A Complete Bibliography of Publications in Nuncius

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

18 October 2024 Version 1.31

Title word cross-reference

$11 [191]. $139.95/£ [2873]. $150 [1944]. $22.50 [2589]. $24.95 [2591]. $26.95 [3281, 345]. $29.95 [2105]. $30.00 [1993]. $35.95 [347]. $36.95 [1993]. $37.50 [2312]. $39 [204, 850]. $39.95 [1911]. $39.95/ [3529]. $39.95/$24.95 [2456]. $40 [3542, 286]. $45 [1845]. $45.60 [3527]. $47.50 [36]. $49 [3543]. $49.50 [851]. $50 [1915, 2585]. $50.00 [2371]. $59 [2375]. $81.00 [26]. $9.95 [613]. ^{∘} [63, 2156, 1391, 2130]. _{2} [2133]. Γ ι [3500]. · [3019]. δ [3500]. ≫ [61, 302, 296, 303, 313, 299, 441, 439, 440, 537, 583, 567, 777, 764, 804, 885, 894, 896, 961, 889, 895, 891, 903, 1005, 995, 1019, 1231, 1288, 1264, 1333, 1365, 1402, 1457, 1406, 1427, 1364, 1454, 1602, 1606, 1587, 1683, 1691, 1657, 1819, 2145, 2159, 2160, 2267, 2421, 2488, 2496, 2495, 1696, 1578, 588, 975, 353, 543, 814, 1235, 2431, 1790, 63, 847, 1552, 72, 85, 58, 309, 443, 442, 444, 446, 447, 445, 449, 448, 576, 574, 966, 887, 964, 873, 886, 890, 830, 915, 991, 996, 1191, 1277, 1278, 1157, 1279, 1280, 1273, 1189, 1187, 1188, 1156, 1359]. ≫ [1358, 1443, 1447, 1445, 1444, 1360, 1448, 1687, 1688, 1685, 1686, 70, 572, 809, 951, 1637, 817, 824, 1147, 1620, 1791, 2161, 2270, 2276, 1401, 334, 926, 367, 547, 408, 386, 1377, 333, 1777, 2126, 343, 1550, 1575, 754, 829, 787, 763, 850, 1190, 1640, 883, 2430, 389, 2258, 2129, 352, 277, 457, 426, 561, 590, 575, 774, 832, 916, 1013, 1015, 1020, 1018, 1033, 1025, 1168, 1161, 1251, 1258, 1162, 1256, 1163, 1241, 1195, 1193, 1160, 1165, 1369, 1355, 1372, 1451, 1452, 1411, 1343, 1424, 1461, 1421, 1435, 1350, 1373, 1423, 1599, 1570]. ≫ [1619, 1665, 2151, 2415, 337, 338, 922, 929, 535, 534, 812, 570, 550, 756, 755, 1414, 1141, 753, 749, 997, 1228, 1229, 1285, 1323, 308, 1230, 2272, 1413, 1579, 1639, 1658, 2132, 2428, 2429, 2499, 1634, 2247, 1153, 2273, 806, 811, 892, 913, 1244, 1202, 1455, 1330, 1803, 2394, 280, 541, 555, 376, 381, 2501, 978, 2138, 51, 1218, 1232, 1233, 1291, 1386, 1363, 1318, 1615, 1565, 1642, 2422, 2498, 1137, 1378, 746, 745, 747, 914, 945, 1224, 1134, 739, 818, 2266, 1294, 3026, 1379, 1135]. ≪ [61, 302, 296, 303, 313, 299, 441, 439, 440, 537, 583, 567, 777, 764, 804, 885, 894, 896, 961, 889, 895, 891, 903, 1005, 995, 1019, 1231, 1288, 1264, 1333, 1365, 1402, 1457, 1406, 1427, 1364, 1454, 1602, 1606, 1587, 1683, 1691, 1657, 1819, 2145, 2159, 2160, 2267, 2421, 2488, 2496, 2495, 1696, 1578, 588, 975, 353, 543, 814, 1235, 2431, 1790, 63, 847, 1552, 72, 85, 58, 309, 443, 442, 444, 446, 447, 445, 449, 448, 576, 574, 966, 887, 964, 873, 886, 890, 830, 915, 991, 996, 1191, 1277, 1278, 1157, 1279, 1280, 1273, 1189, 1187, 1188, 1156, 1359]. ≪ [1358, 1443, 1447, 1442, 1445, 1444, 1360, 1448, 1687, 1688, 1685, 1686, 70, 572, 809, 951, 1637, 817, 824, 1147, 1620, 1791, 2161, 2270, 2276, 1401, 334, 926, 367, 547, 408, 386, 1377, 333, 1777, 2126, 343, 1550, 1575, 754, 829, 787, 763, 850, 1190, 1640, 883, 2430, 389, 2258, 2129, 352, 277, 457, 426, 561, 590, 575, 774, 832, 916, 1013, 1015, 1020, 1018, 1033, 1025, 1168, 1161, 1251, 1258, 1162, 1256, 1163, 1241, 1195, 1193, 1160, 1165, 1369, 1355, 1372, 1451, 1452, 1411, 1343, 1424, 1461, 1421, 1435, 1350, 1373, 1423, 1599]. ≪ [1570, 1619, 1665, 2151, 2415, 337, 338, 922, 929, 535, 534, 812, 570, 550, 756, 755, 1414, 1141, 753, 749, 997, 1228, 1229, 1285, 1323, 308, 1230, 2272, 1413, 1579, 1639, 1658, 2132, 2428, 2429, 2499, 1634, 2247, 1153, 2273, 806, 811, 892, 913, 1244, 1202, 1455, 1330, 1803, 2394, 280, 541, 555, 376, 381, 2501, 978, 2138, 51, 1218, 1232, 1233, 1291, 1386, 1363, 1318, 1615, 1565, 1642, 2422, 2498, 1137, 1378, 746, 745, 747, 914, 945, 1224, 1134, 739, 818, 2266, 1294, 3026, 1379, 1135, 541]. ν [3500]. ω [3500]. [3500]. τ [3500]. ς [2497]. ζ [3500].

* [582].

-Buch [787].

. [1231, 890, 1187, 1401, 2266].

/Diario [1869]. /S [1905]. /The [1671].

0 [2227, 2160, 2312, 2463, 2495, 2464, 2560, 2586, 2601, 2776, 2793, 2953, 2459, 2516, 2717, 2889, 2462, 2606, 2730, 3116, 2669, 2737, 2777, 2538, 2822, 2898, 2707, 2864, 2981, 2664, 2805, 2964, 2580, 3025, 3009, 2850, 2689, 2972, 3268, 2551, 2671, 2791, 2909, 3035, 3030, 2553, 2917, 2371, 2619, 2957, 2311, 2589, 3076, 2824, 1063, 2579, 2834, 2873, 2755, 2962, 2756, 2657, 2594, 3111, 2876, 3003, 2655, 2550, 2774, 2706, 2888, 2751, 2907, 2719, 2827, 2819, 2839, 2600, 2595, 2794, 2893, 2633, 2518, 2800, 2829, 2922, 2947, 2747, 2973, 2783, 2623, 2832, 2656, 2950, 2681, 2854]. 0 [2718, 2862, 2561, 3033, 2709, 2629, 3074, 2736, 2912, 927, 1407, 2966, 2825, 3014, 2365, 3069, 2940, 2790, 2801, 2505, 2775, 2799, 3036, 2310, 2792, 2858, 2686, 2584, 3224, 2745, 2753, 2976, 2617, 2731, 2833, 3008, 2676, 2552, 2749, 2841, 1229, 1285, 2840, 2643, 2678, 2366, 2412, 2506, 2559, 2673, 3039, 2807, 2803, 2316, 2384, 2608, 2585, 2939, 2728, 2786, 2457, 2804, 2609, 2784, 2845, 2626, 2489, 2378, 808, 993, 2806, 2967, 2687, 2591, 2861, 2625, 2587, 2582, 2901, 3527, 2846, 2752, 2796, 2545, 2674, 2739, 2534, 2724, 2266, 2772, 2562, 2848, 2802, 2913, 2688]. 0 [2578, 2692, 2417]. 0-19-282025-7 [1907]. 0-19-516004-5 [2606]. 0-19-516598-5 [2595]. 0-19-517127-6 [2681]. 0-19-517734-7 [2600]. 0-19-726284-8 [2686]. 0-19-815065-2 [2706]. 0-19-815248-5 [2459]. 0-19-823724-3 [1843]. 0-19-850031-9 [1844]. 0-19-851544-8 [2848]. 0-19-852049-2 [1007]. 0-19-856484-8 [2717]. 0-19-856566-6 [2310]. 0-19-924233-X [2378]. 0-19-926326-4 [2687]. 0-19-926988-2 [2582]. 0-19-927117-8 [2609]. 0-19-927136-4 [2587]. 0-19-927668-4 [2861]. 0-19-927905-5 [2830]. 0-19-929644-8 [2947]. 0-226-01055-4 [2783]. 0-226-06860-9 [2747]. 0-226-10921-6 [2850]. 0-226-11378-7 [2739]. 0-226-12111-9 [2669]. 0-226-13681-7 [2559]. 0-226-13948-4 [2862]. 0-226-16056-4 [2424]. 0-226-16472-1 [2845]. 0-226-19994-0 [2800]. 0-226-20487-1 [927]. 0-226-26147-6 [2371]. 0-226-28415-8 [2551]. 0-226-35933-6 [2688]. 0-226-40955-4 [2772]. 0-226-48231-6 [993]. 0-226-48722-9 [2562]. 0-226-48733-4 [2805]. 0-226-54202-5 [2676]. 0-226-57696-5 [2718]. 0-226-57701-5 [2834]. 0-226-61668-1 [2976]. 0-226-62087-5 [2775]. 0-226-64201-1 [2560]. 0-226-64847-8 [2655]. 0-226-73569-9 [2807]. 0-226-75020-5 [1722]. 0-226-75709-9 [2841]. 0-226-76733-7 [2858]. 0-231-13588-2 [2678]. 0-252-03156-3 [2776]. 0-253-21656-7 [2561]. 0-262-10116-5 [2913]. 0-262-11273-6 [2730]. 0-262-12258-8 [2664]. 0-262-12269-3 [2689]. 0-262-16229-6 [2656]. 0-262-19421-X [2120]. 0-262-19516-X [2731]. 0-268-03071-5 [2585]. 0-268-03484-2 [2774]. 0-271-01662-0 [1910]. 0-271-02052-0 [2312]. 0-271-02459-3 [2534]. 0-292-70981-1 [2824]. 0-300-07383-6 [2056]. 0-300-10075-2 [2626]. 0-393-04086-0 [1853]. 0-393-06019-5 [2678]. 0-415-03732-8 [1808]. 0-415-15672-6 [2366]. 0-415-21575-7 [2601]. 0-415-23847-1 [2417]. 0-415-94293-4 [2412]. 0-471-11459-6 [1993]. 0-471-75154-5 [2864]. 0-472-11228-7 [2489]. 0-472-11602-9 [2950]. 0-520-06283-3 [770]. 0-520-20010-1 [1845]. 0-520-20548-0 [1987]. 0-520-20646-0 [1987]. 0-520-23325-5 [2516]. 0-520-23793-5 [2671]. 0-520-24064-2 [2382]. 0-520-24135-5 [2784]. 0-521-08732-5 [2629].

0-521-20072-5 [1154]. 0-521-25586-4 [851]. 0-521-25625-9 [849]. 0-521-30694-9 [803]. 0-521-30984-0 [860]. 0-521-32831-4 [808]. 0-521-33511-6 [848]. 0-521-35168-5 [1670]. 0-521-35371-8 [1997]. 0-521-35733-0 [806]. 0-521-36021-8 [850]. 0-521-38470-2 [1203]. 0-521-38564-4 [1407]. 0-521-40371-5 [1334]. 0-521-43172-7 [1656]. 0-521-44171-4 [1654]. 0-521-45518-9 [1758]. 0-521-47565-1 [2909]. 0-521-53026-1 [2752]. 0-521-55213-3 [1654]. 0-521-55214-1 [1654]. 0-521-56004-7 [2066]. 0-521-56585-5 [2075]. 0-521-57243-6 [2464]. 0-521-58127-3 [2106]. 0-521-58142-7 [1800]. 0-521-58351-9 [1991]. 0-521-59032-9 [2004]. 0-521-62672-2 [2377]. 0-521-63149-1 [2003]. 0-521-64029-6 [2134]. 0-521-64627-8 [1575]. 0-521-65216-2 [2457]. 0-521-77118-8 [2518]. 0-521-79143-X [2152]. 0-521-80602-X [2362]. 0-521-81230-5 [2427]. 0-521-81251-8 [2470]. 0-521-81516-9 [2791]. 0-521-83122-9 [2790]. 0-521-83637-9 [2737]. 0-521-84455-X [2709]. 0-674-01495-2 [2591]. 0-674-01519-3 [2674]. 0-674-01687-4 [2793]. 0-674-01892-3 [2728]. 0-674-02185-1 [2846]. 0-674-02328-5 [3035]. 0-674-02615-2 [2966]. 0-674-09745-9 [2102]. 0-674-58220-9 [1847]. 0-691-00699-7 [2207]. 0-691-01099-4 [1725]. 0-691-01206-7 [2105]. 0-691-01687-9 [1846]. 0-691-05939-X [2061]. 0-691-09685-6 [2463]. 0-691-11398-X [2506]. 0-691-11541-9 [2538]. 0-691-11849-3 [2538]. 0-691-12635-6 [2792]. 0-7123-0894-6 [2617]. 0-7167-3103-7 [2545]. 0-7181-3882-1 [1914]. 0-7190-7373 [2901]. 0-7546-0139-0 [2586]. 0-7546-0506-X [2940]. 0-7546-0960-X [2751]. 0-7546-3718-2 [2633]. 0-7546-3933-9 [2755]. 0-7546-3996-7 [2803]. 0-7546-4102-3 [2801]. 0-7546-5020-0 [2802]. 0-7546-5048-0 [2707]. 0-7546-5156-8 [2962]. 0-7546-5259-9 [2981]. 0-7546-5296-3 [2819]. 0-7546-5568-7 [2873]. 0-7546-5609-8 [2832]. 0-7923-4361-1 [1857]. 0-7923-6782-0 [2462]. 0-8014-8901-6 [2589]. 0-8018-4483-5 [1329]. 0-8018-5223-4 [1837]. 0-8018-5941-7 [2016]. 0-8018-7991-4 [2794]. 0-8018-8055-6 [2584]. 0-8018-8235-4 [2786]. 0-8018-8400-4 [2799]. 0-8018-8401-2 [2799]. 0-8018-8426-8 [2825]. 0-8020-9018-4 [2873]. 0-8020-9064-8 [2953]. 0-8078-5678-9 [2964]. 0-8122-3827-3 [2657]. 0-8133-1575-1 [1763]. 0-8201-1555-X [2957]. 0-8203-1864-7 [1915]. 0-8203-1865-5 [1915]. 0-8223-1204-2 [1144]. 0-8223-1759-1 [1841]. 0-8223-1768-0 [1841]. 0-85115-285-6 [1063]. 0-85115-798-X [2143]. 0-85296-792-6 [2381]. 0-85296-797-7 [2381]. 0-85296-845-0 [1733]. 0-85323-630-5 [1818]. 0-85481-125-7 [2308]. 0-86078-633-1 [2079]. 0-86078-781-8 [2098]. 0-86078-924-1 [2749]. 0-86078-958-6 [2756]. 0-86159-126-7 [1872]. 0-86341-234-3 [847]. 0-86341-250-5 [1732]. 0-86341-251-3 [2080]. 0-87081-722-1 [2505]. 0-87169-220-1 [1819]. 0-87169-235-X [2137]. 0-87169-933-8 [2495]. 0-88135-074-5 [1911]. 0-88135-133-4 [1898]. 0-88135-252-7 [2745]. 0-88135-273-X [2059]. 0-88135-275-6 [2907]. 0-88135-285-3 [2160]. 0-88135-354-X [2917]. 0-88135-356-6 [2839]. 0-88135-365-5 [2972]. 0-88135-372-8 [2553]. 0-88135-376-0 [2594].

0-902205-58-7 [2311]. 0-911221-23-9 [2227]. 0-930405-87-0 [2753]. 0-941901-38-6 [2724]. 0-948382-07-4 [1085]. 0-9536522-4-6 [2365]. 0-9546484-1-2 [2579]. 00 [2488, 2491, 2504, 2492, 1944, 2419, 2499, 2507, 2503, 2416]. 000-year-old-computer [3919]. 0066 [3258]. 0066-6815 [3258]. 0223 [2201]. 0223-5099 [2201]. 0249-7980 [1981]. 0369-6243 [1871]. 0394 [2524]. 0394-4409 [800]. 0394-9001 [2524]. 0418-4459 [1228]. 06 [2541].

1 [714, 1288, 2194, 2228, 2318, 2500, 2560, 2527, 2746, 2986, 2418, 2708, 2710, 3286, 2777, 2936, 3071, 2522, 2770, 3774, 3526, 2909, 2542, 3001, 1151, 2592, 2156, 2605, 2824, 3205, 3127, 3043, 2546, 3087, 2572, 2646, 2719, 3530, 2934, 2946, 3853, 2654, 2740, 2757, 2557, 2835, 2781, 2596, 927, 2945, 3018, 2505, 2684, 2660, 2880, 2785, 2855, 2976, 3114, 2583, 2552, 2749, 2645, 3125, 2618, 2867, 2985, 2380, 2845, 2274, 2856, 2911, 2921, 2549, 2625, 2846, 2752, 2604, 2779, 2628, 2971, 3285, 2417]. 1-288 [2501]. 1-4020-0081-2 [2708]. 1-4020-1631-X [2583]. 1-58465-512-7 [2867]. 1-84046-468-2 [2740]. 1-84046-486-0 [2777]. 1-84046-504-2 [2546]. 1-84046-536-0 [2625]. 1-84046-540-9 [2542]. 1-84046-546-8 [2604]. 1-84046-573-5 [2527]. 1-84171-206-X [2204]. 1-84306-045-0 [2899]. 1-84306-046-9 [2899]. 1-85072-204-8 [1877]. 1-85196-778-8 [2781]. 1-85196-858-X [2956]. 1-85619-604-6 [1945]. 1-85858-269-5 [2779]. 1-85973-872-9 [2746]. 1-86508-658-4 [2549]. 1-873992-99-8 [2710]. 1-878220-76-4 [2971]. 1-884964-29-X [2194]. 1-890951-43-9 [2684]. 1-890951-67-6 [2835]. 1-905375-05-0 [2719]. 1-931968-02-0 [2552]. 1/11 [1279]. 1/5* [720]. 10. [1951, 2022]. 10.24894 [3530]. 10.24894/978 [3530]. 10.24894/978-3-7965-4155-1 [3530]. 10/17 [757]. 10/1996 [2009]. 100.000 [1995]. 103 [1091]. 103-529 [1091]. 109 [1111]. 10th [1429]. 11. [1952, 2023]. 110 [2095]. 110e [243]. 111 [1765]. 1123-2560 [2628]. 113e [836]. 1158-8403 [3280]. 116e [1327, 4042]. 12. [1953, 2024]. 120.000 [970, 1642]. 1200 [2137]. 124 [233]. 1270 [2520]. 129 [416]. 13. [1954, 2025]. 130.000 [1716, 1904]. 13th [1985]. 14 [2433]. 14. [1955, 2026]. 14.95 [1841]. 140 [701]. 1421/2000 [2199]. 145.000 [1431]. 1450 [3906]. 1460 [251]. 1471 [2308]. 1471-5841 [2308]. 14th [1985]. 15-1 [2267]. 15-2 [2267]. 15-II-III [2370]. 15. [1956, 2027]. 15.000 [421, 81]. 150 [1270]. 150-HCI. [1270]. 1536 [1192]. 155 [2461]. 1550 [2819]. 157 [726]. 1575 [3734]. 1575-1639 [3734]. 1576 [2139]. 1578 [3844]. 1597 [2138]. 15th [3557, 2130]. 15th-Century [3557]. 16. [1957, 2028]. 16.50EUR [3426]. 160.- [1838]. 160.000 [2012]. 1600 [3254, 2114, 2369, 4091, 3530]. 1610 [3225]. 1630 [3411]. 1639 [3734]. 165 [1065]. 1666 [2654]. 1669 [2333, 2284]. 168 [950]. 1685 [2947]. 1689-Mars [516]. 1698 [93]. 16th [519, 1860, 3759, 3609, 4017].

17. [1958, 2029]. 170 [1878]. 1700 [2689, 2583]. 1700-talets [1146]. 1700s [1146]. 1720 [3542]. 1726 [180]. 1727 [919]. 1738 [1681]. 1743-1794 [3739]. 1750 [1063, 2098, 3508]. 1774 [2899]. 1777 [2529]. 1780 [3265]. 1780-1820 [3265]. 1785 [2899, 412]. 1789 [1090]. 1789-1914 [3187]. 1789/1790 [565].

1790 [2732]. 1791 [2665]. 1793 [2495]. 1793/94 [1098]. 1794 [2727, 1440]. 17me [210]. 17th [675, 1387, 3205, 26, 1378, 3324, 1213]. 18 [4040]. 18. [1959, 2030]. 18.000 [1115]. 18.50EUR [3426]. 1800 [2675, 3070, 3131, 2828, 1328, 3398]. 1800-tal [2675]. 1800-talets [2500]. 1802 [1495]. 1804 [2534]. 1806 [2609]. 1811 [2605]. 1815 [2375, 847]. 1819 [2899]. 1820 [3265, 3186, 3509, 3691]. 1820-1848 [3691]. 1830 [762, 2956, 728]. 1831 [1006]. 1832 [1912]. 1835 [2957]. 1836 [1408]. 1840 [1731]. 1840/44 [1731, 1731]. 1841 [2332, 4074, 2283]. 1842 [2354, 174]. 1848 [4074, 3691]. 1855 [1020]. 1856 [2976]. 1857 [3030, 519, 851]. 1860 [2071, 1806]. 1865 [3210]. 1870 [3023, 2079, 2866]. 1872 [927]. 1874 [2979]. 1880 [2059, 779]. 1887 [45]. 1890 [415, 3999]. 18th [685, 2228, 1060, 4049, 1067, 2161, 2962, 4072, 3902, 26, 3944, 3616, 4006, 1294]. 18th- [2962]. 18th-Century [1067]. 19. [1960, 2031]. 19.000 [1096]. 19.500 [433]. 19.50EUR [3426]. 1900 [2803, 2861, 3483]. 1900-1925 [3483]. 1901 [4032]. 1914 [3083, 971]. 1918 [1340, 850]. 1920 [2852, 1797, 3977]. 1920s [3839]. 1925 [3263, 3483]. 1930 [1854]. 1930s [3839]. 1931 [3978]. 1934 [2804]. 1939 [2984]. 1940 [2273, 3414, 3675]. 1942 [3369]. 1945 [2670, 2633, 3638]. 1949 [2912]. 1950 [1449, 1422]. 1950s [4054, 4055, 860]. 1952-92 [1255]. 1956 [1372]. 1957 [552]. 1960 [2613, 3639]. 1965 [2856]. 1969 [2075]. 197-227 [35]. 1974 [10]. 1980 [348]. 1983 [176]. 1984 [63, 226]. 1984/85 [492]. 1985 [75, 171, 3967, 12, 177, 2492]. 1985-86 [186]. 1986 [74, 420, 66, 83, 674]. 1987 [386, 333]. 1988 [1094, 713]. 1988-89 [673]. 1989 [620, 643, 839, 999, 751, 694, 915]. 1990 [648, 798]. 1991 [788, 4077, 1018, 1058]. 1992 [2885, 1160]. 1993 [1600, 1488, 1237]. 1994 [4040, 1661, 1265, 1408, 4089, 1608, 1331]. 1995 [1655, 1487]. 1995-96 [1750]. 1996 [4043, 2161, 4066, 1731, 4082, 1654, 1803, 2201, 1916]. 1996-vol [1871]. 1997 [4045, 1984, 4076, 1913, 4078, 2213]. 1998 [2169, 2240, 2093, 4052, 2078, 2118, 2146]. 1999 [2229, 2050, 2113, 2195, 2217, 4058]. 19os [1930]. 19th [3759, 1387, 2295, 2089, 2962, 3104, 3079, 2765, 1936, 545, 3617, 4049, 3472, 4062, 1121, 3488, 4006]. 19th-Century [2962, 3472]. 1er [502]. 1st [824].

2 [2200, 2318, 2496, 2916, 2708, 2843, 2607, 2878, 2942, 3120, 2896, 2735, 2898, 2852, 3773, 3072, 2528, 2837, 2906, 2798, 2309, 2588, 2826, 2974, 2648, 2156, 2579, 3027, 3066, 2519, 2546, 2550, 2774, 2706, 3852, 2633, 2654, 2829, 2740, 2757, 2892, 2859, 2596, 2533, 2977, 2966, 2375, 2468, 3031, 3546, 3251, 2821, 2615, 2693, 2541, 2548, 2624, 2537, 2767, 2678, 3241, 2539, 2729, 2457, 3019, 2626, 2198, 2591, 3086, 2856, 2582, 3278, 4008, 2613, 2748, 2578, 3026]. 2-02-015919-8 [1668]. 2-05-101951-7 [2539]. 2-07-072936-2 [2537]. 2-13-049107-3 [1878]. 2-13-049345-9 [1875]. 2-13-050092-7 [2060]. 2-251-79983-4 [2892]. 2-296-01033-4 [2896]. 2-503-50739-5 [2145]. 2-503-51008-6 [2309]. 2-503-51196-1 [2596]. 2-503-51197-X [2267]. 2-503-51218-6 [2369]. 2-503-51642-4 [2878]. 2-503-52160-6 [2615]. 2-503-52442-7 [3026]. 2-600-00805-5 [2843]. 2-606-01215-1 [2856].

2-7011-2404-2 [2197]. 2-7011-3076-X [2422]. 2-7071-3616-6 [2273]. 2-7116-1064-0 [1667]. 2-7116-1172-8 [1228]. 2-7116-1629-0 [2578]. 2-7116-1754-8 [2648]. 2-7116-1823-4 [2767]. 2-7116-1828-5 [2821]. 2-7116-9755-X [1326]. 2-7132-2015-7 [2613]. 2-7283-0545-5 [2201]. 2-7283-0740-7 [2826]. 2-7288-0356-0 [2898]. 2-7332-1026-2 [2528]. 2-7355-0272-4 [1327]. 2-7384-9903-1 [2318]. 2-7430-0319-7 [2118]. 2-7453-1246-4 [2729]. 2-7453-1363-0 [2829]. 2-7475-7691-4 [2735]. 2-7475-8650-2 [2798]. 2-7606-1822-6 [2375]. 2-8031-0190-4 [2469]. 2-85367-205-0 [1795]. 2-85367-208-5 [2215]. 2-85367-230-1 [2654]. 2-85704-926-9 [2533]. 2-86847-996-0 [2550]. 2-87009-765-4 [2428]. 2-87192-006-0 [295]. 2-88074-571-3 [2519]. 2-88124-400-9 [892]. 2-89448-320-1 [2423]. 2-908327-41-4 [2229]. 2-908965-11-9 [2013]. 2-908965-14-3 [2015]. 2-909344-09-06 [2541]. 2-912450-07-1 [2757]. 2-930175-00-1 [1859]. 2-930175-02-8 [1734]. 2/1 [1435]. 2/3 [1857]. 20 [763, 1914]. 20. [1961, 2032]. 20.000 [340, 336, 1028, 697]. 2000 [2200, 2909, 4059, 4067, 2970, 2363]. 2000s [4038]. 2001 [2581]. 2002 [2376, 3917]. 2003 [2611]. 2005 [2830]. 2007 [3092]. 2008 [3004]. 200th [2724]. 2011 [3156]. 2012 [3194, 3237]. 207-2612-2 [1908]. 2089-1 [955]. 20th [1264, 2228, 4041, 2161, 2156, 4072, 3407, 1936, 2216, 2765, 3902, 3378, 1378]. 20th-Century [1264, 2765]. 21. [1962, 2033]. 213-319 [774]. 215 [1372]. 215-346 [1372]. 21cm [3280]. 21st [3385]. 21st-Century [3385]. 22 [4067].

22. [1963, 2034]. 22.000 [877, 1754, 1650, 346]. 220.000 [2107]. 224 [309]. 23 [1657, 2015]. 23. [2327]. 24 [244]. 24-26 [3967]. 24. [2328]. 24.000 [330, 611]. 241pp [3541]. 243pp [3542]. 246 [791]. 25 [17, 492]. 25. [2329]. 25.00 [25]. 25.000 [493, 953]. 251 [876]. 2519/SFR [1731]. 26 [1489]. 26. [2330]. 26.000 [1148, 2222]. 28 [2497]. 28. [2331]. 28.- [1905]. 28.-/S [1905]. 28.000 [932, 79, 1140]. 29. [2332]. 291 [490]. 295 [2265]. 2ème [868]. 2nd [580, 3448].

3 [2228, 2160, 2776, 2647, 3750, 3094, 3005, 2517, 2822, 2933, 3040, 2851, 3280, 3282, 2653, 3776, 2750, 2682, 2837, 2959, 2689, 3204, 2853, 2908, 3082, 3021, 1151, 2156, 2727, 2789, 2920, 3013, 2831, 2519, 2657, 2646, 3061, 2819, 2742, 3070, 2999, 3253, 2720, 2652, 2268, 2532, 2854, 2859, 3024, 4074, 3091, 2196, 2736, 2598, 2659, 2713, 2525, 2801, 3277, 2574, 2529, 2785, 1074, 3078, 2975, 2725, 2875, 2576, 1285, 3083, 2643, 2663, 2734, 2673, 3130, 2728, 2804, 3528, 2466, 2465, 2544, 2205, 3037, 2674, 3239, 2534, 2857]. 3-03910-299-0 [2804]. 3-05-003848-9 [2576]. 3-11-017966-0 [2822]. 3-335-00500-7 [1992]. 3-406-52891-0 [2643]. 3-447-05337-2 [2837]. 3-484-36589-7 [2831]. 3-486-26308-0 [1862]. 3-515-06096-0 [1285]. 3-515-07079-6 [1912]. 3-515-07252-7 [1838]. 3-515-07561-5 [2055]. 3-515-08185-2 [2859]. 3-518-57688-7 [415]. 3-525-21018-3 [2750]. 3-540-20068-1 [2646]. 3-540-21961-7 [2653]. 3-540-55341-X [1074]. 3-7001-2120-2 [1854]. 3-7643-6584-6 [2647]. 3-7643-7007-6 [2725]. 3-7692-2038-2 [1715].

3-7705-4101-4 [2682]. 3-7728-1483-2 [1202]. 3-7728-2034-4 [2465]. 3-7728-2255-X [2734]. 3-7913-3664-9 [2908]. 3-7930-9448-0 [2854]. 3-7965-1324-7 [2205]. 3-7965-1325-5 [2529]. 3-7965-1327-1 [2785]. 3-8047-1050-6 [1797]. 3-8047-1447-1 [1905]. 3-8353-0159-7 [2853]. 3-88309-030-1 [1151]. 3-89241-022-4 [2161]. 3-89241-024-0 [2161]. 3-89241-029-1 [2228]. 3-89244-629-6 [2532]. 3-89244-649-0 [2673]. 3-89244-893-0 [2736]. 3-89665-174-5 [2517]. 3-925347-73-9 [2544]. 3-928186-51-5 [2196]. 3-928186-56-6 [2466]. 3-929360-70-5 [2598]. 30. [1965, 2036, 2333]. 30.000 [2154, 2220, 435, 165]. 3032 [2080]. 31. [1966, 2037, 2334]. 31.000 [417]. 31.500 [707]. 315 [920]. 31st [3591]. 32. [1967, 2038, 2335]. 32.000 [18, 437, 2108]. 323 [774]. 323-436 [774]. 3249-3792 [994]. 33. [2336]. 336 [419]. 34 [220]. 34. [2337]. 34.000 [1112]. 345/ [1731]. 346 [1372]. 35.000 [1201, 678]. 351 [3243]. 355 [1436]. 36.000 [593, 1903]. 37. [2163]. 373-730 [878]. 3793-4411 [994]. 38. [2164]. 38.000 [300, 2133]. 38/DM [172]. 39. [2165]. 397 [366]. 3rd [824].

4 [2467, 2793, 3164, 2878, 2425, 2896, 2735, 2805, 3080, 2682, 2836, 2979, 3123, 2616, 519, 2958, 3022, 2883, 2873, 3115, 3256, 2874, 2988, 3108, 2794, 3289, 2905, 2690, 2904, 2787, 2783, 2892, 2862, 2494, 2970, 1407, 554, 4077, 2987, 2799, 2554, 3208, 2221, 3932, 997, 2767, 2412, 3226, 2984, 2729, 2679, 2786, 3044, 808, 2465, 2919, 2603, 2687, 3075, 2861, 2914, 2549, 2587, 3259, 2627, 2772, 2971, 3264]. 40 [2106, 411]. 40. [2166]. 40.000 [1190, 360, 1101, 428, 281]. 405 [308]. 407-462 [893]. 41. [2167]. 42. [2326]. 42.000 [1092]. 42.500 [288]. 43. [2386]. 43.000 [924]. 436 [774]. 437 [306]. 44 [1731]. 44. [2387]. 45. [2388]. 45.00 [489]. 450 [993]. 457-505 [720]. 46 [172]. 46. [2389]. 48. [2435]. 48.000 [1075, 1990]. 4e [954, 713].

5 [2527, 3016, 2459, 2516, 2843, 2606, 2517, 2530, 3063, 3059, 2864, 2320, 2471, 2972, 3129, 2671, 3035, 2963, 2535, 2834, 3298, 2536, 3090, 2379, 3012, 2941, 2948, 2595, 2642, 2771, 2773, 2718, 2629, 2196, 2590, 2598, 2275, 2775, 3426, 2821, 2529, 3088, 2676, 2585, 2597, 3096, 2784, 3203, 3261, 2324, 2622, 2779, 2556, 2913, 2571, 3089]. 5* [720]. 50 [2715, 2490, 204]. 50-5356-787-6 [2715]. 50.000 [1200, 930, 2011, 1180, 556, 1159]. ’500 [1]. 500s [4038]. 5099 [2201]. 52.50 [2079]. 529 [1091]. 54.000 [1917, 1809]. 55.000 [947]. 5841 [2308]. 585-754 [12]. 5XE [2683].

6 [2463, 2464, 2715, 3041, 2647, 2730, 3064, 3255, 3254, 3541, 2472, 3545, 3785, 2850, 3023, 2865, 2309, 3122, 2685, 2670, 3544, 2371, 2602, 2589, 1063, 2577, 2978, 2756, 2895, 2907, 2839, 2885, 3131, 2651, 2835, 2656, 2532, 2681, 3017, 2938, 2365, 2375, 3542, 2555, 2615, 3257, 2310, 2732, 2792, 2584, 2617, 2722, 2725, 2366, 2374, 3062, 3206, 3058, 993, 2466, 2724, 3073, 2688]. 6. [1947, 2018]. 6.95 [1085]. 60.000 [200, 2002, 1796]. ’600 [515, 1, 1927, 3906]. 60856 [3068]. 63 [2432]. 65.000 [1718]. 6815 [3258].

7 [2601, 3299, 2573, 2897, 3107, 2612, 3121, 3109, 2653, 2631, 2721, 3015, 2853, 3084, 2826, 2321, 2311, 3263, 3006, 2520, 2831, 2838, 2531, 3060, 57, 2943, 2632, 2600, 2955, 2672, 3288, 2561, 2323, 2965, 2373, 3529, 2741, 2884, 2944, 3267, 3093, 2858, 2677, 3038, 2745, 2844, 3118, 3284, 2559, 2867, 2866, 3085, 2803, 2539, 2951, 2489, 3126, 2982, 2890, 2903, 2820, 2935, 2205, 2810, 2613, 2545, 2739, 2910, 2644, 2367, 2668, 3026]. 7. [1948, 2019]. 70.000 [559, 558, 186]. ’700 [170, 1893, 2174, 3564]. 7373 [2901]. 75.000 [589]. 79* [1731].

8 [2495, 2953, 2710, 3000, 2717, 3065, 3113, 3283, 2894, 3081, 2664, 2667, 2769, 3279, 2523, 2551, 2969, 2648, 2553, 2665, 2658, 2863, 2962, 3209, 2655, 2518, 2712, 2947, 2781, 3067, 2832, 2825, 2754, 2686, 2809, 3010, 2738, 2983, 2609, 3095, 2902, 2604, 2848, 2726, 2960]. 8. [1949, 2020]. 80.000 [2010, 912, 958, 1098]. ’800 [897, 1867, 879]. 80th [4078]. 84 [2808, 2494, 2593]. 84-00-08137-4 [2494]. 84-323-0802-1 [1515]. 84-370-4219-4 [2114]. 84-482-3709-9 [2593]. 84-600-4005-4 [43]. 84-7820-814-3 [2808]. 84-7820-815-1 [2808]. 84-7820-816-X [2808]. 84-87369-03-0 [1860]. 84-88639-32-5 [1849]. 84-89652-87-2 [2068]. 85 [2198]. 85-7291-049-2 [2198]. 85.000 [563, 1665]. 85181 [2879]. 86-82977-02-8 [2069]. 87 [2418, 2658]. 87-7289-815-1 [2418]. 87-7289-817-8 [2658]. 87-fasc [1871]. 88 [2467, 3065, 2425, 2612, 2530, 2522, 2320, 2471, 2780, 2580, 2472, 2543, 2667, 2836, 2631, 2721, 2797, 2865, 2314, 2769, 2523, 2588, 2849, 2685, 2616, 2558, 2670, 2963, 2602, 2619, 2665, 2592, 2232, 2321, 2605, 2727, 2789, 2535, 2883, 2577, 2863, 3027, 3006, 2520, 2536, 2379, 2531, 2572, 2632, 2742, 2672, 2614, 2712, 2642, 2690, 2630, 2651, 2904, 2787, 2557, 2623, 2720, 2652, 2773, 2323, 2590, 2754, 2965, 2275, 2413, 2430, 2373, 2468, 2659, 2713, 2525, 2741, 2743, 2693, 2880, 2221, 2225, 2548, 2574, 2877, 2624, 2855, 2677, 2716, 2809, 2722, 2645, 2272, 2426, 2618]. 88 [2738, 2380, 2316, 2384, 2374, 2608, 2679, 2597, 2603, 2274, 2914, 2269, 2903, 2902, 2820, 2501, 2691, 2324, 2796, 2810, 2714, 2622, 2644, 2628, 2322, 2367, 2266, 2627, 2748, 2726, 2571, 2668, 2692]. 88-00-86088-5 [2642]. 88-04-35111-X [1112]. 88-06-12446-3 [1642]. 88-06-13519-8 [1644]. 88-06-13866-9 [2426]. 88-06-15051-0 [2384]. 88-06-16098-2 [2693]. 88-06-16754-5 [2963]. 88-06-17235-2 [2748]. 88-08-08931-2 [2548]. 88-09-03811-8 [2769]. 88-15-02809-9 [811]. 88-15-03267-3 [913]. 88-15-04157-5 [1909]. 88-15-07665-4 [2315]. 88-15-08676-5 [2471]. 88-15-09730-9 [2543]. 88-15-09759-7 [2721]. 88-15-09926-3 [2789]. 88-15-11277-4 [2836]. 88-17-87169-9 [2503]. 88-203-2762-7 [2156]. 88-203-2789-9 [2156]. 88-203-2790-2 [2156]. 88-203-2791-0 [2156]. 88-204-9708-5 [1917]. 88-204-9853-7 [1820]. 88-204-9854-5 [1820]. 88-205-0798-6 [1869]. 88-207-3704-3 [2659]. 88-209-7242-5 [2379]. 88-218-0950-1 [2713]. 88-218-0963-3 [2713]. 88-220-0543-0 [1836]. 88-220-0559-7 [2810]. 88-222-3743-9 [800]. 88-222-4660-8 [1966]. 88-222-4842-2 [2222]. 88-222-5075-3 [2372]. 88-222-5150-4 [2425].

88-222-5168-7 [2499]. 88-222-5187-3 [2574]. 88-222-5214-4 [2467]. 88-222-5269-1 [2572]. 88-222-5284-5 [2530]. 88-222-5319-1 [2488]. 88-222-5322-1 [2522]. 88-222-5331-0 [2608]. 88-222-5335-3 [2525]. 88-222-5351-5 [2571]. 88-222-5378-7 [2531]. 88-222-5394-9 [2614]. 88-222-5401-5 [2590]. 88-222-5416-3 [2652]. 88-222-5419-8 [2809]. 88-222-5467-8 [2754]. 88-222-5484-8 [2726]. 88-222-5493-7 [2903]. 88-222-5508-9 [2780]. 88-222-5554-2 [3027]. 88-222-5566-6 [2865]. 88-240-3867-0 [1665]. 88-245-0639-9 [2413]. 88-272-0482-2 [1868]. 88-299-1487-8 [2107]. 88-304-1758-0 [2108]. 88-317-8329-7 [2491]. 88-339-1030-X [1998]. 88-339-1055-5 [1921]. 88-339-1089-X [1922]. 88-339-1091-1 [1899]. 88-339-1435-6 [2472]. 88-339-1457-7 [2632]. 88-339-1470-4 [2904]. 88-339-1573-5 [2535]. 88-339-1579-4 [2690]. 88-339-1612-X [2630]. 88-339-1644-8 [2738]. 88-339-5640-7 [2111]. 88-350-9212-4 [1713]. 88-359-5160-7 [2323]. 88-359-5396-0 [2504]. 88-389-1450-8 [2062]. 88-392-0518-7 [2321]. 88-395-8300-9 [2232]. 88-404-5066-1 [2592]. 88-420-5537-9 [1876]. 88-420-5629-4 [2364]. 88-420-5929-3 [2274]. 88-420-5973-0 [2383]. 88-420-6139-5 [2320]. 88-420-6155-7 [2431]. 88-420-6225-1 [2274]. 88-420-7472-1 [2618]. 88-420-7516-7 [2668]. 88-424-9341-4 [1900]. 88-428-0778-8 [2433]. 88-430-1448-X [2206]. 88-430-1546-X [2269]. 88-430-2142-7 [2416]. 88-430-2321-7 [2672]. 88-430-2818-9 [2501]. 88-430-3108-2 [2624]. 88-430-3114-7 [2631]. 88-430-3414-6 [2651]. 88-435-8725-0 [1926]. 88-452-3370-7 [2644]. 88-452-4665-5 [2276]. 88-464-0419-X [1917]. 88-464-1385-7 [1988]. 88-464-2286-4 [2203]. 88-464-3692-X [2430]. 88-464-5449-9 [2716]. 88-464-7132-6 [2722]. 88-464-7290-X [2714]. 88-464-7823-1 [2855]. 88-464-7937-8 [2902]. 88-467-0441-X [2272]. 88-491-1284-X [2322]. 88-491-1650-0 [2213]. 88-491-1651-9 [2213]. 88-491-1940-2 [2468]. 88-491-2126-1 [2460]. 88-495-0256-7 [2370]. 88-495-0414-4 [2370]. 88-495-1060-8 [2712]. 88-498-0769-4 [2616]. 88-500-237-8 [2523]. 88-6074-044-4 [2883]. 88-7078-619-6 [2132]. 88-7078-693-5 [2324]. 88-7088-301-9 [2157]. 88-7088-519-4 [2914]. 88-7145-141-4 [1861]. 88-7145-146-5 [1931]. 88-7145-160-0 [2149]. 88-7145-167-8 [2149]. 88-7145-169-4 [2221]. 88-7145-180 [2414]. 88-7145-188-0 [2414]. 88-7146-356-0 [1851]. 88-7146-437-0 [2316]. 88-7158-132-6 [2670]. 88-7188-593-7 [2429]. 88-7233-016-5 [2275]. 88-7246-211-8 [1768]. 88-7284-597-1 [1850]. 88-7395-012-4 [2603]. 88-7395-094-9 [2691]. 88-7398-001-5 [2622]. 88-7498-028-0 [2619]. 88-7642-146-7 [2677]. 88-7644-298-7 [2367]. 88-7686-977-8 [1906]. 88-7686-980-8 [2461]. 88-7689-192-7 [2741]. 88-7794-091-3 [1745]. 88-7794-272-X [2225]. 88-7804-144-0 [1863]. 88-7970-053-7 [2012]. 88-7989-595-8 [2432]. 88-7989-891-4 [2627]. 88-7997-063-1 [2380]. 88-8086-497-1 [2492]. 88-8086-556-0 [2580]. 88-8091-370-0 [1746]. 88-8091-519-3 [1923]. 88-8096-448-8 [1767]. 88-8114-538-3 [1995]. 88-8147-136-1 [1994]. 88-8219-375-6 [2153]. 88-8231-075-2 [2224].

88-8232-044-8 [2144]. 88-8272-182-5 [2597]. 88-8304-267-0 [2266]. 88-8304-321-9 [2270]. 88-8304-398-7 [2373]. 88-8334-079-5 [2420]. 88-8335-539-3 [2742]. 88-8335-777-9 [2743]. 88-8344-008-0 [2150]. 88-8353-111-6 [2374]. 88-8354-002-6 [2602]. 88-8354-003-4 [2787]. 88-8358-546-1 [2557]. 88-8409-068-7 [2612]. 88-8422-180-3 [2727]. 88-8422-257-5 [2536]. 88-8422-362-8 [2665]. 88-8443-115-8 [3065]. 88-8447-218-0 [2796]. 88-8449-225-4 [2679]. 88-8450-039-7 [2520]. 88-8450-074-5 [2419]. 88-8450-126-1 [2645]. 88-8450-140-7 [2820]. 88-8450-143-1 [2628]. 88-8450-162-8 [2863]. 88-8450-198-9 [2877]. 88-8450-199-7 [3006]. 88-8450-210-1 [2880]. 88-8483-186-5 [2507]. 88-8492-238-0 [2623]. 88-8498-207-3 [2720]. 88-85050-93-X [2014]. 88-85295-51-7 [2965]. 88-85352-12-X [2009]. 88-85962-27-0 [1840]. 88-86143-08-7 [1806]. 88-86413-88-2 [2588]. 88-86685-45-9 [2314]. 88-86739-12-5 [2136]. 88-86830-02-5 [2773]. 88-86977-28-X [2849]. 88-86977-59-X [2849]. 88-86996-10-1 [2605]. 88-87088-88-8 [2113]. 88-87218-98-6 [2685]. 88-87262-14-4 [2490]. 88-87375-00-3 [1919]. 88-87814-30-9 [2797]. 88-88064-22-2 [2458]. 88-88247-09-2 [2497]. 88-88676-19-8 [2667]. 88-88690-02-6 [2577]. 88-88734-09-0 [2692]. 88-900162-1-3 [2078]. 88-900162-2-1 [2498]. 88-900162-3-X [2558]. 89 [2876]. 89-95457-52-0 [2876]. 89/The [3241].

9 [2227, 2746, 2887, 2899, 3543, 3002, 2669, 2737, 2538, 2981, 2891, 2780, 2964, 2543, 2797, 2314, 2908, 2791, 2542, 3124, 2232, 2755, 2949, 3057, 3029, 2614, 3225, 2747, 3110, 3265, 3079, 2950, 2533, 2900, 2413, 2918, 2790, 2743, 2684, 3287, 2195, 2877, 2716, 2886, 2841, 2952, 2576, 3525, 2426, 3034, 2807, 2998, 3547, 2593, 3032, 2544, 3250, 2501, 3092, 2691, 2956, 2562, 2968]. 9. [1950, 2021]. 9/2790 [1645]. 90 [2573, 2607, 2906, 1791, 57, 2771, 2493, 2555, 2554, 2195, 2732, 2833, 304, 997, 1229, 2840, 2937, 2556]. 90-04-09421-0 [1229]. 90-04-09423-7 [914]. 90-04-09464-4 [997]. 90-04-11804-7 [2573]. 90-04-11996-5 [2771]. 90-04-12233-8 [2937]. 90-04-14188-X [2937]. 90-04-15025-0 [2840]. 90-04-15160-5 [2933]. 90-04-15593-7 [2939]. 90-386-2077-2 [2607]. 90-420-1852-6 [2732]. 90-420-1868-2 [2906]. 90-5235-071-X [1791]. 90-5235-103-1 [1792]. 90-5699-103-5 [1944]. 90-5709-007-4 [2067]. 90-5867-293-X [2493]. 90-6203-676-7 [57]. 90-6292-114-0 [1792]. 90-6292-129-9 [2195]. 90-6292-139-6 [2555]. 90-6292-143-4 [2554]. 90-6292-148-5 [2556]. 90-74461-53-0 [2833]. 90.000 [644, 708]. ’900 [897, 1867]. 9001 [2524]. 906844265 [2662]. 91 [2200, 2496, 2500, 2675]. 91-27-07793-4 [2200]. 91-27-08105-2 [2200]. 91-506-1669-2 [2496]. 91-7203-256-1 [2119]. 91-7844-617-1 [2500]. 91-7844-635-X [2675]. 93/94 [1619]. 937 [3011]. 938 [1173]. 94/95 [1619]. 95 [191]. 960-7094-99-9 [1930]. 963 [2421]. 963-641-906-X [2421]. 972-0-34262-5 [2895]. 972-9127-40-9 [1743]. 976 [2922, 2987]. 976-1-59102-475-0 [2922]. 976-1-59102-477-4 [2987]. 978 [2887, 2986, 2916, 3041, 3016, 3299, 2879, 3000, 2899, 2889, 3113, 2942, 3543, 2897, 3116, 3064, 3002, 3107, 3120, 3286, 3283, 3094, 3005, 2894, 2933, 2936, 3063, 3040, 3059, 3071, 3121, 3109, 2852, 2851, 3081, 2770, 2891, 3068, 3255, 3254, 3280, 3282, 3541, 3072, 3080, 3526, 3545, 3025, 2959, 3009, 3023, 3279, 3015, 3204, 2979, 3082, 3084, 3129, 3123, 3122, 3268, 3021, 2974, 2969, 3030, 2958, 3124, 3544, 3076, 2920, 3263, 3022, 3066, 3115, 3209, 3256, 3013, 2978, 3205, 3127, 2949, 3298, 3043, 2838, 3090, 3111, 3087, 2895, 2874, 3061, 3060, 3003, 3012]. 978 [2941, 2888, 2943, 2827, 2948, 2934, 2988, 3057, 2885, 3070, 2955, 3029, 2946, 3108, 2893, 3289, 3288, 3131, 2905, 3225, 2999, 2973, 3253, 3110, 3265, 3067, 3079, 3033, 3024, 3074, 3017, 2970, 3091, 2900, 2977, 2912, 2945, 3014, 2938, 3069, 3018, 3031, 3546, 3542, 3529, 2918, 2884, 2944, 3426, 3267, 3251, 3036, 3277, 3208, 2660, 3287, 3257, 3093, 3088, 3038, 3224, 3078, 2886, 3114, 3008, 2844, 2975, 3118, 2952, 2875, 3010, 3284, 3083, 3125, 3525, 2663, 2866, 2983, 2985, 3039, 3085, 3130, 3226, 3034, 3241, 2984, 2939, 2951, 3096, 3062, 2998, 3044, 3019, 3206]. 978 [3058, 3126, 3528, 3547, 2806, 2919, 2967, 3032, 2982, 3086, 3011, 3203, 3075, 2911, 2890, 2921, 3261, 3095, 3250, 2935, 3092, 3278, 3527, 2956, 3037, 2910, 3239, 3259, 3073, 2857, 2960, 2968, 3089, 3264, 3285]. 978-0-19-280725-0 [3014]. 978-0-19-530573-9 [3124]. 978-0-19-920566-0 [2912]. 978-0-19-921149-4 [3108]. 978-0-19-921530-0 [3025]. 978-0-19-923619-0 [3074]. 978-0-19-924535-2 [3066]. 978-0-19-927820-6 [3131]. 978-0-19-954306-9 [2998]. 978-0-19-956664-8 [3113]. 978-0-19-965498-7 [3263]. 978-0-226-05853-5 [3261]. 978-0-226-20791-9 [3032]. 978-0-226-27650-2 [3033]. 978-0-226-30125-9 [2887]. 978-0-226-43033-1 [2985]. 978-0-226-48118-0 [3033]. 978-0-226-60856 [3068]. 978-0-226-61668-1 [2976]. 978-0-226-68086-6 [3542]. 978-0-226-71287-1 [3043]. 978-0-226-74945-7 [3118]. 978-0-226-76329-3 [3070]. 978-0-226-77056-7 [3126]. 978-0-253-35175-3 [3078]. 978-0-262-01675-9 [3265]. 978-0-262-06254-1 [2911]. 978-0-262-11306-9 [2900]. 978-0-262-13476-7 [2866]. 978-0-262-16234-0 [2806]. 978-0-262-19557-7 [2838]. 978-0-297-64547-4 [2881]. 978-0-300-09887-7 [3038]. 978-0-300-10995-5 [3012]. 978-0-300-11196-5 [2948]. 978-0-300-11796-7 [2944]. 978-0-300-13911-2 [3072]. 978-0-300-15093-3 [3204]. 978-0-306-81742-7 [3060]. 978-0-393-34340-3 [3281]. 978-0-415-34020-5 [3059]. 978-0-472-11602-7 [2950]. 978-0-472-11630-0 [3003]. 978-0-520-24261-6 [2885]. 978-0-520-24376-7 [3084]. 978-0-521-00903-4 [2874]. 978-0-521-47565-5 [2909]. 978-0-521-51484-2 [3086]. 978-0-521-60893-0 [2889]. 978-0-521-61266-1 [2936]. 978-0-521-69200-7 [3093]. 978-0-521-69530-5 [3063]. 978-0-521-80619-0 [3008]. 978-0-521-85181 [2879]. 978-0-521-85779-6 [2938]. 978-0-521-86286-8 [3067]. 978-0-521-86937-9 [3034]. 978-0-521-87087-0 [2827]. 978-0-521-87370-3 [3083]. 978-0-521-87412-0 [3076]. 978-0-521-88766-3 [3037]. 978-0-674-02299-7 [2844]. 978-0-674-02615-5 [2966]. 978-0-674-06663-2 [3278]. 978-0-691-12094-2 [2977]. 978-0-691-12310-3 [2851]. 978-0-691-12825-2 [2852]. 978-0-691-17487-7 [3529].

978-0-7294-0935-3 [3024]. 978-0-7546-3250-4 [2984]. 978-0-7546-4863-5 [3096]. 978-0-7546-5758-3 [2975]. 978-0-7546-5781-1 [2945]. 978-0-7546-5843-6 [3017]. 978-0-7546-5913-6 [3023]. 978-0-7546-5939-6 [3122]. 978-0-7546-6252-5 [3016]. 978-0-7546-6586-1 [3114]. 978-0-8018-8613-3 [2875]. 978-0-8018-8835-9 [3110]. 978-0-8047-4147-7 [2970]. 978-0-8047-4148-4 [2970]. 978-0-8047-5544-3 [2920]. 978-0-8122-3913-3 [2959]. 978-0-86078-996-3 [3005]. 978-0-87071-609-6 [3254]. 978-0-87169-259-7 [2890]. 978-0-87169-965-7 [2982]. 978-0-87220-937 [3011]. 978-0-87580-340-1 [2660]. 978-0-88135-396-9 [2891]. 978-0-88135-397-6 [3041]. 978-0-88135-399-0 [3036]. 978-0-9564591-1-4 [3208]. 978-1-107-00262-3 [3239]. 978-1-4020-2979-0 [2967]. 978-1-4094-2593-9 [3250]. 978-1-4214-0187-4 [3256]. 978-1-4214-0298-7 [3267]. 978-1-4214-0479-0 [3268]. 978-1-4214-2951-9 [3543]. 978-1-58603-832-8 [3081]. 978-1-59102-520-7 [2935]. 978-1-84306-193-9 [2899]. 978-1-84384-178-4 [3044]. 978-1-84628-790-9 [3079]. 978-1-85196-858-9 [2956]. 978-1-878220-76-9 [2971]. 978-1-890951-78-8 [2983]. 978-1-891220-02-0 [3111]. 978-1-905267-48-4 [3289]. 978-1-905375-20-2 [2886]. 978-1-905375-21-9 [2886]. 978-2-13-054657-3 [2663]. 978-2-251-34510-9 [2952]. 978-2-251-34512-3 [3013]. 978-2-251-34513-0 [3013]. 978-2-251-42034-9 [3004]. 978-2-251-42035-6 [3004]. 978-2-251-42047-9 [3287]. 978-2-296-05606-0 [2973]. 978-2-296-07957-1 [3087]. 978-2-35031-368-9 [3257]. 978-2-503-52365-1 [2770]. 978-2-503-54129-7 [3284]. 978-2-503-58107-1 [3526]. 978-2-600-05878-0 [3527]. 978-2-7116-2147-7 [3107]. 978-2-7117-4028-4 [2958]. 978-2-7283-0833-0 [3069]. 978-2-7298-7304-2 [3251]. 978-2-7355-0637-8 [2894]. 978-2-7453-1532-8 [3010]. 978-2-85399-797-3 [3280]. 978-2-87854-466-4 [3075]. 978-2-9600815-0-3 [3091]. 978-3-030-41413-9 [3525]. 978-3-05-004319-7 [2943]. 978-3-05-004496-5 [3129]. 978-3-05-005095-9 [3225]. 978-3-11-018542-3 [2933]. 978-3-406-54403-3 [2857]. 978-3-428-13849-4 [3226]. 978-3-515-08806-0 [3116]. 978-3-515-08862-6 [2978]. 978-3-515-09281-4 [3123]. 978-3-515-09344-6 [3073]. 978-3-515-09842-7 [3288]. 978-3-596-17627-4 [2979]. 978-3-7643-8390-9 [3002]. 978-3-7728-2375-6 [3064]. 978-3-7728-2540-8 [3283]. 978-3-7965-4073-8 [3530]. 978-3-7965-4155-1 [3530, 3530]. 978-3-8258-1635-3 [3253]. 978-3-8258-1958-3 [3130]. 978-3-8442-5258-3 [3282]. 978-3-87909-973-3 [3082]. 978-84-7283-905-2 [3019]. 978-88-02-07715-4 [2919]. 978-88-02-08139-7 [3085]. 978-88-09-05396-0 [3009]. 978-88-15-11889-9 [3029]. 978-88-15-11986-5 [2941]. 978-88-15-12574-3 [3094]. 978-88-15-13873-6 [3206]. 978-88-15-23878-8 [3226]. 978-88-17-01601-8 [2881]. 978-88-17-02334-4 [3022]. 978-88-222-5496-1 [3127]. 978-88-222-5621-8 [2960]. 978-88-222-5658-4 [2905]. 978-88-222-5660-7 [3015]. 978-88-222-5669-0 [2888]. 978-88-222-5686-7 [2955]. 978-88-222-5722-2 [2942]. 978-88-222-5749-9 [2949]. 978-88-222-5859-5 [3088]. 978-88-222-6003-1 [3286]. 978-88-222-6008-6 [3255]. 978-88-222-6122-9 [3258]. 978-88-222-6270-7 [3299]. 978-88-336-7039-3 [3528]. 978-88-339-1746-7 [2910]. 978-88-339-1802-0 [2893]. 978-88-339-1866-2 [3031]. 978-88-399-3411-2 [3546]. 978-88-4001-368-7 [3121]. 978-88-420-8377-1 [2946]. 978-88-430-4270-8 [2969]. 978-88-430-4535-8 [3095]. 978-88-430-9589-6 [3541]. 978-88-430-9620-6 [3545, 3544]. 978-88-430-9827-9 [3547]. 978-88-464-8818-3 [3021]. 978-88-467-1895-2 [2974]. 978-88-467-3603-1 [3285]. 978-88-498-1889-5 [3089]. 978-88-518-0081-9 [2918]. 978-88-568-0023-4 [3115]. 978-88-6036-163-9 [2968]. 978-88-6067-089-2 [3120]. 978-88-6227-097-7 [2951]. 978-88-6227-449-4 [3259]. 978-88-6227-567-5 [3298]. 978-88-6372-275-8 [3279]. 978-88-7038-467-3 [3040]. 978-88-7075-951-8 [3426]. 978-88-7075-952-5 [3426]. 978-88-7075-981-5 [3426]. 978-88-7075-982-2 [3426]. 978-88-7075-986-0 [3426]. 978-88-7088-543-9 [3057]. 978-88-7252-319-3 [3277]. 978-88-7424-216-0 [3039]. 978-88-7493-113-2 [2916]. 978-88-7642-321-5 [3203]. 978-88-7642-331-4 [3080]. 978-88-7794-490-0 [2882]. 978-88-7970-416-8 [3209]. 978-88-8082-681-1 [2921]. 978-88-8342-530-1 [3125]. 978-88-84 [3001]. 978-88-8422-706-5 [3090]. 978-88-8450-239-1 [2934]. 978-88-8450-247-6 [3062]. 978-88-8492-511-4 [2988]. 978-88-8492-564-0 [3030]. 978-88-88649-22-1 [3071]. 978-88-89854-68-6 [3257]. 978-90-04-15268-7 [2897]. 978-90-04-15593-0 [2939]. 978-90-04-15827-6 [3058]. 978-90-04-16186-3 [2999]. 978-90-420-3156-2 [3241]. 978-90-481-2252-3 [3061]. 978-90-481-8644-0 [3224]. 978-90-6554-424-7 [3109]. 978-90-6984-483-1 [3018]. 978-90-6984-506-7 [2884]. 978-90-801833-9-1 [3205]. 978-90-8550-974-1 [2986]. 978-91-7100-803-9 [3092]. 978-960-7916-83-9 [3205]. 978-972-0-45090-6 [2895]. 978-981-270-452-8 [3000]. 978-987-1673-67-4 [3264]. 978030012493 [2961]. 98-102-3920 [2268]. 981-02-2369-2 [2216]. 98338 [3001]. 98338-1 [3001]. 9th [3291].

= [3751, 3763, 420, 3885, 1838, 1133, 1174, 1225, 1391]. ’īsā [735].

Âge [3631, 305, 3127].

A-B [1444]. a-t-on [1041]. A. [195, 1132, 1587, 508, 569, 584, 1682, 29, 279, 1366, 1588, 2579, 1335, 1219, 2214, 2483, 763, 2013, 729, 1099, 2468, 4076, 291, 1671]. A.-J [2644]. A.-L. [1132, 1682]. A.A [1619]. a.C [2363, 2921]. A.D [2363, 735, 3564, 993]. A.D. [3906]. a.M [1058, 2146]. Aa. [1101]. Aa.Vv [1171, 1170]. Aa.Vv. [686, 1155, 1316, 855, 682, 709, 867, 956, 1094, 1115, 1176, 1178, 1275, 1287, 1431, 330, 200, 240, 1458]. ÅKerman [891]. aan [541]. Ångstrom [2304]. aantekenineng [3109]. aantekeningen [802]. år [3852, 3853, 3854]. Abaci [3408]. abacus [2130]. Abade [2895]. Abano [626]. Abb [1918, 1838]. Abbaci [3845, 3593]. Abbacus [3002]. abbandonate [3426]. Abbasal [1358]. Abbatia [1685]. Abboretti [171]. abbozzo [1034]. abbreviations [5]. Abbreviazioni [5]. Abbri [2166, 2237, 2329, 569, 1494, 2117, 2280, 1615, 831]. Aberrations [2292]. Abete [1287]. Abhandlungen [961, 2055, 23, 1191, 542, 1912, 376, 1069]. Abigail [2791]. Abir [440, 2067]. Abir-Am [440, 2067]. Abraham [895, 116, 3918, 3133, 1007]. Abramo [2429, 1906, 3918]. Abridged [3846, 3858, 1987]. Abril [963]. Abruzzese [2667]. Abruzzo [1181]. Absent [1885]. absoluta [250]. absolutism [1400]. Abstruse [2763]. Abt [1202]. Abteilung [217]. Abyssal [3352]. Academia [2517, 1744, 1645, 2808, 1600]. Academiae [302]. Academic [3750, 2462, 3080, 1401, 606, 624, 1570, 1600, 2583, 2939, 3058, 1857, 2850]. Academie [302, 749, 2267, 2495, 3598, 1681, 2894, 2611, 368, 2615, 2118, 2541, 2469, 1795, 3877, 2541]. Academiën [302]. Academies [1]. Academy [1081, 1476, 2338, 3156, 3471, 495, 1701, 1830, 2662, 3877, 1299, 763, 1008, 534, 2686, 3470, 2354, 3469, 3198, 3737]. Acapulco [2727]. Accademia [516, 721, 1091, 4045, 1564, 1689, 2933, 278, 311, 716, 887, 1077, 1603, 1909, 2941, 1741, 2713, 200, 1251, 1461, 1599, 1766, 2112, 2226, 1645, 838, 412, 2581, 2354, 2092, 2815]. Accademica [720, 717, 991, 1279, 1156, 1688, 70, 2536]. Accademiche [1909, 3741, 3750]. Accademie [1]. Access [990, 3282]. accompagna [3873]. accompagnée [2528]. Accompany [2724]. accompanying [3873]. Account [524, 36]. Accursio [1815]. Acedia [248]. Achievements [3469]. Achille [1806]. Achromatic [3422]. Ackerman [1357, 766]. Ackermann [1872]. Acme [1712]. Aconcio [3065]. Acot [1944]. Acqua [734]. Acquaintance [450]. Acquasparta [311]. Acque [2448, 976, 2491, 3255, 2461, 3915]. Acquisiti [1470]. Acquistate [1307]. Acquisto [668]. Acromegaly [3727]. Across [2994, 3723, 3189]. Acta [299, 514, 1055, 2418, 235, 1992, 950, 337, 681]. Actes [1090, 2878, 1440, 260, 32, 444, 836, 1442, 1984, 1327, 314, 868, 540, 2118, 1326, 1734, 2201, 4042, 4047, 4053, 4077, 3952, 4089, 243, 1224, 4043, 954, 713]. Action [2052]. actividad [2101]. activities [1999]. Activity [3080]. actors [1999]. Actualité [1609]. AD [2799, 1303, 1533, 2364, 312, 3433, 1138, 2727, 2531, 1865, 2115, 999, 412, 814, 522, 2833]. Ada [1899]. Adam [772, 218, 787, 1057]. Adams [445]. adapt [1829]. Adattar [1829]. Adda [2921]. Addendum [1690]. Additamenta [92]. Addition [2528]. Addressees [398, 2250]. Adelard [378]. Adeline [3395]. Adelmo [1174]. Adelphi [1248, 1506, 1856]. Adini [2614]. Adivar [2927]. Adler [543, 3111, 2187]. Administration [1588, 4049]. Adnan [2927]. Adolf [173]. Adolfo [832]. Adolph [107, 2544]. Adosa [921]. Adresse [1473]. Adrian [135, 1914]. Adriana [1488, 3040, 2151, 3820]. Adriano [2882, 1644, 2963]. Adriatica [1496]. Adrien [3024]. Adv [2602, 2787]. Advancement [167]. Adventure [3480]. Adventurer [1588]. Adversaries [1683]. Adversity [2899]. Advertising [2295]. Advice [3558]. Aeaomena [2418]. aedificandi [563]. Aegidii [720]. Aegidius [1359]. aerztliche [639]. Aesthetic [3542]. Aesthetics [3857]. Aether [443]. Aevi [307, 720, 717, 991, 1279, 1156, 1688, 534]. Aevii [70]. Affinitatum [2475]. affinité [834]. Affinity [2730]. Afins [430]. afraid [1726]. Africa [841, 3564]. African [1358, 3716, 840]. After [1263, 3334, 3752, 3430, 3868, 3920, 381, 1118, 2231, 3543, 3445]. Afterlives [3434]. Aftermath [2912]. AG [3002, 2205]. Again [3101]. Against [2127, 3769, 1829, 1053]. Agar [3216]. Agassi [1247]. Agassiz [1541]. Agate [3096]. Age [167, 202, 1110, 2424, 2463, 3569, 3328, 3482, 1564, 857, 1187, 1985, 2756, 760, 1057, 3267, 3691, 3210, 2731, 2845, 1986, 2604, 3796, 1273, 1492, 3258, 3947, 3953, 1654, 4018, 2887, 3173, 33, 3834, 2944, 3216, 2792, 3213, 3599, 3600]. Agenda [1231]. Ages [335, 2570, 2766, 2818, 2641, 2364, 30, 1103, 1277, 1446, 1445, 1985, 2997, 2874, 2932, 2705, 2009, 2515, 3106, 3238, 3056, 3747, 3175, 2707, 3006, 3937, 3325, 3595]. Aggiornamenti [1209]. Aggiornamento [162]. Aggiunti [1829]. agli [2616, 1522]. Agnès [1799]. Agnese [930, 71, 1871, 431, 518]. Agnesi [2783]. Agnoletti [2009]. ago [1440]. Agora [3120, 866, 2685]. Agosti [1645]. Agostino [1957, 2028, 3631]. Agraria [1350, 1666, 3710]. agrarie [3873]. Agricola [1317, 1608, 1661, 1147, 754, 1226, 1627, 1628]. agricole [1525, 2153]. agricoltura [170]. Agricultural [3487, 1350, 2804, 3873]. Agrimi [446, 1273, 2419]. Agrippa [1229]. Agro [1641]. Agro-Industries [1641]. agroforestal [2808]. Aguilar [352]. Agustí [2907, 3511, 3675]. Ahmed [2927]. ai [3747, 223, 3807, 2888, 2614, 785, 1092, 500, 875, 20, 778, 1016, 1748, 2009, 2150]. aid [676]. Aids [959, 566]. Aileen [3033]. Ailly [380]. aiónas [1930]. Air [168, 1776, 3392]. Air-Pump [168]. Aires [3264]. Aiton [218]. Aix [3280, 1224]. Aix-en-Provence [3280, 1224]. Ajaloomuuseumis [938]. Ajuntament [1849]. akīm [334]. Akademie [507, 508, 3129, 542, 2943, 3225, 1992, 3018, 2884, 1001, 1610, 2576, 411, 1854, 382, 1133, 1174, 1391]. Akron [2552]. Aktendes [1281]. Akute [1391]. Al [2710, 1974, 1829, 791, 573, 965, 1227, 740, 726, 353, 2199, 686, 1816, 619, 1226, 3138, 3783, 2313, 2298, 817, 1560, 2232, 2490, 1428, 1335, 45, 43, 2808, 2492, 789, 1181, 863, 876, 2257, 3426, 917, 355, 455, 779, 859, 1072, 1251, 1372, 1424, 1926, 2213, 2363, 1978, 3967, 13, 3973, 2083, 2399, 2208, 372, 3673, 1324, 4014, 609, 2266, 4038, 2088, 334, 3370, 3550, 353, 3292]. Al-akīm [334]. al-’amal [353]. Al-Furqän [2710]. Al-Furqn [2199]. al-girāsa [1226]. Al-Haytham [791, 2199, 3292]. al-jabr [609]. Al-Jazari [3370]. Al-Khazini [3550]. al-khmis [2199]. al-khwrisms [609]. Al-kitb [609]. al-muchabala [609]. al-mukhtasar [609]. al-qarn [2199]. al-Riyyt [2199]. al-Samh [353]. al-Sana [1227]. al-tallyah [2199]. al-thlith [2199]. ’Alī [735]. Alabama [174, 285]. Alain [444, 1667, 3013, 2819, 368, 807, 2183, 1414]. Alain-Philippe [3013]. alambicchi [2156]. Alamos [1249]. Alan [520, 2516, 58, 1915, 3716, 1198, 580, 1199, 1333, 2671, 385, 1915, 424, 1694, 1844, 1332]. Alatri [1343]. Alaymo [1648]. alba [331, 330, 3652]. Albanese [1841]. Albergo [1675]. Alberi [265, 1314]. Albert [2852, 2215, 754, 2115, 1386, 156, 2852, 2258]. Albertazzi [2224, 1015]. Albertazzi-Dazzani [1015]. Alberti [1044]. Albertini [955]. Alberto [1502, 1951, 2022, 648, 1794, 2070, 1904, 1399, 2645, 1856, 1658, 904, 1380, 2400, 2665, 1869, 35, 529, 1904, 1491, 1713, 2914]. Albin [1799]. Albrecht [1602, 2529, 127, 1764, 2205]. album [3247]. Alcheandrana [3058]. Alchemia [567]. ’Alchemical [3157, 1439, 3500, 2834, 3731, 3490, 3627, 610, 3964]. alchemici [3277, 3964]. Alchemico [814]. Alchemico-Metallurgiche [814]. Alchemie [1402, 1171, 276]. Alchemische [1402]. Alchemist [3529, 2830]. Alchemists [3157]. Alchemy [301, 1439, 2864, 2891, 2579, 3173, 3590, 3671, 2718, 3236, 4077, 3338, 753, 2840, 3963, 2591, 3311, 3361, 1002, 3808, 749, 3068, 2579, 753, 3313, 3440]. Alchimia [793, 1938, 646, 1601, 1868, 2420, 3808, 752, 682, 1528, 1868]. Alchimia. [165]. Alchimie [4077, 1138, 3173, 1228]. Alchimique [795]. alchimista [995]. Alchymistes [795]. Alcune [2091, 1826, 1558]. Alcuni [1470, 602, 776, 818]. Alden [1406]. Aldershot [2586, 3016, 1393, 3005, 1568, 2707, 3023, 3122, 1278, 1443, 1263, 1366, 2079, 2873, 2755, 2962, 2756, 1669, 2751, 2819, 2633, 2832, 1664, 3017, 2945, 2940, 2801, 835, 1320, 1563, 2098, 3114, 2975, 2749, 2803, 2984, 3096, 3250, 2802]. Aldershot/Burlington [2707]. Aldo [2276, 2933, 214, 1283, 2929]. Aldrovandi [1325, 2882, 3048, 2225, 3989]. Aldrovandian [1378]. Aldrovandiana [1378]. Aleksandar [2069]. Aleksandr [771]. Aleotti [2037, 2461, 4064]. Aleottie [1966]. Alessandra [1589, 1966, 2037, 3286, 2614, 2148, 2491, 715, 2789, 2413, 1482]. Alessandria [2574, 1483]. Alessandrini [2882]. Alessandro [1956, 1962, 2027, 2033, 2334, 2435, 2612, 1488, 1487, 2882, 717, 1869, 2156, 2623, 407, 697, 863, 917, 1343, 1614, 2213, 3687, 3652, 2285, 1483, 2607, 315, 665, 1322, 2836, 3788, 792, 2014, 2156, 3115, 3090, 2988, 2395, 874, 4076, 664, 2271, 3206, 924, 2127]. Alex [2560, 3079]. Alexander [60, 965, 1617, 2822, 3461, 3542, 2801, 534, 3401, 3073, 1462, 770, 1489, 1086, 1203, 4012]. Alexandre [305, 287]. Alexandria [2935, 3164, 2564]. Alexis [3377]. alfabeti [2580, 2968]. Alfabetico [6, 159, 465, 905]. Alfonsi [333]. Alfonso [2333, 915, 3415, 3591, 1782, 1973, 727, 648, 1277, 2058, 503, 1027]. Alfred [1853, 1945, 348, 2911, 2922, 3891, 3579, 3036, 535, 819, 348, 3429]. Alfredo [2429]. Algarotti [894, 312]. algebre [609]. Algorismi [3002]. Algorismus [2228, 2161]. Algra [1983]. alguse [938]. Ali [1358]. alia [2783]. Aliberti [3039]. Alic [77]. Alienationes [28]. alienista [2157, 3927]. Alighieri [711]. alii [1484, 1641, 1407, 2204]. alimentaire [3197]. alimentari [1525]. alimentazione [1314, 1904]. Alina [3258]. Alinari [1519]. Aline [3280]. Alinea [2006]. Alippi [1952, 2023]. Alison [3076, 916, 2975]. Alistair [3480]. Alix [2827]. alk [1987]. alkaline [1987]. Alla [523, 1629, 1241, 1467, 2049, 2399, 1953, 2024, 1947, 2018, 2012, 2354, 726, 18, 612, 2522, 3777, 2144, 486, 716, 718, 949, 1304, 1472, 2091, 314, 1099, 1673, 1677, 1866, 712, 3178, 704, 832, 937, 859, 1056, 2111, 558, 2618, 3136, 957, 1540, 1174, 3822]. Allacci [2085]. All’alba [2140]. all’ascolto [1115]. Alle [89, 802, 1367, 2452, 1574, 3115, 604, 1971, 1068, 1119, 1379, 2233, 2162, 449, 809, 2690, 952, 562, 1255, 2110, 1114, 2905, 1237]. alleanze [2507]. Allegorico [740, 739]. Allegory [3158]. Allemandi [407, 563, 556, 558, 83]. allemands [2539]. Allen [491, 3528, 289, 2549, 2625, 301, 837, 1680, 133, 2828]. allgemeine [557]. all’Hindenburg [2679]. Alliance [3512]. Allianz [1150]. all’Illuminismo [2691]. Alliney [2388]. Allioni [2530, 2531]. Allo [664, 854, 734]. All’origine [2941]. Alma [448]. Almagest [2813]. Almanacs [3612]. Almerico [1961, 2032]. Almqvist [1611]. Almut [1724]. Alnwick [2388]. aló [2497]. Alonso [1197]. Aloud [3477]. Alpenreisen [2853]. Alpes [2085]. Alpha [746, 747]. Alpha-Omega [746, 747]. Alphabetical [398, 399, 2250, 2251]. Alphabets [3100]. Alpi [2965]. Alpo [3440]. Alps [2512, 2085]. Alta [1855]. altera [4028]. altered [4028]. Altertums [1838, 4082]. Altieri [1029]. alto [3961, 3259]. altri [3693, 727, 3088, 3655]. Aluminum [3985, 3333]. Alvar [927]. Alvaro [1197]. Alzani [2650]. Am [440, 2067, 961, 1405, 1212, 2979, 173, 547, 415, 950, 183, 1610, 363]. ’amal [353]. Amaldi [2216, 2216]. Amateur [3701, 3705, 3699, 3702, 3355, 3356]. amazing [2844, 3000]. Amazonian [3579]. Amberes [1860]. Ambientali [809, 778, 844, 1591]. Ambiente [779, 660, 3818, 1599, 2322]. Ambiguous [3055, 3464]. Ambition [3480]. ámbito [2808]. Ambix [714, 2579]. Ambrosiana [947]. Amedeo [1949, 2020, 3029]. Ameghino [3618]. Amelio [561, 562]. America [2075, 1576, 58, 2094, 1935, 1544, 2312, 3800, 816, 2844, 3148, 4034, 1544]. American [242, 441, 885, 1169, 1617, 1819, 2495, 2964, 857, 2824, 174, 286, 829, 2137, 1184, 293, 1193, 2231, 2982, 2890, 1290, 1814, 68, 2971, 2964, 2872]. americanas [2097]. Americhe [861]. Amerigo [3293]. amérindien [3024]. Ames [1362]. Amherst [163, 2935]. Amici [2129, 658, 2354, 3138, 527, 1555, 462, 1895, 2047, 2129, 2255, 2914]. amicorum} [3247]. Amiel [866]. Amma [407, 1168, 558]. Ammazzabambini [346]. Amministrazione [4049, 1016, 2789, 2315]. among [3448]. Amoretti [407]. Amos [335]. Amour [2533]. Amouretti [1224]. Ampère [2859, 1638]. Amphion [1080]. ampliata [2107]. Ampliationes [28]. Ampollini [3496]. Amsterdam [2067, 2715, 3081, 63, 2906, 1359, 2662, 57, 1944, 1275, 3018, 2884, 2732, 3241, 280, 582, 541, 3654, 380]. Amsterdam/Philadelphia [1359]. Amy [2613]. Ana [1516, 2895, 3243]. Anachronisms [3630]. Analecta [302, 1396]. Analgesic [1810]. Analisi [199, 1158]. Analitici [1606]. Analitico [666, 1098]. Analogia [2349]. Analogy [2120, 3994, 3995]. analyse [2654]. Analyser [3104]. Analysis [3437, 1711, 4027, 955, 3064, 4027]. Analytica [30, 58, 890]. Analytical [2403]. analytique [920]. Anand [341]. Anastasios [625]. anastatica [1496, 2605, 788, 1408, 2138, 1197]. Anatomia [2326, 2277, 1951, 2022, 2691, 3364]. anatomica [1175, 3364]. Anatomical [3502, 817, 2699, 3489, 3147, 2556]. Anatomice [1962, 2033]. Anatomiche [817, 2434]. anatomico [3673]. Anatomie [3750, 1909, 2598]. anatomische [2598]. Anatomy [2746, 3750, 3222, 3661, 225, 3455]. Anatrini [3544]. Ancestors [3626]. Ancestry [1443]. ancienne [1224, 707]. anciens [675, 3058, 1224]. Ancient [2459, 2879, 3059, 511, 1358, 3500, 3649, 829, 845, 1640, 3466, 914, 3274, 1683, 707, 1092, 1429, 3465, 3061, 2991, 2076, 4082]. Ancients [4082, 675]. Ancona [2220]. andalusí [1226, 963]. Andalusí. [1227]. Andare [3965, 2206]. Anders [2745]. Andersch [158]. Anderson [414, 1393, 928]. Andorlini [2642]. Andr [1737]. André [516, 1418, 2952]. Andrea [3508, 3569, 1105, 3844, 184, 558, 3741, 3401, 2581, 1064, 1267, 2622, 2266, 3596, 2421, 3286, 402, 3009, 2941, 2614, 704, 2919, 1644]. Andreas [1531, 2978, 2598, 2576, 136, 3173, 2208]. Andreoli [1098]. Andrés [2494]. Andrew [2953, 2962, 14, 1225, 283, 3067, 2790, 1504, 2848]. Andrews [3017]. Anemoscope [2392]. anestesiologia [2414]. Anfänge [1905, 4035]. Angel [3683, 2138, 1226]. Angela [1283, 1029, 3654, 3299]. Angeles [441, 685, 1333, 2516, 1227, 846, 2885, 1987, 191, 770, 1319, 2784]. Angeles-London [770, 1319]. Angeles/London [2885]. Angeletti [1864]. Angeli [515, 897, 339, 588, 18, 456, 1368, 221, 1574, 1693, 2116, 2203, 181, 3021, 449, 1759, 1820, 3115, 1917, 1436, 2430, 810, 88, 922, 2855, 2716, 66, 1788, 2722, 2104, 957, 1988, 2902, 1291, 2714, 178, 546]. Angelica [3494]. Angelika [1918]. Angelini [1909, 3426]. Angelis [2941]. Angelo [1759, 1904, 1728, 2331, 981, 2176, 2010, 3031, 1523, 1895, 1728]. Angerstein [1019]. Angersteins [1019]. Angius [1961, 2032]. anglaise [413]. Anglo [3669]. Anglo-Italian [3669]. anhand [787]. Anhang [565]. Anim [1317]. Anima [734, 1965, 2036, 2963, 2308, 3880]. Animal [2870, 3715, 2985, 3489, 3565, 3955, 2706]. animale [2025, 1369, 3955, 1068, 1954, 2904]. Animali [861, 1956, 2027, 2012, 2334, 3777, 3804, 2323, 2285, 2903]. Animalium [1782]. Animals [3411, 3494, 3412, 816, 3777, 3804]. Anita [1161, 1251, 1877]. Ankunft [1212]. Anlaß [220]. anlässlich [276]. Anmerk [379]. Ann [2957, 3003, 2950, 1192, 194, 3074, 2489]. Anna [3362, 1695, 3737, 1453, 2645, 2650]. Annabel [2378]. Annales [1406, 2229]. Annali [2315, 727, 2009, 1904, 576]. Annalisa [1419]. Annals [2009]. Annalucia [1436]. Annamaria [178, 1909, 2367]. Annarita [3426]. Anne [3495, 3331, 3894, 2529, 1848, 2887, 1669, 1058]. annessi [2468]. Annette [959, 3427]. annexes [2898]. anni [2616, 3839, 2202, 1597, 355, 431, 1423, 1768, 1230, 3085, 3969, 1137, 2960]. Annibale [1790]. Annie [2894]. anniversaire [4089]. Anniversary [2724, 4089]. Anno [1623, 1771, 2039, 2122, 2241, 2340, 2437, 1881, 2010, 1025, 1372, 1599, 1631, 1137]. Annotated [1078, 1660, 1805, 696, 802, 955, 3205, 12, 1907]. Annotations [405]. annoté [2578]. annotée [2898]. annotés [796]. Announcement [205, 2344, 2390, 91, 262, 316, 392]. Annual [1898, 4078, 1749, 3944, 3945]. annum [3920, 800]. annunciata [3496]. Annuncio [2344, 2390]. Annuncio/Announcement [2344, 2390]. Anomalous [3514]. anonime [991, 1156, 70]. Anónimo [1227, 1226]. anoressia [2364]. anorexia [2364]. ANR [798, 994]. ANRW [571]. ans [1440, 410]. Anselmo [715]. Ansichten [2979]. Ansprachenund [296]. Anstalt [93]. antalgica [1810]. ante [3920, 800]. Antenore [621]. anthologie [796, 866]. Anthony [1169, 2095, 2541, 537, 3204, 2757, 762, 387]. Anthropocene [3602]. Anthropologie [1004, 248, 172, 2060, 3829]. anthropologue [416]. Anthropology [2671, 2765, 345, 1786, 3799, 3829, 3954, 3619]. Anti [3375, 819, 539]. Anti-Calvinists [539]. Anti-Vaccination [819]. Anti-Vivisection [3375]. antica [1605, 2517, 442, 1059, 2949, 745, 2135]. Anticamedicinatibetana [1595]. Antiche [1483, 309, 442, 2572, 1233, 1482]. Antichi [268, 1816, 75, 792, 1092, 2693, 1084, 1257, 3416, 607, 1719, 1258]. Antichit [1526]. Antichita [1276, 1604, 447, 797, 1760, 2934, 1683, 1816, 1223]. Anticipations [385]. antico [2933, 3541, 1101, 1062, 1258, 1343, 377, 3314]. Antike [2682, 3857, 706, 1058, 2643]. Antiken [4082, 1838]. Antinewtonian [2761]. Antiphlogistic [2350]. Antipolis [1224]. Antiqua [3298, 2574, 1225]. antiquaria [2665]. Antique [1314, 1217, 332, 2767]. antiques [868, 2820]. Antiquiora [1822, 1399]. Antiquité [3395, 2648, 3924, 638, 1060, 1135]. Antiquity [77, 2570, 2766, 2818, 2641, 1060, 1439, 447, 830, 2997, 2932, 3857, 1865, 2705, 1058, 2515, 3106, 3238, 3056, 3564, 3325, 3650, 3810, 1753, 3944, 2822, 2874, 3924, 3273, 3438, 3733]. Antisemitism [3377]. Antje [706]. Antognazza [3008]. Antoine [1846, 3739, 1440, 2512, 1588, 2423, 3966, 2528, 3881, 2541, 3960, 2603]. Antoine-Laurent [2603]. Antologia [552, 223]. Anton [1042, 412, 201, 1042]. Antonella [2798, 3069]. Antonelli [688]. Antonello [2335, 1065, 2286]. Antoni [802, 1367, 1584, 2099]. Antonietta [3546]. Antonini [958]. Antonino [2954, 2416]. Antonio [722, 1204, 1959, 2030, 1948, 2019, 2164, 2337, 336, 1145, 1648, 1679, 1682, 1693, 74, 2974, 38, 2101, 641, 644, 1166, 1167, 1690, 1745, 969, 2062, 2608, 1803, 1252, 2726, 2960, 3367, 794, 1200, 1616, 2235, 2329, 2462, 2431, 2824, 3027, 2398, 2632, 2955, 522, 2651, 3872, 2373, 2002, 2148, 922, 1506, 957, 1990, 2726, 1743, 1744]. Antony [249]. antropologi [2675, 1729]. Antropologia [943, 284, 300, 882, 780, 864, 2677, 2206, 3285, 931, 3799, 3954, 1767, 1170, 882]. Anziehungskraft [543]. Apa [280, 582]. Apa-Holland [280, 582]. Apennines [2363]. Apes [880, 2275, 3922, 2382]. api [1766]. Apiarium [2713]. Apocalyptic [2840]. Apollo [1654, 3790]. Apollonio [2590, 1973]. Apollonius [3163, 3453, 3684]. apologetic [1464]. Apologetica [2497]. Apologetiche [1464]. apologetico [3426]. Aporie [1976]. Apostolica [281, 1372]. Apotekarsocieteten [1805]. Apothecary [2994]. Apotheker [623, 1715, 1714]. Apother [1905]. app [720, 966, 964, 857, 1105, 996, 1279, 1187, 1446, 1447]. apparati [2523]. Apparaturen [2915]. Apparatus [3680, 3505, 2423, 3966]. Apparecchi [828, 75]. apparecchietto [2974]. appareils [2757]. apparence [625]. Apparenza [384]. appartenant [2757]. Appeal [3558]. Appealing [2303].

Appearance [384, 1294]. Appellationes [28]. Appendice [1553, 984, 83, 983, 2497, 1267]. Appendices [1911, 334, 28]. appendici [28, 31]. appendix [1267]. appendixes [422]. Appennino [2363]. Apperception [890]. Applicata [71, 199, 291, 2005, 2217]. application [2381]. Applications [3764]. Applicazioni [1467, 700]. Applying [2636]. Approaches [1536]. Approche [2229]. Appropriazione [3255, 3915]. Apps [295]. Appunti [1890]. Après [2757]. April [4040, 1749, 276, 753]. aprile [4040, 1741]. Apuano [711]. Aq [1181]. Aquila [2620, 2667, 4059]. araba [1434]. Arabe [333, 352, 3058, 2648]. Arabes [1227, 1226]. Arabic [2055, 1102, 1225, 3896, 3166]. Arabist [378]. Aragno [2308]. Aragon [1392]. araldica [1314]. Araque [3415]. Arato [894]. Arbeiten [32]. Arbeitskreis [1191]. Arbib [934]. arbitrio [2159, 2488]. Arbor [257, 2957, 3003, 2950, 1192]. Arbore [1314]. arbre [3180]. Arcadian [2998]. Arcangelo [1369, 2319]. Arcetri [3323, 3138]. archaeological [3245]. Archaeology [3622, 3624, 3140]. Archaeopress [2204]. Archäologe [173]. Archè [3277]. Archeoastronomia [2113]. Archeologia [736, 575, 1834, 1831, 970]. archeologica [575]. Archeologico [1530]. archéologiques [3245]. Archimede [2881, 2270]. Archimedean [815]. Archimedes [2881, 2731, 3003, 49]. archimetria [4028]. Archinard [1421]. Architectura [393, 1645]. Architectural [1839]. Architecture [2086, 3204, 2687, 3839, 3258, 3953]. architetto [3015, 2577, 2461]. Architettura [561, 562, 1314, 2006]. architetture [3839]. Archiv [1941, 1285]. Archival [3624, 3245]. Archive [4, 2183, 1350, 3335, 2184, 2260, 2187]. Archives [3484, 1401, 1944, 3092, 2188, 2180, 175, 3428, 2067]. archivi [2213]. Archivio [1894, 4, 737, 269, 821, 1383, 2313, 2314, 2376, 1471, 2090, 1124, 1828, 323, 38, 463, 2447, 600, 1302, 1782, 1973, 1979, 663, 2129, 2353, 2397, 953, 1168, 397, 2249, 95, 1633, 528, 1335, 789, 1755, 697, 876, 291, 1350, 3335, 844]. Archivistici [844]. archivistiques [3245]. Arduino [1637, 739]. area [3839, 1260, 1429]. areas [2363]. Arecco [2536]. aree [2363]. Aretino [2112, 635]. Arezzo [2773, 635, 1259, 1874, 2112, 2117, 1615]. Arezzoe [517]. Argenta [2461]. Argentina [3618]. argentinos [3618]. argomentazioni [2141]. Argos [2577]. Argument [805, 3139]. Arguments [387]. aria [1533]. Ariane [3691, 1992, 3095]. Arianna [3410]. Ariosto [1241]. Aristosseno [2246]. Aristotele [1606, 2933, 443, 718, 887, 1189, 2517, 1829, 2518]. Aristoteles [1941]. Aristotelian [1683, 1102, 3366, 3826, 3272]. Aristotelianism [719, 263]. Aristotelica [3064, 1687, 4027]. aristotelico [3279, 3826]. Aristotelis [1687, 2520]. aristotelischen [1941]. Aristotelismus [443, 571]. Aristotle [1688, 2799, 1829, 535, 2812, 3239, 4027]. Arita [499]. Arithmetic [2991]. arithmétique [2596]. Arizona [592]. Ark [418]. arm [1163]. Armando [2073, 2557, 2429, 252]. Armin [751]. Arminianism [539]. Armocida [1431, 1809]. Armonici [2246]. Armonico [2123]. armonie [1658]. Arnaldo [2642, 1000, 644, 712, 810]. Arne [2194]. Arno [695]. Arnold [148, 2940, 157]. Aro [3692]. aromatiche [1268]. Arosio [1355]. Arp [812]. Arpaeolia [355]. Arpago [237]. Arqueològica [353]. Arqueologico [963]. Arquimedes [3167]. Arrangement [3133]. arredi [429]. Arrhenius [1657]. Arrigoni [313]. arrival [1212, 3895]. Arrizabalaga [1720]. Arrow [349]. Ars [867, 1094, 563, 384]. Arsenal [3739]. Arsenale [315]. Art [242, 1818, 2120, 2227, 3509, 699, 1314, 3230, 2908, 3082, 3182, 2356, 3495, 1143, 1566, 2084, 3411, 3494, 2095, 2704, 3357, 3216, 2684, 2761, 3038, 3338, 2357, 3153, 2731, 2985, 2640, 384, 2994, 2815, 840, 1986, 1331, 2338, 1608, 886, 3879, 332, 436, 1865, 2423, 611, 3966, 2786, 3994, 3995, 4090]. Art-Technology [3338]. Arte [2326, 2524, 2356, 2277, 2357, 1394, 2691, 1868, 2810, 1331, 2329, 332, 1434, 1093, 1354, 611, 1565, 2100, 1197, 2491, 3327]. Artemis [193]. Arterial [3148]. Arthur [3369, 1588, 3412, 761, 1203, 3732, 2864, 951, 2961, 2204, 1542, 1854]. Arti [1488, 2882, 1603, 575, 1195, 1343, 1671, 2112, 604, 12, 2650, 2142, 3255, 873, 958, 3915]. Artibus [2145, 2267, 2596, 2369]. articles [681]. Artico [336]. Artiere [1415, 1787]. Artifice [3495]. artifici [1488]. Artificial [820]. artificiale [1459]. Artigas [2839, 2595]. artigiani [2888]. Artillery [2054]. Artioli [726]. Artis [1531]. Artisan [3431, 3909, 3456, 3254]. Artisan/practitioners [3909]. Artisanal [3486, 3459]. Artisans [2098, 3766]. Artista [2769]. Artistarum [1444, 1984]. artisti [750]. Artistic [3382]. Artistiche [1419]. artistici [2002, 2148]. Artium [601]. Arts [722, 2160, 2933, 3195, 3531, 2662, 3189, 3215, 1224, 3459, 3363, 620, 1444, 1984, 3952, 3257, 2767]. Arturo [3412]. Arzneimittel [1724]. Arzneimitteltherapie [1715, 3938]. Arzt [172, 62]. Arzte [690, 557, 1797, 364, 3977]. Arztes [565]. Ärztliche [676]. Asa [2119]. asbl [3091]. Ascarelli [1817]. Ascensioni [1050]. Ascent [3102]. ascesa [3286, 3824]. Asch [126]. Asclepiades [946]. Ash [1575]. Ashbourne [2899]. Ashgate [2586, 3016, 3005, 2707, 2981, 3023, 3122, 2873, 2755, 2962, 2756, 2751, 2819, 2633, 2832, 3017, 2945, 2940, 2801, 1320, 2098, 3114, 2975, 2749, 2803, 2984, 3250, 2802]. Ashgate/Variorum [3023]. Ashmolean [2204, 3913]. Asimov [3151]. asnage [541]. Aspectos [35]. Aspects [1264, 3080, 35, 1812]. Asper [3649]. Aspetti [1789, 859]. Aspm [1681]. Assael [1674, 1867]. Assembly [1249]. assente [1852, 2954]. assenti [2007]. asservatorum [1687]. Assessment [488]. Assessorato [314, 778, 859, 1424]. Assiomatico [2246]. Assiomatico-Deduttivo [2246]. Assisi [1101]. assistance [520]. Assistenza [810, 1335, 879]. Associate [1264]. Associated [491, 68]. Associati [1811, 2743, 635, 1455, 446, 2742]. Association [1944, 4078, 1314, 847, 2204, 2998, 3044]. Associazione [1456]. assoluto [2632]. Asthetik [3123, 2682, 3857]. Astrofili [1013, 1160]. Astrofilo [1013, 1160]. Astrofisica [2290, 240]. Astrolabe [2055, 2441, 3449, 3730, 3707, 1123, 424, 1563, 735, 3669, 1630, 3609, 3896, 3276]. astrolaben [1342]. Astrolabes [3111, 3898, 1342, 3250, 424]. Astrolabi [1160, 353]. Astrolabio [1613, 1581, 1580, 1769]. Astrologer [2442, 2830]. Astrologia [1565, 3259, 2942, 752, 1061, 3961]. astrologica [3920, 800]. astrological [3920]. astrologicis [3165]. astrologiques [3058]. astrologue [1447]. Astrology [649, 1357, 3278, 3756, 275, 3612]. astronauti [3416]. Astronome [2958]. Astronomer [3705, 3740]. Astronomers [1671]. Astronomes [3013, 3890, 3980]. Astronomi [71, 1748, 1671]. Astronomia [1747, 239, 1667, 3496, 430, 633, 1590, 1591, 1514, 2113, 2211, 2213, 2212, 2548, 3939, 790, 1262, 2593, 3782, 3783, 1059, 792, 333, 1919, 2951, 1836, 1862, 2152, 240, 2005, 2070, 999, 3158]. Astronomica [3920, 1748, 1749, 1919, 800]. Astronomical [2442, 2510, 2089, 3435, 3847, 524, 1664, 3920, 631, 1082, 1749, 1308, 3292, 2872, 2404, 2937, 3478, 3451, 406]. astronomiche [2142]. Astronomici [2257]. Astronomico [486, 1220, 1555, 291, 430, 937, 1082, 2110, 3847, 631, 1863]. Astronomie [1171, 542, 2048, 2566, 2396, 1501, 540]. astronomiques [3307, 3071]. astronomisches [751]. Astronomo [688, 1823, 590, 2070]. Astronomy [2516, 2055, 3450, 3005, 511, 3111, 2999, 52, 2970, 1640, 2513, 1749, 2213, 329, 489, 4045, 3783, 1670, 3896, 2396, 4087, 4003, 4004, 2879, 3782, 674]. Astrophysics [2872, 4003, 4004, 1670]. asurlāb [353]. Athanasius [2227, 1610, 3433, 3174, 1770, 1196, 4090]. Atheism [1087, 2873]. Athens [3205, 1915]. Athens-Leuven [3205]. Athesia [948]. Athlone [342, 1717]. Atlanta [1275]. Atlante [233]. Atlantic [1365, 2964, 3342, 3444]. Atlantico [3172]. Atlanticus [3629]. Atlantis [1537]. Atlas [420, 3305]. Atlee [2653]. atolli [2849]. Atom [3263, 2546, 3901]. atomes [796]. Atomi [596]. Atomic [1702, 1446, 3263, 3901, 428]. atomica [1949, 2020]. atomico [428]. atomique [3127]. Atomism [2462, 3162, 1708, 2771]. Atomismo [2348]. atomist [3863]. atomista [3863, 2372]. Atoms [2443, 2718]. Atomtheorien [1446]. Atri [991]. Atteind [1306]. Atti [197, 514, 1091, 2389, 1488, 2522, 2376, 311, 915, 572, 2949, 2492, 709, 1101, 2880, 1813, 2005, 2113, 12, 2581, 1388, 1137, 1270, 2692, 4040, 4059, 4067, 1251, 3967, 721, 1243, 561, 2217, 2809, 838, 831]. Attilio [2754]. Attitudes [651]. Attivit [521, 1880, 1545, 1979]. Attivita [391, 90, 261, 458, 657, 1116, 1292, 1035, 1205, 1374, 1463, 2261, 47, 731, 813, 973, 1372, 818, 1621]. Attractions [1725, 3841]. attraverso [1180, 354, 812, 942]. attrezzi [2009]. attribué [1232]. Attributed [3436, 610]. Attributions [2346]. attuale [566]. attualità [189]. aube [1985, 331]. Aubin [3483]. Auctarium [2531]. Auctores [534]. Audiences [2928, 3033]. Aufkärung [565]. Aufklärer [62, 895]. Aufklärung [220, 542, 2532, 2785]. Aufklärungszeit [62]. Aufnahme [508]. aufrichtige [1095]. Aufsätze [253]. Aufstieg [364, 571, 798]. Aufstiegund [994]. Auftrag [1735]. Auftragdes [2196]. Auge [1918, 300]. augmenté [2767]. August [109, 1191, 2059, 64, 276, 363, 1150, 108]. Augusta [1352, 1191]. Auguste [2444, 2539]. Augustin [2517]. Augustiniennes [574]. Augusto [2612, 2693, 39, 2151]. Aurelianensis [1359]. Aurelianus [1391]. Aurifodina [2138]. Aurivillius [124]. Ausbau [1141].

Auseinandersetzung [343]. Ausejo [1515]. Ausgabe [1731, 1328]. Auslegung [1202, 4033]. auspices [1795, 368, 2118]. Ausstellung [173, 363, 2265]. Austellungsbegleiter [234]. Austin [2824]. Australia [670]. australopithecus [4008]. Austria [1019]. auteurs [2541]. Authenticity [3293]. Authoritative [3321]. Authorities [1683]. Authority [3068, 17, 3108, 3191, 2559]. Authorship [2412, 4048]. Autobiografia [2919, 3757]. Autobiographies [885]. Autobiography [1592, 3757]. Automata [3654]. Automatic [355]. automatiche [355]. Autonoma [1794]. Autora [1584]. Autoren [994]. Autori [1066, 72]. Autour [1984, 1476, 2229]. Aux [2566, 2048, 1170, 3527, 2578, 444]. Auzuox [2556]. av [1019, 1441, 2500, 884]. available [2683]. Avancement [1475]. Avant [3368, 368, 3571]. Avant-Propos [368, 3368]. Avanzini [1867]. Avanzo [1794]. avènement [3895]. aventures [2052]. averne [1899]. Avicenna [381, 2308, 3880]. Avogadro [1949, 2020, 3029, 1830]. avril [4042, 836, 1327, 1421]. Avvento [2871]. avventura [1437]. avventurosa [2910]. Avvertenze [1417]. avvisatori [779]. Awards [3463]. away [3965]. awqāt [1226, 1227]. Axolotl [3554]. Ayelet [3668]. Azafea [352]. Azarquiel [352]. Azevedo [893]. Azienda [1237]. Azzari [1324]. Azzolini [3278].

B [303, 454, 1370, 2516, 2389, 2354, 568, 1911, 3770, 235, 847, 29, 1359, 1444, 1670, 3441, 2589, 2873, 3462, 2907, 2719, 3345, 169, 3412, 1944, 789, 1092, 1431, 1809, 1935, 2152, 462, 36, 26, 862, 1161, 1532, 999, 545, 1645, 3118, 2552, 347, 13, 735, 808, 1329, 380, 2935, 2294, 3445, 2556, 609, 3526, 2371, 407, 3525, 4004, 1717]. B-Natural [2719]. B. [1446, 2091, 1867, 329]. B.C [993, 3906]. b.n [1190, 1335]. B.V [802]. b/n [2371]. b/w [3526, 3525, 1717]. Baasner [895, 542]. Baba [2197]. Babbage [2919]. Babel [3334, 3868]. Babini [1574, 2543, 2104, 178]. Babu [3712]. Babylonian [3000, 1309]. Bacci [1251]. Bach [2734, 2465]. Bache [1617]. Bachelard [1578, 1995, 2429, 3744, 2429, 4021]. Bachi [1890]. Back [3584, 3640, 3696, 2940, 2366, 3101, 772]. background [3842, 204, 1008]. Bacon [167, 679, 918, 1194, 1362, 1649, 1789, 1907, 3016, 2394, 4020]. Bacone [2597]. Baconian [1457]. Baconianism [659]. Bacterial [3601]. Bacteriology [3692, 3185]. Bad [3064, 3283, 173, 343, 787, 344, 1731, 2734, 1202, 2465]. Badaloni [2202]. Bagliani [3631]. Baglivi [3329, 1218]. Baguena [43]. Baillou [1554]. Baima [1426]. Bains [2757, 2757]. Baird [2381, 3797]. Baker [2552]. Baklarish [2998]. bakteriologischen [183]. Balance [3550]. balances [1163]. Baldassarri [3673]. Baldin [3525]. Baldini [1115, 1033, 412, 1040]. Baldinucci [1521]. Baldo [1195, 1532, 3364]. Baliani [1960, 2031]. Balkema [841, 840]. Ball [3229, 2133]. Balla [2212]. Ballesio [2314]. Ballistics [3282, 4015]. Balsamo [877, 882]. Baltimore [1231, 2016, 3543, 3268, 3256, 2794, 1409, 3110, 2825, 3267, 2799, 2584, 3210, 2875, 1458, 2786, 1329, 1837, 1839]. Balugani [840]. Ban [645]. Banca [2368]. Band [303, 567, 961, 895, 1402, 1531, 1602, 2647, 1608, 307, 571, 1001, 1058, 2785, 1731, 1838, 2725, 1797, 364, 411, 1153, 1086, 1328, 751, 1069, 565, 1905]. Bänden [1731]. bandiera [1737]. Bandinelli [3299]. Banfield [843, 1085]. bang [1114, 1114]. Bangham [3573]. Banks [454, 1370, 2156, 2956, 2527]. Baptiste [620, 2203, 3864, 2611]. Bar [2204]. Baraldi [2745]. Barba [1197]. Barbara [2548, 2529, 749, 2433, 1231, 2120, 720, 1143, 2820]. Barbarisi [339]. Barbaro [393]. Barbieri [641, 1678]. Barcan [2746]. Barcellona [686, 494]. Barcellona-Madrid [686, 494]. Barcelona [353, 333, 882, 618, 2099, 3452]. Barcelona-Madrid [618]. Bardi [1295]. Bardinet [707, 1432]. Bargello [777]. Barger [347]. Barghini [558]. Bari [1996, 2320, 2431, 718, 971, 2665, 2224, 2727, 2383, 2536, 3090, 682, 2946, 360, 566, 1030, 1674, 1673, 1675, 1752, 1672, 1753, 1756, 1677, 1676, 1867, 1865, 1866, 330, 1718, 240, 704, 1876, 417, 255, 1075, 2618, 611, 2921, 1836, 2810, 1223, 1651, 1719]. Barigazzi [1174]. Barker [1518, 1225]. Barletti [2476]. Barnabas [609]. Barnes [3148, 3096]. Barnett [3543]. barocca [2142, 2523, 682, 3178, 3614, 2387]. Barocco [756, 1488]. Barocelli [1061]. barocken [1610]. Barometer [843]. Barometers [1085, 3761]. Barometro [1216, 700]. Baron [808, 4032]. Baroncelli [1960, 2031]. Baroque [1365, 2227, 4090, 900, 2086]. Baros [843, 1085]. Baroukh [1674]. Barr [1915]. Barreca [3426, 3717]. Barrera [2824]. Barrera-Osorio [2824]. barroco [3327]. Barrotta [1368]. Barrow [803, 3843, 319]. Barry [3096]. Barsanti [1954, 2025, 453, 655, 522]. Barsantia [522]. Barthlemy [2596]. Bartocci [2883]. Bartoli [1958, 2029]. Bartolomeo [2012, 3777]. Bas [3829, 2060]. Bascule [560]. Base [2051]. based [4054, 4055, 787, 388]. Basel [2647, 3002, 1149, 172, 3530, 2529, 2785, 2725, 1141, 363, 1247, 2205]. Basel/Berlin [3530]. Basel/Boston/Berlin [2647]. Basel/Stuttgart [172]. Bases [990]. Basic [1169]. Basicbooks [1505]. Basil [196, 450, 451]. Basile [2142, 312]. Basilio [1690]. Basilisken [256, 2544]. Basilisken-Presse [256, 2544]. Basler [172, 3530]. Basnage [541]. Bass [3495]. bassa [3710]. Bassetti [2590]. Bassi [3689, 1726, 14, 1906, 3918]. Bassignana [407, 558]. Basso [1745, 1923, 1491, 2363]. bassomedievale [575]. Bassum [1735]. Bate [1161]. Bath [378]. Bathing [3564]. Baths [3564]. Bats [817]. battaglie [2742]. Battaglini [1303, 1302]. Battimelli [2216]. Battista [2370, 1963, 2034, 215, 3138, 3015, 2007, 3457, 2461, 2712, 956, 876, 1895, 2047, 2129, 2255, 929, 3443, 13, 1044, 2208, 12, 2914, 1297]. Battisti [1960, 2031]. Battistini [2941, 2213, 1064]. Bauer [1758, 948]. Baumeister [961]. Bäumer [1058]. Bautz [1151]. Bayern [1501]. Bayle [1908]. bayna [2199]. BC [2646, 2799]. Bce [989]. Bd [23, 251, 188, 443, 798, 994, 1191, 557, 1724, 217, 183, 622, 276, 362, 1202, 1912, 1989, 690]. Be [3467, 2870, 457, 2382, 3451, 2646, 3922]. Be-Ma [457]. Beach [3041, 2891, 2972, 2553, 2917, 2907, 2839, 2828, 3036, 2745]. Beagle [1408]. Beagon [2706]. Bearbeitet [217, 2576, 259]. Bearbeitung [256]. Beast [3661]. Beasts [833]. Beate [1726]. Beauchamp [2381, 1913]. Beaujouan [886]. Beaune [1326]. beautés [3075]. Beautiful [3348, 3415]. Beauval [541]. Became [3035, 3281, 2676, 3870, 3903]. Beccafumi [814]. beccari [4038, 2266, 4038]. Becchi [2384]. Becherucci [3820, 3040]. Beck [543, 706, 2643]. Becocci [277]. Becomes [3055]. Becquemont [1878]. Beddoes [16]. Bede [1278]. Bedeutung [1069]. Bedini [1320, 2098, 1137, 281, 3688, 2930]. Bedside [80]. Beeckman [2372, 3863]. Beer [418]. Bees [2686, 1766]. beffa [2535]. Before [1448, 3104, 3843, 803, 3334, 3463, 3868, 1392, 3470]. Beg [2760]. Begierde [1918]. Beginn [1405]. Beginning [3118, 447, 993]. Beginnings [2586, 2371, 1879, 3855, 3906, 1999, 4035]. begründet [1838]. Behar [2528]. Behavior [1852]. Behaviorism [286, 3382]. Behind [2545]. Beigun [1231]. Beihefte [3116, 1941, 1285]. Beijing [2095]. Being [2418, 2764, 3147, 3840, 4002, 3446]. Beings [2350]. Beispiel [1610, 364]. Beitrag [62, 248, 966]. Beitrage [565, 253, 3530, 1918, 492, 296]. Belcastro [3693]. Belforte [2380]. Belga [2833]. België [302]. Belgioioso [1142]. Belgique [3091, 2428, 2469]. Belgium [3526]. Belgrade [2069]. Belhoste [40, 3330]. Belief [2839]. Beliefs [1725, 2803, 3841]. believer [4089]. Belin [222, 620, 40, 368]. Belknap [347]. Bell [2364]. Bella [1018, 1165, 2504]. Bellagarda [432]. Bellardone [42]. Bellaterra [2099]. ’belle [456]. Belles [3013, 754, 2892, 3287, 216, 2952, 3004, 2118]. belles-lettres [2118]. Bellezza [3426]. Belli [1243, 911]. Bellini [1571]. Bellodi [1243]. Belloli [184]. Bellone [934, 1901, 1265]. Bellosta [2573]. Bellucci [1926, 2414]. Belluzzi [3015]. Belmar [2101]. Belonging [1630]. Belot [3703]. Beltrami [2215, 3768]. Beltrano [1525, 1666]. Bemerkungen. [2374]. ben [2461, 2604]. Bene [1637, 1233]. Benedetta [3402, 2622]. Benedetti [1962, 2033, 12]. Benedetto [2525, 2580]. Benedict [229, 945]. Benedictis [223]. Benedino [3027, 1789, 2372]. Beneficenza [810, 2006]. Benett [1072]. Benevento [2849]. Beni [809, 20, 778, 844, 1599, 1591, 1794, 3693, 2002, 2148, 1645]. Beniamino [1679]. Bénichou [416]. Benin [2197]. Benison [347]. Benito [1062]. Benjamin [283, 3070]. Benkert [3530]. Bennett [1393, 1511, 1662, 489, 3418, 3719, 3732]. Benno [251]. Benoit [1326]. Bensaude [796, 1494, 1848, 2273]. Bensaude-Vincent [796, 1494, 1848, 2273]. Bensi [2490]. Bentham [925]. Bentivoglio [1787]. Benz [1391]. Benzenhöfer [567]. Beonio [718]. ber [508, 1079]. Bérard [893]. Berariu [2267]. Berc [3010]. Berckelmann [106]. Beretta [1055, 3739, 1967, 2038, 1947, 2018, 2436, 1147, 1712, 814]. Berg [2746]. Bergamo [715, 719, 991, 2605, 869, 1456, 2415, 2415]. Bergdolt [950]. Bergenheim [2119]. Berger [149]. Bergman [725]. Bergshistoria [1019]. Bergson [2318, 1578, 2428, 1802, 3805, 4088]. Bergwelten [1661]. Bericht [1328]. Berichte [961]. Berkeley [441, 685, 1333, 2516, 846, 1495, 182, 3084, 2671, 1500, 328, 2885, 1987, 191, 770, 1319, 255, 451, 757, 969, 2382, 2784, 387, 1493, 1845, 388]. Berkeley-Los [1333, 770, 1319]. Berkeley/Los [2885]. Berlin [507, 2647, 3002, 508, 2822, 2933, 2852, 3081, 3282, 2653, 3129, 443, 571, 798, 994, 498, 1149, 2646, 2943, 3530, 3225, 2196, 201, 1001, 1074, 1610, 2576, 3130, 3226, 1247, 411, 348, 2466, 2422, 1133, 1174, 1391, 2146, 695]. Berlin-Diepholz [2466]. Berlin-New [2822, 2653, 443, 571, 798, 994, 348]. Berlin/New [2933, 498]. Berliner [508]. Berlinesi [3497]. Berling [2854]. Berlinguer [883]. Berlioz [1571]. Bermannus [754]. Bern [183, 1764, 2804]. Bernadette [796, 1494, 1848, 2273]. Bernal [2848]. Bernard [1495, 1187, 3637, 3033, 1458, 2316, 281, 1138, 1668, 3904]. Bernardi [61, 1953, 2024, 1904, 1068]. Bernardini [1335]. Bernardino [721, 1963, 2034, 838, 2488]. Bernardo [949, 1045]. Bernardoni [3009, 3596]. Bernasconi [2965]. Bernd [1661, 2962]. Berner [3617]. Bernhard [2161, 107, 1069]. Bernoulli [1141, 2786, 2725, 3235]. Bernoulli-Edition [1141]. Berquist [3316]. Berret [1799]. bersetzung [1531, 256]. Bertani [81]. Berthollet [2898]. Berthon [2528]. Berti [1820, 3394]. Bertola [2458]. Bertoletti [687]. Bertoloni [2825]. Bertomeu [2907, 2101, 3091]. Bertomeu-Sanchez [3091, 2907]. Bertotti [1579]. Bertozzi [2942]. Bertrams [3332]. Bertucci [2893, 2498, 3459]. Berücksichtigung [1724]. Berzelius [1005]. Berzolari [1869]. Besler [1518]. Besomi [2167, 2238]. besonderer [1724]. Bespoke [3534]. Bessarion [1123]. Best [2800]. Besta [1355]. Bestehen [296]. Bestehens [220]. Bestiari [4081, 1642]. bestiary [4081, 833]. Betti [1303, 1302]. Bèttica [1438]. Bettina [1531, 623, 2750]. Betty [1002]. Between [3379, 3137, 2350, 1699, 3053, 817, 1481, 3857, 1246, 1398, 2258, 3623, 3219, 544, 2228, 3760, 3701, 4044, 2161, 1111, 3826, 4072, 3893, 3902, 624, 3904, 3258, 1936, 3925, 778, 3377, 3953, 3956, 2231, 2998, 2804, 808, 1912, 1331, 1989, 768, 4032, 1, 3905]. Beverley [1401]. Bevilacqua [1499, 2156]. Beyond [2671, 3384, 2084, 3617, 3658]. Bezza [3315]. bi [353, 1640]. bi-l-asurlāb [353]. Biagi [1029]. Biagio [1893]. Biagioli [1980, 1400]. Biancastella [2954]. Bianchi [3107, 310, 447, 718, 719, 1908, 1639, 1388, 2398]. Bianchini [2511, 751, 3516]. bianco [1566, 1915]. Biard [1799]. bib [1400]. Bibbia [2141]. bibl [28, 30, 58, 194, 250, 251, 193, 196, 198, 192, 248, 249, 188, 305, 443, 512, 442, 446, 445, 574, 649, 715, 717, 716, 719, 964, 890, 836, 830, 1102, 1103, 1104, 1277, 1279, 1280, 1273, 1187, 1188, 1281, 1446, 1357, 1445, 219, 218, 833, 381, 380, 379, 252, 449]. Bible [582, 524]. biblica [2720]. Biblico [707]. Biblio [3426]. Biblio-grafica [3426]. Bibliografia [2620, 214, 309, 2380, 2260, 86, 576, 3838, 2117, 753, 749, 4083, 85]. Bibliografica [1979, 2261, 3426, 929, 1252, 809]. Bibliografici [69, 1660, 290, 350, 935, 1156, 1250, 1858, 553, 179, 231, 1009, 1338, 1416, 1507]. Bibliografie [989]. Bibliographia [3941, 2114]. Bibliographic [1126, 3245]. Bibliographical [990, 1479, 12]. Bibliographie [1859, 557, 3893, 1079, 893, 4000, 2165, 2236]. Bibliographies [1898, 244, 10, 3944, 3945]. Bibliographique [2267]. bibliographiques [3245]. bibliographische [183]. Bibliographisches [698]. Bibliography [405, 669, 823, 907, 1478, 1556, 1635, 1898, 214, 244, 55, 271, 326, 10, 35, 40, 33, 34, 43, 42, 21, 3838, 422, 3893, 3941, 3944, 3945, 4083, 2260, 4000, 1387, 38, 1660]. Bibliopolis [86, 444, 274, 2157, 3057, 230, 2914]. Bibliosofia [1417]. Biblioteca [39, 61, 517, 583, 617, 764, 804, 995, 1427, 1946, 1964, 2035, 1314, 221, 1235, 2522, 447, 601, 576, 718, 1360, 809, 2453, 2577, 2383, 42, 1219, 2614, 1706, 986, 1979, 1809, 389, 277, 1062, 1011, 1015, 1163, 1136, 1591, 1663, 1876, 2574, 2480, 756, 1994, 2499, 44, 1068, 1252, 607, 1783, 1709, 281, 180, 72, 887, 2773, 2880, 1158, 3613, 2017, 2434, 2248, 2637, 1220, 1307, 2129, 1546, 1417]. biblioteche [991, 1156, 70, 556]. Bibliotek [1078]. Bibliotheca [556]. Bibliothecae [1055, 1157]. bibliothecis [1687]. Bibliothek [1171, 966, 1191, 276, 379]. Bibliothéque [2214, 2894, 1386, 1486, 1447, 2048, 2566, 280]. bibliothèques [1804]. Biblos [2458]. Bicchi [41]. bicentenaire [1090, 4047, 1440, 1326]. Bicentenary [2156, 1440]. Bicentennial [452]. Bicheno [1431]. Bidrag [2119]. Bielcke [100]. Biella [42]. Biémont [2548]. Biennale [165]. Bieri [3086]. Big [2546, 1114, 1114]. big-bang [1114]. Bigazzi [1820]. Bigliazzi [1283, 1461]. bij [2715]. Bijdragen [63, 1791]. Bilance [560, 1163, 1015]. Bilancia [1015, 1163, 1419, 1746, 1806, 1011]. Bilancia/Centro [1806]. Bilancia/Il [1163]. bilanciaio [1746]. bilancio [915]. Bild [3129]. bildende [547]. Bilder [2979, 3129]. Bildning [2200, 3852, 3853, 3854]. Bildungstrieb [687]. Bimbi [2012, 3777]. Bini [3718, 3505]. binnen [541]. binoculars [1584]. biodiversità [3818, 2322]. biodiversity [3818]. Bioelettrici [3103]. Biogeologie [1074, 3981]. biogeology [3981]. Biografia [2164, 2235, 2895, 2667, 2133]. Biografico [554]. Biografie [628, 1586]. Biographica [2576]. Biographical [2291, 689, 57, 3907]. Biographien [303, 2576]. Biographies [303, 1070, 1587, 2377, 2781]. Biographische [183]. Biography [888, 1945, 3047, 3050, 57, 2050, 2781, 3048, 2569, 1562, 218, 3862, 3907, 1843, 1577, 3008]. Biologia [784, 3299, 2043, 78, 241, 865, 912, 359, 637, 1173, 1024, 1272, 24, 957, 944, 1952, 2023, 3760, 1903]. Biologia. [3095]. Biological [1641, 3183, 371]. biologiche [912, 371]. Biologico [1971, 758]. Biologie [172, 1058, 3127, 3571, 2465]. biologiques [2902]. Biologische [3462]. Biology [2161, 3462, 4062, 1641, 2765, 912, 2427, 1058, 2866, 3571, 3760, 4072, 3931, 3975, 2161, 4072, 1847]. biomedica [2780]. Biomedical [2506]. Biomedicine [3081]. Biot [2611]. Biotechnologie [3968, 2857]. biotechnology [1758, 3968, 4046]. Biotti [952]. Biran [2499]. Birds [3412, 1331]. Birgit [3129]. Biringuccio [3047]. Birkhäuser [3002, 1149, 2725, 1247]. Birmingham [2779]. Biroma [2136]. Birth [3508, 1835, 1560, 2697, 3356, 311, 3976, 4089]. Birthing [2660]. bis [567, 2228, 497, 1862, 3902, 2598, 1058, 1141, 2576, 276, 1797, 1153, 1854, 3977, 1086, 1328, 1905, 4035, 4036]. bisher [787]. Bishop [1518, 3316, 719, 3993]. Bisogna [1899]. Bithynia [946]. Bizzarri [72]. Bizzarria [975]. Bizzozero [1551]. Björkman [3092]. Bl [2878, 914]. Black [2408, 68]. Black-Body [2408]. Blackman [2525]. Blackwell [1070, 196, 450, 451]. Blagioli [2412]. Blaise [2594]. Blanc [2826]. Blancard [1386]. Blanchard [1795, 2215, 2654, 1386]. Blanckaert [2843, 416, 1793, 2735]. Blanco [2315]. Blandin [1944]. Blankaert [3365]. Blastema [1210]. Blay [920, 1534, 2422]. Bleichmar [3261]. bless [839]. Bleuler [288]. Blick [1918]. Blinks [3383]. Bloch [1981]. Blondel [2273]. Blood [2863, 3617, 2863, 3573]. Bloomington [2561, 3078, 693]. Blu [2577]. Blu/Saggi [2577]. Blue [2134]. Blunt [222, 1314]. Bnoli [1919]. Bo [1787]. Boas [919]. Boato [1240]. Bobinger [2189]. Bocarro [2473]. Bocconi [1350]. Bocher [1438]. Bochum [1608]. Bochumer [1359, 380]. Bocking [1738]. Bodei [2501]. Bodet [795]. Bodies [2589, 2564, 3716, 3191, 3725, 2640, 2582, 2498, 1199, 3074]. Bodies/Political [2589]. Bodily [3690, 3733]. Bodin [1832]. Body [587, 2746, 3113, 3535, 3482, 3152, 3150, 1143, 3012, 3661, 2655, 2408, 3153, 3231, 2840, 3144, 3801, 2875, 3566]. Bodyworld [2265]. Boel [3617]. Boerhaave [266, 2555, 2554, 1422, 1342, 1792, 2195, 3690, 2556, 4058, 3821]. Boethius [1531, 2055, 23, 1153, 1912, 751, 376, 1069]. Boezio [85]. Bohr [3263, 1008, 1007, 2674, 3137, 3842, 3882, 3901, 3950]. Boi [2215]. Bois [201]. Bois-Reymond [201]. Bokförlag [972, 1023]. Boktryckeri [1441, 884]. Bolis [869]. Bollati [1921, 934, 2063, 2074, 2472, 2535, 2632, 2893, 2690, 2630, 2904, 758, 3031, 3546, 1093, 2111, 1321, 577, 2738, 1658, 1922, 1899, 1998, 765, 1244, 745, 2910]. Bollettino [3408, 1013, 1160, 1372, 1424]. Bollone [1426]. Bologna [89, 202, 1616, 1502, 934, 59, 1902, 336, 493, 4045, 3094, 1066, 2882, 1815, 2155, 2543, 2836, 4052, 2721, 180, 2558, 85, 448, 1105, 955, 170, 257, 519, 792, 1462, 1417, 1869, 2007, 2315, 2789, 1428, 1909, 38, 21, 328, 2941, 3594, 3029, 1148, 912, 331, 641, 1094, 2468, 871, 199, 656, 780, 779, 1062, 1158, 1415, 1411, 1512, 1591, 1745, 1746, 2002, 1923, 2070, 2077, 2078, 2148, 2213, 2153, 2225, 2548, 1712, 570, 550, 755, 684, 969, 1459, 678, 3125, 3226, 1639, 3206, 811, 913, 2603, 2460, 2691, 2498, 2954, 1325, 1530, 1802, 2322, 493]. Bologna [2293, 3783, 1701, 3487, 732, 1039, 3343]. Bologna-Berlin [3226]. Bolognanel [1105]. bolognese [448, 2213]. Bolometer [2408]. Bolt [2152]. Bolzano [1668, 948]. bomb [3974, 2546, 2977]. Bombay [329]. Bompiani [2276, 2644]. bona [1806]. Bonadonna [786]. Bonaparte [2612]. Bonaventura [2328, 447]. Bonaventuras [248]. Bonaventure [447]. Bondi [171, 1015]. Bondio [2682, 2863, 3130]. Bonechi [2337, 2288]. Bonella [1335]. Bonelli [3040, 3820]. Bonera [2156, 1452]. Bones [2103, 3217, 939, 3716]. Bonfigli [1335]. Bonfioli [727]. Bonicalzi [2057]. Bonifacio [871]. Bonifazio [3177]. Bonino [2467, 2208]. Boniolo [1901]. Bonn [173, 173]. Bonner [249]. Bonnet [171, 1700]. Bonoli [633, 2213, 3315]. Bononia [2603, 2691]. Bononiensis [1094]. bons [2015]. Bont [3639]. Bonuzzi [1098]. Book [39, 11, 16, 15, 89, 41, 77, 60, 62, 61, 167, 259, 253, 202, 227, 164, 197, 258, 195, 242, 168, 189, 224, 283, 301, 302, 335, 296, 303, 342, 313, 299, 516, 507, 515, 517, 441, 520, 454, 491, 439, 440, 514, 414, 438, 405, 490, 537, 580, 629, 587, 592, 573, 544, 583, 615, 617, 642, 651, 567, 579, 721, 794, 802, 714, 728, 801, 759, 679, 777, 724, 805, 764, 795, 672, 685, 722, 804, 837, 852, 885, 894, 896, 972, 961, 889, 918, 923, 965, 959, 895, 891, 968, 888, 842, 1002, 1110, 1070, 1005, 1091, 1106, 995, 1055]. Book [1087, 1090, 1078, 1081, 1019, 1023, 1108, 1199, 1200, 1231, 1288, 1234, 1139, 1201, 1169, 1194, 1264, 1146, 1261, 1204, 1333, 1367, 1340, 1365, 1313, 1346, 1362, 1402, 1441, 1457, 1406, 1427, 1364, 1370, 1449, 1404, 1454, 1616, 1531, 1535, 1536, 1528, 1617, 1502, 1537, 1539, 1490, 1562, 1602, 1605, 1606, 1589, 1587, 1680, 1683, 1691, 1657, 1763, 1649, 1725, 1660, 1692, 1819, 1818, 1846, 1853, 1805, 1795, 1945, 2003, 1910, 2016, 2119, 2120, 2075, 2102, 2056, 2067, 2194, 2227, 2143, 2145, 2228, 2159, 2200, 2160, 2267, 2312, 2318, 2421, 2420, 2370, 2424, 2488, 2463, 2467, 2496, 2495, 2500, 2464, 2560, 2586, 2601]. Book [2527, 2715, 2675, 2746, 2776, 2793, 2887, 2953, 2986, 2916, 3041, 3016, 3299, 3347, 3368, 3508, 3569, 3598, 3739, 2418, 2459, 2516, 2573, 2647, 2708, 2710, 2879, 2881, 3000, 2055, 3477, 2843, 772, 2717, 1227, 1393, 2607, 2899, 3534, 1852, 1921, 2878, 3454, 3509, 2889, 3065, 315, 593, 897, 1239, 1578, 1949, 1956, 1957, 1959, 1962, 2020, 2027, 2028, 2030, 2033, 1963, 2034, 1951, 2022, 1952, 1953, 1954, 2023, 2024, 2025, 1955, 1960, 1966, 2026, 2031, 2037, 1967, 2038, 1948, 1950, 1958, 2019, 2021, 2029, 1947, 1965, 2018, 2036, 2012, 1961]. Book [2032, 2164, 2235, 2163, 2168, 2234, 2239, 2233, 2162, 2169, 2240, 2165, 2236, 2167, 2238, 2166, 2237, 2329, 2331, 2337, 2328, 2332, 2326, 2335, 2330, 2333, 2334, 2336, 2338, 2339, 2327, 2434, 2387, 2388, 2389, 2435, 2436, 2386, 3141, 3641, 3659, 3677, 3697, 3720, 3743, 3113, 3395, 3425, 3689, 2142, 339, 496, 552, 653, 588, 807, 726, 934, 2942, 3543, 3499, 1736, 3688, 3328, 3570, 3735, 2425, 2462, 2606, 2730, 2897, 59, 1412, 1576, 1573, 1659, 1801, 1902, 3116, 3064, 3002, 3107, 3120, 3286, 3283, 3562, 228, 222]. Book [340, 336, 282, 416, 493, 421, 675, 688, 931, 1418, 1793, 1903, 2053, 2612, 2896, 3482, 2669, 3326, 3094, 2735, 2737, 3005, 2199, 18, 3631, 2777, 508, 456, 495, 497, 623, 621, 533, 548, 568, 543, 569, 584, 591, 652, 585, 620, 532, 612, 686, 796, 677, 676, 725, 703, 698, 673, 696, 699, 841, 884, 834, 846, 1006, 1097, 1095, 1060, 1066, 1155, 1197, 1171, 1145, 1314, 1316, 1322, 1440, 1317, 1439, 1405, 1368, 1387, 1611, 1488, 1567, 1487, 1564, 1569, 1568, 1518, 1495, 1533, 1608, 1648, 1661, 1681, 1727, 1679, 1643, 1682]. Book [1765, 1689, 1816, 1804, 1811, 1817, 1898, 1911, 1981, 2517, 2530, 2538, 2822, 2894, 2933, 2898, 2882, 2936, 3063, 3040, 3059, 3071, 3121, 3109, 3362, 3475, 3674, 2852, 2851, 3081, 221, 619, 1236, 1235, 1403, 1496, 1842, 2491, 2522, 2707, 2864, 2770, 2981, 2891, 3068, 3255, 3254, 3280, 3282, 3541, 643, 932, 1226, 1613, 1581, 1580, 1574, 1769, 1693, 1815, 1996, 2064, 2063, 2116, 2074, 2073, 2155, 2320, 2431, 2364, 2471, 1004, 3694, 2664, 2780, 1790, 2653, 1835, 166, 287, 278, 3072, 3736, 2805, 2964, 2580, 2203, 2472, 2144, 2543, 3080, 2093, 2750, 2682]. Book [2284, 2528, 223, 2620, 2667, 3526, 3442, 3545, 3025, 3653, 1738, 2836, 2837, 75, 63, 74, 234, 235, 260, 855, 847, 843, 856, 1085, 1112, 1808, 2524, 2631, 2721, 2797, 2959, 2906, 3009, 3510, 2850, 3023, 3316, 2865, 549, 1991, 10, 180, 181, 182, 3476, 2313, 2314, 2376, 2105, 2689, 2106, 2972, 2798, 2769, 2309, 3279, 2523, 2588, 3015, 3204, 3414, 3483, 2853, 2979, 2908, 3082, 3084, 3129, 3123, 3122, 3268, 3369, 2849, 3021, 3512, 2685, 2826, 2616, 2551, 2558, 2671, 2791, 2909, 3035, 2542, 2974, 2969, 3077, 3030, 2670, 28, 72, 29]. Book [85, 30, 23, 31, 58, 86, 32, 194, 243, 252, 250, 251, 193, 196, 198, 192, 248, 249, 188, 312, 311, 307, 310, 306, 305, 309, 443, 512, 442, 444, 446, 511, 447, 445, 449, 448, 571, 576, 648, 574, 649, 715, 798, 720, 717, 716, 718, 719, 966, 887, 964, 873, 886, 890, 836, 857, 830, 915, 994, 991, 1102, 1103, 1105, 1077, 1104, 996, 990, 1191, 1277, 1283, 1278, 1157, 1279, 1280, 1273, 1189, 1187, 1188, 1281, 1156, 1359, 1446, 1358, 1443, 1447, 1357, 1442, 1445, 1444, 1360, 1448, 1687, 1688, 1685, 1684, 1686, 1984, 1985]. Book [1941, 3001, 70, 274, 279, 955, 949, 1059, 170, 257, 519, 572, 792, 809, 951, 971, 1088, 1462, 1417, 2611, 3573, 498, 2283, 2279, 2648, 1726, 2963, 3602, 2958, 2553, 1151, 1500, 2917, 3124, 3544, 3500, 3637, 3672, 3495, 349, 538, 551, 1143, 1566, 1717, 2103, 2371, 2602, 2619, 2665, 2957, 2592, 1067, 1243, 1138, 1238, 1147, 1263, 1366, 1327, 1612, 1516, 1515, 1499, 1572, 1620, 1494, 1603, 1668, 1759, 1758, 1670, 1862, 1869, 1791, 1859, 1820, 1879, 2010, 2007, 2014, 1928, 2080, 2079, 2059, 2066, 2161, 2232, 2216, 2215, 2224, 2229, 2157, 2156, 2270, 2276, 2315, 2321]. Book [2311, 2381, 2504, 2490, 2497, 2605, 2727, 2789, 3441, 3408, 3457, 2589, 3076, 1588, 1667, 2535, 2824, 2920, 3263, 3022, 614, 1063, 1401, 1721, 1789, 2579, 2658, 2834, 2883, 2577, 1428, 1538, 334, 2755, 2863, 2962, 3027, 3066, 3006, 3115, 3209, 3256, 3572, 1111, 1909, 2383, 3397, 3013, 1152, 1335, 2978, 1248, 35, 40, 17, 33, 45, 34, 38, 43, 42, 56, 21, 3205, 3692, 2520, 3127, 2949, 284, 413, 926, 3298, 2536, 2756, 3043, 2662, 2831, 3440, 3630, 2519, 2838, 3461, 3090, 2657, 3714, 2594, 3111, 3435, 2058, 1669, 2379, 2546, 1149, 3087, 3611]. Book [3717, 2531, 2895, 2572, 2646, 2874, 2876, 3061, 3060, 3003, 3012, 3400, 3462, 1908, 2808, 2655, 1164, 2550, 365, 366, 2774, 2941, 2706, 3658, 57, 64, 219, 218, 173, 226, 172, 220, 244, 420, 547, 557, 542, 1715, 2888, 3415, 1917, 3365, 3675, 3418, 2943, 3594, 3249, 3312, 3366, 3491, 2751, 2907, 2719, 2827, 2819, 2948, 3344, 3345, 3409, 3530, 275, 408, 752, 2934, 3742, 1995, 314, 2632, 2839, 2600, 2988, 2742, 682, 3057, 2595, 2885, 3070, 2955, 385, 368, 2672, 3029, 613, 2946, 2461, 3346, 3486, 3511, 333, 3108, 2794, 3411, 2893, 3289, 3288]. Book [2633, 2518, 2470, 2101, 3131, 2614, 2654, 2800, 2829, 2922, 2905, 3497, 1242, 2740, 3225, 3281, 2502, 2712, 2757, 3399, 2642, 2288, 1735, 1505, 1656, 343, 2690, 1260, 2630, 2651, 2999, 3620, 3599, 1492, 2904, 2787, 169, 341, 760, 769, 767, 925, 919, 933, 1246, 1409, 1395, 1575, 3568, 2771, 2947, 754, 2747, 1732, 1398, 1722, 1987, 1543, 2377, 2557, 2973, 3639, 3253, 2308, 2783, 1997, 174, 2828, 3110, 80, 81, 67, 79, 225, 288, 286, 300, 504, 3265, 3331, 2140, 2214, 2202, 2139, 2492, 2623, 2720, 2652, 2835, 2773, 2892, 2268, 2781]. Book [3067, 3079, 415, 2832, 3494, 2656, 2950, 3438, 3649, 3671, 2532, 2681, 2854, 2859, 2915, 3656, 3412, 2718, 2862, 2061, 3670, 2561, 3033, 829, 2709, 2596, 410, 536, 606, 748, 928, 845, 1196, 1315, 1484, 1641, 1664, 1982, 2094, 2095, 2134, 2197, 2137, 2493, 2629, 3718, 2323, 3024, 3074, 422, 1148, 1841, 2494, 3719, 3738, 3017, 2533, 1918, 2970, 3091, 1724, 1714, 217, 2196, 3734, 3455, 2900, 681, 2736, 2977, 3243, 2912, 65, 191, 204, 285, 418, 419, 624, 654, 770, 758, 787, 683, 763, 927, 850, 930, 912, 848, 866, 1407]. Book [1741, 1878, 1875, 1915, 1992, 1901, 2015, 2004, 2013, 1943, 1944, 2362, 2427, 2966, 3458, 2825, 407, 1190, 2590, 3635, 1640, 874, 880, 883, 881, 1183, 2945, 2277, 2419, 2598, 2754, 3014, 3716, 2938, 3597, 2965, 3427, 3480, 3258, 332, 360, 331, 503, 434, 436, 437, 505, 586, 639, 638, 641, 566, 640, 645, 706, 709, 786, 789, 785, 729, 707, 788, 867, 954, 956, 868, 958, 947, 948, 952, 950, 877, 872, 869, 882, 1027, 1094, 1101, 1115, 1096, 1099, 1028, 1026, 1092, 1030, 1098, 1176, 1178, 1275, 1287, 1181, 1177, 1175, 1268, 1173, 1192, 1271, 1179]. Book [1431, 1436, 1438, 1434, 1432, 1595, 1598, 1596, 1597, 1594, 1522, 1674, 1678, 1673, 1675, 1716, 1752, 1755, 1672, 1753, 1756, 1677, 1754, 1676, 1867, 1865, 1812, 1809, 1864, 1866, 1810, 1936, 1925, 1934, 2006, 2011, 1933, 1927, 1935, 2072, 2076, 2071, 2152, 2154, 2220, 2222, 2265, 2275, 2413, 2430, 2373, 2365, 2375, 2468, 3496, 3069, 3018, 2659, 3327, 2287, 3031, 3402, 3481, 3498, 3546, 2713, 3668, 3542, 3439, 3529, 3695, 2940, 14, 2525, 1008, 2741, 499, 432, 433, 500, 435, 644, 564, 554, 712, 713, 810, 710, 697, 708, 711, 871, 863, 879, 870, 876, 875, 1093]. Book [1100, 1029, 1186, 1184, 1274, 1180, 1182, 1433, 1523, 1521, 1652, 839, 1057, 2918, 2450, 2790, 2801, 3636, 3601, 3693, 3713, 2555, 2505, 390, 389, 330, 1913, 2743, 3507, 2775, 2884, 2944, 3426, 2987, 2684, 3267, 3251, 3313, 2821, 2961, 3618, 2207, 344, 917, 2799, 3036, 3619, 3633, 36, 25, 26, 200, 201, 256, 171, 190, 183, 622, 2615, 3638, 616, 920, 1001, 1718, 1843, 3277, 2554, 2683, 3208, 2660, 1849, 1489, 2693, 3287, 3257, 3691, 2880, 88, 71, 20, 240, 187, 184, 185, 238, 237, 199, 233, 356, 354, 291, 355]. Book [293, 295, 352, 277, 424, 431, 518, 452, 453, 457, 430, 406, 427, 428, 502, 429, 409, 426, 425, 455, 561, 562, 540, 560, 655, 563, 581, 631, 589, 633, 556, 590, 575, 635, 656, 559, 704, 762, 730, 705, 761, 700, 776, 780, 779, 701, 702, 782, 727, 774, 778, 844, 970, 858, 878, 893, 939, 835, 832, 862, 916, 861, 963, 937, 864, 941, 953, 854, 859, 938, 1014, 1062, 1080, 1082, 1011, 1016, 1013, 1072, 1012, 1083, 1015, 1020, 1084, 1107, 992, 1056, 998, 1018, 1033, 1161, 1170, 1251, 1258]. Book [1162, 1256, 1257, 1284, 1259, 1163, 1240, 1255, 1136, 1158, 1166, 1167, 1241, 1159, 1185, 1195, 1193, 1160, 1165, 1172, 1453, 1369, 1429, 1430, 1456, 1355, 1354, 1372, 1415, 1451, 1452, 1411, 1422, 1343, 1424, 1461, 1421, 1435, 1350, 1373, 1437, 1342, 1419, 1320, 1423, 1584, 1585, 1609, 1525, 1599, 1512, 1511, 1544, 1510, 1591, 1532, 1614, 1570, 1619, 1563, 1743, 1744, 1690, 1665, 1663, 1730, 1748, 1766, 1768, 1750, 1662, 1666, 1745, 1746, 1749, 1671, 1695, 1877, 1806, 1794, 1874, 1813, 1787, 1861, 1863, 1873, 1792, 1876, 1807, 1860, 1871, 1872, 2002]. Book [2005, 1930, 1919, 1931, 2001, 1923, 2009, 1926, 2069, 2070, 2112, 2115, 2117, 2077, 2078, 2110, 2113, 2118, 2068, 2098, 2114, 2111, 2148, 2151, 2226, 2213, 2153, 2212, 2150, 2149, 2195, 2217, 2221, 2225, 2271, 2317, 2310, 2307, 2319, 2415, 2414, 2369, 2368, 2423, 2363, 2541, 2548, 2574, 558, 1065, 999, 2877, 337, 338, 922, 929, 1058, 1600, 1764, 2624, 2529, 2732, 2785, 417, 545, 626, 2855, 3093, 3492, 1504, 1497, 1655, 2792, 1712, 535, 610, 534, 2858, 3333, 2282, 2686, 3676, 2677, 1074, 1731, 1838, 3088, 3600, 1848, 1904, 2206, 1321, 1399, 3038, 2584, 1645, 2716]. Book [2809, 3224, 3314, 3456, 2280, 66, 1071, 1410, 2745, 1610, 3485, 2753, 3210, 3078, 2976, 833, 1319, 2204, 2886, 3712, 2617, 494, 618, 3690, 3715, 2731, 3114, 3614, 2833, 3008, 2060, 2676, 2844, 216, 2975, 3687, 3118, 812, 1114, 3652, 255, 304, 450, 451, 570, 550, 771, 680, 756, 757, 791, 755, 684, 969, 1052, 1414, 1459, 1788, 2722, 2583, 2552, 2749, 2725, 2841, 2952, 2875, 3010, 1141, 2576, 2537, 3284, 3443, 1855, 165, 276, 753, 749, 799, 800, 678, 997, 1228, 1229, 1285, 1396, 1392, 1397, 2645, 2840, 3083, 2767, 1265, 1249, 1618, 3125, 2643, 3417, 2678]. Book [3525, 1323, 2663, 308, 578, 577, 838, 1061, 1230, 1647, 1646, 2734, 229, 186, 163, 203, 345, 347, 1073, 1075, 1237, 1289, 1458, 1394, 1460, 1506, 1694, 1737, 1653, 1729, 1786, 1799, 1856, 1980, 2052, 2104, 2231, 2272, 2366, 2426, 2382, 2412, 2506, 2618, 2559, 2738, 2673, 2867, 2866, 2983, 2985, 3039, 3085, 3130, 3226, 1413, 1579, 1639, 1658, 1851, 1922, 1899, 1998, 1994, 2107, 2062, 2132, 2428, 2429, 2380, 2469, 2499, 3034, 2807, 611, 3657, 44, 2830, 3446, 3241, 83, 674, 3453, 2803, 2984, 3740, 2316, 2384, 2374, 2432, 2608, 2539, 2729]. Book [625, 921, 1400, 1723, 2585, 1797, 364, 363, 361, 362, 2939, 2951, 2679, 176, 2597, 1503, 1501, 3737, 3096, 1247, 1542, 2728, 3062, 411, 1140, 2099, 2786, 2285, 1847, 2998, 13, 2457, 1153, 3434, 2286, 2804, 2609, 3044, 3019, 2784, 3612, 3444, 3398, 3634, 3396, 2845, 3206, 2626, 3058, 230, 2278, 2489, 3126, 3564, 1840, 3343, 3410, 3460, 3493, 3528, 1326, 3733, 1734, 765, 766, 924, 851, 1154, 1334, 1408, 1541, 2273, 2378, 2507, 1854, 3547, 3741, 3506, 3332, 2372, 3710, 2593, 348, 289, 689, 808, 806, 692, 693, 811, 723, 803, 687, 691, 957, 860]. Book [892, 849, 942, 913, 1007, 1086, 1113, 993, 1034, 1076, 1079, 1109, 1068, 1089, 1203, 1244, 1202, 1144, 1198, 1329, 1328, 1455, 1349, 1425, 1330, 1426, 1348, 1491, 1571, 1713, 1650, 1767, 1654, 1803, 1796, 1988, 1912, 2108, 2466, 2465, 2806, 2919, 2967, 3032, 3330, 3334, 3401, 3429, 2208, 2198, 2603, 2687, 3616, 3617, 12, 2591, 751, 329, 750, 1733, 1844, 1907, 2274, 2982, 3086, 3011, 3203, 387, 388, 384, 371, 373, 372, 370, 412, 3075, 280, 582, 541, 555, 539, 2856, 1720, 2861, 3654, 2911, 1999, 2890, 2921, 2914, 1337, 2544, 3261, 3416, 1993, 2057, 2269, 694, 2903]. Book [3095, 378, 377, 376, 381, 380, 379, 382, 2902, 1837, 1857, 3567, 3655, 3711, 1836, 2136, 2458, 2460, 2549, 2625, 2650, 2820, 2937, 3250, 3325, 3315, 3669, 3673, 2935, 2587, 3632, 3478, 2501, 3092, 2582, 2138, 3004, 3278, 3445, 2901, 2281, 3447, 2581, 2691, 840, 1986, 3527, 3595, 3329, 3364, 3613, 2846, 1003, 1064, 1218, 1232, 1233, 1142, 1291, 1386, 1450, 1371, 1388, 1363, 1318, 1420, 1592, 1615, 1486, 1534, 1483, 1565, 1642, 1868, 2100, 2205, 2201, 2324, 2433, 2422, 2498, 2503, 2752, 2796, 2810, 2954, 2956, 1137, 1267, 1270, 1252, 1269, 1325, 1324, 1352, 1331]. Book [1519, 1540, 1530, 1485, 1644, 607, 608, 746, 745, 747, 946, 914, 945, 1224, 1133, 1174, 1134, 1223, 1225, 1391, 2714, 3413, 3361, 3037, 2096, 2147, 2456, 1990, 2622, 1802, 1906, 1900, 2141, 2613, 2545, 2674, 2739, 2910, 281, 489, 3609, 3239, 1069, 2604, 2534, 2724, 2779, 1920, 1728, 1800, 2556, 1498, 1493, 1651, 1989, 2146, 3259, 2644, 2628, 1332, 1845, 609, 3650, 1850, 2097, 2322, 2367, 3073, 627, 831, 1000, 1150, 1290, 1336, 1577, 1798, 1870, 1814, 2133, 2266, 2627, 690, 175, 177, 178, 346, 565, 546, 768, 2772, 492, 68, 695, 1245, 2562, 3459]. Book [2848, 2802, 2857, 1905, 3565, 2913, 2688, 3363, 2971, 3428, 2748, 3479, 3596, 2726, 2960, 2782, 2571, 3566, 2968, 2668, 2578, 2692, 3026, 3089, 3264, 3285, 3367, 3571, 1914, 1916, 1135, 1389, 1390, 1482, 1719, 1839, 1834, 1983, 2054, 2135, 2416, 2417, 3164]. Books [491, 1078, 1169, 1805, 2527, 1852, 1903, 2777, 696, 699, 1439, 2542, 1448, 3558, 2546, 3100, 2922, 2740, 3420, 2987, 2684, 2683, 3114, 1397, 2983, 2935, 945, 2147, 2545, 2604, 2779, 3419, 3801, 3942]. Bookshelf [2121, 2162, 2325, 2385]. Boole [176]. Borbone [171]. Borchi [700, 776, 1014, 1257, 1807]. Borck [2736]. Border [3219]. Borders [759]. Bordiga [12]. Borelli [2333, 3591, 1782, 1973]. Borgato [2387]. Borgia [1107, 2172]. Boring [3047]. Boringhieri [1921, 934, 2063, 2074, 2472, 442, 2535, 2632, 2893, 2690, 2630, 2904, 67, 758, 3031, 3546, 1093, 2111, 1321, 577, 2738, 1658, 1922, 1899, 1998, 765, 1244, 745, 2910]. Borja [3374]. Bork [2707]. Borla [932]. Borlotti [2415]. Born [3460, 2646]. Bornus [2421]. Boroli [2916]. Borrelli [2164, 2235, 3410, 1990]. Borsa [496]. Bortolotti [38]. Bos [28]. Bosch [3116, 2484]. Bosch-Omori [2484]. boschivo [2009]. Boschung [1764, 2529]. Boscovich [452, 858, 412, 898, 212, 2175]. Bose [2485]. Boselli [593]. Bossi [1570, 452]. Boston [16, 2573, 2647, 2708, 2462, 2897, 3002, 2933, 28, 955, 1149, 218, 2999, 2771, 418, 1570, 1600, 3224, 2840, 2939, 1247, 2937, 2971]. Boswell [573]. Botanic [1370, 1314, 1137]. Botanica [509, 1314, 1809, 82, 1266, 1252, 1351, 1517, 2391, 2421, 686, 83, 3976, 1519]. Botanical [1314, 3501, 3345, 3704, 3261, 1267, 1270, 2391, 3779, 3709]. botaniche [1270]. Botanici [1975, 1267]. Botanico [83, 1270, 1352, 2125, 863, 917, 1137, 1269]. botanico-farmacologiche [1269]. Botanik [2228, 3902]. Botanique [1305]. botanischen [3530]. botanistes [222]. botanists [222]. Botany [2527, 2657, 2477, 3114, 1252, 1225, 2228, 1305, 3902, 3976]. Botey [2114]. Both [1769]. Bots [280, 541]. Bottazzini [1658, 1348]. bottle [2806]. Boudet [1447]. Boudon [3611]. Boudon-Millot [3611]. Boudot [625]. Bougard [2145]. Boulder [1763, 2505, 3052]. Bouley [3455]. Boundaries [2928]. Bourel [1799]. bourgeois [1537, 2761]. Bourguet [3461]. Bovi [330]. Bovolo [3426]. Bowden [2724]. Bowen [2516, 3417]. Bowers [847]. Bowler [2747, 191, 2966]. Box [3468, 3968, 2857]. Boydell [615, 1490, 2143, 1063, 275, 616]. Boyle [60, 168, 515, 1231, 1139, 2143, 1533, 3653, 2873, 3615, 1776, 2873, 3883]. Bozzi [1532, 550]. Bozzolo [1331]. br [1190]. Bracali [2488]. Braccesi [792, 2213]. bracci [1163]. Bracegirdle [2683, 3208]. Bradford [3455]. Bradley [3733]. Bradwardine [1686]. Brahe [888]. Braida [1539]. Braidense [221]. Brain [3379, 3151, 3415, 36, 1072, 3385, 2506, 3381, 2967, 3628, 3874, 3933, 3619]. Brains [3488, 3446, 3874]. Brancacci [2933]. Branching [3668]. Brander [2294]. Brandolini [1502]. Brani [896]. Brasile [588]. Brass [761]. Braun [2529, 860]. Braun-Bucher [2529]. Braunschweig [259, 1724]. Braunschweig/Wiesbaden [259]. Braunschweiger [1724]. Braunwarth [1854]. Bravieri [390, 187]. Bray [580]. Brazier [26, 545]. Brazil [1557]. Breach [1944, 892]. Bredekamp [3129, 2943, 3225]. Breger [1001]. Breidbach [2908, 2734]. Breithaupt [234]. Brenda [2981]. Brenner [625, 1003, 831]. Brenngerte [1083]. Brenni [2168, 2239, 3739, 3592]. Brent [3345]. Brepols [2145, 2267, 2878, 2770, 3526, 2309, 1280, 1273, 1444, 1685, 1984, 2596, 2615, 2369, 2886, 2833, 3284, 3026]. Brera [71, 291, 631, 2212, 462, 631]. Bresadola [2955, 2904, 2078]. Brescia [601, 1115, 782, 970, 1013, 1373, 1713]. Bresciana [1373]. Bresciani [1013, 1160]. bresciano [1373]. Breslau [1731]. Bret [1795, 3368, 1495, 2229, 2118, 3368]. Bretagne [1612, 2060, 3829]. Bretschneider [2741, 1135]. Brett [2731]. Brev [1441, 884]. Brevaglieri [3614]. breve [2503, 2471, 2631, 2624, 2810]. Brevi [1003]. Brewer [3044]. Brewin [2779]. Brian [1194, 1694, 1799, 1907, 1993, 629, 166, 1196, 2775, 2683, 3208]. Brianta [3021]. Bridge [2050]. Brief [293]. Briefe [307, 64, 226, 787, 1001, 1764, 259]. Briefen [64]. Briefwechsel [64, 622, 1001, 411, 751, 201]. Brieux [2183]. Brieven [802, 1367]. Brigati [1533]. Briggs [679]. Brigham [1196]. Brigitte [1187, 3091, 2161]. Brill [891, 1110, 2573, 2897, 2933, 29, 1277, 1188, 1358, 1445, 1686, 2999, 2771, 753, 749, 997, 1229, 2840, 2939, 3058, 2937, 914, 1225, 1983]. Brioschi [2631]. Brisson [2648]. Bristol [772, 218, 1057]. Bristol-New [772]. Britain [651, 923, 1449, 1818, 1568, 1808, 3023, 3637, 2957, 2079, 3714, 1109, 2861, 1490, 3819, 3829, 2751, 3879, 1466, 3935, 4031, 3715]. Britannici [534, 3434]. britannico [2467]. British [454, 642, 1406, 2560, 2607, 1439, 1518, 2964, 2906, 1566, 3079, 927, 1872, 534, 2686, 2617, 3083, 840, 2956, 3837, 2984]. Brittleness [3482]. Broader [3559]. Broberg [972]. Broc [1190]. Brocchieri [718]. brochures [3942]. Bronx [10]. Brook [2134]. Brooke [1395]. Brookline [845]. Brosses [893]. Brotherhood [517]. Brotóns [2114]. Brovarone [407]. Brown [592, 959, 1346, 2848]. Browne [532]. Bruce [840, 1063, 2999, 1618, 2591]. bruchi [1185]. Brückle [3225]. Brüder [1141]. Brugnatelli [1869, 2127]. Brun [1177]. Brun-Hool [1177]. Brunelleschia [1485]. Brunelli [1062]. Brunetta [1794]. Brunfels [2391]. Bruni [1438, 2293]. Bruniana [3259]. Brunn [139]. Bruno [2142, 309, 1186, 578, 1230, 2951, 3330, 1900, 852, 2192, 1375, 312, 40, 3127, 1438, 730, 970, 1162, 2877, 2992, 1230, 1579]. Brunora [3934]. Brunoro [2592, 3934]. Brunschön [2939]. Brunsdon [833]. Brunswick [227, 440, 811]. Brunswick-London [440]. Bruschi [1028]. Brush [1654]. Brussel [302]. Brussels [2833]. Bruxelles [1447, 1879, 295, 2231, 2469]. Bruzzaniti [1240]. Bruzzone [1096]. Bryden [939]. BSB [303, 568]. Bsel [2577]. Bub [1736]. Bubble [820]. Buccellati [2212, 2622]. Bucchioni [1062]. Bucciantini [2968]. Buch [1941, 787, 1391]. Buchanan [2981]. Buchdahl [1499]. Bucher [2529]. Buchges [639]. Buchgesellschaft [188, 1446]. Büchner [363]. Buchwald [1573]. Buck [1191]. Buenos [3264]. Buess [172]. Buffalo [968, 2953, 2139]. Buffon [4047, 518, 1326, 2896, 4047, 2729]. Buffon88 [1326]. bugie [1112, 3831]. Building [2305, 2867]. buildings [2330]. built [2688]. Bulino [560, 1011, 1163]. Bulletin [714]. Bulzoni [1754, 1033]. bumani [1962]. Buniva [2313, 2314, 2376]. Bunsen [2403]. buoi [2726]. buona [3813].

Buonocore [1816]. Burchard [3569]. Burchell [2945]. Burczyk [1256]. Burczyk-Marona [1256]. Bureau [1350, 3657, 2638]. Burghard [695]. Burgos [3722]. Burkhardt [3076, 1407, 2004, 851]. Burleigh [379]. Burlington [2707, 2981]. Burlini [1893]. Burnett [830, 1102, 1358, 916, 2998, 378]. Burns [2381, 3091]. Burton [2561]. Busard [376]. Buschinger [32]. Bush [3402]. business [2745]. Busmann [101]. Bussi [1283]. Bussola [599]. Bussole [2624]. Bussotti [3120]. busti [2142]. Busts [3353]. butchers [4038]. Buti [2088]. Buzzetti [1105, 2213]. Bycroft [1072]. Bynum [224, 532, 2909, 225]. Byron [1113]. byzantine [1669].

C [801, 679, 1367, 1427, 1853, 1945, 2586, 2516, 59, 543, 584, 841, 1643, 1235, 1496, 1226, 2074, 2155, 2528, 235, 1279, 1273, 1156, 1401, 4057, 341, 1982, 2095, 3891, 848, 407, 3635, 1179, 1675, 1652, 3713, 563, 556, 761, 2114, 558, 4079, 2643, 163, 2380, 83, 892, 993, 2108, 2465, 387, 1857, 1142, 1324, 1224, 1728, 3023, 2662, 3468, 850, 3722, 2725, 3957, 735, 539, 1297]. C. [11, 2459, 2974, 1687, 1688, 1640, 434, 1503]. C.I.D.I [189]. C.I.L [1408]. C.N.A.M [702]. C.S.I.C [2114, 2097]. C.T.H.S [728, 243, 836]. C.U.E.C [2113]. c1880 [3638]. c2009 [3061]. CA [2753, 1444, 274, 2520, 829, 3731, 816, 3410]. Ca.Fo.S. [1696]. cabala [752, 1565]. Cabinet [3509, 2700, 3218, 1792, 3634, 3489, 3823, 1422, 4019]. cabinets [2068, 3990, 2940]. Caccioppoli [1706]. Cacucci [2665, 2727, 2536, 3090]. Cadeddu [1693, 957]. cadémie [753]. Cadencia [3264]. Cadet [1476]. Cadice [2727, 2727]. Cadmo [2227, 1034]. Cadrans [322]. caduta [537]. Caelius [1391]. Caelo [1189]. Cafagna [1540]. Cafarella [1166, 1167, 1745]. Cafasso [893]. Cagiano [893]. Cagliano [1673]. Cagliari [2140, 2113, 1399]. Cagliaritana [2140, 2113]. Cahan [1340]. Cahier [1224, 1626]. Cahiers [3280]. Cain [763]. Caizares [2920]. Caizares-Esguerra [2920]. Cajal [45, 45, 3415]. Calabi [2272]. calamita [2722]. Calanchi [780]. Calasanzio [1257]. Calascione [44]. calcolatori [2111]. calcolo [1140]. calcul [886, 681]. Calculus [1141]. Calcutta [1750]. Calder [173]. Calderini [21]. calderone [1753]. Calderoni [2153]. Calegari [1746]. Calendar [2003, 3809, 3836, 3612]. calendari [2548]. Calendario [1227, 2153]. Calendario-lunario. [2153]. Calendars [2879]. Califfati [1434]. California [441, 685, 1333, 2516, 846, 182, 3084, 2671, 1500, 2885, 1987, 191, 770, 1319, 2382, 2784, 1845]. Calisi [937, 2110]. Called [3276, 2870]. Calledda [2113]. Callendar [2971]. Calles [1544]. Callwey [676]. Calmann [653]. Calmann-Lvi [653]. calore [1682]. Calvinists [539]. Calvino [2968]. Cambiagi [1728]. Cambiamenti [3103, 2064]. Cambiamento [898]. Cambridge [60, 164, 801, 679, 837, 918, 923, 888, 1002, 1005, 1199, 1139, 1194, 1261, 1340, 1535, 1536, 1617, 1562, 1587, 1683, 1649, 2003, 2120, 2075, 2102, 2464, 2527, 2793, 2879, 2889, 1736, 2730, 1412, 1576, 2737, 533, 1643, 2936, 3063, 1403, 2664, 166, 856, 1991, 2689, 2106, 2791, 2909, 3035, 2542, 194, 648, 649, 279, 349, 1143, 1758, 1670, 2066, 2311, 3076, 1588, 1721, 1152, 56, 2838, 2546, 2874, 3060, 2827, 2518, 2470, 1656, 1492, 169, 760, 769, 919, 933, 1395, 1575, 2050, 2377, 1997, 3265, 3067, 3079, 2656, 2709, 928, 1315, 1484, 1641, 1982, 2134]. Cambridge [2629, 422, 2900, 3243, 204, 419, 850, 848, 1407, 2004, 2362, 2427, 2966, 1640, 2938, 2076, 2071, 2152, 1274, 1433, 1652, 2790, 406, 409, 762, 1072, 3093, 2731, 3008, 2844, 1392, 3083, 1249, 347, 1694, 2866, 3034, 1503, 1542, 2728, 2457, 3044, 3528, 766, 851, 1154, 1334, 689, 808, 806, 692, 723, 803, 860, 849, 1109, 1203, 1198, 1425, 1654, 2806, 2591, 3086, 2911, 1999, 1337, 3278, 2846, 1592, 2752, 945, 3037, 2096, 2147, 2674, 3239, 1800, 1332, 1577, 1798, 768, 1245, 2913, 3917, 3183, 2482, 3311]. Cambridge-New [1535]. Cambridge/New [3008]. Camerlengo [1794]. Camerota [1955, 2026, 2163, 2234, 2167, 2238, 2339, 2140, 1399, 3687]. Camillo [1951, 2022, 3088, 3588]. Caminer [2787]. cammini [2627]. Campagna [949]. Campanella [2202, 1775]. Campanelliana [3259]. Campani [3688, 2293]. Campania [3390]. Campanile [3402]. Campigotto [2002, 2148]. Campo [2258]. Campogalliano [560, 1015, 1419, 1746, 1806]. Campus [1726]. Camuffo [1088]. Can [1993]. Canadelli [3077, 3285]. Canavero [715]. Cancer [1505, 3676]. Cancrini [288]. Cancro [2233, 2162]. candidates [4000]. Candle [1309]. Candle-Clock [1309]. Candore [2248]. Cannizzaro [764]. Cannocchiale [1555, 1469, 1704]. Cannocchiali [988]. Cannon [347]. Cannstatt [3064, 3283, 343, 344, 2734, 1202, 2465]. Canon [381]. Canone [1157, 1230]. Canova [2612]. Canovai [453]. Canstatt [1731]. Cantagalli [1260, 2597]. cantieri [2523]. Cantini [1283]. Canton [490, 672, 722, 842, 1264, 1204, 1365, 1454, 1657, 2160, 495, 1097, 1898, 1911, 1147, 1494, 2059, 2594, 2456]. Cantoni [697]. Cantor [283, 2737, 2861]. Canziani [181]. Cao [1156]. caos [1998]. Capasso [365, 366, 1181]. Capatti [1904]. Capel [494, 618]. Capelli [1181, 869]. Capitalism [2134, 3780, 3787]. Capitan [1729]. capitoli [1645]. capitolina [2111]. capitolo [2167, 2238, 2140]. capitum [1280]. Capo [3060]. Capocaccia [861]. Capodimonte [1082, 1555, 1082]. capolavori [958]. Cappelletti [299, 1952, 2023, 337]. Capra [1189]. Caprara [1815]. Capshew [2075]. Captain [1408]. Capucci [180]. Caraffe [1968]. Caramanica [2154]. Caramiello [2530, 2531]. caratteri [421]. Carazza [1162]. Carbon [2971]. Carbone [2013]. Cardan [834]. Cardano [2716]. Cardinal [964, 3374, 1123]. Cardinali [727]. Cardini [1754, 1064]. Cardo [1594]. Cardo/Saggi [1594]. Cardot [260]. Care [2933, 2962, 169, 1812]. Career [842, 3501]. Careers [440, 2626]. Carey [551]. Caribbean [3444]. Caring [959]. Cariou [1578]. Carl [1602, 2527, 158, 123, 64, 226, 3905, 2939, 1079, 507, 3905, 3148]. Carl-Wilhelm [507]. Carla [83, 348, 1857]. Carli [1922]. Carlieri [788, 1098]. Carlino [1644]. Carlo [41, 896, 1950, 2021, 2332, 2436, 1145, 2631, 2592, 1759, 2156, 2531, 3412, 3426, 561, 1284, 1355, 1451, 1532, 1794, 1994, 3613, 1651, 1902, 493, 2612, 2530, 1995, 2476, 1895, 2693, 1794, 3140, 2049, 2726, 2416]. Carlos [648, 1860]. Carlotta [2712, 3741]. Carlssons [1023, 2119]. Carmela [2336, 1917, 2287, 1917, 2618]. Carmélia [746]. Carmen [1525, 1666]. Carmichael [194]. Carneiro [2895]. carnet [3461]. carnets [2902]. Carnevale [1186]. Carocci [3541, 3545, 2631, 2969, 3544, 2672, 2651, 2624, 2206, 3547, 2269, 3095, 2501, 2416]. Carol [2707]. Caroli [1394]. Carolina [2964, 1218]. Caroline [3491, 3695, 3650, 3081, 3216]. Carolingian [2999]. Carolus [2884]. Carothers [1814]. Caroti [949]. Carpanetto [2100]. Carpenter [1274]. Carrand [777]. Carrara [887, 3409, 2006, 711, 1768, 3969]. carrés [2519]. Carriole [3645]. Carsten [2917]. Carte [2176, 2614, 2638, 1307, 2129, 1218, 555, 1135, 2314, 305]. Carteggi [171, 799]. Carteggio [1974, 53, 904, 2400, 2398, 2058, 1706, 2449, 2590, 664, 171, 529, 2399, 1783, 1689]. Cartelle [678]. Carter [454]. cartes [754]. Cartesianesimo [1142]. Cartesianism [3508]. Cartesio [755, 1506, 2432]. Cartografia [636, 236, 297, 357, 1917, 238, 858, 3933]. cartografiche [237]. Cartografo [1823, 688, 590]. cartographer [2356]. Cartographies [3385]. Cartography [3600, 3325, 3933]. Carvajal [3214]. Carvallo [2578]. cas [3765, 1878]. Casa [1691, 1817, 1613, 1581, 1580, 947, 948, 1015, 2386]. Casagrande [1156]. Casagrande-Mazzoli [1156]. Casale [1287]. Casalena [2969]. Casalzuigno [1331]. Casamassima [705, 878]. Casanatense [1314]. Casati [2330, 2281]. Case [1778, 1702, 1707, 2303, 79, 3649, 3219, 1294, 54, 3504, 3765, 1721, 1219]. Cases [3649]. Casi [1259, 1437, 2112]. Casino [3590]. Caso [1778, 2473, 1219, 1559, 2487, 89, 2543, 1242, 1148, 2918, 2208]. Casper [1492]. Cassa [863, 917, 858, 1614]. Cassata [2738]. Cassiano [2719, 3345, 3412, 921]. Cassini [1453, 2650, 1705]. Cassino [1230]. Castagnola [752]. Castagnoli [1902]. Castaldo [237]. Castelia [965]. Castell [134]. Castellani [1235, 3140]. Castelli [2525, 2507]. Castello [2523]. Castelnuovo [1658]. Castenaso [1015]. Castile [648]. Catalá [1544]. Catalans [353, 3019]. Catalog [4067, 293, 2265, 1331]. catalogación [2068, 3990]. Catalogazione [1054, 238, 556, 1259, 1301]. Cataloghi [1220, 1975, 1740, 1924, 1661, 1804]. cataloging [2068, 3990]. Catalogo [1696, 1279, 1156, 1360, 809, 2614, 2773, 1741, 867, 1176, 71, 844, 1259, 1794, 83, 1283, 4067, 1167, 1252, 1331, 1530, 221, 70, 929, 1218]. Catalogodei [720]. Catalogodi [991]. Catalogue [1055, 1078, 1805, 2338, 696, 985, 856, 1094, 2555, 1167, 1429, 1662, 1792, 2724, 2757, 1933, 2683, 2369, 4091, 3823, 2554, 2556, 3719]. Catalogues [1448, 2204, 3913]. Catalogus [1687]. Catalonia [3452]. Catania [1648, 1690]. Catanzaro [3089]. Catapano [79, 1522]. catastrofe [39]. Catastrofismo [744]. Catastrophe [3482]. Catastrophes [2103]. Catastrophic [3602]. Catedra [43]. Categorie [1606]. Categories [2518]. Caterina [3412, 3367, 2513]. caterpillars [1185]. Cathedrals [2096]. Catherine [796, 2197, 1249, 1842]. Catheters [3149]. Catholic [58, 3455, 2585]. Catholique [1442]. Cattaneo [2332]. Cattaneo/Giovanni [2332]. cattedre [2468]. cattolici [3426]. Cauchy [40, 400, 40]. cauldron [1753]. Causa [2272, 3798]. Causality [1857, 4057]. cause [3798]. causes [337]. Causis [299, 2571, 337, 945]. Cavalieri [2328]. Cavalli [1856, 3092]. Cavalli-Sforza [1856]. Cavazza [3689]. Cavendish [1819, 2717, 2606, 3892]. Cavicchi [3694]. Cavre [1955, 2026]. Cay [1936]. Cay-Ruediger [1936]. Cazzola [3710]. ccxcix [799]. ccxi [2528]. CD [3286, 2683, 2205]. CD-ROM [3286, 2683, 2205]. CdG [2156]. Ceccarelli [811]. Cecchi [3693, 390, 665]. Cecilia [2934, 2520, 1592]. Ceglia [2144, 2921, 3397]. Celebrating [3466, 311]. celebrative [709]. celebrativo [311]. celebrazioni [2581]. Celebre [740, 2578]. celeste [2934, 1160]. Celestial [3450, 2636, 184, 3325, 233]. Cell [3226, 4062]. Celli [1185]. Cellio [2293]. Cells [3035, 3903]. cellulare [2233, 2162]. Cem [2528]. Cena [1375]. Cenedella [1675]. Censimento [1599, 309]. Censorship [3651]. Censura [889, 3135, 1291]. Census [2259]. cent [2214]. Centaur [3048]. Centenario [299, 709, 311, 448, 572, 2581, 337]. centenary [3762, 4045, 311, 2213, 337]. Center [1837, 840, 1264]. centième [4089]. Centina [3286, 2614]. Cento [2202, 1768, 1061, 3085, 3969]. Central [1110, 3433, 1350, 387, 1783, 1546]. Centrale [2230, 844, 1062, 1350, 1666, 1546, 1783]. centrali [1951, 2022]. Centre [32, 444, 1447, 1612, 1859, 540, 893, 1799, 1734, 329, 1224, 540]. Centres [3170]. Centro [995, 3120, 1648, 3040, 2685, 448, 1360, 2010, 1438, 14, 20, 859, 1259, 1619, 1746, 1806, 1813, 2005, 2078, 999, 1645, 1737, 1424, 2217]. centuria [1098]. Centuries [703, 3469, 1985, 2888, 26, 2802, 3763, 3885, 1936, 544, 2776, 3759, 2350, 4049, 2161, 4072, 4004, 4006]. Century [517, 891, 1110, 1264, 1457, 1404, 1490, 1589, 1725, 1818, 2056, 2228, 2312, 2496, 3739, 4041, 2388, 675, 3535, 2737, 1060, 1439, 3081, 2295, 2906, 3373, 180, 1686, 279, 3637, 2089, 1067, 1670, 2216, 3324, 2962, 40, 2347, 3104, 3511, 1770, 2781, 3656, 3033, 2900, 418, 419, 3236, 3407, 1652, 425, 778, 1193, 1320, 1511, 1860, 2423, 3557, 545, 3600, 3690, 3114, 1618, 2099, 2998, 1039, 3617, 3654, 2911, 378, 3669, 3393, 2901, 1331, 1213, 3609, 1294, 3565, 2913, 3291, 3479, 573, 2464, 2551, 2356, 3205, 3060, 3024, 3919, 938, 1429, 3944, 3448, 3965, 1845]. Century [77, 335, 685, 3744, 2462, 3562, 3471, 1387, 3783, 3149, 63, 3472, 394, 3502, 2105, 2542, 3803, 3431, 3217, 3234, 3198, 4062, 1121, 659, 3841, 2988, 3849, 2470, 3869, 3079, 3879, 3894, 2765, 3900, 3902, 2825, 3523, 3105, 3196, 2961, 3385, 3949, 3488, 3966, 3513, 3378, 735, 3554, 3343, 3474, 2443, 734, 2294, 1378, 2130, 2127, 4034, 3729, 4017, 830, 3714, 3616, 3346]. century-long [3919]. Ceolin [1195, 1532]. C’era [2433]. Ceramica [794]. Ceranski [1726]. Cerchio [1181]. cere [956]. Cerebral [3619, 3375, 3446, 3933]. Cerebrali [1467, 1917, 3933]. Ceregato [2882]. Ceriana [1162]. Ceriana-Mayneri [1162]. Cerletti [2855, 3039, 3505]. Cerletti-Bini [3505]. Cernusco [953]. Ceroplastics [3489]. Cerruti [41, 764, 2504]. certezze [1998]. Cerulus [2725]. cerusici [1598]. Ceruti [371]. cerveau. [1460]. Cervellara [43]. cervelli [3426]. cervello [1965, 2036, 2334, 1917, 2220, 3933, 570, 1052, 1414, 1506, 2618, 2499, 2285, 2336, 2287]. Cesa [338]. Cesare [496, 2491, 2492, 1107, 1813, 3967, 3255, 3802, 2592, 1791, 3061, 1645, 2426, 3085]. Cesariano [1645]. Cesàro [982]. Cesena [70]. Cesi [311, 2941, 3345, 2713]. Ch [378]. Châtelet [1681, 3894, 3331]. Chadwick [85]. Chair [3740]. Challenge [837, 2812]. Challenger [1076]. Cham [3525]. Chamber [3509, 425]. Chamberlin [1654, 3790]. Chambers [964, 2956]. Chambéry [4042, 1327]. Chambre [264]. Chambry [502]. Champ [1080]. Champion [795, 1985, 2829, 3010, 2729]. Chance [2106, 2594, 3830]. Chandelier [3736]. Chang [2681]. Change [1490, 2607, 928, 3210, 2971, 692]. Changed [3686, 2551]. Changes [728]. Changing [2906, 1274, 3215, 2066, 3832, 3981]. Chaoui [2544]. Chapel [3722]. Chapman [1664]. Chapter [3408, 3610, 184]. Chapters [2095, 25, 1489]. characteristics [421]. characters [3664]. Charentes [1609]. Charisma [2850]. Charitable [1652]. charity [1596, 1192]. Charlatanes [3959]. Charlatanes. [3264]. Charlatanism [3066]. Charleroi [855]. Charles [2717, 2538, 2805, 194, 830, 1102, 1358, 3076, 2873, 2755, 385, 1407, 1875, 2004, 171, 190, 851, 1408, 3571, 2538, 3072, 4056, 2849, 1278, 3814, 3816, 3423, 2015, 2761, 893, 1700, 2998, 1571, 2919, 2982, 2044, 413]. Charles-Daubert [1875]. Charlotte [174]. chartes [4043]. Charting [3806, 1844]. Chassagne [2894]. Chatto [1945, 2781, 2956]. Chaucer [3449, 3276]. Che [1052, 3873, 2430, 3693, 3426, 4010, 2503, 3373]. cheat [1112, 3831]. Cheesemaking [3474]. Chemical [242, 405, 488, 837, 885, 2425, 2730, 2897, 1387, 1495, 2891, 3195, 1500, 3173, 2828, 1484, 3635, 36, 2486, 2724, 1290, 1814, 3760, 1381, 2051, 3982]. Chemicus [2517]. Chemie [961, 497, 543, 1151, 1069]. Chemielehrbcher [623]. Chemiker [508, 1336]. Chemikern [508]. cheminement [2973]. chemins [1418]. Chemischen [1381]. Chemist [16, 1991, 4075]. Chemistical [734]. Chemistry [11, 15, 301, 923, 1005, 1333, 1846, 3347, 3739, 2338, 2350, 2462, 1387, 1627, 2262, 3160, 2864, 3770, 1991, 1500, 1147, 2161, 2579, 3821, 3511, 928, 1484, 3243, 3690, 3215, 2591, 2443, 2044, 1290, 1294, 1199, 3784, 3881, 4072, 4075, 2198, 1798, 2496, 2891, 2917, 2907, 3091, 3243, 714, 1494]. Chemists [885, 977, 3236]. Chemnitz [1608]. Chercheur [1700]. Chernin [1203]. Chernoglazov [2822]. Cherubini [1099, 644, 712, 810]. Chester [609]. cheval [1187]. chevaux [1187]. chez [198, 3448]. Chiantera [1029]. Chiara [446, 1273, 2230, 2420]. Chiarugi [901, 788, 1098]. Chiavari [2271]. chiave [1137]. Chicago [1106, 1313, 1404, 2194, 2424, 2560, 2776, 2887, 1573, 1659, 2669, 3068, 2805, 2093, 2850, 2551, 538, 2103, 2371, 2834, 3043, 3111, 2655, 3070, 2800, 2747, 1398, 1722, 2783, 2718, 2862, 3033, 927, 3542, 2775, 1504, 2858, 2976, 2676, 3118, 2841, 2559, 2985, 2807, 1400, 1723, 2845, 3126, 1541, 993, 3032, 3261, 2739, 2772, 2562, 2688, 3323]. Chicago-London [1504]. chickens [1897]. Chiesa [573, 1790]. Chiesura [2849]. Chieti [365, 366, 1033]. Childhood [1477]. Children [3100, 1192]. Chile [3451]. Chilly [2015, 2013]. Chilly-Mazarin [2015, 2013]. Chimeras [35]. Chimica [506, 793, 794, 764, 1091, 1031, 461, 646, 1679, 2223, 960, 1813, 1601, 734, 1868, 2044, 41, 569, 2504, 1730, 3982, 1171, 2425, 831]. chimiche [3299, 3760]. chimici [2208, 2603]. Chimico [3184]. Chimie [1406, 2145, 620, 1090, 2898, 3784, 3665]. chimique [1090, 834, 2051]. Chimiques [1888]. Chimpanzee [3922, 2382]. China [1261, 3205, 2518, 928, 1315, 1484, 1641, 1982, 2629, 1640, 163, 690, 3566, 3780, 3787, 2134, 406]. Chine [748, 2048, 2566, 410]. Chine-Occident [748]. Chinese [1683, 3205, 3599, 3649, 606, 845, 2493, 3535, 3, 3664, 4037]. Chinnici [1350, 1750, 1863]. Chinoise [2048, 2566]. Chinoises [536, 4037, 2197]. Chioggia [1236]. Chiricò [2428]. Chirio [1691, 2467]. Chirurgia [2620, 947, 1176, 3426, 1728]. chirurgiche [2468]. chirurgico [1599]. Chirurgie [2682, 1187, 3857]. chisel [3973]. Chitarrini [1354]. Chitnis [341]. Chiusa [2487]. choisies [866]. choisis [796, 3992]. Chopin [1571]. Chordate [3610]. choses [2821]. Christ [3387]. Christa [2601, 698, 1819]. Christiaan [1323]. Christian [2418, 151, 103, 153, 112, 1188, 1357, 1686, 1229, 2967, 565, 296, 101, 144, 99, 443]. Christiane [201]. Christianidis [2708]. Christianity [573, 2966]. Christianorum [1280]. Christianson [888]. Christie [283, 683, 489]. Christina [891]. Christine [2273, 2915, 3083]. Christoph [1715, 1398, 3253, 2915, 276, 105, 141, 147, 115, 966, 547, 1735, 2196, 1150]. Christopher [414, 1110, 899, 2093, 1836, 1845]. Christos [3572]. Chronicle [1781, 1560, 2131]. Chronische [1391]. Chronology [1822]. Chrysa [2142]. Chrysopoeia [3277]. Chu [2137, 3897]. Chunder [2485]. Church [573, 2774, 3455, 2096]. Churchland [3093]. chymischen [1531]. Chymistry [3598, 2891, 2718]. Chymists [2891]. Ci [747]. Ciamarra [1767]. Ciampoli [3396]. Ciancio [3209, 2796]. Ciappi [1695]. Ciardi [1949, 2020, 3544, 3022, 3029, 1614, 2603, 3416]. ciascuno [2370]. Ciba [332, 503, 434, 436, 639, 640, 785, 1027, 1026, 1177, 1175, 1271, 1434, 1810, 500, 564, 708, 870, 875]. Ciba-Geigy [332, 503, 434, 436, 639, 640, 785, 1027, 1026, 500, 564, 708, 875]. Cic [1098]. cicala [2630, 3956]. Ciccarelli [1156]. Cicle [349]. Cicogna [2257]. Cidado [2895]. Ciel [2638, 1567, 1447, 410]. cieli [2212, 3457]. Cielo [2178, 2942, 1059, 186]. cielos [35]. Ciencia [889, 333, 430, 1855, 3019, 3759, 43, 3427, 3327, 2114, 2097, 2198, 3019, 889, 2494, 1544]. Ciencias [352, 430, 2068, 3990]. Cientifica [2895, 2101, 494, 618]. Cientificas [2494, 889, 1226, 494, 618, 1227]. Cientifico [963, 3990]. Científicos [1849, 1860, 4017, 889]. Cierre [1597]. Cieszkowski [926]. Ciliberto [578]. cilindre [3167]. Cimatti [2323]. Cimento [2338, 1572, 2815]. Cimino [2224, 504]. Cina [832, 1873, 186, 2307]. Cinato [1899]. Cinema [3152]. cinematografiche [185]. cinese [1059]. Cinisello [877, 882]. Cinq [1681]. Cinque [2226, 2307]. cinquecentina [1817]. Cinquecento [515, 278, 2490, 1597, 1029, 922, 2266, 3015, 263, 2140, 2951, 2345, 4038]. Cinzia [1062, 3315]. Cinzio [1236, 2602]. Cioppi [1430]. Cipriani [2797, 3578]. Circa [3100]. Circle [891, 489]. Circles [3340]. Circolazione [3021, 1539]. Circulating [3322, 3622]. Circulation [3253, 3912, 3148, 3835, 3695]. Circumsolar [3133]. Cirene [2574]. Cirillo [3297]. Cirms [1619]. Cis [2498]. Cisalpino [1145, 1869, 1712]. Cisalpino-Istituto [1869]. Ciso [1183, 1180, 1025]. Cité [653, 2229, 701, 2188]. citizen [1329, 4029]. citizen-patient [1329, 4029]. Citt [2314, 42, 1755]. Citta [2523, 1990, 2743, 1597, 1756, 562, 1016, 2115, 1680, 1790, 1279, 1268, 281]. Cittadella [2458]. City [549, 3675, 613, 68, 3479, 3867, 1849, 3597, 3572]. ciudad [3867, 1849]. Ciuffoletti [804]. Civica [42]. Civici [2007]. Civico [2619, 431, 518, 1663, 1530]. civile [180]. Civilisation [444, 1315, 1484, 1641, 1982, 2629, 331]. civilisations [1985, 748]. Civility [1404]. Civilization [3940, 2561, 2937]. Civilizations [2927]. Civilt [1101, 640, 2307, 186]. Civiltà [2810, 3541, 331, 1373, 552]. civitas [446]. Cl [1326]. Clagett [829]. clair [834]. Clara [3497, 556, 2883, 2905]. Clarendon [1087, 192, 2706, 925, 1596, 2609, 1007, 2687, 1844, 539, 946]. Clark [2850, 285, 1934]. Clarke [761]. Class [587]. Classe [2469]. classi [2612]. classica [41, 1134]. Classical [2710, 2879, 945, 606]. Classici [1364, 1606, 1064, 1816, 570]. Classico [1789, 1258, 1241, 706]. Classics [892, 2708]. classification [763]. Classificazione [396]. classique [2648, 3004, 3173]. classiques [866]. Claude [496, 416, 33, 3365, 1228, 2316, 2776, 2843, 416, 1793, 2735, 2770, 512, 1984, 536, 3904, 921, 1135]. Claude-Ren [496]. Claudia [2657, 2616, 2428, 364]. Claudine [300]. Claudio [1967, 2038, 1948, 2019, 1620, 1190, 3426, 3619, 338, 705, 878, 2677]. Claudius [3905]. Claus [3213]. Clave [3665]. Clavio [2172]. Clavis [3064, 3283]. Clavius [1398]. Cleempoel [2369]. Clelia [515]. Clemente [1029]. Cleomedes [2516]. Clercc [63]. Clercq [2195]. clericale [3426]. Clericuzio [2462, 2651]. Clerk [4068, 849]. Clermont [921]. Clermont-Ferrand [921]. Clerq [1792]. Clessidre [3362]. Clifford [347]. Clima [1088, 170]. climat [305]. Climate [2971]. climatiques [305]. Clinica [1467]. Clinical [3220, 1179, 3895, 955]. Clinicals’ [1877, 3928]. clinique [3895, 2375]. Clio [3241]. Clock [2442, 1309, 1215, 428]. Clocks [3370, 1212, 406, 425]. Clockwork [3604, 281, 406]. Clonage [2704]. Cloth [64, 1841, 2312, 1987, 422]. clothbound [1494]. Clouds [1047]. Club [1314, 3208, 1241]. Clueb [1502, 59, 1417, 641, 1094, 2468, 780, 1411, 1746, 1923, 2070, 2213, 969, 678, 2460, 1325, 2322]. Clues [3500]. ClupGuide [2150]. Clusius [2884, 2509]. Cluyt [2509]. Clviii [1599]. clxxxi [3027]. CNRS [1848]. Co [172, 833, 1408, 2205, 54]. Co-Operation [54]. Coast [1617, 2363]. Cocchi [1959, 2030, 2373, 2398, 3872]. Cocheton [638]. Cocoon [1331]. Code [795, 645]. Code-Dd [645]. Codex [2881, 1323, 816, 3629]. codice [2881]. Codici [2257, 3172, 949, 1011]. Codicological [3440]. codicum [1687]. Coelum [51]. coërcendus [3433]. Coercion [3714]. cognitiva [373, 2431, 2618]. cognizione [304]. Cognizioni [1785]. Cogrossi [2726]. Cohen [1313, 2967]. Cohn [143]. Coignet [2291]. Cojannot [2826]. Cojannot-Le [2826]. colaboración [2571]. Colacchi [2667]. Colapietra [2667]. Colavita [3315]. Cole [405, 1872, 3745, 3551].

Colecao [430]. Colección [1544]. Coleridge [2609]. Colin [1315, 1798]. Coll [754]. collaboration [4043, 3280, 2309, 2646]. collaborative [3540]. collaborazione [1957, 2028, 1950, 2021, 2320, 3718, 1998, 83]. Collana [1502, 588, 85, 1243, 1499, 2224, 2614, 2202, 1675, 1672, 1677, 1676, 871, 1033, 1452, 1768, 680, 1988, 1064, 1233, 1318, 1802, 1803]. Collane [1803]. collapse [1398]. Collecio [2593]. Collected [802, 1367, 2852, 886, 1278, 1443, 1669, 2098, 1019]. Collecting [3713, 1319, 3737, 2956, 3918]. Collection [1055, 2145, 2267, 3164, 675, 3276, 1699, 2215, 40, 2719, 314, 866, 3069, 713, 1849, 761, 998, 2069, 2702, 753, 749, 1413, 3704, 2201, 2422, 1135, 3856, 1584, 3948, 4007, 840, 699, 266, 2555, 2147]. Collections [2436, 3641, 3659, 3677, 3697, 3720, 3743, 3184, 1448, 2698, 3217, 2568, 3503, 3695, 2961, 2683, 293, 1080, 1510, 3338, 2204, 3297, 3977, 1928, 2757, 3913, 992, 1270]. Collective [3737]. Collector [1314, 2185, 593]. Collectors [2180]. College [767, 2050, 1299, 2867, 1464, 1464, 2995]. Colleges [2804]. Collegio [2452, 740]. Collegium [1256]. Collezione [777, 828, 2298, 2871, 1561, 864, 1015, 1343, 1491, 75, 2007, 3856, 354, 780, 1666, 1931, 2111, 2149, 2414, 3948, 4007]. Collezioni [1780, 1163, 1510, 2002, 83, 1270, 493, 2619, 1928, 1512, 2622, 486, 2110]. collezionismo [3918, 2226, 1906]. Colliding [693]. Collina [1082]. Collini [2798]. Collins [651, 1571, 2739]. Colloque [1090, 4047, 1440, 32, 243, 1442, 868, 1326, 2878, 4043, 260, 4053, 1984, 314, 954, 713, 4077, 540, 2118, 3952, 1734, 4089, 2201, 1224]. Colloques [728]. Colloquio [2809, 1388, 1488]. Colloquium [2933]. Colnort [795]. Colnort-Bodet [795]. Cologne [3082]. Colomba [1903]. Colombaria [516, 716, 887, 1077]. Colombo [1166, 3364]. Colonial [2964, 3316, 3580, 3993, 2657]. Colonies [3692]. Colonisation [1334, 3887]. Colonna [2565, 3652]. colonne [3209]. color [3528, 1053]. Colorado [1763, 2505]. colorata [203]. colore [2490]. colored [1199]. colori [1566, 2371, 1335, 407, 1796, 840]. Colorless [68]. Colour [2761, 2404, 3526, 3530]. Colouring [3296]. Coltello [970]. colture [3873]. Columbia [1108, 2678]. Columbian [862]. Columbus [899]. columns [57]. combinatus [1610]. Combridge [406]. Come [2323, 1112, 3804, 3831, 3693, 1020]. Comet [3280, 4063]. cometa [3496]. Cometary [3373]. comete [2167, 2238]. Comets [1835, 2636]. Cominale [1142]. coming [1510]. Comino [171]. Comit [2894]. Comite [584, 1060, 1155, 368, 1795, 834]. Command [1408]. Commandino [49]. Commedie [3825, 2370]. commémoratif [4089]. commémoratifs [4039]. Commemorating [2724]. commemoration [452]. commemorative [4089]. Commemorazione [402]. commencant [1444]. Comment [1041, 212]. commentaire [574, 1138, 1432]. Commentari [393]. Commentaries [1102, 1688, 274]. commentarii [1958, 2029]. commentario [3277, 3964]. Commentarius [4023, 2205]. Commentary [167, 1288, 2516, 535, 791, 1611, 3298, 2525, 749, 3964, 1133, 2095]. commenté [754]. commented [1019]. commentées [921]. Commenti [2520]. Commento [764, 223, 3277, 1483, 2644]. comments [2586]. Commerce [2657, 3714, 2944, 914, 3811]. Commercial [3440, 2833]. commerciale [2596]. Commercio [903, 1539, 638, 1029, 1099, 712, 710, 711, 871, 863]. Commissie [802]. Commission [669, 823, 907, 1478, 1556, 1635, 2179, 2299, 55, 271, 326]. Committee [802, 2495]. Commodities [3715]. Common [549, 2751, 4028]. Commonsense [3232]. commoratifs [2067]. communicate [2678]. Communicating [4018, 3328]. Communication [3471, 824, 2867]. Communications [3658]. communissimi [4028]. Communities [3637]. Community [2425, 2406, 3595, 1576, 3982]. Como [1813, 2005, 2217]. Compagni [1229]. Companion [1649, 2889, 2936, 3063, 2095]. Company [16, 1346, 1853, 28, 3281, 624, 3011, 3968]. comparata [537]. Comparative [1541]. comparatives [1224]. Comparetti [171]. Comparison [3225, 214]. Comparisons [405]. Compass [2346, 3290, 3258, 3953, 3204]. Compassi [1214]. Compasso [2163, 2234, 941]. Compendia [2510]. Compendy [2596]. compiler [1805]. compilor [557]. complessa [2581]. complessive [931]. Complete [1880]. Complexity [3460]. comportamento [2276]. comportements [3010]. Composite [645]. Compositori [1902, 2882, 1745, 2225]. Composizione [1774]. Compostela [2068, 3990]. Compotus [1278]. Comprensione [2264]. Compromise [2087]. computer [3757, 3060, 3919, 355, 2919]. Computerized [990]. Computerizzata [602]. Computers [694]. Comte [1681, 257, 2444]. Comunale [1015]. Comune [989, 1979, 953, 859, 1453, 560, 858]. comunicazione [3496, 2903, 2692]. comunit [627]. Comunita [2413, 2425, 3982]. Con [2667, 1077, 563, 2117, 1957, 2028, 1955, 2026, 1950, 2021, 2163, 2234, 2387, 2389, 2386, 2530, 1226, 2320, 3788, 1869, 1538, 3844, 2614, 67, 2590, 1098, 171, 3277, 1895, 2047, 2255, 1926, 753, 749, 1323, 1394, 1737, 1998, 2132, 3964, 83, 2501, 2141, 2571]. Concave [2084]. Concept [204, 26, 2732, 1847, 282, 1184, 1976, 3975]. Conceptions [3477, 214, 1086]. Concepts [1702, 2457, 1983]. Conceptual [1541, 1723]. concernant [2757, 2578]. Concerning [694]. Concetta [3329, 2608]. Concetti [1585, 3095]. Concetto [2445, 660, 2645, 687, 1389]. Concezione [1976, 2142]. Concezioni [214]. Conchology [3647]. Concilio [1291]. Conclusions [2629]. Concordancia [2571]. Concorsi [2092]. condanna [719]. condemnation [719]. Condillac [198]. condizione [809]. conducting [2156, 4022]. conduits [2330]. Conduttori [1824, 2330, 2281]. Confcooperative [1373]. Confederationis [1687]. Conference [311, 4067, 2050, 4082, 753, 4040, 1440, 4077, 4089, 824]. Confidence [3340]. Conflict [1193]. Conforti [3672]. Confronto [214, 3161]. confrontonel [2507]. Confusion [524]. Confutazioni [1606]. Congedo [2580, 2492, 1142]. congiura [2535]. Congrés [243, 4042, 836, 1442, 1327]. Congress [741, 4059, 4071, 4087, 4053]. Congresses [1879]. congressi [2969]. Congresso [741, 4059, 3031, 1813, 2005, 2217, 1710, 809]. Coniche [1973]. Conics [3684]. coniques [2199, 3163]. Conjecture [3229]. conjecturing [2786]. conjunction [10]. Connaissance [198]. connaissances [1232]. connecting [3994, 3995, 2120]. Connection [899, 3449, 1193, 2195, 4058]. Connoisseurs [2815]. conocimiento [2494]. Conoscenza [81, 1619, 934, 310, 996, 371]. conoscenze [2536]. conoscere [3388]. conquêtes [3742]. conquiste [792]. Conrad [101, 114]. Conry [2318, 987]. Conscience [3424]. Consciousness [592, 2120, 890, 3994, 3995]. Conseils [1879]. Consejo [889, 1227, 1226, 2494, 494, 618]. Conselve [2602, 2787]. Consentina [721]. Conservate [2248, 2176, 1218, 1783]. conservateurs [954, 713]. conservati [1252]. conservatinella [44]. conservation [3818]. Conservazione [3103, 3818, 238, 2322]. Conservés [2048, 2566]. Conserving [2664]. considérablement [2767]. Consideration [1724]. consideratione [567]. Considerazioni [1826, 1558, 814, 522]. Consigli [1648]. Consiglio [999, 2274, 3986, 3987, 3988]. Consilia [1273]. consolazione [85]. conspiracy [1798]. Constant [207]. Constantine [763, 1358, 749]. Constantino [1849]. Constantinople [3205, 2760]. Constitution [1667, 2214, 2296, 3930, 3930]. constitutione [2337, 2288]. Construcción [352]. Constructing [3637, 769, 3576]. Construction [1377, 3485, 2199, 184]. constructivas [2876]. Consultation [2906, 1980]. Consultations [955]. Consultatoria [1724, 1724]. Consultazioni [1711]. Consulti [3027]. Consulting [1711]. Consumer [3616]. Consumption [3559]. Contact [2156, 3656, 4022]. Contagi [708]. Contagion [3656]. contagioso [2714, 2726]. Contagium [660]. Contardi [1959, 2030, 2386]. Contatti [2172]. Conte [1658, 2138]. contemporaine [1413]. Contemporaines [1944, 2067]. contemporains [444, 3448]. contemporanea [1871, 3937, 2107, 555, 1615, 2221, 2429, 1348]. contemporaneo [1810]. contemporaries [3448, 362]. Contemporary [3262, 3694, 1738, 3150, 2540, 2907, 2610, 2788, 2847, 2666, 2733, 3028, 3119, 2925, 3669, 3207, 1498, 3242, 1608, 1348]. Contents [3193, 3227, 3269]. Contenuti [3174]. contesa [812]. contest [2783]. Contested [3608]. Context [3437, 3559, 1494, 2658, 1628, 850, 3267, 1683, 3906, 993]. Contexte [2704]. contexts [2745, 3122]. Conti [717, 941]. Continental [348, 3429]. Continenti [744, 656]. Continents [692, 693, 656]. continentur [1687]. contingency [1659]. continu [3540]. Continuatio [1280]. Continuism [1560]. Continuismo [1560]. Continuità [595, 898]. Continuity [595]. continuous [3540]. conto [711]. contraddizione [3457]. contre [2214]. contre-réforme [2214]. Contributi [448, 1794, 372, 2905, 407, 832]. Contribution [305, 3010, 734, 2400, 2767]. Contributions [2193, 2161, 3489, 888, 2528, 2873, 4072, 492]. Contributo [1974, 915, 904, 2400, 1978, 3136, 734, 1926, 2271]. Contributors [2759, 3098, 3155, 3192]. Contro [1829, 1001, 2349, 2247, 1053, 2942, 2132, 3367]. Control [2495, 1126]. controverse [2578]. Controversia [744, 656, 1068]. Controversies [2470]. Controversy [2512, 2732, 3908, 656, 1798, 2781]. Convegni [311]. Convegno [197, 514, 721, 1091, 2522, 215, 2376, 404, 1637, 1243, 2949, 4067, 561, 2113, 827, 1310, 2692, 2389, 915, 572, 2492, 2880, 838, 3967, 1137, 1270, 831, 4040, 311]. Convention [3848]. Convenzione [3848, 1850]. Conversazioni [2384]. Convolutions [1149]. Cook [614, 3599, 2944, 1729, 1844]. Cooper [2827]. Cooperation [2294, 3893, 1912]. Cooperativa [2140, 2113]. Cooperazione [2113]. Coordinamento [859, 2416]. Cope [959]. Copeland [3417]. Copenhagen [1081, 2418, 1659, 235, 2658, 1008, 1363]. Copernic [2201, 4043]. Copernica [2576]. Copernican [35, 2248, 1119]. copernicana [1136, 35]. Copernicane [1119]. Copernicanesimo [1118, 1790]. Copernicanism [1118]. Copernicano [486, 2248, 937, 2110]. Copernico [1136, 2460]. Copernicus [1398, 2799, 2576, 751, 1450, 4043, 3139, 2799, 2812, 2992]. Copernicus-Biographien [2576]. copie [858]. copies [38, 3225]. Coping [1812]. Copley [3468]. copper [1422]. Coppini [2623, 1807]. copti [2644]. copto [2644]. Copy [3225, 49]. Coral [3309]. Corbellini [1925, 2624]. Corbin [807]. Corbo [1189, 1136]. cordis [2205, 4023]. core [3789, 3795, 1654]. Cormack [3441]. Corn [3605]. Cornelis [3244]. Cornelius [2736, 1229]. Cornell [167, 2589, 614, 177]. Corners [3575]. Cornet [1224]. Cornish [3695]. Cornwall [1660]. Corona [965]. Coronelli [2077, 1209, 1779, 2046]. Corpi [3672, 3178, 2836, 1056]. Corpo [1774, 3801, 2754, 2691, 1318, 1644]. corporis [1962, 2033]. Corps [1727, 2952, 175]. Corpus [2635, 307, 720, 717, 991, 1279, 1280, 1156, 1688, 70, 1232, 1133, 1174, 1131, 1391, 3650, 610, 2644]. Corpuscles [60, 2462, 1891]. Corpuscular [3647]. Corrado [1648, 1648]. Correia [2895]. Correios [1743]. correnti [2742, 2525]. Correr [2257]. Correspondance [1795, 368, 3875, 2539, 3785, 4065, 3284, 3368]. Correspondants [2539]. Correspondence [2003, 2717, 4, 1006, 1495, 3475, 2852, 3030, 1238, 2080, 1732, 1770, 1407, 2004, 982, 851, 1783, 4032, 3340, 3785, 4056, 2400, 3809, 3814, 3815, 3816, 3836, 4065, 4073, 4074, 1350, 1671, 808, 3335, 2834]. corretta [1790]. correzioni [2726]. Corriere [867]. Corrispondenza [1894, 4, 737, 269, 821, 1383, 1471, 2090, 1124, 1828, 323, 463, 2447, 600, 1302, 1555, 1782, 1973, 663, 1895, 2047, 2129, 2255, 2353, 2397, 3088, 982, 397, 2249, 95, 1633, 2914, 1783, 528, 2252, 1516, 2590, 71, 1350, 1671, 3335]. corruptione [1359]. Corsi [3030, 758]. Corsini [2448, 747, 1974, 402]. Corso [162, 764, 88, 1615]. Corte [1860, 2049, 2420, 2012, 2522, 3777, 2058, 1874, 4017]. Cortes [1645]. Corti [970, 1343]. Cortina [2132, 2324]. Coruña [2068]. Corvallis [3254]. cosa [1052]. cose [894, 1322, 2002, 3388]. Cosentina [838]. Cosenza [721, 991, 838, 1003, 831]. Cosimo [1212, 2461, 2590, 1906]. Cosmacini [360, 437, 1675, 1716, 1752, 1672, 1865, 929, 2668]. cosmesi [869, 3426]. Cosmetic [3857]. Cosmic [3477, 3555, 2123, 812]. cosmiche [812]. cosmique [3004]. Cosmo [2123]. Cosmografo [2356]. Cosmographia [3017]. cosmologia [3925, 1713]. cosmologie [2820, 3062]. Cosmology [1412, 1835, 511, 2999, 1193, 1398, 3925]. Cosmopolitanism [3414]. Cosmos [281, 3004, 1698, 4036]. Costa [2077, 199]. Costantini [1857].

Costantinopoli [2535]. Coste [3010, 2363]. costituzione [1417]. costruito [2974]. costruttivismo [372]. Costruttori [2045]. costruzione [934, 1414]. Costumi [2009]. Cotes [2683]. Cottingham [196, 755]. Cotugno [2164, 2235, 223, 1990, 53]. Coulston [2538, 2982]. Council [3986, 3987, 3988]. Count [1393]. Counter [2482, 419]. counter-themes [419]. counters [1654]. Countries [14, 3501, 3835, 3253, 3912]. Countway [955]. Coupain [3332]. couple [2052]. Cour [1040]. couronnés [1681]. Cours [2995, 1306, 2898, 1442, 3665]. Course [3160]. courses [3767]. Court [1063, 3494, 1874, 1860, 2841, 2786, 4080, 3777, 1212, 3449, 2185, 4017]. Courtier [1400]. Coury [434]. Coutinho [3504]. Coveney [942]. Covetous [1401]. Cowie [3715]. Coyaud [2150]. Coyne [549, 1920]. cracoviensia [800, 3920]. Craft [3420, 3711]. Craig [3256]. Craik [3079]. Cram [885]. cranio [1244]. cranium [4008]. Crasta [1988]. Crater [3518]. Crawford [1657, 1245]. Creating [3444]. Creation [1449, 2793, 1891, 2638, 1573]. creationis [2138]. creations [1510]. creative [3805]. Creatività [958]. Creativity [15, 722]. créatrice [2318, 3805]. Creaturen [2715]. Creatures [3199, 1225, 3460]. creazioni [1510]. credente [2156]. Credenze [3206]. Credibility [3191]. Credit [2412, 4048]. Crelinsten [2851]. Cremante [180, 2769]. Cremona [70, 1516]. Crepin [32]. Crepuscoli [1050]. Créqui [1681]. Crescenzi [1512]. Crest [2294]. Cresti [561]. Crime [2907, 3770]. crimini [3206]. criminologia [3547]. Crimmins [925]. Crini [1324]. Crisci [1728]. Crisciani [2420, 446, 1273]. crisi [1540]. Crisis [1540]. Crisóstomo [3372]. Crispini [1003, 1233, 1850, 831]. Cristallino [2389]. Cristallo [2815, 1244]. cristiana [2720]. Cristianesimo [573]. Cristiani [2934]. Cristiano [3456]. Cristina [3742, 682, 1456, 1451, 2141, 2517, 2648, 2821, 1325]. Cristóbal [3489]. Cristoforo [1166]. Criteri [2451]. Criterio [2126]. Critica [2083, 1943, 3277, 1712, 969, 2132, 1483, 2205, 1399]. Critical [227, 488, 28, 2050, 2924, 997, 1249, 212, 1907, 10, 3805, 3298, 3863, 3946, 749]. Critiche [826, 327, 1130, 2455, 9, 161, 1385, 273, 531, 742, 1480, 1785]. critici [1581]. Criticism [443, 1143]. criticismo [1691]. critico [3863]. Critique [925, 2083, 198, 512]. critiques [512, 3805, 2318]. Crocologia. [3570]. croisée [1418, 2015]. croisements [2015]. Croisier [1626]. Cronaca [1781, 1560, 2131, 3496]. Crónicas [3264, 3959]. Cronin [2669]. Cronologico [464, 555]. Crookes [1466]. Croom [224, 341, 388]. Cropp [144]. crops [3873]. Crosbie [689]. Crosland [1406, 1568, 3023]. Cross [1110, 2015]. Cross-fertilization [2015]. crosses [2015]. Crossley [2095]. Crossroads [2866]. Crowe [2003, 723]. Crown [1392]. Crows [2549, 2625]. croyant [1734, 4089]. Crucial [1846, 2350, 3881]. Crusade [819]. Crusades [2709]. Crusca [1564, 1645]. Crystal [3579]. CS [886, 1443, 1669]. CSIC [1197]. CT [1660]. CTHS [728, 1327]. CTT [1743]. Cuadernos [43, 352, 2114]. Cuccuini [2230]. cucina [1000]. Cuec [1399]. Cuen [854, 1851, 2316, 1371]. cui [840]. Cullen [1640]. Cult [3217]. Cultivation [2772]. Cultura [1884, 2620, 180, 718, 45, 2536, 2614, 403, 352, 859, 1806, 1876, 1120, 1142, 2100, 745, 1785, 39, 189, 313, 617, 1427, 1691, 3120, 585, 1533, 3040, 1996, 2558, 887, 996, 1360, 572, 2383, 3813, 3820, 3029, 635, 1768, 558, 756, 1658, 687, 1134, 178, 2070, 2151, 517]. culturade [314]. Cultural [1005, 2746, 3137, 3437, 1778, 2957, 3636, 2884, 3740, 1146, 3819]. Culturale [1778, 1372, 2787]. Culturali [809, 20, 354, 778, 844, 1591, 1794, 1645, 1599, 2002, 2148, 3693]. Culture [517, 520, 587, 923, 1617, 1763, 2463, 2560, 2746, 1568, 1835, 2964, 3222, 2671, 2079, 2947, 3494, 1466, 3220, 1935, 3191, 1040, 782, 1411, 1319, 1397, 3148, 229, 2866, 3616, 3261, 3711, 1483, 3233, 3361, 3479, 3753, 3754, 2737, 3779, 32, 1612, 3810, 3813, 2874, 3850, 1575, 2214, 1093, 635, 1168, 3935, 2767, 1799, 1400, 4001, 2459, 3002, 56, 2600, 1411, 3114, 2688]. Culturein [2059]. culturelle [3010]. Cultures [1643, 2964, 3526, 3173, 2867]. Culturing [3903, 3035]. Cummins [2945]. cumulative [1898, 10, 3944, 3945]. Cuneo [2100]. Cunningham [2962]. Cuomo [2874]. cuore [315]. Cupido [2156]. cura [41, 202, 258, 189, 296, 299, 517, 514, 617, 794, 804, 896, 1091, 1200, 1364, 1502, 1606, 2370, 2488, 3065, 46, 1949, 1956, 1957, 1959, 1962, 2020, 2027, 2028, 2030, 2033, 1963, 2034, 1951, 2022, 1952, 1953, 1954, 2023, 2024, 2025, 1955, 1960, 1966, 2026, 2031, 2037, 1948, 1950, 1958, 2019, 2021, 2029, 1947, 2018, 2012, 1961, 2032, 2168, 2239, 2166, 2237, 2387, 2389, 339, 496, 552, 726, 59, 1902, 3120, 421, 1793, 2612, 508, 621, 569, 1006, 1322, 1317, 1488, 1487, 1569, 1648, 1661, 1679, 1682, 1765, 1816, 1804]. cura [1811, 3121, 221, 1235, 1496, 2491, 2522, 643, 1769, 2064, 2063, 2116, 2580, 223, 2620, 74, 260, 847, 180, 181, 2313, 2314, 2376, 2558, 3077, 250, 448, 715, 720, 717, 873, 915, 1105, 1077, 1283, 1157, 1279, 1280, 1189, 1156, 1360, 170, 951, 971, 2602, 1243, 1516, 1499, 1620, 1759, 1820, 2010, 2007, 2014, 1928, 2270, 2276, 2315, 2497, 3027, 1111, 1909, 1335, 1908, 752, 2461, 2614, 1242, 343, 1260, 67, 300, 504, 2202, 2492, 2623, 748, 758, 930, 1741, 1901, 1943, 407, 1190, 1183, 1101, 1115, 1098, 1178, 1287, 1675, 1716, 1755, 1677, 1754]. cura [1866, 1925, 1927, 2072, 14, 554, 710, 879, 1029, 1186, 1523, 1521, 330, 171, 1718, 71, 20, 240, 238, 237, 199, 291, 561, 581, 633, 556, 635, 779, 727, 858, 832, 861, 941, 953, 1056, 1018, 1284, 1259, 1163, 1255, 1136, 1158, 1185, 1369, 1456, 1354, 1411, 1343, 1350, 1373, 1419, 1532, 1665, 1663, 1746, 1806, 1794, 1813, 1861, 1876, 1871, 2005, 1919, 1923, 2115, 2117, 2113, 2212, 2217, 2225, 2319, 2414, 2368, 2574, 558, 1065, 337, 922, 929, 1497, 1712, 1904, 1645, 255, 757, 1459, 1788, 165, 799, 3125, 1061, 1230, 1237, 1289]. cura [1737, 1653, 1729, 3085, 1579, 1658, 2107, 2062, 2132, 2429, 2380, 2499, 83, 2384, 2374, 2432, 1840, 851, 1408, 1034, 1330, 1803, 1796, 2603, 372, 412, 555, 1836, 2460, 2138, 1064, 1388, 1318, 1420, 1615, 1483, 1642, 2205, 2324, 2433, 1137, 1267, 1252, 1269, 1325, 1324, 1519, 1530, 1223, 1990, 2622, 1802, 1900, 2141, 1728, 1651, 2367, 831, 1000, 1870, 2416]. curabile [1677]. curantibus [747]. Cureau [264]. Cures [3706]. Curi [1579]. curieuses [3024]. curiosa [3159, 3319]. curiosit [2893]. curiosità [354]. Curiosities [3412, 3634]. Curiosity [2964, 4051, 3502, 2587, 2801, 2961, 847]. Curious [3053, 2940, 3399]. Current [1898, 3944, 3945]. Curry [3510, 649, 275].

Cursi [3632]. Curti [559]. Curtis [1670, 914]. curve [941]. curvilineorum [2712]. Cusano [1157]. Cusanus [3229]. Cushing [1659]. Cusp [3701]. Cut [3389]. cutting [2009]. Cuvier [620, 2103, 17]. Cv [716, 2608]. CX [746]. Cxiii [576]. cxix [2892]. cxlii [216]. cxlii-321-xv [216]. Cxliv [1280]. Cxviii [887]. CXX [1077]. Cxxi [250, 2952]. cxxv [2590]. Cxxxiv [1401]. Cybernetic [2446]. Cynthia [3442]. Cypess [3399]. Cz [680].

D [16, 1370, 2003, 2424, 1580, 28, 32, 511, 717, 1277, 1156, 408, 933, 1997, 79, 3079, 2709, 927, 2004, 331, 566, 1753, 1866, 499, 25, 1718, 761, 916, 1411, 1563, 2731, 2975, 1414, 1458, 2803, 921, 3044, 1326, 1203, 1198, 1324, 1223, 1225, 2644, 2848, 3611, 2118, 175, 116]. D. [1991, 964, 3712]. D.C [184, 293, 1837, 859]. D.C. [3081, 2921]. d’Abaco [2130]. d’Abano [31]. d’Abbadie [2541, 3960]. Dacosta [332, 434, 436, 638]. d’Adda [1107]. Daedalus [2178]. D’Afrodisia [1483]. Dag [2308]. Dagli [716, 2888, 1092]. Dagognet [2318]. D’Agostino [2319]. d’Agricola [2770]. Dahan [2613]. Dai [1622, 1335, 562, 1452]. d’Ailly [1357]. Dairy [3474]. Dal [899, 726, 2298, 2719, 2690, 1677, 815, 1718, 1437, 1114, 2107, 2679, 957, 2726, 3747, 1957, 2028, 1955, 2026, 2364, 3783, 2490, 3345, 2651, 2492, 3412, 2197, 789, 1181, 863, 876, 917, 779, 859, 1372, 1373, 1510, 1871, 2009, 1926, 2213, 2221, 2363, 558, 3937, 3967, 921, 3741, 2810, 1324, 4014, 2266, 4038]. Dalai [2826]. Dalché [574]. d’alcune [2726]. d’Alexandre [2165, 2236]. Dalholm [1441, 884]. Dalitz [2268]. Dalla [794, 2233, 2162, 3138, 1560, 1307, 1810, 704, 1056, 3757, 1470, 2536, 2614, 3178, 1251, 2153, 2919, 1540, 1119]. Dall’antichità [3426, 1865]. dall’antico [2789]. dall’arpaeolia [355]. Dallas [1617]. Dall’Aste [1502]. dall’Ateneo [3286, 3824]. dall’Autore [1765]. Dalle [18, 3541, 792, 785, 573, 1314, 2616, 500, 875, 861, 1510, 2727]. Dall’et [1866]. dall’età [1816]. Dall’Europa [832]. Dall’infinitamente [315]. Dall’Innesto [1629, 2115]. dall’Istituto [44]. Dallo [3973, 2232]. Dall’Odissea [1428]. dall’orologio [428]. dall’U [1099]. dall’U-nit [1099]. dall’Unit [712, 1748]. Dal‘Rariorum [2531]. d’Altri [3362]. Damasio [1506]. Dame [445, 2774, 613, 2585]. Dameri [1806]. d’America [537, 644]. Dames’ [2638]. Damianaki [2142]. d’Amici [527]. Damn [4038]. Dan [2561]. Dancy [451]. Dangerous [2424, 3834, 3736]. Daniel [113, 846, 2725, 2227, 1921, 2878, 119, 3113, 1576, 850, 1878, 388, 2324]. Daniela [2267, 2491, 561, 1695, 2213, 1691, 3015, 887, 3261, 2581]. Daniele [70, 2335, 393]. Danielle [3631, 1358, 33]. Danilo [688, 340, 453, 655]. d’Anime [264]. Danish [1081, 1008]. D’Annunzio [1033]. Dans [1888, 282, 1567, 260, 32, 1187, 1447, 1442, 1444, 1984, 2085, 2215, 2519, 3461, 3087, 2652, 332, 436, 707, 3251, 3924, 1848, 3540, 3010, 2537, 2767, 1700, 3026, 1135]. Dante [711, 1258, 2307, 3062]. Danti [263, 2392]. d’Antoine [1440]. Danuta [1256]. Danzica [2592, 3934]. Danziger [769]. Dappiano [3065]. d’après [305]. Darás [2114]. d’archivio [1990]. Dare [2963]. Dario [585, 922, 1088, 2888, 522, 2951, 2960]. d’Aristote [2821]. Dark [289]. Darmstadt [188, 1446, 639, 363]. Darnton [846]. Dart [1154]. d’arte [1597, 1011]. Dartmouth [2867]. d’Arturo [2070]. Darwin [1921, 2849, 3602, 418, 770, 927, 1407, 2015, 2004, 2966, 2272, 851, 1408, 4056, 3798, 3084, 3122, 3076, 56, 3814, 3816, 3837, 418, 624, 927, 2799, 4025, 3480]. Darwin. [3082]. Darwinian [228, 2791]. Darwinism [285, 419]. Darwinisme [1073]. Darwins [2979]. Daryn [3460, 2879]. d’Assisi [572]. Daston [2684, 2559, 2983, 3428]. d’astronomie [1626, 1669, 216]. Data [2418, 990, 2291, 1167, 4002, 3840]. Database [2051, 1479]. Date [2253]. d’Athanasius [1293]. d’Athènes [2214]. Dati [989]. Dating [3730, 3292]. Dattelpalme [256]. Daubenton [2898]. Daubert [1875]. Daughters [3361]. d’Autrefois [3362]. Dava [1650]. Davenport [651]. Davenport-Hines [651]. David [1087, 2003, 115, 3025, 3483, 2553, 1588, 334, 1917, 2719, 3345, 2561, 2945, 1768, 3398, 1225, 1905, 4035, 801, 1070, 1340, 2182, 228, 677, 2805, 1717, 2371, 3066, 3111, 2719, 2470, 2977, 624, 2362, 1008, 939, 3933, 1467, 3375, 570, 2552, 674, 2803, 2939, 3058, 808, 993, 2937, 3250, 2562]. Davide [2536]. Davidson [3078]. Davina [3530]. Davis [2873, 1458, 2322, 3669]. Davy [1070, 1588]. Dawn [3456, 1825]. Dawne [1584]. Dawson [2737, 3637]. Day [2361, 3131, 2840, 1865, 2009]. Dazzani [1015]. Dc [242, 885]. Dd [645]. DDR [226, 507]. dea [2748]. Dead [3215]. Deakin [2935]. Dealer [2183, 2185]. Dealers [2180]. Dean [2957, 1337]. Deane [2525]. Dear [805, 2862, 3018, 1723]. Dearborn [2194]. Death [651, 3220, 3309, 3219, 2724, 1440]. Debate [1778, 2781, 3375, 2858, 3610, 723, 1146, 3821]. Débats [2704]. Debora [2662, 1806, 2505]. Deborah [2671, 2948]. Debus [301, 491, 837, 1680, 2579, 2828]. débuts [1879, 3568, 540]. decanis [1280]. December [3194, 2169, 2240, 2852, 519, 4074, 4077, 3591]. décembre [4043, 2229, 4077, 502, 2201]. decennio [2154]. Deception [3196]. Declaration [2978]. Decleva [1759]. declinazione [1166]. Decline [614, 3220, 1210]. Declino [1210]. Decoding [3919, 3060]. decoration [1331]. decorazione [1331]. Découverte [748, 1475, 3768, 2215, 2273]. découvertes [410]. Dedalo [1335, 682, 330, 240, 1836, 2810, 1719, 2178]. Dedeyan [413]. Dedications [3419]. Dedomena [3840, 4002]. deduction [3929]. Deduttivo [2246]. deduzione [3929, 1994]. Deel [802, 1367]. Deep [3627]. Defence [1970]. defenders [624]. defense [2248, 2357, 2356]. Defenso [2357]. Defesa [2356]. Definition [1147]. definizione [1693]. degeneration [850, 3957]. Degli [1054, 1691, 1894, 2421, 2467, 159, 1951, 2022, 1953, 2024, 2334, 4, 737, 1784, 336, 2246, 269, 1314, 821, 1383, 2580, 75, 2313, 2376, 1471, 2090, 2969, 72, 991, 1102, 1156, 70, 792, 809, 741, 1124, 1499, 2224, 2789, 1828, 323, 38, 396, 463, 2447, 600, 3838, 3856, 1302, 522, 1477, 1782, 1973, 1192, 2468, 663, 710, 2129, 2353, 2693, 2397, 71, 291, 780, 727, 832, 1033, 1259, 1195, 1372, 1452, 1423, 1671, 1861, 2005, 1931, 2009, 2117, 2151, 2212, 2149, 2217, 2271, 2415, 2414, 1712, 1230, 2380, 4083, 397, 2249, 2285, 95]. degli [1426, 1633, 3416, 4007, 1318, 1615, 528, 3335, 1870, 1719]. d’Egypte [2118, 4053]. Dehtleef [140, 140]. d’Einstein [4089]. Dejan [2069]. DeKalb [2660]. Dekker [3325]. Del [1441, 2200, 1300, 459, 3286, 2452, 884, 1778, 1629, 828, 321, 2309, 2346, 2356, 2473, 2248, 1823, 904, 1380, 2400, 3103, 1780, 3852, 3853, 2614, 2041, 50, 596, 1975, 2178, 2173, 1774, 393, 1893, 2174, 664, 1976, 2258, 453, 953, 1871, 211, 529, 1210, 1551, 1978, 660, 1634, 3019, 2815, 2910, 2130, 1832, 1000, 39, 202, 197, 189, 313, 517, 514, 721, 794, 777, 852, 1091, 995, 1427, 1616, 1539, 1680, 1691, 2159, 2488, 3744, 736, 4040, 740, 1239, 1951, 2022, 1966, 2037, 2163, 2234, 2167, 2238, 2333, 2334, 2336, 2338, 2389, 3689, 807]. del [726, 2942, 3757, 340, 336, 493, 1903, 456, 612, 686, 1488, 1811, 3040, 3121, 221, 1236, 1496, 2522, 1996, 2116, 2780, 1790, 278, 2284, 3788, 180, 2376, 486, 3279, 1970, 3015, 2849, 2685, 3077, 3801, 72, 250, 312, 311, 447, 576, 716, 718, 719, 915, 1279, 3001, 70, 949, 572, 809, 2619, 1243, 1620, 1869, 1928, 2232, 2224, 2276, 2504, 2490, 3457, 1428, 2863, 3006, 3209, 38, 2520, 2949, 3820, 4059, 2572, 1917, 3839, 2934, 2742, 3057, 3029, 4064, 2461, 3851, 2893, 4067, 1706, 522, 1260, 2651, 3620, 80, 300, 2492, 2494, 758, 1190, 2419]. del [789, 1101, 1030, 1178, 1438, 1675, 1672, 1866, 1927, 2222, 2275, 2430, 2468, 2287, 876, 1186, 2743, 3426, 1718, 3277, 2129, 2880, 431, 428, 429, 425, 561, 581, 575, 730, 776, 780, 782, 858, 937, 864, 854, 1162, 1257, 1284, 1163, 1255, 1241, 1159, 1453, 1369, 1343, 1424, 1373, 1591, 1730, 1746, 1813, 1860, 2005, 2001, 2112, 2115, 2110, 2113, 2111, 2153, 2217, 2271, 2415, 2548, 999, 2877, 929, 3933, 618, 3949, 2081, 1414, 678, 1323, 578, 838, 186, 1075, 2107, 2429, 2499, 3967, 2951, 3062, 1140, 2285, 3181, 3972, 3973, 4084, 2399, 3710, 687]. del [942, 1349, 2108, 2919, 12, 2274, 3986, 3987, 3988, 1836, 2650, 2820, 3632, 734, 4010, 3329, 1233, 1318, 2100, 2796, 1137, 1270, 1644, 1134, 1223, 2714, 1783, 1990, 2622, 1900, 281, 1728, 2628, 2322, 831, 2627, 178, 2726, 2968, 2692, 1389, 4017]. del’500 [181]. del’600 [181]. Delalande [3362]. Delambre [620]. Delbourgo [2844]. Dele [1050]. d’électricité [260]. Deleuze [756]. Delfina [897]. Delfino [912, 789, 644, 876]. Delfo [970]. Deli [1050]. Delia [2426]. Delicatissime [2815]. delinquenti [3547]. Della [740, 1035, 1205, 1374, 1463, 1221, 3772, 215, 1820, 2041, 2525, 47, 731, 813, 973, 827, 3443, 1890, 3410, 1546, 744, 2394, 3329, 89, 41, 61, 189, 246, 299, 517, 583, 794, 777, 773, 781, 896, 1091, 1017, 1200, 1201, 1364, 1606, 1691, 1894, 2421, 2370, 2488, 3299, 736, 2043, 46, 315, 521, 593, 1946, 2017, 1948, 2019, 2169, 2240, 2166, 2237, 2338, 2339, 3099, 4, 934, 2942, 3760, 737, 975, 594, 1037, 459, 1902, 3286, 336, 265, 421, 688, 2612, 4044, 3094, 18, 1887, 269, 621, 532, 677, 1322, 1564, 1648, 1679, 1811]. della [821, 221, 1236, 1235, 643, 1711, 1574, 1815, 2064, 2063, 2073, 2431, 2471, 1383, 2293, 2620, 828, 1112, 2631, 2797, 2865, 1215, 214, 989, 2313, 2314, 2376, 1829, 2826, 2616, 1471, 2090, 2670, 85, 250, 311, 442, 447, 716, 910, 1105, 1077, 1360, 572, 809, 2248, 1701, 1124, 1243, 1516, 1499, 1560, 1513, 2010, 2224, 2490, 1828, 2444, 2883, 3115, 3209, 323, 1335, 1248, 3807, 2172, 3812, 2536, 3818, 463, 3090, 2447, 3824, 3831, 1164, 2348, 2941, 600, 1917, 3839, 1995, 1377, 4067, 2614, 1036, 2502, 1302, 1706, 2712, 522, 1550, 986, 2557, 81, 300, 1782]. della [1973, 2140, 2720, 2773, 599, 909, 1901, 2590, 162, 874, 881, 1183, 1561, 1979, 2261, 360, 331, 503, 505, 639, 638, 709, 867, 947, 948, 952, 1027, 1096, 1092, 1030, 1098, 1176, 1431, 1674, 1675, 1716, 1752, 1672, 1677, 1676, 1865, 1809, 1864, 1810, 2072, 2413, 3496, 1382, 1554, 3402, 663, 3911, 3481, 979, 433, 435, 564, 554, 710, 697, 711, 871, 1093, 661, 391, 200, 2129, 2353, 2397, 2880, 90, 88, 71, 261, 238, 199, 291, 294, 458, 428, 657, 556, 590, 575, 656, 634, 778, 858, 862, 859, 1116, 1062, 1011, 1015, 1056]. della [1018, 1025, 1292, 1168, 1251, 1259, 1163, 1136, 1166, 1165, 1452, 1343, 1435, 1350, 1419, 1423, 1599, 1746, 1749, 1806, 1794, 1813, 2002, 2005, 1923, 2009, 2115, 2117, 2078, 2148, 2226, 2150, 2217, 2225, 2307, 2319, 999, 211, 1897, 1344, 1712, 2677, 318, 1645, 3314, 658, 3388, 3949, 2833, 3652, 1210, 304, 319, 570, 1788, 2722, 678, 1323, 186, 1075, 1237, 1289, 1394, 1579, 1998, 2107, 2062, 2132, 2264, 397, 2249, 83, 3136, 1971, 1697, 2597, 3062, 3976, 3343, 924, 1408, 3547, 95, 3710, 942, 1034, 1068, 1244, 1455, 1348, 1491, 1803, 1988, 1633, 3203]. della [371, 373, 3989, 2914, 3190, 1119, 2269, 3655, 734, 2501, 2581, 1565, 2433, 2503, 2810, 1267, 1270, 1331, 1540, 1530, 1783, 1802, 2141, 2044, 528, 739, 818, 3335, 1651, 4021, 2322, 831, 2960, 2088, 2391, 2692, 1834, 788, 2461]. della’astronomia [2113]. dell’abisso [2384]. dell’Accademia [2905, 1251, 1873, 2226, 46]. dell’acqua [2133]. dell’acque [2525]. dell’alchimia [3544]. dell’alimentazione [872]. dell’allestimento [1411]. dell’America [434]. dell’anima [1867, 3178]. Dell’Anschaulichkeit [2567]. dell’antico [1241]. dell’archeologia [3209]. Dell’Architettura [783]. dell’Archivio [844, 558, 407]. dell’‘Archivum [2522]. Dell’Argomento [2349]. dell’aria [1947, 2018]. dell’aristotelismo [719]. dell’armonia [2434]. dell’Arte [14, 432, 3209, 2461, 1373, 1485]. dell’Astronomia [832, 1813, 2005, 2217]. dell’Ateneo [72, 310, 306]. dell’Atomismo [1789, 456, 2933]. dell’atomo [456]. dell’Atto [1971]. dell’attuale [1352]. dell’audizione [203]. dell’ed [788]. Dell’Eden [2125]. dell’Edizione [2608, 2726, 1496, 2605, 288, 2138]. dell’elettricità [89]. dell’elettroshock [3039]. dell’Embrione [61]. dell’empirio [1691]. dell’empirismo [2467]. dell’Enciclopedia [337]. dell’eruzione [2333, 2284]. dell’esercito [1436]. dell’esperienza [2324]. dell’esperimento [1533]. Dell’Et [604]. dell’Età [2797]. Dell’Etere [1976]. dell’Etna [2333, 2284]. dell’etnologia [537]. dell’Europa [562]. dell’i [2748]. dell’i-dea [2748]. dell’iconografia [1175, 1926, 4014]. dell’Imperatore [2233, 2162]. dell’individuale [1754]. dell’industria [449, 971]. Dell’Infanzia [1477]. dell’infinito [2951]. dell’influenza [2430, 4010]. Dell’Informazione [1343, 2321, 3402]. dell’Inghilterra [2721]. dell’innovazione [1020]. dell’insegnamento [1183]. Dell’Interpretazione [1606]. dell’inventore [3757, 2919]. dell’Irritabilita [1701]. Dell’Istituto [2248, 1979, 521, 1035, 1205, 1374, 1463, 1477, 599, 2261, 391, 47, 90, 261, 458, 657, 731, 813, 973, 1116, 1292, 2168, 2239, 75, 2315, 1677, 14, 238, 1033, 1512, 12]. Dello [2473, 1970, 951, 926, 2202, 705, 878, 1158, 1665, 2213, 2429]. dell’occulto [765]. dell’Oceania [864]. dell’Officina [44]. dell’Orologeria [1510]. dell’Orso [1483]. dell’Orto [83, 1137]. dell’Ospedale [2637, 2871, 709, 789, 1756, 953, 876]. Dell’Osservatorio [1220, 1555, 291, 631, 1082]. dell’Ottocento [1240, 1718, 3545]. dell’Ufficio [1350, 1666]. dell’ultimo [589]. dell’Umanit [708]. dell’UNESCO [2131]. dell’Unione [1013, 1160]. dell’unità [1260]. Dell’Università [1780, 1706, 493, 2313, 2376, 250, 1105, 1869, 2010, 697, 727, 864, 1861, 1931, 2414, 2100, 1252, 448, 1316, 448, 873, 1243, 1195, 1665, 2149, 2414, 83, 1270, 4019]. dell’universo [1114]. Dell’Uomo [2351, 537, 2334, 2285, 745, 178]. dell’URSS [46]. Delpino [1952, 2023]. dels [3019]. Demaitre [2875]. dementi [1335]. Dementia [288]. Democrat [16]. Democrito [2933, 3277]. Democritus [2933, 2933]. Demoiatria [640]. demon [1874]. Demonstration [1763, 2360, 2423, 3966, 4001]. Dendrochronology [2483]. Denecke [108]. Denis [2108, 1993]. Denise [1841]. Dennet [1921]. Dennis [576, 2957]. d’enseignement [1442]. Dentaria [432]. dentition [4008]. D’Entrecasteaux [1304, 1472]. Dents [707]. Déodat [2965]. Deparcieux [2528]. Department [648]. Depicting [3648]. d’Epicure [3865, 3861, 2892, 3312]. d’épidémie [3010]. d’epoca [1510]. Depots [3484]. Depressed [337]. Depressive [645]. depressivi [645]. depuis [620]. Deputación [2068]. Deputazione [2667, 2100]. d’erbari [1270]. Derek [406, 219, 2187, 2188]. Deriva [744, 656]. Desargues [206]. Descartes [1364, 3735, 1843, 2307, 450, 1228, 3203, 3673, 2772, 1379, 842, 196, 3524, 3256, 3862, 3923, 2057, 3180, 1609]. Descendenztheorie [1042, 1043]. Descent [3102, 1043, 1042]. Described [2565]. Describing [3017, 2775]. descriptio [1324]. Description [2267, 574, 3663, 3244, 2552]. Descriptions [405]. Descriptive [696, 2555, 2554, 1792, 2556, 3823]. descritta [1160]. d’Esculape [332]. Desedibus [337]. desiderata [4061]. Desiderate [4061, 1285]. Desideri [1330]. desiderio [3614]. Design [885, 1970, 3663, 2966]. Designing [3033, 2485, 2594]. Designs [3201]. Desirable [3451]. desire [1918]. Desmond [1914]. d’espèce [282]. Dessì [1148]. Destinatari [6, 905]. Destiné [1475]. Destiny [2800]. Destruction [1891]. d’Estudis [353, 3019]. Detail [3411]. Detailed [405, 3570]. Detective [3726]. determined [2728]. d’Étienne [3665]. Detleef [140]. d’Etnografia [864]. detta [1091, 200, 412]. Detti [546]. detto [2461]. d’étude [1447, 1859, 2229, 1734]. d’Études [512, 444, 1442, 32, 3952]. d’Eugenio [2215, 3768]. d’Europa [644]. deutsch [3893, 1912, 567]. deutsch-französischen [3893]. Deutsche [1992, 4078, 690, 547, 3186]. Deutschen [1735, 2196, 690, 542, 2532, 1074, 1396, 3968, 3981, 2857]. Deutscher [497, 623, 557, 1715, 1714, 1797, 3977, 1336, 1905]. Deutsches [961, 1862, 2409]. Deutschland [2228, 415, 3902, 3999]. Deutschsprachige [623]. deutschsprachigen [361]. devant [582]. Development [2635, 3282, 3422, 2067, 3968, 1999, 235]. Developments [11, 2470, 1810]. Deventer [2555]. Devices [3291]. devil [1874]. Dévoilé. [795]. devoir [1232]. DeVun [3633]. d’Henri [3805]. d’Henry [2318]. d’Histoire [2145, 675, 1418, 1327, 1401, 1304, 1472, 2995, 1421, 753, 749, 2856, 1486, 2267, 4042, 2898, 3280, 260, 243, 444, 836, 954, 713, 3287, 1080, 2118, 3992, 1224, 1473]. Dhombers [620]. Dhombres [653, 1612, 40]. d’Hygiène [1040]. Diabete [1438]. diaeta [1145]. Diagnoses [3706]. diagnosi [1693, 645]. diagnosis [676, 3454]. Diagnostik [676]. Diagrams [3668, 3533, 3363, 3272]. Dial [2346]. dialettica [377]. Dialoghi [255]. Dialogo [3858, 2395]. Dialogue [3258, 754, 3953, 2139, 748]. Dials [2347]. Diamond [3351]. Diana [1538, 1598, 2072]. Diari [3497]. Diario [3788, 1869, 1521, 2374]. Diary [3788]. Diasio [2060]. Diatetikvorlesung [565]. diavoletto [1874]. diavolo [1874]. Dibattista [2557]. Dibattito [1778, 1629, 1824, 211, 2167, 2238, 250]. d’Ibn [353]. Dibner [2731]. dicano [3373]. Dicembre [1871, 1360, 572, 1176, 199, 1137, 1530]. Dichter [363]. diciembre [45]. Dick [951, 2779]. Dickson [2106]. Dictionary [532, 1097, 57, 2362, 3907, 2752]. Did [3133]. Didactic [3487, 2686]. Didattica [2476, 13, 2557, 1424]. didattici [1452]. Didattico [859]. Diderot [413, 3160]. Diderota [257]. Didier [3671]. Didn’t [1208]. died [3817]. Diego [965]. Diepholz [2196, 2466]. Diet [3398]. Dieter [567, 1715]. diététique [3197]. Dietrich [98, 3650, 137, 307, 557, 1731, 1336]. Dietro [1120]. Difendere [3057]. Difesa [1970, 2248]. Difference [177]. Differences [3582]. Different [3378, 2156, 4022]. Differential [3104]. différentiel [681]. Differing [3226]. Difficile [2567, 979]. Difficult [1699, 1198]. difformi [66]. Diffusion [1795, 582, 2214]. Dig [3484]. Digital [3629]. Digiuno [2364]. Digne [2757, 2757]. Digne-les-Bains [2757, 2757]. Dijon [4047, 1326]. Dilemma [3656, 2804]. Dilettante [3702]. Dilg [1285]. Dillon [1283]. diluvio [493]. d’Image [527]. dimension [1469]. Dimensioni [1469]. Dimensions [419]. Dimenticata [2264]. DiMeo [3615]. Dimitri [1225]. d’immagini [457]. dinamica [2155]. d’influences [728]. d’Ingegneria [1062]. dingen [3109]. Dinges [2365]. Dini [2334, 874, 2285, 924]. Dino [1105, 2213, 390, 187, 2100]. Dinosauri [1113]. d’insegnare [1174]. d’intërieur [425]. dintorni [2951]. d’inventions [410]. Dio [250, 953, 2432]. Diodati [2652]. Diognostic [645]. Diogo [2895]. Dioptrics [2710, 2347]. Diorama [3538]. Dioxide [2971]. Dip [1648]. Dipartimenti [1780]. Dipartimento [250, 1706, 575, 1258, 1195, 1423, 1591, 1532, 2117, 2078, 2271, 2415, 1712, 1318, 1615, 2498, 2954, 2622]. Diplomacy [3512]. Diplomat [3073]. Diplomats [3639]. Diptych [998]. Diputacio [2101]. Dirac [1296]. directed [40]. Directeurs [2648]. Directing [3469]. direction [2067, 620, 1612, 1080, 1848, 1410, 1228, 1073, 1799, 1326, 1486]. Directions [1126]. directory [3905, 992]. diretta [588, 2202, 2107]. direttamente [1765, 44]. Direzione [523, 2416]. Direzioni [304]. dirigée [3368, 866]. Dirigibili [2679]. diritti [3718]. Dirk [3328]. d’Isabelle [2652]. Disability [3536]. disasters [4082]. dischetto [1572]. disciplina [2836, 3057, 4028]. Discipline [1723, 2557, 4028]. Disciplines [1500, 2967, 3999]. disco [502]. Discontinuità [595]. discontinuity [595]. Discorsi [1829, 1128, 2869]. Discorso [1605, 4059, 3426]. Discourse [2840, 4059]. discourses [759, 1829]. discover [3060, 3919, 840]. Discovered [2175, 778]. discoveries [2551]. Discovering [3086]. Discovery [885, 421, 1564, 349, 3399, 36, 3768, 2215, 2678]. Discussion [3373]. discussione [2159]. Discussioni [826, 327, 1130, 2455, 9, 161, 1148, 1385, 273, 531, 742, 1480, 1785]. Discussions [3213]. Disease [1152, 1222, 3656, 3185, 1720, 3864, 2071, 2752]. disegni [2590]. Disegno [278, 1970, 2232, 3973]. Disgrace [2085]. Disgraceful [3937]. disingannata [1957, 2028]. disorder [850, 3957]. Disorders [645]. dispersione [2614]. Display [2700, 3717, 2303, 1411]. Displayed [3579]. Displaying [3622]. dispute [812, 2678]. dissection [3801, 2835]. Dissections [3220]. Dissent [414, 2633]. Dissertatio [1145, 2337, 2288]. dissertations [1829]. Dissertazioni [1829]. dissezione [3801]. dissezioni [1644]. dissimulata [2659]. Dissociazione [979]. dissonanti [718]. Distance [1377]. distances [812]. Distant [3245]. Distanza [1547]. distanze [812]. Distillateurs [795]. Distillling [2591]. distribution [2833, 3026]. disturbi [645]. Disziplinen [415, 3999]. d’Italia [1752, 1748, 1787, 1904, 555, 3752, 2789, 3089]. diteurs [1879]. Divagazioni [2630]. divazioni [3956]. Diversibus [2145]. Diversis [2267, 2596, 2369]. Diversità [758]. Diversity [3602]. Divide [2671]. Divided [489]. Divination [2094]. divinatoria [2942]. Divine [1188]. divisioni [2499]. divisions [3956]. divulgazione [2672]. Divulging [1231]. Dizionario [532, 912, 1173, 1431, 1716, 554]. DM [64, 173, 172, 220, 1918, 950, 622, 1731, 1838, 1797, 361, 362, 690, 1905]. DM/SFR [1905]. Dmitrij [1870]. DN36 [2683]. Do [2356, 2357, 1305, 2895, 1744, 2068, 3990]. Dobbs [1002]. Doc [2118]. d’Occidente [1434]. Doce [1544]. docente [3121]. Docenti [2849, 2468]. Doctor [1476, 787, 3972, 2959]. doctoral [4000]. doctorands [1859, 4000]. Doctors [80, 34, 2609, 34, 1275]. Doctrine [25, 1210, 388, 2763, 1100]. Doctrines [2648]. Docu [3505]. Docu-Museum [3505]. Documenta [3221]. Documentación [2114]. Documentaire [2051]. Documentales [1855, 2097]. documentaria [2497, 1741, 844]. Documentary [2051]. Documentation [1478, 1556, 1635, 675]. Documentazione [1746, 1806, 1424]. Documenti [313, 736, 487, 252, 200, 556, 1990, 1802, 407, 1251, 1453, 1373, 558, 1230, 3088]. Documentie [1512]. Documento [1634]. Documentos [430]. Documents [2048, 2566, 584]. Dogma [551, 2132]. Dogmatische [217]. Dogmatismo [1368]. Dohrn [201, 1042]. Dohrns [1042]. DOI [3530]. Doing [3288, 4069]. Dokumente [420, 787]. Dokumenten [64]. Dolby [2066]. Dollo [1648]. Dolnikowski [1686]. Dolomieu [2965, 2782]. dolore [3426]. Dolza [897, 3094, 1020]. Dom [2525]. Domain [3664]. Domains [3849, 1845]. domato [1242]. Domenico [2164, 2235, 2326, 2434, 1705, 1145, 1453, 2650, 3315, 1990, 814, 4050, 53, 223, 2825, 880, 1704, 3297, 1857]. Domergue [1135]. dominio [81]. Dominique [3362, 1799, 2613]. Domninus [3298]. Domus [4040, 4040]. Don [2345]. Donald [652, 592, 885, 1535, 2064, 36]. Donata [2428, 3021]. Donatella [1251, 1748, 2221]. Donati [2049, 2399]. Donativi [2126]. Donato [1522, 3443, 1369]. Donazione [2479]. Done [3334, 3868]. Doni [555]. Donna [2351, 1980, 2104, 178]. Donne [3689, 897, 1029, 3741, 4067, 786, 2367, 3545, 3547]. Donnés [2051]. D’Onofrio [377]. Donovan [1588]. Don’t [3537, 1505]. Donzelli [1605, 1729, 2432, 2627, 2968]. Doom [2840]. door [802, 3109, 541]. doors [1367]. Dopo [1118, 48, 3752, 3089, 3085]. doppio [611]. d’optique [3071]. Dora [1911, 1329]. Dorandi [3277]. Dordrecht [16, 2708, 2462, 28, 1401, 606, 1570, 1600, 3224, 2583, 2967, 1857, 1498]. Dordrecht/Boston/London [2708, 1570]. Dore [1062]. d’organo [1092]. Doria [861, 1142]. d’Oriente [1434]. d’origine [3058]. Dorigny [2229]. Dorinda [440, 587, 1536, 17]. Dorland [2731]. Dormant [2184]. Dorna [1671]. Dorner [3714]. d’oro [1540]. Dorothy [642, 16, 1409]. D’Orsi [2010]. Dossier [2148]. Dottor [641]. dottori [2460]. Dottrina [1210, 3179]. dottrine [2720]. Double [3050, 3107, 611]. Douglass [440, 2483]. Doutes [2578]. Douthwaite [2424]. d’outil [3424]. Dover [341]. Downfall [3538]. Doxographica [798, 994]. Dr [855, 2556]. Dr. [2765]. Drake [968]. Draw [3661]. Drawing [427, 735, 699, 3973]. Drawings [3200, 3053, 3415, 840, 2886]. Dream [2730, 2446]. Dreams [885, 3333, 3985]. Drebbel [3244]. Dresden [1012, 1083, 860]. Dress [2746]. Drift [348, 656, 348, 3429]. Drifting [693, 692]. Dritte [1001]. dritter [1001]. Driver [3695, 2845]. Droescher [2468]. Droscher [1992, 3095, 3691]. Droysen [154]. Droz [2843, 1567, 198, 3527]. drug [3938]. Drugman [1411]. Drugs [3736, 1724, 1433]. druk [541]. Dry [3700]. Dubinsky [3415]. Dublin [916, 176]. Duc [3701]. Ducale [726]. Ducato [726, 581, 858]. Duccio [1820]. Duch [425]. Duchet [2896]. Duckworth [833, 691, 387]. Ducos [1985]. Ducth [802]. Dudley [2786]. Due [897, 2354, 3763, 250, 2248, 1782, 3885, 268, 1298, 1745, 1464, 2394, 668, 2720, 453, 186, 2132]. Duecento [1423]. Duels [3467]. Duemila [2266, 4038]. Dugald [209]. Duhem [1859, 1560, 2469, 4000]. Duhem. [625]. Duillier [3490]. Duke [1212, 1841, 1144, 3278, 3756]. d’Ulm [2898]. Dumit [2506]. Dumont [1004, 2958, 1138, 362]. d’un [1187, 1626, 3365, 2652, 2052]. Duncan [3091]. Duncker [3226]. d’une [3127, 2214, 2762, 3004]. d’Ungheria [1372]. Dunham [1089]. Dünkel [3129]. Duntze [157]. Dunze [157]. Duplication [527]. Dupré [3253, 3526]. Duprez [2770]. Duprez-Masson [2770]. dura [416, 211]. durabilità [1373]. Durand [3280]. durante [880, 2154, 1464]. Durata [1802]. d’Urbino [425]. durch [1141]. Dure [2169, 2240]. durée [2528]. Durham [1841, 1144]. During [1495, 3214, 2443, 3037, 1825, 2607, 703, 3691, 1464, 1249, 1408, 3274]. Duris [2053]. Dürkheim [787]. Dussert [2013]. Dussert-Carbone [2013]. Dutch [2887, 3495, 57, 2944, 1422, 541, 3644, 3907, 3561, 3508]. Duxford [2777, 2740]. DVD [2363, 2662]. Dyck [2003]. Dyer [3616]. Dyes [3666]. Dyestuffs [2059, 3850]. Dynamics [1702, 1500, 1857, 4057]. dynamique [516].

E-Book [3641, 3659, 3677, 3697, 3720, 3743]. E. [2055, 1808, 31, 1686, 1901, 1096, 1177, 1756, 854, 1014, 1749, 2054]. Eagle [174]. Eahmh [1936, 1934, 1933, 1935]. Eahmh-Research [1934]. Eamon [1397]. earlier [1275]. Early [491, 615, 837, 1108, 3240, 2717, 3112, 3543, 3328, 3002, 2882, 1842, 2891, 3007, 2964, 3653, 1703, 3050, 2649, 30, 649, 3558, 3651, 3441, 2589, 2824, 2579, 1481, 2514, 3066, 34, 2657, 3198, 2827, 2823, 3070, 2470, 3860, 341, 52, 3253, 3871, 2521, 2781, 3342, 2832, 2656, 3671, 2768, 3731, 3244, 3455, 2945, 3597, 3159, 2940, 3492, 3252, 1319, 2204, 3712, 3114, 3202, 2583, 3231, 1397, 2711, 2585, 3737, 1542, 3612, 3444, 3398, 3493, 1733, 2575, 2404, 378, 3711, 3669, 3519, 2587, 2582, 3233, 3361, 2816, 3587, 3419, 2127, 3363, 3422, 1261, 2228, 3074, 3902, 763]. early [3018, 3913, 2844, 3968, 4018, 3175, 3522, 3501, 3835, 3857, 3644, 2781, 3407, 3912, 2554, 3513, 3274, 2556]. Early-Fourteenth-Century [3669]. Early-Modern [3363]. Earrings [3683]. Earth [551, 1717, 2756, 656, 2976, 289, 1654, 2802, 3789, 1076, 2777, 3789, 3792, 3795, 3796, 3430, 3176, 3431, 2259, 2361, 3159, 3623, 3183]. Earthly [2636]. earthquake [519]. earthquakes [1429]. Eastern [2879, 3267, 3111, 3947]. Eastwood [2999]. easy [1199]. Ebbesen [1684]. Eberhard [725]. eBook [3530]. eccellenti [1271]. Echt [3344]. echten [2265]. Ecig [1096, 354]. Ecig-Edizioni [354]. Eckart [1838, 690, 1838]. Eclipse [3292]. Ecole [4043, 3069, 2782, 1090, 1727, 287, 2826, 243, 305, 314, 2214, 425, 2201, 2613, 1135, 3069]. Écoles [868, 2214]. Ecologia [1350, 1666]. Ecologists [1738]. Ecology [1852, 1738, 1350, 3746, 204, 1944]. Economia [1502]. Economica [2364]. Economics [2807]. economista [2916]. Economy [209, 267]. Écrire [1543]. écrit [1232]. Écrits [3287, 3992]. Écriture [582, 3087, 2902]. ECT [3505]. Ed [520, 490, 672, 495, 532, 63, 2473, 990, 1975, 1973, 909, 2196, 1774, 332, 331, 434, 436, 639, 640, 706, 785, 707, 958, 947, 948, 872, 869, 882, 1027, 1115, 1099, 1026, 1287, 1181, 1175, 1173, 1522, 1810, 2154, 1410, 818, 1498, 1831, 41, 242, 439, 438, 580, 1090, 1078, 1108, 1139, 1454, 1649, 1692, 2194, 2227, 2160, 2464, 2916, 2708, 2055, 772, 2878, 3509, 2889, 496, 2942, 1576, 228, 2896, 2669, 652, 673, 696, 699, 1368, 2517, 2898, 2882, 3063, 3475, 3081, 2981, 2891, 3280, 3282, 166, 287, 3526, 2689, 3129, 2849, 2685, 873, 830, 1102]. ed [1277, 1278, 1358, 1685, 1984, 1462, 2648, 2619, 1758, 2080, 2161, 2229, 3408, 3022, 1063, 2579, 2883, 2863, 2531, 2572, 2808, 2774, 2706, 64, 220, 3344, 275, 682, 2955, 3029, 3289, 2633, 2905, 3225, 2787, 767, 1409, 1732, 300, 2139, 2892, 2596, 1196, 2629, 419, 1915, 1944, 2590, 881, 3427, 360, 789, 868, 1028, 1812, 1864, 1936, 2071, 2365, 3481, 2713, 712, 710, 711, 871, 863, 876, 875, 1029, 1184, 1182, 2790, 2684, 622, 2660, 779, 844, 1251, 1195, 1749, 1872, 2153, 2195, 2855, 534, 3088, 2206, 2716, 2809, 1071, 2753, 1114, 3652, 1414]. ed [1459, 2722, 2552, 2725, 2841, 3284, 1855, 997, 1229, 229, 163, 1786, 1980, 2231, 2867, 3965, 2608, 3096, 2786, 2998, 3044, 1734, 765, 3710, 2593, 289, 808, 692, 803, 892, 2919, 12, 3011, 2921, 378, 2902, 2581, 2691, 2752, 2714, 2456, 2779, 1493, 2644, 1290, 3428, 2726, 2578, 2054, 1442, 3069]. ed. [3442, 3441, 3630, 3462, 3497, 3599, 3649, 3670, 3445, 3613]. Ed.s [491, 440]. edd [2273]. Eddington [1503]. Eddy [2803]. Edicin [2876]. Edición [1226]. Ediciones [1544]. edicone [1620]. édifiantes [3024]. Edificando [2523].

Edifir [2012, 1488, 3209]. edifizj [2330, 2281]. Edimond [2523]. Edinburgh [652, 3289, 624, 169]. edita [14]. Edité [1224, 1765, 444]. Edited [802, 1367, 2329, 2331, 2337, 2328, 2332, 2335, 2330, 2333, 2334, 2336, 2338, 2339, 2327, 2436, 1611, 1495, 1898, 1911, 2284, 10, 249, 2283, 2279, 2288, 2287, 2282, 2280, 2285, 2286, 2278, 1408, 2281, 945, 3368, 1967, 2038, 3395, 3328, 3362, 1238, 1603, 2873, 334, 1164, 3365, 3366, 933, 848, 1275, 2076, 1001, 753, 1229, 1458, 1907, 3315, 3329, 3364, 1592, 1225, 3363]. editées [921]. édités [540]. Editeur [795, 1795, 2428, 855, 1985, 3010, 2375]. Edith [2786]. Editiemme [638]. Editime [864]. Edition [567, 1661, 1681, 3282, 28, 58, 1448, 1941, 2156, 3060, 3593, 3225, 2139, 1841, 566, 338, 1141, 997, 1907, 1133, 3684, 4015, 3205, 226, 244, 2362, 2525, 406, 1731, 3946, 749, 4024, 3282, 3368, 834, 2898, 1327, 3251, 1060, 2013, 1848, 2767]. Editions [2894, 2898, 2613, 2716, 728, 282, 1418, 584, 620, 287, 305, 836, 1612, 1668, 2215, 368, 748, 419, 2015, 1944, 540, 701, 1080, 1170, 1609, 2118, 1848, 2052, 921, 2273, 1386, 2422, 608, 2067, 295, 2423]. editor [11, 227, 805, 1264, 1152, 2695]. Editora [2895, 3264]. Editore [61, 516, 517, 573, 583, 794, 764, 804, 1605, 2488, 2142, 726, 456, 1316, 1322, 1487, 1679, 1765, 1816, 221, 619, 2073, 2580, 3015, 252, 447, 1283, 1157, 1189, 70, 949, 2665, 2010, 2883, 3127, 413, 926, 3298, 2536, 2955, 2557, 1183, 3258, 789, 788, 2525, 644, 389, 171, 277, 455, 589, 878, 1056, 1136, 1451, 1435, 1748, 1695, 1863, 2077, 2574, 626, 680, 799, 2132, 2429, 377, 2138, 1233, 1142, 1388, 1420, 2324, 1269, 3259, 1000, 2133, 2692, 3089, 39, 223, 752, 2308, 930, 407, 874, 554, 876, 1180, 776, 1082, 1033, 1730, 1768]. editore [800, 3039, 1034, 2136, 1906]. Editores [686, 1515, 35]. Editori [852, 1200, 807, 2491, 2431, 2364, 2014, 2504, 2323, 429, 635, 704, 730, 859, 1585, 2115, 771, 578, 2380, 750, 1565, 607, 1201]. Editoria [1539]. Editorial [965, 2508, 3143, 3548, 3660]. Editoriale [1869, 1186, 2220, 1663]. Editoriali [1994]. Editors [1264, 846, 164, 224, 722, 1370, 519, 1494, 2990]. Editrice [1691, 3065, 315, 1902, 621, 1817, 1613, 1581, 1580, 1790, 2588, 250, 312, 2140, 645, 2275, 3426, 457, 562, 780, 782, 861, 1018, 1240, 1165, 1343, 1666, 1745, 1813, 2001, 2070, 2113, 2225, 929, 165, 2679, 13, 1796, 1318, 1252, 1134, 85, 880, 697, 1861, 1931, 2149, 1653, 1802]. Editrici [2882]. Ediz [877]. Edizione [2370, 2192, 1569, 1765, 2172, 1260, 1973, 2590, 638, 2451, 171, 2107, 2644, 1852, 1790, 1098, 3277, 1323, 1483, 2205, 2914]. Edizioni [197, 313, 896, 2227, 2370, 593, 2012, 496, 1317, 1682, 3121, 1496, 2064, 3080, 2667, 3279, 2849, 3030, 72, 310, 306, 448, 915, 996, 3001, 274, 2602, 2619, 2007, 2224, 2156, 2270, 2490, 2605, 1538, 2863, 3006, 1335, 21, 2520, 2934, 1995, 2988, 2502, 2712, 2787, 504, 2623, 2720, 1190, 2419, 503, 434, 436, 882, 1177, 1271, 1434, 2413, 432, 564, 708, 330, 2880, 240, 354, 452, 561, 631, 575, 953, 1369, 1512, 1690, 1663, 2877, 2677, 757, 165, 3125, 2272, 3062, 1140, 1803, 2274, 3203, 384, 2820, 1003, 1483, 1868, 2810, 2628, 627, 831, 1870, 2266]. edizioni [3120, 340, 74, 181, 2685, 2497, 435, 871, 870, 560, 1011]. Edizionitalia [1415, 1787]. Edmé [1232]. Edmond [3806, 1844]. Edoardo [2216, 2113, 2232]. Edocere [446]. Edreira [2068]. Eds [414, 724, 279, 1515, 1113, 1857, 283, 685, 1005, 1087, 1204, 1562, 2601, 2986, 3041, 2459, 3477, 2843, 2717, 1393, 1643, 2933, 2882, 2936, 3059, 2707, 2805, 2093, 2750, 3025, 3510, 549, 1991, 2972, 2798, 3082, 2826, 2551, 2671, 2791, 194, 648, 1445, 2553, 2059, 2161, 2216, 2727, 2589, 3076, 2834, 2755, 2962, 2978, 2949, 2662, 2831, 2657, 2895, 3061, 2550, 2941, 2888, 3415, 2907, 2719, 2819, 2934, 3070, 385, 3288, 2712, 3253, 2783, 225, 2268, 2781, 2656, 3412, 3033, 2134, 3091, 2977, 683, 1407, 2004, 2945, 2754, 2938, 2965, 1934, 1935, 3018, 2741, 2801, 2743]. eds [2884, 616, 2880, 452, 862, 1570, 1563, 2078, 2213, 2423, 2785, 2745, 2886, 2731, 2583, 2952, 1249, 2734, 1694, 2366, 2412, 2559, 2866, 3226, 2428, 2803, 2939, 2845, 3410, 860, 1198, 2967, 329, 2274, 2856, 2911, 3092, 3527, 2498, 2954, 2622, 2613, 1845, 2571, 2692, 3013, 3572]. eds. [3570, 3631, 3674, 3694, 3653, 3637, 3611, 3675, 3742, 3671, 3718, 3695, 3693, 3687, 3737, 3634, 3733, 3741, 3650]. Eduard [1203]. Educare [1115]. éducatif [1486]. Education [699, 1277, 1445, 3594, 3852, 3853, 3854, 3170, 2767, 3694]. Educational [1486]. Edvige [75]. Edward [857, 2873, 244, 1798, 2953, 80, 2799]. Edwin [843, 1085]. Eerdmans [624]. eerste [541]. Effect [1657, 2971]. effects [1573]. efter [1146]. Efthymios [3205, 3267]. Egbert [28]. egg [4023]. Eggs [3520]. Egidio [1279, 2946, 595]. Egitto [869]. Egmond [2884, 3411]. Egnatio [263]. Egnazio [2392]. Egypt [707, 3157]. Egypte [707, 1432]. Egyptian [829, 2076, 4053, 94]. Ehrung [507]. Eigenschaften [379]. Eight [916]. Eighteenth [517, 544, 1110, 1725, 1818, 2312, 2496, 2464, 3739, 2906, 180, 1670, 3714, 2888, 2781, 2900, 3236, 1652, 778, 2423, 3690, 4004, 2901, 1331, 1845, 2350, 394, 3841, 3849, 3879, 3894, 3105, 3196, 3966, 1039, 2294, 4034, 3729]. Eighteenth-Century [1110, 1725, 1818, 2312, 2496, 3739, 2906, 2900, 1652, 2423, 2901, 2464, 1845, 3841, 3849, 3879, 3105, 3196, 4034, 3729]. Eileen [1986]. Eilhard [961]. Einarson [945]. Einaudi [39, 537, 617, 1427, 1528, 2963, 883, 1716, 2693, 1904, 756, 186, 2426, 2384, 1349, 1330, 1420, 1642, 1644, 346, 2748, 1834]. Eindhoven [2607]. einem [420, 3905, 2576, 362, 565]. Einf [379]. Einfluß [547]. Einführung [201]. eingeladen [4078]. eingeleitet [362]. Einheit [1381]. Einleitung [307, 571, 256]. einsegnamento [189]. Einselne [994]. Einstein [2852, 811, 1993, 1802, 2674, 2851, 3786, 2061, 3882, 3889, 2977, 3908, 2207, 2258, 3125, 3974, 3126, 3998, 4089, 2763]. Einsteinian [1976]. Einsteiniano [1976]. Einthoven [2556]. Einzelschriften [173, 217]. Eire [992]. Eirene [13]. Ekeland [2800]. Ekmeleddin [2749]. Ekström [699]. Elaboration [2259, 2403]. Elaine [3493]. Elbert [1202]. Elder [2706, 1408]. Eldorado [2872]. Electa [1927, 879, 20, 635, 1016, 1172, 1419, 1510, 1926, 165, 1331, 1530]. Electa/Editori [635]. Electric [2435, 526, 1825, 1573, 2498]. Electrical [1006, 856, 2542, 1238, 2080, 2381, 1732, 1913, 1733]. électricité [260]. Electricity [2168, 2239, 2512, 1067, 2156, 190, 2844, 971, 3955, 4022, 1913, 2498]. electrodynamics [1542]. Electromagnetic [2478]. Electronic [1880]. Elektrizität [2532]. Elektrodynamik [2859]. Elektroenzephalographie [2736]. élémentaire [1090]. Elemente [3130]. Elementi [1903, 2246, 1870]. Elementorum [2712]. Elements [1822, 376, 3796, 1654, 2462, 32]. Elena [35, 3738, 1252, 3077, 1515, 563, 1056, 1435, 3285]. elenco [1748]. elettrica [971]. Elettrici [1824]. elettricit [2904]. Elettricità [2168, 2239, 75, 3955, 1423]. Elettrolitica [979]. elettrometri [1258]. Elettrotecnica [2622]. Élie [2652]. Elin [347]. Elio [1082, 1615, 926]. Eliot [418]. Elisabeth [1657, 2854, 1245]. Elisabetta [2949, 3718, 780, 1430]. elite [3874]. Elitegehirnforschung [3874, 2673]. élites [3021]. elixir [3808]. Elizabeth [3694, 1769]. Elizabethan [2948, 1872, 2830, 3745, 2310, 4013]. Elizabetta [2787]. Elkins [2970]. Ellegård [927]. Ellen [2707, 3971, 2378, 2724]. Ellenberger [675]. Ellenistica [2264, 1785]. Elliot [2678]. Elliott [1370, 3345]. Ellipses [3251]. Ellissi [792, 1462]. Elly [3325]. Éloge [530]. Elogio [1728]. Elsheikh [1434]. Elusive [2730]. Elwert [64]. Emanuel [156, 3267, 1988, 725, 1611]. Emanuela [519, 199, 1519, 1429]. Emanuele [2415, 1851, 2049, 2449, 2659]. Ematopoietica [1551]. Embalming [3216, 3148]. Emblematic [3349]. Emblembuch [93]. Embodiments [2093]. Embriologia [1774]. Embroidering [3539]. Embryo [3055]. Embryos [3400]. Emergence [2947, 3999, 3430, 2762]. emergenti [2116, 304]. emigration [3413]. Emil [201]. Emile [1681, 3703, 2548]. Emilia [2002, 2148, 2363, 3039, 2368]. Emilia-Romagna-Grafis [2148]. Emiliae [180]. Emiliani [2826]. Emiliano [1028]. Emilie [184, 3894, 3894, 3331]. Emilio [1807, 700, 776, 1257, 3088]. Emily [3316]. Eminent [885]. Emma [3510]. Emmanuel [1495, 1727]. Emotions [2655, 3405]. Emozione [1506]. Empereur [620]. Emperor [694]. Empire [1340, 2527, 2824, 2920, 3692, 3456, 2749, 689, 3261, 3625, 3779, 3068, 3198, 3216, 2845]. Empires [1072]. Empirical [3552, 1381]. Empiricism [2656]. Empiricists [2771]. Empirischen [1381]. empirisme [3865]. Emploi [210]. Empoli [1257]. empreinte [3448]. Empresa [3618]. en-counters [1654]. Enchantment [2560]. Encheiridion [3970, 3298]. Enciclopaedica [337]. Enciclopedia [299, 257, 862, 2319]. enciclopedico [912]. Enclosing [3678]. Encounter [3349, 624]. encounters [3790, 3794]. Encyclopaedica [299]. Encyclopedia [2227, 3059, 1786, 4090]. Encyclopédies [1804]. End [579, 2538, 3072, 2970, 4010]. Endersby [3032]. endings [2678]. Endocasts [1154, 4009]. Endocrinology [1179]. Endoscopy [3146]. Energia [2622]. Energy [689, 1274]. enfermedad [1855]. Eng [235]. Engelbert [256]. Engelhardt [557, 1731]. Engelmann [3572]. Engine [2604]. Engineer [3703, 3374, 3061, 3224]. Engineering [2664, 3479]. Engineers [1006, 1238, 2080, 2381, 1732, 1913, 1733, 3452, 303, 1966, 4049, 4064]. England [342, 438, 580, 1404, 2586, 649, 2883, 385, 2633, 3131, 2071, 1877, 1872, 3114, 1798, 1725, 2056, 3745, 3204, 3841, 3869, 2945, 2582, 1213, 3493]. Englightenment [1143]. Englische [1212]. English [1085, 192, 2156, 3298, 420, 341, 2139, 2940, 2525, 3638, 1167, 1419, 3434, 2609, 3044, 3612, 892, 539, 378, 3669, 3761, 985, 1212, 3449, 3868, 3888, 2087, 3334]. Engraved [3049]. Enhancement [2193]. Enigk [361]. Enigma [3529, 3505]. enjeux [1543]. enlarged [1448]. Enlightened [3738, 2534]. Enlightenment [258, 759, 724, 923, 1110, 1087, 1146, 1536, 1910, 2601, 3117, 2435, 3425, 2526, 652, 1568, 3260, 3674, 2664, 3222, 1147, 385, 341, 3265, 3020, 2661, 2842, 2778, 2599, 1144, 2723, 3459, 2844, 3979, 4001, 4011, 4030, 1005, 1763, 2160, 2424, 2463, 2953, 3767, 3779, 3834, 2893, 3074, 2801, 2731, 3118, 2784, 4086, 3261, 3347]. Enlightenment.Collectors [2961]. Enlightment [169, 760]. Enough [1814]. enquêtes [1793]. Enrica [2226]. Enrico [2328, 31, 1077, 2279, 3408, 2965, 1173, 171, 858, 941, 2062, 1303, 1759, 2270, 1302, 581, 631, 1284, 1456, 2319, 1321, 2745, 1265, 1140, 2507]. Enrique [678]. Enriques [3120, 2685, 1658, 2380, 4083, 3838, 2306, 4083, 2763]. Enriquese [3120]. Ensayo [2494, 3618]. enseignement [2973]. Entanglements [3531]. entangling [3244]. Ente [185]. enterprise [4053]. Enterprises [2719]. Entralgo [882]. Entre [2048, 2566, 728, 3742, 3069, 3197, 2578]. entreprise [4053, 2118]. Entstehung [415, 3999]. entwicklungsgeschichtliche [1989]. Environment [3543, 3818, 2553]. Environmental [1738, 3512]. Environnement [2762]. Envisioning [3316, 3993]. Epicureanism [3063]. Epicuri [1896]. epidemia [566]. epidemie [1181]. Epikureismus [798]. epilessia [1867]. Epilogue [420, 3275]. Episode [492]. Episodes [2498]. Episodio [211]. Epistemic [3282, 4015, 4021]. epistemica [1995, 4021]. Epistemologia [909, 374, 1693]. Epistemologica [1897]. Epistemological [35, 1897]. epistemologiche [1148]. epistemológicos [35]. Epistemology [3328, 2640, 1498, 4018, 3038, 3486]. epistémonika [1930]. Epistolari [2450, 658]. Epistolario [922]. Epistolarum [799]. epistole [1251]. Epistolography [3328, 4018]. Eponyms [3464]. époque [456, 40, 3010]. époques [3527]. epos [2135]. Eppur [2171]. épreuve [3075]. Epulario [1000]. equal-arm [1163]. Equations [3532]. Equatorial [3308]. equestre [3009]. equipment [2556]. Era [1005, 1169, 3886, 2417]. Eradication [499]. Erasistrati [607]. Erasmo [2159, 577, 2415]. Erasmus [1791, 1792]. Erbari [435, 83]. Erbario [2886]. eredità [758, 926]. eresia [1368]. Eresus [2939, 1225]. Erfahrung [1202]. Erffordiensis [251]. Erfurt [251]. Erfurter [251]. Ergänzungsband [1328]. Ergon [1918]. Ergon-Verlag [1918]. Erhard [3064, 3283]. Eric [3415, 1799, 3124, 1428, 218, 3362]. Erich [1412, 251]. Eriksson [1365]. Erinnerung [1731]. Erkenntnis [1857, 1381]. Erlangen [3116, 1405]. erlebte [1731]. Erluterungen [226]. Erma [1135]. Ermanno [1926]. Ermete [1280]. ermetici [2644]. Ernan [2774]. Ernest [1683, 547]. Ernesto [2616, 1011, 982]. Ernst [303, 698, 2144, 2908, 2202, 201, 1838, 3332, 111, 131, 420, 758]. Ernsting [1661]. Eröffnung [173]. Eros [3502]. Erratic [3052]. Error [3504]. errore [1506]. errori [1462]. Erster [420, 379, 1001]. Ertraege [188]. érudit [2492, 3967]. Erudition [2656]. Eruditorum [681]. Érudits [582, 3803]. Erudizione [1823, 1973]. Erwin [1579]. Esame [988]. Esami [1189]. esattezza [3652]. Escatologia [2473]. Escuela [1227, 1226, 1860, 4017].

Esculapio [332]. Esedra [2588]. esempio [1729]. Esercizio [1369]. esfera [3167]. Esguerra [2920]. esorcismo [3178]. Esortazione [1174]. esoterismo [765]. esotiche [1268]. España [1515, 35, 965, 2808]. español [3327]. Española [889, 1860, 2097, 4017]. especially [2228, 3902]. Esperia [1064]. Esperiences [959]. Esperienza [3486]. Esperienze [1953, 2024, 2726]. Esperimenti [906]. Esperto [264]. Espiritismo [3427]. Esplorando [792]. Esplorazioni [1158, 3022]. Esposizione [2743]. esprit [1460, 2428, 2856, 4088]. Esq [1408]. essä [1146]. Essai [2528, 2757, 3026, 2528, 1232]. essais [1413]. Essay [2586, 1146, 3864]. Essays [227, 164, 520, 439, 1562, 4041, 1393, 3442, 511, 1263, 2216, 2755, 3630, 613, 225, 848, 1563, 229, 3044, 230, 1198, 1857, 3762, 4057, 4079, 1267, 2982, 275]. esse [814, 522]. Essen [1661]. essentiae [567]. Essential [3011]. Essere [897, 44]. Est [370]. Established [2971]. Establishing [615]. establishment [1406]. Estácio [1743, 1744]. Estate [642]. esteem [1989]. Estense [2129, 1011]. estensi [3813, 1011, 3813]. Ester [1538, 1598]. Estes [1097]. esteta [2916]. estetica [2142]. Esther [2072]. Estonian [938]. estr [233, 2112]. Estrange [1393, 3237]. estratto [2009]. estrelas [3833, 1743]. estructura [494, 618]. Estudi [353]. estudio [1226, 3427]. Estudios [889, 1227, 3759, 1226, 43, 1855, 3867, 1849, 2097]. etä [555, 794, 3672, 2536, 958, 3937, 3961, 3655, 1318, 3259]. établi [754, 893, 2539]. Eterna [3264]. Eternal [3479]. Eternity [1686]. Ethan [3509]. Ether [1976]. Ethics [1225, 1275, 2978, 1275]. Ethik [248]. éthique [3197]. Ethnicity [3377]. ethnographie [2067, 4039]. ethnography [4039, 3965]. Ethnological [3579]. ethnosciences [1170, 1170]. Etica [883, 1864]. Etienne [2898, 762]. Etimologico [1173]. Etnografi [2675]. etnografia [2206, 3965]. Etnologia [931]. étonnant [1727]. Etrusca [640]. ETS [3285, 2974, 2272, 1134]. Ettore [1364, 2415, 3080]. Étude [2267]. Études [287, 305, 574, 1442, 1444, 1984, 1080, 308, 3058, 3280, 2309, 2229, 4063, 1669]. ètudiant [1984]. Étudier [295, 3876]. Euclid [2418, 376, 3840, 3273, 4002, 4027]. Euclide [3064, 4027]. Euclidean [2215, 3768]. Euclides [1321]. euclidienne [3768, 2215]. Eudes [912]. Eufrasia [3412]. Eugen [288]. Eugene [1246, 3701]. Eugenetica [2616, 2738]. Eugeni [494, 618]. Eugenia [2430]. Eugenics [972]. Eugenio [612, 718, 2215, 752, 862, 3768, 250, 1157, 1230, 750]. Euklid [2647]. Euler [1492, 2905, 2041, 1681]. EUR [3541, 3526, 3545, 3544, 3530, 3546, 3525, 3528, 3547, 2196]. Euro [2150, 2206, 2266]. Eurobase [41]. Euroclio [2231]. Euroma [1343, 1666]. Europa [2590, 727, 861, 280, 2097, 3021]. europäischen [276]. Europalia [855]. Europe [491, 1110, 1234, 2228, 847, 1991, 192, 305, 1067, 2059, 3070, 2652, 2832, 1935, 2375, 1570, 3398, 3528, 1720, 1999, 2846, 3175, 4051, 1277, 3522, 3552, 2589, 2514, 2962, 2827, 3850, 2656, 2048, 2566, 3895, 3898, 4075, 3902, 3407, 2775, 2489, 2994, 3250, 3170, 3543, 3441, 3455, 3711]. europea [1926, 4014, 3120, 2558, 1230]. European [1536, 2160, 3156, 857, 1515, 1494, 1063, 2838, 547, 654, 850, 1944, 1812, 2077, 2732, 993, 3746, 3755, 2193, 3906, 4080, 3957, 1230, 3474, 4014]. européen [444]. Europëens [728]. Europennes [2231]. Europeo [310, 306, 1157, 1388, 2571, 1488, 2077, 1061]. Europischer [2785]. Eva [3279]. Evaluation [645]. Evan [3734, 1005]. evangelical [624]. Evangelista [2368, 1000, 1469]. Evans [1103, 3076, 2801]. Eve [3369, 541]. Evelyn [3241]. Evemarie [1905]. Even [3668]. Even-Ezra [3668]. Events [2636]. Everett [1980]. Evers [156]. Every [3302]. Everyday [3493]. Evidence [899, 3160, 1703, 1468, 3564]. Evo [27, 84, 1810]. Evolution [2793, 1852, 3084, 1468, 3400, 191, 927, 1122, 2966, 3958, 2994, 1914, 3796, 3799, 719, 3805, 3827, 848, 4079, 1114, 3973, 1654, 3268, 3837, 3805]. Evolution-Creation [2793]. Evolution. [3076]. Evolutionary [418, 1847, 3554, 3102, 419, 624, 3975, 1989]. Evolutionstheorie [2979]. Evoluzione [2232, 1244, 719, 758, 1114, 3973]. evoluzionismo [3799, 3285]. Evolving [3091, 3635, 3480]. ex [3064, 355, 4027]. ex-machina [355]. exact [1198]. exacten [23]. exactes [2757]. Exakten [376]. Examination [36, 1217]. Example [2485, 2396]. Examples [3216]. Exartim [1256]. Excellence [2545, 3170]. excellent [1271]. Exception [3219, 3102]. Exchange [2944, 2228, 3902]. excited [2156, 4022]. exemplar [1744]. Exemple [2396, 1187]. Exhibit [2724]. Exhibiting [1913]. Exhibition [2338, 2311, 3717, 1072, 1662, 2265, 3917, 4090, 1331, 3579]. Exhibitions [3472, 3196]. Exile [3201]. Existence [2388, 1308]. existing [3905]. Exotic [3494]. expand [4012, 1203]. Expanding [439, 3086]. Expedición [2727, 686]. Expedition [1408, 3261, 3451, 4053, 4053, 816]. Expeditions [2802, 3779]. ’Experience [2869, 615, 642, 1723]. experienced [1731]. Experiences [3033]. Experiencing [3271, 2824]. Experiential [3444]. Experiment [919, 2828, 2196, 26, 4078, 1753, 3921, 1800]. Experimental [168, 1067, 3462, 919, 2718, 835, 2069, 3375, 2081, 2626, 2404]. Experimentation [3149, 2978, 2338]. Experimente [2859]. Experiments [2424, 3200, 2594, 3288, 2732, 1069, 3834, 4069]. Experten [364]. Expertise [3431]. explain [371]. explained [2418]. explanation [1499]. Explanatory [2568]. Exploitation [734]. Exploration [15, 841, 1566, 1334, 3429, 3887]. Explorations [1852, 2920, 1683]. Explorative [2859]. Exploring [3151, 2505]. Explosives [2981]. Exposed [2794]. expositionesque [1687]. expressing [4028]. expressioris [4028]. expsicion [45]. Extant [1448]. Extase [1293]. extended [1167]. external [60]. Extinction [3602]. extracted [2009]. Extraterrestrial [723]. Eye [1703, 2371, 1918, 3246, 3855, 3411]. Eye-Witnesses [1703]. Eyes [3053]. Eyewitness [2794]. Eystettensis [1518]. Ezio [1317, 75]. Ezra [3133, 3668].

F [224, 852, 1393, 726, 532, 3475, 234, 260, 2909, 252, 1279, 1188, 1445, 2215, 2229, 3115, 1152, 1335, 225, 788, 1099, 1436, 2006, 2152, 344, 730, 701, 337, 2584, 255, 2576, 163, 921, 3444, 1455, 2465, 2294, 840, 1615, 2752, 1989, 831, 178, 609, 1668, 1667, 1647]. F. [1407, 2004, 958, 499, 1647, 806, 1034]. F.lli [1354]. F.R.S [1408]. f.t [2163, 2234, 2386, 1190, 1746, 1836, 424, 1766]. F/18 [2229]. fī [1227, 1226]. Fa [1710, 559]. Fabbri [429, 2221, 687, 3203]. fabbrica [3801, 1644]. Fabbroni [619, 455]. Fabiani [2114]. Fabio [1499, 2156, 3652, 2565, 1137, 3121, 863, 917, 88, 2886]. fable [656]. fables [1080]. Fabre [1138]. Fabri [2420, 921, 3808]. Fabricius [1868]. Fabrique [2085, 3127, 2952]. Fabrizia [1352]. Fabrizio [2163, 2234, 3298, 2213, 2951, 1330, 3315, 3673, 3259, 546, 2346, 2619, 2949, 633, 1919, 2213]. Fabulous [2922]. Fac [868, 2528]. fac-simile [2528]. Faccia [740]. face [3769, 2273]. Faces [654, 3957, 1002, 850]. Facets [2873]. Fachwissenschaft [1151]. Facing [3626]. Facini [281, 1045]. Facolt [2467, 2620, 250, 2468]. Facoltà [75, 1516, 1128, 72]. facsimile [2525]. Facsimiles [2957]. Fact [1457]. factory [3801]. Facts [1449]. Faculté [1984, 1224, 1444]. Faculty [68]. Faenza [951]. Faes [720]. Failed [3452]. Faille [2833]. Fairbairn [1580]. Faire [1305, 1475]. Fairs [3369]. Faith [167]. Faiths [2998]. Falco [2580]. Falconi [2442]. Fall [2595, 3617, 3295]. Fallacy [3384]. Falsi [1462]. Famiglia [2174, 2773]. familiarum [799]. familières [893]. Families [959, 3425, 34]. Family [3771, 3332]. fantasia [1948, 2019]. fantasmagoria [684]. fantasmi [1965, 2036]. Fantastic [3151]. fantastica [684]. Fantini [1963, 2034]. Fantoli [1790]. Fara [1725, 2527, 561, 562, 3499]. Faracovi [2380, 3259]. Faraday [1006, 847, 2542, 1238, 2080, 1732, 2859, 1733, 4051, 3860, 526, 4073, 4074]. Farben [3037]. Farbenlehre [1053]. Farber [164]. Farbtafeln [1797]. Farbzentrenforschung [1153]. fare [1746]. Farinelli [2605]. farmaceutica [1354]. Farmacia [277]. farmaco [1030]. farmacologia [435]. farmacologiche [1269]. Farmers [3359]. Farnese [233, 281]. Fasc [172, 814, 2524, 1273, 522, 782, 1871]. Fascicule [368, 2214]. fascination [2265]. Fascism [3413, 3851]. fascismo [3851, 2208, 2275]. Fascist [3717]. fascista [2670]. Fashion [3715]. fast [543]. Fasting [2364]. Faszination [2265]. Faszinationsgeschichte [2854]. Fate [511]. Fatebenefratelli [953]. Father [348]. Fathers [3267, 3947]. fatica [2331, 2282]. Fatio [3490]. fatto [1955, 2026]. Fattori [1388]. Faubourg [1727]. Fauna [2612]. Fausti [2581]. Fausto [1437, 1259, 2112]. Favalia [1352]. Favino [3545, 3396]. favola [656]. favoloso [1674]. Favours [2496]. Favrholdt [1008]. Fay [3363]. Fayard [548, 1432, 1232]. fazione [81]. Fazzi [776, 1768]. FB [1734]. fear [4010]. Febbraio [2354, 1373, 83]. febbre [1483]. febrero [1860]. February [2354]. fece [2430, 4010]. fécondation [2015]. fede [688, 655]. Federale [720]. Federica [3545, 2428, 3396]. Federici [1061]. Federico [299, 1952, 2023, 4043, 2522, 311, 2941, 3345, 1706, 832, 3492, 1540, 1112, 2713, 827]. Federicum [1138]. Federigo [1658, 3120, 3838, 2380, 4083, 2763]. Feel [3537]. Feeling [1457]. Feingold [2600, 803]. Feinmechanik [234]. Felice [2420, 1954, 2025, 1712, 2931, 2698, 2697, 3808, 2323]. Feliciano [1663]. Felipe [2101]. Felix [152, 307, 3695, 379, 1989, 3390, 2845]. Fellowship [2686]. Felsch [2853]. Feltre [2001]. Feltrinelli [195, 80, 373, 372]. Female [2377, 3722, 2482, 3943]. femmes [2973, 3395, 3742]. femminile [178]. femminili [3040, 3820]. femminismo [2104]. Fend [830]. Feng [599]. Fenner [499]. Fenomeni [3103, 661]. Fenomenologia [550]. fenomenologo [1995, 4021]. fenómenos [3427]. Feraboli [1280]. Ferber [740, 739]. Ferdinand [248, 3205, 3205, 2493]. Ferdinandea [2332]. Ferdinandea. [2283]. Ferdinando [2166, 2237, 2329, 569, 2667, 1494, 2280, 2117, 1615]. ferenza [3179]. Fergola [2449]. Fergus [2161]. Fermat [3568]. Fermentation. [2897]. Fermi [2319, 2669]. Fermilab [4054, 4055]. Fernanda [1817, 178]. Fernando [2559, 3446]. Ferno [899]. Ferrand [921]. Ferrante [2226]. Ferrara [873, 1189, 2461, 1190, 727, 1136, 1241, 2002, 2460, 2461]. Ferrara-Poggio [1241]. Ferrara-Roma [1189]. Ferraresi [2789]. Ferrari [1962, 2033, 519, 199, 779]. Ferraris [2479, 1889]. Ferri [1840, 1267, 1269, 3439]. Ferriani [1105]. Ferrier [1917, 3933, 1467, 3375]. Ferriolo [1811]. Ferro [643, 788]. Ferroni [2592]. ferrovia [2332, 2283]. Ferruccio [3061, 407, 558]. fertilization [2015]. fervor [3433]. Fery [3423, 3423]. Festa [2946]. Festina [1173]. Festkörperphysik [1153]. Festschrift [234, 23]. Festschriften [1797, 3977]. Festschrifts [3977]. Festugière [2644]. feu [1681]. Feuer [811]. Feyerabend [2132]. FF [416, 1878, 1875, 2015, 2013, 920, 1944, 748, 1981]. Ff. [681]. Fi [2227]. Fibonacci [3408, 2922, 3593]. Fibre [2870]. Fichte [1288]. Ficino [799]. Fickel [105]. Fiction [3053, 418]. Field [2302, 2258, 3434, 1999]. Fields [3678]. Fiera [185]. Fiesc [1681]. Fiesole [2227]. Fifteenth [3431]. Fifteenth-Century [3431]. fig [593, 708, 711, 871, 502, 563, 556, 575, 559, 701, 970, 916, 953, 859, 1015, 1056, 1161, 1258, 1256, 1259, 1163, 1158, 1159, 1193, 1746, 558, 840, 1283, 713, 870, 875, 1100, 1029, 1186, 1184, 1274, 1182, 518, 562, 540, 560, 655, 581, 631, 633, 590, 635, 762, 705, 700, 776, 780, 779, 702, 778, 844, 858, 878, 893, 939, 835, 862, 861, 963, 937, 864, 941, 854, 938, 1062, 1082, 1011, 1016, 1013, 1072, 1012, 1083, 1084, 1107, 992, 998, 1170, 1257, 1284, 1240, 1255, 1136, 1158, 1166, 1167, 1241, 1185]. fig [1195, 1165, 1172, 1342, 1000, 1150, 1389, 1719, 410]. fig. [782, 1014]. figg [2163, 2234, 2387, 2389, 1656, 1713, 1654, 952]. figli [1092]. figlie [897]. Figlio [2081, 2313]. figs [829]. Figura [2326, 2277]. Figural [283]. figurative [2142, 958]. Figure [2742, 2802, 1394, 1135]. Figures [1152, 3363, 422]. Figuring [3376, 3735, 3549]. Filantropismo [1295]. Filare [1631]. Filing [2259]. Filippi [951]. Filippine [2727]. Filippo [1955, 2026, 2163, 2339, 390, 2714, 665, 2234, 643, 900, 930, 1521, 1040, 3246, 384]. Filo [976, 1343, 1540]. filogenia [3618]. Filologia [1831]. Filologiche [1318]. Filopanti [1591]. Filosofi [995, 719, 2507]. ’Filosofia [1036, 515, 1364, 1606, 2467, 898, 1648, 72, 250, 274, 949, 1513, 2276, 2444, 1789, 1207, 50, 162, 3135, 727, 2078, 922, 1344, 1712, 2124, 969, 1459, 2306, 678, 2107, 2062, 4084, 3203, 1233, 1615, 2498, 1900, 764, 896, 1201, 1691, 621, 677, 1314, 74, 85, 448, 715, 716, 719, 915, 2949, 3812, 1917, 3838, 81, 1901, 3904, 1672, 1864, 3925, 1033, 626, 3933, 3949, 2380, 4083, 3961, 3062, 3396, 1034, 1455, 1713, 1988, 2501, 1651, 3259, 2742, 978, 2316, 1061]. filosofiada [61]. filosofica [612, 2721, 2865, 1943, 756, 1330]. filosofiche [1954, 2025, 1364]. Filosofici [1829, 446, 238, 2117, 44, 1455, 1615, 1990, 4043, 181, 991, 1156, 70]. filosofico [2062, 2374]. filosofie [1200]. Filosofo [1207, 1638, 12, 39, 3121, 31, 1111, 3872, 1579, 2373]. Filze [1783]. Fimiani [757]. Fin [210, 3013, 3890, 3980, 3004]. fina [1951, 2022]. final [3798]. finale [3798, 2272]. finde [695]. Finding [2175]. Findlen [2783, 3412, 1319]. Fine [3171, 486, 2552, 716, 718, 1360, 949, 2140, 2430, 879, 1998, 4010, 178, 346]. Finger [2959]. finium [1688]. Fink [2682]. Finmeccanica [552]. Finocchiaro [2885, 1987, 3011]. Finzi [170]. finzione [2167, 2238]. Fiocca [1966, 2037, 3286, 2491, 2614]. Fiona [1818]. Fiore [2522]. Fiorella [1922]. Fiorentina [1973]. Fiorentinagas [704]. fiorentini [576, 453, 2266, 4038]. Fiorentino [252, 4040, 2224, 1430, 2126]. Fiorini [426]. Fiorino [709]. Fire [3433, 3173, 3244, 3635, 3529]. Firenza [585]. Firenze [61, 189, 313, 516, 517, 583, 777, 804, 1539, 1691, 2159, 2488, 2467, 3299, 521, 593, 1239, 1966, 2012, 2386, 2942, 2425, 1784, 3002, 3286, 340, 688, 931, 1488, 1564, 1533, 1817, 2530, 3040, 3071, 619, 1235, 2522, 3255, 2780, 278, 2797, 3009, 2865, 2769, 3015, 31, 252, 250, 309, 447, 576, 720, 717, 716, 887, 873, 991, 1077, 1283, 1157, 1279, 1156, 1360, 1688, 70, 741, 2592, 2270, 1789, 1538, 2863, 3027, 3006, 3209, 1111, 2520, 3127, 2949, 413, 2058, 2531, 2572, 2888, 752, 2934, 2955, 2614, 2905, 2652, 1943, 2590, 874, 881, 2419]. Firenze [2754, 3258, 706, 709, 729, 788, 952, 872, 1028, 1098, 1178, 1287, 1268, 1173, 1598, 1678, 1755, 2006, 2072, 2222, 2373, 14, 2525, 1186, 1523, 1521, 390, 389, 2880, 187, 277, 453, 455, 561, 655, 589, 590, 575, 700, 844, 953, 1014, 1107, 1258, 1284, 1255, 1430, 1354, 1424, 1461, 1437, 1748, 1807, 2574, 1065, 2877, 3088, 2809, 304, 799, 800, 2645, 1323, 1289, 2499, 2608, 3062, 924, 2372, 687, 1068, 2903, 1388, 1270, 1519, 1485, 1728, 1493, 2628, 1000, 2266, 2726, 2960, 2571, 1389, 1035, 1205, 1374, 1463, 1949, 1956, 1957, 1959, 1962, 2020]. Firenze [2027, 2028, 2030, 2033, 1963, 2034, 1951, 2022, 1952, 1953, 1954, 2023, 2024, 2025, 1955, 1960, 2026, 2031, 2037, 1967, 2038, 1948, 1950, 1958, 2019, 2021, 2029, 1947, 1965, 2018, 2036, 1961, 2032, 2164, 2235, 2163, 2168, 2234, 2239, 2233, 2162, 2169, 2240, 2165, 2236, 2167, 2238, 2166, 2237, 2329, 2331, 2337, 2328, 2332, 2326, 2335, 2330, 2333, 2334, 2336, 2339, 2327, 2434, 2387, 2388, 2389, 2435, 2436, 2386, 2354, 594, 1711, 1778, 3138, 2284, 2248, 2283, 2279, 2230, 2697, 2637, 2871, 2288, 1477, 599, 2277]. Firenze [1307, 2287, 391, 47, 90, 261, 458, 657, 731, 813, 973, 1116, 1292, 2282, 2280, 2285, 2286, 2278, 1546, 2281, 1783, 2088, 1252]. Firenze-Milano [2769]. Firm [3771, 3332]. Firmata [1704]. Firpo [1539]. First [3499, 28, 3319, 3717, 3555, 3225, 3599, 1550, 3523, 2987, 36, 3434, 2198, 2443, 734, 3565, 421, 675, 3932]. Fischer [62, 1405, 1004, 2979, 420, 67, 866, 2013, 622, 361, 362, 3650, 690, 565]. Fish [2435, 3353]. Fisher [2771]. Fisica [1902, 1887, 75, 828, 940, 2697, 2476, 2918, 2258, 71, 291, 1258, 1162, 1240, 1241, 1195, 1423, 1532, 1665, 1813, 2005, 1937, 2211, 2271, 1634, 1697, 1262, 978, 89, 1954, 2025, 2168, 2239, 2386, 18, 1243, 1572, 2605, 1084, 1257, 1343, 1424, 2217, 626, 1737, 2062, 2429, 4019, 1679, 1813, 2271]. Fisiche [76, 1516, 595, 3744, 1947, 2018, 2429]. fisici [2603]. Fisico [22, 321, 2871, 1369, 3029, 1343, 2910]. Fisico-Matematiche [22]. Fisico-Matematici [2871]. Fisico-matematico [1369]. Fisiocritici [200, 1251, 2581]. Fisiognomia [2326, 2277]. fisiognomica [3748, 2419, 1394, 1774]. Fisiologia [662, 1956, 2027, 2144, 1917, 874, 3933, 924, 3822]. Fiss [3369]. Fissell [2582]. Fits [1199]. Fitzpatrick [2601]. Fitzroy [2194, 1408]. Fiumi [1461, 3710]. Five [703, 2392]. Flaccovio [1863]. Flage [388]. Flammarion [1804, 3051]. Flannery [3713]. flauto [1348]. Flavia [3457]. Flavio [3061, 1665, 1394]. Fleming [2971]. Flemish [1511]. Flood [3543]. flora [3873, 2530, 2886, 3345]. Floram [2531]. Florence [2169, 2240, 2338, 741, 3717, 1192, 1596, 893, 2820, 1252, 3413, 2169, 2240, 2354, 1711, 2700, 1778, 1212, 194, 2248, 2245, 2698, 34, 1477, 3590, 2701, 2702, 3241, 2345, 1546, 3393, 1783, 2088, 3494, 3597]. Florentine [14, 4038, 4040, 899, 2181, 3684, 1297, 3147]. Florentino [3618]. Florenz’ [1212]. Florey [201]. Florian [2750]. Florianopolis [2198]. Florike [3411, 2884]. Floris [1313]. Flou [834]. Flowers [3389]. flu [4010]. Fludd [579]. fluidi [75, 1068]. Fluids [3690, 3148, 3733]. flute [3956, 1348]. Fluviale [1300]. Focusing [3219]. Foder [1748, 1863]. Foelix [152]. Foes [1366]. Fogel [3687]. folclore [442]. Fold [3605]. Folge [253, 961, 542, 492]. Folger [491]. Folgori [1824]. folio [1413]. folio/essais [1413]. Folk [242]. Folkerts [2055, 23]. folkhemmet [972]. folletos [2114, 3942]. Folli [1158]. Follia [2411, 1335, 952]. followed [4089]. Fond [954]. fondamentali [2321, 639, 564]. Fondamenti [1091, 1697, 3911, 3481, 1803, 831]. Fondation [282, 713]. Fondations [2615]. Fondatore [1352]. Fondazione [1539, 3071, 986, 2965, 355, 1373, 2368, 1737, 709, 1251, 1323, 1140, 1470, 3136, 2168, 2239]. fondements [2820]. fondi [2614, 1696]. Fondo [1691, 2467, 221, 2614, 1350, 929, 1706, 2257, 2129, 1783]. Fons [377]. Fontana [1954, 2025, 2002, 1530, 2698, 2697]. Fonti [1125, 1624, 1772, 1882, 2297, 1046, 1784, 1811, 2177, 2256, 738, 2355, 7, 1829, 1384, 1869, 1472, 1335, 822, 1975, 1477, 325, 270, 96, 401, 237, 1056, 1456, 1435, 929, 555, 734, 1802, 2130, 667, 31, 3174, 165, 2100]. Food [3398, 3528]. Footprints [3917, 2311]. Footsteps [3589]. Foramen [3628]. Forbidden [3651]. Force [1999]. forces [3026]. forest [2009]. Foresta [1016]. Forestale [736]. forestali [1525]. forestali-Ucea [1525]. foreste [3747, 2009]. Forestry [1641]. forests [3747, 2009]. Foreword [227, 440, 1791, 3354, 43, 927]. forg’d [342]. Forgery [3671]. Forgotten [2264, 2646, 624, 845, 3567]. förhistoria [2675]. Forl [2002]. Forlag [1363, 2675, 2500]. Forlati [426]. Forma [3620]. formacin [618]. formación [494]. Formalismo [2996]. Forman [2830]. Formation [1702, 2698, 3650, 2550, 2308, 3880, 4023]. formatione [2205, 4023]. Formato [989]. Formazione [459, 2473, 2363, 4049, 2849, 2315, 2789, 687, 2100, 2960, 2672]. Forme [1884, 996, 1201]. formed [1898, 10, 3944, 3945]. Formes [1543]. Fornaci [563]. Fornander [1019, 1019]. Fornari [2368]. Forneris [2530, 2531, 83]. Forni [1000]. Fors [2496, 2745, 3347]. Forscherleben [1074, 3981]. Forschung [188, 571, 798, 994, 1285, 4061]. Forschungen [1402, 1004, 622, 276, 362, 1086, 1989, 565]. Forschungsergebnisse [188]. Fortenbaugh [1225]. Forti [1436, 2738]. fortificata [562]. Fortino [1850]. Fortis [2665, 171, 904, 1380, 2400, 529]. Forts [571, 994]. fortuna [3839]. Fortunati [1530]. Fortune [3566, 3839]. fortunes [654]. fortunosa [2574]. forza [1948, 2019]. Foschi [2116]. Fossati [1352]. Fossil [2103]. fossili [493]. Fossils [3353, 3681, 3432, 2976, 3215, 3030, 3434]. Fossombroni [171]. fotocopia [1258]. fotocopie [1259]. Fotografiae [1519]. fotografie [1343]. Fotolito [1901]. Foucault [1605]. Foundation [2710, 2168, 2239, 2891, 235, 538, 3205, 3024, 1355, 2724, 1290, 1814]. Foundations [2216, 3139, 1100, 2231, 2247, 1541, 2268, 3911, 1723]. founded [3968]. Fountains [3654]. Four [2392, 3490, 1245, 2802, 4059, 14]. Fourcroy [1476]. Fournier [2554]. fourteen [1164]. Fourteenth [2388, 1686, 3669, 573]. Fourth [642, 63, 1448, 311]. Fox [2354, 2059, 1431, 1845, 3414]. Fr [1343, 497, 1941, 1859]. Fra [2172, 1, 1975, 503, 1467, 2128, 662, 250, 915, 1759, 2156, 2504, 1917, 2934, 1268, 1867, 3925, 453, 3933, 3614, 3961, 1713, 3259]. Fragility [3371]. fragmenta [607]. Frame [511]. Framework [3435]. Francês [2473]. Franca [862, 1525]. francais [1985]. Francaise [2826, 314, 3069, 2396, 1232, 1135, 4043, 2229, 3087, 2214, 2201]. francaises [2214]. France [438, 837, 1578, 496, 584, 846, 1567, 1568, 2538, 260, 1808, 3023, 1187, 3803, 1612, 17, 33, 3692, 314, 385, 1246, 1543, 285, 1878, 1875, 920, 762, 2118, 2060, 3010, 2663, 1073, 582, 2587, 1486, 1534, 175, 3286, 3171, 3788, 3803, 1041, 3824, 3829, 1121, 2995, 3236, 1464, 2901, 3609, 4031, 3425, 3459]. Francesca [1260, 712, 697, 1251, 1437, 1861, 1931, 2149, 2414, 1729, 1267, 72, 1377, 1183, 710, 879, 2115, 1653, 611, 1988, 2057]. Francescanesimo [572]. Franceschelli [2432]. Francesco [894, 965, 4040, 1953, 2024, 2620, 3009, 964, 955, 572, 3397, 3717, 752, 300, 2492, 2773, 3426, 426, 1768, 2112, 2415, 2475, 1061, 3969, 2266, 2726, 4040, 2522, 2144, 3080, 873, 2224, 2172, 2511, 2572, 1782, 2590, 1186, 2112, 2738, 2921, 2458, 3516]. Francese [2536, 2154, 1464]. Franchini [1084]. Francia [496, 2969, 257, 3544, 1869, 3788]. Franciosi [2685]. Francis [167, 679, 918, 1194, 1362, 2586, 3016, 3570, 955, 1789, 1563, 441, 1907, 4020]. Francisco [190, 3446, 371, 648, 2172, 304]. Franco [515, 897, 339, 588, 18, 456, 1368, 221, 1574, 1693, 2116, 2203, 3021, 449, 1283, 1499, 1759, 1820, 2605, 1917, 2461, 1287, 2154, 2275, 2430, 554, 810, 88, 941, 1355, 922, 2716, 66, 1788, 1851, 83, 3710, 957, 2902, 1291, 2141, 2910, 1850, 546, 894, 3772, 1093, 457, 1016, 1355, 2317, 1003, 1233]. francobolli [1026]. Francoforte [2138]. Francois [2318, 2165, 3160, 598, 2236, 675, 1859, 1554, 2423, 1922, 2469, 1734, 3527, 582, 1386]. Francoise [3611, 1875]. Frangsmyr [685, 1005, 1146, 1454]. Frank [1006, 1238, 1164, 761, 1071, 2080, 2751, 1732, 1786]. Frankel [1203]. Franken [1402]. Frankfurt [2979, 1726, 547, 67, 415, 183, 1058, 363, 2146, 2184]. Frankfurt/New [1726]. Frankland [1798]. Franklin [2959]. Frankreich [3186, 3073]. Frans [2725]. Franz [296, 567, 1531, 2055, 3116, 1004, 3123, 23, 1941, 2978, 173, 3288, 2859, 217, 93, 1285, 1396, 362, 1501, 1153, 751, 376, 1069, 609, 3073, 389, 2255, 589]. französischen [3893, 1912, 565]. Fraser [163]. Fratelli [1816, 859]. Fratenita [517]. Frattarolo [1158]. Fraud [3400]. Frazer [766]. freccia [942]. Fred [2549]. fredda [3718, 1552, 2918]. Freddy [1501]. Frederic [15, 1846, 2626]. Frederick [851, 3076]. Frederik [2715]. Fredi [1411]. Free [345, 913]. Freedberg [2371, 2719]. Freedheim [2064]. Fregionaia [876]. Fregionata [789]. Fregonese [2156]. Freiberg [307, 818]. Freiburg [1687, 948]. frenastenici [1574]. French [587, 837, 2067, 222, 282, 675, 1060, 1440, 1403, 2664, 1985, 279, 1879, 2215, 2156, 40, 1149, 760, 2015, 1944, 707, 1275, 1170, 3118, 1912, 1720, 1486, 4039, 4042, 1476, 4043, 3765, 3767, 3768, 3769, 977, 2051, 985, 3778, 3781, 3784, 3559, 3785, 4053, 2085, 3803, 1211, 3805, 1305, 3217, 3829, 4063, 3423, 1121, 4065, 3875, 3876, 3877, 3731, 3893, 3894, 3895, 2396, 4077, 3924, 3930, 3932, 3936, 3540, 3245, 1464, 3448, 2872, 2443, 3992, 4088, 4089, 4000, 2127, 4035, 4037]. Fresco [460, 2360]. Fresquet [1855]. Freud [1613, 1581, 3182, 1246, 67, 1394]. Freundeskreis [64]. Frewer [2978]. Frezzotti [2149]. Friberg [3000]. Friburg [1687]. Friburgense [1687]. Friederich [127]. Friedman [2152, 2583, 2911, 2545]. Friedmann [1203, 4012]. Friedr [259]. Friedrich [62, 1602, 144, 108, 125, 64, 173, 1715, 2915, 1724, 1079, 896, 154, 117, 173, 2859, 1328, 1916]. Friend [2786]. Friends [959, 2371, 1366, 1196, 3521, 3855]. Frigessi [2426]. Frignano [726]. Friis [840]. Fringe [224, 3198]. Frisi [339]. Fritscher [2161, 1069]. Fritz [354, 1141, 1336, 1577, 2061]. Friuli [70, 980]. Frizzoni [1415]. Frlagsaktiebolaget [673]. Frngsmyr [672, 2200, 495]. fro [1654]. Frogs [3955]. Fromman [3283]. Fromman-Holzboog [3283]. Frommann [3064, 343, 344, 1731, 1202, 1328, 2465]. Frommann-Holzboog [3064, 343, 1731, 1202, 1328, 2465]. Front [3585, 3642, 3698]. fronte [812, 2208]. Frontier [1999, 1211]. Frontière [1211]. Frontiers [3414]. Frontispiece [3049]. Frosini [1782]. Frsen [146]. Fruehhumanismus [950]. Frühe [895]. frühen [567, 2228, 966, 3902, 1396, 3968, 2857]. früher [2682, 3857]. Frühparacelsismus [2831]. Fruit [3603]. Fruitful [3790, 3794, 1654]. frumento [170]. Frumer [3478]. Fruton [2897]. fu [3426]. Fuchs [2391]. Fujani [2548]. Fuller [2740]. fulmini [2330, 2281]. Fulvia [1599]. Fumagalli [718]. Function [3371, 3419, 1381]. Functions [3375, 1309]. Fund [1783]. Fundacin [2876, 1860]. Fundacion [965, 1860]. Fundament [961]. Fundamental [2161, 1503]. fundamentals [4072]. Funding [3469]. Funerary [3216, 3722]. Fungi [2719]. Funkenstein [335]. Funktion [1381]. Funktionen [2837]. funzionali [2336, 2287]. funzionario [619, 455]. Funzione [1551]. funzioni [2334, 2285]. fuori [2214, 930, 2004, 638, 1099, 712, 710, 711, 871, 863, 1029]. fürchtet [1726]. furnishings [429]. Furqän [2710]. Furqn [2199]. Fürsten [3344]. Fusione [1552]. Fusoris [2441]. future [1934, 1137]. Futures [3428]. Futurinsieme. [185]. futuro [1113, 1137]. futurocomune [185]. füüsika [938]. Fyfe [3033].

G [301, 303, 491, 440, 414, 837, 965, 1370, 1680, 2067, 2953, 1393, 568, 1387, 1496, 2074, 2155, 1004, 2528, 181, 194, 1156, 1445, 1243, 1620, 1759, 1928, 2066, 2579, 64, 2614, 1575, 2828, 2950, 748, 2134, 2629, 1901, 3716, 788, 1431, 1716, 1867, 1809, 1184, 1913, 344, 291, 452, 1062, 1033, 999, 929, 3226, 13, 766, 1654, 2466, 381, 382, 1142, 1324, 1224, 1225, 1920, 1728]. G. [516, 759, 2389, 2354, 1103, 522, 681, 436, 586, 786, 789, 947, 948, 952, 1675, 1716, 1752, 1865, 1925, 462, 1001, 1013, 1065, 411, 1455, 1733, 2294, 2501]. G.S [1408]. Gāyat [334]. GA [1275]. Gabetti [563]. gabinetes [2068, 3990]. Gabinetto [1162, 1665, 4019, 1258, 1257]. Gabriel [2053, 808, 4032, 1417]. Gabriele [1189, 1136, 165, 814, 2001]. Gabrieli [1689]. Gabriella [277, 2314, 1354, 1730]. Gabrietti [1746]. Gadebusch [2682, 2863, 3130]. Gaetana [2783, 3088]. Gaetano [1186, 453]. Gago [3504]. Gahn [1441, 884]. Gaia [2472]. Gail [959, 2655]. Gal [1399]. Galan [2059, 3675, 2907, 3511]. galassie [1114]. Galatea [3435]. Galatina [2580, 2492]. Galavotti [1857]. Galaxien [1862, 4036]. galaxies [1114, 4036]. Galaxy [1412]. Galbani [2367]. Galeazzi [1178]. Galen [274, 3110, 1133]. Galenic [3650]. Galenica. [3650]. Galeno [887, 1134, 3057, 2220, 1174]. Galera [2494]. Galician [2068, 3990]. Galien [3491]. Galilaeana [4040, 4040, 92]. Galile [2201, 3778]. Galilean [2338, 2, 1699, 2248, 2124, 3540, 3956, 2869, 2260]. Galilée [2422, 4043, 2085, 2652]. galiléen [3540]. Galilei [2653, 312, 3486, 581, 1195, 1451, 1423, 1532, 1399, 1064, 2141, 2943, 1822, 2081, 2247, 2083, 1310, 1119]. Galileiana [2338, 1968, 2260, 2630, 2140, 1321, 3652]. Galileiane [2248, 2170, 3956, 2124, 2289]. Galileiani [1790, 613]. ’Galileiano [596]. ’Galileo [2479, 968, 2142, 1790, 2588, 2371, 1791, 2885, 613, 1036, 3225, 3399, 1398, 2720, 3507, 581, 1195, 1452, 1423, 1532, 1874, 1399, 3525, 1400, 2687, 1986, 1291, 4043, 2167, 2238, 3099, 3156, 1375, 2142, 975, 2653, 3778, 1118, 1699, 1969, 2085, 1880, 2774, 2084, 3248, 2839, 2595, 3846, 2946, 3855, 3858, 3225, 3232, 1207, 524, 1987, 988, 1208, 3258, 3914, 2993, 1451, 3224, 1559, 2171, 2081, 3953, 3685, 2869, 1120, 1697, 2082, 2247, 1117, 1889, 1310, 2487, 2569, 3011, 3203, 1119, 1064, 2141, 1630, 1886, 2292, 3589, 1376, 3719, 1874]. Galileo. [3326]. Galileos [207]. Galileothek [2260, 2260]. Galison [1980, 2412, 2983]. Gallaccini [3258, 3953]. Gallais [2528]. Gallais-Hamonno [2528]. Galli [2616, 38, 3620]. Gallica [1280]. Gallimard [2537, 1413]. Gallo [1255, 94]. Galloni [1599]. Galluzzi [2338, 2436, 1252]. Galluzzo [2863, 3006, 2520, 2934, 2419, 2880, 2877, 3062, 2820, 2628]. Galvani [4052, 2904, 2078, 3955]. Galvanismus [3186]. galvanometridel [1258]. Galzigna [81]. Gamba [581, 858, 2910]. Gambaro [999]. Gambaruto [407, 558]. Gambi [588, 1437]. Game [3229, 2489, 2088].

Games [1852, 2594]. Gans [926]. Ganz [1794]. Gaposchkin [1592]. Gara [1482]. Garbari [863, 917, 2886, 1137]. Garber [3113]. Garbugli [2349]. Garcia [2652, 2101, 965]. García-Diego [965]. Garden [1314, 1518, 2553, 3305, 3704, 1137]. Gardens [1370, 1267]. Gardner [373]. Gareth [1439]. Garfagnana [726]. Garfagnini [555]. Gargam [3395]. Garin [612, 752, 750]. Gario [1658]. Garland [1786]. Garofalo [607]. Garratt [1733]. Garrigó [3489]. Garrod [1222]. Garsault [3230]. Garten [193]. Garum [914]. Garun [3811]. Garwin [2551]. Gary [1106]. Garzanti [1088, 433, 1491, 371]. Garzoni [2722]. Gasca [1516]. Gascoigne [760]. Gaskell [767]. Gaslight [3265, 4011]. Gaspare [1289, 3394]. Gasparo [2592, 3934]. Gassendi [2757, 2771, 2892, 2178, 1896]. Gaston [1995, 2429, 3744, 1639, 2429, 4021]. gastroenterologia [638]. gastroentrologie [638]. Gastronomie [2877]. Gathorne [1945]. Gathorne-Hardy [1945]. Gattei [2132]. Gattinara [2507]. Gaudant [2782]. Gaudenzi [586, 1716]. Gaukroger [2947, 1843]. Gauld [1332]. Gautier [574]. Gauvin [2423]. Gavagna [2712]. Gavin [3038]. Gavinelli [1156]. Gavitt [1192]. Gavroglu [1577, 3243]. Gay [580]. Gayon [1326]. gaze [1918, 3448]. Gazette [2354, 2354]. Gazzaniga [1864]. GB [1857]. Gdansk [3934]. gear [2009]. Geber [997, 3946]. Gedichten [3905]. Geert [2715]. gehalten [173, 276]. Gehirne [3874, 2673]. Geigy [332, 503, 434, 436, 639, 640, 785, 1027, 1026, 500, 564, 708, 870, 875]. geïllustreerd [802]. Geistesgeschichte [29, 535]. Geleerden [802]. Gelehrsamkeit [3530]. gelehrten [364]. Gelehrtenbriefwechsel [2785]. Gelehrter [1916]. geletterd [280]. Gembillo [2062]. Gemelli [2933, 1789, 3027, 2231, 2372]. Geminiano [1705, 3655]. genannt [217]. Gender [544, 1763, 1735, 2835, 2832, 3738, 3633, 440, 4001]. Gendered [3653]. Gendering [743]. Genealogies [2134, 3780, 3787]. Genealogy [2730]. Geneeskunde [63, 1791]. General [2744, 3211, 2980, 2860, 3042, 2680, 2629, 3266, 2505, 3128, 2795, 2621, 2547, 3837, 927, 788, 4004, 2920, 1546]. Generale [1502, 71, 291, 1546, 1864, 2005, 2217]. Generali [585, 2888, 922, 2960, 621, 522]. Generalization [3134]. Generation [2835, 3126, 1571, 3460]. generatione [1359]. Generations [2105, 3900, 811]. Généraux [2267]. generazione [1953, 2024, 811]. genere [3689, 788, 1098, 3547]. Genes [3922, 2382]. Genèse [3004]. Genesi [740, 660]. Genesis [3633, 2230, 2814, 2087]. Genetic [1458, 2671]. Genetica [1890, 421, 3090, 2677]. Genetics [3140, 421, 3573]. Geneva [499, 1510, 3527]. Genève [2843, 2539, 2856, 1567, 1421, 1421]. Genevieve [2267]. geni [1856]. Geniale [3874, 2673]. génie [855]. genio [2157, 729, 3927, 3632, 2503]. Genius [3717, 1337, 1002, 3927, 1089, 2595, 3688]. Gennaio [1619, 572, 2809, 1388, 1137]. Gennaro [339, 1815, 3397]. genom [1019]. Genova [894, 2490, 1096, 354, 562, 861, 1240, 1730, 2271, 999, 203, 2141]. Genovesi [2337]. genre [3331, 3894]. Gent [410, 1342]. Gente [2009].

Gentilcore [3066, 3398]. Gentile [799]. Gentili [1343]. Gentleman [3726, 2186]. Gentlemen [3359]. Genuine [3140]. Genuth [1835]. Genve [198]. géo [3890, 3980]. géo-héliocentrique [3890, 3980]. geocartografia [1324]. Geochemie [1069]. Geochemistry [2161, 4066]. Geodesy [2802]. geodeta [688, 590]. Geoffrey [283, 1683, 1763, 2737, 442, 1334, 2861]. Geofisica [519, 199, 779, 1166, 1167, 1429]. geografi [1663]. Geografia [588, 3939, 1190, 3918, 1856, 1906, 2605, 2593]. geografica [688, 590]. geografiche [1564, 1135]. Geograph [1916]. geographer [787]. Geographia [2574]. Geographica [1838]. Geographical [1408]. Geographie [1838, 4082, 3087]. Geographies [2562]. Geography [4082, 3918, 2805]. géohéliocentrique [3013]. Geologia [493, 2007, 1524, 1939, 818, 3209, 2007]. Geological [349, 2103]. geologiche [2849]. Geologici [739]. géologie [675, 2782]. geologo [2849]. Geology [675, 538, 1717, 2957, 2756, 2087, 3819, 692]. Geomagnetic [1167, 3763, 3885]. Geomagnetici [1301, 3763, 3885, 1745]. geomagnetico [1167]. Geomagnetism [293]. Geomantica [599]. Geometer [3374]. Geometra [1207]. Geometria [2444, 1236, 941, 4028]. geometriadello [2685]. Geometric [3708, 1037]. Geometrica [1037]. géométrie [2199, 3768, 2829, 3568]. Geometries [3519]. géométriques [2199]. Geometry [3522, 206, 2991, 1662, 3646, 3271, 3768, 2215, 4028, 2710]. géomtrie [2573, 2215]. Georg [126, 100, 151, 147, 131, 114, 1608, 2144, 234, 420, 2375, 1489, 338, 565, 303, 155, 128, 754, 363, 3822]. Georg-Agricola-Gesellschaft [1608]. George [549, 418, 518, 833, 969, 808, 945, 1493, 4032, 520, 1288, 214, 1443, 2834, 2838, 3515, 3377, 255, 757, 229, 345, 2433]. George-Louis [518]. Georges [620, 3280, 17, 1734, 4089, 2103, 4063]. Georgia [3059, 1915, 2417]. Georgius [1661, 1627, 1628]. Georgofili [1461, 1766, 2354]. Georgy [2662]. Geoscience [693]. Gerald [833, 2674]. Gerard [1393, 1420, 3237, 839, 1057, 835, 2310]. Gerardo [1663]. Gerbino [3204]. Gerd [3462, 1499]. Geredigeerd [1367]. Gerhard [1391, 123, 787, 1916]. Gericke [23]. Gerl [751]. Germ [3656]. Germagnana [575]. German [253, 303, 2228, 2312, 676, 1608, 1446, 1726, 2156, 420, 1735, 1575, 2061, 1918, 1724, 2912, 787, 2265, 1058, 1731, 362, 3978, 806, 1912, 382, 1989, 2146, 768, 492, 3571, 3753, 3754, 985, 1212, 4061, 1043, 4062, 1977, 3857, 3874, 3893, 2765, 3902, 3905, 4078, 3938, 1042, 3338, 4082, 3968, 3977, 3981, 3998, 3999, 2130, 1381, 54, 4033, 4034, 4036]. Germana [2162, 2202, 2233, 765]. Germania [720, 3103]. Germaniae [1448]. germanico [2315]. Germanistik [32]. Germany [2228, 385, 3288, 3226, 2804, 3902, 2585, 3999, 2587, 2913]. Germs [3638]. Geroulanos [3482]. Gerrit [57]. Gershom [1528]. Gertschaft [1531]. gesammelt [64]. Gesammelte [253]. Geschichte [296, 507, 2228, 2055, 3116, 23, 251, 571, 798, 994, 1684, 2161, 172, 420, 557, 1715, 1735, 3874, 1918, 1724, 2196, 639, 256, 4078, 338, 2673, 1797, 1153, 1086, 1079, 1912, 2466, 376, 1450, 1069, 690, 695, 1905, 1916, 1058, 3938, 361]. Geschichteder [183]. Geschichtsschreibung [188].

Geschichtswissenschaft [3530, 364]. Geschiedenis [63, 1791]. Geschlechterverhältnisse [1735]. Gesellschaft [1608, 966, 1151, 226, 557, 1735, 2196, 4078, 1141, 492]. Gesellschaften [253]. Gesenius [132]. Gestalt [3753, 3754, 1575]. Gestation [3571]. gesuita [2577]. Gesuiti [2387, 1033]. Gesuitica [898]. Gesundheit [193]. Getting [3636]. Geyer [1275]. Geyer-Kordesch [1275]. Geymonat [3121]. Gheverghese [2433]. Ghiberti [1958, 2029]. Ghilarza [355]. Ghulām [735]. Giacardi [2215, 556]. Giacomo [740, 1768, 1244, 1631]. Giacomoni [1201, 3065, 1318]. Giambattista [1966, 2037, 3410, 1352, 1966, 2387, 4064]. Giampaolo [1255]. Giampiero [1330]. Gian [1815, 74, 2631, 2007, 1186, 1794, 1705, 1693]. Giancarlo [1283, 2692]. Gianfranco [912, 811, 1323]. Gianguido [1145]. Gianluigi [3364]. Gianna [2656]. Giannetto [2062]. Gianni [2156, 2383, 862, 2107, 1091, 2313, 1452, 1137, 1389]. Giannichedda [1834]. Giannini [1456]. Giants [913]. giapponese [2197, 500]. Giard [1486, 1486]. Giardini [1267, 607]. Giardino [2270, 1975, 1268, 863, 917, 1488, 1811, 433, 2125]. Giarre [1648]. Gibertoni [1419]. Gibin [1236, 1496, 2602]. Gideon [3442]. Gidlunds [972, 2500, 2675]. Giffard [2679]. Gift [3214]. giganti [913]. Gigli [1869]. Giglio [575]. Gigliola [2149, 2414, 1034, 1260, 1861, 1931]. Giglioni [2203]. Gilbert [444]. Gilberto [2941, 2624]. Giles [595]. Gilles [756, 1639]. Gilles-Gaston [1639]. Gillespie [2936]. Gillian [418]. Gillispie [2538, 2982]. Gilman [177]. Gimma [1142]. ginecologia [2414]. Ginevra [1510]. ginevrina [72]. Gingerich [1193]. Ginnasio [1620, 1928, 1343]. Gino [2580, 2780]. Ginori [668]. Ginori-Venturi [668]. Gio [2461, 1453, 2650]. Gioacchino [320]. Gioco [2088]. Gioia [2077, 2954, 1491]. Giora [3410]. Giordano [309, 789, 876, 578, 2951, 2192, 2992, 1230]. Giorgi [1437]. Giorgia [1748, 1863]. Giorgio [1955, 2026, 2522, 971, 1672, 3693, 861, 1185, 2151, 1851, 1218, 1317, 360, 437, 432, 1372, 1663, 338, 2679, 3062, 1906, 2668, 3329]. Giornale [1599]. Giornali [1235, 180]. giornate [709]. giorni [3747, 4040, 792, 785, 500, 875, 1748, 2009]. Gioseffo [2083]. Giovacchini [3312]. Giovan [2370, 38, 956, 3443, 13, 2370, 215, 3138, 2712, 2714, 1297]. Giovane [1955, 2026, 629, 2849, 72]. Giovanili [2091]. Giovann [1960, 2031]. Giovanna [1962, 2033, 180, 3693]. Giovanni [258, 313, 965, 1956, 2027, 1960, 2031, 2332, 2942, 688, 619, 3541, 1693, 74, 3015, 2670, 1360, 1928, 2216, 3457, 3115, 3497, 2129, 455, 655, 590, 893, 953, 1160, 1456, 2368, 929, 3314, 3396, 1840, 2208, 2274, 2914, 740, 2333, 336, 621, 1145, 2063, 2116, 2797, 3053, 1637, 2216, 2398, 3591, 1782, 1973, 912, 1741, 883, 2741, 876, 2450, 1895, 2047, 2255, 1240, 1415, 1787, 2153, 12, 2136, 1540, 739, 1000]. Giovannini [1438, 277]. Giovanninni [1920]. Giovannozzi [688, 655]. girāsa [1226]. Girard [206]. Girardi [817]. giri [502]. Girl [2424, 3834]. Girolamo [1295, 417, 2716]. Giudice [1499]. giugno [583, 4059, 1101]. Giulia [980, 1011, 1419, 1806, 1163, 1142]. Giuliana [2530, 2531, 83, 680, 2231]. Giuliani [1572]. Giuliano [861, 2078, 2498, 2463, 2558, 697]. Giulio [39, 1954, 2025, 1420, 522, 1551, 3967, 3588, 377, 2492]. Giunta [1267, 1252, 1324]. Giunti [1949, 1956, 1957, 1959, 1962, 2020, 2027, 2028, 2030, 2033, 1963, 2034, 1951, 2022, 1952, 1953, 1954, 2023, 2024, 2025, 1955, 1960, 2026, 2031, 1948, 1950, 1958, 2019, 2021, 2029, 1947, 2018, 1961, 2032, 2329, 2331, 2337, 2328, 2332, 2335, 2330, 2333, 2334, 2336, 2339, 2327, 1235, 2284, 3009, 2769, 2283, 2279, 2288, 706, 2287, 1107, 1284, 2282, 2280, 1729, 2285, 2286, 2278, 2281, 1485]. Giuntini [2266]. Giura [2727]. Giuseppe [2330, 2336, 2327, 1689, 1236, 223, 2313, 2602, 1243, 2224, 1248, 2941, 2139, 2773, 2754, 1523, 1521, 631, 556, 1013, 1240, 1671, 2062, 2278, 412, 3315, 2954, 1269, 3450, 1303, 1496, 911, 2376, 525, 1572, 1777, 1302, 2754, 332, 708, 870, 982, 1039, 2691, 3688]. Giuseppina [893]. Giussani [2965]. Giusti [2328, 1488, 2279, 2270, 3408, 941, 1284, 1321, 1297]. Giusto [1390]. given [3840, 4002, 2418]. Givens [2819]. Giving [3318, 3214]. Gjertsen [219]. Glance [3553]. Glass [761]. glassmaking [3934]. glassware [3948]. Glatigny [2826]. Glaxo [638]. Glennie [3131]. Global [1370, 3543, 2356, 3658, 3599, 3334, 3868, 1540, 2357]. globale [1428]. globalization [3947, 3267]. Globally [3445]. Globes [3450, 160, 980, 184, 3516]. Globi [268, 1298, 1779, 2046]. Globo [2356, 2357]. Globusfreund [233]. glosario [1226]. Glossario [3172]. Glosses [1102]. Glückauf [1661]. GmbH [567, 376, 251]. Gmunden [382]. Gnostics [3448]. gnostiques [3448]. Gnt [1618, 1450]. Gnt-Verlag [1618]. God [3175, 839, 3214]. Goddu [29]. Godula [2576]. Goeppingen [32]. Goeppinger [32]. Goerke [676]. Goethe [1201, 226, 622, 1796, 2567, 403, 2704, 1053, 226]. Goethe-Gesellschaft [226]. Goethes [543]. Gogh [729]. Gohau [2053]. Going [3965]. Goldammer [253, 217]. Golden [2887, 2944, 1540]. Goldgar [2887]. Golfond [871]. Golgi [1951, 2022, 3088]. Goliardica [1801, 1572, 1343, 1666]. Golinski [923]. Golvers [3205, 2493]. Gömez [1849, 1849, 3444]. Gonnelli [313]. Gonzaga [964]. Gonzlez [2876]. Good [293, 2772, 3813, 2015]. Gooday [2737]. Goodman [2671]. Goods [964]. Goodwin [1194]. Gordin [3334]. Gordon [3674, 1944, 892, 1100]. Görel [3092]. Gorge [1154, 4008, 4009]. Gorgonean [3215]. Gori [2320, 2224]. Gori-Savellini [2224]. Gorilla [2966]. Got [3152, 3447]. Gotshalk [2094]. Gottfried [121]. Gotthelf [895, 100]. Gotthold [111]. Göttingen [2750, 2853, 966, 220, 2532, 1764, 542, 2673, 364]. Göttingen. [542]. Göttinger [2736]. Gottlieb [1441, 884, 173, 565]. Gouk [2056, 830, 409, 998]. Gould [349]. Goupil [1090, 834, 1495, 368]. Gouzévitch [2550]. Govanni [12]. governo [389]. Govoni [2672]. Gowan [2737, 3637]. Grabmann [248]. Grabmann-Institutes [248]. grace [3826]. Gracia [30]. Grada [585]. Graeca [3298]. Graecorum [1133, 1174]. Graeme [2737, 654]. Graf [1402]. grafica [3426]. Grafiche [2502, 2773, 426, 1343, 1671, 2650]. grafici [305]. Grafis [2007, 1512, 2148]. Graham [1164, 1071, 3515]. Gramophones [356].

Gramsci [514, 2208]. Granada [1227, 1226, 2138]. Granafei [1369]. Grand [624, 1158, 3201, 3702, 692, 1212]. Grande [3483, 2060, 2916, 315, 3077, 1540, 3829]. Grande-Bretagne [2060, 3829]. Grandezze [595]. Grandi [1469, 1779, 2046, 1287, 942, 1974]. Grandin [2972]. Granduca [1538, 589]. granducale [389]. Granducato [844]. Grandy [2561]. Granger [1639]. Grani [1890]. Grant [2799]. grasshopper [3956]. Grassi [1963, 2034, 2577]. Grassina [2642]. Grattan [1149]. Grattan-Guinness [1149]. Gratton [240, 1114]. Grau [353]. Gravesande [3508, 266]. Gravitational [2739]. Gravity [2739]. gravium [1960, 2031]. gravures [609]. Grayling [387]. Grazi [1034]. grazia [3826, 3279]. Graziano [519, 2934, 199, 779]. Grazie [952, 1255]. Graziella [1061, 2212, 2622]. Great [2227, 3482, 3471, 886, 2751, 1057, 1720, 1540, 406, 1089, 4090, 3829, 3453]. Greatest [2881]. Greci [2011, 792, 1092]. Grecia [1605, 442, 745, 1390, 2135, 2416, 4085]. Greco [3438, 1851, 2692, 2644, 3157, 94, 3277]. Greco-Roman [3438]. grècque [2648]. grecs [1224]. Greece [3465, 4085, 3466, 1051, 1294, 2127]. Greek [1683, 2418, 2459, 2708, 3000, 3059, 3205, 2874, 3649, 2938, 3267, 1225, 1983, 3810, 3886, 3157, 3533, 3947, 2926, 2054, 2417]. Greeks [1092]. Green [2467, 3108]. Greenberg [2864]. Greenblatt [3281]. Greene [848, 4079, 3429]. Greenhouse [1657]. Greening [3512]. Gregorian [2176]. Gregoriana [2176]. Gregorio [65, 1185, 3525]. Gregory [2865, 996, 3001, 293, 3676, 370, 2427]. Greiner [508]. Grell [2962]. Grelon [2550]. Gren [1602]. Grendler [2584]. Grene [419]. Greppi [1190]. Greuter [1298]. Grgoire [1879]. Gribbin [2625]. Grieco [3528]. Grifo [1369]. Grignani [3546]. Grimaudo [3057, 2135]. Grimsby [2683]. Grison [1495, 1727]. Grmek [331, 566, 1753, 1866, 1718, 1223]. Grob [2556]. Groeben [201]. Gronda [180]. Groningen [3109, 753]. Groom [3494]. Gross [2556]. Grosseteste [192, 1685]. Grossherzog [1212]. Grossi [1015, 2129]. Grössing [382]. Grotesque [3199]. Group [1179, 43, 3440]. Growth [192, 2617]. Grubrich [67]. Grubrich-Simitis [67]. Grumbrecht [145]. Grünbaum [1581]. Grundstoffindustrie [497]. Grundzuge [639]. Grüner [1359, 380]. Gruppo [97, 1214, 288]. Gruyter [2822, 2933, 443, 571, 798, 994, 498]. Gualtiero [2414]. guarigione [1093]. guarire [1178, 2100]. Guarnieri [881, 66, 346, 3413]. Guatteri [1352]. Gudrun [1862, 1450]. Guedj [2108]. Guenter [1935, 169]. Guerini [1811, 446, 2742, 2743, 1455]. Guerra [3080, 1099, 1678, 2670, 3718, 880, 712, 2918]. Guerraggio [3031]. Guerre [3483, 3013, 638, 3890, 3980, 2613]. Guerrini [3345, 2590, 2373, 2713]. Guglielmo [1589, 2420, 2614, 1741, 3286, 3808, 3824, 2995]. Guicciardini [3630, 752, 3476]. Guicciardinie [752]. Guida [1201, 45, 3807, 284, 20, 937, 1016, 2110, 2111, 2429, 2150]. Guide [2194, 182, 1984, 3846, 1934, 1933, 1072, 1419, 3434, 3858, 4070, 3807, 1987, 3417]. Guido [2388, 2849, 2224, 504, 407, 71, 1746, 1658, 1252, 1519, 2467, 1974, 18, 2203, 2849, 2070]. Guidobaldo [2346, 815]. Guidoboni [519, 199, 1429]. Guillaume [3160]. Guillerme [1080, 921]. Guinness [1149]. Guintini [704]. Gunnar [972, 1365, 86, 274]. Gunpowder [2981]. Gunter [1684, 698, 1004]. Günther [173, 2265]. Gurst [698]. Gustav [62, 420, 622, 361, 362, 690, 565]. Gustavian [1537]. gustavianska [1537]. Gustavo [932]. Gutas [1225]. Gutenberg [2753]. Guy [540, 2971, 886]. Guyon [2898]. Guyton [1495]. Gymnasium [1928]. Gynaecology [3108]. Gyung [2730].

H [579, 1019, 1853, 2075, 2467, 2715, 2573, 2354, 508, 543, 1727, 3475, 1991, 3369, 2671, 1445, 1687, 1688, 172, 3070, 3108, 1656, 343, 767, 1395, 65, 787, 638, 948, 1001, 88, 406, 1342, 570, 1229, 2643, 280, 382, 3037, 1800, 3571, 1916, 2133]. H. [1313, 280, 376]. H.M.S [1408]. Hôpitaux [496]. Haag [3344]. Haase [571, 798, 994]. Haber [163, 1336]. Habilis [1154, 4009]. Habits [3564]. Habsburg [3552]. Hachette [496]. Hachler [2529]. Hackensack [3000]. Hackett [3011]. Hacking [1242]. Hackmann [1563]. Haddad [3600]. Hadot [2767, 2748]. Haeckel [420, 2765, 2908, 3400]. Haeckels [547]. Haefliger [1177]. Hagelin [1078, 1805, 696, 699]. Hagens [2265]. Hagia [2760]. Hagner [2673]. Hahn [3225, 2537, 3284, 2728]. Hakfoort [1492]. Half [2443, 734, 675, 3949]. Hall [195, 1449, 1587, 1366, 1588, 919, 1537]. Halle [1992]. Haller [221, 1764, 2732, 2205, 1897]. Hallerian [1701]. Halleriana [2529, 2785, 2205, 1701]. Hallers [2529, 2785]. Halleux [2267, 754, 295]. Halley [1844, 3806]. Halton [812]. Hamann [382]. Hambledon [511]. Hambly [427]. Hamburg [296, 307, 966, 3905, 379, 1989, 3905]. Hamende [1737]. Hamesse [1442]. Hamilton [176]. Hamm [1944]. Hammerschmidt [136]. Hammond [2505]. Hamonno [2528]. Hampshire [341, 835]. Hand [3130, 3246, 3513, 3346, 3018, 1342, 2994, 2943]. Handbook [219, 608]. Handbuch [498, 948]. Handling [3681]. Hands [2482]. Hankins [2889]. Hannaford [1837]. Hannah [3651, 3506, 3035]. Hannaway [3081]. Hanns [2784]. Hanover [2867]. Hans [1611, 2938, 1764, 2866, 541, 2544, 961, 3226, 2384, 2967]. Hans-Jörg [2866]. Hans-Werner [961]. Hantaro [2305]. Hapke [1151]. Happened [2987]. Harald [1405, 3137]. Haramundanis [1592]. Harassowitz [1191]. Haraway [1980]. Harbsmeier [1982]. hardback [683, 2427, 1800]. Hardie [2936]. Hardy [1945, 2909]. Harkness [2948]. Harman [1263, 849, 1198]. Harmattan [2318, 1793, 2896, 2735, 2798, 3087, 2973]. Harmonia [819, 2934]. Harmonie [2814, 3203]. Harmonious [3477, 2913]. Harmonisch [2986]. harmony [2123, 3477]. Harmut [1150]. Harold [3599, 454, 614, 2944]. Harper [1571]. Harrassowitz [2837, 276]. Harriet [2873, 939]. Harris [3565]. Harrison [3971, 2378]. Harry [3086, 2846, 2739, 1800]. Hartmut [1285]. Hartwig [104]. Harvard [679, 2102, 2793, 3509, 1576, 3035, 648, 349, 422, 2966, 998, 2844, 347, 2728, 1847, 3528, 2591, 3278, 2846, 945, 2096, 2674]. Harvester [544]. Harvey [1194, 1403, 2004]. Harwood [2804]. Haskell [2686]. Haslam [1818]. Hasok [2681]. Hasse [2308]. Hassenfratz [1727]. hatte [4078]. Haugeland [1459]. Haupt [623]. Hautes [287, 305, 2613]. Haven [587, 1234, 1449, 2056, 3072, 1738, 3204, 3012, 2948, 65, 2944, 2961, 2683, 3038, 2626, 1720, 840]. Haven-London [587]. Hawaii [1729]. Hawkes [1370]. Haworth [990]. Hay [2898]. Haytham [2199, 791, 3292]. Hazen [627]. HCI. [1270]. heads [3905]. Healers [3567, 3594]. Healing [3653, 2514, 3012, 1178, 1935, 3444, 3110]. Health [642, 2755, 3714, 1812, 1934, 499, 3231, 3398, 3567, 768, 3851, 3867, 435, 1849, 2962, 3653]. Hearing [830]. Heart [3149, 4023]. Heath [2671]. heaven [35, 2493]. Heavenly [3303, 406]. Heavens [3806, 2999, 2178, 1844, 2937, 1986, 3060, 3919, 2516]. Heavy [2564, 2031]. Hebrew [3707, 1048, 3731, 1047]. Heckmann [343]. Hedemora [2500, 2675]. Hedin [1264]. Heering [3694, 3288]. Heft [296, 173]. Hegel [1455, 344]. Hegeliana [1455]. hegeliane [926]. hegemony [1659]. Heidelberg [2265, 1074]. Heidelberger [567, 1602, 2915]. Heijenoort [230]. Heilbron [89, 685, 2096, 3326]. Heilkult [706]. Heilkunst [706]. Heinemann [945]. Heiner [1989]. Heinrich [139, 1573, 2198, 110, 193, 433]. Heinz [1141, 676, 2643]. Heinz-Jürgen [1141]. Heirs [2731, 3059]. Heisenberg [2062, 2674, 3882]. Heisters [3116]. Helander [1611]. Helbing [2140]. Held [1819, 2495, 2338, 4040, 2169, 2240, 4045, 4054, 2195, 4058]. Hele [1370]. Helen [3510, 3715, 3037, 2553]. Héléne [548, 1647, 1646]. Helge [533, 2105, 3263]. Heliconius [1611]. Heliocentric [3139, 2812]. héliocentrique [3890, 3980]. Helisaeus [2138]. Hellenic [3205]. Hellénikes [1930]. Hellenistic [2417, 3886, 2264]. Heller [1920]. Hellyer [2585]. Helm [224, 341, 388]. Helmholtz [2567]. Helmholtz. [2690]. Helminthologie [361]. Helmont [2203, 3821, 3864]. Helmut [2471]. Helmuth [23, 2502, 2276]. Helsingfors [673]. Helsinki [673, 2978]. Helvetica [1687]. Helveticae [1687]. Hemel [544, 1342]. Hempstead [544]. Henderson [2161, 3572, 3012, 1596, 499, 1720, 3597]. Henk [14]. Henley [418]. Henri [3368, 541, 1850, 3805, 3848, 1289, 1799, 2428, 4088, 1802, 2186, 541]. Henrich [1731, 1731]. Henrici [150]. Henrietta [3412]. Henriksen [1249]. Henrique [430]. Henry [1587, 2606, 3762, 182, 2185, 1915, 85, 1489]. Hensing [36]. Hentschel [2912]. hepato [638]. hepato-gastroentrologie [638]. Hepper [840]. Her [891]. Heraus [344]. Herausgegeben [567, 698, 307, 1359, 226, 362, 1391, 259, 23, 1191, 1684, 64, 173, 201, 1838, 1141, 1854, 382]. Herbal [3409, 3561]. herbals [435]. Herbarium [2230, 3713, 1270, 2230]. Herbarum [1811]. Hercules [3305]. Herder [2269]. Heredity [2866]. Heresies [2791]. Heritage [77, 2710, 228, 2193, 2891, 209, 3717, 2583, 2724, 1290, 1814, 3990]. Hermann [1333, 1905, 4035]. Hermanos [3618]. Hermaphrodites [2832]. Hermes [1280, 173, 1814, 3166]. Hermeticism [491]. Hermeticum [2644]. Hernandez [816]. Heroes [3083]. Heron [2564]. Herschel [2003, 1862, 4036, 3809, 3836, 4036]. Hertz [1573]. Hertzka [948]. hervorragender [303]. Herzberg [1151]. Herzen [1956, 2027]. Herzlich [300]. Herzog [1191, 276]. Hes [424]. Hess [1141]. Hesse [934]. Hessen [64]. Hessenbruch [2194]. Hessische [64]. Heurkamp [364]. Heusser [172]. heute [296]. Hevelius [2849, 1826]. Hfl [57]. Hg [3130]. Hicks [1877, 3928]. Hidden [3218, 3713]. Hierarchy [592]. Hierholzer [201]. Higgitt [2781]. High [1928, 3827, 4076]. Highfield [942]. Highly [3451]. Hight [1914]. Hilbert [1348]. Hilbertian [3911]. hilbertiana [3911, 3481]. Hildegard [948]. Hildegard-Medizin [948]. Hildesheim [1489, 338, 746, 747]. Hildesheim-Zürich-New [1489]. Hilfsmittel [676, 1448]. Hilger [772, 218, 1057]. Hill [1660, 1082]. Hillenaar [280]. Hills [2899]. Hilversum [1589, 2986]. Himsel [129]. Hines [651]. Hints [3339]. hipnotismo [3427]. Hippocrate [1131]. Hippocrates [1092]. Hippocratic [608]. Hippocratica [446]. Hippocraticum [1131]. Hirai [3671]. Hirnströme. [2736]. Hiro [3671]. hisb [609]. Hispania [2876]. Hispanic [3261, 3779, 3725, 3294]. hispanica [2114, 3941]. Hispano [352]. Hispano-Arabe [352]. Histoire [834, 2051, 243, 198, 2648, 2215, 3611, 3365, 536, 2048, 2566, 566, 872, 1799, 1386, 2422, 2188, 3107, 548, 305, 1859, 2757, 1543, 3876, 748, 2197, 436, 295, 3010, 2767, 3657, 2729, 1734, 582, 541, 2856, 4037, 3571, 675, 2053, 638]. Histoires [2401, 1981]. Historia [1962, 2033, 2639, 966, 43, 2876, 333, 1544, 2114, 1645, 3687, 3019, 430, 2198, 2656, 2119]. historian [3824]. Historic [3111, 2361, 212]. Historica [2418, 235, 1448, 1992, 2337, 2288, 1838]. Historical [592, 805, 1264, 4041, 1393, 2093, 28, 511, 2553, 2216, 3630, 275, 769, 1395, 2134, 2900, 1812, 1935, 3191, 3267, 2993, 916, 1082, 998, 4082, 1308, 2763, 4039, 1659, 3780, 3787, 2747, 1275, 3910, 1184, 4078, 1167, 2068, 3990, 2752, 3772, 3694]. historical-scientific [2068, 3990]. Històrico [353, 2068, 3990]. histórico-científico [3990]. histórico-scientífico [2068]. historicos [2097]. Historie [282]. Histories [2300, 3675, 2928, 1271, 3378, 3445, 2863]. Historiographic [1702, 233]. Historiographical [1313]. Historiography [1106, 3188, 2926, 533]. Historique [280, 2067, 4039, 796]. Historiques [834, 1060, 2894, 584, 1155, 2229]. Historisch [1328]. Historisch-kritische [1328]. Historische [1171, 3109, 571, 2196, 497, 443, 4078, 382]. Historischen [64, 1838, 4082, 201]. History [11, 15, 77, 224, 301, 454, 491, 439, 490, 592, 714, 672, 722, 891, 842, 1110, 1005, 1055, 1081, 1288, 1194, 1264, 1204, 1333, 1365, 1406, 1404, 1454, 1536, 1657, 2016, 2194, 2160, 2464, 3041, 2708, 669, 823, 907, 1478, 1556, 1635, 1967, 2038, 3562, 282, 675, 3764, 495, 532, 1097, 1387, 1495, 1643, 1898, 1911, 2051, 2262, 2822, 2990, 3464, 2891, 3184, 3282, 1711, 3771, 2964, 3442, 1738, 847, 856, 1991, 4054, 1699, 3502, 2105, 2972, 3182, 3268, 2558, 194, 444, 1188, 1357, 1686, 971, 2248, 2245, 2553, 2917, 551, 1717, 3430, 1147, 1494, 1560, 2059]. History [2161, 2698, 2920, 2579, 1481, 2863, 1152, 328, 3043, 2662, 3630, 2594, 3717, 1164, 2706, 2907, 3100, 2719, 2827, 2819, 3345, 2839, 3852, 3853, 3854, 3288, 3131, 3860, 3188, 2999, 169, 767, 1395, 1575, 2050, 174, 2828, 80, 225, 1477, 3412, 3468, 4071, 2928, 1479, 3017, 3244, 2927, 3900, 3243, 1944, 2938, 1092, 1179, 1865, 2365, 1126, 2568, 3191, 3636, 2775, 2884, 3036, 2409, 3619, 26, 55, 190, 271, 326, 2701, 4078, 3208, 356, 293, 409, 635, 762, 939, 938, 1072, 1511, 1665, 1662, 1749, 1058, 4079, 545, 3676, 2924, 2753, 3078, 2552, 753, 1229]. History [1392, 1249, 1618, 3958, 3623, 1786, 3183, 3241, 974, 3058, 3474, 3332, 892, 1348, 1733, 2982, 2925, 4087, 1127, 1837, 743, 3278, 2565, 2498, 2147, 2456, 489, 2180, 1290, 1814, 1636, 2088, 283, 342, 685, 772, 1576, 4044, 1608, 643, 1808, 3791, 3792, 3793, 3794, 10, 182, 2381, 1111, 43, 3818, 3847, 3855, 4066, 1656, 933, 3874, 3876, 3878, 4069, 4070, 4072, 4075, 191, 1810, 1934, 1182, 3925, 3926, 25, 184, 631, 1750, 1897, 3944, 3945, 833, 2731, 3956, 3973, 1198, 1654, 2198, 3986, 3987, 3988, 541, 3992, 1119, 3997, 4004, 4010, 2433, 1332, 4031, 4035]. history [4037, 4038, 4023, 2981, 2371, 2978, 3812, 408, 2950, 848, 2152, 1976, 1600, 3938, 974, 1839, 1433, 3510, 3740, 1071]. historyofscience.com [2753]. Histrico [2895]. Histrico-Naturis [2895]. Histricos [1855]. Hjalmar [3347, 2745, 2496]. Hmso [992]. Hndler [1095]. Hoare [2802]. Hobbes [168, 1533, 1550, 1708, 2394, 3525]. Hochadel [3675, 2532]. Höchler [2785]. Hoddeson [1249, 860]. Hodge [3122]. Hodierna [2389, 13]. Hoechst [3037]. Hoefer [2475]. Hoel [2215]. Hoenen [1188, 1445]. Hoepli [2156]. Hofe [1212]. Høffding [3137]. Hoffmann [1724]. Hoffmanns [1724]. Hofmann [1150]. Hoftijzer [2884]. Hogarth [1818, 3879]. Hohenheims [1396]. Hohenhein [217]. Hohenlohe [1402]. Holbrook [992]. Holden [2687]. Holes [1296]. Holger [1838]. Holism [1575, 3753, 3754]. Holland [280, 582, 541]. Hollein [3082]. Hollister [1164, 1071]. Hollister-Short [1164, 1071]. Holmes [15, 1846, 2626]. Hologenesis [3377]. Holstenius [2058]. Holton [2674]. Holy [3068, 2364, 3214]. Holzboog [3064, 3283, 343, 344, 1731, 1202, 1328, 2465]. Homberg [3598]. Homburg [173, 3332]. Home [3475, 1067, 972]. Homeopathic [1935]. Homer [692]. Homme [3365, 419, 314]. hommes [3280, 4063, 1224]. Homo [1154, 4009]. Homoeopathy [2365]. Homosexualities [1692]. Homosexuality [580, 573]. homosexuals [573]. Homunculus [3199]. Hon [3410]. Hong [2216, 2268]. Honkapohja [3440]. Honor [164, 2887, 3442, 3738, 3010, 3772]. Honoré [1985, 2829, 2729, 795]. Honour [2436, 1857, 4057, 1198]. Honourable [3472]. Hooke [616, 1776, 3432]. Hooker [3032]. Hool [1177]. Hopeful [304]. Hopkin [3079]. Hopkins [1231, 2016, 3543, 3268, 3256, 2794, 1409, 3110, 2825, 3267, 2799, 2584, 3210, 2875, 1458, 2786, 1329, 1837, 1839]. Hoppe [2161]. Hopwood [3400]. Hoquet [2729]. Horacio [494, 618]. Horeau [368]. Horizon [820]. horizons [253]. Horizonten [253]. horlogérie [425]. Hormigón [1515]. Horn [2070]. Horoscopes [408]. Horst [725, 3129, 2943, 3225]. Hortus [1518]. hospices [1933]. Hospital [3012, 1652, 3343, 169]. Hospitals [1911, 1778, 4005, 1778]. Hoüel [3768]. Houltz [2745]. Hourglasses [3362]. Hourya [1668]. House [672, 2948]. Household [3493]. Housewife [3726]. Howard [373]. Howells [1184]. Howie [2470]. Hoyer [698]. Høyrup [3002]. Hrsg [571, 1191, 1735, 1838, 1450, 1150, 29, 798, 994, 1281, 172, 2529, 1731, 276, 1285, 1396, 1854, 379].

Hrsgs [1004]. Hs [1076]. Hsi [2137, 3897]. Huard [500]. Hubble [1862, 4036]. Hubel [570]. Huber [1764]. Hubert [2529, 2732]. Huby [1225]. Huff [1261]. Huffman [579]. Hugeniorum [1323]. Hugford [1212, 1212]. Hugh [3658, 1617]. Hughes [2546, 2688]. Hugo [109]. Hugues [574, 609]. Huldrych [639, 564]. Hulin [2973]. Hull [927]. Hulme [724]. Hulton [840]. humain [2952]. humaine [2528, 1232]. humaines [1793, 2663, 1224]. Human [592, 3175, 3482, 3150, 3573, 1152, 2978, 3661, 2706, 3849, 1409, 2835, 3656, 3296, 3148, 2985, 3102, 1225, 2628, 3144, 3751, 3804, 371, 1845, 2688]. Human-built [2688]. humani [2033]. Humaniora [950]. Humanism [1627, 3409, 3940]. humanisme [754, 2877]. humanity [848, 4079]. Humberto [371]. Humblot [3226]. Humboldt [3461, 3401, 3073]. Humfrey [3551]. Humoring [2655]. Humours [3666]. Humphrey [1872, 3745]. Humphry [1070, 1588]. Hundert [1151]. Hundred [699, 2392, 3605, 3969]. hundredth [4089]. Hünemörder [296]. Hungarian [3976]. Hungary [1019, 3976]. Hunter [615, 1087, 1139, 1490, 2143, 2873, 616, 3635]. Hunters [3676]. Hutchins [2984]. Huxley [65, 1915, 3762, 3827, 3884, 1915]. Huxley. [1914]. Huygens [1323, 3645]. Hyde [2765]. Hydraulic [2478]. hygiene [422, 3937, 3197]. Hylas [255]. Hypatia [77, 2935]. hypnotism [1332]. Hypothesis [3139]. Hysteria [2016, 3916].

I-xix [2501]. I. [1059, 2550, 2011]. i.e [57]. Ian [3412, 2362, 2076, 1242, 2716]. Iannaccone [1059, 832, 1873, 2307]. Iaria [1335]. Ib [840]. Ibc [2148]. Iberian [3653, 2920]. ibérique [1135]. Ibn [2199, 216, 3133, 2573, 791, 3292, 2998, 1047]. Ibnal [1358]. Ibnal-Abbasal-Magusi [1358]. Ibrhm [2573]. ich [1731, 695]. ichnographia [3434]. Icmai [970]. Icmai-Il [970]. Icon [2527, 2777, 2542, 2546, 2740, 2604]. Iconografia [83, 1378, 2612]. Iconographic [395, 3624]. Iconography [862, 1378, 184, 4014]. Ida [2416]. ide [2496]. ide- [2496]. Idea [680, 819, 723, 2394, 1837, 2726, 1921, 3065, 3878, 191]. idealistic [1086]. Ideas [60, 1683, 2776, 1439, 1717, 2658, 2988, 3852, 3853, 3854, 174, 225, 850, 1274, 806, 2967, 2802, 4024, 3041]. Idee [1539, 1345, 3426, 2132, 1022, 41, 995, 1648, 2624, 680, 1802]. Idées [1401]. idehistoria [3854, 1023, 3852, 3853]. identification [3844]. identificazione [3844]. identità [2558]. Identities [3624]. Identity [2075, 3091, 3480, 3800, 3083, 2506]. ideologia [442]. Ideologiae [339]. ideologie [2135]. Ideology [2686]. ides [198]. idhistoria [2200]. idraulica [1461, 3710, 782]. idrauliche [736]. IEE [847]. Ier [746]. I’Europa [1236]. if [3191, 3199]. IG [3037]. Igiene [3057, 2221, 3937]. Ignacio [2876]. Ignaz [172]. Ignazio [1212, 1212]. Ignorance [3229]. ihr [547, 2146]. ihre [2837]. ihrer [1918, 639, 4078]. Ihsanoglu [2749]. II [1091, 1402, 1441, 2267, 2370, 4043, 339, 620, 221, 1769, 3015, 243, 307, 446, 571, 798, 873, 994, 1516, 2010, 2727, 3013, 1335, 2808, 173, 2719, 3345, 3225, 1706, 343, 225, 2892, 2713, 171, 1489, 589, 727, 832, 2529, 2725, 2107, 362, 3019, 289, 892, 12, 3004, 1291, 1391, 2692, 1370, 748, 3163, 467, 3111, 2955, 1048, 984, 827]. IIe [3448]. III [2370, 2495, 2878, 2199, 2898, 221, 1212, 1688, 2010, 1909, 2808, 173, 3345, 1641, 217, 2590, 360, 171, 727, 186, 945, 1391, 831, 2015, 3163, 221, 1212, 522, 2049, 1049, 468]. Ilardi [2890]. Ilaria [3496, 2269, 2644]. Ildegarda [948]. Ileana [1350, 1863, 1750]. Iliffe [2781]. Ilimitada [3618]. Ill [1314, 1648, 1911, 3254, 1320, 766, 303, 777, 1110, 1005, 1055, 1201, 1264, 1763, 1725, 1819, 1818, 1910, 2120, 2056, 2200, 2160, 2421, 2500, 2601, 2715, 2675, 2746, 2012, 2606, 2612, 2669, 2737, 1811, 2530, 2522, 2707, 3255, 3280, 2364, 2653, 2682, 2528, 3526, 2524, 2631, 2313, 2523, 2588, 2685, 2671, 2542, 31, 32, 512, 511, 449, 448, 576, 648, 966, 964, 886, 1077, 1191, 1157, 1280, 1187, 1281, 1358, 1442, 1360, 1717, 2371, 2602, 2592, 1147, 1327, 1862, 1869, 1791, 1928, 2080, 2059, 2232, 2216, 2270, 2321, 2311, 2381, 2504, 2490, 2497, 2605, 2589, 1588, 2535, 2658]. ill [2577, 3209, 1335, 3205, 2536, 2662, 2657, 2546, 2531, 2572, 2646, 2808, 2655, 2774, 542, 1715, 2751, 2719, 2632, 2600, 2672, 2614, 2654, 2712, 2642, 2690, 2630, 2651, 2747, 1732, 2783, 1997, 2623, 2652, 2773, 2656, 2681, 2718, 2561, 2709, 2197, 2137, 2493, 2736, 2015, 2004, 407, 1190, 2598, 2555, 622, 2554, 3208, 1452, 1411, 1343, 1350, 1419, 1544, 1591, 1743, 1665, 1768, 1745, 1746, 1877, 1806, 1863, 1873, 1792, 1860, 1871, 1872, 2002, 1930, 1919, 1931, 1923, 2009, 1926, 2069, 2112, 2078, 2118, 2068, 2098, 2111, 2148, 2226, 2213, 2153, 2149, 2195, 2221]. ill [2225, 2310, 2415, 2414, 2369, 2423, 2363, 2541, 2529, 2732, 2792, 2686, 2753, 833, 2617, 2731, 2676, 2749, 2725, 2678, 3525, 2618, 2673, 2803, 1400, 2585, 2679, 2597, 3206, 1734, 1154, 1334, 1541, 2378, 2593, 1007, 993, 1796, 1988, 2603, 582, 378, 381, 382, 1836, 2136, 2458, 2460, 2549, 2625, 2650, 3250, 2587, 2582, 2691, 1642, 1868, 2433, 2422, 2503, 2796, 1644, 1990, 2622, 2674, 2604, 2534, 2724, 2779, 2556, 2266, 2627, 690, 2562, 2688, 2748, 1006, 1097, 1155, 1197, 1171, 1145, 1314, 1316, 1322, 1440, 1317, 1439, 1569, 1727, 1320, 1324, 197, 794, 1090, 1108]. ill [1234, 1261, 1589, 1736, 1573, 612, 686, 676, 725, 703, 698, 673, 696, 699, 841, 884, 834, 846, 1095, 1060, 1488, 1487, 1564, 1518, 1495, 1533, 1608, 1661, 1681, 1643, 1682, 1689, 1804, 1817, 1835, 1738, 243, 193, 249, 188, 311, 574, 836, 857, 830, 1447, 2103, 1243, 1238, 1366, 1612, 1620, 1879, 557, 1315, 1484, 1640, 761, 1453, 1429, 1430, 1456, 1355, 1354, 1415, 1422, 1343, 1421, 1437, 1423, 1584, 1585, 1609, 1525, 1599, 1512, 1511, 1510, 1591, 1532, 1614, 1563, 1744, 1690, 1663, 1730, 1662, 1666, 1749, 1695, 1794, 1874, 1813, 1787, 1807]. ill [2001, 2009, 2115, 2077, 2110, 2151, 2212, 2271, 2317, 2319, 2368, 1764, 1113, 1076, 1109, 1089, 1720, 1999, 2581, 1986, 1371, 1267, 1270, 1252, 1269, 1352, 1331, 1519, 1530, 1485, 745, 1336, 1577, 1798, 1814, 1390, 1482, 1839, 1834, 2054]. ill. [1420]. ill.n.n [1861]. Illinois [2776, 2660]. Illness [2906]. illuminata [1427, 1653]. Illumination [2377, 3943]. illuminato [585]. Illuminismo [621]. illuminista [3178]. illuministica [258]. illus [3121, 2456, 218]. illustr [172, 1407]. Illustrata [432, 503, 638, 1027]. Illustrated [1078, 1805, 3434, 1044, 3451, 3561, 802, 1494, 761, 174]. Illustrating [2478, 3319, 3325]. Illustration [1314, 2565, 2391, 3682, 3078]. illustrations [64, 173, 3530, 489, 21]. illustrato [2954]. Illustrazione [2173, 2391, 688, 1816, 590]. illustrazioni [1566, 1641, 930, 1915, 1992]. illustré [638]. Ilse [67, 1074]. Ilsetraut [2767]. Ilustracin [2101]. Ilya [1998]. Im [2196, 2853, 193, 571, 798, 994, 1862, 220, 547, 4062, 1735, 1992, 950, 4078, 1396, 364, 361, 3981, 690, 565, 4036]. Image [2160, 3055, 2770, 3705, 3048, 1511, 3051, 3294, 3054, 2146, 2391, 2066, 3832, 4086]. Images [544, 1439, 3322, 3372, 3321, 3400, 2988, 3411, 2970, 3432, 2152, 3554, 3451, 3363, 3997, 2820, 1230]. Imaginaire [1700]. imaginandi [384]. Imagination [335, 1457, 3433, 384, 3003, 3864]. Imaging [3200, 1143]. imaginibus [3165]. Imagining [3543, 862]. Imago [1926, 1378, 1324]. Imbach [915]. imbrogliano [3831, 1112]. imitations [2195, 4058]. immaginario [1140]. Immaginazione [3864, 3426, 2203]. Immagine [48, 1120, 2391, 1159, 1861, 1931, 2149, 2414, 1653]. Immagini [2488, 265, 3135, 1011, 1117, 1318, 2796, 178, 250, 844, 1695, 1230]. Immanuel [2605, 2374]. Immoderatus [3433]. Immortality [3148]. immunditia [2221, 3937]. immunology [1656, 3926]. Immutability [2262]. Imola [2007]. Impact [774]. impatto [1769]. Imperato [2226]. Imperial [3649, 1299, 1329, 4029, 2698, 3032]. Imperiale [2386, 2126, 3541]. Implications [2505, 2972]. impone [1033]. Importance [2418, 3147, 3840, 4002]. Important [1078, 696]. importante [2140]. Impossibile [3544]. imposture [2693]. Imprenditori [1676]. Imprenta [2593]. imprese [2113]. impression [3448]. imprimè [3010]. Improve [3339]. Improvement [3316, 3993]. Impulses [1409]. Impulso [687]. impure [3829, 2060, 3829]. Imschoot [410]. IMSS [2162, 2121, 2325, 2385]. Inaugurali [1128]. Incarnate [2093]. inchiesta [2903]. incipit [991, 1156, 70, 1102]. incisioni [814]. include [1728]. including [422]. Incomparable [3615]. incomprensioni [2383]. inconscio [932, 1613, 2057]. incubated [4023]. incubato [4023, 2205]. Incunabuli [1077]. Incurabili [1522]. incurables [3959, 3264]. Indagine [3103, 1159, 1779, 2046, 2009]. Indagini [1467]. Indeed [2702]. Index [2267, 213, 398, 399, 2250, 2251, 1688, 746, 86, 35, 40, 33, 34, 38, 43, 21, 218, 422, 2013, 1618, 1546, 42]. indexing [10]. India [1750, 2485, 3712]. Indian [329]. Indiana [1288, 445, 613, 2561, 3078, 693]. Indianapolis [693, 3011]. indiano [537]. Indians [441, 3579]. indicati [2415]. indicato [1901, 2424, 1992, 2004]. indicazione [2423]. Indice [6, 5, 159, 324, 465, 464, 666, 905, 1624, 1625, 1772, 1773, 1882, 1883, 2242, 2243, 2244, 2253, 2254, 2341, 2342, 2343, 2438, 2439, 2440, 1546, 2571, 72, 70, 407, 753]. Indices [334, 219]. Indici [1623, 1771, 2039, 2122, 2241, 2340, 2437, 1881, 1190, 2590, 994]. indicie [749]. Indigenous [2827]. indipendente [2472]. indirizzi [1815]. Individualità [66]. individualitet [2500]. Individuality [1222]. Individuation [30]. Indivisibles [2099]. Indizes [746, 747]. industria [2808, 1030, 635, 1540]. Industrial [2607, 2059, 2975, 3850, 1020]. industriale [558, 970]. industriali [1020]. industrialization [3018]. Industrie [653, 701, 2188, 1150, 2229]. industrielle [1612]. industriellen [497]. Industries [1641]. Industriezeitalters [1405]. Industry [2972, 3265, 2686, 1808, 971, 635, 4011, 1540, 4031]. Ined [2528]. inédit [4089, 574]. Inedita [1037, 1829, 3221, 3088]. Inedite [1782, 2349, 3838, 4083, 2380]. Inediti [2450, 211, 313]. Inedito [2176, 2398, 2178, 739, 2167, 2238, 67]. inédits [516]. infanticidio [2963]. infanzia [2057]. infini [3781]. infinita [715]. infinitamente [315]. Infinitely [2099]. Infinitesimale [2171, 1140]. infinitésimales [2199]. Infinitesimals [2099]. Infinito [2171, 2503]. infinity [3781]. Infirmary [169]. Infirmity [3595]. Influence [1085, 2957, 2658, 2749, 1225, 3761, 3819, 2757]. influencia [2097]. Influenza [818, 4010]. Informatiche [1645]. Information [990, 2753, 3293, 4046, 1758]. Informatione [2714]. Informazione [2363]. Infrastructure [3479]. Ingaliso [2714]. Inganni [1462]. Ingeborg [64, 1391, 64]. ingegnere [3402, 3596]. ingegneri [1966, 2037, 4049, 2315, 2789, 4064]. Ingegneria [2415]. ingégnieurs [2550]. ingeleid [3109]. Ingeniera [2876]. ingeniería [2808]. ingenieríaen [2808]. Ingeniero [965]. ingenieros [494, 618]. ingénieurs [1612]. Ingeniosa [3748, 2419]. Ingenious [3748, 3874]. Inghen [28, 1188, 1281]. Inghilterra [2883]. Inghirami [688, 590]. inglese [2721, 912, 348]. inglese-italiano [912]. inglesi [2520]. Ingmar [2922]. Ingrassia [2714]. Ingrid [491, 568]. inheritance [421]. Initiatives [3512]. initiorum [1688]. initiorum-Index [1688]. inizi [716, 3090]. iniziatico [2849]. inizio [1186, 1718]. Inken [2857]. Inlay [3484]. Inner [3152]. innesto [1674]. Innocent [1477]. Innocenti [1784, 1477, 1192, 1755]. Innovation [3210, 1020, 3036]. Innovazione [1759, 1975, 1020]. Innsbruck [3449]. inondazioni [1461]. inquieto [417]. Inquiry [1313, 3018, 3621]. Inquisición [889]. Inquisition [2993]. insanity [788]. insbesondere [2228, 3902]. Inscriptions [3050, 3707, 2118]. Insect [3520, 3495]. insectes [2015]. insects [2015]. insegna [575]. insetti [1953, 2024]. Insights [2928]. institucional [494, 618]. Institut [507, 2228, 3116, 3286, 353, 698, 512, 1442, 2161, 2265, 2118, 3019, 3824]. Institute [224, 1078, 1340, 2168, 2239, 2964, 445, 964, 830, 1102, 2248, 334, 3205, 408, 2308, 1182, 1355, 2068, 610, 2731, 791, 378, 2245, 3824, 1477, 3969, 3990]. Institute-University [791]. Institutes [1078, 248]. Institution [3471, 1006, 1238, 2080, 2381, 2697, 2751, 1732, 1913, 184, 293, 1733]. institutional [3906, 993]. Institutionalization [672]. Institutionen [2496]. Institutions [3023, 1063, 225, 1972, 2130, 1829, 3895, 1749, 4080, 2749]. Instituto [333, 352, 2068, 2114, 1855, 2097, 3990, 3990]. Instituts [296, 1687]. Instructif [1305]. Instruction [939, 1474]. Instructions [1793, 2798]. Instructive [1305]. Instrument [2179, 2299, 2442, 63, 2295, 3472, 2185, 317, 3320, 3342, 4078, 761, 916, 1161, 2913, 1808, 762, 2310, 4013, 2196, 4078]. instrument-making [762]. Instrumenta [1685, 1343]. Instrumental [3399]. Instrumentalizing [3555]. Instrumentation [631, 2937]. Instrumente [1012, 2915]. instrumenti [1955, 2026]. Instrumentos [1860, 4017]. Instruments [1393, 2169, 2240, 2436, 3171, 3282, 985, 3694, 63, 856, 3360, 2183, 2089, 3111, 2084, 3288, 1664, 3719, 2483, 2152, 427, 835, 1084, 998, 1320, 2098, 2975, 1308, 1044, 1972, 2485, 2815, 2147, 489, 2180, 3587, 4015, 3763, 3766, 2757, 3871, 4069, 3885, 761, 992, 1256, 1860, 2069, 2195, 2369, 3997, 4091, 4031, 4058, 4017, 2556]. Int [868]. Integraphs [3104]. Integrated [2260]. Integrating [2161, 4072]. Integration [3482, 3930, 1667, 3930]. Integrato [2260]. Intellectual [491, 891, 1110, 1490, 2658, 3008, 2412, 3058, 4048, 3862, 1843, 2077]. Intellectuals [2991]. intellectuel [1444]. Intellettuale [310, 306, 1157, 1388, 2571, 619, 1236, 455, 2077, 1003]. intellettuali [897]. Intelligent [2966]. intelligenza [1459]. Intelligibility [2862]. intend [1305]. Interaction [190, 362]. Interactive [3458]. interamente [2107]. Interazioni [2429]. Interdipartimentale [1619]. interdisciplinaire [2229]. Interesse [1054, 1259, 2002, 2415]. Interest [1308]. Interests [3414]. interfacultaire [1859, 1734]. Interferential [2407]. Interferometer [2406]. Interlinea [2156]. International [583, 1611, 2933, 4052, 4054, 648, 1494, 2161, 2594, 2839, 4066, 1770, 4071, 2745, 2204, 753, 4087, 3092, 1137, 3684, 2878, 4043, 4047, 260, 4053, 1442, 1984, 3893, 4077, 2078, 2118, 1799, 1326, 1912, 2201, 1224, 669, 823, 907, 1478, 1556, 1635, 328, 4059, 1479, 55, 271, 326]. Internationale [2145, 1563, 753, 749, 2267, 3893, 1912]. Internationalen [1281, 343]. Internationales [1401]. Internationalis [1600]. Internationalism [3639, 1245]. Internazionale [721, 1488, 2522, 215, 1360, 404, 3031, 711, 2078, 2113, 2809, 1310, 1388, 915, 572, 4059, 1943, 2743, 838, 1737, 1137]. Internazionali [354, 1994, 1025]. Internet [2753, 2244, 2343, 2440, 1711]. interpersonale [2324]. Interplanétaire [1293]. Interpretation [2662, 3630, 1656, 3926, 4033, 2821]. Interpretations [2106, 2103, 3830]. Interpretazione [578, 199]. Interpreting [1204, 1736, 2470]. Interrelations [2161, 4072]. Interuniversitaires [2231]. Interventi [1373]. intervento [2667]. Intimate [440]. intitulé [1232]. Intorno [1037, 817, 1424, 2399, 1953, 2024, 1947, 2018, 1673]. intra [3433]. intrigo [2535]. introd [1414]. Introducción [3939, 2097]. Introduccióna [2593]. Introduction [2516, 1911, 3463, 2402, 3300, 28, 30, 58, 3430, 1859, 334, 3317, 3298, 2662, 3336, 368, 754, 3046, 3721, 893, 451, 791, 749, 3355, 3386, 3270, 3212, 533, 1611, 2528, 574, 1668, 681, 763, 1944, 14, 4024, 3822, 3418]. Introductions [998, 2898, 512]. Introductory [606, 955]. introduit [2578]. introduttivo [563]. introduttore [2376]. Introduzione [764, 612, 2620, 1417, 2497, 2605, 2202, 788, 1411, 1663, 929, 13, 1540, 1728, 3822, 2692, 1521, 854, 684, 1994, 2499, 1269, 2141, 2144]. introduzioni [2530]. invasi [3178]. Inventaire [1379]. Inventario [2176, 904, 2400, 38, 1260, 697, 664, 291, 953, 1162, 2068, 529, 555, 789, 876, 3990, 2314, 2399, 1783]. inventata [372]. Inventing [3849, 2827, 2681, 1845]. Invention [1842, 3670, 3401, 2339, 1037, 3018, 701, 1563, 3083, 2604]. Inventions [3399, 3061]. Inventor [3688, 3757, 1237, 1814]. inventore [1690, 1237, 2769]. Inventories [1448]. Inventory [964, 1618, 2176, 2400, 14, 2068, 3990, 1783]. Invenzione [1037, 2339, 807, 312, 505, 1491]. invenzioni [559].

Investigaciones [1227, 1226, 2494, 494, 618]. Investigating [2404, 3406]. Investigation [2350, 58, 3054, 1198]. Investigationes [889]. Investigations [1683, 2318, 30, 3805, 3366]. Investigative [2626]. Inviato [664, 2590]. Invisible [1842]. Ion [235]. Ionic [1657]. IOS [3081]. Iowa [1362]. ipotesi [2233, 2162]. Ippocrate [1092]. Ippolito [955]. Iranian [2564]. Irby [3059, 2417]. Irby-Massie [3059]. Ireland [1933, 916]. Irene [3225, 1158, 303, 1797]. Irina [3618, 3264]. Irish [916]. Irma [862]. Irrelevant [3301]. irresistibile [3286, 3824]. irresistible [3824]. Irritability [2732, 1701]. Irritating [2732]. Irwin [1694, 1334]. Isaac [2586, 2311, 1588, 2600, 2781, 803, 143, 3843, 3863, 3490, 3917, 2372, 3476]. Isabel [3475, 2876, 3025]. Isabella [2642]. Isabelle [2013, 1998, 1567, 1899]. Isaia [832, 1873, 2307]. ISBN [1763, 1725, 1819, 1818, 1846, 1853, 1795, 1945, 2003, 1910, 2016, 2119, 2120, 2075, 2102, 2056, 2067, 2194, 2227, 2143, 2145, 2228, 2200, 2160, 2267, 2312, 2318, 2421, 2420, 2370, 2424, 2488, 2463, 2467, 2496, 2495, 2500, 2464, 2560, 2586, 2601, 2527, 2715, 2675, 2746, 2776, 2793, 2887, 2953, 2986, 2916, 3041, 3016, 3299, 2418, 2459, 2516, 2573, 2647, 2708, 2710, 2879, 2881, 3000, 2055, 2843, 2717, 2607, 2899, 1852, 1921, 2878, 2889, 3065, 1966, 2012, 3113, 2942, 3543, 2425, 2462, 2606, 2730, 2897, 3116, 3064, 3002, 3107, 3120, 3286, 3283, 2612, 2896, 2669, 3094, 2735, 2737, 3005, 2777]. ISBN [1898, 1911, 1981, 2517, 2530, 2538, 2822, 2894, 2933, 2898, 2882, 2936, 3063, 3040, 3059, 3071, 3121, 3109, 2852, 2851, 3081, 2491, 2522, 2707, 2864, 2770, 2981, 2891, 3068, 3255, 3254, 3280, 3282, 3541, 2320, 2431, 2364, 2471, 2664, 2780, 2653, 3072, 2805, 2964, 2580, 2203, 2472, 2144, 2543, 3080, 2750, 2682, 2528, 2667, 3526, 3545, 3025, 2836, 2837, 847, 1085, 1112, 1808, 2631, 2721, 2797, 2959, 2906, 3009, 3023, 2865, 1991, 2314, 2105, 2689, 2106, 2972, 2798, 2769, 2309, 3279, 2523, 2588, 3015, 3204, 2853, 2979, 2908, 3082, 3084, 3129, 3123, 3122]. ISBN [3268, 2849, 3021, 2685, 2826, 2616, 2551, 2558, 2671, 2791, 2909, 3035, 2542, 2974, 2969, 3030, 2670, 3001, 2648, 2963, 2958, 2553, 1151, 2917, 3124, 3544, 2371, 2602, 2619, 2665, 2957, 2592, 1327, 1515, 1668, 1758, 1670, 1862, 1869, 1791, 1859, 1820, 2007, 2014, 1928, 2080, 2079, 2059, 2066, 2161, 2232, 2216, 2215, 2224, 2229, 2157, 2156, 2270, 2276, 2315, 2321, 2311, 2381, 2504, 2490, 2497, 2605, 2727, 2789, 2589, 3076, 1667, 2535, 2824, 2920, 3263, 3022, 1063, 2579, 2658, 2834, 2883, 2873, 2577, 2755, 2863, 2962, 3027, 3066, 3006, 3115, 3209, 3256, 1909, 2383, 3013]. ISBN [2978, 43, 3205, 2520, 3127, 2949, 3298, 2536, 2756, 3043, 2662, 2831, 2519, 2838, 3090, 2657, 2594, 3111, 2379, 2546, 3087, 2531, 2895, 2572, 2646, 2874, 2876, 3061, 3060, 3003, 3012, 1908, 2808, 2655, 2550, 2774, 2941, 2706, 57, 1715, 2888, 1917, 2943, 2751, 2907, 2719, 2827, 2819, 2948, 3530, 2934, 1995, 2632, 2839, 2600, 2988, 2742, 3057, 2595, 2885, 3070, 2955, 2672, 3029, 2946, 2461, 3108, 2794, 2893, 3289, 3288, 2633, 2518, 2470, 3131, 2614, 2654, 2800, 2829, 2922, 2905, 2740, 3225, 3281, 2502, 2712, 2757, 2642, 1656, 2690, 2630, 2651, 2999, 2904, 2787, 1575, 2771]. ISBN [2947, 2747, 1732, 1722, 1987, 2377, 2557, 2973, 3253, 2308, 2783, 1997, 3110, 3265, 2492, 2623, 2720, 2652, 2835, 2773, 2892, 2268, 2781, 3067, 3079, 415, 2832, 2656, 2950, 2532, 2681, 2854, 2859, 2718, 2862, 2061, 2561, 3033, 2709, 2596, 2134, 2197, 2137, 2493, 2629, 2323, 3024, 3074, 1841, 2494, 3017, 2533, 2970, 3091, 2196, 2900, 2736, 2977, 2912, 770, 683, 927, 850, 848, 1407, 1878, 1875, 1915, 1992, 1901, 2015, 2004, 2013, 1944, 2362, 2427, 2966, 2825, 2590, 2945, 2419, 2598, 2754, 3014, 2938, 2965, 3258, 2152, 2222, 2275, 2413, 2430, 2373]. ISBN [2365, 2375, 2468, 3069, 3018, 2659, 3031, 3546, 2713, 3542, 3529, 2940, 2525, 2741, 2918, 2790, 2801, 2555, 2505, 2743, 2775, 2884, 2944, 3426, 2987, 2684, 3267, 3251, 2821, 2961, 2207, 2799, 3036, 2615, 1843, 3277, 2554, 3208, 2660, 1849, 2693, 3287, 3257, 2880, 295, 1743, 1665, 1768, 1745, 1746, 1877, 1806, 1861, 1863, 1792, 1876, 1860, 1871, 1872, 1930, 1919, 1931, 1923, 2009, 1926, 2069, 2078, 2113, 2118, 2068, 2098, 2114, 2111, 2213, 2153, 2150, 2149, 2195, 2221, 2225, 2310, 2415, 2414, 2369, 2423, 2541, 2548, 2574, 2877, 2624, 2529, 2732, 2785, 2855]. ISBN [3093, 2792, 2858, 2686, 2677, 1074, 1838, 3088, 2206, 3038, 2584, 2716, 2809, 3224, 2745, 2753, 3078, 2976, 2204, 2886, 2617, 2731, 3114, 2833, 3008, 2060, 2676, 2844, 2975, 3118, 2722, 2583, 2552, 2749, 2725, 2841, 2952, 2875, 3010, 2576, 2537, 3284, 800, 997, 1228, 1229, 1285, 2645, 2840, 3083, 2767, 3125, 2643, 2678, 3525, 2663, 2734, 2272, 2366, 2426, 2382, 2412, 2506, 2618, 2559, 2738, 2673, 2867, 2866, 2983, 2985, 3039, 3085, 3130, 3226, 1851, 1922, 1899, 1998, 1994, 2107, 2062, 2132, 2428, 2429, 2380, 2469, 2499, 3034, 2807, 2830, 3241, 2803, 2984, 2316, 2384]. ISBN [2374, 2432, 2608, 2539, 2729, 2585, 1797, 2939, 2951, 2679, 2597, 3096, 2728, 3062, 2786, 1847, 2998, 2457, 2804, 2609, 3044, 3019, 2784, 2845, 3206, 2626, 3058, 2489, 3126, 1840, 3528, 1326, 1734, 851, 1154, 1334, 2273, 2378, 2507, 1854, 3547, 2372, 2593, 808, 806, 811, 803, 860, 892, 849, 913, 1007, 993, 1203, 1202, 1144, 1329, 1713, 1767, 1654, 1988, 1912, 2108, 2466, 2465, 2806, 2919, 2967, 3032, 2198, 2603, 2687, 2591, 1733, 1844, 1907, 2274, 2982, 3086, 3011, 3203, 3075, 582, 2856, 2861, 2911, 2890, 2921, 2914, 2544, 3261, 1993, 2269, 2903, 3095, 2902, 1837]. ISBN [1857, 1836, 2136, 2458, 2460, 2549, 2625, 2820, 2937, 3250, 2935, 2587, 2501, 3092, 2582, 3004, 3278, 2901, 2691, 3527, 2846, 1642, 1868, 2205, 2201, 2324, 2433, 2422, 2498, 2503, 2752, 2796, 2810, 2954, 2956, 1644, 914, 2714, 3037, 2622, 1906, 1900, 2613, 2545, 2674, 2739, 2910, 3239, 2604, 2534, 2724, 2779, 1800, 2556, 3259, 2644, 2628, 1845, 1850, 2322, 2367, 3073, 2266, 2627, 2772, 2562, 2848, 2802, 2857, 1905, 2913, 2688, 2971, 2748, 2726, 2960, 2571, 2968, 2668, 2578, 2692, 3026, 3089, 3264, 3285, 1914, 2416, 2417, 2882, 2915, 2782]. ISBN-10 [2822, 2909, 2781, 2950, 2966, 2880, 2939, 2956, 2971]. ISBN-13 [2899, 3071, 2959, 2909, 2838, 2950, 2977, 2966, 2938, 3267, 2976, 2939, 2956]. ISBN13 [2971]. Isegni [1620]. Isenburg [588]. Iside [2748]. Isidore [3290]. ISIS [10, 1898, 3944, 3945, 10]. Isisi [1898]. Isla [3363]. Islam [1261, 2710, 52]. Islamic [2710, 2937, 2838, 3707, 674]. Islamicate [184]. islamiques [2519]. Isnb [2850]. Isola [304, 952]. ispirazioni [1233]. Israel [505]. ISSN [2524, 3258, 800, 1228, 2201, 2628]. Istituti [2006, 1994]. Istituto [299, 514, 4043, 1784, 2882, 72, 309, 446, 1105, 1869, 1909, 752, 2461, 2202, 1190, 707, 712, 291, 832, 862, 1257, 1158, 1166, 1167, 1195, 1429, 1355, 1665, 1768, 1873, 2148, 2217, 2319, 337, 2833, 799, 800, 44, 1455, 2208, 12, 555, 1990, 2005, 2169, 2240, 519, 1561, 71, 238, 779]. Istituzioni [358, 632, 1021, 1739, 1125, 2297, 1046, 2177, 2256, 738, 2355, 7, 1829, 1384, 1472, 822, 1975, 325, 270, 96, 401, 19, 1339, 2130, 667, 2670, 1876, 2164, 2235, 1784]. Istoria [1119]. Istrumenti [1730]. Istruzione [1424]. Istruzioni [1304, 1472]. ital [1273]. Italia [852, 1239, 2425, 1801, 1533, 1679, 1574, 2616, 2969, 250, 1279, 971, 1572, 1759, 2270, 2315, 1908, 288, 1741, 874, 880, 881, 360, 1030, 2430, 14, 697, 730, 1033, 1159, 1745, 1876, 1919, 1065, 1114, 1265, 1323, 577, 1289, 44, 924, 1482, 258, 2043, 3763, 4049, 3782, 2620, 1118, 3090, 3866, 50, 3885, 2128, 662, 1890, 2738, 3181, 3982, 4010, 818, 1564]. Italian [202, 258, 517, 629, 573, 1966, 2168, 2239, 2237, 2332, 2330, 2338, 2339, 2425, 3002, 421, 1564, 643, 2364, 1085, 1112, 180, 311, 447, 719, 971, 741, 1499, 1928, 2156, 1111, 3717, 786, 788, 1026, 1092, 1271, 1753, 1865, 1810, 435, 355, 428, 429, 425, 631, 635, 656, 778, 1014, 1082, 1020, 1107, 1185, 1172, 1411, 1350, 1510, 1766, 1750, 1749, 1874, 2009, 2077, 2363, 1065, 337, 2584, 812, 1265, 1237, 611, 1348, 384, 371, 381, 3669, 3278, 2433, 1267, 1252, 1269, 1331, 1540, 3744, 3747, 3748, 4040, 3750, 5, 665, 2354, 3752, 1375, 4]. Italian [3467, 3757, 3760, 3761, 1705, 1037, 4044, 2635, 3055, 1711, 985, 3772, 1778, 3777, 4049, 3782, 3783, 1781, 3788, 1215, 214, 1118, 3798, 1829, 1970, 3799, 3801, 2356, 2248, 817, 1701, 741, 2089, 2176, 2400, 1560, 3804, 3807, 3808, 3812, 3431, 3813, 3818, 2091, 4059, 3824, 3826, 3831, 3838, 3839, 2123, 1219, 3844, 4064, 3847, 3848, 3851, 3856, 4067, 1, 3863, 3864, 3866, 1477, 2178, 3872, 3873, 3223, 3904, 3908, 1382, 3910, 3911, 4076, 2396, 3918, 3925, 1976, 2258, 3927, 2047, 3929, 1897, 3933, 3934, 4081, 3937, 3948, 2124, 3949, 2081, 982, 1114]. Italian [1210, 319, 1469, 1704, 1464, 3954, 3955, 3956, 1230, 2131, 2264, 4083, 1308, 3964, 3965, 3969, 3972, 3973, 2260, 4084, 4085, 2083, 3976, 1546, 1053, 3982, 3986, 3987, 3988, 3989, 1049, 1047, 1119, 734, 4006, 4007, 4010, 1378, 1783, 595, 2044, 3335, 4014, 4019, 4021, 4038, 2088, 2391, 3822, 3828, 3413]. Italiana [299, 1946, 931, 1235, 1829, 2042, 1748, 1871, 1919, 2319, 337, 2100, 2425, 1817, 1996, 2116, 576, 288, 1176, 862, 1749, 3982, 1064, 178, 2017, 2319, 2543]. italiana-Journal [1749]. Italiane [2370, 1995, 2191, 1128, 384, 991, 1156, 70, 2712, 2468]. Italiani [1894, 4, 737, 269, 461, 821, 1383, 1552, 1471, 2090, 741, 1124, 1828, 323, 463, 2447, 600, 1302, 1782, 1973, 2045, 663, 2129, 2353, 2397, 397, 2249, 1301, 95, 1633, 1064, 528, 818, 588, 809, 3426, 1372, 1350, 1748, 3335]. Italianistica [2954]. Italiano [995, 605, 446, 1824, 712, 238, 211, 1978, 44, 1455, 1990, 4043, 912, 3908, 425, 1167, 3125, 750, 555, 515, 1901]. italiano-inglese [912]. Italica [2612]. italico [1436]. Italicum [516, 2230]. Italie [3461, 314, 2060, 3307, 3829, 1888]. Italien [4062, 1992]. italienischen [950]. italienne [425]. Italiens [210, 2396, 2539]. Italo [2421, 1415, 2315, 3976, 2968]. italo-germanico [2315]. Italo-Ungheresi [2421, 3976]. Italy [258, 2425, 1564, 971, 3409, 980, 3439, 1320, 1749, 1319, 3564, 381, 3752, 3763, 1778, 4049, 3782, 3373, 1118, 3558, 3651, 1481, 3066, 3829, 4062, 3866, 3885, 2477, 3226, 3982, 4010, 3717, 3399, 3528, 3567, 3361]. Itc [1512]. Iter [516]. Ithaca [167, 2589, 614, 177]. itinéraire [2537]. Itinerari [1538, 457, 3125, 3908]. itineraries [3908]. itinerario [1268]. Itinerrios [2895]. Iurilli [223]. IV [1488, 2517, 311, 3412, 2821, 1411, 1923, 2539, 945, 1091, 307, 1280, 368, 171, 565, 1264, 182, 2740, 1781, 469]. Ivan [1411, 1837, 607]. Ivana [999]. Ivano [2726]. Ivanovi [1870]. Ivar [2800]. I’ve [3152]. Ivor [1149]. Ivory [3502, 409, 998]. Ivv [1255]. Iwan [2747, 2542, 2676]. IX [313, 717, 2785, 1645, 1137, 587, 3059, 949, 3003, 275, 2794, 2947, 2061, 2709, 2095, 419, 654, 791, 1330, 2901, 607, 474, 393, 791]. ix-246 [791]. ix-367 [791]. IXe [2199]. Izquierdo [3412].

J [164, 685, 885, 891, 1110, 1288, 1201, 1333, 1370, 2003, 2459, 1393, 1576, 584, 532, 1006, 1727, 1643, 3475, 2981, 3770, 2805, 2528, 549, 3122, 2791, 29, 30, 512, 1277, 1188, 1358, 1442, 1445, 2103, 1238, 2080, 614, 2873, 1428, 40, 2756, 2662, 1164, 218, 2751, 2907, 385, 2757, 3599, 2747, 1732, 191, 763, 2004, 2966, 638, 868, 1275, 1177, 1274, 2801, 2944, 2821, 3928, 88, 406, 762, 939, 998, 1563, 1877, 2858, 3676, 1071, 753, 749, 997, 1228, 1229, 2767, 163, 1460, 1980, 212, 625, 921, 3528, 1326, 766, 806, 723, 1034, 2465, 280]. J [2856, 3445, 2324, 2810, 914, 1225, 2044, 1920, 2556, 2644, 2578]. J. [2647, 772, 1097, 1447, 1685, 408, 928, 1484, 1641, 1982, 638, 1596, 1812, 3928, 535, 3118, 186, 2294, 946, 489, 2848]. J.-Cl [1326]. J.-F [921]. J.-H [1727]. J.-P. [1447]. Jaargang [302]. Jabr [609]. jäbrigen [296]. Jaca [1431, 812, 2057]. Jack [3417, 2079, 1997, 2617]. jacket [2470]. Jackson [3034, 2913]. Jacob [1427, 102, 1489, 1922, 892, 1086, 152, 142, 138, 2473, 2786, 3235]. Jacobi [2074]. Jacobs [1566]. Jacomien [3477, 2986]. Jacopo [3693, 1269, 3065, 59, 3002]. Jacquart [3631, 1358, 33]. Jacqueline [2717, 2952]. Jacques [81, 1080, 1911, 2519, 3197, 175]. Jacyna [2909]. Jaeger [3003]. Jaffe [862]. Jagadis [2485]. Jagiellonian [1256]. Jagow [2750]. Jahncke [567, 1715]. Jahre [676, 234, 1151, 3905, 93]. Jahrestagung [4078]. Jahrhundert [4062, 364, 2228, 2853, 3902, 1992]. Jahrhunderts [567, 64, 2576]. Jahrhundertwende [547]. jährigen [220]. Jaisson [2615]. Jaka [421]. Jaki [1859]. Jakob [2668]. James [1346, 2003, 2267, 1006, 3510, 1238, 2080, 1164, 2751, 1732, 848, 1071, 849, 840, 1290, 1853, 2075, 2495, 2899, 2889, 1659, 2669, 3063, 2472, 925, 1398, 4068, 2970, 2844, 3241, 3732, 2971]. Jami [2197]. Jamie [1346, 2586]. Jammer [2207]. Jan [1402, 1441, 884, 2203, 923, 1537, 3864, 1934]. Jane [2378, 3078, 3971]. Janeiro [430]. Janello [3456]. Janet [3415, 3704]. Janiak [3067, 2790]. Janine [2544]. Janis [2664]. Janni [2574, 1719]. Janssen [2715, 280, 2302]. January [3194, 3237, 2080]. Janus [1002]. Japan [3492, 3478]. Japanese [2305]. Jar [3291]. Jarcho [955]. Jardin [1418]. Jardine [1643, 3510, 3013, 3363]. Jardinier [1475]. jaws [707]. Jay [349]. Jayawardene [1660]. Jazari [3370]. Je [1731]. Jean [2165, 620, 444, 2648, 1138, 1612, 866, 2013, 586, 540, 893, 2423, 1228, 2469, 2708, 2236, 2512, 2441, 1859, 2819, 536, 2533, 3197, 701, 2423, 2541, 1460, 2052, 2469, 3966, 230, 1734, 1832, 2782, 2611]. Jean-Baptiste [620, 2611]. Jean-Claude [1228, 536]. Jean-Francois [2236, 1859, 2469, 1734]. Jean-Jacques [3197]. Jean-Louis [866, 2013]. Jean-Noël [1460]. Jean-Paul [1138, 586, 2541]. Jean-Pierre [2533]. Jeanneret [1543]. Jed [1573]. Jeff [2546]. Jeffrey [1736, 2851, 4057, 1857]. Jeffreys [1654, 3789, 3795]. Jehl [248]. Jena [420, 1042]. jenese [1455]. Jenner [244]. Jenneriana [1629]. Jenni [3595]. Jennifer [2794]. Jenny [3573]. Jens [3002]. Jensen [3612]. Jèrôme [2958, 3450]. Jeremias [103]. Jeremy [2579, 925, 543, 2753]. Jerez [2101]. Jernkontorets [1019]. Jerónimo [2593]. Jerry [167]. Jersey [1846, 1357, 2216, 2268, 654, 1397]. Jersey-London-Hong [2216]. Jersey/London/Hong [2268]. Jerusalem [3413]. Jervis [2063]. Jesse [3154]. Jesuit [3205, 1196, 2686, 2585, 3433, 1770]. jésuite [2214]. jésuites [1486]. Jesuits [1486]. Jesus [2146]. Jeu [3234, 3231]. Jeux [3234]. Jewel [2948]. Jewish [3413, 3576]. Jews [2861]. Jezek [499]. Jhan [598]. Jim [3418, 3719, 3732, 1072, 1511, 1662, 3032]. jing [1640]. Jinghwa [1076]. Jo [1002, 1500, 3060]. Joachim [567, 142, 966, 2831, 1012, 146, 118, 3216, 1396]. Joachim-Napoleon [3216]. Joan [2589, 494, 618, 3019]. Joannes [3736, 2833]. Joelle [1985]. Johan [1441, 884]. Johanivisson [703]. Johann [62, 127, 139, 149, 141, 117, 99, 155, 135, 138, 120, 137, 121, 128, 787, 1873, 148, 113, 787, 1796]. Johannes [156, 152, 145, 136, 110, 1074, 680, 382, 1868, 1905, 4035, 567, 122, 3158, 717, 3523, 3981]. John [283, 888, 2003, 2435, 3326, 652, 1898, 2864, 10, 511, 3441, 2873, 3572, 3809, 3836, 1409, 2095, 848, 2362, 3597, 1934, 1935, 3267, 344, 356, 406, 2584, 2366, 1720, 1993, 3669, 1225, 4035, 3571, 89, 573, 679, 1234, 1362, 1502, 2916, 2717, 3797, 196, 443, 511, 910, 2381, 2883, 3012, 760, 1395, 286, 683, 3911, 1184, 25, 356, 916, 4079, 2124, 450, 755, 1459, 693, 1202, 1720, 2625, 2096, 3481]. Johns [1231, 2016, 3543, 3268, 3256, 2794, 3110, 2825, 2799, 3210, 2875, 1458, 2786, 1329, 1837, 1839]. Johnston [3204, 1662]. Joint [2990]. Jol [3010]. Jolande [2074]. Jole [446, 1273, 2658, 2419]. Jolivet [444]. Joly [1138]. Jonas [2384]. Jonathan [2737, 3637, 3417, 1457, 1945, 451, 2382, 2804]. Jones [1853, 2601, 652, 1278, 2772]. Jöran [3000]. Jordanova [227, 544, 724]. Jörg [2866, 3226]. Jorge [2920, 1849, 30]. José [43, 2907, 2101, 889, 965, 3005, 45, 43, 2895, 2101, 3091, 3327, 1849, 1544, 1855, 3554, 2097]. Josef [1237]. Joseph [454, 414, 896, 1370, 1910, 2527, 1368, 2156, 1328, 3032, 2433, 2534, 2724, 1914, 130, 2897, 416, 3025, 748, 2629, 3236, 406, 2506, 1247, 3979, 2956, 2724, 2779]. Joshua [3635]. Josiane [1410]. Josselyn [2883]. Jost [1402]. Journal [439, 3322, 367, 1466, 1749, 3978]. journalist [1589]. Journals [2038, 2990, 1515, 3320, 3755, 180, 1967]. Journées [2229]. Journey [3293, 1089, 3147, 3788, 1107]. jours [638]. Jouvence [2011]. Jovilabe [2]. Joyal [921]. Jr [1362, 356, 229, 345, 68, 1839]. Juan [3634]. Juanel [2876]. Juanelo [965]. Jude [1336]. Judgement [830]. Judith [2660]. Juette [1812, 1935]. Juice [3603]. juin [4047, 444, 2118, 1326]. Jules [2215, 2302, 3768]. Julia [2424, 2979]. Julian [918]. Julie [3312]. Julien [701, 416]. Julien-Joseph [416]. Juliet [2945]. Julius [154, 148, 120]. Jullien [2829]. July [3237, 4071, 4087]. June [4045, 4059]. Jung [2074]. jungen [343]. Jungius [966]. Jungnickel [1819]. Juni [1141]. Junio [963]. Jürgen [1141, 1153, 2198]. Jury [2851]. Juste [3058]. Justice [3306]. Justine [2841]. Justus [149, 259, 145].

K. [950, 1913, 2556]. kabbalah [1528]. kabinet [1422]. Kaempfer [256]. Kagan [3368]. Kahn [3671]. Kai [1079, 1912]. Kaiserfeld [1264]. Kaiserreich [690]. Kampfplatz [2854]. Kangshen [2095]. Kanigel [2503]. Kant [2605, 284, 2374, 723, 1498, 1989, 1043, 1042, 2663, 2687]. Kant-Forschungen [1989]. Kantian [2911]. Kantiana [621]. Kantiani [1994]. Kants [1043, 1042, 1989]. Kanz [1079, 1912]. Kaoru [2197]. Kaplan [1231]. Karachalios [2208]. Karen [3369, 2819, 1438]. Kargon [3369]. Kari [3656]. Karin [703]. Karl [158, 2466, 1852, 2972, 3123, 2792, 2643, 2432, 361]. Karl-Heinz [2643]. Karolinska [1078]. Karp [1411]. Karr [3458]. kart [1797]. Kartoniert [1905]. Karvar [2550]. Kassel [234]. Kassell [2830]. Kassler [2586]. Kastner [2466, 568, 895]. Katalog [2265, 1083, 2576, 363]. Kate [1072]. Katharine [34, 2835]. Katherine [1592]. Kathie [3412]. Kathleen [2832, 1144]. Katholieke [3205]. Kathrin [2529]. Kathryn [2377]. Katrin [2374]. Kaunitz [212]. Kavalier [336, 522]. Kawashima [3331]. Keaney [1225]. Keel [2375]. Kegan [219]. Keiko [3331]. Keimpe [1983]. Keith [1337]. Keller [965, 3737]. Kelvin [689, 808, 4032]. Kemp [3014, 2985, 1485]. Ken [2940]. Kennedy [2055]. Kennet [2752]. Kenneth [3332, 856, 2381, 1152, 1274, 1076]. Kenny [2587]. Kepler [382, 3158, 2636, 1118, 3134, 2123, 3049, 3523, 2393, 382]. Keplero [1118]. Kepos [1811]. Kept [3656]. Kern [2655]. Keuren [2553]. Kevin [2003, 1999]. Kevles [1576]. Kew [1370, 1314]. key [1137]. Keynes [1502, 2916, 2781]. Keyser [3059, 2417]. Kg [948]. Kh [2576]. Khafīf [735]. Khazini [3550]. khmis [2199]. khwrisms [609]. Kielmeyer [1079, 2465]. Kielmeyer-Bibliographie [1079]. Kilmister [1503]. Kim [2730, 2137]. kinds [4022, 2156]. King [1370, 3570, 648, 1877, 674, 2937, 3250, 3928, 2676]. King-Hele [1370]. Kingdon [992]. Kinnaird [840]. Kinsey [1853, 1945, 3891]. Kiple [1152, 2752]. Kira [3594]. Kircher [1196, 2227, 1293, 3433, 3174, 1770, 999, 4090]. Kirchers [1610]. Kirchhoff [2403]. Kirschke [2466]. Kirschner [1941, 2576]. Kirsten [1838]. Kirwan [1495]. Kistemaker [2662]. Kitāb [1226]. Kitàb [353]. kitb [609]. Kites [3295]. Kiyosi [2197]. Klaerich [125]. Klare [508]. Klasse [302, 542, 382]. klassische [173]. klassischen [746, 747]. Klaus [201, 3650, 498, 3289, 2912, 1012, 1083]. Klaus-Dietrich [3650]. Klein [1769, 2900]. kleine [1714]. Kleineidam [251]. Klemme [1989]. Klestinec [3442]. Kline [1349]. Klinge [673]. Kluwer [2708, 2462, 1401, 606, 1570, 1600, 2583, 1857, 1498]. Knellwolf [2601].

Knight [1070, 1991, 1588, 2873, 3114, 2803]. Knobloch [1001]. Knoppers [2589]. know [1505]. Knowing [2227, 4090]. Knowledge [1449, 1819, 2495, 2887, 3282, 2636, 394, 1188, 3651, 3522, 3501, 3552, 2066, 3441, 3622, 3675, 2827, 3404, 3253, 829, 3738, 1935, 3708, 3912, 3490, 1652, 3556, 3492, 2867, 3444, 3634, 3493, 3474, 3588, 2591, 3711, 3479, 3748, 3780, 3787, 3832, 3835, 3070, 3895, 1170, 3540, 371, 1486, 1381, 1989, 2093, 2134, 3096, 2562, 2783, 2749, 3492]. known [424]. Knoxville [438]. Knudsen [2967]. Knut [2304]. Knuuttila [1104]. Københaven [753, 997]. København [914]. Koch [3185]. Kochina [591]. Kockerbeck [547]. Koelbing [639, 564]. Koelliker [3276]. Koenraad [2369]. Koerner [2102]. Kohen [143]. Kohlmann [2831]. Kohn [228, 1462]. Kolbe [1333]. Kolloquien [507]. Kolloquium [1838, 4082]. Köln [2573, 29, 1277, 1358, 914, 1188, 1445, 1686, 753, 749, 997]. Kommentar [1941]. kommentarer [1441, 884]. kommenterade [1019]. kommentiert [64]. Kommentierung [1531]. Kommission [64]. Kommunikationsmedium [1151]. Kompendium [3129]. kön [2500]. Kong [2216, 2268]. Kongresses [1281]. Koninklijke [302, 3018, 2884]. Konkordanzen [747]. Konkordanzes [746]. Konrad [114]. Konsiliarkorrespondenz [3116]. Kontinente [348]. Kontroverse [1069]. Kopaneva [2662]. koperen [1422]. Kordesch [1275]. Kórperwelte. [2265]. Korrespondenz [2529]. Kors [385]. Korshin [385]. Korst [280]. Kort [3082]. Kosmetik [2682, 3857]. Kosmos [1862, 4036, 2986]. Kosmosophie [1086]. Kostas [3243, 1577]. Kostic [2069]. Kounelis [796]. Kovalevskaya [591]. Kováts [1387]. Koyr [2236]. Koyré [287, 2165]. Kragh [533, 1991, 2105, 3263]. Krakow [3920, 1256]. Krämer [1448]. Krankheit [950]. Krankheiten [1391]. Krings [343]. Kristeller [1448]. Kritik [950]. Kritische [364, 1328]. kritischer [567]. kroppen [2500]. Krug [706]. kt [1838]. Kuemmerle [32]. Kuhlemann [141]. Kuhn [2132, 2740]. Kühnken [173]. Kuksewicz [1359]. Kul [1281]. Kullmann [173]. Kulstad [890]. Kultur [62, 2200, 571, 798, 994, 276, 690]. Kultur- [62, 276]. Kultureller [2854]. Kulturgebiet [1396]. Kulturgeschichte [2736, 3130]. Kulturkreislehre [3377]. Kulturmission [690]. Kundiger [725]. Kunitzsch [2055]. Kunkel [2228]. Kunst [1531, 547, 379]. Kunsten [302]. Kunstformen [547]. Künstler [2943]. Kunstverlagsgesellschaft [1012]. Kunze [379]. Kurdyla [356]. Kurt [217, 253, 769]. kurz [751]. Kusukawa [1721]. Kuuliala [3595]. Kyle [3409]. Kynismus [798].

L [89, 438, 685, 804, 1200, 1367, 1818, 2899, 3065, 593, 2012, 340, 336, 493, 421, 1903, 3326, 18, 569, 584, 1006, 3059, 932, 2780, 2580, 2203, 2144, 3025, 847, 1112, 2797, 447, 1188, 1238, 1859, 2010, 2080, 2232, 2224, 2157, 2156, 1909, 1335, 2058, 1908, 1164, 1917, 2751, 2614, 1732, 80, 81, 67, 79, 288, 300, 2773, 1196, 1148, 285, 927, 850, 912, 1943, 2754, 360, 437, 641, 645, 706, 707, 958, 947, 877, 882, 1101, 1115, 1096, 1028, 1092, 1030, 1098, 1431, 1598, 1716, 1809, 1864, 1925, 2006, 2011, 2154, 2220, 2222, 2275, 433, 435, 644]. L [697, 708, 1180, 1718, 237, 428, 563, 589, 556, 575, 559, 970, 953, 1020, 1033, 1241, 1159, 1665, 1671, 1806, 2002, 1923, 2009, 2114, 2150, 558, 929, 417, 1904, 2206, 1071, 771, 2583, 165, 678, 3148, 186, 347, 1075, 1851, 1922, 1899, 1998, 1994, 2107, 2062, 2132, 2429, 2380, 611, 1140, 13, 924, 851, 1988, 370, 280, 555, 376, 2136, 1388, 2100, 1324, 2096, 1990, 1906, 1900, 281, 1728, 1850, 2133, 177, 346, 2772, 353, 1682, 1226, 2620, 1907, 1162]. L. [1132, 841, 1682, 234, 2551, 1908, 640, 1181, 2597]. l’A-cadémie [753]. l’A.H.E.F. [260]. Lab [3499]. Laberinti [2488]. Labirinti [1318, 2159]. Labor [635, 2131, 3968, 2857]. laboratoire [3365]. Laboratoires [2856]. Laboratorien [2915]. Laboratories [3739]. Laboratorio [3184, 2128, 4040, 2224, 1241, 1343, 1424, 1730, 2115]. Laboratory [2302, 3160, 3157, 3185, 2069, 2834, 3968, 2556, 54, 1603]. Laborlandschaften. [2853]. labyrint [1537, 2986]. Labyrinth [1537]. labyrinthe [3180]. Labyrinths [3654]. l’Academie [1476, 1795, 2145]. Lacaita [223, 1759]. l’Accademia [3750]. Lacchini [1417]. lacitt [2461]. Lactea [2458]. Ladies [3738]. Ladnyi [499]. Ladrire [2469]. Lady [3615]. Lægeforeningens [1363]. Lafayette [1288]. Lafleur [512, 1984]. Laghi [1695]. Lagier [2663]. Lagrange [3778, 3101, 2422]. laici [517]. Laico [264]. l’Aigle [2611]. Lain [2601, 882]. Laird [2139]. Laissus [620]. Laity [517]. Läkaresällskapet [696]. Lalande [3450, 2958]. L’albero [371, 2269]. l’Alchimie [2648]. Lalli [1695]. Lama [2104]. Lamarck [3760, 404, 1211, 3136, 3299]. Lamarckien [1041]. Lamarckismo [404]. Lamb [3617]. Lambeck [3341]. Lamberini [2491, 3015, 561]. Lambert [88]. Lamberti [2111]. Lamberto [2111]. Lament [3560]. l’Amérique [434]. Lamp [1792]. l’analyse [3075]. Lancaster [16, 28]. Lancastrian [3449]. Lancelot [1613]. Lanciotti [912]. land [1111, 3199]. Landecker [3035]. Landes [2589]. Landesbibliothek [725]. Landmark [2899]. Landmarks [1561]. Landriani [1947, 2018]. Landscapes [2957, 3431, 3296, 3819]. Landucci [1956, 2027]. Lane [2683]. Lang [547, 183, 1058, 2231, 2804, 2146]. Langage [2428, 198, 4088, 2428]. Langhans [119]. Langins [2664]. Language [30, 58, 1494, 2518, 1982, 1119, 283, 1829, 1563, 4088, 1381, 1119]. Languages [227, 648]. Langue [1232, 3742]. Langues [1224]. Lanmarco [711, 1180]. L’Antica [1134, 2368]. l’antiquité [3251, 2767]. Lanz [1724]. Lanzarini [38]. Laplace [1148, 2537, 3284, 2728, 1654, 2350, 3789, 3795, 3875]. Laplanche [582]. L’apprendista [1565]. Laquidara [711, 1180]. l’Archivio [2614]. lärdomshistoria [2496]. large [1469]. Largs [808, 4032]. L’Arioste [3952, 3257]. Larissa [3298]. L’Aristotelisme [308]. Larry [3674, 1109]. Larsson [2500, 3036]. L’arte [1865, 2338, 2280]. lasciapassare [1107]. Lasser [3509]. Last [3591, 1997, 3223, 3962, 1375]. Laszlo [2627]. lat [1134]. Late [77, 1490, 3162, 1188, 3409, 3656, 3191, 1193, 3564, 3567, 3337, 3595, 3479, 1596, 3018, 3273]. Lateinischen [256, 1446]. Later [2392, 1213]. Laterza [1996, 2320, 2431, 2364, 718, 971, 2224, 2383, 2946, 360, 566, 1030, 1674, 1673, 1675, 1752, 1672, 1753, 1756, 1677, 1676, 1867, 1865, 1866, 1718, 704, 1876, 417, 255, 1075, 2618, 611, 2274, 1223, 1651, 2668, 1390]. Latin [2031, 1611, 1102, 1446, 1448, 1688, 334, 2308, 337, 2686, 384, 376, 1225, 609, 1138, 3748, 3880, 3920, 4026, 4027, 4028, 4023]. latina [1134]. Latino [995, 2644]. Latinorum [1391, 1687]. latins [868, 3058, 1224]. Latinus [1280]. Latitude [2777]. Latitudini [2113]. l’Atlantico [2317]. Lattis [1398]. Laudan [538]. Laude [4060]. Laura [2933, 1726, 3638, 1591, 2221, 1357, 2589, 1084, 2858, 3689]. Laureates [1290]. Lauren [2830]. Laurence [3733]. Laurent [3739, 1440, 3280, 1878, 2603]. Laurenza [2326, 2434]. Laurenziana [1360]. Laurie [860]. Laurioux [2877]. Lausanne [2519, 868]. L’avènement [2375]. L’Aventure [702]. Lavine [1411]. Lavinia [3620, 1530]. Lavoisier [15, 1406, 1846, 1795, 3299, 3368, 3739, 2425, 675, 569, 584, 1440, 1682, 1147, 1588, 368, 3331, 3894, 2198, 2603, 1090, 1132, 3437, 2425, 2350, 3760, 671, 977, 2051, 3160, 3200, 3785, 4065, 3881, 2533, 3894, 2198, 3982, 1381, 1494]. Lavoisiers [1381]. Lavorata [1469]. lavoratori [1186]. Lavori [1615, 1452]. Lavoro [2128, 2220, 1186, 782, 1663, 635]. Law [1613, 3134, 1999]. Lawrence [15, 164, 414, 1846, 3598, 2093, 2834, 2873, 1652, 3313, 163, 2626, 2891, 1839]. Laws [694, 421]. Laylin [1290]. Lazzara [2316]. Lazzaro [1765, 1111, 1569, 1235, 602, 171]. l’cartista [2826]. Lck [2471]. Leading [2605]. Leah [3633, 3114]. Learned [648, 3431, 1972, 3229]. Learning [167, 1081, 703, 3414, 1445, 3205, 919, 3420, 4069, 1563, 3272, 2749, 3288, 2950]. Leary [1362, 933]. Leben [362]. Lecce [2144, 2492, 1369, 2138, 1142, 3967]. Lecce-Taurisano [2492]. Leclerc [518]. l’École [2898]. Lecons [2898]. Lecourt [1414]. lecture [1042]. Lectures [2516, 1608, 492, 4054, 1977, 767]. le‘Disputatione [2942]. Leed [1428]. Leen [2497]. Leeuwenhoek [802, 1367, 3686]. Lefanu [244]. Lefèbvre [295]. Lefèvre [2689, 2900]. Left [3101]. Legacy [2182, 1699, 3394, 2967, 2911]. Legado [963]. legale [3115, 1183]. Legami [3299, 3760]. legazione [858]. Legem [1033]. Legge [346]. Legged [3048]. leggi [421, 1435, 1998]. Legislation [3851]. Legislazione [3851, 1314, 2275]. Legitimacy [3703]. legitimist [40]. légitimiste [40]. Legname [1300]. legno [449]. Lehmann [143, 2922]. Lehoux [2879, 3460]. Lehrer [2466]. Leibnitiana [1141]. Leibniz [516, 583, 1535, 310, 890, 218, 681, 1001, 756, 1141, 411, 1140, 2772, 2578, 3113, 2474, 3646]. Leibniz-Gesellschaft [1141]. Leibniz. [3008]. Leiden [891, 1110, 2573, 2897, 2933, 63, 29, 1277, 1188, 1358, 1445, 1686, 2999, 3734, 2555, 2554, 1422, 1342, 1792, 2195, 535, 753, 749, 997, 1229, 2840, 2939, 3058, 2937, 914, 2556, 4058, 1983, 266, 3823]. Leiden-Boston [2840]. Leiden-Köln [29]. Leiden-New [1277, 1445, 1686, 753, 749, 997]. Leiden/Boston [2933, 2939, 2937]. Leiden/Boston/Köln [2573]. Leinen [1731]. Leinkauf [1610]. Leipzig [303, 497, 568, 698, 1095, 251, 226, 1012]. L’élite [3483]. l’elixir [2420]. Lemaître [4089]. Lemaitre [1734]. Lemery [2145, 3784]. L’Empirisme [3312]. Lena [2119]. l’Encyclopédie [2843]. Lennerhed [2119]. Lenoir [762]. lens [1469, 1704]. l’Enseignement [1984]. lenses [319]. Lente [1173, 1469, 1704, 354]. Lenti [319]. Leo [61, 516, 583, 804, 2488, 2467, 2942, 2425, 3286, 3015, 252, 309, 576, 720, 717, 716, 887, 873, 1157, 1279, 1156, 1360, 1688, 70, 3127, 413, 752, 2955, 874, 881, 3258, 2525, 455, 2574, 799, 800, 1388, 1493, 1389]. Leon [1044]. Leonard [3733, 3727]. Leonardi [1156]. Leonardiano [3172]. Leonardo [573, 1950, 2021, 340, 3544, 2598, 858, 1107, 1159, 1424, 1394, 3632, 2691, 1324, 2326, 688, 3195, 3009, 3372, 2769, 3593, 1774, 3014, 3432, 590, 2722, 1767, 3628, 1783, 3629, 2277, 3613, 3596]. Leoncini [697]. Leone [2712, 2085]. Leonello [764]. Leong [3493]. Leonhard [2905, 2041]. Leoniceno [887]. Leonid [2822]. Leopardi [1157, 2501]. leopardiani [2501]. Leopoldina [1992]. Leopoldo [589, 197]. Lepenies [1066]. Leprosy [2875]. l’Erbario [2882]. l’Ermeneutica [2720]. Lerner [3004]. Lesley [1449, 3441]. Leslie [3265]. l’Esperienza [2429]. L’esprit [1460]. L’esprit-cerveau. [1460]. Lessico [459, 310, 306, 1157, 1388, 2571]. Lessons [2684]. L’età [202, 1564]. Letizia [893, 3216]. Letouzey [1418]. Letter [1037, 2156, 212, 2786, 2695]. Lettera [400, 1037, 2172, 1451, 2497, 2141]. letterari [180]. letteraria [2167, 2238]. letterario [2632]. Letteratura [894, 3279, 3001, 2157, 2720, 4085, 1064, 764, 1314, 312, 915, 996, 1101, 2416]. Lettere [516, 1691, 2467, 665, 1303, 2168, 2239, 2169, 2240, 2253, 2254, 2354, 1705, 1817, 2882, 72, 250, 716, 1077, 2248, 1603, 2091, 752, 1302, 3168, 1782, 1523, 1521, 1195, 2112, 982, 1658, 2380, 12, 1119, 1728, 887, 3838, 999, 2132, 4083, 2602, 2787, 799, 412]. Letteree [555]. Letteren [302, 541]. Letters [802, 1081, 1367, 2354, 3328, 2906, 955, 1707, 2080, 3076, 2091, 2259, 2361, 1732, 4068, 3490, 1008, 982, 541, 1705, 3762, 3768, 2248, 2215, 3838, 787, 1618, 4083, 849, 1119, 4018, 665, 3875, 2047, 1915]. lettore [3863, 2372]. Lettori [1622, 3783, 2213]. Lettres [3013, 386, 754, 2892, 3024, 868, 3287, 216, 2952, 2469, 3004, 1224, 3768, 620, 2215, 2118, 516, 893, 921]. Lettura [1884, 1499]. Letture [2170, 2289, 1452]. leur [3026]. leurs [3448]. Leuven [3205]. Levante [2049, 2399]. Level [820]. Leven [2643]. Leviatano [1533]. Leviathan [168]. Levine [520]. L’évolution [2318]. Levrotto [1018, 1165]. Levy [2646]. Lewis [2423, 811]. Lexicon [698]. Lexika [746, 747]. Lexika-Indizes-Konkordanzen [747]. Lexika-Indizes-Konkordanzes [746]. Lexikon [2750, 2643]. Lexis [1411, 1923]. L’expédition [2118]. Leyde [1225]. Leyden [2771]. Leyden-Boston [2771]. Lezioni [931, 1569, 2307]. l’homme [2896]. Lhote [921]. li [1679]. Lia [3670, 1794]. Liana [1268]. Liba [3438]. Libavius [1531, 3173]. Liber [567, 59, 3408, 3845, 609, 799, 2520, 2571, 3593]. Libera [444]. liberalism [3083]. Liberi [3181, 2468]. Libero [2488, 677, 1998, 2159]. Libert [1737]. libertà [688, 655]. libertinage [3936, 2309]. Libertine [891]. Libertines [3803]. ’libertinisme [2492, 3967]. libertins [3803, 1875]. Liborio [2557]. Libr [2767]. Libra [1011]. Librairie [1795, 198, 512, 2821, 1228, 3527, 2578]. Librandi [2880]. Libraria [3257, 2107, 3528, 1816]. Libraria-Paris [3257]. Librarie [795]. Libraries [2227, 4090, 14]. librario [1062]. Library [491, 1055, 1078, 1805, 696, 699, 1439, 1518, 3282, 28, 955, 2245, 2311, 2600, 2999, 1196, 1841, 3905, 3917, 2617, 3062, 2998, 1546, 1252, 945, 1783, 180, 1219, 2248, 3905]. Library/Oxford [2600]. libraux [2767]. Libre [353, 1879]. Libreria [1790, 2667, 1077, 2001, 2307]. Libres [1386]. Libri [903, 1589, 552, 3286, 1823, 2614, 1973, 2202, 13, 1644, 59, 3801, 72, 887, 951, 2712, 1015, 1645, 1229, 3632, 3364, 3824, 2995, 2614]. libris [2716]. Libro [1950, 2021, 393, 1539, 2881, 2517, 2954]. Libros [889, 2114, 3942, 494, 618]. license [1944]. Liceo [828, 1620, 1928, 4076, 776, 1258, 1241, 1343, 1424, 1241, 1343, 1424]. Lidia [2014]. Lidio [3578]. Liebersohn [2846]. Liebig [259]. Liebigs [2466]. Liège [4071, 4087]. Lienhard [2785]. Lies [1112, 3831]. Lieshout [280]. Life [15, 168, 1234, 1449, 1589, 1853, 1910, 1852, 3757, 3615, 3055, 703, 3268, 3035, 1401, 2751, 3191, 1161, 2792, 3958, 347, 1698, 2345, 3219, 766, 2378, 723, 1993, 2549, 3447, 1225, 2534, 3052, 2772, 2971, 1537, 3760, 282, 3786, 3843, 3844, 169, 3891, 3903, 3928, 1877, 611, 362, 3969, 3971, 3972, 3979, 803, 3981, 1655]. lifetime [1814]. Lifter [2564]. Light [1776, 3318, 3555, 3405, 732, 3189, 2296, 3208, 3333, 3985, 1999]. lighting [541]. Lightman [3637, 3033]. lightning [2330]. Lights [4053]. Ligozzi [1269, 1269]. Liguori [1908, 79, 1522, 2659, 417, 377]. Liguri [912]. Liguria [2363]. Lii [1417, 1289]. Likeness [3723]. Lili [1745, 1166, 1167]. Lilia [861]. Liliana [2224]. Lille [582]. Lillian [860, 1249]. l’Image [1155]. Limited [405, 916, 1320, 1877, 2098]. límites [3427]. limiti [2276]. Limits [3347, 1997, 1038, 3962]. Linati [2156]. Linceans [3198]. Lincee [3687]. Lincei [1689, 311, 1741, 2713, 2741, 1873, 2941]. linceo [1689]. Lincoln [2551]. Linda [3817, 285]. Lindberg [801, 993]. Lindeboom [57]. Lindee [2671]. Lindgren [23]. Lindner [3037]. Lindqvist [1264, 3092, 2456]. Lineamenti [2321, 1335]. Linear [206, 2360]. Lineare [1785]. Linee [2865]. Lines [3575, 3668]. L’inferno [3498]. l’Infini [1668]. L’Influsso [515]. Ling [406]. Lingua [1829]. linguaggi [969, 3062]. Linguaggio [2996, 1119, 569, 2269]. Linguerri [3125]. Linguistic [3440, 1564]. linguistici [1564]. L’inizio [447]. Link [945]. Linked [3449]. links [3760]. Linn [222]. Linnaeus [2527, 2102, 3899]. Linnaeus. [1454]. Linnean [1121]. Linnéenne [1121]. Linneo [2125]. L’insegnamento [1105]. Lint [74, 448]. lion [2311, 3917]. Lions [1241]. Lipomo [1813]. liquidorum [1960, 2031]. liquids [2031]. Lire [1447, 2308, 200, 930, 2269]. Lis [2155]. Lisa [3636, 2724]. Lisbet [2102]. Lisboa [1744]. Lissa [3018]. Lisse [802, 1367]. List [1448, 160, 5]. L’Istituto [3750, 594, 3969]. Lists [3641, 3659, 3677, 3697, 3720, 3743]. Liszt [1571]. LIT [3253, 1201, 1190, 330, 3130, 1713, 1650, 1767, 1796, 2108, 1642, 1868]. Literal [283]. Literary [805, 1457, 2247, 180, 2354, 2354]. Literatur [1079, 695, 2750]. Literature [227, 520, 405, 1457, 3205, 3578, 2945, 2737, 10, 4085, 4000, 2906, 3044, 3656]. Lithophylacii [3434]. Litografia [809, 1084]. Litten [1501]. litteraria [966]. litteratorum [1145]. Litteratura [1918]. littérature [1859, 3491, 3010, 4000, 3742]. Little [1346, 2198]. Littleton [2873]. littraire [198]. Liv [1873, 1537, 217, 1870]. live [1744]. Liverpool [1818]. Lives [440, 3616, 1562, 3434]. Livia [2215, 1158, 556, 1158]. Liviana [645]. Living [1225, 1753, 2350]. Livingstone [2805, 624, 2562]. Livio [240, 1114]. Livorno [2338, 2380]. Livre [2814, 2821, 1646]. Livres [3163, 2952]. Livro [2198]. Ljungström [2675]. Llorente [2571]. Lloyd [1683, 442, 998]. Llull [249]. Lob [542]. Lobacevskiana [2444]. Lobeiras [2146]. Local [2827, 2749]. locale [1260]. Localismo [744]. Locality [2106, 3830]. Localization [3375, 3933]. Localizzazioni [1467, 2336, 1917, 2287, 3933]. Locations [405]. loci [2321]. Locke [60]. Locus [1660]. Locy [1133]. Lodovica [1539]. Lodovisi [1806]. Loeb [945]. Loeber [111]. Loffredo [1767]. Loge [8]. Loggia [2616, 1975]. Logic [30, 58, 444, 1982, 3581, 1225, 2146]. Logica [58, 444, 85, 3443]. logicaa [1105]. Logical [1102]. logiche [1148]. Logie [3797, 2381]. Logik [379, 2146]. Logique [2573, 1232, 444]. logisch [443]. Logistics [3218]. Lohr [1687, 1688]. lointaines [3245]. Lojacono [1364]. Lomagno [1271]. Lombardia [1373]. Lombardo [2116, 1871]. Lombrichi [2349]. Lombroso [897, 2157, 3546, 3085, 2351, 3802, 3927, 2426]. Lonard [175]. Londa [2657]. London [11, 77, 167, 224, 301, 342, 454, 491, 439, 440, 414, 405, 580, 587, 642, 651, 579, 724, 959, 968, 1106, 1108, 1231, 1234, 1313, 1346, 1449, 1404, 1692, 1853, 1945, 2016, 2120, 2102, 2056, 2194, 2424, 2560, 2601, 2793, 2887, 2953, 2708, 2710, 2881, 2462, 1573, 1659, 2199, 532, 1006, 1439, 1518, 3059, 3068, 3072, 2805, 847, 1808, 10, 3204, 3030, 194, 511, 964, 830, 1102, 1103, 1104, 990, 349, 538, 1143, 1566, 1717, 2103, 1067, 1238, 2080, 2216, 2381, 614, 334, 3043, 3012, 2655, 219, 2719, 2948, 408, 2885, 3070, 2800, 3281, 341, 767, 1246, 1409, 1732]. London [1398, 1722, 2308, 2783, 225, 286, 2139, 2268, 2781, 3079, 2862, 3033, 1841, 65, 191, 418, 770, 927, 2966, 3542, 1182, 1913, 2775, 2944, 2961, 427, 939, 992, 1161, 1570, 1872, 2310, 1600, 1504, 610, 2858, 3038, 2584, 1071, 2976, 833, 1319, 2886, 2617, 2676, 2844, 3118, 791, 2841, 347, 1786, 1980, 2366, 2412, 2559, 2985, 2807, 2830, 674, 1400, 1723, 2784, 2626, 3126, 1408, 1541, 289, 691, 1144, 1571, 3032, 1733, 387, 388, 3261, 378, 1837, 3278, 2846, 2956, 945, 2096, 2739, 1577, 177, 2772, 68, 2562, 1914, 1839, 2417, 2303, 3196, 1652]. London [2683, 4013, 2816]. London-Bronx [10]. London-New [405, 724, 2417]. London/Chicago [2194]. London/New [3059, 2366]. London/Turin [2308]. Londra [1164]. Long [3254, 3714, 2832, 1306, 2131, 3060, 2973, 1865, 3919, 3214, 3479]. Longanesi [2108]. Longinos [3218]. Longitudine [1650]. Longo [2588, 2727, 1248]. Longobardo [740]. Lonie [279]. lontano [2206, 3965]. look [772, 2068, 3990]. Looking [3151, 2940]. Lopez [45, 1226, 43, 2101, 1849, 2114, 2097, 3327]. Lorch [2055]. Lord [888, 689]. Lorella [2773, 1015]. Lorena [804, 619, 1538, 389, 455, 589, 844, 1451, 2141]. lorenese [517, 340]. Lorenz [3116]. Lorenzo [1958, 2029, 1028, 1215, 2395, 1028, 3951, 2272]. Lorgna [412, 1219]. L’Origine [486, 2580]. Loris [307, 716, 1175]. lornetki [1584]. loro [1956, 2027]. Lorraine [517, 3428, 2684, 2559, 2983]. L’Ortobotanicodi [1487]. Losano [2313]. L’ospedale [1335, 1597]. L’Osservatorio [3847]. l’Ossesione [2677]. lost [946]. lotta [2946]. Louis [1495, 598, 866, 2013, 518, 2556]. Louis-Bernard [1495]. Louis-Francois [598]. Louise [3715]. Louvain [1442, 1859, 3091, 1860, 2369, 1734, 4089, 4091, 4089, 4017]. Louvain-La-Neuve [1734, 1442, 1859, 3091, 4089, 4089]. Louvre [3257]. Lovaina [4017, 1860]. Love [591, 3475, 348]. Loved [959]. Lovelock [2472]. Lovers [959, 1081, 3357, 3358, 3359]. Low [3369, 3501, 3835, 3253, 3912, 14, 761, 916, 2975]. Lowell [2450, 723]. Lowith [2432]. Lowood [182]. Loyola [2686]. Lozeran [1681]. Lt [2154]. Ltd [2777, 652, 2864, 847, 939, 833, 1408]. Lublin [1281]. Lubrina [715, 719]. Luc [2648, 2785, 2867]. Luca [1284, 1856, 2380, 3107, 447, 718, 719, 3209, 1923, 2796]. Lucas [1370, 2058]. Lucca [70, 776, 1768]. lucchese [1768]. Lucchesini [171]. Luccio [2320]. Luce [1550, 2041, 1486, 2810, 2622]. luces [2808, 2808]. Lucia [2012, 2530, 448, 1481, 863, 917, 953, 1614, 2886, 1269, 1755, 1461, 2368]. Luciana [1283, 1461, 2845]. Luciani [2336]. Luciani/Giuseppe [2336]. Luciano [2612, 1540, 1996, 1189, 2215, 877, 1026, 1434, 1075]. Lucio [588, 2156, 2646]. Lucky [3381]. Lucretius [2936, 3109, 3863]. Lucrezio [1322, 3863, 2372]. Ludmilla [724, 227, 544]. Ludolph [99]. Ludovico [1948, 2019, 3121, 2088]. Ludus [1611]. Ludwig [629, 106, 144, 109, 691, 153, 104, 112]. Luft [1405]. luglio [844]. Luigi [764, 1234, 1539, 2336, 310, 1516, 1869, 2270, 3345, 752, 288, 407, 1521, 727, 1532, 558, 1426, 840, 41, 2334, 4052, 2315, 2504, 365, 366, 2965, 641, 2373, 2713, 1136, 2078, 969, 1856, 2460, 2714, 1900, 2127]. Luigina [1642]. Luis [3005, 1855]. Luisa [3040, 1820, 3820, 1029, 2104, 2324, 3094, 3820, 1020, 3214]. Luise [2014, 1585]. Lukas [3572]. Luke [2875]. Lukoschik [2787]. Lull [610]. lulliana [1565]. Lumen [3146]. Lumi [2536, 2690, 1491]. lumière [1460]. Lumières [1985, 2958, 2118, 2843, 3767, 4053, 3069, 3180, 3330]. Lun [2095]. Luna [1279, 525]. Lunar [1468, 3518, 3600, 2248]. Lunare [2248]. lunario. [2153]. Lund [1441, 884]. Lundbeck [235]. lunettes [3071]. lunety [1584]. lunga [1865]. Lunger [2589]. Lungo [2131]. Luni [1796]. L’università [3741]. l’Université [2878]. Lunwerg [686]. Luoghi [2254, 658, 1016]. luogo [2057]. l’Uomo [2503]. Lupacchini [3481]. Lupato [2136]. Luperini [2974]. Luppi [1011, 1163, 1419, 1806]. Lurija [771]. Lussana [2415]. Lustig [2791]. lustració [3019]. Lusverg [1631]. Lute [3643]. Lutero [2159]. Lüthy [3253]. Luxograph [514]. Luz [2101, 2097]. lv [1238, 2202, 727]. LVI [2114, 653, 808]. Lvii [1228, 3283, 2937]. Lviii [1157]. Lx [1001]. Lxi [1388]. Lxiv [407, 2388, 1904]. lxix [2852, 955]. LXV [1686]. lxvi [1000]. lxvii [3064, 3283]. lxx [746]. lxx-viii [746]. lxxii [1731]. lxxix [1486]. lxxvi [2937]. lxxx [2781]. Lxxxv [12]. lxxxviii [2429]. Lyall [856]. Lydia [3543, 3086]. Lydon [1933]. Lyell [1654, 3796]. Lynceorum [3687]. Lynn [2901]. Lynne [2841]. Lynteris [3572]. Lynx [2371, 3855]. Lyon [954, 713, 3245]. Lyonnaise [704].

M [959, 1005, 3598, 1239, 726, 934, 2933, 3475, 2891, 643, 2364, 235, 2671, 990, 1278, 1188, 1156, 1445, 279, 2553, 1263, 2834, 2873, 45, 56, 173, 3415, 2819, 1219, 2614, 1656, 767, 1398, 174, 286, 2004, 639, 788, 869, 1674, 1867, 2152, 564, 1100, 3313, 452, 1584, 1655, 534, 2753, 753, 1737, 2803, 13, 689, 860, 849, 1198, 2967, 1733, 1142, 1388, 1324, 747, 1224, 1225, 2130, 1920, 627, 2974, 2130]. M. [1476, 934, 814, 3122, 1188, 1445, 38, 1475, 785, 872, 1594, 1753, 1866, 2222, 616, 1718, 237, 1497, 800, 13, 803, 1388]. M.A [1408]. M.D [16]. M.F.N [75]. MA [1264, 1365, 2160, 1576, 3035, 2553, 1494, 2966, 2844, 2866, 3278, 2096, 2456, 2913, 722, 1657, 495, 1097, 2900, 457, 2911, 2631, 2313, 1227]. maîtres [1444]. Maar [2198]. Maarssen [582]. Maccagni [1284, 1532]. Macchia [893]. Macchie [2178, 1119]. Macchina [1710, 449, 2651, 1389]. Macchine [2523, 1777, 1827, 552, 3172, 3541, 407, 355, 1259, 558, 2148, 1634]. Maccioni [1589]. macellai [2266, 4038]. Macerata [1101]. Macgregor [2961, 2204]. Machiavelli [776]. machina [355, 459]. Machina. [2809]. Machinae [1544]. Machine [3047, 2553, 1162]. Machines [2689, 3280, 4063, 355]. machoires [707]. macht [2715]. Macii [700, 776, 1014, 1257, 1807]. Macini [1317]. MacIntosh [2873]. Mackensen [234]. Maclean [2716]. Macleod [3083]. MacMillan [532]. Maconi [947]. Macquer [3236]. macroseismic [519]. Madame [2533, 3200]. Maddalena [2503]. Madders [1999]. Maddison [1563]. Made [735, 1203, 4010, 840, 4012, 3447]. Madeddu [2324]. Madhouse [3505]. Madison [15, 520, 1457, 229, 348]. Madness [225, 342, 177]. Madre [2091, 211]. Madrid [889, 965, 686, 1197, 1515, 35, 2876, 2494, 352, 963, 1544, 1860, 1645, 494, 618, 3172, 1299, 3489]. maestranze [2523]. Maestri [2368, 873]. maestro [3121, 2130]. Maffeo [2379]. Maffi [546]. Maffioli [2491, 3255, 1791, 1519]. Magalotti [668]. magazine [2737]. Magellan [3475]. Maggio [514, 1430, 4067, 844, 1813, 2005, 2217, 83]. maggio-31 [844]. Maggioli [2014]. Maggiori [3280]. Magherini [952]. Magi [2064]. Magia [2192, 604, 310, 2949]. Magic [842, 1106, 2056, 3072, 356, 2830, 2816, 3291, 3869, 2561, 3924]. magica [3443]. Magiche [2192]. magie [3251, 3924, 1646]. Magiotti [1874]. magiques [2519]. magische [543]. Magister [601]. Magistra [2639]. Magliabechi [555]. Magnetic [1725, 2777, 856, 3841]. magnetica [1166]. Magnetism [2168, 2239]. magnetismo [75, 2168, 2239]. Magneto [526]. Magneto-Electric [526]. Magnificent [1518]. Magnifico [1028, 3951]. Magnifying [3105]. Magnino [1145]. Magno [2645]. Magnolfi [958]. magrūsāt [1226]. Magusi [1358]. Mahmoud [1434]. mai [2878, 4042, 1476, 1440, 1327, 314, 1224]. Maiden [3145]. Maiello [300]. Maierù [1277, 915]. Main [2979, 547, 415, 183, 363, 1999]. Maine [2499]. Maines [2016]. Maintain [3133]. Mainzer [498]. Maio [2151]. Maison [314, 419]. Maistre [1707]. Maius [1256]. Maja [2500]. Major [1907, 3884, 1915]. Majorana [3080]. Maker [2442, 3320, 1161, 1309]. Makers [985, 63, 843, 3472, 3176, 2888, 3616, 2913, 1808, 2310, 4013, 4031]. Makes [3225, 2862]. Making [3187, 1393, 2607, 3322, 1991, 3573, 3522, 2838, 3675, 3594, 2600, 52, 3342, 761, 3333, 3446, 2485, 4075, 762, 3985, 4013, 3070, 3108, 2747, 2975, 3241, 2310, 3734, 3478]. Malacrida [1028]. Maladetti [2266]. maladies [331]. Malady [2875]. Malanima [449]. Malaquias [3475]. malaria [1963, 1925, 2034]. Malaspina [2727]. malattia [2203, 2721, 2797, 3864, 300, 1099, 1677, 712, 2624, 3951, 729, 1028]. malattie [365, 366, 331, 1099, 712, 1523, 1728]. Malcolm [2717, 2779]. Male [3108, 300, 1677, 1867, 2726]. Maledetti [4038]. malerbe [3873, 1177]. Malet [2099]. Malinowski [229]. Mallet [519]. Malpighi [61]. Malvezzi [727]. mammelle [1523, 1728]. Man [1454, 2424, 1044, 1203, 3834, 1847, 3975, 4012, 3014]. Manacles [342]. mancante [1928, 1901]. Mancha [3005]. Manchester [283, 759, 17, 683, 2901]. Mancini [277, 1354]. Mancosu [1451]. Mandelbrote [2311, 2755]. Mandrino [71, 291]. Manduria [223]. Manetti [2622]. Manfred [622]. Manfredini [956]. Manfredonia [237]. Manganaro [284]. Mangani [2773, 1663, 1906]. Mangianti [1525]. Mangle [1504]. Manhattan [2546]. Manhood [1169]. Manicomi [3181]. maniera [1174]. Manlio [1746]. Mann [1004, 1918]. Manna [3213]. Mannelli [2616, 680, 1850, 3089]. Manners [80]. Manning [3442]. Mannoni [1834]. Manns [1918]. Mano [2126]. Manoscritte [1624, 1772, 1882, 31, 165]. manoscrittee [1314]. Manoscritti [2242, 2341, 2438, 602, 1546, 1252, 668, 2974, 28, 720, 991, 1102, 1279, 1156, 70, 1056, 44]. Manoscritto [1976, 2258, 2592, 67, 3934, 1645]. Manovra [641]. Mansell [405, 10, 1164, 1071]. Mantegazza [931]. Mantovani [2616, 1343]. Manual [1489, 3485, 939]. Manuale [2155, 948]. Manuel [3475, 2473, 2808]. Manuela [555]. Manuels [1442]. Manufacture [2085]. Manuscript [1448, 3154, 3440, 2814, 3236, 3913, 2204, 3205, 1976, 2258, 3934]. Manuscripts [1439, 990, 1830, 2873, 2781, 2181, 1546, 1252, 3345]. Manuscriptum [1138]. manuscrits [1447, 609]. Manuzio [1283]. Many [2873, 3059]. Manzi [3693]. Manziana [2142, 949]. Manzini [1111]. Manzoni [2220]. Manzuoli [189]. Map [3583]. Maple [1814]. Mappe [265, 574, 1641, 2493, 844]. Mapping [3054]. Maps [3385, 3745, 2176, 1872, 1152, 3600]. Mara [2436, 2101, 1849, 777]. Marage [1879]. Marat [586]. Marazzini [1801]. Marble [3353]. Marburg [64, 256, 2544]. Marburger [183]. Marc [2888, 681, 2545, 3641, 3659, 3677, 3697, 3720, 3743, 300, 2856, 2539]. Marc-Auguste [2539]. Marcacci [3457]. Marcaida [3327]. Marcard [102]. Marcel [713, 1663]. Marcello [1606]. March [2075, 2338, 2161, 2156, 2311, 4066, 4076, 3917, 3800]. Marchant [3060]. Marche [20, 1663]. Marchetti [1322]. marchi [1151]. Marchini [1056, 1435]. Marchis [1961, 2032, 407, 1020, 1018, 558]. Marci [680]. marciani [1056]. Marciano [2933]. Marco [1589, 2488, 3739, 1949, 1957, 2020, 2028, 1967, 2038, 1947, 1965, 2018, 2036, 2166, 2237, 2435, 2436, 2620, 3544, 2904, 1599, 2117, 2499, 2603, 3416, 3632, 1615, 1055, 2942, 814, 2685, 2670, 1147, 3022, 2955, 3029, 2654, 2630, 2904, 3590, 656, 2078, 1712, 2809].

Marcon [3492]. Marcone [2642]. Marconi [2523, 1741, 1733, 3860]. Marcovecchio [1173]. Marcus [680, 3651, 2585]. Mardaga [2428]. mare [807, 1719]. Maréchal [3564]. Marelli [1756]. Margaret [440, 3653, 2362, 77, 164, 1427, 2755, 190, 2807, 2457, 892]. Margarete [259]. Margarida [1421]. Margine [1379]. Margolin [1228]. Maria [2120, 2387, 3570, 1784, 3116, 1903, 1488, 3040, 221, 1778, 4050, 223, 74, 307, 951, 3672, 1143, 3177, 1335, 2895, 2637, 2871, 2965, 641, 709, 3546, 554, 3693, 277, 776, 1062, 1185, 1354, 1525, 1768, 1695, 2115, 1712, 2104, 2429, 2380, 3343, 1988, 1857, 3315, 1064, 2205, 2954, 1325, 2101, 3554, 2097, 2969, 2665, 1820, 43, 3820, 2783, 1029, 1185, 1666, 1745, 2077, 3008, 1788, 2429, 2903, 2902, 1352, 745, 2146]. Maria-José [43]. Mariacarla [2682, 2863, 3130]. Marian [2554]. Mariangela [1754]. Marianne [1019, 2826]. Mariano [1515, 2839, 2595, 1794, 2130, 2130]. Mariapaola [757]. Mariateresa [718]. Marie [2933, 3461, 3331, 3894, 1934, 3010, 3737, 959, 1578, 2770, 919, 2615]. Marie-Anne [3331, 3894]. Marie-Claude [2770]. Marie-Noëlle [3461]. Marietti [189, 894, 2141]. Marietti-Manzuoli [189]. Marika [1264]. Mariken [2986]. Marilyn [445]. Marin [264]. Marina [1528, 2154, 700, 1591]. Marine [1045, 2553]. Marinella [937, 2110]. Mariner [424]. Marinha [1744]. Marini [1794]. Marinus [2418]. Mario [2389, 1322, 2140, 412, 1903, 2321, 81, 65, 1980, 2412, 2107, 1400]. Marion [3116]. Mariotte [1232]. Mariotti [1268]. Marisa [2826, 3495]. Marja [575]. Marjorie [419]. Mark [3695, 3417, 3634, 3733, 1854, 890, 3727]. Marketing [3241]. Marketplace [3033]. Markets [2888]. Markey [3670]. Markku [1649]. Markowski [800]. Marks [2382, 1800]. Markus [3649, 1602]. Marne [1681]. Marona [1256]. Marquez [648]. Marquez-Villanueva [648]. Marr [2801]. Marri [1028]. Mars [516, 32, 502, 3053]. Marsala [929]. Marsden [2604, 2054]. Marshall [829, 1729]. Marsigli [1234, 1234]. Marsilio [794, 852, 1947, 2018, 807, 1487, 2491, 958, 730, 1064, 1565, 1324, 1540, 799]. Marsilius [1188, 1281, 28, 1281]. Marsilius-von-Inghen-Kongresses [1281]. Marta [3689, 1516, 2934]. Martelli [2590, 3277]. Martels [753]. Martin [918, 3256, 3251, 1016, 2529, 3687, 2601, 2103, 1758, 2756, 3014, 2365, 2785, 2985, 348, 2466, 1485, 565]. Martine [2856]. Martínez [3372]. Martini [2773, 3741]. Martino [180]. Martins [2845]. martiri [1956, 2027]. Martoglio [83]. Martullo [237]. Martyr [2935]. Martzloff [536]. Marvellous [3014]. Marvels [1510]. Marvin [1248]. Mary [2625, 1500, 3003, 2706, 2377, 26, 992, 545, 3943, 1541, 2582, 2724]. Maryland [2794]. Maryvonne [2596]. Marzetti [1502]. marzo [572, 1741, 4076, 2581]. marzo-30 [1741]. Masaccio [2339, 460, 2360]. Masani [1713]. Mascagni [1653]. Mascarino [2631]. Mascheroni [2395]. Masculine [3108]. Masculinity [1169, 3727]. Masedu [2620]. Mass [679, 672, 842, 1204, 1454, 2120, 2102, 2793, 1898, 1911, 2664, 1552, 349, 1143, 1147, 2059, 2838, 3060, 422, 3243, 1274, 2731, 347, 2728, 2806, 2591, 2846, 945, 2674]. Mass.-London [1143]. Massa [726, 711, 2670, 726]. Massachusetts [2730, 2656, 845, 163]. Massie [3059, 2417]. Massimi [1375, 3858]. Massimo [2461, 912, 1115, 2002, 1811, 310, 635, 3096, 2968]. Masson [2770, 1581, 536, 1809, 1864]. Massoneria [1427, 1491]. master [2130]. Mastroianni [1693]. Mastrorosa [2416]. Matassi [926, 1615]. Matematica [2043, 940, 1347, 2444, 1780, 1706, 1207, 1973, 3481, 2258, 298, 859, 1514, 3286, 2270, 3824, 3866, 556, 1284, 186, 1348, 1667, 3426, 2433, 2503]. Matematiche [22, 1516, 3408, 2191, 2380, 1379, 3838, 3915, 4083, 3255]. Matematici [2871, 1780, 3031, 1311, 1350, 2881]. Matematico [2356, 76, 12, 3286, 2577, 3824, 1369, 1349]. Matematico-Fisiche [76]. Mater [1811, 448]. Matera [2880]. Materia [1805, 3562, 1550, 914, 1200, 315, 3811, 2645, 2198, 3527]. Material [2262, 3222, 3531, 3253, 3468, 3912, 3536, 3835, 2716, 3616]. materiale [1101]. Materiali [778, 1861, 1931, 2149, 2414, 1456]. Materialien [183, 338]. Materialismo [417]. Materialist [1095]. materialista [2668]. Materials [3484, 3360, 778, 2076, 2900]. materie [29, 30]. Mateu [2907, 3770]. Mathemata. [23]. Mathematica [3283, 3298, 556, 4026, 2114, 3941, 2647]. mathematicae [4028]. Mathematical [2086, 3204, 1515, 3630, 2084, 2095, 2258, 3533, 674, 1723, 2457, 1213, 2816, 3587, 3755, 3941, 3441]. Mathematician [2935, 2356, 40, 3824, 2717]. mathematicians [1350]. mathématicien [40]. mathematico [1647]. Mathematics [2459, 3171, 1443, 1515, 3630, 1149, 2800, 3079, 1299, 1640, 2987, 1511, 2833, 3711, 3235, 1225, 4026, 3755, 591, 3824, 3838, 4083, 1089, 1348, 2433, 4028, 4037, 2708, 3000, 3866, 3712]. Mathematik [296, 2055, 1912, 1069]. mathématique [2214]. Mathématiques [620, 2199, 3483, 2197, 4037, 536]. Mathematisch [542, 1083, 1012]. Mathematisch-physikalische [542]. mathematisch-physikalischen [1012]. Mathematisch-Physikalischer [1083]. mathematische [1001]. Mathematischer [411, 1001]. Mathematizations [2099]. matheseos [4028]. Mathildenhöhe [363]. matière [2821]. Matire [1981]. Matocchi [456]. Matricon [701]. Matteo [3277, 2150, 3436, 3282, 3224]. Matter [3347, 1500, 1147, 3525, 2687, 3584, 3640, 3585, 3696, 3642, 3698]. Mattern [3110]. matters [35, 2944]. Matteucci [1895]. Matthäus [1298]. Matthew [3635, 2803, 3017, 1814, 2772]. Matthews [3016]. Matthias [106, 3905, 3905]. Matti [79, 1522, 3373, 673]. Mattias [972]. Mattioli [3844, 1840, 2581, 1267, 1267]. Matton [1280, 1228]. Maturana [371]. Maude [3477]. Maugini [1252]. Maupertuis [338]. Maura [3713]. Maurice [625, 1406, 1568, 3023, 2885, 1987, 3011]. Maurizio [1952, 2023, 2337, 2389, 1105, 2288, 1570, 1371]. Mauro [2009, 371, 2141, 2726, 2721, 2720, 2009]. Mauroliciano [2635]. Maurolico [2172]. Max [895, 3282, 3082, 2831, 860, 2207, 684]. maxillary [4008]. Maximilian [2189]. Maximus [1826]. Maxwell [849, 4068]. May [1476, 1440, 2852, 4054, 4067, 3092]. Maya [427]. Mayhew [3366]. Maynard [1502, 2916]. Mayneri [1162]. Mayr [758]. Mazarin [2015, 2013]. Mazumdar [1656]. Mazzarello [2157, 2535, 3546, 3498]. Mazzini [1101, 2011]. Mazzoli [1156]. Mazzolini [1928, 3226]. Mazzotta [1029].

Mazzotti [3096]. MBCA [1176]. mbH [1012, 348, 1095]. McBurney [3412]. McCalman [2601]. McClellan [2495, 2267]. Mcconnell [1161, 1877]. McCord [445]. Mccormick [14]. McCormmach [2606, 1819]. McEvoy [1685]. McGregor [3412]. McGuinness [629]. Mcintosh [1110, 204]. McKnight [1108]. Mclean [3017]. McMullin [2774]. McOuat [3674]. Mcvaugh [3006, 1392]. Meade [1249]. Meadows [2617]. Meaning [490, 2763, 440, 388]. Meanings [3504]. Means [3922, 2382]. Measure [1945, 428, 2069]. Measured [2625]. Measurement [857, 2681, 3478, 3745, 3423, 2525, 1872]. Measurements [1044]. Measurers [1511]. Measures [857, 3712, 3813]. Measuring [1743, 3833]. Mecacci [1996, 771, 1075]. mécanique [920]. mécaniques [502]. mécanismes [855]. Meccanica [75, 1369, 18, 2140, 3911, 3481, 1323]. meccaniche [854]. meccanismo [371]. Mechanical [2478, 1776, 3210, 3459, 226, 1703]. mechanick [916]. Mechanics [2139, 2970, 2825, 2296, 3038, 3272, 1659, 3830, 3911, 2106, 2139]. mechanism [371]. Mechanisms [3524]. Meckel [117]. med [1441, 884]. Medal [3468]. Medals [3467, 3472]. medea [1753]. Médecin [1476, 1040]. Medecine [2658, 3491, 2878, 836, 1187, 3829, 3895, 332, 434, 2375, 2060]. Mededelingen [302]. MedGG [3116]. MedGG-Beihefte [3116]. Media [1552, 1312, 2454, 1979, 2190]. Mediaeval [376]. Mediaevalis [1280]. Médiathèque [2188]. médiation [2550]. Mediävistischen [1687]. Medica [1805, 3562, 1811, 639, 867, 882, 1809, 432, 3241, 567, 2620, 2721, 3811, 1434, 1752, 564, 1056, 3527, 882, 914]. Medical [224, 759, 837, 1078, 1231, 1903, 2512, 3149, 2909, 279, 3558, 614, 2514, 2978, 57, 3594, 3185, 3556, 3727, 2998, 3343, 381, 2498, 3170, 676, 3821, 3907, 707, 1182, 33, 3759, 3066, 3857, 4006]. Médicale [243, 1040, 2578, 707, 3197]. Médicales [868, 954, 713]. Medicalium [235]. médicaments [3527]. médicaux [1273, 868, 1432]. Medicea [593, 278, 1360, 340, 1538, 2058]. Mediche [1128, 912, 2100]. Medici [2012, 2348, 3494, 2590, 1173, 1752, 1648, 877, 1598, 1861, 4006, 3777, 644, 3027, 869]. Medicina [784, 2219, 1353, 2164, 2235, 2836, 78, 241, 446, 2298, 1724, 865, 360, 359, 437, 640, 637, 706, 867, 947, 948, 1096, 1026, 1431, 1594, 1593, 1675, 1751, 1932, 554, 712, 1024, 1029, 1272, 24, 1599, 2114, 660, 944, 1520, 2327, 612, 2620, 2631, 873, 949, 3115, 45, 43, 912, 1183, 332, 434, 1092, 1434, 1673, 1672, 1677, 1754, 1676, 1865, 1809, 1864, 1866, 2011, 433, 500, 871, 870, 3426, 3178, 1025, 1065, 2278, 957, 1174, 3759, 2642, 786, 1173, 1672, 1809, 1864, 2006, 2154, 2413, 3673, 2524, 3672]. Medicinae [787, 448, 747]. médicinales [436]. medicinali [436, 1268, 1271]. Medicinalia [1775]. Medicinalium [2418]. Medicine [15, 242, 224, 301, 414, 438, 544, 1818, 2312, 3187, 676, 696, 2990, 3072, 3442, 86, 194, 1358, 274, 955, 3651, 1143, 1063, 2755, 1222, 34, 3440, 2907, 2819, 3108, 2999, 3599, 169, 1997, 2950, 3671, 1479, 422, 1944, 786, 1026, 1934, 1935, 1126, 1182, 2944, 3267, 1557, 190, 4078, 1750, 3557, 1065, 3690, 2875, 3231, 1392, 3241, 3398, 3493, 3506, 1144, 3617, 3337, 608, 914, 3770, 45, 43, 3829, 1735, 3879, 606, 3895, 912, 877, 1092, 1865, 2071, 4080, 3962, 4030, 4034, 2906, 601, 2709, 845, 2830, 2959, 3409]. Medicine-by-Post. [2906]. Medicines [3714, 3241, 1271]. Medico [2049, 31, 887, 2642, 80, 1098, 1672, 1866, 2222, 2468, 1186, 1718, 1599, 3972, 2902, 3329, 1223, 546, 2668]. medico-analitico [1098]. medico-biologiques [2902]. medico-chirurgiche [2468]. medico-paziente [80]. Medicorum [1133, 1174, 1391]. medicos [446]. Medieval [3005, 3162, 192, 511, 445, 1102, 1104, 990, 1277, 1188, 1443, 2441, 3594, 2819, 3409, 3707, 2709, 3668, 3338, 2583, 2937, 3250, 3565, 3748, 1596, 406, 833, 1187, 3449, 263, 3898, 4081, 1047]. Medievale [910, 827, 444, 3748, 449, 715, 719, 915, 996, 32, 2820, 2517, 3001, 2419]. Médiévales [32, 512, 1442, 1444, 1984, 3527, 444]. Medievali [2628, 572, 4081, 1642]. Medievalia [1187]. médiévaux [886]. Medii [307, 720, 717, 991, 1279, 1156, 1688, 70, 534]. Medio [27, 84]. Medioeval [2489]. medioevale [870, 1868]. Medioevo [995, 1276, 1816, 2364, 247, 510, 647, 797, 716, 718, 962, 1356, 1940, 2009, 558, 1527, 377, 375, 3259, 433, 1871, 2221, 1223, 3747, 1187, 1684, 2934, 3937, 3961, 2136]. Medir [3833, 1743]. Mediterranean [1429]. Méditerranée [1868, 1224]. méditerranéennes [1224]. Mediums [3407]. Medizin [62, 3116, 2750, 193, 172, 420, 1735, 2196, 706, 950, 4078, 2643, 3130, 690, 565, 676, 948]. Medizin- [62, 3130]. Mediziner [303]. Medizingeschichte [183]. Medizinische [2682, 3857]. Medizintheorie [172]. Medusen [420]. Medvei [1179]. Mees [2938]. meeting [4045, 4078, 1749, 1270]. meetings [3767]. Megas [2571]. Meghan [3425]. Meier [99]. Meijers [2662]. Meine [1714]. Meinel [966, 1735, 2915, 2196, 276, 1150]. Meiner [307, 379, 1989]. Meis [3315]. Meitzner [1531]. Melancholie [248]. Melanchton [1721]. melanconia [2721]. Melania [2631]. Melanie [1769]. Melbourne [194, 418]. Melchior [127]. Melchisedech [820]. Melencolia [3454]. Melhado [1005]. Meli [2825]. Melissa [3735]. Melodi [1419]. Melon [2055, 3896]. Melon-Shaped [2055, 3896]. Meloni [2012, 1166, 1167, 1745]. Melons [3560]. Meltemi [2374]. Memoir [691, 1305, 1731]. Mémoire [2541, 2469, 2067, 4039, 2902, 1305]. Mémoires [1681]. Memoirs [1819, 1911, 829, 2137, 4005, 1464]. Memoria [1682, 1620, 2614, 2754, 505, 1159, 2313]. memorial [4039]. Memoriam [2931]. Memorie [1952, 2023, 569, 709, 200, 844, 1749, 1464, 1949, 2020, 1105, 930, 1251]. Memories [3722, 353]. Memory [3182, 3140, 3732, 2067]. Memosciences [3091]. Men [580, 1707, 2988, 3079, 3623, 2625, 4063, 2660]. Menagerie [3344]. Menchi [577]. Mendel [421]. Mendeleev [1870]. Mendelism [3377]. Mendelsohn [1980]. Mendelssohn [1571]. Mendera [575, 1695]. Mendolicchio [2503]. Menghini [1516]. Mengoli [596]. Menozzi [1856]. Menschen [1004]. Mensing [2195, 4058]. Menso [2055, 23]. Mental [2717, 2912]. Mentalities [166, 974]. Mentality [2912]. Mente [2128, 643, 2431, 3804, 1428, 1248, 1917, 2323, 3546, 680, 1459, 2618, 373, 370, 2269, 2324]. Mentions [3472]. Mentor [2081]. Mentore [2081]. mentori [2849]. Meo [794, 1200, 1679, 1682, 1803]. meraviglia [4044, 1923]. meraviglie [2893]. meravigliosa [1185]. Meravigliose’ [2815]. Merc [2571]. mercati [2888]. Mercatore [1663]. Mercè [353]. Merchant [3736]. Merchant-Poet [3736]. Merchants [3714]. Mercury [3133]. mere [2156, 4022]. Meredith [3361]. Merian [1185]. meridian [3245]. Meridiana [2760]. Meridiane [1415, 1787, 1013]. méridiennes [3245]. Meridione [514]. Merieux [713, 954]. Merit [3738]. merito [2387]. Merkel [491]. Meroi [2949]. Merril [1196]. Mersenne [3203, 2814, 2869]. Merton [913]. Merveilles [1510]. Merz [3308]. Merzagora [2150]. Meschiari [2914]. Mesini [1317]. Meskens [2833]. mesopotamiche [3541]. Messbarger [2783]. Messenger [1118]. Messengers [3755, 3253, 3912, 3835, 1515]. Messina [1436, 2429]. Messzirkel [2544]. mesure [3423]. Metà [2399, 734, 3949]. ’Metafisica [1968, 2272, 1053, 3798, 88, 2432]. Metafisiche [61]. Metafisici [1697, 3798, 2272]. Metafora [459]. Métailié [748]. Metal [2707]. metales [1197]. Metallica [1317, 2707, 2770]. Metallico [740]. Metallurgical [1627]. Metallurgiche [814]. Metallurgy [4015, 3282]. Metamatematica [3911]. metamathematics [3911]. metamorfosi [1185]. metamorphosis [1185]. Metaphor [349, 56, 2764, 3931, 2427]. Metaphors [933]. metaphysical [3798]. Metaphysics [1489, 2583, 1225, 1053, 1425]. Meteore [525, 170]. météorite [2611]. Meteorografo [981]. Meteorologia [93, 1524, 2218, 3136, 2333, 2284, 170, 2271]. Meteorologica [2517]. Meteorological [2971]. météorologie [1985]. Météorologiques [2821]. Meteorology [1985, 3256, 3431, 2513, 3923]. Metheora [535]. Method [1846, 511, 209, 3525, 1199, 3881, 1377]. méthode [548]. Methodique [2843]. methodis [59]. Methods [2917, 1]. Metodi [1777, 1]. Metodo [569, 1693, 181, 3673]. Metodologia [2065, 1619]. Metodologici [909]. métre [702]. Metrics [3306]. Metrologies [2405]. metropolización [1544]. metropolization [1544]. Mettrie [1144, 4030]. Metze [2576]. Metzger [548]. Meucci [1690]. Meulders [2690]. Meyer [99, 343]. Meyerbeer [1571]. Meyers [3482]. Meyerson [1994]. Mezzagora [1925]. Mezzi [1777]. Mezzo [2041]. Mezzogiorno [3089, 3752]. Mgr [1734]. Mgr. [4089]. Mi [877, 2730]. Michael [1006, 2106, 2791, 830, 1238, 2080, 2873, 1732, 225, 2268, 2938, 344, 1854, 3334, 1225, 1914, 615, 1087, 1139, 1490, 1562, 2003, 2143, 2793, 1405, 2542, 1566, 3006, 4073, 4074, 1392, 2673, 1847, 723, 1425, 2911, 2935, 2901, 1920, 2802, 3251]. Micheau [3611]. Michel [2843, 496, 1667, 2140, 1799, 2578, 1605, 2145, 2690, 920, 3257, 3004, 1534, 2422, 2186]. Michel-Pierre [3004]. Michel/Centre [1799]. Michela [745, 995, 2383, 610]. Michelangelo [1020, 1037]. Michele [2167, 2238, 2331, 2896, 612, 2313, 432, 1162, 2009, 2282, 3687, 2376, 817, 2140, 875, 2009, 1399, 578, 2130]. Micheletti [932]. Micheli [2294, 1252, 1389]. Michelle [3615, 1495, 368, 1090, 834]. Michelon [1091]. Michelson [2407, 2406]. Michiel [2291]. Michigan [3003, 2950, 624, 1192, 2489]. Michl [2896]. Micoli [869]. Micoli-M [869]. Micro [974]. Micro-History [974]. Microbi [708]. microfiche [1618, 1618]. Micrologu [3062]. Micrologus. [2628]. Micrométre [527]. Micrometro [1631]. Microscope [1842, 3686, 3208, 839]. Microscopes [2888, 2554]. Microscopi [2888, 854]. microscopia [854]. Microscopical [839, 3208]. Microscopy [2970, 3105, 2683, 3246]. Microstorie [346]. Mid [3236, 1466]. Mid-Eighteenth [3236]. Mid-Victorian [1466]. Middle [335, 2570, 2766, 2818, 2641, 2707, 2364, 30, 1103, 1277, 1446, 1445, 1985, 3006, 2997, 2874, 2932, 2705, 2009, 2515, 3106, 3238, 3056, 3325, 3595, 3747, 3175, 3937]. Midhat [2927]. Midollo [1551]. Midwife [2841]. Mie [930]. Mieli [214, 2929]. Miert [3328]. Miglietta [3367]. Miglioramenti [2726]. Mignini [384]. Migration [3041]. Miguel [2138]. Mikrobiologie [183]. Mikulàs [1433]. Milan [2332, 1355, 3528, 3669, 3278, 3756, 3276, 3839, 4067, 1382]. milanese [3839]. Milani [2332]. Milano [41, 195, 515, 629, 573, 2916, 2881, 897, 2332, 339, 496, 552, 588, 421, 18, 456, 585, 612, 677, 1368, 1811, 3121, 221, 643, 1574, 1693, 2116, 2203, 1112, 181, 2769, 3021, 250, 449, 991, 1088, 2283, 1759, 1820, 2156, 2276, 2605, 3022, 3115, 1248, 1917, 2742, 4067, 1242, 80, 300, 1901, 437, 505, 586, 786, 867, 947, 1092, 1431, 1436, 1678, 1677, 1809, 1864, 1810, 1927, 2430, 1554, 433, 500, 564, 554, 713, 810, 879, 870, 875, 2918, 2743, 3426, 3277, 3257, 88, 71, 20, 185, 291, 431, 518, 452, 457, 428, 429, 631]. Milano [559, 953, 1011, 1016, 1159, 1172, 1355, 1350, 1373, 1419, 1510, 1749, 1871, 2005, 2212, 2217, 922, 929, 2855, 1712, 1321, 2716, 66, 812, 1788, 2722, 1394, 1506, 1856, 2104, 2132, 2507, 957, 942, 1113, 1455, 1491, 1650, 1796, 1988, 2108, 371, 373, 372, 370, 2057, 2902, 1291, 2324, 2433, 2503, 1252, 1331, 1530, 2714, 2622, 1900, 2644, 2367, 627, 2133, 178, 546, 3077, 1382, 2743, 221, 954]. Milano-Trento [1796]. Milchstraße [1862, 4036]. Mile [1994]. milieu [2877, 338]. militare [3015, 880, 561]. militares [494, 618]. Military [2664, 1484]. Milky [1412, 4036]. Milla [1532, 1195]. Millar [2362]. Millàs [333]. Mille [2202]. Millén [1516]. millénaire [3611]. Millenni [1642]. millennium [4059]. Miller [1542]. Millot [3611]. Mills [2579]. Millwood [68]. Milner [3084, 684]. Milo [2781]. Mimesis [2507]. Mimi [3333]. Minazzi [3121]. Mind [3379, 192, 3231, 3246, 694, 643, 3804, 2069, 3943, 342, 2377, 3668]. Mind-forg’d [342]. mindful [3018]. Minds [511, 694]. Mineccia [2266]. Minelli [2612, 1487]. Mineral [734]. Minerale [734]. Mineralogy [538, 2161, 4066]. Mineraria [818, 3021]. Minerva [432, 494, 618]. Mines [1727, 754, 2782, 1135]. mineur [754]. Ming [500, 1873, 3566]. Miniati [777, 2436]. Miniato [2886, 2345]. Minici [1794]. minières [1060]. Mining [3347, 1060]. Ministero [809, 844, 1525, 1591]. Mino [165, 814]. Minor [1873]. minori [2531]. Minsky [1248]. mint [1872, 3745]. Minturno [2154]. minute [1598]. Minuz [178]. Miotto [631]. Mir [591]. Mirabili [1794]. Mirabilia [2530]. Miracles [524]. miracoli [1927, 2693]. miracolo [3397]. Mirandola [2942, 1360, 1565]. Mirella [1850]. mires [3245]. Mirhady [1225]. Mirko [1223, 331, 566]. miroir [3024]. Mirror [1439, 2084, 2146, 1020]. Mirrors [3436, 3182]. Mirto [2058]. Miscellanea [3319, 42, 2100, 1225]. Miscrope [1057]. mise [2067, 4039]. Missa [1460].

Missed [2833]. Missiles [1484]. Mission [3072]. Missouri [1108]. misteri [942]. Misterioso [208]. mistero [3416]. Misticismo [2083, 1906, 2364, 3918]. Misunderstanding [1694]. Misura [428, 2525, 2153]. Misurare [1059, 1955, 2026, 2135]. misure [3813, 1806]. Mitchell [2709, 1575]. Miti [2670, 2627]. mito [2467, 503, 957]. mitraille [3483]. Mitri [2780, 2580]. Mitscherlich [961]. Mittelalter [193, 1684]. Mittelalters [29, 535, 1446]. Mittenti [6, 905]. Mittler [3073]. Mitwirkung [1854]. mixte [2578]. mobile [656, 3695]. Mobilising [3708]. Möbius [3123]. Moby [951]. Mocchi [680]. Modal [2583]. Modalità [1380]. Modalities [1104]. Model [1702, 3263, 1468, 3901, 1296, 1608, 1781]. Modèles [2550]. Modelli [1917, 1780, 2191, 2394, 1011, 1499]. Modello [1781]. Models [2478, 3502, 1499, 3518, 3703, 3579, 2699, 3513, 2556]. Modem [2610, 2521, 2666, 2575]. Modena [726, 1322, 1569, 1765, 2461, 2502, 1190, 1595, 435, 171, 560, 1011, 1163, 1906, 2129]. Modern [167, 491, 837, 1108, 1169, 2586, 3187, 3240, 3262, 2717, 3112, 3543, 3328, 2606, 3175, 2882, 1842, 2891, 3007, 1835, 3322, 3653, 2649, 649, 3558, 3651, 2917, 3522, 3501, 2371, 2540, 1560, 3441, 2589, 2579, 1481, 2514, 3066, 2978, 3205, 2657, 3835, 2827, 2823, 2600, 3070, 3108, 3857, 3399, 3644, 2788, 3253, 3342, 2832, 2656, 3671, 2768, 3455, 2945, 2847, 3597, 3912, 3191, 2660, 3492, 3038, 2733, 3252, 1319, 3712, 3028, 3114, 3202, 2583, 2926, 3231, 1397, 3148, 3183, 3146, 2711, 2585, 3737, 3612, 3444, 3398, 3493, 3119, 3711, 3519, 2587, 2582, 3445, 3233, 3361, 2456]. Modern [2816, 3587, 3207, 3274, 3419, 1290, 1814, 3363, 3242, 759, 1261, 1576, 3791, 3792, 3793, 3794, 1879, 3839, 3855, 2747, 3870, 3895, 3930, 692, 1348, 1654, 3999, 4018, 2560, 2907, 3281, 3216, 3213]. Moderna [72, 718, 1560, 1876, 604, 89, 2865, 2670, 2383, 2949, 2941, 562, 1788, 3961, 3019, 1348, 555, 1233, 1318, 2797, 3136, 2836, 2721, 1866, 3426, 2062, 2597, 3206, 3259, 3672, 3655]. Moderne [1918, 1879, 1667, 3895, 2375, 3069, 3930, 3010]. modernen [415, 3999]. modernes [3568, 3527]. moderni [2693]. Modernist [1409]. Modernit [2192, 2117]. Modernity [3187, 3369, 2947, 3439, 3985, 2861, 3333]. modernizzazione [1759]. Moderno [2871, 1239, 3839]. modernorum [444]. Modesty [2833]. Modica [13]. mœurs [3861, 2892]. Moggi [3693, 1252, 1519]. moitié [675, 1567]. moléculaire [3127]. Molella [3369]. Moleschott [2668]. Moletti [2139]. Moliner [2101]. Molland [1443]. Molti [658, 2738]. Moment [2600]. Momenti [639, 564]. mon [2013, 3932]. Monaco [720, 1671]. Monad [3975, 3113, 1847]. Monarchy [3294]. Moncada [494, 618]. Mond [2943]. Mondadori [1112, 1092, 1394, 1113, 370, 1900]. mondana [2934]. Monde [1476, 1700, 3004, 1804, 260, 1187, 3461, 3087, 3024, 2537]. Mondella [1712, 2931]. Mondi [2132, 3614]. Mondiale [1099, 880, 712, 1540]. Mondo [1493, 3541, 447, 2651, 706, 1101, 2430, 3918, 1159, 1373, 1766, 2002, 3314, 771, 2432, 1426, 4010, 1906, 2572, 2968, 3457]. mondoantico [2642]. Money [2887]. Monferrato [1287]. Mongolfiere [2536]. Monica [3257, 1519, 3108, 2722, 3278]. Monika [1735]. Monkey [2966]. Monnaie [2229]. Monnier [2642]. Monografas [43]. Monograph [1161]. Monographie [420]. Monographs [1406, 2686, 2609]. Monro [3628]. Monsignor [4089]. Monster [2424, 304, 3834]. Monstrosities [2589]. Monstrous [2589]. Monsummano [1766]. Montacchini [83]. montagne [4042, 1327]. Montaldo [3085, 3547]. Montanari [3655, 1705, 2293]. Montandon [3377]. Montbard [4047]. Monte [2368, 2346, 815]. Montebard [1326]. Montebelli [581, 858]. Montecatini [197]. Montefeltro [2522]. Montero [678]. Montessori [2104]. Montevarchi [1874]. Montevarchialla [1874]. Montgomery [2707]. Monti [221, 2965, 1712, 1788, 2903, 2902, 2205, 1039]. Montmartre [1727]. Montpellier [2878, 243]. Montral [512]. Montral-Paris [512]. Montréal [512, 2375]. Montres [1510]. monument [3540]. Monumenta [1448]. monumento [3009]. Monza [1756]. Moon [3600, 1654, 3790]. Moore [848, 2427]. Moquin [2013]. Moquin-Tandon [2013]. Morabito [2336, 1917, 2287, 2618]. Moral [3191, 2559]. morale [872]. Morales [3412]. Morales-Muñiz [3412]. moralisation [3197]. Moralische [93]. moraliska [2500]. Morals [3504]. Moran [1063, 2591]. Morano [926, 1233]. morbo [2714]. Mordechai [2600]. Mordente [2163, 2234, 2346]. Morel [2933, 3711]. Morello [2333, 2284]. Morelon [216]. Moreniana [2614]. Moretti [2592, 1330]. Morfologia [1201]. Morgagni [299, 337]. Morgan [3676]. Mori [971]. Morini [1242, 1642]. Moritz [101]. Morize [430]. Morlacchi [2883]. Morley [3542]. Mormino [1323]. Morosi [1777]. Morpurgo [678]. Morrell [2079]. Morris [3369, 3716, 1349]. Morrison [761, 916, 2975, 2457]. Morrison-Low [761, 916, 2975]. mort [4047, 1440, 1326]. Morte [3343]. Morus [2542, 2747, 2676]. Morveau [1495]. mosca [1922]. Moscati [211]. Moscow [591]. Mosini [1603]. Mosso [2331]. most [2844, 4028]. Mostert [1589]. Mostra [911, 4067, 1741, 867, 1176, 83, 809, 844, 1411, 1331, 1530]. mot [2770]. Mother [2091]. Mothu [2309]. Motion [842, 1880, 2031, 3778, 3525]. Motor [2870]. Mott [3429]. Motta [2141]. Mottoni [720]. Motu [1822, 1782, 1960, 2031, 2140, 1399, 757]. motus [1323]. Mount [916]. mountain [4042]. mouse [2232, 3973]. mouvement [3778, 920, 3540, 2422]. Movement [2870, 3540]. Movies [3586]. movimento [2622]. Moving [3304, 3654]. Moyen [3631, 305, 1187, 1985, 33, 1273]. Moyer [2489]. Mozart [877]. Mr [2765, 3079]. Ms [1399, 1984, 955]. mss [991, 1156]. Mucchi [1322, 1569, 1765, 171]. Mucchielli [1878]. Mucciarelli [2221]. Much [3537]. muchabala [609]. Mud [3460]. Mu’Dh’s [1047]. Mudry [868]. Mugnai [887, 255]. Muir [289]. mukhtasar [609]. Mulino [89, 202, 1616, 934, 3094, 1066, 1815, 2155, 2471, 2543, 2836, 2721, 180, 85, 170, 257, 2315, 2789, 1428, 1909, 2941, 3029, 1148, 331, 656, 550, 755, 684, 1459, 3226, 1639, 3206, 811, 913]. Mulino-Duncker [3226]. Müllberger [3427]. Muller [567, 1715, 3462, 787, 2866, 1916]. Muller-Jahncke [567, 1715]. Müller-Wille [2866]. Multa [3159]. Multhauf [293]. Multimedia [2144]. Multinational [3771, 3332]. Multiply [3605].

Mummies [3216]. München [2228, 676, 1387, 2682, 2908, 30, 193, 248, 1862, 2161, 2598, 706, 2643, 2857, 961, 543, 58, 890]. München-Wien [30]. München-Zürich [193]. Münchener [248]. Mundana [996]. Mundi [459, 574, 2934]. Mundo [2895]. Mundus [1610, 3293]. Münehen [1448]. Munich [2161, 4066]. Muñiz [3412]. Munksgaard [1081]. Muñoz [2593]. Münster [3253, 3017]. Muntschick [256]. muqbala [609]. mural [1172]. murali [1172]. muranesi [2592, 3934]. Murano [3934]. Muratori [1948, 2019]. Murdoch [910]. Murphy [3506]. Murray [767]. Murri [2543]. Mursia [586]. Musa [2436]. Musaei [2436]. Muscle [2870]. Muse [1488, 1065]. Museale [1373, 1411]. Museali [2191]. Musealization [3452]. Musée [3365, 2757, 3257, 502, 702, 2856, 3257, 1421, 3365]. musée-laboratoire [3365]. musées [954, 713, 1421]. Musei [2007, 21, 3818, 20, 19, 2213, 4044, 3807, 1016, 1923, 2150, 2322]. Museo [777, 521, 1035, 1205, 1374, 1463, 2169, 2240, 2386, 594, 493, 2452, 1711, 486, 2248, 2697, 38, 2126, 599, 3719, 1561, 1979, 2261, 1438, 391, 47, 90, 261, 458, 431, 518, 428, 657, 633, 731, 813, 780, 973, 963, 937, 864, 1116, 1082, 1011, 1015, 1292, 1163, 1159, 1343, 1419, 1423, 1510, 1665, 1663, 1746, 1806, 2009, 2110, 1270, 1530, 593, 2619, 2270, 854, 1373, 4019, 2322, 3099, 3156, 2293, 2257, 3685, 2088]. Museo/The [1665]. Museologia [1253, 1508, 2109, 2209, 1582, 1137, 1159, 2150]. MuseoPoli [1411, 1923]. Muses [1065]. Museu [430]. Museum [454, 414, 961, 2418, 3156, 1711, 2700, 847, 856, 2248, 2245, 2510, 1566, 1862, 2698, 2658, 2662, 3111, 3345, 3289, 3412, 1299, 2568, 2555, 2409, 2701, 2554, 2683, 293, 409, 762, 761, 938, 1082, 1072, 1256, 1422, 1511, 1510, 1665, 1662, 1792, 1872, 2069, 2195, 2204, 2702, 3709, 3505, 3092, 2956, 2147, 3452, 2556, 4058, 1418, 2118, 2719, 3913, 1411, 1342, 266, 1541, 2088]. Museums [3150, 3695, 761, 1319, 2817, 3092, 4044, 3807, 3818, 3818, 2940, 2456]. Music [1106, 1346, 2586, 3437, 3486, 3399, 3869, 356, 355, 1225, 2123, 2056, 2776, 3905]. Musica [1616, 1887, 2123, 2934, 50, 596, 1491, 1615, 85, 3203, 355]. Musical [830]. musicale [355]. musicali [355]. Musicians [2913]. Musick [2586]. Musikbibliothek [3905]. Musique [3437]. Musiques [502]. Musis [1603, 1343]. Musitelli [1431]. Musschenbroek [1792]. Mussini [2433]. Must [3661]. Musurgia [3174]. mutamento [2796]. Mutio [3201]. Mutual [2496]. Mutualismo [810]. muziek [2986]. my [3152, 3932]. Myles [2913]. Myology [2870]. Myra [3213]. Mystery [2561]. Mysticism [2828, 2083, 2364, 3918]. Mystik [1086]. mystique [287]. Myth [1699, 349]. Mythologie [1202, 4033]. mythology [4033].

N [283, 439, 2805, 1105, 1277, 64, 1505, 2095, 422, 624, 761, 953, 1749, 163, 3092, 2562, 61, 714, 777, 1536, 2016, 2467, 2496, 1902, 2313, 572, 2371, 1516, 2010, 367, 386, 787, 763, 200, 727, 1025, 1165, 1372, 1350, 1619, 2009, 1600, 308, 2231, 2679, 2201, 1137, 914, 346, 407, 63]. N. [1643]. N.F [248]. N.H [2867]. N.J [249, 34]. n.n [1259, 1369, 1423, 558]. N.S [2145, 2267, 1033]. n.t [2163, 2234, 2387, 2389]. N.Y [10]. n/bl [914]. Nacci [2383]. nach [4078, 2734]. Nachleben [1396]. Nacional [963, 3618]. Nadia [3080, 2271]. Naess [2653]. Nagaoka [2305]. Nagel [1141]. Naissance [2652, 920, 681, 866, 4089]. naixement [3019]. Nakladem [1584]. Nallet [502]. Names [213, 3464, 34, 43, 21, 35, 33]. Naming [833]. Nancy [3442, 2656, 2950, 381]. Nani [2331, 2282]. Nanni [1813]. Nannoni [1523, 1728]. Nantes [868]. Napier [939]. Naples [893, 1040, 4014, 3729]. Napoleon [3216]. Napoleone [3412]. Napoleonic [2538, 3487, 3216]. Napoleonica [2536]. napoléonienne [314]. napoletane [1084]. napoletani [1230]. Napoletano [1629, 3397]. Napoli [896, 1201, 1680, 2370, 4043, 75, 86, 444, 446, 274, 1059, 3672, 2665, 2157, 2497, 284, 926, 1908, 1995, 3057, 386, 2712, 79, 504, 880, 1522, 2659, 238, 237, 832, 854, 1082, 1084, 1873, 1919, 1926, 2226, 2307, 1052, 1414, 1851, 2429, 44, 2316, 230, 1767, 384, 2914, 377, 1233, 1142, 1371, 1990, 3367, 1706, 4014]. Napolip [2154]. narcotics [1433]. Nardini [2592]. Narr [1684]. Narrative [3318, 418]. Narratives [3152, 305, 3649]. Nascita [2043, 1560, 811, 2421, 311, 572, 3402, 3976, 3547, 2581]. Nasim [3346, 3740]. Nassau [2715]. Nastasi [514, 3031, 412]. Natacha [3203]. Natalie [2228, 2544]. Natalya [2662]. Natasa [3742]. Nathalie [3075]. Nathan [2053, 1410]. Nation [2920, 3899, 2102]. National [1418, 699, 3414, 243, 836, 2510, 1327, 3205, 3717, 3289, 2928, 1299, 3624, 293, 540, 761, 702, 1355, 3092, 1783, 768, 4042, 2118, 3986, 3987, 3988, 1783]. nationale [4043]. Nationalism [2977, 1912, 3974, 3893, 1245]. Nationalismus [3893, 1912]. Nationalmusei [3092]. nativi [1729]. NATO [3512]. Natur [2200, 1004, 498, 547, 343, 93]. Natur-und [498]. Natura [265, 1066, 1871, 658, 1119, 2628, 896, 1201, 1903, 4044, 1322, 1236, 3279, 715, 3826, 2754, 1923, 2151, 2226, 2225, 3388, 678, 3062, 1086, 1034, 1455, 1988, 3989, 2501, 1331, 2882, 2558, 1998, 2384, 2748, 3443, 3652]. naturae [1086, 819]. Natural [454, 918, 1346, 2056, 2424, 3735, 2606, 1643, 3059, 3230, 2964, 2093, 3510, 3476, 511, 2371, 1263, 3441, 2698, 2873, 3043, 2706, 3100, 2719, 2827, 2819, 3345, 3850, 174, 3671, 3412, 2137, 65, 763, 2775, 2701, 833, 2807, 2803, 2585, 3393, 2565, 1213, 4026, 2031, 1403, 1721, 3818, 3834, 3855, 3869, 3893, 3897, 4082, 1109, 1912, 2059, 3614]. Naturale [898, 597, 2697, 1207, 50, 431, 518, 2996, 537, 2386, 3001, 3397, 2949, 3818, 872, 3396, 2322, 978]. naturaleza [2494]. Naturali [903, 1516, 554, 1871, 1960, 2031, 2002, 2148]. Naturalis [3283, 2394, 4026]. naturalismo [2665]. Naturalist [1408, 3872, 3932, 2884]. naturalista [1236, 3872, 2373, 2602]. naturaliste [416, 2013, 3932]. Naturalistes [1474]. naturalisti [588, 2888]. Naturalistic [2698]. naturalistica [1871]. Naturalistico [1380, 934]. Naturalium [2418, 235]. Nature [227, 592, 679, 1535, 2102, 3175, 620, 2908, 2551, 2671, 3495, 2824, 3723, 56, 3100, 2814, 732, 2862, 3033, 2704, 683, 3014, 3351, 3542, 3196, 3529, 3492, 1319, 1397, 3387, 2784, 3528, 819, 1541, 3032, 3401, 1119, 2565, 3748, 4044, 2735, 3826, 3899, 2419, 806, 1086, 3989, 3075, 1331, 3026, 4024, 2920, 2794, 267, 2559, 2737, 3639]. Nature/Culture [2671]. Naturelle [1418, 598, 1473, 2829, 872, 2118, 2729, 1304, 1472, 2898]. naturforschenden [1141]. Naturforscher [725, 557, 1992, 1797, 3977, 1079]. Naturforschung [3186]. Naturis [2895]. Naturkatastrophen [4082, 1838]. Naturkontraktet [1023]. Naturkunde [567, 787]. Naturphilosophie [343, 1086, 2734]. naturphilosophischen [1328]. naturumgängets [1023]. Naturwissenschaft [307, 420, 1735, 2196, 950, 4078, 382, 565, 695]. Naturwissenschaften [296, 2228, 2161, 547, 1735, 3893, 1724, 2196, 1079, 1912, 2466, 1450, 1069, 1381, 1916, 2055, 344]. Naturwissenschaftler [303, 1916]. naturwissenschaftlichen [2466]. Naturwissenschaftlicher [411, 3893, 1001]. Naturwissenschaftliches [93]. Naturwissenschaftsgeschichte [1602]. Natuur [3109]. Natuurwetenschappen [63, 1791]. Naudé [1417]. Nauenberg [2268]. Navarro [1227, 2101, 1849, 2114, 2593]. navegación [2808]. Navegar [2357, 2356]. Navicula [1421]. Navigating [2357, 2356]. Navigation [489, 3418]. navigazione [2574]. Navy [3451]. Nawa [2915]. Nazionale [777, 1091, 2370, 1569, 1689, 221, 311, 309, 720, 717, 991, 1279, 1156, 1688, 70, 519, 1603, 752, 1741, 1307, 729, 1677, 2713, 428, 779, 1166, 1167, 1195, 1429, 1355, 1813, 2217, 2363, 999, 799, 800, 44, 2274, 412, 2914, 1783, 2726, 1765, 809, 171, 2005, 2608, 3986, 3987, 3988, 831, 2692]. Nazionale-Ufficio [2363]. Nazione [2969]. Nazioni [1737]. Nazis [422]. Nazism [768]. Neapolitan [519, 1084, 1230]. Near [2879]. Near-Eastern [2879]. Nebel [1862, 4036]. Nebula [3051, 4036]. nebulae [3346]. Nebulous [3789, 3792, 3795, 1654]. Necessary [3350]. Necessità [2258]. Nederlandse [802, 63, 1791, 3018, 2884]. Need [2258]. Needham [3225, 748, 1484, 1641, 1982, 2629, 406, 186]. Neeley [2377]. Neeraja [3601]. negli [3813, 3839, 429, 1619, 1806, 2501]. Negociating [1494]. Nehru [329]. Nei [1552, 393, 4040, 1564, 3172, 2524, 446, 2888, 407, 952, 869, 1026, 1434, 1678, 1755, 1011, 558, 44]. Neighbours [3091]. Neil [2609, 2587, 2956]. Neither [3243]. Nel [2178, 2002, 299, 515, 517, 1200, 3747, 315, 2434, 3752, 496, 1784, 688, 1066, 1488, 2491, 3541, 2780, 278, 3788, 180, 3801, 250, 448, 716, 170, 1824, 2665, 3022, 1789, 1111, 2536, 2893, 2642, 79, 1992, 706, 789, 1101, 1598, 2154, 3496, 2659, 979, 433, 876, 1029, 3918, 2257, 3426, 1976, 1020, 1255, 1159, 1373, 1512, 2009, 337, 922, 2677, 2206, 3314, 2124, 658, 733, 1061, 1394, 1653, 2618, 1851, 3965, 1140, 1426, 3329, 1868, 2954, 1644, 1783, 1906, 2125, 3089, 3367, 3388, 2339]. nel‘De [3057]. Nella [1221, 2473, 3103, 1220, 50, 1311, 660, 1120, 1053, 1218, 1785, 195, 313, 1539, 1605, 1691, 2421, 340, 569, 1368, 1648, 2517, 3040, 1236, 1996, 2116, 3545, 2721, 2797, 442, 996, 1462, 1789, 1538, 3209, 3818, 3820, 682, 2140, 436, 1096, 1752, 3546, 708, 871, 879, 2918, 3925, 2129, 238, 631, 575, 1240, 1591, 1873, 2001, 1321, 982, 678, 1265, 2132, 2429, 2432, 3976, 687, 1713, 370, 3655, 734, 1252, 1324, 745, 2322, 1000, 178, 2135]. nell’Accademia [2338, 1701]. Nell’Alta [1300]. nell’antica [332, 867]. nell’antichità [1753]. nell’antico [869]. Nell’Archivio [2260, 237]. nell’arte [436, 1927]. Nelle [2351, 1829, 1128, 197, 991, 1156, 70, 1243, 2965, 958, 1093, 556, 844, 1663, 2002, 2363, 999, 2141, 178, 1482]. nell’et [2387, 612, 1491]. nell’età [1616, 3426, 1788, 765, 1615]. Nell’Europa [2473, 1427, 2949, 561]. nell’evoluzione [365]. Nelli [1056, 1435]. Nelli-Elena [1056, 1435]. Nell’Italia [1778, 3689, 2836, 2670, 2742, 2672, 2893, 2741, 3426, 3206]. Nello [1550, 909, 2942]. nell’Occidente [377]. Nell’Opera [1637, 2567, 1785, 339, 1917, 3933, 2645]. nell’Osservatorio [700]. nell’ottavo [572]. nell’Ottocento [588, 711, 2271, 710]. nell’Università [1183]. Nelson [1934]. Neohellenic [3205, 1825]. Neolitico [794]. Neos [2490]. neotranscendental [1499]. neotrascendentale [1499]. Nepi [2230]. Neptune [2553]. Neptunismus [1069]. Nequam [534]. Neri [2329, 1056, 1435, 969]. nero [1915, 1566]. Nerve [2678]. Nerves [25, 3054, 2678]. nervoso [1951, 2022, 2334, 2285]. Nesbitt [3695]. Nest [174, 2549, 2625]. Neswald [2854]. Netherlands [2662, 3829, 659, 2928, 3918, 908]. Network [2312, 3341, 1936, 1935, 4034]. Networks [3303, 3490, 3036]. Netz [2881, 2785]. Neu [1898, 10]. Neudruck [1731]. Neue [961]. neuen [253]. neueren [571, 798, 994]. Neuman [2707]. Neumann [3481, 3911]. Neuro [3384]. Neuro-Realism [3384]. Neurobiologia [2324]. Neurological [1355]. Neurologico [1355]. neurology [25]. neuron [1100]. Neurophilosophy [3093]. Neurophysiology [26, 545]. neuropsicologia [2618]. Neuroscience [2764, 3378]. Neurosciences [3385, 1460]. neuroscienze [2690]. neurotransmitters [2678]. Neuve [1442, 1859, 3091, 1734, 4089, 4089]. Neuzeit [567, 895, 966, 1396, 565, 2682, 3857]. Neve [2321]. Never [3191, 3101]. nevrosi [67]. Newbury [2916]. Newly [2175]. Newman [2834, 3415, 2718, 3529, 997, 3638]. Newnham [2050]. news [447]. Newsletter [3097, 2989, 3132, 2811, 2868, 2923, 3045, 2563, 2634, 2694, 2758]. Newspapers [3055]. Newton [617, 1002, 1199, 2586, 3476, 447, 1263, 2311, 1588, 219, 2600, 2781, 3118, 803, 1198, 447, 1366, 3843, 2654, 2651, 3067, 3490, 661, 3917, 2790, 1053, 1534, 3529]. Newtonian [2600, 2268, 1109, 1051]. Newtonians [892, 3888]. Next [1846, 3881]. Nexus [2972]. Niarcel [2229]. Niccolò [3476, 1829, 3630]. Nice [1224]. Nice-Sophia [1224]. Nicholas [3510, 3363, 1518, 3013, 539, 3213]. Nicholson [2076]. Nick [3400]. Nico [2692]. Nicola [2202]. Nicolaidis [3205, 3267]. Nicolas [2145, 3332, 3784, 198, 3490, 921]. Nicolas-Claude [921]. Nicolaus [129, 1941, 1450, 102, 2870, 2992]. Nicole [653, 2973]. Nicoletta [2333, 2284, 2523, 2149, 2060, 1134]. Nicolini [2149]. Nicolò [4015, 3282, 887]. Nicolodi [2796]. Nicolson [2881]. Nicomp [2797]. Niedergang [571, 798, 994]. niedergeschrieben [1731]. Niels [3137, 3263, 3842, 3901, 1008, 3950, 1007]. Nielsen [235, 2583]. niemandem [1726]. Niemeyer [895, 2831]. Nieto [2059, 3675, 2907, 3511]. Nieto-Galan [2059, 3675, 2907, 3511]. Nietzsche [2427, 3931, 2432]. Nieuwe [63, 1791]. Nievole [197]. Nigel [3131, 840, 3565]. Nikolaus [2308]. Nile [840]. Nina [2972]. Nine [2379, 2095]. Nineteenth [77, 1457, 1589, 2737, 3637, 1670, 40, 2988, 2781, 3656, 3033, 418, 2961, 778, 2911, 4004, 1331, 2913, 3346, 3965, 2350, 3471, 63, 3217, 3198, 3554, 2443, 734, 2127]. Nineteenth-Century [1457, 1589, 2737, 3637, 2781, 3656, 3033, 418, 2911, 2913, 63, 3217, 3554, 2127]. Nino [2308]. Nis [3489]. nit [1099]. Nixon [3656]. NJ [1725, 3000]. Nlg [1857]. nn [774, 1803]. No [3467, 1041, 613, 2882, 40, 33, 38, 42, 21, 2470, 2915, 2782]. No-Lamarckien [1041]. Noakes [2737]. Nobel [672, 3036, 3092, 2456, 2545, 1290, 1245, 3463, 3470]. Nobelpreisträger [1336]. Nobiltà [926]. nobody [1726]. Nocciolo [1672]. Nock [2644]. nodale [2597]. Noel [2881, 2717, 3034, 3205, 2493, 1460]. Noëlle [3461]. Nollet [2512, 2423, 3966]. nombres [886, 2596]. Nombreux [516]. Nomenclatura [3172]. Nomenklatur [183]. Nomi [1625, 1773, 1883, 2243, 2342, 2439, 28, 29, 30, 31, 58, 991, 1102, 1156, 70, 407]. noms [1444]. Non [1055, 2106, 824, 2424, 3768, 2215, 3804, 3830, 2323, 1992, 1901, 2004, 3496, 2415]. non-Euclidean [2215, 3768]. non-human [3804]. Non-Locality [2106, 3830]. Non-Printed [1055]. Non-Verbal [824]. Nonbinary [3633]. nónio [1744]. nono [448]. nor [3243]. Norci [893]. Nordman [2911]. Noreen [122]. Normal [3470]. Normale [3080, 941, 2677, 1645, 3203, 2898]. normaliens [3483]. Norman [2753]. North [2312, 2586, 2964, 511, 408, 1935, 2683, 3564, 4034]. Northern [182, 2660]. Norton [1853, 3281, 1497, 2678, 689]. Nos [1857, 638]. Nose [3371]. nosocomio [2072]. nostri [3747, 792, 785, 500, 875, 1748, 2009]. Nota [1470, 1704, 764]. Note [459, 1549, 1189, 2374, 1051, 3142, 620, 1417, 2497, 3905, 1704, 1379]. notebooks [3798, 2834, 435]. Notenhälsen [3905]. Notes [1476, 2759, 3098, 3155, 3192, 1911, 28, 58, 1775, 1880, 2013, 1475, 620, 2898, 2528, 422, 681, 763, 233, 893, 3932, 1647, 1476, 3437]. ’Nothing [819]. Notions [388]. Notizie [336, 522, 1424, 1777]. Notomi [2349]. Notre [445, 2774, 613, 2585]. Notturno [902]. nourriture [872]. Nouvelle [3075, 3127]. Nouvelles [386, 2869]. Nova [3282, 1312, 2454, 3670, 1979, 2190, 1363, 3158, 4015]. Novalis [1330]. Novam [51]. Novant’anni [1572]. Novara [1156, 2156, 4050, 3315]. Novato [2753]. Nove [2002]. Novecento [3040, 3121, 1996, 2116, 915, 1759, 2504, 2742, 662, 1075, 1851, 3181, 1519, 456, 3744, 3820, 3426, 3949, 2107, 2429, 2507, 3544]. novel [654]. Novella [277, 1354]. November [648, 173, 4089, 1331]. Novembre [561, 1360, 1734, 4089, 2581, 1331]. noviembre [1860]. novità [447]. Novohispanic [3218]. Novum [51]. Novus [3293]. Nr [302, 542, 233]. NSW [2625]. Nuclear [1758, 4046]. Nucleare [1195, 2918]. Nucleus [1702]. Nude [3382]. Nudity. [2746]. nueva [2808]. Nuland [1092]. Numa [1190]. Number [15, 184, 293, 916, 2433]. Numbers [438, 842, 1655, 3712, 2922]. numeri [330, 2580]. Numerical [2083, 233]. Numerico [2083]. numero [2433]. Nummedal [3068]. Nuncius [1964, 2035, 2040, 2434, 1481, 3225, 1064, 814, 523, 522, 1118]. Nunes [1744, 2356]. Nuova [1239, 1533, 250, 3177, 2883, 2172, 2637, 2871, 1036, 2041, 288, 709, 697, 1976, 3135, 727, 1861, 1931, 2149, 2414, 1065, 1114, 1323, 1289, 1653, 2107, 2264, 1330, 745, 2726, 1482, 569, 435, 1265, 2104, 373, 3773, 3774, 3775, 3776, 1778, 2156]. Nuove [1132, 1801, 910, 1, 1373]. Nuovedizioni [31]. Nuovi [1778, 1, 570, 3614]. Nuovo [1572, 902, 2062, 377, 371, 3138]. Nuremberg [409, 3506]. Nussbaumer [3086]. Nutrition [1274]. Nuvolari [2607]. Nuvole [1050]. Nuzzo [1922]. Ny [68]. Nye [1500]. nylon [1814]. Nyomda [2421]. Nyomda-Roma [2421].

O [3254, 1448, 3225, 683, 706, 1178, 771, 2583, 2432, 2133, 3479, 288, 2391, 1744, 2068, 3990, 3500]. O-7190-2567-2 [683]. O. [1678]. O.P [252]. Oberbibliothekar [173]. Oberman [1229]. Oberta [2593]. Obituary [272, 2703]. Object [2684, 3295, 3635]. Objectifs [210]. Objectivity [3859, 1655, 3753, 3754, 1575, 2983]. Objects [3679, 3531, 3070, 3253, 2825, 3912, 3725, 3621, 3532, 3634, 3835]. Obrist [749, 2820]. obscuris [3146]. Observateur [1700]. Observational [1468, 2812, 1308]. Observatione [3652]. Observations [3328, 817, 2578, 4018]. Observatoire [210, 2048, 2566, 3245, 2638]. Observatoires [3307, 2396]. Observatories [2089, 1308, 2984, 2396, 1749, 2096]. Observatorio [430]. Observatory [2302, 2379, 613, 3847, 631, 1082, 1920, 1750, 3154, 462, 3245]. Observed [2839]. Observing [3346, 3363]. Obstetricia [1094]. occasion [1090, 1440, 4090]. Occasional [1872]. occasione [809]. occasioni [811, 913]. Occhio [570, 1052]. Occident [748]. occidental [1273]. occidentale [305, 331, 1866, 1718, 2644, 2631, 1223]. Occidente [165]. Occult [491, 166, 2561, 3769]. occulta [1229]. occulte [3769, 752, 2273]. Occultism [2560]. occulto [1965, 2036]. Occupation [3725]. Oceania. [864]. Océanique [2762]. Ochino [2488]. Ockham [445, 29]. O’Connor [2976]. Octavo [1283]. October [63, 4052, 311, 824, 2080, 2311, 3917, 2078, 2195, 1137, 4058, 2701, 3932]. octobre [2013, 3932]. oculistica [3856, 2149]. Oddi [3201]. Oddone [2588]. Odoardo [2448, 1974]. O’Donnell [286, 176]. odontoiatria [1180]. oecumenique [748]. Oeder [151]. Oeuvres [1795, 3785, 4065, 1232, 368, 216, 3368]. Offenbarung [1202, 4033]. Öffentliche [2532]. offered [4063]. offerta [2888]. offertes [3280, 4063]. Office [182]. Official [3294]. Officina [3257, 1354, 2128, 3528]. officine [1030]. Oficyna [1256]. oggetti [1585]. oggetto [1580]. oggi [1769, 559, 1452, 2364, 1865, 2153, 2810, 3741]. Ogilvie [2775]. Ohya [500]. okkulten [1191]. Ökonomien [3530]. Olaf [2734, 1081, 2908, 380]. Olandese [14]. Olaus [1365]. Old [1805, 2538, 1778, 614, 160, 980, 1084, 1584, 2817, 3459, 3060, 3919, 1665, 2069, 2068, 3990, 4019, 816]. Oldenbourg [961]. Oldroyd [677, 1717]. Olduvai [1154, 4008, 4009]. Ole [2967, 2962]. Olff [1410]. Olff-Nathan [1410]. Olga [306, 1444]. Oliver [3675, 3225, 2532]. Oliverio [1951, 2022]. Olivi [1236, 1496, 2602]. Olivier [866, 1981, 3087].

Olivieri [1532]. ollada [2068, 3990]. Olmi [2941, 2754, 2691, 2954, 1269]. Olmo [1813, 2005, 2217]. Olms [1489, 338, 746, 747]. Olms-Weidmann [746, 747]. Olof [2675]. Ologenesi [2335, 2286]. Olschki [61, 516, 517, 583, 804, 1539, 2159, 2488, 2467, 3299, 1966, 2037, 1967, 2038, 1965, 2036, 2164, 2235, 2163, 2234, 2233, 2162, 2165, 2236, 2167, 2238, 2166, 2237, 2326, 2434, 2387, 2388, 2389, 2435, 2436, 2386, 2942, 2425, 3286, 2530, 619, 2522, 3255, 2780, 278, 2865, 3015, 252, 309, 447, 576, 720, 717, 716, 887, 873, 991, 1077, 1157, 1279, 1156, 1360, 1688, 70, 1789, 3027, 1111, 3127, 2949, 413, 2058, 2531, 2572, 2888, 752, 2955, 2614, 2905, 2652, 1943, 2590, 874, 881, 2277, 2754, 3258, 1178, 2222, 2525, 455, 2574, 3088]. Olschki [2809, 799, 800, 2499, 2608, 924, 2372, 687, 1068, 2903, 1388, 1493, 2726, 2960, 2571, 1389]. Olshausen [1838]. Olson [2799]. Oltre [941]. Olympiodore [2821]. Om [1023, 1146]. Omaggio [448, 1591, 2472]. Omar [3346, 494, 618, 3740]. Ombre [3546]. Omega [746, 747]. Omer [2509]. Omnès [1413]. Omnia [720, 1279, 307]. omnium [567]. Omodeo [2453]. Omohundro [2964]. Omologia [2352]. Omori [2484]. omosessuali [573]. omosessualita [573]. once [2433]. oncogenetiche [2233, 2162]. ondulatoria [18]. One [520, 959, 3499, 3471, 699, 180, 3661, 1654, 1225, 2551, 1744, 3969, 4010, 1814]. Onere [2348]. Ongaro [2327, 2278]. Onno [3412]. onore [3772, 1820, 2117]. Ontological [2388]. Ontology [30, 58, 2900]. Oocyte [3055]. Oönskade [972]. Open [3678, 3282]. oper [1270]. Opera [307, 720, 1279, 1330, 619, 2826, 455]. operaio [810]. Operation [54]. Opere [2192, 223, 1823, 1209, 504, 1136, 2914, 1379, 2370, 1314, 1569, 1765, 3015, 2849, 991, 1156, 70, 1243, 2531, 3844, 2222, 171, 2608, 13, 3972, 2138, 2726, 1364, 2815, 1840, 2202]. operum [1687]. Ophthalmologica [1600]. ophthalmology [3856, 1600]. Opificio [2169, 2240]. opinion [3813, 2901]. opinione [3813, 1806, 1233]. Oppenheimer [1249]. Opportunities [1126, 2833]. Opsomer [746]. Optical [2635, 2292, 2088]. optician [1382]. Optics [2293, 3234, 2084, 1492, 3349, 3410, 3273, 4087, 2175, 3005, 791, 3526, 263]. Opus [3165, 1381, 1381]. Opuscoli [2448, 2603]. Orain [3087]. Orange [3604]. orarie [2148]. oration [1829]. Orationis [4059]. Orazio [2577, 1307, 559]. Orazione [1829]. Orbs [3517]. Orchestrating [2302]. Orchidées [2015]. orchids [2015]. Order [1346, 1535, 829, 2094, 3096, 3528, 3506]. Ordering [2999]. Ordinando [1903]. ordine [1148, 953]. ordre [3180]. Oregon [3254]. Oresmes [1941]. Oreste [3687]. Orfila [2907, 3770]. Organ [2870, 1092]. organa [1930]. organi [1951, 2022]. Organic [1333]. Organica [2502]. organiche [2499]. Organicisme [2735]. Organisation [507]. organisations [1933]. Organise [3700, 4043, 1440, 260, 314, 2118]. organisme [2578]. organismo [874, 924, 687]. Organization [672, 499, 3343]. organized [1440]. Organizzatore [2043]. organizzatrice [2787]. Organon [1606]. Orgasm [2016, 3916]. Oriele [355]. Oriental [1792]. Orientale [832, 1873]. Orientamenti [909]. Orig [639, 1576, 948, 1414, 1459]. Origin [3000, 2822, 286, 573, 3789, 3790, 3795, 1654, 4014, 1114]. original [226, 420]. originale [496, 532, 332, 331, 434, 436, 638, 566]. originals [2195, 4058]. origine [2337, 3013, 366, 2288, 3890, 1926, 2597, 3980, 4014, 1114]. origines [444, 2654]. Origini [2452, 2635, 604, 1971, 2044, 89, 537, 573, 3299, 1239, 3760, 18, 1574, 2616, 449, 971, 1620, 3115, 2905, 874, 785, 1176, 500, 875, 203, 924, 1068, 1389, 2865]. Origins [592, 2635, 2636, 2850, 971, 3176, 2697, 3658, 526, 3265, 2835, 2718, 2926, 2807, 3126, 3516, 2044, 3746, 3760, 769, 1944, 2310, 4011, 4013, 3082]. Orinoco [3073]. Orio [2667]. Orittologo [740]. Orlandi [1158, 1343]. Orléans [1187]. Ornaments [2086]. Ornavasso [2965]. Ornella [2380, 3259]. Oro [995]. Orologi [425, 2148, 1415, 1787, 2001, 429]. orologiai [425, 2148]. orologiaio [250]. Orologio [1215, 902, 428]. Orologist [3701]. Oronce [3171]. oroscopo [752]. Orphica [798]. Orrery [3515]. Orsini [861]. Ørsted [2967]. Ortega [3446]. Ortelio [1906, 3918]. Ortelius [3918]. Orthodox [3448]. orthodoxes [3448]. Orthodoxy [224, 1490, 3267, 3947]. Orti [1267]. Orto [917, 1352, 863, 1137]. Orville [116]. Oscariansk [2675]. Oschwald [110]. oscuro [1677]. Osiride [2619]. Oskar [226]. Osler [164]. Osorio [2824]. Ospedale [1784, 1192, 1335, 1436, 1755, 2871]. Ospedali [1778, 496, 710, 1599]. Ospedaliere [1778]. Ospedaliero [953]. ospizio [2072]. Osseo [1551]. osseruatorio [1863]. Osservatore [740]. Osservatorio [688, 3138, 390, 187, 453, 655, 590, 700, 937, 1014, 1430, 2110]. Osservazioni [779, 2955, 1098, 2608, 4016, 2726, 1235, 817, 2499]. Ossietzky [3905]. Österreichische [382]. Österreichischen [1854, 382]. Österreichs [253, 492]. Österrike [1019]. ostetriche [956]. ostetricia [2414]. ostetrico [38]. Otava [673]. Otero [2571]. Other [955, 3394, 3225, 3412, 3199, 3513, 3354, 3291, 1611, 2886, 1592, 3682]. Othmar [2375]. Ott [2529]. Ottaviani [3687, 3652]. Ottaviano [2252]. Ottavio [2167, 2238]. Otte [535]. ottenuto [44]. Ottica [790]. ottiche [854, 1794]. Ottici [2174, 75]. Ottico [2635, 988, 1382, 1554, 1893, 2088]. Otto [915, 1191, 1759, 2140, 132]. Ottobre [1015, 2389, 2522, 2376, 311, 809, 2949, 2492, 2880, 3967, 1137, 1530]. ottocentesche [2616]. Ottocento [3544, 662, 1331, 1519, 818, 346, 4049, 2315, 3103, 3029, 2206, 3965, 778]. Ottoman [2749]. Ottoson [86]. Ottosson [274]. Ouest [1612, 175]. Ouest-France [1612]. Our [3626, 3636]. outils [3280, 4063]. Outram [440, 587, 1536, 17]. outstanding [303]. ouvrages [541]. Ove [1078, 1805, 696, 699]. Oven [3244]. Overlooked [3485]. Overseas [1944]. Overture [3184]. Overview [2924]. ovo [2205, 4023]. Owen [2560, 1193]. Owens [3653]. Own [3499]. Oxford [1070, 1087, 1406, 1763, 2463, 2746, 2459, 2717, 1852, 3113, 2606, 2538, 2852, 2851, 3081, 3025, 196, 192, 3124, 3263, 3066, 2706, 2600, 2595, 3108, 3131, 925, 2947, 2681, 2095, 3024, 3074, 2912, 3014, 1596, 1100, 839, 25, 1843, 1511, 1662, 2310, 534, 2686, 2204, 450, 451, 2506, 2830, 2998, 2609, 2378, 1007, 2687, 1844, 1907, 539, 2861, 694, 2587, 2582, 946, 489, 2848, 2054, 209, 3913, 1636].

Oxford-New [2746].

P [242, 852, 2016, 2055, 3065, 688, 1682, 287, 2543, 3369, 28, 2962, 3110, 2773, 2832, 204, 758, 1915, 1944, 868, 14, 390, 2660, 184, 291, 293, 452, 655, 590, 730, 1062, 1813, 3078, 255, 1541, 379, 1142, 1324, 747, 178, 2688, 1002, 3280, 3526, 3079, 2222, 812, 3284, 1868, 1989]. P. [1589, 235, 1447, 1448, 869, 1192, 1677, 608, 1071, 1391, 1245]. P.-S [1682]. P.Ri.St.Em [1350]. PA [1814]. Pablo [3444]. Pabst [1446]. Pacciarelli [2007]. Pacific [3887, 1334, 2846]. Pacini [1316, 917, 776, 1768, 1269, 3246]. Pacinotti [2974]. Pacioli [1284]. Paderborn [248]. Padova [621, 1487, 1236, 1496, 74, 2588, 448, 645, 356, 864, 1056, 1195, 1423, 1532, 1061, 1579, 2107, 2136]. Padre [981]. Padua [3691, 1195, 3487, 3709, 3409]. paese [2893]. Paesi [14, 1906, 3918]. pag [1619]. Pagan [3166]. Paganini [181, 2107]. Pagden [537]. Pagel [612]. pages [2274]. Pagnoni [307]. Pagnoni-Sturlese [307]. págs [353]. Painted [1566]. painter [3777]. Painting [2970, 3393, 1986, 3635]. Pais [1007]. Països [3019]. Pal [1134]. Palace [3579, 429]. Paladini [2974]. Palaeontology [1039]. Palas [494, 618]. Palatinate [787]. Palazzi [1268]. Palazzo [429, 1325]. Paleis [302]. Paleologo [2151]. Paleontologia [493, 3618]. Paleontology [3078]. Paleopathology [367]. Palermo [258, 514, 764, 1903, 1156, 572, 930, 1863, 2062, 2208, 2135, 3154, 3308, 1220, 3847]. Palestine [3564]. Palimpsest [2881]. Palisca [2776]. Palladio [2001]. Palletta [929]. Palm [1367]. Palma [330, 859, 2111]. Palmieri [1134]. Paloma [2571]. Palombi [1816, 859]. Paloscia [1287]. Paltrinieri [1013, 1160, 1415, 1787, 2153]. Palumbo [258]. Pamela [1225, 3479, 3254, 3082, 3070]. Pamphilj [2086]. Pancaldi [2463, 2558, 2078, 2498]. Pancino [74]. Panconesi [1028]. Pandolfini [1077]. Panicali [425]. Panini [2461, 1190, 435, 1906]. Panisperna [2910]. Panopolis [2704]. Panopticon [2051]. Pansini [729]. Pansophia [2159]. Pansophiae [3064, 3283]. Pantegni [1358]. Panti [2934]. Pantin [1567, 2652]. Panza [2654]. Panzeri [929]. Paola [1574, 951, 1111, 2893, 1029, 2743, 1794, 1658, 3459, 1201, 2351, 2672, 1148, 1318, 1565, 2498]. Paolantonio [1801]. Paoletti [2271]. Paoli [3257, 1663]. Paoline [877, 882]. Paolo [202, 3739, 1957, 2028, 2168, 2239, 2338, 2436, 339, 1317, 2522, 1693, 2144, 1360, 2157, 3397, 2492, 2623, 2659, 3498, 3546, 1614, 2117, 1653, 1856, 3085, 2384, 2921, 1252, 1802, 3194, 899, 3120, 931, 3592, 449, 2490, 2535, 3115, 284, 2742, 3426, 426, 1437, 1532, 550, 1291, 1498]. Paoloni [764, 2216, 1741, 2741, 2274]. Papa [2461, 1399, 2198, 2420, 3808]. Papa-Livro [2198]. Papagno [2754]. Pape [1391]. Paper [2662, 173, 172, 3345, 3412, 1841, 1872, 2067, 2424, 64, 2719, 1987]. paperback [1818, 2075, 613, 624, 2362, 1843, 2456]. Paperbacks [537, 2471, 418, 755, 1639]. Papers [1495, 2852, 63, 182, 2579, 2873, 2185, 4068, 2189, 849]. papi [2420, 1335]. pappagallo [2108]. Pappelbaum [147]. Pappus [3164]. Papyri [3500]. papyrus [1432]. Para [2356, 2357]. Paracelsian [2658]. paracelsiana [1680]. Paracelsianism [3590]. Paracelsians [837]. Paracelsica [1396]. Paracelsisten [492]. Paracelsists [492]. Paracelso [1680, 612, 310, 951]. Paracelsus [253, 568, 217, 1285, 492, 253, 4061, 492]. Paracelsus-Forschung [1285, 4061]. Paracelsus. [3072, 3671]. Paracelsusforschung [253, 492]. Paradigm [845]. Paradigmatic [3225]. Paradigme [1187]. Paradigms [693]. Paradoxes [1668, 3781, 3781]. Paragraph [212]. Parallèles [2401]. Parallels [3601]. paranormales [3427]. Parapegmata [2879]. Parassiti [708]. Paratexts [3358]. Paravia [2232]. Paravicini [3631]. Parcours [3330, 3767, 1727]. Pardo [889]. Parés [3019]. Pareti [2233, 2162, 765]. Parigi [2197, 638, 2415, 1493, 2298]. parigina [719]. Parigino [1523, 1521, 1728]. Parini [1691, 2467]. Parini-Chirio [1691, 2467]. Paris [728, 795, 1090, 1795, 2267, 2318, 2495, 1578, 496, 653, 3107, 222, 282, 416, 675, 1418, 1793, 2053, 2896, 4047, 2735, 548, 584, 620, 796, 834, 1060, 1155, 1440, 1681, 1727, 1804, 1911, 1981, 2894, 2933, 2898, 287, 2528, 260, 2798, 3030, 243, 305, 512, 574, 719, 836, 1444, 1985, 2611, 2648, 2958, 1138, 1327, 1668, 2215, 2229, 1667, 3013, 40, 33, 3087, 314, 368, 2654, 2829, 754, 1543, 2973, 2892, 536, 748, 2533, 681, 419, 866, 1878, 1875, 2013, 332, 331, 434, 436, 566, 872, 1432, 3251, 2821, 920, 3277, 3287, 3257, 540, 701]. Paris [702, 1170, 1563, 2118, 3932, 1848, 1410, 2060, 216, 2952, 3010, 2537, 1228, 2767, 2663, 1647, 1646, 1073, 1460, 1799, 2052, 1413, 2729, 625, 1326, 2273, 3075, 3004, 1232, 1386, 1486, 1534, 2422, 2613, 2782, 2578, 3767, 3469, 2183, 3341, 210, 3877, 2048, 2566, 3932, 1473, 1329, 4005, 2188, 4029, 3330]. Parisi [700]. Parisini [2148]. Park [1910, 2312, 34, 2835, 2534]. Parkinson [3038, 2125]. parla [3908, 3125]. Parlatore [930]. Parma [2368, 1352]. Parmentier [681]. parnaso [2167, 2238]. Parodi [2413]. Parola [2445]. Paroxysms [1199]. Parrini [2742, 1251, 1498]. Parrish [2964]. pars [1280, 4028]. Part [802, 1288, 2495, 1670, 2719, 3345, 1484, 1489, 1193, 2204, 1225, 1641, 2198, 4028, 3852, 3853, 4004]. Parte [1569, 1765, 221, 171, 12, 1091, 1982, 2198]. partécipation [314]. Partenza [2254]. parthénogénèse [3448]. parthenogenesis [3448]. Parthenon [1179]. participation [4043]. Particle [2970, 860, 4054, 4055]. Partielle [1474]. partigiano [3121]. Parts [3412]. Parva [58]. Pascal [2826, 295, 2772, 2053, 3043, 2594, 1944, 3646]. Pascale [1456]. Paschetto [31]. Paschoud [3024]. Pascolini [315]. Pasini [1140]. Pasquale [3541, 2727, 3497, 71, 631, 1745, 2005, 12, 2217, 3314]. Pasquinelli [1502]. pass [1107]. Passages [3757]. Passaggi [3757, 2919]. passaggio [2597]. Passato [1113, 1925, 854, 1137]. passi [421]. Passione [2855, 3039]. Passioni [643]. Passions [1199, 643]. PäßLer [3073]. Past [439, 3362, 3439, 3793, 3796, 1812, 1934, 1654, 3996, 1137, 3289]. Pasta [619, 455]. Paster [2655]. Pasteur [957, 3692]. Pasticci [2349]. Pastore [2836, 3115, 3206]. Pasts [3428]. Paterna [2593]. Path [2658]. Pathologies [3319]. Pathology [1936, 177]. Pathos [3502]. Pathways [885, 2626]. patient [1329, 4029]. Patientenwege. [3116]. Patients [80, 1392, 2365]. Patologia [869, 3179]. Patria [2667, 2100, 688, 655]. Patrice [3368, 1495, 1795, 2229, 2118]. Patricia [3499, 1725, 2527]. Patrick [275, 1944, 1073, 574, 649]. Patrimonio [605, 1861, 1931, 2149, 2414, 1260, 1062, 2068, 3990]. patriot [1589]. Patristica [1685]. Patrizi [1647]. Patrizia [2148, 3413, 42, 881, 66, 346]. Patron [336]. Patronage [1617, 2715, 4080, 1063]. Patroni [1382]. Patrons [3766, 2953, 2098]. Pattaro [1325]. Patterns [2626, 3245]. Patterson [1654, 3796]. Paul [164, 441, 2878, 1412, 1138, 219, 172, 244, 385, 3225, 586, 2884, 356, 1249, 2132, 2417, 3059, 1581, 2633, 3131, 606, 845, 1933, 2076, 2541, 3093, 2584, 372, 840, 768]. Paul-Valry-Montpellier [2878]. Paula [2895, 2783, 3412, 1319]. Pauline [1656]. Paulus [58]. Paume [3234, 3231]. paumier [3230]. paura [1899]. Pauwels [2867]. Pavese [1801, 1572]. Pavia [1801, 1145, 1156, 1243, 1499, 1572, 1869, 2156, 713, 1452, 3487, 2476, 954]. Pavia-Milano [954]. Pavone [2389, 13]. Payne [3258, 1592]. Payne-Gaposchkin [1592]. Payot [331, 566]. pays [728, 2519, 3829, 2060]. Pays-Bas [3829, 2060]. paziente [80]. pazienti [1598]. pazzi [1335]. pazzia [788, 1098]. pbk [1987]. PD [2602, 2458]. PDF [3530]. Peace [2977, 3974]. Peano [982]. Pearl [3683]. Pearn [3076]. Pearson [2792, 68]. peces [3618]. Pecs [2421]. Pedemontana [2530]. Pedemontanam’ [2531]. Pedemontii [2531]. Pedersen [1081, 2044]. pediatria [2414]. Pedretti [1950, 2021]. Pedro [1744, 2356, 882]. Pedull [2202]. Peep [3150]. Pegler [2719, 3345]. Peirce [1148]. Peiresc [921]. Pelageya [591]. Pell [2717]. pelle [3751, 2628]. Pellegrini [726]. Pelling [2755]. Peltonen [1649]. pendant [314]. Pendragon [3125]. Pendulum [3706]. Pendulums [1208]. Penelope [830, 998, 2056, 409]. Penetralia [1323]. Penetration [1051]. Penfield [3199]. Penguin [1852]. péninsule [1135]. Pennsylvania [805, 1819, 1910, 2312, 2959, 2657, 385, 763, 2534]. Pennuto [2608, 3329]. Penrose [694]. Pensa [2144]. pensano [3804, 2323]. Pensare [2383]. pensé [2767]. pensée [413, 1121, 338]. Pensées [2869]. Penseur [1211]. Pensieri [932]. Pensiero [643, 1691, 3065, 1239, 1903, 3001, 758, 1866, 1718, 1140, 13, 1349, 2269, 1223, 2644, 2668]. Penso [332, 436, 708, 870]. People [225, 3191, 3922, 2382, 972]. Pepe [2270, 727, 1136, 2460]. Peperoni [1591]. Pépin [2648]. Pequeña [2198]. Pequignot [638]. Per-Gunnar [86, 274]. perception [1918]. Perceptions [1458]. Perceptual [450]. percezione [1788, 1485]. Percezioni [3062]. Perché [1112, 3831]. Percival [2450]. Percorsi [3001, 3544, 1579, 952, 2117]. Percorso [185, 2632]. perduto [2881, 771]. peregrinatio [1230]. Peregrinus [847]. Pereira [995, 610]. Peressut [1923]. Perfect [3302, 2604]. Perfected [1044]. Perfectionis [997, 3946]. perfectly [4051, 847]. Performance [3153]. Perga [1973]. Pergame [3491]. Perge [3163]. Perhaps [3301]. Periccioli [197, 561]. Pericle [1569, 171, 1765]. pericolosa [1921]. Perillo [2934]. Perinaldo [1453]. Period [3274]. période [314]. Periodic [3124]. periodica [180]. periodical [2737, 180, 3837, 927]. Periodicals [3637]. Periodici [1159, 2261]. periodico [1870]. periodiek [280]. periods [3944]. peripatetic [2308, 3880]. Peripatoi [443]. Periphery [1264]. Pérouse [1304]. Perreiah [58]. Perron [2073]. Perrone [1229]. Perrot [584]. Pers [280, 541]. Persia [3291]. Persicus [256]. Persistent [3048]. personale [2468]. Personalising [3561]. personalità [1815]. Personality [229]. Personenregister [362]. Personhood [2506]. Persons [213]. perspectiva [1493]. Perspective [1370, 206, 460, 1377, 2993, 611, 2360, 2339, 2550, 1812, 3653, 3526]. Perspectives [394, 1685, 2553, 1395, 1812, 693, 2925, 3684, 1935, 1184, 3565]. Persuasion [2360]. Pertlik [1854]. Peru [686, 3993, 3316]. Perugia [991, 2883, 635, 1840, 1267]. Peruta [1820, 3772]. Pes [1558]. pesare [2135]. Pesaro [858]. Pesaro-Cassa [858]. pesatura [1011]. Pesce [2720, 2141]. Pesendorfer [389, 589]. Pesic [2806]. Pesole [1084]. peste [1096, 1181, 3010]. pestifero [2714]. Pestre [2613]. Peter [2601, 130, 3694, 847, 3298, 172, 547, 3018, 183, 1792, 1058, 1285, 1980, 2231, 2412, 2983, 2804, 1854, 380, 2146, 1905, 4035, 60, 439, 724, 805, 134, 140, 652, 1095, 63, 1263, 3341, 3288, 2747, 2862, 191, 2966, 2987, 1792, 2195, 1723, 2784, 849, 942, 1198, 2806, 2322]. Petersburgh [2662]. Petit [3650, 3491]. Petite [675]. Petits [1888]. Petrarca [950, 2112]. Pétrarque [1228]. Petri [1685]. Petri-Brepols [1685]. Petrifaction [3215]. Petrology [2161, 4066]. Petronio [1013]. Petrus [2658]. Petry [344]. Pettenati [83]. Peutingeriana [2572]. Peyrouse [1476, 1475]. Pfalz [787]. Pfalzmuseum [787]. Pfinzing [3521]. Pflanzengeograph [787]. Pflugg [173]. Phaenomen [188]. Phaenomena [3325]. Phaidon [489]. Phaidon-Christie [489]. phanerogamic [2230]. pharaonique [1432]. Phares [1447]. Pharmaceutical [1805, 2312, 4034]. Pharmacie [1602]. Pharmacist [3644]. pharmacology [435]. pharmacopée [746]. pharmacy [2228, 3902, 4035]. Pharmakon [705, 878]. Pharmazie [2228, 1724, 1714, 3902, 1905]. Pharmaziegeschichte [1714]. Pharmaziegeschichtsschreibung [1905, 4035]. Phebe [3612]. Phenomena [2478]. phenomenology [4021]. phénomnalisme [2469]. Philadelphia [805, 1819, 2495, 2959, 857, 1359, 2657, 385, 829, 2137, 763, 1193, 2982, 2890, 2724]. Philip [2936, 1721, 3527, 2938, 2133]. Philipp [2853]. Philippe [3013, 1133, 3071]. Phillip [1154]. Philo [443]. Philo-logisch-historische [443]. Philologe [173]. Philologie [173, 746, 747]. Philonous [255]. Philoponu [443]. Philosoph [172]. Philosopher [3757, 3025, 2489, 1111, 3872, 3067]. Philosophers [1213, 719, 1603]. Philosophia [30, 58, 890, 1896, 2394, 1225, 4026, 1229, 3283]. Philosophiae [4028]. philosophica [2138]. Philosophical [891, 1819, 2495, 2586, 3425, 857, 2658, 3366, 3139, 829, 2137, 2790, 1193, 2982, 2890, 2763, 1829, 3842, 3906, 1008, 993, 2586]. Philosophico [93]. Philosophico-Politica [93]. Philosophie [571, 798, 994, 1359, 3173, 172, 338, 1228, 2428, 1202, 2465, 380, 1232, 1989, 4033, 512, 1984, 4077, 2767, 1460, 3571, 2829, 1646, 1413, 1799, 2729]. Philosophies [1875, 1534]. Philosophique [512, 2821, 1228, 2578, 548, 338, 2767]. philosophiques [2648, 308]. Philosophisch [382]. Philosophisch-historische [382]. Philosophische [379, 380]. Philosophorum [307, 720, 717, 1279, 1688, 70, 991, 1156]. Philosophy [301, 918, 1288, 2586, 3508, 2708, 3509, 669, 823, 907, 1478, 1556, 1635, 2606, 652, 3230, 1842, 3162, 549, 30, 58, 86, 511, 1103, 1104, 1445, 274, 1500, 1263, 1776, 2873, 209, 2771, 2828, 2681, 2137, 55, 271, 326, 835, 1570, 2858, 3224, 2124, 2803, 2585, 4084, 892, 1144, 3571, 283, 1721, 3812, 3838, 2308, 3880, 3897, 3904, 1934, 4077, 3925, 3933, 3949, 3950, 4083, 806, 1007, 1109, 1989, 4030, 4033, 4024, 2889, 4026, 4028, 1403, 3735, 3476, 3671]. Phoenix [256, 3729]. Phonographs [356]. Phosphorus [36]. Photo [226]. Photo-mechanical [226]. Photograph [3050, 3052]. Photographing [3681]. Photographs [2303, 3578]. Photography [3182, 2794, 2970, 3635, 3208, 3215]. Phsychology [1225]. Phylogeny [3610]. Physica [3366]. Physical [2917, 2258, 1786, 2457, 595, 3941]. Physical-mathematical [3941]. physichen [1004]. Physician [16, 2994, 2830]. Physicians [3442, 57, 3594, 3734, 3977, 3506, 303, 3907, 955]. physicist [1726]. Physicists [772, 2912, 2913, 1576]. physico [2337, 2288, 2114, 3941, 1647]. physico-historica [2337, 2288]. physico-mathematica [2114, 3941]. Physicorum [2520]. Physics [1199, 4041, 2168, 2239, 2700, 549, 2105, 29, 1443, 1067, 2161, 2216, 2698, 2305, 3232, 2970, 3900, 3243, 2207, 2701, 1072, 1084, 1792, 2676, 1618, 2585, 1225, 1920, 1051, 3744, 772, 4054, 4055, 3791, 3792, 3793, 3794, 1879, 4072, 3889, 938, 2423, 3950, 3966, 860, 1007, 1425, 1654, 3999, 4089, 4019, 1231, 1264, 694, 3239, 3823]. Physik [1941, 415, 3999]. Physikalisch [1340]. Physikalisch-Technische [1340]. physikalische [1151, 542]. physikalischen [1012]. Physikalischer [1083]. Physikerin [1726]. Physiognomy [654, 3748, 3566]. physiologique [3026]. Physiologische [2853]. Physiology [1225, 3571, 3933, 3822]. physique [2898, 1879, 4089]. Physis [2499]. phytonymes [1224]. Pia [2969, 1062, 1325]. piacere [1580]. Piaget [1510, 1510]. Pianeta [2511]. Pianeti [1215]. piano [1160]. Pianta [2391, 435]. Piante [1926, 4014, 2012, 3777, 3844, 436, 1268, 1271, 861, 1269]. pianto [2276]. pianura [3710]. Piaskowsky [1584]. Piave [1300, 736]. Piazza [1856, 2499]. Piazzette [1968]. Pica [1767]. Picard [540]. Picardie [32]. Picatrix. [334]. Piccin [2107]. Piccolino [2435, 2630, 2904]. piccolo [315]. Pichiani [1360]. Pick [850]. Pickering [2781, 1504, 2956]. Pico [2942, 1360, 1565, 1360]. Picolet [540]. Picta [2882]. Pictet [2539]. Picture [2864, 3047, 3052]. Pictures [3735, 1628]. Picturing [3662, 3607, 2689, 2506]. Pieces [2175]. pied [3087]. piedi [449]. Piedmont [778, 1020, 2363]. piega [756]. Piemonte [2376, 2789, 778, 1020, 2363]. piemontesi [1599]. Pier [2276, 407, 1512, 558, 1252, 407, 558, 1426]. Pieraccini [1186]. Pierangelo [1271]. Pierleone [2222, 3972]. Pierluigi [715, 719, 2213, 1368]. Piero [223, 1283, 45, 1794, 1483, 258, 2230, 1377, 678]. Pierre [2933, 1357, 1138, 1859, 1908, 2892, 2537, 3284, 2428, 2469, 305, 2359, 1560, 1879, 2771, 3875, 2533, 3236, 500, 3287, 1896, 2728, 3704, 4000, 3004, 2627, 2748]. Pierre-Jean [1138]. Pierre-Joseph [3236]. Pierre-Marie [2933]. Pierre-Simon [3875]. Piers [2709]. Piersandro [2156]. Piet [3109]. Pietà [1335]. Pietrasanta [1488]. Pietre [2169, 2240, 563]. Pietro [197, 1967, 2038, 2332, 1569, 1765, 2116, 31, 2283, 2590, 3031, 171, 20, 1159, 2574, 412, 2581, 1267, 2692, 514, 3030, 2453, 3844, 596, 1382, 2743, 1373, 211, 1851, 1840, 1651, 1719]. Pietsch [3268]. Piety [1596]. Pigeaud [868]. Pighetti [515]. pilgrimage [1230]. Piliarvu [2213]. Pilotto [2001]. Pimentel [3634]. Pina [1943]. Piñero [43, 2101, 1849, 2114, 2097]. Pinerolo [2376, 2650, 1453]. Pingree [334, 3111]. Pino [2113]. Pio [1675]. Pioneer [968, 3797, 2381]. Pions [860, 4054, 4055]. Pious [4034, 2312, 3455]. Piovan [626]. Pipistrelli [817]. Piramide [1874]. Pirella [1730]. Piro [2880]. Pisa [4040, 1316, 3080, 2974, 3030, 809, 3298, 2988, 2623, 863, 917, 941, 1614, 2677, 1645, 749, 2272, 1994, 2951, 3203, 1137, 1269, 607, 1134, 3259, 3285, 4040, 809, 3824, 2623]. Pisa-Roma [1994]. Pisa/Roma [2951]. Pisano [3286, 3824]. Pistacchio [871]. Pistoia [991]. Pitagora [493, 1802]. Pitrelli [2692]. Pitti [2345]. pittore [2012, 3777]. Pittura [1950, 2021, 2142, 2490]. Più [3799, 3285]. Pizzigalli [2415]. Pizzorusso [684]. pl [914]. Placanica [39]. Place [2560, 1682, 56, 65, 2745, 2180, 2562, 3762, 1187, 1915]. Placucci [1677]. Plague [3559, 3597, 1392, 194, 3572]. Plain [1267]. Plana [2047]. planches [608]. Planck [3282]. Planes [1208]. Planetarium [3111, 2187]. Planetary [3513, 4003, 4004, 3516, 1215, 3791, 3792, 3793, 3794, 1654, 1670]. Planimeters [3104]. Planisferologio [281]. Plant [2391, 787]. plantarum [945]. plantas [2097]. plantes [1418, 436, 1224]. Plants [3411, 816, 3114, 840, 3777, 3844, 1271, 1269, 4014, 3570, 3713, 3691]. Plastic [3857, 3385]. Plastica [794]. Plastination [2265]. plastische [2682, 3857]. plates [2595, 2590, 3528, 2628, 2650]. Platone [1594]. Platonismo [680]. Platonismus [571]. Players [3230]. Pleasure [3679, 762]. plein [3087]. plein-pied [3087]. Plessner [2276, 2502]. Plight [526]. Pliniiconcordantiae [747]. Pliny [2706]. plots [418]. Pluchinotta [1027]. Plurality [723]. Plus [3030, 2988, 2623, 3058]. Pluta [380]. Pneumatics [3274]. Pnina [440, 2067]. poche [1646]. Pocket [796, 866]. Podgorny [3618, 3264, 3618]. Poem [2247]. poems [1611, 3905]. Poesia [2395]. poésie [1567]. Poet [3736]. Poetiche [1411]. Poetics [1411, 2976, 1225]. Poetry [2686, 3034]. Poets [2991]. Poggi [1570]. Poggibonsi [1695]. Poggio [1241, 1052]. Pogliano [1967, 2038, 1948, 2019, 3426, 705, 878, 2677, 3619]. Poincaré [1289, 1850, 3848]. Point [1377]. Poirier [2533, 2541]. Poitiers [444, 1609]. Poitou [1609]. Poitou-Charentes [1609]. Polaris [1481]. Polarlichts [188]. Polemical [3173]. polemiche [3908, 3125]. polemische [217]. Poleni [74]. Policies [3081, 1411, 1999]. Policy [2972, 3851]. Polignac [282]. Poligrafici [1994]. Poligrafico [2202, 1158, 1665]. Poligrafo [1236, 1061, 1579]. Polish [1584]. Polistampa [2270, 2373, 2266]. Polite [1763, 4001]. politecniche [4067, 2367]. Politecnico [1780, 2622, 3077]. Politica [1427, 2473, 2092, 93, 1311, 197, 619, 2670, 3851, 2154, 2275, 2918, 455]. Political [587, 2589, 209]. politiche [1411]. politico [1648]. politico-medici [1648]. Politics [167, 1362, 3081, 1738, 2079, 3692, 2657, 2839, 3511, 3738, 2977, 3691, 2866, 3528, 2582, 3278, 1225, 2545, 3452, 768, 1194, 3756, 3765, 1505, 760, 3895, 3935, 3974, 3669]. politique [3765, 1878]. politiques [3895, 582]. Polity [2538, 649, 3950, 1007]. Poliziano [1360]. Polizzi [1289, 2501]. Polla [1181]. Polli [1890]. Pollich [787]. Pollichia [787]. Polskich [1584]. Polytechnic [4067]. polytechnique [1090]. polytechniques [2519]. Pomar [816]. Pomata [2656]. Pomet [1095]. pompa [1533]. Pompei [1214]. Pompeii [1839]. Pompeo [2380, 3259]. Pomponazzi [3256, 3923]. Pontaniana [2226]. Ponte [952, 1255, 1840]. Pontecorvo [785, 875, 2918]. Pontifical [2176]. Pontificia [2176]. Pontificio [707]. Poor [2962, 194]. Pope [3808]. Popes [2379]. Popper [2740]. Poppies [3607]. populaire [1848]. populaires [1170]. Popular [3450, 1835, 3558, 45, 3578, 1170, 1855, 2901]. Popularising [3561]. Popularizers [3033]. populärmedicin [2500]. Population [1245]. Porcia [336, 522]. Porret [2843]. Porretta [1343]. Porro [1671]. Porta [2370, 2712, 3443, 3410, 1331, 215]. Porta-Bozzolo [1331]. Porta-Cocoon [1331]. Portable [1342, 3]. Portal [3404, 3727]. Portar [2247]. Porter [224, 342, 642, 1449, 2464, 1239, 532, 2631, 174, 225, 2004, 2071, 554, 1182, 1433, 1655, 2792, 1845]. Porto [2895]. Portoghese [2356]. Portrait [176, 2391]. Portraits [2576, 1269, 2988]. Portugal [3562, 160, 1557, 2925]. Portuguese [1744, 2356, 2357, 3833]. Porzio [3279, 3826]. Posamentier [2922]. positive [3799]. Positivism [202, 3701]. positivisme [2843]. positivismo [1616, 3426, 202]. positivisti [3799, 3285]. Positivistica [1120]. positivists [3799]. positivo [3799, 3285]. Positron [1296]. Possessing [1319]. possessioni [3178]. possibile [1901]. possibili [2132]. Possibilità [2258]. Possibilities [2568]. Possibility [2258]. Possible [2800]. Post [1778, 17, 759, 81, 1382]. post-fazione [81]. post-modern [759]. Post-Revolutionary [17]. Post-Spagnola [1382]. Post-Unitaria [1778]. Post-Unitarian [1778]. Post. [2906]. Postal [3680]. Postdoctoral [2686]. postfazione [771, 1728]. posthume [338]. Postmortems [3455]. Postume [1379]. Postumum} [1381, 1381]. potente [2504]. potentia [250]. potenza [2670]. poteri [1899]. Potter [608]. Poul [235]. Poulets [1897]. Poultry [1161]. pounds [616]. Pour [2067, 748, 2197, 1475, 2396, 1474, 4039, 4047, 2798, 1305, 1326, 2613, 3107]. pouvoir [3491]. poveri [1335]. Powell [959]. Powell-Cope [959]. Power [1070, 2607, 649, 1588, 3580, 3511, 4046, 1758]. Pox [1720]. Pozza [1056, 1435]. Pozzi [726]. Pozzo [899, 2719, 3345, 3412, 921, 1111]. Pozzuoli [3209, 2796]. pp [39, 15, 61, 197, 224, 302, 296, 303, 313, 299, 516, 515, 517, 441, 520, 491, 440, 514, 490, 537, 583, 617, 567, 794, 728, 777, 764, 672, 685, 804, 885, 894, 961, 889, 895, 891, 1110, 1070, 1005, 1091, 995, 1055, 1090, 1108, 1200, 1288, 1234, 1201, 1264, 1261, 1333, 1365, 1402, 1457, 1406, 1427, 1364, 1454, 1531, 1536, 1528, 1502, 1539, 1602, 1606, 1589, 1587, 1683, 1691, 1657, 1725, 1819, 1818, 1846, 1853, 1795, 1945, 2003, 1910, 2016, 2119, 2120, 2075, 2102, 2056, 2067, 2194, 2227, 2143, 2145, 2228, 2159, 2200, 2160, 2267, 2312, 2318, 2421, 2420, 2370, 2424, 2488, 2463]. pp [2467, 2496, 2495, 2500, 2464, 2560, 2586, 2601, 2527, 2715, 2675, 2746, 2776, 2793, 2887, 2953, 2986, 2916, 3041, 3016, 3299, 2418, 2459, 2516, 2573, 2647, 2708, 2710, 2879, 2881, 3000, 2055, 2843, 2717, 2607, 2899, 1852, 1921, 2878, 2889, 3065, 593, 1578, 1949, 1956, 1957, 1959, 1962, 2020, 2027, 2028, 2030, 2033, 1963, 2034, 1951, 2022, 1952, 1953, 1954, 2023, 2024, 2025, 1955, 1960, 1966, 2026, 2031, 2037, 1967, 2038, 1948, 1950, 1958, 2019, 2021, 2029, 1947, 1965, 2018, 2036, 2012, 1961, 2032, 2164, 2235, 2163, 2168, 2234, 2239, 2233]. pp [2162, 2169, 2240, 2165, 2236, 2167, 2238, 2166, 2237, 2329, 2331, 2337, 2328, 2332, 2326, 2335, 2330, 2333, 2334, 2336, 2338, 2339, 2327, 2434, 2387, 2388, 2389, 2435, 2436, 2386, 3113, 2142, 339, 496, 588, 726, 934, 2942, 3543, 1736, 2425, 2462, 2606, 2730, 2897, 1573, 3116, 3064, 3002, 3107, 3120, 3286, 3283, 340, 336, 282, 416, 493, 421, 675, 688, 931, 1903, 2612, 2896, 2669, 3094, 2735, 2737, 3005, 18, 2777, 620, 612, 686, 676, 725, 703, 698, 673, 696, 699, 841, 884, 834, 846, 1006, 1097, 1095, 1060]. pp [1155, 1197, 1171, 1145, 1314, 1316, 1322, 1440, 1317, 1439, 1488, 1487, 1564, 1569, 1518, 1495, 1533, 1608, 1648, 1661, 1681, 1727, 1643, 1682, 1765, 1689, 1804, 1811, 1817, 1898, 1911, 1981, 2517, 2530, 2538, 2822, 2894, 2933, 2898, 2882, 2936, 3063, 3040, 3059, 3071, 3121, 3109, 2852, 2851, 3081, 2491, 2522, 2707, 2864, 2770, 2981, 2891, 3068, 3255, 3254, 3282, 932, 1769, 2320, 2431, 2364, 2471, 1004, 2664, 2780, 2653, 1835, 3072, 2805, 2964, 2580, 2203, 2472, 2144, 2543, 3080, 2093, 2750, 2682, 2284, 2528, 2667, 3545, 3025, 1738, 2836, 2837, 847, 1085]. pp [1112, 1808, 2524, 2631, 2721, 2797, 2959, 2906, 3009, 2850, 3023, 2865, 1991, 180, 182, 2313, 2314, 2376, 2105, 2689, 2106, 2972, 2798, 2769, 2309, 3279, 2523, 2588, 3015, 3204, 2853, 2979, 2908, 3082, 3084, 3129, 3123, 3122, 3268, 2849, 3021, 2685, 2826, 2616, 2551, 2558, 2671, 2791, 2909, 3035, 2542, 2974, 2969, 2670, 28, 72, 29, 85, 30, 23, 31, 58, 86, 32, 194, 243, 252, 250, 251, 193, 196, 198, 192, 248, 249, 188, 312, 311, 307, 310, 306, 305, 309, 443, 512, 442, 444, 446, 511, 447, 445, 449, 448, 576, 648]. pp [574, 649, 715, 720, 717, 716, 718, 719, 966, 887, 964, 886, 890, 836, 857, 830, 915, 994, 991, 1102, 1103, 1105, 1077, 1104, 996, 990, 1191, 1277, 1283, 1278, 1157, 1279, 1280, 1273, 1189, 1187, 1188, 1281, 1156, 1359, 1446, 1358, 1443, 1447, 1357, 1442, 1445, 1360, 1687, 1688, 1685, 1686, 1941, 3001, 70, 792, 951, 1462, 1417, 2283, 2279, 2648, 2963, 2958, 2553, 1151, 2917, 3124, 3544, 1566, 1717, 2103, 2371, 2602, 2619, 2665, 2957, 2592, 1243, 1238, 1147, 1366, 1327, 1612, 1516, 1515, 1572, 1620, 1494, 1668, 1758, 1670, 1862, 1869, 1791, 1859, 1820, 1879]. pp [2010, 2007, 2014, 1928, 2080, 2079, 2059, 2066, 2161, 2232, 2216, 2215, 2224, 2229, 2157, 2156, 2270, 2276, 2315, 2321, 2311, 2381, 2504, 2490, 2497, 2605, 2727, 2789, 2589, 3076, 1588, 1667, 2535, 2824, 2920, 3263, 3022, 1063, 1401, 2579, 2658, 2834, 2883, 2873, 2577, 2755, 2863, 2962, 3027, 3066, 3006, 3115, 3209, 3256, 1111, 1909, 2383, 3013, 1152, 1335, 2978, 35, 40, 17, 33, 34, 38, 43, 42, 21, 3205, 2520, 3127, 2949, 926, 3298, 2536, 2756, 3043, 2662, 2831, 2519, 2838, 3090, 2657, 2594, 3111, 1669, 2379, 2546, 3087, 2531, 2895, 2572, 2646, 2874]. pp [2876, 3061, 3060, 3003, 3012, 1908, 2808, 2655, 2550, 366, 2774, 2941, 2706, 57, 64, 219, 218, 173, 172, 220, 244, 547, 557, 542, 1715, 2888, 1917, 2943, 2751, 2907, 2719, 2827, 2819, 2948, 3530, 408, 752, 2934, 1995, 314, 2632, 2839, 2600, 2988, 2742, 3057, 2595, 2885, 3070, 2955, 2672, 3029, 613, 2946, 2461, 3108, 2794, 2893, 3289, 3288, 2633, 2518, 2470, 2101, 3131, 2614, 2654, 2800, 2829, 2922, 2905, 2740, 3225, 3281, 2502, 2712, 2757, 2642, 2288, 1656, 2690, 2630, 2651, 2999, 2904, 2787, 1395, 1575, 2771, 2947, 754, 2747, 1732, 1987, 1543, 2377]. pp [2557, 2973, 3253, 2308, 2783, 1997, 174, 3110, 80, 81, 67, 79, 225, 288, 286, 300, 2214, 2202, 2492, 2623, 2720, 2652, 2835, 2773, 2892, 2268, 2781, 3067, 2832, 2656, 2950, 2532, 2681, 2854, 2859, 2915, 2718, 2862, 2061, 2561, 3033, 829, 2709, 2596, 410, 606, 748, 1315, 1484, 1641, 2095, 2134, 2197, 2137, 2493, 2629, 2323, 3024, 3074, 422, 1148, 1841, 2494, 3017, 2533, 1918, 2970, 3091, 1724, 2900, 2736, 2977, 3243, 2912, 191, 204, 419, 624, 770, 758, 787, 683, 763, 927, 850, 930, 912, 848, 1407, 1878, 1875]. pp [1915, 1992, 1901, 2015, 2004, 2013, 1944, 2362, 2427, 2966, 2825, 407, 1190, 2590, 1640, 2945, 2277, 2419, 2598, 2754, 3014, 2938, 2965, 3258, 332, 360, 331, 434, 436, 437, 639, 638, 641, 566, 645, 706, 707, 867, 958, 947, 948, 952, 950, 877, 872, 882, 1101, 1115, 1096, 1099, 1092, 1030, 1098, 1177, 1173, 1431, 1436, 1675, 1716, 1672, 1677, 1754, 1676, 1812, 1809, 1936, 1934, 2011, 1933, 1935, 2071, 2152, 2154, 2220, 2265, 2275, 2413, 2430, 2373, 2365, 2375, 2468, 3069, 3018, 2659, 2287, 3031, 3546, 2713, 3529, 2940, 14, 2525, 2741, 433, 435, 644, 554, 712]. pp [713, 710, 697, 708, 711, 871, 863, 870, 876, 875, 1100, 1029, 1186, 1184, 1274, 1180, 1182, 2918, 2790, 2801, 2555, 2505, 330, 2743, 2775, 2884, 2944, 3426, 2987, 2684, 3267, 3251, 2821, 2961, 2207, 917, 2799, 3036, 36, 25, 26, 200, 622, 2615, 616, 920, 1001, 1718, 1843, 3277, 2554, 3208, 2660, 1849, 2693, 3287, 3257, 2880, 88, 184, 233, 293, 295, 277, 424, 518, 457, 430, 428, 502, 562, 540, 560, 655, 563, 581, 631, 589, 633, 556, 590, 575, 635, 559, 762, 705, 761, 700, 776, 780, 779]. pp [701, 702, 727, 774, 778, 844, 970, 858, 878, 893, 939, 835, 862, 916, 861, 963, 937, 864, 941, 953, 854, 859, 938, 1014, 1062, 1080, 1082, 1011, 1016, 1013, 1072, 1012, 1083, 1015, 1020, 1084, 1107, 992, 1056, 998, 1033, 1161, 1170, 1251, 1258, 1256, 1257, 1284, 1259, 1163, 1240, 1255, 1136, 1158, 1166, 1167, 1241, 1159, 1185, 1195, 1193, 1160, 1165, 1172, 1453, 1369, 1429, 1430, 1456, 1355, 1354, 1372, 1415, 1451, 1452, 1411, 1422, 1343, 1424, 1461, 1421, 1435, 1350, 1373, 1437, 1342, 1419, 1320, 1423, 1584, 1585, 1609, 1525]. pp [1599, 1512, 1511, 1544, 1510, 1591, 1532, 1614, 1570, 1563, 1743, 1744, 1690, 1665, 1663, 1730, 1766, 1768, 1662, 1666, 1745, 1746, 1749, 1695, 1877, 1806, 1794, 1874, 1813, 1787, 1861, 1863, 1873, 1792, 1876, 1807, 1860, 1871, 1872, 2002, 1930, 1919, 1931, 2001, 1923, 2009, 1926, 2069, 2112, 2115, 2077, 2078, 2110, 2113, 2118, 2068, 2098, 2114, 2111, 2148, 2151, 2226, 2213, 2153, 2212, 2150, 2149, 2195, 2221, 2225, 2271, 2317, 2310, 2319, 2415, 2414, 2369, 2368, 2423, 2363, 2541, 2548, 2574, 558, 1065, 2877, 338, 922, 1764, 2624, 2529, 2732, 2785, 417]. pp [2855, 3093, 2792, 1712, 535, 610, 534, 2858, 2282, 2686, 2677, 1074, 3088, 1848, 1904, 2206, 3038, 2584, 1645, 2716, 2809, 3224, 2280, 2745, 2753, 3078, 2976, 833, 2204, 2886, 2617, 2731, 3114, 2833, 3008, 2060, 2676, 2844, 2975, 3118, 304, 680, 756, 791, 755, 684, 1414, 2722, 2583, 2552, 2749, 2725, 2841, 2952, 2875, 3010, 1141, 2576, 2537, 165, 276, 753, 749, 800, 678, 997, 1229, 1285, 1392, 2645, 2840, 3083, 2767, 1618, 3125, 2643, 2678, 3525, 2663, 308, 1230, 1647, 2734, 186, 203, 345, 347, 1075, 1394, 2231, 2272, 2366, 2426, 2382, 2412, 2506, 2618]. pp [2559, 2738, 2673, 2867, 2866, 2983, 2985, 3039, 3085, 3130, 3226, 1413, 1579, 1639, 1658, 1851, 1922, 1899, 1998, 1994, 2107, 2062, 2132, 2428, 2429, 2380, 2469, 2499, 3034, 2807, 611, 44, 2830, 3241, 2803, 2984, 2316, 2384, 2374, 2608, 2539, 2729, 1400, 1723, 2585, 361, 362, 2939, 2951, 2679, 2597, 3096, 2728, 3062, 411, 1140, 2786, 2285, 1847, 2998, 13, 2457, 1153, 2286, 2804, 2609, 3044, 3019, 2784, 2845, 3206, 2626, 3058, 2278, 2489, 3126, 1840, 3528, 1326, 1734, 924, 851, 1154, 1334, 1541, 2273, 2378, 2507, 3547, 2372, 2593, 348, 808, 806, 803]. pp [860, 892, 849, 913, 1007, 1086, 1113, 993, 1076, 1079, 1109, 1068, 1089, 1203, 1244, 1202, 1144, 1329, 1455, 1330, 1491, 1713, 1650, 1767, 1654, 1796, 1988, 1912, 2108, 2466, 2465, 2806, 2919, 2967, 3032, 2198, 2603, 2687, 12, 2591, 1733, 1844, 1907, 2982, 3086, 3011, 3203, 3075, 582, 555, 2856, 1720, 2861, 2911, 1999, 2890, 2921, 2914, 2544, 3261, 1993, 2269, 2903, 3095, 378, 377, 381, 380, 379, 382, 2902, 1837, 1836, 2136, 2458, 2460, 2549, 2625, 2650, 2820, 2937, 3250, 2935, 2587, 2501, 3092, 2582, 3004, 3278, 2901, 2281, 2581, 2691, 840, 1986, 3527, 2846, 1064]. pp [1232, 1233, 1291, 1386, 1450, 1371, 1388, 1318, 1420, 1615, 1486, 1534, 1483, 1565, 1642, 2100, 2205, 2201, 2324, 2433, 2422, 2498, 2503, 2752, 2796, 2810, 2954, 2956, 1267, 1270, 1252, 1269, 1324, 1352, 1331, 1519, 1530, 1485, 1644, 607, 608, 746, 745, 747, 914, 945, 1224, 1133, 1174, 1134, 1225, 2714, 3037, 2456, 1990, 2622, 1906, 1900, 2613, 2545, 2674, 2739, 2910, 281, 489, 3239, 1069, 2604, 2534, 2724, 2779, 1728, 1800, 2556, 3259, 2644, 2628, 1845, 609, 1850, 2322, 2367, 3073, 1000, 1150, 1290, 1336, 1577, 1798, 1814, 2133, 2266, 2627, 690, 346, 2772, 2562]. pp [2848, 2802, 2857, 1905, 2913, 2688, 2971, 2748, 2726, 2960, 2782, 2968, 2668, 2578, 2692, 3026, 3089, 3264, 3285, 1135, 1389, 1390, 1482, 1719, 1839, 1834, 2054, 2416, 2417, 2620]. pp. [11, 16, 89, 41, 77, 60, 62, 167, 259, 253, 202, 227, 164, 258, 195, 242, 168, 189, 283, 301, 335, 342, 507, 454, 439, 414, 438, 405, 580, 629, 587, 592, 573, 544, 615, 642, 651, 579, 721, 802, 801, 759, 679, 724, 805, 795, 722, 837, 852, 896, 972, 918, 923, 965, 959, 968, 888, 842, 1002, 1106, 1087, 1078, 1081, 1019, 1023, 1199, 1231, 1139, 1169, 1194, 1146, 1204, 1367, 1340, 1313, 1346, 1362, 1441, 1370, 1449, 1404, 1616, 1535, 1617, 1537, 1490, 1562, 1605, 1680, 1649, 1660, 1692, 1805, 772, 1227, 1393, 315, 897, 1239, 339, 552, 653]. pp. [807, 934, 59, 1412, 1576, 1659, 1801, 1902, 228, 222, 688, 1418, 1793, 2053, 2199, 508, 456, 495, 497, 623, 621, 533, 548, 568, 543, 569, 584, 591, 652, 585, 532, 796, 677, 1066, 1405, 1368, 1611, 1567, 1568, 1648, 1679, 1816, 619, 1236, 1403, 1496, 1842, 1226, 1613, 1581, 1580, 1574, 1769, 1693, 1815, 1996, 2064, 2063, 2116, 2074, 2073, 2155, 1790, 166, 287, 278, 234, 260, 855, 843, 856, 549, 181, 571, 873, 1444, 1448, 1684, 1984, 1985, 274, 279, 955, 949, 1059, 170, 257, 572, 809, 971, 1088, 498, 1726, 1500, 349]. pp.

[538, 551, 1143, 1067, 1138, 1263, 1516, 1499, 1603, 1759, 2156, 1588, 614, 1721, 1789, 1428, 1538, 334, 1248, 45, 56, 284, 413, 2058, 1149, 1164, 365, 275, 682, 385, 368, 2101, 1242, 1505, 343, 1260, 1492, 341, 760, 769, 767, 925, 919, 933, 1246, 1409, 1398, 225, 504, 2140, 2139, 536, 928, 845, 1196, 1664, 1982, 2094, 1714, 217, 65, 285, 418, 654, 866, 1741, 1943, 874, 880, 881, 1183, 503, 505, 586, 640, 709, 786, 789, 785, 729, 954, 956, 868, 869, 1027, 1026, 1176, 1178, 1275, 1287, 1181, 1175, 1268, 1192]. pp. [1271, 1179, 1438, 1434, 1432, 1595, 1598, 1596, 1597, 1594, 1522, 1674, 1678, 1673, 1752, 1755, 1753, 1756, 1867, 1865, 1812, 1864, 1866, 1810, 1925, 2006, 1927, 2072, 2076, 1008, 499, 432, 500, 564, 810, 879, 1093, 1433, 1523, 1521, 1652, 1057, 390, 389, 344, 201, 256, 190, 183, 1001, 1489, 71, 20, 240, 187, 185, 238, 237, 199, 356, 354, 291, 355, 352, 431, 452, 453, 406, 427, 429, 409, 426, 425, 455, 561, 656, 704, 730, 832, 1162, 1748, 1750, 1671, 2005, 2070, 2117, 2217, 2307, 337, 929, 545, 626, 1504, 1497, 1321, 1399]. pp. [1410, 1610, 1319, 494, 618, 216, 1114, 255, 450, 451, 570, 550, 771, 757, 969, 1052, 1459, 1788, 1855, 1228, 1396, 1397, 1249, 1323, 578, 577, 838, 1061, 1646, 229, 163, 1073, 1237, 1289, 1458, 1460, 1506, 1694, 1737, 1653, 1729, 1786, 1799, 1856, 1980, 2052, 2104, 625, 921, 364, 363, 1503, 1501, 1247, 1542, 2099, 230, 765, 766, 289, 689, 692, 693, 723, 687, 691, 957, 942, 1034, 1198, 1328, 1349, 1425, 1426, 1348, 1571, 2208, 329, 750, 387, 388, 384, 371, 373, 372, 370, 412, 280, 539, 1337, 2057, 376, 2138, 1003, 1142, 1363]. pp. [1592, 1137, 1325, 1540, 946, 1223, 2096, 2147, 1802, 2141, 1920, 1498, 1493, 1651, 1332, 2097, 627, 831, 1870, 175, 178, 565, 546, 768, 492, 68, 695, 1983, 2135]. Pra [2107]. Practical [3557]. Practice [1617, 2075, 2496, 3526, 3023, 3204, 2755, 3503, 1504, 2804, 3504, 3800, 1936, 1400]. Practices [1725, 1097, 3081, 3442, 2514, 2888, 3070, 3189, 3624, 2867, 974, 3032, 3841, 707, 2967]. Practitioners [3254, 3441, 1392, 1213, 3909, 1652]. Praecisio [3229]. praeclarum [3165]. praecox [288]. Praga [844]. Praktiken [3530]. Prati [1620, 1928]. Pratica [1380, 1769, 1056, 1461, 3314]. pratiche [2888, 2651]. praticque [2596]. pratique [707, 2199]. Prattico [1851]. Pre [759, 3108, 1888, 1020]. Pre- [759]. Pre-Modern [3108]. pre-unification [1020]. Pre-Unitaire [1888]. Precept [2993]. precipitazioni [1525]. Precision [1808, 1497, 762, 2310, 4013, 4031, 540]. precolombiana [434]. précolombienne [434]. precursors [3061]. pref [3277]. Preface [2652, 2696, 2318, 893, 3010, 2469, 625, 2578, 921, 1154, 1138]. Prefazione [2616, 576, 1438, 678, 2501, 718, 1185, 1399, 771, 1644]. Prehistoric [1334, 3887]. prehistory [1810]. Preis [1838]. Preiss [1113]. Preistoria [1821, 1810, 2153]. Preliminary [38, 659, 3049]. Premessa [1156, 752, 2659]. Premi [2092]. premier [260, 2013, 3932]. première [675]. Premodern [3338, 2875]. Premuda [1175]. prép [868]. Preparation [3214]. Present [2361, 3747, 1865, 1812, 1934, 2009, 3937, 1999, 441, 3061]. Present-Day [2361]. Presentate [2354]. Presentation [2896, 3736, 1647, 620]. Presentazione [2142, 223, 407, 969, 371, 2383, 288, 1850]. presentazioni [1252]. Presente [740, 1925, 1113, 2013]. Presented [1393, 2354, 1123, 63, 1563]. présentés [584, 866]. Presents [3428]. Presenze [3820, 3040]. preservare [2330, 2281]. Preservation [3717, 3639]. preserved [2176, 992, 1252]. Preserving [3713, 2330]. présntés [796]. Press [15, 77, 60, 167, 227, 164, 168, 283, 335, 342, 441, 520, 439, 440, 438, 580, 587, 592, 615, 801, 759, 679, 805, 685, 837, 918, 923, 959, 968, 888, 1002, 1005, 1106, 1087, 1108, 1199, 1231, 1288, 1234, 1139, 1194, 1261, 1333, 1340, 1313, 1362, 1457, 1406, 1449, 1404, 1535, 1536, 1617, 1490, 1562, 1587, 1683, 1763, 1649, 1725, 1660, 1818, 1846, 1805, 1910, 2016, 2120, 2075, 2102, 2056, 2143, 2312, 2424, 2463, 2464, 2560, 2715, 2776, 2793, 2887, 2953, 2418, 2459, 2516, 2879, 3000, 2717, 2607, 1852, 2889, 3113, 3543, 1736, 2606, 2730, 1412, 1576, 1573, 1659, 228, 2669]. Press [2737, 533, 532, 796, 703, 846, 1643, 2538, 2936, 3063, 2852, 2851, 3081, 1403, 1842, 3068, 3254, 2664, 1835, 166, 3072, 2805, 2964, 2093, 3025, 1738, 2959, 2850, 1991, 2105, 2689, 2106, 3204, 3084, 3268, 2551, 2671, 2791, 2909, 3035, 3030, 58, 194, 192, 249, 511, 445, 649, 990, 1357, 279, 1500, 3124, 349, 538, 551, 1143, 1566, 2103, 2371, 1758, 1670, 2066, 2589, 3076, 1588, 2824, 2920, 3263, 614, 1063, 1721, 2834, 2873, 3066, 3256, 1152, 17, 34, 56, 3043, 2838, 2657, 2874, 3060, 3003, 3012, 2655, 2774, 2706, 2827, 2948, 275, 2600, 2595]. Press [2885, 3070, 385, 613, 3108, 2794, 2518, 2470, 3131, 2800, 1656, 1492, 169, 760, 769, 925, 919, 933, 1246, 1409, 1395, 1575, 2947, 2747, 1398, 1722, 1987, 2377, 2783, 1997, 174, 3110, 286, 3265, 2139, 2835, 3067, 2656, 2950, 2681, 3700, 2718, 2862, 2061, 2561, 3033, 2709, 928, 1315, 1484, 1641, 1982, 2094, 2095, 2134, 2493, 2629, 3074, 422, 1841, 2970, 2900, 2977, 3243, 2912, 65, 191, 204, 285, 419, 624, 654, 770, 683, 850, 848, 1407, 1915, 2004, 2362, 2427, 2966, 2825, 1640, 3014, 2938, 1192, 1596, 2076, 2071, 2152, 3542]. Press [3529, 1100, 1184, 1274, 1433, 1652, 2505, 2775, 2944, 3267, 2961, 2207, 2799, 36, 25, 26, 190, 616, 1843, 2660, 184, 293, 406, 916, 2310, 545, 3093, 1504, 1497, 1655, 2792, 2858, 2686, 3038, 2584, 3078, 2976, 1319, 2731, 3008, 2676, 2844, 3118, 2552, 2841, 2875, 3083, 1249, 2678, 229, 163, 345, 347, 1458, 1694, 2382, 2506, 2559, 2867, 2866, 2985, 3034, 2807, 2830, 1400, 1723, 2585, 176, 1503, 1542, 2728, 2786, 1847, 2998, 2457, 2609, 2784, 2845, 2626, 2489, 3126, 3528, 766, 851, 1154, 1334, 1541, 2378, 689, 808, 806, 692, 693, 723]. Press [803, 860, 849, 913, 1007, 993, 1076, 1109, 1203, 1144, 1198, 1329, 1425, 1654, 2806, 3032, 2603, 2687, 2591, 1844, 1907, 3086, 582, 1720, 2861, 2911, 1999, 1337, 3261, 694, 381, 1837, 2587, 2582, 3278, 2901, 2691, 1986, 2846, 1592, 2752, 946, 945, 3037, 2096, 2674, 2739, 3239, 2534, 1800, 1332, 1845, 1577, 1798, 1814, 177, 768, 2772, 68, 1245, 2562, 2848, 2913, 2688, 1839, 2054, 2658, 3837, 927, 2790, 3210]. Presse [2095, 256, 2544, 1848]. Presses [491, 1578, 1727, 3280, 260, 33, 2519, 2550, 1543, 866, 1878, 1875, 2375, 920, 2060, 2663, 1073, 2231, 3075, 582, 1486, 1534]. Pressman [1997]. Presso [2248, 2176, 539, 1218, 4040, 3783, 1599, 2213]. Prestel [2908]. Prestige [223]. Prete [2309, 2726]. Preti [3257]. Preunitario [711, 1020]. Preußen [3073]. Preußens. [364]. prevedibilità [1619]. Previdenza [1099, 712]. previously [787]. Prévot [1187]. Prez [1681]. Prez-sur-Marne [1681]. prezzo [2424, 2007, 1992, 1901, 2004, 2415, 2423]. Price [2105, 406, 2274, 2187, 2470, 2188]. priest [3827, 1914]. Priestley [414, 1910, 1368, 3025, 2534, 2724, 2779, 977, 3979, 2724]. Prigogine [834, 1998]. Prima [2476, 2173, 734, 1852, 1613, 2116, 880, 1099, 1098, 712, 171, 3961, 3655, 1836, 3259]. primaire [1859, 4000]. Primary [2103, 4000]. primatologische [2544]. Prime [2128, 2960]. Primeira [2198]. Primi [2172, 1606, 421]. primitifs [3058]. Primo [1606, 1034, 456, 180, 809, 3029, 1435, 1748, 593, 1379]. primordiale [1620]. primórdios [2198]. prince [222, 212]. Princeton [168, 335, 1725, 1846, 2463, 228, 2538, 2852, 2851, 1842, 1835, 2105, 249, 1357, 34, 2061, 2977, 654, 3529, 2207, 1497, 1655, 2792, 1397, 2506, 1076, 381, 1986]. Principat [571, 798, 994]. Principe [2891, 2834, 2873, 527, 237, 3598, 3313]. principes [1232]. Principia [1534]. Principio [2082]. Principles [2462, 1550]. principum [1279]. Prins [2986, 3477]. Print [846, 3171, 3557, 3114, 2147]. Printed [1055, 3559, 1448, 3530]. Printer [49]. Printing [3534, 3342, 3556]. Printmaking [3458]. Prior [824]. Priority [1193]. Prisma [2062]. Prismi [386]. privata [2916, 2221, 3937]. Private [1853, 3160, 2781, 2977, 1510, 3891, 2002, 3937, 3974]. Privileges [1020]. privilegi [1020]. privo [2502]. Prix [2615]. Prize [2545, 3156]. Prizes [3471, 3463, 3472, 3469, 3470]. probabilit [1502]. probabilità [1148]. probabilités [2528]. Probability [2106, 2470, 3830, 4057, 1857]. Probevorlesung [1042]. Problem [30, 3933]. Problema [1831, 1917, 1754, 3933]. Problemata [3366, 343]. problèmatique [2469]. Probleme [1004, 188]. Problemi [181, 1373, 1788, 1579]. Problems [283, 974]. Proceedings [2169, 2240, 2481, 1440, 2933, 4052, 4053, 2161, 452, 2078, 2195, 753, 3092, 1137, 1270, 4040, 4045, 4059, 4066, 4067, 4077, 4087, 4089, 4058, 4071]. Process [2897]. processo [346]. Processus [1888]. Prochiantz [1414]. procreation [3448, 3448]. proctologia [1027]. Proctor [1505, 422]. Prodigy [176]. prodotti [1695]. Produced [3191, 2866, 43, 1469]. Producer [3705]. Production [3500, 3440, 2699, 3711, 914, 3811, 3839, 1486]. Products [3474]. produzione [3839, 575, 1834]. Produzioni [903]. professeur [3010]. profession [4049, 2617]. professional [2075, 3800, 1275]. professionalità [1373]. professione [4049, 2315, 1752]. professionellen [364]. professioni [2100]. Professions [2514]. professore [1728]. Professorenkarriere [2466]. Professors [3442]. profeti [1462]. profilassi [3367]. profile [3910]. Profiles [885]. profilo [2320, 3910]. profit [762]. profumo [1354]. profumo-farmaceutica [1354]. Progettare [1459]. Progetto [1696, 1546, 4044, 1923, 782]. Programma [1054]. programmes [1442, 1080]. Programmi [2092]. Progrès [1306, 620]. Progress [1847, 3921, 3975, 1270, 1800, 2681]. Progressive [4011, 3265]. Progresso [1737, 1270, 2042]. Proiettive [1785]. Project [2546, 1770, 4044, 1546]. Projective [206]. Projetté [1475]. Prolusione [2692]. Prolusioni [1431]. Prometheus [2922, 2987, 2935]. Promise [2828]. Promoting [1819, 2495, 3552, 919]. Promozione [1237]. Pronostici [3315, 4050]. Prontera [2572]. Proof [3225]. propagation [1681]. Properties [28]. Property [2412, 4048]. Prophecy [649, 1357]. Prophet [3166]. proporzione [1321, 2163, 2234]. Proporzioni [733]. Propos [368, 3368, 1386]. Proposent [1305]. Proposito [744, 1802]. Proposta [1559]. Proposte [910]. Proprietario [1295]. propriis [2716]. Prosa [1918]. prose [3884, 1918, 1915]. Prospects [1458]. Prosperi [2963, 1644]. Prospettiva [1485, 1785, 611, 2685, 2826, 2339]. Prospettive [1132, 910, 809, 2116, 181, 709, 1871]. Protagon [2115]. Protein [1274]. Protestant [2840]. protestants [582]. protettori [1956, 2027]. Protokolle [343]. Proton [1296]. Protopharmacology [1097]. Prototype [3505]. Prouse [1472]. Prova [1380, 2348]. Provence [3280, 1224]. Provenienti [1214, 1510]. Provenienza [668]. Proverbio [1749, 2113]. Providence [3491]. province [2002]. Provinces [2555]. Provincia [710, 1456, 1372, 1424, 1512, 1794, 1415]. Provincial [2068]. Provinciale [1016]. Provo [1196]. Prowincjonalna [1256]. Prucha [441]. Pruell [1936]. Pryds [1277]. Ps. [1483]. Pseudo [1108, 997, 3946, 3671, 3277, 3650]. Pseudo-Democrito [3277]. Pseudo-Galenica. [3650]. Pseudo-Geber [997, 3946]. Pseudo-Paracelsus. [3671]. Pseudo-Science [1108]. pseudosphere [3768, 2215, 3768, 2215]. psichiatri [3181]. psichiatria [881, 710]. Psichiatric [1184]. psichiatrici [710]. Psichology [1246]. psicoanalisi [1769, 1996, 2063, 2073]. psicoanalitica [1693]. Psicologia [1815, 1996, 245, 369, 4040, 1574, 2116, 2074, 2155, 2320, 3799, 1917, 3910, 1931, 3933, 1459, 1075, 1394, 4007, 3285, 1815, 2155, 2471, 2224]. Psicoterapia [2064, 2431, 1693]. Psicotropismo [705, 878]. psychiatrist [3927]. Psychiatrists [767]. Psychiatry [767, 225]. Psychological [3680, 769]. Psychologie [380, 2856]. Psychologists [3800, 2075]. Psychology [699, 286, 2404, 3413, 4040, 3753, 3754, 3799, 933, 1575, 3933, 1337, 4007, 3910, 2069, 2552]. Psychopathologist [3704]. Psychosurgery [1997, 3962]. Ptolemaic [1398]. Ptolemy [1468, 2813, 2812]. pu [1041, 44]. Pubblica [2476, 1136, 1424, 2916, 810, 879, 2221, 3937]. Pubblications [1147, 2379]. Pubblicazioni [1206, 250, 873, 3120]. Pubblichiamo [814, 522]. Pubblico [1637, 1325, 2672]. Public [923, 1853, 2700, 2600, 2781, 2977, 1655, 1458, 3867, 3891, 1849, 3937, 1694, 3974, 1109, 2901, 3564]. Public/Private [1853]. pública [3867, 1849]. Publicacions [2099]. Publication [1442, 38, 845, 540, 1090, 3657, 2828]. Publication/Watson [2828]. Publications [15, 490, 672, 722, 842, 1264, 1204, 1365, 1454, 1657, 2160, 2267, 2495, 3041, 3471, 495, 1097, 1898, 1911, 2891, 2972, 512, 445, 1447, 2553, 2917, 1494, 2059, 2594, 2907, 2839, 613, 225, 868, 3036, 427, 1224, 2456, 1920, 3877, 3877]. Publications/U.S.A [490]. Publications/USA [2160, 3041, 1898, 1911, 2553]. Publications/Watson [2839]. publico [2461]. Publié [1795]. Publiée [368]. publiés [32]. Publique [1040]. Publish [1208]. Published [1898, 10, 420, 1944, 4090]. Publisher [652, 3142]. Publishers [1589, 2194, 2462, 591, 235, 1401, 606, 1944, 424, 1570, 1600, 2833, 2583, 2939, 348, 1571, 1857]. Publishing [16, 405, 2899, 28, 1494, 1791, 2216, 2579, 2594, 2839, 624, 1179, 1320, 2098, 2745, 1786, 3058, 3011, 1792]. Puccini [2206]. pudore [852, 730]. PUF [2060]. Puglia [2921]. Pugliese [1598]. Puig [352]. Pulchinotta [503]. Pumfrey [2777]. Pump [168]. Punishment [651]. Punta [1279]. Punti [1821]. pupils [1652]. purae [4028]. Purdue [1288]. Pure [3191, 4028]. Purification [3722]. Purposes [2751]. Pursuit [1655, 3567]. Putting [3583, 2562]. Pyenson [2423]. Pygmalion [2533]. pyrheliometric [3423]. pyrhéliométrique [3423]. Pyrheliometry [2304].

Q [1343]. qarn [2199]. quad [1516]. Quaderni [1902, 1811, 2685, 309, 1516, 2315, 2955, 575, 1018, 1025, 1165, 1350, 2226, 999, 1712, 4016, 3798, 2272, 2010, 2608]. Quaderno [1619, 1455]. Quadrans [3485]. Quadrant [1045]. Quadratura [318]. quadrimestrale [1372]. Quaestio [717, 319]. Quaestionen [1941]. Quaestiones [1359]. Quakers [2861]. Quale [2391]. qualit [1672]. qualità [854]. Qualities [60]. qualunque [1995, 4021]. Quando [2263]. quante [2632]. quanti [1801]. Quantifying [685, 2304]. quantistica [18, 3911, 3481]. Quantitative [1846, 3462, 3881]. quantities [595, 2099]. Quantum [441, 1736, 1659, 2106, 3263, 3830, 2970, 3900, 3901, 3243, 2296, 3038, 1247, 3911, 1542, 2105, 2106]. Quarenghi [2415]. quarks [4054, 4055, 860]. Quarta [1765]. Quarterly [1466]. Quarto [252]. Quasar [956, 812]. quasars [812]. Quatre [512, 3742, 32, 3665]. Quattro [425, 2612, 4059]. Quattrocento [1360]. Quattroemme [1840, 1267]. QuattroVenti [2321, 581, 858]. quatuor [59]. Que [1306]. Québec [2423, 608]. Queen [891]. Quekett [3208]. Quel [3693]. Quellen [1079, 1905]. querelle [3890, 3980]. querelleau [3013]. Quest [549, 3524, 3703, 3529, 2802, 3753, 3754, 1575]. question [319]. Questione [660, 2487, 852, 1574, 730, 1136]. Questioni [2520, 883]. Questions [1578, 2388, 2296, 836]. Qui [1305, 814, 522]. quibus [1687]. Quiet [1333]. quimeras [35]. química [2198]. quintae [567]. Quintino [42]. Quirico [1591]. Quodlibet [715, 719]. Quotidiani [1552]. quotidiano [1598]. quotidienne [496]. Qurra [216].

R [283, 490, 722, 961, 888, 842, 1055, 1019, 1204, 1370, 2003, 2160, 2055, 2737, 532, 63, 549, 58, 442, 1103, 1359, 279, 3124, 3637, 1717, 2957, 2658, 2834, 2594, 3658, 3409, 2839, 2595, 343, 767, 2139, 2854, 2718, 3734, 848, 2004, 332, 434, 436, 586, 638, 1275, 1716, 1812, 1864, 3529, 2525, 2701, 452, 854, 1014, 1161, 771, 753, 997, 1392, 1618, 1506, 1733, 380, 2501, 2901, 1324, 747, 2200]. R. [1370, 2573, 1393, 1067, 2066, 343, 1810, 1342, 534, 2583, 212]. R.N [1408]. Raccolta [915, 996, 3001, 2129]. Raccolte [2191, 1077]. Raccolto [1955, 2026]. Race [2851, 3574, 3658, 3878, 3619, 3377, 2131, 2366, 1837, 177, 768, 3648, 3758, 3991, 3716]. races [2663]. Rachel [2016, 538]. Racial [422, 3488, 54]. racines [2648]. Racism [2366, 3758, 3991]. Raction [2052].

Radiation [1247]. radiazioni [1801]. Radical [3337, 2724]. Radicali [3103]. radici [435, 371]. Radio [3860, 1733, 593, 457]. radiologo [1437]. Radium [3310]. Raf [3639]. Rafael [2068]. Raffaele [2370, 585, 2667, 2727]. Raffaella [3041, 752, 1741, 2741, 3125, 2425, 1876, 2225, 2274]. Raffaelli [2014, 1585]. Raffaello [858, 1874, 2324]. Raffigurazioni [1591]. Rafinesque [763]. Ragazzini [1252]. Ragazzo [2910]. Raggi [1437]. Ragione [2488, 715, 2797, 1506, 1233]. ragioni [2501]. Ragland [3734]. Ragozzino [75, 854]. Ragusa [2389, 13]. railway [2332]. Raimondi [2492]. Raimondo [2497]. Rainer [895, 248, 542]. Raison [3234, 886]. raisonné [2369, 4091]. Ralf [2598]. Ralph [2976]. RaM [3092]. Ramandi [3570]. Ramanujan [2503]. Rambaldi [1730]. Ramelli [2644]. Ramn [249]. Ramon [3019, 2907, 2101, 3091, 3415, 3327]. Ramsden [3154]. Rand [2802]. Randazzo [1748]. Randi [250, 718]. Rane [3955, 2904]. Ranelagh [3615]. Ranieri [1331]. Raphael [1874, 3507, 2663]. Rapids [624]. Rapport [2678]. Rapporti [817, 658, 3976, 2421, 448]. Rapporto [1221, 80, 2384]. Rapports [2048, 2566, 620]. Rappresentare [2691]. rappresentazione [3873, 384]. Rappresentazioni [2434, 1120]. Rare [1078, 1805, 696, 1510]. Rariora [2530]. Rarit [2883]. Rarities [2883]. Raschi [3742]. Rasera [2619]. Rasetti [2910]. rasforskning [2675]. Rashed [2573, 2710, 2199, 3568]. Rashygien [972]. Rasini [2276, 2502]. Rasmussen [1848]. Raspadori [873]. Rassenkunde [1004]. Rastatt [1171]. Ratcliff [2888, 2856]. Rather [2870]. Rathschüler [354]. Ratio [1388]. Rational [1535, 3006]. Rationalité [1138]. rationality [3848]. Ratzel [1916]. Rauch [520]. Raum [361]. Raven [36, 26, 545]. Ravenna [2002]. Ravenni [1186]. raxionalismo [2429]. Ray [2435, 235, 3361]. Raymond [227, 610, 1154]. rayonnement [2878]. razionalit [1850]. razionalità [3848]. razza [2677]. Razzismo [1978]. rben [2576]. RCS [867]. Re [1317, 2707, 3244, 2770, 1756, 3364]. Re-entangling [3244]. Reach [3513]. Reactions [3433]. Read [3646]. Reader [439, 4070, 1980, 2869, 3758, 3837, 3863, 927, 3991, 2194, 3535, 2366]. Readers [3783]. Reading [1499, 1628, 3338, 3387, 2640, 2737, 1541, 3507]. readings [606]. Real [2808, 1645, 2265, 2697]. Realdo [3364]. Reale [2386, 429, 1140]. Reali [79]. Realism [3384, 3391]. réalisme [3087]. réalité [625]. Réalités [3395]. realizzata [638]. Realizzazione [165]. realizzazioni [562]. Realm [2262]. Realtà [2502, 934, 559, 372]. Reappraisal [1708]. Reappraisals [801]. Reardon [928]. Reardon-Anderson [928]. Rearranging [2917]. Reason [1110]. Reassessment [3188, 3410]. Rebecca [3399, 2783]. Rebekah [2781]. Rebentrost [2857]. reborn [2646]. Rebuilding [1219]. Recami [2415]. recensione [522]. Recentemente [1470]. récentes [1131]. Reception [3559, 3502, 2699, 1825, 1701, 3837, 927]. Recezione [1701, 1691]. Recherche [540, 1700, 3540]. Recherches [1305, 3245, 1224, 1131]. Rechnen [751]. Recht [1405]. Recipe [3338]. Recipes [3934, 3493]. Recollections [4041, 2216, 1592]. Reconcilement [2761]. Reconciling [3232]. Reconfiguring [3175]. reconsidered [1139]. Reconstructing [2193]. Reconstruction [2484, 3629, 1694, 388].

Reconstructions [3996, 3289]. Record [3292]. Recordati [1676]. Recording [3338]. Records [3641, 3659, 3677, 3697, 3720, 3743, 1933]. Recounting [3517]. Recreating [3289, 3996]. Rectilignes [322]. Recuperi [1003]. Recupero [48]. Recycling [3587]. red [2119]. Redakcja [1281]. Redhead [1425]. Redi [1953, 2024, 2773, 2112, 1782, 2726]. Rediscovered [1896]. Rediscovering [3484]. rediscovery [421]. Rediviva [1218]. Redondi [1967, 2038, 2332, 287, 2283, 2743]. Redpath [691]. réédition [2528]. Reeds [2819]. Reeves [1986]. Referate [343]. référence [2550]. reflecting [1014]. Reflection [890, 3134, 1199, 1381]. Reflections [2568, 2629]. reflexes [1564]. Reflexion [1381]. Reform [2755, 3079, 1800, 918, 3921]. Reformation [1087, 3017, 3506, 3337, 3017, 2840, 3214]. reformatus [1321]. réforme [2214]. Reforming [2858]. Refractive [2346, 2347]. Refractometer [2486]. Refugee [2231]. Refusés [3473]. Regame [2538]. regard [3448]. Regarding [3160]. Regency [342]. Regensburg [2915, 4078]. Reggio [2368, 3039]. Regi [3177]. Regia [1868]. Regime [614, 3459, 2789, 3197]. Regiment [3244]. Regimine [1279]. Regio [2126]. Regiomontanus [1123, 751]. Regiomontanus-Bianchini [751]. region [778]. Regional [540]. regionale [1267, 1252, 1324]. Regione [1648, 778, 2148, 2002, 20]. Régis [216, 3135, 333]. Registro [1379]. Regius [3508]. Regno [2665, 79, 2154, 3367]. Regny [1851]. Regogliosi [1754]. Regolamento [3177]. regolare [2461]. regole [2836, 2111]. regoli [2111]. regolo [2111]. regressu [59]. regulating [3244]. Rehn [307]. Reich [3037, 1410]. Reichsanstalt [1340]. Reid [450]. Reidel [16, 28]. Reig [1849]. Reihe [1001, 746, 747]. Reill [2784]. Reimers [173]. Reinhard [1854]. Reinhardt [2917]. Reinheit [379]. Reinhold [1019, 1019]. Reinzer [93]. Reis [1743, 1744]. Reitzel [235]. Related [2879, 1358, 3719, 1933]. Relatifs [2048, 2566]. Relations [3573, 1481, 1735, 3893, 1912]. Relationship [1110, 3137, 1043, 1936, 1042, 1989]. Relationships [817]. Relative [1822, 1476]. Relativement [1476]. Relativistic [3929]. relativistica [3929, 1323, 1994]. Relativit [2082]. relatività [1801, 2632]. Relativitätstheorie [1977]. Relativity [3038, 3126, 2851, 1977, 2763]. relativo [2632]. relazionale [2324]. Relevance [2361, 490]. Relics [3214]. Relief [2962]. religieuses [707]. Religion [164, 925, 1395, 3074, 1878, 3267, 2207, 2799, 1225, 3765, 760, 3889, 3925]. religione [688, 1314, 3925, 655, 1713]. Religions [2201, 4043]. Religionsphilosophische [217]. religious [3906, 707, 993, 2861]. remède [3611]. Remedies [2998]. remedios [3959, 3264]. Remembered [2669]. Reminisciences [1734]. Remises [1476]. Remmert [2837]. Remondini [1794]. Remus [131]. Ren [496]. Renacimiento [2808]. Renaissance [491, 580, 583, 651, 579, 1106, 3477, 2889, 1966, 698, 166, 3526, 194, 964, 990, 1191, 1277, 1445, 279, 1670, 34, 3012, 2988, 2214, 2832, 2950, 2598, 1192, 3069, 3018, 2801, 2775, 2884, 3923, 1058, 3940, 3241, 2489, 3528, 1720, 381, 3567, 4004, 3278, 1486, 3240, 3112, 3756, 3467, 2442, 2882, 3007, 3171, 2649, 3801, 3173, 2838, 2823, 4064, 2999, 2521, 2768, 3731, 3918, 4077, 1058, 2584, 3252, 3202, 1646, 2711, 3972, 2890, 2575, 3337, 4003, 3256, 3458, 3633, 1228, 3650, 3670]. Renaissanceforschung [1191]. Renate [2312]. Renatico [1241]. Renato [2116, 1706, 1438, 2149, 1796, 619, 1928, 455, 3226]. rencontres [3767, 3330]. Rendiconti [1091]. René [842, 2539, 1670, 1364]. Renée [2662, 3507]. Renewal [1778, 1121, 2396]. Renneberg [1735]. Rennes [1612, 2550, 175]. Renouveau [1121, 2396]. Renzi [2147]. Renzo [180, 700, 776, 1257, 1807, 421]. Reperta [3670]. Repertoire [180, 3624, 1444]. Repertorio [180, 2520, 2468, 2045]. Repertorium [2529]. replica [814, 522]. répliques [2578]. Report [824, 670, 1618, 3474, 908, 2050, 1636]. Reports [1225, 3976]. Represent [3700]. Representation [3430, 1880, 2970, 3153, 3873, 384, 3004]. Representational [2867]. representations [707, 707, 3010, 3395]. Representing [3189, 3542]. Reprint [2957, 226, 420, 1731, 1857]. Reprinted [12]. Reprints [301, 674]. Reproduction [2582]. Repubblica [720, 1435]. Republic [1933, 541]. Republicain [1727]. Republiek [541]. République [516, 386]. reputation [2757]. Reputations [3030]. rerum [2337, 310, 2288, 567, 86]. rerum-deutsch [567]. Resarch [54]. résaux [2550]. rescue [1781]. Research [2435, 3282, 3469, 2850, 3205, 3487, 2305, 1770, 1934, 1933, 3601, 3226, 3986, 3987, 3988, 3772, 1305, 4061, 769, 3874, 778, 3540, 1270, 492, 1546, 3092]. Researcher [2185, 3981]. réseau [2652]. residences [429]. residenze [429]. Resistance [1758, 4046]. Resistenza [2474, 1678]. resor [1019]. Resort [1997, 3962]. Resources [1127]. Respighi [1671]. Respiration [3200]. responsable [855]. Response [2990]. responses [2861]. Ressel [1237]. rest [3528]. resta [3693]. Restauration [3893, 1912]. restaurazione [2376, 1540]. Restauro [828, 1632, 1298]. resti [3693]. restituta [4027, 3064]. Restoration [342, 2169, 2240, 3176, 760, 1912, 1540, 3893]. restored [4027]. Restrictiones [28]. Resultate [4061, 1285]. résultats [2015]. results [2015, 4061]. Résumé [2952]. Resurrection [3372]. Retailers [843]. Retained [1783, 1933]. Retaining [3708]. Rete [3730, 3077, 2903]. Rethinking [2867, 3096]. Retorica [1053]. Retrying [2885]. réunies [2309]. réunis [1326]. Revealing [2881]. revelation [4033]. Revelations [1548]. Reversion [2404]. Reviel [2881]. Review [39, 11, 16, 15, 89, 41, 77, 60, 62, 61, 167, 259, 253, 202, 227, 164, 197, 258, 195, 242, 168, 189, 224, 283, 301, 302, 335, 296, 303, 342, 313, 299, 516, 507, 515, 517, 441, 520, 454, 491, 439, 440, 514, 414, 438, 405, 490, 537, 580, 629, 587, 592, 573, 544, 583, 615, 617, 642, 651, 567, 579, 721, 794, 802, 714, 728, 801, 759, 679, 777, 724, 805, 764, 795, 672, 685, 722, 804, 837, 852, 885, 894, 896, 972, 961, 889, 918, 923, 965, 959, 895, 891, 968, 888, 842, 1002, 1110, 1070, 1005, 1091, 1106, 995, 1055]. Review [1087, 1090, 1078, 1081, 1019, 1023, 1108, 1199, 1200, 1231, 1288, 1234, 1139, 1201, 1169, 1194, 1264, 1146, 1261, 1204, 1333, 1367, 1340, 1365, 1313, 1346, 1362, 1402, 1441, 1457, 1406, 1427, 1364, 1370, 1449, 1404, 1454, 1616, 1531, 1535, 1536, 1528, 1617, 1502, 1537, 1539, 1490, 1562, 1602, 1605, 1606, 1589, 1587, 1680, 1683, 1691, 1657, 1763, 1649, 1725, 1660, 1692, 1819, 1818, 1846, 1853, 1805, 1795, 1945, 2003, 1910, 2016, 2119, 2120, 2075, 2102, 2056, 2067, 2194, 2227, 2143, 2145, 2228, 2159, 2200, 2160, 2267, 2312, 2318, 2421, 2420, 2370, 2424, 2488, 2463, 2467, 2496, 2495, 2500, 2464, 2560, 2586, 2601]. Review [2527, 2715, 2675, 2746, 2776, 2793, 2887, 2953, 2986, 2916, 3041, 3016, 3299, 3347, 3368, 3508, 3569, 3598, 3739, 2418, 2459, 2516, 2573, 2647, 2708, 2710, 2879, 2881, 3000, 2055, 3477, 2843, 772, 2717, 1227, 1393, 2607, 2899, 1852, 1921, 2878, 3454, 3509, 2889, 3065, 315, 593, 897, 1239, 1578, 1949, 1956, 1957, 1959, 1962, 2020, 2027, 2028, 2030, 2033, 1963, 2034, 1951, 2022, 1952, 1953, 1954, 2023, 2024, 2025, 1955, 1960, 1966, 2026, 2031, 2037, 1967, 2038, 1948, 1950, 1958, 2019, 2021, 2029, 1947, 1965, 2018, 2036, 2012, 1961, 2032]. Review [2164, 2235, 2163, 2168, 2234, 2239, 2233, 2162, 2169, 2240, 2165, 2236, 2167, 2238, 2166, 2237, 2329, 2331, 2337, 2328, 2332, 2326, 2335, 2330, 2333, 2334, 2336, 2338, 2339, 2327, 2434, 2387, 2388, 2389, 2435, 2436, 2386, 3113, 3395, 3425, 3689, 2142, 339, 496, 552, 653, 588, 807, 726, 934, 2942, 3543, 3499, 1736, 3688, 3328, 3570, 3735, 2425, 2462, 2606, 2730, 2897, 59, 1412, 1576, 1573, 1659, 1801, 1902, 3116, 3064, 3002, 3107, 3120, 3286, 3283, 228, 222, 340, 336, 282, 416, 493, 421, 675, 688, 931]. Review [1418, 1793, 1903, 2053, 2612, 2896, 3482, 2669, 3326, 3094, 2735, 2737, 3005, 2199, 18, 3631, 2777, 508, 456, 495, 497, 623, 621, 533, 548, 568, 543, 569, 584, 591, 652, 585, 620, 532, 612, 686, 796, 677, 676, 725, 703, 698, 673, 696, 699, 841, 884, 834, 846, 1006, 1097, 1095, 1060, 1066, 1155, 1197, 1171, 1145, 1314, 1316, 1322, 1440, 1317, 1439, 1405, 1368, 1387, 1611, 1488, 1567, 1487, 1564, 1569, 1568, 1518, 1495, 1533, 1608, 1648, 1661, 1681, 1727, 1679, 1643, 1682, 1765, 1689, 1816, 1804, 1811, 1817, 1898, 1911, 1981]. Review [2517, 2530, 2538, 2822, 2894, 2933, 2898, 2882, 2936, 3063, 3040, 3059, 3071, 3121, 3109, 3362, 3475, 3674, 2852, 2851, 3081, 221, 619, 1236, 1235, 1403, 1496, 1842, 2491, 2522, 2707, 2864, 2770, 2981, 2891, 3068, 3255, 3254, 3280, 3282, 3541, 643, 932, 1226, 1613, 1581, 1580, 1574, 1769, 1693, 1815, 1996, 2064, 2063, 2116, 2074, 2073, 2155, 2320, 2431, 2364, 2471, 1004, 3694, 2664, 2780, 1790, 2653, 1835, 166, 287, 278, 3072, 3736, 2805, 2964, 2580, 2203, 2472, 2144, 2543, 3080, 2093, 2750, 2682, 2284, 2528, 223, 2620, 2667, 3526, 3442, 3545, 3025]. Review [3653, 1738, 2836, 2837, 75, 63, 74, 234, 235, 260, 855, 847, 843, 856, 1085, 1112, 1808, 2524, 2631, 2721, 2797, 2959, 2906, 3009, 3510, 2850, 3023, 3316, 2865, 549, 1991, 10, 180, 181, 182, 3476, 2313, 2314, 2376, 2105, 2689, 2106, 2972, 2798, 2769, 2309, 3279, 2523, 2588, 3015, 3204, 3414, 3483, 2853, 2979, 2908, 3082, 3084, 3129, 3123, 3122, 3268, 3369, 2849, 3021, 3512, 2685, 2826, 2616, 2551, 2558, 2671, 2791, 2909, 3035, 2542, 2974, 2969, 3077, 3030, 2670, 28, 72, 29, 85, 30, 23, 31, 58, 86, 32, 194, 243, 252]. Review [250, 251, 193, 196, 198, 192, 248, 249, 188, 312, 311, 307, 310, 306, 305, 309, 443, 512, 442, 444, 446, 511, 447, 445, 449, 448, 571, 576, 648, 574, 649, 715, 798, 720, 717, 716, 718, 719, 966, 887, 964, 873, 886, 890, 836, 857, 830, 915, 994, 991, 1102, 1103, 1105, 1077, 1104, 996, 990, 1191, 1277, 1283, 1278, 1157, 1279, 1280, 1273, 1189, 1187, 1188, 1281, 1156, 1359, 1446, 1358, 1443, 1447, 1357, 1442, 1445, 1444, 1360, 1448, 1687, 1688, 1685, 1684, 1686, 1984, 1985, 1941, 3001, 70, 274, 279, 955, 949, 1059, 170, 257]. Review [519, 572, 792, 809, 951, 971, 1088, 1462, 1417, 2611, 3573, 498, 2283, 2279, 2648, 1726, 2963, 3602, 2958, 2553, 1151, 1500, 2917, 3124, 3544, 3637, 3672, 3495, 349, 538, 551, 1143, 1566, 1717, 2103, 2371, 2602, 2619, 2665, 2957, 2592, 1067, 1243, 1138, 1238, 1147, 1263, 1366, 1327, 1612, 1516, 1515, 1499, 1572, 1620, 1494, 1603, 1668, 1759, 1758, 1670, 1862, 1869, 1791, 1859, 1820, 1879, 2010, 2007, 2014, 1928, 2080, 2079, 2059, 2066, 2161, 2232, 2216, 2215, 2224, 2229, 2157, 2156, 2270, 2276, 2315, 2321, 2311, 2381, 2504, 2490, 2497, 2605, 2727, 2789, 3441, 3408, 3457]. Review [2589, 3076, 1588, 1667, 2535, 2824, 2920, 3263, 3022, 614, 1063, 1401, 1721, 1789, 2579, 2658, 2834, 2883, 2577, 1428, 1538, 334, 2755, 2863, 2962, 3027, 3066, 3006, 3115, 3209, 3256, 3572, 1111, 1909, 2383, 3397, 3013, 1152, 1335, 2978, 1248, 35, 40, 17, 33, 45, 34, 38, 43, 42, 56, 21, 3205, 3692, 2520, 3127, 2949, 284, 413, 926, 3298, 2536, 2756, 3043, 2662, 2831, 3440, 3630, 2519, 2838, 3461, 3090, 2657, 3714, 2594, 3111, 3435, 2058, 1669, 2379, 2546, 1149, 3087, 3611, 3717, 2531, 2895, 2572, 2646, 2874, 2876, 3061, 3060, 3003]. Review [3012, 3400, 3462, 1908, 2808, 2655, 1164, 2550, 365, 366, 2774, 2941, 2706, 3658, 57, 64, 219, 218, 173, 226, 172, 220, 244, 420, 547, 557, 542, 1715, 2888, 3415, 1917, 3365, 3675, 3418, 2943, 3594, 3312, 3366, 3491, 2751, 2907, 2719, 2827, 2819, 2948, 3344, 3345, 3409, 3530, 275, 408, 752, 2934, 3742, 1995, 314, 2632, 2839, 2600, 2988, 2742, 682, 3057, 2595, 2885, 3070, 2955, 385, 368, 2672, 3029, 613, 2946, 2461, 3346, 3486, 3511, 333, 3108, 2794, 3411, 2893, 3289, 3288, 2633, 2518, 2470, 2101, 3131, 2614, 2654, 2800, 2829, 2922, 2905, 3497]. Review [1242, 2740, 3225, 3281, 2502, 2712, 2757, 3399, 2642, 2288, 1735, 1505, 1656, 343, 2690, 1260, 2630, 2651, 2999, 3620, 3599, 1492, 2904, 2787, 169, 341, 760, 769, 767, 925, 919, 933, 1246, 1409, 1395, 1575, 3568, 2771, 2947, 754, 2747, 1732, 1398, 1722, 1987, 1543, 2377, 2557, 2973, 3639, 3253, 2308, 2783, 1997, 174, 2828, 3110, 80, 81, 67, 79, 225, 288, 286, 300, 504, 3265, 3331, 2140, 2214, 2202, 2139, 2492, 2623, 2720, 2652, 2835, 2773, 2892, 2268, 2781, 3067, 3079, 415, 2832, 3494, 2656, 2950, 3438, 3649, 3671, 2532, 2681]. Review [2854, 2859, 2915, 3656, 3412, 2718, 2862, 2061, 3670, 2561, 3033, 829, 2709, 2596, 410, 536, 606, 748, 928, 845, 1196, 1315, 1484, 1641, 1664, 1982, 2094, 2095, 2134, 2197, 2137, 2493, 2629, 3718, 2323, 3024, 3074, 422, 1148, 1841, 2494, 3719, 3738, 3017, 2533, 1918, 2970, 3091, 1724, 1714, 217, 2196, 3734, 3455, 2900, 681, 2736, 2977, 3243, 2912, 65, 191, 204, 285, 418, 419, 624, 654, 770, 758, 787, 683, 763, 927, 850, 930, 912, 848, 866, 1407, 1741, 1878, 1875, 1915, 1992, 1901, 2015, 2004, 2013, 1943, 1944]. Review [2362, 2427, 2966, 3458, 2825, 407, 1190, 2590, 3635, 1640, 874, 880, 883, 881, 1183, 2945, 2277, 2419, 2598, 2754, 3014, 3716, 2938, 3597, 2965, 3427, 3480, 3258, 332, 360, 331, 503, 434, 436, 437, 505, 586, 639, 638, 641, 566, 640, 645, 706, 709, 786, 789, 785, 729, 707, 788, 867, 954, 956, 868, 958, 947, 948, 952, 950, 877, 872, 869, 882, 1027, 1094, 1101, 1115, 1096, 1099, 1028, 1026, 1092, 1030, 1098, 1176, 1178, 1275, 1287, 1181, 1177, 1175, 1268, 1173, 1192, 1271, 1179, 1431, 1436, 1438, 1434, 1432, 1595, 1598, 1596, 1597, 1594, 1522, 1674, 1678]. Review [1673, 1675, 1716, 1752, 1755, 1672, 1753, 1756, 1677, 1754, 1676, 1867, 1865, 1812, 1809, 1864, 1866, 1810, 1936, 1925, 1934, 2006, 2011, 1933, 1927, 1935, 2072, 2076, 2071, 2152, 2154, 2220, 2222, 2265, 2275, 2413, 2430, 2373, 2365, 2375, 2468, 3496, 3069, 3018, 2659, 3327, 2287, 3031, 3402, 3481, 3498, 3546, 2713, 3668, 3542, 3439, 3529, 3695, 2940, 14, 2525, 1008, 2741, 499, 432, 433, 500, 435, 644, 564, 554, 712, 713, 810, 710, 697, 708, 711, 871, 863, 879, 870, 876, 875, 1093, 1100, 1029, 1186, 1184, 1274, 1180, 1182, 1433, 1523, 1521, 1652]. Review [839, 1057, 2918, 2450, 2790, 2801, 3636, 3601, 3693, 3713, 2555, 2505, 390, 389, 330, 1913, 2743, 3507, 2775, 2884, 2944, 2987, 2684, 3267, 3251, 3313, 2821, 2961, 3618, 2207, 344, 917, 2799, 3036, 3619, 3633, 36, 25, 26, 200, 201, 256, 171, 190, 183, 622, 2615, 3638, 616, 920, 1001, 1718, 1843, 3277, 2554, 2683, 3208, 2660, 1849, 1489, 2693, 3287, 3257, 3691, 2880, 88, 71, 20, 240, 187, 184, 185, 238, 237, 199, 233, 356, 354, 291, 355, 293, 295, 352, 277, 424, 431, 518, 452, 453, 457, 430, 406, 427]. Review [428, 502, 429, 409, 426, 425, 455, 561, 562, 540, 560, 655, 563, 581, 631, 589, 633, 556, 590, 575, 635, 656, 559, 704, 762, 730, 705, 761, 700, 776, 780, 779, 701, 702, 782, 727, 774, 778, 844, 970, 858, 878, 893, 939, 835, 832, 862, 916, 861, 963, 937, 864, 941, 953, 854, 859, 938, 1014, 1062, 1080, 1082, 1011, 1016, 1013, 1072, 1012, 1083, 1015, 1020, 1084, 1107, 992, 1056, 998, 1018, 1033, 1161, 1170, 1251, 1258, 1162, 1256, 1257, 1284, 1259, 1163, 1240, 1255, 1136, 1158, 1166, 1167, 1241]. Review [1159, 1185, 1195, 1193, 1160, 1165, 1172, 1453, 1369, 1429, 1430, 1456, 1355, 1354, 1372, 1415, 1451, 1452, 1411, 1422, 1343, 1424, 1461, 1421, 1435, 1350, 1373, 1437, 1342, 1419, 1320, 1423, 1584, 1585, 1609, 1525, 1599, 1512, 1511, 1544, 1510, 1591, 1532, 1614, 1570, 1619, 1563, 1743, 1744, 1690, 1665, 1663, 1730, 1748, 1766, 1768, 1750, 1662, 1666, 1745, 1746, 1749, 1671, 1695, 1877, 1806, 1794, 1874, 1813, 1787, 1861, 1863, 1873, 1792, 1876, 1807, 1860, 1871, 1872, 2002, 2005, 1930, 1919, 1931, 2001, 1923, 2009, 1926, 2069, 2070, 2112, 2115, 2117]. Review [2077, 2078, 2110, 2113, 2118, 2068, 2098, 2114, 2111, 2148, 2151, 2226, 2213, 2153, 2212, 2150, 2149, 2195, 2217, 2221, 2225, 2271, 2317, 2310, 2307, 2319, 2415, 2414, 2369, 2368, 2423, 2363, 2541, 2548, 2574, 558, 1065, 999, 2877, 337, 338, 922, 929, 1058, 1600, 1764, 2624, 2529, 2732, 2785, 417, 545, 626, 2855, 3093, 3492, 1504, 1497, 1655, 2792, 1712, 535, 610, 534, 2858, 3333, 2282, 2686, 3676, 2677, 1074, 1731, 1838, 3088, 3600, 1848, 1904, 2206, 1321, 1399, 3038, 2584, 1645, 2716, 2809, 3224, 3314, 3456, 2280, 66, 1071, 1410, 2745, 1610, 3485, 2753, 3210]. Review [3078, 2976, 833, 1319, 2204, 2886, 3712, 2617, 494, 618, 3690, 3715, 2731, 3114, 3614, 2833, 3008, 2060, 2676, 2844, 216, 2975, 3687, 3118, 812, 1114, 3652, 255, 304, 450, 451, 570, 550, 771, 680, 756, 757, 791, 755, 684, 969, 1052, 1414, 1459, 1788, 2722, 2583, 2552, 2749, 2725, 2841, 2952, 2875, 3010, 1141, 2576, 2537, 3284, 3443, 1855, 165, 276, 753, 749, 799, 800, 678, 997, 1228, 1229, 1285, 1396, 1392, 1397, 2645, 2840, 3083, 2767, 1265, 1249, 1618, 3125, 2643, 3417, 2678, 3525, 1323, 2663, 308, 578, 577, 838, 1061, 1230, 1647, 1646, 2734]. Review [229, 186, 163, 203, 345, 347, 1073, 1075, 1237, 1289, 1458, 1394, 1460, 1506, 1694, 1737, 1653, 1729, 1786, 1799, 1856, 1980, 2052, 2104, 2231, 2272, 2366, 2426, 2382, 2412, 2506, 2618, 2559, 2738, 2673, 2867, 2866, 2983, 2985, 3039, 3085, 3130, 3226, 1413, 1579, 1639, 1658, 1851, 1922, 1899, 1998, 1994, 2107, 2062, 2132, 2428, 2429, 2380, 2469, 2499, 3034, 2807, 611, 3657, 44, 2830, 3446, 3241, 83, 674, 3453, 2803, 2984, 3740, 2316, 2384, 2374, 2432, 2608, 2539, 2729, 625, 921, 1400, 1723, 2585, 1797, 364, 363, 361, 362, 2939, 2951, 2679]. Review [176, 2597, 1503, 1501, 3737, 3096, 1247, 1542, 2728, 3062, 411, 1140, 2099, 2786, 2285, 1847, 2998, 13, 2457, 1153, 3434, 2286, 2804, 2609, 3044, 3019, 2784, 3612, 3444, 3398, 3634, 3396, 2845, 3206, 2626, 3058, 230, 2278, 2489, 3126, 3564, 1840, 3410, 3460, 3493, 3528, 1326, 3733, 1734, 765, 766, 924, 851, 1154, 1334, 1408, 1541, 2273, 2378, 2507, 1854, 3547, 3741, 3506, 3332, 2372, 3710, 2593, 348, 289, 689, 808, 806, 692, 693, 811, 723, 803, 687, 691, 957, 860, 892, 849, 942, 913, 1007, 1086, 1113, 993, 1034, 1076, 1079, 1109, 1068, 1089]. Review [1203, 1244, 1202, 1144, 1198, 1329, 1328, 1455, 1349, 1425, 1330, 1426, 1348, 1491, 1571, 1713, 1650, 1767, 1654, 1803, 1796, 1988, 1912, 2108, 2466, 2465, 2806, 2919, 2967, 3032, 3330, 3334, 3401, 3429, 2208, 2198, 2603, 2687, 3616, 3617, 12, 2591, 751, 329, 750, 1733, 1844, 1907, 2274, 2982, 3086, 3011, 3203, 387, 388, 384, 371, 373, 372, 370, 412, 3075, 280, 582, 541, 555, 539, 2856, 1720, 2861, 3654, 2911, 1999, 2890, 2921, 2914, 1337, 2544, 3261, 3416, 1993, 2057, 2269, 694, 2903, 3095, 378, 377, 376, 381, 380, 379, 382, 2902, 1837, 1857, 3567, 3655]. Review [3711, 1836, 2136, 2458, 2460, 2549, 2625, 2650, 2820, 2937, 3250, 3325, 3315, 3669, 3673, 2935, 2587, 3632, 3478, 2501, 3092, 2582, 2138, 3004, 3278, 3445, 2901, 2281, 3447, 2581, 2691, 840, 1986, 3527, 3595, 3329, 3364, 3613, 2846, 1003, 1064, 1218, 1232, 1233, 1142, 1291, 1386, 1450, 1371, 1388, 1363, 1318, 1420, 1592, 1615, 1486, 1534, 1483, 1565, 1642, 1868, 2100, 2205, 2201, 2324, 2433, 2422, 2498, 2503, 2752, 2796, 2810, 2954, 2956, 1137, 1267, 1270, 1252, 1269, 1325, 1324, 1352, 1331, 1519, 1540, 1530, 1485, 1644, 607, 608, 746, 745, 747, 946, 914, 945, 1224]. Review [1133, 1174, 1134, 1223, 1225, 1391, 2714, 3413, 3361, 3037, 2096, 2147, 2456, 1990, 2622, 1802, 1906, 1900, 2141, 2613, 2545, 2674, 2739, 2910, 281, 489, 3239, 1069, 2604, 2534, 2724, 2779, 1920, 1728, 1800, 2556, 1498, 1493, 1651, 1989, 2146, 3259, 2644, 2628, 1332, 1845, 609, 3650, 1850, 2097, 2322, 2367, 3073, 627, 831, 1000, 1150, 1290, 1336, 1577, 1798, 1870, 1814, 2133, 2266, 2627, 690, 175, 177, 178, 346, 565, 546, 768, 2772, 492, 68, 695, 1245, 2562, 3459, 2848, 2802, 2857, 1905, 3565, 2913, 2688, 3363, 2971, 3428, 2748, 3479, 3596, 2726]. Review [2960, 2782, 2571, 3566, 2968, 2668, 2578, 2692, 3026, 3089, 3264, 3285, 3367, 3571, 1914, 1916, 1135, 1389, 1390, 1482, 1719, 1839, 1834, 1983, 2054, 2135, 2416, 2417]. Reviews [3141, 3249, 3426, 2982, 1466]. revised [1448]. Revisited [672, 3551, 671, 3380, 2176, 3236, 3588, 3422, 753].

Revival [815]. Revolt [3495]. revolución [35]. Revolution [488, 587, 801, 1169, 1333, 1313, 1587, 1660, 2425, 497, 846, 1495, 3228, 857, 1588, 760, 3893, 3691, 3936, 1458, 2264, 1723, 3474, 693, 892, 1912, 2640, 3983, 3984, 3445, 1381, 565, 2772, 1090, 2229, 3760, 35, 2646, 1237, 692, 4018, 2607, 2730, 977, 2051, 2664, 2805, 2837, 2824, 2756, 2948, 1722, 2718, 3888, 3893, 2975, 1464, 1247, 3126, 3982, 2591, 2051, 3936, 3569, 3328, 3456, 2309]. Revolutionär [363, 1450]. Revolutionary [441, 2538, 17, 1329, 4029]. revolutionized [1076]. Revolutions [2527, 2542, 2546, 2740, 2604, 490]. revue [2767, 3657]. Rewarded [3472]. Rewarding [3469, 3465]. Rewards [3464]. Reymond [201]. Reynolds [3635]. Rezeption [3186]. Rhades [142]. Rheinberger [2866, 3226]. Rhetoric [679, 3110, 2945, 1109, 1053, 1225]. rhétorique [3491]. Rhodes [576]. Rhys [2542, 2747, 2676]. Ribémont [1187]. Ribera [1855]. Ricasoli [1307]. Riccardo [2320]. Ricci [554, 1768, 1745, 1037]. Riccioli [2387, 3457, 999]. Ricco [453]. Ricerca [1380, 2092, 1304, 1472, 1880, 2476, 1545, 1619, 1876, 550, 1546, 1621, 2434, 1815, 2276, 778, 1456, 1373, 1871, 999, 1270, 2865, 181]. Ricerche [894, 2025, 1947, 2018, 456, 3772, 252, 448, 715, 719, 1820, 1943, 1369, 1372, 2002, 999, 2206, 1645, 2128, 449, 1291, 1615, 1954, 2274, 3986, 3987, 3988, 2960]. Ricerche/Antropologia [2206]. Ricette [3934, 2592]. Ricevute [1206]. Ricezione [1890]. Richard [164, 1617, 1562, 1683, 2899, 2737, 1495, 1791, 4057, 2799, 1857, 242, 617, 651, 2916, 2899, 3084, 192, 2577, 613, 2268, 2094, 2678, 3126, 370, 2810]. Richards [2791]. Richardson [1839]. Richers [148]. richesse [584]. Richet [3043]. Ricomposizione [1559]. Ricordi [1746]. Ricordo [2410, 825, 1465, 2358, 1129]. Ricossa [407]. Ricostruire [1219]. Ricupero [2567]. Rider [685, 182]. Ridolfi [576]. Riduano [740]. Rieder [3527]. Riedlberger [3298]. Riefolo [788]. Riemann [978]. Rietsch [2528]. riflessi [1564]. riflessione [1014]. Riflessioni [1833, 1424]. rifondazione [1693]. Riforma [88, 2394]. Riforme [804, 197]. riformismo [585]. Rigenerare [2616]. rigenerazioni [2903]. Righini [3040, 3820]. Right [2764]. Rigo [3088]. Rihll [2459]. riistu [938]. rilegate [1259]. rilegato [1915]. riminese [4076, 2014]. Rimini [2014, 4076, 1585, 2002, 4076]. Riminie [199]. rinacentista [965]. Rinascimentali [2461, 1190, 1873]. Rinascimento [1282, 1761, 2037, 612, 27, 84, 247, 309, 510, 647, 797, 962, 1187, 1356, 1684, 752, 2651, 2492, 1190, 1718, 425, 1284, 799, 800, 1527, 750, 375, 2810, 1644, 1906, 1832, 1966, 2434, 2491, 3801, 4064, 2222, 3918, 3426, 3967, 3972]. Rindi [1145]. Rinman [3160]. Rinnovamento [1778, 2620]. Rinzivillo [854]. Rio [430]. Ripoll [1984]. Riposte [1658]. Riproduzione [2605, 1258]. Risāla [1227]. riscoperta [421]. riscritta [2107]. Rise [2606, 3254, 2595, 3108, 3506, 3617, 539, 3295, 1261, 1670, 3824, 3909, 1109, 4003, 4004]. Rising [1047]. riso [2276]. Risorgimento [517, 688, 2477, 2513]. risorse [1525]. Risparmio [863, 917, 858, 1614]. Risposta [740, 212, 739]. Risse [169, 1935]. rist [1408, 2054]. ristampa [1197, 1496, 788, 1098, 2138]. Rita [223, 307, 2429, 309, 2787, 2880]. Rites [3216, 2067, 4039]. Rithmomachia [2489]. Ritmi [2548]. ritr [3280]. ritratti [2142, 1269]. Ritratto [2391]. Ritrovamento [2449]. ritrovati [778]. Ritrovato [1216, 2881, 771]. Ritter [926]. Riuniti [1200, 2504, 2323, 771, 578, 750]. riveduta [1790]. Rivers [3025, 229]. Rivieccio [1851]. Rivincita [2929]. Rivisitata [2352]. Rivisitato [2176]. Rivista [1516, 1871, 552]. rivoluzionario [586]. Rivoluzione [3299, 461, 1295, 2536, 2264, 195, 1691, 2425, 3760, 585, 619, 2376, 455, 558, 1237, 3982, 373, 1464]. Riyyt [2199]. Rizzi [2965, 1810]. Rizzoli [2881, 3022, 437, 505, 786, 867, 559, 942, 1650, 2503, 2133]. Road [3146, 3981]. Roark [3239]. Robert [515, 579, 801, 1231, 1139, 2516, 150, 1412, 3116, 855, 3414, 3369, 2791, 192, 1685, 2873, 3366, 754, 2977, 2884, 616, 1072, 1193, 3690, 1113, 2967, 1225, 1845, 414, 1910, 2143, 2267, 3615, 846, 2707, 549, 2059, 1776, 56, 2518, 1505, 2781, 3883, 422, 204, 1935, 2801, 3185, 293, 295, 766, 289, 913, 2967, 2503, 914, 2545, 2534, 609, 627]. Roberta [1419, 2221, 2855, 3039]. Roberto [1533, 576, 563, 1343, 2159, 3041, 456, 170, 1754, 425, 1614]. Roberts [1439, 3018, 3425]. Roberval [1232]. robespierristes [2229]. Robin [685, 3417, 182]. Robinet [516]. Robinson [1581, 2629, 2378]. Robison [3594]. Robotti [3080, 2271]. Roche [772, 846]. Rochelle [194]. Rocke [1333]. Roderick [3475]. Rodger [2971]. Rodolfi [2645]. Rodopi [63, 2906, 57, 1275, 2732, 3241]. Roeck [2962]. Roero [2905, 3497, 556]. Roger [2586, 81, 1249, 1326, 942, 1720, 1403, 2073, 2537, 3284, 2984, 2728, 694]. Roi [1040]. Roland [1804, 3288, 1413]. Rolando [2334]. Role [3464, 3503, 3554, 3355, 1002, 2237, 3818, 1270]. Rolfe [3412]. Rollin [103]. Rolof [153]. Roloff [153]. ROM [3286, 2683, 2363, 2205]. Roma [299, 1091, 1200, 1605, 2421, 2420, 2142, 569, 532, 1314, 1682, 1689, 1816, 3541, 932, 1613, 1581, 1580, 1769, 1996, 2064, 2073, 2320, 2431, 3545, 828, 2631, 3279, 2523, 2826, 2969, 72, 312, 311, 310, 306, 915, 991, 996, 1189, 3001, 949, 971, 3544, 1516, 2504, 2577, 2383, 1335, 3298, 2379, 682, 2672, 2712, 2651, 2557, 288, 504, 2214, 2202, 2720, 2323, 1741, 1943, 880, 332, 789, 707, 867, 956, 1099, 1098, 1176, 1754, 1925, 2011, 2275, 2713, 2741, 644, 712, 876, 1718, 355, 844, 862, 937, 859, 1016, 1033, 1158, 1166, 1167, 1354, 1451, 1343]. Roma [1525, 1619, 1690, 1665, 1666, 1876, 2110, 2151, 2319, 2363, 2624, 2206, 2809, 3614, 2833, 1114, 255, 771, 1052, 1414, 1265, 578, 1729, 2618, 1994, 2374, 2432, 2951, 3547, 1034, 1803, 750, 412, 555, 2269, 3095, 2501, 1388, 1868, 1223, 1651, 1870, 2627, 2968, 2668, 2692, 1135, 1390, 1482, 2416, 486, 3031, 4085, 2416, 1350]. Roma-La [1343]. Roma-Napoli [504, 1052, 1414]. Roma/Bari [1996, 2320, 2431, 971, 2383, 255, 2618, 1651]. Romae [3688]. Romagna [180, 1107, 2002, 2148, 2363, 2148, 1107]. romagnola [199]. romaine [332, 1135]. Roman [3068, 2874, 3003, 2999, 3438, 2938, 3157, 2054, 2114, 3810, 3271]. romana [1816, 2876, 332]. Romance [648]. romandes [2519]. romane [956]. Romani [720, 2348, 1214, 2011]. Romanicum [2522]. Romano [2452, 1279, 2214, 3069, 2623, 595, 94]. Romans [2596]. Romantic [1005, 3819, 3034, 2967, 1571, 2957]. Romantica [1833, 1866, 1614]. Romanticism [3701, 1570]. Romantik [226, 1086]. romantischer [3186]. Romanus [1558]. Romanzesca [740, 739]. romanzo [3039]. Rombai [804, 340, 688, 590, 1324]. Romby [1324]. Rome [4043, 2826, 2356, 314, 2595, 2214, 332, 868, 3069, 1429, 1350, 2877, 2833, 2274, 2201, 1135, 2086, 3324, 3465, 3341, 2991, 3644, 2513, 3069, 2357, 4085, 3466, 595, 3491, 3479]. Romelio [601]. römischen [571, 798, 994]. Roms [571, 798, 994]. Ronald [438, 1580, 857]. Ronan [1315]. Ronceverte [511]. Ronchi [3071]. ronde [1224]. Ronne [235]. Ronny [3736]. Rookmaaker [841]. Rooms [4044]. Roos [3737]. Roosth [3447]. Roots [167, 435, 371]. Roriz [899]. Rosa [2335, 2880, 1082, 1020, 3008]. Rosales [2473]. Rosanna [2530, 2531]. Rosario [1692]. Rose [1110]. Roselló [3412]. Roselló-Izquierdo [3412]. Rosemberg [194]. Rosen [1571]. Rosenberg [1185]. Rosenfield [505]. Roser [352]. Roshdi [3568, 2710, 2199]. Rosicrucianism [1110]. Röslin [2138]. Rosner [2724]. Ross [1409, 3210]. Rossana [2215]. Rossella [3481]. Rosselló [2114]. Rossetti [448]. Rossi [202, 3194, 1957, 2028, 3115, 3061, 2461, 1927, 2659, 1369, 2070, 2117, 2319, 1651, 2692, 2117]. Rossini [223]. Rossiter [440]. Roter [363]. Rotoli [869]. Rotta [3655]. Rotterdam [841, 1791, 1792, 840]. Rotundo [1169]. Rotzoll [2222]. Rouelle [3160]. Roumette [701]. Roussau [1065]. Rousseau [759, 198, 3197]. Routes [3226]. Routledge [579, 724, 1692, 2601, 3059, 1103, 1104, 219, 1980, 2366, 2412, 2417]. Rovai [1768]. Rovatti [2276]. Rovereto [2619, 2796, 2619]. Roverino [2346]. Rovida [1243, 2232]. Rovigati [862]. Rowan [176]. Rowbottom [190]. Rowe [2977]. Rowlandson [1818, 3879]. Roy [224, 1476, 225, 1845, 342, 642, 1449, 2464, 2631, 2071, 554, 1182, 1433]. Royal [11, 615, 1081, 1370, 495, 1314, 2698, 2662, 2751, 2719, 169, 767, 919, 3542, 1008, 839, 429, 761, 916, 1408, 3664, 1476, 3471, 2700, 2303, 3877, 3468, 1299, 2702, 3489, 3474]. Royale [2267, 2495, 3598, 1681, 2894, 3877, 2469]. Royaume [584]. Royston [2604]. Rozwadowski [2553]. Rubbettino [2616, 680, 1850, 3089]. Rucellai [1307]. Ruchat [2856]. Rudbeck [1365]. rudits [1875]. Rudman [2987]. Rudolf [307, 220, 415]. Rudolph [1285, 106, 2364, 362]. Rudolphinae [3049]. Rudwick [2103, 2756]. rue [2898]. Ruedi [915]. Ruediger [1936]. Ruellius [1905, 4035]. Ruffini [400]. Ruggero [2597]. Ruggiero [412]. Ruggiu [1900]. Ruins [3431, 3439]. Ruisinger [3116]. Ruju [1589]. Rule [3204, 3219]. Rumao [2356, 2357]. Run [3627]. Runge [155]. running [2525]. ruolo [2166, 2237, 3818, 2322]. ruota [782, 578]. Rupert [195, 1587, 1366, 1588]. Rupescissa [567]. Ruprecht [2750, 966, 220, 542, 364]. Ruse [2793, 2791, 1847]. russa [1691]. Russel [3579, 819]. Russell [11, 1819, 549, 2381, 1343, 1798, 2606]. Russia [2228, 3902]. Russian [2822, 770, 4025]. Russland [2228, 3902]. Russo [2646, 3061, 1386]. Ruston [3044]. Rutgers [227, 440]. Ruth [2746]. Rutherford [1535]. Rutledge [1808]. rvolution [653]. Ryan [1393].

S [61, 164, 516, 583, 617, 801, 759, 804, 1660, 2488, 2467, 2055, 2942, 2897, 3286, 1682, 3255, 2155, 3015, 3122, 2909, 252, 309, 447, 576, 720, 717, 716, 887, 964, 873, 1105, 1157, 1279, 1280, 1281, 1156, 1360, 1448, 1688, 1685, 70, 2103, 1515, 1791, 1481, 3127, 413, 2756, 2058, 752, 2955, 613, 2922, 2999, 415, 1943, 874, 881, 2754, 3258, 958, 1431, 2154, 2525, 2987, 616, 1001, 455, 1013, 1193, 2541, 2574, 1065, 1731, 1838, 799, 800, 2678, 2132, 1797, 3044, 1326, 2967, 3092, 1388, 1920, 2556, 1493, 1905, 1389, 2313, 953, 3448, 974]. S. [551, 1735, 1673, 1927, 1843, 1840]. s.a [199]. s.a.s [1354]. s.d [1167]. S.E.A [711]. s.i.p [2080, 2268, 2428, 44, 2458]. s.i.p. [282, 931, 1928, 1669, 1712, 1848, 276, 203, 13]. S.J [549, 2379, 2493, 93, 2833]. S.J. [2176]. s.l [2579, 452, 878, 1195, 1743, 1671]. s.n.c [1084, 1343]. s.n.t [633, 963, 1162, 2317]. S.p.A [41]. S.P.E.S [1755]. s.p.i [582]. s.p.i. [541]. s.r.l [1159, 1874, 1836]. s/N [953]. Sa [1187, 2757, 4089]. Saale [1992]. sabaude [429]. Sabin [2938, 2115]. Sabina [2592, 3614]. Sabine [3344]. Sabino [2379]. Sabix [1090]. Sabliers [3362]. Sabra [791]. Sabrina [3439, 3057, 2135]. Saccenti [1322]. Sachiko [1721]. Sackmann [183]. Sacks [771]. sacré [582]. Sacred [3214]. Sacrificial [2840]. Sadi [3564]. Saemtliche [1001]. Saetti [2368]. Saffron [3570]. Sagamore [3041, 2891, 2972, 2553, 2917, 2907, 2839, 2828, 3036, 2745]. Sage [3166, 2848]. Sagep [562, 861, 1240]. Saggi [1528, 1949, 2020, 312, 2577, 1190, 1594, 2117, 684, 1922, 1998, 1244, 1565, 1267]. Saggiatore [629, 612, 677, 1242, 300, 2433]. Saggio [995, 2334, 934, 684, 2203, 1077, 3864, 67, 88, 563, 2285]. Sahlins [1729]. Saint [574, 2509]. Saints [3595]. sajandi [938]. Sala [1544, 3435]. Salandin [74]. Salavert [2114]. sale [2627]. Salem [1434]. Salento [3329]. Salernitana [1811, 678]. Salerno [312, 2592, 757]. Saliba [2838]. Salis [208]. Sally [3570, 2737, 3637, 683]. Salmoiraghi [3504]. Salomone [2386]. Salon [3473, 1083]. Salons [1012]. Salsament [914]. salsamenta [3811]. salubrit [1947]. salubrita [2018]. Salud [1855, 3867, 1849]. Salus [916]. Salute [2797, 365, 3672, 3057, 1028, 1030, 1716, 433, 1186, 2624, 3951, 883]. Salvat [882]. Salvataggio [1781]. Salvatore [2667, 3655, 2620, 504, 2316]. Salvemini [611]. Salviati [900]. Salvo [3315, 2319]. Salzburg [492]. Salzburger [253, 492]. Samantha [3076]. Samat [872]. Samh [353]. Sámi [3583]. Samling [1218]. sammanställda [1019]. Sammlung [3116]. Sampaolo [612]. Samuel [763, 362, 811, 158, 124, 362]. San [572, 3590, 190, 953, 2009, 2368, 1840, 3397]. Sana [1227, 435]. Sanchez [3091, 2907]. Sanctity [3455]. Sander [177]. Sandglasses [395]. Sandra [1325, 277, 2206, 3125]. Sandri [1755, 953]. Sandro [727]. Sands [2683]. Sangro [2497]. sani [2738]. Sanit [1056, 2376, 711]. Sanità [1675, 1677, 1676, 880, 1436, 435, 1025, 1435, 360, 1598]. sanitari [1678]. sanitaria [3851, 1260, 2154, 2275]. sanitate [3057]. Sankaran [3601]. Sankt [2517]. Sanmarino [3015]. Sansoni [872, 389, 589]. sant [175]. Santa [1784, 3177, 1335, 2637, 2871, 709, 948, 277, 1354, 2115, 2364, 1778, 3343, 3213]. Santha [1372]. Santiago [3415, 2068, 3990, 2068, 3990]. Santinello [621].

Santini [2416]. Santis [1167]. Santit [3178]. santo [1756]. Santoni [1781]. Santorio [2327]. Santucci [969]. Sant’Uffizio [2348]. Saper [1746]. Sapere [3161, 88, 3614]. saperi [2789, 3206, 2627]. sapientia [996]. Sapienza [2421, 1516, 2623, 1619, 2151, 2949]. Sara [2337, 2288, 1267, 1269, 1835]. Sarah [3653, 3409, 2578, 2578]. Saranno [503]. Sarasohn [3636]. Sareshel [535]. Sarlo [4040, 2224, 4040]. Sarma [329]. Saronno [332, 434, 436, 785, 1027, 1026, 1177, 1175, 1271, 1434]. Saros [525]. Sarti [493]. Sarton [214]. Sartorio [1528]. Sarzana [2685]. Sassi [745]. Satellite [3658]. Satirical [2247]. satisfaction [3916, 2016]. satisfied [847]. satisfyed [4051]. Satolli [586, 1716]. saubere [1405]. Saul [955, 2771, 347]. Saur [1387]. sauvage [1170]. Savage [184]. Savage-Smith [184]. savans [541]. Savant [1700, 3767, 40, 1734, 3330, 4089]. Savantes [243, 836, 1327, 4042]. Savants [3425, 653, 3769, 1612, 2273]. Savellini [2320, 2224]. Savigny [64]. Savini [1671]. Savino [1283]. Savoia [2421, 3426]. savoir [33, 3540, 1486]. Savoirs [3527, 1804, 886, 3895, 1170]. Savoisi [502]. Savona [2683]. Savoy [429]. Sawmills [2009, 3747]. say [4028]. scaffale [2880]. Scaino [3230]. Scala [2115]. Scale [265, 3680]. scalpello [2232, 3973]. Scambi [2450]. Scambio [989]. Scampoli [187]. Scandali [1036]. Scano [1693]. Scans [2506]. Scapecchi [1283]. scapigliato [3546]. Scapparone [2949]. Scappini [1325]. Scarabelli [2007]. Scarantino [2380]. Scarpelli [1244]. Scarpellini [2513]. scavo [575]. Scazzieri [3041]. scelti [896, 1959, 2030]. scenario [1901]. scene [2844]. Scerri [3124]. Sceuastica [1531]. Schabas [2807]. Schaer [1804]. Schaffer [168, 1533, 3018, 616]. Schaller [1918]. Schardin [1012]. Schatkamers [1422]. Schechner [1835]. Scheda [1632, 814]. Schede [603, 572, 3687, 1794, 1269]. Scheele [507]. Scheele-Ehrung [507]. Scheifele [1797]. Scheintod [3219]. Scheiwiller [552, 2367]. Scheiwiller-Finmeccanica [552]. Scheletri [3693]. Schelling [896, 343, 1034, 1328, 2734, 2465, 4033, 3571]. Schelling-Tagung [343]. Schellings [1202, 1328, 4033]. Schematic [3533]. Scheme [460]. Schettino [75, 854]. Schianchi [2368]. Schiaparelli [2450, 3053, 3497]. Schicksal [565]. Schiebinger [2657]. Schillinger [1012, 1083]. Schio [1961, 2032]. Schipperges [193, 433]. schizofrenie [288]. schizophrenia [1184]. Schloss [1402]. Schmalz [763]. Schmidt [114, 2978, 3070, 1905, 4035, 98, 3458]. Schmitt [3026, 114]. Schnack [64]. Schneider [259, 3129, 1445, 1714]. Schnitzler [951, 1854]. Schóbinger [115]. Schoening [248]. Schofield [1910, 2534]. Schola [678]. Scholar [2183, 2957, 2186, 1589, 4089, 1196]. Scholarly [3767]. Scholars [2231, 2180, 3803]. Scholarship [1685, 2662, 2268]. Scholastic [86, 274, 719]. Scholem [1528]. Scholz [1150]. Schönbeck [2647]. Schone [302]. School [1928, 328, 425, 1860, 3489, 4091, 4076, 2369, 4017, 3690]. Schools [1278, 1084]. Schoop [3037]. Schopenhauer [1965, 2036, 1767, 3954]. Schreck [1873]. Schriften [1602, 248, 307, 226, 217, 183, 1001, 1328, 787]. Schriftenreihe [557, 1797]. Schrijvers [3109]. Schrödinger [1579]. Schroeder [188, 1172, 1172]. Schrotflinte [2544]. Schubert [1571].

Schüddekopf [226]. Schulmann [2977]. Schultz [2544]. Schumann [1571]. Schuster [3441, 80]. Schütt [961]. Schwabe [172, 3530, 2529, 2205]. Schwarzbach [348]. Schweiz [1687]. Schweizer [172]. Science [15, 77, 167, 227, 164, 242, 301, 520, 440, 414, 490, 544, 615, 672, 722, 923, 842, 1070, 1005, 1055, 1108, 1264, 1204, 1333, 1365, 1346, 1362, 1457, 1404, 1454, 1617, 1587, 1657, 1692, 2075, 2056, 2160, 2463, 2464, 2527, 3041, 3016, 3187, 3240, 3508, 2708, 3262, 3112, 3509, 2038, 2168, 2239, 3499, 3482, 495, 652, 532, 1097, 1568, 1495, 1898, 1911, 2193, 2822, 2933, 2882, 2990, 3475, 3464, 3463, 2891, 3007, 3184, 1711, 847, 856, 1112, 3222, 394, 214, 2972, 2086, 2649, 2908, 3053, 3150, 3369, 3512, 2558, 2671, 3574, 2542, 857, 2248, 2245, 2510, 3651, 2553]. Science [2917, 3421, 538, 824, 2540, 3433, 1147, 1366, 1494, 1560, 1707, 1791, 2059, 1588, 2920, 1063, 3324, 2755, 17, 328, 3205, 3465, 2756, 2838, 2657, 2594, 2546, 3717, 3198, 3675, 2084, 2751, 2907, 275, 2823, 3487, 2839, 2988, 3511, 3849, 3852, 3853, 3854, 3411, 3289, 3288, 1, 2740, 3188, 2999, 760, 925, 1395, 2771, 2947, 754, 2610, 2050, 2377, 3639, 2788, 2828, 1477, 3331, 2521, 2681, 2862, 2561, 3033, 4071, 2928, 1479, 829, 1315, 2095, 2197, 3578, 2768, 3074, 3223, 1841, 1299, 2533, 2927, 1466, 2900, 65, 1944, 2847, 3539, 3407, 3523, 2762, 1126, 2568, 3542]. Science [3191, 3529, 2505, 2775, 2944, 3166, 3267, 1557, 3036, 4078, 2761, 2683, 409, 916, 1072, 992, 1320, 1511, 1544, 1750, 1662, 2069, 2068, 2098, 1655, 3935, 2666, 2858, 3038, 2733, 3224, 3252, 3947, 2924, 2976, 2731, 3028, 3202, 2926, 1397, 3525, 163, 1980, 2867, 3183, 2264, 2711, 2803, 625, 1723, 2247, 974, 2804, 3044, 3493, 1541, 348, 811, 892, 2967, 3334, 3429, 2640, 4086, 3119, 2925, 4087, 2575, 382, 1127, 3393, 4001, 734, 743, 3466, 1986, 2147, 2456, 2674, 2604, 3207, 1800, 1294, 2127, 1636, 68, 54, 2562, 3459, 3470, 2088, 3213, 3242, 2417, 283]. science [685, 1194, 1261, 1406, 1562, 1683, 2067, 4039, 3749, 2237, 3759, 3762, 3765, 3766, 533, 287, 4048, 3778, 10, 182, 2551, 3800, 2066, 1111, 43, 3807, 3127, 3813, 2646, 3831, 3829, 3832, 3742, 1735, 2747, 3868, 3869, 1543, 3876, 4069, 4070, 3882, 3883, 3886, 3894, 1878, 1915, 3904, 3258, 3906, 3069, 3914, 3921, 3925, 920, 762, 774, 1170, 1422, 1570, 2541, 1897, 4080, 1848, 3943, 3944, 3945, 1410, 2745, 2060, 3953, 1265, 3956, 1694, 1799, 1413, 1400, 993, 1076, 1109, 1053, 2861, 1337, 3992, 3996, 2901, 2422, 1381, 1845, 1245, 4024, 672, 1490, 1763]. Science [2194, 2143, 2160, 2464, 2459, 669, 823, 907, 1478, 1556, 1635, 1967, 3752, 2606, 2737, 2538, 2079, 2794, 2633, 3033, 1484, 1641, 1982, 2629, 2945, 1561, 2684, 3267, 2799, 55, 271, 326, 2660, 2124, 2749, 2412, 2985, 3034, 2803, 3032, 2982, 3990, 2911, 2498, 2545, 2067, 1155, 2740, 2537, 2848, 3674, 3363, 1609, 3414, 3637, 3438, 1466, 2555, 3428]. Science-Industry [2972]. Sciences [1081, 2267, 2495, 4043, 653, 675, 2053, 495, 1681, 2894, 3254, 287, 305, 1985, 551, 3430, 1327, 2215, 3531, 2756, 2662, 368, 2259, 2361, 2995, 1409, 3189, 419, 763, 2015, 2013, 1475, 3159, 1008, 2615, 701, 1421, 2541, 753, 749, 2446, 3183, 2273, 2856, 1386, 1486, 2422, 1224, 2188, 1290, 1814, 1795, 2145, 4042, 3395, 282, 1793, 2735, 620, 243, 836, 2611, 1859, 2229, 314, 2757, 2973, 3876, 3877, 3893, 954, 3909, 713, 295, 2118, 611, 2539, 1734, 1198, 1912, 2861, 3992, 2613, 1476, 3471, 2093, 3469, 1701, 1830, 2229, 2201, 3470, 548, 748, 3287]. Sciences [3598, 3507, 3445, 3631]. scientia [3748, 3282, 2419, 4015]. Scientiae [2639, 377]. Scientiarum [2418, 235, 257]. Scientific [15, 335, 801, 805, 837, 842, 1313, 1587, 1660, 2228, 2496, 3569, 3000, 1393, 2169, 2240, 2179, 2299, 2436, 3467, 3328, 2425, 3055, 1495, 2193, 3228, 985, 3694, 166, 4048, 3322, 3080, 63, 2295, 3469, 3050, 2183, 1702, 3030, 511, 3637, 2161, 2216, 2824, 1481, 2697, 3111, 3717, 2888, 3366, 2948, 3703, 2259, 2361, 317, 3139, 2947, 3868, 1722, 2268, 4068, 2681, 2718, 3357, 3503, 3191, 761, 835, 916, 998, 1161, 1256, 2792, 3936, 3456, 1319, 2975, 1458, 1723, 2457, 3554, 3474, 1972, 2591, 2640, 3983, 3984, 2485, 1337, 3295, 2815, 3445, 3361, 2180, 3419, 1825, 2772]. Scientific [2562, 759, 1562, 3746, 1576, 1573, 1829, 1499, 1928, 3848, 4072, 3893, 1944, 3906, 4076, 778, 992, 1860, 2069, 2068, 2369, 3948, 1546, 692, 849, 993, 1912, 3982, 3990, 3999, 1270, 4091, 1577, 4018, 3023, 3902, 2195, 2412, 4085, 4058, 4017, 3150, 3334]. Scientifica [1549, 2092, 2871, 403, 288, 1343, 1137, 1482, 189, 2425, 1574, 3077, 312, 2672, 3029, 3848, 871, 2880, 778, 1619, 3948, 1114, 1658, 4085, 3982, 1270, 1850, 178, 2416, 2167, 2238, 2149]. Scientifiche [358, 632, 1021, 2370, 2064, 486, 1829, 1304, 1472, 1995, 19, 1339, 384, 2164, 2235, 569, 2670, 1928, 2888, 2712, 2651, 1876, 811, 1499, 2557, 2141]. Scientifici [777, 1696, 1054, 396, 1220, 2045, 1546, 181, 3022, 1260, 1259, 2112, 2415, 969, 1922, 1998, 1244]. scientificidi [1016]. Scientifico [2043, 1629, 605, 2313, 2376, 2449, 664, 1424, 2399, 515, 1691, 3065, 2387, 2632, 4076, 88, 453, 1861, 2002, 1931, 2149, 2414, 2104, 2068, 2468]. scientifico-letterario [2632]. Scientifique [540, 1700, 1386, 1795, 2309, 1612, 2652, 3936, 3657]. Scientifiques [834, 1060, 2894, 2048, 2566, 728, 4042, 584, 1155, 2798, 886, 1327, 2757]. Scientist [968, 3322, 3025, 3699, 2617, 2081, 347, 2345, 378, 2971, 1589, 1219, 4076, 1195, 1874, 2728]. Scientists [4042, 4, 3059, 741, 3176, 3716, 3977, 68, 303, 802, 1112, 3831, 541, 3335, 3769, 2362, 2626]. Scienza [246, 583, 773, 894, 46, 521, 1035, 1205, 1374, 1463, 1946, 2168, 2239, 2169, 2240, 1221, 2448, 594, 688, 619, 1235, 1711, 1295, 1112, 1470, 1552, 214, 989, 2376, 2473, 442, 447, 910, 1637, 2248, 1969, 1243, 1516, 1499, 1560, 1513, 1757, 2224, 682, 2893, 2395, 986, 2263, 599, 909, 162, 1979, 2261, 1927, 391, 47, 90, 261, 294, 458, 428, 455, 657, 731, 813, 973, 1116, 1016, 1056, 1292, 1343, 2078, 922, 1897, 1344, 1712, 827, 2124, 304, 604, 2306, 1851, 2264, 2316, 1889, 765, 1068, 1053, 1833, 2394, 734, 1832, 2088, 61, 195]. scienza [515, 1616, 2017, 1965, 2036, 2166, 2237, 3752, 3689, 339, 532, 677, 1368, 2491, 3545, 2524, 2313, 448, 809, 1462, 3544, 2504, 2789, 3807, 3813, 3831, 2941, 2742, 2672, 2461, 2630, 2557, 2140, 1901, 883, 3904, 503, 3496, 2659, 2743, 3426, 3925, 200, 354, 1033, 1452, 1532, 1806, 2150, 3314, 3652, 570, 680, 3956, 1061, 1653, 1579, 1639, 2062, 2132, 2597, 942, 1426, 1713, 1803, 3203, 373, 370, 3190, 3655, 2581, 2810, 1485, 745, 346, 2968, 2692, 3089, 1691, 3099, 2670, 2157, 1561, 186, 2100, 2946, 1265, 1289, 611, 1111]. Scienze [61, 22, 516, 514, 1091, 3750, 46, 2351, 4045, 1648, 2882, 2431, 75, 76, 716, 1077, 274, 1701, 1516, 1603, 2092, 3408, 1909, 2536, 1, 1477, 2042, 554, 1251, 1195, 1619, 1871, 2112, 827, 1737, 1899, 2132, 12, 412, 2501, 1318, 1420, 1709, 2628, 442, 887, 752, 81, 912, 200, 1873, 186, 1639, 178, 189]. Scienziati [1894, 4, 737, 269, 461, 821, 1383, 1471, 2090, 741, 1124, 1828, 323, 463, 2447, 600, 1302, 1782, 1973, 663, 2129, 2353, 2397, 397, 2249, 95, 1633, 1371, 528, 818, 1066, 1112, 2969, 809, 3831, 581, 2507, 750, 3335, 2921, 3718]. Scienziato [3326, 2473, 2014, 1219, 586, 2081, 2650, 313, 3757, 340, 2472, 2769, 2156, 4076, 1453, 1874, 1579, 2919]. Scilla [1957, 2028, 3432]. Scimmie [597, 2125]. Scintillation [2482]. scintille [2904, 3955]. scolastica [446]. scolastico [719]. Scolastique [3173]. Scolopi [1372, 453]. Scolopica [1372]. Scolopio [664, 688, 655]. Scoperta [421, 1166, 1241, 627]. Scoperte [2142, 1564]. Scopoli [1145]. Scorci [3190]. Scossiroli [421]. Scot [3247]. Scotland [2899, 3289, 169, 761, 3100]. Scotland.s.l [761]. Scott [2707, 2964, 2755, 2311]. Scottish [652, 341, 624]. scratch [435]. Scribal [3507]. scrigno [1451]. Scrinium [1687]. Script [3050]. Scriptorium [1447]. Scripture [1119]. Scritti [1959, 2030, 1580, 223, 2283, 3277, 2855, 3964, 1142, 3015, 3838, 390, 727, 1399, 2380, 2501, 1324, 2332, 2588, 4083]. Scritto [739]. scrittori [644]. Scrittura [1119, 1066]. Scritture [3632]. Scrupulosity [2143, 3883]. scudo [3367]. Sculptor [3489]. Sculptural [3458]. Scuola [1811, 3080, 1115, 782, 941, 1645, 1713, 425, 1512, 1321, 818]. scuole [1084]. se [1305]. sea [424, 1076]. Sean [176]. Search [3082, 3176, 2739, 3664, 3729, 1146, 3060, 3919, 1503]. searches [3245]. Seas [3806, 1844]. Seat [1690]. Seattle [959, 1246]. Seattle-London [1246]. Seawall [2683]. Sebastian [3017]. Sebastiano [1343, 799]. Sebi [1901]. Sebregondi [2006]. Sec [1408, 1300, 949]. Secc [572, 1335, 1695]. Secchi [2176, 1671, 981, 1895]. secció [353]. Sección [1645]. Secco [1456]. Secoli [87, 513, 650, 967, 1032, 1361, 1607, 1762, 2158, 1648, 254, 1529, 1942, 37, 1286, 2921, 383, 514, 3763, 3021, 446, 2888, 3885, 952, 1755, 1180, 2880, 559, 1354, 1745, 1861, 4006, 1194, 1689, 1817, 1872, 2524]. Secolo [2389, 486, 2473, 1823, 979, 658, 1978, 660, 734, 1378, 2130, 894, 1066, 2064, 2063, 3783, 2523, 250, 576, 716, 1105, 2383, 2690, 881, 789, 863, 876, 1186, 389, 917, 453, 561, 859, 1251, 1730, 678, 12, 178, 1784, 2356, 2399, 573, 496, 2116, 2780, 2213, 2677, 3496, 3329, 1324]. Second [830, 3151, 244, 406, 2345, 2362, 3949, 3965, 4028, 4024, 2400]. seconda [1569, 1790, 3949, 2107]. secondaire [1859, 4000]. secondary [4000]. Seconde [2767, 1567]. Secondi [1606]. Secondo [1606, 2370, 2400, 878, 621, 2220, 2206, 1645, 165, 3965, 3103]. Secord [1643, 3510]. Secret [3529, 3704]. Secrets [2881, 3313, 1397, 3963, 3919, 749, 2835, 3060]. Section [4042, 243, 836, 1327, 2230]. Sections [3471]. Sectis [1134]. Secular [925]. Secularisation [394]. século [430]. secunda [4028]. security [762]. sede [238]. Sedibus [299]. Sedley [1225]. Seed [3603]. Seeds [3607]. Seeing [3553, 3051, 3387]. Seemann [1012]. Segala [1957, 2028, 1965, 2036, 2166, 2237, 2620, 656, 1373, 2117, 1615]. Seget [3247]. Segherie [3747, 736, 2009]. Segnali [2317]. Segnalibro [2010]. Segnatempo [426]. Segni [310, 2117]. segno [2339]. Segoloni [747]. Segons [3013]. Segrate [1317]. segreti [2881, 3443, 1000]. segreto [3397]. seguaci [1565]. segue [814, 522]. seguente [2974]. Séguin [3200]. seguito [814, 522]. Séha [3277]. Séha-Milano [3277]. sehend [1918]. Sei [180, 1426, 2100]. Seibold [1074]. seicentesca [1000]. Seicento [515, 1789, 50, 1453, 922, 733, 3741, 2659, 2226, 2650]. Seidel [1288, 577]. Seils [1602]. Seine [172, 1918]. seinen [362]. seiner [1602, 543]. seismic [2363]. Seismograph [2484]. Seiten [1731, 1854]. seizième [1567]. Séjour [893, 2013, 3932]. Seklernas [2119]. Selbstbewustsein [1989]. Selbstdenker [1602]. Selbsterkenntnis [1989]. Selbstverlag [787]. Selbstzeugnisse [1764]. Select [1055]. selecta [1343]. selectae [1157]. Selected [3076, 230, 249, 3992]. Selection [4041, 699, 2216]. Selezione [2445]. Self [3736, 1246, 3244, 2360, 1989]. self-esteem [1989]. self-knowledge [1989]. Self-Persuasion [2360]. Self-Presentation [3736]. Self-regulating [3244]. Seligardi [2425]. Sella [42]. Sellerio [764, 930, 2062]. Sellier [1185]. selon [2663, 3180]. Selva [2174, 1704]. sémantique [444]. Semen [1133]. Semerari [2431]. Seminar [1481]. Seminari [1619]. Seminario [2224, 1943, 1101, 1373, 2271]. Sempione [2743]. Semplici [1268, 863, 917, 2581, 1267, 1811]. Senders [399, 2251]. senese [1260, 200]. Senesio [3964]. senologia [503]. Sensation [3727, 3034]. Sensazionale [2263]. Sense [3050, 830, 3302, 2862, 300, 2069]. Sensi [2269]. Sensibility [2732]. senso [1957, 2028]. sent [2675]. Senthil [3712]. Sentimental [3425]. Senza [906, 2423]. Separating [3173]. Sepkoski [3602]. Seppilli [2336]. September [2933, 2050, 4078, 452, 3932, 363, 1331, 2701, 4078]. septembre [1442, 2013, 954, 868, 713, 3932]. Septentrion [2423]. septième [444]. Sera [867]. Serafina [2874]. Serafini [617]. Serapide [3209, 2796]. Serbal [494, 618]. Serbal-Consejo [618]. Serena [3616]. Serge [496]. Sergio [407, 2415]. Serie [2522, 1516, 43, 14, 727, 2114, 2062, 2100, 1091, 1019]. Series [224, 1288, 1587, 847, 886, 1278, 2381, 1669, 547, 2719, 3345, 613, 767, 1395, 3412, 1936, 1934, 1933, 1935, 1873, 2204, 1392, 3241, 1363, 440, 2216, 2098]. Serio [1748, 1863]. Sermons [2966]. Serpieri [2014, 4076, 665, 4076, 664]. Serra [3298, 2895, 2951, 3259, 2895]. Serrano [3738]. Serravezza [1616]. Serres [2578]. Servei [2099]. Service [2379, 1040, 1224]. Services [990, 3839]. servicio [965]. servizi [3839, 1436, 1678, 2918]. Servizio [1794, 2113, 2363, 619, 455]. ses [444, 2648, 1859, 338, 2952, 4000]. Sesiano [2519]. Sessions [1877]. sessuale [852, 730]. Sesti [1519]. seta [1540]. setensi [1806]. Setim [13]. Setim-Editrice [13]. Sets [2786, 457]. Sette [2315, 1354, 4049]. settecentesche [493]. Settecento [39, 313, 517, 1200, 1427, 1539, 726, 340, 1236, 180, 72, 1824, 2536, 778, 1331, 818, 1651, 2490, 3022, 2893, 2100, 2954, 3689]. Settembre [1015, 1488, 915, 1943, 1168, 2113, 1331]. Settembre/31 [1015].

Settentrionale [818]. setting [1275, 1005]. settings [3905]. Settle [1532]. seu [4028]. Seuchengeschichte [62]. Seuil [1668, 1170, 1410, 2052]. Seurat [1485]. Seven [3345]. Seventeenth [335, 891, 1404, 1490, 2056, 2776, 2462, 3535, 1439, 830, 2888, 2825, 1320, 3600, 2099, 3654, 3393, 3562, 3373, 3502, 3803, 3234, 659, 1770, 3869, 3474]. Seventeenth-Century [891, 1404, 1490, 2056, 3535, 1320, 2099, 3654, 3393, 3562, 3234, 659, 3869, 3474]. Seventh [1749, 4045, 2213]. Severinus [2658]. Sex [587, 651, 1945, 3633, 2119, 1280, 2527]. Sexual [1449, 2016, 3916, 544]. Sexualité [33]. sexualitetens [2119]. Sexuality [177, 651]. Seyssel [1080]. Sez [2217, 2005]. Sezione [517, 250, 291, 727, 2608, 2726, 697, 1802]. Sezioni [575]. sfera [1160]. sfida [2826]. sfondo [2062]. Sforza [1856, 3009]. SFR [172, 547, 1731, 1905]. Sfruttamento [734]. SGA [199, 779, 519]. Shackelford [2658]. Shadow [1406, 2739]. Shadows [724, 3304]. Shadowy [3391]. Shaftesbury [1233]. Shakespeare [491]. Shakespearean [2655]. Shank [3118, 3445]. Shape [1490, 3620, 3633, 1541]. Shaped [2055, 3896]. shapes [1505]. Shapin [168, 1404, 1533, 2093, 1722]. Shaping [2947, 3131]. Shapiro [1199, 1198]. Sharon [1980, 3567, 3044]. Sharpe [717]. Sharples [1225]. Shaw [2076, 3241]. Shea [490, 722, 842, 1204, 2160, 2594, 2839, 2595]. Sheffield [1812, 1936, 1934, 1933, 1935, 2365]. Sheller [3333]. Shelling [806]. Shen [2095]. Shepherd [225, 1100]. Sherwin [1092]. Shifting [692, 2917]. ship [1076]. Short [1164, 3418, 1550, 1071, 608]. Shortell [2707]. Shorter [80, 1315]. Shortland [1562]. Should [2870, 3451]. Shoulders [913]. Show [4067, 2976, 3150]. Shows [3196]. Shuie [599]. Shuttleworth [2737, 3637, 683]. Siacci [1706]. Sib [582]. Sibylla [1185]. Sic [1054]. sich [1726]. Sicilia [1648, 572]. Siciliana [1648, 764]. Siciliano [514, 2429, 2208]. Sick [3343]. Sickness [642, 1812]. sicurezza [2918]. Sida [566]. Side [289]. Sider [2939]. Sidereus [3225, 1064, 1481, 1118]. Sidney [194]. Siebente [1001]. Siècle [1626, 210, 1700, 2199, 260, 512, 425, 3395, 1793, 1567, 1804, 2798, 306, 1187, 1447, 1984, 3803, 1138, 2973, 3890, 3894, 1875, 540, 1799, 3657, 3058, 3980, 582, 746, 175, 1121, 675, 2573, 496, 2735, 2894, 3013, 40, 3087, 2550, 2829, 2652, 2596, 2877, 308, 1060, 3331]. Siécles [2048, 2566, 2878, 1985, 3173, 3742, 920, 1848]. Siege [3597]. Siegel [2324]. Siegemund [2841]. Sieges [1484]. Siegfried [698]. Siena [197, 991, 3844, 1260, 1183, 710, 697, 879, 1180, 200, 561, 575, 1251, 1861, 1931, 2115, 2117, 2149, 2221, 2414, 1653, 2597, 1840, 2581, 1615, 2115, 3937]. Siena-Dipartimento [2117]. sienambente [2221]. Sierotowicz [1920]. Sigem [2502]. sight [1918]. Siglo [1515, 35, 2808, 1860, 494, 618, 4017]. Siglos [889, 43, 3759]. Sigmaringen [1402]. Sigmund [1852, 67]. Sign [3099, 2339, 1792]. Signatura [310]. signed [1704]. significa [1052]. Significance [3124]. significato [1270]. Signor [2726, 740]. Signs [2939]. Sigrist [2539]. Sik [2137]. Silent [3835, 3912, 3804, 3253]. silenziosa [3804, 2323]. silk [1540]. Silke [1872]. Sillabe [2338]. Sillery [2423]. Silva [2808]. Silvana [83, 577]. Silvano [1200, 3547, 1870, 2275, 3085]. Silver [3403]. Silverberg [1113]. Silvestri [2605, 2622]. Silvia [1502, 2905, 3497, 556, 3088, 2147, 2012, 2798, 1695, 1493]. Silvio [3688, 2646, 1320, 2098, 281, 2930]. simbolismo [1868]. simbologia [1314]. simile [2528]. Simili [3041, 1741, 2741, 1876, 2225, 3125, 2274]. Simitis [67]. Simo [1104]. Simões [2895, 3243]. Simon [168, 1533, 1447, 80, 3018, 2537, 3284, 2830, 2728, 3650, 3875]. Simona [1242, 2769]. Simonazzi [2721]. Simoncelli [1291]. Simone [1959, 2030, 2386, 3512, 1671, 3279, 2958, 3826, 2918]. Simonetta [2320, 2224]. Simonton [1337]. Simples [2998]. Simplicio [2933]. Simpson [761]. simultaneità [1802]. simultaneous [1810]. Sin [3480]. Sinaceur [1668]. Since [857, 651, 1568]. Sinesio [3277, 2574]. sinestesie [203]. Sing [3477]. Singapore [2216, 2268, 2258]. Singapore-New [2216]. Singapore/New [2268]. Singer [282]. Singer-Polignac [282]. Single [2391]. Singola [2391]. Singular [3]. Sinn [2573]. Sinologia [909]. Sinological [2134, 3780, 3787]. sinottica [3873]. Sintesi [67]. Sinthese [28]. Sir [454, 1370, 2003, 2717, 2156, 3635, 808, 4032, 3809, 3836, 4032]. Siraisi [2656, 2950, 381, 3442]. Sironi [1030, 1676, 2918]. Sirri [2370]. Sis [1161]. Sismel [2863, 3006, 2520, 2934, 2419, 2880, 2877, 2645, 3062, 2820, 2628]. Sismel-Edizioni [2520, 3062]. sismiche [2363]. Sismico [2363]. sismografi [779]. sismologia [2271]. Sismondi [72]. Sistema [2246, 2248, 1373, 1951, 2022, 2334, 3457, 562, 2285, 1034, 1870]. sistematico [1903]. Sistemi [3161, 1375, 257, 3858]. Sister [3615]. Sisto [2068]. Site [94]. Sites [3033]. Siti [2244, 2343, 2440]. Sitoni [965]. Situated [3191]. Situation [2015, 2013]. Sitzungsberichte [382]. sive [1962, 2033]. Six [3736, 2970]. sixiéme [2214]. Sixteenth [2776, 180, 279, 2347, 2961, 425, 1193, 1511, 3114, 3479, 2356, 3513, 735, 3343]. Sixteenth-Century [1193, 3114, 3479, 735, 3343]. Sj [1610]. Skepticism [3735]. Skeptizismus [798]. sketching [3346]. Skin [3152, 3636, 3751, 2628]. skizziert [2915]. Skrifter [2496]. skriftserie [3092]. Skull [3581]. Skulls [3217, 1154, 4009]. Sky [2806]. Slatkine [2539]. Slaviero [2313, 2376, 556]. Slime [3460]. Sloane [3440]. Slotten [1617, 3658]. Sm [2009]. Small [2099]. Smallpox [499]. Smet [855]. Smith [1457, 3070, 184, 851, 1408, 689]. Smithsonian [184, 293]. Smoller [1357]. Smtliche [217]. Sn1054 [2136]. Snakes [3662]. Snchez [494, 618]. Snowflakes [3520]. Snyder [2858]. So-Called [3276]. Sobel [1650]. Sobre [3167, 889, 3759, 43, 3867, 2494, 1849, 1855, 2097]. Soc [711]. Social [439, 592, 1404, 2496, 3055, 1991, 925, 1246, 285, 3636, 2505, 3096, 3528, 2237, 4075]. Sociale [1169, 712, 2166, 2237, 1674, 1372]. Sociales [287, 305, 2735, 2613]. Sociali [2117, 1615, 1099, 712]. socialista [546]. Societ [2235, 2042, 1748, 1749, 1871, 1919, 1482]. Società [2164, 199, 2628, 517, 2616, 1248, 2642, 3402, 804, 931, 85, 442, 2504]. Société [2229, 1304, 1472, 2273, 2735, 774, 2877, 1563]. Sociétés [243, 836, 1327, 4042]. Societies [2879]. Society [11, 242, 441, 615, 885, 1204, 1406, 1763, 1819, 1805, 2495, 696, 1898, 10, 3352, 857, 1277, 1445, 1366, 2579, 275, 2303, 919, 225, 3468, 829, 2137, 3542, 3191, 839, 916, 1161, 1193, 1749, 3712, 163, 3398, 1408, 2982, 2890, 1290, 1814, 68, 2971, 517, 2751, 2071, 4001, 341, 3468, 2858, 3474]. Sociologia [246, 300]. Sociology [3096]. Socit [1473, 1543, 1073]. sodalizio [453]. Soemmering [362]. Soemmering-Forschungen [362]. Soemmerring [1004, 622, 362, 565]. Soemmerring-Forschungen [1004, 622, 565]. Soemmerrings [362]. Sofia [72]. Sofistiche [1606]. Sofya [591]. soggetti [58]. soggettivita [1995, 4021]. soggetto [1995, 3057, 4021]. soglie [809, 1237]. sognate [559]. Sohn [259, 234]. soil [2009]. soins [32]. Sökandet [1146]. Solaires [322]. Solar [3292, 1654, 3789, 3790, 3795, 2096]. solare [428, 1083]. Solari [2178, 1119, 1415, 1787, 2001]. Soldano [2882]. Soldato [3279, 1436]. Soldier [1234]. Sole [1680, 1990]. Soleil [3423]. Solida [4026, 3283]. solidariet [2072]. Solidarietà. [810]. solidarum [1960, 2031]. solids [2031]. Solla [406]. Sollenberger [1225]. Solomon [672]. Solomos [2366]. Solution [2812]. Solvay [1879, 3771]. Solvay. [3332]. sombra [2494]. Some [394, 1395, 2291, 3053, 2091]. Somenzi [626]. Somerville [2377, 3943]. Somewhere [3248]. Son [2757, 338, 2878, 2821, 2081, 2782]. Sonda [1140]. sondaggio [2002]. Sonde [1414]. Sonderband [1687]. Sonderheft [1141]. Sonnabend [1838]. Sonne [2943]. Sons [2864, 2814, 1092]. Sophia [3447, 1224, 2760]. Sopra [740, 1954, 2025, 2334, 2317, 2285]. Sorba [3741]. Sorbonne [3075]. sordi [932]. Soresini [457, 2317]. SÖRlin [1023]. Sosio [677, 1998]. Sostenute [2298]. Sotgiu [355]. Sotheby [427]. sotto [3672, 2317]. Souffrin [3287, 2359]. Soul [2388, 2933, 3074, 3724, 734, 2740, 3880, 3012, 2308]. Soule [3316]. Soups [3380, 2678]. Source [829, 2296, 329, 1542, 840]. Sourcebook [2753, 3886]. Sources [1846, 3282, 3205, 524, 1477, 1126, 1557, 2409, 3140, 1618, 1972, 1127, 734, 1225, 2130, 1636, 182, 305, 1273, 3881, 14, 338, 1829]. sous [1795, 2067, 620, 3483, 1612, 368, 1080, 2118, 1848, 1410, 1228, 1073, 1799, 1486]. South [2512, 3716, 3712, 3752]. Southern [841, 192, 2962, 2363]. Southgate [1401]. Soveria [2616, 680, 1850, 3089]. Soviet [54]. Sovietici [46]. Sovrani [2126]. sovrano [250]. sowie [1941]. Spa [638]. Spaans [3736]. Space [3658, 3191, 3600, 1265, 1983, 282, 1999]. Spaces [3184, 2568, 3674]. spacio [1647]. Spadafora [722, 1204]. Spaetmittelalter [251]. Spagnesi [1077]. spagnola [2430, 4010, 1382]. Spain [438, 3624, 2924, 2998, 3511, 3738]. Spallanzani [61, 1569, 1765, 1235, 1111, 171, 602, 817, 2535, 2349, 2903]. spalle [913]. Spanish [2824, 35, 45, 1544, 1743, 1860, 3456, 2198, 3759, 2510, 3867, 3941, 3942, 4010, 4017]. Spark [526]. Sparking [2512]. Sparks [3380, 3955, 2678]. Spary [1643, 3510]. Sparzani [2632]. spas [1182]. Spätankite [1684]. spatial [2745]. Spatializing [3582]. spaziale [2321]. spazio [2429, 1265, 2685]. speaks [3908]. Specchio [2942, 777, 2328, 2279, 1020, 3961, 3632, 3259]. Specerey [1095]. Specerey-Hndler [1095]. Special [1977, 613, 1724]. Spécialis [2188, 4028]. Specialist [2495]. Specie [2391, 1098, 680, 1244, 788]. Species [1656, 3926, 2391, 788]. Specificity [1656, 3926]. specimen [1744]. Specola [1790, 3138, 2699, 2293]. Spectacles [2890]. Spectacular [2700]. Spectra [2303]. Spectro [2408]. Spectro-Bolometer [2408]. Spectroscope [2302, 2300, 2305, 2406, 2404]. Spectroscopic [2405, 2089, 2406]. Spectroscopies [2301]. Spectroscopy [2403, 2301, 2407, 2409]. Spectrum [2408]. Speculation [3480]. Speculationi [596]. Speculationum [534]. Speculative [2606]. speculazione [1957, 2028]. Speculum [534, 3001]. Spedali [3177]. Spekulant [1602]. Spekulation [1202]. Spencer [1878, 1786, 3765]. Sperienze [1235]. Sperimentale [1969, 2476, 2081, 74, 874, 1532, 550, 924]. Sperimentali [1467]. sperimentazione [2338, 1753]. Spezia [3120, 2317]. Speziali [1676]. Spezielle [1977]. Spezieria [2368]. sphere [233]. Spheres [1346, 3477, 3004]. Spherical [2084]. Sphinx [608]. spiegare [371]. spiegazione [1499]. Spiegelbild [2146]. Spiegelder [571, 798, 994]. Spiesser [2596]. Spillane [25]. Spillius [1769]. Spinoza [384, 384]. Spinoziana [1943]. Spiral [3051, 1037]. Spirale [1037]. Spirit [685, 3392, 4088]. spiriti [1426]. spirito [926, 1873]. ’Spiritu [596]. Spiritual [3606]. spirituale [1521]. spiritualismo [1244]. Spiritualismus [1086]. Spiritus [208]. Spitz [695]. Spoglio [2261]. Spoleto [995, 2222, 3972]. Sponsel [3480]. Spontaneous [3460]. spora [2334, 2285]. Spot [3379]. Sprache [1381]. Sprachtheorie [1684]. Sprachtheorien [1684]. Spread [1627]. Sprevak [3417]. Sprimont [2428]. Spring [1776]. Springer [2653, 2646, 3061, 3079, 201, 1074, 3224, 3525, 348, 2967]. Springer-Verlag [348]. Sproegel [135]. spropositi [3373]. Spruit [2497]. Sptmittelalter [1086]. Spunti [1452]. Spurious [3298, 3970]. Sputnik [3658]. Squibb [767]. Squillacciotti [635]. Srinivasa [2503]. srl [1901, 1840]. St [251, 2662, 244, 3017]. Staatlicher [1083]. Staats [3905]. Staats- [3905]. Staatsmann [62]. Stabile [3062]. stabilimenti [2468]. Staccioli [1237]. Stackfreed [1703]. Stadera [1831]. Stadere [560]. Stadt [1171]. Stafford [2120, 1143]. Stage [2655]. Stagecraft [2816]. Stages [2626]. Stahl [2578, 3571, 303, 2144, 3179, 3822]. stahliennes [2578]. Staley [3126]. Stampa [2173, 729, 1779, 2046, 3426, 88, 165]. stampadella [1314]. Stampatori [607]. stampe [309]. stamperia [1794]. stamps [1026]. Stancati [2428]. Stand [364]. Standardisation [2304]. Standards [2406]. Stands [2977, 3974]. Stanford [2227, 551, 2920, 2970, 4090]. Stanislao [453, 764]. Stanley [1859, 2959]. Stansfield [16]. Stanze [4044, 1923]. Star [1885]. Stargazer [2549]. Starkey [2834]. Starobinski [2052]. Starry [1118]. Stars [511, 1357, 3730, 1743, 1698, 3278, 3756, 3833]. Start [3577]. Statale [1258, 4076]. State [918, 1362, 1910, 2312, 2981, 3254, 549, 1970, 3905, 2534, 4076]. Statecraft [3535]. Statements [2131]. States [3210, 1800, 3813, 3921]. stati [3813, 1806]. statica [2327, 2278]. Statistical [1412, 2792]. statistico [1239]. Statistique [2267]. Stato [1970, 2969, 1823, 2202, 237, 844, 1158, 1665, 2789]. Statua [1044]. Statues [3215, 1044, 3654]. Status [2388, 2993]. Staubermann [3289]. stay [3932]. Steam [2607, 2604]. Stearn [1314].

Stearns [439]. Stedall [2717]. Steding [123]. Steele [2731]. Steenbrugis [1685]. Stefan [2785, 2576, 62, 1941]. Stefani [912, 1437]. Stefania [1252]. Stefano [2330, 2592, 304, 2132, 2281, 949, 1570, 687]. Stefanos [3482]. Steffan [2866]. Steffens [1731]. Steger [2750]. Steila [1691]. Steiner [242, 296, 567, 1531, 2055, 3116, 3123, 23, 1941, 2978, 173, 3288, 2859, 217, 1838, 1285, 1396, 1501, 1153, 1086, 1912, 751, 376, 1069, 609, 3073]. Steinheil [2403]. Steinke [2529, 2732]. Steinle [2859]. Stella [1903, 2155, 1481]. Stellare [2290]. stelle [1851]. Stem [3226]. stems [3905]. Sten [1684]. Stendardo [2226]. Stengers [1899, 1998]. Steno [2870]. Stenoniana [1363]. Stensen [3524]. Stephan [64, 2529, 150, 3037]. Stéphane [2952, 2652, 3026]. Stephen [3204, 1662, 1108, 2777, 1738, 349, 3281, 2947, 1654]. Stephenson [2152]. Steps [2408, 421]. Stereocartografo [1781]. stereocartograph [1781]. Stereotypes [177]. sterilisering [972]. sterilization [972]. Stern [2061, 363]. Sternkunst [542]. Sterpellone [640, 877, 1026, 1434]. Stevan [496]. Steve [2740]. Steven [2707, 2093, 3441, 1411, 168, 1404, 3016, 1533, 1722, 998]. Stevens [990, 1278]. Stewart [3674, 693, 1109, 2971, 209]. Stiaffini [1695]. Stichweh [415]. Stiftung [3116, 173]. Stile [1380]. Stilgeschichte [3129]. Still [2653]. Stillman [968]. Stimson [424]. Stinche [952]. Stirpium [2531]. Stirtz [130]. Stockholm [1078, 1019, 1023, 1146, 1537, 1805, 2119, 2200, 696, 699, 3092, 1379]. Stocking [229, 345]. Stoeger [549]. Stoffel [2165, 2236, 1859, 2469, 1734]. Stoizismus [798]. Stokes [808, 4032]. Stolberg [1405]. Stoltzenberg [1336]. Stolzenberg [2227]. stomaco [871]. Stone [3215, 732]. Stones [3520]. Stood [2653]. storage [4039]. stored [2248]. Storia [583, 794, 773, 781, 852, 1091, 1017, 1691, 46, 521, 1035, 1205, 1374, 1463, 2169, 2240, 2333, 3099, 2411, 1221, 594, 1902, 597, 18, 621, 677, 1316, 1679, 1711, 1693, 1815, 2667, 989, 3279, 72, 447, 448, 716, 910, 3001, 783, 1345, 2248, 2963, 2619, 1243, 1516, 1499, 1560, 2010, 2224, 3408, 2697, 3397, 3812, 1477, 2720, 599, 909, 758, 1561, 1979, 2261, 360, 503, 639, 566, 789, 785, 947, 872, 1027, 1096, 1092, 1030, 1178, 1175, 1431, 1674, 1675, 1756, 1677, 1676, 1865, 1809, 1864, 2006, 2072, 2430, 14, 432, 554, 708, 876, 875, 391, 47]. Storia [90, 71, 261, 354, 291, 294, 458, 431, 518, 657, 575, 635, 634, 731, 813, 730, 705, 700, 779, 973, 878, 1116, 1056, 1292, 1430, 1423, 1665, 1813, 2078, 2113, 1897, 827, 1904, 2445, 203, 1075, 1394, 1856, 83, 1068, 1022, 1349, 1348, 373, 555, 3095, 1291, 2100, 1802, 739, 1651, 627, 2266, 2748, 2088, 1831, 61, 517, 764, 894, 2881, 2163, 2234, 2386, 1903, 4044, 532, 1314, 1648, 2471, 3772, 2543, 2631, 2313, 2376, 2523, 309, 915, 1105, 996, 1869, 1820, 2232, 3209, 1111, 1335, 3818, 3847, 80, 81, 300, 2140, 881, 1183]. storia [436, 437, 1716, 1752, 710, 711, 871, 3925, 200, 631, 556, 832, 1020, 1018, 1025, 1255, 1165, 1354, 1452, 1435, 1768, 1863, 2005, 1923, 2217, 999, 2624, 1712, 680, 3956, 1579, 2132, 3181, 3973, 3710, 687, 957, 1713, 1796, 2274, 3986, 3987, 3988, 3190, 3655, 4010, 1003, 2433, 2810, 4019, 2322, 831, 4038, 2135, 1169, 3094, 2064, 2073, 2431, 2780, 1164, 2557, 1901, 638, 1866, 1810, 564, 1718, 1871, 2110, 2113, 2307, 2107, 2627, 643, 2630, 2284, 971, 1420, 1223]. Storia-Geofisica-Ambiente [779]. Storica [517, 804, 236, 389, 297, 357, 39, 617, 764, 1427, 697, 238, 2618, 1491]. Storiche [1648, 1318, 449, 1291, 2622]. Storici [2507]. Storicità [189]. Storico [1054, 605, 2314, 38, 1173, 953, 1168, 2117, 2833, 1615, 3286, 2315, 1335, 3824, 1741, 407, 789, 3910, 876, 185, 1082, 1259, 1167, 1861, 1871, 1931, 2149, 2415, 2414, 558, 555, 2320, 1399]. storico-critica [1399]. storico-documentaria [1741]. storico-naturalistica [1871]. Storico-Scientifico [605, 1861, 1931, 2149, 2414]. Storico-Sociali [2117, 1615]. Storie [1824, 1598, 621, 2632, 1271, 2413, 3426]. Stories [2970]. Storiografia [910, 1096]. storiografico [915]. Storm [977]. Story [3124, 1589, 2527, 2864]. Stoye [1234]. strada [726]. Stradanus [3670]. strade [2720, 3314]. Strane [2507]. Strange [3617]. Strasbourg [836]. Strasser [3127]. Strategie [2092, 1178]. Strazzoni [3508, 3569]. stregone [1565]. Strehlow [948]. Strings [1783]. Strocchia [3567]. Stromenti [1634]. Strömfeld [363]. Strömfeld/Roter [363]. Strong [3468, 4026]. Strozzi [2181]. Strube [303, 497]. Structure [805, 3263, 2230, 3901, 4036]. Structures [3210, 2457]. Struensee [62]. struggle [2740, 2793]. Struggling [3191]. Struktur [1610]. Strukturen [1862, 4036]. Strumentaria [1892, 1509, 2871, 351, 73, 232, 292, 423, 501, 630, 775, 853, 936, 1010, 1241, 1254, 1341, 1583, 1742, 1860, 1929, 2000, 2210, 1826, 1779, 1972, 1664, 1160, 1807, 1860, 2147]. strumentazione [631]. Strumenti [69, 777, 894, 1221, 1054, 3138, 1470, 2313, 2376, 290, 350, 935, 1250, 1858, 396, 2871, 553, 1220, 2476, 2045, 179, 231, 1084, 1009, 1259, 1240, 1338, 1416, 1507, 1861, 1645, 1301, 4006, 2168, 2239, 3763, 75, 2523, 2974, 715, 719, 3847, 3856, 1260, 3885, 390, 3426, 776, 780, 782, 1011, 1257, 1456, 1343, 1666, 1745, 1863, 1931, 2112, 2110, 2149, 2415, 2414, 929, 4007, 1156, 1420, 1660, 778]. Struttura [321, 1550, 2334, 1159, 2285]. Strutture [1778, 1695]. struzzi [883]. Stuart [614, 2936]. Stuber [2529, 2785]. Student [691]. studentesse [3741]. Studenti [97, 159, 2460]. Students [3594]. ’Studi [1077, 516, 721, 894, 995, 1539, 2421, 46, 1963, 3120, 1648, 2933, 2522, 1790, 2667, 75, 2313, 2523, 2685, 312, 309, 446, 449, 720, 716, 887, 1105, 1360, 1637, 1499, 2224, 38, 613, 2461, 522, 1190, 2419, 71, 238, 291, 561, 727, 832, 1033, 1195, 1452, 1423, 1861, 2117, 2151, 2212, 2149, 2217, 2271, 2415, 1712, 799, 800, 1230, 1994, 44, 1455, 750, 1291, 1318, 1615, 1483, 2100, 1990, 3259, 197, 1691, 2467, 3748, 4043, 336, 2376, 85, 717, 915, 996, 1279, 3001, 1869, 752, 2492, 881, 2754, 2005, 2414, 838, 3967, 3655, 607, 1134]. Studi [2034]. Studia [1444, 1984, 2529, 2785, 1141, 2205]. Studiare [1827]. Studie [547, 787]. Studied [3716]. Studien [567, 1402, 29, 251, 443, 1359, 2196, 4078, 338, 535, 1610, 1396, 364, 380, 1905]. Studier [1537]. Studies [301, 805, 891, 1110, 1005, 1264, 1333, 1365, 1454, 1537, 1657, 2016, 2160, 2708, 772, 1568, 2891, 3122, 2558, 445, 648, 886, 830, 1278, 1188, 1358, 1443, 1686, 1147, 2156, 334, 3205, 2756, 1669, 547, 1575, 2681, 3024, 3017, 3159, 3191, 616, 1849, 184, 293, 916, 1570, 1750, 2098, 3224, 2552, 1229, 1980, 2099, 2998, 3044, 3058, 2404, 2937, 3278, 2498, 914, 1245, 685, 3748, 3759, 2737, 3773, 3774, 3775, 43, 4063, 4078, 2731, 541, 1989, 2436, 3005, 3867, 3776]. Studii [251]. Studio [2010, 909, 238, 2213, 658, 1469, 2460, 1325, 1214, 3783, 252, 1280, 572, 1101, 1677, 854, 2226, 999, 3329]. Studiolo [277]. Studley [2779]. Study [1910, 2607, 3570, 2462, 395, 274, 3440, 3661, 3049, 928, 3432, 1557, 1469, 997, 163, 3564, 689, 2180, 3239, 2534, 68, 3783, 787, 1274, 3946, 3979, 1086, 4024]. Studying [3876]. Stumpo [1521, 1728]. Sturlese [307, 309, 716]. Sturloni [2692]. Stuttgart [62, 296, 567, 1531, 1602, 2055, 3116, 3064, 3283, 623, 725, 1004, 3123, 23, 1941, 2978, 173, 172, 557, 1715, 3288, 1735, 343, 2859, 1714, 344, 622, 1731, 1838, 1285, 1396, 1618, 2734, 1797, 361, 362, 1501, 1153, 348, 1086, 1079, 1202, 1328, 1912, 2465, 751, 376, 1450, 1069, 609, 3073, 690, 565, 1905, 1916, 4082]. Stuttgart-Bad [3283, 2734, 1202, 2465]. Stuttgart-New [1004, 361, 362, 565]. Stuttgarter [1838, 4082]. Sua [2356, 734, 2614, 2630, 2773, 437, 641, 729, 1030, 1665, 3956, 811, 4019]. suan [1640]. Suardo [1456]. Suárez [2808]. subactis [1033]. Subalpina [2100, 2313, 2376]. Subbarayappa [329]. subject [769, 1989, 4021, 3446]. subjectivity [4021]. Subjects [3944]. Subjektivität. [343]. Subjekts [1989]. sublime [3029]. submitted [1476]. Subsidia [991, 1156, 1688, 70, 2372]. Subskript [1838]. Subskript.-Preis [1838]. Substance [3113]. Substances [2156, 4022]. Subventions [2615]. successives [3665]. Sudhoffs [1941, 1285]. Sue [3135, 858]. Suenjel [3583]. Suffering [3727]. Sufficient [3350]. Suffolk [1314, 1063]. Suffrage [3499]. Sugarcane [1641]. Suhrkamp [415]. Sui [1824, 3103, 1779, 2046, 2276]. suisses [2539]. Suivi [1232, 4089]. sujet [3013, 3890, 3980]. Sul [1221, 1778, 2449, 1782, 976, 1298, 1826, 1558, 1682, 309, 752, 1753, 635, 1619, 799, 800, 2384, 750, 2136, 1014]. Sulla [1951, 2022, 459, 162, 733, 1890, 744, 1833, 1502, 3748, 1949, 2020, 1963, 2034, 2332, 934, 456, 814, 2933, 1769, 250, 1077, 2283, 522, 1148, 2419, 3498, 3426, 656, 1159, 2009, 999, 1483]. sull’Alto [995]. Sulle [1780, 1663, 2191, 913, 1119, 1379, 995, 2167, 2238, 1523, 812, 2499, 1728]. sull’elet [1068]. sull’elet-tricit [1068]. sull’elettroshock [2855]. sull’ereditarietà [421]. Sull’Illustrazione [1549]. Sull’Insensibilita [211]. Sull’orlo [2384]. sull’ottica [684]. Sull’ottima [1174]. Sulloway [1581]. sull’uso [2141]. Sumer [1804]. Summa [3946, 997]. summary [1167]. Summer [328]. Sun [3392, 2096, 3423, 3514, 3304]. Sundial [3, 428]. Sundials [409, 998]. Sundin [1934]. Sunshine [2304]. sunspot [2178]. Sunspots [3169, 1119]. Sunto [764]. Suo [2049, 1769, 2588, 448, 1335, 2140, 789, 729, 876, 1690, 1665, 2112, 2077, 1653, 611, 4019]. suoi [2881, 1564, 1581]. suolo [2009]. Suoni [596, 2151]. Sup [2422]. super [717, 1359]. Superconduttori [627]. Supérieures [444]. Superintended [1408]. Superior [889, 1227, 1226, 2494, 494, 618]. Superiore [3080, 941, 1645]. Superiori [2191]. supernova [2136]. Superstition [3373]. Supper [1375]. suppl [2581]. Supplement [3815, 1407, 406]. Supplementary [2852, 2716]. Supplementi [3259]. Suppositiones [28]. Supramolecolare [2044]. Supramolecular [2044]. Surface [3667]. Surfaces [3348]. Surgeon [2709]. Surgeons [57, 3907]. Surgery [3006, 3857]. surgical [3871, 2556]. Surrealism [3038]. Surrounding [3505]. Survey [1617, 990, 659, 2009, 1636, 519, 424, 2747]. Surveying [3719, 489]. Surveyor [3271]. Surveyors [2991]. Surveys [964, 830, 1102, 408, 610, 378]. Surviving [3597, 424]. Susan [2671, 3267, 2964, 3110, 1652]. Susanna [891]. Susanne [1854]. Susskind [190]. Sutton [1763]. Suzanne [3458, 795]. Svante [3092, 1264, 2456]. SVEC [3024]. Sven [3526, 2972, 3253]. Svensk [2200, 3852, 3853, 3854]. Svenska [696]. svenskakulturdebatt [1146]. Sverige [972]. Sverker [1023]. Svezia [682, 740]. Sviluppo [2635, 811]. Svizzera [1869, 3788]. Swain [3650]. Swainston [133]. Swan [2657]. Sweden [972, 891, 1537, 495, 1481]. Swedenborg [725, 1611, 1988]. swedenborgarnas [1537]. Swedenborgs [1537]. Swedish [972, 1019, 1264, 1146, 1537, 1805, 2496, 3471, 495, 696, 3852, 3853, 3854, 3470]. Swerve [3281, 3870]. Swets [802, 1367]. Switzerland [3788]. Sydney [341, 851, 388]. Sylla [2786]. sylloges [1930]. Sylvain [1228]. Sylvie [2150, 2892]. Symbolism [1725, 3841]. Symbolistes [795]. Symmetrien [498]. Symons [847]. Sympathetic [3841, 1725]. Symposion [1141]. Symposions [276]. Symposium [672, 63, 648, 2161, 4066, 3092, 1137, 2456, 4054, 444, 2195, 4058, 4055]. synapse [2678]. Synchrony [2937]. synoptic [3873]. synthèse [1799]. Synthesis [2505, 3215, 3102]. Synthetic [3447]. Syriac [1102]. System [420, 3148, 1654, 3789, 3790, 3795, 2248, 3999, 1486, 1381, 1381]. Systematic [1989]. systématique [3026]. Systematische [1989]. Système [2537, 1486, 3013, 3890, 3980, 3135]. Systems [3846, 415, 2505, 3999, 1375, 3858, 1987]. Szulakowska [2840].

T [2267, 2459, 1659, 1314, 3059, 1496, 2551, 2974, 1063, 767, 65, 2076, 3636, 163, 1198, 3429, 2591, 946, 1920, 2145, 1316, 1041, 2369, 753, 749]. T. [2467, 2134, 2220, 761, 203, 1834]. Tabarroni [1105]. tabelle [1992]. Tabellen [4062]. Table [3124, 1224]. Tables [1152, 3575, 4062, 1172, 3873, 34]. tablié [1859, 2013, 216]. Tabula [2475, 2572]. Tabulae [3049]. Tacchi [809]. Tacchini [1350]. taccuini [435]. Taddei [320]. Taddia [2603]. Tafeln [420]. Tagebuch [1854, 3978]. Tagliaferri [18, 71, 291, 631]. Tagliagambe [1200, 1870]. Taglianti [2612, 2151]. Tagliavini [178]. taglio [2009]. Tagliolini [1488]. Tagung [173, 343]. Tahiti [3725]. Taisbak [2418]. Taking [2745]. tal [2675]. Talbot [2354]. Tale [3601]. talets [1146, 2500]. Talismans [3453]. Talk [2684, 3537]. tallyah [2199]. Tamagno [563]. Tamassia [2148]. Tamburello [832]. Tamera [2731]. Taming [2435]. Tampellini [2597]. Tandon [2013]. Tang [3664]. Tanganelli [2614]. Tani [2269]. tänkande [1537]. Tansey [2909]. tape [3905]. Tappero [3085]. Tara [3068]. Tarabochia [715]. tarantismo [2780]. tarantola [3329]. Tardo [1785, 515, 1816, 2536, 2492, 3967, 377]. tardo-antico [377]. Tardo-Ellenistica [1785]. Targioni [313, 2252]. Taroni [1802]. tartīb [1226]. Tartaglia [3282, 4015]. Tartu [938]. Taschenbuch [2979]. Tascon [2876]. Tassinari [1483]. Tasso [1189]. Tatasciore [896]. Tatlock [2841]. Taton [1670]. Tatti [3278]. Taub [3438]. Taube [113, 104]. Taurisano [2492, 3967]. Taussig [2892]. tav [424, 1369, 1766, 840, 867]. Tavistock [225]. Tavole [3873, 305, 2214, 407, 1172, 1614, 1230, 1658, 1135, 1177, 1641]. tavolo [3673]. Tavoni [2077, 2954]. tavv [1461, 1836, 2054, 2142, 917]. taw [2163, 2234, 2389, 2386]. Tazzioli [2215]. Teacher [3486]. Teaching [3509, 3734, 3503, 381, 4045, 169, 4069, 1256, 2213, 2423, 3966, 3288]. teatro [1314, 74, 2225, 3989, 2796]. Teatro. [2370]. Tebtynis [3484]. Tec [2118]. Tech [348]. Teche [3685]. Techne [782]. Technical [2607, 3047, 1778, 3503, 1249, 3969, 2054, 3973, 4085]. Technik [296, 676, 1608, 571, 798, 420, 1735, 2196, 4078, 1079, 2466, 1450, 695, 1916, 1735, 2196]. Technik.- [1735]. Technikentwicklung [3968, 2857]. Techniker [303]. Technikgeschichte [1608]. Technique [675, 1795, 1612, 2770]. Techniques [1327, 702, 728, 1060, 3280, 243, 836, 1442, 1985, 2229, 3876, 1080, 2539, 1224, 4042, 295]. Technische [1340, 3129]. Technischer [411, 1001]. Techno [2762]. Techno-Science [2762]. Technological [2981, 3210, 3382]. Technologie [2725]. Technologies [3606, 3507, 3903, 3035]. Technology [2016, 3187, 2607, 2168, 2239, 3764, 1387, 2990, 847, 3369, 2510, 2553, 1830, 1063, 3810, 1164, 3289, 2928, 1479, 1484, 1641, 1299, 1944, 1126, 1557, 293, 916, 2069, 3338, 3083, 3968, 3452, 4046, 676, 1608, 182, 1758, 2381, 3807, 1735, 3916, 1544, 2009, 4080, 2745, 2731, 2749, 1237, 1694, 1109, 3996, 2688, 2874, 4078, 2753, 2804, 3452, 2076, 1071]. Tecnica [781, 1017, 2168, 2239, 1637, 1757, 986, 428, 1482, 1544, 2383, 3807, 1168, 999, 3314, 4085, 1390, 2416, 1470, 2808, 2150]. Técnicas [1544, 2876]. Tecniche [2490, 2431, 2536, 999]. Tecnico [1778, 1512, 1768, 3969, 2002, 3973, 2232]. Tecnico-Culturale [1778]. tecnico-scientifico [2002]. Tecnologia [2008, 1164, 634, 3314, 1889, 1314, 1018, 1165, 2009, 304, 1851, 1482, 3094]. tecnologica [1237]. Tecnologico [2313, 2376, 3021]. tecnologie [185]. tecnologo [3402]. Tecnology [184]. tedesca [716, 687]. Teeter [1002]. Teeth [707, 1154, 4009]. Teeuwen [2986]. Tega [257]. Tegni [274]. Teich [1433]. Teichmann [1153]. Teil [251, 571, 798, 994, 2831, 420, 1391]. Teilband [571, 798]. Teknos [1682, 1803, 1870]. Telegraphy [3764, 2381]. Teleologia [2272, 3798]. teleology [3798]. Telescope [2293, 1217, 1548, 1886, 2393, 3422, 3423, 3258, 1014, 1469, 1704, 3953, 3423, 3308, 2549]. Telescopes [462, 1584, 2890]. Telescopic [3248]. telescopio [1014, 1836]. Telesio [721, 838]. Television [3152, 3797]. televisive [185]. Telle [567, 2831, 1396]. Telling [1562]. Tema [1376, 1619]. temere [2430, 4010]. Temi [1899, 1815]. Temperature [2681, 3244, 3423, 3423]. Tempi [739, 447, 3362]. ’Tempio [2796, 3209]. Tempo [321, 3209, 3620, 428, 1620, 1343, 1512, 1690, 2153, 2548, 1265, 578, 1653, 1998, 4084, 942, 2057, 1900, 2588, 2112]. Temporally [3635]. temps [3767, 701, 3010, 3330, 2782]. Tenenbaum [2374]. Tennessee [438]. Tennis [3228, 3230, 3232, 3235, 3233, 2786]. Tenon [1911]. Tentative [3292]. tentazione [765]. tenu [1090, 1734, 1224]. tenuti [1619]. tenuto [4040]. Teofilo [3258, 3953]. teologia [85, 2720, 3178]. Teologische [251]. teorema [2108]. Teoria [2233, 2162, 1701, 1550, 2041, 979, 2258, 733, 1971, 744, 2135, 1949, 2020, 1769, 1693, 2321, 656, 1321, 1802]. teorica [1737]. teorici [1815]. Teorie [1200, 1221, 906, 2651, 1788, 2431, 1499, 3426, 779, 811, 2502]. Teramo [70]. terapeutica [1855]. Terapia [639, 869, 1810, 564, 3179]. Terenna [1260, 1861, 1931, 2149, 2414]. Teresa [2387, 3570, 221, 1712, 1788, 2903, 2902, 2205, 588]. Teresiana [1554]. Terme [197, 1766, 626]. Termini [1173, 379, 28, 30]. terminologia [2865, 3174, 1943]. Terminologie [306]. termologia [75]. Termometri [1807]. Termometro [2173]. termoscopi [1807]. Terms [28]. Terni [70]. Terra [740, 739, 1111, 656, 858, 2001, 186]. Terrada [2101, 2097]. terrain [1793]. Terram [51]. Terre [563, 410]. terremoto [39, 199, 1430]. Terrentius [1873]. territoires [314]. territori [1373]. territoriale [584]. Territories [3091]. Territorio [804, 1970, 1871, 635]. Territory [1970, 635]. Terrors [3310]. Terse [3139]. Tertius [2520]. Terzo [2370, 3457]. Tesi [2298]. Tesnière [3657]. Tessicini [2951]. Test [2851, 319]. Testi [1539, 1145, 1648, 720, 717, 1279, 752, 2461, 1453, 1369, 558, 922, 800, 1483, 1134, 1816, 223, 2667, 181, 85, 915, 996, 3001, 869, 1645, 1579, 2644, 1230]. testimonianze [709]. Testimony [2702]. Testo [2713, 1174, 2214, 930, 1992, 2004, 3277, 1394, 2644]. testuale [1943]. Teubner [303, 568]. Teutonicorum [307]. Texas [2824]. Text [2418, 2085, 334, 1048, 1419, 1049, 3298, 420, 4089]. Textbooks [2443]. Texte [2055, 584, 29, 23, 574, 2085, 893, 535, 216, 2576, 1086, 1912, 376, 1069, 2578, 1138, 754, 4089]. Textes [567, 548, 796, 512, 1442, 866, 3277, 2820, 516, 1444, 868]. Texts [2879, 3535, 964, 830, 1102, 1358, 2103, 408, 3070, 2781, 610, 3271, 806, 378, 1230, 3630]. Texture [2552]. Th [14]. Thabit [216]. theater [3989]. Theatre [3588]. Theatrum [1544]. Their [490, 3558, 2361, 1664, 3922, 2382, 3554, 2180, 63, 3122, 34, 190, 4078, 184, 1392, 3448]. Themes [419]. Theodicy [267]. Theodolites [3504]. Theodor [1239]. Theodore [3268, 1655, 2792, 691]. Theologia [3283, 4026]. Theologian [3025]. Theologische [217]. Theology [2388, 3162, 549, 1103, 1225, 1920, 4026, 3889, 624, 335, 3016, 2207]. Theophrast [1396, 217]. Theophrastus [945, 2939, 945, 1225]. theorems [1089]. Theoretical [2606, 1500, 3462]. Theoria [569, 532, 504, 1052, 1414]. Theorie [507, 2199, 2578, 1667, 3930]. Theories [3758, 1702, 3122, 551, 3189, 3432, 3140, 2457, 692, 3991, 1825, 1199, 1446, 1499, 2366]. Theorist [3122]. Theory [1657, 3373, 443, 1701, 1830, 1222, 2978, 2084, 3656, 927, 3503, 2258, 2761, 2583, 1247, 3429, 388, 946, 2971, 1043, 204, 1936, 3930, 656, 1042, 1977, 3837, 2505, 419, 1503]. Therapeutic [1097, 1800, 3921]. Therapie [1724, 639, 950, 676]. therapy [676, 1810, 3938]. There [3467, 2433]. Therein [36]. Theriac [3563]. Thériaque [3611]. Thermodynamik [2854]. Thermometer [3244]. Thermometers [2294]. Thevenot [820]. Thick [2552]. Thief [3683, 1589]. Thierry [707, 1432, 2729]. Things [3376, 1945, 3486, 3537, 3339, 1198, 2684]. Think [2602, 2787, 3804, 2688]. Thinker [2724, 1211]. Thinking [3602, 1717, 3554, 3445, 3272, 2825, 3649]. Thinks [3383]. Third [3293, 3037]. Thirteenth [3565]. Thirteenth-Century [3565]. thlith [2199]. Thologie [598]. Thomas [1264, 126, 1686, 1401, 1918, 1915, 1532, 1794, 3600, 362, 3711, 16, 3762, 1151, 2185, 3247, 1550, 1708, 3884, 1915, 645, 1610, 2734, 2132, 2465, 2687, 2688]. Thomasset [33]. Thomaz [3475]. Thomson [3074, 534, 3210, 808, 4032]. Thorbecke [1402]. Thorburn [3091]. Thoren [888]. Thought [1108, 3016, 1188, 1685, 1686, 1588, 925, 3668, 1229, 1825, 1983, 1002, 1537, 1121, 3074, 770, 4025, 624]. Thought& [1225]. Thoughts [3151, 2977, 3974]. Thouin [1418]. Thousand [3852, 3853, 3854]. thread [1540]. Threat [2990]. Three [2388, 2998, 1654, 1705]. threshold [1237]. Thrift [3131]. Thurner [3634]. tiempos [45]. Tierisch [3344]. Tierwelt [3123]. Tiggelen [3091]. Tihon [1669]. till [2119]. Timber [1213]. Time [2899, 1686, 349, 3620, 3191, 3944, 163, 3219, 3478, 3239, 447, 45, 428, 1265, 4084, 2388, 3072, 3043, 163]. Timekeeping [3131, 2937]. Times [1234, 1589, 1161, 347, 1007, 2549, 2545, 3770, 3773, 3774, 3775, 3776, 3843, 3928, 1877, 3950, 803, 2156, 2907, 3857]. Timpanaro [1064]. Tina [330]. Tipografia [514, 1260, 710, 1813, 1807, 1817]. tipografici [576]. tipografo [1283]. tipologia [1752]. Tisci [3367]. Tit [639, 948, 882]. Titelbilderund [2837]. Title [3641, 3659, 3677, 3697, 3720, 3743]. Tito [3285, 3799]. Titolo [638, 566]. Tiziano [3277, 313]. Toaldo [2330, 3450, 525]. Tobias [1154]. Toby [1261]. Toccafondi [1324]. Toccare [1851]. Today [3502, 2364]. Todd [3482, 2516]. Todesco [88]. Todorovic [2069]. Tofani [1343]. Toga [2247]. Together [2786, 2525, 1225]. Tognotti [2430]. Toil [2899]. Tokugawa [3478]. Tokyo [3081]. tolerance [573]. Tolkningar [2500]. tolleranza [573, 3065]. Tolomeo [2212, 1785]. Tolstoj [2157, 3927]. Tolstoy [3927]. Tomàs [889]. Tomasi [2012, 2530, 863, 917, 1614, 2886, 1269]. Tomaso [1052]. Tome [675, 243, 2539, 3004, 3163]. tomi [621, 252, 186, 2107]. Tomlinson [1106]. Tommasi [2620, 2592, 2667]. Tommaso [715, 1639, 1775, 1782, 546, 2202]. Tomo [2370, 12]. Tomory [3265]. Tomus [307, 1280]. Tonelli [338]. Tongiorgi [2012, 2530, 863, 917, 1614, 2886, 1269]. Tonköpfen [3905]. Tony [3239]. Tool [2050]. Tools [3360, 3421, 778, 2345, 2168, 2239, 3847, 3856, 4006, 4007, 4063]. Topham [2737, 3637, 2004]. Topici [1606]. Topics [1104]. topo [1922]. Topografica [44]. Topography [3479]. Tore [1005, 672, 685, 1146, 1454, 2200, 495]. torinesi [556]. Torino [39, 537, 617, 1427, 1364, 1528, 1539, 1606, 1691, 2467, 1921, 934, 2063, 2074, 2472, 2313, 2314, 2376, 2670, 442, 2963, 2010, 2232, 2535, 2632, 2893, 2690, 2630, 2904, 67, 758, 407, 1438, 1716, 2413, 3546, 432, 1093, 1029, 2693, 429, 563, 556, 778, 1020, 1018, 1162, 1185, 1165, 1599, 2111, 2150, 558, 1904, 756, 577, 186, 2426, 2738, 3085, 1658, 1922, 1899, 1998, 83, 2384, 1140, 765, 1244, 1349, 1330, 1426, 1348, 2919, 1420, 1642, 2100, 1644, 745, 2910, 346, 2748, 1834, 2010, 2092, 1780, 3031, 1162, 83]. tormentata [80]. Tornitore [203]. Toronto [491, 968, 2953, 2873, 2139]. Toronto-Buffalo-London [968, 2139]. Toronto/Buffalo/London [2953]. torpedini [2904, 3955]. torpedoes [3955]. Torquato [1189]. torre [2148]. Torriani [3456]. Torricell [1469]. Torricelli [1062, 1325, 1037, 1469]. Torrini [2337, 2389, 2288, 1371]. Torsten [1079, 1912]. Tort [1073]. Torti [955]. Torzini [2159]. Tosca [1186]. Toscana [313, 516, 517, 804, 340, 688, 2797, 716, 1077, 1538, 710, 1186, 389, 589, 590, 575, 844, 1695, 2363, 1267, 1324, 887, 1252, 1000]. Toscana-Marsilio [1324]. Toscanelli [899]. Toscani [2115, 1696]. Toscani-Santa [2115]. Toscano [2168, 2239, 1295, 2665, 1025, 346]. Tosi [2988, 2623, 863, 917, 1614]. tot [63, 1791]. Totaro [1943]. Totelin [3733]. Touching [3728, 3722, 3724]. Toulouse [2638]. Tour [1158]. tournant [920]. Tournhout [2878]. Tours [4077]. Tous [1804]. Toussaint [872]. Toussaint-Samat [872]. Touwaide [2819]. Tower [36]. Tozzetti [313, 2252]. Tozzi [2002]. Tr [1920, 2881, 1277, 706, 811, 2324, 2433, 2503]. Tra [1221, 1295, 989, 1118, 887, 817, 1823, 2949, 2398, 3103, 2395, 3718, 2258, 3135, 626, 660, 1889, 1378, 818, 2159, 3299, 897, 2167, 2238, 3760, 688, 4044, 585, 4049, 2376, 3279, 2523, 2849, 448, 3544, 2315, 1111, 3826, 2632, 2630, 81, 3904, 877, 1925, 655, 778, 1923, 2226, 3179, 255, 3956, 611, 2384, 1233, 2100, 1331, 1519, 2316]. Trabucco [3687]. tracce [1663]. Traces [3000, 3536]. Track [2764]. Tract [3559, 1550]. Tractato [1050]. Tractatus [3002, 2140]. Trad [1533, 677, 2631, 1273, 1248, 80, 67, 3277, 1639, 1922, 1899, 1998]. Trade [916, 2310, 4013]. Traders [2312, 4034]. Tradition [837, 2909, 2441, 263, 2305, 3349, 258, 3059, 2648, 3906, 993]. Traditions [2950, 845, 3734, 2812, 1102]. Tradizione [898, 2041, 1975, 258, 1680, 1811, 2517, 2721, 165]. tradotto [2197]. Traducció [353]. traducción [1226]. traduction [1138, 1668, 681, 1647]. Traductionintégrale [1432]. traduit [216, 2578]. Traduzione [1528, 1145, 348, 612, 1417, 2548, 304, 1994, 2499, 1064, 1483, 1174, 2713]. Traina [1390]. Training [3230, 1277, 4049]. Trait [1090]. traité [1187]. traités [3058]. Traits [1040]. Trajectories [3601]. Traktat [1531, 379]. Tramite [319]. Tramontano [3040, 3820]. Tramonti [2006]. Transaction [811]. Transactions [2495, 1193, 2586]. Transazioni [2429]. Transcribed [2873]. transfer [2749]. Transferring [3501]. transfigured [683]. Transformation [2917, 3441, 3173, 2407, 2825, 3273, 891, 1721, 762]. Transformations [1169, 394, 635]. Transformed [3274]. Transfusion [3573, 3617]. transl [3011, 3037]. Translated [2822, 249, 955, 791, 2418, 1944, 3267, 945, 1225]. Translatio [1280]. Translation [2516, 28, 58, 3846, 2139, 3731, 49, 997, 376, 609, 1611, 3205, 3298, 3858, 1987, 763, 2525, 3946, 749, 1133, 2518, 3599]. Translations [2103, 3893]. Translator [3450, 2646]. Transmutations [3598].

Transmuted [3793, 3796, 1654]. Transnational [3471, 3248]. Transparency [3372]. Transpiration [3200]. transplants [1092]. Trapani [70]. trapianti [1092]. trasferimento [3021]. trasformazione [3915]. trasformazioni [3255, 635]. trasgressivo [2916]. traslazione [67]. trasmissione [2933]. Trasporto [1300]. Tratado [2356, 2357, 1226]. Tratados [352]. Trattati [2722, 1000, 562]. trattatista [3015]. Trattato [3230, 1098, 1523, 1483, 1728, 1502]. trattazione [2167, 2238]. tratte [1452]. tratti [1003]. Traugott [1151]. Travaglio [2131]. travail [1444]. Travaux [2145, 584, 834, 1060, 2894, 3277, 753, 749, 2267, 1155]. Travel [2302, 1566, 2363, 3170, 1476, 3965]. Traveler [2846]. Travellers [2702, 1305]. Travelling [3322]. Travels [3828, 1019, 4051, 847, 840]. Travi [2474]. Traweek [1980]. Tre [1705, 1041, 1061]. Treasures [2184]. Treasuries [1422]. Treatise [3436, 3731]. Treatises [28, 2054]. treatment [1724]. Treaty [2356, 2357]. Trecento [995]. Tree [3603, 371]. Trees [3958, 3268]. trefpunt [280]. Trends [2994]. Trenta [2616, 3839]. Trent’Anni [1547, 1710, 1737]. Trentina [2009]. Trentino [2009, 3747]. Trento [3065, 824, 1620, 1928, 2315, 3844, 1794, 2009, 1840, 1796, 1318, 1928]. tres [1229, 2712]. Treviso [2612]. Trew [3116]. Triads [2913]. Trial [613, 1042]. Trials [3734, 2966]. Triarico [2436]. tricentenaire [540]. tricit [1068]. Tricky [3577]. Trieste [448, 1237, 1737]. Trifogli [2520]. triginta [1280]. Trigonometrisch [751]. Trigonometrisch-astronomisches [751]. Trinchero [3742]. Trinchese [504]. trinit [715]. Trinity [460, 2360]. triomphe [3004]. trionfo [2797]. trip [3932]. Tripode [896]. Trismegisto [1280]. Triton [3403]. Triumph [2899, 3563]. Triune [3381]. Trivulzio [1675]. Trkulja [2012]. Trocchio [299, 1112, 337]. Trofast [1441, 884]. trois [2214, 609]. Troisième [1410]. Tromometri [779]. Troncon [1796]. Tropical [2845]. Tropp [1203]. Troubled [80]. Troublesome [2595]. Trowbridge [843, 1085]. Troxler [172]. True [3351, 2802, 2433]. Trumereien [3186]. Trust [414, 1655]. Trust/ [414]. Trustees [1872]. Truth [1404, 2606, 1401]. Truths [1231]. Tryonia [763]. Tube [235]. Tübingen [2831, 895, 1684]. Tubus [1826]. Tucci [3497, 71, 291, 452, 631, 1745, 1813, 2005, 2217]. Tucker [2794]. Tucson [592]. tude [512]. tudes [2613]. tuenda [3057]. Tugnoli [1620, 1325]. Tulipmania. [2887]. Tullio [2865, 996, 3001]. Tumhout [2309]. Tumor [3601, 3676]. tumori [1677]. tuo [1672]. Tuplin [2459]. Turchetti [2918, 3512]. Turco [1000]. Turin [2308, 429, 1830, 3503, 2817, 3101]. Turing [3417]. turismo [1428]. Turistica [1237]. Turn [3565]. Turner [1393, 3237, 2757, 767, 839, 1057, 762, 835, 998, 1563, 2310, 2541, 1420]. Turnholt [1280]. Turnhout [2145, 2267, 2770, 3526, 1273, 1444, 1685, 1984, 2596, 2615, 2369, 3284, 3026]. Turriano [965, 2876]. Tuscan [2239]. Tuscany [2363, 1267, 14]. Tusciae [1324]. Tusculanum [2418, 2658]. tusen [2200, 3852, 3853, 3854]. tutte [3388]. tutti [3181]. TV [2381]. Twee [1342]. Twelfth [3408, 2998, 378, 3783]. Twelfth-Century [2998]. Twentieth [544, 3081, 2105, 3511, 2470, 419, 1618, 3513, 3744, 3783, 3149, 3900, 3949]. Twentieth-Century [3511, 419]. Twenty [2392, 3565, 2551, 3523]. Twenty-First-Century [3565, 3523]. Twenty-Five [2392]. twenty-one [2551]. Twilight [1047]. Twist [3379]. Two [2354, 3200, 2346, 2671, 2510, 3048, 3601, 3507, 1342, 3605, 2702, 2231, 1654, 2198, 1225, 2248, 226, 420, 1464, 3763, 3469, 3885]. Two-Culture [2671]. Two-Hundred-Fold [3605]. Two-Legged [3048]. Tyacke [539]. Tyana [3453]. Tycho [888]. Tychonico [2167, 2238]. Tydén [972]. type [3894, 3576]. Typologie [1273]. Tytler [654].

U [1370, 606, 845, 2006, 83, 690]. U.S [1617]. U.S. [3451]. U.S.A [490, 3000, 1652]. Überblick [251]. übers [379]. Übersetzt [1391]. Übersetzungen [3893, 2576]. Übersicht [557, 67]. Uberto [2475]. Übertragungsneurosen [67]. Überwelt [725]. Ubrizsy [2421]. Uccelli [1331]. Ucea [1525]. Udine [70]. Udo [567]. Ufficio [844, 2363]. Uffizi [2338]. Ugo [3039, 412, 1033, 2855]. Uguaglia [2722]. uguali [1163]. uguali/equal [1163]. uguali/equal-arm [1163]. Uhren [1212, 1012]. Uitgegeven [802, 3109, 541]. Uitgeverij [3109]. UK [1085, 1933, 2625, 2683, 1993]. UK£ [606]. Ulf [2978, 3036]. Ülikooli [938]. Ulisse [2882, 2225, 3989]. Ullmann [698]. Ulrich [3073]. Ultima [400]. Ulugh [2760]. Ulyssis [1378]. Umana [2326, 2277, 1948, 2019, 3751, 588, 809, 365, 2628, 371]. umana/The [2628]. umane [1093]. Umanesimo [1761, 1360, 1754]. umani [3804, 3693, 1856, 2323, 3718]. umanista [887, 3326]. Umanisti [750]. umanistici [799]. Umanit [223]. umano [453, 1506, 2276]. Umberto [407, 563, 556, 558, 1579, 1658, 1348]. Umbri [635]. Umbruchs [2466]. Umweltkonflikte [1405]. Unacceptables [2231]. Un’Analisi [975, 1711, 1399]. un’antica [640]. Un’Aporia [319]. unbekannter [787]. Uncertain [3832, 2066]. uncleanliness [3937]. Uncommon [3688]. unconscious [1086]. Unconventional [2594, 3488]. Unconvincing [2407]. Underground [3711]. Underlying [3435]. Understanding [3175, 549, 2873, 2872, 2264]. Uneasy [440]. Un’Edizione [602]. UNESCO [2131]. Unfer [2787]. Unfinished [2139]. Ungern [1019]. Ungheresi [2421, 3976]. Ungheria [2421, 3976]. Unha [2068, 3990]. Único [1744]. Unicopli [496, 3121, 452, 631]. Unification [768, 1020, 3752]. Unificato [2258]. Unified [2258]. Unifying [2457]. unigenito [2313]. Unimaginable [3686]. Un’Infanzia [979]. Un’introduzione [612, 2618]. Union [669, 823, 907, 1478, 1556, 1635, 55, 271, 326]. Unione [720, 717, 991, 1279, 1156, 1688, 70]. Unit [2789, 1381]. Unità [3089, 3752, 2741]. Unitaire [1888]. Unitaria [1778]. Unitarian [1778]. Unite [1737]. United [992, 3210, 1800, 3921]. Unity [2388]. Univ [3108, 2152, 1274, 2626]. Universal [511]. Universale [2471, 755, 1639, 1330, 493, 1064, 1437, 3632]. universalia [717]. Universalis [3174]. Universalism [68]. Universalwissenschaft [1610]. Universe [1346, 3084, 551, 1342, 1203, 1038, 2152, 2077, 1114, 4012, 3086, 3997, 2625, 1896]. Universelle [2814, 280]. universes [4059]. universi [4059]. Universidad [333]. Universit [1499, 1879, 203]. Universita [2078, 2498, 1691, 2467, 4043, 75, 2313, 2376, 2176, 1516, 2224, 38, 2642, 727, 1033, 2005, 2117, 2078, 2151, 2212, 2217, 2271, 1712, 1230, 2460, 2558, 2421, 1648, 75, 72, 448, 21, 71, 291, 575, 832, 1452, 1350, 1373, 1423, 1619, 2415, 1318, 1615]. Universitad [2097]. Universitaire [2550, 1875, 2519]. Universitaires [1578, 260, 33, 1543, 1878, 920, 2060, 2663, 1073, 582, 1486, 1534, 3280]. Universitaria [809, 2140, 1343, 1591, 1666, 2113, 3065, 252, 2100]. Universitario [1869, 14, 832, 1873]. Universitas [251]. Universitat [2114, 2099, 3019, 251, 1074, 296, 1687, 3981]. Universitatis [1055]. Universitäts [248]. Universitäts-Schriften [248]. Universitätsbibliothek [173, 3905]. Universitätsverlag [1687]. Université [2878, 32, 1442, 2375, 1224, 1224]. Universiteé [512]. Universiteit [3205]. Universiteits [280, 541]. Universités [1984, 306, 1442, 1444]. Universitet [2496, 673]. Universitetsbiblioteket [1218]. Universitetstryckeriet [2496]. Universities [1445, 3487, 2584, 381]. Universitria [45]. Universitt [2915]. University [15, 60, 167, 227, 164, 168, 283, 335, 441, 520, 491, 439, 440, 438, 587, 592, 801, 759, 679, 805, 685, 837, 918, 923, 959, 968, 888, 1002, 1005, 1106, 1055, 1108, 1199, 1231, 1288, 1234, 1139, 1194, 1261, 1333, 1340, 1313, 1362, 1457, 1449, 1404, 1535, 1536, 1617, 1562, 1587, 1683, 1649, 1725, 1818, 1846, 1910, 2016, 2075, 2102, 2056, 2227, 2312, 2424, 2463, 2464, 2560, 2715, 2776, 2793, 2887, 2953, 2459, 2516, 2879, 2717, 2607, 1852, 2889, 3113, 3543, 1736, 2606, 1412, 1576, 1573, 1659, 228, 2669, 2737, 533, 703, 846, 1643, 2538, 2936, 3063, 2852, 2851, 1403]. University [1842, 3068, 3254, 1835, 166, 3072, 2805, 2964, 2093, 3025, 1738, 2959, 2850, 1991, 182, 2105, 2106, 3204, 3084, 3268, 2551, 2671, 2791, 2909, 3035, 3030, 58, 194, 249, 445, 648, 964, 1102, 1357, 279, 955, 1500, 3124, 349, 538, 551, 2103, 2371, 1758, 1670, 2066, 2311, 2589, 3076, 1588, 2824, 2920, 3263, 614, 1721, 2834, 2873, 334, 3066, 3256, 1152, 17, 34, 43, 56, 3043, 2657, 2874, 3003, 3012, 2655, 2774, 547, 2827, 2948, 2600, 2595, 2885, 3070, 385, 613, 2794, 2518, 2470, 3131, 2800, 1656, 1492, 169, 760, 769, 919, 933, 1246, 1409]. University [1395, 1575, 2947, 2747, 1398, 1722, 1987, 2377, 2783, 1997, 174, 3110, 286, 2139, 3067, 2950, 2681, 2718, 2862, 2061, 2561, 3033, 2709, 928, 1196, 1315, 1484, 1641, 1982, 2094, 2095, 2134, 2493, 2629, 3074, 422, 1841, 2970, 3734, 2977, 2912, 65, 191, 204, 285, 419, 654, 770, 683, 927, 850, 848, 1407, 1915, 2004, 2362, 2427, 2966, 2825, 1640, 3014, 2938, 3905, 1192, 2076, 2071, 3542, 3529, 1100, 1433, 1652, 3917, 2505, 2775, 2944, 3267, 2961, 2207, 2799, 25, 1843, 2660, 406, 938, 998, 1665, 2069, 2310, 3093, 1504, 1497, 1655]. University [2792, 610, 2858, 2686, 3038, 2584, 3210, 3078, 2976, 1319, 3008, 2676, 2844, 3118, 791, 2552, 2841, 2875, 753, 3083, 1249, 2678, 3709, 229, 163, 347, 1458, 1694, 2382, 2506, 2559, 2985, 3034, 2807, 2830, 2984, 1400, 1723, 2585, 1503, 1542, 2728, 2786, 1847, 2998, 2457, 2784, 2845, 2489, 3126, 3528, 766, 851, 1154, 1334, 1541, 2378, 689, 808, 806, 692, 693, 723, 803, 860, 849, 993, 1076, 1109, 1203, 1144, 1198, 1329, 1425, 1654, 3032, 2603, 2591, 1907, 3086, 582, 1720, 2861, 1999, 1337, 3261, 694, 381, 1837, 4090, 2587, 2582, 3278, 2901, 2691]. University [1986, 2846, 1592, 2752, 1252, 945, 3037, 2096, 2674, 2739, 3239, 2534, 1800, 1332, 1845, 1577, 1798, 177, 768, 2772, 1245, 2562, 2848, 2688, 1839, 2054, 2790, 3981, 2184, 2850, 1277, 2176, 3824, 4019]. universo [792, 1437, 2077]. Unknown [3200, 1972, 787]. Unpublished [1037, 1448, 2781, 3236, 3490, 1829, 2176, 3838, 2178, 4083, 4089]. Unruly [3531, 3532]. Unschuld [606, 845]. Unseen [1143]. Unternehmengründung [2857]. Unternehmensgründung [3968]. Untersuchungen [1202, 1989]. until [3902]. Untold [3293]. Unusual [3048]. unveiled [2152, 3997]. Unveiling [3196]. Unwanted [972]. Unwin [289, 2549, 2625]. Uomini [41, 2720, 2619, 1373, 3710]. Uomo [1710, 1088, 2156, 2070, 1922, 2384, 2432]. update [12]. upon [3339]. Upplysningen [1146]. Uppsala [685, 1005, 1055, 1264, 1365, 1454, 1657, 2496, 703, 1611, 1147, 1218]. Upsala [328]. Upsaliensis [1055]. Urania [1082, 1591]. Uraniae [1591]. Uraniborg [888]. Urbach [1854]. urban [4014, 3675]. Urbana [2776, 1926, 4014]. Urbanae [2085, 2085]. Urbania [1663]. urbinate [425]. Urbino [2522, 2321, 425, 581, 858, 1665, 3201, 4019]. Urbis [1926]. Urs [2529, 1764]. Ursprünge [2859]. URSS [240]. Ursula [2900]. Urszula [2840]. Urzeit [497]. US$34.95 [422]. US$50.00 [2362]. US$54.00 [606]. US$55.00 [2427]. US$60 [829]. US$80.00 [1991]. US$88 [1857]. USA [2160, 3041, 1898, 1911, 2553, 1993]. Usage [1475]. usages [3527, 1224]. Uschmann [420]. use [1829, 184]. Useful [1231, 1819, 2495, 3738, 819, 1305]. Users [2888, 1664]. Uses [2707, 2699, 2587, 4028]. Usi [2009]. Using [2485]. Usl [1259]. Uso [1829, 352]. ustorio [2328, 2279]. usus [4028]. Uta [23]. Utah [1196]. utensils [938]. Utet [1364, 1606, 1029, 3085, 1348, 2919]. Utgivna [1441, 884]. Utiles [1305]. Utilità [2476]. Utilitarianism [925]. utilité [3203]. Utility [2700, 3371]. Utopia [3316, 3993]. Utopian [167, 3522]. Utopianism [1108]. Utrecht [424]. utskott [1019].

V [714, 728, 1091, 1763, 2420, 2776, 814, 549, 444, 70, 38, 2662, 3412, 952, 1098, 1864, 171, 622, 237, 337, 1154, 329, 2921, 2581, 2100, 945, 1133, 1174, 1391, 1920, 2653, 2895, 2808, 547, 2642, 1732, 2660, 1065, 1348, 2627, 2968, 3808, 1973, 470, 163]. V-Vi [945]. V. [1098, 1179]. Vaccination [819]. Vaccinazione [1629, 2376, 1674]. vaccinazioni [785, 875]. Vaccinogeno [1784]. Vaetterli [138]. Vai [2007]. Vaiolo [1629, 3367]. vajuolo [2115]. Val [197]. valdarnese [2072]. Valdarno [1255]. Valdelsa [575]. Valencia [45, 43, 2101, 1849, 2114, 1855, 3019, 2097, 1849, 3019, 3867]. Valencia-C.S.I.C [2114]. Valencian [3759]. Valenciana [45, 2101]. valencianas [3759, 43]. Valencianos [43, 1849, 2114]. Valenstein [2678]. Valentin [1387]. Valentini [1324]. Valentino [1869, 1159]. Valeria [1603, 1574, 2543, 2104, 2910, 178]. Valérie [3657]. Valerio [238, 237, 233, 1926, 1237, 44]. Vallance [946]. Valle [1300, 736]. Vallecchi [31]. Valleriani [3282, 3224]. Vallicrosa [333]. Vallisneri [336, 3027, 2955, 922, 2608, 2902, 2726, 2960, 522, 3159]. Vallon [1080]. Vallori [2276, 2502]. valore [1289]. Valry [2878]. Valsalva [641]. Valturio [407]. Value [3602, 2361]. Values [1497]. Valutazione [645]. vana [1957, 2028]. Vandelli [726]. Vandenhoeck [966, 220, 542, 364]. Vandenhoek [2750]. Vangelista [585, 1678]. Vanhaelen [3477, 3654]. Vanin [2001]. Vanini [2492, 3967]. Vannevar [3402]. Vanni [552]. Vannoni [2798]. Vannozzi [1260, 1183, 1099, 712, 710, 697, 879, 1251, 1437, 1861, 1931, 2115, 2149, 2414, 1653, 1267]. Vanzan [2156, 1056, 1435]. Vaprio [1107]. varî [1000]. Varela [304, 371]. Varese [1331]. Vari [658]. Varia [223]. varia. [1619]. Variations [2812, 305]. Variazioni [1376]. Varied [3355]. varii [1955, 2026]. Variorum [301, 1393, 1568, 3023, 886, 1278, 1443, 1067, 1263, 1366, 2079, 1669, 1664, 835, 1320, 1563, 2098, 674, 3250]. Various [1225]. Värnamo [972]. Varni [1904]. Varro [2140]. Vasari [1955, 2026]. Vasoli [59]. Vassallo [2151]. Vasta [1020]. Vatican [549, 2379, 613, 1920]. Vaticana [1790, 281]. Vaticano [1790, 1279, 1291, 281]. Vattimo [2383]. Vauban [2664]. Vazquez [2146]. VCH [1095, 950, 1336]. VEB [698, 497, 420]. vec [2878]. Vecce [1950, 2021, 3613]. Vecchi [1778, 1675, 3614]. Vecchia [3138]. vecchiaia [958]. Vecchiarelli [2142, 949, 788]. Vecchio [1156]. Vedere [1816, 1052]. Védrine [1646, 1647]. Vedute [1794, 1926, 4014]. Veen [14]. Vega [648]. vegetale [1952, 2023]. Vegetti [1594]. Vegvar [2707]. Vein [3149]. Veith [2915]. Velasco [3554]. Veleno. [3206]. velenose [1268]. vello [2068, 3990]. Velmet [3692]. velo [2748]. Velocità [318]. Velsquez [2876]. Velut [2952]. Venere [2511]. Veneta [2679, 2490]. Venetiis [1421]. Veneto [2882, 1195, 12]. Venetus [58]. Venezia [794, 852, 315, 2332, 807, 1487, 2882, 2491, 2283, 980, 958, 730, 1435, 1064, 1565, 1324, 1540, 2257]. Veneziani [2809]. Veneziano [2174, 12, 1893]. Venice [2332, 3225, 1195, 12, 3341]. Vent’Anni [48]. Venti [425, 3839]. Ventrice [12]. Ventricles [3628]. Ventura [1759, 1158]. Venturi [1811, 668]. Venus [3133]. Venuses [3147]. Vera [3129, 3737, 1530, 2334, 2285, 2433]. Verardi [3443]. Veratti [3088]. Verbal [824]. Verbali [1251]. Verband [253, 492]. Verbiest [3205, 2493]. Vercillo [2428]. Verde [252]. Vergata [2335, 1065, 2286]. Vergognosa [3937, 2221]. Verhaltnis [1042, 1043, 547, 1989]. veridica [2631]. verit [718]. verità [2501, 1605]. vérité’ [3107]. Verkehr [362]. Verkennigen [541]. Verlag [62, 296, 567, 961, 895, 1402, 1531, 2647, 2055, 3116, 3064, 3002, 3283, 508, 497, 623, 543, 1405, 2517, 1004, 2682, 2837, 234, 2853, 2979, 2908, 3082, 3129, 3123, 30, 23, 58, 32, 251, 193, 307, 890, 1684, 1941, 1726, 1151, 2978, 2831, 1149, 173, 420, 1715, 2943, 3530, 3288, 3225, 1735, 3253, 67, 415, 2854, 2859, 1918, 1714, 217, 2196, 2736, 2598, 344, 201, 183, 622, 1001, 1489, 338, 1764, 2785, 1610, 2725, 2576, 1285, 1396, 1618, 2734, 2673, 3130, 361, 362, 1501, 1247, 411, 1153, 1854, 348, 1079, 1912, 2466, 751, 376, 380, 379]. Verlag [382, 1450, 2205, 1133, 1174, 1391, 1069, 2146, 609, 3073, 690, 565, 695, 2857, 1905, 1916]. Verlages [220]. Verlagsgesellschaft [303, 1095, 557, 1797, 348]. Verlangsgesellschaft [1602]. verlichting [541]. Verloren [1589, 2986]. Vermin [3636]. Vermisste [1012]. Vernacular [3559, 3558, 2582]. Vernet [3019]. vernier [1744]. Vernon [1692]. vero [312]. Veroeffentlichungen [248]. Veröffentlichung [966]. Veröffentlichungen [64, 172, 1724, 1687]. Véron [3071]. Verona [81, 1597, 426, 705, 1137, 3730, 3707, 1597]. veronensis [426]. Veronica [2712]. Véronique [3611]. verreinato [686]. Verres [3071]. Versammlungen [557, 1797, 3977]. versehen [362]. version [334, 3408]. versione [1134]. versiones [1687]. Versions [1047, 3665]. Verso [1979, 1693]. Versuchsanstalt [3462]. vertaald [3109]. Vertebrati [2612]. Vertical [1703]. Vertonungen [3905]. Verwaal [3690]. Very [3418, 2087]. Verzeichnis [3905, 1079]. Vesal [2598]. Vésale [2952]. Vescovile [2271]. Vescovini [1061]. vescovo [719]. Vespucci [3293]. Vesubio [2494]. vetenskap [2200, 3852, 3853, 3854]. vétérinaire [1187]. Veterinaria [21]. Veterinary [3489]. Vetrano [1665]. vetraria [2329, 2280]. vetrarie [2592, 3934]. vetreria [575, 2149, 3948]. vetro [575, 1255, 1695]. vetta [3498]. Vetustissimus [3485]. Vexation [2674, 3882]. Vezzosi [3718]. Vhc [1150]. VI [1795, 1004, 309, 386, 1484, 1641, 859, 2205, 1391, 895, 2710, 1852, 2167, 2238, 2166, 2237, 2942, 1495, 1689, 2689, 1572, 3263, 2962, 2657, 2546, 3024, 850, 912, 2884, 2617, 997, 692, 2911, 1003, 945, 1973, 471]. Via [2910, 726, 3255, 715, 3915, 2458]. Viaggi [2965, 2206, 3828, 3022, 3178, 3965, 2126]. Viaggiatore [740, 1428]. Viaggiatori [588, 1304, 1472, 1287, 1663]. Viaggio [315, 1380, 1304, 1472, 1673, 727, 2049, 2399, 942, 3788, 2849, 1869, 1538, 2590, 1107, 2893, 2363]. Vial [3586]. viale [44]. Viano [2517, 2648, 2821, 2693, 1651]. Vibrator [2016, 3916]. vicenda [1236]. Vicente [2114]. Vicenza [1594, 1435, 2679]. Vichiana [1821]. Vickers [166, 2609, 1907]. Vico [581, 858]. Victims [3715]. Victor [574, 2101, 2114, 888, 2593, 2114]. Victoria [2114, 3733]. Victorian [56, 2794, 341, 2858, 2617, 3715, 345, 3032, 1798, 3033, 1466]. Victory [3882, 2674]. Vidal [2559, 3446, 2571]. vide [2503]. videlicet [4028]. videotape [1094]. Vidi [51]. Vie [3861, 2052, 496, 2528, 2572, 866, 2578, 2892, 2053, 282]. Viella [2420]. vienen [3618]. Vienna [3462, 492, 3341, 2482]. vier [1941]. vierges [3448]. Vierhaus [220]. vierzig [420]. View [1686, 3248]. Vieweg [259, 259]. Views [3488, 4014]. vigilia [2536]. Vigna [2612, 2151]. Vigni [2221]. Vignola [1679]. Vignoli [3799, 3285]. Vignotto [1599]. VII [3368, 991, 1077, 2629, 709, 2529, 1086, 1388, 805, 959, 652, 1518, 2093, 1494, 21, 752, 1492, 2720, 2548, 3088, 2663, 347, 1334, 945, 2726, 472, 1973]. vii-397 [752]. vii-95 [21]. VII/1 [2529]. VIII [1105, 1156, 2100, 242, 283, 313, 516, 520, 794, 801, 685, 1110, 1087, 1090, 1365, 1846, 2120, 2160, 2370, 1967, 2038, 2436, 3120, 495, 533, 2822, 2805, 2791, 2958, 35, 2895, 173, 2719, 385, 2747, 3074, 2427, 3208, 535, 1074, 3038, 746, 747, 68, 473]. viii-237 [35]. viii-488 [2436]. Viktor [2502, 1203, 1971]. Viladrich [353]. Vilibankhov [2662]. Villa [1813, 2005, 2217, 2919, 1331, 2086, 1975]. Villanueva [648]. ville [314]. Villes [314]. Villette [2188]. Vinassa [1851]. Vincent [796, 1494, 1848, 2273, 2829, 2890]. Vincentiis [2415]. Vincenzi [3374]. Vincenzio [901, 788]. Vincenzo [3486, 2077, 299, 2871, 1209, 3374, 2081, 2083, 1779, 2046]. Vinci [1950, 2021, 2769, 2598, 3014, 1424, 3372, 3628]. vintage [1510]. Vinti [1995, 1994]. Viola [1273, 862]. Violani [1145, 3412]. Viollet [3701]. Viollet-le-Duc [3701]. Virey [416]. Virgilio [2415]. Virginia [2964]. Virginio [3497, 2450]. virgins [3448]. Virology [3676]. Virtue [2772]. Virtuosi [3357]. Virtuoso [900, 1234]. Virtus [2321]. Virus [3676]. Viruses [3601, 3601]. Visconti [828, 930, 431, 518, 1343, 1871]. Visconti-Roma [1343]. visibile [1493]. Visible [3779, 3524, 3261]. Vision [1365, 1686, 3608, 611, 2890, 2404, 3406, 319]. Visione [559, 319, 570, 1788, 611]. visioni [1794]. Visions [544, 3728, 2845, 3625, 2908]. visit [3927]. visita [2157, 3927, 937]. Visitors [2701]. visiva [1788, 1485]. Visser [2884]. vista [1955, 2026, 185]. Visual [2160, 3268, 3501, 3430, 3468, 3153, 3958, 3740, 3261, 3994, 3995, 3233, 3779, 1918, 4086, 2120, 2867]. Visualizing [3555, 3580, 2819]. visuelle [1918]. Vita [2647, 585, 643, 1209, 522, 874, 1268, 2222, 661, 1311, 924, 1291, 546, 61, 2916, 3299, 3757, 3760, 336, 2472, 180, 3015, 252, 442, 2490, 3844, 81, 883, 589, 1768, 3969, 13, 3972, 1840, 1068, 2919, 2503, 2910, 2668, 2743]. Vital [172, 2897]. Vitale [407]. Vitale-Brovarone [407]. vitales [3026]. Vitaliano [2049, 2399]. vitalisme [3571]. vitaliste [3197]. Vitality [3054]. Vitalizing [2784]. Viti [1360]. Vito [2043]. Vitrometrum [2175]. Vitruvian [3456]. Vitruvio [1645, 393]. Vittoria [1417, 1229, 456]. vittoriana [765]. Vittorio [1961, 2032, 1815, 407, 1020, 1018, 2415, 558, 1961, 2032, 79, 1030, 1522, 1676, 626]. Vivarium. [3462]. vivente [1201, 1753]. Viviani [2871]. Vivisection [3375]. vivo [1744]. Vivoli [1324]. Vivum [660]. Vladimiro [238, 233, 1926, 44]. vm [2656, 1497]. vocabulaire [1224]. Vocation [17, 3191]. voi [2156]. Voice [2906]. Voigts [3440]. Vol [583, 1657, 2199, 1006, 1648, 1765, 235, 3792, 3793, 3794, 1670, 3027, 1335, 2808, 547, 3345, 613, 3225, 2781, 829, 1484, 2713, 554, 1671, 2529, 3944, 2886, 186, 3183, 1857, 2937, 4004, 1223, 1055, 2418, 2055, 339, 621, 620, 1316, 2522, 1769, 1790, 2524, 198, 1447, 1442, 70, 2156, 2605, 2727, 1401, 334, 367, 64, 420, 2892, 1315, 1641, 1982, 2629, 2004, 2590, 1866, 171, 727, 1193, 1570, 1750, 1749, 1871, 922, 3224, 1071, 1229, 849, 3987, 3988, 1225]. Vol. [3750, 3773, 3774, 3776, 3785, 4068, 4074]. Volder [3569]. Volgar [1829]. volgare [2880]. Volgarizzamento [1645]. Volker [2837]. Voll [519, 3013, 617, 1364, 2200, 1950, 2021, 2332, 3283, 931, 497, 686, 673, 1569, 1898, 221, 1235, 3015, 249, 445, 2283, 1516, 1820, 2156, 1149, 1944, 1595, 2011, 1690, 791, 1786, 13, 1154, 1408, 808, 1349, 1330, 541, 746]. voll. [1387, 10]. Vollmuth [2598]. Volpaia [1215]. Volpone [3090]. Vols [278, 2612, 2530, 2882, 2522, 2605, 2949, 2719, 2101, 225, 2892, 2590, 2938, 2615, 2886, 3284, 2274, 2914, 2937, 3089]. Volta [1869, 2156, 2904, 1813, 2005, 2217, 2433, 2435, 3773, 3774, 3775, 3776, 3788, 2395, 2476, 3955, 2127]. Volta.

[2463]. Voltaggio [1093]. Voltaire [1681, 3024]. Volterra [2043]. Voltiana [3776, 2156, 3773, 3774, 3775, 2156]. Volume [1367, 1606, 1819, 1795, 2464, 3368, 2899, 3193, 3227, 3269, 675, 1608, 2852, 180, 252, 991, 1156, 1238, 2080, 3111, 2955, 1732, 1407, 1058, 2107, 2608, 851, 1154, 1654, 4032, 2726, 313, 2495, 1516, 1164, 71, 878, 1435, 44, 2274, 4008, 4009]. Volumen [2114]. volumes [226, 420]. volvelle [1160]. Vom [364, 2228, 173, 3902, 276, 1086]. Voniangi [740]. Vons [2952]. voor [302]. vooravond [541]. voorzien [802, 3109]. vor [1726, 751]. Vorbildmoderner [1608]. vorhandenen [3905]. Vorlesungen [1977]. Vorstandes [1735, 2196]. Vorträge [1608, 173, 276, 492, 296, 557]. Voss [2979]. Vosse [118]. Voten [343]. Voyage [1476, 3151, 1566, 3145, 1475, 3153, 3148, 1408, 2013, 3932]. Voyages [1306]. Voyageurs [1305, 1474, 2798]. Voyaging [3149, 3144]. Vraie [2578]. Vrin [3107, 416, 1981, 512, 2648, 1138, 1667, 2821, 1228, 2767, 1647, 1460, 625, 1326, 2578]. Vs [2262, 2740]. VT [2707, 2981]. Vtterli [138]. Vucinich [770]. Vuibert [2958]. Vuillemin [3075]. vulcani [2849]. vulcano [3672]. Vulgar [916, 1829]. vulgarisation [1543]. Vulkanismus [1069]. Vulkanismus-Neptunismus-Kontroverse [1069]. Vulkanismusstreit [1069]. Vultus [1591]. Vv [1101].

W [516, 440, 414, 1201, 1393, 3509, 2669, 3475, 2805, 234, 235, 1808, 3268, 192, 798, 994, 1278, 1067, 2381, 3346, 2740, 343, 2095, 681, 948, 2801, 2775, 1001, 450, 753, 1249, 229, 345, 3740, 1503, 411, 689, 806, 1034, 1455, 2466, 2465, 945, 1225, 2556, 2913, 2054, 3526, 1584, 3525, 1717]. W. [2354, 497, 532, 2106, 2909, 2139, 1563]. W.-C [2095]. W.W [1853, 3281]. wa [2199, 1226]. wa-al-qarn [2199]. wa-l-magrūsāt [1226]. Wabster [194]. Waer [279]. Wagner [362]. Wahlvervandtschaften [543]. Wahlvorschlge [508]. Wahrnehmung [1918]. wa’l [609]. wa’l-muqbala [609]. Walbaum [120]. Wäldern [3073]. Wales [3131]. Walker [1836]. Wallace [307, 1814, 3579, 819]. Wallenborn [1879]. Waller [1055, 1218]. Wallin [1441, 884]. Wallstein [2853, 2532, 2736, 2673]. Walsdorf [137]. Walsh [2435]. Walstorf [137]. Walstorff [137]. Walter [1953, 2024, 2822, 2933, 443, 571, 798, 994, 2202, 1001, 347, 61, 612, 257, 1068, 379]. Walther [1074, 3981]. Walton [2707, 3441]. Walzel [226]. Wandel [1074, 3981]. Wandering [3247]. Wang [406, 3566]. Wangji [3664]. Wanrooij [852, 730]. War [441, 3158, 3369, 2977, 1122, 1662, 2731, 2446, 2678, 3974, 3499, 3482]. Warburg [964, 830, 1102, 334, 408, 2308, 610, 791, 378]. ward [1918]. Wardy [2518]. Warfare [2709, 2938]. Warren [3063, 551]. Wars [3118, 2231, 1505]. Warszawa [1584]. Was [3191, 1731, 3334, 3451, 2646, 3868, 356, 2433, 4028]. Washington [242, 491, 885, 959, 3081, 1246, 1184, 184, 293, 1837, 1290, 58]. Wassermann [871]. Watches [429, 1510]. Watchmaking [1510]. Watelet [1663]. Water [734, 3370, 2525, 1182, 734]. Waters [1791, 3627, 3914]. Watson [1494, 2594, 2839, 2745, 2549, 2828]. Watt [2899, 2604]. Watzlawick [372]. waves [1573, 2739]. Wax [3222, 3147]. Waxworks [3654]. Way [1412, 1723, 382, 3412, 4036]. Wayne [2906]. Ways [2764, 3144].

Waźbiński [278]. Wear [1275]. Weather [2879, 2939]. Webb [1246]. Weber [565]. Webster [3072, 2755]. Wechselwirkung [1151]. Wedekind [565]. Wedekinds [565]. weeds [3873]. Weg [497, 1074, 3981, 382]. Wegener [348, 3429]. Wehefritz [1387]. Weidenfeld [2881]. Weidmann [746, 747]. Weigel [3064, 3283]. Weights [857]. Weijers [306, 1444]. Weikersheim [1402]. Weimar [226]. Weinberger [167]. Weindling [768]. Weiner [1911, 1329]. Weinheim [1095, 950, 1150, 1336]. Weiss [2810, 695]. Weizig [1854]. Weizsäcker [2502, 1971]. Welch [3241]. Welcker [173]. Welcker-Tagung [173]. Welker [173]. well [1111]. Wellcome [224, 414, 1182, 3241]. Wellman [1144]. Welt [571, 798, 994, 4082, 1838]. Weltbildes [1086]. Welterklärung [2837]. Wenderoth [234]. Went [3101]. Wenzel [622]. Werdegang [1714, 565]. Were [3467, 3472, 3533]. wereld [2986]. Werk [1281, 2544, 1150, 173]. Werke [217, 2725, 1328, 3893]. Werner [173, 961, 183, 2062, 1854]. Werner-Reimers-Stiftung [173]. Wesley [990]. West [1288, 1261, 1660, 2308, 163, 1837, 3878, 3880]. Western [2228, 2909, 857, 2561, 2749, 2985, 993, 3902, 3906]. westerse [2986]. Westeuropa [2228, 3902]. Westfall [164, 617, 1791, 613, 1249]. Westman [801, 1193]. Westview [1763]. wetenschap [1422]. Wetenschappen [302, 3018, 2884]. Weyer [1402]. whatever [4021]. Wheatland [2182]. Wheaton [1618]. Wheatsheaf [544]. Whewell [3051]. which [3530]. Whipple [856, 409, 762, 1072, 2147]. Whitby [2938]. White [1401, 2296, 68]. whitenes [2248]. Whiteside [1198]. Who [2625, 2971, 1305, 1203, 4012]. Whole [2875]. Whyte [1613]. Wi [348]. wider [1450]. Widmalm [2972]. Widmung [2837]. Wie [695]. Wiederabdruck [1531]. Wieland [1445]. Wielgus [1281]. Wien [253, 30, 58, 248, 233, 1854, 382]. Wienand [3082]. Wiesbaden [259, 567, 2837, 1191, 217, 276, 376, 609]. Wijnands [3412]. Wiken [1146]. Wiksell [1611]. Wild [2424, 2906, 1170, 3834]. Wildberg [443]. Wiley [2864, 1089, 1993]. Wilfrid [1314, 222]. Wilfried [188]. Wilhelm [896, 3598, 104, 1328, 1150, 507, 125, 98, 2831]. Wilhelminischen [547]. Wilkins [2124]. Wilks [1686]. Will [964, 3591]. Wille [2866]. Willem [2715, 2556]. Willen [1450]. William [2881, 2388, 1314, 549, 29, 307, 445, 173, 624, 3529, 1343, 1877, 808, 1225, 4032, 224, 405, 490, 579, 722, 842, 1204, 2160, 1403, 2850, 2834, 2594, 244, 2839, 2595, 225, 2718, 1466, 997, 1397, 176, 1089]. Williams [227, 1808]. Williamsburg [2964]. Willich [112]. Willie [68]. Willig [112]. Wilma [330, 859, 2111, 833]. Wilson [2312, 1842, 1670, 808, 1202, 1837, 3417]. Winchester [244]. Winckler [107]. Wind [3392]. Winding [3146]. Windus [1945]. Winkelmann [64]. Winkle [62]. Winkler [107]. Winsor [1541]. Wireless [457]. Wirkung [1281, 173, 1150]. Wisconsin [15, 520, 1457, 229]. Wisdom [3737, 3643, 3550]. Wise [1497, 689]. Wiss [639]. Wissens [3530]. Wissenschaft [296, 507, 2532, 1150, 1381]. Wissenschaften [507, 23, 571, 798, 966, 1191, 220, 542, 1854, 376, 382]. Wissenschaftler [363]. Wissenschaftliche [1602, 188, 1446, 557, 1918, 1797, 348, 296]. Wissenschaftlichen [2837, 253, 492]. wissenschaftlicher [3129, 415, 3999].

Wissenschaftsaustausch [2228, 3902]. Wissenschaftsbeziehungen [3893, 1912]. Wissenschaftsgeschichte [296, 1941, 557, 276, 1079]. Wissenschaftshistorischer [1328]. Wissenschaftslegitimierung [2837]. Wissenschaftslehre [1288]. Wissenschaftsphilosophie [498]. Withers [2805]. within [3437, 3513]. without [3414]. Witnesses [1703]. Wittgenstein [629, 691]. Wittgenstein. [629]. Wittich [1193]. Wittje [3288]. Wokler [1845]. Wolf [567, 1905, 1066, 1715]. Wolf-Dieter [567]. Wolfe [347, 3571]. Wolfenbütteler [1191, 276]. Wolff [2902]. Wolfgang [259, 1402, 571, 173, 2900, 256, 1796, 2689, 1714, 690]. Wolfram [2939]. Wolfschmidt [1862, 1450]. Women [77, 440, 2016, 3108, 2050, 2835, 2832, 786, 2477, 2660, 3616, 3567, 3361, 4067, 3916]. Won [3467]. Wonder [3196, 2801, 3460, 4044]. Wonderful [2702, 1185]. wonders [2844]. Wong [500]. Wood [2633, 289]. Woodblock [3501]. Woodbridge [615, 1490, 2143, 1314, 1063, 616]. Woodcut [3047, 3048]. Woodcuts [1628]. Woodrow [1837]. Woodward [1812, 1934, 1935]. Wootton [1087]. Word [3223]. Wordly [964]. Work [1910, 1401, 1880, 3599, 3157, 3702, 766, 2378, 2534, 2971, 2311, 2178, 3917, 3933, 3971, 3979, 1454, 3093]. Workplace [3638]. Works [249, 3298, 2765, 3014, 2815, 2294, 3844, 3893, 763, 3970, 3972, 541, 3785, 4065, 1907]. Workshop [583, 3484, 4052, 3699, 1126, 1297, 4040, 2169, 2240, 1792, 2078]. Workshops [761]. World [438, 1204, 2881, 3000, 2717, 3229, 3499, 1736, 3175, 1842, 2653, 3653, 3414, 2216, 3176, 1152, 3846, 3281, 1770, 2268, 3686, 1841, 3017, 499, 3713, 862, 3634, 3294, 2846, 2956, 3363, 3147, 60, 3749, 1476, 1375, 447, 3832, 3858, 1987, 3870, 3918, 1766, 3528, 3998, 4010, 2688, 2601, 2879, 2964, 2920, 2657, 2564, 2832, 2862, 2061, 816, 4082, 2551, 2066, 3692, 3401, 3369, 2124]. Worlds [3171, 2800, 3513, 723, 3510]. Worldview [164]. Wormald [1194]. Wormbs [2972]. Woronoff [1326]. Worth [1097]. Wounds [2709]. Wragge [3542]. Wragge-Morley [3542]. Writing [3438, 3581, 699, 4035]. Writings [4041, 2332, 1627, 2216, 1225, 3838, 4083, 3964, 3992, 2790]. written [3347, 3326, 3316, 3369, 3397, 3400, 3312, 3344, 3345, 3346, 3399, 3331, 3327, 3402, 3313, 3333, 1731, 3314, 3398, 3396, 3332, 3330, 3334, 3401, 3325, 3361, 3367]. Wroclaw [1731]. Wrong [2764]. Wulf [3401]. Württembergisch [1402]. Württembergische [725]. Würzburg [1918]. Wvg [1905]. Wydawnictw [1281]. Wydawnicza [1256]. Wykes [3025]. Wynne [1694]. Wynter [939].

X [77, 2120, 2267, 235, 1112, 2849, 648, 720, 1447, 1688, 2014, 2059, 1917, 2061, 2137, 2362, 2152, 1437, 2009, 2206, 1645, 2204, 1922, 1998, 1326, 2422, 2956, 441, 439, 592, 1200, 1457, 1617, 1683, 1692, 2143, 2776, 3041, 2573, 2647, 2708, 2607, 2434, 2389, 2606, 3005, 3063, 3081, 2580, 2865, 3035, 1278, 3001, 2917, 1366, 1670, 1669, 1164, 2595, 2905, 1395, 2832, 2094, 2900, 770, 1878, 616, 1320, 2785, 2858, 2686, 450, 1618, 3125, 2866, 1723, 2585, 2939, 2728, 2784, 3126, 766, 289, 1203, 1426, 1654, 2465, 1733, 1337, 2549, 2625, 2582, 3004, 1483, 2147, 2604, 1150, 1905]. x [2688, 2571, 2668, 2194, 2421, 2675, 2849, 475, 2558, 2917, 2957, 1791, 2808, 2751, 2630, 2709, 2493, 2430, 2940, 2225, 1074, 2731, 2583, 2734, 2272, 2506, 2378, 2269, 2937, 2714, 2322]. x-109 [1905]. x-248 [450]. x-346 [2389]. X-ray [235]. Xaver [2255]. Xavier [1707]. XC [516]. XCIII [2571]. xcvi [2488]. Xe [2573, 3058, 746]. Xelmírez [2068, 3990, 3990]. Xhayet [2267]. XI [534, 2809, 454, 414, 490, 2462, 2737, 1440, 498, 2589, 2873, 2838, 2946, 3131, 760, 767, 1246, 1987, 2214, 2134, 1841, 285, 418, 862, 3093, 1075, 1501, 3096, 3044, 2274, 2861, 3004, 946, 1498, 1623, 476]. xi-739 [498]. XIe [2199]. XII [802, 1204, 23, 716, 886, 1279, 2213, 678, 301, 335, 342, 579, 923, 888, 1002, 1231, 1234, 1194, 1362, 1406, 1427, 1449, 1490, 1819, 1910, 2075, 2056, 2495, 2746, 2881, 2164, 2235, 3002, 2669, 841, 1533, 1727, 2538, 2933, 3254, 3025, 1443, 538, 1067, 1263, 1366, 2010, 2824, 3066, 34, 2949, 2662, 2831, 2774, 408, 3289, 1409, 1575, 286, 2862, 1664, 624, 848, 2362, 2940, 25, 704, 835, 1320, 2676, 2844, 2734, 2807, 2539, 1503, 803, 3032, 387, 1800, 1245, 2562, 2802, 1771, 3783, 477]. xii-218 [387]. xii-232 [408]. xii-298 [34]. XIII [1367, 250, 446, 572, 1462, 408, 204, 2880, 1695, 2628, 629, 544, 842, 1763, 1945, 623, 2891, 2959, 349, 57, 2999, 919, 933, 2718, 1915, 2945, 1071, 2372, 693, 1089, 1425, 2919, 2846, 2752, 1845, 2772, 1881, 478]. XIII-XVI [2880]. XIIIe [2878, 306, 512, 1984, 1985]. Ximenes [340, 1783]. Ximeniano [688, 390, 187, 453, 655, 590, 700, 1014, 1430, 700]. Xing [3566]. XIV [573, 250, 1105, 949, 952, 1372, 378, 168, 615, 672, 918, 1070, 1005, 1169, 1261, 1454, 1535, 1562, 1587, 1725, 2067, 2312, 2560, 3016, 2710, 2879, 2165, 2236, 2435, 1736, 3072, 2105, 2920, 3043, 219, 218, 2751, 2633, 1722, 3079, 3243, 191, 1640, 2754, 2938, 2944, 26, 2976, 833, 969, 3034, 2803, 1144, 1349, 2324, 2956, 2674, 1336, 2848, 2960, 1032, 2039, 479]. xiv-342 [1144]. xiv-410 [1261]. xiv-665 [219]. XIVe [1985]. XIVeme [1187]. XIX [514, 1300, 496, 3021, 43, 3205, 752, 789, 876, 453, 938, 1730, 1861, 2217, 535, 1324, 1346, 2159, 1393, 2106, 3013, 2379, 1398, 928, 1944, 1542, 2982, 2501, 87, 1032, 3759, 486, 37, 979, 484, 734, 4006, 1378, 938]. xix-429 [2379]. XIXe [3395, 496, 2735, 2798, 40, 2973, 2048, 1848, 3657, 175, 2566, 1121]. Xl [1091, 412, 2530, 2004, 1001]. Xliii [310, 1330]. xliv [2781, 201, 329]. xlvi [2955]. Xlviii [1229, 2529]. xm [2424]. xn [1400]. xn-402 [1400]. XV [2882, 446, 576, 376, 3364, 1324, 164, 440, 651, 759, 795, 891, 1657, 2586, 2897, 846, 1316, 2664, 2671, 1721, 2579, 2755, 3061, 2888, 368, 2654, 1656, 1997, 1315, 758, 1411, 1319, 216, 3083, 1249, 2643, 1328, 1571, 2603, 1720, 2890, 2587, 2138, 2581, 914, 2545, 2534, 1332, 2322, 2971, 2122, 480, 2130]. XVe [2878, 1447, 2596, 2877, 1626]. XVI [514, 889, 1091, 1194, 1648, 1817, 2523, 29, 572, 1335, 43, 863, 917, 2880, 561, 1813, 1860, 1872, 12, 679, 837, 1106, 1108, 1333, 1536, 1818, 2459, 2516, 2889, 493, 1488, 2964, 3023, 3279, 3015, 279, 2103, 1147, 2010, 2655, 172, 244, 2600, 2614, 2377, 2801, 2775, 36, 1369, 2098, 545, 2753, 3078, 3114, 2975, 3118, 2875, 2618, 3085, 3058, 2378, 1198, 1329, 3203, 280, 3250, 967, 1607, 2241, 2158, 3759, 481, 2356, 658, 1378, 4017]. xvi-160 [244]. xvi-396 [1147]. XVIe [3173, 3013, 425, 308, 3890, 3980]. XVII [1648, 1689, 2888, 1153, 3329, 1097, 2528, 2972, 1500, 771, 345, 1007, 3086, 2748, 87, 513, 650, 889, 1361, 1762, 2340, 2473, 254, 1529, 1942, 37, 482, 1286, 660, 383]. xvii-186 [771]. XVIIe [675, 2798, 1138, 3173, 2550, 2829, 2652, 2048, 1875, 920, 540, 582, 2566, 3803]. XVIII [1784, 1793, 2199, 1066, 2780, 2523, 3021, 3331, 3024, 952, 3496, 453, 1251, 1861, 2005, 610, 259, 722, 968, 1199, 1313, 1531, 1795, 2387, 1790, 551, 2756, 2268, 3267, 2799, 2678, 1786, 3130, 2830, 361, 1541, 993, 1837, 1290, 2437, 1823, 2888, 483, 494, 618, 660, 2399, 4006]. xviii-484 [2387]. xviii-540 [1790]. XVIIIe [1060, 2894, 920, 3894, 1700]. xvm [1139, 1683, 1649, 1660, 2551]. XX [1194, 1793, 1689, 1817, 964, 547, 863, 1186, 917, 1695, 1872, 2213, 2677, 800, 617, 801, 1340, 1853, 2016, 2463, 2601, 2887, 3000, 2386, 1643, 619, 1791, 1152, 56, 2819, 3108, 2652, 1641, 2970, 1177, 2713, 1843, 929, 2584, 2725, 2645, 2867, 2985, 2786, 2626, 348, 1455, 2967, 3037, 1798, 513, 650, 967, 1361, 1607, 1762, 3783, 254, 1529, 1942, 1335, 485, 1286, 1978, 383]. xx-1176 [1152]. xx-viii [801]. XXe [2550, 2973, 1848, 1799, 3657]. XXème [3087]. XXI [3286, 1515, 35, 751, 2921, 885, 621, 1143, 65, 1461, 2158]. xxi-253 [65]. XXIe [1804]. XXII [830, 1605, 3113, 3286, 612, 1568, 3124, 3076, 2706, 2745, 2552, 1856, 723, 555, 1592, 2739]. xxii-651 [555]. XXIII [512, 1102, 3525, 405, 2864, 2981, 707, 534, 2984]. XXIV [1280, 889, 1816, 2522, 2780, 2605, 3013, 1982, 2713, 1369, 1577]. xxix [1911, 2851, 2059, 2977, 3014]. xxn [1915]. XXth [4071, 4087]. XXV [1280, 1069, 2873]. XXVI [445, 537, 2370, 1966, 2037, 2907, 1484, 689, 806]. XXVI/1 [445]. XXVII [1685, 2469, 1502, 1322, 3018, 3008]. XXVIII [43, 59, 2553, 2771, 1407, 849]. xxx [2194, 2464, 1717, 1242, 1343, 255, 2576, 1642]. xxx-148 [255]. XXXI [1055, 57, 1076]. XXXII [3259, 1135, 1404, 1239, 1317, 2783, 2841, 851, 860]. XXXIII [868, 3127]. xxxiv [2883, 3012, 2712, 2013, 1109]. XXXIX [306, 172, 2790, 380]. xxxv [2605, 800, 1154]. XXXVI [714, 353, 2834, 582]. xxxvi-1017 [582]. xxxvii [1196].

Yaakov [3410]. Yabuuti [2197]. Yale [587, 1234, 1449, 2056, 3072, 1738, 3204, 3012, 2948, 65, 2944, 2961, 3038, 2626, 1720, 840]. Yans [754]. Yapp [833]. Yarmouth [886]. Yasmin [2686]. Year [1846, 2350, 3881, 3919, 3920]. Years [2899, 2852, 3319, 3852, 3853, 3854, 2392, 1195, 1249, 1408, 676, 1440, 3060, 3905, 355, 1230, 3969, 1137, 2538, 2987]. Yeo [1562]. Yolton [450]. York [439, 405, 579, 759, 724, 1169, 1535, 1692, 1853, 2601, 2746, 772, 2606, 2822, 2933, 3059, 2864, 1004, 2653, 1808, 2906, 194, 443, 571, 798, 994, 1103, 1104, 990, 1277, 1188, 1358, 498, 1726, 3061, 219, 2827, 2600, 2922, 3225, 3281, 1505, 919, 933, 2947, 80, 225, 286, 2835, 683, 1179, 1100, 2987, 2684, 36, 26, 622, 1489, 862, 1877, 545, 2686, 1071, 3008, 451, 2678, 345, 1786, 1980, 2366, 2412, 2983, 3241, 361, 362, 851, 348, 860, 892, 913, 1089, 2861, 1993, 2587, 2582, 746, 747, 914, 2545, 1814, 690, 565, 68, 2417, 1445]. York [1686, 753, 749, 997, 3413]. York-London [439, 1692, 990, 3281]. Young [56, 1196, 347, 629]. Youth [2091]. Yulia [3478]. Yung [2137]. Yunus [1309]. Yves [620, 3010, 1543]. Yves-Marie [3010]. Yvette [987, 2318]. Yvonne [1418].

Z [1573, 499, 753, 3084]. Z. [804]. Zabarella [59]. Zacchia [3115]. Zaccolini [3436]. Zach [2255]. Zachary [3714]. Zadigroma [2692]. Zagli [2266]. zahlr [1838]. Zambelli [1565]. Zammito [3571]. Zamorani [2670]. Zampetti [20]. Zan [2726]. Zanatta [1606]. Zanca [2754]. Zanetti [3527, 3456]. Zanfi [1595]. Zanichelli [792, 1462, 912, 2548, 570]. Zanobio [1809]. Zanrosso [2679]. Zanzi [2965]. Zaragoza [2808]. Zarlino [2083]. Zavattini [2155]. zbiorach [1584]. Zdzisław [1359]. Zecca [2202, 1665]. Zeit [1602, 543, 2785, 1150]. Zeitalter [220, 547, 2598, 565]. Zeiten [2466]. Zeitgenossen [362]. Zeitlinger [802, 1367]. Zeitschrift [1941, 1151, 173]. Zelig [951]. Zellbiologie [4062, 1992]. Zenone [1821]. Zensetsou [500]. Zentralantiquariat [226]. Zhmud [2822]. Zhou [1640]. Zhouyi [2094]. Zik [3410]. Zimmermann [698, 29, 128]. Zina [1599]. Zinato [2659]. Zinn [121]. Zinner [2184]. Zinner-Archive [2184]. Zinsser [2660]. Zionism [2977, 3974]. Zocchi [2743]. Zodiacs [94]. Zoltan [1387]. Zone [2684, 2983]. Zoologia [1496]. Zoological [841]. Zoologie [1058, 1306]. zoologo [1963, 2034]. Zoology [1058, 1408, 1541]. Zosime [2704]. Zotti [1794]. Zuccari [1555]. Zuccoli [1591]. zufolo [2630, 3956]. Zukunft [1714]. zum [296, 2228, 443, 547, 3902, 1396, 364, 1989]. Zumbo [2416]. Zunino [1903]. Zupko [857]. Zur [1043, 343, 2854, 1042, 2673, 3186, 1202, 62, 253, 296, 567, 1602, 2055, 508, 497, 1171, 29, 23, 32, 307, 966, 1191, 1359, 1941, 498, 173, 172, 420, 3530, 3874, 1918, 1724, 2265, 183, 338, 1058, 2785, 535, 1074, 1838, 1610, 4082, 1797, 364, 1086, 1912, 3981, 376, 380, 746, 747, 1069, 565, 492, 695, 4033, 1905, 415, 1153, 3999]. Zürich [193, 248, 1489, 746, 747, 343]. Zusammen [1731]. Zusammenarbeit [1191, 3893, 201, 1912]. Zweiten [1281, 690]. Zweiter [2831, 411]. Zwink [725]. zwischen [2228, 2682, 1151, 64, 3857, 3893, 3902, 3186, 1912, 3073, 3905]. Zygmunt [278].

References

Garin:1986:FSN

[1] Eugenio Garin. Fra ’500 e ’600: Scienze nuove, metodi nuovi, nuove accademie. (Italian) [Between 500 and 600: New science, new methods, new academies]. Nuncius, 1(1):3–23, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

      brillonline.com/content/10.1163/182539186x00016.

Bedini:1986:GJ

[2] Silvio A. Bedini. The Galilean jovilabe. Nuncius, 1(1):25–46, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

      tronic). URL http://booksandjournals.brillonline.com/content/
      10.1163/182539186x00025.

Proverbio:1986:SCP

[3] Edoardo Proverbio and Giuliano Bertuccioli. On a singular Chinese portable sundial. Nuncius, 1(1):47–58, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539186x00034.

Arrigoni:1986:PAD

[4] Tiziano Arrigoni. Per un archivio della corrispondenza degli scienziati italiani. (Italian) [towards an archive of correspondence of Italian scientists]. Nuncius, 1(1):59–61, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00043.

Anonymous:1986:IDA

[5] Anonymous. Indice delle abbreviazioni. (Italian) [List of abbreviations]. Nuncius, 1(1):62–131, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00052.

Anonymous:1986:IAM

[6] Anonymous. Indice alfabetico dei mittenti e dei destinatari. Nuncius, 1 (1):132–139, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00061.

Brenni:1986:IF

[7] Paolo Brenni and Massimo Misiti. Istituzioni e fonti. Nuncius, 1(1): 141–184, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00070.

Turner:1986:L

[8] Anthony J. Turner. Loge. Nuncius, 1(1):185–187, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539186x00089.

Finocchiaro:1986:DC

[9] Maurice A. Finocchiaro. Discussioni critiche. Nuncius, 1(1):189–202, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00098.

Bucciantini:1986:BRI

[10] Massimo Bucciantini. Book review: Isis cumulative bibliography 1966– 1975. A bibliography of the history of science formed from ISIS critical bibliographies 91–100 indexing literature published from 1965 through 1974. Edited by John Neu. London-Bronx, N.Y., Mansell in conjunction with the History of Science Society 1980–1985, 2 voll. Nuncius, 1 (1):203–204, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00106.

Abbri:1986:BRC

[11] Ferdinando Abbri. Book review: C. A. Russell (editor), Recent Developments in the History of Chemistry, London, the Royal Society of Chemistry 1985, 333 pp. Nuncius, 1(1):205–207, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00115.

Settle:1986:BRG

[12] Thomas B. Settle. Book review: Giovanni Bordiga, Govanni Battista Benedetti, Filosofo e Matematico Veneziano del secolo XVI (Reprinted from the Atti dell’Istituto Veneto di Scienze, Lettere ed Arti. Tomo Lxxxv, 1925–1926, Parte II, pp. 585-754), with an annotated bibliographical update by Pasquale Ventrice, Venice, Istituto Veneto di Scienze, Lettere ed Arti, 1985. Nuncius, 1(1):207–209, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00124.

Romano:1986:BRM

[13] Renato Romano. Book review: M. Pavone, Introduzione al pensiero di G. B. Hodierna, 2 voll., Modica, Setim-Editrice 1982, 419 pp. L. 30.000. M. Pavone, La vita e le opere di Giovan Battista Hodierna, Ragusa, Didattica Libri Eirene Editrice 1986, 436 pp. s.i.p. Nuncius, 1(1):209–213, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00133.

Maffioli:1986:BRH

[14] Cesare Maffioli. Book review: Henk Th. Van Veen and Andrew P. Mccormick, Tuscany and the Low Countries: an introduction to the sources and an inventory of four Florentine libraries, Firenze, Centro di 1985, 182 pp. (Italia e i Paesi Bassi. Serie edita a cura dell’Istituto Universitario Olandese di Storia dell’Arte di Firenze, 2). Nuncius, 1(1):213–216, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00142.

Abbri:1986:BRF

[15] Ferdinando Abbri. Book review: Frederic Lawrence Holmes, Lavoisier and the Chemistry of Life. An Exploration of Scientific Creativity, Madison, The University of Wisconsin Press 1985, 565 pp. (Wisconsin Publications in the History of Science and Medicine Number 4). Nuncius, 1 (1):216–220, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00151.

Abbri:1986:BRD

[16] Ferdinando Abbri. Book review: Dorothy A. Stansfield, Thomas Beddoes M.D. 1760–1808. Chemist, Physician, Democrat, Dordrecht–Boston– Lancaster, D. Reidel Publishing Company 1984, 306 pp. Nuncius, 1 (1):220–222, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00160.

Corsi:1986:BRDa

[17] Pietro Corsi. Book review: Dorinda Outram, Georges Cuvier. Vocation, Science and Authority in Post-Revolutionary France, Manchester, Manchester University Press 1984, 299 pp., £25. Nuncius, 1(1):222–225, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00179.

Belloni:1986:BRG

[18] Lanfranco Belloni. Book review: Guido Tagliaferri, Storia della fisica quantistica: Dalle origini alla meccanica ondulatoria, Milano, Franco Angeli 1985, 459 pp., L. 32.000. Nuncius, 1(1):226–227, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00188.

Miniati:1986:ISM

[19] Mara Miniati. Istituzioni scientifiche e musei. Nuncius, 1(1):229–230, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00197.

Miniati:1986:BRR

[20] Mara Miniati. Book review: Regione Marche — Centro Per I Beni Culturali, Guida ai Musei delle Marche, a cura di Pietro Zampetti, Milano, Electa 1985, 132 pp. Nuncius, 1(1):230–231, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00205.

Corsi:1986:BRU

[21] Pietro Corsi. Book review: Università di Bologna, Bologna. Musei di Veterinaria, Bologna, Edizioni Calderini 1985, vii-95 pp., no bibliography, no index of names, illustrations. Nuncius, 1(1):231, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00214.

Abbri:1986:SFM

[22] Ferdinando Abbri. Scienze fisico-matematiche. Nuncius, 1(1):232–233, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00223.

Caroti:1986:BRM

[23] Stefano Caroti. Book review: Mathemata. Festschrift für Helmuth Gericke, herausgegeben von Menso Folkerts und Uta Lindgren, Stuttgart, Franz Steiner Verlag 1985, 742 pp. (Boethius. Texte und Abhandlungen zur Geschichte der exacten Wissenschaften, Bd. XII). Nuncius, 1 (1):233–234, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00232.

Mazzolini:1986:MB

[24] Renato G. Mazzolini. Medicina e biologia. Nuncius, 1(1):235, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00241.

Mazzolini:1986:BRJ

[25] Renato G. Mazzolini. Book review: John D. Spillane, the Doctrine of the Nerves: Chapters in the history of neurology, Oxford, Oxford University Press 1981, xii + 468 pp., £25.00. Nuncius, 1(1):235–236, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00250.

Mazzolini:1986:BRM

[26] Renato G. Mazzolini. Book review: Mary A. B. Brazier, A History of Neurophysiology in the 17th and 18th Centuries: From Concept to Experiment, New York, Raven Press 1984, xiv + 230 pp., $81.00. Nuncius, 1(1):236, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00269.

Caroti:1986:MERa

[27] Stefano Caroti. Medio evo e rinascimento. Nuncius, 1(1):237, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00278.

Caroti:1986:BRE

[28] Stefano Caroti. Book review: Egbert P. Bos, Marsilius of Inghen: Treatises on the Properties of Terms. A First Critical Edition of the Suppositiones, Ampliationes, Appellationes, Restrictiones and Alienationes with Introduction, Translation, Notes and Appendices, Dordrecht–Boston– Lancaster, D. Reidel Publishing Company 1983, 274 pp., ind. dei nomi, dei termini e dei manoscritti, bibl., appendici (Sinthese Historical Library, 22). Nuncius, 1(1):238, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00287.

Caroti:1986:BRG

[29] Stefano Caroti. Book review: A. Goddu, The Physics of William Ockham, Leiden-Köln, E. J. Brill 1984, 243 pp., ind. nomi e materie (Studien und Texte zur Geistesgeschichte des Mittelalters, hrsg. von A. Zimmermann, B. XVI). Nuncius, 1(1):238–239, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00296.

Caroti:1986:BRJ

[30] Stefano Caroti. Book review: Jorge J. E. Gracia, Introduction to the Problem of Individuation in the Early Middle Ages, München-Wien, Philosophia Verlag 1984, 302 pp., bibl., ind. dei nomi, delle materie e dei termini (Analytica, Investigations in Logic, Ontology and the Philosophy of Language). Nuncius, 1(1):239–240, ????

1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00304.

Caroti:1986:BRPa

[31] Stefano Caroti. Book review: E. Paschetto, Pietro d’Abano medico e filosofo, Firenze, Nuovedizioni Enrico Vallecchi 1984, 358 pp., ind. dei nomi, delle fonti manoscritte, ill., appendici. Nuncius, 1(1):240, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00313.

Caroti:1986:BRQ

[32] Stefano Caroti. Book review: Les quatre éléments dans la culture médiévale. Actes du Colloque des 25, 26, 27 mars 1982, publiés par les soins de D. Buschinger et A. Crepin, Goeppingen, Kuemmerle Verlag 1983, 308 pp., ill. (Goeppinger Arbeiten zur Germanistik, Université de Picardie, Centre d’études Médiévales, 386). Nuncius, 1(1):241, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00322.

Corsi:1986:BRDb

[33] Pietro Corsi. Book review: Danielle Jacquart, Claude Thomasset, Sexualité et savoir médical au Moyen Age, Paris, Presses Universitaires de France 1985, 269 pp., no bibliography, no index of names. Nuncius, 1(1): 242, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00331.

Corsi:1986:BRK

[34] Pietro Corsi. Book review: Katharine Park, Doctors and Medicine in Early Renaissance Florence, Princeton, N.J., Princeton University Press, xii-298 pp., bibliography, index of names, index of doctors and their families, tables. Nuncius, 1(1):242–243, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00340.

Corsi:1986:BRA

[35] Pietro Corsi. Book review: Alberto Elena, Las quimeras de los cielos. Aspectos epistemológicos de la revolución copernicana, (Spanish) [Chimeras of heaven. Epistemological aspects of the Copernican revolution]. Madrid, Siglo XXI de España Editores 1985, viii-237 pp., bibliography pp. 197-227, index of matters and of names. Nuncius, 1(1):

243, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00359.

Mazzolini:1986:BRD

[36] Renato G. Mazzolini. Book review: Donald B. Tower, Hensing, 1719: An Account of the First Chemical Examination of the Brain and the Discovery of Phosphorus Therein, New York, Raven Press 1983, xvi + 407 pp., $47.50. Nuncius, 1(1):243–244, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00368.

Corsi:1986:SXX

[37] Pietro Corsi. Secoli XVII–XIX. Nuncius, 1(1):244, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539186x00377.

Corsi:1986:BRMa

[38] Pietro Corsi. Book review: M. Bortolotti, V. Lanzarini, Inventario del Museo ostetrico Giovan Antonio Galli, Bologna, Università degli Studi di Bologna, Archivio Storico 1985, Preliminary Publication in 300 copies, 41 pp., no index, no bibliography. Nuncius, 1(1):244–245, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00386.

Abbri:1986:BRA

[39] Ferdinando Abbri. Book review: Augusto Placanica, Il filosofo e la catastrofe. un terremoto del Settecento, Torino, Giulio Einaudi editore 1985, 244 pp. (Biblioteca di cultura storica 155). Nuncius, 1(1):245–246, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00395.

Corsi:1986:BRB

[40] Pietro Corsi. Book review: Bruno Belhoste, Cauchy 1789–1857. Un mathématicien légitimiste au XIXe siècle, (French) [Cauchy, 1789– 1857: a legitimist mathematician in the Nineteenth Century] Paris, Belin 1985, 224 pp., index, no bibliography (Collection Un savant, une époque, directed by J. Dhombres). Nuncius, 1(1):246–247, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00403.

Abbri:1986:BRL

[41] Ferdinando Abbri. Book review: Luigi Cerruti, Uomini ed idee della chimica classica, a cura di Carlo Bicchi, Milano, Eurobase S.p.A. 1985, 88 pp. Nuncius, 1(1):247–248, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00412.

Corsi:1986:BRP

[42] Pietro Corsi. Book review: Patrizia Bellardone, Miscellanea Quintino Sella. Citt di Biella, Biblioteca Civica 1984, 216 pp., no bibliography, no index. Nuncius, 1(1):248, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00421.

Corsi:1986:BRMb

[43] Pietro Corsi. Book review: Maria-José Baguena Cervellara et al., Estudios sobre la medicina y la ciencia valencianas. Siglos XVI–XIX, foreword by José López Piñero, Valencia, Catedra de Historia de la Medicina 1985, pp. 193, bibliography and index of names; pp. 179–193, bibliography of studies in the history of science and medicine produced by the history of medicine group at The University of Valencia (Cuadernos Valencianos de Historia de la Medicina y de la Ciencia, XXVIII Serie A — Monografas) [ISBN 84-600-4005-4]. Nuncius, 1(1):248–249, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00430.

Principe:1986:BRV

[44] Ilario Principe. Book review: Vladimiro Valerio, l’Italia nei manoscritti dell’Officina Topografica conservatinella Biblioteca Nazionale di Napoli, Napoli, Istituto Italiano per gli Studi Filosofici 1985, 112 pp., s.i.p. (il volume pu essere ottenuto direttamente dall’Istituto, viale Calascione 7, 80132 Napoli). Nuncius, 1(1):249–250, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00449.

Corsi:1986:BRJ

[45] Pietro Corsi. Book review: José M. Lopez Piero et. al., Cajal y la medicina Valenciana de su tiempos, 1833–1887. (Spanish) [Cajal and the medicine of his time, Valencia, 1833–1887]. Guida de la expsicion 5–21 diciembre 1983, Valencia, Cultura Universitria Popular, 40 pp. Nuncius, 1(1):250–251, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00458.

Anonymous:1986:SSS

[46] Anonymous. Studi sovietici di storia della scienza, 1985 (a cura dell’Accademia delle Scienze dell’URSS). Nuncius, 1(1):251–256, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00467.

Miniati:1986:ADMa

[47] Mara Miniati. Attività dell’Istituto e Museo di Storia Della Scienza di Firenze. Nuncius, 1(1):257–xii, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00476.

Miniati:1986:VDI

[48] Mara Miniati. Vent’anni dopo: Immagine di un recupero. Nuncius, 1(2):iii–5, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00485.

Neville:1986:PCC

[49] Pamela Neville. The Printer’s copy of Commandino’s translation of Archimedes, 1558. Nuncius, 1(2):7–12, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539186x00494.

Gozza:1986:MNF

[50] Paolo Gozza. La musica nella filosofia naturale del seicento in Italia. Nuncius, 1(2):13–47, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00502.

Torrini:1986:VCN

[51] Maurizio Torrini. ≪ et vidi coelum novum et terram novam ≫. Nuncius, 1(2):49–77, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00511.

Goldstein:1986:MAE

[52] Bernard R. Goldstein. The making of astronomy in early Islam. Nuncius, 1(2):79–92, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00520.

Borrelli:1986:CDC

[53] Antonio Borrelli. Carteggio di Domenico Cotugno. Nuncius, 1(2):93– 101, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00539.

Weindling:1986:GSC

[54] Paul Weindling. German–Soviet co-operation in science: the case of the Laboratory for Racial Resarch, 1931–1938. Nuncius, 1(2):103–109, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00548.

Mazzolini:1986:CBI

[55] Renato G. Mazzolini. Commission on bibliography of the International Union of the History and Philosophy of Science. Nuncius, 1(2):111–137, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00557.

Corsi:1986:BRR

[56] Pietro Corsi. Book review: Robert M. Young, Darwin’s Metaphor. Nature’s Place in Victorian Culture, Cambridge, Cambridge University Press 1985, xx + 341 pp. Nuncius, 1(2):139–146, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00566.

Dougherty:1986:BRG

[57] F. W. P. Dougherty. Book review: Gerrit A. Lindeboom, Dutch Medical Biography. A biographical Dictionary of Dutch Physicians and Surgeons 1475–1975. Amsterdam, Rodopi 1984, xiii + 2243 columns (i.e. 1122 pp. + xxxi), Hfl. 295, ISBN 90-6203-676-7. Nuncius, 1(2):147–148, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00575.

Caroti:1986:BRPb

[58] Stefano Caroti. Book review: Paulus Venetus, Logica Parva, Translation of the 1472 Edition with Introduction and Notes by Alan R. Perreiah, the Catholic University of America Press, Philosophia Verlag, Washington– München, Wien 1984 (≪ Analytica. Investigation in Logic, Ontology and the Philosophy of Language ≫ ), 372 pp., bibl. ind. nomi e soggetti. Nuncius, 1(2):148, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00584.

Baroncini:1986:BRJ

[59] Gabriele Baroncini. Book review: Jacopo Zabarella, De methodis libri quatuor. Liber de regressu, a cura di C. Vasoli, Bologna, Clueb 1985, xxviii + 193 pp. Nuncius, 1(2):149–153, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539186x00593.

Abbri:1986:BRP

[60] Ferdinando Abbri. Book review: Peter Alexander, Ideas, Qualities and Corpuscles. Locke and Boyle on the external world, Cambridge, Cambridge University Press 1985, 336 pp. Nuncius, 1(2):153–155, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00601.

Abbri:1986:BRW

[61] Ferdinando Abbri. Book review: Walter Bernardi, Le Metafisiche dell’Embrione. Scienze della vita e filosofiada Malpighi a Spallanzani (1672–1793), Firenze, Leo S. Olschki Editore 1986, 503 pp. (≪ Biblioteca di storia della scienza n. 24 ≫ ). Nuncius, 1(2):156–158, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00610.

Abbri:1986:BRS

[62] Ferdinando Abbri. Book review: Stefan Winkle, Johann Friedrich Struensee. Arzt, Aufklärer und Staatsmann. Beitrag zur Kultur-, Medizin- und Seuchengeschichte der Aufklärungszeit, Stuttgart, Gustav Fischer Verlag 1983, 655 pp. Nuncius, 1(2):158–160, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00629.

Brenni:1986:BRP

[63] Paolo Brenni. Book review: Peter R. De Clercc (Ed.), Nineteenth-Century Scientific Instruments and their Makers, Papers presented at the Fourth Scientific Instrument Symposium, Amsterdam 23–24 October 1984. Leiden 1985, Rodopi Amsterdam 1985. (≪ Nieuwe Nederlandse Bijdragen tot de Geschiedenis der Geneeskunde en der Natuurwetenschappen N^{∘} 17 ≫ ). Nuncius, 1(2):161–163, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00638.

Dougherty:1986:BRI

[64] F. W. P. Dougherty. Book review: Ingeborg Schnack (ed.), Der Briefwechsel zwischen Friedrich Carl von Savigny und Stephan August Winkelmann (1800–1804) mit Dokumenten und Briefen aus dem Freundeskreis gesammelt, herausgegeben und kommentiert von Ingeborg Schnack. (Veröffentlichungen der Historischen Kommission für Hessen,

23. Hessische Briefe des 19. Jahrhunderts, vol. 3) Marburg, N. G. Elwert 1984, 533 pp. + 4 illustrations. Cloth: DM 95,–; paper: DM 90,–. Nuncius, 1(2):163–165, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00647.

LaVergata:1986:BRM

[65] Antonello La Vergata. Book review: Mario A. di Gregorio, T. H. Huxley’s Place in Natural Science, New Haven and London, Yale University Press 1984, xxi-253 pp. Nuncius, 1(2):165–167, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00656.

Nicasi:1986:BRP

[66] Stefania Nicasi. Book review: Patrizia Guarnieri, Individualità difformi, Milano, Franco Angeli 1986. Nuncius, 1(2):167–170, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00665.

Guarnieri:1986:BRS

[67] Patrizia Guarnieri. Book review: Sigmund Freud, Sintesi delle nevrosi di traslazione, un manoscritto inedito a cura e con un saggio di Ilse Grubrich-Simitis, Torino, Boringhieri 1968, 121 pp., L. 20.000 (trad. it. di Übersicht der Übertragungsneurosen, Frankfurt, Fischer Verlag 1985). Nuncius, 1(2):170–172, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00674.

Weindling:1986:BRW

[68] Paul Weindling. Book review: Willie Pearson, Jr., Black Scientists, White Society, and Colorless Science: A Study of Universalism in American Science, Millwood, Ny, New York City and London: Associated Faculty Press 1985, viii + 201 pp. Nuncius, 1(2):173–174, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00683.

Abbri:1986:SB

[69] Ferdinando Abbri. Strumenti bibliografici. Nuncius, 1(2):175, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00692.

Carotin:1986:BRC

[70] Stefano Carotin. Book review: Catalogo di manoscritti filosofici nelle biblioteche italiane, vol. 5: Cesena, Cremona, Lucca, S. Daniele del Friuli, Teramo, Terni, Trapani, Udine, Firenze, Leo S. Olschki Editore 1985 (Unione Accademica Nazionale. ≪ Corpus Philosophorum Medii Aevii, Subsidia ≫, V), 365 pp., indice degli incipit, nomi e opere anonime. Nuncius, 1(2):176, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00700.

Miniati:1986:BRI

[71] Mara Miniati. Book review: Istituto di Fisica Generale Applicata — Storia della Fisica, 1726–1799 Catalogo della corrispondenza degli Astronomi di Brera, volume I, a cura di Agnese Mandrino, Guido Tagliaferri, Pasquale Tucci. Milano, Università degli Studi di Milano 1986, 427 pp. Nuncius, 1(2):176–177, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00719.

Caroti:1986:BRF

[72] Stefano Caroti. Book review: Francesca Sofia, Una biblioteca ginevrina del Settecento: I libri del giovane Sismondi, Roma, Edizioni dell’Ateneo e Bizzarri 1983 (Istituto di Storia Moderna. Facoltà di Lettere e Filosofia. Università di Roma. ≪ Storia ≫, 11), 297 pp., indice degli autori. Nuncius, 1(2):177–178, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00728.

Miniati:1986:S

[73] Mara Miniati. Strumentaria. Nuncius, 1(2):178, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539186x00737.

Brenni:1986:BRT

[74] Paolo Brenni. Book review: Il teatro di filosofia sperimentale di Giovanni Poleni, a cura di Gian Antonio Salandin e Maria Pancino, Padova, edizioni Lint 1986. Nuncius, 1(2):178–179, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00746.

Brenni:1986:BRE

[75] Paolo Brenni. Book review: Ezio Ragozzino, Edvige Schettino, La collezione degli antichi apparecchi dell’Istituto di Fisica: Gli strumenti ottici (1840–1890), Napoli, Università degli Studi di Napoli, Facoltà di Scienze M.F.N. 1984 Meccanica dei fluidi e termologia (1840–1900), Napoli, Università degli Studi di Napoli, Facoltà di Scienze M.F.N. 1984 Elettricità e magnetismo (1835–1900), Napoli, Università degli Studi di Napoli, Facoltà di Scienze M.F.N. 1985. Nuncius, 1(2):179–180, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00755.

Caroti:1986:SMF

[76] Stefano Caroti. Scienze matematico-fisiche. Nuncius, 1(2):180–181, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00764.

Abbri:1986:BRM

[77] Ferdinando Abbri. Book review: Margaret Alic, Hypatia’s Heritage. A History of Women in Science from Antiquity to the Late Nineteenth Century, London, the Women’s Press 1986, X + 230 pp. Nuncius, 1 (2):181–182, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00773.

Caroti:1986:MB

[78] Stefano Caroti. Medicina e biologia. Nuncius, 1(2):182–183, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00782.

Guarnieri:1986:BRV

[79] Patrizia Guarnieri. Book review: Vittorio D. Catapano, Le Reali Case de’ Matti nel Regno di Napoli, Napoli, Liguori 1986, 351 pp., L. 28.000. Nuncius, 1(2):183, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00791.

Guarnieri:1986:BRE

[80] Patrizia Guarnieri. Book review: Edward Shorter, La tormentata storia del rapporto medico-paziente, Milano, Feltrinelli 1986, 231 pp., L. 27.000 (trad. it. di Bedside Manners. The Troubled History of Doctors and Patients, New York, Simon and Schuster 1985). Nuncius, 1(2):184, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00809.

Guarnieri:1986:BRM

[81] Patrizia Guarnieri. Book review: Mario Galzigna, Conoscenza e dominio. Le scienze della vita tra filosofia e storia, post-fazione di Jacques Roger, Verona, Bertani 1985, 226 pp., L. 15.000. Nuncius, 1(2):185, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00818.

Ragazzini:1986:B

[82] Stefania Ragazzini. Botanica. Nuncius, 1(2):186, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539186x00827.

Ragazzini:1986:BRE

[83] Stefania Ragazzini. Book review: Erbari e Iconografia botanica. Storia delle Collezioni dell’Orto Botanico dell’Università di Torino. A cura di Franco Montacchini con la collaborazione di Giuliana Forneris e Carla Martoglio. Appendice a cura di Silvana Pettenati, Torino, U. Allemandi e C. 1986, Catalogo della Mostra, Torino, febbraio–maggio 1986. Nuncius, 1(2):187, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539186x00836.

Caroti:1986:MERb

[84] Stefano Caroti. Medio evo e rinascimento. Nuncius, 1(2):188, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00845.

Caroti:1986:BRH

[85] Stefano Caroti. Book review: Henry Chadwick, Boezio. La consolazione della musica, della logica, della teologia, della filosofia, Bologna, Società editrice il Mulino 1986 (≪ Collana di testi e studi. Filosofia ≫ ), 358 pp., bibliografia. Nuncius, 1(2):189, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00854.

Caroti:1986:BRPc

[86] Stefano Caroti. Book review: Per-Gunnar Ottoson, Scholastic Medicine and Philosophy, Napoli, Bibliopolis 1984, 322 pp., bibliografia, index rerum. Nuncius, 1(2):189–190, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00863.

Abbri:1986:SXX

[87] Ferdinando Abbri. Secoli XVII–XIX. Nuncius, 1(2):190–191, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00872.

Miniati:1986:BRF

[88] Mara Miniati. Book review: Fabio Todesco, Riforma della metafisica e sapere scientifico: saggio su J. H. Lambert (1728–1777). Milano, Franco Angeli, 330 pp. (in corso di stampa). Nuncius, 1(2):191, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539186x00881.

Abbri:1986:BRJ

[89] Ferdinando Abbri. Book review: John L. Heilbron, Alle origini della fisica moderna. Il caso dell’elettricità, Bologna, il Mulino 1984, 454 pp. Nuncius, 1(2):192–198, ???? 1986. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539186x00890.

Miniati:1986:ADMb

[90] Mara Miniati. Attivita dell’Istituto e Museo di Storia della Scienza di Firenze. Nuncius, 1(2):199–215, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539186x00908.

Scriba:1986:A

[91] C. J. Scriba. Announcement. Nuncius, 1(2):216–xii, ???? 1986. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539186x00917.

Baldini:1987:AG

[92] Ugo Baldini. Additamenta Galilaeana. Nuncius, 2(1):3–36, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00179.

Meinel:1987:NMA

[93] Christoph Meinel. Natur Als Moralische Anstalt Die Meteorologia Philosophico-Politica Des Franz Reinzer, S.J., Ein Naturwissenschaftliches Emblembuch Aus Dem Jahre 1698. Nuncius, 2(1):37–94, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00188.

Turner:1987:GEZ

[94] A. J. Turner. Greco–Egyptian zodiacs from a Gallo–Romano site. Nuncius, 2(1):95–110, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539177x00197.

Schettino:1987:PAD

[95] Edvige Schettino. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 2(1):111–124, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539177x00205.

Mazzolini:1987:IF

[96] Renato G. Mazzolini. Istituzioni e fonti. Nuncius, 2(1):125–134, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00214.

Anonymous:1987:GS

[97] Anonymous. I gruppo di studenti. Nuncius, 2(1):134, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00223.

Anonymous:1987:SDW

[98] Anonymous. 2. Schmidt, Dietrich Wilhelm. Nuncius, 2(1):135, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00232.

Anonymous:1987:MMJ

[99] Anonymous. 3. Meier (Meyer), Johann Ludolph Christian. Nuncius, 2(1):135, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x00241.

Anonymous:1987:BGG

[100] Anonymous. 4. Bielcke, Georg Gotthelf. Nuncius, 2(1):136, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00250.

Anonymous:1987:BCM

[101] Anonymous. 5. Busmann, Conrad Moritz Christian. Nuncius, 2(1):136, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00269.

Anonymous:1987:MJN

[102] Anonymous. 6. Marcard, Jacob Nicolaus. Nuncius, 2(1):137, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00278.

Anonymous:1987:RCJ

[103] Anonymous. 7. Rollin, Christian Jeremias. Nuncius, 2(1):137–138, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00287.

Anonymous:1987:THW

[104] Anonymous. 8. Taube, Hartwig Wilhelm Ludwig. Nuncius, 2(1):138, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00296.

Anonymous:1987:FC

[105] Anonymous. 9. Fickel, Christoph. Nuncius, 2(1):139, ???? 1987. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539177x00304.

Anonymous:1987:BML

[106] Anonymous. 10. Berckelmann, Matthias Ludwig Rudolph. Nuncius, 2(1):139, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x00313.

Anonymous:1987:WWA

[107] Anonymous. 11. Winkler (Winckler), Adolph Bernhard. Nuncius, 2(1): 140, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00322.

Anonymous:1987:DFA

[108] Anonymous. 12. Denecke, Friedrich August. Nuncius, 2(1):140–141, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00331.

Anonymous:1987:HAL

[109] Anonymous. 13. Hugo, August Ludwig Von. Nuncius, 2(1):141, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00340.

Anonymous:1987:OJH

[110] Anonymous. 14. Oschwald, Johannes Heinrich. Nuncius, 2(1):141–142, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00359.

Anonymous:1987:LGE

[111] Anonymous. 15. Loeber, Gotthold Ernst. Nuncius, 2(1):142, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00368.

Anonymous:1987:WWC

[112] Anonymous. 16. Willig (Willich), Christian Ludwig. Nuncius, 2(1):142– 143, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00377.

Anonymous:1987:TDJ

[113] Anonymous. 17. Taube, Daniel Johann. Nuncius, 2(1):143, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00386.

Anonymous:1987:SSG

[114] Anonymous. 18. Schmitt (Schmidt), Georg Conrad (Konrad). Nuncius, 2(1):144, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x00395.

Anonymous:1987:SDC

[115] Anonymous. 19. Schóbinger, David Christoph. Nuncius, 2(1):144–145, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00403.

Anonymous:1987:OA

[116] Anonymous. 20. Orville, Abraham D’. Nuncius, 2(1):145–146, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00412.

Anonymous:1987:MJF

[117] Anonymous. 22. Meckel, Johann Friedrich. Nuncius, 2(1):146–147, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00421.

Anonymous:1987:VJ

[118] Anonymous. 23. Vosse, Joachim. Nuncius, 2(1):147, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539177x00430.

Anonymous:1987:LD

[119] Anonymous. 24. Langhans, Daniel. Nuncius, 2(1):147–148, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00449.

Anonymous:1987:WJJ

[120] Anonymous. 25. Walbaum, Johann Julius. Nuncius, 2(1):148–149, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00458.

Anonymous:1987:ZJGa

[121] Anonymous. 26. Zinn, Johann Gottfried. Nuncius, 2(1):149–150, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00467.

Anonymous:1987:NJ

[122] Anonymous. 27. Noreen, Johannes. Nuncius, 2(1):150, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00476.

Anonymous:1987:SCG

[123] Anonymous. 28. Steding, Carl Gerhard. Nuncius, 2(1):150–151, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00485.

Anonymous:1987:AS

[124] Anonymous. 29. Aurivillius, Samuel. Nuncius, 2(1):151, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00494.

Anonymous:1987:KFW

[125] Anonymous. 30. Klaerich, Friedrich Wilhelm. Nuncius, 2(1):151–152, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00502.

Anonymous:1987:AGT

[126] Anonymous. 31. Asch, Georg Thomas Von. Nuncius, 2(1):152–153, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00511.

Anonymous:1987:AJM

[127] Anonymous. 32. Albrecht, Johann Melchior Friederich. Nuncius, 2(1): 153, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00520.

Anonymous:1987:ZJGb

[128] Anonymous. 33. Zimmermann, Johann Georg. Nuncius, 2(1):153–154, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00539.

Anonymous:1987:HN

[129] Anonymous. 34. Himsel, Nicolaus Von. Nuncius, 2(1):154–155, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00548.

Anonymous:1987:SPJ

[130] Anonymous. 35. Stirtz, Peter Joseph. Nuncius, 2(1):155, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00557.

Anonymous:1987:RGE

[131] Anonymous. 36. Remus, Georg Ernst. Nuncius, 2(1):156, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00566.

Anonymous:1987:GO

[132] Anonymous. 37. Gesenius, Otto. Nuncius, 2(1):156, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539177x00575.

Anonymous:1987:SA

[133] Anonymous. 38. Swainston, Allen. Nuncius, 2(1):157, ???? 1987. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539177x00584.

Anonymous:1987:CP

[134] Anonymous. 39. Castell, Peter. Nuncius, 2(1):157–158, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00593.

Anonymous:1987:SJA

[135] Anonymous. 40. Sproegel, Johann Adrian. Nuncius, 2(1):158, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00601.

Anonymous:1987:HJA

[136] Anonymous. 41. Hammerschmidt, Johannes Andreas. Nuncius, 2(1): 158–159, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x00610.

Anonymous:1987:WWW

[137] Anonymous. 42. Walstorff (Walstorf, Walsdorf), Johann Dietrich. Nuncius, 2(1):159, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x00629.

Anonymous:1987:VVJ

[138] Anonymous. 43. Vaetterli (Vtterli), Johann Jacob. Nuncius, 2(1):159– 160, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00638.

Anonymous:1987:BJH

[139] Anonymous. 44. Brunn, Johann Heinrich Von. Nuncius, 2(1):160, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00647.

Anonymous:1987:DDD

[140] Anonymous. 45. Dehtleef (Detleef, Dehtleef), Peter. Nuncius, 2(1):161, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00656.

Anonymous:1987:KJC

[141] Anonymous. 46. Kuhlemann, Johann Christoph. Nuncius, 2(1):161–162, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00665.

Anonymous:1987:RJJ

[142] Anonymous. 47. Rhades, Joachim Jacob. Nuncius, 2(1):162–163, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00674.

Anonymous:1987:KCL

[143] Anonymous. 2. Kohen (Cohn), Lehmann Isaac. Nuncius, 2(1):163, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00683.

Anonymous:1987:CFL

[144] Anonymous. 3. Cropp, Friedrich Ludwig Christian. Nuncius, 2(1):164, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00692.

Anonymous:1987:GJJ

[145] Anonymous. 4. Grumbrecht, Johannes Justus. Nuncius, 2(1):164–165, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00700.

Anonymous:1987:FJ

[146] Anonymous. 5. Frsen, Joachim. Nuncius, 2(1):165, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539177x00719.

Anonymous:1987:PGC

[147] Anonymous. 6. Pappelbaum, Georg Christoph. Nuncius, 2(1):166, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00728.

Anonymous:1987:RAJ

[148] Anonymous. 7. Richers, Arnold Julius Johann. Nuncius, 2(1):166–167, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00737.

Anonymous:1987:BJJ

[149] Anonymous. 8. Berger, Johann Justus Von. Nuncius, 2(1):167–168, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00746.

Anonymous:1987:HRS

[150] Anonymous. 10. Henrici, Robert Stephan. Nuncius, 2(1):168, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00755.

Anonymous:1987:OGC

[151] Anonymous. 11. Oeder, Georg Christian Von. Nuncius, 2(1):168–169, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00764.

Anonymous:1987:FFJ

[152] Anonymous. 12. Foelix (Felix), Johannes Jacob. Nuncius, 2(1):169, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00773.

Anonymous:1987:RRC

[153] Anonymous. 13. Roloff (Rolof), Christian Ludwig. Nuncius, 2(1):170, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00782.

Anonymous:1987:DJF

[154] Anonymous. 14. Droysen, Julius Friedrich. Nuncius, 2(1):170–171, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00791.

Anonymous:1987:RJG

[155] Anonymous. 15. Runge, Johann Georg. Nuncius, 2(1):171, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00809.

Anonymous:1987:EEJ

[156] Anonymous. 16. Evers, Emanuel Johannes Albert. Nuncius, 2(1):171, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00818.

Anonymous:1987:DDA

[157] Anonymous. 17. Duntze (Dunze), Arnold. Nuncius, 2(1):172, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00827.

Anonymous:1987:AKC

[158] Anonymous. 19. Andersch, Karl (Carl) Samuel. Nuncius, 2(1):173–192, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00836.

Anonymous:1987:IAD

[159] Anonymous. Indice alfabetico degli studenti. Nuncius, 2(1):194, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00845.

EstacioDosReis:1987:LOG

[160] Antonio Estacio Dos Reis. List of old globes in Portugal. Nuncius, 2 (1):195–201, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x00854.

Hackmann:1987:DC

[161] Willem D. Hackmann. Discussioni critiche. Nuncius, 2(1):203–213, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00863.

Lanfredini:1987:CAS

[162] Roberta Lanfredini and Stefano Casati. Corso di aggiornamento sulla filosofia della scienza. Nuncius, 2(1):215–219, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539177x00872.

Pogliano:1987:BRTa

[163] Claudio Pogliano. Book review: Time, Science, and Society in China and the West. The Study of Time V, ed. by J. T. Fraser, N. Lawrence and F. C. Haber, Amherst, The University of Massachusetts Press 1986, 262 pp. Nuncius, 2(1):221–223, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539177x00881.

Abbri:1987:BRM

[164] Ferdinando Abbri. Book review: Margaret J. Osler, Paul Lawrence Farber (editors), Religion, Science, and Worldview. Essays in Honor of Richard S. Westfall, Cambridge, Cambridge University Press 1985, xv + 350 pp. Nuncius, 2(1):224–226, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539177x00890.

Pereira:1987:BRA

[165] Michela Pereira. Book review: Alchimia. La tradizione in Occidente secondo le fonti manoscritte e a stampa, a cura di Mino Gabriele, Edizioni La Biennale — Realizzazione Electa Editrice 1986, 90 pp., L. 30.000. Nuncius, 2(1):226–230, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539177x00908.

Biagioli:1987:BRB

[166] Mario Biagioli. Book review: Brian Vickers (ed.), Occult and Scientific Mentalities in the Renaissance, Cambridge, Cambridge University Press 1984, 408 pp. Nuncius, 2(1):230–236, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539177x00917.

Abbri:1987:BRJa

[167] Ferdinando Abbri. Book review: Jerry Weinberger, Science, Faith, and Politics. Francis Bacon and the Utopian Roots of the Modern Age. A Commentary on Bacon’s Advancement of Learning, Ithaca and London, Cornell University Press 1985, 342 pp. Nuncius, 2(1):236–241, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00926.

Abbri:1987:BRSa

[168] Ferdinando Abbri. Book review: Steven Shapin, Simon Schaffer, Leviathan and the Air-Pump. Hobbes, Boyle, and the Experimental Life, Princeton, Princeton University Press 1985, xiv + 440 pp. Nuncius, 2 (1):241–244, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x00935.

Giuntini:1987:BRG

[169] Chiara Giuntini. Book review: Guenter B. Risse, Hospital life in Enlightment Scotland. Care and teaching at the Royal Infirmary of Edinburgh, Cambridge University Press 1986 (Cambridge History of Medicine). Nuncius, 2(1):245–247, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539177x00944.

Casati:1987:BRR

[170] Stefano Casati. Book review: Roberto Finzi (a cura di), Le meteore e il frumento: clima, agricoltura, meteorologia a Bologna nel ’700, Bologna, il Mulino 1986, 387 pp. Nuncius, 2(1):247–249, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00953.

Mazzolini:1987:BRL

[171] Renato G. Mazzolini. Book review: Lazzaro Spallanzani, Edizione nazionale delle opere. Parte prima Carteggi, vol. I, Carteggi con Abboretti … Bondi, a cura di Pericle di Pietro, Modena, Enrico Mucchi Editore 1984; vol. II, Carteggio con Charles Bonnet …, 1984; vol. III, Carteggi con Borbone … Comino …, 1985; vol. Iv, Carteggi con Comparetti … Fortis …, 1985; vol. V, Carteggi con Fossombroni … Lucchesini …, 1985. Nuncius, 2(1):249–252, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00962.

Dougherty:1987:BRH

[172] F. W. P. Dougherty. Book review: Heusser Peter, Der Schweizer Arzt und Philosoph Ignaz Paul Vital Troxler (1780–1866). Seine Philosophie, Anthropologie und Medizintheorie (Basler Veröffentlichungen zur Geschichte der Medizin und der Biologie, hrsg. von H. Buess, Fasc. XXXIX), Basel/Stuttgart, Schwabe Co. 1984, xvi + 350 pp., 1 illustr. Paper SFR 38/DM 46. Nuncius, 2(1):252–253, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00971.

Dougherty:1987:BRF

[173] F. W. P. Dougherty. Book review: Friedrich Gottlieb Welker, Werk und Wirkung. Vorträge, gehalten auf der Welcker-Tagung in der Werner-Reimers-Stiftung in Bad Homburg vom 5.–7. November 1984 (I) und zur Eröffnung der Ausstellung Friedrich Gottlieb Welcker (1784–1868): Philologe, Archäologe und Oberbibliothekar in Bonn in der Universitätsbibliothek Bonn am 5. November 1984 (II) herausgegeben von William M. Calder III, Adolf Kühnken, Wolfgang Kullmann, Günther Pflugg. (Hermes. Zeitschrift für klassische Philologie, Einzelschriften, Heft 49). Stuttgart, Franz Steiner Verlag 1986, viii + 293 pp. + illustrations. Paper DM 58,-. Nuncius, 2(1):253–255, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00980.

Greene:1987:BRC

[174] John C. Greene. Book review: Charlotte M. Porter, The Eagle’s Nest. Natural History and American Ideas, 1812–1842, Alabama, University of Alabama Press 1986, 251 pp. illustrated. Nuncius, 2(1):255–257, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x00999.

Wanrooij:1987:BRJ

[175] Bruno P. F. Wanrooij. Book review: Jacques Lonard, Archives du corps. La sant au XIXe siècle, Rennes, d. Ouest France 1986, 329 pp. Nuncius, 2(1):257–258, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x01006.

Richards:1987:BRS

[176] Joan L. Richards. Book review: Sean O’Donnell, William Rowan Hamilton: Portrait of a Prodigy, Boole Press Dublin 1983. Nuncius, 2(1): 258–259, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x01015.

Wanrooij:1987:BRS

[177] Bruno P. F. Wanrooij. Book review: Sander L. Gilman, Difference and Pathology. Stereotypes of Sexuality, Race, and Madness, Ithaca and London, Cornell University Press 1985. Nuncius, 2(1):259–262, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01024.

Wanrooij:1987:BRV

[178] Bruno P. F. Wanrooij. Book review: Valeria P. Babini, Fernanda Minuz, Annamaria Tagliavini, La donna nelle scienze dell’uomo. Immagini del femminile nella cultura scientifica italiana di fine secolo, Milano, F. Angeli 1986, 175 pp. Nuncius, 2(1):263, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01033.

Mazzolini:1987:SB

[179] Renato G. Mazzolini. Strumenti bibliografici. Nuncius, 2(1):265–266, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01042.

Bucciantini:1987:BRB

[180] Massimo Bucciantini. Book review: La biblioteca periodica. Repertorio dei giornali letterari del Sei–Settecento in Emiliae in Romagna. Volume primo: 1668–1726, (Italian) [The periodical library. Repertoire of literary journals of the Sixteenth–Eighteenth Century in Emiliae in Romagna. Volume One: 1668–1726], a cura di Martino Capucci, Renzo Cremante e Giovanna Gronda, Bologna, Il Mulino 1985, 550 pp. (Cultura e vita civile nel Settecento). Nuncius, 2(1):266–267, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01051.

Bucciantini:1987:BRE

[181] Massimo Bucciantini. Book review: Le edizioni dei testi filosofici e scientifici del’500 e del’600. Problemi di metodo e prospettive di ricerca, a cura di G. Canziani e G. Paganini, Milano, Angeli 1986, 228 pp. Nuncius, 2(1):267–268, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539177x01060.

Bucciantini:1987:BRR

[182] Massimo Bucciantini. Book review: Robin E. Rider — Henry E. Lowood, Guide to sources in Northern California for history of science and technology, Berkeley, Office for History of Science and Technology, University of California 1985, iv, 194 pp. (Berkeley Papers in History of Science, 10). Nuncius, 2(1):268–269, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539177x01079.

Mazzolini:1987:BRW

[183] Renato G. Mazzolini. Book review: Werner Sackmann, Biographische und bibliographische Materialien zur Geschichteder Mikrobiologie un zur bakteriologischen Nomenklatur (Marburger Schriften zur Medizingeschichte), Bd. 16, Frankfurt am Main, Bern, Verlag Peter Lang 1985, 412 pp. Nuncius, 2(1):269, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539177x01088.

Miniati:1987:BRE

[184] Mara Miniati. Book review: Emilie Savage-Smith, Islamicate Celestial Globes: their history, construction and use — with a chapter on iconography by Andrea P. A. Belloli, Washington, D.C., Smithsonian Institution Press, 1985, 354 pp. (Smithsonian Studies in History and Tecnology — Number 46). Nuncius, 2(1):270, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539177x01097.

Miniati:1987:BRF

[185] Mara Miniati. Book review: Futurinsieme. Percorso storico di tecnologie cinematografiche e televisive in vista di un futurocomune, Milano, Ente Fiera Milano 1986, 63 pp. Nuncius, 2(1):270–271, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01105.

Pogliano:1987:BRJ

[186] Claudio Pogliano. Book review: J. Needham, Scienza e civilt in Cina. Vol. III: la matematica e le scienze del cielo e della terra, Torino, Einaudi 1985-86, pp. 1115, in due tomi, L. 60.000 + 70.000. Nuncius, 2(1): 271, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01114.

Miniati:1987:BRD

[187] Mara Miniati. Book review: Dino Bravieri, Scampoli, Firenze, Osservatorio Ximeniano 1986, 235 pp. Nuncius, 2(1):272, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01123.

Caroti:1987:BRW

[188] Stefano Caroti. Book review: Wilfried Schroeder, Das Phaenomen des Polarlichts. Geschichtsschreibung, Forschungsergebnisse und Probleme, Wissenschaftliche Buchgesellschaft, Darmstadt 1984 (Ertraege der Forschung, Bd. 218), 156 pp., bibl., ind., ill. Nuncius, 2(1):272–273, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01132.

Abbri:1987:BRSb

[189] Ferdinando Abbri. Book review: Storicità e attualità della cultura scientifica einsegnamento delle scienze, a cura del C.I.D.I. di Firenze, Firenze, Marietti-Manzuoli 1986, 260 pp. Nuncius, 2(1):273, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01141.

Mazzolini:1987:BRM

[190] Renato G. Mazzolini. Book review: Margaret Rowbottom and Charles Susskind, Electricity and Medicine: History of their Interaction, San Francisco, San Francisco Press 1984, 303 pp. Nuncius, 2(1):274, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01150.

LaVergata:1987:BRP

[191] Antonello La Vergata. Book review: Peter J. Bowler, Evolution: the history of an idea, Berkeley, Los Angeles, London, University of California Press 1984, xiv, 412 pp., $11,95. Nuncius, 2(1):274–275, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01169.

Caroti:1987:BRRa

[192] Stefano Caroti. Book review: Richard W. Southern, Robert Grosseteste. the Growth of an English Mind in Medieval Europe, Oxford, Clarendon Press 1986, 337 pp., bibl., ind. Nuncius, 2(1):275–276, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01178.

Caroti:1987:BRH

[193] Stefano Caroti. Book review: Heinrich Schipperges, Der Garten Der Gesundheit. Medizin im Mittelalter, München-Zürich, Artemis Verlag 1985, 296 pp., bibl., ill. Nuncius, 2(1):276–277, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01187.

Caroti:1987:BRAa

[194] Stefano Caroti. Book review: Ann G. Carmichael, Plague and the poor in Renaissance Florence, Cambridge–London–New York–New Rochelle– Melbourne–Sidney, Cambridge University Press 1986 (Cambridge History of Medicine, eds. Charles Wabster and Charles Rosemberg), 180 pp., bibl., ind. Nuncius, 2(1):277–278, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539177x01196.

Abbri:1987:BRRa

[195] Ferdinando Abbri. Book review: A. Rupert Hall, La rivoluzione nella scienza. 1500–1750, Milano, Feltrinelli 1986, 438 pp. Nuncius, 2(1):278, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01204.

Caroti:1987:BRJ

[196] Stefano Caroti. Book review: John Cottingham, Descartes, Oxford, Basil Blackwell 1986, 171 pp., bibl., ind. Nuncius, 2(1):278–279, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01213.

Abbri:1987:BRPa

[197] Ferdinando Abbri. Book review: Una politica per le Terme: Montecatini e la Val di Nievole nelle riforme di Pietro Leopoldo. Atti del Convegno di studi, Siena, Edizioni Periccioli 1985, 275 pp., ill. Nuncius, 2(1): 279–280, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539177x01222.

Caroti:1987:BRN

[198] Stefano Caroti. Book review: Nicolas Rousseau, Connaissance et langage chez Condillac, Genve, Librairie Droz 1986 (Histoire des ides et critique littraire, vol. 242), 467 pp., bibl., ind. Nuncius, 2(1):280–281, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01231.

Miniati:1987:BRT

[199] Mara Miniati. Book review: Il terremoto di Riminie della costa romagnola: 25 dicembre 1786 — Analisi e interpretazione. A cura di Emanuela Guidoboni, Graziano Ferrari, Bologna, SGA (Società di Geofisica Applicata, s.a.) 1986, 293 pp. Nuncius, 2(1):281, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01240.

Mazzolini:1987:BRA

[200] Renato G. Mazzolini. Book review: Aa.Vv., Documenti per una storia della scienza senese, Siena, Accademia delle scienze di Siena detta dei Fisiocritici, Memorie n. 2, 1985, 939 pp. Lire 60.000. Nuncius, 2(1): 282, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01259.

Mazzolini:1987:BREa

[201] Renato G. Mazzolini. Book review: Emil du Bois-Reymond (1818–1896) Anton Dohrn (1840–1909) Briefwechsel, herausgegeben von Christiane Groeben in Zusammenarbeit mit Klaus Hierholzer mit einer historischen Einführung von Ernst Florey, Berlin, Springer Verlag 1985, xliv + 322 pp. Nuncius, 2(1):282–283, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539177x01268.

Abbri:1987:BRLa

[202] Ferdinando Abbri. Book review: L’età del positivismo, (Italian) [The Age of Positivism], a cura di Paolo Rossi, Bologna, Il Mulino 1986, 491 pp. Nuncius, 2(1):283–284, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539177x01277.

Pogliano:1987:BRTb

[203] Claudio Pogliano. Book review: T. Tornitore, Storia delle sinestesie. Le origini dell’audizione colorata, Genova, Universit di Genova 1986, 63 pp., s.i.p. Nuncius, 2(1):284, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539177x01286.

LaVergata:1987:BRR

[204] Antonello La Vergata. Book review: Robert P. Mcintosh, the background of ecology. Concept and theory, Cambridge, Cambridge University Press 1985, XIII, 383 pp., £30,00, $39,50. Nuncius, 2(1):285–297, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539177x01295.

Anonymous:1987:A

[205] Anonymous. Announcement. Nuncius, 2(1):299–xii, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539177x01303.

Field:1987:LPP

[206] J. V. Field. Linear perspective and the projective geometry of Girard Desargues. Nuncius, 2(2):3–40, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539187x00015.

Drake:1987:GC

[207] Stillman Drake. Galileos constant. Nuncius, 2(2):41–54, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00024.

Abbri:1987:MSS

[208] Ferdinando Abbri. Il misterioso spiritus salis. Nuncius, 2(2):55–88, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00033.

Corsi:1987:HDS

[209] Pietro Corsi. The heritage of Dugald Stewart: Oxford philosophy and the method of political economy. Nuncius, 2(2):89–144, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00042.

Grillot:1987:LOI

[210] Solange Grillot. L’emploi des objectifs italiens l’observatoire de Paris la fin du 17me siècle. Nuncius, 2(2):145–155, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539187x00051.

Monti:1987:IPM

[211] Maria Teresa Monti. Inediti di Pietro Moscati. Un episodio del dibattito italiano sull’insensibilita della Dura Madre. Nuncius, 2(2):157–170, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00060.

Proverbio:1987:HCC

[212] Edoardo Proverbio. Historic and critical comment on the risposta of R. J. Boscovich to a paragraph in a letter by Prince Kaunitz. Nuncius, 2 (2):171–226, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539187x00079.

Anonymous:1987:INP

[213] Anonymous. Index of names of persons. Nuncius, 2(2):227, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00088.

Bucciantini:1987:GSA

[214] Massimo Bucciantini. George Sarton e Aldo Mieli: Bibliografia e concezioni della scienza a confronto. (Italian) [George Sarton and Aldo Mieli: Bibliography and conceptions of science in comparison]. Nuncius, 2(2): 229–239, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539187x00097.

Bianchi:1987:CIG

[215] Lorenzo Bianchi. Convegno internazionale su Giovan Battista Della Porta. Nuncius, 2(2):241–251, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539187x00105.

Panella:1987:BRT

[216] Emilio Panella. Book review: Thabit Ibn Qurra, Oeuvres d’astronomie. Texte tablié et traduit par Régis Morelon, Paris, Les Belles Lettres 1987, cxlii-321-xv pp. Nuncius, 2(2):253–255, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539187x00114.

Klatt:1987:BRT

[217] Norbert Klatt. Book review: Theophrast Von Hohenhein, genannt Paracelsus, Theologische und Religionsphilosophische Schriften. Bd. III: Dogmatische und polemische Einzelschriften. Bearbeitet von Kurt Goldammer (Smtliche Werke, 2. Abteilung, Bd. III), Wiesbaden, Franz Steiner Verlag 1986, liv + 299 pp. Nuncius, 2(2):255–256, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00123.

Dougherty:1987:BRE

[218] F. W. P. Dougherty. Book review: Eric J. Aiton, Leibniz. A biography, Bristol Boston, Adam Hilger 1985, xiv + 307 pp., bibl., index, illus. Nuncius, 2(2):256–257, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539187x00132.

Dougherty:1987:BRD

[219] F. W. P. Dougherty. Book review: Derek Gjertsen, the Newton Handbook, London and New York, Routledge Kegan Paul 1986, xiv-665 pp., bibl., indices. Nuncius, 2(2):258–259, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00141.

Dougherty:1987:BRR

[220] F. W. P. Dougherty. Book review: Rudolf Vierhaus (ed.), Wissenschaften im Zeitalter der Aufklärung. Aus Anlaß des 250 jährigen Bestehens des Verlages Vandenhoeck Ruprecht. Göttingen, Vandenhoeck Ruprecht 1985, 282 pp., DM 34. Nuncius, 2(2):259–262, ????

1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00150.

Bernardi:1987:BRC

[221] Walter Bernardi. Book review: Catalogo del Fondo Haller della Biblioteca Nazionale Braidense di Milano, a cura di Maria Teresa Monti, Milano, Franco Angeli Editore 1983–1986, Parte I voll. I–III, Parte II voll. I–III. Nuncius, 2(2):262–266, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539187x00169.

Barsanti:1987:BRW

[222] Giulio Barsanti. Book review: Wilfrid Blunt, Linn. Le prince des botanistes, (French) [The prince of botanists]. Paris, Belin 1986, 350 pp. Nuncius, 2(2):266–268, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539187x00178.

Borrelli:1987:BRD

[223] Antonio Borrelli. Book review: Domenico Cotugno, Opere. Antologia a cura di Antonio Iurilli. Commento ai testi di Rita Maria Rossini. Presentazione di Giuseppe De Benedictis, Manduria, Piero Lacaita editore 1986 (Prestige. Scritti di Varia Umanit, 2). Nuncius, 2(2):268–270, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00187.

Abbri:1987:BRW

[224] Ferdinando Abbri. Book review: William F. Bynum, Roy Porter (editors), Medical Fringe Medical Orthodoxy 1750–1850, London, Croom Helm 1987, 274 pp. (the Wellcome Institute Series in the History of Medicine). Nuncius, 2(2):271–272, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539187x00196.

Guarnieri:1987:BRW

[225] Patrizia Guarnieri. Book review: William F. Bynum, Roy Porter, Michael Shepherd (eds.), The Anatomy of Madness. Essays in the History of Psychiatry, London and New York, Tavistock Publications 1985, 2 vols., I. People and Ideas, 316 pp. II. Institutions and Society, 283 pp. Nuncius, 2(2):273–276, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539187x00204.

Dougherty:1987:BRG

[226] F. W. P. Dougherty. Book review: Goethe und die Romantik. Briefe mit Erluterungen. Herausgegeben von Carl Schüddekopf und Oskar Walzel. (Schriften der Goethe-Gesellschaft, 13) Photo-mechanical Reprint of the original edition Weimar, 1898 in two volumes. Leipzig, Zentralantiquariat der DDR 1984. Nuncius, 2(2):276–277, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00213.

Abbri:1987:BRLb

[227] Ferdinando Abbri. Book review: Ludmilla Jordanova (editor), Languages of Nature. Critical Essays on Science and Literature. Foreword by Raymond Williams, New Brunswick, Rutgers University Press 1986, 351 pp. Nuncius, 2(2):277–280, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539187x00222.

Barsanti:1987:BRD

[228] Giulio Barsanti. Book review: David Kohn (ed.), The Darwinian Heritage. Princeton, Princeton University Press 1985, 1139 pp. Nuncius, 2 (2):280–282, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539187x00231.

Pogliano:1987:BRG

[229] Claudio Pogliano. Book review: George W. Stocking, Jr. (ed.), Malinowski, Rivers, Benedict and Others. Essays on Culture and Personality, Madison, University of Wisconsin Press 1986, 259 pp. Nuncius, 2 (2):282–285, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539187x00240.

Santambrogio:1987:BRJ

[230] Marco Santambrogio. Book review: Jean Van Heijenoort, Selected Essays. Napoli, Bibliopolis 1985, 166 pp. Nuncius, 2(2):286–289, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00259.

Miniati:1987:SB

[231] Mara Miniati. Strumenti bibliografici. Nuncius, 2(2):291, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00268.

Miniati:1987:S

[232] Mara Miniati. Strumentaria. Nuncius, 2(2):292, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539187x00277.

Miniati:1987:BRV

[233] Mara Miniati. Book review: Vladimiro Valerio, Historiographic and numerical notes on the Atlante Farnese and its celestial sphere. estr. da Der Globusfreund nr. 35–37, Wien 1987, pp. 97–124. Nuncius, 2(2):292, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00286.

Brenni:1987:BRM

[234] Paolo Brenni. Book review: L. Von Mackensen, Feinmechanik aus Kassel 225 Jahre F. W. Breithaupt Sohn (Festschrift und Austellungsbegleiter), Kassel, Georg Wenderoth Verlag 1987, 85 pp. Nuncius, 2(2): 293, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00295.

Brenni:1987:BRP

[235] Paolo Brenni. Book review: P. Ronne, A. B. W. Nielsen, Development of the Ion X-ray Tube. Copenhagen, the Foundation of Poul Lundbeck, Eng. M. Sc., C. A. Reitzel Publishers 1986 (Acta Historica Scientiarum Naturalium et Medicalium, Vol. 35). Nuncius, 2(2):293–294, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00303.

Caroti:1987:CS

[236] Stefano Caroti. Cartografia storica. Nuncius, 2(2):294–295, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00312.

Miniati:1987:BRM

[237] Mara Miniati. Book review: M. A. Martullo Arpago, L. Castaldo Manfredonia, I. Principe, V. Valerio (a cura di), Fonti cartografiche nell’Archivio di Stato di Napoli. Napoli 1987, 164 pp. Nuncius, 2(2):

295, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00321.

Miniati:1987:BRI

[238] Mara Miniati. Book review: Istituto Italiano Per Gli Studi Filosofici, Catalogazione, Studio e conservazione della cartografia storica, a cura di Vladimiro Valerio. Napoli, nella sede dell’Istituto 1987, 117 pp. Nuncius, 2(2):296, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539187x00330.

Caroti:1987:A

[239] Stefano Caroti. Astronomia. Nuncius, 2(2):296–297, ???? 1987. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539187x00349.

Miniati:1987:BRA

[240] Mara Miniati. Book review: Aa.Vv., Astronomia e astrofisica in URSS. A cura di Livio Gratton. Bari, Edizioni Dedalo 1987, 166 pp. Nuncius, 2(2):297–298, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539187x00358.

Caroti:1987:MB

[241] Stefano Caroti. Medicina e biologia. Nuncius, 2(2):298–299, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00367.

Abbri:1987:BRRb

[242] Ferdinando Abbri. Book review: Richard P. Steiner (ed.), Folk Medicine. the Art and the Science, Washington Dc, American Chemical Society 1986, viii + 223 pp. Nuncius, 2(2):299–300, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00376.

Caroti:1987:BRAb

[243] Stefano Caroti. Book review: Actes du 110e Congrés National des Sociétés Savantes. (Montpellier, 1985). Section d’histoire des sciences et des techniques. Tome II. Histoire de l’École Médicale de Montpellier. Colloque, Paris, C.T.H.S. 1985, 365 pp, ill. Nuncius, 2(2):300–302, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00385.

Dougherty:1987:BRW

[244] F. W. P. Dougherty. Book review: William Lefanu. A Bibliography of Edward Jenner. Second edition, Winchester, St Paul’s Bibliographies 1985, xvi-160 pp., £24. Nuncius, 2(2):303, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00394.

Pogliano:1987:P

[245] Claudio Pogliano. Psicologia. Nuncius, 2(2):304, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539187x00402.

Abbri:1987:SDS

[246] Ferdinando Abbri. Sociologia della scienza. Nuncius, 2(2):304–306, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00411.

Caroti:1987:MR

[247] Stefano Caroti. Medioevo e rinascimento. Nuncius, 2(2):306, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00420.

Caroti:1987:BRRb

[248] Stefano Caroti. Book review: Rainer Jehl, Melancholie und Acedia. Ein Beitrag zu Anthropologie und Ethik Bonaventuras, Paderborn — München — Wien — Zürich, Ferdinand Schoening 1985 (Münchener Universitäts-Schriften, Veroeffentlichungen des Grabmann-Institutes, N.F. 32), 323 pp., bibl., ind. Nuncius, 2(2):306–307, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00439.

Caroti:1987:BRS

[249] Stefano Caroti. Book review: Selected Works of Ramn Llull (1232–1316), Edited and Translated by Antony Bonner, N.J., Princeton University Press 1985, 2 voll., 1329 pp., bibl., ind., ill. Nuncius, 2(2):307–308, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00448.

Caroti:1987:BREa

[250] Stefano Caroti. Book review: Eugenio Randi, Il sovrano e l’orologiaio. Due immagini di Dio nel dibattito sulla potentia absoluta fra XIII e XIV secolo, Firenze, La Nuova Italia Editrice 1987, 191 pp., bibl. ind. (Pubblicazioni della Facolt di Lettere e Filosofia dell’Università di Milano, Cxxi. Sezione a cura del Dipartimento di Filosofia, 11). Nuncius, 2(2):308–309, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00457.

Caroti:1987:BREb

[251] Stefano Caroti. Book review: Erich Kleineidam, Universitas Studii Erffordiensis. Überblick über die Geschichte der Universität Erfurt. Teil I: Spaetmittelalter 1392–1460, Leipzig, St. Benno Verlag Gmbh 1985, 480 pp., bibl., ind. (Erfurter Teologische Studien, Bd. 14). Nuncius, 2(2): 309–310, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539187x00466.

Caroti:1987:BRAc

[252] Stefano Caroti. Book review: Armando F. Verde O.P., Lo studio Fiorentino 1473–1503. Ricerche e Documenti. Volume Quarto. La vita universitaria, Firenze, Leo S. Olschki Editore 1985, 3 tomi, 1546 pp., bibl. Nuncius, 2(2):310–312, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539187x00475.

Abbri:1987:BRK

[253] Ferdinando Abbri. Book review: Kurt Goldammer, Paracelsus in neuen Horizonten. Gesammelte Aufsätze, (German) [Paracelsus new horizons]. Wien, Verband der wissenschaftlichen Gesellschaften Österreichs 1986 (Salzburger Beiträge zur Paracelsusforschung; Folge 24), 400 pp. Nuncius, 2(2):312–313, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539187x00484.

Caroti:1987:SXX

[254] Stefano Caroti. Secoli XVII–XX. Nuncius, 2(2):313–314, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00493.

Parigi:1987:BRG

[255] Silvia Parigi. Book review: George Berkeley, Dialoghi tra Hylas e Philonous, a cura di P. F. Mugnai, Roma/Bari, Laterza 1987, xxx-148 pp. Nuncius, 2(2):314, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539187x00501.

Mazzolini:1987:BREb

[256] Renato G. Mazzolini. Book review: Engelbert Kaempfer, Phoenix Persicus. Die Geschichte der Dattelpalme. Einleitung, bersetzung aus dem Lateinischen und Bearbeitung von Wolfgang Muntschick, Marburg, Basilisken-Presse 1987, 227 pp. Nuncius, 2(2):315, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00510.

Casati:1987:BRW

[257] Stefano Casati. Book review: Walter Tega, Arbor scientiarum. Enciclopedia e sistemi in Francia da Diderota Comte. Bologna, il Mulino 1984, 406 pp. Nuncius, 2(2):316, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539187x00529.

Abbri:1987:BRPb

[258] Ferdinando Abbri. Book review: Piero di Giovanni (a cura di), La tradizione illuministica in Italia. (Italian) [The tradition of Enlightenment in Italy]. Palermo, Palumbo 1986, 294 pp. Nuncius, 2(2):317, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00538.

Abbri:1987:BRJb

[259] Ferdinando Abbri. Book review: Justus Von Liebig, Briefe an Vieweg, herausgegeben und bearbeitet von Margarete und Wolfgang Schneider, Braunschweig/Wiesbaden, Friedr. Vieweg Sohn 1986, xviii + 437 pp. Nuncius, 2(2):317–318, ???? 1987. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539187x00547.

Brenni:1987:BRS

[260] Paolo Brenni. Book review: Un siécle d’électricité dans le monde. Actes du premier colloque international d’histoire de l’électricité, organisé par l’A.H.E.F. A cura di F. Cardot. Paris, Presses Universitaires de France 1987, 444 pp. Nuncius, 2(2):318–319, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539187x00556.

Miniati:1987:ADM

[261] Mara Miniati. Attivita dell’Istituto e Museo di Storia della Scienza di Firenze. Nuncius, 2(2):321–334, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539187x00565.

Scriba:1987:A

[262] C. J. Scriba. Announcement. Nuncius, 2(2):337–xii, ???? 1987. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539187x00574.

Frangenberg:1988:EDO

[263] Thomas Frangenberg. Egnatio Danti’s optics. Cinquecento Aristotelianism and the Medieval tradition. Nuncius, 3(1):3–38, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00014.

Pogliano:1988:ELD

[264] Claudio Pogliano. Un esperto laico d’anime: Marin Cureau de la chambre (1596–1669). Nuncius, 3(1):39–54, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539188x00023.

Barsanti:1988:IDN

[265] Giulio Barsanti. Le immagini della natura: Scale, mappe, alberi 1700– 1800. Nuncius, 3(1):55–125, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539188x00032.

DeClercq:1988:GCM

[266] Peter De Clercq. The ’s Gravesande Collection in the Museum Boerhaave, Leiden. Nuncius, 3(1):127–137, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539188x00041.

LaVergata:1988:TNE

[267] Antonello La Vergata. Theodicy and Nature’s economy. Nuncius, 3 (1):139–152, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00050.

Miniati:1988:DAG

[268] Mara Miniati. Due antichi globi. Nuncius, 3(1):153–155, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00069.

Beretta:1988:PAD

[269] Marco Beretta. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 3(1):157–161, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539188x00078.

Mandrino:1988:IF

[270] A. Mandrino, G. Tagliaferri, and P. Tucci. Istituzioni e fonti. Nuncius, 3(1):163–176, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00087.

Mazzolini:1988:CBIa

[271] Renato G. Mazzolini. Commission on bibliography of the International Union of the History and Philosophy of Science. Nuncius, 3(1):177–188, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00096.

Bedini:1988:O

[272] Silvio A. Bedini. Obituary. Nuncius, 3(1):189–191, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539188x00104.

Renn:1988:DC

[273] Jürgen Renn. Discussioni critiche. Nuncius, 3(1):193–241, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00113.

Carrara:1988:BRP

[274] Daniela Mugnai Carrara. Book review: Per-Gunnar Ottosson, Scholastic Medicine and Philosophy. A Study of Commentaries on Galen’s Tegni (ca. 1300–1400), Napoli, Bibliopolis, Edizioni di Filosofia e Scienze 1984, 322 pp. Nuncius, 3(1):243–244, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539188x00122.

Ernst:1988:BRA

[275] Germana Ernst. Book review: Astrology, Science and Society. Historical Essays, ed. by Patrick Curry, the Boydell Press 1987, ix + 302 pp. Nuncius, 3(1):244–250, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539188x00131.

Pereira:1988:BRA

[276] Michela Pereira. Book review: Die Alchemie in der europäischen Kultur- und Wissenschaftsgeschichte [Vorträge gehalten anlässlich des 16 Wolfenbütteler Symposions vom 2 bis 5 April 1984 in der Herzog August Bibliothek] hrsg. von Christoph Meinel. Wiesbaden, Harrassowitz 1986 (Wolfenbütteler Forschungen Bd 32), 356 pp., s.i.p. Nuncius, 3 (1):250–254, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00140.

Miniati:1988:BRS

[277] Mara Miniati. Book review: Sandra Giovannini — Gabriella Mancini, La Farmacia di Santa Maria Novella. Firenze, Becocci Editore 1987, 32 pp. (Biblioteca de ≪ Lo Studiolo ≫ ). Nuncius, 3(1):255–256, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00159.

Biagioli:1988:BRZ

[278] Mario Biagioli. Book review: Zygmunt Waźbiński, l’Accademia Medicea del Disegno a Firenze nel Cinquecento, 2 Vols., Firenze, Olschki 1987, 557 pp. Nuncius, 3(1):256–259, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539188x00168.

Carrara:1988:BRW

[279] Daniela Mugnai Carrara. Book review: A. Waer, R. K. French, I. M. Lonie (Eds), the Medical Renaissance of the Sixteenth Century, Cambridge, University Press 1985, xvi + 349 pp. Nuncius, 3(1):259–262, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00177.

Simonutti:1988:BRH

[280] Luisa Simonutti. Book review: H. Bots, H. Hillenaar, J. Janssen, J. Van Der Korst & L. Van Lieshout, De ≪ Bibliothèque Universelle et Historique ≫ (1686–1693). Een periodiek als trefpunt van geletterd Europa. Amsterdam, Apa-Holland Universiteits Pers 1981, xvi, 399 pp. Nuncius, 3(1):262–264, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539188x00186.

Turner:1988:BRS

[281] Gerard L’E. Turner. Book review: Silvio A. Bedini, Clockwork Cosmos: Bernard Facini and the Farnese Planisferologio, Città del Vaticano, Biblioteca Apostolica Vaticana 1985, 221 pp., L. 40.000. Nuncius, 3(1): 264–266, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00195.

Barsanti:1988:BRH

[282] Giulio Barsanti. Book review: Historie du concept d’espèce dans les sciences de la vie, (French) [History of the concept of space in the life sciences], Paris, Éditions de la Fondation Singer-Polignac 1987, 324 pp. s.i.p. Nuncius, 3(1):266–269, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539188x00203.

Abbri:1988:BRAa

[283] Ferdinando Abbri. Book review: Andrew E. Benjamin, Geoffrey N. Cantor, John R. R. Christie (eds), the Figural and the Literal. Problems of language in the history of science and philosophy 1630–1800, Manchester, Manchester University Press 1987, viii + 229 pp. Nuncius, 3(1):269–273, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00212.

Crispini:1988:BRP

[284] Ines Crispini. Book review: Paolo Manganaro, l’Antropologia di Kant, Napoli, Guida 1983, 290 pp. Nuncius, 3(1):274–276, ????

1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00221.

LaVergata:1988:BRL

[285] Antonello La Vergata. Book review: Linda L. Clark, Social Darwinism in France, The University of Alabama Press 1984, xi + 261 pp. Nuncius, 3(1):277–280, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00230.

Guarnieri:1988:BRJ

[286] Patrizia Guarnieri. Book review: John M. O’Donnell, the Origin of Behaviorism. American Psychology, 1870–1920, New York and London, New York University Press 1985, xii + 299 pp., $40. Nuncius, 3(1): 280–281, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00249.

Biagioli:1988:BRA

[287] Mario Biagioli. Book review: Alexandre Koyré, De la mystique à la science, P. Redondi (ed.). Paris, Éditions de l’École des Hautes Études en Sciences Sociales 1986, 227 pp. Nuncius, 3(1):282–284, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00258.

Guarnieri:1988:BRE

[288] Patrizia Guarnieri. Book review: Eugen Bleuler, Dementia praecox o il gruppo delle schizofrenie, presentazione dell’edizione italiana di Luigi Cancrini, Roma, La Nuova Italia Scientifica 1985, 365 pp., L. 42.500. Nuncius, 3(1):284–285, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539188x00267.

Segala:1988:BRR

[289] Marco Segala. Book review: Robert Muir Wood, the Dark Side of the Earth, London, Allen & Unwin 1985, II ed., 1986, x + 246 pp. Nuncius, 3(1):286–291, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00276.

Caroti:1988:SBa

[290] Stefano Caroti. Strumenti bibliografici. Nuncius, 3(1):293–294, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00285.

Miniati:1988:BRMa

[291] Mara Miniati. Book review: A. Mandrino, G. Tagliaferri, P. Tucci (a cura di), Inventario di archivio dell’Osservatorio Astronomico di Brera 1726–1917. Milano, Università degli Studi di Milano, Istituto di Fisica Generale Applicata, Sezione di Storia della Fisica 1987, 270 pp. Nuncius, 3(1):294, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00294.

Miniati:1988:S

[292] Mara Miniati. Strumentaria. Nuncius, 3(1):295, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539188x00302.

Miniati:1988:BRRa

[293] Mara Miniati. Book review: Robert P. Multhauf — Gregory Good, A Brief History, of Geomagnetism and a Catalog of the Collections of the National Museum of American History. Washington, D.C., Smithsonian Institution Press 1987, 87 pp. (Smithsonian Studies in History and Technology. Number 48). Nuncius, 3(1):295–296, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00311.

Miniati:1988:SDS

[294] Mara Miniati. Storia della scienza. Nuncius, 3(1):296–297, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00320.

Miniati:1988:BRRb

[295] Mara Miniati. Book review: Robert Halleux et Pascal Lefèbvre, Étudier l’histoire des sciences et des techniques. Bruxelles, Apps éditions 1987, 153 pp. [ISBN 2-87192-006-0]. Nuncius, 3(1):297, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00339.

Abbri:1988:BRC

[296] Ferdinando Abbri. Book review: Christian Hünemörder (a cura di), Wissenschaftsgeschichte heute. Ansprachenund wissenschaftliche Vorträge zum 25 jäbrigen Bestehen des Instituts für Geschichte der Naturwissenschaften, Mathematik und Technik der Universität Hamburg. Stuttgart, Franz Steiner Verlag 1987, 123 pp. (≪ Beiträge zur Geschichte der Wissenschaft und der Technik, Heft 20 ≫ ). Nuncius, 3(1):297–298, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00348.

Miniati:1988:CSa

[297] Mara Miniati. Cartografia storica. Nuncius, 3(1):298–299, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00357.

Miniati:1988:M

[298] Mara Miniati. Matematica. Nuncius, 3(1):299, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539188x00366.

Abbri:1988:BRV

[299] Ferdinando Abbri. Book review: Vincenzo Cappelletti, Federico Di Trocchio (a cura di), De Sedibus, et Causis. Morgagni nel Centenario. Roma, Istituto della Enciclopedia Italiana 1986, 393 pp. (≪ Acta Encyclopaedica 6 ≫ ). Nuncius, 3(1):300–301, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539188x00375.

Guarnieri:1988:BRM

[300] Patrizia Guarnieri. Book review: Marc Augè e Claudine Herzlich, Il sense del male. Antropologia, storia e sociologia della malattia, ed. it. a cura di Francesco Maiello. Milano, Il Saggiatore 1986, 272 pp., L. 38.000. Nuncius, 3(1):301–302, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539188x00384.

Abbri:1988:BRAb

[301] Ferdinando Abbri. Book review: Allen G. Debus, Chemistry, Alchemy and the New Philosophy, 1550–1700. Studies in the History of Science and Medicine. London, Variorum Reprints 1987, xii + 320 pp. Nuncius, 3(1):302–304, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00393.

Abbri:1988:BRAc

[302] Ferdinando Abbri. Book review: Academiae Analecta. Brussel, Paleis der Academiën 1986, 128 pp. (≪ Mededelingen van de Koninklijke Academie voor Wetenschappen, Letteren en Schone Kunsten van België, Klasse der Wetenschappen, Jaargang 48, Nr. 4, 1986 ≫ ). Nuncius, 3(1):304–305, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00401.

Abbri:1988:BRI

[303] Ferdinando Abbri. Book review: Irene Strube, Georg Ernst Stahl. Leipzig, BSB B. G. Teubner Verlagsgesellschaft 1984, 82 pp., ill. (≪ Biographien hervorragender Naturwissenschaftler (German) [Biographies of outstanding scientists, engineers and physicians], Techniker und Mediziner Band 76 ≫ ). Nuncius, 3(1):305–306, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00410.

Parigi:1988:BRF

[304] Silvia Parigi. Book review: Francisco Varela, Scienza e tecnologia della cognizione. Direzioni emergenti (traduzione di Stefano Isola). Firenze, Hopeful Monster 1987, pp. 90. Nuncius, 3(1):306–307, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00429.

Caroti:1988:BRP

[305] Stefano Caroti. Book review: Pierre Alexandre, Le climat en Europe au Moyen Âge. Contribution à l’histoire des variations climatiques de 1000 à 1425, d’après les sources narratives de l’Europe occidentale. Paris Éditions de l’École des Hautes Études en Sciences Sociales 1987, pp. 827, bibl., grafici, tavole e carte. Nuncius, 3(1):307–308, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00438.

Caroti:1988:BRO

[306] Stefano Caroti. Book review: Olga Weijers, Terminologie des universités au XIIIe siècle. Roma, Edizioni dell’Ateneo 1987 (Lessico Intellettuale Europeo, XXXIX), pp. 437. Nuncius, 3(1):308–309, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00447.

Caroti:1988:BRD

[307] Stefano Caroti. Book review: Dietrich Von Freiberg, Opera omnia. Tomus Iv. Schriften zur Naturwissenschaft. Briefe, mit einer Einleitung von Loris Sturlese. Herausgegeben von Maria Rita Pagnoni-Sturlese, Rudolf Rehn, Loris Sturlese und William A. Wallace. Hamburg, Felix Meiner Verlag 1985 (Corpus Philosophorum Teutonicorum Medii Aevi, Band II, 4), pp. 380, ind. Nuncius, 3(1):309–311, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00456.

Plastina:1988:BRL

[308] Sandra Plastina. Book review: L’Aristotelisme au XVIe siècle, ≪ Études philosophiques ≫, 1986, n. 3, pp. 281–405. Nuncius, 3(1):312–313, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00465.

Caroti:1988:BRR

[309] Stefano Caroti. Book review: Rita Sturlese, Bibliografia, censimento e storia delle antiche stampe di Giordano Bruno. Firenze, Leo S. Olschki 1987 (Istituto Nazionale di Studi sul Rinascimento, Quaderni di ≪ Rinascimento ≫, VI), pp. 224. Nuncius, 3(1):314–315, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00474.

Caroti:1988:BRM

[310] Stefano Caroti. Book review: Massimo Luigi Bianchi, Signatura rerum. Segni, magia e conoscenza da Paracelso a Leibniz. Roma, Edizioni dell’Ateneo 1987 (Lessico Intellettuale Europeo, Xliii), pp. 199, ind. Nuncius, 3(1):315–316, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539188x00483.

Caroti:1988:BRC

[311] Stefano Caroti. Book review: Convegno celebrativo del IV centenario della nascita di Federico Cesi (Acquasparta, 7–9 ottobre 1985). (Italian)

[Conference celebrating the fourth centenary of the birth of Federico Cesi (Acquasparta 7–9 October 1985)]. Roma, Accademia Nazionale dei Lincei 1986 (Atti dei Convegni lincei, 78), pp. 352, ill. Nuncius, 3(1): 316–319, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539188x00492.

Caroti:1988:BRB

[312] Stefano Caroti. Book review: Bruno Basile, l’invenzione del vero. La letteratura scientifica da Galilei ad Algarotti. Roma, Salerno Editrice 1987, 245 pp., ind. (Studi e Saggi, 5). Nuncius, 3(1):319–321, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00500.

Abbri:1988:BRT

[313] Ferdinando Abbri. Book review: Tiziano Arrigoni, Uno scienziato nella Toscana del Settecento. Giovanni Targioni Tozzetti. Firenze, Edizioni Gonnelli 1987, viii + 174 pp. (≪Documenti inediti di cultura toscana volume IX ≫ ). Nuncius, 3(1):321–322, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00519.

Falzoni:1988:BRV

[314] Mauro Falzoni. Book review: Villes et territoires pendant le période napoléonienne (France et Italie). Actes du colloque organisé par l’École Française de Rome et l’Assessorato alla culturade la ville de Rome, avec la partécipation de la Maison des sciences de l’homme (Paris) (Rome 3, 4 et 5 mai 1987). Rome, École française de Rome 1987, 556 pp. (Collection de l’École française de Rome, 96). Nuncius, 3(1):322–324, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00528.

Anonymous:1988:BRA

[315] Anonymous. Book review: Alessandro Pascolini, Dall’infinitamente grande all’infinitamente piccolo. Viaggio nel cuore della materia. Venezia, Arsenale Editrice 1987, 108 pp. Nuncius, 3(1):325–336, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539188x00537.

Scriba:1988:Aa

[316] C. J. Scriba. Announcement. Nuncius, 3(1):338–354, ???? 1988. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539188x00546.

Field:1988:WSA

[317] J. V. Field. What is scientific about a scientific instrument? Nuncius, 3(2):3–26, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x00303.

Nardi:1988:QDV

[318] Antonio Nardi. La quadratura della velocità. Nuncius, 3(2):27–64, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00312.

Parigi:1988:QBU

[319] Silvia Parigi. La quaestio di Barrow: Un’aporia della visione tramite lenti. (Italian) [The question of Barrow: a test of vision through lenses]. Nuncius, 3(2):65–69, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00321.

Antoniotti:1988:GT

[320] Paola Antoniotti. Gioacchino taddei, 1792–1860. Nuncius, 3(2):71–100, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00330.

Bruzzaniti:1988:STF

[321] Giuseppe Bruzzaniti. La struttura del tempo fisico. Nuncius, 3(2):101– 147, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00349.

Archinard:1988:CSR

[322] Margarida Archinard. Les cadrans solaires rectilignes. Nuncius, 3(2): 149–181, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x00358.

Contardi:1988:PAD

[323] Simone Contardi. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 3(2):183–187, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539178x00367.

Anonymous:1988:I

[324] Anonymous. Indice. Nuncius, 3(2):188–191, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539178x00376.

LopezPinero:1988:IF

[325] José M. López Piñero. Istituzioni e fonti. Nuncius, 3(2):193–207, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00385.

Mazzolini:1988:CBIb

[326] Renato G. Mazzolini. Commission on bibliography of the International Union of the History and Philosophy of Science. Nuncius, 3(2):209–225, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00394.

Alborn:1988:DC

[327] Timothy L. Alborn. Discussioni critiche. Nuncius, 3(2):227–250, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00402.

Corsi:1988:ISS

[328] Pietro Corsi. International Summer School in History of Science (Berkeley–Bologna–Upsala). Nuncius, 3(2):251–252, ???? 1988. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539178x00411.

Shea:1988:BRB

[329] William R. Shea. Book review: B. V. Subbarayappa and K. V. Sarma (eds.), Indian Astronomy. A Source Book. Bombay, Nehru Centre 1985, xliv + 338 pp. Nuncius, 3(2):253–254, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00420.

Maracchia:1988:BRA

[330] Silvio Maracchia. Book review: Aa.Vv., l’alba dei numeri, a cura di Tina Bovi e Wilma di Palma. Bari, Edizioni Dedalo 1987, 152 pp., Lit. 24.000. Nuncius, 3(2):254–259, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00439.

Lippi:1988:BRM

[331] Donatella Lippi. Book review: Mirko D. Grmek, Le malattie all’alba della civiltà occidentale, Bologna, Il Mulino 1985, 599 pp. (Ed. originale: Les maladies à l’aube de la civilisation occidentale, Paris, Payot 1983). Nuncius, 3(2):259–261, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00448.

Lippi:1988:BRGa

[332] Donatella Lippi. Book review: Giuseppe Penso, La medicina romana. l’arte di Esculapio nell’antica Roma, Saronno, Ciba-Geigy Ed. 1985, 607 pp. (Ed. originale: La médecine romaine. l’art d’Esculape dans la Rome antique, Paris, R. Dacosta 1984). Nuncius, 3(2):261–262, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00457.

ForcadaNogues:1988:BRA

[333] Miquel Forcada Nogués. Book review: De astronomia Alfonsi Regis. Barcelona, Universidad de Barcelona, Instituto ≪ Millàs Vallicrosa ≫ de Historia de la Ciencia Arabe 1987. Nuncius, 3(2):262–266, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00466.

Compagni:1988:BRP

[334] Vittoria Perrone Compagni. Book review: Picatrix. The Latin version of the Gāyat Al-akīm edited by David Pingree (Text, Introduction, Appendices, Indices), the Warburg Institute, University of London 1986 (≪ Studies of the Warburg Institute ≫, vol. 39), 326 pp. Nuncius, 3 (2):266–268, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x00475.

Abbri:1988:BRAd

[335] Ferdinando Abbri. Book review: Amos Funkenstein, Theology and the Scientific Imagination from the Middle Ages to the Seventeenth Century, Princeton, Princeton University Press 1986, xii + 421 pp. Nuncius, 3 (2):268–272, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x00484.

Barsanti:1988:BRG

[336] Giulio Barsanti. Book review: Giovanni Artico di Porcia, Notizie della vita, e degli studi del Kavalier Antonio Vallisneri, Bologna, Patron 1986, 244 pp., L. 20.000. Nuncius, 3(2):272–275, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00493.

Monti:1988:BRD

[337] Maria Teresa Monti. Book review: Desedibus, et causis. Morgagni nel centenario. (Latin and Italian) [Depressed, and the causes. The centenary of Morgagni]. A cura di V. Cappelletti e F. Di Trocchio, Istituto dell’Enciclopedia Italiana 1986. (≪ Acta Enciclopaedica ≫, 6), 393 pp. Nuncius, 3(2):275–279, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00501.

Monti:1988:BRG

[338] Maria Teresa Monti. Book review: Giorgio Tonelli, La pensée philosophique de Maupertuis. Son milieu et ses sources. Edition posthume par Claudio Cesa, Hildesheim, Georg Olms Verlag 1987, 144 pp. (≪ Studien und Materialien zur Geschichte der Philosophie ≫, 25). Nuncius, 3(2):279–282, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00510.

Arrigoni:1988:BRI

[339] Tiziano Arrigoni. Book review: Ideologiae scienza nell’opera di Paolo Frisi (1728–1784), a cura di Gennaro Barbarisi, Milano, Franco Angeli 1987, vol. I, 473 pp.; vol. II, 706 pp. Nuncius, 3(2):282–284, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00529.

Barsanti:1988:BRD

[340] Giulio Barsanti. Book review: Danilo Barsanti, Leonardo Rombai, Leonardo Ximenes, uno scienziato nella Toscana lorenese del Settecento, Firenze, edizioni medicea 1987, 241 pp., L. 20.000. Nuncius, 3(2):285– 287, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00538.

Giuntini:1988:BRA

[341] Chiara Giuntini. Book review: Anand C. Chitnis, The Scottish Enlightenment & Early Victorian English Society, London–Sydney–Dover (New Hampshire), Croom Helm 1986, 201 pp. Nuncius, 3(2):287–289, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00547.

Abbri:1988:BRR

[342] Ferdinando Abbri. Book review: Roy Porter, Mind-forg’d Manacles. A history of madness in England from the Restoration to the Regency, London, the Athlone Press 1987, xii + 412 pp. Nuncius, 3(2):289–293, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00556.

Giacche:1988:BRR

[343] Vladimiro Giacché. Book review: R. Heckmann, H. Krings, R. W. Meyer (a cura di), Natur und Subjektivität. Zur Auseinandersetzung mit der Naturphilosophie des jungen Schelling. Referate, Voten und Protokolle der II. Internationalen Schelling-Tagung Zürich 1983, Stuttgart– Bad Cannstatt, Frommann-Holzboog 1985 (≪ Problemata ≫ 106), 391 pp. Nuncius, 3(2):293–298, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539178x00565.

Matassi:1988:BRH

[344] Elio Matassi. Book review: Hegel und die Naturwissenschaften (Heraus. von Michael John Petry), Stuttgart–Bad Cannstatt, F. Frommann Verlag–G. Holzboog 1987, 562 pp. Nuncius, 3(2):298–301, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00574.

Pogliano:1988:BRG

[345] Claudio Pogliano. Book review: George W. Stocking Jr., Victorian Anthropology, New York, the Free Press 1987, xvii + 429 pp., $26.95. Nuncius, 3(2):301–304, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00583.

Wanrooij:1988:BRP

[346] Bruno P. F. Wanrooij. Book review: Patrizia Guarnieri, l’Ammazzabambini. Legge e scienza in un processo toscano di fine Ottocento, Torino, Einaudi (Microstorie, n. 15) 1988, 225 pp., L. 22.000. Nuncius, 3(2):304–306, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00592.

Pogliano:1988:BRS

[347] Claudio Pogliano. Book review: Saul Benison, A. Clifford Barger, Elin L. Wolfe, Walter B. Cannon. The Life and Times of a Young Scientist, Cambridge (Mass.) and London, the Belknap Press of Harvard University Press 1987, vii + 520 pp., $35.95. Nuncius, 3(2):306–309, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00600.

Segala:1988:BRM

[348] Marco Segala. Book review: Martin Schwarzbach, Alfred Wegener, the Father of Continental Drift, Science Tech Publishers, Madison, Wi, Springer-Verlag, Berlin-New York 1986, pp. xx + 241. Traduzione inglese di Carla Love di: Alfred Wegener und die Drift der Kontinente, Wissenschaftliche Verlagsgesellschaft mbH, Stuttgart 1980. Nuncius, 3 (2):310–315, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x00619.

Ciancio:1988:BRS

[349] Luca Ciancio. Book review: Stephen Jay Gould, Time’s Arrow, Time’s Cicle. Myth and Metaphor in the Discovery of Geological Time, Cambridge Mass. and London, Harvard University Press 1987, xiii + 222 pp. Nuncius, 3(2):315–318, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00628.

Caroti:1988:SBb

[350] Stefano Caroti. Strumenti bibliografici. Nuncius, 3(2):319, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00637.

Labarta:1988:S

[351] E. Calvo Labarta. Strumentaria. Nuncius, 3(2):320–321, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00646.

Miniati:1988:BRRc

[352] Mara Miniati. Book review: Roser Puig Aguilar, Los Tratados de Construcción y Uso de la Azafea de Azarquiel. Madrid, Instituto Hispano-Arabe de Cultura 1987 (≪ Cuadernos de Ciencias ≫ 1), 132 pp. Nuncius, 3(2):321–322, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00655.

Bassols:1988:MVG

[353] Sergi Bassols. Mercè Viladrich I Grau, El ≪ Kitàb al-’amal bi-l-asurlāb ≫ (Libre de l’ús de l’astrolabi) d’Ibn al-Samh. Estudi i Traducció, Barcelona, Institut d’Estudis Catalans. Memòries de la secció Històrico–Arqueològica XXXVI, 1986, 158 págs. Nuncius, 3(2):322–323, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00664.

Miniati:1988:BRL

[354] Mara Miniati. Book review: La lente. Storia, scienza, curiosità attraverso la collezione Fritz Rathschüler. Genova, Ecig-Edizioni Culturali Internazionali Genova 1988, 198 pp. Nuncius, 3(2):323, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00673.

Miniati:1988:BRMb

[355] Mara Miniati. Book review: Musica ex-machina. 3000 anni di macchine musicali automatiche dall’arpaeolia al computer musicale. (Italian) [Music ex-machina. 3000 years of machines. Automatic music to computer music of Arpaeolia] Roma, Fondazione Oriele Sotgiu di Ghilarza 1987, 75 pp. Nuncius, 3(2):324, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00682.

Miniati:1988:BRJ

[356] Mara Miniati. Book review: John Paul Kurdyla Jr., When Music was Magic. History, Phonographs and Gramophones from 1879 to 1939. Padova, John Paul Kurdyla Jr. 1987, 215 pp. Nuncius, 3(2):324, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00691.

Miniati:1988:CSb

[357] Mara Miniati. Cartografia storica. Nuncius, 3(2):325, ???? 1988. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539178x00709.

Abbri:1988:IS

[358] Ferdinando Abbri and Stefano Casati. Istituzioni scientifiche. Nuncius, 3(2):325–327, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x00718.

Lippi:1988:MB

[359] Donatella Lippi. Medicina e biologia. Nuncius, 3(2):327–328, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00727.

Lippi:1988:BRGb

[360] Donatella Lippi. Book review: Giorgio Cosmacini, Storia della Medicina e della Sanità in Italia, Bari, Laterza 1987 (III ed.), 448 pp., L. 40.000. Nuncius, 3(2):328–329, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00736.

Renato:1988:BRK

[361] G. Mazzolini Renato. Book review: Karl Enigk, Geschichte der Helminthologie im deutschsprachigen Raum. Stuttgart-New York, Gustav Fischer Verlag 1986, xviii + 356 pp., DM. 58,-. Nuncius, 3(2):329, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00745.

Renato:1988:BRR

[362] G. Mazzolini Renato. Book review: Rudolph Wagner, Samuel Thomas von Soemmerrings Leben und Verkehr mit seinen Zeitgenossen. (German) [Samuel Thomas Soemmerring’s life and interaction with contemporaries]. Herausgegeben, eingeleitet und mit einem Personenregister versehen von Franz Dumont. (Soemmering-Forschungen, Bd. II) Stuttgart-New York, Gustav Fischer Verlag 1986, 747 pp., DM. 118,-. Nuncius, 3(2):330, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00754.

Renato:1988:BRG

[363] G. Mazzolini Renato. Book review: Georg Büchner: 1813– 1837; Revolutionär, Dichter, Wissenschaftler. Katalog der Ausstellung Mathildenhöhe, Darmstadt, 2. August–27. September 1987. Basel– Frankfurt am Main, Strömfeld/Roter Stern 1987, 432 pp. Nuncius, 3 (2):330–331, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x00763.

Renato:1988:BRC

[364] G. Mazzolini Renato. Book review: Claudia Heurkamp, Der Aufstieg Der Ärzte im 19. Jahrhundert. Vom gelehrten Stand zum professionellen Experten: Das Beispiel Preußens. (Kritische Studien zur Geschichtswissenschaft, Band 68) Göttingen, Vandenhoeck & Ruprecht 1985, 409 pp. Nuncius, 3(2):331–332, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00772.

Donatella:1988:BRLa

[365] Lippi Donatella. Book review: Luigi Capasso, Salute e malattie nell’evoluzione umana, Chieti 1987, 87 pp. Nuncius, 3(2):332–333, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00781.

Donatella:1988:BRLb

[366] Lippi Donatella. Book review: Luigi Capasso, l’origine delle malattie, Chieti 1985, pp. 397. Nuncius, 3(2):333–334, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539178x00790.

Donatella:1988:JPV

[367] Lippi Donatella. ≪ journal of paleopathology ≫, 1987, vol. 1, n. 1. Nuncius, 3(2):335, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00808.

Ferdinando:1988:BRO

[368] Abbri Ferdinando. Book review: Oeuvres de Lavoisier — Correspondance. Publiée sous les auspices du Comité Lavoisier de l’Académie des Sciences. Avant-Propos de Alain Horeau, Introduction de Michelle Goupil. Fascicule Iv 1784–1786, Paris, Éditions Belin 1986, xv + 351 pp. Nuncius, 3(2):335–336, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00817.

Silvia:1988:P

[369] Parigi Silvia. Psicologia. Nuncius, 3(2):337–338, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539178x00826.

Silvia:1988:BRR

[370] Parigi Silvia. Book review: Richard L. Gregory, La mente nella scienza, Milano, Est Mondadori 1985, 421 pp. Nuncius, 3(2):338–339, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00835.

Silvia:1988:BRHa

[371] Parigi Silvia. Book review: Humberto Maturana, Francisco Varela, L’albero della conoscenza. Un nuovo meccanismo per spiegare le radici biologiche della conoscenza umana, (Italian) [The tree of knowledge. A new mechanism to explain the biological roots of human knowledge]. Presentazione di Mauro Ceruti, Milano, Garzanti 1987, 207 pp. Nuncius, 3(2): 339, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00844.

Silvia:1988:BRP

[372] Parigi Silvia. Book review: Paul Watzlawick (a cura di), La realtà inventata. Contributi al costruttivismo, Milano, Feltrinelli 1988, 281 pp. Nuncius, 3(2):340, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00853.

Silvia:1988:BRHb

[373] Parigi Silvia. Book review: Howard Gardner, La nuova scienza della mente. Storia della rivoluzione cognitiva, Milano, Feltrinelli 1988, 472 pp. Nuncius, 3(2):341, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00862.

Silvia:1988:E

[374] Parigi Silvia. Epistemologia. Nuncius, 3(2):341–342, ???? 1988. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539178x00871.

Stefano:1988:MR

[375] Caroti Stefano. Medioevo e rinascimento. Nuncius, 3(2):342–343, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00880.

Stefano:1988:BRH

[376] Caroti Stefano. Book review: H. L. L. Busard, the Mediaeval Latin Translation of Euclid’s ≪Elements≫, Stuttgart, Franz Steiner Verlag Wiesbaden GmbH 1987 (Boethius. Texte und Abhandlungen zur Geschichte der Exakten Wissenschaften, XV), 411 pp. Nuncius, 3(2): 343–344, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x00899.

Stefano:1988:BRG

[377] Caroti Stefano. Book review: Giulio D’Onofrio, Fons scientiae. La dialettica nell’Occidente tardo-antico, Napoli, Liguori Editore 1986 (Nuovo Medioevo, 31), 346 pp., ind. Nuncius, 3(2):344–345, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00907.

Stefano:1988:BRA

[378] Caroti Stefano. Book review: Adelard of Bath. An English Scientist and Arabist of the Early Twelfth Century, ed. by Ch. Burnett, London, the Warburg Institute 1987 (Warburg Institute Surveys and Texts, XIV), 208 pp., ill., ind. Nuncius, 3(2):345–347, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00916.

Stefano:1988:BRWa

[379] Caroti Stefano. Book review: Walter Burleigh, Von der Reinheit der Kunst der Logik. Erster Traktat. Von den Eigenschaften der Termini, übers. und mit Einf. und Anmerk. hrsg. von P. Kunze, Hamburg, Felix Meiner Verlag 1988 (Philosophische Bibliothek, 401), 269 pp., bibl., ind. Nuncius, 3(2):347–348, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00925.

Stefano:1988:BRO

[380] Caroti Stefano. Book review: Olaf Pluta, Die philosophische Psychologie des Peter von Ailly, Amsterdam, Verlag B. R. Grüner 1987 (Bochumer Studien zur Philosophie, 6), xxxix + 150 pp., bibl., ind. Nuncius, 3 (2):348–349, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x00934.

Stefano:1988:BRN

[381] Caroti Stefano. Book review: Nancy G. Siraisi, Avicenna in Renaissance Italy. The ≪ Canon ≫ and Medical Teaching in Italian Universities after 1500, Princeton, Princeton University Press 1987, 410 pp., ill., bibl., ind. Nuncius, 3(2):349–350, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00943.

Stefano:1988:BRWb

[382] Caroti Stefano. Book review: Der Weg der Naturwissenschaft von Johannes von Gmunden zu Johannes Kepler, (German) [The Way of Science from Johannes of Gmunden to Johannes Kepler], herausgegeben von G. Hamann und H. Grössing, Wien, Verlag der Österreichischen Akademie der Wissenschaften 1988 (Österreichische Akademie der Wissenschaften, Philosophisch-historische Klasse, Sitzungsberichte, 497), 214 pp., ill., ind. Nuncius, 3(2):351–352, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539178x00952.

Stefano:1988:SXX

[383] Casati Stefano. Secoli XVII–XX. Nuncius, 3(2):352–353, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00961.

Silvia:1988:BRF

[384] Parigi Silvia. Book review: Filippo Mignini, Ars imaginandi. Apparenza e rappresentazione in Spinoza, (Latin and Italian) [Art of imagination. Appearance and representation in Spinoza], Napoli, Edizioni Scientifiche Italiane 1981, 438 pp. Nuncius, 3(2):353–354, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00970.

Ferdinando:1988:BRA

[385] Abbri Ferdinando. Book review: Alan Charles Kors and Paul J. Korshin (eds.), Anticipations of the Enlightenment in England, France and Germany, Philadelphia, University of Pennsylvania Press 1987, viii + 290 pp. Nuncius, 3(2):354–356, ???? 1988. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539178x00989.

Ferdinando:1988:NRL

[386] Abbri Ferdinando. ≪ Nouvelles de la République des Lettres ≫, VI, 1986, n. 2, Napoli, Prismi 1987. Nuncius, 3(2):356–357, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x00998.

Silvia:1988:BRA

[387] Parigi Silvia. Book review: Anthony C. Grayling, Berkeley: the Central Arguments, London, Duckworth 1986, xii-218 pp. Nuncius, 3(2):357– 358, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x01005.

Silvia:1988:BRD

[388] Parigi Silvia. Book review: Daniel E. Flage, Berkeley’s Doctrine of Notions: A reconstruction based on his Theory of Meaning, London & Sydney, Croom Helm 1987, 226 pp. Nuncius, 3(2):358–359, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x01014.

Mara:1988:BRF

[389] Miniati Mara. Book review: Franz Pesendorfer, La Toscana dei Lorena. un secolo di governo granducale. Firenze, Sansoni Editore 1987 (≪ Biblioteca Storica ≫ ), 191 pp. Nuncius, 3(2):359, ????

1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539178x01023.

Mara:1988:BRD

[390] Miniati Mara. Book review: Dino Bravieri, P. Filippo Cecchi: scritti e strumenti. Firenze, Osservatorio Ximeniano 1988, 85 pp. Nuncius, 3 (2):359–360, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539178x01032.

Mara:1988:ADM

[391] Miniati Mara. Attivita dell’Istituto e Museo di Storia della Scienza di Firenze. Nuncius, 3(2):361–377, ???? 1988. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539178x01041.

Scriba:1988:Ab

[392] C. J. Scriba. Announcement. Nuncius, 3(2):378–390, ???? 1988. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539178x01050.

Losito:1989:ILA

[393] Maria Losito. Il IX libro del de architectura di Vitruvio nei commentari di Daniele Barbaro (1556–1567). Nuncius, 4(1):3–42, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00013.

Brooke:1989:SSK

[394] John Hedley Brooke. Science and the secularisation of knowledge: Perspectives on some Eighteenth Century transformations. Nuncius, 4(1): 43–65, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00022.

Boullin:1989:ISS

[395] David J. Boullin. An iconographic study of sandglasses. Nuncius, 4 (1):67–85, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00031.

DAgostino:1989:CDS

[396] Salvo D’Agostino. La classificazione degli strumenti scientifici. Nuncius, 4(1):87–92, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00040.

Proverbio:1989:PAD

[397] Edoardo Proverbio. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 4(1):93–150, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00059.

Anonymous:1989:AIA

[398] Anonymous. Alphabetical index of addressees. Nuncius, 4(1):151–153, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00068.

Anonymous:1989:AIS

[399] Anonymous. Alphabetical index of senders. Nuncius, 4(1):154–159, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00077.

Barbieri:1989:LLR

[400] Francesco Barbieri and Carla Fiori. L’ultima lettera di ruffini a Cauchy. Nuncius, 4(1):161–163, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00086.

Mazzolini:1989:IF

[401] Renato G. Mazzolini. Istituzioni e fonti. Nuncius, 4(1):165–176, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00095.

Belloni:1989:CAC

[402] Luigi Belloni. Commemorazione di Andrea Corsini. Nuncius, 4(1):177– 190, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00103.

Giacomoni:1989:GCS

[403] Paola Giacomoni. Goethe e la cultura scientifica. Nuncius, 4(1):191– 203, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00112.

Casati:1989:CIL

[404] Stefano Casati, Silvia Collini, and Antonella Vannoni. Convegno internazionale su Lamarck e il Lamarckismo. Nuncius, 4(1):205–215, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00121.

Abbri:1989:BRWa

[405] Ferdinando Abbri. Book review: William A. Cole, Chemical Literature 1700–1860. A Bibliography with Annotations, Detailed Descriptions, Comparisons and Locations, London-New York, Mansell Publishing Limited 1988, xxiii + 582 pp. Nuncius, 4(1):217–218, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00130.

Miniati:1989:BRJ

[406] Mara Miniati. Book review: Joseph Needham, Wang Ling, Derek J. De Solla Price, Heavenly clockwork. The great astronomical clocks of medieval China. Second edition with Supplement by John H. Combridge. Cambridge, Cambridge University Press 1986, 266 pp. Nuncius, 4(1): 218–220, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00149.

Lamberini:1989:BRP

[407] Daniela Lamberini. Book review: Pier Luigi Bassignana (a cura di), Le macchine di Valturio nei documenti dell’archivio storico Amma. Presentazione di Sergio Ricossa, contributi di Guido Amoretti, Pier Luigi Bassignana, Ferruccio Gambaruto, Vittorio Marchis, Alessandro Vitale-Brovarone, Torino, Umberto Allemandi & C. editore 1988, 176 pp., indice dei nomi, Lxiv tavole a colori, 50 ill. b. e n. Nuncius, 4(1):220–225, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00158.

Ernst:1989:BRJ

[408] Germana Ernst. Book review: J. D. North, Horoscopes and History. London, The Warburg Institute 1986, pp. xii-232 (≪ Warburg Institute Surveys and Texts ≫, XIII). Nuncius, 4(1):225–226, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00167.

Miniati:1989:BRPa

[409] Mara Miniati. Book review: Penelope Gouk, the Ivory Sundials of Nuremberg 1500–1700. Cambridge, Whipple Museum of the History of Science 1988, 144 pp. Nuncius, 4(1):226–228, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00176.

Iannaccone:1989:BRC

[410] Isaia Iannaccone. Book review: Chine ciel et terre: 5000 ans d’inventions et de découvertes. Gent, Imschoot 1988, 438 pp, fig. Nuncius, 4(1):229– 230, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00185.

Roero:1989:BRG

[411] Clara Silvia Roero. Book review: G. W. Leibniz, Mathematischer Naturwissenschaftlicher und Technischer Briefwechsel, Zweiter Band 1676– 1679, Berlin, Akademie Verlag 1987, pp. 1013 + 40. Nuncius, 4(1): 230–233, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00194.

Silvia:1989:BRR

[412] Parigi Silvia. Book review: Ruggiero Giuseppe Boscovich, Lettere ad Anton Mario Lorgna 1765–1785, a cura di Ugo Baldini e Pietro Nastasi, Roma, Accademia Nazionale delle Scienze detta dei Xl 1988, 142 pp. Nuncius, 4(1):233–237, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00202.

Crispini:1989:BRC

[413] Franco Crispini. Book review: Charles Dedeyan, Diderot et la pensée anglaise, Firenze, Leo S. Olschki Editore 1987, 365 pp. Nuncius, 4(1):

237–242, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00211.

Abbri:1989:BRRa

[414] Ferdinando Abbri. Book review: Robert G. W. Anderson, Christopher Lawrence (Eds.), Science, Medicine and Dissent: Joseph Priestley (1733– 1804), London, Wellcome Trust/ Science Museum 1987, xi + 105 pp. Nuncius, 4(1):242–245, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00220.

Guntau:1989:BRR

[415] Martin Guntau. Book review: Rudolf Stichweh, Zur Entstehung des modernen Systems wissenschaftlicher Disziplinen. Physik in Deutschland 1740–1890, Suhrkamp Verlag Frankfurt am Main 1984, 559 S. (ISBN 3-518-57688-7). Nuncius, 4(1):245–247, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00239.

Barsanti:1989:BRCa

[416] Giulio Barsanti. Book review: Claude Bénichou e Claude Blanckaert (a dura di), Julien-Joseph Virey, naturaliste et anthropologue. Paris, Vrin 1988, pp. 291, FF 129. Nuncius, 4(1):247–250, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00248.

Morabito:1989:BRG

[417] Carmela Morabito. Book review: Girolamo De Liguori, Materialismo inquieto. Bari, Laterza 1988, 238 pp., L. 31.000. Nuncius, 4(1):250–252, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00257.

LaVergata:1989:BRG

[418] Antonello La Vergata. Book review: Gillian Beer, Darwin’s plots. Evolutionary narrative in Darwin, George Eliot and Nineteenth-Century fiction, London, Boston, Melbourne and Henley, Ark Paperbacks 1983, xi, 303 pp. Nuncius, 4(1):253–254, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00266.

LaVergata:1989:BRM

[419] Antonello La Vergata. Book review: Marjorie Grene (ed.), Dimensions of Darwinism. Themes and counter-themes in Twentieth-Century evolutionary theory, Cambridge, Cambridge University Press; Paris, Éditions de la Maison des Sciences de l’Homme 1983, pp. ix, 336. Nuncius, 4 (1):254–256, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00275.

Dougherty:1989:BRE

[420] F. W. P. Dougherty. Book review: Ernst Haeckel, Das System der Medusen. Erster Teil einer Monographie der Medusen. Mit einem Atlas von vierzig Tafeln. Reprint of the original two volumes (vol. 1 = text, vol. 2 = atlas) published in Jena: 1880, with an Epilogue (in German and English) by Georg Uschmann (Dokumente zur Geschichte von Naturwissenschaft, Medizin und Technik, 9) Jena: Veb Gustav Fischer Verlag 1986. Nuncius, 4(1):256–259, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00284.

Barsanti:1989:BRR

[421] Giulio Barsanti. Book review: Renzo E. Scossiroli (a cura di), I primi passi della genetica. Scoperta e riscoperta delle leggi di Mendel sull’ereditarietà dei caratteri, (Italian) [The first steps of genetics. Discovery and rediscovery of the laws of Mendel’s inheritance of characteristics]. Milano, Jaka Book 1987, 108 pp., L. 15.000. Nuncius, 4(1):259–260, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00293.

Ipsen:1989:BRR

[422] Carl Ipsen. Book review: Robert N. Proctor, Racial hygiene: Medicine under the Nazis. Cambridge, Mass., Harvard University Press 1988, 414 pp. including appendixes, bibliography, notes, index + 43 figures, US$34.95 (cloth). Nuncius, 4(1):261–265, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00301.

Miniati:1989:Sa

[423] Mara Miniati. Strumentaria. Nuncius, 4(1):267, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00310.

Miniati:1989:BRAa

[424] Mara Miniati. Book review: Alan Stimson, The Mariner’s Astrolabe. A survey of known, surviving sea astrolabes. Utrecht, Hes Publishers 1988, 191 pp., 1 tav. f.t. Nuncius, 4(1):267–268, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00329.

Miniati:1989:BRRa

[425] Mara Miniati. Book review: Roberto Panicali, Orologi e orologiai del Rinascimento italiano. La scuola urbinate / l’horlogérie d’intërieur italienne au XVIe siécle et l’École du Duch d’Urbino / Sixteenth Century Italian Chamber clocks and the Urbino School. Urbino, Quattro Venti 1988, 145 pp. Nuncius, 4(1):268–269, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00338.

Miniati:1989:BRPb

[426] Mara Miniati. Book review: Paolo Francesco Forlati, Segnatempo ≪ veronensis ≫. Verona, Grafiche Fiorini 1987, 163 pp. Nuncius, 4(1): 269–270, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00347.

Miniati:1989:BRMa

[427] Mara Miniati. Book review: Maya Hambly, Drawing Instruments 1580– 1980. London, Sotheby’s Publications 1988, 206 pp. Nuncius, 4(1):270, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00356.

Miniati:1989:BRMb

[428] Mara Miniati. Book review: La Misura del Tempo dall’orologio solare all’orologio atomico. (Italian) [The measure of time from the sundial to the atomic clock]. Milano, Museo Nazionale della Scienza e della Tecnica 1988, 98 pp., L. 40.000. Nuncius, 4(1):270–271, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00365.

Miniati:1989:BRO

[429] Mara Miniati. Book review: Orologi negli arredi del Palazzo Reale di Torino e delle residenze sabaude. (Italian) [Watches in the furnishings of the Royal Palace of Turin and the Savoy residences] Milano, Fabbri Editori 1988, 367 pp. Nuncius, 4(1):271, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00374.

Miniati:1989:BRH

[430] Mara Miniati. Book review: Henrique Morize, Observatorio Astronomico — um século de historia (1827–1927). Rio de Janeiro, Museu de Astronomia e Ciencias Afins 1987, 179 pp. (Coleçao Documentos de Historia da Ciencia, 1). Nuncius, 4(1):272, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00383.

Miniati:1989:BRAb

[431] Mara Miniati. Book review: Agnese Visconti, I 150 anni del Museo Civico di Storia Naturale di Milano (1838–1988). Milano, Museo Civico di Storia Naturale 1988, 51 pp. Nuncius, 4(1):272–273, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00392.

Malacrida:1989:BRG

[432] Lorenzo Marri Malacrida. Book review: Giorgio e Michele Bellagarda, Storia Illustrata dell’Arte Dentaria. Torino, Edizioni Minerva Medica 1987, 367 pp. Nuncius, 4(1):273, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00400.

Malacrida:1989:BRH

[433] Lorenzo Marri Malacrida. Book review: Heinrich Schipperges, Il giardino della salute. La medicina nel medioevo. Milano, Garzanti 1988, 273 pp., L. 19.500. Nuncius, 4(1):273–274, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00419.

Lippi:1989:BRC

[434] Donatella Lippi. Book review: C. Coury, La medicina dell’America precolombiana, Saronno, Ciba-Geigy Edizioni 1987 (Ed. originale: La médecine de l’Amérique précolombienne, Paris, R. Dacosta 1969 e 1982, 305 pp.). Nuncius, 4(1):274–276, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00428.

Malacrida:1989:BRS

[435] Lorenzo Marri Malacrida. Book review: Di sana pianta. Erbari e taccuini di sanità. Le radici della nuova farmacologia. (Italian) [From scratch: herbals and notebooks on health. The roots of the new pharmacology]. Modena, edizioni Panini 1988, 159 pp., L. 30.000. Nuncius, 4(1):276– 277, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00437.

Lippi:1989:BRGa

[436] Donatella Lippi. Book review: G. Penso, Le piante medicinali nell’arte e nella storia, Saronno, Ciba-Geigy Edizioni 1986 (Ed. originale: Les plantes médicinales dans l’art et l’histoire, Paris, Ed. R. Dacosta 1986, 300 pp.). Nuncius, 4(1):277–278, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00446.

Lippi:1989:BRGb

[437] Donatella Lippi. Book review: Giorgio Cosmacini, La Medicina e la sua storia, Milano, Rizzoli 1989, 236 pp., L. 32.000. Nuncius, 4(1):278–279, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00455.

Abbri:1989:BRRb

[438] Ferdinando Abbri. Book review: Ronald L. Numbers (ed.), Medicine in the New World. New Spain, New France, and New England, Knoxville, The University of Tennessee Press 1987, 175 pp. Nuncius, 4(1):279–281, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00464.

Abbri:1989:BRPa

[439] Ferdinando Abbri. Book review: Peter N. Stearns (ed.), Expanding the Past: A Reader in Social History. Essays from the ≪ Journal of Social History ≫, New York-London, New York University Press 1988, x + 394 pp. Nuncius, 4(1):281–283, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00473.

Abbri:1989:BRPb

[440] Ferdinando Abbri. Book review: Pnina G. Abir-Am, Dorinda Outram (ed.s), Uneasy Careers and Intimate Lives. Women in Science 1789–1979, with a Foreword by Margaret W. Rossiter, New Brunswick-London, Rutgers University Press 1987, xv + 365 pp. (≪ the Douglass series on women’s lives and the meaning of gender ≫ ). Nuncius, 4(1):283–285, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00482.

Abbri:1989:BRF

[441] Ferdinando Abbri. Book review: Francis Paul Prucha, The Indians in American Society. From the Revolutionary War to the Present, Berkeley– Los Angeles, University of California Press 1988, x + 127 pp. (≪ A Quantum Book ≫ ). Nuncius, 4(1):286, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00491.

Caroti:1989:BRGa

[442] Stefano Caroti. Book review: Geoffrey E. R. Lloyd, Scienza folclore ideologia. Le scienze della vita nella Grecia antica. Torino, Boringhieri 1987 (≪ Società antiche ≫ ), 225 pp., bibl., ind. Nuncius, 4(1):287–288, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00509.

Caroti:1989:BRCa

[443] Stefano Caroti. Book review: Christian Wildberg, John Philoponus’ Criticism of Aristotele’s Theory of Aether. Berlin-New York, Walter de Gruyter 1988 (≪ Peripatoi. Philo-logisch-historische Studien zum Aristotelismus ≫, Bd. 16), 274 pp., bibl., ind. Nuncius, 4(1):289–290, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00518.

Caroti:1989:BRGb

[444] Stefano Caroti. Book review: Gilbert de Poitiers et ses contemporains. Aux origines de la≪Logica modernorum≫. Actes du septième symposium européen d’histoire de la logique et de la sémantique médiévales. Centre d’Études Supérieures de Civilisation Médiévale de Poitiers. Poitiers 17– 22 juin 1985, edité par Jean Jolivet et Alain de Libera. Napoli, Bibliopolis 1987 (≪ History of Logic ≫, V), 481 pp., ind. Nuncius, 4(1):291–294, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00527.

Caroti:1989:BRM

[445] Stefano Caroti. Book review: Marilyn McCord Adams, William Ockham. Notre Dame, Indiana, University of Notre Dame Press 1987 (≪ Publications in Medieval Studies. The Medieval Institute. University of Notre Dame ≫, XXVI/1–2), 2 voll., 1402 pp., bibl., ind. Nuncius, 4 (1):294–296, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00536.

Caroti:1989:BRJa

[446] Stefano Caroti. Book review: Jole Agrimi, Chiara Crisciani, Edocere medicos. Medicina scolastica nei secoli XIII–XV. Napoli, Guerini e associati 1988 (Istituto Italiano per gli Studi Filosofici, ≪ Hippocratica civitas ≫, II), 309 pp., bibl., ind. Nuncius, 4(1):296–297, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00545.

Caroti:1989:BRL

[447] Stefano Caroti. Book review: Luca Bianchi, L’inizio dei tempi. Antichità e novità del mondo da Bonaventura a Newton, (Italian) [The beginning of time. Antiquity and news of the world from Bonaventure to Newton]. Firenze, L. S. Olschki Editore 1987 (≪ Biblioteca di Storia della Scienza ≫, 26), 113 pp., ind. Nuncius, 4(1):298–299, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00554.

Caroti:1989:BRR

[448] Stefano Caroti. Book review: Rapporti tra le Università di Padova e Bologna. Ricerche di filosofia, medicinae scienza. Omaggio dell’Università di Padova all’≪Alma Mater≫ bolognese nel suo nono centenario, a cura di Lucia Rossetti. Trieste, Edizioni Lint 1988 (Centro per la Storia dell’Università di Padova. ≪ Contributi ≫, 20), 439 pp., ill., ind. Nuncius, 4(1):299–302, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00563.

Caroti:1989:BRP

[449] Stefano Caroti. Book review: Paolo Malanima, I piedi di legno. Una macchina alle origini dell’industria medievale, Milano, Franco Angeli 1988 (≪ Studi e ricerche storiche ≫, 110), 149 pp., ill., bibl. Nuncius, 4(1):302–303, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00572.

Parigi:1989:BRJa

[450] Silvia Parigi. Book review: John W. Yolton, Perceptual Acquaintance from Descartes to Reid, Oxford, Basil Blackwell 1984, x-248 pp. Nuncius, 4(1):303–304, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00581.

Parigi:1989:BRJb

[451] Silvia Parigi. Book review: Jonathan Dancy, Berkeley: An Introduction, Oxford and New York, Basil Blackwell 1987, 165 pp. Nuncius, 4(1): 304–305, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00590.

Miniati:1989:BRB

[452] Mara Miniati. Book review: Bossi M.–Tucci P. (eds.), Bicentennial commemoration of R. G. Boscovich. Milano, September 15–18, 1987. Proceedings. s.l., Edizioni Unicopli 1988, 281 pp. Nuncius, 4(1):305–306, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00608.

Miniati:1989:BRD

[453] Mara Miniati. Book review: Danilo Barsanti, Stanislao Canovai e Gaetano Del Ricco. Il sodalizio umano e scientifico di due scolopi fiorentini fra XVIII e XIX secolo. Firenze, Osservatorio Ximeniano 1988, 105 pp. Nuncius, 4(1):306, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00617.

Abbri:1989:BRH

[454] Ferdinando Abbri. Book review: Harold B. Carter, Sir Joseph Banks 1743–1820, London, British Museum (Natural History) 1988, xi + 671 pp. Nuncius, 4(1):307–308, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00626.

Miniati:1989:BRRb

[455] Mara Miniati. Book review: Renato Pasta, Scienza politica e rivoluzione. l’opera di Giovanni Fabbroni (1752–1822) intellettuale e funzionario al servizio dei Lorena, Firenze, Leo S. Olschki Editore 1989, 639 pp. Nuncius, 4(1):308–309, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00635.

Beretta:1989:BRR

[456] Marco Beretta. Book review: Roberto Matocchi, La ’belle époque ’ dell’atomo. Ricerche sulla vittoria dell’atomismo del primo novecento, Milano, Franco Angeli Editore 1988, 315 pp. Nuncius, 4(1):309–310, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00644.

Miniati:1989:BRF

[457] Mara Miniati. Book review: Franco Soresini, La radio — Wireless sets. Milano, Be-Ma Editrice 1988, 141 pp. (≪ Itinerari d’immagini ≫, 14). Nuncius, 4(1):311, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00653.

Miniati:1989:ADM

[458] Mara Miniati. Attivita dell’Istituto e Museo di Storia della Scienza di Firenze. Nuncius, 4(1):313–350, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00662.

Baroncini:1989:NSF

[459] Gabriele Baroncini. Note sulla formazione del lessico della metafora Machina mundi. Nuncius, 4(2):3–30, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00671.

Field:1989:PSM

[460] J. V. Field, R. Lunardi, and T. B. Settle. The perspective scheme of Masaccio’s Trinity fresco. Nuncius, 4(2):31–118, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00680.

Beretta:1989:SIR

[461] Marco Beretta. Gli scienziati italiani e la rivoluzione chimica. Nuncius, 4(2):119–146, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00699.

Mandrino:1989:GBA

[462] A. Mandrino, G. Tagliaferri, and P. Tucci. G. B. Amici’s telescopes for Brera Observatory. Nuncius, 4(2):147–164, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00707.

DeGregorio:1989:PAD

[463] Mario De Gregorio. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 4(2):165–211, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00716.

Anonymous:1989:IC

[464] Anonymous. Indice cronologico. Nuncius, 4(2):212–214, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00725.

Anonymous:1989:IA

[465] Anonymous. Indice alfabetico. Nuncius, 4(2):215–217, ???? 1989. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00734.

Scheffer:1989:I

[466] Carl Fredrick Scheffer. I. Nuncius, 4(2):218–219, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00743.

Catellacci:1989:I

[467] Antonio Catellacci. II. Nuncius, 4(2):219–220, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00752.

Thomson:1989:I

[468] William Thomson. III. Nuncius, 4(2):220, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00761.

Dandolo:1989:I

[469] Vincenzo Dandolo. IV. Nuncius, 4(2):221, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00770.

Spallanzani:1989:Va

[470] Lazzaro Spallanzani. V. Nuncius, 4(2):221–222, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00789.

Spallanzani:1989:Vb

[471] Lazzaro Spallanzani. VI. Nuncius, 4(2):222–223, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00798.

Barzellotti:1989:V

[472] Giacomo Barzellotti. VII. Nuncius, 4(2):223–225, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00806.

Fabbroni:1989:V

[473] Giovanni Fabbroni. VIII. Nuncius, 4(2):225–227, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00815.

Fabbroni:1989:I

[474] Giovanni Fabbroni. IX. Nuncius, 4(2):227–230, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00824.

Capelle:1989:X

[475] Guillaume Antoine Benot Capelle. X. Nuncius, 4(2):231, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00833.

DeLasteyrieDuSaillant:1989:X

[476] Charles Philibert De Lasteyrie Du Saillant. XI. Nuncius, 4(2):231–232, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00842.

DeSaintFond:1989:X

[477] Barthelemy Faujas De Saint Fond. XII. Nuncius, 4(2):232–233, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00851.

Thouin:1989:X

[478] Andr Thouin. XIII. Nuncius, 4(2):233–234, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00860.

Bodard:1989:X

[479] Pierre Henry Hippolyte Bodard. XIV. Nuncius, 4(2):234–237, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00879.

Hauy:1989:X

[480] Rene Just Hauy. XV. Nuncius, 4(2):237–238, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00888.

Borson:1989:X

[481] Stefano Borson. XVI. Nuncius, 4(2):238–239, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00897.

Mazzoni:1989:Xa

[482] Giovanni Battista Mazzoni. XVII. Nuncius, 4(2):240–241, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00905.

Mazzoni:1989:Xb

[483] Giovanni Battista Mazzoni. XVIII. Nuncius, 4(2):241–243, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00914.

Savi:1989:X

[484] Paolo Savi. XIX. Nuncius, 4(2):243–244, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00923.

DeCandolle:1989:X

[485] Augustin Pyramus De Candolle. XX. Nuncius, 4(2):244–248, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00932.

Calisi:1989:LDC

[486] Marinella Calisi. L’Origine delle Collezioni Scientifiche del Museo Astronomico e Copernicano in Roma alla fine del XIX secolo. Nuncius, 4 (2):249–263, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x00941.

Anonymous:1989:D

[487] Anonymous. Documenti. Nuncius, 4(2):264–302, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x00950.

Abbri:1989:CRC

[488] Ferdinando Abbri. The chemical revolution: a critical assessment. Nuncius, 4(2):303–315, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00969.

Turner:1989:BRJ

[489] Gerard L’E. Turner. Book review: J. A. Bennett, the Divided Circle: A History of Instruments for Astronomy, Navigation and Surveying. Oxford, Phaidon-Christie’s 1987, 224 pp., 260 illustrations, £45.00. Nuncius, 4(2):317–318, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x00978.

Abbri:1989:BRWb

[490] Ferdinando Abbri. Book review: William R. Shea (Ed.), Revolutions in Science. Their Meaning and Relevance. Canton, Science History Publications/U.S.A. 1988, pp. xi + 291. Nuncius, 4(2):318–321, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x00987.

Abbri:1989:BRI

[491] Ferdinando Abbri. Book review: Ingrid Merkel, Allen G. Debus (Ed.s), Hermeticism and the Renaissance. Intellectual History and the Occult in Early Modern Europe. Washington, the Folger Shakespeare Library, London and Toronto, Associated University Presses 1988, 438 pp. (Folger Books). Nuncius, 4(2):322–326, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x00996.

Webster:1989:BRP

[492] Charles Webster. Book review: Paracelsus und Paracelsisten: Vorträge 1984/85. Salzburger Beiträge zur Paracelsusforschung, Folge 25, (German) [Paracelsus and Paracelsists: Lectures 1984/85. Salzburg Paracelsus contributions to research, Episode 25], Verband der wissenschaftlichen Gesellschaft Österreichs, Vienna 1987, 190 pp. Nuncius, 4(2):326–327, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x01003.

Barsanti:1989:BRCb

[493] Giulio Barsanti. Book review: Carlo Sarti, I fossili e il diluvio universale. Le collezioni settecentesche del Museo di Geologia e Paleontologia dell’Università di Bologna. Bologna, Pitagora 1988, xvi + 189 pp., L. 25.000. Nuncius, 4(2):327–330, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x01012.

Ordoez:1989:BRH

[494] Javier Ordoez and Alberto Elena. Book review: Horacio Capel, Joan Eugeni Snchez, Omar Moncada, De Palas a Minerva. La formación científica y la estructura institucional de los ingenieros militares en el siglo XVIII. Barcellona-Madrid, Libros de Serbal Consejo Superior de Investigaciones Científicas 1988, 390 pp. Nuncius, 4(2):330–331, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01021.

Beretta:1989:BRT

[495] Marco Beretta. Book review: Tore Frngsmyr (Ed.), Science in Sweden, the Royal Swedish Academy of Sciences 1739–1989. Canton, Ma, Science History Publications 1989, viii + 291 pp. Nuncius, 4(2):332–333, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01030.

Arrigoni:1989:BRS

[496] Tiziano Arrigoni. Book review: Serge Borsa, Claude-Ren Michel, Gli ospedali in Francia nel XIX secolo, a cura di Cesare Stevan. Milano, Edizioni Unicopli 1987, 209 pp. (ed. originale: La vie quotidienne des Hôpitaux en France au XIXe siècle, Paris, Hachette 1985). Nuncius, 4 (2):333–335, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x01049.

Beretta:1989:BRW

[497] Marco Beretta. Book review: W. Strube, Der historische Weg der Chemie, von der Urzeit bis zur industriellen Revolution. Leipzig, Veb Deutscher Verlag fr Grundstoffindustrie 1984–1986, 2 voll., 188 e 219 pp. Nuncius, 4(2):335–336, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x01058.

Castellani:1989:BRK

[498] Elena Castellani. Book review: Klaus Mainzer, Symmetrien der Natur. Ein Handbuch zur Natur-und Wissenschaftsphilosophie. Berlin/New York, De Gruyter 1988, xi-739 pp. Nuncius, 4(2):336–343, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01067.

Malacrida:1989:BRF

[499] Lorenzo Marri Malacrida. Book review: F. Fenner, D. A. Henderson, I. Arita, Z. Jezek, I. D. Ladnyi, Smallpox and its Eradication. Geneva, World Health Organization 1988, 1464 pp. Nuncius, 4(2):343– 345, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01076.

Malacrida:1989:BRP

[500] Lorenzo Marri Malacrida. Book review: Pierre Huard, Zensetsou Ohya, Ming Wong, La medicina giapponese dalle origini ai nostri giorni. Milano, Ciba-Geigy 1988, 312 pp. Nuncius, 4(2):345–346, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01085.

Miniati:1989:Sb

[501] Mara Miniati. Strumentaria. Nuncius, 4(2):347, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x01094.

Miniati:1989:BRMc

[502] Mara Miniati. Book review: Musiques mécaniques. Musée Savoisi en Chambry 1er décembre 1988–1er mars 1989. Chambry, Nallet 1988, 198 pp., fig. + 1 disco 45 giri. Nuncius, 4(2):347–348, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01102.

Lippi:1989:BRA

[503] Donatella Lippi. Book review: Alfonso Pulchinotta, Storia illustrata della senologia. Fra scienza e mito. Saranno, Ciba-Geigy Edizioni 1989, 573 pp. Nuncius, 4(2):348–349, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x01111.

Guarnieri:1989:BRS

[504] Patrizia Guarnieri. Book review: Salvatore Trinchese, Opere, a cura di Guido Cimino. Roma-Napoli, Edizioni Theoria 1989, 326 pp. Nuncius, 4(2):349–350, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x01120.

Lippi:1989:BRI

[505] Donatella Lippi. Book review: Israel Rosenfield, l’invenzione della memoria. Milano, Rizzoli 1989, 248 pp. Nuncius, 4(2):351, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01139.

Abbri:1989:C

[506] Ferdinando Abbri. Chimica. Nuncius, 4(2):352–353, ???? 1989. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x01148.

Abbri:1989:BRCa

[507] Ferdinando Abbri. Book review: Carl-Wilhelm Scheele-Ehrung 1986, Akademie der Wissenschaften der Ddr. Institut fur Theorie, Geschichte und Organisation der Wissenschaft. Berlin 1987 (Kolloquien 62), 231 pp. Nuncius, 4(2):354–356, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x01157.

Beretta:1989:BRG

[508] Marco Beretta. Book review: A. Greiner, H. Klare (a cura di), Chemiker ber Chemiker, Wahlvorschlge zur Aufnahme von Chemikern in die Berliner Akademie 1822–1925. Berlin, Akademie Verlag 1986, 253 pp.

Nuncius, 4(2):356–357, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x01166.

Barsanti:1989:B

[509] Giulio Barsanti. Botanica. Nuncius, 4(2):357–358, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539189x01175.

Caroti:1989:MR

[510] Stefano Caroti. Medioevo e rinascimento. Nuncius, 4(2):358–359, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01184.

Caroti:1989:BRJb

[511] Stefano Caroti. Book review: John D. North, Stars, Minds and Fate. Essays in Ancient and Medieval Cosmology. London and Ronceverte, the Hambledon Press 1989, 429 pp., ill. ind. John D. North, the Universal Frame. Historical Essays in Astronomy Natural Philosophy and Scientific Method. London and Ronceverte, the Hambledon Press 1989, 384 pp., ill. ind. Nuncius, 4(2):359–360, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x01193.

Caroti:1989:BRCb

[512] Stefano Caroti. Book review: Claude Lafleur, Quatre introductions la philosophie au XIIIe siécle. Textes critiques et tude critique. Montral-Paris, Institut d’Études Médiévales de Montréal — Librairie Philosophique J. Vrin 1988, 428 pp., ill., bibl., ind. (Universiteé de Montréal. Publications de l’Institut d’Études Médiévales, XXIII). Nuncius, 4(2):360–361, ???? 1989. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539189x01201.

Abbri:1989:SXX

[513] Ferdinando Abbri. Secoli XVII–XX. Nuncius, 4(2):362–363, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01210.

Abbri:1989:BRPc

[514] Ferdinando Abbri. Book review: Pietro Nastasi (a cura di), Atti del Convegno Il Meridione e le Scienze (secoli XVI–XIX). Palermo, 14–16 Maggio 1985. Palermo, Tipografia Luxograph 1988, 621 pp. (Istituto Gramsci Siciliano, Acta 1). Nuncius, 4(2):363–364, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01229.

Abbri:1989:BRCb

[515] Ferdinando Abbri. Book review: Clelia Pighetti, L’Influsso scientifico di Robert Boyle nel tardo ’600 italiano. Milano, Franco Angeli 1988, 275 pp. (Filosofia e scienza nel Cinquecento e nel Seicento). Nuncius, 4(2): 364–365, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539189x01238.

Abbri:1989:BRA

[516] Ferdinando Abbri. Book review: André Robinet, G. W. Leibniz Iter Italicum (Mars 1689-Mars 1690). La dynamique de la République des Lettres. Nombreux textes inédits. Firenze, Leo S. Olschki Editore 1988, viii + 493 pp. (Accademia Toscana di Scienze e Lettere ‘La Colombaria’, Studi XC). Nuncius, 4(2):365–366, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x01247.

Abbri:1989:BRCc

[517] Ferdinando Abbri. Book review: Cultura e società nel Settecento lorenese. Arezzoe la Fratenita dei laici. (Italian) [Culture and society in Eighteenth Century Lorraine. Arezzoe of the Brotherhood of Laity]. Firenze, Olschki Editore 1988, 310 pp. (Biblioteca Storica Toscana. Sezione di storia del Risorgimento, a cura della Società Toscana di Storia del Risorgimento 14). Nuncius, 4(2):367–368, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x01256.

Miniati:1989:BRAc

[518] Mara Miniati. Book review: Agnese Visconti, George-Louis Leclerc de Buffon (1707–1788). Milano, Museo Civico di Storia Naturale di Milano 1988, 35 pp., fig. Nuncius, 4(2):368–369, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01265.

Casati:1989:BRI

[519] Stefano Casati. Book review: Istituto Nazionale di Geofisica, Mallet’s macroseismic survey on the Neapolitan earthquake of 16th December 1857. Emanuela Guidoboni and Graziano Ferrari editors. Bologna, Sga 1987. Voll. 4. Nuncius, 4(2):369, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x01274.

Abbri:1989:BRG

[520] Ferdinando Abbri. Book review: George Levine (Ed.), One Culture. Essays in Science and Literature. With the assistance of Alan Rauch. Madison, The University of Wisconsin Press 1987, viii + 359 pp. (Science and Literature). Nuncius, 4(2):370–372, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x01283.

Anonymous:1989:ADM

[521] Anonymous. Attivit dell’Istituto e Museo di Storia della Scienza di Firenze I. Nuncius, 4(2):373–382, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539189x01292.

Generali:1989:PQS

[522] Dario Generali. Pubblichiamo qui di seguito le considerazioni di Dario Generali sulla recensione di Giulio Barsantia G. A. di Porcia, notizie della vita, e degli studi del kavalier Antonio Vallisneri (Nuncius III, 1988, fasc. 2). Ad esse segue la replica di Giulio Barsanti. Nuncius, 4(2):383–387, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01300.

Barsanti:1989:ADN

[523] Giulio Barsanti. Alla direzione di Nuncius. Nuncius, 4(2):387–398, ???? 1989. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539189x01319.

Goldstein:1990:GAA

[524] Bernard R. Goldstein. Galileo’s account of astronomical miracles in the Bible: a confusion of sources. Nuncius, 5(1):3–16, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00651.

Casati:1990:GTL

[525] Stefano Casati. Giuseppe Toaldo: la luna, il Saros e le meteore. Nuncius, 5(1):17–42, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00660.

Gee:1990:FPO

[526] Brian Gee. Faraday’s plight and the origins of the magneto-electric spark. Nuncius, 5(1):43–69, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00679.

Fauque:1990:APD

[527] Danielle Fauque. Amici et le principe de duplication d’image: Le micrométre d’Amici. Nuncius, 5(1):71–77, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00688.

Vaccari:1990:PAD

[528] Ezio Vaccari. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 5(1):79–126, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00697.

Muljacic:1990:PIC

[529] Zarko Muljacić. Per un inventario del carteggio di Alberto Fortis. Nuncius, 5(1):127–203, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00705.

Mazzolini:1990:E

[530] Renato G. Mazzolini. Éloge. Nuncius, 5(1):205–238, ???? 1990. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00714.

Segala:1990:DC

[531] Marco Segala. Discussioni critiche. Nuncius, 5(1):239–256, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00723.

Beretta:1990:BRWa

[532] Marco Beretta. Book review: W. F. Bynum, E. J. Browne e R. Porter, Dizionario di storia della scienza, Roma, Theoria 1988 (Ed. originale Dictionary of the History of Science, London, MacMillan Press 1981), 566 pp. Nuncius, 5(1):257–258, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00732.

Beretta:1990:BRHa

[533] Marco Beretta. Book review: Helge Kragh, An introduction to the historiography of science, Cambridge, Cambridge University Press 1987, viii + 235 pp. Nuncius, 5(1):258–260, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00741.

Morpurgo:1990:BRR

[534] Piero Morpurgo. Book review: R. M. Thomson (ed.), Alexander Nequam Speculum Speculationum, Oxford, the British Academy 1988, xxiii + 507 pp. (≪ Auctores Britannici Medii Aevi ≫, XI). Nuncius, 5(1): 260–262, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00750.

Morpurgo:1990:BRJ

[535] Piero Morpurgo. Book review: J. K. Otte, Alfred of Sareshel’s Commentary on the Metheora of Aristotle, Leiden 1988, viii + 131 pp. (≪ Studien und Texte zur Geistesgeschichte des Mittelalters ≫, XIX). Nuncius, 5(1):262–265, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00769.

Iannaccone:1990:BRJ

[536] Isaia Iannaccone. Book review: Jean-Claude Martzloff, Histoire des Mathématiques Chinoises, Paris, Masson 1988, 375 pp. Nuncius, 5(1): 265–268, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00778.

Abbri:1990:BRAa

[537] Ferdinando Abbri. Book review: Anthony Pagden, La caduta dell’uomo naturale. L’indiano d’America e le origini dell’etnologia comparata, Torino, Einaudi 1989, xxvi + 300 pp. (≪ Einaudi Paperbacks 197 ≫ ). Nuncius, 5(1):269–271, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00787.

Ciancio:1990:BRR

[538] Luca Ciancio. Book review: Rachel Laudan, From Mineralogy to Geology. The Foundation of a Science, 1650–1830, Chicago and London, The University of Chicago Press 1987, xii + 278 pp. Nuncius, 5(1):271–275, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00796.

Simonutti:1990:BRN

[539] Luisa Simonutti. Book review: Nicholas Tyacke, Anti-Calvinists. The Rise of English Arminianism, c. 1590–1640, Oxford, Clarendon Presso 1987, 305 pp. Nuncius, 5(1):275–276, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00804.

Miniati:1990:BRCa

[540] Mara Miniati. Book review: Centre Regional De Publication De Paris, Jean Picard et les débuts de l’astronomie de précision au XVIIe siècle. Actes du colloque du tricentenaire édités par Guy Picolet, Paris, Éditions du Centre National de la Recherche Scientifique 1987, 382 pp., fig. Nuncius, 5(1):277–279, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00813.

Simonutti:1990:BRH

[541] Luisa Simonutti. Book review: Henri Basnage de Beauval en de ≪ histoire des ouvrages des savans≫, 1687–1709. Verkennigen binnen de Republiek der Letteren aan de vooravond van de verlichting, (Dutch) [asnage Henri de Beauval and the llHistoire des ouvrages des savans≫ [history of the works of scientists]: 1687–1709: studies in the Republic of Letters on the eve of lighting] uitgegeven door Hans Bots, Amsterdam, Holland Universiteits Pers, eerste druk 1976, 2 voll., s.p.i. Nuncius, 5(1):279–282, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00822.

Dougherty:1990:BRR

[542] F. W. P. Dougherty. Book review: Rainer Baasner, Das Lob der Sternkunst. Astronomie in der deutschen Aufklärung, Göttingen, Vandenhoeck & Ruprecht 1987, 239 pp., ill. (Abhandlungen der Akademie der Wissenschaften in Göttingen. Mathematisch-physikalische Klasse,

3. Folge Nr. 40). Nuncius, 5(1):283–284, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00831.

Beretta:1990:BRJ

[543] Marco Beretta. Book review: Jeremy Adler, ≪ Eine fast magische Anziehungskraft ≫. Goethes Wahlvervandtschaften und die Chemie seiner Zeit, München, Verlag C. H. Beck 1987, 260 pp. Nuncius, 5(1): 284–286, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00840.

Abbri:1990:BRLa

[544] Ferdinando Abbri. Book review: Ludmilla Jordanova, Sexual Visions. Images of Gender in Science and Medicine between the Eighteenth and Twentieth Centuries, Hemel Hempstead, Harvester Wheatsheaf 1989, xiii + 207 pp. Nuncius, 5(1):287–289, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00859.

Morabito:1990:BRM

[545] Carmela Morabito. Book review: Mary A. B. Brazier, A History of Neurophysiology in the 19th Century, New York, Raven Press 1988, xvi + 265 pp. Nuncius, 5(1):289–292, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00868.

Wanrooij:1990:BRT

[546] Bruno P. F. Wanrooij. Book review: Tommaso Detti, Fabrizio Maffi. Vita di un medico socialista, Milano, Franco Angeli 1987, 311 pp. Nuncius, 5(1):292–295, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00877.

Dougherty:1990:BRC

[547] F. W. P. Dougherty. Book review: Christoph Kockerbeck, Ernest Haeckels ≪ Kunstformen der Natur ≫ und ihr Einfluß auf die deutsche bildende Kunst der Jahrhundertwende. Studie zum Verhältnis von Kunst und Naturwissenschaften im Wilhelminischen Zeitalter, Frankfurt am Main, Peter Lang 1986, v + 246 pp., SFR 56,00 (≪ European University Studies, Series XX ≫, Vol. 194). Nuncius, 5(1):295–296, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00886.

Beretta:1990:BRHb

[548] Marco Beretta. Book review: Héléne Metzger, La méthode philosophique en histoire des sciences. Textes 1914–1939, Paris, Fayard 1987, 277 pp. Nuncius, 5(1):296–298, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00895.

Brooke:1990:BRR

[549] John Hedley Brooke. Book review: Robert J. Russell, William R. Stoeger S.J., George V. Coyne S.J. (eds.), Physics, Philosophy and Theology: a Common Quest for Understanding, Vatican City State 1988, 419 pp. Nuncius, 5(1):299–302, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00903.

Parigi:1990:BRP

[550] Silvia Parigi. Book review: Paolo Bozzi, Fenomenologia sperimentale, Bologna, Il Mulino 1989 (≪ Ricerca ≫ ), 337 pp. Nuncius, 5(1):303– 304, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00912.

Ciancio:1990:BRW

[551] Luca Ciancio. Book review: S. Warren Carey, Theories of the Earth and Universe. A History of Dogma in the Earth Sciences, Stanford, Stanford University Press 1988, xviii + 413 pp. Nuncius, 5(1):304–306, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00921.

Arrigoni:1990:BRC

[552] Tiziano Arrigoni. Book review: Civiltà delle macchine. Antologia di una rivista 1953–1957, a cura di Vanni Scheiwiller, Milano, Libri Scheiwiller-Finmeccanica 1988, 529 pp. Nuncius, 5(1):306–308, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00930.

Dougherty:1990:SB

[553] F. W. P. Dougherty. Strumenti bibliografici. Nuncius, 5(1):309, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00949.

Malacrida:1990:BRR

[554] Lorenzo Marri Malacrida. Book review: Roy Porter (a cura di), Dizionario Biografico della Storia della Medicina e delle Scienze Naturali. Milano, Franco Maria Ricci editore 1989, 243 pp. Vol. 4. Nuncius, 5(1):309–310, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00958.

Simonutti:1990:BRL

[555] Luisa Simonutti. Book review: Letteree carte Magliabechi. Inventario cronologico a cura di Manuela Doni Garfagnini, Roma, Istituto storico italiano per l’etä moderna e contemporanea 1988, xxii-651 pp., L. 45.000 (≪ Fonti per la Storia d’Italia ≫ ). Nuncius, 5(1):310–311, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00967.

Miniati:1990:BRLa

[556] Mara Miniati. Book review: Livia Giacardi, Clara Silvia Roero (a cura di), Bibliotheca Mathematica. Documenti per la storia della matematica nelle biblioteche torinesi. Catalogazione di Giuseppe Slaviero. Torino, Umberto Allemandi & C. 1987, 222 pp., fig., L. 50.000. Nuncius, 5(1): 312, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00976.

Dougherty:1990:BRD

[557] F. W. P. Dougherty. Book review: Dietrich Von Engelhardt (compilor), Wissenschaftsgeschichte auf den Versammlungen der Gesellschaft Deutscher Naturforscher und Ärzte 1822–1972. Bibliographie der Vorträge und allgemeine Übersicht (Schriftenreihe zu Geschichte der Versammlungen der Gesellschaft Deutscher Naturforscher und Ärzte, Bd. 4). Stuttgart, Wissenschaftliche Verlagsgesellschaft 1987, 206 pp., ill. Nuncius, 5(1):312–313, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00985.

Mintati:1990:BRP

[558] Mara Mintati. Book review: Pier Luigi Bassignana (a cura di), La cultura delle macchine dal Medioevo alla rivoluzione industriale nei documenti dell’Archivio storico Amma. Testi di Andrea Barghini, Pier Luigi Bassignana, Ferruccio Gambaruto, Vittorio Marchis. Torino, Umberto Allemandi & C. 1989, 109 pp. + 77 n.n., fig., L. 70.000. Nuncius, 5 (1):313–314, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00994.

Miniati:1990:BRO

[559] Mara Miniati. Book review: Orazio Curti, Visione e realtà. Le invenzioni di oggi sognate secoli fa. Milano, Rizzoli 1989, 143 pp., fig., L. 70.000. Nuncius, 5(1):314–315, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01001.

Miniati:1990:BRCb

[560] Mara Miniati. Book review: Comune di Campogalliano, Bilance Bascule Stadere. Modena, edizioni il Bulino 1987, 69 pp., fig. Nuncius, 5(1): 315, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01010.

Miniati:1990:BRAa

[561] Mara Miniati. Book review: Atti del Convegno di Studi ≪Architettura militare nell’Europa del XVI secolo≫ Firenze, 25–28 Novembre 1986, a cura di Carlo Cresti, Amelio Fara, Daniela Lamberini. Siena, Edizioni Periccioli 1988, 448 pp. Nuncius, 5(1):316–317, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01029.

Miniati:1990:BRAb

[562] Mara Miniati. Book review: Amelio Fara, Il sistema e la città. Architettura fortificata dell’Europa moderna dai trattati alle realizzazioni 1464– 1794. Genova, Sagep Editrice 1989, 267 pp., fig. Nuncius, 5(1):317–318, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01038.

Miniati:1990:BREa

[563] Mara Miniati. Book review: Elena Tamagno, Fornaci. Terre e pietre per l’ars aedificandi. Con un saggio introduttivo di Roberto Gabetti. Torino, Umberto Allemandi & C. 1987, 174 pp., fig., L. 85.000. Nuncius, 5(1):318, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x01047.

Malacrida:1990:BRH

[564] Lorenzo Marri Malacrida. Book review: Huldrych M. Koelbing, Storia della terapia medica. Momenti fondamentali. Milano, Edizioni Ciba-Geigy 1989, 221 pp. Nuncius, 5(1):318–319, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x01056.

Wanrooij:1990:BRM

[565] Bruno P. F. Wanrooij. Book review: Martin Weber, Georg Christian Gottlieb Wedekind 1761–1831. Werdegang und Schicksal eines Arztes im Zeitalter der Aufkärung und der Französischen Revolution (Mit einem Anhang: Wedekinds Diatetikvorlesung von 1789/1790), Stuttgart-New York, Gustav Fischer Verlag (Soemmerring-Forschungen. Beitrage zur Naturwissenschaft und Medizin der Neuzeit, Band Iv), 1988, 465 pp. Nuncius, 5(1):319–320, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01065.

Lippi:1990:BRM

[566] Donatella Lippi. Book review: Mirko D. Grmek, Aids. Storia di una epidemia attuale, Bari, Laterza 1989, 330 pp. (Titolo originale: Histoire du Sida, Edition Payot, Paris 1989). Nuncius, 5(1):321–322, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01074.

Abbri:1990:BRU

[567] Ferdinando Abbri. Book review: Udo Benzenhöfer, Johannes de Rupescissa Liber de consideratione quintae essentiae omnium rerumdeutsch. Studien zur Alchemia medica des 15. bis 17. Jahrhunderts mit kritischer Edition des Textes, Stuttgart, Franz Steiner Verlag Wiesbaden GmbH 1989, 213 pp. (≪ Heidelberger Studien zur Naturkunde der frühen Neuzeit, Herausgegeben von Wolf-Dieter Muller-Jahncke und Joachim Telle. Band 1 ≫ ). Nuncius, 5(1):322–323, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01083.

Beretta:1990:BRI

[568] Marco Beretta. Book review: Ingrid KÄStner, Paracelsus. Leipzig, Bsb B. G. Teubner 1985, 92 pp. Nuncius, 5(1):324, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x01092.

Beretta:1990:BRLa

[569] Marco Beretta. Book review: A. L. Lavoisier, Memorie scientifiche. Metodo e linguaggio nella nuova chimica (a cura di Ferdinando Abbri). Roma, Theoria 1986, 204 pp. Nuncius, 5(1):325, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01100.

Parigi:1990:BRD

[570] Silvia Parigi. Book review: David H. Hubel, Occhio, cervello e visione, Bologna, Zanichelli 1989 (≪ Nuovi classici della scienza ≫ ), 250 pp. Nuncius, 5(1):326, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01119.

Caroti:1990:BRA

[571] Stefano Caroti. Book review: Aufstieg und Niedergang der römischen Welt (ANRW). Geschichte und Kultur Roms im Spiegelder neueren Forschung. Teil II: Principat, Band 36: Philosophie, Wissenschaften, Technik. 1. Teilband: Philosophie (Historische Einleitung, Platonismus).

2. Teilband: Philosophie (Platonismus [Forts.]; Aristotelismus), Hrsg. von Wolfgang Haase, Berlin-New York, Walter de Gruyter 1987, 1322 pp. Nuncius, 5(1):327–330, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x01128.

Casati:1990:BRF

[572] Stefano Casati. Book review: Francescanesimo e cultura in Sicilia (Secc. XIII–XVI). Atti del convegno internazionale di studio nell’ottavo centenario della nascita di San Francesco d’Assisi. Palermo, 7–12 marzo 1982, ≪ Schede medievali ≫, n. 12–13, gennaio–dicembre 1987, 491 pp. Nuncius, 5(1):331–332, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01137.

Abbri:1990:BRJ

[573] Ferdinando Abbri. Book review: John Boswell, Cristianesimo, tolleranza, omosessualita. La Chiesa e gli omosessuali dalle origini al XIV secolo, (Italian) [Christianity, tolerance, homosexuality. The Church and homosexuals from its origin to the fourteenth century] Milano, Leonardo Editore 1989, 512 pp. Nuncius, 5(1):332–334, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01146.

Caroti:1990:BRPa

[574] Stefano Caroti. Book review: Patrick Gautier Dalché, La ≪ Description Mappe Mundi ≫ de Hugues de Saint Victor. Texte Inédit avec introduction et commentaire, Paris, Études Augustiniennes 1988, 228 pp., bibl., ind., ill. Nuncius, 5(1):334–335, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x01155.

Miniati:1990:BRMa

[575] Mara Miniati. Book review: Marja Mendera, La produzione di vetro nella Toscana bassomedievale. Lo scavo della vetreria di Germagnana in Valdelsa. Firenze, Edizioni all’insegna del Giglio 1989, 90 pp., fig., L. 25.000 (≪ Quaderni del Dipartimento di Archeologia e Storia delle Arti Sezioni archeologica–Università di Siena ≫, 15). Nuncius, 5(1): 336–337, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x01164.

Caroti:1990:BRD

[576] Stefano Caroti. Book review: Dennis E. Rhodes, Gli annali tipografici fiorentini del XV secolo, Prefazione di Roberto Ridolfi, Firenze, Leo S. Olschki 1988 (≪ Biblioteca di bibliografia italiana ≫, Cxiii), 149 pp., ill., ind. Nuncius, 5(1):337, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01173.

Plastina:1990:BRS

[577] Sandra Plastina. Book review: Silvana Seidel Menchi, Erasmo in Italia 1520–1580, Torino, Bollati Boringhieri 1987, 530 pp. Nuncius, 5(1):338– 341, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01182.

Plastina:1990:BRM

[578] Sandra Plastina. Book review: Michele Ciliberto, La ruota del tempo. Interpretazione di Giordano Bruno, Roma, Editori Riuniti 1986, 249 pp. Nuncius, 5(1):341–345, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01191.

Abbri:1990:BRW

[579] Ferdinando Abbri. Book review: William H. Huffman, Robert Fludd and the End of the Renaissance, London and New York, Routledge 1988, xii + 252 pp. Nuncius, 5(1):345–346, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x01209.

Abbri:1990:BRAb

[580] Ferdinando Abbri. Book review: Alan Bray, Homosexuality in Renaissance England, 2nd ed., London, the Gay Men’s Press 1988, 156 pp. Nuncius, 5(1):346–347, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01218.

Miniati:1990:BREb

[581] Mara Miniati. Book review: Enrico Gamba, Vico Montebelli (a cura di), Galileo Galilei e gli scienziati del Ducato di Urbino, Urbino, QuattroVenti 1989, 79 pp., fig. Nuncius, 5(1):348, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x01227.

Simonutti:1990:BRF

[582] Luisa Simonutti. Book review: François Laplanche, l’Écriture, le sacré et l’histoire. Érudits et politiques protestants devant la Bible en France au XVIIe siècle, Amsterdam & Maarssen, Apa-Holland University Press, Diffusion en France: Presses Universitaires de Lille 1986, xxxvi-1017 pp., 10 ill. (ISBN * 90 302 1012 5) (Sib 12), s.p.i. Nuncius, 5(1):348–350, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01236.

Abbri:1990:BRLb

[583] Ferdinando Abbri. Book review: The Leibniz Renaissance. International Workshop (Firenze, 2–5 giugno 1986), Firenze, Leo S. Olschki Editore 1989, 305 pp. (≪ Biblioteca di Storia della Scienza ≫, Vol. 28). Nuncius, 5(1):350–351, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x01245.

Beretta:1990:BRLb

[584] Marco Beretta. Book review: A. L. Lavoisier, De la richesse territoriale du Royaume de France, (Texte et documents présentés par J. C. Perrot), Paris, Éditions du Comite des Travaux historiques et scientifiques 1988, 269 pp. Nuncius, 5(1):351–353, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x01254.

Beretta:1990:BRRa

[585] Marco Beretta. Book review: Raffaele De Grada, Vita Firenza, Dario Generali, La cultura a Milano tra riformismo illuminato e rivoluzione, Milano, Vangelista 1989, 149 pp. Nuncius, 5(1):353–354, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01263.

Lippi:1990:BRG

[586] Donatella Lippi. Book review: G. Gaudenzi — R. Satolli, Jean-Paul Marat. Scienziato e rivoluzionario, Milano, Mursia 1989, 410 pp. Nuncius, 5(1):354–355, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01272.

Abbri:1990:BRDa

[587] Ferdinando Abbri. Book review: Dorinda Outram, the Body and the French Revolution. Sex, Class and Political Culture, New Haven-London, Yale University Press 1989, ix + 197 pp. Nuncius, 5(1):355– 357, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01281.

Arrigoni:1990:BRT

[588] Tiziano Arrigoni. Book review: Teresa Isenburg, Viaggiatori naturalisti italiani in Brasile nell’Ottocento, Milano, Franco Angeli 1989, 119 pp. (≪ Geografia umana. Collana diretta da Lucio Gambi ≫ ). Nuncius, 5 (1):357–359, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x01290.

Miniati:1990:BRFa

[589] Mara Miniati. Book review: Franz Pesendorfer, Leopoldo II di Lorena. La vita dell’ultimo Granduca di Toscana (1824–1859), Firenze, Sansoni Editore 1989, 446 pp., L. 75.000. Nuncius, 5(1):359, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x01308.

Miniati:1990:BRLb

[590] Mara Miniati. Book review: Leonardo Rombai, P. Giovanni Inghirami. Astronomo, geodeta e cartografo. ≪ l’illustrazione geografica della Toscana ≫, Firenze, Osservatorio Ximeniano 1989, 183 pp., fig. Nuncius, 5(1):359–360, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01317.

Beretta:1990:BRPa

[591] Marco Beretta. Book review: Pelageya Kochina, Love and mathematics: Sofya Kovalevskaya, Moscow, Mir Publishers Moscow 1985, 340 pp. Nuncius, 5(1):360, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01326.

Abbri:1990:BRDb

[592] Ferdinando Abbri. Book review: Donald E. Brown, Hierarchy, History, and Human Nature. The Social Origins of Historical Consciousness, Tucson, The University of Arizona Press 1988, x + 384 pp. Nuncius, 5(1): 360–362, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x01335.

Anonymous:1990:BRP

[593] Anonymous. Book review: Primo Boselli, Il museo della radio / the radio collector, Firenze, Edizioni Medicea 1989, 345 pp., fig., L. 36.000. Nuncius, 5(1):362–382, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x01344.

Baroncelli:1990:LMS

[594] Giovanna Baroncelli and Massimo Bucciantini. L’Istituto e Museo di Storia della Scienza di Firenze. Nuncius, 5(2):5–52, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00011.

Trifogli:1990:CDD

[595] Cecilia Trifogli. Continuità e discontinuità delle grandezze fisiche in Egidio Romano. (Italian) [Continuity and discontinuity of physical quantities in Giles of Rome]. Nuncius, 5(2):53–73, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00020.

Gozza:1990:ASS

[596] Paolo Gozza. Atomi, ’spiritus’, suoni: Le speculationi di musica (1670) del ’Galileiano’ Pietro Mengoli. Nuncius, 5(2):75–98, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00039.

Barsanti:1990:SND

[597] Giulio Barsanti. Storia naturale delle scimmie 1600–1800. Nuncius, 5 (2):99–165, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00048.

Blanckaert:1990:TNL

[598] Claude Blanckaert. La thologie naturelle de louis-françois jhan (1803– 1871). Nuncius, 5(2):167–204, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00057.

Iannaccone:1990:FSB

[599] Isaia Iannaccone. Il feng shuie la bussola geomantica dell’Istituto e Museo di Storia della Scienza di Firenze. Nuncius, 5(2):205–219, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00066.

Drake:1990:PAD

[600] Stillman Drake. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 5(2):221–228, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00075.

Caroti:1990:BRB

[601] Stefano Caroti. La Biblioteca di Romelio Da Brescia Magister Artium Et Medicine. Nuncius, 5(2):229–248, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00084.

Castellani:1990:UCA

[602] Carlo Castellani. Un’edizione computerizzata di alcuni manoscritti di Lazzaro Spallanzani. Nuncius, 5(2):249–260, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00093.

Anonymous:1990:S

[603] Anonymous. Schede. Nuncius, 5(2):261–265, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00101.

Parigi:1990:AOD

[604] Silvia Parigi. Alle origini dell’et moderna: Arti, magia, scienza. Nuncius, 5(2):267–276, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00110.

Brenni:1990:PSS

[605] Paolo Brenni. Il patrimonio storico-scientifico italiano. Nuncius, 5(2): 277–280, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00129.

Iannaccone:1990:BRP

[606] Isaia Iannaccone. Book review: Paul U. Unschuld, Introductory readings in classical Chinese medicine. Dordrecht, Kluwer Academic Publishers 1988, pp. 463, US$54.00, UK£47.00 . Nuncius, 5(2):281–283, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00138.

Touwaide:1990:BRI

[607] Alain Touwaide. Book review: Ivan Garofalo, Erasistrati fragmenta, Pisa, Giardini Editori e Stampatori 1988, ix + 215 pp. (Biblioteca di studi antichi, 62). Nuncius, 5(2):284–288, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00147.

Touwaide:1990:BRP

[608] Alain Touwaide. Book review: P. Potter, Short Handbook of Hippocratic Medicine, Québec, Les Éditions du Sphinx 1988, 59 pp. 4 planches. Nuncius, 5(2):288–290, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00156.

Vilaplana:1990:BRB

[609] Dulce Liars Vilaplana. Book review: Barnabas B. Hugues, Robert of Chester’s Latin Translation of al-khwrisms al Jabr. (Al-kitb al-mukhtasar f hisb al-jabr wa’l-muqbala / Liber algebre et al-muchabala), Wiesbaden Stuttgart, Franz Steiner Verlag 1989, 67 pp. + 6 pp. de gravures des trois manuscrits. Nuncius, 5(2):290–294, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00165.

Morpurgo:1990:BRM

[610] Piero Morpurgo. Book review: Michela Pereira, the alchemical corpus attributed to Raymond Lull, London, the Warburg Institute — University of London 1989, 118 pp. (Warburg Institute Surveys and Texts, XVIII). Nuncius, 5(2):294–297, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00174.

Potest:1990:BRF

[611] Paola Potest. Book review: Francesca Salvemini, La visione e il suo doppio. La prospettiva tra arte e scienza, (Italian) [Vision and its double. Perspective through art and life sciences], Bari, Laterza 1990, 182 pp., L. 24.000. Nuncius, 5(2):297–298, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00183.

Beretta:1990:BRWb

[612] Marco Beretta. Book review: Walter Pagel, Paracelso. Un’introduzione alla medicina filosofica nell’et del Rinascimento. Introduzione di Eugenio Garin, traduzione di Michele Sampaolo, Milano, Il Saggiatore 1989, xxii + 404 pp., ill. Nuncius, 5(2):298–301, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00192.

Finocchiaro:1990:BRR

[613] Maurice A. Finocchiaro. Book review: Richard S. Westfall, Essays on the Trial of Galileo, Vatican City, Vatican Observatory Publications (Special Series, Studi Galileiani, Vol. 1, No. 5); and Notre Dame (Indiana), University of Notre Dame Press 1989, 103 pp., paperback, $9.95. Nuncius, 5(2):301–307, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00200.

Clericuzio:1990:BRH

[614] Antonio Clericuzio. Book review: Harold J. Cook, the Decline of the Old Medical Regime in Stuart London, Ithaca and London, Cornell University Press 1986, 310 pp. Nuncius, 5(2):307–311, ????

1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00219.

Abbri:1990:BRM

[615] Ferdinando Abbri. Book review: Michael Hunter, Establishing the New Science. The Experience of the Early Royal Society, Woodbridge, the Boydell Press 1989, xiv + 382 pp. Nuncius, 5(2):311–313, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00228.

Meli:1990:BRM

[616] Domenico Bertoloni Meli. Book review: M. Hunter — S. Schaffer (eds.), Robert Hooke. New Studies, Woodbridge, the Boydell Press 1989, x + 310 pp., 39.50 pounds. Nuncius, 5(2):313–316, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00237.

Abbri:1990:BRRa

[617] Ferdinando Abbri. Book review: Richard S. Westfall, Newton, a cura di A. Serafini, Torino, Einaudi 1989, 2 voll., xx + 959 pp. (Biblioteca di cultura storica 179). Nuncius, 5(2):316–318, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00246.

Ordoez:1990:BRH

[618] Javier Ordoez and Alberto Elena. Book review: Horacio Capel, Joan Eugeni Snchez, Omar Moncada, De Palas a Minerva. La formacin científica y la estructura institucional de los ingenieros militares en el siglo XVIII, Barcelona-Madrid, Libros del Serbal-Consejo Superior de Investigaciones Científicas 1988, 390 pp. Nuncius, 5(2):318–320, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00255.

Bernardi:1990:BRR

[619] Walter Bernardi. Book review: Renato Pasta, Scienza politica e rivoluzione. l’opera di Giovanni Fabbroni (1752–1822) intellettuale e funzionario al servizio dei Lorena, Firenze, Olschki Editore 1989, xx + 639 pp. Nuncius, 5(2):320–323, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00264.

Beretta:1990:BRRb

[620] Marco Beretta. Book review: Rapports à l’Empereur sur les progrés des sciences, des lettres et des arts depuis 1789: vol. I, Jean-Baptiste Delambre, Sciences Mathématiques, présentation et notes de Jean Dhombers, 333 pp.; vol. II, Georges Cuvier, Chimie et Sciences de la Nature, présentation et note sous la direction de Yves Laissus, 333 pp., Paris, Éditions Belin 1989. Nuncius, 5(2):323–325, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00273.

Beretta:1990:BRG

[621] Marco Beretta. Book review: Giovanni Santinello (a cura di), Storia delle storie generali della filosofia, vol. 3, Il secondo Illuminismo e l’et Kantiana, Padova, Editrice Antenore 1988, 2 tomi, xxi + 1057 pp. Nuncius, 5(2):326–328, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00282.

Mazzolini:1990:BRM

[622] Renato G. Mazzolini. Book review: Manfred Wenzel (ed.), Goethe und Soemmerring. Briefwechsel 1784–1828. (Soemmerring-Forschungen, Bd. V), Stuttgart — New York, Gustav Fischer Verlag 1988, 179 pp., 23 ill., DM. 58,-. Nuncius, 5(2):329–330, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00291.

Beretta:1990:BRB

[623] Marco Beretta. Book review: Bettina Haupt, Deutschsprachige Chemielehrbcher (1775–1850), Stuttgart, Deutscher Apotheker Verlag 1987, xiii + 424 pp. Nuncius, 5(2):330–331, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00309.

LaVergata:1990:BRD

[624] Antonello La Vergata. Book review: David N. Livingstone, Darwin’s forgotten defenders. The encounter between evangelical theology and evolutionary thought, Grand Rapids, Michigan, William Eerdmans Publishing Company; Edinburgh, Scottish Academic Press 1987, xii + 210 pp., £7,95 (paperback). Nuncius, 5(2):332–336, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00318.

Redondi:1990:BRA

[625] Pietro Redondi. Book review: Anastasios Brenner, Duhem. Science, réalité et apparence, Préface de Maurice Boudot, Paris, J. Vrin 1990, 249 pp. Nuncius, 5(2):337–340, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00327.

Morabito:1990:BRV

[626] Carmela Morabito. Book review: Vittorio Somenzi, Tra fisica e filosofia, Abano Terme, Piovan Editore 1989, 338 pp. Nuncius, 5(2):340–343, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00336.

Vista:1990:BRR

[627] Michele Vista. Book review: Robert M. Hazen, Superconduttori. Storia di una scoperta, Milano, Edizioni di comunit 1989, 269 pp. Nuncius, 5 (2):343–345, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00345.

Abbri:1990:B

[628] Ferdinando Abbri. Biografie. Nuncius, 5(2):347–348, ???? 1990. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00354.

Abbri:1990:BRB

[629] Ferdinando Abbri. Book review: Brian McGuinness, Wittgenstein. Il giovane Ludwig (1889–1921), (Italian) [Wittgenstein. The young Ludwig (1889–1921)]. Milano, Il Saggiatore 1990, xiii + 473 pp. Nuncius, 5(2): 348–350, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00363.

Miniati:1990:S

[630] Mara Miniati. Strumentaria. Nuncius, 5(2):350, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00372.

Miniati:1990:BREc

[631] Mara Miniati. Book review: Enrico Miotto, Giuseppe Tagliaferri, Pasquale Tucci, La strumentazione nella storia dell’Osservatorio astronomico di Brera, (Italian) [Instrumentation in the history of the Astronomical Observatory of Brera], Milano, Edizioni Unicopli 1990, 121 pp., fig. Nuncius, 5(2):351, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00381.

Abbri:1990:IS

[632] Ferdinando Abbri. Istituzioni scientifiche. Nuncius, 5(2):352–353, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00390.

Miniati:1990:BRFb

[633] Mara Miniati. Book review: Fabrizio Bonoli (a cura di), Il Museo di Astronomia, s.n.t., 8 pp., fig. Nuncius, 5(2):353, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00408.

Miniati:1990:SDT

[634] Mara Miniati. Storia della tecnologia. Nuncius, 5(2):354, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00417.

Miniati:1990:BRMb

[635] Mara Miniati. Book review: Massimo Squillacciotti (a cura di), Lavoro, industria e cultura. Storia delle trasformazioni sul territorio aretino (Italian) [Labor, industry and culture. History of transformations in the territory of Arezzo], Perugia, Electa/Editori Umbri Associati 1990, 171 pp., fig. Nuncius, 5(2):354, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00426.

Beretta:1990:C

[636] Marco Beretta. Cartografia. Nuncius, 5(2):355–356, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539190x00435.

Lippi:1990:MB

[637] Donatella Lippi. Medicina e biologia. Nuncius, 5(2):356–358, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00444.

Lippi:1990:BRJ

[638] Donatella Lippi. Book review: J. J. Cocheton, J. Guerre, H. Pequignot, Storia illustrata della gastroenterologia, Editiemme 1989, 526 pp. (Edizione fuori commercio realizzata per la Glaxo Spa. Titolo originale Histoire illustré de l’hepato-gastroentrologie de l’Antiquitè à nos jours, Parigi, R. Dacosta 1987). Nuncius, 5(2):358–359, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00453.

Lippi:1990:BRH

[639] Donatella Lippi. Book review: Huldrych M. Koelbing, Storia della Terapia Medica. Momenti fondamentali, Ciba-Geigy Ed., 1989, 221 pp. (Tit. Orig. Die aerztliche Therapie: Grundzuge ihrer Geschichte, Darmstadt, Wiss. Buchges 1985). Nuncius, 5(2):360–361, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00462.

Lippi:1990:BRS

[640] Donatella Lippi. Book review: L. Sterpellone, La Medicina Etrusca. Demoiatria di un’antica civilt, Ciba-Geigy Ed. 1990, 198 pp. Nuncius, 5(2):361–363, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539190x00471.

Lippi:1990:BRL

[641] Donatella Lippi. Book review: Luigi L. Barbieri, Il Dottor Antonio Maria Valsalva e la sua Manovra, Bologna, Clueb 1990, 166 pp. (L. 25.000). Nuncius, 5(2):363–364, ???? 1990. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539190x00480.

Abbri:1990:BRRb

[642] Ferdinando Abbri. Book review: Roy Porter — Dorothy Porter, in Sickness and in Health. The British Experience 1650–1850, London, Fourth Estate 1988, 324 pp. Nuncius, 5(2):364–365, ????

1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00499.

Berni:1990:BRF

[643] Stefano Berni. Book review: Filippo M. Ferro (a cura di), Passioni della mente e della storia, (Italian) [Passions of the mind and history]. Milano, Vita e Pensiero 1989. Nuncius, 5(2):365–367, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00507.

Malacrida:1990:BRA

[644] Lorenzo Marri Malacrida. Book review: Arnaldo Cherubini, Medici scrittori d’Europa e d’America, Roma, Antonio Delfino Editore 1989, 344 pp., L. 90.000. Nuncius, 5(2):367–368, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00516.

Lippi:1990:BRT

[645] Donatella Lippi. Book review: Thomas A. Ban, Code-Dd. Valutazione e diagnosi dei disturbi depressivi, Padova, Liviana Editrice 1989, 153 pp. (L. 28.000) (Code-Dd. Composite Diognostic Evaluation of Depressive Disorders, 1929). Nuncius, 5(2):368–369, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00525.

Beretta:1990:AC

[646] Marco Beretta. Alchimia e chimica. Nuncius, 5(2):369–370, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00534.

Caroti:1990:MR

[647] Stefano Caroti. Medioevo e rinascimento. Nuncius, 5(2):370–372, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00543.

Caroti:1990:BRF

[648] Stefano Caroti. Book review: Francisco Marquez-Villanueva, Carlos Alberto Vega (eds.), Alfonso X of Castile the Learned King (1221–1284).

An International Symposium, Harvard University, 17 November 1984, Cambridge, Department of Harvard University 1990, 165 pp., ill., ind. (Harvard Studies in Romance Languages, 43). Nuncius, 5(2):372–373, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00552.

Caroti:1990:BRPb

[649] Stefano Caroti. Book review: Patrick Curry, Prophecy and Power. Astrology in Early Modern England, Cambridge, Polity Press 1989, 238 pp., bibl., ind. Nuncius, 5(2):373–375, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00561.

Abbri:1990:SXX

[650] Ferdinando Abbri. Secoli XVII–XX. Nuncius, 5(2):375–377, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00570.

Abbri:1990:BRRc

[651] Ferdinando Abbri. Book review: Richard Davenport-Hines, Sex, Death and Punishment. Attitudes to sex and sexuality in Britain since the Renaissance, London, Collins 1990, xv + 439 pp. Nuncius, 5(2):377–379, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00589.

Beretta:1990:BRPb

[652] Marco Beretta. Book review: Peter Jones (ed.), Philosophy and Science in the Scottish Enlightenment, Edinburgh, John Donald Publisher Ltd. 1988, vii + 230 pp. Nuncius, 5(2):379, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00598.

Arrigoni:1990:BRN

[653] Tiziano Arrigoni. Book review: Nicole Dhombres, Les savants en rvolution 1789–1799, Cité des Sciences et de l’Industrie, Paris, Calmann-Lvi 1989, 180 pp. Nuncius, 5(2):380, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00606.

LaVergata:1990:BRG

[654] Antonello La Vergata. Book review: Graeme Tytler, Physiognomy in the European novel. Faces and fortunes, Princeton, New Jersey, Princeton University Press 1982, ix + 436 pp. Nuncius, 5(2):381, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00615.

Miniati:1990:BRD

[655] Mara Miniati. Book review: Danilo Barsanti, P. Giovanni Giovannozzi. Uno scolopio tra fede e libertà, religione e patria (1860–1928), Firenze, Osservatorio Ximeniano 1990, 217 pp., fig. Nuncius, 5(2):382, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00624.

Miniati:1990:BRMc

[656] Mara Miniati. Book review: Marco Segala, La favola della terra mobile. La controversia sulla teoria della deriva dei continenti, (Italian) [The fable of mobile Earth. The controversy over the drift theory of the continents]. Bologna, Il Mulino 1990, 340 pp. Nuncius, 5(2):382–383, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539190x00633.

Miniati:1990:ADM

[657] Mara Miniati. Attivita dell’Istituto e Museo di Storia della Scienza di Firenze. Nuncius, 5(2):385–394, ???? 1990. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539190x00642.

Olmi:1991:MAV

[658] Giuseppe Olmi. Molti amici in vari luoghi: Studio della natura e rapporti epistolari nel secolo XVI. Nuncius, 6(1):3–31, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00010.

Elena:1991:BSC

[659] Alberto Elena. Baconianism in the Seventeenth-Century Netherlands: a preliminary survey. Nuncius, 6(1):33–47, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539191x00029.

Potest:1991:QCV

[660] Paola Potest. La questione del contagium vivum e la genesi del concetto di ambiente nella medicina tra XVII e XVIII secolo. Nuncius, 6 (1):49–67, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00038.

Mamiani:1991:NFD

[661] Maurizio Mamiani, Emanuela Trucco, and Emanuela Trucco. Newton e I fenomeni della vita. Nuncius, 6(1):69–96, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539191x00047.

Pogliano:1991:FIF

[662] Claudio Pogliano. La fisiologia in Italia fra ottocento e novecento. Nuncius, 6(1):97–121, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x00056.

Luigi:1991:PAD

[663] Pepe Luigi. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 6(1):123–134, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x00065.

Mantovani:1991:ICS

[664] Roberto Mantovani and Flavio Vetrano. Inventario del carteggio scientifico inviato allo scolopio Alessandro Serpieri. Nuncius, 6(1):135–160, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00074.

Anonymous:1991:LAS

[665] Anonymous. Lettere di Alessandro Serpieri a Filippo Cecchi. (Italian) [Letters from Alessandro Serpieri to Filippo Cecchi]. Nuncius, 6(1):161, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00083.

Anonymous:1991:IA

[666] Anonymous. Indice analitico. Nuncius, 6(1):162–166, ???? 1991. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539191x00092.

Wrakberg:1991:IF

[667] Urban Wrakberg. Istituzioni e fonti. Nuncius, 6(1):167–170, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00100.

Truci:1991:ADM

[668] Isabella Truci. Acquisto di due manoscritti magalotti di provenienza ginori-venturi. Nuncius, 6(1):171–174, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539191x00119.

Anonymous:1991:CBI

[669] Anonymous. Commission on bibliography of the International Union of the History and Philosophy of Science. Nuncius, 6(1):175, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00128.

Home:1991:RA

[670] R. W. Home. Report on Australia (1989). Nuncius, 6(1):176–190, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00137.

Beretta:1991:LR

[671] Marco Beretta. Lavoisier revisited. Nuncius, 6(1):191–203, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00146.

Abbri:1991:BRTa

[672] Ferdinando Abbri. Book review: Tore Frngsmyr (Ed.), Solomon’s House Revisited. The Organization and Institutionalization of Science, Canton Mass., Science History Publications 1990, xiv + 350 pp. (Nobel Symposium 75). Nuncius, 6(1):205–208, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00155.

Beretta:1991:BRM

[673] Marco Beretta. Book review: Matti Klinge (ed.), Helsingfors Universitet (1640–1990), 2 voll., Helsinki, Frlagsaktiebolaget Otava, 1988-89, 740 + 931 pp., ill. Nuncius, 6(1):209–211, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00164.

Ragep:1991:BRD

[674] F. Jamil Ragep. Book review: David A. King, Islamic Mathematical Astronomy, London, Variorum Reprints, 1986. Nuncius, 6(1):211–213, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00173.

Barsanti:1991:BRF

[675] Giulio Barsanti. Book review: François Ellenberger, Histoire de la géologie. Tome I: Des anciens à la première moitié du XVIIe siécle, (French) [History of Geology. Volume I. From the ancients to the first half of the 17th Century]. Paris, Technique et Documentation — Lavoisier, 1988, 352 pp. (Petite Collection d’Histoire des Sciences). Nuncius, 6 (1):213–214, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00182.

Beretta:1991:BRHa

[676] Marco Beretta. Book review: Heinz Goerke, Medizin und Technik. 3000 Jahre Ärztliche Hilfsmittel für Diagnostik und Therapie, (German) [Medicine and technology. 3000 years of medical aid for diagnosis and therapy]. München, Callwey, 1988, 284 pp., ill. Nuncius, 6(1):215–216, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00191.

Beretta:1991:BRD

[677] Marco Beretta. Book review: David Oldroyd, Storia della filosofia della scienza, trad. it. di Libero Sosio, Milano, Il Saggiatore, 1989, 540 pp. Nuncius, 6(1):216–218, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x00209.

Pereira:1991:BRP

[678] Michela Pereira. Book review: Piero Morpurgo, Filosofia della natura nella Schola Salernitana del secolo XII, Prefazione di Enrique Montero Cartelle, Bologna, Clueb, 1990, 245 pp., L. 35.000. Nuncius, 6(1):218– 221, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00218.

Abbri:1991:BRJ

[679] Ferdinando Abbri. Book review: John C. Briggs, Francis Bacon and the Rhetoric of Nature, Cambridge Mass., Harvard University Press, 1989, xvi + 285 pp. Nuncius, 6(1):221–223, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00227.

Parigi:1991:BRGa

[680] Silvia Parigi. Book review: Giuliana Mocchi, Idea, mente, specie. Platonismo e scienza in Johannes Marcus Marci (1595–1667), Soveria Mannelli (Cz), Rubbettino Editore, 1990, 314 pp. (Collana di storia delle idee). Nuncius, 6(1):223–226, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x00236.

Knobloch:1991:BRG

[681] Eberhard Knobloch. Book review: G. W. Leibniz, La naissance du calcul différentiel, 26 articles des Acta Eruditorum, introduction, traduction et notes par Marc Parmentier. Paris, 1989. 180 Ff. Nuncius, 6(1):226–228, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00245.

Farinella:1991:BRA

[682] Calogero Farinella. Book review: Aa.Vv., Cristina di Svezia. Scienza ed alchimia nella Roma barocca, Bari, Dedalo, 1990, 272 pp. Nuncius, 6 (1):228–233, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00254.

LaVergata:1991:BRJa

[683] Antonello La Vergata. Book review: John Christie, Sally Shuttleworth, (eds.), Nature transfigured, 1700–1900, Manchester and New York, Manchester University Press, 1989, ISBN O-7190-2567-2, 225 pp., £29.95 (hardback). Nuncius, 6(1):233–236, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00263.

Parigi:1991:BRM

[684] Silvia Parigi. Book review: Max Milner, La fantasmagoria. Saggio sull’ottica fantastica, introduzione di A. Pizzorusso, Bologna, Il Mulino, 1989, 263 pp. (Saggi 363). Nuncius, 6(1):236–239, ????

1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00272.

Abbri:1991:BRTb

[685] Ferdinando Abbri. Book review: Tore Frangsmyr, J. L. Heilbron, Robin E. Rider (eds.), The Quantifying Spirit in the 18th Century, Berkeley–Los Angeles, University of California Press, 1990, viii + 411 pp. (Uppsala studies in history of science 7). Nuncius, 6(1):239–242, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00281.

Beretta:1991:BRAa

[686] Marco Beretta. Book review: Aa.Vv., La expedicion botanica al verreinato del Peru (1777–1788), Barcellona-Madrid, Lunwerg Editores, 1988, 2 voll., 179 + 181 pp., ill. Nuncius, 6(1):243–246, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00290.

Segala:1991:BRS

[687] Marco Segala. Book review: Stefano Fabbri Bertoletti, Impulso formazione e organismo. Per una storia del concetto di Bildungstrieb nella cultura tedesca, Firenze, Olschki, 1990, 245 pp. Nuncius, 6(1):246–249, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00308.

Barsanti:1991:BRL

[688] Giulio Barsanti. Book review: Leonardo Rombai, P. Giovanni Inghirami. Astronomo, geodeta e cartografo. l’illustrazione geografica della Toscana, Firenze, Osservatorio Ximeniano, 1989, 185 pp.; Danilo Barsanti, P. Giovanni Antonelli. Scienza e fede nel Risorgimento, Firenze, Osservatorio Ximeniano, 1989, 127 pp.; Danilo Barsanti, P. Giovanni Giovannozzi. Uno scolopio tra fede e libertà, religione e patria (1860–1929), Firenze, Osservatorio Ximeniano, 1990, 218 pp. Nuncius, 6(1):249–251, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00317.

Segala:1991:BRC

[689] Marco Segala. Book review: Crosbie W. Smith — M. Norton Wise, Energy and Empire. A Biographical Study of Lord Kelvin, Cambridge, Cambridge University Press, 1989, xxvi + 866 pp. Nuncius, 6(1):251– 254, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00326.

Volkmar:1991:BRW

[690] Barbara Volkmar. Book review: Wolfgang U. Eckart, Deutsche Arzte in China 1897–1914. Medizin als Kulturmission im Zweiten Deutschen Kaiserreich, Stuttgart, New York, Gustav Fischer Verlag, 1989, 230 pp., ill., DM 58. (Medizin in Geschichte und Kultur, Bd. 16). Nuncius, 6 (1):255–257, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00335.

Segala:1991:BRT

[691] Marco Segala. Book review: Theodore Redpath, Ludwig Wittgenstein. A Student’s Memoir, London, Duckworth, 1990, 109 pp. Nuncius, 6 (1):257–260, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00344.

Segala:1991:BRH

[692] Marco Segala. Book review: Homer E. Le Grand, Drifting Continents and Shifting Theories. The modern revolution in geology and scientific change, Cambridge, Cambridge University Press, 1990 (I ed. 1988), vi + 313 pp. Nuncius, 6(1):260–264, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00353.

Segala:1991:BRJ

[693] Marco Segala. Book review: John A. Stewart, Drifting Continents and Colliding Paradigms. Perspectives on the Geoscience Revolution, Bloomington Indianapolis, Indiana University Press, 1990, xiii + 285 pp. Nuncius, 6(1):264–269, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x00362.

Stanzione:1991:BRR

[694] Massimo Stanzione. Book review: Roger Penrose, The Emperor’s New Mind. Concerning Computers, Minds, and the Laws of Physics, Oxford, Oxford University Press, 1989. Nuncius, 6(1):269–272, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00371.

Weindling:1991:BRB

[695] Paul Weindling. Book review: Burghard Weiss, Wie finde ich Literatur zur Geschichte der Naturwissenschaft und der Technik, Berlin, Berlin Verlag, Arno Spitz, 1990, 454 pp. Nuncius, 6(1):273, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00380.

Beretta:1991:BROa

[696] Marco Beretta. Book review: Ove Hagelin (ed.), Rare and Important Books in the Library of the Swedish Society of Medicine. A Descriptive and Annotated Catalogue, Stockholm, Svenska Läkaresällskapet, 1989, 176 pp., ill. Nuncius, 6(1):273–274, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00399.

Malacrida:1991:BRGa

[697] Lorenzo Marri Malacrida. Book review: Giuliano Cantoni, Alessandro Leoncini, Francesca Vannozzi, l’archivio dell’Università di Siena. Inventario della sezione storica, Siena, La Nuova Italia editrice, 1990, 311 pp., L. 20.000. Nuncius, 6(1):275, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00407.

Beretta:1991:BRL

[698] Marco Beretta. Book review: Lexicon der Renaissance, (Herausgegeben von Günter Gurst, Siegfried Hoyer, Ernst Ullmann und Christa Zimmermann), Leipzig, VEB Bibliographisches Institut, 1989, 798 pp., ill. Nuncius, 6(1):275–276, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x00416.

Beretta:1991:BROb

[699] Marco Beretta. Book review: Ove Hagelin (ed.), The Art of writing and drawing. A Selection of One Hundred Books from the Ekström Collection, Stockholm, the National Library for Psychology and Education, 1987, 232 pp., ill. Nuncius, 6(1):276–277, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00425.

Miniati:1991:BREa

[700] Mara Miniati. Book review: Emilio Borchi, Renzo Macii, Marina Parisi, Il barometro. Storia e applicazioni nell’Osservatorio Ximeniano.

Firenze, Osservatorio Ximeniano 1990, 140 pp., fig. Nuncius, 6(1):277, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00434.

Miniati:1991:BRJ

[701] Mara Miniati. Book review: Jean Matricon, Julien Roumette, l’invention du temps, Paris, Éditions de la Cité des Sciences et de l’Industrie, 1989, 125 pp., fig., F. F. 140. Nuncius, 6(1):278, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00443.

Miniati:1991:BRL

[702] Mara Miniati. Book review: L’Aventure du métre. Paris, C.N.A.M. — Musée National des Techniques, 1989, 103 pp., fig. Nuncius, 6(1):278– 279, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00452.

Beretta:1991:BRK

[703] Marco Beretta. Book review: Karin Johanivisson, A Life of Learning. Uppsala University during Five Centuries, Uppsala, Uppsala University Press, 1989, 133 pp., ill. Nuncius, 6(1):279–280, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00461.

Miniati:1991:BRAa

[704] Mara Miniati. Book review: Andrea Guintini, Dalla Lyonnaise alla Fiorentinagas 1839–1989. Bari, Editori Laterza, 1990, xii + 268 pp. Nuncius, 6(1):280, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00470.

Miniati:1991:BRCa

[705] Mara Miniati. Book review: Claudio Pogliano, Pharmakon. Storia dello Psicotropismo. Verona, Casamassima, 1990, 362 pp., fig. Nuncius, 6 (1):280–281, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00489.

Lippi:1991:BRAa

[706] Donatella Lippi. Book review: Antje Krug, Medicina nel mondo classico, Firenze, Giunti Ed., 1990, 270 pp., L. 35.000 (tr. it. di Heilkunst und Heilkult. Medizin in der Antike, München, O. Beck, 1985). Nuncius, 6 (1):281–282, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00498.

Lippi:1991:BRT

[707] Donatella Lippi. Book review: Thierry Bardinet, Dents et machoires dans les représentations religieuses et la pratique médicale de l’Egypte ancienne, (French) [Teeth and jaws in religious representations and medical practices of ancient Egypt], Roma, Ed. Pontificio Istituto Biblico, 1990, xxiii + 280 pp., L. 31.500. Nuncius, 6(1):282–283, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00506.

Malacrida:1991:BRGb

[708] Lorenzo Marri Malacrida. Book review: Giuseppe Penso, Parassiti, Microbi e Contagi nella Storia dell’Umanit. Ciba-Geigy Edizioni, 1990, 547 pp., fig. L. 90.000. Nuncius, 6(1):283–284, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00515.

Lippi:1991:BRAb

[709] Donatella Lippi. Book review: Aa.Vv., Santa Maria Nuova in Firenze. Memorie, testimonianze, prospettive, VII Centenario della fondazione dell’Ospedale. Atti delle giornate celebrative, Firenze, Il Fiorino, 1990, 300 pp. Nuncius, 6(1):284–285, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00524.

Malacrida:1991:BRF

[710] Lorenzo Marri Malacrida. Book review: Francesca Vannozzi (a cura di), La storia della psichiatria e degli ospedali psichiatrici in Toscana nell’Ottocento, Siena, Tipografia della Provincia, 1990, 243 pp., ed. fuori commercio. Nuncius, 6(1):285–286, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00533.

Malacrida:1991:BRL

[711] Lorenzo Marri Malacrida. Book review: Lanmarco Laquidara, Per una storia della sanit nell’Ottocento Preunitario Apuano. Massa Carrara, S.E.A., (per conto della Soc. Internazionale Dante Alighieri), 181 pp., fig., ed. fuori commercio. Nuncius, 6(1):286–287, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00542.

Malacrida:1991:BRAa

[712] Lorenzo Marri Malacrida. Book review: Arnaldo Cherubini, Francesca Vannozzi, Previdenza di malattia e malattie sociali dall’Unit alla prima guerra mondiale, Roma, Istituto Italiano di Medicina Sociale, 1990, 257 pp., ed. fuori commercio. Nuncius, 6(1):287–288, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00551.

Malacrida:1991:BRAb

[713] Lorenzo Marri Malacrida. Book review: Actes du 4e colloque des conservateurs des musées d’histoire des sciences médicales. 7 au 10 septembre 1988: Pavia–Milano, Lyon, Collection Fondation Marcel Merieux, 1990, 327 pp., fig. Nuncius, 6(1):288–290, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00560.

Abbri:1991:BRB

[714] Ferdinando Abbri. Book review: Bulletin for the History of Chemistry, V, 1989. Ambix, XXXVI, 1989, n. 1. Nuncius, 6(1):290–291, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00579.

Caroti:1991:BRAa

[715] Stefano Caroti. Book review: Alessandra Tarabochia Canavero, (a cura di), l’infinita via. Ragione, natura e trinit da Anselmo a Tommaso, Bergamo, Pierluigi Lubrina, 1990, 272 pp., bibl., ind. (Quodlibet. Ricerche e strumenti di filosofia medievale, 7). Nuncius, 6(1):291–292, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00588.

Caroti:1991:BRLa

[716] Stefano Caroti. Book review: Loris Sturlese, Storia della filosofia tedesca nel Medioevo. Dagli inizi alla fine del XII secolo, Firenze, Leo S. Olschki, 1990, 241 pp., bibl., ind. (Accademia Toscana di Scienze e Lettere ’La Colombaria’, Studi, Cv). Nuncius, 6(1):292–293, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00597.

Caroti:1991:BRJ

[717] Stefano Caroti. Book review: Johannes Sharpe, Quaestio super universalia, a cura di Alessandro D. Conti, Firenze, Leo S. Olschki, 1990, 375 pp., bibl., ind. (Unione Accademica Nazionale. Corpus Philosophorum Medii Aevi. Testi e studi IX). Nuncius, 6(1):293–294, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00605.

Caroti:1991:BRLb

[718] Stefano Caroti. Book review: Luca Bianchi, Eugenio Randi, Le verit dissonanti. Aristotele alla fine del Medioevo, prefazione di Mariateresa Fumagalli Beonio Brocchieri, Bari, Laterza, 1990, 194 pp., ind. (Biblioteca di Cultura Moderna, 991). Nuncius, 6(1):294–295, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00614.

Caroti:1991:BRLc

[719] Stefano Caroti. Book review: Luca Bianchi, Il vescovo e i filosofi. La condanna parigina del 1277 e l’evoluzione dell’aristotelismo scolastico, (Italian) [The bishop and the philosophers. The condemnation of Paris of 1277 and the scholastic evolution of Aristotelianism]. Bergamo, Pierluigi Lubrina, 1990, 280 pp., bibl., ind. (Quodlibet. Ricerche e strumenti di filosofia medievale, 6). Nuncius, 6(1):295–296, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00623.

Caroti:1991:BRB

[720] Stefano Caroti. Book review: Barbara Faes De Mottoni, (a cura di), Aegidii Romani Opera Omnia. I. Catalogodei manoscritti (457-505). 1/5*. Repubblica Federale di Germania (Monaco), Firenze, Leo S. Olschki, 1990, 350 pp., ind., app. (Unione Accademica Nazionale. Corpus Philosophorum Medii Aevi, Testi e Studi, X). Nuncius, 6(1):296–298, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00632.

Abbri:1991:BRAa

[721] Ferdinando Abbri. Book review: Atti del Convegno Internazionale di Studi su Bernardino Telesio, Cosenza, Accademia Consentina, 1990, 221 pp. Nuncius, 6(1):298–299, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00641.

Abbri:1991:BRW

[722] Ferdinando Abbri. Book review: William R. Shea, Antonio Spadafora (editors), Creativity in the Arts and Science, Canton Ma, Science History Publications, 1990, xviii + 187 pp. Nuncius, 6(1):299–301, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00650.

Segala:1991:BRMa

[723] Marco Segala. Book review: Michael J. Crowe, the Extraterrestrial Life Debate 1750–1900. The Idea of Plurality of Worlds from Kant to Lowell, Cambridge, Cambridge University Press, 1988, xxii + 680 pp. Nuncius, 6(1):301, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00669.

Abbri:1991:BRPa

[724] Ferdinando Abbri. Book review: Peter Hulme, Ludmilla Jordanova (Eds.), the Enlightenment and its Shadows, London-New York, Routledge, 1990, 232 pp. Nuncius, 6(1):302–303, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00678.

Beretta:1991:BRHb

[725] Marco Beretta. Book review: Horst Bergman und Eberhard Zwink, Emanuel Swedenborg — 1688–1772. Naturforscher und Kundiger der Überwelt, Stuttgart, Württembergische Landesbibliothek, 1988, 168 pp., ill. Nuncius, 6(1):303–304, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00687.

Arrigoni:1991:BRV

[726] Tiziano Arrigoni. Book review: La via Vandelli strada Ducale del Settecento da Modena a Massa — Dal Frignano alla Garfagnana e al Ducato di Massa, a cura di M. Pellegrini e F. M. Pozzi, Modena, Artioli Editore 1989, pp. 157. Nuncius, 6(1):304–305, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00696.

Miniati:1991:BRSa

[727] Mara Miniati. Book review: Sandro Cardinali, Luigi Pepe (a cura di), Alfonso Bonfioli Malvezzi. Viaggio in Europa e altri scritti, Ferrara, Università degli Studi di Ferrara, 1988, lv + 321 pp. (Annali dell’Università di Ferrara (Nuova Serie) Sezione III — Filosofia, vol. II, n. 1). Nuncius, 6(1):305, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00704.

Abbri:1991:BRC

[728] Ferdinando Abbri. Book review: Changes d’influences scientifiques et techniques entre pays Europëens de 1780 à 1830, Paris, Éditions du CTHS, 1990, 171 pp. (Colloques du C.T.H.S., V). Nuncius, 6(1):305– 307, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00713.

Lippi:1991:BRP

[729] Donatella Lippi. Book review: A. Pansini, Van Gogh. Il suo genio e la sua malattia, Firenze, Stampa Nazionale, 1990, 111 pp. Nuncius, 6 (1):307–308, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00722.

Miniati:1991:BRB

[730] Mara Miniati. Book review: Bruno P. F. Wanrooij, Storia del pudore. La questione sessuale in Italia 1860–1940. Venezia, Marsilio Editori, 1990, 234 pp. Nuncius, 6(1):308–309, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00731.

Miniati:1991:ADMa 7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00740.

Gomez:1991:BSN

[732] Susana Gomez. The Bologna Stone and the nature of light. Nuncius, 6(2):v–32, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00759.

Palladino:1991:STD

[733] Franco Palladino. Sulla teoria delle proporzioni nel seicento. Nuncius, 6(2):33–81, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00768.

Taiani:1991:LSA

[734] Rodolfo Taiani. L’acqua e la sua anima: Il contributo della scienza chimica allo sfruttamento delle fonti di acqua minerale nella prima metà del XIX secolo. (Italian) [Water and its soul: The contribution of chemistical science to the exploitation of mineral water sources in the first half of the Nineteenth Century]. Nuncius, 6(2):83–107, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00777.

Saliba:1991:SCD

[735] George Saliba. A Sixteenth-Century drawing of an astrolabe made by Khafīf Ghulām ’Alī B. ’īsā (c. 850 A.D.). Nuncius, 6(2):109–119, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00786.

Agnoletti:1991:DAF

[736] Mauro Agnoletti. Documenti di archeologia forestale: Le segherie idrauliche della Valle del Piave. Nuncius, 6(2):121–143, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00795.

Barbieri:1991:PAD

[737] Francesco Barbieri and Marina Zuccoli. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 6(2):145–155, ????

1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00803.

Biagioli:1991:IF

[738] Mario Biagioli. Istituzioni e fonti. Nuncius, 6(2):157–169, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00812.

Vaccari:1991:SDT

[739] Ezio Vaccari. Storia della terra e tempi geologici in uno scritto inedito di Giovanni Arduino: la ≪Risposta Allegorico–Romanzesca≫ a Ferber. Nuncius, 6(2):171–188, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x00821.

Anonymous:1991:RAR

[740] Anonymous. Risposta allegorico romanzesca di Voniangi Riduano, Osservatore Longobardo, al celebre orittologo viaggiatore Signor Giovanni — Giacomo Ferber del Collegio Metallico di Svezia, sopra la genesi della presente faccia della terra. Nuncius, 6(2):189–212, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00830.

Cecconi:1991:CDS

[741] Aldo Cecconi. Il Congresso degli Scienziati Italiani a Firenze (1841). (Italian) [The Congress of Italian Scientists in Florence (1841)]. Nuncius, 6(2):213–278, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00849.

Touwaide:1991:DC

[742] Alain Touwaide. Discussioni critiche. Nuncius, 6(2):279–293, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00858.

Tanio:1991:GHS

[743] Nadine Tanio. Gendering the history of science. Nuncius, 6(2):295–305, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00867.

Segala:1991:CLP

[744] Marco Segala. Catastrofismo e localismo: a proposito della controversia sulla teoria della deriva dei continenti. Nuncius, 6(2):307–323, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00876.

Touwaide:1991:BRMa

[745] Alain Touwaide. Book review: Maria Michela Sassi, La scienza dell’uomo nella Grecia antica (≪Nuova Cultura≫, 6), Torino, Bollati Boringhieri, 1988, 250 pp., ill. Nuncius, 6(2):325–330, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00885.

Touwaide:1991:BRC

[746] Alain Touwaide. Book review: Carmélia Opsomer, Index de la pharmacopée du Ier au Xe siècle, Hildesheim, Zürich et New York (Olms-Weidmann), 1989, 2 voll., lxx-viii + 824 pp. (≪Alpha-Omega, Reihe A: Lexika-Indizes-Konkordanzes zur klassischen Philologie≫, CX). Nuncius, 6(2):330–333, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x00894.

Touwaide:1991:BRMb

[747] Alain Touwaide. Book review: Medicinae Pliniiconcordantiae curantibus A. R. Corsini et M. P. Segoloni Hildesheim, Zürich et New York, Olms-Weidmann, 1989, viii + 351 pp. (≪Alpha-Omega, Reihe A: Lexika-Indizes-Konkordanzen zur klassischen Philologie≫, Ci). Nuncius, 6(2): 334–336, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00902.

Iannaccone:1991:BRJ

[748] Isaia Iannaccone. Book review: Joseph Needham, Dialogue des civilisations Chine-Occident. Pour une histoire oecumenique des sciences (a cura di G. Métailié), Éditions La Découverte, Paris, 1991, ii + 367 pp., 165 Ff. Nuncius, 6(2):337–340, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00911.

Pereira:1991:BRC

[749] Michela Pereira. Book review: Constantine of Pisa, the book of the secrets of alchemy, Introduction, critical edition, translation and commentary by Barbara Obrist, Leiden-New York–Köln, E. J. Brill, 1990, 339 pp., con indicie bibliografia. (≪Collection des Travaux de l’Academie Internationale d’Histoire des Sciences≫, t. 34). Nuncius, 6(2):340–342, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00920.

Shea:1991:BRE

[750] William R. Shea. Book review: Eugenio Garin, Umanisti artisti scienziati. Studi sul Rinascimento italiano. Roma, Editori Riuniti, 1989, 333 pp. Nuncius, 6(2):342–345, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x00939.

Shank:1991:BRA

[751] Michael H. Shank. Book review: Armin Gerl, Trigonometrischastronomisches Rechnen kurz vor Copernicus: Der Briefwechsel Regiomontanus-Bianchini, Boethius, Band XXI, Stuttgart, Franz Steiner Verlag, 1989. Nuncius, 6(2):345–348, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539191x00948.

Ernst:1991:BRG

[752] Germana Ernst. Book review: I Guicciardinie le scienze occulte. L’oroscopo di Francesco Guicciardini. Lettere di alchimia, astrologia e cabala a Luigi Guicciardini, a cura di Raffaella Castagnola, premessa di Eugenio Garin, Firenze, Leo S. Olschki editore, 1990, vii-397 pp. (Istituto Nazionale di studi sul Rinascimento, Studi e Testi, XIX). Nuncius, 6(2):348–351, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00957.

Pereira:1991:BRA

[753] Michela Pereira. Book review: Alchemy revisited. Proceedings of the International Conference on the History of Alchemy at The University of Groningen, 17–19 April 1989, edited by Z. R. W. M. von Martels, Leiden-New York–Københaven–Köln, E. J. Brill, 1990, 284 pp., con bibliografia e indice. (≪Collection des Travaux de l’A-cadémie Internationale d’Histoire des Sciences≫, t. 33). Nuncius, 6(2):351–353, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00966.

Gohau:1991:BRG

[754] Gabriel Gohau. Book review: Georg Agricola, Bermannus (le mineur) — un dialogue sur les mines, Introduction, texte établi et commenté par Robert Halleux et Albert Yans, Paris, Les Belles Lettres, 1990, 185 pp. + 2 cartes. (Coll. ≪Science et humanisme≫ ). Nuncius, 6(2):353–355, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x00975.

Parigi:1991:BRJ

[755] Silvia Parigi. Book review: John Cottingham, Cartesio, Bologna, Il Mulino, 1991, 217 pp. (≪Universale Paperbacks≫, 253). Nuncius, 6 (2):355–357, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x00984.

Parigi:1991:BRGb

[756] Silvia Parigi. Book review: Gilles Deleuze, La piega. Leibniz e il Barocco, Torino, Einaudi, 1990, 206 pp. (≪Biblioteca di cultura filosofica≫ 52). Nuncius, 6(2):357–359, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x00993.

Parigi:1991:BRGc

[757] Silvia Parigi. Book review: George Berkeley, De motu, a cura di Mariapaola Fimiani, Salerno, Edizioni 10/17, 1989, 211 pp. Nuncius, 6(2): 359–361, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01000.

LaVergata:1991:BRE

[758] Antonello La Vergata. Book review: Ernst Mayr, Storia del pensiero biologico. Diversità, evoluzione, eredità, a cura di P. Corsi, Torino, Bollati Boringhieri, 1990, xv + 932 pp. a. Nuncius, 6(2):361–364, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01019.

Abbri:1991:BRG

[759] Ferdinando Abbri. Book review: G. S. Rousseau, Enlightenment Borders. Pre- and post-modern discourses. Medical, scientific, Manchester and New York, Manchester University Press, 1991, xv + 359 pp. Nuncius, 6(2):364–369, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01028.

Giuntini:1991:BRJ

[760] Chiara Giuntini. Book review: John Gascoigne, Cambridge in the Age of the Enlightment. Science, religion and politics from the Restoration to the French Revolution, Cambridge University Press, 1989, xi + 358 pp. Nuncius, 6(2):369–373, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x01037.

Miniati:1991:BRCb

[761] Mara Miniati. Book review: T. N. Clarke, A. D. Morrison-Low, A. D. C. Simpson, Brass & Glass. Scientific Instrument Making Workshops in Scotland as illustrated by instruments from the Arthur Frank Collection at the Royal Museum of Scotland.s.l., National Museums of Scotland, 1989, 320 pp., ill. Nuncius, 6(2):373–374, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01046.

Miniati:1991:BRAb

[762] Mara Miniati. Book review: Anthony J. Turner, From pleasure and profit to science and security: Étienne Lenoir and the transformation of precision instrument-making in France, 1760–1830. Cambridge, the Whipple Museum of the History of Science, 1989, 107 pp., fig. Nuncius, 6(2):374–376, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01055.

LaVergata:1991:BRJb

[763] Antonello La Vergata. Book review: A. J. Cain, Constantine Samuel Rafinesque Schmalz on classification. A translation of the early works by Rafinesque with introduction and notes, Philadelphia, Pennsylvania, the Academy of Natural Sciences of Philadelphia, 1990, 240 pp., ≪Tryonia≫, n. 20. Nuncius, 6(2):376–378, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01064.

Abbri:1991:BRSa

[764] Ferdinando Abbri. Book review: Stanislao Cannizzaro, Sunto di un corso di filosofia Chimica. Commento e nota storica di Luigi Cerruti. Introduzione di Leonello Paoloni, Palermo, Sellerio Editore, 1991, 286 pp. (≪Biblioteca siciliana di storia e letteratura≫ 27). Nuncius, 6(2):378– 379, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01073.

Scarpelli:1991:BRG

[765] Giacomo Scarpelli. Book review: Germana Pareti, La tentazione dell’occulto. Scienza ed esoterismo nell’età vittoriana, Torino, Bollati Boringhieri, 1990, 297 pp. Nuncius, 6(2):380–381, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539191x01082.

Scarpelli:1991:BRR

[766] Giacomo Scarpelli. Book review: Robert Ackerman, J. G. Frazer. His Life and Work, Cambridge, Cambridge University Press, 1987, Ill., x + 348 pp. Nuncius, 6(2):381–383, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x01091.

Giuntini:1991:BRL

[767] Chiara Giuntini. Book review: Lectures on the History of Psychiatry (The Squibb Series), ed. by R. M. Murray, T. H. Turner, London, Gaskell

— the Royal College of Psychiatrists, 1990, xi + 223 pp. Nuncius, 6 (2):383–387, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01109.

Wanrooij:1991:BRP

[768] Bruno P. F. Wanrooij. Book review: Paul Weindling, Health, Race and German Politics between National Unification and Nazism 1870–1945, Cambridge, Cambridge University Press, 1989, 641 pp. Nuncius, 6(2): 387–389, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01118.

Giuntini:1991:BRK

[769] Chiara Giuntini. Book review: Kurt Danziger, Constructing the subject. Historical origins of psychological research, Cambridge, Cambridge University Press, 1990, 249 pp. Nuncius, 6(2):390–393, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01127.

LaVergata:1991:BRA

[770] Antonello La Vergata. Book review: Alexander Vucinich, Darwin in Russian thought, Berkeley-Los Angeles-London, University of California Press, 1988, x + 468 pp. (ISBN 0-520-06283-3). Nuncius, 6(2):393–396, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01136.

Parigi:1991:BRA

[771] Silvia Parigi. Book review: Aleksandr R. Lurija, un mondo perduto e ritrovato, prefazione di O. Sacks, postfazione di L. Mecacci, Roma, Editori Riuniti, 19912, (19731), xvii-186 pp. Nuncius, 6(2):396–399, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01145.

Alessandro:1991:BRJ

[772] Bruno Alessandro. Book review: J. Roche (ed.), Physicists look back. Studies in the history of physics, Bristol-New York, Adam Hilger, 1990, 393 pp. Nuncius, 6(2):399–403, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x01154.

Abbri:1991:SDS

[773] Ferdinando Abbri. Storia della scienza. Nuncius, 6(2):405–407, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01163.

Miniati:1991:BRSb

[774] Mara Miniati. Book review: ≪Impact science et société≫, nn. 159 e 160, 1990, pp. 213-319; 323-436. Nuncius, 6(2):407–408, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01172.

Miniati:1991:S

[775] Mara Miniati. Strumentaria. Nuncius, 6(2):408–409, ???? 1991. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539191x01181.

Miniati:1991:BREb

[776] Mara Miniati. Book review: Emilio Borchi, Renzo Macii, Alcuni strumenti del Liceo Machiavelli di Lucca, Lucca, Maria Pacini Fazzi editore, 1991, 61 pp., fig. Nuncius, 6(2):409–410, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01190.

Abbri:1991:BRM

[777] Ferdinando Abbri. Book review: Mara Miniati, Strumenti Scientifici della Collezione Carrand, Firenze, Museo Nazionale del Bargello, 1991, 70 pp., ill. (≪Lo specchio del Bargello≫ n. 31). Nuncius, 6(2):410–411, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01208.

Miniati:1991:BRSc

[778] Mara Miniati. Book review: Strumenti ritrovati. Materiali della ricerca scientifica in Piemonte tra Settecento e Ottocento. (Italian) [Tools discovered. Materials of scientific research in the Piedmont region between the Eighteenth and Nineteenth Century]. Torino, Regione Piemonte — Assessorato ai Beni Culturali e Ambientali, 1991, 77 pp., fig. Nuncius, 6(2): 411, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01217.

Miniati:1991:BRI

[779] Mara Miniati. Book review: Istituto Nazionale di Geofisica, Tromometri, avvisatori, sismografi. Osservazioni e teorie dal 1850 al 1880. A cura di Graziano Ferrari. Bologna, ed. SGA Storia-Geofisica-Ambiente, 1991, 139 pp., fig. Nuncius, 6(2):411–412, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x01226.

Miniati:1991:BREc

[780] Mara Miniati. Book review: Elisabetta Calanchi, La collezione degli strumenti del Museo di Antropologia. Bologna, Editrice Clueb, 1991, 127 pp., fig. Nuncius, 6(2):412, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x01235.

Abbri:1991:SDT

[781] Ferdinando Abbri. Storia della tecnica. Nuncius, 6(2):413–415, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01244.

Miniati:1991:BRP

[782] Mara Miniati. Book review: Progetto Techne, Culture e strumenti del lavoro. La ruota idraulica. Brescia, Editrice la Scuola, 1989, 5 fasc., fig. Nuncius, 6(2):415, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x01253.

Casati:1991:SD

[783] Stefano Casati. Storia dell’architettura. Nuncius, 6(2):416, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01262.

Abbri:1991:MB

[784] Ferdinando Abbri. Medicina e biologia. Nuncius, 6(2):416–417, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01271.

Lippi:1991:BRM

[785] Donatella Lippi. Book review: M. Pontecorvo, Storia delle vaccinazioni. Dalle origini ai giorni nostri, Saronno, Ciba-Geigy Ed., 1991, 151 pp. Nuncius, 6(2):417–418, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x01280.

Lippi:1991:BRGa

[786] Donatella Lippi. Book review: G. Bonadonna, Donne in Medicina, (Italian) [Women in Medicine], Milano, Rizzoli, 1991, 323 pp. Nuncius, 6 (2):418–419, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01299.

LaVergata:1991:BRG

[787] Antonello La Vergata. Book review: Gerhard H. Müller, Johann Adam Pollich (1741–1780), ‘Medicinae Doctor’ und Pflanzengeograph der Pfalz. Eine Studie anhand bisher unbekannter Briefe. Dokumente und Schriften, (German) [Johann Adam Pollich (1741–1780), ‘Medicinae Doctor’ and plant geographer of the Palatinate. A study based on previously unknown letters], Bad Dürkheim, Pfalzmuseum für Naturkunde, Selbstverlag der ≪Pollichia≫, 1990, 176 pp. (≪Pollichia≫-Buch n. 21). Nuncius, 6(2):419–420, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x01307.

Lippi:1991:BRV

[788] Donatella Lippi. Book review: Vincenzio Chiarugi, Della pazzia in genere, e in specie, (Italian) [On insanity in general, and in species], ristampa anastatica dell’ed. Carlieri, 1793–94. Introduzione di F. M. Ferro e G. Riefolo. Firenze, Vecchiarelli Editore, 1991. Nuncius, 6(2):420–421, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01316.

Lippi:1991:BRGb

[789] Donatella Lippi. Book review: G. B. Giordano, Storia dell’Ospedale di Fregionata nel secolo XIX ed inventario del suo archivio storico dal 1813 al 1942, Roma, A. Delfino Editore, 1991, 251 pp. Nuncius, 6(2): 421–422, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01325.

Parigi:1991:AO

[790] Silvia Parigi. Astronomia e ottica. Nuncius, 6(2):422–423, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01334.

Parigi:1991:BRI

[791] Silvia Parigi. Book review: Ibn Al-Haytham, Optics, Translated with Introduction and Commentary by A. I. Sabra, London, the Warburg Institute-University of London, 1989, 2 voll., pp. ix-367 e ix-246. Nuncius, 6(2):423–424, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01343.

Casati:1991:BRA

[792] Stefano Casati. Book review: Alessandro Braccesi, Esplorando l’universo. Dalle conquiste degli antichi greci all’astronomia dei nostri giorni. Bologna, Zanichelli, 1988, pp. 292, (Le Ellissi). Nuncius, 6(2): 425, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01352.

Abbri:1991:AC

[793] Ferdinando Abbri. Alchimia e chimica. Nuncius, 6(2):425–427, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01361.

Abbri:1991:BRAb

[794] Ferdinando Abbri. Book review: Antonio di Meo (a cura di), Storia della Chimica. Dalla Ceramica del Neolitico all’età della Plastica, Venezia, Marsilio Editore, 1990, viii + 335 pp., ill. Nuncius, 6(2):427– 428, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01370.

Abbri:1991:BRSb

[795] Ferdinando Abbri. Book review: Suzanne Colnort-Bodet, Le Code Alchimique Dévoilé. Distillateurs. Alchymistes et Symbolistes, Paris, Librarie Honoré Champion, Editeur, 1989, xv + 388 pp. Nuncius, 6(2): 428–429, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01389.

Beretta:1991:BRAb

[796] Marco Beretta. Book review: Les atomes. Une anthologie historique, Textes choisis, présntés et annotés par Bernadette Bensaude-Vincent et Catherine Kounelis, Paris, Press Pocket, 1991, 334 pp. Nuncius, 6(2): 430, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01398.

Caroti:1991:AMR

[797] Stefano Caroti. Antichità, medioevo e rinascimento. Nuncius, 6(2):430– 431, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01406.

Caroti:1991:BRAb

[798] Stefano Caroti. Book review: Aufstieg und Niedergang der römischen Welt (ANR W) Geschichte und Kultur Roms im Spiegelder neueren Forschung. Teil II: Principat. Bd. 36: Philosophie, Wissenschaften, Technik. 3. Teilband: Philosophie (Stoizismus) e 4. Teilband: Philosophie (Epikureismus, Skeptizismus, Kynismus, Orphica; Doxographica), hrsg. W. Haase, Walter de Gruyter, Berlin-New York, 1989 e 1990. Nuncius, 6(2):431–434, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x01415.

Pereira:1991:BRMa

[799] Michela Pereira. Book review: Marsilio Ficino, Lettere — (I) Epistolarum familiarum liber I, a cura di Sebastiano Gentile, Firenze, Leo. S. Olschki Editore, 1990, ccxcix + 322. (Istituto Nazionale di Studi sul Rinascimento, Carteggi umanistici). Nuncius, 6(2):434–435, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01424.

Pereira:1991:BRMb

[800] Michela Pereira. Book review: M. Markowski, Astronomica et astrologica cracoviensia ante annum 1550, Leo S. Olschki editore, Firenze 1990, xxxv + 377 pp. ISBN 88-222-3743-9; ISSN 0394-4409 (Istituto Nazionale di Studi sul Rinascimento, Studi e Testi, XX). Nuncius, 6(2):435–436, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01433.

Abbri:1991:BRD

[801] Ferdinando Abbri. Book review: David C. Lindberg, Robert S. Westman, Reappraisals of the Scientific Revolution, Cambridge, Cambridge University Press, 1990, xx-viii + 551 pp. Nuncius, 6(2):436–438, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01442.

Abbri:1991:BRAc

[802] Ferdinando Abbri. Book review: Antoni Van Leeuwenhoek, Alle de Brieven. Uitgegeven, geïllustreerd en van aantekeningen voorzien door een Commissie van Nederlandse Geleerden (The Collected Letters. Edited, illustrated and annotated by a Committee of Ducth scientists), Deel (Part) XII, Lisse, Swets & Zeitlinger B.V., 1989, 441 pp. Nuncius, 6(2):438–439, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x01451.

Segala:1991:BRMb

[803] Marco Segala. Book review: M. Feingold (ed.), Before Newton. The life and times of Isaac Barrow, Cambridge, Cambridge University Press, 1990, xii + 380 pp. (ISBN 0-521-30694-9). Nuncius, 6(2):439–440, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01460.

Abbri:1991:BRZ

[804] Ferdinando Abbri. Book review: Z. Ciuffoletti, L. Rombai (a cura di), La Toscana dei Lorena. Riforme, Territorio, Società, Firenze, Leo S. Olschki Editore, 1989, 670 pp. (≪Biblioteca Storica Toscana≫ 19). Nuncius, 6 (2):440–442, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01479.

Abbri:1991:BRPb

[805] Ferdinando Abbri. Book review: Peter Dear (editor), the Literary Structure of Scientific Argument. Historical Studies, Philadelphia, University of Pennsylvania Press, 1991, vii + 211 pp. Nuncius, 6(2):442–444, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01488.

Segala:1991:BRF

[806] Marco Segala. Book review: F. W. J. Shelling, Ideas for a philosophy of nature, Cambridge, Cambridge University, Press, 1988, xxvi + 294 pp. (≪Texts in German Philosophy≫ ) (ISBN 0-521-35733-0). Nuncius, 6(2):444, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01497.

Arrigoni:1991:BRA

[807] Tiziano Arrigoni. Book review: Alain Corbin, l’invenzione del mare, Venezia, Marsilio Editori, 1990, 365 pp. Nuncius, 6(2):444–446, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01505.

Segala:1991:BRD

[808] Marco Segala. Book review: David B. Wilson (ed.), The correspondence between Sir George Gabriel Stokes and Sir William Thomson, Baron Kelvin of Largs, Cambridge, Cambridge University Press, 2 voll., 1990, lvi + 783 pp. ISBN 0-521-32831-4. Nuncius, 6(2):446–447, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01514.

Casati:1991:BRP

[809] Stefano Casati. Book review: Pisa, Biblioteca Universitaria — Ministero Per I Beni Culturali e Ambientali, Pisa, ottobre 1839. Il primo congresso degli scienziati italiani. Catalogo della mostra bibliografica in occasione del congresso nazionale ≪Prospettive della scienza e condizione umana alle soglie del 2000≫, Pisa, Litografia Tacchi, 1989, 172 pp. Nuncius, 6 (2):447–448, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539191x01523.

Malacrida:1991:BRAc

[810] Lorenzo Marri Malacrida. Book review: Arnaldo Cherubini, Beneficenza e Solidarietà. Assistenza pubblica e Mutualismo operaio. 1860– 1900. Milano, Franco Angeli, 1991, 400 pp. Nuncius, 6(2):448–449, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01532.

Segala:1991:BRL

[811] Marco Segala. Book review: Lewis Samuel Feuer, Einstein e la sua generazione. Nascita e sviluppo di teorie scientifiche, Bologna, Il Mulino, 1990 (≪Le occasioni≫ 35) ISBN 88-15-02809-9 (tr. it. di Gianfranco Ceccarelli di Einstein and the Generations of Science, New Brunswick, Transaction, Inc., 1982). Nuncius, 6(2):449–450, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x01541.

Papini:1991:BRH

[812] Paola Papini. Book review: Halton Arp, La contesa sulle distanze cosmiche e le quasar. (Italian) [The dispute over cosmic distances and quasars] Milano, Jaca Book, 1989, 218 p. (≪Di fronte e attraverso≫ ). Nuncius, 6(2):450–451, ???? 1991. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539191x01550.

Miniati:1991:ADMb

[813] Mara Miniati. Attività dell’Istituto e Museo di Storia Della Scienza di Firenze. Nuncius, 6(2):453–465, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539191x01569.

Beretta:1991:PQS

[814] Marco Beretta. Pubblichiamo qui di seguito le considerazioni di Mino Gabriele sulla scheda di Marco Beretta a M. Gabriele, le incisioni alchemico-metallurgiche di Domenico Beccafumi (≪nuncius≫, v, fasc. 2). Ad esse segue la replica di Marco Beretta. Nuncius, 6(2):467–493, ???? 1991. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539191x01578.

Meli:1992:GDM

[815] Domenico Bertoloni Meli. Guidobaldo Dal Monte and the Archimedean revival. Nuncius, 7(1):3–34, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00019.

LopezPinero:1992:PCC

[816] José M. López Piñero. The Pomar Codex (ca. 1590): Plants and animals of the Old World and from the Hernandez Expedition to America. Nuncius, 7(1):35–52, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00028.

Castellani:1992:RTS

[817] Carlo Castellani. I rapporti tra Spallanzani e Girardi e le ≪Osservazioni Anatomiche Intorno al Pipistrelli≫ di Michele Girardi. (Italian) [The relationships between Spallanzani and Girardi and the ≪anatomical observations of bats≫ by Michele Girardi]. Nuncius, 7(1):53–91, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00037.

Vaccari:1992:GAM

[818] Ezio Vaccari. Geologia e attività mineraria in Italia settentrionale tra settecento ed ottocento: l’influenza della ≪Scuola di Freiberg≫ su alcuni scienziati italiani 1. Nuncius, 7(1):93–107, ????

1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00046.

Scarpelli:1992:NNU

[819] Giacomo Scarpelli. ’nothing in nature that is not useful’ the antivaccination crusade and the idea of Harmonia Naturae in Alfred Russel Wallace. Nuncius, 7(1):109–130, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00055.

Turner:1992:MTB

[820] A. J. Turner. Melchisedech Thevenot, the bubble level, and the artificial horizon. Nuncius, 7(1):131–145, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00064.

Bernabeo:1992:PAD

[821] Raffaele A. Bernabeo, Stefano Arieti, Donatella Lippi, and Bonifacio Pistacchio. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 7(1):147–166, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00073.

Folkerts:1992:IF

[822] Menso Folkerts, Andreas Kühne, and Michael Segre. Istituzioni e fonti. Nuncius, 7(1):167–171, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00082.

Anonymous:1992:CBIa

[823] Anonymous. Commission on bibliography of the International Union of the History and Philosophy of Science. Nuncius, 7(1):173, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00091.

Ciancio:1992:RCV

[824] Luca Ciancio. Report on the Conference ≪Non-Verbal Communication in Science Prior to 1900≫ (Trento, 1st–3rd October 1991). Nuncius, 7 (1):174–179, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00109.

LaVergata:1992:R

[825] Antonello La Vergata. Ricordo. Nuncius, 7(1):181–194, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00118.

Abbri:1992:DC

[826] Ferdinando Abbri. Discussioni critiche. Nuncius, 7(1):195–205, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00127.

Morpurgo:1992:FIS

[827] Piero Morpurgo. Federico II e le scienze: un convegno di storia della scienza medievale. Nuncius, 7(1):207–217, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539192x00136.

Brenni:1992:RDC

[828] Paolo Brenni. Il restauro della collezione di apparecchi di fisica del liceo visconti di roma. Nuncius, 7(1):219–222, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539192x00145.

Hornung:1992:BRM

[829] Erik Hornung. Book review: Marshall Clagett, Ancient Egyptian Science. A Source Book, Vol. I: Knowledge and Order, American Philosophical Society, Philadelphia, 1989 (≪Memoirs of the American Philosophical Society≫, Vol. 184), 15 + 736 pp., ca. 150 figs., US$60. Nuncius, 7 (1):223–225, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00154.

Caroti:1992:BRSa

[830] Stefano Caroti. Book review: The Second Sense. Studies in Hearing and Musical Judgement from Antiquity to the Seventeenth Century, ed. by Charles Burnett, Michael Fend and Penelope Gouk, the Warburg Institute, London, 1991 (≪Warburg Institute Surveys and Texts≫, XXII), 245 pp., ind., bibl., ill. Nuncius, 7(1):225–227, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00163.

Vista:1992:BRA

[831] Michele Vista. Book review: Atti del III convegno nazionale di storia e fondamenti della chimica. F. Abbri — F. Crispini (a cura di), Cosenza, Edizioni Brenner, 1991, 402 pp. Nuncius, 7(1):227–231, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00172.

Miniati:1992:BRIa

[832] Mara Miniati. Book review: Isaia Iannaccone, Adolfo Tamburello (a cura di), Dall’Europa alla Cina: contributi per una storia dell’Astronomia, Napoli, Università degli Studi ≪Federico II≫, Istituto Universitario Orientale Napoli, 1990, 256 pp. Nuncius, 7(1):231–238, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00181.

Olmi:1992:BRW

[833] Giuseppe Olmi. Book review: Wilma George — Brunsdon Yapp, the Naming of the Beasts. Natural history in the medieval bestiary, London, Gerald Duckworth & Co Ltd., 1991, xiv + 231 pp., ill., bibl., ind. Nuncius, 7(1):238–240, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00190.

Beretta:1992:BRM

[834] Marco Beretta. Book review: Michelle Goupil, Du Flou au clair? Histoire de l’affinité chimique de Cardan à Prigogine, Paris, Édition du Comité des Travaux Historiques et Scientifiques, 1991, 348 pp., ill. Nuncius, 7 (1):240–242, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00208.

Miniati:1992:BRG

[835] Mara Miniati. Book review: Gerard l’E. Turner, Scientific Instruments and Experimental Philosophy 1550–1850, Aldershot, Hampshire, Variorum, 1990, xii + 329 pp., fig. Nuncius, 7(1):242–243, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00217.

Caroti:1992:BRQ

[836] Stefano Caroti. Book review: Questions d’histoire de la médecine. Actes du 113e Congrès National des Sociétés Savantes (Strasbourg, 5–9 avril 1988). Section d’histoire des sciences et des techniques, Éditions du C.T.H.S., Paris, 1991, 113 pp., bibl., ill. Nuncius, 7(1):244–245, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00226.

Abbri:1992:BRA

[837] Ferdinando Abbri. Book review: Allen G. Debus, the French Paracelsians. the Chemical Challenge to Medical and Scientific Tradition in Early Modern France, Cambridge, Cambridge University Press, 1991, xvi + 247 pp. Nuncius, 7(1):246–248, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00235.

Plastina:1992:BRA

[838] Sandra Plastina. Book review: Atti del convegno internazionale di studi su Bernardino Telesio, Cosenza, Accademia Cosentina, 1990, 220 pp. Nuncius, 7(1):248–253, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00244.

Maltese:1992:BRG

[839] Giulio Maltese and Giorgio Dragoni. Book review: Gerard l’E. Turner, God bless the microscope, Oxford, Royal Microscopical Society, 1989. Nuncius, 7(1):253–255, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00253.

Tomasi:1992:BRP

[840] Lucia Tongiorgi Tomasi. Book review: Paul Hulton, F. Nigel Hepper & Ib Friis, Luigi Balugani’s Drawings of African Plants. From the collection made by James Bruce of Kinnaird on his travels to discover the source of the Nile 1767–1773. New Haven, the Yale Center for British Art — Rotterdam, A. A. Balkema, 1991, 140 pp., 16 fig., 330 tav. (di cui 48 a colori). Nuncius, 7(1):255–258, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00262.

Beretta:1992:BRC

[841] Marco Beretta. Book review: L. C. Rookmaaker, the Zoological Exploration of Southern Africa 1650–1790. Rotterdam, A. A. Balkema, 1989, xii + 368 pp., ill. Nuncius, 7(1):258–260, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00271.

Abbri:1992:BRW

[842] Ferdinando Abbri. Book review: William R. Shea, The Magic of Numbers and Motion. The Scientific Career of René Descartes, Canton Mass., Science History Publications, 1991, xiii + 371 pp. Nuncius, 7(1):260–262, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00280.

Brenni:1992:BRE

[843] Paolo Brenni. Book review: Edwin Banfield, Barometer Makers and Retailers 1660–1900, Trowbridge, Baros Book, 246 pp. Nuncius, 7(1):263, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00299.

Miniati:1992:BRA

[844] Mara Miniati. Book review: Archivio di Stato di Firenze, La Toscana dei Lorena nelle mappe dell’Archivio di Stato di Praga. Memorie ed immagini di un Granducato. Catalogo e mostra documentaria Firenze, 31 maggio-31 luglio 1991. Roma, Ministero per i Beni Culturali e Ambientali, Ufficio Centrale per i Beni Archivistici, 1991, 430 pp., fig. Nuncius, 7(1):264–267, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00307.

Iannaccone:1992:BRP

[845] Isaia Iannaccone. Book review: Paul U. Unschuld, Forgotten Traditions of Ancient Chinese Medicine, Brookline Massachusetts, Paradigm Publication, 1990, 403 pp. Nuncius, 7(1):267–270, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00316.

Beretta:1992:BRR

[846] Marco Beretta. Book review: Robert Darnton and Daniel Roche (Editors), Revolution in Print. The Press in France 1775–1800, Berkeley and Los Angeles, University of California Press, 1989, xv + 351 pp., ill. Nuncius, 7(1):270–272, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00325.

Brenni:1992:BRC

[847] Paolo Brenni. Book review: Curiosity perfectly satisfied. Faraday’s travels in Europe 1813–1815. A cura di B. Bowers e L. Symons. IEE History of Technology Series 16, Peter Peregrinus Ltd in association with the Science Museum, London, 1991, 168 pp. (ISBN 0-86341-234-3) (≪IEE History of Technology Series≫, 16). Nuncius, 7(1):272–274, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00334.

LaVergata:1992:BRH

[848] Antonello La Vergata. Book review: History, humanity and evolution. Essays for John C. Greene, edited by James R. Moore, Cambridge, Cambridge University Press, 1989, xii + 429 pp., (ISBN 0-521-33511-6). Nuncius, 7(1):274–278, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00343.

Segala:1992:BRPa

[849] Marco Segala. Book review: Peter M. Harman, The scientific letters and papers of James Clerk Maxwell, 1846–1862, vol. I, Cambridge, Cambridge University Press, 1990, xxviii + 748 pp. (ISBN 0-521-25625-9). Nuncius, 7(1):278–281, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00352.

LaVergata:1992:BRD

[850] Antonello La Vergata. Book review: Daniel Pick, Faces of degeneration. A European disorder, c. 1848–1918, Cambridge, Cambridge University Press (≪Ideas in Context≫ ), 1989, vi + 275 pp., L. 27,50 ($39,50) (ISBN 0-521-36021-8). Nuncius, 7(1):281–284, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00361.

Scarpelli:1992:BRC

[851] Giacomo Scarpelli. Book review: The Correspondence of Charles Darwin, Volume 6, 1856–1857, a cura di Frederick Burkhardt e Sydney Smith, Cambridge & New York, Cambridge University Press, 1990, xxxii + 673 pp., L. 35.00 / $49.50 (ISBN 0-521-25586-4). Nuncius, 7(1):284– 286, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00370.

Abbri:1992:BRB

[852] Ferdinando Abbri. Book review: Bruno P. F. Wanrooij, Storia del pudore. La questione sessuale in Italia 1860–1940, Venezia, Marsilio Editori, 1990, 234 pp. Nuncius, 7(1):286–289, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539192x00389.

Miniati:1992:Sa

[853] Mara Miniati. Strumentaria. Nuncius, 7(1):291, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539192x00398.

Miniati:1992:BRR

[854] Mara Miniati. Book review: E. Ragozzino, R. Rinzivillo, E. Schettino, I microscopia del passato. Una introduzione allo studio delle qualità ottiche e meccaniche dei microscopi del museo, Napoli, Cuen, 1991, 72 pp., fig. Nuncius, 7(1):291–292, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00406.

Brenni:1992:BRA

[855] Paolo Brenni. Book review: Aa.Vv., Les mécanismes du génie, Editéur responsable Europalia Dr. Robert De Smet, Charleroi, 1991, 426 pp. Nuncius, 7(1):292–293, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00415.

Brenni:1992:BRK

[856] Paolo Brenni. Book review: Kenneth Lyall, Electrical and Magnetic Instruments, Catalogue 8, Cambridge, the Whipple Museum of the History of Science, 1991, 136 pp. Nuncius, 7(1):293–294, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00424.

Caroti:1992:BRR

[857] Stefano Caroti. Book review: Ronald Edward Zupko, Revolution in Measurement: Western European Weights and Measures Since the Age of Science, Philadelphia, the American Philosophical Society, 1990, 547 pp., ind., app., ill. Nuncius, 7(1):294, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00433.

Miniati:1992:BRCa

[858] Mara Miniati. Book review: Comune di Pesaro-Cassa di Risparmio Di Pesaro, La terra e le sue copie. Leonardo, Raffaello, Boscovich, la cartografia del ducato e della legazione di Urbino, a cura di Enrico Gamba e Vico Montebelli, Urbino, Quattroventi, 1991, 95 pp., fig. Nuncius, 7(1): 295, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00442.

Miniati:1992:BRW

[859] Mara Miniati. Book review: Wilma di Palma, Aspetti della Cultura Matematica a Roma dal I al VI secolo d.C., Roma, Fratelli Palombi Editori, 1991, 27 pp., fig. (Comune di Roma, Assessorato alla Cultura Centro di Coordinamento Didattico, 40). Nuncius, 7(1):295–296, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00451.

Segala:1992:BRL

[860] Marco Segala. Book review: Laurie M. Braun — Max Dresden — Lillian Hoddeson (eds.), Pions to quarks. Particle physics in the 1950s, New York, Cambridge University Press, 1989, xxxii + 734 pp. (ISBN 0-521-30984-0). Nuncius, 7(1):296–297, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00460.

Miniati:1992:BRLa

[861] Mara Miniati. Book review: Lilia Capocaccia Orsini, Giorgio Doria, Giuliano Doria (a cura di), 1492–1992 Animali e piante dalle Americhe all’Europa, Genova, Sagep Editrice, 1991, 323 pp., fig. Nuncius, 7(1): 297–298, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00479.

Miniati:1992:BRIb

[862] Mara Miniati. Book review: Irma B. Jaffe, Gianni Eugenio Viola, Franca Rovigati (eds.), Imagining the New World. Columbian Iconography, Roma, New York, Istituto della Enciclopedia italiana, 1991, xi + 234 pp. fig. Nuncius, 7(1):298–299, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00488.

Malacrida:1992:BRFa

[863] Lorenzo Marri Malacrida. Book review: Fabio Garbari, Lucia Tongiorgi Tomasi, Alessandro Tosi, Giardino dei Semplici, l’orto botanico di Pisa dal XVI al XX secolo, Pisa, Cassa di Risparmio di Pisa, 1991, 397 pp., ed. fuori commercio. Nuncius, 7(1):299–300, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00497.

Miniati:1992:BROa

[864] Mara Miniati. Book review: Oceania. Collezione dell’Oceania del Museo di Antropologia e d’Etnografia dell’Università di Padova, Padova, Editime, 1991, 58 pp., fig. Nuncius, 7(1):300, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00505.

LaVergata:1992:BM

[865] Antonello La Vergata. Biologia e medicina. Nuncius, 7(1):301–302, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00514.

LaVergata:1992:BRN

[866] Antonello La Vergata. Book review: La naissance de la vie. Une anthologie, Textes choisies et présentés par Jean-Louis Fischer, Paris, Presses Pocket (Agora. Les classiques, Collection dirigée par Olivier Amiel), 1991, 296 pp. Nuncius, 7(1):302–303, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00523.

Lippi:1992:BRAa

[867] Donatella Lippi. Book review: Aa.Vv., Ars Medica. La Medicina nell’antica Roma, Catalogo della Mostra, Milano, RCS Rizzoli Corriere della Sera, 1991, 64 pp., tav. Nuncius, 7(1):303–304, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00532.

Lippi:1992:BRE

[868] Donatella Lippi. Book review: Les Écoles Médicales à Rome — Actes du 2ème Colloque Int. sur les textes médicaux latins antiques, Lausanne, septembre 1986, ed. prép. par. P. Mudry et J. Pigeaud, Publications de la Fac. des Lettres — Un. Lausanne —, XXXIII, Nantes, 1991, 317 pp. Nuncius, 7(1):304, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00541.

Lippi:1992:BRPa

[869] Donatella Lippi. Book review: P. Micoli-M. Rotoli, I capelli nell’antico Egitto. — Patologia, terapia, cosmesi nei testi medici, Bergamo, Bolis Ed., 1991, 115 pp. Nuncius, 7(1):304–305, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00550.

Malacrida:1992:BRGa

[870] Lorenzo Marri Malacrida. Book review: Giuseppe Penso, La medicina medioevale, Milano, Ciba Geigy edizioni, 1991, 608 pp., fig. Nuncius, 7 (1):305–306, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00569.

Malacrida:1992:BRB

[871] Lorenzo Marri Malacrida. Book review: Bonifacio Pistacchio, Lo stomaco nella storia della medicina, Bologna, Collana scientifica Wassermann, edizioni Golfond, 1991, 139 pp., fig., ed. fuori commercio. Nuncius, 7 (1):306–307, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00578.

Lippi:1992:BRM

[872] Donatella Lippi. Book review: M. Toussaint-Samat, Storia naturale e morale dell’alimentazione, Firenze, Sansoni, 1991, pp. 901 (Ed. or. Histoire naturelle & morale de la nourriture, Paris, 1987). Nuncius, 7(1): 307–308, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00587.

Caroti:1992:BRF

[873] Stefano Caroti. Book review: Francesco Raspadori (a cura di), I maestri di medicina ed arti dell’Università di Ferrara 1390–1950, Firenze, Leo S. Olschki, 1991 (≪Pubblicazioni dell’Università di Ferrara≫, II), 295 pp. Nuncius, 7(1):308–309, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00596.

Laquidara:1992:BRA

[874] Lanmarco Laquidara. Book review: Alessandro Dini, Vita e organismo. Le origini della fisiologia sperimentale in Italia, Firenze, Leo S. Olschki editore, 1991, 207 pp. Nuncius, 7(1):309–310, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00604.

Malacrida:1992:BRM

[875] Lorenzo Marri Malacrida. Book review: Michele Pontecorvo, Storia delle vaccinazioni, dalle origini ai nostri giorni, Milano, Ciba-Geigy ed., 1991, 151 pp., fig. Nuncius, 7(1):310, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00613.

Malacrida:1992:BRGb

[876] Lorenzo Marri Malacrida. Book review: Giovanni Battista Giordano, Storia dell’ospedale di Fregionaia, nel secolo XIX ed inventario del suo archivio storico dal 1813 al 1942, Roma, Delfino editore, 1991, pp. 251. Nuncius, 7(1):311, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00622.

Lippi:1992:BRL

[877] Donatella Lippi. Book review: Luciano Sterpellone, Mozart tra medici e medicine, Cinisello Balsamo — Mi — Ediz. Paoline, 1991, pp. 177, L. 22.000. Nuncius, 7(1):312, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00631.

Miniati:1992:BRCb

[878] Mara Miniati. Book review: Claudio Pogliano, Pharmakon. Storia dello Psicotropismo. Secondo volume, s.l., Casamassima Editore, 1991, 373-730 pp., fig. Nuncius, 7(1):313, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00640.

Malacrida:1992:BRFb

[879] Lorenzo Marri Malacrida. Book review: Francesca Vannozzi (a cura di), l’assistenza pubblica nella Siena di fine ’800, Milano, Electa, 1991, 81 pp. Nuncius, 7(1):313, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00659.

Laquidara:1992:BRD

[880] Lanmarco Laquidara. Book review: Domenico De Napoli, La sanità militare in Italia durante la prima guerra mondiale, Roma, Apes editrice, 1989, 321 pp. Nuncius, 7(1):314, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00668.

Laquidara:1992:BRP

[881] Lanmarco Laquidara. Book review: Patrizia Guarnieri, La storia della psichiatria. un secolo di studi in Italia, Firenze, Leo S. Olschki ed., 1991, 53 + 102 pp. Nuncius, 7(1):314–315, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00677.

Lippi:1992:BRPb

[882] Donatella Lippi. Book review: Pedro Laín Entralgo, Antropologia Medica, Cinisello Balsamo, Edizioni Paoline, 1988, 444 pp., L. 22.000 (tit. or. Antropología médica, Barcelona, Salvat. Ed., 1984). Nuncius, 7(1):315– 316, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00686.

Laquidara:1992:BRG

[883] Lanmarco Laquidara. Book review: Giovanni Berlinguer, Questioni di vita. Etica, scienza, salute, Einaudi, 1991, (≪Gli struzzi≫ ). Nuncius, 7(1):316–317, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00695.

Beretta:1992:BRJ

[884] Marco Beretta. Book review: Johan Gottlieb Gahn, Brev, Utgivna med kommentarer av Jan Trofast. Del I., Lund, Wallin & Dalholm Boktryckeri, 1991, 307 pp., ill. Nuncius, 7(1):317–318, ????

1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00703.

Abbri:1992:BRD

[885] Ferdinando Abbri. Book review: Donald J. Cram, From Design to Discovery, Washington Dc, American Chemical Society, 1990, xxi + 146 pp. (≪Profiles, Pathways, and Dreams. Autobiographies of Eminent Chemists, 3≫ ). Nuncius, 7(1):318–319, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00712.

Caroti:1992:BRG

[886] Stefano Caroti. Book review: Guy Beaujouan, Par raison de nombres. l’art du calcul et les savoirs scientifiques médiévaux, Great Yarmouth, Variorum, 1991, XII + 300 pp., ind., ill. (≪Collected Studies Series≫ CS 344). Nuncius, 7(1):319–321, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00721.

Caroti:1992:BRDa

[887] Stefano Caroti. Book review: Daniela Mugnai Carrara, La biblioteca di Nicolò Leoniceno. Tra Aristotele e Galeno: cultura e libri di un medico umanista, Leo S. Olschki, Firenze, 1991 (≪Accademia toscana di scienze e lettere ‘La Colombaria’. Studi≫, Cxviii), 247 pp., ind. Nuncius, 7(1): 321–322, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00730.

Abbri:1992:BRV

[888] Ferdinando Abbri. Book review: Victor E. Thoren, the Lord of Uraniborg. A Biography of Tycho Brahe, with contributions by John R. Christianson, Cambridge, Cambridge University Press, 1990, xii + 523 pp. Nuncius, 7(1):322–324, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00749.

Abbri:1992:BRJa

[889] Ferdinando Abbri. Book review: José Pardo Tomàs, Ciencia y Censura. La Inquisición Española y los Libros Científicos en los Siglos XVI y XVII, Madrid, Consejo Superior de Investigationes Científicas, 1991, xxiv + 390 pp. (≪Estudios sobre la Ciencia, 13≫ ). Nuncius, 7(1):325– 326, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00758.

Caroti:1992:BRM

[890] Stefano Caroti. Book review: Mark Kulstad, Leibniz on Apperception, Consciousness, and Reflection, München, Philosophia Verlag, 1991, 183 pp., ind., bibl. (≪Analytica≫ ). Nuncius, 7(1):326–327, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00767.

Abbri:1992:BRSa

[891] Ferdinando Abbri. Book review: Susanna ÅKerman, Queen Christina of Sweden and Her Circle. The transformation of a Seventeenth-Century Philosophical Libertine, Leiden, E. J. Brill, 1991, xv + 399 pp. (≪Brill’s Studies in Intellectual History, 21≫ ). Nuncius, 7(1):327–330, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00776.

Segala:1992:BRM

[892] Marco Segala. Book review: Margaret C. Jacob, the Newtonians and the English Revolution 1689–1720, II ed., New York, Gordon and Breach, 1990, 288 pp. (≪Classics in the History and Philosophy of Science≫ ), 7 (ISBN 2-88124-400-9). Nuncius, 7(1):330–331, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00785.

Miniati:1992:BRCc

[893] Mara Miniati. Book review: Charles De Brosses, Lettres familières. Texte établi par Giuseppina Cafasso. Introduction, notes et bibliographie par Letizia Norci Cagiano de Azevedo. Préface de Giovanni Macchia. Séjour à Florence. Naples, Centre Jean Bérard, 1991, pp. 407-462, 1 fig. Nuncius, 7(1):331, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00794.

Abbri:1992:BRFa

[894] Ferdinando Abbri. Book review: Franco Arato, Il secolo delle cose. Scienza e storia in Francesco Algarotti, Genova, Marietti, 1991, 173 pp.

(≪Ricerche Studi Strumenti — Letteratura, 25≫ ). Nuncius, 7(1):331– 332, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00802.

Abbri:1992:BRR

[895] Ferdinando Abbri. Book review: Rainer Baasner, Abraham Gotthelf Kästner, Aufklärer (1719–1800), Tübingen, Max Niemeyer Verlag, 1991, vi + 683 pp. (≪Frühe Neuzeit, Band 5≫ ). Nuncius, 7(1):332–334, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00811.

Abbri:1992:BRFb

[896] Ferdinando Abbri. Book review: Friedrich Wilhelm Joseph Schelling, La filosofia della natura (1797–1801). Brani scelti a cura di Carlo Tatasciore, Napoli, Edizioni ≪Il Tripode≫, 1990, 144 pp. Nuncius, 7(1):334, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00820.

Anonymous:1992:BRD

[897] Anonymous. Book review: Delfina Dolza, Essere figlie di Lombroso. Due donne intellettuali tra ’800 e ’900, Milano, Franco Angeli, 1990, 239 pp. Nuncius, 7(1):334–350, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00839.

Baldini:1992:BTG

[898] Ugo Baldini. Boscovich e la tradizione gesuitica in filosofia naturale: continuità e cambiamento. Nuncius, 7(2):3–68, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539192x00848.

Apfelstadt:1992:CCP

[899] Eric Apfelstadt. Christopher Columbus, Paolo Dal Pozzo Toscanelli and Ferno De Roriz: New evidence for a Florentine connection. Nuncius, 7(2):69–80, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00857.

Biagioli:1992:FSB

[900] Mario Biagioli. Filippo Salviati: a Baroque virtuoso. Nuncius, 7(2):81– 96, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00866.

Berni:1992:VC

[901] Stefano Berni. Vincenzio chiarugi. Nuncius, 7(2):97–113, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00875.

Miniati:1992:NON

[902] Mara Miniati. Un nuovo orologio notturno. Nuncius, 7(2):115–117, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00884.

Abbri:1992:CLP

[903] Ferdinando Abbri. ≪commercio di libri e produzioni naturali≫. Nuncius, 7(2):119–139, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00893.

Ciancio:1992:CAC

[904] Luca Ciancio. Contributo all’inventario del carteggio di Alberto Fortis. Nuncius, 7(2):141–158, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x00901.

Anonymous:1992:IAM

[905] Anonymous. Indice alfabetico dei mittenti e dei destinatari. Nuncius, 7(2):159, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00910.

Ciardi:1992:TSE

[906] Marco Ciardi. Teorie senza esperimenti. Nuncius, 7(2):161–193, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00929.

Anonymous:1992:CBIb

[907] Anonymous. Commission on bibliography of the International Union of the History and Philosophy of Science. Nuncius, 7(2):195, ????

1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00938.

Vermij:1992:RN

[908] R. H. Vermij. Report on the Netherlands (1989–1992). Nuncius, 7(2): 196–208, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x00947.

Iannaccone:1992:SEE

[909] Isaia Iannaccone. Sinologia ed epistemologia: Orientamenti metodologici nello studio della storia della scienza. Nuncius, 7(2):209–230, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00956.

Caroti:1992:NPD

[910] Stefano Caroti. Nuove prospettive della storiografia di storia della scienza medievale: Le proposte di John E. Murdoch. Nuncius, 7(2):231–252, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00965.

Brenni:1992:MGB

[911] Paolo Brenni. Una mostra su Giuseppe Belli. Nuncius, 7(2):253–255, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x00974.

LaVergata:1992:BRG

[912] Antonello La Vergata. Book review: Giovanni Delfino, Eudes Lanciotti, Gianfranco Liguri e Massimo Stefani, Biologia e medicina. Biology & medicine. Dizionario enciclopedico di scienze biologiche e mediche, italiano-inglese, inglese-italiano, Bologna, Zanichelli, 1990, vi + 1719 pp., L. 80.000. Nuncius, 7(2):257–260, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00983.

Segala:1992:BRR

[913] Marco Segala. Book review: Robert K. Merton, Sulle spalle dei giganti, Bologna, Il Mulino, 1991, 294 pp. (≪Le occasioni≫, 42) ISBN 88-15-03267-3 (On the Shoulders of Giants, New York, the Free Press, 1965, 19852). Nuncius, 7(2):260–263, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x00992.

Touwaide:1992:BRR

[914] Alain Touwaide. Book review: Robert I. Curtis, Garum and Salsament. Production and Commerce in Materia Medica, Leiden, New York, København, Köln, E. J. Brill, 1991, xv + 226 pp., 8 pl., n/bl. (≪Studies in Ancient Medicine≫, 3) (ISBN 90-04-09423-7). Nuncius, 7(2):263–265, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01009.

Caroti:1992:BRSb

[915] Stefano Caroti. Book review: Gli studi di filosofia medievale fra Otto e Novecento. Contributo a un bilancio storiografico. Atti del convegno internazionale Roma, 21–23 settembre 1989, a cura di Ruedi Imbach e Alfonso Maierù, Roma, Edizioni di storia e letteratura, 1991, 425 pp., ind. (≪Storia e letteratura. Raccolta di studi e testi≫, 179). Nuncius, 7(2):265–268, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x01018.

Miniati:1992:BRJ

[916] Mara Miniati. Book review: John E. Burnett, Alison D. Morrison-Low, ≪Vulgar and mechanick≫. The Scientific Instrument Trade in Ireland 1650–1921, Dublin, Mount Salus Press Limited, 1989, 166 pp., fig. (≪Royal Dublin Society Historical Studies in Irish Science and Technology≫, Number Eight). Nuncius, 7(2):268–270, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01027.

Maugini:1992:BRF

[917] Elena Maugini. Book review: Fabio Garbari — Lucia Tongiorgi Tomasi — Alessandro Tosi, Giardino dei Semplici. l’Orto botanico di Pisa dal XVI al XX secolo, Pisa, Cassa di Risparmio di Pisa, Pacini, 1991, 397 pp., 323 tavv. Nuncius, 7(2):270–273, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01036.

Abbri:1992:BRJb

[918] Ferdinando Abbri. Book review: Julian Martin, Francis Bacon, the State, and the reform of Natural Philosophy, Cambridge, Cambridge University Press, 1992, xiv + 236 pp. Nuncius, 7(2):273–275, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01045.

Giuntini:1992:BRMa

[919] Chiara Giuntini. Book review: Marie Boas Hall, Promoting Experimental Learning. Experiment and the Royal Society, 1660–1727, Cambridge– New York, Cambridge University Press, 1991, xiii + 207 pp. Nuncius, 7(2):275–278, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x01054.

Meli:1992:BRM

[920] Domenico Bertoloni Meli. Book review: Michel Blay, La Naissance de la mécanique analytique. La science du mouvement au tournant des XVIIe et XVIIIe siècles, Paris, Presses Universitaires de France, 1992, 414 pp., Ff 315. Nuncius, 7(2):278–280, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01063.

Redondi:1992:BRN

[921] Pietro Redondi. Book review: Nicolas-Claude Fabri De Peiresc, Lettres à Cassiano dal Pozzo (1626–1637), editées et commentées par J.-F. Lhote et D. Joyal, preface de J. Guillerme, Clermont-Ferrand, Éditions Adosa, 1989, 277 pp. Nuncius, 7(2):280–285, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01072.

Monti:1992:BRA

[922] Maria Teresa Monti. Book review: Antonio Vallisneri, Epistolario. 1679– 1710, vol. I, a cura di Dario Generali, Milano, Franco Angeli, 1991, 639 pp. (≪Filosofia e Scienza nel Cinquecento e nel Seicento. Testi≫, 9). Nuncius, 7(2):285–290, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x01081.

Abbri:1992:BRJc

[923] Ferdinando Abbri. Book review: Jan Golinski, Science as Public Culture. Chemistry and Enlightenment in Britain, 1760–1820, Cambridge, Cambridge University Press, 1992, xii + 342 pp. Nuncius, 7(2):290–294, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01090.

Scarpelli:1992:BRA

[924] Giacomo Scarpelli. Book review: Alessandro Dini, Vita e organismo. Le origini della fisiologia sperimentale in Italia, Firenze, Olschki, 1991, 213 pp., L. 43.000. Nuncius, 7(2):294–296, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01108.

Giuntini:1992:BRJ

[925] Chiara Giuntini. Book review: James E. Crimmins, Secular Utilitarianism. Social Science and the Critique of Religion in the Thought of Jeremy Bentham, Oxford, Clarendon Press, 1990, 348 pp. Nuncius, 7 (2):296–299, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x01117.

Crispini:1992:BRE

[926] Ines Crispini. Book review: Elio Matassi, Eredità hegeliane. Da Cieszkowski e Gans a Ritter, Napoli, Morano Editore, 1991, 313 pp. (≪Nobiltà dello spirito≫, 29). Nuncius, 7(2):299–302, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01126.

LaVergata:1992:BRA

[927] Antonello La Vergata. Book review: Alvar Ellegård, Darwin and the general reader. The reception of Darwin’s Theory of Evolution in the British periodical press, 1859–1872. With a new foreword by D. L. Hull, Chicago and London, The University of Chicago Press, 1990, 394 pp., ISBN 0-226-20487-1. Nuncius, 7(2):302–304, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01135.

Iannaccone:1992:BRJ

[928] Isaia Iannaccone. Book review: J. Reardon-Anderson, the Study of Change. Chemistry in China, 1840–1949, Cambridge, Cambridge University Press, 1991, xix + 444 pp. Nuncius, 7(2):304–309, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01144.

Monti:1992:BRC

[929] Maria Teresa Monti. Book review: Catalogo del fondo Giovanni Battista Palletta. A cura di G. Marsala e L. Panzeri. Introduzione di G. Cosmacini, Milano, Editrice Bibliografica, 1991. (≪Fonti e strumenti≫, 17), xx + 155 pp. Nuncius, 7(2):310–312, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01153.

LaVergata:1992:BRF

[930] Antonello La Vergata. Book review: Filippo Parlatore, Mie memorie, a cura di Agnese Visconti, Palermo, Sellerio editore, 1992, 482 pp., 32 illustrazioni fuori testo, lire 50.000. Nuncius, 7(2):312–314, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01162.

Barsanti:1992:BRP

[931] Giulio Barsanti. Book review: Paolo Mantegazza, Lezioni di antropologia (1870–1910), Firenze, Società Italiana di Antropologia e Etnologia, 1989, 2 voll. di complessive 916 pp., s.i.p. Nuncius, 7(2):314–316, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01171.

Berni:1992:BRG

[932] Stefano Berni. Book review: Gustavo Micheletti, I Pensieri sordi e l’inconscio, Roma, Borla, 1991, 357 pp., L. 28.000. Nuncius, 7(2):316– 319, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01180.

Giuntini:1992:BRMb

[933] Chiara Giuntini. Book review: Metaphors in the history of psychology, edited by D. E. Leary, Cambridge–New York, Cambridge University Press, 1990, xiii + 380 pp. Nuncius, 7(2):319–323, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01199.

Artuso:1992:BRM

[934] Paolo Artuso. Book review: M. Arbib — M. Hesse, La costruzione della realtà, Bologna, Il Mulino, 1992, 420 pp. E. Bellone, Saggio naturalistico sulla conoscenza, Torino, Bollati Boringhieri, 1992, 135 pp. Nuncius, 7 (2):323–330, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x01207.

Caroti:1992:SB

[935] Stefano Caroti. Strumenti bibliografici. Nuncius, 7(2):331–332, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01216.

Miniati:1992:Sb

[936] Mara Miniati. Strumentaria. Nuncius, 7(2):332–333, ???? 1992. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539192x01225.

Miniati:1992:BRM

[937] Mara Miniati. Book review: Marinella Calisi, Guida alla visita del Museo Astronomico e Copernicano di Roma, Roma, Osservatorio Astronomico e Museo Astronomico e Copernicano, 1991, 145 pp., fig. Nuncius, 7(2): 333, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01234.

Miniati:1992:BRX

[938] Mara Miniati. Book review: XIX sajandi alguse füüsika riistu Tartu Ülikooli Ajaloomuuseumis. (Estonian) [XIX century physics utensils. Tartu University History Museum]. Tartu, 1989, 78 pp., fig. Nuncius, 7(2):333–334, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x01243.

Miniati:1992:BRD

[939] Mara Miniati. Book review: David J. Bryden, Napier’s Bones. A History and Instruction Manual, London, Harriet Wynter Ltd., 1992, 23 pp., fig. Nuncius, 7(2):334, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x01252.

Caroti:1992:MF

[940] Stefano Caroti. Matematica e fisica. Nuncius, 7(2):334–335, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01261.

Miniati:1992:BROb

[941] Mara Miniati. Book review: Oltre il compasso. La geometria delle curve, a cura di Franco Conti e Enrico Giusti, Pisa, Scuola Normale Superiore, 1992, 73 pp., fig. Nuncius, 7(2):335, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01270.

Segala:1992:BRPb

[942] Marco Segala. Book review: Peter Coveney — Roger Highfield, La freccia del tempo. Viaggio attraverso uno dei grandi misteri della scienza, Milano, Rizzoli, 1991, 455 pp. Nuncius, 7(2):336, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01289.

Abbri:1992:A

[943] Ferdinando Abbri. Antropologia. Nuncius, 7(2):337–338, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01298.

Touwaide:1992:BM

[944] Alain Touwaide. Biologia e medicina. Nuncius, 7(2):338–339, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01306.

Touwaide:1992:BRT

[945] Alain Touwaide. Book review: Theophrastus, De causis plantarum, Books III–IV, Edited and translated by Benedict Einarson & George K. K. Link, Cambridge, Mass., Harvard University Press, London, W. Heinemann, 1990, vi + 361 pp. (≪Loeb Classical Library≫, 474). Theophrastus, De causis plantarum, Books V-Vi, Edited and translated by Benedict Einarson & George K. K. Link, Cambridge, Mass., Harvard University Press, London, W. Heinemann, 1990, vii + 465 pp. (≪Loeb Classical Library≫, 472). Nuncius, 7(2):340–341, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01315.

Touwaide:1992:BRJ

[946] Alain Touwaide. Book review: J. T. Vallance, the lost Theory of Asclepiades of Bithynia, Oxford, Clarendon Press, 1990, xi + 162 pp. Nuncius, 7(2):341–343, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x01324.

Lippi:1992:BRGa

[947] Donatella Lippi. Book review: G. Maconi, Storia della Medicina e della Chirurgia, Milano, Casa Ed. Ambrosiana, 1991, 374 pp. L. 55.000. Nuncius, 7(2):343–344, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x01333.

Lippi:1992:BRGb

[948] Donatella Lippi. Book review: G. Hertzka e W. Strehlow, Manuale della Medicina di Santa Ildegarda, Bolzano, Casa Ed. Athesia, 1992, 342 pp. (Tit. orig.: Handbuch der Hildegard-Medizin, Freiburg, H. Bauer Kg, 1987). Nuncius, 7(2):344–345, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01342.

Carrara:1992:BRS

[949] Daniela Mugnai Carrara. Book review: Stefano Caroti, I codici di Bernardo Campagna. Filosofia e medicina alla fine del sec. XIV, Manziana (Roma), Vecchiarelli Editore 1991, ix + 275 pp. Nuncius, 7(2):345–347, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x01351.

Lippi:1992:BRK

[950] Donatella Lippi. Book review: K. Bergdolt, Krankheit und Therapie bei Petrarca, Die Kritik am Medizin und Naturwissenschaft im italienischen Fruehhumanismus, Weinheim, VCH, Acta Humaniora, 1992, 364 pp., DM 168. Nuncius, 7(2):348, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01360.

Casati:1992:BRA

[951] Stefano Casati. Book review: Arthur Schnitzler, Paracelso a cura di Paola Maria Filippi, Faenza, Moby Dick, 1991, 124 pp. (≪I libri dello Zelig≫, 14). Nuncius, 7(2):349, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01379.

Lippi:1992:BRGc

[952] Donatella Lippi. Book review: G. Magherini — V. Biotti, l’isola delle Stinche e i percorsi della follia a Firenze nei secoli XIV–XVIII, Firenze, Ponte alle Grazie, 1992, 187 pp., figg. Nuncius, 7(2):349–351, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01388.

Miniati:1992:BROc

[953] Mara Miniati. Book review: Ordine Ospedaliero di San Giovanni di Dio Fatebenefratelli, Archivio Storico Del Comune di Firenze, l’Archivio dell’Ospedale di San Giovanni di Dio di Firenze (1604–1890), Inventario a cura di Lucia Sandri, Cernusco s/N (Milano), Edizioni Fatebenefratelli, 1991, 306 pp., fig., L. 25.000. Nuncius, 7(2):351–352, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01397.

Lippi:1992:BRAb

[954] Donatella Lippi. Book review: Actes du 4e colloque des conservateurs des musées d’histoire des sciences medicales, 7 au 10 septembre 1988, Pavia-Milano, Lyon, Fond. Mérieux, 1990, 327 pp. Nuncius, 7(2):352– 353, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01405.

Carrara:1992:BRC

[955] Daniela Mugnai Carrara. Book review: Clinical Consultations and Letters by Ippolito Francesco Albertini, Francesco Torti and Other Physicians. University of Bologna ms. 2089-1. Translated and annotated with an introductory analysis by Saul Jarcho, Boston, the Francis A. Countway Library of Medicine, 1989, lxix + 356 pp. Nuncius, 7(2):353–355, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01414.

Lippi:1992:BRAc

[956] Donatella Lippi. Book review: Aa.Vv., Le cere ostetriche romane di Giovan Battista Manfredini, Roma, Quasar, 1991, 101 pp. Nuncius, 7 (2):355–356, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x01423.

Segala:1992:BRA

[957] Marco Segala. Book review: Antonio Cadeddu, Dal mito alla storia. Biologia e medicina in Pasteur, Milano, Franco Angeli, 1991, 296 pp. Nuncius, 7(2):356–357, ???? 1992. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539192x01432.

Lippi:1992:BRF

[958] Donatella Lippi. Book review: F. Antonini — S. Magnolfi, l’età dei capolavori. Creatività e vecchiaia nelle arti figurative, Venezia, Marsilio Ed., 1991, 243 pp., L. 80.000. Nuncius, 7(2):357–358, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01441.

Abbri:1992:BRM

[959] Ferdinando Abbri. Book review: Marie Annette Brown, Gail M. Powell-Cope, Caring for a Loved One with Aids. The Esperiences of Families, Lovers, and Friends, Seattle and London, The University of Washington Press 1992, vii + 64 pp. Nuncius, 7(2):358–359, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01450.

Miniati:1992:C

[960] Mara Miniati. Chimica. Nuncius, 7(2):359–360, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539192x01469.

Abbri:1992:BRH

[961] Ferdinando Abbri. Book review: Hans-Werner Schütt, Eilhard Mitscherlich. Baumeister am Fundament der Chemie, München, R. Oldenbourg Verlag, 1992, 198 pp. (≪Deutsches Museum. Abhandlungen und Berichte, Neue Folge≫, Band 8). Nuncius, 7(2):360–361, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01478.

Caroti:1992:MR

[962] Stefano Caroti. Medioevo e rinascimento. Nuncius, 7(2):361–362, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01487.

Miniati:1992:BRLb

[963] Mara Miniati. Book review: El Legado Cientifico Andalusi. Museo Arqueologico Nacional, Madrid Abril Junio, 1992, s.n.t., 340 pp., fig. Nuncius, 7(2):362, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x01496.

Caroti:1992:BRDb

[964] Stefano Caroti. Book review: D. S. Chambers, A Renaissance Cardinal and His Wordly Goods: the Will and Inventory of Francesco Gonzaga (1444–1483), London, the Warburg Institute, University of London, 1992, 257 pp., app., ill., bibl., ind. (≪Warburg Institute Surveys and Texts≫, XX). Nuncius, 7(2):363, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01504.

Abbri:1992:BRJd

[965] Ferdinando Abbri. Book review: José A. García-Diego, Alexander G. Keller, Giovanni Francesco Sitoni. Ingeniero rinacentista al servicio de la Corona de España, Madrid, Fundacion Juanelo Turriano — Editorial Castelia, 1990, 231 pp. Nuncius, 7(2):363–364, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01513.

Caroti:1992:BRC

[966] Stefano Caroti. Book review: Christoph Meinel, Die Bibliothek des Joachim Jungius. Ein Beitrag zur Historia litteraria der frühen Neuzeit, Göttingen, Vandenhoeck & Ruprecht, 1992, 224 pp., app., ill., ind. (≪Veröffentlichung der Joachim Jungius Gesellschaft der Wissenschaften, Hamburg≫, 67). Nuncius, 7(2):364–365, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01522.

Abbri:1992:SXX

[967] Ferdinando Abbri. Secoli XVI–XX. Nuncius, 7(2):365–368, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01531.

Abbri:1992:BRSb

[968] Ferdinando Abbri. Book review: Stillman Drake, Galileo: Pioneer Scientist, Toronto-Buffalo-London, University of Toronto Press, 1990, xviii + 261 pp. Nuncius, 7(2):368–369, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01540.

Parigi:1992:BRL

[969] Silvia Parigi. Book review: Luigi Neri, George Berkeley. Filosofia e critica dei linguaggi scientifici, Presentazione di Antonio Santucci, Bologna, Clueb, 1991, xiv + 241 pp. Nuncius, 7(2):369–370, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01559.

Miniati:1992:BRB

[970] Mara Miniati. Book review: Bruno Corti, Archeologia Industriale, Brescia, Icmai-Il Coltello di Delfo, 1991, 298 pp., fig., L. 120.000. Nuncius, 7(2):370–372, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539192x01568.

Casati:1992:BRS

[971] Stefano Casati. Book review: Storia dell’industria elettrica in Italia. 1. Le origini. 1882–1914, (Italian) [History of the electricity industry in Italy. 1. Origins. 1882–1914] a cura di Giorgio Mori, Roma/Bari, Laterza, 1992, 1022 pp. Nuncius, 7(2):372–373, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01577.

Abbri:1992:BRG

[972] Ferdinando Abbri. Book review: Gunnar Broberg och Mattias Tydén, Oönskade i folkhemmet. Rashygien och sterilisering i Sverige, (Swedish) [Unwanted people in the home. Eugenics and sterilization in Sweden], Värnamo, Gidlunds Bokförlag, 1991, 219 pp. Nuncius, 7(2):373–376, ???? 1992. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539192x01586.

Miniati:1992:ADM 7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539192x01595.

Roger:1993:HSP

[974] Jacques Roger. History of science: Problems and practices. History of science (S), history of mentalities, micro-history. Nuncius, 8(1):3–26, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00019.

Barcaro:1993:UDB

[975] Umberto Barcaro. Un’analisi della ≪Bizzarria≫ di Galileo. Nuncius, 8(1):27–39, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00028.

Maffioli:1993:SFD

[976] Cesare S. Maffioli. Sul filo delle acque. Nuncius, 8(1):41–74, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00037.

Beretta:1993:CSL

[977] Marco Beretta. Chemists in the storm: Lavoisier, Priestley and the French Revolution. Nuncius, 8(1):75–104, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539183x00046.

Tazzioli:1993:FFN

[978] Rossana Tazzioli. Fisica e ≪Filosofia Naturale≫ in Riemann. Nuncius, 8(1):105–120, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00055.

Maiocchi:1993:UDT

[979] Roberto Maiocchi. Un’infanzia difficile: la teoria della dissociazione elettrolitica nel XIX secolo. Nuncius, 8(1):121–167, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00064.

Franco:1993:OGF

[980] Loredana Franco. The old globes in Friuli–Venezia Giulia (Italy). Nuncius, 8(1):169–195, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00073.

Brenni:1993:MPA

[981] Paolo Brenni. Il meteorografo di padre Angelo Secchi. Nuncius, 8(1): 197–247, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00082.

Palladino:1993:LGP

[982] Franco Palladino. Le lettere di Giuseppe Peano nella corrispondenza di Ernesto Cesàro. (Italian) [The letters of Giuseppe Peano in the correspondence of Ernesto Cesàro]. Nuncius, 8(1):249–281, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00091.

Ziwet:1993:A

[983] Alexander Ziwet and Giuseppe Peano. Appendice I. Nuncius, 8(1): 282–283, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00109.

Macfarlane:1993:AI

[984] Alexander Macfarlane and Giuseppe Peano. Appendice II. Nuncius, 8 (1):284–285, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00118.

Bertini:1993:CEF

[985] Silvia Bertini. A catalogue of English, French, Italian and German scientific instruments makers from 1750 to 1900. Nuncius, 8(1):287–294, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00127.

Gori:1993:BDF

[986] Guido Gori and Massimo Misiti. La biblioteca della fondazione scienza e tecnica. Nuncius, 8(1):295–298, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00136.

Abbri:1993:YC

[987] Ferdinando Abbri and Antonello La Vergata. Yvette Conry (1930–1992). Nuncius, 8(1):299–304, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00145.

Greco:1993:EOC

[988] Vincenzo Greco, Giuseppe Molesini, and Franco Quercioli. Esame ottico dei cannocchiali di Galileo. Nuncius, 8(1):305–311, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00154.

Bucciantini:1993:BFC

[989] Massimo Bucciantini, Anna C. Citernesi, and Fabio Guidi. Bce un formato comune per lo scambio di dati tra bibliografie di storia della scienza. Nuncius, 8(1):313–320, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00163.

Caroti:1993:BRW

[990] Stefano Caroti. Book review: Wesley M. Stevens (Ed.), Bibliographical Access to Medieval and Renaissance Manuscripts: A Survey of Computerized Data Bases and Information Services, New York-London, the Haworth Press Inc., 1992, 196 pp., ind. Nuncius, 8(1):321–322, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00172.

Caroti:1993:BRC

[991] Stefano Caroti. Book review: Catalogodi manoscritti filosofici nelle biblioteche italiane. Volume 6. Atri, Bergamo, Cosenza, Milano, Perugia, Pistoia, Roma, Siena, Firenze, Olschki, 1992, 381 pp., ind. dei mss., degli incipit, dei nomi e delle opere anonime (≪Unione Accademica Nazionale. Corpus philosophorum Medii Aevi. Subsidia, VII≫ ). Nuncius, 8(1):322– 323, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00181.

Miniati:1993:BRM

[992] Mara Miniati. Book review: Mary Holbrook, Science preserved. A directory of scientific instruments in collections in the United Kingdon and Eire. London, Hmso, 1992, 271 pp., fig. Nuncius, 8(1):323–324, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00190.

Segala:1993:BRD

[993] Marco Segala. Book review: David C. Lindberg, The beginning of Western science. The European scientific tradition in philosophical, religious, and institutional context, 600 B.C. to A.D. 450, Chicago, The University of Chicago Press, 1992, xviii + 455 pp., ill., ISBN 0-226-48231-6. Nuncius, 8(1):324–327, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00208.

Caroti:1993:BRA

[994] Stefano Caroti. Book review: Aufstiegund Niedergang der Römischen Welt (Anr W). Geschichte und Kultur Roms im Spiegelder neueren Forschung. Teil II: Principat, Bd. 36, 5: Philosophie (Einselne Autoren; Doxographica); Bd. 36, 6: Philosophie (Doxographica [Forts.]), hrsg. von W. Haase, Berlin-New York, Walter De Gruyter, 1992, 3249-3792 e 3793-4411 pp., indici. Nuncius, 8(1):327–328, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00217.

Abbri:1993:BRMa

[995] Ferdinando Abbri. Book review: Michela Pereira, l’Oro dei Filosofi. Saggio sulle idee di un alchimista del Trecento, Spoleto, Centro Italiano di Studi sull’Alto Medioevo, 1992, 265 pp. (≪Biblioteca di Medioevo Latino≫, 7). Nuncius, 8(1):328–330, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539183x00226.

Caroti:1993:BRT

[996] Stefano Caroti. Book review: Tullio Gregory, Mundana sapientia. Forme di conoscenza nella cultura medievale, Roma, Edizioni di storia e letteratura, 1992, 480 pp., app. ind., (≪Storia e letteratura. Raccolta di studi e testi, 181≫ ). Nuncius, 8(1):330–334, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00235.

Pereira:1993:BRW

[997] Michela Pereira. Book review: William R. Newman (ed.), They ≪Summa Perfectionis≫ of Pseudo-Geber. A Critical Edition, Translation and Study, E. J. Brill, Leiden-New York–Københaven–Köln, 1991, vi + 785 pp., ISBN 90-04-09464-4. Nuncius, 8(1):334–341, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00244.

Miniati:1993:BRSa

[998] Mara Miniati. Book review: Steven A. Lloyd, Ivory Diptych Sundials 1570–1750, with Introductions by Penelope Gouk and A. J. Turner, Harvard, the Collection of Historical Scientific Instruments, Harvard University, 1992, 169 pp., fig. Nuncius, 8(1):341–342, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00253.

Mocchi:1993:BRI

[999] Giuliana Mocchi. Book review: Ivana Gambaro, Astronomia e tecniche di ricerca nelle lettere di G. B. Riccioli ad A. Kircher, Quaderni del Centro di studio sulla storia della tecnica del Consiglio Nazionale delle Ricerche, 15, Genova, 1989. Nuncius, 8(1):342–345, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00262.

Vista:1993:BRG

[1000] Michele Vista. Book review: Giovanni Del Turco, Epulario e segreti varî. Trattati di cucina toscana nella Firenze seicentesca (1602–1636), a cura di A. Evangelista, Arnaldo Forni Editore, 1992, lxvi + 170 pp., fig. Nuncius, 8(1):346–347, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00271.

Meli:1993:BRG

[1001] Domenico Bertoloni Meli. Book review: G. W. Leibniz, Saemtliche Schriften und Briefe, Siebente Reihe, erster Band, mathematische Schriften, edited by E. Knobloch and Walter S. Contro, Berlin, Akademie Verlag, 1990, xl + 954 pp. — Dritte Reihe, dritter Band, mathematischer naturwissenschaftlicher und technischer Briefwechsel, edited by H. Breger, Berlin, Akademie Verlag, 1991, Lx + 896 pp. Nuncius, 8(1):347– 351, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00280.

Abbri:1993:BRB

[1002] Ferdinando Abbri. Book review: Betty Jo Teeter Dobbs, the Janus faces of genius. The role of alchemy in Newton’s thought, Cambridge, Cambridge University Press 1991 (p. 1992), xii + 359 pp. Nuncius, 8 (1):351–355, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00299.

Toscano:1993:BRF

[1003] Anna Toscano. Book review: Franco Crispini, Recuperi. Brevi tratti di storia intellettuale, Cosenza, Edizioni Brenner 1992, vi + 169 pp. Nuncius, 8(1):355–357, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00307.

Bertoletti:1993:BRG

[1004] Stefano Fabbri Bertoletti. Book review: Gunter Mann — Franz Dumont (Hrsgs.) Die Natur des Menschen. Probleme der physichen Anthropologie und Rassenkunde (1750–1850), Stuttgart-New York, G. Fischer Verlag 1990, 457 pp., (Soemmerring-Forschungen, VI). Nuncius, 8(1):357–360, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00316.

Abbri:1993:BRE

[1005] Ferdinando Abbri. Book review: Evan M. Melhado & Tore Frängsmyr (eds.), Enlightenment Science in the Romantic Era. The Chemistry of Berzelius and its Cultural Setting, Cambridge, Cambridge University Press 1992, xiv + 246 pp., ill. (≪Uppsala Studies in History of Science≫, 10). Nuncius, 8(1):360–363, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00325.

Beretta:1993:BRC

[1006] Marco Beretta. Book review: The Correspondence of Michael Faraday, Vol. 1, 1811–1831, a cura di Frank A. J. L. James, London, Institution of Electrical Engineers, 1991, 673 pp. Ill. Nuncius, 8(1):363–364, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00334.

Segala:1993:BRAa

[1007] Marco Segala. Book review: Abraham Pais, Niels Bohr’s Times, in physics, philosophy, and polity, Oxford, Clarendon Press, 1991, xvii + 565 pp. ill. (ISBN 0-19-852049-2). Nuncius, 8(1):364–367, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00343.

Maiocchi:1993:BRD

[1008] Roberto Maiocchi. Book review: David Favrholdt, Niels Bohr’s philosophical background, Copenhagen, the Royal Danish Academy of Sciences and Letters, 1992, 147 pp. Nuncius, 8(1):367–370, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00352.

Miniati:1993:SB

[1009] Mara Miniati. Strumenti bibliografici. Nuncius, 8(1):371, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00361.

Miniati:1993:S

[1010] Mara Miniati. Strumentaria. Nuncius, 8(1):371–372, ???? 1993. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539183x00370.

Miniati:1993:BRE

[1011] Mara Miniati. Book review: Ernesto Milano, Giulia Luppi, Libra. La bilancia nei codici estensi. Immagini e modelli di strumenti di pesatura, Modena, Biblioteca Estense, Museo della Bilancia, Il Bulino edizioni d’arte, 1991, 57 pp., fig. Nuncius, 8(1):372–373, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00389.

Miniati:1993:BRKa

[1012] Mara Miniati. Book review: Klaus Schillinger, Joachim Schardin, Vermisste Instrumente und Uhren des mathematisch-physikalischen Salons Dresden, Leipzig, E. A. Seemann Kunstverlagsgesellschaft mbH, 1992, 44 pp., fig. Nuncius, 8(1):373, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00398.

Miniati:1993:BRG

[1013] Mara Miniati. Book review: G. Paltrinieri, Le meridiane: da S. Petronio a S. Giuseppe di Brescia, ≪L’Astrofilo. Bollettino dell’Unione Astrofili Bresciani≫, Brescia, 18–1991, 32 pp., fig. Nuncius, 8(1):373–374, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00406.

Miniati:1993:BRBa

[1014] Mara Miniati. Book review: E. Borchi, R. Macii, Sul telescopio a riflessione, (Italian) [On the reflecting telescope], Firenze, Osservatorio Ximeniano, 1992, 126, pp., fig. Nuncius, 8(1):374, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00415.

Miniati:1993:BRLa

[1015] Mara Miniati. Book review: Lorella Grossi, Collezione Albertazzi-Dazzani. Biblioteca Comunale di Castenaso Casa Bondi 11 Settembre/31 Ottobre 1992, Campogalliano, Museo della Bilancia, 1992, 63 pp., fig. (≪I libri delle bilance≫ 1). Nuncius, 8(1):374–375, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00424.

Miniati:1993:BRF

[1016] Mara Miniati. Book review: Franco Foresta Martin, Scienza in città. Guida ai luoghi e ai musei scientificidi Roma, Milano, Amministrazione Provinciale di Roma, Electa, 1992, 151 pp., fig. Nuncius, 8(1):375, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00433.

Abbri:1993:SDT

[1017] Ferdinando Abbri. Storia della tecnica. Nuncius, 8(1):375–376, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00442.

Miniati:1993:BRSb

[1018] Mara Miniati. Book review: ≪Quaderni di storia della tecnologia≫, a cura di Vittorio Marchis, Torino, Editrice Levrotto & Bella, 1991. Nuncius, 8(1):376–377, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00451.

Abbri:1993:BRR

[1019] Ferdinando Abbri. Book review: Reinhold R. Angersteins resor genom Ungern och Österrike 1750, sammanställda och kommenterade av Marianne Fornander, (Swedish) [Reinhold R. Angerstein’s travels through Hungary and Austria 1750, collected and commented by Marianne Fornander], Stockholm, Jernkontorets Bergshistoria utskott 1992 (≪ serie H≫, 50) 290 pp. Nuncius, 8(1):377–378, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00460.

Miniati:1993:BRLb

[1020] Mara Miniati. Book review: Luisa Dolza, Vittorio Marchis, Michelangelo Vasta, I privilegi industriali come specchio dell’innovazione nel Piemonte Preunitario 1814–1855, (Italian) [Privileges as a mirror of industrial innovation in the pre-unification Piedmont 1814–1855]. Torino, la Rosa, 1992, 143, pp., L. 22.000 (≪Innovazione e storia≫ ). Nuncius, 8(1): 378, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00479.

Abbri:1993:IS

[1021] Ferdinando Abbri. Istituzioni scientifiche. Nuncius, 8(1):378–380, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00488.

Segala:1993:SDI

[1022] Marco Segala. Storia delle idee. Nuncius, 8(1):380–381, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00497.

Abbri:1993:BRSa

[1023] Ferdinando Abbri. Book review: Sverker SÖRlin, Naturkontraktet. Om naturumgängets idéhistoria, Stockholm, Carlssons Bokförlag, 1991, 293 pp. Nuncius, 8(1):381–383, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00505.

Malacrida:1993:BM

[1024] Lorenzo Marri Malacrida. Biologia e medicina. Nuncius, 8(1):383–384, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00514.

Miniati:1993:QIS

[1025] Mara Miniati. ≪quaderni internazionali di storia della medicina e della sanità≫, ciso toscano, anno I, n. 1, 1992. Nuncius, 8(1):384–385, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00523.

Lippi:1993:BRL

[1026] Donatella Lippi. Book review: Luciano Sterpellone, La Medicina nei francobolli, (Italian) [Medicine in stamps], Saronno, Ciba-Geigy Ed., 1992, 323 pp. Nuncius, 8(1):385–386, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539183x00532.

Lippi:1993:BRAa

[1027] Donatella Lippi. Book review: Alfonso Pluchinotta, Storia illustrata della proctologia, Saronno, Ciba-Geigy Ed., 1992, 321 pp. Nuncius, 8 (1):386–387, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00541.

Lippi:1993:BRE

[1028] Donatella Lippi. Book review: Emiliano Panconesi e Lorenzo Marri Malacrida, Lorenzo il Magnifico in salute e in malattia. Firenze, A. Bruschi ed., 1992, L. 20.000. Nuncius, 8(1):387, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00550.

Malacrida:1993:BRM

[1029] Lorenzo Marri Malacrida. Book review: Maria Luisa Altieri Biagi, Clemente Mazzotta, Angela Chiantera, Paola Altieri (a cura di), Medicina per le donne nel Cinquecento, Torino, Utet, 1992, 124 pp., ed. fuori commercio, fig. Nuncius, 8(1):388, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00569.

Lippi:1993:BRVa

[1030] Donatella Lippi. Book review: Vittorio A. Sironi, Le officine della salute. Storia del farmaco e della sua industria in Italia. Bari, Laterza, 1992, 274 pp. (L. 45.000). Nuncius, 8(1):388–390, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00578.

Abbri:1993:C

[1031] Ferdinando Abbri. Chimica. Nuncius, 8(1):390–392, ???? 1993. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539183x00587.

Abbri:1993:SXX

[1032] Ferdinando Abbri. Secoli XIV–XIX. Nuncius, 8(1):392–393, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00596.

Miniati:1993:BRU

[1033] Mara Miniati. Book review: Ugo Baldini, Legem impone subactis. Studi su filosofia e scienza dei Gesuiti in Italia. Roma, Bulzoni editore, 1992, 601 pp., L. 78.000 (≪Università degli Studi G. D’Annunzio di Chieti, Collana dell’Istituto di Filosofia≫, n.s. 3). Nuncius, 8(1):393–394, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00604.

Segala:1993:BRF

[1034] Marco Segala. Book review: F. W. J. Schelling, Primo abbozzo di un sistema di filosofia della natura, a cura di Gigliola Grazi, Roma, Cadmo editore, 1989, 382 pp. Nuncius, 8(1):394, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00613.

Anonymous:1993:ADM

[1035] Anonymous. Attività dell’Istituto e Museo di Storia Della Scienza di Firenze. Nuncius, 8(1):395–408, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00622.

Garin:1993:GSD

[1036] Eugenio Garin. Galileo: Gli scandali della nuova ’filosofia’. Nuncius, 8(2):v–430, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00631.

Baroncelli:1993:IAD

[1037] Giovanna Baroncelli. Intorno all’invenzione della spirale geometrica. Una lettera inedita di Torricelli a Michelangelo Ricci. (Italian) [On the invention of the geometric spiral. An unpublished letter of Torricelli to Michelangelo Ricci]. Nuncius, 8(2):14–606, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00721.

Woltjer:1993:LU

[1038] Lodewijk Woltjer. The limits of the universe. Nuncius, 8(2):431–441, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00640.

Sarti:1993:GMP

[1039] Carlo Sarti. Giuseppe Monti and palaeontology in Eighteenth Century Bologna. Nuncius, 8(2):443–455, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00659.

Marin:1993:TDP

[1040] Brigitte Marin. Les traits d’hygiène publique (1784–1797) de Filippo Baldini, médecin à la Cour De Naples: Culture médicale et service du roi. Nuncius, 8(2):457–486, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00668.

Conry:1993:CPT

[1041] Yvette Conry. Comment a-t-on pu tre no-Lamarckien en France (1843– 1930)? Nuncius, 8(2):487–520, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00677.

Muller:1993:KVD

[1042] Gerhard H. Müller and Klaus Wenig. Kants Verhaltnis Zur Descendenztheorie: Anton Dohrns Probevorlesung 1868 in Jena. (German) [Kant’s relationship to the theory of descent: Anton Dohrn’s 1868 trial lecture in Jena]. Nuncius, 8(2):521–528, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00686.

Dohrn:1993:KVD

[1043] Anton Dohrn. Kants Verhältnis Zur Descendenztheorie. (German) [Kant’s relationship to the theory of descent]. Nuncius, 8(2):529–553, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00695.

Scaglia:1993:IPM

[1044] Gustina Scaglia. Instruments perfected for measurements of man and statues illustrated in Leon Battista Alberti’s De Statua. Nuncius, 8(2): 555–596, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00703.

Bedini:1993:MQB

[1045] Silvio A. Bedini. A marine quadrant by Bernardo Facini. Nuncius, 8 (2):597–599, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00712.

Arrigoni:1993:IF

[1046] Tiziano Arrigoni. Istituzioni e fonti. Nuncius, 8(2):607–610, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00730.

Smith:1993:MHI

[1047] A. Mark Smith and Bernard R. Goldstein. The Medieval Hebrew and Italian versions of Ibn Mu’Dh’s on twilight and the rising of clouds. Nuncius, 8(2):611–613, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00749.

Goldstein:1993:IHT

[1048] Bernard R. Goldstein. II. The Hebrew text. Nuncius, 8(2):614–626, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00758.

Smith:1993:IIT

[1049] A. Mark Smith. III. The Italian text. Nuncius, 8(2):628–630, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00767.

Anonymous:1993:TDC

[1050] Anonymous. [tractato deli crepuscoli e dele ascensioni dele nuvole]. Nuncius, 8(2):631–643, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00776.

Vlahakis:1993:NPN

[1051] George N. Vlahakis. A note on the penetration of Newtonian physics in Greece. Nuncius, 8(2):645–656, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00785.

Parigi:1993:BRT

[1052] Silvia Parigi. Book review: Tomaso Poggio, l’occhio e il cervello. Che cosa significa Vedere. Roma-Napoli, Theoria, 1991, 124 pp. Nuncius, 8 (2):657–669, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00794.

Segala:1993:CNR

[1053] Marco Segala. Contro Newton: Retorica, scienza e metafisica nella Farbenlehre di Goethe. (Italian) [Against Newton: Rhetoric, science and metaphysics in Goethe’s science of color]. Nuncius, 8(2):671–688, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00802.

Berni:1993:SPP

[1054] M. Berni, P. Brenni, F. Guidi, M. Miniati, G. di Pasquale, F. Principe, A. Giatti, and F. Vannozzi. Sic, un programma per la catalogazione degli strumenti scientifici di interesse storico. Nuncius, 8(2):689–696, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00811.

Abbri:1993:BRMb

[1055] Ferdinando Abbri. Book review: Marco Beretta, A History of Non-Printed Science. A Select Catalogue of the Waller Collection, Uppsala, Uppsala University Library, 1993, 200 pp., ill. (Acta Bibliothecae R. Universitatis Upsaliensis, vol. XXXI). Nuncius, 8(2):699–701, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00820.

Miniati:1993:BRN

[1056] Mara Miniati. Book review: Nelli-Elena Vanzan Marchini (a cura di), Dalla Scienza medica alla pratica dei corpi. Fonti e manoscritti marciani per la Storia della Sanit, Padova, Neri Pozza, Editore, 1993, 192 pp., fig. (Fonti per la Storia della Sanit, 1). Nuncius, 8(2):701–703, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00839.

Maltese:1993:BRG

[1057] Giulio Maltese. Book review: Gerard l’E. Turner the Great Age of the Miscrope, Bristol, Adam Hilger, 1989, 379 pp. Nuncius, 8(2):703–705, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00848.

Monti:1993:BRA

[1058] Maria Teresa Monti. Book review: Änne Bäumer, Geschichte der Biologie, Band 1. Biologie der Antike bis zur Renaissance, Band 2. Zoologie der Renaissance, (German) [History of Biology. Volume 1. Biology of Antiquity through the Renaissance. Volume 2. Zoology of the Renaissance] Frankfurt a.M., Peter Lang, 1991. Nuncius, 8(2):705–709, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00857.

Casanovas:1993:BRI

[1059] Juan Casanovas. Book review: I. Iannaccone, Misurare il cielo: l’antica astronomia cinese, Napoli, 1991, 218 pp. Nuncius, 8(2):709–710, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00866.

Beretta:1993:BRT

[1060] Marco Beretta. Book review: Les techniques minières de l’antiquité au XVIIIe siécle, (French) [Mining techniques from Antiquity to the 18th Century], Paris, édition du Comité des Travaux Historiques et Scientifiques 1992, 601 pp., ill. Nuncius, 8(2):710–712, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00875.

Plastina:1993:BRF

[1061] Sandra Plastina. Book review: Filosofia scienza e astrologia nel tre cento europeo, a cura di Graziella Federici Vescovini e Francesco Barocelli, Padova, Il Poligrafo, 1992, 222 pp. Nuncius, 8(2):712–717, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00884.

Miniati:1993:BRBb

[1062] Mara Miniati. Book review: Benito Brunelli, Cinzia Bucchioni, Maria Pia Torricelli, Il patrimonio librario antico della Biblioteca d’Ingegneria, Bologna, Biblioteca Centrale G. P. Dore, [1992], 210 pp., fig. Nuncius, 8(2):717–718, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00893.

Clericuzio:1993:BRB

[1063] Antonio Clericuzio. Book review: Bruce T. Moran (ed.), Patronage and Institutions. Science, Technology and Medicine at the European Court, 1500–1750. Woodbridge, Suffolk, the Boydell Press, 1991, 261 pp. ISBN 0-85115-285-6. Nuncius, 8(2):718–720, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00901.

Toscano:1993:BRG

[1064] Anna Toscano. Book review: Galileo Galilei, Sidereus Nuncius, a cura di Andrea Battistini, traduzione italiana di Maria Timpanaro Cardini, Venezia, Letteratura universale Marsilio, 1993, 247 pp. (Esperia, Collana dei Classici Italiani). Nuncius, 8(2):720–722, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00910.

Mocchi:1993:BRG

[1065] Giuliana Mocchi. Book review: G. S. Roussau, La medicina e le Muse, (Italian) [Medicine and the Muses], a cura di Antonello la Vergata, Firenze, la Nuova Italia, 1993, pp. v + 165. Nuncius, 8(2):722–728, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00929.

Beretta:1993:BRW

[1066] Marco Beretta. Book review: Wolf Lepenies, Natura e scrittura. Autori e scienziati nel XVIII secolo, Bologna, Il Mulino, 1992, 131 pp. Nuncius, 8(2):728–730, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00938.

Ciardi:1993:BRR

[1067] Marco Ciardi. Book review: R. W. Home, Electricity and Experimental Physics in 18th-Century Europe, London, Variorum, 1992, xii + 396 pp. Nuncius, 8(2):730–733, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x00947.

Segala:1993:BRWa

[1068] Marco Segala. Book review: Walter Bernardi, I fluidi della vita. Alle origini della controversia sull’elet-tricit animale, Firenze, Olschki, 1992, 388 pp. (Biblioteca di Storia della Scienza, 33). Nuncius, 8(2):733–736, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00956.

Vaccari:1993:BRB

[1069] Ezio Vaccari. Book review: Bernhard Fritscher, Vulkanismusstreit und Geochemie. Die Bedeutung der Chemie und des Experiments in der Vulkanismus-Neptunismus-Kontroverse, Stuttgart, Franz Steiner Verlag, 1991, 346 pp. (Boethius Texte und Abhandlungen zur Geschichte der Mathematik und der Naturwissenschaften, Band XXV). Nuncius, 8(2): 737–739, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00965.

Abbri:1993:BRD

[1070] Ferdinando Abbri. Book review: David Knight, Humphry Davy. Science Power, Oxford, Blackwell, 1992, xiv + 218 pp. (Blackwell Science Biographies). Nuncius, 8(2):739–742, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x00974.

Nicolini:1993:BRH

[1071] Nicoletta Nicolini. Book review: History of Technology, vol. 13, ed. Graham Hollister-Short e Frank A. J. L. James, Mansell, London e New York, 1991, xiii + 242 p. Nuncius, 8(2):742–745, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x00983.

Miniati:1993:BRJ

[1072] Mara Miniati. Book review: Jim Benett, Robert Brain, Kate Bycroft et al., Empires of Physics. A Guide to the Exhibition. Cambridge, Whipple Museum of the History of Science, 1993, 110 pp., fig. Nuncius, 8(2): 745–748, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x00992.

Pogliano:1993:BRD

[1073] Claudio Pogliano. Book review: Darwinisme et socit, sous la direction de Patrick Tort, Paris, Presses Universitaires de France, 1992, 690 pp. Nuncius, 8(2):748–750, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x01009.

Muller:1993:BRI

[1074] Gerhard H. Müller. Book review: Ilse Seibold, Der Weg zur Biogeologie. Johannes Walther 1860–1937. Ein Forscherleben in Wandel der deutschen Universität, Berlin, Heidelberg, Springer, 1992, viii + 196 pp., ISBN 3-540-55341-X. Nuncius, 8(2):750–752, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01018.

Pogliano:1993:BRL

[1075] Claudio Pogliano. Book review: Luciano Mecacci, Storia della psicologia del Novecento, Bari, Laterza, 1992, xi + 554 pp., L. 48.000. Nuncius, 8 (2):752–754, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x01027.

Segala:1993:BRKa

[1076] Marco Segala. Book review: Kenneth Jinghwa Hs, Challenger at sea: a ship that revolutionized earth science, Princeton, Princeton University Press, 1992, xxxi + 417 pp., ill. Nuncius, 8(2):754–757, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01036.

Caroti:1993:BRR

[1077] Stefano Caroti. Book review: Le raccolte della Colombaria. I. Incunabuli. Con un saggio sulla Libreria Pandolfini, a cura di Enrico Spagnesi, Firenze, Olschki 1993 (Accademia Toscana di Scienze e Lettere ‘La Colombaria’. ’Studi’, CXX–VII), pp. 315, ill., ind. Nuncius, 8(2):759– 760, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01045.

Abbri:1993:BROa

[1078] Ferdinando Abbri. Book review: Ove Hagelin (ed.), Rare and Important Medical Books in the Library of the Karolinska Institute. An Illustrated and Annotated Catalogue, Stockholm, Karolinska Institutes Bibliotek, 1992, 212 pp. Nuncius, 8(2):760, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x01054.

Segala:1993:BRKb

[1079] Marco Segala. Book review: Kai Torsten Kanz, Kielmeyer-Bibliographie. Verzeichnis der Literatur von und ber den Naturforscher Carl Friedrich Kielmeyer (1765–1844), Stuttgart, Verlag für Geschichte der Naturwissenschaften und der Technik, 1991, 161 pp. (Quellen der Wissenschaftsgeschichte, 1). Nuncius, 8(2):760–761, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539183x01063.

Miniati:1993:BRCa

[1080] Mara Miniati. Book review: Les Collections: fables et programmes sous la direction de Jacques Guillerme. Seyssel, Éditions Champ Vallon, 1993, 364 pp. (Amphion, 4, Études d’histoire des techniques). Nuncius, 8(2): 761–763, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x01072.

Abbri:1993:BROb

[1081] Ferdinando Abbri. Book review: Olaf Pedersen, Lovers of Learning. A History of the Royal Danish Academy of Sciences and Letters 1742– 1992, Copenhagen, Munksgaard, 1992, 348 pp. Nuncius, 8(2):763–765, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01081.

Miniati:1993:BRCb

[1082] Mara Miniati. Book review: La Collina di Urania. Il Museo storico dell’Osservatorio Astronomico di Capodimonte. (Italian) [The Hill of Urania. The Historical Museum of the Astronomical Observatory of Capodimonte]. Napoli, Elio de Rosa editore, 1992, 111 pp., fig. Nuncius, 8(2):765, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x01090.

Miniati:1993:BRKb

[1083] Mara Miniati. Book review: Klaus Schillinger, Solare Brenngerte. Katalog. Dresden, Staatlicher Mathematisch-Physikalischer Salon, 1992, 28 pp., fig. Nuncius, 8(2):765–766, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x01108.

Miniati:1993:BRLc

[1084] Mara Miniati. Book review: Laura Franchini, Strumenti antichi di fisica di scuole napoletane. Old Physics Instruments from Neapolitan Schools. Napoli, Litografia Pesole s.n.c., 1993, 46 pp., fig. Nuncius, 8(2):766, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01117.

Brenni:1993:BRE

[1085] Paolo Brenni. Book review: Edwin Banfield, the Italian Influence on English Barometers from 1780, Baros Book, Trowbridge (UK), 1993, 160 pp., ISBN 0-948382-07-4, £6.95. Nuncius, 8(2):766–768, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01126.

Segala:1993:BRAb

[1086] Marco Segala. Book review: Alexander Jacob, De naturae natura. A study of idealistic conceptions of nature and the unconscious, Stuttgart, Steiner, 1992, 181 pp. (Kosmosophie. Forschungen und Texte zur Geschichte des Weltbildes der Naturphilosophie, der Mystik und des Spiritualismus vom Sptmittelalter bis zur Romantik, Band VII). Nuncius, 8 (2):768–769, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x01135.

Abbri:1993:BRMc

[1087] Ferdinando Abbri. Book review: Michael Hunter and David Wootton (eds.), Atheism from the Reformation to the Enlightenment, Oxford, Clarendon Press, 1992, viii + 307 pp. Nuncius, 8(2):769–771, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01144.

Casati:1993:BRD

[1088] Stefano Casati. Book review: Dario Camuffo, Clima e uomo, Milano, Garzanti, 1990, 207 pp. Nuncius, 8(2):772, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539183x01153.

Segala:1993:BRWb

[1089] Marco Segala. Book review: William Dunham, Journey through genius. the great theorems of mathematics, New York, Wiley, 1990, xiii + 300 pp., ill. Nuncius, 8(2):772–773, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x01162.

Abbri:1993:BRMd

[1090] Ferdinando Abbri. Book review: Michelle Goupil (ed.), Lavoisier et la Révolution chimique. Actes du Colloque tenu l’occasion du bicentenaire de la publication du Trait élémentaire de chimie 1789, Paris, Sabix — École polytechnique, 1992, viii + 372 pp., ill. Nuncius, 8(2):773–776, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01171.

Abbri:1993:BRGa

[1091] Ferdinando Abbri. Book review: Gianni Michelon (a cura di), Atti del Iv Convegno Nazionale di Storia e Fondamenti della Chimica, Roma, Accademia Nazionale delle Scienze detta dei Xl, 1993 (Rendiconti, serie V, XVI, parte II, 1992), pp. 103-529. Nuncius, 8(2):776–777, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01180.

Lippi:1993:BRS

[1092] Donatella Lippi. Book review: Sherwin B. Nuland, I figli di Ippocrate. Storia della medicina dagli antichi greci ai trapianti d’organo, (Italian) [The sons of Hippocrates. History of medicine from the ancient Greeks to organ transplants]. Milano, Mondadori, 1992, pp. 498, L. 42.000. Nuncius, 8(2):777–778, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x01199.

Malacrida:1993:BRF

[1093] Lorenzo Marri Malacrida. Book review: Franco Voltaggio, l’arte della guarigione nelle culture umane, Torino, Bollati Boringhieri, 1992, 741 pp. Nuncius, 8(2):778, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x01207.

Lippi:1993:BRAb

[1094] Donatella Lippi. Book review: Aa.Vv., Ars Obstetricia Bononiensis (Catalogue and videotape), Bologna, Clueb, 1988. Nuncius, 8(2):779, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01216.

Beretta:1993:BRP

[1095] Marco Beretta. Book review: Peter Pomet, Der aufrichtige Materialist und Specerey-Hndler (Leipzig, 1717), VCH Verlagsgesellschaft mbh, Weinheim, 1987, 21 + 901 pp., ill. Nuncius, 8(2):780, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01225.

Lippi:1993:BRB

[1096] Donatella Lippi. Book review: E. Bruzzone, Storia e Medicina nella storiografia della peste, Genova, Ecig, 1987, pp. 189, L. 19.000. Nuncius, 8(2):780–781, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x01234.

Beretta:1993:BRJ

[1097] Marco Beretta. Book review: J. Worth Estes, Dictionary of Protopharmacology. Therapeutic Practices, 1700–1850, Science History Publications, Canton, Ma., 1990, xvii + 215 pp. Ill. Nuncius, 8(2):781–782, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01243.

Lippi:1993:BRVb

[1098] Donatella Lippi. Book review: V. Chiarugi, della pazzia in genere, e in specie. Trattato medico-analitico con una centuria di osservazioni., ristampa della prima edizione (Firenze, Carlieri 1793/94), a cura di V. Andreoli e L. Bonuzzi, Roma, Cic, 1991, pp. 353, L. 80.000. Nuncius, 8(2):782, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x01252.

Lippi:1993:BRC

[1099] Donatella Lippi. Book review: A. Cherubini — F. Vannozzi, Previdenza di malattia e malattie sociali dall’U-nit alla prima Guerra Mondiale, Roma, 1990, pp. 257. Ed. fuori commercio. Nuncius, 8(2):783, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01261.

Malacrida:1993:BRG

[1100] Lorenzo Marri Malacrida. Book review: Gordon M. Shepherd, Foundations of the neuron doctrine, 1991, New York, Oxford University Press, 338 pp., fig. Nuncius, 8(2):783–784, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x01270.

Lippi:1993:BRAc

[1101] Donatella Lippi. Book review: Aa. Vv., Civilt materiale e letteratura nel mondo antico, Atti del Seminario di studio, Macerata 28–29 giugno 1991, a cura di I. Mazzini, Assisi, 1992, pp. 173, L. 40.000. Nuncius, 8 (2):784–785, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x01289.

Caroti:1993:BRGa

[1102] Stefano Caroti. Book review: Glosses and Commentaries on Aristotelian Logical Texts. The Syriac, Arabic and Medieval Latin Traditions, ed. by Charles Burnett, London, the Warburg Institute University of London, 1993 (Warburg Institute Surveys and Texts, XXIII), pp. 192, bibl. ind. dei nomi, dei manoscritti e degli incipit. Nuncius, 8(2):785–786, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01298.

Caroti:1993:BRGb

[1103] Stefano Caroti. Book review: G. R. Evans, Philosophy and Theology in the Middle Ages, London and New York, Routledge, 1993, pp. 135, bibl., ind. Nuncius, 8(2):786–787, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x01306.

Caroti:1993:BRS

[1104] Stefano Caroti. Book review: Simo Knuuttila, Modalities in Medieval Philosophy, London and New York, Routledge, 1993 (Topics in Medieval Philosophy), pp. 236, bibl., ind. Nuncius, 8(2):787, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01315.

Caroti:1993:BRL

[1105] Stefano Caroti. Book review: L’insegnamento della logicaa Bolognanel XIV secolo, a cura di Dino Buzzetti, Maurizio Ferriani e Andrea Tabarroni, Bologna, Istituto per la storia dell’Università di Bologna, 1992 (Studi e memorie per la storia dell’Università di Bologna, N. S. VIII), pp. 645, app., ind. Nuncius, 8(2):788–789, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01324.

Abbri:1993:BRGb

[1106] Ferdinando Abbri. Book review: Gary Tomlinson, Music in Renaissance Magic. Toward a Historiography of Others, Chicago and London, The University of Chicago Press, 1993, xvi + 291 pp. Nuncius, 8(2):789–791, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01333.

Miniati:1993:BRLd

[1107] Mara Miniati. Book review: Il lasciapassare di Cesare Borgia a Vaprio d’Adda e il viaggio di Leonardo in Romagna, (Italian) [The pass of Cesare Borgia in Vaprio d’Adda and the journey of Leonardo in Romagna], Firenze, Giunti, 1993, 60 pp., fig. Nuncius, 8(2):791–792, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01342.

Abbri:1993:BRSb

[1108] Ferdinando Abbri. Book review: Stephen A. McKnight (ed.), Science, Pseudo-Science, and Utopianism in Early Modern Thought, Columbia and London, University of Missouri Press, 1992, xvi + 221 pp., ill. Nuncius, 8(2):792–794, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x01351.

Segala:1993:BRL

[1109] Marco Segala. Book review: Larry Stewart, the rise of public science. Rhetoric, technology, and natural philosophy in Newtonian Britain, 1660–1750, Cambridge, Cambridge University Press, 1992, xxxiv + 453 pp., ill. Nuncius, 8(2):794–795, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x01360.

Abbri:1993:BRC

[1110] Ferdinando Abbri. Book review: Christopher Mcintosh, the Rose Cross and the Age of Reason. Eighteenth-Century Rosicrucianism in Central Europe and its Relationship to the Enlightenment, Leiden, E. J. Brill, 1992, viii + 200 pp., ill., (Brill’s Studies in Intellectual History, 29). Nuncius, 8(2):795–797, ???? 1993. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539183x01379.

Contardi:1993:BRF

[1111] Simone Contardi. Book review: Il filosofo nel pozzo. La terra di Lazzaro Spallanzani tra storia e scienza (Italian) [The philosopher in the well. The land of Lazzaro Spallanzani between history and science], a cura di Paola Manzini, Firenze, Olschki, 1990, pp. 109. Nuncius, 8(2):797–798, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01388.

Brenni:1993:BRF

[1112] Paolo Brenni. Book review: Federico di Trocchio, Le bugie della Scienza. Perché e come gli scienziati imbrogliano, (Italian) [The Lies of Science. Why and how scientists cheat], Milano, A. Mondadori, 1993, pp. 364, ISBN 88-04-35111-X, L. 34.000. Nuncius, 8(2):798–802, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01397.

Segala:1993:BRB

[1113] Marco Segala. Book review: Byron Preiss — Robert Silverberg (Eds.), Dinosauri. Passato, presente, futuro, Milano, Mondadori, 1992, 351 pp., ill. Nuncius, 8(2):802–803, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x01405.

Papini:1993:BRL

[1114] Paola Papini. Book review: Livio Gratton, Origine ed evoluzione dell’universo. Dal big-bang alle galassie, (Italian) [Origin and evolution of the universe. From the Big Bang to galaxies]. Roma, La Nuova Italia scientifica, 1992, 189 pp. Nuncius, 8(2):803, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539183x01414.

Lippi:1993:BRAd

[1115] Donatella Lippi. Book review: Aa.Vv., Educare all’ascolto, a cura di Massimo Baldini, Brescia, La Scuola Ed., 1988, 217 pp., L. 18.000. Nuncius, 8(2):804, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539183x01423.

Miniati:1993:ADM

[1116] Mara Miniati. Attivita dell’Istituto e Museo di Storia della Scienza di Firenze. Nuncius, 8(2):805–837, ???? 1993. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539183x01432.

Rossi:1994:IG

[1117] Paolo Rossi. Immagini di Galileo. Nuncius, 9(1):3–14, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00018.

Bucciantini:1994:DSN

[1118] Massimo Bucciantini. Dopo il Sidereus Nuncius: Il copernicanesimo in Italia tra Galileo e Keplero. (Italian) [After the Starry Messenger:

Copernicanism in Italy from Galileo to Kepler]. Nuncius, 9(1):15–35, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00027.

Stabile:1994:LDN

[1119] Giorgio Stabile. Linguaggio della natura e linguaggio della scrittura in Galilei. Dalla istoria sulle macchie solari alle lettere copernicane. (Italian) [Language of nature and language of scripture in Galileo. From the history of sunspots to Copernican letters]. Nuncius, 9(1):37–64, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00036.

Redondi:1994:DLR

[1120] Pietro Redondi. Dietro l’immagine. Rappresentazioni di Galileo nella cultura positivistica. Nuncius, 9(1):65–116, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x00045.

Duris:1994:RPL

[1121] Pascal Duris. Le renouveau de la pensée Linnéenne en France au XIXe Siècle. (French) [The renewal of Linnean thought in France in the 19th Century]. Nuncius, 9(1):117–142, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00054.

LaVergata:1994:EW

[1122] Antonello La Vergata. Evolution and war, 1871–1918. Nuncius, 9(1): 143–163, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00063.

King:1994:APR

[1123] David A. King and Gerard L’E. Turner. The astrolabe presented by Regiomontanus to Cardinal Bessarion in 1462. Nuncius, 9(1):165–206, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00072.

Ceranski:1994:PAD

[1124] Beate Ceranski. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 9(1):207–231, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00081.

Anonymous:1994:IF

[1125] Anonymous. Istituzioni e fonti. Nuncius, 9(1):233, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x00090.

Lowood:1994:WBC

[1126] Henry Lowood. Workshop on bibliographic control of sources for the history of science, technology, and medicine: New directions and opportunities. Nuncius, 9(1):234–238, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00108.

Stevens:1994:SRH

[1127] Wesley M. Stevens. Sources and resources for history of science to 1600. Nuncius, 9(1):239–264, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00117.

Pogliano:1994:DIN

[1128] Claudio Pogliano. Discorsi inaugurali nelle facoltà mediche italiane (1875–1925). Nuncius, 9(1):265–294, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x00126.

Shea:1994:R

[1129] William R. Shea. Ricordo. Nuncius, 9(1):295–297, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x00135.

Bonsanti:1994:DC

[1130] Giorgio Bonsanti. Discussioni critiche. Nuncius, 9(1):299–303, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00144.

Touwaide:1994:RRH

[1131] Alain Touwaide. Recherches récentes sur Hippocrate et le Corpus Hippocraticum. Nuncius, 9(1):305–319, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00153.

Abbri:1994:LNP

[1132] Ferdinando Abbri. A.-L. Lavoisier: Nuove prospettive. Nuncius, 9(1): 321–332, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00162.

Touwaide:1994:BRGa

[1133] Alain Touwaide. Book review: Galen, on Semen, Edition, translation and commentary by Philippe de Locy, Berlin, Akademie Verlag, 1992, 291 pp. (= Corpus Medicorum Graecorum, V, 3, 1). Nuncius, 9(1): 333–335, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00171.

Touwaide:1994:BRN

[1134] Alain Touwaide. Book review: Nicoletta Palmieri, L’Antica versione latina del ≪De Sectis≫ di Galeno (Pal. lat. 1090), Pisa, Ets Editrice, 1992, 233 pp. (≪Testi e studi di cultura classica≫, 7). Nuncius, 9(1): 335–338, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00180.

diPasquale:1994:BRC

[1135] Giovanni di Pasquale. Book review: Claude Domergue, Les mines de la péninsule ibérique dans l’antiquité Romaine, Roma, l’≪Erma≫ di Bretschneider, 1990, 625 pp., XXXII tavole, 53 figure, 6 carte geografiche (≪Collection de l’École Française de Rome≫, 127). Nuncius, 9(1):338– 342, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00199.

Miniati:1994:BRLa

[1136] Mara Miniati. Book review: Luigi Pepe (a cura di), Copernico e la questione copernicana. Opere della Pubblica Biblioteca di Ferrara, Ferrara, Gabriele Corbo Editore, 1993, 203 pp., fig. Nuncius, 9(1):342–344, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00207.

Tosi:1994:BRGa

[1137] Alessandro Tosi. Book review: Gianni Bedini, Fabio Garbari (a cura di), I 400 anni dell’Orto botanico di Pisa. l’orto botanico: il passato chiave per il futuro?, Atti del convegno internazionale, Pisa, 11–12–13 ottobre 1991 (The 400 years of the Pisa Botanic Garden. The botanic garden: is its past the key to its future?, Proceedings of the International Symposium, Pisa, 11–12–13 October 1991), ≪Museologia Scientifica≫, anno IX, n. 1–4, gennaio–dicembre 1992, Verona, 1993, 460 pp. Nuncius, 9(1):344–346, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00216.

Ciardi:1994:BRB

[1138] Marco Ciardi. Book review: Bernard Joly, Rationalité de l’alchimie au XVIIe siècle, avec le texte latin, la traduction et le commentaire du Manuscriptum ad Federicum de Pierre-Jean Fabre, Préface de Jean-Paul Dumont, Paris, Vrin, 1992, 408 pp. Nuncius, 9(1):346–348, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00225.

Abbri:1994:BRM

[1139] Ferdinando Abbri. Book review: Michael Hunter (ed.), Robert Boyle reconsidered, Cambridge, Cambridge University Press, 1994, xvm + 231 pp. Nuncius, 9(1):348–352, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00234.

Roero:1994:BRE

[1140] Clara Silvia Roero. Book review: Enrico Pasini, Il reale e l’immaginario. La fondazione del calcolo infinitesimale nel pensiero di Leibniz, Torino, Edizioni Sonda, 1993, 224 pp., L. 28.000. Nuncius, 9(1):352–355, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00243.

Pepe:1994:BRA

[1141] Luigi Pepe. Book review: Der Ausbau des Calculus durch Leibniz und die Brüder Bernoulli, Symposion der Leibniz-Gesellschaft und der Bernoulli-Edition der naturforschenden Gesellschaft in Basel, 15 bis 17 Juni 1987, herausgegeben von Heinz-Jürgen Hess und Fritz Nagel, 1989, 175 pp. (≪Studia Leibnitiana≫, Sonderheft 17). Nuncius, 9(1):355–356, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00252.

Toscano:1994:BRG

[1142] Anna Toscano. Book review: Giulia Belgioioso, Cultura a Napoli e Cartesianesimo. Scritti su G. Gimma, P. M. Doria, C. Cominale, Lecce, Congedo Editore, 1992, 459 pp. Nuncius, 9(1):356–358, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00261.

Ciancio:1994:BRB

[1143] Luca Ciancio. Book review: Barbara Maria Stafford, Body Criticism. Imaging the Unseen in Englightenment Art and Medicine, Cambridge Mass.-London, The MIT Press, 1991, xxi + 587 pp. Nuncius, 9(1): 359–363, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00270.

Segala:1994:BRK

[1144] Marco Segala. Book review: Kathleen Wellman, La Mettrie. Medicine, Philosophy, and Enlightenment, Durham and London, Duke University Press, 1992, xiv-342 pp. (ISBN 0-8223-1204-2). Nuncius, 9(1):364–367, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00289.

Beretta:1994:BRG

[1145] Marco Beretta. Book review: Giovanni Antonio Scopoli, Dissertatio de diaeta litteratorum. Traduzione di Domenico Magnino. Testi di Gianguido Rindi e Carlo Violani, Pavia, Cisalpino, 1991, 141 pp., Ill. Nuncius, 9(1):367–368, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00298.

Abbri:1994:BRTa

[1146] Ferdinando Abbri. Book review: Tore Frängsmyr, Sökandet efter Upplysningen. En essä om 1700-talets svenskakulturdebatt, (Swedish) [The search for Enlightenment. An essay on the Swedish cultural debate in the 1700s], Stockholm, Wiken, 1993, 200 pp. Nuncius, 9(1):369–371, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00306.

Ciardi:1994:BRM

[1147] Marco Ciardi. Book review: Marco Beretta, the Enlightenment of Matter: the Definition of Chemistry from Agricola to Lavoisier, Canton, Mass., Science History Pubblications, 1993, xvi-396 pp., ill. (≪Uppsala Studies in the History of Science≫, 15). Nuncius, 9(1):371–377, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00315.

Israel:1994:BRP

[1148] Giorgio Israel. Book review: Paola Dessì, l’ordine e il caso. Discussioni epistemologiche e logiche sulla probabilità da Laplace a Peirce, Bologna, Il Mulino, 1989, 230 pp., L. 26.000. Nuncius, 9(1):377–379, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00324.

Dhombres:1994:BRI

[1149] Jean Dhombres. Book review: Ivor Grattan-Guinness, Convolutions in French Mathematics 1800–1840, Basel, Boston, Berlin, Birkhäuser Verlag, 1990, 3 voll., 1602 pp. Nuncius, 9(1):379–386, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00333.

Vista:1994:BRC

[1150] Michele Vista. Book review: Christoph Meinel, Harmut Scholz (Hrsg.), Die Allianz von Wissenschaft und Industrie — August Wilhelm Hofmann: (1818–1892); Zeit, Werk, Wirkung, Weinheim, Vhc, 1992, x + 375 pp., fig. Nuncius, 9(1):386–389, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00342.

Cerruti:1994:BRT

[1151] Luigi Cerruti. Book review: Thomas Hapke, Die Zeitschrift für physikalische Chemie. Hundert Jahre Wechselwirkung zwischen Fachwissenschaft, Kommunikationsmedium und Gesellschaft, Herzberg, Verlag Traugott Bautz, 1990, 146 pp., 58 marchi, ISBN 3-88309-030-1. Nuncius, 9(1):389–393, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00351.

Corbellini:1994:BRK

[1152] Gilberto Corbellini. Book review: Kenneth F. Kiple (editor), the Cambridge World History of Human Disease, Cambridge, Cambridge University Press, 1993, xx-1176 pp., with Tables, Figures and Maps. Nuncius, 9(1):393–395, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00360.

Ronchi:1994:BRJ

[1153] Lucia Rositani Ronchi. Book review: Jürgen Teichmann, Zur Geschichte der Festkörperphysik. Farbzentrenforschung bis 1940, Stuttgart, Franz Steiner Verlag, 1988, 162 pp. (≪Boethius≫, Band XVII). Nuncius, 9 (1):395–400, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00379.

Scarpelli:1994:BRP

[1154] Giacomo Scarpelli. Book review: Phillip V. Tobias, Olduvai Gorge. Volume 4. The Skulls, Endocasts and Teeth of Homo Habilis, preface by Raymond A. Dart, Cambridge, Cambridge University Press, 1991, 2 voll., xxxv + 921 pp., ill. (ISBN 0-521-20072-5). Nuncius, 9(1):400–403, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00388.

Beretta:1994:BRAa

[1155] Marco Beretta. Book review: Aa.Vv., l’Image et la science, Paris, Comité des travaux historiques et scientifiques, 1992, 425 pp., Ill. Nuncius, 9 (1):403–404, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00397.

Caroti:1994:BRS

[1156] Stefano Caroti. Book review: Strumenti Bibliografici Catalogo di manoscritti filosofici nelle biblioteche italiane. Volume 7. Novara, Palermo, Pavia, a cura di G. M. Cao, C. Casagrande, M. A. Casagrande-Mazzoli, D. Ciccarelli, S. Gavinelli, S. Vecchio. Premessa di C. Leonardi, Firenze, Leo S. Olschki, 1993, 319 pp., ind. dei mss., ind. degli incipit, ind. dei nomi e delle opere anonime (≪Unione Accademica Nazionale. Corpus philosophorum Medii Aevi. Subsidia≫, VIII). Nuncius, 9(1):405–406, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00405.

Caroti:1994:BREa

[1157] Stefano Caroti. Book review: Eugenio Canone (a cura di), Bibliothecae selectae. Da Cusano a Leopardi, Firenze, Leo S. Olschki Editore, 1993, 631 pp., ill. (≪Lessico Intellettuale Europeo≫, Lviii). Nuncius, 9(1): 406–407, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00414.

Miniati:1994:BRLb

[1158] Mara Miniati. Book review: Livia Orlandi Frattarolo (a cura di), Bologna e il Grand Tour, Roma, Istituto Poligrafico dello Stato, 1991, 46 pp. fig. Livia Orlandi Frattarolo, Irene Ventura Folli, Esplorazioni in biblioteca, Bologna, Analisi, 1991, 114 pp. fig. Nuncius, 9(1):407–408, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00423.

Miniati:1994:BRMa

[1159] Mara Miniati. Book review: Museologia Pietro A. Valentino, l’immagine e la memoria. Indagine sulla struttura del Museo in Italia e nel mondo, Milano, Leonardo Periodici s.r.l., 1993, 302 pp., fig., L. 50.000. Nuncius, 9(1):408–409, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00432.

Miniati:1994:BRSa

[1160] Mara Miniati. Book review: Strumentaria Giovanni Paltrinieri, Astrolabi e volvelle. La sfera celeste descritta su un piano, 39 pp., in ≪L’Astrofilo. Bollettino dell’Unione Astrofili Bresciani≫, 20, 1992. Nuncius, 9(1):409, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00441.

Miniati:1994:BRAa

[1161] Mara Miniati. Book review: Anita Mcconnell, R. B. Bate of the Poultry 1782–1847. The Life and Times of a Scientific Instrument Maker, London, Scientific Instrument Society, 1993, 74 pp. fig. (≪Sis Monograph≫, 1). Nuncius, 9(1):409–410, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00450.

Miniati:1994:BRB

[1162] Mara Miniati. Book review: Bruno Carazza, Michele Ceriana-Mayneri, l’≪Inventario delle machine≫ del Gabinetto di Fisica di Torino, s.n.t. [Torino, 1993], [84] pp. Nuncius, 9(1):410–411, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00469.

Miniati:1994:BRGa

[1163] Mara Miniati. Book review: Giulia Luppi (a cura di), Bilance a bracci uguali/equal-arm balances, Modena, Museo della Bilancia/Il Bulino, 1993, 205 pp., fig. (≪Biblioteca del Museo della Bilancia. Collezioni del Museo≫, 1). Nuncius, 9(1):411–412, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00478.

Dolza:1994:BRS

[1164] Luisa Dolza. Book review: Storia della Tecnologia History of Technology, volume fourteen (14) edited by Graham Hollister-Short and Frank A. J. L. James, Londra, Mansell, 1992, x + 182 pp. Nuncius, 9(1):412–414, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00487.

Miniati:1994:BRSb

[1165] Mara Miniati. Book review: ≪Quaderni di storia della tecnologia≫, n. 3. Torino, Editrice Levrotto & Bella, 1993, 193 pp., fig. Nuncius, 9 (1):414–415, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00496.

Miniati:1994:BRLc

[1166] Mara Miniati. Book review: Lili Cafarella, Antonio Meloni, Cristoforo Colombo e la scoperta della declinazione magnetica, Roma, Istituto Nazionale di Geofisica, 1993, 21 pp., fig. Nuncius, 9(1):415–416, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00504.

Miniati:1994:BRLd

[1167] Mara Miniati. Book review: Lili Cafarella, A. De Santis, Antonio Meloni, Il catalogo geomagnetico storico italiano. The historical Geomagnetic Data Catalogue. With an extended summary in English, [Roma], Istituto Nazionale di Geofisica, s.d., 160 pp., fig. Nuncius, 9(1):416–417, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00513.

Miniati:1994:ASA

[1168] Mara Miniati. Archivio storico amma, ≪le culture della tecnica≫, settembre 1993. Nuncius, 9(1):417, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00522.

Abbri:1994:BRSa

[1169] Ferdinando Abbri. Book review: Storia Sociale E. Anthony Rotundo, American Manhood. Transformations in Masculinity from the Revolution to the Modern Era, New York, Basic Books, 1993, xiv + 382 pp. Nuncius, 9(1):418–420, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00531.

Miniati:1994:BRAb

[1170] Mara Miniati. Book review: Antropologia Aa.Vv., La science sauvage. Des savoirs populaires aux ethnosciences, (French) [Wild science. Popular knowledge to ethnosciences]. Paris, Éditions du Seuil, 1993, 212 pp., fig. Nuncius, 9(1):420–421, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00540.

Beretta:1994:BRC

[1171] Marco Beretta. Book review: Chimica Aa.Vv., Von der Astronomie zur Alchemie, Rastatt, Historische Bibliothek der Stadt Rastatt, 1991, 273 pp. Ill. Nuncius, 9(1):421–422, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00559.

Miniati:1994:BRT

[1172] Mara Miniati. Book review: Le tavole murali di von Schroeder, (Italian) [The mural tables of von Schroeder], Milano, Electa, 1992, 89 pp., fig. Nuncius, 9(1):422, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00568.

Lippi:1994:BRM

[1173] Donatella Lippi. Book review: Medicina e Biologia Enrico Marcovecchio, Dizionario Etimologico Storico dei Termini Medici, Firenze, Ed. Festina Lente, 1993, pp. 938. Nuncius, 9(1):422–423, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00577.

Touwaide:1994:BRGb

[1174] Alain Touwaide. Book review: Galeno, Sull’ottima maniera d’insegnare — Esortazione alla medicina, Testo e traduzione di Adelmo Barigazzi, Berlin, Akademie Verlag, 1991, 172 pp. (= Corpus Medicorum Graecorum V, 1, 1). Nuncius, 9(1):423–425, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x00586.

Lippi:1994:BRLa

[1175] Donatella Lippi. Book review: Loris Premuda, Storia dell’iconografia anatomica, Saronno, Ed. Ciba, 1993, 385 pp. Nuncius, 9(1):425–426, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00595.

Lippi:1994:BRAa

[1176] Donatella Lippi. Book review: Aa.Vv., Le origini della chirurgia italiana, Catalogo della Mostra (Roma, dicembre 1993), Roma, MBCA, 1993, 135 pp. Nuncius, 9(1):427–428, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00603.

Lippi:1994:BRH

[1177] Donatella Lippi. Book review: E. Haefliger e J. Brun-Hool, Tavole delle malerbe, Saronno, Ciba Edizioni, 1991, xx pp., tavole. Nuncius, 9(1): 428–429, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00612.

Lippi:1994:BRAb

[1178] Donatella Lippi. Book review: Aa.Vv., Healing. Storia e strategie del guarire, a cura di O. Galeazzi, Firenze, Olschki, 1993, 362 pp. Nuncius, 9(1):429–430, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00621.

Lippi:1994:BRV

[1179] Donatella Lippi. Book review: V. C. Medvei, the History of Clinical Endocrinology, New York, Parthenon Publishing Group, 1993, 551 pp. Nuncius, 9(1):431–432, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00630.

Malacrida:1994:BRL

[1180] Lorenzo Marri Malacrida. Book review: Lanmarco Laquidara, l’odontoiatria attraverso i secoli, Siena, Ciso editore, 1993, 400 pp., L. 50.000. Nuncius, 9(1):432–433, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00649.

Lippi:1994:BRC

[1181] Donatella Lippi. Book review: L. Capasso, A. Capelli, Le epidemie di peste in Abruzzo, dal 1348 al 1702, Cerchio (Aq), A. Polla Ed., 1993, 102 pp. Nuncius, 9(1):433–434, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00658.

Malacrida:1994:BRR

[1182] Lorenzo Marri Malacrida. Book review: Roy Porter (ed.), The medical history of water and spas, London, Wellcome Institute for the History of Medicine, 1990, 150 pp., fig. Nuncius, 9(1):434, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00667.

Laquidara:1994:BRF

[1183] Lanmarco Laquidara. Book review: Francesca Vannozzi (a cura di), Per una storia dell’insegnamento della medicina legale nell’Università di Siena, Siena, Ciso Editore, 1993, 200 pp. Nuncius, 9(1):434–435, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00676.

Malacrida:1994:BRJ

[1184] Lorenzo Marri Malacrida. Book review: John G. Howells (ed.), The concept of schizophrenia: historical perspectives, Washington, American Psichiatric Press, 1991, 211 pp., fig. Nuncius, 9(1):436, ????

1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00685.

Miniati:1994:BRMb

[1185] Mara Miniati. Book review: Maria Sibylla Merian, La meravigliosa metamorfosi dei bruchi, (Italian) [The wonderful metamorphosis of caterpillars], a cura di Maria Gregorio, prefazione di Giorgio Celli, Torino, Rosenberg & Sellier, 1993, 212 pp., fig. Nuncius, 9(1):436–437, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00694.

Malacrida:1994:BRF

[1186] Lorenzo Marri Malacrida. Book review: Francesco Carnevale, Gian Bruno Ravenni (a cura di), Gaetano Pieraccini medico del lavoro. La salute dei lavoratori in Toscana all’inizio del XX secolo, Firenze, Editoriale Tosca, 301 pp. fig. Nuncius, 9(1):437–438, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00702.

Caroti:1994:BRMa

[1187] Stefano Caroti. Book review: Medioevo e Rinascimento Brigitte Prévot, Bernard Ribémont, Le cheval en France au Moyen Age. Sa place dans le monde médiéval; sa médecine: l’exemple d’un traité vétérinaire du XIVeme siècle, la ≪Chirurgie des chevaux≫, Orléans, Paradigme, 1994, 522 pp., ill., bibl., app. (≪Medievalia≫ ). Nuncius, 9(1):438–439, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00711.

Caroti:1994:BRMb

[1188] Stefano Caroti. Book review: M. J. F. M. Hoenen, Marsilius of Inghen. Divine Knowledge in Late Medieval Thought, Leiden — New York — Köln, E. J. Brill, 1993, 287 pp., ind. bibl. (≪Studies in the History of Christian Thought≫, L). Nuncius, 9(1):439–440, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00720.

Caroti:1994:BRL

[1189] Stefano Caroti. Book review: Luciano Capra (a cura di), Note di Torquato Tasso a ≪De Caelo≫ di Aristotele, Ferrara-Roma, Gabriele Corbo Editore, 1993, 194 pp., ind. (≪Esami≫, 2). Nuncius, 9(1):440, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00739.

Lamberini:1994:BRN

[1190] Daniela Lamberini. Book review: Numa Broc, La geografia del Rinascimento, a cura di Claudio Greppi, Modena, Edizioni Panini, 1989, 172 pp., ill. b.n., f.t. 15 br. Indici, (≪Istituto di Studi Rinascimentali, Ferrara≫, Saggi), Lit. 40.000. Nuncius, 9(1):441, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00748.

Caroti:1994:BRAa

[1191] Stefano Caroti. Book review: Augusta Buck (Hrsg.), Die okkulten Wissenschaften in der Renaissance, Wiesbaden, Otto Harassowitz, 1992, 293 pp., ill. ind. (≪Wolfenbütteler Abhandlungen zur Renaissanceforschung. in Zusammenarbeit mit dem Wolfenbütteler Arbeitskreis für Renaissanceforschung herausgegeben von der Herzog August Bibliothek≫, Bd. 12). Nuncius, 9(1):442–443, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00757.

Lippi:1994:BRPa

[1192] Donatella Lippi. Book review: P. Gavitt, Charity and children in Renaissance Florence. The Ospedale degli Innocenti, 1410–1536, Ann Arbor, The University of Michigan Press, 1990, 330 pp. Nuncius, 9(1):443–444, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00766.

Miniati:1994:BRO

[1193] Mara Miniati. Book review: Owen Gingerich, Robert S. Westman, the Wittich Connection: Conflict and Priority in Late Sixteenth-Century Cosmology, Philadelphia, the American Philosophical Society, 1988, 148 pp., fig. (≪Transactions of the American Philosophical Society≫ vol. 78, Part 7). Nuncius, 9(1):444–445, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00775.

Abbri:1994:BRSb

[1194] Ferdinando Abbri. Book review: Secoli XVI–XX Brian Harvey Goodwin Wormald, Francis Bacon. History, politics and science. 1561–1626, Cambridge, Cambridge University Press, 1993, xii + 409 pp. Nuncius, 9(1):445–447, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00784.

Miniati:1994:BRMc

[1195] Mara Miniati. Book review: Milla Baldo Ceolin (ed.), Galileo scientist. His Years at Padua and Venice, s.l., Istituto Nazionale di Fisica Nucleare, Istituto Veneto di Scienze, Lettere ed Arti, Dipartimento di Fisica ≪Galileo Galilei≫ dell’Università degli Studi di Padova, [1993], 140 pp., fig. Nuncius, 9(1):447–448, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00793.

Iannaccone:1994:BRB

[1196] Isaia Iannaccone. Book review: Brian L. Merril (ed.), Athanasius Kircher (1602–1680) Jesuit Scholar, Provo, Utah, Friends of the Brigham Young University Library, 1989, xxxvii + 74 pp. Nuncius, 9(1):448–450, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00801.

Beretta:1994:BRAb

[1197] Marco Beretta. Book review: Alvaro Alonso Barba, Arte de los metales (1640), ristampa anastatica, Madrid, CSIC, 1992, 228 pp. Ill. Nuncius, 9(1):450, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x00810.

Segala:1994:BRP

[1198] Marco Segala. Book review: Peter M. Harman — Alan E. Shapiro (eds.), The investigation of difficult things. Essays on Newton and the history of the exact sciences in honour of D. T. Whiteside, Cambridge, Cambridge University Press, 1992, xvi + 531 pp. Nuncius, 9(1):451–454, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00829.

Abbri:1994:BRAa

[1199] Ferdinando Abbri. Book review: Alan E. Shapiro, Fits, Passions, and Paroxysms. Physics, method, and chemistry and Newton’s theories of colored bodies and fits of easy reflection, Cambridge, Cambridge University Press, 1993, xviii + 400 pp. Nuncius, 9(1):454–456, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00838.

Abbri:1994:BRAb

[1200] Ferdinando Abbri. Book review: Antonio di Meo, Silvano Tagliagambe (a cura di), Teorie e filosofie della materia nel Settecento, Roma, Editori Riuniti, 1993, x + 230 pp. L. 50.000. Nuncius, 9(1):457–458, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00847.

Abbri:1994:BRP

[1201] Ferdinando Abbri. Book review: Paola Giacomoni, Le forme e il vivente. Morfologia e filosofia della natura in J. W. Goethe, Napoli, Guida editori, 1993, 284 pp. ill. Lit. 35.000. Nuncius, 9(1):458–459, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00856.

Segala:1994:BRJ

[1202] Marco Segala. Book review: John Elbert Wilson, Schellings Mythologie. Zur Auslegung der Philosophie der Mythologie und der Offenbarung, Stuttgart-Bad Cannstatt, Frommann-Holzboog, 1993, 292 pp. (≪Spekulation und Erfahrung≫, Abt. 2, Untersuchungen; Bd. 31) (ISBN 3-7728-1483-2). Nuncius, 9(1):459–460, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00865.

Segala:1994:BRE

[1203] Marco Segala. Book review: Eduard A. Tropp, Viktor Y. Frankel, and Arthur D. Chernin, Alexander A. Friedmann: the Man who Made the Universe Expand, Cambridge, Cambridge University Press, 1993, x + 267 pp. (ISBN 0-521-38470-2). Nuncius, 9(1):460–461, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00874.

Abbri:1994:BRW

[1204] Ferdinando Abbri. Book review: William R. Shea, Antonio Spadafora (eds.), Interpreting the World. Science and Society, Canton, Mass., Science History Publications, 1992, XII + 218 pp. Nuncius, 9(1):461–462, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00883.

Anonymous:1994:ADM

[1205] Anonymous. Attività dell’Istituto e Museo di Storia Della Scienza di Firenze. Nuncius, 9(1):463–468, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00892.

Anonymous:1994:PR

[1206] Anonymous. Pubblicazioni ricevute. Nuncius, 9(1):469–476, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00900.

Giusti:1994:FGM

[1207] Enrico Giusti. Il filosofo geometra. matematica e filosofia naturale in Galileo. Nuncius, 9(2):485–498, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00919.

Hill:1994:PPW

[1208] David K. Hill. Pendulums and planes: What Galileo didn’t publish. Nuncius, 9(2):499–515, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00928.

Franco:1994:VCV

[1209] Loredana Franco. Vincenzo Coronelli: Vita e opere. Aggiornamenti. Nuncius, 9(2):517–541, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x00937.

Pareti:1994:DDD

[1210] Germana Pareti. Il declino della dottrina del blastema. (Italian) [The decline of the doctrine of blastema]. Nuncius, 9(2):543–558, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00946.

Conry:1994:LPF

[1211] Yvette Conry. Lamarck, penseur de frontière. (French) [Lamarck, frontier thinker]. Nuncius, 9(2):559–592, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x00955.

Bohr:1994:EUG

[1212] Michael Bohr. Englische Uhren Für Grossherzog Cosimo III. Und Die Ankunft von Ignazio Hugford am Hofe in Florenz’. (German) [English clocks for Grand Duke Cosimo III. And the arrival of Ignazio Hugford at the Court in Florence]. Nuncius, 9(2):593–618, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00964.

Turner:1994:NPM

[1213] Anthony J. Turner. Natural philosophers, mathematical practitioners and timber in later 17th Century England. Nuncius, 9(2):619–634, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00973.

diPasquale:1994:SGC

[1214] Giovanni di Pasquale. Studio su un gruppo di compassi romani provenienti da Pompei. Nuncius, 9(2):635–644, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x00982.

Brusa:1994:LPL

[1215] Giuseppe Brusa. L’orologio dei pianeti di Lorenzo della Volpaia. (Italian) [the planetary clock of Lorenzo della Volpaia]. Nuncius, 9(2):645–669, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x00991.

Brenni:1994:BR

[1216] Paolo Brenni. Il barometro ritrovato. Nuncius, 9(2):670–676, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01008.

Miniati:1994:EAT

[1217] Mara Miniati, Vincenzo Greco, Giuseppe Molesini, and Franco Quercioli. Examination of an antique telescope. Nuncius, 9(2):677–682, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01017.

Toscano:1994:BRC

[1218] Anna Toscano. Book review: Catalogo Delle Carte di Giorgio Baglivi Conservate Nella Waller Samling, Presso Universitetsbiblioteket ≪Carolina Rediviva≫ di Uppsala. Nuncius, 9(2):683–738, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01026.

Farinella:1994:RBS

[1219] Calogero Farinella. Ricostruire la biblioteca di uno scienziato: Il caso di A. M. Lorgna. (Italian) [rebuilding the library of a scientist: the case of A. M. Lorgna]. Nuncius, 9(2):739–758, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01035.

Fodera:1994:CSS

[1220] Giorgia Foderá, Francesca Martines, and Donatella Randazzo. Cataloghi di strumenti scientifici nella Biblioteca Dell’Osservatorio Astronomico di Palermo. Nuncius, 9(2):759–796, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01044.

Barbara:1994:SRT

[1221] Brusa Maria Barbara. Sul rapporto tra strumenti e teorie nella storia della scienza. Nuncius, 9(2):797–807, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x01053.

Corbellini:1994:IDG

[1222] Gilberto Corbellini. Individuality and disease in Garrod’s theory of medicine. Nuncius, 9(2):809–817, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01062.

Touwaide:1994:BRS

[1223] Alain Touwaide. Book review: Storia del pensiero medico occidentale, a cura di Mirko D. Grmek, Vol. 1: Antichità e medioevo, Roma & Bari, Laterza, 1993, 577 pp. Nuncius, 9(2):819–821, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01071.

Touwaide:1994:BRA

[1224] Alain Touwaide. Book review: Actes du colloque international ≪Les phytonymes grecs et latins≫, tenu à Nice les 14, 15 et 16 mai 1992 à la Faculté des Lettres, Arts et Sciences humaines, Université de Nice-Sophia Antipolis, Nice, Université, 1993, 324 pp. (≪Université de Nice Sophia Antipolis, Centre de Recherches comparatives sur les Langues de la Méditerranée ancienne≫, 12). Des hommes et des plantes — Plantes méditerranéennes, vocabulaire et usages anciens, Table ronde Aix-en-Provence, mai 1992, Edité par M. C. Amouretti & G. Cornet, Aix-en-Provence, Université, Service des Publications, 1993, 174 pp. (≪Cahier d’histoire des techniques≫, 2). Nuncius, 9(2):822–824, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01080.

Touwaide:1994:BRT

[1225] Alain Touwaide. Book review: Theophrastus of Eresus, Sources for his Life, Writings, Thought& Influence, edited and translated by William W. Fortenbaugh, Pamela M. Huby, Robert W. Sharples (Greek and Latin) and Dimitri Gutas (Arabic), together with Andrew D. Barker, John J. Keaney, David D. Mirhady, David Sedley and Michael G. Sollenberger. 2 vol. Part One: Life, Writings, Various Reports, Logic, Physics, Metaphysics, Theology, Mathematics. Part Two: Phsychology, Human Physiology, Living Creatures, Botany, Ethics, Religion, Politics, Rhetoric and Poetics, Music, Miscellanea, Leyde, Brill, 1992, 465 & 705 pp. (= Philosophia antiqua, 54, 1 & 2). Nuncius, 9(2):825–827, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01099.

Berrani:1994:BRA

[1226] Hamid Berrani. Book review: Angel C. López y López, Kitāb fī tartīb awqāt al-girāsa wa-l-magrūsāt. un tratado agrícola andalusí anónimo. Edición, traducción y estudio con glosario de… Granada, Consejo Superior de Investigaciones Científicas. Escuela de Estudios Arabes, 1990, 410 pp. Nuncius, 9(2):828–829, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01107.

Alkuwaifi:1994:BRM

[1227] Ahmad Alkuwaifi. Book review: Ma Angeles Navarro, Risāla fī Awqāt al-Sana. un Calendario Anónimo Andalusí. Granada, Consejo Superior de Investigaciones Científicas, Escuela de Estudios Arabes, 1990, 284 pp.

Nuncius, 9(2):829–831, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x01116.

Pereira:1994:BRA

[1228] Michela Pereira. Book review: Alchimie et Philosophie à la Renaissance, sous la direction de Jean-Claude Margolin et Sylvain Matton, Paris, Librairie Philosophique J. Vrin, 1993, (≪De Pétrarque à Descartes≫ Lvii), ISSN 0418-4459, ISBN 2-7116-1172-8, 478 pp. Nuncius, 9(2):831–835, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01125.

Pereira:1994:BRC

[1229] Michela Pereira. Book review: Cornelius Agrippa, De occulta philosophia libri tres, edited by Vittoria Perrone Compagni, Leiden, E. J. Brill, 1992, 659 pp., (≪Studies in the History of Christian Thought≫, ed. by H. A. Oberman, vol. Xlviii), ISBN 90-04-09421-0. Nuncius, 9(2):835–837, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01134.

Plastina:1994:BRG

[1230] Sandra Plastina. Book review: Giordano Bruno. Gli anni napoletani e la ≪ peregrinatio≫ europea. Immagini. Testi (Italian) [Giordano Bruno. The years and the Neapolitan European pilgrimage. Images. Texts]. Documenti a cura di Eugenio Canone, Cassino, Università degli Studi di Cassino, 1992, 246 pp. + 8 tavole. Nuncius, 9(2):838–840, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01143.

Abbri:1994:BRB

[1231] Ferdinando Abbri. Book review: Barbara Beigun Kaplan, ≪Divulging of Useful Truths in Physics≫. the Medical Agenda of Robert Boyle, Baltimore and London, the Johns Hopkins University Press, 1993, xii + 216 pp. Nuncius, 9(2):840–842, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01152.

Toscano:1994:BRE

[1232] Anna Toscano. Book review: Edmé Mariotte, Essai de Logique. Suivi de l’écrit intitulé ’Les principes du devoir et des connaissances humaines’ attribué a Roberval, Paris, Fayard, 1992, 206 pp. (≪Corpus des Oeuvres de Philosophie en Langue Française≫ ). Nuncius, 9(2):843–845, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01161.

Toscano:1994:BRF

[1233] Anna Toscano. Book review: Franco Crispini, l’opinione del bene. A. Shaftesbury tra ispirazioni antiche e ragione moderna, Napoli, Morano Editore, 1994, 166 pp. (≪Collana di Filosofia≫ ). Nuncius, 9(2):846– 847, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01170.

Abbri:1994:BRJ

[1234] Ferdinando Abbri. Book review: John Stoye, Marsigli’s Europe 1680– 1730. The Life and Times of Luigi Ferdinando Marsigli, Soldier and Virtuoso, New Haven and London, Yale University Press, 1994, xii + 356 pp., ill. Nuncius, 9(2):847–849, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01189.

Bernardi:1994:BRL

[1235] Walter Bernardi. Book review: Lazzaro Spallanzani, I Giornali delle Sperienze e Osservazioni, a cura di C. Castellani, Firenze, Giunti, 1994, voll. 5 in 6 tt. (≪Biblioteca della Scienza Italiana≫, 1). Nuncius, 9 (2):849–856, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x01198.

Bernardi:1994:BRC

[1236] Walter Bernardi. Book review: Cinzio Gibin, La geometria della natura. Chioggia e I’Europa nella vicenda intellettuale di Giuseppe Olivi naturalista del Settecento, Padova, Il Poligrafo, 1994, 173 pp. Nuncius, 9 (2):856–859, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x01206.

Pogliano:1994:BRA

[1237] Claudio Pogliano. Book review: Alle soglie della rivoluzione tecnologica. Josef Ressel: un inventore a Trieste 1773–1993, (Italian) [On the threshold of the technology revolution. Josef Ressel: an inventor in Trieste 1773–1993], a cura di Valerio Staccioli, Trieste, Azienda di Promozione Turistica, 1994, 184 pp. Nuncius, 9(2):859–862, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01215.

Ciardi:1994:BRC

[1238] Marco Ciardi. Book review: The Correspondence of Michael Faraday. Volume 2, 1832–1840, edited by Frank A. J. L. James, London, the Institution of Electrical Engineers, 1993, lv + 807 pp., ill. Nuncius, 9 (2):862–865, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x01224.

Anonymous:1994:BRT

[1239] Anonymous. Book review: Theodor M. Porter, Le origini del moderno pensiero statistico (1820–1900), Firenze, La Nuova Italia, 1993, xxxii + 349 pp. Nuncius, 9(2):865–868, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01233.

Miniati:1994:BRGb

[1240] Mara Miniati. Book review: Giovanni Boato, Giuseppe Bruzzaniti, Strumenti nella Fisica dell’Ottocento, Genova, Sagep Editrice, 1993, 142 pp. fig. Nuncius, 9(2):868, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x01242.

Miniati:1994:BRLe

[1241] Mara Miniati. Book review: Liceo Classico ≪L. Ariosto≫ Ferrara, Strumentaria. Alla scoperta dell’antico laboratorio di Fisica del Liceo Ariosto (1860–1924). Ferrara-Poggio Renatico, Lions Club, 1993, 117 pp., fig. Nuncius, 9(2):869–870, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x01251.

Garbolino:1994:BRI

[1242] Paolo Garbolino. Book review: Ian Hacking, Il caso domato, a cura di Simona Morini, Milano, Il Saggiatore, 1994, xxx + 335 pp. Nuncius, 9 (2):870–873, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x01260.

Ciardi:1994:BRA

[1243] Marco Ciardi. Book review: Atti del Convegno: La fisica a Pavia nelle opere di Giuseppe Belli, a cura di G. Bellodi e E. Rovida, Pavia, Collana di Storia della Scienza dell’Università di Pavia, 1994, 144 pp., ill. Nuncius, 9(2):873–877, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x01279.

Segala:1994:BRG

[1244] Marco Segala. Book review: Giacomo Scarpelli, Il cranio di cristallo. Evoluzione della specie e spiritualismo, Torino, Bollati Boringhieri, 1993, 219 pp. (≪Saggi scientifici≫ ). Nuncius, 9(2):877–880, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01288.

Weindling:1994:BRE

[1245] Paul Weindling. Book review: Elisabeth Crawford, Nationalism and internationalism in science, 1880–1939. Four Studies of the Nobel Population, Cambridge, Cambridge University Press, 1992, xii + 157 p. Nuncius, 9(2):880–881, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x01297.

Giuntini:1994:BRE

[1246] Chiara Giuntini. Book review: Eugene Webb, the Self Between. From Freud to the New Social Psichology of France, Seattle-London, University of Washington Press, 1993, xi + 268 pp. Nuncius, 9(2):881–883, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01305.

Robotti:1994:BRJ

[1247] Nadia Robotti. Book review: Joseph Agassi, Radiation Theory and the Quantum Revolution, Basel–Boston–Berlin, Birkhäuser Verlag, 1993, 178 pp. Nuncius, 9(2):883–885, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01314.

Cordeschi:1994:BRM

[1248] Roberto Cordeschi. Book review: Marvin Minsky, La società della mente, trad. it. di Giuseppe Longo, Milano, Adelphi, 1989, 674 pp. Nuncius, 9 (2):885–888, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x01323.

Peruzzi:1994:BRL

[1249] Giulio Peruzzi. Book review: Lillian Hoddeson, Paul W. Henriksen, Roger A. Meade, Catherine Westfall (eds.), Critical Assembly: A Technical History of Los Alamos during the Oppenheimer Years, 1943–1945, Cambridge, Cambridge University Press, 1993, xv + 509 pp. Nuncius, 9(2):888–892, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x01332.

Caroti:1994:SB

[1250] Stefano Caroti. Strumenti bibliografici. Nuncius, 9(2):893–894, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01341.

Miniati:1994:BRAc

[1251] Mara Miniati. Book review: Anita Bacci, Donatella Parrini, Francesca Vannozzi, I documenti dell’Accademia I. Verbali, memorie, epistole ed atti dalla fondazione al secolo XVIII. Siena, Accademia delle Scienze di Siena della de’ Fisiocritici, 1994, 261 pp. (≪Memorie≫, 5). Nuncius, 9(2):894, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x01350.

Tosi:1994:BRM

[1252] Alessandro Tosi. Book review: I Manoscritti di Pier Antonio Micheli conservati nella Biblioteca Botanica dell’Università di Firenze, (Italian) [The manuscripts of Pier Antonio Micheli preserved in the Library of Botany at The University of Florence], catalogo a cura di Stefania Ragazzini, presentazioni di Paolo Galluzzi, Elena Maugini e Guido Moggi, Milano, Giunta regionale toscana & Editrice Bibliografica, 1993, 99 pp., ill. Nuncius, 9(2):895–896, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01369.

Miniati:1994:M

[1253] Mara Miniati. Museologia. Nuncius, 9(2):896, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x01378.

Miniati:1994:S

[1254] Mara Miniati. Strumentaria. Nuncius, 9(2):897, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x01387.

Miniati:1994:BRGc

[1255] Mara Miniati. Book review: Giampaolo Gallo (a cura di), Per una storia del vetro nel Valdarno. Ivv 1952-92. Firenze, Ponte alle Grazie, 1994, 187 pp., fig. Nuncius, 9(2):897–898, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01396.

Miniati:1994:BRD

[1256] Mara Miniati. Book review: Danuta Burczyk-Marona, Scientific and teaching instruments in the Jagiellonian Museum. Kraków, Prowincjonalna Oficyna Wydawnicza Exartim, 1993, 32 pp., fig. (≪Collegium Maius≫, 3). Nuncius, 9(2):898, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01404.

Miniati:1994:BREa

[1257] Mara Miniati. Book review: Emilio Borchi, Renzo Macii, Gli antichi strumenti del gabinetto di fisica. Empoli, Istituto Calasanzio, 1993, 204 pp., fig. Nuncius, 9(2):899, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01413.

Miniati:1994:BRAd

[1258] Mara Miniati. Book review: Antichi elettrometri e galvanometridel Dipartimento di Fisica di Firenze. Firenze, 1992, 50 pp., fig. (riproduzione in fotocopia). l’antico gabinetto di Fisica di un Liceo Classico. Firenze, Liceo Classico Statale ≪Dante≫, 1993, 57 pp., fig. (riproduzione in fotocopia). Nuncius, 9(2):899–900, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01422.

Miniati:1994:BRF

[1259] Mara Miniati. Book review: Fausto Casi (a cura di), Strumenti scientifici e macchine di interesse storico della Usl di Arezzo. Catalogo. Arezzo, Centro di Catalogazione degli strumenti scientifici di interesse storico, 1994, 182 pp. n.n., fig. (fotocopie rilegate). Nuncius, 9(2):900– 901, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01431.

Giatti:1994:BRG

[1260] Anna Giatti. Book review: Gigliola Terenna, Francesca Vannozzi (a cura di), Inventario del patrimonio dell’unità sanitaria locale 30 — area senese. Gli strumenti scientifici. Siena, Tipografia Edizione Cantagalli, 1994, 98 pp. Nuncius, 9(2):901, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x01440.

Abbri:1994:BRTb

[1261] Ferdinando Abbri. Book review: Toby E. Huff, the rise of early modern science. Islam, China, and the West, Cambridge, Cambridge University Press, 1993, xiv-410 pp., ill. Nuncius, 9(2):902–904, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01459.

Segala:1994:AF

[1262] Marco Segala. Astronomia e fisica. Nuncius, 9(2):904–905, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01468.

Ciardi:1994:BRP

[1263] Marco Ciardi. Book review: Peter M. Harman, After Newton: Essays on Natural Philosophy, Aldershot, Variorum, 1993, xii + 315 pp. Nuncius, 9(2):905–907, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x01477.

Abbri:1994:BRSc

[1264] Ferdinando Abbri. Book review: Svante Lindqvist (editor), Center on the Periphery. Historical Aspects of 20th-Century Swedish Physics. Marika Hedin and Thomas Kaiserfeld Associate Editors, Canton MA, Science History Publications, 1993, iv + 516 pp. ill. (≪Uppsala Studies in History of Science≫, 17). Nuncius, 9(2):907–909, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01486.

Peruzzi:1994:BRE

[1265] Giulio Peruzzi. Book review: Enrico Bellone, Spazio e tempo nella nuova scienza, (Italian) [Space and time in the new science], Roma, la Nuova Italia, 1994. Nuncius, 9(2):909–911, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01495.

Tosi:1994:B

[1266] Alessandro Tosi. Botanica. Nuncius, 9(2):911–912, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x01503.

Tosi:1994:BRGb

[1267] Alessandro Tosi. Book review: I Giardini dei Semplici e gli Orti Botanici della Toscana. In appendice saggi su Pietro Andrea Mattioli, (Italian) [The Gardens of the Plain and Botanical Gardens of Tuscany. An appendix of essays by Pietro Andrea Mattioli], a cura di Sara Ferri e Francesca Vannozzi, Perugia, Giunta regionale toscana & Quattroemme, 1993, 214 pp., ill. Nuncius, 9(2):912–914, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01512.

Lippi:1994:BRLb

[1268] Donatella Lippi. Book review: Liana Palazzi Mariotti, Il Giardino dei Semplici, un itinerario fra le piante, aromatiche, medicinali, velenose, esotiche, Firenze, Città di Vita, 1993, 325 pp. Nuncius, 9(2):914–915, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01521.

Tosi:1994:BRR

[1269] Alessandro Tosi. Book review: I ritratti di piante di Jacopo Ligozzi, (Italian) [The portraits of plants by Jacopo Ligozzi], a cura di Lucia Tongiorgi Tomasi, introduzione di Giuseppe Olmi, schede botanico-farmacologiche di Sara Ferri, Pisa, Pacini Editore, 1993, 198 pp., ill. Nuncius, 9(2):915– 916, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01530.

Tosi:1994:BRH

[1270] Alessandro Tosi. Book review: 150-HCI. Collezioni botaniche e ricerca scientifica. Atti del convegno. Il significato delle collezioni d’erbari oper.

Il progresso della ricerca scientifica (Botanical collections and scientific research. Proceedings of the meeting. The role of the herbarium collections for the progress of the botanical research), Firenze, Museo Botanico dell’Università, 1993, 849 pp., ill. Nuncius, 9(2):916–918, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01549.

Lippi:1994:BRPb

[1271] Donatella Lippi. Book review: Pierangelo Lomagno, Storie di piante medicinali eccellenti, (Italian) [Histories of plants: excellent medicines], Saronno, Ciba Edizioni, 1994, 199 pp. Nuncius, 9(2):918–919, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01558.

Malacrida:1994:MB

[1272] Lorenzo Marri Malacrida. Medicina e biologia. Nuncius, 9(2):919–920, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01567.

Caroti:1994:BRJ

[1273] Stefano Caroti. Book review: Jole Agrimi — Chiara Crisciani, Les ≪Consilia≫ médicaux, trad. de l’ital. par C. Viola, Turnhout, Brepols, 1994, 106 pp., bibl., ind. (≪Typologie des sources du moyen age occidental, fasc. 69≫ ). Nuncius, 9(2):920–921, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01576.

Malacrida:1994:BRK

[1274] Lorenzo Marri Malacrida. Book review: Kenneth J. Carpenter, Protein and energy. A study of Changing Ideas in Nutrition, Cambridge, Mass., Cambridge Univ. Press., 1994, 280 pp., fig. Nuncius, 9(2):921– 922, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01585.

Lippi:1994:BRAc

[1275] Donatella Lippi. Book review: Aa.Vv., Doctors and ethics: the earlier historical setting of professional ethics, edited by A. Wear, J. Geyer-Kordesch and R. French, Amsterdam—Atlanta, GA, Rodopi, 1993, 303 pp. Nuncius, 9(2):922–924, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01594.

Abbri:1994:AM

[1276] Ferdinando Abbri. Antichita e medioevo. Nuncius, 9(2):924–925, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01602.

Caroti:1994:BRAb

[1277] Stefano Caroti. Book review: Alfonso Maierù, University Training in Medieval Europe, tr. and ed. by D. N. Pryds, Leiden-New York– Köln, E. J. Brill, 1994, 188 pp., bibl., ind. (≪Education and Society in the Middle Ages and Renaissance, 3≫ ). Nuncius, 9(2):925–927, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01611.

Caroti:1994:BRC

[1278] Stefano Caroti. Book review: Charles W. Jones, Bede, the Schools and the ≪Compotus≫, ed. by W. M. Stevens, Aldershot, Variorum, 1994, x + 356 pp., ind. (≪Collected Studies Series, 436≫ ). Nuncius, 9(2): 927, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01620.

Caroti:1994:BREb

[1279] Stefano Caroti. Book review: Egidio Romano, Opera Omnia. I. Catalogo dei manoscritti (1001–1075) De Regimine principum. 1/11 Città del Vaticano — Italia, a cura di F. Del Punta e C. Luna, Firenze, Leo S. Olschki, 1993, 420 pp., bibl., app., ind. (≪Unione Accademica Nazionale. Corpus Philosophorum Medii Aevi. Testi e studi, XII≫ ). Nuncius, 9 (2):927–928, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539184x01639.

Caroti:1994:BREc

[1280] Stefano Caroti. Book review: Ermete Trismegisto, De triginta sex decanis, cura et studio S. Feraboli, Translatio Gallica capitum XXIV–XXV, cura et studio S. Feraboli et S. Matton, Turnholt, Brepols, 1994, 386 pp., bibl., ill., ind. (≪Corpus Christianorum Continuatio Mediaevalis, Cxliv. Hermes Latinus, Tomus Iv, pars I≫ ). Nuncius, 9(2):928–929, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01648.

Caroti:1994:BRMc

[1281] Stefano Caroti. Book review: Marsilius von Inghen. Werk und Wirkung. Aktendes Zweiten Internationalen Marsilius-von-Inghen-Kongresses, hrsg. von S. Wielgus, Lublin, Redakcja Wydawnictw Kul, 1993, 291 pp., ill., bibl., ind. Nuncius, 9(2):929–931, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01657.

Abbri:1994:R

[1282] Ferdinando Abbri. Rinascimento. Nuncius, 9(2):931–933, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01666.

Caroti:1994:BRAc

[1283] Stefano Caroti. Book review: Aldo Manuzio tipografo. 1494–1515, catalogo a cura di Luciana Bigliazzi, Angela Dillon Bussi, Giancarlo Savino, Piero Scapecchi, Firenze, Octavo Franco Cantini Editore, 1994, 239 pp., fig. Nuncius, 9(2):934, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x01675.

Miniati:1994:BREb

[1284] Mara Miniati. Book review: Enrico Giusti, Carlo Maccagni (a cura di), Luca Pacioli e la matematica del Rinascimento. Firenze, Giunti, 1994, 108 pp., fig. Nuncius, 9(2):934–936, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01684.

Pereira:1994:BRR

[1285] Michela Pereira. Book review: Resultate und Desiderate der Paracelsus-Forschung, hrsg. von Peter Dilg und Hartmut Rudolph, Franz Steiner Verlag, Stuttgart, 1993, 204 pp. (≪Sudhoffs Archiv Beihefte≫ 31), ISBN 3-515-06096-0. Nuncius, 9(2):936–937, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01693.

Miniati:1994:SXX

[1286] Mara Miniati. Secoli XVII–XX. Nuncius, 9(2):937, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539184x01701.

Lippi:1994:BRAd

[1287] Donatella Lippi. Book review: Aa.Vv., Firenze dei grandi viaggiatori, a cura di Franco Paloscia, Casale Monferrato, Ed. Abete, 1994, 191 pp. Nuncius, 9(2):937–938, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x01710.

Abbri:1994:BRG

[1288] Ferdinando Abbri. Book review: George J. Seidel, Fichte’s Wissenschaftslehre of 1794. A Commentary on Part 1, West Lafayette, Indiana, Purdue University Press, 1993, 126 pp. (≪Purdue University Series in the History of Philosophy≫ ). Nuncius, 9(2):938–939, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01729.

Pogliano:1994:BRH

[1289] Claudio Pogliano. Book review: Henri Poincaré, Il valore della scienza, a cura di Gaspare Polizzi, Firenze, La Nuova Italia, 1994, lii + 204 pp. Nuncius, 9(2):939–940, ???? 1994. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539184x01738.

Vista:1994:BRL

[1290] Michele Vista. Book review: Laylin K. James (ed.), Nobel Laureates in Chemistry, 1901–1992. Washington, American Chemical Society and the Chemical Heritage Foundation, 1993, xviii + 798 pp. (History of Modern Chemical Sciences). Nuncius, 9(2):940–942, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01747.

Toscano:1994:BRP

[1291] Anna Toscano. Book review: Paolo Simoncelli, Storia di una censura. ≪Vita di Galileo≫ e Concilio Vaticano II, Milano, Franco Angeli, 1992, 153 pp. (≪Studi e ricerche storiche≫ ). Nuncius, 9(2):942–943, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539184x01756.

Miniati:1994:ADM

[1292] Mara Miniati. Attivita dell’Istituto e Museo di Storia della Scienza di Firenze. Nuncius, 9(2):945–958, ???? 1994. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539184x01765.

Camenietzki:1995:LIA

[1293] Carlos Ziller Camenietzki. L’extase interplanétaire d’Athanasius Kircher. Nuncius, 10(1):3–32, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x00017.

Vlahakis:1995:ANS

[1294] George N. Vlahakis. The appearance of a ≪New≫ science in 18th Century Greece: the case of chemistry. Nuncius, 10(1):33–50, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00026.

Bonechi:1995:PTT

[1295] Simone Bonechi. Un proprietario toscano tra scienza, rivoluzione e filantropismo: Girolamo De’ Bardi (1777–1829). Nuncius, 10(1):51–97, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00035.

Monti:1995:DHM

[1296] Dalida Monti. Dirac’s holes model: From proton to positron. Nuncius, 10(1):99–130, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00044.

Turner:1995:FWG

[1297] Gerard L’E. Turner. The Florentine workshop of Giovan Battista Giusti, 1556–c. 1575. Nuncius, 10(1):131–172, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539185x00053.

Miniati:1995:SRD

[1298] Mara Miniati, Valeria Cocchetti, Caterina Toso, and Sergio Boni. Sul restauro di due globi di Matthäus Greuter. Nuncius, 10(1):173–178, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00062.

Jimenez:1995:RAM

[1299] Josefa Jiménez, Martinez Manuela, and Amparo Sebastian. The Royal Academy of Mathematics and the Imperial College in the National Museum of Science and Technology of Madrid. Nuncius, 10(1):179–192, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00071.

Agnoletti:1995:TFL

[1300] Mauro Agnoletti. Il trasporto fluviale del legname nell’alta valle del piave (xix sec.). Nuncius, 10(1):193–209, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00080.

Ricci:1995:CSG

[1301] Maria Basso Ricci and Pasquale Tucci. La catalogazione di strumenti geomagnetici italiani. Nuncius, 10(1):211–215, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539185x00099.

Gatto:1995:PAD

[1302] Romano Gatto. Per un archivio della corrispondenza degli scienziati italiani lettere di Giuseppe Battaglini a Enrico Betti. Nuncius, 10(1): 217–221, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00107.

Anonymous:1995:LGB

[1303] Anonymous. Lettere di Giuseppe Battaglini ad Enrico Betti. Nuncius, 10(1):222–256, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00116.

Collini:1995:SHN

[1304] Silvia Collini and Antonella Vannoni. La Société d’Histoire Naturelle e il viaggio di D’Entrecasteaux alla ricerca di la Pérouse: Le istruzioni scientifiche per i viaggiatori. Nuncius, 10(1):257–274, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00125.

Delamarck:1995:MIP

[1305] M. Delamarck. Mémoire instructif pour les voyageurs qui se proposent de faire des recherches utiles à la botanique. (French) [Instructive memoir for travellers who intend to do useful research on botany]. Nuncius, 10 (1):275–286, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00134.

Pinel:1995:PQZ

[1306] M. Pinel. Sur le progrès que la zoologie atteind des voyages de long cours. Nuncius, 10(1):287–291, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00143.

Lazzi:1995:COR

[1307] Giovanna Lazzi and Isabella Truci. Le carte di Orazio Ricasoli Rucellai acquistate dalla Biblioteca Nazionale di Firenze. Nuncius, 10(1):293– 305, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00152.

Proverbio:1995:OIH

[1308] Edoardo Proverbio. Observational instruments of historical interest in existence in Italian astronomical observatories. Nuncius, 10(1):307–320, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00161.

Turner:1995:YCC

[1309] A. J. Turner. Yunus the candle-clock maker and Babylonian functions. Nuncius, 10(1):321–323, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00170.

Segre:1995:CIG

[1310] Michael Segre. Convegno internazionale su Galileo Galilei. Nuncius, 10 (1):325–326, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00189.

Pasta:1995:MNV

[1311] Renato Pasta. I matematici nella vita politica, 1789–1848. Nuncius, 10 (1):327–329, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00198.

Bucciantini:1995:NM

[1312] Massimo Bucciantini, Fabio Guidi, and Alessandra Lenzi. Nova media. Nuncius, 10(1):331–343, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00206.

Abbri:1995:BRH

[1313] Ferdinando Abbri. Book review: H. Floris Cohen, the Scientific Revolution. A Historiographical Inquiry, Chicago & London, The University of Chicago Press, 1994, xviii + 662 pp. Nuncius, 10(1):345–347, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00215.

Beretta:1995:BRAa

[1314] Ilva Beretta. Book review: De Arbore. Botanica, alimentazione, architettura, teatro, storia, legislazione, filosofia, simbologia, araldica, religione, letteratura, tecnologia degli alberi dalle opere manoscrittee a stampadella Biblioteca Casanatense. Roma, Biblioteca Casanatense, 1991, 811 pp., Ill. Wilfrid Blunt, and T. William Stearn, the Art of Botanical Illustration, Woodbridge, Suffolk, Antique Collectors’ Club in association with the Royal Botanic Garden, Kew, 1994, 368 pp., Ill. Nuncius, 10(1):347– 351, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00224.

Iannaccone:1995:BRC

[1315] Isaia Iannaccone. Book review: Colin A. Ronan, the Shorter Science and Civilisation in China, vol. 4, Cambridge, Cambridge University Press, 1994, xv + 334 pp., 142 ill. Nuncius, 10(1):351–354, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00233.

Beretta:1995:BRAb

[1316] Marco Beretta. Book review: Aa.Vv., Storia dell’Università di Pisa — vol. 1 1343–1737, 2 t. Pisa, Pacini Editore, 1993, xv + 739 pp., Ill. Nuncius, 10(1):355–358, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x00242.

Beretta:1995:BRGa

[1317] Marco Beretta. Book review: Giorgio Agricola, De Re Metallica, a cura di Paolo Macini e Ezio Mesini, Segrate, Edizioni Anim, 1994, xxxii + 542 pp., Ill. Nuncius, 10(1):358–359, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00251.

Toscano:1995:BRP

[1318] Anna Toscano. Book review: Paola Giacomoni (a cura di), Immagini del corpo in età moderna, Trento, Editrice Università degli Studi di Trento, 1994, 282 pp. (≪Labirinti≫ Collana del Dipartimento di Scienze Filologiche e Storiche). Nuncius, 10(1):359–362, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00260.

Olmi:1995:BRP

[1319] Giuseppe Olmi. Book review: Paula Findlen, Possessing Nature. Museums, Collecting and Scientific Culture in Early Modern Italy, Berkeley-Los Angeles-London, University of California Press, 1994, xv + 449 pp. Nuncius, 10(1):363–365, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00279.

Miniati:1995:BRS

[1320] Mara Miniati. Book review: Silvio A. Bedini, Science and Instruments in Seventeenth-Century Italy, Aldershot, Variorum Ashgate Publishing Limited, 1994, x + 338 pp., Ill. Aldershot, Variorum Ashgate Publishing Limited, 1994, xii + 338 pp., Ill. Nuncius, 10(1):365–368, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00288.

Nardi:1995:BRE

[1321] Antonio Nardi. Book review: Enrico Giusti, Euclides reformatus: la teoria delle proporzione nella scuola galileiana, Milano, Bollati Boringhieri, 1993, 348 pp. Nuncius, 10(1):368–374, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00297.

Beretta:1995:BRAc

[1322] Marco Beretta. Book review: Alessandro Marchetti, della natura delle cose di Lucrezio, a cura di Mario Saccenti, Modena, Mucchi Editore, 1992, xxvii + 275 pp., Ill. Nuncius, 10(1):374–375, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00305.

Pighetti:1995:BRG

[1323] Clelia Pighetti. Book review: Gianfranco Mormino, ≪Penetralia motus≫. La fondazione relativistica della meccanica in Christiaan Huygens, con l’edizione del ≪Codex Hugeniorum 7 A≫, Firenze, La Nuova Italia, 1993, 355 pp. Nuncius, 10(1):375–377, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00314.

Tosi:1995:BRI

[1324] Alessandro Tosi. Book review: Imago et descriptio Tusciae. La Toscana nella geocartografia dal XV al XIX secolo, a cura di Leonardo Rombai, scritti di M. Azzari, P. Crini, L. Rombai, G. C. Romby, D. Toccafondi, R. Valentini, C. Vivoli, Venezia, Giunta regionale Toscana-Marsilio, 1993, 399 pp., Ill. Nuncius, 10(1):377–379, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00323.

Tosi:1995:BRC

[1325] Alessandro Tosi. Book review: Cristina Scappini, Maria Pia Torricelli, Lo Studio Aldrovandi in Palazzo Pubblico (1617–1742), a cura di Sandra Tugnoli Pattaro, Bologna, Clueb, 1993, 180 pp. Nuncius, 10(1):379–381, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00332.

Savorelli:1995:BRB

[1326] Mirella Brini Savorelli. Book review: Buffon88, Actes du Colloque international pour le bicentenaire de la mort de Buffon (Paris, Montebard, Dijon, 14–22 juin 1988), réunis par J.-Cl. Beaune, S. Benoit, J. Gayon, J. Roger, D. Woronoff, direction J. Gayon, Paris, Vrin, 1992, 771 pp. (ISBN 2-7116-9755-X). Nuncius, 10(1):381–388, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00341.

Ciardi:1995:BRS

[1327] Marco Ciardi. Book review: Les scientifiques et la montagne. Actes du 116e Congrès National des Sociétés Savantes, section d’Histoire des Sciences et des Techniques (Chambéry, 29 avril–4 mai 1991), Paris, Édition du CTHS, 1993, 296 pp., ill [ISBN 2-7355-0272-4]. Nuncius, 10(1):388– 391, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00350.

Segala:1995:BRF

[1328] Marco Segala. Book review: Friedrich Wilhelm Joseph Schelling, Historisch-kritische Ausgabe. Ergänzungsband zu Werke Band 5 bis 9. Wissenschaftshistorischer Bericht zu Schellings naturphilosophischen Schriften 1797–1800, Stuttgart, Frommann-Holzboog, 1994, xv + 845 pp. Nuncius, 10(1):392–395, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00369.

Segala:1995:BRD

[1329] Marco Segala. Book review: Dora B. Weiner, The citizen-patient in revolutionary and imperial Paris, Baltimore, The Johns Hopkins University Press, 1993, xvi + 444 pp. (ISBN 0-8018-4483-5). Nuncius, 10 (1):395–398, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00378.

Segala:1995:BRN

[1330] Marco Segala. Book review: Novalis, Opera filosofica, a cura di Fabrizio Desideri e Giampiero Moretti, 2 voll., Torino, Einaudi, 1993, xliii + 620 e ix + 965 pp. (≪Nuova Universale Einaudi≫, 214). Nuncius, 10 (1):398–400, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00387.

Tosi:1995:BRU

[1331] Alessandro Tosi. Book review: Gli Uccelli dei Ranieri. Arte, natura e decorazione tra Settecento e Ottocento, catalogo della mostra, Casalzuigno (Varese), Villa della Porta-Bozzolo, 25 settembre–1 novembre 1994, (Italian) [The Birds of Ranieri. Art, nature and decoration between the Eighteenth and Nineteenth Century, exhibition catalog, Casalzuigno (Varese), Villa della Porta-Cocoon, September 25–November 1 1994]. Milano, Electa, 1994, 134 pp., ill. Nuncius, 10(1):400–401, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00396.

Vidal:1995:BRA

[1332] Fernando Vidal. Book review: Alan Gauld, A history of hypnotism, Cambridge, Cambridge University Press, 1992, xv + 738 pp. Nuncius, 10(1):401–402, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00404.

Abbri:1995:BRAa

[1333] Ferdinando Abbri. Book review: Alan J. Rocke, the Quiet Revolution. Hermann Kolbe and the Science of Organic Chemistry, Berkeley-Los Angeles, University of California Press, 1993, xvi + 502 pp. (≪California Studies in the History of Science≫ ). Nuncius, 10(1):402– 405, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00413.

Scarpelli:1995:BRG

[1334] Giacomo Scarpelli. Book review: Geoffrey Irwin, The Prehistoric Exploration and Colonisation of the Pacific, Cambridge, Cambridge University Press, 1992, vii + 240 pp. ill [ISBN 0-521-40371-5]. Nuncius, 10 (1):405–407, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00422.

Corbellini:1995:BRB

[1335] Gilberto Corbellini. Book review: A. L. Bonella (a cura di), L’ospedale dei pazzi di Roma dai papi al 900. Vol. I, Fonti per la storia della follia: Santa Maria della Pietà e il suo archivio storico — secc. XVI–XX, Roma, Edizioni Dedalo, 296 pp., 32 ill. F. F. Bernardini, A. Iaria, A. Bonfigli (a cura di), l’ospedale dei pazzi di Roma dai papi al 900. Vol. II, Lineamenti di assistenza e cura a poveri e dementi, Roma, Edizioni Dedalo, 448 pp., 36 ill. a colori e 191 b.n. Nuncius, 10(1):407–410, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00431.

Vista:1995:BRD

[1336] Michele Vista. Book review: Dietrich Stoltzenberg, Fritz Haber: Chemiker, Nobelpreisträger, Deutscher Jude, Weinheim, VCH, 1994, xiv + 670 pp., ill. Nuncius, 10(1):410–413, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539185x00440.

Somenzi:1995:BRD

[1337] Vittorio Somenzi. Book review: Dean Keith Simonton, Scientific Genius (A psychology of science), Cambridge, Cambridge University Press, 1988, x + 229 pp. Nuncius, 10(1):413–414, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00459.

Miniati:1995:SBa

[1338] Mara Miniati. Strumenti bibliografici. Nuncius, 10(1):415–416, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00468.

Miniati:1995:IS

[1339] Mara Miniati. Istituzioni scientifiche. Nuncius, 10(1):416–418, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00477.

Abbri:1995:BRD

[1340] Ferdinando Abbri. Book review: David Cahan, An Institute for an Empire. The Physikalisch-Technische Reichsanstalt 1871–1918, Cambridge, Cambridge University Press, 1989, xx + 315 pp. Nuncius, 10(1):418– 419, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00486.

Miniati:1995:S

[1341] Mara Miniati. Strumentaria. Nuncius, 10(1):419–420, ???? 1995. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539185x00495.

Miniati:1995:BRRa

[1342] Mara Miniati. Book review: R. H. Van Gent, De hemel in de hand. Twee astrolaben van het museum Boerhaave. The Portable Universe. Two Astrolabes of the Museum Boerhaave, Leiden, Museum Boerhaave, 1994, 44 pp., fig. Nuncius, 10(1):420, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00503.

Miniati:1995:BRLa

[1343] Mara Miniati. Book review: Liceo Ginnasio ≪Corti Gentili≫ Alatri (Fr), Il filo del tempo: l’antico laboratorio fisico. Instrumenta selecta, a cura di Roberto Mantovani. Alatri, Arti Grafiche Tofani s.n.c., 1994, xxx + 88 pp., ill. Musis Museo della Scienza e Dell’Informazione Scientifica, Collezione strumenti di fisica Liceo E. Q. Visconti-Roma, a cura di Alessandro Orlandi e William Russell, fotografie di Sebastiano Porretta, Roma, Euroma Editrice Universitaria di Roma-La Goliardica, 1994, 82 pp., ill. Nuncius, 10(1):421–422, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00512.

Morini:1995:FDS

[1344] Simona Morini. Filosofia della scienza. Nuncius, 10(1):422–423, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00521.

Casati:1995:SDI

[1345] Stefano Casati. Storia delle idee. Nuncius, 10(1):423–426, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00530.

Abbri:1995:BRJa

[1346] Ferdinando Abbri. Book review: Jamie James, the Music of the Spheres. Music, Science, and the Natural Order of the Universe, London, Little, Brown and Company, 1994, xix + 262 pp. Nuncius, 10(1):426–427, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00549.

Caroti:1995:M

[1347] Stefano Caroti. Matematica. Nuncius, 10(1):427–428, ???? 1995. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539185x00558.

Segala:1995:BRU

[1348] Marco Segala. Book review: Umberto Bottazzini, Il flauto di Hilbert. Storia della matematica moderna e contemporanea, (Italian) [Hilbert’s flute. History of modern and contemporary mathematics], Torino, Utet, 1990, v + 455 pp. Nuncius, 10(1):428, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00567.

Segala:1995:BRMa

[1349] Marco Segala. Book review: Morris Kline, Storia del pensiero matematico, 2 voll., Torino, Einaudi, 1991, xiv + 1470 pp. Nuncius, 10(1): 428–429, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00576.

Miniati:1995:BRPa

[1350] Mara Miniati. Book review: Per l’archivio della corrispondenza dei matematici italiani. Il fondo ≪Tacchini≫ dell’Ufficio Centrale di Ecologia Agraria in Roma, (Italian) [On the archive of correspondence of Italian mathematicians. Il fondo ≪Tacchini≫ of the Central Bureau of Agricultural Ecology in Rome], a cura di Ileana Chinnici, Milano, Università Bocconi, 1994, 111 pp., ill. (Quaderni P.Ri.St.Em. n. 7). Nuncius, 10 (1):429–430, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00585.

Tosi:1995:B

[1351] Alessandro Tosi. Botanica. Nuncius, 10(1):430, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539185x00594.

Tosi:1995:BRM

[1352] Alessandro Tosi. Book review: Maria Augusta Favalia, Fabrizia Fossati, Giambattista Guatteri. Fondatore dell’attuale Orto Botanico di Parma, Parma, 1993, 139 pp., ill. Nuncius, 10(1):431, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00602.

Anonymous:1995:M

[1353] Anonymous. Medicina. Nuncius, 10(1):432–433, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539185x00611.

Miniati:1995:BRGa

[1354] Mara Miniati. Book review: Gabriella Mancini (a cura di), l’Officina profumo-farmaceutica di Santa Maria Novella in Firenze: sette secoli di storia e di arte, Roma, F.lli Chitarrini s.a.s., 1994, 220 pp., ill. Nuncius, 10(1):433–434, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00620.

Miniati:1995:BRF

[1355] Mara Miniati. Book review: Franco Arosio, Carlo Besta and the Foundation of the National Neurological Institute in Milan, Milan, ≪Carlo Besta≫ National Neurological Institute, 1993, 47 pp. Franco Arosio, Carlo Besta 1876–1940, Milano, Istituto Nazionale Neurologico ≪Carlo Besta≫, 1993, 111 pp., ill. Nuncius, 10(1):434–435, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00639.

Caroti:1995:MR

[1356] Stefano Caroti. Medioevo e rinascimento. Nuncius, 10(1):435–436, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00648.

Caroti:1995:BRL

[1357] Stefano Caroti. Book review: Laura Ackerman Smoller, History, Prophecy and the Stars. The Christian Astrology of Pierre d’Ailly, 1350– 1420, Princeton, New Jersey, Princeton University Press, 1994, 233 pp., bibl., ind. Nuncius, 10(1):436–437, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00657.

Caroti:1995:BRC

[1358] Stefano Caroti. Book review: Constantine the African and Ali Ibnal-Abbasal-Magusi. The ≪Pantegni≫ and Related Texts, ed. by Charles Burnett and Danielle Jacquart, Leiden New York–Köln, E. J. Brill, 1994, 364 pp., ill., ind. (≪Studies in Ancient Medicine≫, 10). Nuncius, 10 (1):437–439, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00666.

Caroti:1995:BRA

[1359] Stefano Caroti. Book review: Aegidius Aurelianensis, Quaestiones super ≪De generatione et corruptione≫, Herausgegeben von Zdzisław Kuksewicz, Amsterdam/Philadelphia, B. R. Grüner, 1993, 237 pp., ind. (≪Bochumer Studien zur Philosophie≫, 18). Nuncius, 10(1):439–440, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00675.

Caroti:1995:BRPa

[1360] Stefano Caroti. Book review: Pico, Poliziano e l’Umanesimo di fine Quattrocento. Biblioteca Medicea Laurenziana 4 novembre–31 dicembre 1994, Catalogo a cura di Paolo Viti, Firenze, Leo S. Olschki, 1994, 376 pp., ill. ind. (≪Centro Internazionale di cultura Giovanni Pico della Mirandola. Studi Pichiani≫, 2). Nuncius, 10(1):440–441, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00684.

Abbri:1995:SXX

[1361] Ferdinando Abbri. Secoli XVII–XX. Nuncius, 10(1):441–442, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00693.

Abbri:1995:BRJb

[1362] Ferdinando Abbri. Book review: John Leary, Jr., Francis Bacon and the Politics of Science, Ames, Iowa State University Press, 1994, xii + 297 pp. Nuncius, 10(1):442–444, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00701.

Toscano:1995:BRN

[1363] Anna Toscano. Book review: ≪Stenoniana. Nova Series≫, I, Copenhagen, Lægeforeningens Forlag, 1991, 159 pp. Nuncius, 10(1):444–446, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00710.

Abbri:1995:BRRa

[1364] Ferdinando Abbri. Book review: Renè Descartes, Opere filosofiche, a cura di Ettore Lojacono, Torino, Utet, 1994, 2 voll., 920 + 777 pp. (≪Classici della Filosofia≫ ). Nuncius, 10(1):446–448, ????

1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00729.

Abbri:1995:BRG

[1365] Ferdinando Abbri. Book review: Gunnar Eriksson, the Atlantic Vision. Olaus Rudbeck and Baroque Science, Canton MA, Science History Publications, 1994, viii + 196 pp. (≪Uppsala Studies in History of Science≫, 19). Nuncius, 10(1):448–450, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00738.

Ciardi:1995:BRR

[1366] Marco Ciardi. Book review: A. Rupert Hall, Newton, his Friends and his Foes, Aldershot, Variorum, 1993, xii + 330 pp. A. Rupert Hall, Science and Society, Aldershot, Variorum, 1994, x + 324 pp., ill. Nuncius, 10 (1):450–452, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00747.

Abbri:1995:BRAb

[1367] Ferdinando Abbri. Book review: Antoni Van Leeuwenhoek, Alle de Brieven (The Collected Letters), Geredigeerd doors (Edited by) L. C. Palm, Deel (Volume) XIII (1700–1701), Lisse, Swets & Zeitlinger, 1994, 417 pp. Nuncius, 10(1):452–453, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00756.

Beretta:1995:BRP

[1368] Marco Beretta. Book review: Pierluigi Barrotta, Dogmatismo ed eresia nella scienza: Joseph Priestley, Milano, Franco Angeli, 1994, 183 pp. Nuncius, 10(1):453–455, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00765.

Miniati:1995:BRD

[1369] Mara Miniati. Book review: Donato Granafei, Meccanica animale. Esercizio Fisico-matematico a cura di Arcangelo Rossi, Lecce, Edizioni del Grifo, 1993, xxiv + xvi + 38 pp., 1 tav. n.n. (≪Testi e Ricerche≫, 9). Nuncius, 10(1):455, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x00774.

Abbri:1995:BRRb

[1370] Ferdinando Abbri. Book review: R. E. R. Banks, B. Elliott, J. G. Hawkes, D. King-Hele, G. U. Lucas (editors), Sir Joseph Banks: A Global Perspective, Kew, the Royal Botanic Gardens, 1994, ii + 235 pp. Nuncius, 10(1):455–456, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00783.

Toscano:1995:BRMa

[1371] Anna Toscano. Book review: Maurizio Torrini, Scienziati a Napoli. 1830–1845, Napoli, Cuen, 1989, 166 pp., ill. Nuncius, 10(1):457–458, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00792.

Miniati:1995:BRGb

[1372] Mara Miniati. Book review: Giorgio Santha, La Provincia Scolopica d’Ungheria. Attività apostolica culturale e sociale dal 1642 al 1956, in ≪Ricerche. Bollettino quadrimestrale degli Scolopi italiani≫, Anno XIV (1994), n. 3, pp. 215-346. Nuncius, 10(1):458–459, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00800.

Miniati:1995:BRPb

[1373] Mara Miniati. Book review: Pietro Segala (a cura di), Università, ricerca e nuove professionalità per la durabilità dell’arte nel museo del mondo. Interventi e documenti dal seminario bresciano del febbraio 1992, Milano, Brescia, Confcooperative Lombardia, Fondazione Civiltà Bresciana, 1994, 191 pp. (≪Sistema Museale: problemi, uomini, territori≫, 1). Nuncius, 10(1):459–461, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00819.

Anonymous:1995:ADMa

[1374] Anonymous. Attività dell’Istituto e Museo di Storia Della Scienza di Firenze. Nuncius, 10(1):463–476, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00828.

Aquilecchia:1995:MSG

[1375] Giovanni Aquilecchia. I massimi sistemi di Galileo e la cena di Bruno. (Italian) [The world systems of Galileo and the Last Supper of Bruno].

Nuncius, 10(2):485–496, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00837.

diBrozolo:1995:VTG

[1376] Luigi A. Radicati di Brozolo. Variazioni su un tema di Galileo. Nuncius, 10(2):497–508, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00846.

Field:1995:PDF

[1377] J. V. Field. Piero della Francesca and the ≪Distance Point Method≫ of perspective construction. Nuncius, 10(2):509–530, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00855.

Tosi:1995:UII

[1378] Alessandro Tosi. ≪Ulyssis Imago≫: Iconografia Aldrovandiana tra XVI e XIX secolo. (Italian) [≪Ulyssis Imago≫: Aldrovandian iconography from the 17th to 20th Century]. Nuncius, 10(2):531–550, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00864.

diLoreto:1995:ISP

[1379] Mario di Loreto. L’≪Inventaire De Stockholm≫ e il ≪Primo Registro≫ di descartes. Note in margine alle opere postume sulle matematiche. Nuncius, 10(2):551–615, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x00873.

Ciancio:1995:AFP

[1380] Luca Ciancio. Alberto Fortis e la pratica del viaggio naturalistico. Stile di ricerca e modalità di prova. Nuncius, 10(2):617–644, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00882.

Vasconi:1995:SEN

[1381] Paola Vasconi. System Der Empirischen Naturwissenschaften Und Einheit Der Erkenntnis: Die Funktion Der Chemischen Revolution Lavoisiers in Der Reflexion Des Opus Postumum Über Die Sprache Der Wissenschaft. (German) [The system of empirical science, and unit of knowledge: The function of the chemical revolution. Lavoisier in the reflection of Opus Postumum on the language of science]. Nuncius, 10(2):645–669, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00891.

Lualdi:1995:PPO

[1382] Alberto Lualdi. Pietro Patroni, un ottico della Milano Post-Spagnola. (Italian) [Pietro Patroni, an optician of the Milan Post-Spagnola]. Nuncius, 10(2):671–689, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x00909.

Betti:1995:PAD

[1383] Gian Luigi Betti. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 10(2):691–714, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00918.

Cavazza:1995:IF

[1384] Marta Cavazza. Istituzioni e fonti. Nuncius, 10(2):715–753, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00927.

Mocchi:1995:DC

[1385] Giuliana Mocchi. Discussioni critiche. Nuncius, 10(2):755–767, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00936.

Toscano:1995:BRF

[1386] Anna Toscano. Book review: François Russo, Libres propos sur l’Histoire des Sciences, Paris, Éditions Albert Blancard, 1995, 283 pp. (≪Bibliothèque Scientifique Albert Blanchard≫ ). Nuncius, 10(2):769– 770, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00945.

Beretta:1995:BRV

[1387] Marco Beretta. Book review: Valentin Wehefritz & Zoltan Kováts, Bibliography on the History of Chemistry and Chemical Technology. 17th to the 19th Century. K. G. Saur, München, 1994, 3 voll. Nuncius, 10 (2):771–772, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00954.

Toscano:1995:BRMb

[1388] Anna Toscano. Book review: M. Fattori e M. L. Bianchi (a cura di), Ratio. VII Colloquio Internazionale. Roma 9–11 gennaio 1992. Atti, Firenze, Leo S. Olschki Editore, 1994, 574 pp. (Lessico Intellettuale Europeo Lxi). Nuncius, 10(2):772–775, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x00963.

diPasquale:1995:BRGa

[1389] Giovanni di Pasquale. Book review: Gianni Micheli, Le origini del concetto di macchina, Firenze, Leo S. Olschki, 1995, 184 pp., fig. Nuncius, 10(2):776–779, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x00972.

diPasquale:1995:BRGb

[1390] Giovanni di Pasquale. Book review: Giusto Traina, La tecnica in Grecia e a Roma, Roma, Laterza 1994, 155 pp., ill. Nuncius, 10(2):779–781, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00981.

Touwaide:1995:BRC

[1391] Alain Touwaide. Book review: Caelius Aurelianus, Akute Krankheiten Buch I–III. Chronische Krankheiten Buch I–V. Herausgegeben von Gerhard Benz. Übersetzt von Ingeborg Pape. Teil I: Akute Krankheiten I–III; Chronische Krankheiten I–II (= Corpus Medicorum Latinorum, VI, 1). Berlin, Akademie Verlag, 1990, in 8^{∘}, 677 p. Nuncius, 10(2):782–786, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x00990.

Pereira:1995:BRM

[1392] Michela Pereira. Book review: Michael R. Mcvaugh, Medicine before the Plague. Practitioners and their Patients in the Crown of Aragon 1285–1345, Cambridge, Cambridge Up, 1993, 280 pp. (Cambridge History of Medicine Series). Nuncius, 10(2):786–788, ????

1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01007.

Amelin:1995:BRR

[1393] Olov Amelin. Book review: R. G. W. Anderson, J. A. Bennett, W. F. Ryan (eds.), Making Instruments Count: Essays on Historical Scientific Instruments. Presented to Gerard l’Estrange Turner, Aldershot, Variorum, 1993, xix + 492 pp. Nuncius, 10(2):789–790, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01016.

Pogliano:1995:BRF

[1394] Claudio Pogliano. Book review: Flavio Caroli, Storia della fisiognomica. Arte e psicologia da Leonardo a Freud, Milano, Leonardo (Mondadori), 1995, 287 pp. (con 133 figure nel testo). Nuncius, 10(2):790– 793, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01025.

Giuntini:1995:BRJ

[1395] Chiara Giuntini. Book review: John H. Brooke, Science and Religion. Some Historical Perspectives, Cambridge, Cambridge University Press, 1991, x + 422 pp. (The Cambridge History of Science Series). Nuncius, 10(2):793–797, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01034.

Pereira:1995:BRJ

[1396] Michela Pereira. Book review: Joachim Telle (hrsg.), Analecta Paracelsica. Studien zum Nachleben Theophrast von Hohenheims im deutschen Kulturgebiet der frühen Neuzeit, Stuttgart, Franz Steiner Verlag, 1994, 590 pp. Nuncius, 10(2):797–799, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01043.

Pereira:1995:BRW

[1397] Michela Pereira. Book review: William Eamon, Science and the Secrets of Nature. Books of Secrets in Early Modern Culture, Princeton (New Jersey), Princeton Up, 1994, 490 pp. Nuncius, 10(2):799–803, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01052.

Gorman:1995:BRJ

[1398] Michael John Gorman. Book review: James M. Lattis, Between Copernicus and Galileo: Christoph Clavius and the collapse of Ptolemaic cosmology. Chicago and London, University of Chicago Press, 1994, xix + 293 pp. Nuncius, 10(2):803–808, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01061.

Nardi:1995:BRM

[1399] Antonio Nardi. Book review: Michele Camerota, Gli scritti De motu antiquiora di Galileo Galilei: il Ms Gal 71. Un’analisi storico-critica, prefazione di Alberto Papa, Cagliari, Cuec, 1992, 149 pp. Nuncius, 10 (2):808–812, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01070.

Redondi:1995:BRM

[1400] Pietro Redondi. Book review: Mario Biagioli, Galileo Courtier. The practice of science in the culture of absolutism, Chicago and London, The University of Chicago Press, 1993, xn-402 pp., bib., ill., ind. Nuncius, 10(2):812–817, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01089.

Clericuzio:1995:BRB

[1401] Antonio Clericuzio. Book review: Beverley C. Southgate, ≪Covetous of Truth≫. the Life and Work of Thomas White, 1593–1676, Dordrecht, Kluwer Academic Publishers, 1993, 189 pp. (Archives Internationales d’Histoire des Idées, vol. Cxxxiv). Nuncius, 10(2):817–819, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01098.

Abbri:1995:BRJc

[1402] Ferdinando Abbri. Book review: Jost Weyer, Graf Wolfgang II. von Hohenlohe und die Alchemie. Alchemische Studien in Schloss Weikersheim 1587–1610, Sigmaringen, Jan Thorbecke Verlag, 1992, 516 pp. (≪Forschungen aus Württembergisch Franken≫, Band 39). Nuncius, 10(2):819–824, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01106.

Bernardi:1995:BRR

[1403] Walter Bernardi. Book review: Roger French, William Harvey’s natural philosophy, Cambridge, Cambridge University Press, 1994, 386 pp. Nuncius, 10(2):824–827, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x01115.

Abbri:1995:BRS

[1404] Ferdinando Abbri. Book review: Steven Shapin, A Social History of Truth. Civility and Science in Seventeenth-Century England, Chicago and London, The University of Chicago Press, 1994, xxxii + 483 pp. Nuncius, 10(2):828–830, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x01124.

Beretta:1995:BRM

[1405] Marco Beretta. Book review: Michael Stolberg, Ein Recht auf saubere Luft? Umweltkonflikte am Beginn des Industriezeitalters, Erlangen, Harald Fischer Verlag, 1994, 365 pp. Nuncius, 10(2):831–832, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01133.

Abbri:1995:BRMa

[1406] Ferdinando Abbri. Book review: Maurice Crosland, in the Shadow of Lavoisier: the Annales de Chimie and the establishment of a new science, Oxford, the Alden Press, 1994, xii + 354 pp. (≪British Society for the History of Science Monographs≫, 9). Nuncius, 10(2):832–834, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01142.

LaVergata:1995:BRF

[1407] Antonello La Vergata. Book review: F. Burkhardt et alii (eds.), The Correspondence of Charles Darwin. Volume 7, 1858–1859. Supplement to the Correspondence, 1821–1857, Cambridge, Cambridge University Press, 1991, xxviii + 671 pp. (14 illustr.), ISBN 0-521-38564-4. Nuncius, 10(2):834–836, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01151.

Scarpelli:1995:BRZ

[1408] Giacomo Scarpelli. Book review: The Zoology of the Voyage of H.M.S. Beagle, under the Command of Captain Fitzroy R.N., during the Years 1832 to 1836, Edited and Superintended by Charles Darwin Esq., M.A., F.R.S., Sec. G.S., Naturalist to the Expedition [London, Smith, Elder & Co. 1839–1843], 4 voll., rist. anastatica a cura della Royal Geographical Society, London, C.I.L. Ltd., 1994. Nuncius, 10(2):836–838, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01160.

Giuntini:1995:BRD

[1409] Chiara Giuntini. Book review: Dorothy Ross (ed.), Modernist Impulses in the Human Sciences, 1870–1930, Baltimore & London, John Hopkins University Press, 1994, xii + 379 pp. Nuncius, 10(2):838–841, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01179.

Nicolini:1995:BRS

[1410] Nicoletta Nicolini. Book review: La science sous le Troisième Reich, sous la direction de Josiane Olff-Nathan, Paris, Ed. du Seuil, 1993, 336 pp. Nuncius, 10(2):841–844, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01188.

Miniati:1995:BRI

[1411] Mara Miniati. Book review: Ivan Karp, Steven D. Lavine (a cura di), Culture in mostra. Poetiche e politiche dell’allestimento museale, (Italian) [Culture on display. Poetics and policies of the museum]. Introduzione di Fredi Drugman, Bologna, Clueb, 1995, xv + 183 pp., ill. (Lexis, IV ≪MuseoPoli≫, 1). Nuncius, 10(2):844–847, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01197.

Barone:1995:BRE

[1412] Vincenzo Barone. Book review: Erich Robert Paul, the Milky Way Galaxy and Statistical Cosmology, 1890–1924, Cambridge, Cambridge University Press, 1994, 262 pp. Nuncius, 10(2):847–854, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01205.

Polizzi:1995:BRR

[1413] Gaspare Polizzi. Book review: Roland Omnès, Philosophie de la science contemporaine, Paris, Gallimard, 1994, 426 pp. (≪Collection folio/essais≫ ). Nuncius, 10(2):854–859, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539185x01214.

Parigi:1995:BRA

[1414] Silvia Parigi. Book review: Alain Prochiantz, La costruzione del cervello (introd. di D. Lecourt), Roma-Napoli, Theoria, 1992 (ed. orig. 1989), 105 pp. (≪Sonde≫, 36). Nuncius, 10(2):859–861, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01223.

Miniati:1995:BRGc

[1415] Mara Miniati. Book review: Giovanni Paltrinieri, Italo Frizzoni, Meridiane e orologi solari di Bologna e provincia, Bologna, l’Artiere Edizionitalia, 1995, 497 pp., ill. Nuncius, 10(2):861–863, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01232.

Miniati:1995:SBb

[1416] Mara Miniati. Strumenti bibliografici. Nuncius, 10(2):865, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01241.

Casati:1995:BRG

[1417] Stefano Casati. Book review: Gabriel Naudé, Avvertenze per la costituzione di una biblioteca. Introduzione, traduzione e note di Vittoria Lacchini, Bologna, Clueb, 1992, Lii, 124 pp. (Bibliosofia, 1). Nuncius, 10(2):866–867, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01250.

Barsanti:1995:BRY

[1418] Giulio Barsanti. Book review: Yvonne Letouzey, Le Jardin des plantes à la croisée des chemins avec André Thouin (1747–1824), Paris, Éditions du Muséum National d’Histoire Naturelle, 1989, 678 pp. Nuncius, 10 (2):867–868, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01269.

Miniati:1995:BRRb

[1419] Mara Miniati. Book review: Roberta Gibertoni, Annalisa Melodi, Giulia Luppi (a cura di), Il Museo della Bilancia a Campogalliano. With English Text, Milano, Electa, 1995, 87 pp., ill. (Guide Artistiche Electa). Nuncius, 10(2):868–869, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x01278.

Toscano:1995:BRS

[1420] Anna Toscano. Book review: Storia delle Scienze. Gli Strumenti, a cura di Gerard l’E. Turner, Torino, Giulio Einaudi Editore, 1991, 594, pp., ill. Nuncius, 10(2):869–871, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01287.

Miniati:1995:BRM

[1421] Mara Miniati. Book review: Margarida Archinard, La ≪Navicula de Venetiis≫ du Musée d’Histoire des Sciences de Genève, in ≪Musées de Genève≫, 334, avril 1995, pp. 2–8, ill. Nuncius, 10(2):871, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01296.

Miniati:1995:BRK

[1422] Mara Miniati. Book review: Het koperen kabinet. Schatkamers van de wetenschap, 1550–1950, (Dutch) [The copper cabinet. Treasuries of science, 1550–1950], Leiden, Museum Boerhaave, 1994, 64 pp., ill. Nuncius, 10(2):872, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01304.

Miniati:1995:BRU

[1423] Mara Miniati. Book review: Università degli Studi di Padova. Dipartimento di Fisica ≪Galileo Galilei≫, Duecento anni di elettricità, Padova, Museo di Storia della Fisica, 1995, 298 pp. n.n., ill. Nuncius, 10(2):872– 873, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01313.

Miniati:1995:BRLb

[1424] Mara Miniati. Book review: Liceo Scientifico Leonardo Da Vinci Firenze, Intorno al laboratorio di fisica di un liceo scientifico. Notizie e riflessioni, Firenze, Provincia di Firenze, Assessorato Pubblica Istruzione, 1995, 39 pp. (≪Bollettino del centro di documentazione didattica≫, 8). Nuncius, 10(2):873–874, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01322.

Segala:1995:BRMb

[1425] Marco Segala. Book review: Michael Redhead, From physics to metaphysics, Cambridge, Cambridge University Press, 1995, xiii + 92 pp. Nuncius, 10(2):874–875, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01331.

Segala:1995:BRP

[1426] Marco Segala. Book review: Pier Luigi Baima Bollone, La scienza nel mondo degli spiriti, Torino, Sei, 1995, x + 298 pp. Nuncius, 10(2): 875–877, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01340.

Abbri:1995:BRMb

[1427] Ferdinando Abbri. Book review: Margaret C. Jacob, Massoneria illuminata. Politica e cultura nell’Europa del Settecento, Torino, Einaudi, 1995, xii + 390 pp. (≪Biblioteca di cultura storica≫, 206). Nuncius, 10 (2):877–878, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01359.

Collini:1995:BRE

[1428] Silvia Collini. Book review: Eric J. Leed, La mente del viaggiatore. Dall’Odissea al turismo globale, Bologna, Il Mulino, 1992, 386 pp. Nuncius, 10(2):878–879, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x01368.

Miniati:1995:BREa

[1429] Mara Miniati. Book review: Emanuela Guidoboni, Catalogue of ancient earthquakes in the Mediterranean area up to the 10th century, Rome, Istituto Nazionale di Geofisica, 1994, 504 pp., ill. Nuncius, 10(2):879–880, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01377.

Miniati:1995:BREb

[1430] Mara Miniati. Book review: Elisabetta Cioppi, 18 Maggio 1895. Storia di un terremoto fiorentino, Firenze, Osservatorio Ximeniano, 1995, 305 pp. ill. Nuncius, 10(2):880–881, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01386.

Lippi:1995:BRAa

[1431] Donatella Lippi. Book review: Aa.Vv., Storia della Medicina. Prolusioni di G. Armocida, E. Bicheno, B. Fox. Dizionario di S. Musitelli, Milano, Jaca Book, 1993, pp. 396, L. 145.000. Nuncius, 10(2):881–882, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01395.

Lippi:1995:BRT

[1432] Donatella Lippi. Book review: Thierry Bardinet, Les papyrus médicaux de l’Egypte pharaonique. Traductionintégrale et commentaire, Paris, Fayard, 1995, 591 pp. Nuncius, 10(2):882–884, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01403.

Malacrida:1995:BRR

[1433] Lorenzo Marri Malacrida. Book review: Roy Porter, Mikulàs̊ Teich, Drugs and narcotics in history, Cambridge, Cambridge University Press, 1995, 227 pp. Nuncius, 10(2):884–886, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539185x01412.

Lippi:1995:BRL

[1434] Donatella Lippi. Book review: Luciano Sterpellone, Mahmoud Salem Elsheikh, La medicina araba. l’arte medica nei Califfati d’Oriente e d’Occidente, Saronno, Ciba Edizioni, 1995, 344 pp. Nuncius, 10(2):886– 887, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01421.

Miniati:1995:BRN

[1435] Mara Miniati. Book review: Nelli-Elena Vanzan Marchini, Le leggi di sanità della Repubblica di Venezia, volume primo, Vicenza, Neri Pozza Editore, 1995, 503 pp. (≪Fonti per la storia della sanità≫, 2/1). Nuncius, 10(2):887–888, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x01430.

Lippi:1995:BRAb

[1436] Donatella Lippi. Book review: Annalucia Forti Messina, Il soldato in ospedale. I servizi di sanità dell’esercito italico (1796–1814), Milano, F. Angeli, 1991, pp. 355. Nuncius, 10(2):888–890, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01449.

Miniati:1995:BRPc

[1437] Mara Miniati. Book review: Paolo Stefani, Francesca Vannozzi, Fausto Casi, l’avventura dei Raggi X (1895–1995). Dal radiologo universale… all’universo dei Raggi X, Firenze, Giorgi & Gambi, 1995, 238 pp., ill. Nuncius, 10(2):890–892, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01458.

Lippi:1995:BRB

[1438] Donatella Lippi. Book review: Bruno Bruni, Il Museo del Diabete. Prefazione di Renato Bèttica Giovannini. Torino, Centro Karen Bruni Bocher, 1995, 153 pp. Nuncius, 10(2):892–893, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01467.

Beretta:1995:BRGb

[1439] Marco Beretta. Book review: Gareth Roberts, the Mirror of Alchemy. Alchemical Ideas and Images in Manuscripts and Books from Antiquity to the Seventeenth Century, London, the British Library, 1994, 128 pp., Ill. Nuncius, 10(2):893, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x01476.

Beretta:1995:BRAd

[1440] Marco Beretta. Book review: Il y a 200 ans Lavoisier. Actes du Colloque organisé à l’occasion du bicentenaire de la mort d’Antoine Laurent Lavoisier, le 8 mai 1794, (French) [Lavoisier 200 years ago. Proceedings of the conference organized on the occasion of the bicentenary of the death of Antoine Laurent Lavoisier, 8 May 1794], Paris, Lavoisier, 1994, xi + 335 pp. Ill. Nuncius, 10(2):893–894, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01485.

Abbri:1995:BRJd

[1441] Ferdinando Abbri. Book review: Johan Gottlieb Gahn, Brev. Utgivna med kommentarer av Jan Trofast. Del II, Lund, Wallin & Dalholm Boktryckeri, 1994, 347 pp. Nuncius, 10(2):894–895, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01494.

Caroti:1995:BRMa

[1442] Stefano Caroti. Book review: Manuels, programmes de cours et techniques d’enseignement dans les universités médiévales. Actes du Colloque international de Louvain-la-Neuve (9–11 septembre 1993), éd. par J. Hamesse, Louvain-la-Neuve, Institut d’Études Médiévales, 1994 (≪Université Catholique de Louvain. Publication de l’Institut d’Études Médiévales. Textes, Études, Congrès, vol. 16), 477 pp., ill., ind. Nuncius, 10(2):895–896, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x01502.

Caroti:1995:BRG

[1443] Stefano Caroti. Book review: George Molland, Mathematics and the Medieval Ancestry of Physics, Aldershot, Variorum, 1995, xii + 336 pp. (≪Collected Studies, Cs 481≫ ). Nuncius, 10(2):896–897, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01511.

Caroti:1995:BRO

[1444] Stefano Caroti. Book review: Olga Weijers, Le travail intellectuel à la Faculté des arts de Paris: textes et maîtres (ca. 1200– 1500). Répertoire des noms commençant par A-B, Turnhout, Brepols, 1994 (≪Studia Artistarum. Études sur la Faculté des arts dans les Universités médiévales, 1≫ ), 92 pp. Nuncius, 10(2):898, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01520.

Caroti:1995:BRMb

[1445] Stefano Caroti. Book review: M. J. F. M. Hoenen, J. H. J. Schneider, G. Wieland (eds.), Philosophy and Learning. Universities in the Middle Ages, Leiden-New York–Köln, E. J. Brill, 1995 (≪Education and Society in the Middle Ages and Renaissance, 6≫ ), 435 pp., bibl., ind. Nuncius, 10(2):898–900, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01539.

Caroti:1995:BRB

[1446] Stefano Caroti. Book review: B. Pabst, Atomtheorien des lateinischen Mittelalters, (German) [Atomic theories of the Latin Middle Ages], Darmstadt, Wissenschaftliche Buchgesellschaft, 373 pp., app., bibl., ind. Nuncius, 10(2):900–901, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01548.

Caroti:1995:BRJ

[1447] Stefano Caroti. Book review: J.-P. Boudet, Lire dans le ciel. La bibliothèque de Simon de Phares, astrologue du XVe siècle, Bruxelles, Centre d’étude des manuscrits, 1994 (≪Les Publications de Scriptorium, vol. X≫ ), 220 pp., app., ind., ill. Nuncius, 10(2):901, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01557.

Caroti:1995:BRPb

[1448] Stefano Caroti. Book review: P. O. Kristeller, Latin Manuscript Books Before 1600. A List of the Printed Catalogues and Unpublished Inventories of Extant Collections. Fourth revised and enlarged Edition by S. Krämer, Münehen, Monumenta Germaniae Historica 1993 (≪Monumenta Germaniae Historica. Hilfsmittel, 13≫ ), 941 pp. Nuncius, 10(2):902, ???? 1995. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539185x01566.

Abbri:1995:BRRc

[1449] Ferdinando Abbri. Book review: Roy Porter, Lesley Hall, The Facts of Life. The Creation of Sexual Knowledge in Britain, 1650–1950, New Haven and London, Yale University Press, 1995, xii + 415 pp. Nuncius, 10(2):903–904, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01575.

Toscano:1995:BRG

[1450] Anna Toscano. Book review: Gudrun Wolfschmidt (Hrsg.), Nicolaus Copernicus. Revolutionär wider Willen, Stuttgart, Gnt Verlag, 1994, 350 pp. [Verlag für Geschichte der Naturwissenschaften und der Technik]. Nuncius, 10(2):904–906, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01584.

Miniati:1995:BRGd

[1451] Mara Miniati. Book review: Galileo Galilei, Lettera a Cristina di Lorena, Roma, Carlo Mancosu Editore, 1993, 80 pp. (≪Lo scrigno≫ 9). Nuncius, 10(2):906, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01593.

Miniati:1995:BRGe

[1452] Mara Miniati. Book review: Gianni Bonera, Galileo oggi. Spunti didattici e letture tratte dai lavori di Galileo, Pavia, Università degli Studi di Pavia, 1995, 207 pp., ill. (≪Collana di storia della scienza≫ ). Nuncius, 10(2):907, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01601.

Miniati:1995:BRA

[1453] Mara Miniati. Book review: Anna Cassini, Gio: Domenico Cassini. Uno scienziato del Seicento. Testi e documenti, Piñerolo, Comune di Perinaldo, 1994, 315 pp., ill. Nuncius, 10(2):907–908, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01610.

Abbri:1995:BRT

[1454] Ferdinando Abbri. Book review: Tore Frängsmyr (ed.), Linnaeus. The Man and His Work, Canton Mass., Science History Publications, 1994, xiv + 206 pp. (≪Uppsala Studies in History of Science≫, 18). Nuncius, 10(2):908–909, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01629.

Segala:1995:BRG

[1455] Marco Segala. Book review: G. W. F. Hegel, La filosofia della natura. Quaderno jenese 1805–1806, Milano, Guerini e Associati, 1994, xx + 153 pp. (≪Istituto Italiano per gli Studi Filosofici. Hegeliana≫, 7). Nuncius, 10(2):909–910, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01638.

Miniati:1995:BREc

[1456] Mara Miniati. Book review: Enrico De Pascale, Cristina Giannini (a cura di), Giovanni Secco Suardo (1798–1873). Fonti strumenti materiali di ricerca, Bergamo, Provincia di Bergamo, Associazione Giovanni Secco Suardo, 1995, 177 pp., ill. Nuncius, 10(2):910–911, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01647.

Abbri:1995:BRJe

[1457] Ferdinando Abbri. Book review: Jonathan Smith, Fact and Feeling, Baconian Science and the Nineteenth-Century Literary Imagination, Madison, The University of Wisconsin Press, 1994, x + 278 pp. (≪Science and Literature≫ ). Nuncius, 10(2):911–913, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01656.

Pogliano:1995:BRA

[1458] Claudio Pogliano. Book review: Aa.Vv., the Genetic Revolution. Scientific Prospects and Public Perceptions, edited by Bernard D. Davis, Baltimore, The Johns Hopkins University Press, 1991, 296 pp. Nuncius, 10(2):913, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539185x01665.

Parigi:1995:BRJ

[1459] Silvia Parigi. Book review: John Haugeland (a cura di), Progettare la mente. Filosofia, psicologia, intelligenza artificiale, Bologna, Il Mulino, 1989 (ed. orig. 1981), 396 pp. Nuncius, 10(2):913–915, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01674.

Pogliano:1995:BRJ

[1460] Claudio Pogliano. Book review: Jean-Noël Missa, L’esprit-cerveau. La philosophie de l’esprit à la lumière des neurosciences, Paris, J. Vrin, 1993, 264 pp. Nuncius, 10(2):916, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01683.

Miniati:1995:BRLc

[1461] Mara Miniati. Book review: Lucia Bigliazzi, Luciana Bigliazzi, Fiumi, inondazioni e ≪ idraulica pratica≫, Firenze, Accademia dei Georgofili, 1995, 21 pp. + xxi tavv., ind. Nuncius, 10(2):916–917, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539185x01692.

Casati:1995:BRA

[1462] Stefano Casati. Book review: Alexander Kohn, Falsi profeti. Inganni ed errori nella scienza, Bologna, Zanichelli, 1991, XIII, 318 pp. (Le Ellissi). Nuncius, 10(2):917–918, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01700.

Anonymous:1995:ADMb

[1463] Anonymous. Attività dell’Istituto e Museo di Storia Della Scienza di Firenze. Nuncius, 10(2):919–936, ???? 1995. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539185x01719.

Pepe:1996:CFD

[1464] Luigi Pepe. Il Collége de France durante la Rivoluzione Francese due memorie apologetiche. (Italian) [The College of France during the French Revolution: two apologetic memoirs]. Nuncius, 11(1):3–41, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00808.

Mazzolini:1996:R

[1465] Renato G. Mazzolini. Ricordo. Nuncius, 11(1):22–318, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00952.

Knight:1996:SCM

[1466] David Knight. Science and culture in Mid-Victorian Britain: the Reviews, and William Crookes’ Quarterly Journal of Science. Nuncius, 11 (1):43–54, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00817.

Morabito:1996:LCD

[1467] Carmela Morabito. Le localizzazioni cerebrali di David Ferrier, fra indagini sperimentali e applicazioni alla clinica. Nuncius, 11(1):55–91, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00826.

DelSanto:1996:OEE

[1468] Paolo Del Santo and Giorgio Strano. Observational evidence and the evolution of Ptolemy’s lunar model. Nuncius, 11(1):93–122, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00835.

Paternoster:1996:SLP

[1469] G. Paternoster, R. Rinzivillo, and E. Schettino. Studio di una lente per cannocchiale di grandi dimensioni lavorata da Evangelista Torricelli. (Italian) [study of telescope lens of large dimension produced by Evangelista Torricell]. Nuncius, 11(1):123–134, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00844.

Brenni:1996:NAS

[1470] Paolo Brenni. Nota su alcuni strumenti recentemente acquisiti dalla Fondazione Scienza e Tecnica. Nuncius, 11(1):135–149, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00853.

Carbone:1996:PAD

[1471] Luciano Carbone, Pietro Nastasi, and Franco Palladino. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 11(1):151–225, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00862.

Collini:1996:IFS

[1472] Silvia Collini and Antonella Vannoni. Istituzioni e fonti la Société d’Histoire Naturelle e il viaggio di d’Entrecasteaux alla ricerca di la Prouse: Le istruzioni scientifiche per i viaggiatori. Nuncius, 11(1):227– 250, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00871.

Richard:1996:ASD

[1473] Louis-Claude-Marie Richard. Adresse la socit D’histoire naturelle de Paris. Nuncius, 11(1):251–253, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00880.

Richard:1996:IPP

[1474] Louis-Claude-Marie Richard. Instruction partielle pour les voyageurs naturalistes. Nuncius, 11(1):254–260, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x00899.

Lezerme:1996:NLD

[1475] Lezerme. Notes l’usage du jardinier destiné faire le voyage projetté pour l’avancement des sciences et la découverte de M. De la Peyrouse. Nuncius, 11(1):261–264, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00907.

Anonymous:1996:NRP

[1476] Anonymous. Notes remises par M. De Fourcroy médecin de l’Academie Roy. Des Sciences, M. Cadet, relativement au voyage de M. De la Peyrouse autour du monde, du 20 mai 1785. (French) [Notes submitted by M. De Fourcroy, doctor of the Royal Academy of Sciences, M. Cadet, relative to the travel of M. De la Peyrouse around the world, of 20 may 1785]. Nuncius, 11(1):265–275, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00916.

Guarnieri:1996:PSD

[1477] Patrizia Guarnieri. Per una storia delle scienze dell’infanzia: Le fonti dell’Istituto degli Innocenti di Firenze. (Italian) [toward a history of childhood science: Sources of the Institute of the Innocent in Florence]. Nuncius, 11(1):277–307, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00925.

Anonymous:1996:CBDa

[1478] Anonymous. Commission on bibliography and documentation of the International Union of the History and Philosophy of Science. Nuncius, 11(1):309, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00934.

Home:1996:TIB

[1479] Roderick W. Home. Towards an International bibliographical database in the history of science, technology and medicine. Nuncius, 11(1):310–315, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00943.

Touwaide:1996:DC

[1480] Alain Touwaide. Discussioni critiche. Nuncius, 11(1):319–336, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00961.

Conforti:1996:SNS

[1481] Maria Conforti. Sidereus Nuncius Stella Polaris the scientific relations between Italy and Sweden in early modern history, the 1995 S. Lucia Seminar. Nuncius, 11(1):337–342, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00970.

diPasquale:1996:BRA

[1482] Giovanni di Pasquale. Book review: Alessandra Gara, Tecnica e tecnologia nelle societ antiche, Roma, La Nuova Italia Scientifica, 1994, 149 pp., ill. Nuncius, 11(1):343–347, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00989.

Toscano:1996:BRPa

[1483] Anna Toscano. Book review: Ps. Alessandro D’Afrodisia, Trattato sulla febbre, edizione critica, traduzione e commento a cura di Piero Tassinari, Alessandria, Edizioni dell’Orso, 1994, x + 140 pp. (Culture Antiche. Studi e Testi). Nuncius, 11(1):347–349, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00998.

Iannaccone:1996:BRJ

[1484] Isaia Iannaccone. Book review: J. Needham et alii, Science and Civilisation in China, Vol. 5 (Chemistry and Chemical Technology), Part VI (Military Technology: Missiles and Sieges), Cambridge, Cambridge University Press, 1994, xxvi + 601 pp. ill. Nuncius, 11(1):349–352, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01005.

Tosi:1996:BRM

[1485] Alessandro Tosi. Book review: Martin Kemp, La scienza dell’arte. Prospettiva e percezione visiva da Brunelleschia Seurat, Firenze, Giunti, 1994, 423 pp., ill. Nuncius, 11(1):352–354, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x01014.

Toscano:1996:BRJ

[1486] Anna Toscano. Book review: Les jésuites à la Renaissance. Système éducatif et production du savoir, sous la direction de Luce Giard, (French) [The Jesuits in the Renaissance. Educational system and knowledge production, under the direction of Luce Giard]. Paris, Presses Universitaires de France, 1995, lxxix + 336 pp. (Bibliothèque d’Histoire des Sciences). Nuncius, 11(1):354–357, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01023.

Beretta:1996:BRLa

[1487] Ilva Beretta. Book review: L’Ortobotanicodi Padova 1545–1995, a cura di Alessandro Minelli, Venezia, Marsilio Editore, 1995, 311 pp., ill. Nuncius, 11(1):357–359, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x01032.

Beretta:1996:BRG

[1488] Ilva Beretta. Book review: Il giardino delle Muse Arti e artifici nel barocco europeo. Atti del IV colloquio Internazionale, Pietrasanta 8–10 settembre 1993, a cura di Maria Adriana Giusti and Alessandro Tagliolini, Firenze, Edifir, 1995, xvi + 277 pp., ill. Nuncius, 11(1):359–361, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01041.

Micheletti:1996:BRA

[1489] Mario Micheletti. Book review: Alexander Jacob, Henry More’s Manual of Metaphysics, Part I: chapters 1–10 and 27–28, Part II: chapters 11– 26, Hildesheim-Zürich-New York, Georg Olms Verlag, 1995, 160; 277 pp. Nuncius, 11(1):361–365, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01050.

Abbri:1996:BRMa

[1490] Ferdinando Abbri. Book review: Michael Hunter, Science and the Shape of Orthodoxy. Intellectual Change in Late Seventeenth-Century Britain, Woodbridge, the Boydell Press, 1995, xii + 345 pp. Nuncius, 11(1): 366–367, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x01069.

Segala:1996:BRA

[1491] Marco Segala. Book review: Alberto Basso, l’invenzione della gioia. Musica e massoneria nell’et dei Lumi, Milano, Garzanti, 1994, 733 pp. (Collezione storica). Nuncius, 11(1):367–370, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01078.

Giudice:1996:BRC

[1492] Franco Giudice. Book review: Casper Hakfoort, Optics in the age of Euler, Cambridge, Cambridge University Press, 1995, vii + 243 pp. Nuncius, 11(1):370–375, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x01087.

Vasconi:1996:BRS

[1493] Paola Vasconi. Book review: Silvia Parigi, Il Mondo visibile. George Berkeley e la perspectiva, Firenze, Leo S. Olschki ed., 1995, 226 pp. Nuncius, 11(1):375–377, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x01096.

Ciardi:1996:BRL

[1494] Marco Ciardi. Book review: Lavoisier in European Context. Negociating a New Language for Chemistry, Ferdinando Abbri — Bernadette Bensaude-Vincent editors, Canton MA, Watson Publishing International, Science History Publications, 1995, vii + 303 pp., illustrated, clothbound. Nuncius, 11(1):377–379, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01104.

Beretta:1996:BRSa

[1495] Marco Beretta. Book review: A Scientific Correspondence During the Chemical Revolution. Louis-Bernard Guyton de Morveau, Richard Kirwan, 1782–1802. Edited by Emmanuel Grison, Michelle Goupil and Patrice Bret, Berkeley, Berkeley Papers in History of Science 17, 1994, vi + 257 pp., ill. Nuncius, 11(1):379–381, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01113.

Bernardi:1996:BRG

[1496] Walter Bernardi. Book review: Giuseppe Olivi, Zoologia Adriatica, a cura di C. Gibin, ristampa anastatica dell’edizione del 1792, Padova, T G Edizioni, 1995, 334 pp. Nuncius, 11(1):381–384, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01122.

Morini:1996:BRM

[1497] Simona Morini. Book review: M. Norton Wise (a cura di), the Values of Precision, Princeton, Princeton University Press, 1995, vm + 372 pp. Nuncius, 11(1):385–388, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01131.

Vasconi:1996:BRP

[1498] Paola Vasconi. Book review: Paolo Parrini (Ed.), Kant and Contemporary Epistemology, Dordrecht, Kluwer, 1994, xi + 376 pp. Nuncius, 11 (1):388–390, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x01140.

Ciardi:1996:BRGa

[1499] Marco Ciardi. Book review: Gerd Buchdahl, Modelli di spiegazione. Per una lettura neotrascendentale delle teorie scientifiche, (Italian) [Models of explanation. Reading neotranscendental scientific theories], a cura di Fabio Bevilacqua e Franco Giudice, Pavia, Universit degli Studi di Pavia — Collana di Storia della Scienza, 1995, 279 pp. Nuncius, 11(1):390– 392, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01159.

Cerruti:1996:BRM

[1500] Luigi Cerruti. Book review: Mary Jo Nye, From Chemical Philosophy to Theoretical Chemistry. Dynamics of Matter and Dynamics of Disciplines, 1800–1950, Berkeley, University of California Press, 1993, xvii + 328 pp. Nuncius, 11(1):392–395, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01168.

Rigutti:1996:BRF

[1501] Mario Rigutti. Book review: Freddy Litten, Astronomie in Bayern 1914– 1945, Stuttgart, Franz Steiner Verlag, 1992, xi + 329 pp. Nuncius, 11 (1):395–399, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x01177.

Abbri:1996:BRJa

[1502] Ferdinando Abbri. Book review: John Maynard Keynes, Trattato sulla probabilit. A cura di Alberto Pasquinelli e Silvia Marzetti Dall’Aste Brandolini, Bologna, Clueb, 1994, xxvii + 504 pp. (Collana di Economia Generale, 9). Nuncius, 11(1):400–403, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01186.

Rigutti:1996:BRC

[1503] Mario Rigutti. Book review: C. W. Kilmister, Eddington’s search for a fundamental theory, Cambridge, Cambridge University Press, 1994, xii + 256 pp. Nuncius, 11(1):403–410, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01195.

Morini:1996:BRA

[1504] Simona Morini. Book review: Andrew Pickering, the Mangle of Practice, Chicago-London, The University of Chicago Press, 1995, 281 pp. Nuncius, 11(1):410–412, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x01203.

Ghia:1996:BRR

[1505] Paolo Ghia. Book review: Robert N. Proctor, Cancer wars. How politics shapes what we know and don’t know about cancer, New York, Basicbooks, 1995, 356 pp. Nuncius, 11(1):413–416, ????

1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01212.

Pogliano:1996:BRA

[1506] Claudio Pogliano. Book review: Antonio R. Damasio, l’errore di Cartesio. Emozione, ragione e cervello umano, Milano, Adelphi, 1995, 408 pp. Nuncius, 11(1):416–419, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x01221.

Miniati:1996:SB

[1507] Mara Miniati. Strumenti bibliografici. Nuncius, 11(1):421–422, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01230.

Miniati:1996:M

[1508] Mara Miniati. Museologia. Nuncius, 11(1):422–423, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x01249.

Brenni:1996:S

[1509] Paolo Brenni. Strumentaria. Nuncius, 11(1):423–425, ???? 1996. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x01258.

Miniati:1996:BRMa

[1510] Mara Miniati. Book review: Montres et Merveilles. 200 rare creazioni d’epoca provenienti dal Museo dell’Orologeria di Ginevra e dalle Collezioni Private Piaget. (Italian) [Watches and Marvels. 200 rare vintage creations coming from the Watchmaking Museum of Geneva and from Piaget Private Collections]. Milano, Electa, 1992, 165 pp., ill. Nuncius, 11(1):425–426, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x01267.

Miniati:1996:BRJa

[1511] Mara Miniati. Book review: Jim Bennett, the Measurers. A Flemish Image of Mathematics in the Sixteenth Century, Oxford, the Museum of the History of Science, 1995, 56 pp., ill. Nuncius, 11(1):426–427, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01276.

Miniati:1996:BRI

[1512] Mara Miniati. Book review: Itc Pier Crescenzi, Provincia di Bologna, Una scuola nel tempo. Documentie collezioni dell’Istituto Tecnico Pier Crescenzi, Bologna, Grafis Edizioni, 1995, 120 pp., ill. Nuncius, 11(1): 427–428, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x01285.

Ciardi:1996:FDS

[1513] Marco Ciardi. Filosofia della scienza. Nuncius, 11(1):429–430, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01294.

Miniati:1996:MA

[1514] Mara Miniati. Matematica e astronomia. Nuncius, 11(1):431, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01302.

Ciardi:1996:BRE

[1515] Marco Ciardi. Book review: Elena Ausejo and Mariano Hormigón (Eds.), Messengers of Mathematics: European Mathematical Journals (1800–1946), Madrid, Siglo XXI de España Editores, S. A., 1993, 297 pp. [ISBN 84-323-0802-1]. Nuncius, 11(1):431–432, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01311.

Ciardi:1996:BRC

[1516] Marco Ciardi. Book review: La corrispondenza di Luigi Cremona, 2 voll., Roma, Università di Roma La Sapienza, Facoltà di Scienze Matematiche, Fisiche e Naturali, Serie di Quaderni della Rivista di Storia della Scienza: I volume, quad. n. 1 (1992) a cura di Ana Millén Gasca, 181 pp., II volume, quad. n. 3 (1994) a cura di Marta Menghini, 175 pp. Nuncius, 11(1):432–433, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x01320.

Tosi:1996:B

[1517] Alessandro Tosi. Botanica. Nuncius, 11(1):433–434, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x01339.

Beretta:1996:BRN

[1518] Ilva Beretta. Book review: Nicholas Barker, Hortus Eystettensis. the Bishop’s Garden and Besler’s Magnificent Book, London, the British Library, 1995, vii + 80 + 80 pp., ill. Nuncius, 11(1):434–435, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01348.

Tosi:1996:BRF

[1519] Alessandro Tosi. Book review: Fotografiae botanica tra Ottocento e Novecento, a cura di Guido Moggi, Monica Maffioli e Emanuela Sesti, Firenze, Alinari, 1994, 162 pp., ill. Nuncius, 11(1):435–437, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01357.

diPasquale:1996:M

[1520] Giovanni di Pasquale. Medicina. Nuncius, 11(1):437–438, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01366.

Malacrida:1996:BRF

[1521] Lorenzo Marri Malacrida. Book review: Filippo Baldinucci, Diario spirituale, a cura di Giuseppe Parigino, (introduzione di Luigi Stumpo), Firenze, Le Lettere, 1995, 226 pp. Nuncius, 11(1):438–439, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01375.

Lippi:1996:BRV

[1522] Donatella Lippi. Book review: Vittorio Donato Catapano, Matti agli Incurabili di Napoli, Napoli, Liguori Ed., 1995, 117 pp. Nuncius, 11 (1):439–440, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x01384.

Malacrida:1996:BRA

[1523] Lorenzo Marri Malacrida. Book review: Angelo Nannoni, Trattato sulle malattie delle mammelle (a cura di Giuseppe Parigino), Firenze, le Lettere, 1995, 132 pp. Nuncius, 11(1):440–441, ????

1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01393.

Miniati:1996:GM

[1524] Mara Miniati. Geologia e meteorologia. Nuncius, 11(1):441–442, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01401.

Miniati:1996:BRFa

[1525] Mara Miniati. Book review: Franca Mangianti, Maria Carmen Beltrano, Le precipitazioni a Roma, Roma, Ministero delle risorse agricole, alimentari e forestali-Ucea, 1995, 47 pp., ill. Nuncius, 11(1):442–443, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01410.

Abbri:1996:Aa

[1526] Ferdinando Abbri. Antichit. Nuncius, 11(1):443–444, ???? 1996. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x01429.

Pereira:1996:MR

[1527] Michela Pereira. Medioevo e rinascimento. Nuncius, 11(1):444–445, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01438.

Abbri:1996:BRG

[1528] Ferdinando Abbri. Book review: Gershom Scholem, Alchimia e kabbalah. Traduzione di Marina Sartorio, Torino, Einaudi, 1995, 96 pp. (Saggi, 800). Nuncius, 11(1):445–446, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01447.

Ciardi:1996:SXX

[1529] Marco Ciardi. Secoli XVII–XX. Nuncius, 11(1):446–447, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01456.

Tosi:1996:BRL

[1530] Alessandro Tosi. Book review: Lavinia Fontana (1552–1614), catalogo della mostra a cura di Vera Fortunati, Bologna, Museo Civico Archeologico, 1 ottobre–4 dicembre 1994, Milano, Electa, 1994, 224 pp., ill. Nuncius, 11(1):447–448, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x01465.

Abbri:1996:BRB

[1531] Ferdinando Abbri. Book review: Bettina Meitzner, Die Gertschaft der chymischen Kunst. Der Traktat De Sceuastica Artis des Andreas Libavius von 1606. bersetzung, Kommentierung und Wiederabdruck, Stuttgart, Franz Steiner Verlag, 1995, xviii + 308 + 179 pp. (Boethius, Band 34). Nuncius, 11(1):449, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x01474.

Miniati:1996:BRP

[1532] Mara Miniati. Book review: Paolo Bozzi, Carlo Maccagni, Luigi Olivieri, Thomas B. Settle, Galileo e la scienza sperimentale, a cura di Milla Baldo Ceolin, Padova, Dipartimento di Fisica Galileo Galilei, 1995, 132 pp., ill. Nuncius, 11(1):450–451, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01483.

Beretta:1996:BRSb

[1533] Marco Beretta. Book review: Steven Shapin, Simon Schaffer, Il Leviatano e la pompa ad aria. Hobbes, Boyle e la cultura dell’esperimento. Trad. it. di Roberto Brigati, Firenze, La Nuova Italia, 1994, xii + 531 pp., ill. Nuncius, 11(1):451–452, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01492.

Toscano:1996:BRM

[1534] Anna Toscano. Book review: Michel Blay, Les Principia de Newton, Paris, Presses Universitaires de France, 1995, 124 pp. (Philosophies). Nuncius, 11(1):452–453, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01500.

Abbri:1996:BRDa

[1535] Ferdinando Abbri. Book review: Donald Rutherford, Leibniz and the Rational Order of Nature, Cambridge-New York, Cambridge University Press, 1995, xiv + 301 pp. Nuncius, 11(1):453–455, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01519.

Abbri:1996:BRDb

[1536] Ferdinando Abbri. Book review: Dorinda Outram, the Enlightenment, Cambridge, Cambridge University Press, 1995, xvi + 141 pp. (New Approaches to European History, n. 7). Nuncius, 11(1):455–456, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01528.

Abbri:1996:BRJb

[1537] Ferdinando Abbri. Book review: Jan Häll, I Swedenborgs labyrint. Studier i de gustavianska swedenborgarnas liv och tänkande, (Swedish) [In Sweden’s Labyrinth: Studies in Gustavian Sweden’s bourgeois life and thought] Stockholm, Atlantis, 1995, 512 pp. Nuncius, 11(1):456– 458, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01537.

Collini:1996:BRE

[1538] Silvia Collini. Book review: Ester Diana, in viaggio con il Granduca. Itinerari nella Toscana dei Lorena, Firenze, Edizioni medicea, 1994, 275 pp. Nuncius, 11(1):458–459, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01546.

Abbri:1996:BRL

[1539] Ferdinando Abbri. Book review: Lodovica Braida, Il Commercio delle Idee. Editoria e circolazione del libro nella Torino del Settecento, Firenze, Olschki, 1995, 404 pp. (Fondazione Luigi Firpo. Studi e Testi, 2). Nuncius, 11(1):459–460, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x01555.

Tosi:1996:BRG

[1540] Alessandro Tosi. Book review: Giovanni Federico, Book Review: Il filo d’oro. L’industria mondiale della seta dalla restaurazione alla grande crisi, (Italian) [The golden thread. The global industry of silk from the Restoration to the Great Crisis]. Introduzione di Luciano Cafagna, Venezia, Marsilio, 1994, 573 pp. Nuncius, 11(1):460–461, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01564.

Scarpelli:1996:BRM

[1541] Giacomo Scarpelli. Book review: Mary P. Winsor, Reading the Shape of Nature. Comparative Zoology at the Agassiz Museum, Chicago London, The University of Chicago Press, 1991, xviii + 324 pp., 49 ill. (Science and its Conceptual Foundations). Nuncius, 11(1):461–462, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01573.

Robotti:1996:BRA

[1542] Nadia Robotti. Book review: Arthur I. Miller, Early quantum electrodynamics: a source book, Cambridge, Cambridge University Press, 1994, xix + 259 pp. Nuncius, 11(1):462–463, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01582.

Govoni:1996:BRY

[1543] Paola Govoni. Book review: Yves Jeanneret, Écrire la science. Formes et enjeux de la vulgarisation, Paris, Presses Universitaires de France, 1994, 398 pp. (Science, histoire et socit). Nuncius, 11(1):463–465, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01591.

Miniati:1996:BRJb

[1544] Mara Miniati. Book review: José Sala Catalá, Ciencia y Técnica en la metropolización de América, (Spanish) [Science and technology in the metropolization of America], Madrid, Ediciones Doce Calles, 1994, 346 pp., ill. (Theatrum Machinae. Colección de Historia de las Técnicas). Nuncius, 11(1):465–467, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01609.

Gorman:1996:AR

[1545] Michael John Gorman. Attivit di ricerca. Nuncius, 11(1):469–473, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x01618.

Scotti:1996:IGM

[1546] Andrea Scotti. Indice generale dei manoscritti scientifici della Biblioteca Centrale di Firenze. Progetto di ricerca. (Italian) [General index of the scientific manuscripts of the Central Library of Florence. Research project]. Nuncius, 11(1):473–492, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x01627.

Miniati:1996:DT

[1547] Mara Miniati. A distanza di trent’anni. Nuncius, 11(2):iii–506, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00015.

Shea:1996:RT

[1548] William R. Shea. The revelations of the telescope. Nuncius, 11(2):507– 526, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00024.

Baroncini:1996:NSS

[1549] Gabriele Baroncini. Note sull’illustrazione scientifica. Nuncius, 11(2): 527–543, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00033.

Giudice:1996:TDL

[1550] Franco Giudice. Teoria della luce e struttura della materia nello ≪short tract on first principles≫ di Thomas Hobbes. Nuncius, 11(2):545–561, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00042.

Pareti:1996:GBF

[1551] Germana Pareti. Giulio Bizzozero e la funzione ematopoietica del Midollo Osseo. Nuncius, 11(2):563–580, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00051.

Bucchi:1996:SMM

[1552] Massimiano Bucchi. La scienza e I mass media: la ≪Fusione Fredda≫ nei quotidiani italiani. Nuncius, 11(2):581–600, ????

1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00060.

Anonymous:1996:A

[1553] Anonymous. Appendice 1. Nuncius, 11(2):601–611, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x00079.

Lualdi:1996:FBO

[1554] Alberto Lualdi. François De Baillou, un Ottico della Milano Teresiana. Nuncius, 11(2):613–630, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00088.

Gatto:1996:CAD

[1555] Romano Gatto. Il cannocchiale Amici dell’Osservatorio Astronomico di Capodimonte e la corrispondenza Amici–Zuccari. Nuncius, 11(2):631– 653, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00097.

Anonymous:1996:CBDb

[1556] Anonymous. Commission on bibliography and documentation of the International Union of the History and Philosophy of Science. Nuncius, 11(2):655, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00105.

Martins:1996:SSS

[1557] Roberto De Andrade Martins. Sources for the study of science, medicine and technology in Portugal and Brazil. Nuncius, 11(2):656–667, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00114.

diPasquale:1996:ACS

[1558] Giovanni di Pasquale and Vittorio Marchis. Alcune considerazioni sul pes romanus. Nuncius, 11(2):669–675, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00123.

Nonnoi:1996:CGP

[1559] Giancarlo Nonnoi. Il caso Galileo. Una proposta di ricomposizione. Nuncius, 11(2):677–686, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x00132.

Ciardi:1996:DCA

[1560] Marco Ciardi. Dalla cronaca al continuismo. Pierre Duhem e la nascita della storia della scienza moderna. (Italian) [From chronicle to continuism. Pierre Duhem and the birth of a history of modern science]. Nuncius, 11(2):687–698, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x00141.

Lenzi:1996:CLS

[1561] Alessandra Lenzi and Fabio Guidi. La collezione Landmarks of Science all’Istituto e Museo di Storia della Scienza. Nuncius, 11(2):699–704, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00150.

Abbri:1996:BRMb

[1562] Ferdinando Abbri. Book review: Michael Shortland and Richard Yeo (eds.), Telling lives in science. Essays on scientific biography, Cambridge, Cambridge University Press, 1996, xiv + 295 pp. Nuncius, 11(2):705– 708, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00169.

Miniati:1996:BRW

[1563] Mara Miniati. Book review: W. D. Hackmann, A. J. Turner (eds.), Learning, language and invention: Essays presented to Francis Maddison, Aldershot & Paris, Variorum and the Société Internationale de l’Astrolabe, 1994, 333 pp., ill. Nuncius, 11(2):708–711, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00178.

Beretta:1996:BRLb

[1564] Marco Beretta. Book review: L’età delle scoperte geografiche nei suoi riflessi linguistici in Italia, (Italian) [The Age of Discovery and its linguistic reflexes in Italy], Firenze, Accademia della Crusca, 1994, 372 pp., ill. Nuncius, 11(2):711–713, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00187.

Toscano:1996:BRPb

[1565] Anna Toscano. Book review: Paola Zambelli, L’apprendista stregone. Astrologia, cabala e arte lulliana in Pico della Mirandola e seguaci, Venezia, Marsilio Editori, 1995, 230 pp. (≪Saggi Marsilio≫ ). Nuncius, 11(2): 713–715, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00196.

Ciancio:1996:BRM

[1566] Lucia Ciancio. Book review: Michael Jacobs, the Painted Voyage. Art, Travel and Exploration 1564–1875, London, the British Museum Press, 1995, 160 pp., 64 illustrazioni a colori, 42 in bianco e nero. Nuncius, 11 (2):715–719, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00204.

Beretta:1996:BRI

[1567] Marco Beretta. Book review: Isabelle Pantin, La poésie du ciel en France dans la seconde moitié du seizième siècle, Genève, Droz, 555 pp. Nuncius, 11(2):719–721, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00213.

Beretta:1996:BRM

[1568] Marco Beretta. Book review: Maurice Crosland, Studies in the Culture of Science in France and Britain since the Enlightenment, Aldershot, Variorum, 1995, xxii + 291 pp. Nuncius, 11(2):721–722, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00222.

Beretta:1996:BRLc

[1569] Marco Beretta. Book review: Lazzaro Spallanzani, Edizione Nazionale delle opere. Parte seconda: Lezioni, a cura di Pericle di Pietro, Modena, E. Mucchi, 1994, 2 voll., 353 + 323 pp., Ill. Nuncius, 11(2):723–724, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00231.

Miniati:1996:BRS

[1570] Mara Miniati. Book review: Stefano Poggi, Maurizio Bossi (eds.), Romanticism in science. Science in Europe, 1790–1840, Dordrecht/Boston/London, Kluwer Academic Publishers, 1994, 245 pp. (≪Boston Studies in the Philosophy of Science≫, vol. 152). Nuncius, 11(2):724–727, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00240.

Segala:1996:BRC

[1571] Marco Segala. Book review: Charles Rosen, the romantic generation. Schubert Mendelssohn Meyerbeer Schumann Liszt Bellini Chopin Berlioz, London, Harper Collins Publishers, 1996, xv + 723 pp. Nuncius, 11(2):727–730, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x00259.

Ciardi:1996:BRGb

[1572] Marco Ciardi. Book review: Giuseppe Giuliani, Il Nuovo Cimento. Novant’anni di fisica in Italia (1855–1944), Pavia, La Goliardica Pavese, vi + 159 pp. + dischetto. Nuncius, 11(2):730–732, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00268.

Barone:1996:BRJ

[1573] Vincenzo Barone. Book review: Jed Z. Buchwald, the creation of scientific effects: Heinrich Hertz and electric waves, Chicago and London, The University of Chicago Press, 1994, 482 pp., 85 ill. Nuncius, 11 (2):733–735, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00277.

Bertini:1996:BRV

[1574] Silvia Bertini. Book review: Valeria Paola Babini, La questione dei frenastenici. Alle origini della psicologia scientifica in Italia (1870– 1910), Milano, Franco Angeli, 1996, 179 pp. Nuncius, 11(2):735–737, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00286.

Giuntini:1996:BRM

[1575] Chiara Giuntini. Book review: Mitchell G. Ash, Gestalt psychology in German culture, 1890–1967. Holism and the quest for objectivity, Cambridge University Press (≪Cambridge Studies in the History of Psychology≫ ), 1995, xii + 513 pp. [ISBN 0-521-64627-8]. Nuncius, 11(2):737–740, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00295.

Barone:1996:BRD

[1576] Vincenzo Barone. Book review: Daniel J. Kevles, the physicists: the history of a scientific community in modern America, Cambridge (MA), Harvard University Press, 1995 (ed. orig. 1977), 489 pp. Nuncius, 11 (2):740–743, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00303.

Vista:1996:BRK

[1577] Michele Vista. Book review: Kostas Gavroglu, Fritz London: a scientific biography, Cambridge, Cambridge University Press, 1995, xxiv + 299 pp., ill. Nuncius, 11(2):743–746, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00312.

Anonymous:1996:BRM

[1578] Anonymous. Book review: Marie Cariou, Bergson et Bachelard, Paris, Presses Universitaires de France, 1995, 117 pp. (≪Questions≫, 52). Nuncius, 11(2):746–750, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x00321.

Polizzi:1996:BRB

[1579] Gaspare Polizzi. Book review: Bruno Bertotti — Umberto Curi (a cura di), Erwin Schrödinger scienziato e filosofo, Padova, Il poligrafo 1994, 198 pp. (≪Percorsi della scienza storia testi problemi≫, 3). Nuncius, 11 (2):750–755, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00330.

Bertini:1996:BRR

[1580] Silvia Bertini. Book review: Ronald D. Fairbairn, Il piacere e l’oggetto. Scritti 1953–1963, Roma Casa Editrice Astrolabio, (1992), 174 pp. Nuncius, 11(2):756–760, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x00349.

Bertini:1996:BRP

[1581] Silvia Bertini. Book review: Paul Robinson, Freud e i suoi critici. Sulloway — Masson — Grünbaum, Roma, Casa Editrice Astrolabio, 1995, 200 pp. Nuncius, 11(2):760–764, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00358.

Toscano:1996:M

[1582] Anna Toscano. Museologia. Nuncius, 11(2):765–767, ???? 1996. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x00367.

Miniati:1996:S

[1583] Mara Miniati. Strumentaria. Nuncius, 11(2):767–768, ???? 1996. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x00376.

Miniati:1996:BRA

[1584] Mara Miniati. Book review: Antoni M. Piaskowsky, Dawne lunety i lornetki w zbiorach Polskich, (Polish) [Old telescopes and binoculars in the Polish collection]. Warszawa, Nakladem Autora, 1996, 194 pp., ill. Nuncius, 11(2):768–769, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x00385.

Miniati:1996:BRC

[1585] Mara Miniati. Book review: Concetti e oggetti, Rimini, Raffaelli e Luisè Editori, 1996, 159 pp., ill. Nuncius, 11(2):769–770, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00394.

Anonymous:1996:B

[1586] Anonymous. Biografie. Nuncius, 11(2):770–771, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x00402.

Abbri:1996:BRR

[1587] Fedinando Abbri. Book review: A. Rupert Hall, Henry More and the Scientific Revolution, Cambridge, Cambridge University Press, 1996, xiv + 304 pp. (≪Cambridge Science Biographies Series≫ ). Nuncius, 11 (2):771–772, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00411.

Clardi:1996:BRR

[1588] Marco Clardi. Book review: A. Rupert Hall, Isaac Newton. Adventurer in Thought, Cambridge, Cambridge University Press, 1996, 468 pp. Arthur Donovan, Antoine Lavoisier. Science, Administration and Revolution, Cambridge, Cambridge University Press, 1996, 351 pp., ill. David Knight, Humphry Davy. Science and Power, Cambridge, Cambridge University Press, 1996, 218 pp. Nuncius, 11(2):772–773, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00420.

Abbri:1996:BRP

[1589] Ferdinando Abbri. Book review: P. Alessandra Maccioni Ruju and Marco Mostert, the Life and Times of Guglielmo Libri (1802–1869) scientist, patriot, scholar, journalist and thief. A Nineteenth-Century story, Hilversum, Verloren Publishers, 1995, 448 pp. ill. Nuncius, 11(2):773–775, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00439.

Miniati:1996:A

[1590] Mara Miniati. Astronomia. Nuncius, 11(2):775–776, ???? 1996. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x00448.

Miniati:1996:BRMb

[1591] Maria Miniati. Book review: Ministero Per I Beni Culturali e Ambientali, Omaggio a Quirico Filopanti, Bologna, Biblioteca Universitaria, 1995, 41 pp., ill. Laura Peperoni, Marina Zuccoli, Vultus Uraniae. Raffigurazioni di Urania nella Biblioteca del Dipartimento di Astronomia, Bologna, Biblioteca Universitaria, 1996, 32 pp., ill. Nuncius, 11(2):776– 777, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00457.

Toscano:1996:BRC

[1592] Anna Toscano. Book review: Cecilia Payne-Gaposchkin, An Autobiography and other Recollections, edited by Katherine Haramundanis, Cambridge, Cambridge University Press, 1996, xxii + 277 pp. Nuncius, 11(2):777–778, ???? 1996. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539196x00466.

Lippi:1996:M

[1593] Donatella Lippi. Medicina. Nuncius, 11(2):778–779, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x00475.

Lippi:1996:BRM

[1594] Donatella Lippi. Book review: M. Vegetti, La Medicina in Platone, Vicenza, Il Cardo/Saggi, 1995, 122 pp. Nuncius, 11(2):779–781, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00484.

Lippi:1996:BRA

[1595] Donatella Lippi. Book review: Anticamedicinatibetana, Modena, Zanfi, 1995, 2 voll., 180 + 172 pp. Nuncius, 11(2):781, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00493.

Lippi:1996:BRJ

[1596] Donatella Lippi. Book review: J. Henderson, Piety and charity in late medieval Florence, Oxford, Clarendon Press, 1994, 533 pp. Nuncius, 11 (2):782–783, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00501.

Lippi:1996:BRL

[1597] Donatella Lippi. Book review: L’Ospedale e la città. Cinquecento anni d’arte a Verona, Verona, Cierre, 1996, 306 pp. Nuncius, 11(2):783–784, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00510.

Lippi:1996:BRE

[1598] Donatella Lippi. Book review: Ester Diana, Sanità nel quotidiano. Storie minute di medici, cerusici e pazienti, Firenze, L. Pugliese, 1995, 256 pp. Nuncius, 11(2):784–785, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00529.

Miniati:1996:BRFb

[1599] Mara Miniati. Book review: Fulvia Zina Vignotto, Marco Galloni, Beni culturali in ambiente medico chirurgico. Censimento presso gli ospedali piemontesi, ≪Giornale della Accademia di Medicina di Torino≫, Anno Clviii, 1995, 257 pp., ill. Nuncius, 11(2):785–786, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00538.

Monti:1996:BRA

[1600] Maria Teresa Monti. Book review: Academia Ophthalmologica Internationalis, History of Ophthalmology, n. 5, Dordrecht–Boston–London, Kluwer Academic Publishers, 1993. Nuncius, 11(2):786–787, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00547.

Pereira:1996:AC

[1601] Michela Pereira. Alchimia e chimica. Nuncius, 11(2):787–790, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00556.

Abbri:1996:BRMc

[1602] Ferdinando Abbri. Book review: Markus Seils, Friedrich Albrecht Carl Gren in seiner Zeit: 1760–1798. Spekulant oder Selbstdenker?, Stuttgart, Wissenschaftliche Verlangsgesellschaft, 1995, 261 pp. (≪Heidelberger Schriften zur Pharmacie und Naturwissenschaftsgeschichte≫, Band 14). Nuncius, 11(2):790–791, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00565.

Ciardi:1996:BRP

[1603] Marco Ciardi. Book review: Philosophers in the Laboratory, edited by Valeria Mosini, Accademia Nazionale di Scienze Lettere e Arti, Musis, 1996, 244 pp. Nuncius, 11(2):791–792, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00574.

Abbri:1996:Ab

[1604] Ferdinando Abbri. Antichità. Nuncius, 11(2):792–794, ???? 1996. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x00583.

Abbri:1996:BRMd

[1605] Ferdinando Abbri. Book review: Michel Foucault, Discorso e Verità nella Grecia antica, Roma, Donzelli Editore, 1996, xxii + 120 pp. Nuncius, 11(2):794–795, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00592.

Abbri:1996:BRMe

[1606] Ferdinando Abbri. Book review: Marcello Zanatta (a cura di), Organon di Aristotele. Volume Primo: Categorie — Dell’Interpretazione

— Analitici primi, Torino, Utet, 1996, 504 pp. Volume Secondo: Analitici Secondi — Topici — Confutazioni Sofistiche, Torino, Utet, 1996, 653 pp. (≪Classici della Filosofia≫ ). Nuncius, 11(2):795–796, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00600.

Abbri:1996:SXX

[1607] Ferdinando Abbri. Secoli XVI–XX. Nuncius, 11(2):797–799, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00619.

Beretta:1996:BRT

[1608] Marco Beretta. Book review: Die Technikgeschichte als Vorbildmoderner Technik. Band 19. Agricola Vorträge Chemnitz 1994, (German) [The history of technology as a model for contemporary art. Volume 19 Agricola Lectures at Chemnitz 1994]. Bochum, Georg-Agricola-Gesellschaft, 1995, 156 pp., ill. Nuncius, 11(2):799, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00628.

Miniati:1996:BRD

[1609] Mara Miniati. Book review: Descartes et la science, Poitiers, Éditions de l’Actualité Poitou-Charentes, 1996, 95 pp. ill. Nuncius, 11(2):799–800, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00637.

Nocenti:1996:BRT

[1610] Luca Nocenti. Book review: Thomas Leinkauf, Mundus combinatus: Studien zur Struktur der barocken Universalwissenschaft am Beispiel Athanasius Kirchers Sj (1602–1680), Berlin, Akademie Verlag, 1993, 434 pp. Nuncius, 11(2):800–801, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00646.

Beretta:1996:BRE

[1611] Marco Beretta. Book review: Emanuel Swedenborg, Ludus Heliconius and other Latin poems. Edited with introduction, translation and commentary by Hans Helander, Uppsala, Almqvist & Wiksell International, 1995, 255 pp. Nuncius, 11(2):801–802, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00655.

Ciardi:1996:BRB

[1612] Marco Ciardi. Book review: La Bretagne des savants et des ingénieurs (1825–1900), sous la direction de Jean Dhombres, Rennes, Éditions Ouest-France, Centre de culture scientifique, technique et industrielle, 1994, 384 pp., ill. Nuncius, 11(2):802–803, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00664.

Bertini:1996:BRL

[1613] Silvia Bertini. Book review: Lancelot Law Whyte, l’inconscio prima di Freud, Roma, Casa Editrice Astrolabio, (1995), 166 pp. Nuncius, 11 (2):803–804, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00673.

Miniati:1996:BRR

[1614] Mara Miniati. Book review: Roberto Paolo Ciardi, Lucia Tongiorgi Tomasi, Alessandro Tosi, Pisa romantica, Pisa, Cassa di Risparmio, 1995, 38 pp., [3] + 117 tavole, ill. Nuncius, 11(2):804–806, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00682.

Toscano:1996:BRF

[1615] Roberto Toscano. Book review: Ferdinando Abbri, Elio Matassi, Marco Segala, Filosofia e Musica nell’Età contemporanea. Studi e ricerche, a cura di F. Abbri, Arezzo, Dipartimento di Studi Storico-Sociali e Filosofici, Università degli Studi di Siena, 1995, 65 pp. (≪Lavori in corso≫ ). Nuncius, 11(2):806–807, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00691.

Abbri:1996:BRA

[1616] Ferdinando Abbri. Book review: Antonio Serravezza, Musica e scienza nell’età del positivismo, Bologna, il Mulino, 1996, 330 pp. Nuncius, 11 (2):807–808, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539196x00709.

Abbri:1996:BRH

[1617] Ferdinando Abbri. Book review: Hugh Richard Slotten, Patronage, Practice, and the Culture of American Science. Alexander Dallas Bache and the U.S. Coast Survey, Cambridge, Cambridge University Press, 1994, x + 228 pp. Nuncius, 11(2):809–810, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00718.

Peruzzi:1996:BRB

[1618] Giulio Peruzzi. Book review: Bruce R. Wheaton, Inventory of Sources for History of Twentieth–Century Physics: Report and microfiche index to 700,000 letters, Stuttgart, Gnt-Verlag, 1993, x + 294 pp. e 61 microfiche. Nuncius, 11(2):810–811, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00727.

Miniati:1996:BRU

[1619] Mara Miniati. Book review: Università di Roma ≪La Sapienza≫ — Cirms Centro Interdipartimentale di Ricerca in Metodologia Delle Scienze, Quaderno n. 1. Seminari tenuti negli A.A. 93/94 e 94/95 sul tema: Conoscenza scientifica e prevedibilità, Roma, Gennaio 1996, pag. varia. Nuncius, 11(2):811–812, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539196x00736.

Ciardi:1996:BRI

[1620] Marco Ciardi. Book review: Isegni del tempo. Memoria delle origini edicone del primordiale, a cura di Claudio Tugnoli, Trento, Liceo Ginnasio ≪G. Prati≫, 1993, 539 pp., ill. Nuncius, 11(2):812–813, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00745.

diPasquale:1996:AR

[1621] Giovanni di Pasquale. Attività di ricerca. Nuncius, 11(2):815–817, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00754.

Bonoli:1996:DL

[1622] Fabrizio Bònoli, Marinella Calisi, Salvatore Pilloni, and Piero Ranfagni. Dai lettori. Nuncius, 11(2):819–821, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539196x00763.

Anonymous:1996:IAX

[1623] Anonymous. Indici (anno XI, 1996). Nuncius, 11(2):823–824, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00772.

Anonymous:1996:IDF

[1624] Anonymous. Indice delle fonti manoscritte. Nuncius, 11(2):825, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00781.

Anonymous:1996:IN

[1625] Anonymous. Indice dei nomi. Nuncius, 11(2):827–842, ???? 1996. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539196x00790.

Chabas:1997:CAC

[1626] José Chabás. Le cahier d’astronomie d’un croisier du XVe siècle. Nuncius, 12(1):3–16, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00014.

Beretta:1997:HCS

[1627] Marco Beretta. Humanism and chemistry: the spread of Georgius Agricola’s metallurgical writings. Nuncius, 12(1):17–47, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x00023.

Hannaway:1997:RPC

[1628] Owen Hannaway. Reading the pictures: the context of Georgius Agricola’s woodcuts. Nuncius, 12(1):49–66, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x00032.

Borrelli:1997:DVA

[1629] Antonio Borrelli. Dall’innesto del vaiolo alla vaccinazione jenneriana: Il dibattito scientifico napoletano. Nuncius, 12(1):67–85, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00041.

Turner:1997:ABG

[1630] Gerard L’E. Turner. An astrolabe belonging to Galileo? Nuncius, 12 (1):87–92, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00050.

Todesco:1997:MFG

[1631] Piero Todesco. Il micrometro filare di Giacomo Lusverg (anno 1677). Nuncius, 12(1):93–103, ???? 1997. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539197x00069.

Grilli:1997:SR

[1632] Antonio Grilli. Scheda di restauro. Nuncius, 12(1):104–107, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00078.

Segre:1997:PAD

[1633] Michael Segre. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 12(1):109–122, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00087.

Rambaldi:1997:SMF

[1634] Gabriella Rambaldi. ≪Stromenti e Macchine≫ di fisica in un documento del 1785. Nuncius, 12(1):123–130, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00096.

Anonymous:1997:CBD

[1635] Anonymous. Commission on bibliography and documentation of the International Union of the History and Philosophy of Science. Nuncius, 12(1):131, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00104.

Watson:1997:SHS

[1636] Katherine D. Watson. Sources for the history of science in Oxford Survey Report. Nuncius, 12(1):132–136, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00113.

Casati:1997:CSS

[1637] Stefano Casati. Convegno di studi: Scienza, tecnica e ≪pubblico bene≫ nell’opera di Giovanni Arduino (1714–1795). Nuncius, 12(1):137–143, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00122.

Segala:1997:AF

[1638] Marco Segala. Ampère filosofo. Nuncius, 12(1):145–160, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00131.

Polizzi:1997:BRG

[1639] Gaspare Polizzi. Book review: Gilles-Gaston Granger, La scienza e le scienze, trad. it. di Tommaso Bianchi, Bologna, Il Mulino, 1996, 122 pp. (≪Universale Paperbacks il Mulino≫ 306). Nuncius, 12(1):161–165, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00140.

Lannaccone:1997:BRC

[1640] Isaia Lannaccone. Book review: C. Cullen, Astronomy and Mathematics in Ancient China: the ≪Zhou bi suan jing≫, Cambridge University Press, Cambridge, 1996, xiv + 241 pp., ill. Nuncius, 12(1):165–168, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00159.

Iannaccone:1997:BRJ

[1641] Isaia Iannaccone. Book review: J. Needham et alii, Science and Civilisation in China, vol. VI, (Biology and Biological Technology), part III (Agro-Industries and Forestry; Agro-Industries: Sugarcane Technology), Cambridge University Press, Cambridge, 1996, xx + 740 pp.; 135 illustrazioni, 6 mappe, 22 tavole. Nuncius, 12(1):168–171, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00168.

Toscano:1997:BRL

[1642] Anna Toscano. Book review: Luigina Morini (a cura di), Bestiari medievali, Torino, Einaudi, xxx + 644 pp., ill. (≪I Millenni≫ ) (ISBN 88-06-12446-3). Lit. 120.000. Nuncius, 12(1):171–173, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00177.

Beretta:1997:BRJ

[1643] Marco Beretta. Book review: N. Jardine — J. A. Secord — E. C. Spary (eds.) Cultures of Natural History, Cambridge, Cambridge University Press, 1996, xx + 501 pp., ill. Nuncius, 12(1):173–175, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00186.

Tosi:1997:BRA

[1644] Alessandro Tosi. Book review: Andrea Carlino, La fabbrica del corpo. Libri e dissezioni nel Rinascimento, prefazione di Adriano Prosperi, Torino, Einaudi, 1994, 267 pp., ill. (ISBN 88-06-13519-8). Nuncius, 12 (1):175–178, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00195.

Nenci:1997:BRC

[1645] Elio Nenci. Book review: Cesare Cesariano, Volgarizzamento dei libri IX (capitoli 7 e 8) e X di Vitruvio, De Architectura, secondo il manoscritto 9/2790 Sección De Cortes della Real Academia de la Historia, Madrid, a cura di B. Agosti, Pisa, Scuola Normale Superiore, 1996, 267 pp. (Centro di Ricerche Informatiche per i beni Culturali. Accademia della Crusca. Strumenti e testi 1). Nuncius, 12(1):178–180, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00203.

Plastina:1997:BRH

[1646] Sandra Plastina. Book review: Hélène Védrine, Philosophie et magie à la Renaissance, Paris, Le livre de poche, 1996, 119 pp. Nuncius, 12 (1):180–184, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00212.

Plastina:1997:BRF

[1647] Sandra Plastina. Book review: F. Patrizi, De spacio physico et mathematico, presentation, traduction et notes par Hélène Védrine, Paris, Vrin, 1996, 93 pp., 150 F. Nuncius, 12(1):184–187, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00221.

Beretta:1997:BRCa

[1648] Marco Beretta. Book review: Corrado Dollo (a cura di), Filosofia e Scienze nella Sicilia dei Secoli XVI e XVII. Vol. I — Le idee. Giarre, Regione Siciliana, Università di Catania (Dip. di Scienze Storiche) — Centro Studi per la storia della filosofia in Sicilia, 1996, 252 pp., Ill. Corrado Dollo (a cura di), Filosofia e Scienze nella Sicilia dei Secoli XVI e XVII. Vol. 2 — Testi. Marco Antonio Alaymo — Consigli politico-medici. Giarre, Regione Siciliana — Università di Catania (Dip. di Scienze Storiche)

— Centro Studi per la storia della filosofia in Sicilia, 1996, 258 pp. Nuncius, 12(1):187–189, ???? 1997. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539197x00230.

Abbri:1997:BRM

[1649] Ferdinando Abbri. Book review: Markku Peltonen, (ed.), The Cambridge Companion to Bacon, Cambridge, Cambridge University Press, 1996, xvm+372 pp. Nuncius, 12(1):189–191, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00249.

Segala:1997:BRD

[1650] Marco Segala. Book review: Dava Sobel, Longitudine, Milano, Rizzoli, 1996, 153 pp. Lit. 22.000. Nuncius, 12(1):191–193, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00258.

Vasconi:1997:BRP

[1651] Paola Vasconi. Book review: Pietro Rossi — Carlo A. Viano (a cura di), Storia della filosofia, 4. Il Settecento, Roma/Bari, Laterza, 1996, 776 pp. Nuncius, 12(1):193–195, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00267.

Malacrida:1997:BRS

[1652] Lorenzo Marri Malacrida. Book review: Susan C. Lawrence, Charitable Knowledge. Hospital pupils and practitioners in Eighteenth-Century London, U.S.A., Cambridge University Press, 1996, 390 pp. Nuncius, 12 (1):195–196, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00276.

Pogliano:1997:BRF

[1653] Claudio Pogliano. Book review: Francesca Vannozzi (a cura di), La scienza illuminata. Paolo Mascagni nel suo tempo (1755–1815), Siena, Nuova immagine editrice, 1996, 254 pp. Nuncius, 12(1):197–199, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00285.

Segala:1997:BRS

[1654] Marco Segala. Book review: Stephen G. Brush, A history of modern planetary physics. Volume One: Nebulous Earth. The origin of the Solar System and the core of the Earth from Laplace to Jeffreys, x + 313 pp., figg.; (ISBN 0-521-44171-4). Volume Two: Transmuted past. The age of the Earth and the evolution of the elements fro Lyell to Patterson, x + 134 pp., figg.; (ISBN 0-521-55213-3). Volume Three: Fruitful en-counters. The origin of the Solar System and of the Moon from Chamberlin to Apollo, x + 354 pp., figg., (ISBN 0-521-55214-1) Cambridge, Cambridge University Press, 1996. Nuncius, 12(1):199–202, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00294.

Morini:1997:BRT

[1655] Simona Morini. Book review: Theodore M. Porter, Trust in Numbers. The Pursuit of Objectivity in Science and Public Life, Princeton, Princeton University Press, 1995. Nuncius, 12(1):202–204, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00302.

Ghia:1997:BRP

[1656] Paolo Ghia. Book review: Pauline M. H. Mazumdar, Species and Specificity. An interpretation of the history of immunology, Cambridge, Cambridge University Press, 1995, xv + 457 pp., figg. (ISBN 0-521-43172-7). Nuncius, 12(1):204–207, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00311.

Abbri:1997:BRE

[1657] Ferdinando Abbri. Book review: Elisabeth Crawford, Arrhenius. From Ionic Theory to the Greenhouse Effect, Canton, Ma, Science History Publications, 1996, xv + 320 pp. ≪Uppsala Studies in History of Science≫, Vol. 23. Nuncius, 12(1):207–209, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00320.

Polizzi:1997:BRU

[1658] Gaspare Polizzi. Book review: Umberto Bottazzini Alberto Conte, Paola Gario (a cura di), Riposte armonie. Lettere di Federigo Enriques a Guido Castelnuovo, Torino. Bollati Boringhieri, 1996, 722 pp. + 8 tavole (≪La cultura scientifica≫ ). Nuncius, 12(1):209–215, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00339.

Barone:1997:BRJ

[1659] Vincenzo Barone. Book review: James T. Cushing, Quantum mechanics: historical contingency and the Copenhagen hegemony, Chicago and London, The University of Chicago Press, 1994, 317 pp. Nuncius, 12 (1):215–220, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00348.

Abbri:1997:BRS

[1660] Ferdinando Abbri. Book review: Strumenti Bibliografici S. A. Jayawardene, the Scientific Revolution. An Annotated Bibliography, West Cornwall CT, Locus Hill Press, 1996, xvm + 383 pp. Nuncius, 12(1):221–222, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00357.

Beretta:1997:BRCb

[1661] Marco Beretta. Book review: Cataloghi Bernd Ernsting, (a cura di), Georgius Agricola Bergwelten 1494–1994, Essen, Edition Glückauf, 1994, 350 pp., ill. Nuncius, 12(1):222–223, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x00366.

Miniati:1997:BRJ

[1662] Mara Miniati. Book review: Jim Bennett, Stephen Johnston, the Geometry of War 1500–1750. Catalogue of Exhibition. Oxford, Museum of the History of Science, 1996, 85 pp., ill. Nuncius, 12(1):224, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00375.

Miniati:1997:BRGa

[1663] Mara Miniati. Book review: Giorgio Mangani, Feliciano Paoli (a cura di), Gerardo Mercatore. Sulle tracce di geografi e viaggiatori nelle Marche. Introduzione di Marcel Watelet, Urbania, Edizioni Biblioteca e Civico Museo di Urbania Il lavoro editoriale, 1996, 191 pp., ill. Nuncius, 12 (1):224–225, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00384.

Iannaccone:1997:BRS

[1664] Isaia Iannaccone. Book review: Strumentaria A. Chapman, Astronomical Instruments and Their Users, Aldershot Variorum 1996, xii + 320 pp. Nuncius, 12(1):225–228, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00393.

Miniati:1997:BRF

[1665] Mara Miniati. Book review: Flavio Vetrano, (a cura di), Il Gabinetto di Fisica dell’Università di Urbino: la sua Storia, Il suo Museo/ The old ≪Gabinetto di Fisica≫ of the Urbino University: its History, its Museum, Roma, Istituto Poligrafico e Zecca dello Stato, 1996, 263 pp., ill. (ISBN 88-240-3867-0), L. 85.000. Nuncius, 12(1):228–229, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00401.

Miniati:1997:BRMa

[1666] Mara Miniati. Book review: Maria Carmen Beltrano, La collezione di strumenti dell’Ufficio Centrale di Ecologia Agraria, Roma, Euroma Editrice Universitaria di Roma — La Goliardica, 1996, 61 pp., ill. Nuncius, 12(1):229, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00410.

Clardi:1997:BRM

[1667] Marco Clardi. Book review: Matematica e Astronomia Alain Michel, Constitution de la théorie moderne de l’intégration, Paris, Vrin, 1992, 338 pp. (ISBN 2-7116-1064-0), 212 F. Nuncius, 12(1):230, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00429.

Ciardi:1997:BRB

[1668] Marco Ciardi. Book review: Bernard Bolzano, Les Paradoxes de l’Infini, introduction et traduction de Hourya Sinaceur, Paris, Éditions du Seuil, 1993, 192 pp. (ISBN 2-02-015919-8), 140 F. Nuncius, 12(1):230–231, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00438.

DelSanto:1997:BRA

[1669] Paolo Del Santo. Book review: Anne Tihon, Études d’astronomie byzantine, Aldershot Variorum (Collected Studies Series; Cs 454), 1994, x + 324 pp., s.i.p. Nuncius, 12(1):231–232, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x00447.

Ciardi:1997:BRR

[1670] Marco Ciardi. Book review: René Taton, Curtis Wilson, Planetary astronomy from the Renaissance to the rise of astrophysics. Vol. 2, Part B: The Eighteenth Century and Nineteenth Century, Cambridge, Cambridge University Press, 1995, x + 281 pp. (ISBN 0-521-35168-5). Nuncius, 12(1):232, ???? 1997. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539197x00456.

Miniati:1997:BRSa

[1671] Mara Miniati. Book review: A. De Simone, Giuseppe Monaco, La corrispondenza degli astronomi /The correspondence of the Astronomers. Vol. I. (I. Porro, A. Secchi, L. Respighi, A. Dorna), s.l., Arti Grafiche Savini, 1996, 249 pp. Nuncius, 12(1):232–233, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00465.

Lippi:1997:BRMa

[1672] Donatella Lippi. Book review: Medicina Giorgio Cosmacini, La qualit del tuo medico. Per una filosofia della medicina. Bari, Laterza, 95 pp. (Collana Il Nocciolo). Nuncius, 12(1):233, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00474.

Lippi:1997:BRC

[1673] Donatella Lippi. Book review: S. Cagliano, Viaggio intorno alla medicina, Bari, Laterza, 1993, 268 pp. Nuncius, 12(1):233–234, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00483.

Lippi:1997:BRBa

[1674] Donatella Lippi. Book review: Baroukh M. Assael, Il favoloso innesto. Storia sociale della vaccinazione, Bari, Laterza, 1995, 302 pp. Nuncius, 12(1):234–235, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00492.

Lippi:1997:BRGa

[1675] Donatella Lippi. Book review: G. Cosmacini — C. Cenedella, I vecchi e la cura. Storia del Pio Albergo Trivulzio, Bari, Laterza, 1994, 214 pp. (Collana di Storia della Medicina e della Sanità). Nuncius, 12(1): 235, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00500.

Lippi:1997:BRV

[1676] Donatella Lippi. Book review: Vittorio Sironi, Da Speziali a Imprenditori. Storia dei Recordati, Bari, Laterza, 1995, 233 pp. (Collana di Storia della medicina e della Sanità). Nuncius, 12(1):235–236, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00519.

Lippi:1997:BRP

[1677] Donatella Lippi. Book review: P. Placucci, Dal male oscuro alla malattia curabile. Storia dell’Istituto Nazionale per lo studio e la cura dei tumori di Milano, Bari, Laterza, 1995, 192 pp. (Collana di Storia della medicina e della Sanità). Nuncius, 12(1):236, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00528.

Lippi:1997:BRBb

[1678] Donatella Lippi. Book review: O. Barbieri, Guerra e resistenza nei servizi sanitari a Firenze, Milano, Vangelista, 1993, 362 pp. Nuncius, 12(1):237, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00537.

Beretta:1997:BRF

[1679] Marco Beretta. Book review: Fisica e Chimica Antonio di Meo, (a cura di), Storia della chimica in Italia, Beniamino Vignola Editore, 1996, li + 639 pp. Nuncius, 12(1):237–238, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00546.

Abbri:1997:BRAa

[1680] Ferdinando Abbri. Book review: Allen G. Debus, Paracelso e la tradizione paracelsiana, Napoli, La Città del Sole, 1996, 126 pp. Nuncius, 12(1):239, ???? 1997. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539197x00555.

Beretta:1997:BREa

[1681] Marco Beretta. Book review: Euler, Lozeran De Fiesc, Comte De Créqui, Emile Du Châtelet, Voltaire, De la propagation du feu. Cinq Mémoires couronnés par l’Académie Royale des Sciences. Paris 1738, Prez-sur-Marne, Edition Aspm, 1994, 232 pp., ill. Nuncius, 12(1):239– 240, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00564.

Beretta:1997:BRL

[1682] Marco Beretta. Book review: A.-L. Lavoisier, P.-S. De La Place, Memoria sul calore, a cura di Antonio di Meo, Roma. Edizioni Teknos, 1995. l + 59 pp. ill. Nuncius, 12(1):240–241, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00573.

Abbri:1997:BRAb

[1683] Ferdinando Abbri. Book review: Antichità Geoffrey Ernest Richard Lloyd, Aristotelian explorations, Cambridge, Cambridge University Press, 1996, x + 242 pp. Adversaries and Authorities. Investigations into ancient Greek and Chinese science, Cambridge, Cambridge University Press, 1996, xvm + 250 pp. (≪Ideas in context≫ ). Nuncius, 12 (1):242–243, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00582.

Caroti:1997:BRM

[1684] Stefano Caroti. Book review: Medioevo e Rinascimento Sten Ebbesen, herausgegeben von, Geschichte der Sprachtheorie. 3. Sprachtheorien in Spätankite und Mittelalter, Tübingen, Gunter Narr Verlag 1995, 408 pp. Nuncius, 12(1):244, ???? 1997. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539197x00591.

Caroti:1997:BRJ

[1685] Stefano Caroti. Book review: J. McEvoy (ed.), Robert Grosseteste: New Perspectives on His Thought and Scholarship, Steenbrugis, in Abbatia S. Petri-Brepols, Turnhout 1995, 438 pp. (≪Instrumenta Patristica, XXVII≫ ). Nuncius, 12(1):245, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00609.

Caroti:1997:BRW

[1686] Stefano Caroti. Book review: E. Wilks Dolnikowski, Thomas Bradwardine. A View of Time and a Vision of Eternity in Fourteenth Century Thought, Leiden-New York–Köln, Brill, 1995, 250 pp. (≪Studies in the History of Christian Thought, LXV≫ ). Nuncius, 12(1):246, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00618.

Caroti:1997:BRCa

[1687] Stefano Caroti. Book review: C. H. Lohr, Aristotelica Helvetica. Catalogus codicum latinorum in bibliothecis Confederationis Helveticae asservatorum, quibus versiones expositionesque operum Aristotelis continentur, Universitätsverlag Friburg Schweiz, Friburg, 1994, 387 pp. (≪Scrinium Friburgense. Veröffentlichungen des Mediävistischen Instituts der Universität Freiburg Schweiz. Sonderband, 6≫ ). Nuncius, 12(1):246–247, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00627.

Caroti:1997:BRCb

[1688] Stefano Caroti. Book review: C. H. Lohr, Latin Aristotle Commentaries. III Index initiorum-Index finium, Leo S. Olschki, Firenze, 1995, 289 pp. (≪Unione Accademica Nazionale, Corpus Philosophorum Medii Aevi. Subsidia X≫ ). Nuncius, 12(1):247–248, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00636.

Beretta:1997:BRS

[1689] Marco Beretta. Book review: Secoli XVII–XX Giuseppe Gabrieli, Il carteggio linceo, Roma, Accademia Nazionale dei Lincei, 1996, vi + 1446 pp., ill. Nuncius, 12(1):248–249, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00645.

Miniati:1997:BRB

[1690] Mara Miniati. Book review: Basilio Catania, Antonio Meucci, l’inventore e il suo tempo, Roma, Edizioni Seat, 1994, 2 voll. + 1 Addendum, 1292 pp., ill. Nuncius, 12(1):249–250, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00654.

Abbri:1997:BRD

[1691] Ferdinando Abbri. Book review: Daniela Steila, Scienza e rivoluzione. La recezione dell’empirio criticismo nella cultura russa (1877–1910), Firenze, Casa Editrice Le Lettere, 1996, 530 pp. (≪Università degli studi di Torino. Fondo di studi Parini-Chirio. Storia della filosofia e del pensiero scientifico, 3≫ ). Nuncius, 12(1):250–251, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00663.

Abbri:1997:BRV

[1692] Ferdinando Abbri. Book review: Vernon A. Rosario (ed.), Science and Homosexualities. New York-London, Routledge, 1997, x + 308 pp. Nuncius, 12(1):251–252, ???? 1997. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539197x00672.

Bertini:1997:BRG

[1693] Silvia Bertini. Book review: Gian Paolo Scano, Antonio Mastroianni e Giovanni Cadeddu, Psicoterapia psicoanalitica: verso una rifondazione. Storia, definizione, metodo, diagnosi, teoria, epistemologia, Milano, Franco Angeli, 1995, 316 pp. Nuncius, 12(1):253, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00681.

Pogliano:1997:BRAa

[1694] Claudio Pogliano. Book review: Alan Irwin and Brian Wynne (eds.), Misunderstanding science? the public reconstruction of science and technology, Cambridge, Cambridge University Press, 1996, 232 pp. Nuncius, 12(1):254, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00690.

Miniati:1997:BRSb

[1695] Mara Miniati. Book review: Silvia Ciappi, Anna Laghi, Maria Mendera, Daniela Stiaffini, Il vetro in Toscana. Strutture prodotti immagini (secc. XIII–XX). Poggibonsi, Lalli Editore, 1995, 189 pp., ill. Nuncius, 12 (1):254–256, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00708.

Andreini:1997:PCF

[1696] Francesca Andreini. Il progetto ≪Ca.Fo.S.≫ catalogo dei Fondi Scientifici Toscani. Nuncius, 12(1):257–258, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x00717.

Redondi:1997:FMD

[1697] Pietro Redondi. I fondamenti metafisici della fisica di Galileo. Nuncius, 12(2):267–289, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00726.

Rees:1997:SLC

[1698] Martin J. Rees. Stars, life and the Cosmos. Nuncius, 12(2):291–309, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00735.

Bucciantini:1997:DLG

[1699] Massimo Bucciantini. A difficult legacy: Galileo and the Galilean collection between myth and history. Nuncius, 12(2):311–328, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00744.

Ratcliff:1997:COM

[1700] Marc J. Ratcliff and Marino L. Buscaglia. Chercheur, observateur et monde savant au XVIIIe siècle: la recherche scientifique dans l’imaginaire de Charles Bonnet (1720–1793). Nuncius, 12(2):329–358, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00753.

Cavazza:1997:RDT

[1701] Marta Cavazza. La recezione della teoria Halleriana dell’irritabilita nell’Accademia Delle Scienze di Bologna. (Italian) [The reception of the Hallerian Theory of Irritability in the Bologna Academy of Sciences]. Nuncius, 12(2):359–377, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00762.

Bruzzaniti:1997:HMD

[1702] Giuseppe Bruzzaniti. A historiographic model of the dynamics of scientific theories and the formation of concepts: the case of the atomic nucleus. Nuncius, 12(2):379–410, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00771.

Brusa:1997:VSM

[1703] Giuseppe Brusa and John H. Leopold. The vertical stackfreed. Mechanical evidence and early eye-witnesses. Nuncius, 12(2):411–426, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00780.

Paternoster:1997:NLP

[1704] G. Paternoster, R. Rinzivillo, and E. Schettino. Nota su una lente per cannocchiale firmata Domenico Selva. (Italian) [A note on a telescope lens signed by Domenico Selva]. Nuncius, 12(2):427–431, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00799.

Barbieri:1997:TLG

[1705] Francesco Barbieri and Franca Cattelani Degani. Tre lettere di Geminiano Montanari a Gian Domenico Cassini. (Italian) [Three letters from Geminiano Montanari to Gian Domenico Cassini]. Nuncius, 12(2):433– 441, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00807.

Gatto:1997:CFS

[1706] Romano Gatto and Luciano Carbone. Il carteggio del Fondo Siacci della Biblioteca del Dipartimento di Matematica Renato Caccioppoli: Dell’Università Federico II di Napoli. Nuncius, 12(2):443–446, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00816.

Ciardi:1997:MLS

[1707] Marco Ciardi. Men of letters and science. The case of Xavier De Maistre (1763–1852). Nuncius, 12(2):447–469, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x00825.

Giudice:1997:THA

[1708] Franco Giudice. Thomas Hobbes and atomism: a reappraisal. Nuncius, 12(2):471–485, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00834.

Trombetta:1997:SB

[1709] Vincenzo Trombetta. Le scienze in biblioteca. Nuncius, 12(2):487–495, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00843.

Urbani:1997:LMC

[1710] Riccardo Urbani. L’uomo e la macchina. un congresso di trent’anni fa. Nuncius, 12(2):497–524, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00852.

Berni:1997:MSD

[1711] Marco Berni and Luisa Barattin. Il Museo di Storia della Scienza di Firenze su Internet un’analisi delle consultazioni. (Italian) [The Museum of the History of Science in Florence: An analysis of consulting on the Internet]. Nuncius, 12(2):525–531, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00861.

Mormino:1997:BRM

[1712] Gianfranco Mormino. Book review: Marco Beretta, Felice Mondella e Maria Teresa Monti (a cura di), Per una storia critica della Scienza, Quaderni di Acme 26, Università degli Studi di Milano, Dipartimento di Filosofia, Cisalpino, Bologna, 1996, 518 pp. s.i.p. Nuncius, 12(2): 533–536, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00870.

Segala:1997:BRA

[1713] Marco Segala. Book review: Alberto Masani La cosmologia nella storia fra scienza, religione e filosofia, Brescia, La Scuola, 1996, 410 pp., figg., Lit. 50.000 (ISBN 88-350-9212-4). Nuncius, 12(2):536–539, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00889.

Kirschke:1997:BRW

[1714] Martin Kirschke. Book review: Wolfgang Schneider, Werdegang und Zukunft der Pharmazie. Meine kleine Pharmaziegeschichte, Stuttgart, Deutscher Apotheker Verlag, 1995, 92 pp. Nuncius, 12(2):539–541, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00898.

Drbek:1997:BRW

[1715] P. Drbek. Book review: Wolf Dieter Muller-Jahncke and Christoph Friedrich, Geschichte der Arzneimitteltherapie, Stuttgart, Deutscher Apotheker Verlag, 1996, 295 pp., 48 ill. (ISBN 3-7692-2038-2). Nuncius, 12(2):541–542, ???? 1997. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539197x00906.

Lippi:1997:BRGb

[1716] Donatella Lippi. Book review: G. Cosmacini, G. Gaudenzi, R. Satolli, (a cura di), Dizionario di storia della salute, Torino, Einaudi, 1997, 678 pp., L. 130.000. Nuncius, 12(2):542–546, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00915.

Ciancio:1997:BRD

[1717] Luca Ciancio. Book review: David R. Oldroyd, Thinking about the Earth. A History of Ideas in Geology, London, Athlone, 1996, xxx + 410 pp., ill. b/w. Nuncius, 12(2):546–550, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x00924.

Meli:1997:BRM

[1718] Domenico Bertoloni Meli. Book review: M. D. Grmek, (a cura di), Storia del pensiero medico occidentale. 2. Dal Rinascimento all’inizio dell’Ottocento, Roma–Bari, Laterza, 1996, 613 pp., L. 65.000. Nuncius, 12(2):550–551, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00933.

diPasquale:1997:BRP

[1719] Giovanni di Pasquale. Book review: Pietro Janni, Il mare degli antichi, Bari, Dedalo, 1996, 500 pp., 79 fig. Nuncius, 12(2):551–554, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00942.

Siraisi:1997:BRJ

[1720] Nancy G. Siraisi. Book review: John Arrizabalaga, John Henderson, and Roger French, the Great Pox: the French Disease in Renaissance Europe, New Haven, Yale University Press, 1997, xv + 352 pp., 30 ill. Nuncius, 12(2):554–558, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x00951.

Clericuzio:1997:BRS

[1721] Antonio Clericuzio. Book review: Sachiko Kusukawa, the transformation of natural philosophy. The case of Philip Melanchton, Cambridge, Cambridge University Press, 1995, xv + 245 pp. Nuncius, 12(2):559–561, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00960.

Gorman:1997:BRS

[1722] Michael John Gorman. Book review: Steven Shapin, The Scientific Revolution, Chicago and London, University of Chicago Press, 1996, xiv + 218, (ISBN 0-226-75020-5). Nuncius, 12(2):561–564, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00979.

Redondi:1997:BRP

[1723] Pietro Redondi. Book review: Peter Dear, Discipline and Experience. the Mathematical Way in the Scientific Revolution, Chicago and London, the Chicago University Press, x + 290 pp., (Science and its conceptual foundations). Nuncius, 12(2):564–567, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x00988.

Kirschke:1997:BRA

[1724] Martin Kirschke. Book review: Almut Lanz, Arzneimittel in der Therapie Friedrich Hoffmanns (1660–1742) unter besonderer Berücksichtigung der Medicina Consultatoria (1721–1723), (German) [Drugs in the treatment of Friedrich Hoffmann (1660–1742), with special Consideration of Medicina Consultatoria (1721–1723)]. Braunschweig, 1995, 241 pp., (Braunschweiger Veröffentlichungen zur Geschichte der Pharmazie und der Naturwissenschaften Bd. 35). Nuncius, 12(2):567–570, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x00997.

Abbri:1997:BRP

[1725] Ferdinando Abbri. Book review: Patricia Fara, Sympathetic Attractions. Magnetic Practices, Beliefs, and Symbolism in Eighteenth-Century England, Princeton, NJ, Princeton University Press, 1996, xiv + 327 pp., ill. (ISBN 0-691-01099-4). Nuncius, 12(2):571–573, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01004.

Cavazza:1997:BRB

[1726] Marta Cavazza. Book review: Beate Ceranski, Und sie fürchtet sich vor niemandem. Die Physikerin Laura Bassi (1711–1778), (German) [And she is afraid of nobody: the physicist Laura Bassi (1711–1778)], Frankfurt/New York, Campus Verlag, 1996, 291 pp. Nuncius, 12(2):573–576, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01013.

Beretta:1997:BREb

[1727] Marco Beretta. Book review: Emmanuel Grison, Du Faubourg Montmartre au Corps des Mines. L’étonnant parcours du Republicain J.-H. Hassenfratz (1755–1827), Paris, École des Mines — Les Presses, 1996, xii + 387 pp. Ill. Nuncius, 12(2):576–578, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01022.

Vannini:1997:BRA

[1728] Fabrizio Vannini. Book review: Angelo Nannoni, Trattato sulle malattie delle mammelle, a cura di G. Parigino, Firenze, Le Lettere, 1995, 132 pp. (Introduzione di E. Stumpo; postfazione di C. Crisci; include L. Nannoni, Elogio del professore di chirurgia Angelo Nannoni, Firenze, Cambiagi, 1790). Nuncius, 12(2):578–582, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x01031.

Pogliano:1997:BRM

[1729] Claudio Pogliano. Book review: Marshall Sahlins, Capitan Cook, per esempio. Le Hawaii, gli antropologi, i nativi, a cura di Francesca Giunti, Roma, Donzelli, 1997, 288 pp. Nuncius, 12(2):582–584, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01040.

Miniati:1997:BRGb

[1730] Mara Miniati. Book review: Gabriella Rambaldi, Istrumenti di chimica. un laboratorio del XIX secolo, Genova, Pirella editore, 1996, 157 pp. ill. Nuncius, 12(2):585–586, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x01059.

Muller:1997:BRD

[1731] Gerhard H. Müller. Book review: Dietrich Von Engelhardt (hrsg.), Henrich Steffens: Was ich erlebte. Aus des Erinnerung niedergeschrieben. Neudruck der Ausgabe Breslau 1840/44. (German) [Henrich Steffens: What I experienced. From the written memoir. Reprint of the Wroclaw edition 1840/44]. 10 in 5 Bänden, Zusammen 79*, lxxii, 3898 Seiten, Leinen. Je Band DM 345/ S 2519/SFR 308. Stuttgart–Bad Canstatt: Frommann-Holzboog, 1995–1996. Nuncius, 12(2):586–590, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01068.

Gooday:1997:BRF

[1732] Graeme Gooday. Book review: Frank A. J. L. James (ed.), The Correspondence of Michael Faraday, Volume 3, 1841–1848 Letters 1334–2145, London, Institution of Electrical Engineers, 1996, v + 835 pp., 12 ill. (ISBN 0-86341-250-5). Nuncius, 12(2):590–593, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01077.

Shea:1997:BRG

[1733] William Shea. Book review: G. R. M. Garratt, the Early History of Radio From Faraday to Marconi, London, the Institution of Electrical Engineers, 1994, x + 93 pp. £19.00 (ISBN 0-85296-845-0). Nuncius, 12 (2):593–594, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x01086.

Scarantino:1997:BRJ

[1734] Luca M. Scarantino. Book review: Jean-François Stoffel (ed.), Mgr Georges Lemaitre, savant et croyant. Actes du colloque tenu Louvain-La-Neuve le 4 novembre 1994, Louvain-la-Neuve, Centre interfacultaire d’étude en histoire des sciences, 1996, 371 pp., 30 ill., (Reminisciences, 3) 1350 FB, (ISBN 2-930175-02-8). Nuncius, 12(2):594–599, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01095.

Gerstengarbe:1997:BRC

[1735] Sybille Gerstengarbe. Book review: Christoph Meinel und Monika Renneberg (Hrsg.), Geschlechterverhältnisse in Medizin, Naturwissenschaft und Technik. (German) [Gender relations in medicine, science and technology]. Hrsg. im Auftrag des Vorstandes der Deutschen Gesellschaft für Geschichte der Medizin, Naturwissenschaft und Technik.- Bassum, Stuttgart, Verlag für Geschichte der Naturwissenschaften und der Technik, 1996, 349 S. Nuncius, 12(2):599–605, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01103.

Bacciagaluppi:1997:BRJ

[1736] Guido Bacciagaluppi. Book review: Jeffrey Bub, Interpreting the Quantum World, Cambridge, Cambridge University Press, 1997, xiv + 298 pp., 21 ill. Nuncius, 12(2):606–610, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x01112.

Pogliano:1997:BRAb

[1737] Claudio Pogliano. Book review: Andr M. Hamende, (a cura di), Trent’anni di fisica con la bandiera delle Nazioni Unite a Trieste. Il Centro internazionale di fisica teorica (1964–1994), Trieste, Fondazione Internazionale per il Progresso e la Libert delle Scienze, 1996, 528 pp. Nuncius, 12(2):610–613, ???? 1997. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539197x01121.

Bowler:1997:BRS

[1738] Peter J. Bowler. Book review: Stephen Bocking, Ecologists and Environmental Politics: A History of Contemporary Ecology, New Haven, Yale University Press, 1997, 271 pp. 9 ill. Nuncius, 12(2):613–615, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01130.

Abbri:1997:I

[1739] Ferdinando Abbri. Istituzioni. Nuncius, 12(2):617, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x01149.

Miniati:1997:C

[1740] Mara Miniati. Cataloghi. Nuncius, 12(2):618, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x01158.

LaVergata:1997:BRG

[1741] Antonello La Vergata. Book review: Giovanni Paoloni e Raffaella Simili (a cura di), Guglielmo Marconi e l’Italia. Mostra storicodocumentaria (Roma, 30 marzo-30 aprile 1996). Catalogo, Roma, Accademia Nazionale dei Lincei, 1996, 240 pp. Nuncius, 12(2):619, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01167.

Miniati:1997:S

[1742] Mara Miniati. Strumentaria. Nuncius, 12(2):619–620, ???? 1997. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x01176.

Miniati:1997:BRAa

[1743] Mara Miniati. Book review: António Estácio Dos Reis, Medir Estrelas. (Spanish) [Measuring Stars], s.l., CTT Correios, 1997, 168 pp., ill. (ISBN 972-9127-40-9). Nuncius, 12(2):620–621, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01185.

Miniati:1997:BRAb

[1744] Mara Miniati. Book review: António Estácio Dos Reis, O Único exemplar vivo do nónio de Pedro Nunes?, (Portuguese) [The one live specimen of the vernier of Pedro Nunes?], Lisboa, Academia de Marinha, 1995, 44 pp., ill. Nuncius, 12(2):621–622, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x01194.

Miniati:1997:BRMb

[1745] Mara Miniati. Book review: Maria Basso Ricci, Lili Cafarella, Antonio Meloni, Pasquale Tucci, Due secoli di strumenti geomagnetici in Italia (1740–1971), Bologna, Editrice Compositori, 1997, 234 pp., ill. (ISBN 88-7794-091-3). Nuncius, 12(2):622–623, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01202.

Miniati:1997:BRMc

[1746] Mara Miniati. Book review: Manlio Calegari, Guido Gabrietti, Saper fare. Ricordi di un bilanciaio di Campogalliano, a cura del Centro Documentazione del Museo della Bilancia, Bologna, Clueb, 1996, 179 pp. + 19 fig. f.t., ill. (ISBN 88-8091-370-0). Nuncius, 12(2):623–624, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01211.

Abbri:1997:A

[1747] Ferdinando Abbri. Astronomia. Nuncius, 12(2):624–625, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01220.

Miniati:1997:BRGc

[1748] Mara Miniati. Book review: Giorgia Foder Serio, Donatella Randazzo, Astronomi italiani dall’Unit d’Italia ai nostri giorni: un primo elenco, Firenze, Societ Astronomica Italiana Editore, 1997, 116 pp. Nuncius, 12 (2):625–626, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x01239.

Miniati:1997:BRP

[1749] Mara Miniati. Book review: E. Proverbio (ed.), Astronomical observatories and institutions in Italy, Seventh annual meeting on the History of Astronomy, Milano, April 21–22 1995, Memorie della Societ Astronomica italiana-Journal of the Italian Astronomical Society, vol. 66, N. 4, 1995, pp. 751–936, ill. Nuncius, 12(2):626–627, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01248.

Miniati:1997:BRI

[1750] Mara Miniati. Book review: Ileana Chinnici, An Italian observatory in India: the history of the Calcutta Observatory, Studies in History of Medicine and Science, 1995-96, vol. 14, 21 pp. Nuncius, 12(2):627, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01257.

Lippi:1997:M

[1751] Donatella Lippi. Medicina. Nuncius, 12(2):628–629, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x01266.

Lippi:1997:BRGc

[1752] Donatella Lippi. Book review: G. Cosmacini, Medici nella storia d’Italia. Per una tipologia della professione medica, Bari, Laterza, 1996, 218 pp. Nuncius, 12(2):629–630, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x01275.

Lippi:1997:BRMb

[1753] Donatella Lippi. Book review: M. D. Grmek, Il calderone di medea. La sperimentazione sul vivente nell’antichità, (Italian) [The cauldron of Medea. The experiment of living in antiquity], Bari, Laterza, 1996, 144 pp. Nuncius, 12(2):630–631, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x01284.

Lippi:1997:BRR

[1754] Donatella Lippi. Book review: Roberto Cardini e Mariangela Regogliosi (a cura di), Umanesimo e medicina. Il problema dell’individuale, Roma, Bulzoni, 1996, 99 pp., £22.000. Nuncius, 12(2):631–632, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01293.

Lippi:1997:BRL

[1755] Donatella Lippi. Book review: Lucia Sandri (a cura di), Gli Innocenti e Firenze nei secoli. un ospedale, un archivio, una citt, Firenze, S.P.E.S., 1996, 199 pp. Nuncius, 12(2):632–633, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01301.

Lippi:1997:BRMc

[1756] Donatella Lippi. Book review: E. Marelli, Un santo, un re, una città. Storia dell’Ospedale di Monza, Bari, Laterza, 1996, 291 pp. Nuncius, 12(2):633, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x01310.

Ciardi:1997:ST

[1757] Marco Ciardi. Scienza e tecnica. Nuncius, 12(2):634, ???? 1997. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x01329.

Ciardi:1997:BRM

[1758] Marco Ciardi. Book review: Martin Bauer (ed.), Resistance to new technology. Nuclear power, information technology and biotechnology, Cambridge, Cambridge University Press, 1995, 422 pp. (ISBN 0-521-45518-9). Nuncius, 12(2):635, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x01338.

Ciardi:1997:BRE

[1759] Marco Ciardi. Book review: Enrico Decleva, Carlo G. Lacaita e Angelo Ventura (a cura di), Innovazione e modernizzazione in Italia fra Otto e Novecento, Milano, Franco Angeli, 1995, 658 pp. Nuncius, 12 (2):636–637, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x01347.

Casati:1997:A

[1760] Stefano Casati. Antichità. Nuncius, 12(2):637–638, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x01356.

Abbri:1997:UR

[1761] Ferdinando Abbri. Umanesimo e rinascimento. Nuncius, 12(2):638–640, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01365.

Abbri:1997:SXX

[1762] Ferdinando Abbri. Secoli XVII–XX. Nuncius, 12(2):640–641, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01374.

Abbri:1997:BRG

[1763] Ferdinando Abbri. Book review: Geoffrey V. Sutton, Science for a Polite Society. Gender, Culture, and the Demonstration of Enlightenment, Boulder, Colorado and Oxford, Westview Press, 1995, xiii + 391 ill. (ISBN 0-8133-1575-1). Nuncius, 12(2):641–643, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01383.

Monti:1997:BRU

[1764] Maria Teresa Monti. Book review: Urs Boschung, Albrecht von Haller in Göttingen. Briefe und Selbstzeugnisse, Bern, Verlag Hans Huber, 1994, 110 pp., ill. Nuncius, 12(2):643–644, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x01392.

Beretta:1997:BRP

[1765] Marco Beretta. Book review: Pericle di Pietro (a cura di), Edizione nazionale delle opere di Lazzaro Spallanzani. Parte Quarta. Opere edite direttamente dall’Autore. Vol. 1 (1760–1768), Modena, Mucchi Editore, 1996, 259 pp. 111. Nuncius, 12(2):644–645, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01400.

Miniati:1997:BRGd

[1766] Mara Miniati. Book review: I Georgofili e il mondo delle api, (Italian) [The Georgofili and the world of bees], Monsummano Terme, Accademia dei Georgofili, 1997, 15 pp. + 13 tav. f.t. Nuncius, 12(2):645– 646, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01419.

Segala:1997:BRL

[1767] Marco Segala. Book review: Leonardo Pica Ciamarra, l’antropologia di Schopenhauer, Napoli, Loffredo, 1996, 191 pp., Lit. 22.000 (ISBN 88-8096-448-8). Nuncius, 12(2):646–647, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539197x01428.

Miniati:1997:BRGe

[1768] Mara Miniati. Book review: Giacomo Ricci, David Rovai, l’Istituto Tecnico Francesco Carrara 1896–1996. Cento anni di vita, Lucca, Maria Pacini Fazzi editore, 1996, 246 pp., ill. (ISBN 88-7246-211-8) (Collana di cultura e storia lucchese, 18). Nuncius, 12(2):647–648, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01437.

Bertini:1997:BRE

[1769] Silvia Bertini. Book review: Elizabeth Both Spillius (a cura di), Melanie Klein e il suo impatto sulla psicoanalisi oggi, Roma, Astrolabio, 1995, vol. I. La teoria, 381 pp., vol. II La pratica, 354 pp. Nuncius, 12(2): 648–649, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539197x01446.

Gorman:1997:CAK

[1770] Michael John Gorman. The correspondence of Athanasius Kircher: the world of a Seventeenth Century Jesuit. An international research project. Nuncius, 12(2):651–658, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539197x01455.

Anonymous:1997:IAX

[1771] Anonymous. Indici (anno XII, 1997). Nuncius, 12(2):659–660, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01464.

Anonymous:1997:IDF

[1772] Anonymous. Indice delle fonti manoscritte. Nuncius, 12(2):661–662, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01473.

Anonymous:1997:IN

[1773] Anonymous. Indice dei nomi. Nuncius, 12(2):663–682, ???? 1997. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539197x01482.

Laurenza:1998:CCF

[1774] Domenico Laurenza. La composizione del corpo. Fisiognomica ed embriologia in Leonardo. Nuncius, 13(1):3–37, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539198x00013.

Couzinet:1998:NMT

[1775] Marie-Dominique Couzinet. Notes sur les medicinalia de Tommaso Campanella. Nuncius, 13(1):39–67, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00022.

Clericuzio:1998:MPS

[1776] Antonio Clericuzio. The mechanical philosophy and the spring of air. New light on Robert Boyle and Robert Hooke. Nuncius, 13(1):69–75, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00031.

Funaro:1998:MMM

[1777] Liana Elda Funaro. ≪mezzi, metodi e macchine≫. Notizie su Giuseppe Morosi. Nuncius, 13(1):77–137, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00040.

Biagini:1998:OVN

[1778] Carlo Biagini. Ospedali vecchi e nuovi: Il dibattito tecnico-culturale sul rinnovamento delle strutture ospedaliere nell’Italia post-unitaria. Il caso del Santa Maria Nuova a Firenze. (Italian) [Hospitals old and new: The debate on technical and cultural renewal of hospitals of post-unitarian Italy. The case of Santa Maria Nuova in Florence]. Nuncius, 13(1):139– 150, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00059.

Scianna:1998:SIS

[1779] Nicolangelo Scianna. Strumentaria indagine sui grandi globi a stampa di Vincenzo Coronelli. Nuncius, 13(1):151–168, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539198x00068.

Ferrarese:1998:SCM

[1780] Giorgio Ferrarese and Franco Palladino. Sulle collezioni di modelli matematici dei Dipartimenti di Matematica Dell’Università e Del Politecnico di Torino. Nuncius, 13(1):169–185, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539198x00077.

Brenni:1998:SSM

[1781] Paolo Brenni and Anna Giatti. Lo stereocartografo Santoni Modello IV: Cronaca di un salvataggio. (Italian) [The Santoni stereocartograph Model IV: Chronicle of a rescue]. Nuncius, 13(1):187–191, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00086.

Guerrini:1998:PAD

[1782] Luigi Guerrini. Per un archivio della corrispondenza degli scienziati italiani due lettere inedite di Tommaso Frosini a Francesco Redi sul de motu animalium di Giovanni Alfonso Borelli. Nuncius, 13(1):193–208, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00095.

Triarico:1998:CLX

[1783] Carlo Triarico. La corrispondenza di Leonardo Ximenes. Inventario delle filze del carteggio conservate nel Fondo Nazionale della Biblioteca Nazionale Centrale di Firenze. (Italian) [The correspondence of Leonardo Ximenes. Inventory of the correspondence strings retained by the National Fund of the National Central Library of Florence]. Nuncius, 13 (1):209–246, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00103.

Barchielli:1998:IFI

[1784] Silvia Barchielli. Istituzioni e Fonti l’Istituto Vaccinogeno All’Ospedale di Santa Maria degli Innocenti di Firenze nel XVIII Secolo. Nuncius, 13 (1):247–263, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00112.

Valerio:1998:DCC

[1785] Vladimiro Valerio. Discussioni critiche cognizioni proiettive e prospettiva lineare nell’opera di Tolomeo e nella cultura tardo-ellenistica. Nuncius, 13(1):265–298, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00121.

Pogliano:1998:BRF

[1786] Claudio Pogliano. Book review: Frank Spencer (ed.), History of Physical Anthropology. An Encyclopedia, New York & London, Garland Publishing, Inc., 1997, 2 voll., xviii + 1196 pp. Nuncius, 13(1):299–301, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00130.

Miniati:1998:BRGa

[1787] Mara Miniati. Book review: Giovanni Paltrinieri, Meridiane e orologi solari d’Italia, Bentivoglio (Bo), l’Artiere Edizionitalia, 1997, 263 pp., ill. Nuncius, 13(1):301–302, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00149.

Parigi:1998:BRM

[1788] Silvia Parigi. Book review: Maria Teresa Monti (a cura di), Teorie della visione e problemi di percezione visiva nell’età moderna, Milano, Franco Angeli, 1995, 228 pp. Nuncius, 13(1):302–305, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00158.

Clericuzio:1998:BRB

[1789] Antonio Clericuzio. Book review: Benedino Gemelli, Aspetti dell’Atomismo Classico nella Filosofia di Francis Bacon e nel Seicento, Firenze, Olschki, 1996, 433 pp. Nuncius, 13(1):306–307, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00167.

Betti:1998:BRA

[1790] Gian Luigi Betti. Book review: Annibale Fantoli, Galileo per il copernicanesimo e per la Chiesa, seconda edizione riveduta e corretta. ≪Studi Galileiani, vol. 2≫, Città del Vaticano, Specola Vaticana — Libreria Editrice Vaticana, 1997, xviii-540 pp. Nuncius, 13(1):307–309, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00176.

Ciardi:1998:BRC

[1791] Marco Ciardi. Book review: Cesare Maffioli, Out of Galileo. The Science of Waters, 1628–1718. Foreword by Richard S. Westfall, Rotterdam, Erasmus Publishing, 1994, xx + 509 pp., ill., (≪Nieuwe Nederlandse Bijdragen tot de Geschiedenis der Geneeskunde en der Natuurwetenschappen, 49≫ ) ISBN 90-5235-071-X. Nuncius, 13(1):309–311, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00185.

Miniati:1998:BRP

[1792] Mara Miniati. Book review: Peter De Clerq, at the sign of the Oriental Lamp. The Musschenbroek workshop in Leiden, 1660–1750, Rotterdam, Erasmus publishing, 1997, 326 pp., ill. (ISBN 90-5235-103-1). Peter De Clerq, the Leiden Cabinet of Physics. A descriptive Catalogue, Leiden, Museum Boerhaave, 1997, 198 pp., ill. (ISBN 90-6292-114-0). Nuncius, 13(1):311–313, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00194.

Barsanti:1998:BRC

[1793] Giulio Barsanti. Book review: Claude Blanckaert (a cura di), Le terrain des sciences humaines. Instructions et enquêtes (XVIII–XX siècle), Paris, l’Harmattan, 1996, 412 pp. Nuncius, 13(1):313–316, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00202.

Miniati:1998:BRC

[1794] Mara Miniati. Book review: Carlo Alberto Zotti Minici (a cura di), Mirabili visioni. Vedute ottiche della stamperia Remondini. Catalogo. Contributi e schede di Mariano Avanzo, Gian Piero Brunetta, Lia Camerlengo, Thomas Ganz, Paola Marini, Carlo Alberto Zotti Minici, Trento, Provincia Autonoma di Trento, Servizio Beni Culturali, 1996, 159 pp., ill. Nuncius, 13(1):316–318, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00211.

Abbri:1998:BRP

[1795] Ferdinando Abbri. Book review: Patrice Bret (Editeur), Oeuvres de Lavoisier. Correspondance. Volume VI. 1789–1791. Publié sous les auspices du Comité Lavoisier de l’Académie des sciences, Paris, Académie des sciences, 1997, xviii + 494 pp. Diffusion: Librairie scientifique et technique A. Blanchard, Paris (ISBN 2-85367-205-0). Nuncius, 13(1): 318–319, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00220.

Segala:1998:BRJ

[1796] Marco Segala. Book review: Johann Wolfgang Goethe, La storia dei colori, a cura di Renato Troncon, Milano-Trento, Luni Editrice, 1997, ill., 494 pp., Lit. 60.000. Nuncius, 13(1):320–322, ????

1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00239.

Reinhardt:1998:BRI

[1797] Carsten Reinhardt. Book review: Irene Scheifele, Festschriften der Versammlungen Deutscher Naturforscher und Ärzte 1822 bis 1920. Stuttgart: Wissenschaftliche Verlagsgesellschaft 1996 (Schriftenreihe zur Geschichte der Versammlungen Deutscher Naturforscher und Ärzte, Band 6). 106 S., 4 Farbtafeln, kart., DM 48 (ISBN 3-8047-1050-6). Nuncius, 13(1):322–323, ???? 1998. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539198x00248.

Vista:1998:BRC

[1798] Michele Vista. Book review: Colin A. Russell, Edward Frankland: chemistry, controversy, and conspiracy in Victorian England, Cambridge, Cambridge University Press, 1996, xx + 536 pp., ill. Nuncius, 13(1): 323–326, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00257.

Pogliano:1998:BRH

[1799] Claudio Pogliano. Book review: Henri Berret, La culture du XXe siècle. Histoire, science, philosophie, sous la direction de Agnès Biard, Dominique Bourel, Eric Brian, Paris, Albin Michel/Centre international de synthèse, 1997, 366 pp. Nuncius, 13(1):326–329, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00266.

Vannozzi:1998:BRH

[1800] Francesca Vannozzi. Book review: Harry H. Marks, the progress of experiment. Science and Therapeutic Reform in the United States, 1900–1990, Cambridge, Cambridge University Press, 1997, xii + 258 pp. (ISBN 0-521-58142-7 hardback). Nuncius, 13(1):330–335, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00275.

Barone:1998:BRP

[1801] Vincenzo Barone. Book review: Paolantonio Marazzini, Nuove radiazioni, quanti e relatività in Italia, 1896–1925, Pavia, La Goliardica Pavese, 1996, 199 pp. Nuncius, 13(1):335–338, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539198x00284.

Tugnoli:1998:BRH

[1802] Claudio Tugnoli. Book review: Henri Bergson, Durata e simultaneità (a proposito della teoria di Einstein), a cura di Paolo Taroni, Pitagora editrice, Collana di Storia delle idee, sezione Fonti e Documenti, Bologna, 1997, 209 pp. Nuncius, 13(1):338–341, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00293.

Segala:1998:BRC

[1803] Marco Segala. Book review: Collane ≪I fondamenti della scienza≫, collana a cura di Antonio di Meo, nn. 1–20, Roma, Edizioni Teknos, 1994– 1996. Nuncius, 13(1):343–344, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00301.

Beretta:1998:BRC

[1804] Marco Beretta. Book review: Cataloghi Roland Schaer (a cura di), Tous les savoirs du monde. Encyclopédies et bibliothèques, de Sumer au XXIe siècle, Paris, Flammarion, 1996, 495 pp., ill. Nuncius, 13(1):344–345, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00310.

Abbri:1998:BRO

[1805] Ferdinando Abbri. Book review: Ove Hagelin (compiler), Old and Rare Books on Materia Medica in the Library of the Swedish Pharmaceutical Society (Apotekarsocieteten). An Illustrated and Annotated Catalogue, Stockholm, Swedish Pharmaceutical Press, 1997, 223 pp. Nuncius, 13 (1):345–346, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00329.

Miniati:1998:BRB

[1806] Mara Miniati. Book review: La bona opinione. Cultura, scienza e misure negli stati e setensi, 1598–1860, a cura di Debora Dameri, Achille Lodovisi, Giulia Luppi, Campogalliano, Museo della Bilancia/Centro di Documentazione, 1997, 292 pp., ill., L. 60.000 (ISBN 88-86143-08-7). Nuncius, 13(1):346–347, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00338.

Miniati:1998:BRSa

[1807] Mara Miniati. Book review: Strumentaria Emilio Borchi, Renzo Macii, Termometri e termoscopi, Firenze, Tipografia Coppini, 1997, 159 pp., ill. Nuncius, 13(1):347–348, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00347.

Brenni:1998:BRW

[1808] Paolo Brenni. Book review: E. W. Williams, the Precision makers, A history of the instrument industry in Britain and France, 1870–1939, London and New York, Rutledge, 1994, 216 pp. (ISBN 0-415-03732-8). Nuncius, 13(1):348–349, ???? 1998. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539198x00356.

Lippi:1998:BRMa

[1809] Donatella Lippi. Book review: Medicina e Botanica B. Zanobio, G. Armocida, Storia della medicina, Milano, Biblioteca Medica Masson, 1997, 339 pp., L. 54.000. Nuncius, 13(1):349–350, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00365.

Lippi:1998:BRR

[1810] Donatella Lippi. Book review: R. Rizzi, Storia della terapia antalgica. Dalla preistoria all’evo contemporaneo, (Italian) [Analgesic therapy history. From prehistory to simultaneous developments] Milano, Ciba Ed., 1996, 275 pp. Nuncius, 13(1):350–351, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00374.

Beretta:1998:BRM

[1811] Ilva Beretta. Book review: Massimo Venturi Ferriolo (a cura di), Mater Herbarum. Fonti e tradizione del giardino dei semplici della Scuola Medica Salernitana, Milano, Guerini e Associati, 1995, 299 pp., ill. (Kepos, Quaderni 6). Nuncius, 13(1):351–352, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00383.

Lippi:1998:BRJ

[1812] Donatella Lippi. Book review: J. Woodward and R. Juette (ed. by), Coping with sickness, Historical aspects of Health care in a European perspective. Sheffield, 1995, 225 pp. Coping with sickness, Perspectives on Health care, past and present, Sheffield, 1995, 221 pp. Nuncius, 13 (1):352–353, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00392.

Miniati:1998:BRF

[1813] Mara Miniati. Book review: Fisica e Chimica P. Tucci (a cura di), Atti del XVI Congresso Nazionale di Storia della Fisica e dell’Astronomia. Centro Volta, Villa Olmo, Como 24–25 maggio 1996, Lipomo (Como), Tipografia Editrice Cesare Nanni, 1997, 715 pp., ill. Nuncius, 13(1): 353–354, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00400.

Vista:1998:BRM

[1814] Michele Vista. Book review: Matthew E. Hermes, Enough for one lifetime: Wallace Carothers, inventor of nylon, History of Modern Chemical Sciences, American Chemical Society and the Chemical Heritage Foundation, 1996, Maple Press, York, PA, 345 pp., ill. Nuncius, 13(1):355–356, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00419.

Bertini:1998:BRP

[1815] Silvia Bertini. Book review: Psicologia Gian Vittorio Caprara, Accursio Gennaro, Psicologia della personalità. Storia, indirizzi teorici e temi di ricerca, Bologna, Il Mulino, 1994, 563 pp. Nuncius, 13(1):357–359, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00428.

Beretta:1998:BRA

[1816] Marco Beretta. Book review: Antichità e Medioevo Marco Buonocore (a cura di), Vedere i classici. l’illustrazione libraria dei testi antichi dall’età romana al tardo medioevo, Roma, Fratelli Palombi Editore, 1996, xxiv + 596 pp. Nuncius, 13(1):359–360, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00437.

Beretta:1998:BRS

[1817] Marco Beretta. Book review: Secoli XVI–XX Fernanda Ascarelli, La tipografia cinquecentina italiana, Firenze, Casa Editrice Le Lettere, 1996, 259 pp., ill. Nuncius, 13(1):361, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00446.

Abbri:1998:BRFa

[1818] Ferdinando Abbri. Book review: Fiona Haslam, From Hogarth to Rowlandson. Medicine in Art in Eighteenth-Century Britain, Liverpool, Liverpool University Press, 1996, xvi + 336 pp., ill. L. 39.99 (paperback L. 18.99) (ISBN 0-85323-630-5). Nuncius, 13(1):361–363, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00455.

Abbri:1998:BRC

[1819] Ferdinando Abbri. Book review: Christa Jungnickel and Russell Mccormmach, Cavendish, Philadelphia, Pennsylvania, the American Philosophical Society, 1996, xii + 414 pp., ill. (ISBN 0-87169-220-1) (≪Memoirs of the American Philosophical Society Held at Philadelphia for Promoting Useful Knowledge≫, Volume 220). Nuncius, 13(1):363–366, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00464.

Ciardi:1998:BRM

[1820] Marco Ciardi. Book review: Maria Luisa Berti, Duccio Bigazzi (a cura di), Ricerche di storia in onore di Franco Della Peruta, 2 voll., Milano, Franco Angeli, 1996, 602 e 629 pp. (ISBN 88-204-9854-5 e 88-204-9853-7). Nuncius, 13(1):366–368, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00473.

Rossi:1998:PZP

[1821] Paolo Rossi. I punti di zenone: Una preistoria vichiana. Nuncius, 13 (2):378–425, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00482.

Giusti:1998:ERC

[1822] Enrico Giusti. Elements for the relative chronology of Galilei’s de motu antiquiora. Nuncius, 13(2):427–460, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539198x00491.

Cavagna:1998:OLA

[1823] Anna Giulia Cavagna. Opere e libri di un astronomo cartografo del XVIII secolo: Tra erudizione e stato. Nuncius, 13(2):461–491, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00509.

Casati:1998:SFD

[1824] Stefano Casati. Storie di folgori: Il dibattito italiano sui conduttori elettrici nel settecento. Nuncius, 13(2):493–512, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539198x00518.

Vlahakis:1998:RET

[1825] George N. Vlahakis. The reception of electric theories during the dawn of NeoHellenic scientific thought. Nuncius, 13(2):513–532, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00527.

Monaco:1998:SAC

[1826] Giuseppe Monaco. Strumentaria alcune considerazioni sul Maximus Tubus di Hevelius. Nuncius, 13(2):533–550, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539198x00536.

Misiti:1998:MPS

[1827] Massimo Misiti. Macchine per studiare. Nuncius, 13(2):551–565, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00545.

Cicenia:1998:PAD

[1828] Salvatore Cicenia. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 13(2):567–593, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00554.

Camerota:1998:IFA

[1829] Michele Camerota. Istituzioni e fonti adattar la volgar lingua al filosofici discorsi. Una inedita orazione di Niccolo Aggiunti contro Aristotele e per l’uso della lingua italiana nelle dissertazioni scientifiche. (Italian) [Sources and institutions adapt the vulgar language in philosophical discourses. An unpublished oration of Niccolo Aggiunti against Aristotle and for the use of the Italian language in scientific dissertations]. Nuncius, 13 (2):595–623, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00563.

Ciardi:1998:TTA

[1830] Marco Ciardi. Theory and technology: the Avogadro manuscripts at the Turin Academy of Sciences. Nuncius, 13(2):625–656, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00572.

diPasquale:1998:SPF

[1831] Giovanni di Pasquale. La stadera: un problema di filologia, storia ed archeologia. Nuncius, 13(2):657–666, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539198x00581.

Vasoli:1998:JBS

[1832] Cesare Vasoli. Jean Bodin e la scienza del rinascimento. Nuncius, 13 (2):667–675, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00590.

Segala:1998:RSS

[1833] Marco Segala. Riflessioni sulla scienza romantica. Nuncius, 13(2):677– 691, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00608.

diPasquale:1998:BRM

[1834] Giovanni di Pasquale. Book review: T. Mannoni, E. Giannichedda, Archeologia della produzione, Torino, Einaudi, 1996, 352 pp., ill. Nuncius, 13(2):693–694, ???? 1998. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539198x00617.

Bettini:1998:BRS

[1835] Amalia Bettini. Book review: Sara Schechner Genuth, Comets, Popular Culture, and the Birth of Modern Cosmology, Princeton, Princeton University Press, 1997, 365 pp., 53 ill. Nuncius, 13(2):694–696, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00626.

Strano:1998:BRC

[1836] Giorgio Strano. Book review: Christopher Walker (a cura di), l’astronomia prima del telescopio, Bari, Dedalo s.r.l., 1997, 504 pp., ill. + 14 tavv. f.t., £50.000 (ISBN 88-220-0543-0). Nuncius, 13(2):696–699, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00635.

Stocking:1998:BRI

[1837] George W. Stocking. Book review: Ivan Hannaford, Race: the History of an Idea in the West. Washington, D.C., Baltimore, and London, Woodrow Wilson Center Press and The Johns Hopkins University Press, 1996, xviii + 448 pp. (ISBN 0-8018-5223-4). Nuncius, 13(2):699–702, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00644.

Muller:1998:BRE

[1838] Gerhard H. Müller. Book review: Eckart Olshausen und Holger Sonnabend (Hrsg.) Naturkatastrophen in der antiken Welt. Stuttgarter Kolloquium zur Historischen Geographie des Altertums 6, 1996. (= Geographica historica, begründet von Ernst Kirsten herausgegeben von Eckart Olshausen, Band 10). Stuttgart, Steiner, 1998. 485 S. mit zahlr. Abb., kt., DM 198. (Subskript.-Preis DM 160.-) (ISBN 3-515-07252-7). Nuncius, 13(2):702–704, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00653.

diPasquale:1998:BRL

[1839] Giovanni di Pasquale. Book review: Lawrence Richardson, Jr., Pompeii: An Architectural History, The Johns Hopkins University Press, Baltimore and London, 1997, 445 pp., ill. Nuncius, 13(2):704–707, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00662.

Savoia:1998:BRF

[1840] Andrea Ubrizsy Savoia. Book review: S. Ferri, (a cura di), Pietro Andrea Mattioli (Siena 1501–Trento 1578), La vita, le opere, Ponte San Giovanni, Perugia, Quattroemme srl, 1997, 405 pp. (ISBN 88-85962-27-0). Nuncius, 13(2):707–709, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00671.

Jacob:1998:BRD

[1841] James R. Jacob. Book review: Denise Albanese, New Science, New World. Durham and London: Duke University Press, 1996. xi + 244 pp. Library Cloth Edition (ISBN 0-8223-1759-1). £43.95. Paper (ISBN 0-8223-1768-0). £14.95. Nuncius, 13(2):709–710, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00680.

Bernardi:1998:BRC

[1842] Walter Bernardi. Book review: Catherine Wilson, the Invisible World. Early Modern Philosophy and the Invention of the Microscope, Princeton, Princeton University Press, 1995, 280 pp. Nuncius, 13(2):710–712, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00699.

Meli:1998:BRG

[1843] Domenico Bertoloni Meli. Book review: S. Gaukroger, Descartes. An intellectual biography, Oxford, University Press, 1995 (paperback 1997), xx + 500 pp. (ISBN 0-19-823724-3). Nuncius, 13(2):712–714, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00707.

Shea:1998:BRA

[1844] William Shea. Book review: Alan Cook, Edmond Halley. Charting the Heavens and the Seas. Oxford: Clarendon Press, 1998, 540 pp. (ISBN 0-19-850031-9). Nuncius, 13(2):714–716, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539198x00716.

Vidal:1998:BRC

[1845] Fernando Vidal. Book review: Christopher Fox, Roy Porter and Robert Wokler (eds.), Inventing human science. Eighteenth-century domains. Berkeley, University of California Press, 1995. xiii + 357 pp. $45. (ISBN 0-520-20010-1). Nuncius, 13(2):716–720, ????

1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00725.

Abbri:1998:BRFb

[1846] Ferdinando Abbri. Book review: Frederic Lawrence Holmes, Antoine Lavoisier — the Next Crucial Year or the Sources of His Quantitative Method in Chemistry, Princeton, New Jersey, Princeton University Press, 1998, viii + 184 pp. (ISBN 0-691-01687-9). Nuncius, 13(2):720– 722, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00734.

Roll-Hansen:1998:BRM

[1847] Nils Roll-Hansen. Book review: Michael Ruse, Monad to man. the Concept of Progress in Evolutionary biology, Harvard University Press, 1996, pp. 628 (ISBN 0-674-58220-9). Nuncius, 13(2):722–725, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00743.

Nani:1998:BRB

[1848] Michele Nani. Book review: Bernadette Bensaude-Vincent et Anne Rasmussen (sous la direction de), La science populaire dans la presse et l’édition. XIXe et XXe siècles, Paris, CNRS Éditions, 1997, 299 pp., s.i.p. Nuncius, 13(2):725–728, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00752.

Mestre:1998:BRJ

[1849] Josep Bernabeu Mestre. Book review: José Mara López Piñero, Jorge Navarro López. Los estudios sobre la salud pública en la ciudad de Valencia 1880–1890. Constantino Gómez Reig (Studies in public health in the city of Valencia, 1880–1890. Constantino Gömez Reig). Valéncia: Ajuntament de Valéncia (Collection Científicos Valencianos), 1994, 235 pp. (ISBN 84-88639-32-5). Nuncius, 13(2):728–729, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00761.

Vinti:1998:BRM

[1850] Carlo Vinti. Book review: Mirella Fortino, Convenzione e razionalit scientifica in Henri Poincaré, presentazione di Franco Crispini, Rubbettino, Soveria Mannelli, 1997, 268 pp. L 30.000 (ISBN 88-7284-597-1). Nuncius, 13(2):729–732, ???? 1998. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539198x00770.

Polizzi:1998:BRP

[1851] Gaspare Polizzi. Book review: Pietro Greco, Franco Prattico, Giorgio Rivieccio, Emanuele Vinassa De Regny, Toccare le stelle. Scienza e tecnologia nel Novecento, Napoli, Cuen, 1997, 245 pp., L. 25.000 (ISBN 88-7146-356-0). Nuncius, 13(2):732–736, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00789.

Andreozzi:1998:BRK

[1852] Luciano Andreozzi. Book review: Karl Sigmund, Games of Life. Explorations in Ecology, Evolution and Behavior, Penguin Books 1995 (prima edizione Oxford University Press, 1993), vi + 244 pp., £7.99, ISBN assente. Nuncius, 13(2):736–739, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00798.

Abbri:1998:BRJ

[1853] Ferdinando Abbri. Book review: James H. Jones, Alfred C. Kinsey. A Public/Private Life, New York — London, W.W. Norton Company, 1997, xx + 937 pp. £28 (ISBN 0-393-04086-0). Nuncius, 13(2):739–742, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00806.

Schaller:1998:BRW

[1854] Angelika Schaller. Book review: Werner Weizig (hrsg.) Arthur Schnitzler: Tagebuch 1927 bis 1930. Unter Mitwirkung von Peter Michael Braunwarth, Susanne Pertlik und Reinhard Urbach, herausgegeben Verlag der Österreichischen Akademie der Wissenschaften, Wien, 1997, 510 Seiten, 67 Mark (ISBN 3-7001-2120-2). Nuncius, 13(2):742–745, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00815.

Perdiguero:1998:BRJ

[1855] Enrique Perdiguero. Book review: José Luis Fresquet (ed.), Salud, enfermedad y terapeutica Popular en la Ribera Alta. Valencia: Instituto de Estudios Documentales e Histricos sobre la Ciencia, 1995, 322 pp. Nuncius, 13(2):746–748, ???? 1998. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539198x00824.

Pogliano:1998:BRL

[1856] Claudio Pogliano. Book review: Luigi Luca Cavalli-Sforza, Paolo Menozzi, Alberto Piazza, Storia e geografia dei geni umani, Milano, Adelphi, 1997, xxii + 794 pp. Nuncius, 13(2):748–751, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00833.

Stoltzner:1998:BRD

[1857] Michael Stoltzner. Book review: Domenico Costantini, Maria Carla Galavotti (Eds.), Probability, Dynamics and Causality. Essays in Honour of Richard C. Jeffrey, Dordrecht, Kluwer Academic Publishers, 1997 (reprint from Erkenntnis Vol. 45, Nos. 2/3, 1996), Nlg 135, US$88, GB £59 (ISBN 0-7923-4361-1). Nuncius, 13(2):751–755, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00842.

Casati:1998:SB

[1858] Stefano Casati. Strumenti bibliografici. Nuncius, 13(2):757–760, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00851.

Ciardi:1998:BRJ

[1859] Marco Ciardi. Book review: Jean-François Stoffel (tablié par), Pierre Duhem et ses doctorands. Bibliographie de la littérature primaire et secondaire. Introduction de Stanley L. Jaki, Louvain-la-Neuve, Centre interfacultaire d’étude en histoire des sciences, 1996, 326 pp., 1200 fr. (ISBN 2-930175-00-1). Nuncius, 13(2):760–761, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00860.

Miniati:1998:BRSb

[1860] Mara Miniati. Book review: Strumentaria Instrumentos cientificos del siglo XVI. La Corte española y la escuela de Lovaina. (Spanish) [Strumentaria: 16th Century scientific instruments. The Spanish Court and Louvain School]. Fundación Carlos de Amberes Madrid 26 de noviembre de 1997 a 1 de febrero de 1998, Madrid, Fundacin Carlos de Amberes, 1997, 248 pp., ill (ISBN 84-87369-03-0). Nuncius, 13(2):761, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00879.

Miniati:1998:BRGb

[1861] Mara Miniati. Book review: Gigliola Terenna, Francesca Vannozzi (a cura di), Strumenti medici dei secoli XVIII–XIX, Siena, Nuova immagine editrice, 1997, 78 pp., 23 ill.n.n., (ISBN 88-7145-141-4) (Patrimonio storico-scientifico dell’Università degli Studi di Siena Materiali 1). Nuncius, 13(2):762, ???? 1998. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539198x00888.

Ciardi:1998:BRAa

[1862] Marco Ciardi. Book review: Astronomia Gudrun Wolfschmidt, Milchstraße, Nebel, Galaxien. Strukturen im Kosmos von Herschel bis Hubble, München, Deutsches Museum, 1995, 186 pp. ill. (ISBN 3-486-26308-0). Nuncius, 13(2):762–763, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00897.

Miniati:1998:BRGc

[1863] Mara Miniati. Book review: Giorgia Foder Serio, Ileana Chinnici, l’osseruatorio astronomico di Palermo, la storia e gli strumenti, Palermo, Flaccovio Editore, 1997, 175 pp., ill., (ISBN 88-7804-144-0). Nuncius, 13 (2):763–764, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00905.

Lippi:1998:BRMb

[1864] Donatella Lippi. Book review: Medicina L. R. Angeletti — V. Gazzaniga, Storia, filosofia ed etica generale della medicina, Milano, Masson, 1998, 177 pp. Nuncius, 13(2):764–765, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00914.

Lippi:1998:BRG

[1865] Donatella Lippi. Book review: G. Cosmacini, L’arte lunga. Storia della medicina dall’Antichità ad oggi, (Italian) [The long art. History of medicine from Antiquity to the present day], Bari, Laterza, 1997, 418 pp. Nuncius, 13(2):765–766, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00923.

Lippi:1998:BRMc

[1866] Donatella Lippi. Book review: M. D. Grmek (a cura di), Storia del pensiero medico occidentale. 3 vol. Dall’et romantica alla medicina moderna. Bari, Laterza, 1998, 648 pp. Nuncius, 13(2):766–767, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00932.

Lippi:1998:BRB

[1867] Donatella Lippi. Book review: B. M. Assael — G. Avanzini, Il male dell’anima. l’epilessia fra ’800 e ’900. Bari, Laterza, 1997, 190 pp. Nuncius, 13(2):767–768, ???? 1998. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539198x00941.

Toscano:1998:BRA

[1868] Anna Toscano. Book review: Alchimia e Chimica Johannes Fabricius, Alchimia. l’Arte Regia nel simbolismo medioevale, Roma, Edizioni Méditerranée, 1997, 248 p.+ill., Lit. 90.000 (ISBN 88-272-0482-2). Nuncius, 13(2):768–769, ???? 1998. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539198x00950.

Ciardi:1998:BRAb

[1869] Marco Ciardi. Book review: Alberto Gigli Berzolari, Luigi Valentino Brugnatelli /Diario del viaggio in Svizzera e in Francia con Alessandro Volta, Bologna, Cisalpino-Istituto Editoriale Universitario, 301 pp., ill. (Fonti e studi per la storia dell’Università di Pavia, 28) (ISBN 88-205-0798-6). Nuncius, 13(2):769–770, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x00969.

Vista:1998:BRD

[1870] Michele Vista. Book review: Dmitrij Ivanovi Mendeleev, Il sistema periodico degli elementi, a cura di Silvano Tagliagambe, Roma, Edizioni Teknos, 1994, liv + 182 pp. Nuncius, 13(2):770–771, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00978.

Miniati:1998:BRSc

[1871] Mara Miniati. Book review: Storia Del Territorio Agnese Visconti (a cura di), Il territorio lombardo: prospettive di ricerca storico-naturalistica dal medioevo all’et contemporanea, Milano, Societ Italiana di Scienze Naturali, 1998, 135 pp., ill., (ISBN 0369-6243) (Natura. Rivista di Scienze Naturali Milano Dicembre 1996-vol. 87-fasc. 2). Nuncius, 13(2):771–772, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x00987.

Miniati:1998:BRSd

[1872] Mara Miniati. Book review: Secoli XVI–XX Silke Ackermann (ed.), Humphrey Cole: mint, measurement and maps in Elizabethan England. London, the Trustees of the British Museum, 1998, 106 pp., ill. (ISBN 0-86159-126-7) (British Museum, Occasional Paper 126). Nuncius, 13 (2):772–773, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x00996.

Miniati:1998:BRI

[1873] Mara Miniati. Book review: Isaia Iannaccone, Johann Schreck Terrentius. Le scienze rinascimentali e lo spirito dell’Accademia dei Lincei nella Cina dei Ming, Napoli, Istituto Universitario Orientale, 1998, 147 pp., ill. (Series Minor, Liv). Nuncius, 13(2):773–774, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x01003.

Miniati:1998:BRD

[1874] Mara Miniati. Book review: Il diavolo e il diavoletto. Raffaello Magiotti, uno scienziato di Montevarchialla corte di Galileo, (Italian) [The devil and the demon. Raphael Magiotti, a scientist of Montevarchi at the Court of Galileo] Arezzo, La Piramide s.r.l., 1997, 189 pp., ill. Nuncius, 13(2): 774–775, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x01012.

LaVergata:1998:BRF

[1875] Antonello La Vergata. Book review: Françoise Charles-Daubert, Les libertins rudits en France au XVIIe siècle, Paris, Presses Universitaire de France (Philosophies, 106), 1998, 132 pp. FF 48 (ISBN 2-13-049345-9). Nuncius, 13(2):775–776, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x01021.

Miniati:1998:BRR

[1876] Mara Miniati. Book review: Raffaella Simili (a cura di), Ricerca e istituzioni scientifiche in Italia, Roma, Bari, Laterza, 1998, 244 pp. (ISBN 88-420-5537-9), (Biblioteca di Cultura Moderna, 1142). Nuncius, 13(2): 776, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x01030.

Miniati:1998:BRA

[1877] Mara Miniati. Book review: Anita Mcconnell, The life and times of J. J. Hicks (1837–1916) ‘King of the Clinicals’, York, England, William Sessions Limited, 1998, 150 pp., ill. (ISBN 1-85072-204-8). Nuncius, 13(2):777, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x01049.

LaVergata:1998:BRD

[1878] Antonello La Vergata. Book review: Daniel Becquemont, Laurent Mucchielli, Le cas Spencer. Religion, science et politique, Paris, Presses Universitaires de France, 1998, x + 358 pp., (ISBN 2-13-049107-3), FF 170. Nuncius, 13(2):777–778, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539198x01058.

Ciardi:1998:BRP

[1879] Marco Ciardi. Book review: Pierre Marage, Grgoire Wallenborn (diteurs), Les Conseils Solvay et les débuts de la physique moderne, (French) [The Solvay Congresses and the beginnings of modern physics], Bruxelles, Universit Libre, 265 pp., ill. Nuncius, 13(2):778–779, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x01067.

Damerow:1998:ARG

[1880] Peter Damerow and Jürgen Renn. Attivit di ricerca Galileo at work: His complete notes on motion in an electronic representation. Nuncius, 13 (2):781–790, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539198x01076.

Ciardi:1998:IAX

[1881] M. Ciardi. Indici (anno XIII, 1998). Nuncius, 13(2):791–795, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x01085.

Anonymous:1998:IDF

[1882] Anonymous. Indice delle fonti manoscritte. Nuncius, 13(2):797–801, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x01094.

Anonymous:1998:IN

[1883] Anonymous. Indice dei nomi. Nuncius, 13(2):803–832, ???? 1998. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539198x01102.

Baroncini:1999:FLF

[1884] Gabriele Baroncini. Forme di lettura, forme di cultura. Nuncius, 14 (1):3–18, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00733.

Strano:1999:AS

[1885] Giorgio Strano. The absent star. Nuncius, 14(1):19–29, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00742.

Zik:1999:GT

[1886] Yaakov Zik. Galileo and the telescope. Nuncius, 14(1):31–67, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00751.

Bensa:1999:FDM

[1887] Elisa Bensa and Gianni Zanarini. La fisica della musica. Nuncius, 14 (1):69–111, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00760.

Nicolini:1999:PPC

[1888] Nicoletta Nicolini. Petits processus chimiques dans l’Italie pre-unitaire. Nuncius, 14(1):113–120, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00779.

Rossi:1999:GFT

[1889] Arcangelo Rossi. Galileo Ferraris (1847–1897) tra scienza e tecnologia. Nuncius, 14(1):121–132, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00788.

Pogliano:1999:BPG

[1890] Claudio Pogliano. Bachi, polli e grani. Appunti sulla ricezione della genetica in Italia (1900–1953). Nuncius, 14(1):133–168, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00797.

Monti:1999:CDC

[1891] Dalida Monti. Creation and destruction of corpuscles. Nuncius, 14(1): 169–200, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00805.

Bnoli:1999:S

[1892] Fabrizio Bnoli and Marina Zuccoli. Strumentaria. Nuncius, 14(1):201– 212, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00814.

Lualdi:1999:BBO

[1893] Alberto Lualdi. Biagio Burlini, un ottico del ’700 Veneziano. Nuncius, 14(1):213–220, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00823.

Abbri:1999:PAD

[1894] Ferdinando Abbri. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 14(1):221–231, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00832.

Meschiari:1999:CGB

[1895] Alberto Meschiari. Corrispondenza di Giovanni Battista Amici con Carlo Matteucci e Angelo Secchi. Nuncius, 14(1):233–261, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00841.

Palmerino:1999:PGP

[1896] Carla Rita Palmerino. Pierre Gassendi’s De Philosophia Epicuri Universe rediscovered. Nuncius, 14(1):263–294, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x00850.

Monti:1999:PHS

[1897] Maria Teresa Monti. I poulets di Haller e la storia epistemologica della scienza. (Italian) [Haller’s chickens and the history of epistemological science]. Nuncius, 14(1):295–304, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00869.

Beretta:1999:BRI

[1898] Marco Beretta. Book review: Isisi cumulative Bibliography 1986–95. A bibliography of the History of Science formed from the Annual Isis Current Bibliographies. Edited by John Neu, Published for the History of Science Society, Canton Mass., Science History Publications/USA, 1997, 4 voll. 935 pp. + 1362 pp. (ISBN 0-88135-133-4). Nuncius, 14 (1):305–306, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00878.

Polizzi:1999:BRIa

[1899] Gaspare Polizzi. Book review: Isabelle Stengers, Scienze e poteri. Bisogna averne paura?, trad. di Ada Cinato, Bollati Boringhieri, Torino, 1998, 95 pp., L. 24.000 (ISBN 88-339-1091-1) (Temi 75). Nuncius, 14 (1):306–309, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00887.

Tugnoli:1999:BRL

[1900] Claudio Tugnoli. Book review: Luigi Ruggiu (a cura di), Filosofia del tempo, Milano, Bruno Mondadori, 1998, pp. 383, L. 42.000 (ISBN 88-424-9341-4). Nuncius, 14(1):309–314, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x00896.

LaVergata:1999:BRB

[1901] Antonello La Vergata. Book review: E. Bellone, G. Boniolo (a cura di), Storia e filosofia della scienza. Un possibile scenario italiano, Milano, Fotolito Sebi srl, 1998, 190 pp., prezzo non indicato (ISBN mancante). Nuncius, 14(1):314–315, ???? 1999. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539199x00904.

Barone:1999:BRC

[1902] Vincenzo Barone. Book review: Carlo Castagnoli (a cura di), Quaderni di Storia della Fisica, n. 1, 1997, Bologna, Editrice Compositori, 357 pp. Nuncius, 14(1):315–320, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00913.

Barsanti:1999:BRM

[1903] Giulio Barsanti. Book review: Mario Zunino e Maria Stella Colomba, Ordinando la natura. Elementi di storia del pensiero sistematico in biologia, Palermo, Medical Books, 1997, 160 pp., L. 36.000. Nuncius, 14 (1):320–321, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00922.

Nani:1999:BRA

[1904] Michele Nani. Book review: Alberto Capatti, Alberto De Bernardi, Angelo Varni (a cura di), l’alimentazione (Storia d’Italia — Annali 13), Torino, Einaudi, 1998, lxiv + 1052 pp., L. 130.000. Nuncius, 14(1):322– 325, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00931.

Wilson:1999:BRE

[1905] Renate Wilson. Book review: Evemarie Wolf, Über die Anfänge der Pharmaziegeschichtsschreibung. Von Johannes Ruellius (1529) bis David Peter Hermann Schmidt (1835), Stuttgart, Deutscher Apother Verlag, 1996, pp. x-109, Kartoniert. (Quellen und Studien zur Geschichte der Pharmazie, Band 72), DM/SFR 28.-/S 204 (ISBN 3-8047-1447-1) (Wvg). Nuncius, 14(1):325–327, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00940.

Tugnoli:1999:BRG

[1906] Claudio Tugnoli. Book review: Giorgio Mangani, Il mondo di Abramo Ortelio. Misticismo, geografia e collezionismo nel Rinascimento dei Paesi Bassi, Modena, Cosimo Panini editore, 1998, 333 pp., L. 60.000 (ISBN 88-7686-977-8). Nuncius, 14(1):328–332, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00959.

Shea:1999:BRF

[1907] William R. Shea. Book review: Francis Bacon, A Critical Edition of the Major Works, edited and annotated by Brian Vickers, Oxford University Press, l + 811 pp., L. 13.99 (ISBN 0-19-282025-7). Nuncius, 14 (1):332–333, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00968.

Dini:1999:BRB

[1908] Alessandro Dini. Book review: L. Bianchi (a cura di), Pierre Bayle e l’Italia, Napoli, Liguori, 1996, 249 pp., L. 25.000 (ISBN 207-2612-2). Nuncius, 14(1):333–338, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00977.

Contardi:1999:BRA

[1909] Simone Contardi. Book review: Annamaria Angelini (a cura di), Anatomie Accademiche. III. l’Istituto delle Scienze e l’Accademia, Bologna, Il Mulino, 1993, 573 pp., L. 60.000 (ISBN 88-15-04157-5). Nuncius, 14(1):338–340, ???? 1999. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539199x00986.

Abbri:1999:BRRa

[1910] Ferdinando Abbri. Book review: Robert E. Schofield, the Enlightenment of Joseph Priestley. A Study of His Life and Work from 1733 to 1773, University Park, Pennsylvania, the Pennsylvania State University Press, 1997, xii + 305 pp., ill., £40.50 (ISBN 0-271-01662-0). Nuncius, 14 (1):340–342, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00995.

Beretta:1999:BRJ

[1911] Marco Beretta. Book review: Jacques Tenon, Memoirs on Paris Hospitals, Edited with an Introduction, Notes and Appendices by Dora B. Weiner, Canton Mass., Science History Publications/USA, 1997, xxix + 407 pp., Ill. (ISBN 0-88135-074-5 $39.95). Nuncius, 14(1):342–343, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01002.

Segala:1999:BRK

[1912] Marco Segala. Book review: Kai Torsten Kanz, Nationalismus und internationale Zusammenarbeit in den Naturwissenschaften. Die deutsch–französischen Wissenschaftsbeziehungen zwischen Revolution und Restauration, 1789–1832, (German) [Nationalism and international cooperation in the natural sciences. The German–French scientific relations between Revolution and Restoration, 1789–1832]. Stuttgart, Steiner Verlag, 1997, 352 pp. (ISBN 3-515-07079-6) (Boethius. Texte und Abhandlungen zur Geschichte der Mathematik und der Naturwissenschaften, Bd. 39). Nuncius, 14(1):344–346, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x01011.

Marchis:1999:BRK

[1913] Vittorio Marchis. Book review: K. G. Beauchamp, Exhibiting Electricity, London, the Institution of Electrical Engineers, 1997. Nuncius, 14(1): 346–349, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01020.

diGregorio:1999:BRA

[1914] Mario A. di Gregorio. Book review: Adrian Desmond, Huxley. Evolution’s Hight Priest, London, Michael Joseph, 1997 (ISBN 0-7181-3882-1). £20. Nuncius, 14(1):349–350, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01039.

LaVergata:1999:BRAa

[1915] Antonello La Vergata. Book review: Alan P. Barr (ed.), The major prose of Thomas Henry Huxley, Athens, Georgia, The University of Georgia Press, 1997, xxn+366 pp., illustrazioni in bianco e nero (ISBN 0-8203-1864-7), $50,00 (rilegato); Alan P. Barr (ed.), Thomas Henry Huxley’s place in science and letters, Athens, Georgia, The University of Georgia Press, 1997, pp. xiii + 353, illustrazioni in bianco e nero (ISBN 0-8203-1865-5), $50,00 (rilegato). Nuncius, 14(1):350–354, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01048.

diGregorio:1999:BRG

[1916] Mario A. di Gregorio. Book review: Gerhard H. Müller, Friedrich Ratzel (1844–1904): Naturwissenschaftler, Geograph, Gelehrter, Stuttgart, Verlag für Geschichte der Naturwissenschaften und Technik, 1996. Nuncius, 14(1):354–356, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01057.

Drscher:1999:BRC

[1917] Ariane Drscher. Book review: Carmela Morabito, La cartografia del cervello. Il problema delle localizzazioni cerebrali nell’opera di David Ferrier, fra fisiologia, psicologia e filosofia, Milano, Franco Angeli, 1996, 234 pp. (ISBN 88-204-9708-5) L. 45.000. Carmela Morabito, Modelli della mente, modelli del cervello, Milano, Franco Angeli, 1998, 303 pp. (ISBN 88-464-0419-X) L. 54.000. Nuncius, 14(1):356–358, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01066.

Keil:1999:BRA

[1918] Gundolf Keil. Book review: Angelika Schaller, Und seine Begierde ward sehend. Auge, Blick und visuelle Wahrnehmung in der Prosa Thomas Manns, (German) [And his desire for sight. Eye gaze and visual perception in the prose of Thomas Mann], Würzburg. Ergon-Verlag, 1997, 407 pp., 4 Abb., DM 68,- — (Litteratura. Wissenschaftliche Beiträge zur Moderne und ihrer Geschichte, 5). Nuncius, 14(1):358–361, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01075.

Miniati:1999:BRF

[1919] Mara Miniati. Book review: Fabrizio Bnoli (a cura di), l’astronomia in Italia, Napoli, Societ Astronomica Italiana, 1998, 156 pp., ill. (ISBN 88-87375-00-3). Nuncius, 14(1):361–362, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x01084.

VanDerMeer:1999:BRM

[1920] Jitse M. Van Der Meer. Book review: Michael Heller, the New Physics and a New Theology, Tr. G. V. Coyne, S. J., S. Giovanninni, T. M. Sierotowicz. Vatican Observatory Publications, 1996, 103 pp. Nuncius, 14(1):362–365, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01093.

Andreozzi:1999:BRD

[1921] Luciano Andreozzi. Book review: Daniel Dennet, l’idea pericolosa di Darwin, Torino, Bollati Boringhieri, 1997, 723 pp. (ISBN 88-339-1055-5). Nuncius, 14(1):365–369, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01101.

Polizzi:1999:BRF

[1922] Gaspare Polizzi. Book review: François Jacob, Il topo, la mosca e l’uomo, trad. di Fiorella Nuzzo De Carli, Bollati Boringhieri, Torino, 1998, 147 pp., L. 35.000 (ISBN 88-339-1089-X), (Saggi scientifici). Nuncius, 14 (1):369–373, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01110.

Miniati:1999:BRL

[1923] Mara Miniati. Book review: Luca Basso Peressut (a cura di), Stanze della meraviglia. I musei della natura tra storia e progetto, Bologna, Clueb, 1997, 363 pp., ill. (ISBN 88-8091-519-3), L. 43.000 (Lexis, MuseoPoli, IV). Nuncius, 14(1):375–377, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01129.

Miniati:1999:C

[1924] Mara Miniati. Cataloghi. Nuncius, 14(1):377, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x01138.

Lippi:1999:BRG

[1925] Donatella Lippi. Book review: G. Corbellini e L. Mezzagora (a cura di), La malaria tra passato e presente, Roma, 1998, 109 pp. Nuncius, 14(1):378, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01147.

Miniati:1999:BRV

[1926] Mara Miniati. Book review: Vladimiro Valerio, Piante e vedute di Napoli dal 1486 al 1599. l’origine dell’iconografia urbana europea, con il contributo di Ermanno Bellucci, Napoli, Electa, 1998, 95 pp., ill., (ISBN 88-435-8725-0), (Imago Urbis, 1). Nuncius, 14(1):378–379, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01156.

Lippi:1999:BRRa

[1927] Donatella Lippi. Book review: S. Rossi (a cura di), Scienza e miracoli nell’arte del ’600 Milano, Electa, 1998, 406 pp. Nuncius, 14(1):379–380, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01165.

Ciardi:1999:BRR

[1928] Marco Ciardi. Book review: Renato G. Mazzolini (a cura di), Le collezioni scientifiche del Ginnasio Liceo ‘Giovanni Prati’ di Trento, (Italian) [The scientific collections of the ‘Giovanni Prati’ Gymnasium (High School) of Trento], Trento, Ginnasio Liceo Giovanni Prati, 1997, 358 pp., ill. (ISBN mancante), s.i.p. Nuncius, 14(1):380–381, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01174.

Miniati:1999:Sa

[1929] Mara Miniati. Strumentaria. Nuncius, 14(1):381–382, ???? 1999. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x01183.

Miniati:1999:BRE

[1930] Mara Miniati. Book review: Ta epistémonika organa, 19os aiónas hoi Hellénikes sylloges 1997, 48 pp., ill. (ISBN 960-7094-99-9). Nuncius, 14 (1):382–383, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01192.

Miniati:1999:BRGa

[1931] Mara Miniati. Book review: Gigliola Terenna, Francesca Vannozzi, La collezione degli strumenti di psicologia, Siena, Nuova immagine editrice, 1998, 78 pp., ill. (ISBN 88-7145-146-5), (Patrimonio storicoscientifico dell’Università di Siena Materiali 2). Nuncius, 14(1):383, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01200.

Lippi:1999:M

[1932] Donatella Lippi. Medicina. Nuncius, 14(1):384, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x01219.

Lippi:1999:BRP

[1933] Donatella Lippi. Book review: Paul Lydon, A catalogue of records retained by hospices and related organisations in the UK and the Republic of Ireland, Sheffield, 1998, 124 pp. (Eahmh, Research Guide Series 3). Nuncius, 14(1):384–385, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01228.

Lippi:1999:BRJ

[1934] Donatella Lippi. Book review: Jan Sundin, Marie Clark Nelson and John Woodward (eds.), The history and philosophy of Medicine and Health: past, present, future, Sheffield, 1998, 30 pp. (Eahmh-Research Guide Series 4). Nuncius, 14(1):385, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01237.

Lippi:1999:BRRb

[1935] Donatella Lippi. Book review: Robert Juette, Guenter B. Risse and John Woodward (eds.), Culture, Knowledge, Healing. Historical perspectives of Homeopathic Medicine in Europe and North America, Sheffield, 1998, 338 pp. (Eahmh Network Series 3). Nuncius, 14(1):385–386, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01246.

Lippi:1999:BRC

[1936] Donatella Lippi. Book review: Cay-Ruediger Pruell (ed.), Pathology in the 19th and 20th centuries. The relationship between theory and practice, Sheffield, 1998, 142 pp. (Eahmh Network Series 2). Nuncius, 14(1):386–387, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01255.

Miniati:1999:F

[1937] Mara Miniati. Fisica. Nuncius, 14(1):387–388, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x01264.

Abbri:1999:A

[1938] Ferdinando Abbri. Alchimia. Nuncius, 14(1):388–389, ???? 1999. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x01273.

Miniati:1999:G

[1939] Mara Miniati. Geologia. Nuncius, 14(1):389–390, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x01282.

Caroti:1999:M

[1940] Sstefano Caroti. Medioevo. Nuncius, 14(1):390–391, ???? 1999. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x01291.

Caroti:1999:BRS

[1941] Stefano Caroti. Book review: Stefan Kirschner, Nicolaus Oresmes Kommentar zur Physik des Aristoteles. Kommentar mit Edition der Quaestionen zu Buch 3 und 4 der aristotelischen Physik sowie von vier Quaestionen zu Buch 5, Stuttgart, Franz Steiner Verlag, 1997, 491 pp. (Sudhoffs Archiv. Zeitschrift fr Wissenschaftsgeschichte, Beihefte 39). Nuncius, 14 (1):391–392, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01309.

Ciardi:1999:SXX

[1942] Marco Ciardi. Secoli XVII–XX. Nuncius, 14(1):392–393, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01318.

LaVergata:1999:BRPa

[1943] Antonello La Vergata. Book review: Pina Totaro (a cura di), Spinoziana. Ricerche di terminologia filosofica e critica testuale, Seminario internazionale, Roma, 29–30 settembre 1995, Firenze, L. S. Olschki, 1997, 278 pp. Nuncius, 14(1):393–394, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01327.

LaVergata:1999:BRPb

[1944] Antonello La Vergata. Book review: Pascal Acot (ed.), The European origins of scientific ecology (1800–1901), introduction by Patrick Blandin, translated from the French by B. P. Hamm, Amsterdam, Overseas Publishers Association (published under license of Gordon and Breach Publishers and Éditions des Archives Contemporaines), 2 voll., xix + 932 pp., (ISBN 90-5699-103-5), (The History of Science, Technology and Medicine, 8), ff. 884, $150,00, £95,00. Nuncius, 14(1):394–395, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01336.

Abbri:1999:BRJ

[1945] Ferdinando Abbri. Book review: Jonathan Gathorne-Hardy, Alfred C. Kinsey. Sex the Measure of All Things. A Biography, London, Chatto Windus, 1998, xiii + 514 pp. (ISBN 1-85619-604-6). Nuncius, 14(1): 395–396, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01345.

Anonymous:1999:BDSa

[1946] Anonymous. Biblioteca della scienza italiana. Nuncius, 14(1):397–400, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01354.

Anonymous:1999:BRMa

[1947] Anonymous. Book review: 6. Marsilio Landriani Ricerche fisiche intorno alla salubrit dell’aria a cura di Marco Beretta 1995, 126 pp., Giunti, Firenze. Nuncius, 14(1):400–401, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01363.

Anonymous:1999:BRLa

[1948] Anonymous. Book review: 7. Ludovico Antonio Muratori della forza della fantasia umana a cura di Claudio Pogliano 1995, 172 pp., Giunti, Firenze. Nuncius, 14(1):401–402, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01372.

Anonymous:1999:BRAa

[1949] Anonymous. Book review: 8. Amedeo Avogadro Saggi e memorie sulla teoria atomica (1811–1838) a cura di Marco Ciardi 1995, 149 pp., Giunti, Firenze. Nuncius, 14(1):402, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01381.

Anonymous:1999:BRLb

[1950] Anonymous. Book review: 9. Leonardo Da Vinci Libro di Pittura a cura di Carlo Pedretti con la collaborazione di Carlo Vecce 1995, 2 voll., 543 pp., Giunti, Firenze. Nuncius, 14(1):402–403, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01390.

Anonymous:1999:BRCa

[1951] Anonymous. Book review: 10. Camillo Golgi Sulla fina anatomia degli organi centrali del sistema nervoso a cura di Alberto Oliverio 1995, 264 pp., Giunti, Firenze. Nuncius, 14(1):403, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01408.

Anonymous:1999:BRFa

[1952] Anonymous. Book review: 11. Federico Delpino Memorie di biologia vegetale a cura di Maurizio Alippi Cappelletti 1996, 427 pp., Giunti, Firenze. Nuncius, 14(1):403–404, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01417.

Anonymous:1999:BRFb

[1953] Anonymous. Book review: 12. Francesco Redi Esperienze intorno alla generazione degli insetti a cura di Walter Bernardi 1996, 233 pp., Giunti, Firenze. Nuncius, 14(1):404, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01426.

Anonymous:1999:BRFc

[1954] Anonymous. Book review: 13. Felice Fontana Ricerche filosofiche sopra la fisica animale a cura di Giulio Barsanti 1996, 134 pp., Giunti, Firenze. Nuncius, 14(1):404–405, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01435.

Anonymous:1999:BRGa

[1955] Anonymous. Book review: 14. Giorgio Vasari Il Giovane Raccolto fatto dal Cavre. Giorgio Vasari: di varii instrumenti per misurare con la vista a cura di Filippo Camerota 1996, 423 pp., Giunti, Firenze. Nuncius, 14(1):405, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01444.

Anonymous:1999:BRAb

[1956] Anonymous. Book review: 15. Alessandro Herzen Gli animali martiri, I loro protettori e la fisiologia a cura di Giovanni Landucci 1996, 173 pp., Giunti, Firenze. Nuncius, 14(1):406, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01453.

Anonymous:1999:BRAc

[1957] Anonymous. Book review: 16. Agostino Scilla La vana speculazione disingannata dal senso a cura di Paolo Rossi con la collaborazione di Marco Segala 1996, 142 pp., Giunti, Firenze. Nuncius, 14(1):406–407, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01462.

Anonymous:1999:BRLc

[1958] Anonymous. Book review: 17. Lorenzo Ghiberti I commentarii a cura di Lorenzo Bartoli 1998, 316 pp., Giunti, Firenze. Nuncius, 14(1):407, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01471.

Anonymous:1999:BRAd

[1959] Anonymous. Book review: 18. Antonio Cocchi Scritti scelti a cura di Simone Contardi 1998, 134 pp., Giunti, Firenze. Nuncius, 14(1):407–408, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01480.

Anonymous:1999:BRGb

[1960] Anonymous. Book review: 19. Giovanni Battisti Baliani De motu naturali gravium solidarum et liquidorum a cura di Giovann Baroncelli 1998, 289 pp., Giunti, Firenze. Nuncius, 14(1):408, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01499.

Anonymous:1999:BRVa

[1961] Anonymous. Book review: 20. Vittorio Angius Almerico da Schio a cura di Vittorio Marchis 1998, 132 pp., Giunti, Firenze. Nuncius, 14 (1):408–409, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x01507.

Anonymous:1999:BRAe

[1962] Anonymous. Book review: 21. Alessandro Benedetti Historia corporis bumani sive Anatomice a cura di Giovanna Ferrari 1998, 364 pp., Giunti, Firenze. Nuncius, 14(1):409, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01516.

Anonymous:1999:BRBa

[1963] Anonymous. Book review: 22. Battista Grassi Studi di uno zoologo sulla malaria a cura di Bernardino Fantini 1998, 462 pp., Giunti, Firenze. Nuncius, 14(1):410, ???? 1999. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539199x01525.

Anonymous:1999:BNa

[1964] Anonymous. Biblioteca di nuncius. Nuncius, 14(1):410–411, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01534.

Anonymous:1999:BRMb

[1965] Anonymous. Book review: 30. Marco Segala I fantasmi, il cervello, l’anima. Schopenhauer, l’occulto e la scienza 1998, 248 pp., Olschki, Firenze. Nuncius, 14(1):411–412, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x01543.

Anonymous:1999:BRGc

[1966] Anonymous. Book review: 31. Giambattista Aleottie gli ingegneri del Rinascimento, (Italian) [Giambattista Aleottie and engineers of the Renaissance], a cura di Alessandra Fiocca 1998, xxvi + 466 pp., Olschki, Firenze [ISBN 88-222-4660-8]. Nuncius, 14(1):412, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01552.

Anonymous:1999:BRJa

[1967] Anonymous. Book review: 32. Journals and History of Science edited by Marco Beretta, Claudio Pogliano and Pietro Redondi 1998, viii + 272 pp., Olschki, Firenze. Nuncius, 14(1):412–413, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x01561.

Ossola:1999:PCM

[1968] Carlo Ossola. Piazzette e caraffe: ’metafisica’ galileiana. Nuncius, 14 (2):423–441, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00012.

Ceolin:1999:GSS

[1969] Massimilla Baldo Ceolin. Galileo e la scienza sperimentale. Nuncius, 14 (2):443–454, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00021.

Camerota:1999:DTD

[1970] Filippo Camerota. Il disegno del territorio e la difesa dello stato. (Italian) [design of the territory and the defence of the state]. Nuncius, 14(2): 455–469, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00030.

Rasini:1999:AOD

[1971] Vallori Rasini. Alle origini della teoria dell’atto biologico di Viktor Von Weizsäcker. Nuncius, 14(2):471–489, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x00049.

Sebastin:1999:SLI

[1972] Amparo Sebastin and Maria J. Jimenez. Strumentaria learned institutions: Sources for unknown scientific instruments. Nuncius, 14(2):491– 504, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00058.

Guerrini:1999:PAD

[1973] Luigi Guerrini. Per un archivio della corrispondenza degli scienziati italiani matematica ed erudizione. Giovanni Alfonso Borelli e l’edizione fiorentina dei libri V, VI e VII delle coniche di Apollonio di Perga. Nuncius, 14(2):505–568, ???? 1999. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539199x00067.

Barbieri:1999:CAC

[1974] Francesco Barbieri and Marina Zuccoli. Un contributo al carteggio di Guido Grandi e Odoardo Corsini. Nuncius, 14(2):569–582, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00076.

Grifoni:1999:IFC

[1975] Tiziana Grifoni. Istituzioni e fonti I cataloghi botanici del giardino di Villa la loggia fra tradizione ed innovazione. Nuncius, 14(2):583–628, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00085.

Mastrobisi:1999:ANC

[1976] Giorgio Jules Mastrobisi. Aporie di una nuova concezione dell’etere nel manoscritto einsteiniano del 1920. (Italian) [History of a new concept in ether. The 1920 Einsteinian manuscript]. Nuncius, 14(2):629–649, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00094.

Einstein:1999:VSR

[1977] Albert Einstein. Vorlesungen Über Spezielle Relativitätstheorie. (German) [Lectures on the Theory of Special Relativity]. Nuncius, 14(2): 650–661, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00102.

Pogliano:1999:CIA

[1978] Claudio Pogliano. Il contributo italiano al razzismo del XX secolo. Nuncius, 14(2):663–669, ???? 1999. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539199x00111.

Lenzi:1999:NMB

[1979] Alessandra Lenzi. Nova media biblioteca e attivit bibliografica dell’istituto e museo di storia della scienza: Verso un archivio comune. Nuncius, 14(2):671–673, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00120.

Pogliano:1999:BRM

[1980] Claudio Pogliano. Book review: Mario Biagioli (ed.), The Science Studies Reader, (in consultation with Peter Galison, Donna J. Haraway, Everett Mendelsohn, Sharon Traweek), New York and London, Routledge, 1999, 590 pp. Nuncius, 14(2):675–677, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00139.

Beretta:1999:BRO

[1981] Marco Beretta. Book review: Olivier Bloch, Matire histoires, Paris, Vrin, 1997, 464 pp. (ISBN 0249-7980), 340 ff. Nuncius, 14(2):677–679, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00148.

Iannaccone:1999:BRJ

[1982] Isaia Iannaccone. Book review: J. Needham (C. Harbsmeier), Science and Civilisation in China, vol. 7, parte I, (Language and Logic), Cambridge University Press, 1988, 480 + xxiv, pp. Nuncius, 14(2):679–683, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00157.

diPasquale:1999:BRK

[1983] Giovanni di Pasquale. Book review: Keimpe Algra, Concepts of Space in Greek Thought, Leiden, Brill, 1995, 366 pp. Nuncius, 14(2):683–685, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00166.

Caroti:1999:BRC

[1984] Stefano Caroti. Book review: Claude Lafleur (ed.), l’Enseignement de la philosophie au XIIIe siècle. Autour du ‘Guide de l’ètudiant’ du ms. Ripoll 109. Actes du colloque international, Brepols, Turnhout, 1997 (Studia Artistarum. Études sur la Faculté des arts dans les Universités médiévales, 5), 719 pp. Nuncius, 14(2):685–690, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00175.

Caroti:1999:BRJ

[1985] Stefano Caroti. Book review: Joelle Ducos, La météorologie en français au Moyen Age (XIIIe–XIVe siècles), (French) [Meteorology in French in the Middle Ages (13th–14th Centuries)], Honoré Champion Éditeur, Paris, 1998 (Sciences, techniques et civilisations du Moyen Age à l’aube des Lumières, 2), 493, pp. Nuncius, 14(2):690–692, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00184.

Tomasi:1999:BRE

[1986] Lucia Tongiorgi Tomasi. Book review: Eileen Reeves, Painting the Heavens. Art and Science in the Age of Galileo, Princeton University Press, 1997, 310 pp., ill. Nuncius, 14(2):692–695, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00193.

Gorman:1999:BRM

[1987] Michael John Gorman. Book review: Maurice A. Finocchiaro, Galileo on the world systems: a new abridged translation and guide, Berkeley and Los Angeles, California, University of California Press; 1997, xi + 425 pp., £40.00, cloth; £15.95, paper (ISBN 0-520-20548-0 alkaline Paper; ISBN 0-520-20646-0 pbk.: alk. paper). Nuncius, 14(2):695–700, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00201.

Segala:1999:BRF

[1988] Marco Segala. Book review: Francesca Maria Crasta, La filosofia della natura di Emanuel Swedenborg, Milano, Angeli, 1999, 335 pp., ill., L. 42.000 (Collana di filosofia, 65) (ISBN 88-464-1385-7). Nuncius, 14(2): 700–702, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00210.

Vasconi:1999:BRH

[1989] Paola Vasconi. Book review: Heiner F. Klemme, Kants Philosophie des Subjekts (Systematische und entwicklungsgeschichtliche Untersuchungen zum Verhältnis von Selbstbewustsein und Selbsterkenntnis), (German) [Kant’s philosophy of the subject (Systematic and evolutionary studies on the relationship between self-esteem and self-knowledge)], Hamburg, Felix Meiner, 1996, 432 p. (Kant-Forschungen Bd. 7). Nuncius, 14(2):

702–704, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00229.

Trombetta:1999:BRA

[1990] Vincenzo Trombetta. Book review: Antonio Borrelli (a cura di), Domenico Cotugno. Documenti d’archivio 1766–1833. Napoli, Istituto Italiano per gli Studi Filosofici, La Citta del Sole, 1977, pp. 238, ill. 30, L. 48.000. Nuncius, 14(2):704–709, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00238.

Brooke:1999:BRD

[1991] John Hedley Brooke. Book review: D. Knight and H. Kragh (eds.), The Making of the Chemist. The Social History of Chemistry in Europe, 1789–1914, Cambridge: Cambridge University Press, 1998, 347 pp. (ISBN 0-521-58351-9), £50.00 (US$80.00). Nuncius, 14(2):709–712, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00247.

LaVergata:1999:BRAb

[1992] Antonello La Vergata. Book review: Ariane Droscher, Die Zellbiologie in Italien im 19. Jahrhundert, Halle (Saale), Deutsche Akademie der Naturforscher Leopoldina (Acta Historica Leopoldina, 26), 184, pp. 18 illustrazioni e 18 tabelle nel testo (ISBN 3-335-00500-7), prezzo non indicato. Nuncius, 14(2):712–713, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00256.

Staley:1999:BRD

[1993] Richard Staley. Book review: Denis Brian, Einstein: A Life, New York, John Wiley, 1996, 509 pp. (ISBN 0-471-11459-6) $30.00 USA, $36.95 Can, £18.99 UK. Nuncius, 14(2):713–715, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00265.

Polizzi:1999:BRM

[1994] Gaspare Polizzi. Book review: Mile Meyerson, La deduzione relativistica, introduzione e traduzione di Carlo Vinti, Pisa-Roma, Istituti Editoriali e Poligrafici Internazionali, 1998, (Biblioteca di Studi Kantiani) 403 pp. L. 60.000, (ISBN 88-8147-136-1). Nuncius, 14(2):715–718, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00274.

Fabriziani:1999:BRC

[1995] Anna Fabriziani. Book review: Carlo Vinti, Il soggetto qualunque. Gaston Bachelard fenomenologo della soggettivita epistemica, Napoli, Edizioni Scientifiche Italiane, 1997, 1010 pp. (ISBN 88-8114-538-3), £100.000. Nuncius, 14(2):718–723, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00283.

Bertini:1999:BRL

[1996] Silvia Bertini. Book review: Luciano Mecacci, Psicologia e psicoanalisi nella cultura italiana del Novecento, Roma/Bari, Laterza, 1998, 106 pp. Nuncius, 14(2):723–725, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00292.

Greenblatt:1999:BRJ

[1997] Samuel H. Greenblatt. Book review: Jack D. Pressman, Last Resort. Psychosurgery and the Limits of Medicine, Cambridge, Cambridge University Press, xv + 555 pp., ill. (ISBN 0-521-35371-8). Nuncius, 14(2): 725–727, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00300.

Polizzi:1999:BRIb

[1998] Gaspare Polizzi. Book review: Ilya Prigogine, La fine delle certezze. Il tempo, il caos e le leggi della natura, in collaborazione con Isabelle Stengers, trad. di Libero Sosio, Torino, Bollati Boringhieri, 1997, (Saggi scientifici) 189 pp., L. 50.000 (ISBN 88-339-1030-X). Nuncius, 14(2): 727–732, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00319.

Smith:1999:BRK

[1999] Robert W. Smith. Book review: Kevin Madders, A New Force at a New Frontier. Europe’s development in the space field in the light of its main actors, policies, law and activities from its beginnings up to the present (Cambridge University Press, Cambridge, 1997), 604 pp., ill. Nuncius, 14(2):732–733, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00328.

Miniati:1999:Sb

[2000] Mara Miniati. Strumentaria. Nuncius, 14(2):735, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x00337.

Miniati:1999:BRGb

[2001] Mara Miniati. Book review: Gabriele Vanin, Orologi solari nella terra del Palladio, Feltre, Libreria Pilotto Editrice, 1998, 119 pp., ill. Nuncius, 14(2):736, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00346.

Miniati:1999:BRAa

[2002] Mara Miniati. Book review: Antonio Campigotto, Massimo Tozzi Fontana, Nel mondo delle cose. Collezioni private di interesse tecnicoscientifico. un sondaggio nelle province di Bologna, Ferrara, Forl, Ravenna e Rimini, Bologna, Il Nove, 1997, 502 pp., ill., (Ricerche per i beni artistici culturali e naturali della regione Emilia Romagna 18) L. 60.000. Nuncius, 14(2):736–737, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00355.

Abbri:1999:BRM

[2003] Ferdinando Abbri. Book review: Michael J. Crowe, David D. Dyck, James R. Kevin, A Calendar of the Correspondence of Sir John Herschel, Cambridge, Cambridge Up., 1998, 828 pp. (ISBN 0-521-63149-1). Nuncius, 14(2):738–739, ???? 1999. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539199x00364.

LaVergata:1999:BRF

[2004] Antonello La Vergata. Book review: F. Burkhardt, D. M. Porter, J. Harvey, J. R. Topham (eds.), The Correspondence of Charles Darwin, vol. 10: 1862, Cambridge, Cambridge University Press, 1997, xl + 936 pp., 16 ill. fuori testo, (ISBN 0-521-59032-9) prezzo non indicato. Nuncius, 14(2):739–740, ???? 1999. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539199x00373.

Miniati:1999:BRAb

[2005] Mara Miniati. Book review: Astronomia e Fisica Pasquale Tucci (a cura di), Atti del XVIII Congresso nazionale di storia della fisica e dell’astronomia. Centro Volta, Villa Olmo, Como 15–16 maggio 1998, Milano, Università degli studi di Milano istituto di fisica generale applicata sez. di storia della fisica, 1999, 294 pp. Nuncius, 14(2):740–741, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00382.

Lippi:1999:BRMa

[2006] Donatella Lippi. Book review: Medicina F. Carrara — L. Sebregondi — U. Tramonti, Gli Istituti di beneficenza a Firenze. Storia e architettura, Firenze, Alinea, 1999, 191 pp. Nuncius, 14(2):741–742, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00391.

Ciardi:1999:BRG

[2007] Marco Ciardi. Book review: Geologia Musei Civici di Imola, La collezione Scarabelli. I. Geologia, a cura di Marco Pacciarelli e Gian Battista Vai, Bologna, Grafis Edizioni, 1995, 407 pp., ISBN e prezzo assenti. Nuncius, 14(2):742–743, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00409.

Carusi:1999:T

[2008] Paola Carusi. Tecnologia. Nuncius, 14(2):743–744, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539199x00418.

Miniati:1999:BRM

[2009] Mara Miniati. Book review: Mauro Agnoletti, Segherie e foreste nel Trentino dal Medioevo ai giorni nostri, (Italian) [Sawmills and forests in Trentino from the Middle Ages to the present day], Trento, Museo degli Usi e Costumi della Gente Trentina, 1998, 233 pp., ill. (ISBN 88-85352-12-X), L. 32.000. Mauro Agnoletti, Indagine sulla tecnologia degli attrezzi da taglio di suolo boschivo in Trentino, estratto da Sm Annali di San Michele, (Italian) [Survey on gear technology for cutting of forest soil in Trentino, extracted from Sm Annals of San Michele], n. 9–10/1996–97, 74 pp., ill. Nuncius, 14(2):744–745, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00427.

Ciardi:1999:BRA

[2010] Marco Ciardi. Book review: Angelo D’Orsi (a cura di), Quaderni di Storia dell’Università di Torino, Torino, Centro di Studio della Storia dell’Università di Torino, Il Segnalibro Editore. Anno I, 1996, n. 1, xvi + 390 pp., L. 45.000. Anno II–III, 1997–1998, n. 2, xii + 564, L. 80.000. Nuncius, 14(2):745–747, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00436.

Lippi:1999:BRMb

[2011] Donatella Lippi. Book review: I. Mazzini, La medicina dei Greci e dei Romani, 2 voll., Roma, Jouvence, 1997, 476 pp., L. 50.000. Nuncius, 14 (2):747–748, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00445.

Anonymous:1999:BRS

[2012] Anonymous. Book review: Silvia Meloni Trkulja, Lucia Tongiorgi Tomasi (a cura di), Bartolomeo Bimbi. un pittore di piante e animali alla corte dei Medici, Firenze, Edifir Edizioni Firenze, 1998, 223 pp., ill. (ISBN 88-7970-053-7) L. 160.000. Nuncius, 14(2):748–749, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00454.

LaVergata:1999:BRM

[2013] Antonello La Vergata. Book review: A. Moquin-Tandon, un naturaliste à Paris, 1834. Notes sur mon séjour à Paris. Mon premier voyage à Paris, en 1834, du 4 septembre au 23 octobre, édition présenté par Jean-Louis Fischer, index tablié par Isabelle Dussert-Carbone, Chilly-Mazarin, Sciences en Situation, 1999, xxxiv + 163 pp. (ISBN 2-908965-11-9), Ff. 129, 20 E. Nuncius, 14(2):749–750, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00463.

Ciardi:1999:BRL

[2014] Marco Ciardi. Book review: Lidia Maggioli (a cura di), Alessandro Serpieri, Scienziato riminese, Rimini, Raffaelli e Luise Editori, 1998, 158 pp. (ISBN 88-85050-93-X). Nuncius, 14(2):750–751, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00472.

LaVergata:1999:BRC

[2015] Antonello La Vergata. Book review: Charles Darwin, De la fécondation croisée des Orchidées par les insectes et des bons résultats de croisements, (French) [Cross-fertilization of orchids by insects and good results of crosses], Chilly-Mazarin, Éditions Sciences en Situation, 1999, 17, iii, 332 pp., 34 ill. (ISBN 2-908965-14-3), FF 149, E 23. Nuncius, 14(2): 751–752, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00481.

Abbri:1999:BRRb

[2016] Ferdinando Abbri. Book review: Rachel P. Maines, The Technology of Orgasm. Hysteria, the Vibrator and Women’s Sexual Satisfaction, Baltimore London, The Johns Hopkins University Press, 1999, xx + 182 pp. (Johns Hopkins Studies in the History of Technology n. 24) (ISBN 0-8018-5941-7). Nuncius, 14(2):752–753, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00490.

Anonymous:1999:BDSb

[2017] Anonymous. Biblioteca della scienza Italiana. Nuncius, 14(2):755–758, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00508.

Anonymous:1999:BRMc

[2018] Anonymous. Book review: 6. Marsilio Landriani Ricerche fisiche intorno alla salubrita dell’aria a cura di Marco Beretta 1995, 126 pp., Giunti, Firenze. Nuncius, 14(2):758–759, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00517.

Anonymous:1999:BRLd

[2019] Anonymous. Book review: 7. Ludovico Antonio Muratori della forza della fantasia umana a cura di Claudio Pogliano 1995, 172 pp., Giunti, Firenze. Nuncius, 14(2):759–760, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00526.

Anonymous:1999:BRAf

[2020] Anonymous. Book review: 8. Amedeo Avogadro Saggi e memorie sulla teoria atomica (1811–1838) a cura di Marco Ciardi 1995, 149 pp., Giunti, Firenze. Nuncius, 14(2):760, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00535.

Anonymous:1999:BRLe

[2021] Anonymous. Book review: 9. Leonardo Da Vinci Libro di Pittura a cura di Carlo Pedretti con la collaborazione di Carlo Vecce 1995, 2 voll., 543 pp., Giunti, Firenze. Nuncius, 14(2):760–761, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00544.

Anonymous:1999:BRCb

[2022] Anonymous. Book review: 10. Camillo Golgi Sulla fina anatomia degli organi centrali del sistema nervoso a cura di Alberto Oliverio 1995, 264 pp., Giunti, Firenze. Nuncius, 14(2):761, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00553.

Anonymous:1999:BRFd

[2023] Anonymous. Book review: 11. Federico Delpino Memorie di biologia vegetale a cura di Maurizio Alippi Cappelletti 1996, 427 pp., Giunti, Firenze. Nuncius, 14(2):761–762, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00562.

Anonymous:1999:BRFe

[2024] Anonymous. Book review: 12. Francesco Redi Esperienze intorno alla generazione degli insetti a cura di Walter Bernardi 1996, 233 pp., Giunti, Firenze. Nuncius, 14(2):762, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00571.

Anonymous:1999:BRFf

[2025] Anonymous. Book review: 13. Felice Fontana Ricerche filosofiche sopra la fisica animale a cura di Giulio Barsanti 1996, 134 pp., Giunti, Firenze. Nuncius, 14(2):762–763, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00580.

Anonymous:1999:BRGd

[2026] Anonymous. Book review: 14. Giorgio Vasari Il Giovane Raccolto fatto dal Cavre. Giorgio Vasari: di varii instrumenti per misurare con la vista a cura di Filippo Camerota 1996, 423 pp., Giunti, Firenze. Nuncius, 14(2):763, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00599.

Anonymous:1999:BRAg

[2027] Anonymous. Book review: 15. Alessandro Herzen Gli animali martiri, I loro protettori e la fisiologia a cura di Giovanni Landucci 1996, 173 pp., Giunti, Firenze. Nuncius, 14(2):764, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00607.

Anonymous:1999:BRAh

[2028] Anonymous. Book review: 16. Agostino Scilla La vana speculazione disingannata dal senso a cura di Paolo Rossi con la collaborazione di Marco Segala 1996, 142 pp., Giunti, Firenze. Nuncius, 14(2):764–765, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00616.

Anonymous:1999:BRLf

[2029] Anonymous. Book review: 17. Lorenzo Ghiberti I commentarii a cura di Lorenzo Bartoli 1998, 316 pp., Giunti, Firenze. Nuncius, 14(2):765, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00625.

Anonymous:1999:BRAi

[2030] Anonymous. Book review: 18. Antonio Cocchi Scritti scelti a cura di Simone Contardi 1998, 134 pp., Giunti, Firenze. Nuncius, 14(2):765–766, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00634.

Anonymous:1999:BRGe

[2031] Anonymous. Book review: 19. Giovanni Battisti Baliani, De motu naturali gravium solidarum et liquidorum, (Latin) [On the motion of natural heavy solids and liquids], a cura di Giovann Baroncelli 1998, 289 pp., Giunti, Firenze. Nuncius, 14(2):766, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00643.

Anonymous:1999:BRVb

[2032] Anonymous. Book review: 20. Vittorio Angius Almerico da Schio a cura di Vittorio Marchis 1998, 132 pp., Giunti, Firenze. Nuncius, 14 (2):766–767, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539199x00652.

Anonymous:1999:BRAj

[2033] Anonymous. Book review: 21. Alessandro Benedetti Historia corporis humani sive Anatomice a cura di Giovanna Ferrari 1998, 364 pp., Giunti, Firenze. Nuncius, 14(2):767, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00661.

Anonymous:1999:BRBb

[2034] Anonymous. Book review: 22. Battista Grassi, Studi di uno zoologo sulla malaria a cura di Bernardino Fantini 1998, 462 pp., Giunti, Firenze. Nuncius, 14(2):768, ???? 1999. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539199x00670.

Anonymous:1999:BNb

[2035] Anonymous. Biblioteca di nuncius. Nuncius, 14(2):768–769, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00689.

Anonymous:1999:BRMd

[2036] Anonymous. Book review: 30. Marco Segala I fantasmi, il cervello, l’anima. Schopenhauer, l’occulto e la scienza 1998, 248 pp., Olschki, Firenze. Nuncius, 14(2):769–770, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00698.

Anonymous:1999:BRGf

[2037] Anonymous. Book review: 31. Giambattista Aleotti e gli ingegneri del Rinascimento a cura di Alessandra Fiocca 1998, xxvi + 466 pp., Olschki, Firenze. Nuncius, 14(2):770, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539199x00706.

Anonymous:1999:BRJb

[2038] Anonymous. Book review: 32. Journals and History of Science edited by Marco Beretta, Claudio Pogliano and Pietro Redondi 1998, viii + 272 pp., Olschki, Firenze. Nuncius, 14(2):770–771, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00715.

Anonymous:1999:IAX

[2039] Anonymous. Indici (anno XIV, 1999). Nuncius, 14(2):773–796, ???? 1999. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539199x00724.

Anonymous:2000:N

[2040] Anonymous. Nuncius. Nuncius, 15(1):2, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539100x00434.

Giudice:2000:TMN

[2041] Franco Giudice. La tradizione del mezzo e la nuova teoria della luce di Leonhard Euler. Nuncius, 15(1):3–49, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539100x00443.

Linguerri:2000:SIP

[2042] Sandra Linguerri. La societ italiana per il Progresso delle scienze:1. Nuncius, 15(1):51–78, ???? 2000. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539100x00452.

Andreozzi:2000:VVO

[2043] Luciano Andreozzi. Vito Volterra organizzatore scientifico e la nascita della biologia matematica in Italia. Nuncius, 15(1):79–109, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00461.

Turco:2000:CJP

[2044] Francesca Turco and Luigi Cerruti. Charles J. Pedersen e le origini della chimica supramolecolare. (Italian) [Charles J. Pedersen and the origins of supramolecular chemistry]. Nuncius, 15(1):111–168, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00470.

Lualdi:2000:RCI

[2045] Alberto Lualdi. Repertorio dei costruttori italiani di strumenti scientifici. Nuncius, 15(1):169–234, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00489.

Scianna:2000:ISG

[2046] Nicolangelo Scianna. Indagine sui grandi globi a stampa di Vincenzo Coronelli. Nuncius, 15(1):235–257, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00498.

Meschiari:2000:CGB

[2047] Alberto Meschiari. Corrispondenza di Giovanni Battista Amici con Giovanni Plana. (Italian) [Letters from Giovanni Battista Amici to Giovanni Plana]. Nuncius, 15(1):259–323, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00506.

Iannaccone:2000:DRL

[2048] Isaia Iannaccone. Documents relatifs à l’histoire de l’astronomie chinoise et aux rapports scientifiques entre l’Europe et la Chine (xviie–xixe siécles) conservés à la Bibliothèque de l’Observatoire de Paris. Nuncius, 15(1): 325–364, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00515.

Scalva:2000:MAC

[2049] Giuse Scalva. Un medico alla corte di Carlo Emanuele III: Vitaliano Donati e il suo viaggio in Levante (1759–1762). Nuncius, 15(1):365–397, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00524.

Govoni:2000:BCT

[2050] Paola Govoni. Biography: a critical tool to bridge the history of science and the history of women in science. Report on a conference at Newnham College, Cambridge, 10–12 September 1999. Nuncius, 15(1):

399–409, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00533.

Beretta:2000:PLB

[2051] Marco Beretta. Panopticon Lavoisier: Base de donnés et histoire documentaire de la Révolution Chimique. (French) [Panopticon Lavoisier: Database and documentary history of the Chemical Revolution]. Nuncius, 15(1):411–425, ???? 2000. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539100x00542.

Pogliano:2000:BRJ

[2052] Claudio Pogliano. Book review: Jean Starobinski, Action et Raction. Vie et aventures d’un couple, Paris, Éditions du Seuil, 1999, 458 pp. Nuncius, 15(1):427–430, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00551.

Barsanti:2000:BRP

[2053] Giulio Barsanti. Book review: Pascal Duris, Gabriel Gohau, Histoire des Sciences de la vie, Paris, Nathan, 1997, 416 pp. Nuncius, 15(1): 430–432, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00560.

diPasquale:2000:BRW

[2054] Giovanni di Pasquale. Book review: E. W. Marsden, Greek and Roman Artillery. Technical Treatises, Oxford, Oxford University Press, 1990 (rist. ed. 1971), 277 pp., 14 tavv. ill. Nuncius, 15(1):432–435, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00579.

Ackermann:2000:BRK

[2055] Silke Ackermann. Book review: E. S. Kennedy, P. Kunitzsch, and R. P. Lorch, The Melon-Shaped Astrolabe in Arabic Astronomy, Boethius — Texte und Abhandlungen zur Geschichte der Mathematik und der Naturwissenschaften, ed. by Menso Folkerts, vol. 43, Stuttgart, Franz Steiner Verlag: 1999, 235 pp. (ISBN 3-515-07561-5). Nuncius, 15(1):435–436, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00588.

Abbri:2000:BRPa

[2056] Ferdinando Abbri. Book review: Penelope Gouk, Music, Science and Natural Magic in Seventeenth-Century England, New Haven and London, Yale University Press, 1999, xii + 308 pp. ill, (ISBN 0-300-07383-6). Nuncius, 15(1):436–438, ???? 2000. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539100x00597.

Stancati:2000:BRF

[2057] Claudia Stancati. Book review: Francesca Bonicalzi, A tempo e luogo. l’infanzia e l’inconscio in Descartes, Milano, Jaca Book, 1998, 203 pp. Nuncius, 15(1):439–442, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00605.

DelNero:2000:BRA

[2058] Valerio Del Nero. Book review: Alfonso Mirto, Lucas Holstenius e la corte medicea. Carteggio (1629–1660). Firenze, L. S. Olschki, 1999, 327 pp. Nuncius, 15(1):442–444, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00614.

Ciardi:2000:BRRa

[2059] Marco Ciardi. Book review: Robert Fox, August Nieto-Galan (eds.), Natural Dyestuffs and Industrial Culturein Europe, 1750– 1880, Canton Mass., Science History Publications, 1999, xxix + 354 pp., ill. (ISBN 0-88135-273-X). Nuncius, 15(1):444–446, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00623.

Panarari:2000:BRN

[2060] Massimiliano Panarari. Book review: Nicoletta Diasio, La science impure. Anthropologie et médecine en France, Grande-Bretagne, Italie, Pays-Bas, Paris, Presses Universitaires de France (PUF), 1999, 286 pp. (ISBN 2-13-050092-7). Nuncius, 15(1):446–449, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00632.

Hentschel:2000:BRF

[2061] Klaus Hentschel. Book review: Fritz Stern, Einstein’s German World, Princeton, Princeton University Press, 1999, ix + 335 pp. (ISBN 0-691-05939-X). Nuncius, 15(1):449–452, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00641.

Polizzi:2000:BRW

[2062] Gaspare Polizzi. Book review: Werner Heisenberg, Lo sfondo filosofico della fisica moderna, a cura di, Giuseppe Gembillo e Enrico Antonio Giannetto, Palermo, Sellerio, 1999, 210 pp., L. 35.000 (ISBN 88-389-1450-8) (Nuovo Prisma, Serie Filosofia della scienza, 19). Nuncius, 15 (1):452–455, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00650.

Bertini:2000:BRGa

[2063] Silvia Bertini. Book review: Giovanni Jervis (a cura di), Il secolo della psicoanalisi, Torino, Bollati Boringhieri, 1999, 248 pp. Nuncius, 15(1): 455–458, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00669.

Bertini:2000:BRD

[2064] Silvia Bertini. Book review: Donald K. Freedheim (a cura di), Storia della Psicoterapia. un secolo di cambiamenti, Roma, Edizioni Scientifiche Magi, 1998, 1122 pp. Nuncius, 15(1):458–459, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00678.

Abbri:2000:M

[2065] Ferdinando Abbri. Metodologia. Nuncius, 15(1):461–462, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00687.

Ciardi:2000:BRRb

[2066] Marco Ciardi. Book review: R. G. A. Dolby, Uncertain Knowledge. An image of science for a changing world, Cambridge, Cambridge University Press, 1996,365 pp. (ISBN 0-521-56004-7). Nuncius, 15(1):463, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00696.

Abbri:2000:BRPb

[2067] Ferdinando Abbri. Book review: Pnina G. Abir-Am (sous la direction de), La mise en mémoire de la science. (French) [The development of memory of science]. Pour une ethnographie historique des rites commoratifs, Amsterdam, éditions des archives contemporaines, 1998, xiv + 321 pp. paper (ISBN 90-5709-007-4). Nuncius, 15(1):464–465, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00704.

Miniati:2000:BRR

[2068] Mara Miniati. Book review: Rafael Sisto Edreira, O patrimonio históricoscientífico do Instituto Xelmírez I (Santiago de Compostela). Inventario e catalogación. Unha ollada ós gabinetes de ciencias do vello Instituto de Santiago. (Galician) [The historical-scientific Xelmírez I Institute (Santiago de Compostela) inventory and cataloging: a look at the cabinets of the old Science Institute of Santiago]. A Coruña, Deputación Provincial da Coruña, 1999, 191 pp., ill. (ISBN 84-89652-87-2). Nuncius, 15 (1):465–466, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00713.

Miniati:2000:BRAa

[2069] Mara Miniati. Book review: Aleksandar Kostic, Dejan Todorovic, Sense, mind measure. Collection of old scientific instruments of the Laboratory for Experimental Psychology, University of Belgrade, Belgrade, Museum of Science and Technology, 1997, 103 pp., ill. (ISBN 86-82977-02-8). Nuncius, 15(1):467–468, ???? 2000. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539100x00722.

Miniati:2000:BRAb

[2070] Mara Miniati. Book review: Astronomia Alberto Rossi, Guido Horn d’Arturo: astronomo e uomo di cultura, Bologna, Editrice Clueb, 1994, 85 pp. Nuncius, 15(1):468–469, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00731.

Lippi:2000:BRR

[2071] Donatella Lippi. Book review: Roy Porter, Disease, medicine and society in England, 1550–1860, Cambridge, Cambridge University Press, 1999, 78 pp. (I ed. Cambridge 1995). Nuncius, 15(1):469, ????

2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00740.

Lippi:2000:BRE

[2072] Donatella Lippi. Book review: Esther Diana (a cura di), Da ospizio a nosocomio. Storia della solidariet valdarnese, Firenze, 2000, 181 pp. Nuncius, 15(1):470, ???? 2000. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539100x00759.

Bertini:2000:BRR

[2073] Silvia Bertini. Book review: Roger Perron, Storia della psicoanalisi, Roma, Armando Editore, 1997, 112 pp. Nuncius, 15(1):470–472, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00768.

Bertini:2000:BRJ

[2074] Silvia Bertini. Book review: Jolande Jacobi, La psicologia di C. G. Jung, Torino, Bollati Boringhieri, 1997, 238 pp. Nuncius, 15(1):472, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00777.

Abbri:2000:BRJ

[2075] Ferdinando Abbri. Book review: James H. Capshew, Psychologists on the March. Science, Practice, and professional Identity in America, 1929–1969, Cambridge, Cambridge University Press, 1999, xii + 276 pp., paperback, £14.95 (ISBN 0-521-56585-5). Nuncius, 15(1):472–474, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00786.

Lippi:2000:BRP

[2076] Donatella Lippi. Book review: Paul T. Nicholson and Ian Shaw (edited by), Ancient Egyptian materials and technology, Cambridge, Cambridge University Press, 2000, 702 pp. Nuncius, 15(1):474–475, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00795.

Miniati:2000:BRMa

[2077] Mara Miniati. Book review: Maria Gioia Tavoni, Un intellettuale europeo e il suo universo. Vincenzo Coronelli (1650–1718). (Italian) [A European intellectual and his universe. Vincenzo Coronelli (1650–1718)]. Bologna, Costa Editore, 1999, 266 pp., ill. Nuncius, 15(1):475–476, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00803.

Miniati:2000:BRMb

[2078] Mara Miniati. Book review: Marco Bresadola, Giuliano Pancaldi (eds.), Luigi Galvani international workshop Proceedings. Bologna, 9 October 1998. Bologna, Università di Bologna Dipartimento di Filosofia Centro Internazionale per la Storia delle Universita e della Scienza, 1999, 215 pp., ill. (ISBN 88-900162-1-3). Nuncius, 15(1):477, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00812.

Ciardi:2000:BRJ

[2079] Marco Ciardi. Book review: Jack Morrell, Science, Culture and Politics in Britain, 1750–1870, Aldershot, Variorum, 1997, 350 pp. (ISBN 0-86078-633-1), £52.50. Nuncius, 15(1):477–478, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00821.

Ciardi:2000:BRF

[2080] Marco Ciardi. Book review: Frank A. J. L. James (ed.), The Correspondence of Michael Faraday. Volume 4, January 1849–October 1855, Letters 2146–3032, London, the Institution of Electrical Engineers, 1999, 1003 pp., 12 ill., s.i.p. (ISBN 0-86341-251-3). Nuncius, 15(1):478–487, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00830.

Palisca:2000:VGS

[2081] Claude V. Palisca. Vincenzo Galilei, scienziato sperimentale, mentore del figlio Galileo (Italian) [Vincenzo Galilei, experimental scientist, mentor of his son Galileo]. Nuncius, 15(2):497–514, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00010.

Regge:2000:GPR

[2082] Tullio Regge. Galileo e il principio di relativit. Nuncius, 15(2):515–518, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00029.

Sargolini:2000:CVG

[2083] Federica Sargolini. La critica di Vincenzo Galilei al misticismo numerico di Gioseffo Zarlino (Italian) [the critique of Vincenzo Galilei to the numerical mysticism of Gioseffo Zarlino]. Nuncius, 15(2):519–550, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00038.

Dupre:2000:MIT

[2084] Sven Dupre. Mathematical instruments and the theory of the concave spherical mirror: Galileo’s optics beyond art and science. Nuncius, 15 (2):551–588, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00047.

Cerbu:2000:DGA

[2085] Thomas Cerbu and Michel-Pierre Lerner. La disgrace de Galilée dans les alpes urbanae: Sur la fabrique du texte de Leone Allacci. (French) [The disgrace of Galileo in the Urbanae Alps: On the manufacture of text of Leone Allacci]. Nuncius, 15(2):589–610, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00056.

Camerota:2000:ASB

[2086] Filippo Camerota. Architecture and science in Baroque Rome. the mathematical ornaments of Villa Pamphilj. Nuncius, 15(2):611–638, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00065.

Knight:2000:GGV

[2087] David Knight. Genesis geology: a very English compromise. Nuncius, 15(2):639–664, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00074.

Zanieri:2000:GOL

[2088] Stefania Zanieri. Un gioco ottico di Ludovico Buti al Museo di Storia della Scienza di Firenze. (Italian) [An optical game of Ludovico Buti in the Museum of History of Science in Florence]. Nuncius, 15(2):665–670, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00083.

Chinnici:2000:CSI

[2089] Ileana Chinnici. 19th century spectroscopic instruments in Italian astronomical observatories. Nuncius, 15(2):671–679, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539100x00092.

Carbone:2000:PAD

[2090] Luciano Carbone, Romano Gatto, and Franco Palladino. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 15(2):681–719, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00100.

DeFinetti:2000:ALG

[2091] Fulvia De Finetti. Alcune lettere giovanili di B. De Finetti alla madre. (Italian) [Some letters of youth: B. De Finetti to his mother]. Nuncius, 15(2):721–740, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00119.

Ciardi:2000:CPA

[2092] Marco Ciardi. Concorsi e premi all’Accademia Delle Scienze di Torino: Programmi di ricerca e strategie di politica scientifica. Nuncius, 15(2): 741–749, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00128.

Borck:2000:BRC

[2093] Cornelius Borck. Book review: Christopher Lawrence, Steven Shapin (eds.), Sciences Incarnate: Historical Embodiments of Natural Knowledge, vii + 342 pp. Chicago, University of Chicago Press, 1998. Nuncius, 15(2):751–754, ???? 2000. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539100x00137.

Iannaccone:2000:BRR

[2094] Isaia Iannaccone. Book review: Richard Gotshalk, Divination, Order and the Zhouyi, University Press of America, 1999, x + 418 pp. Nuncius, 15(2):754–757, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00146.

Iannaccone:2000:BRS

[2095] Isaia Iannaccone. Book review: Shen Kangshen, John N. Crossley, Anthony W.-C. Lun, The Nine Chapters on the Mathematical Art. Companion and Commentary, Oxford University Press, Science Presse Beijing, 1999, ix + 596 pp., £110. Nuncius, 15(2):757–759, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00155.

Triarico:2000:BRJ

[2096] Carlo Triarico. Book review: John L. Heilbron, The Sun in the Church: Cathedrals as solar observatories, Cambridge, MA — London, Harvard University Press, 1999, 366 pp. Nuncius, 15(2):760–765, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00164.

Visconti:2000:BRJ

[2097] Agnese Visconti. Book review: José María López Piñero, María Luz López Terrada, La influencia española en la introducción en Europa de las plantas americanas (1493–1623), Valencia, Instituto de estudios documentales e historicos sobre la ciencia, Universitad de Valencia —- C.S.I.C., 1997, 214 pp. Nuncius, 15(2):766–768, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00173.

Miniati:2000:BRS

[2098] Mara Miniati. Book review: Silvio A. Bedini, Patrons, Artisans and Instruments of Science, 1600–1750, Aldershot, Ashgate Publishing Limited, 1999, xvi + 336 pp., ill. (ISBN 0-86078-781-8), Variorum Collected Studies Series. Nuncius, 15(2):768–770, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00182.

Roero:2000:BRA

[2099] Clara Silvia Roero. Book review: Antoni Malet, From Indivisibles to Infinitesimals. Studies on Seventeenth-Century Mathematizations of Infinitely Small Quantities, Universitat de Barcelona, Servei de Publicacions Bellaterra, 1996, 163 pp. Nuncius, 15(2):770–772, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00191.

Toscano:2000:BRD

[2100] Anna Toscano. Book review: Dino Carpanetto, Scienza e arte del guarire. Cultura, formazione universitaria e professioni mediche a Torino tra Sei e Settecento, Cuneo 1998, 318 pp., L. 40.000 (Deputazione Subalpina di Storia Patria. Miscellanea di Storia Italiana. Serie V. Studi e fonti per la storia dell’Università di Torino, VIII). Nuncius, 15(2):772– 775, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00209.

Galera:2000:BRJ

[2101] Andrs Galera. Book review: José María López Piñero, Victor Navarro, Mara Luz López Terrada, José Ramón Bertomeu, Antonio García Belmar, Felipe Jerez Moliner, La actividad científica valenciana de la Ilustracin, Valencia, Diputacio de Valencia, 1998, 2 vols., 254 pp. + 320 pp. Nuncius, 15(2):775–776, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00218.

Abbri:2000:BRL

[2102] Ferdinando Abbri. Book review: Lisbet Koerner, Linnaeus: Nature and Nation, Cambridge Mass. and London, Harvard University Press, 1999, 298 pp. (ISBN 0-674-09745-9). Nuncius, 15(2):776–778, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00227.

Ciancio:2000:BRM

[2103] Luca Ciancio. Book review: Martin J. S. Rudwick, Georges Cuvier, Fossil Bones, and Geological Catastrophes. New Translations Interpretations of the Primary Texts, Chicago and London, The University of Chicago Press, 1997, xvi + 301 pp., ill. Nuncius, 15(2):778–781, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00236.

Pogliano:2000:BRV

[2104] Claudio Pogliano. Book review: Valeria Babini, Luisa Lama, Una donna nuova. Il femminismo scientifico di Maria Montessori, Milano, Angeli, 2000, 320 pp. Nuncius, 15(2):781–784, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00245.

Butterfield:2000:BRH

[2105] Jeremy Butterfield. Book review: Helge Kragh, Quantum Generations: a History of Physics in the Twentieth Century, Princeton University Press, 1999, xiv + 494 pp. Price: $29.95 £18.95. (ISBN 0-691-01206-7). Nuncius, 15(2):784–787, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00254.

Cachro:2000:BRW

[2106] Jacek Cachro. Book review: W. Michael Dickson, Quantum Chance and Non-Locality: Probability and Non-Locality in the Interpretations of Quantum Mechanics. Cambridge, Cambridge University Press, 1998, xix + 244 pp. (ISBN 0-521-58127-3), £40. Nuncius, 15(2):787–791, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00263.

Polizzi:2000:BRM

[2107] Gaspare Polizzi. Book review: Mario Dal Pra (diretta da), Storia della Filosofia, Volume 11: La filosofia contemporanea: la seconda met del Novecento (2 tomi), II Edizione interamente riscritta e ampliata a cura di Gianni Paganini, Padova, Piccin Nuova Libraria, 1998, 1762 pp., (ISBN 88-299-1487-8), L. 220.000. Nuncius, 15(2):791–794, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00272.

Segala:2000:BRD

[2108] Marco Segala. Book review: Denis Guedj, Il teorema del pappagallo, Milano, Longanesi C., 2000, 562 pp., (ISBN 88-304-1758-0), Lit. 32.000. Nuncius, 15(2):795–796, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00281.

Miniati:2000:M

[2109] Mara Miniati. Museologia. Nuncius, 15(2):797, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539100x00290.

Miniati:2000:BRMc

[2110] Mara Miniati. Book review: Marinella Calisi, Storia e strumenti del Museo Astronomico e Copernicano di Roma. Guida alle collezioni, Roma, Osservatorio Astronomico di Roma, 2000, 167 pp., ill. Nuncius, 15(2): 798, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00308.

Miniati:2000:BRW

[2111] Mara Miniati. Book review: Wilma di Palma, Lamberto Lamberti, Le regole del regolo. Guida alla collezione capitolina di regoli calcolatori, Torino, Bollati Boringhieri, 2000, 135 pp., ill. (ISBN 88-339-5640-7). Nuncius, 15(2):798–799, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00317.

Miniati:2000:BRFa

[2112] Mara Miniati. Book review: Fausto Casi, Francesco Redi e gli strumenti scientifici del suo tempo, Arezzo, Accademia Petrarca di Lettere Arti e Scienze, 1999, 231–284; 391–407 pp., ill. (estr. da Francesco Redi Aretino). Nuncius, 15(2):799–800, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00326.

Miniati:2000:BRPa

[2113] Mara Miniati. Book review: Pino Calledda, Edoardo Proverbio (a cura di), Storia del Servizio Internazionale delle Latitudini e delle imprese di Cooperazione internazionale (1850–1950) Astronomia e Archeoastronomia. Atti del Convegno di Storia della’astronomia Cagliari 24–25 settembre 1999, Cagliari, Cooperativa Universitaria Editrice Cagliaritana C.U.E.C., 2000, 373 pp. (ISBN 88-87088-88-8). Nuncius, 15(2):800–802, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00335.

Miniati:2000:BRV

[2114] Mara Miniati. Book review: Víctor Navarro Brotóns, Vicente L. Salavert Fabiani, Victoria Rosselló Botey, Víctor Darás Román, Bibliographia physico-mathematica hispanica (1474–1900). Volumen I Libros y folletos, 1475–1600, Valencia, Instituto di Historia de la ciencia y Documentación López Piñero, Universitat de Valencia-C.S.I.C., 1999, 374 pp. (ISBN 84-370-4219-4) (Cuadernos Valencianos de Historia de la Medicina y de la Ciencia LVI — Serie C). Nuncius, 15(2): 802, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00344.

Miniati:2000:BRFb

[2115] Mara Miniati. Book review: Francesca Vannozzi (a cura di), Siena la città laboratorio. Dall’innesto del vajuolo ad Albert Sabin, Siena, Protagon Editori Toscani-Santa Maria della Scala, 1999, 108 pp., ill. Nuncius, 15(2):803–804, ???? 2000. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539100x00353.

Bertini:2000:BRGb

[2116] Silvia Bertini. Book review: Giovanni Pietro Lombardo, Renato Foschi (a cura di), La psicologia italiana e il Novecento. Le prospettive emergenti nella prima met del secolo, Milano, Franco Angeli, 1997, 337 pp. Nuncius, 15(2):804–805, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00362.

Miniati:2000:BRFc

[2117] Mara Miniati. Book review: Ferdinando Abbri e Marco Segala (a cura di), Segni e percorsi della modernit. Saggi in onore di Paolo Rossi. Con la bibliografia di Paolo Rossi, Arezzo, Università degli Studi di Siena-Dipartimento di Studi Storico-Sociali e Filosofici, 2000, 229 pp. Nuncius, 15(2):805–806, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539100x00371.

Miniati:2000:BRPb

[2118] Mara Miniati. Book review: Patrice Bret (d.), L’expédition d’Egypte, une entreprise des Lumières, 1798–1801. Actes du colloque international organisé par l’Académie des inscriptions et belles-lettres et l’Académie des sciences, sous les auspices de l’Institut de France et du Muséum national d’histoire naturelle, 8–10 juin, 1998, Paris, Éditions Tec Doc, 1999, 436 pp., ill. (ISBN 2-7430-0319-7). Nuncius, 15(2):806–807, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00380.

Abbri:2000:BRA

[2119] Ferdinando Abbri. Book review: Asa Bergenheim, Lena Lennerhed (red.), Seklernas sex. Bidrag till sexualitetens historia, Stockholm, Carlssons, 1997, 264 pp. (ISBN 91-7203-256-1). Nuncius, 15(2):807–808, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00399.

Abbri:2000:BRB

[2120] Ferdinando Abbri. Book review: Barbara Maria Stafford, Visual Analogy. Consciousness as the Art of Connecting, Cambridge Mass. — London, The MIT Press, 1999, viii + 219 pp., ill. (ISBN 0-262-19421-X). Nuncius, 15(2):808–809, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539100x00407.

Anonymous:2000:IB

[2121] Anonymous. The IMSS bookshelf. Nuncius, 15(2):811–817, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00416.

Anonymous:2000:IAX

[2122] Anonymous. Indici (anno XV, 2000). Nuncius, 15(2):821–886, ???? 2000. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539100x00425.

Fabbri:2001:KCA

[2123] Natacha Fabbri. Kepler: Il cosmo armonico e la musica. (Italian) [Kepler: cosmic harmony and music]. Nuncius, 16(1):3–48, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00019.

Nonnoi:2001:SFG

[2124] Giancarlo Nonnoi. La scienza e la filosofia Galileiane nel New World di John Wilkins. (Italian) [Science and Galilean philosophy in the New World by John Wilkins]. Nuncius, 16(1):49–84, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00028.

Zucchi:2001:LPB

[2125] Luca Zucchi. Linneo e Parkinson: Il botanico e le scimmie nel Giardino Dell’Eden. Nuncius, 16(1):85–152, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00037.

Funaro:2001:CMV

[2126] Liana Elda Funaro. ≪il criterio e la mano≫. Viaggi e donativi sovrani all’Imperiale e Regio Museo Fiorentino. Nuncius, 16(1):153–189, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00046.

Vlahakis:2001:AFS

[2127] George N. Vlahakis. Against French science: Alessandro Volta and Luigi Brugnatelli in early Nineteenth-Century Greece. Nuncius, 16(1):191–210, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00055.

Passione:2001:MLP

[2128] Roberta Passione. Mente e lavoro. Le prime ricerche in Italia fra laboratorio e officina. Nuncius, 16(1):211–235, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539101x00064.

Meschiari:2001:PAD

[2129] Alberto Meschiari. Per un archivio della corrispondenza degli scienziati italiani le carte del ≪Fondo Giovanni Battista Amici≫ e della ≪Raccolta Amici Grossi≫ nella Biblioteca Estense di Modena. Nuncius, 16(1):237–300, ???? 2001. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539101x00073.

Ulivi:2001:IFM

[2130] Elisabetta Ulivi. Istituzioni e fonti mariano del M^{∘} Michele, un Maestro d’abaco del XV secolo. (German) [institutions and sources for Mariano Del M^{∘} Michele, a master of the abacus in the 15th Century]. Nuncius, 16(1):301–346, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00082.

Pogliano:2001:SRD

[2131] Claudio Pogliano. Statements on race dell’UNESCO: Cronaca di un lungo travaglio (1949–1953). (Italian) [UNESCO statements on race: Chronicle of a long labor (1949–1953)]. Nuncius, 16(1):347–399, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00091.

Polizzi:2001:BRT

[2132] Gaspare Polizzi. Book review: Thomas S. Kuhn, Dogma contro critica. Mondi possibili nella storia della scienza, con due lettere di Paul K. Feyerabend, a cura di Stefano Gattei (≪Scienze e Idee≫, 61), Milano, Cortina Editore, 2000, 400 pp., L. 59.000 (ISBN 88-7078-619-6). Nuncius, 16(1):401–405, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00109.

Vista:2001:BRP

[2133] Michele Vista. Book review: Philip Ball, H_{2}O — Una biografia dell’acqua, Milano, Rizzoli Editore 2000, 432 pp., L. 38.000. Nuncius, 16(1):405–407, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00118.

Iannaccone:2001:BRB

[2134] Isaia Iannaccone. Book review: T. Brook, G. Blue (eds.), China and Historical Capitalism. Genealogies of Sinological Knowledge, Cambridge, Cambridge University Press, 1999, 291 pp. + xi (ISBN 0-521-64029-6). Nuncius, 16(1):407–409, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00127.

diPasquale:2001:BRS

[2135] Giovanni di Pasquale. Book review: Sabrina Grimaudo, Misurare e pesare nella Grecia antica. Teoria, storia, ideologie, Palermo, l’epos, 1998, 206 pp. Nuncius, 16(1):409–411, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00136.

Strano:2001:BRG

[2136] Giorgio Strano. Book review: Giovanni Lupato, Sn1054: una supernova sul Medioevo, Padova, Biroma editore, 1997, 208 pp. ill., L. 35.000 (ISBN 88-86739-12-5). Nuncius, 16(1):411–415, ????

2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00145.

Iannaccone:2001:BRY

[2137] Isaia Iannaccone. Book review: Yung Sik Kim, The Natural Philosophy of Chu Hsi. 1130–1200, Philadelphia, Memoirs of the American Philosophical Society, 2000, 380 pp. + ill. (ISBN 0-87169-235-X). Nuncius, 16(1):415–417, ???? 2001. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539101x00154.

Tessicini:2001:BRH

[2138] Dario Tessicini. Book review: Helisaeus Röslin, De opere Dei creationis, ristampa anastatica dell’edizione Francoforte 1597, a cura di Miguel Angel Granada, ≪Aurifodina philosophica≫, Conte Editore, Lecce 2000, xv + 55 pp. Nuncius, 16(1):418–420, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00163.

Guerrini:2001:BRW

[2139] Luigi Guerrini. Book review: W. R. Laird (ed.), The Unfinished Mechanics of Giuseppe Moletti. An Edition and English Translation of his ‘Dialogue on Mechanics’, 1576, University of Toronto Press, Toronto-Buffalo-London, 2000, 222 pp. Nuncius, 16(1):420–424, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00172.

Guerrini:2001:BRM

[2140] Luigi Guerrini. Book review: Michele Camerota e Mario Otto Helbing, All’alba della scienza galileiana. Michel Varro e il suo ‘De motu tractatus’. Un importante capitolo nella storia della meccanica di fine Cinquecento, Cagliari, Cooperativa Universitaria Editrice Cagliaritana, 2000, 367 pp. Nuncius, 16(1):424–427, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00181.

Tugnoli:2001:BRG

[2141] Claudio Tugnoli. Book review: Galileo Galilei, Lettera a Cristina di Lorena sull’uso della Bibbia nelle argomentazioni scientifiche, a cura di Franco Motta e con introduzione di Mauro Pesce, Marietti 1820, Genova, 2000, 173 pp. Nuncius, 16(1):427–431, ????

2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00190.

Aquilecchia:2001:BRC

[2142] Giovanni Aquilecchia. Book review: Chrysa Damianaki, Galileo e le arti figurative. I ritratti e i busti di Galileo. Scoperte astronomiche e pittura barocca. La concezione estetica di Galileo. Presentazione di Bruno Basile, Manziana, Roma, Vecchiarelli Editore, 2000, 103 pp. + 46 (51) tavv. Nuncius, 16(1):431–432, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00208.

Abbri:2001:BRMa

[2143] Ferdinando Abbri. Book review: Michael Hunter, Robert Boyle (1627– 91). Scrupulosity and Science, Woodbridge, the Boydell Press, 2000, x + 293 pp. (ISBN 0-85115-798-X). Nuncius, 16(1):432–434, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00217.

Bondi:2001:BRF

[2144] Roberto Bondì. Book review: Francesco Paolo De Ceglia, Introduzione alla fisiologia di Georg Ernst Stahl, Lecce, Pensa Multimedia, 2000, 236 pp., L. 28.000 (ISBN 88-8232-044-8). Nuncius, 16(1):434–435, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00226.

Abbri:2001:BRMb

[2145] Ferdinando Abbri. Book review: Michel Bougard, La Chimie de Nicolas Lemery, Turnhout, Brepols, 1999, 524 pp. ≪De Diversibus Artibus. Collection de Travaux de l’Académie Internationale d’Histoire des sciences≫, t. 40 N.S. 3 (ISBN 2-503-50739-5). Nuncius, 16(1):436–438, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00235.

Vasconi:2001:BRM

[2146] Paola Vasconi. Book review: Maria Jesus Vazquez Lobeiras, Die Logik und ihr Spiegelbild, (German) [Logic and its Mirror Image]. Frankfurt a.M., Berlin, Peter Lang Verlag, 1998. Nuncius, 16(1):438–440, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00244.

Triarico:2001:BRS

[2147] Carlo Triarico. Book review: Strumentaria Silvia De Renzi, Instruments in Print. Books from the Whipple Collection, Cambridge, Whipple Museum of the History of Science, 2000, x + 108 pp. Nuncius, 16(1):441–442, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00253.

Miniati:2001:BRAa

[2148] Mara Miniati. Book review: Antonio Campigotto, Alessandra Parisini, Patrizia Tamassia, Macchine orarie. Orologi da torre e orologiai in Emilia Romagna, Bologna, Istituto per i beni artistici, culturali e naturali della Regione Emilia-Romagna-Grafis, 2000, 267 pp., ill. (Ibc Dossier 38). Nuncius, 16(1):442–443, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00262.

Miniati:2001:BRN

[2149] Mara Miniati. Book review: Nicoletta Nicolini, Gigliola Terenna, La collezione di vetreria scientifica, Siena, Nuova immagine editrice, 1999, 192 pp. (ISBN 88-7145-160-0) (Materiali 3 — Patrimonio storicoscientifico dell’Università degli Studi di Siena). Renato Frezzotti, Gigliola Terenna, Francesca Vannozzi, La collezione degli strumenti di oculistica, Siena, Nuova immagine editrice, 2000, 157 pp. ill. (ISBN 88-7145-167-8) (Materiali 4 — Patrimonio storico-scientifico dell’Università degli Studi di Siena). Nuncius, 16(1):443–445, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00271.

Miniati:2001:BRM

[2150] Mara Miniati. Book review: Museologia Sylvie Coyaud, Matteo Merzagora, Guida ai musei della scienza e della tecnica. Torino, ClupGuide, 2000, 285 pp., L. 36.000 (Euro 18,59) (ISBN 88-8344-008-0). Nuncius, 16(1):445–446, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00280.

Miniati:2001:BRAb

[2151] Mara Miniati. Book review: Augusto Vigna Taglianti, Giorgio Di Maio, Adriana Vassallo Paleologo, Suoni natura e cultura, Roma, Università degli Studi ≪La Sapienza≫, 2000, 111 pp., ill. Nuncius, 16(1):446–447, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00299.

Lippi:2001:BRA

[2152] Donatella Lippi. Book review: Astronomia B. Stephenson, M. Bolt, A. F. Friedman, The universe unveiled. Instruments and Images through History, Cambridge, Cambridge Univ. Press 2000, 152 pp., (ISBN 0-521-79143-X). Nuncius, 16(1):447–448, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00307.

Miniati:2001:BRGa

[2153] Mara Miniati. Book review: Giovanni Paltrinieri, Calendario-lunario. La misura del tempo dalla preistoria a oggi, Bologna, Calderoni ed agricole, 2000, 329 pp., ill. (ISBN 88-8219-375-6). Nuncius, 16(1):449–450, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00316.

Lippi:2001:BRMa

[2154] Donatella Lippi. Book review: Medicina S. Franco, La politica sanitaria durante il decennio francese nel Regno di Napolip, Marina di Minturno (Lt), A. Caramanica Ed., 2000, 197 pp., L. 30.000. Nuncius, 16(1): 450–451, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00325.

Bertini:2001:BRP

[2155] Silvia Bertini. Book review: Psicologia A. Lis, S. Stella, G. C. Zavattini, Manuale di psicologia dinamica, Bologna, Il Mulino, 1999, 536 pp. Nuncius, 16(1):452–453, ???? 2001. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539101x00334.

Ciardi:2001:BRV

[2156] Marco Ciardi. Book review: Voltiana Alessandro Volta, on the Electricity excited by the mere Contact of conducting Substances of different kinds. Bicentenary Edition in French, English, German and Italian of the Letter to Sir Joseph Banks of the 20th of March 2000, Milano, Hoepli, 1999, 102 pp., L. 10.000 (ISBN 88-203-2762-7). Fabio Bevilacqua, Lucio Fregonese, Nuova Voltiana. Studies on Volta and his Times, Milano, Hoepli, 2000– 2001, 3 voll., 1^{∘} vol., 157 pp., L. 30.000 (ISBN 88-203-2789-9); 2^{∘} voi., 113 pp., L. 30.000 (ISBN 88-203-2790-2); 3^{∘} voi., 151 pp. (ISBN 88-203-2791-0). Gianni Bonera, Piersandro Vanzan, Alessandro Volta: l’uomo, lo scienziato, il credente, Pavia, Edizioni CdG, 1999, 152 pp., L. 10.000. Carlo Linati, Cupido fra gli alambicchi, Novara, Interlinea Edizioni, 2000, 124 pp. Nuncius, 16(1):454–455, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00343.

Ciardi:2001:BRS

[2157] Marco Ciardi and Ferdinando Abbri. Book review: Scienza e Letteratura. Paolo Mazzarello, Il genio e l’alienista. La visita di Lombroso a Tolstoj, Napoli, Bibliopolis, 1998, 94 pp., L. 25.000 (ISBN 88-7088-301-9). Nuncius, 16(1):456–457, ???? 2001. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539101x00352.

Anonymous:2001:SXX

[2158] Anonymous.  Secoli XVI–XXI.  Nuncius, 16(1):457–458, ????

2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00361.

Abbri:2001:BRR

[2159] Ferdinando Abbri. Book review: Roberto Torzini, I labirinti del libero arbitrio. La discussione tra Erasmo e Lutero, Firenze, Olschki, 2000, xix + 267 pp. (≪Pansophia≫, 4). Nuncius, 16(1):458–459, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00370.

Abbri:2001:BRW

[2160] Ferdinando Abbri. Book review: William R. Shea (ed.), Science and the Visual Image in the Enlightenment, Canton MA, Science History Publications/USA, 2000, viii + 232 pp., ill. (≪European Studies in Science History and the Arts≫, 4) [ISBN] 0-88135-285-3. Nuncius, 16(1):459– 461, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00389.

Ciardi:2001:BRB

[2161] Marco Ciardi. Book review: Brigitte Hoppe (ed.), Biology Integrating Scientific Fundamental. Contributions to the History of Interrelations between Biology, Chemistry, and Physics from the 18th to the 20th Centuries, München, Institut für Geschichte der Naturwissenschaften, 1997, 434 pp. (ISBN 3-89241-022-4) (≪Algorismus≫, 21). Bernhard Fritscher, Fergus Henderson (eds.), Toward a History of Mineralogy, Petrology, and Geochemistry. Proceedings of the International Symposium on the History of Mineralogy, Petrology, and Geochemistry, Munich, March 8–9, 1996, München, Institut für Geschichte der Naturwissenschaften, 1998, 424 pp. (ISBN 3-89241-024-0). (≪Algorismus≫, 23). Nuncius, 16(1): 461–463, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00398.

Anonymous:2001:BRIa

[2162] Anonymous. Book review: The IMSS Bookshelf 36. Germana Pareti, Il Cancro dell’Imperatore Dalla Teoria cellulare alle ipotesi oncogenetiche 2000, 220 pp., Olschki, Firenze. Nuncius, 16(1):465–466, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00406.

Anonymous:2001:BRFa

[2163] Anonymous. Book review: 37. Filippo Camerota, Il Compasso di Fabrizio Mordente Per la storia del Compasso di proporzione 2000, 300 pp. con 21 taw. f.t e 134 figg. n.t. Olschki, Firenze. Nuncius, 16(1):466–467, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00415.

Anonymous:2001:BRAa

[2164] Anonymous. Book review: 38. Antonio Borrelli, Istituzioni scientifiche Medicina e Società Biografia di Domenico Cotugno (1736–1822) 2000, xii + 272 pp., Olschki, Firenze. Nuncius, 16(1):467–468, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00424.

Anonymous:2001:BRJa

[2165] Anonymous. Book review: 39. Jean–François Stoffel Bibliographie d’Alexandre Koyré 2000, xiv + 196 pp., Olschki, Firenze. Nuncius, 16(1): 468, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00433.

Anonymous:2001:BRRa

[2166] Anonymous. Book review: 40. Il ruolo sociale della scienza (1789–1830) a cura di Ferdinando Abbri e Marco Segala 2000, vi + 178 pp., Olschki, Firenze. Nuncius, 16(1):469, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00442.

Anonymous:2001:BROa

[2167] Anonymous. Book review: 41. Ottavio Besomi e Michele Camerota, Galileo e il parnaso Tychonico un capitolo inedito del dibattito sulle comete tra finzione letteraria e trattazione scientifica, 2000, vi + 276 pp., Olschki, Firenze. Nuncius, 16(1):469–470, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00451.

Anonymous:2001:BRFb

[2168] Anonymous. Book review: Fondazione Scienza e Tecnica Gli strumenti di fisica dell’Istituto Toscano Elettricità e magnetismo, (Italian) [Science and Technology Foundation: The tools of the Toscano Institute of Physics: Electricity and Magnetism], a cura di Paolo Brenni 2000, 319 pp. Le Lettere, Firenze. Nuncius, 16(1):470, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00460.

Anonymous:2001:BRIb

[2169] Anonymous. Book review: Istituto e Museo di Storia della Scienza, Florence Opificio delle Pietre Dure, Florence The Restoration of Scientific Instruments. Proceedings of the workshop held in Florence December 14–15, 1998, 2000, 113 pp., Le Lettere, Firenze. Nuncius, 16(1): 471, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00479.

Renn:2001:LG

[2170] Jürgen Renn and Matteo Valleriani. Letture galileiane. Nuncius, 16 (2):481–503, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00497.

Odifreddi:2001:GIE

[2171] Piergiorgio Odifreddi. Galileo: Infinitesimale, eppur infinito. Nuncius, 16(2):505–510, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00505.

DAlessandro:2001:PCF

[2172] Paolo D’Alessandro and Pier Daniele Napolitani. I primi contatti fra Maurolico e Clavio: Una nuova edizione della lettera di Francesco Maurolico a Francisco Borgia. Nuncius, 16(2):511–522, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00514.

Helbing:2001:PIS

[2173] Mario Otto Helbing. Una prima illustrazione a stampa del termometro. Nuncius, 16(2):523–529, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00523.

Lualdi:2001:FSO

[2174] Alberto Lualdi. La famiglia selva, ottici del ’700 veneziano. Nuncius, 16 (2):531–546, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00532.

Triarico:2001:AND

[2175] Carlo Triarico. About newly discovered pieces by Boscovich on optics and on the finding of a vitrometrum. Nuncius, 16(2):547–569, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00541.

Chinnici:2001:CAS

[2176] Ileana Chinnici and Wiktor Gramatowski. Le carte di Angelo Secchi S.J. (1818–1878) conservate presso la Pontificia Università Gregoriana. Un inventario inedito rivisitato. (Italian) [The maps of Angelo Secchi S.J. (1818–1878) preserved at the Pontifical Gregorian University. An unpublished inventory revisited]. Nuncius, 16(2):571–627, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00550.

Beretta:2001:IF

[2177] Francesco Beretta. Istituzioni e fonti. Nuncius, 16(2):629–641, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00569.

Guerrini:2001:NDC

[2178] Luigi Guerrini. Nel Dedalo del cielo. Gassendi e le macchie solari in un inedito del 1633. (Italian) [From Daedalus of the heavens. Gassendi and the sunspot in an unpublished work of 1633]. Nuncius, 16(2):643–672, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00578.

Anonymous:2001:SIC

[2179] Anonymous. Scientific instrument commission. Nuncius, 16(2):673, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00587.

Turner:2001:ASC

[2180] Anthony Turner. The archives of scholars, collectors, and dealers: Their place in the study of the history of scientific instruments. Nuncius, 16 (2):675–679, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00596.

Miniati:2001:AFS

[2181] Mara Miniati. About the Florentine Strozzi manuscripts. Nuncius, 16 (2):681–686, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00604.

Andrewes:2001:LDW

[2182] William J. H. Andrewes. The legacy of David Wheatland. Nuncius, 16 (2):687–701, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00613.

Bruyns:2001:ABD

[2183] Willem F. J. Mrzer Bruyns. Alain Brieux, dealer and scholar in Paris: His archive on scientific instruments. Nuncius, 16(2):703–709, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00622.

Ackermann:2001:DTZ

[2184] Silke Ackermann. Dormant treasures: the Zinner-Archive at Frankfurt University. Nuncius, 16(2):711–722, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539101x00631.

DeClercq:2001:PID

[2185] Peter De Clercq. The papers of the instrument dealer, collector and researcher Thomas Henry Court (1868–1951). Nuncius, 16(2):723–731, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00640.

VanCleempoel:2001:HMG

[2186] Koenraad Van Cleempoel. Henri Michel, a gentleman–scholar. Nuncius, 16(2):733–738, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00659.

Stephenson:2001:DPA

[2187] Bruce Stephenson. The Derek Price Archive at the Adler Planetarium. Nuncius, 16(2):739–746, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00668.

Turner:2001:ADP

[2188] A. J. Turner. The archives of Derek Price in the médiathèque spécialis en histoire des sciences, Cité des Sciences et de l’Industrie at La Villette in Paris. Nuncius, 16(2):747–761, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00677.

Keill:2001:PMB

[2189] Inge Keill. The papers of Maximilian Bobinger. Nuncius, 16(2):763–765, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00686.

Tolmasquim:2001:NM

[2190] Alfredo Tiomno Tolmasquim, Alexandre Magno Da Costa, and Lucia Alves Lino. Nova media. Nuncius, 16(2):767–779, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00695.

Palladino:2001:SRM

[2191] Franco Palladino and Nicla Palladino. Sulle raccolte museali italiane di modelli per le matematiche superiori. Nuncius, 16(2):781–790, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00703.

Angelini:2001:MML

[2192] Annarita Angelini. Magia e modernit. l’edizione delle opere magiche di Giordano Bruno. Nuncius, 16(2):791–798, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539101x00712.

Beretta:2001:RSC

[2193] Marco Beretta. Reconstructing science: Contributions to the enhancement of European scientific heritage. Nuncius, 16(2):799–816, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00721.

Abbri:2001:BRA

[2194] Ferdinando Abbri. Book review: Arne Hessenbruch (ed.), Reader’s Guide to the History of Science, London/Chicago, Fitzroy Dearborn Publishers, 2000, xxx + 934 pp., ISBN 1-884964-29-X. Nuncius, 16(2):817–819, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00730.

Miniati:2001:BRPa

[2195] Mara Miniati. Book review: Peter De Clercq (ed.), Scientific instruments: originals and imitations. The Mensing connection. Proceedings of a symposium, held at the Museum Boerhaave, Leiden, 15–16 October 1999, Leiden, Museum Boerhaave, 2000, 141 pp., ill., ISBN 90-6292-129-9. Nuncius, 16(2):819–821, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00749.

Kleinert:2001:BRC

[2196] Andreas Kleinert. Book review: Christoph Meinel (Ed.): Instrument — Experiment. Historische Studien. Im Auftragdes Vorstandes der Deutschen Gesellschaft für Geschichte der Medizin, Naturwissenschaft und Technik, Berlin, Diepholz: Verlag für Geschichte der Naturwissenschaften und der Technik, 2000, Eur 34,00, ISBN 3-928186-51-5. Nuncius, 16(2):822–823, ???? 2001. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539101x00758.

Iannaccone:2001:BRK

[2197] Isaia Iannaccone. Book review: Kiyosi Yabuuti, Une histoire des mathématiques chinoises (tradotto dal giapponese da Kaoru Baba e Catherine Jami), Parigi, Benin, 2000, 191 pp. + ill. (ISBN 2-7011-2404-2), (Pour la Science). Nuncius, 16(2):823–825, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00767.

Seligardi:2001:BRJ

[2198] Raffaella Seligardi. Book review: Jürgen Heinrich Maar, Pequeña história da química. Primeira parte — Dos primórdios a Lavoisier / una história da ciéncia da matéria, (Spanish) [A little history of chemistry. First part. Two first to Lavoisier]. Florianopolis: Papa-Livro, 1999, 848 pp., ISBN 85-7291-049-2. Nuncius, 16(2):825–827, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00776.

Bellino:2001:BRR

[2199] Francesca Bellino. Book review: Roshdi Rashed, Les mathématiques infinitésimales du IXe au XIe siécle (al-Riyyt al-tallyah bayna al-qarn al-thlith wa-al-qarn al-khmis). Vol. III: Ibn al-Haytham. Théorie des coniques, construction géométriques et géométrie pratique, Al-Furqn, London 1421/2000, XVIII, 1034 pp. Nuncius, 16(2):827–831, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00785.

Abbri:2001:BRT

[2200] Ferdinando Abbri. Book review: Tore Frngsmyr, Svensk idhistoria. Bildning och vetenskap under tusen r. Del 1 1000–1809. Del 2 1809–2000, Stockholm, Natur och Kultur, 2000, 2 voll., 448 + 422 pp., ill., ISBN 91-27-07793-4. 91-27-08105-2. Nuncius, 16(2):831–833, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00794.

Toscano:2001:BRS

[2201] Anna Toscano. Book review: Sciences et Religions. De Copernic à Galile, Actes du colloque international Rome 12–14 décembre 1996, École française de Rome, 1999 (Collection de l’école française de Rome n. 260), pp. 550, ISBN 2-7283-0545-5, ISSN 0223-5099. Nuncius, 16(2):833–836, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00802.

Guerrini:2001:BRT

[2202] Luigi Guerrini. Book review: Tommaso Campanella, Opere, a cura di Germana Ernst. Introduzione di Nicola Badaloni, Istituto Poligrafico e Zecca dello Stato, Roma 1999, pp. lv + 1027 (Cento Libri per Mille anni. Collana diretta da Walter Pedull). Nuncius, 16(2):836–839, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00811.

Bond:2001:BRG

[2203] Roberto Bond. Book review: Guido Giglioni, Immaginazione e malattia: saggio su Jan Baptiste van Helmont, Milano, Franco Angeli, 2000, 187 pp., L. 32.000 (ISBN 88-464-2286-4). Nuncius, 16(2):839–843, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00820.

Olmi:2001:BRA

[2204] Giuseppe Olmi. Book review: Arthur Macgregor et alii, Manuscript Catalogues of the Early Museum Collections 1683–1886 (Part I), Oxford, Archaeopress in association with the Ashmolean Museum, 2000, pp. 451, £50.00, ISBN 1-84171-206-X (Bar International Series, 907). Nuncius, 16(2):843–844, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00839.

Toscano:2001:BRA

[2205] Anna Toscano. Book review: Albrecht Von Haller, Commentarius de formatione cordis in ovo incubato, edizione critica a cura di Maria Teresa Monti, Schwabe Co. Ag Verlag, Basel 2000 (Studia Halleriana VI), pp. 677 + CD-ROM, ISBN 3-7965-1324-7. Nuncius, 16(2):845–847, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00848.

Nani:2001:BRS

[2206] Michele Nani. Book review: Sandra Puccini, Andare lontano. Viaggi ed etnografia nel secondo Ottocento, Roma, Carocci, 1999 (Ricerche/ Antropologia, 63), pp. 297, ISBN 88-430-1448-X, L. 37.000 [22,21 Euro]. Nuncius, 16(2):847–849, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x00857.

Mastrobisi:2001:BRM

[2207] Giorgio Jules Mastrobisi. Book review: Max Jammer, Einstein and Religion: Physics and Theology, Princeton, Princeton University Press, 1999, pp. 279 (ISBN 0-691-00699-7). Nuncius, 16(2):850–854, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00866.

Seligardi:2001:BRA

[2208] Raffaella Seligardi. Book review: Andreas Karachalios, I chimici di fronte al fascismo. Il caso di Giovanni Battista Bonino (1899–1985), Palermo, Istituto Gramsci Siciliano, 2001, 219 pp. Nuncius, 16(2):854– 855, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00875.

Miniati:2001:Ma

[2209] Mara Miniati. Museologia. Nuncius, 16(2):857–858, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539101x00884.

Miniati:2001:S

[2210] Mara Miniati. Strumentaria. Nuncius, 16(2):858–859, ???? 2001. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539101x00893.

Miniati:2001:AF

[2211] Mara Miniati. Astronomia e fisica. Nuncius, 16(2):859–860, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00901.

Miniati:2001:BRGb

[2212] Mara Miniati. Book review: Graziella Buccellati (a cura di), I cieli di Brera. Astronomia da Tolomeo a Balla, Milano, Università degli Studi di Milano, 2000, 207 pp., ill. Nuncius, 16(2):860–861, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00910.

Miniati:2001:BRF

[2213] Mara Miniati. Book review: Fabrizio Bonoli, Daniela Piliarvu, I lettori di Astronomia presso lo Studio di Bologna dal XII al XX secolo, Bologna, Clueb, 2001, 282 pp., ill., ISBN 88-491-1651-9 (Musei e archivi dello studio bolognese, 7). Pierluigi Battistini, Fabrizio Bonoli, Alessandro Braccesi, Dino Buzzetti (eds.), Seventh centenary of the teaching of Astronomy in Bologna 1297–1997, Bologna, Clueb, 2001, 212 pp., ill., ISBN 88-491-1650-0 (Musei e archivi dello studio bolognese, 8). Nuncius, 16(2):861–863, ???? 2001. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539101x00929.

Guerrini:2001:BRR

[2214] Luigi Guerrini. Book review: A. Romano, La contre-réforme mathématique. Constitution et diffusion d’une culture mathématique jésuite la Renaissance, Bibliothéque des écoles françaises d’Athènes et de Rome (fascicule trois cent sixiéme). Roma, École française de Rome, 1999 (pp. xi + 691 + 21 tavole fuori testo). Nuncius, 16(2):863–864, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00938.

Ciardi:2001:BRLa

[2215] Marco Ciardi. Book review: Luciano Boi, Livia Giacardi, Rossana Tazzioli, La découverte de la géomtrie non euclidienne sur la pseudosphère. Les lettres d’Eugenio Beltrami à Jules Hoel (1868–1881), (French) [The discovery of non-Euclidean geometry on the pseudosphere. The letters of Eugenio Beltrami to Jules Hoel (1868–1881)]. Paris, Éditions Albert Blanchard, 1998, 278 pp., 480 F, ISBN 2-85367-208-5 (Collection Sciences dans l’Histoire). Nuncius, 16(2):865, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00947.

Ciardi:2001:BRG

[2216] Marco Ciardi. Book review: Giovanni Battimelli, Giovanni Paoloni (eds.), 20th Century Physics: Essays and Recollections. A Selection of Historical Writings by Edoardo Amaldi, Singapore-New Jersey-London-Hong Kong, World Scientific Publishing, 1998, 747 pp., ill., ISBN 981-02-2369-2 (Edoardo Amaldi Foundations series, 3). Nuncius, 16(2):865–866, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00956.

Miniati:2001:BRPb

[2217] Mara Miniati. Book review: Pasquale Tucci (a cura di), Atti del XIX Congresso Nazionale di storia della fisica e dell’astronomia, centro Volta, Villa Olmo, Como 28–29 maggio 1999, Milano, Università degli Studi di Milano, Istituto di Fisica generale applicata, Sez. di storia della fisica, 2000, 312 pp. Nuncius, 16(2):866–867, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x00965.

Miniati:2001:Mb

[2218] Mara Miniati. Meteorologia. Nuncius, 16(2):867–868, ???? 2001. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539101x00974.

Abbri:2001:M

[2219] Ferdinando Abbri. Medicina. Nuncius, 16(2):868–869, ???? 2001. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539101x00983.

Lippi:2001:BRMb

[2220] Donatella Lippi. Book review: T. Manzoni, Il cervello secondo Galeno, Ancona, Il lavoro editoriale, 2001, pp. 138, L. 30.000. Nuncius, 16(2): 869–871, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x00992.

Miniati:2001:BRRa

[2221] Mara Miniati. Book review: Roberta Mucciarelli, Laura Vigni, Donatella Fabbri, Vergognosa immunditia. Igiene pubblica e privata a Siena dal medioevo all’et contemporanea, Siena, sienambente, 2000, 223 pp., ill., ISBN 88-7145-169-4. Nuncius, 16(2):871–872, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01009.

Lippi:2001:BRMc

[2222] Donatella Lippi. Book review: M. Rotzoll, Pierleone da Spoleto. Vita e opere di un medico del Rinascimento, Firenze, Olschki, 2000, p. 126, ISBN 88-222-4842-2, L. 26.000. Nuncius, 16(2):872–873, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01018.

Ciardi:2001:C

[2223] Marco Ciardi. Chimica. Nuncius, 16(2):874, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539101x01027.

Ciardi:2001:BRLb

[2224] Marco Ciardi. Book review: Liliana Albertazzi, Guido Cimino, Simonetta Gori-Savellini, Francesco De Sarlo e il laboratorio fiorentino di psicologia, Bari, Edizioni Giuseppe Laterza di Giuseppe Laterza, 1999, 467 pp., L. 70.000, ISBN 88-8231-075-2 (Università degli Studi di Bari. Collana del Seminario di Storia della Scienza, 5). Nuncius, 16(2):874–875, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01036.

Miniati:2001:BRRb

[2225] Mara Miniati. Book review: Raffaella Simili (a cura di), Il teatro della natura di Ulisse Aldrovandi, Bologna, Editrice Compositori, 2001, 143 pp., ill., ISBN 88-7794-272-X. Nuncius, 16(2):875–876, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01045.

Miniati:2001:BRE

[2226] Mara Miniati. Book review: Enrica Stendardo, Ferrante Imperato: collezionismo e studio della natura a Napoli tra Cinque e Seicento, Napoli, Accademia Pontaniana, 2001, 155 pp., ill. (Quaderni dell’accademia pontaniana 31). Nuncius, 16(2):876–877, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01054.

Abbri:2001:BRD

[2227] Ferdinando Abbri. Book review: Daniel Stolzenberg (ed.), The Great Art of Knowing. The Baroque Encyclopedia of Athanasius Kircher, Stanford, Stanford University Libraries, Fiesole Fi, Edizioni Cadmo, 2001, pp. 160. ISBN 0-911221-23-9. Nuncius, 16(2):877–879, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01063.

Abbri:2001:BRN

[2228] Ferdinando Abbri. Book review: Natalie Kunkel, Wissenschaftsaustausch zwischen Russland und Westeuropa, insbesondere Deutschland, in der Botanik und Pharmazie vom 18. bis zum frühen 20. Jahrhundert, (German) [Scientific exchange between Russia and Western Europe, especially Germany, in botany and pharmacy in the 18th to the early 20th Century], München, Institut für Geschichte der Naturwissenschaften, 1999, 277 pp. (Algorismus, 28). ISBN 3-89241-029-1. Nuncius, 16(2):879–880, ???? 2001. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539101x01072.

Ciardi:2001:BRP

[2229] Marco Ciardi. Book review: Patrice Bret, Niarcel Dorigny (ed.), Sciences et techniques autour de la Révolution française. Approche interdisciplinaire. Journées d’étude des 10 et 11 décembre 1999, Cité des sciences et de l’industrie — Monnaie de Paris, Paris, Société des études robespierristes, 2000, 230 pp., 120 F/18,29, ISBN 2-908327-41-4 (Annales historiques de la Révolution française, 320). Nuncius, 16(2):880–881, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01081.

Contardi:2001:PCC

[2230] Simone Contardi. Piero Cuccuini Chiara Nepi, Herbarium Centrale Italicum (the phanerogamic section): the Genesis and structure of a herbarium, Firenze, 1999. Nuncius, 16(2):881–882, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01090.

Pogliano:2001:BRG

[2231] Claudio Pogliano. Book review: Giuliana Gemelli (ed.), The Unacceptables. American Foundations and Refugee Scholars between the Two Wars and after, Bruxelles, Presses Interuniversitaires Europennes — Peter Lang, 2000, 306 pp. (Euroclio, n. 18). Nuncius, 16(2):882–883, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01108.

Ciardi:2001:BRE

[2232] Marco Ciardi. Book review: Edoardo Rovida, Dallo scalpello al mouse. Evoluzione e storia del disegno tecnico, Torino, Paravia, 1999, 141 pp., ill., L. 40.000, ISBN 88-395-8300-9. Nuncius, 16(2):883, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01117.

Anonymous:2001:BRG

[2233] Anonymous. Book review: Germana Pareti, Il Cancro dell’Imperatore Dalla Teoria cellulare alle ipotesi oncogenetiche 2000, 220 pp., Olschki, Firenze. Nuncius, 16(2):885–886, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x01126.

Anonymous:2001:BRFc

[2234] Anonymous. Book review: Filippo Camerota, Il Compasso di Fabrizio Mordente Per la storia del Compasso di proporzione 2000, 300 pp. con 21 taw. f.t e 134 figg. n.t. Olschki, Firenze. Nuncius, 16(2):886–887, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01135.

Anonymous:2001:BRAb

[2235] Anonymous. Book review: Antonio Borrelli, Istituzioni scientifiche Medicina e Societ Biografia di Domenico Cotugno (1736–1822) 2000, xii + 272 pp., Olschki, Firenze. Nuncius, 16(2):887–888, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01144.

Anonymous:2001:BRJb

[2236] Anonymous. Book review: Jean-François Stoffel, Bibliographie d’Alexandre Koyr 2000, xiv + 196 pp., Olschki, Firenze. Nuncius, 16 (2):888–889, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x01153.

Anonymous:2001:BRRb

[2237] Anonymous. Book review: Il ruolo sociale della scienza (1789–1830). (Italian) [The social role of science (1789–1830)], a cura di Ferdinando Abbri e Marco Segala 2000, vi + 178 pp., Olschki, Firenze. Nuncius, 16(2):889, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539101x01162.

Anonymous:2001:BROb

[2238] Anonymous. Book review: Ottavio Besomi e Michele Camerota, Galileo e il parnaso Tychonico: un capitolo inedito del dibattito sulle comete tra finzione letteraria e trattazione scientifica 2000, vi + 276 pp., Olschki, Firenze. Nuncius, 16(2):889–890, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x01171.

Anonymous:2001:BRFd

[2239] Anonymous. Book review: Fondazione Scienza e Tecnica Gli strumenti di fisica dell’Istituto Toscano Elettricità e magnetismo, (Italian) [Science and Technology Foundation: The tools of the Tuscan Institute of Physics: Electricity and magnetism], a cura di Paolo Brenni 2000, 319 pp. Le Lettere, Firenze. Nuncius, 16(2):890, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539101x01180.

Anonymous:2001:BRIc

[2240] Anonymous. Book review: Istituto e Museo di Storia della Scienza, Florence Opificio delle Pietre Dure, Florence The Restoration of Scientific Instruments. Proceedings of the workshop held in Florence December 14–15, 1998, 2000, 113 pp., Le Lettere, Firenze. Nuncius, 16(2): 891, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01199.

Anonymous:2001:IAX

[2241] Anonymous. Indici (anno XVI, 2001). Nuncius, 16(2):893–896, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01207.

Anonymous:2001:IM

[2242] Anonymous. Indice dei manoscritti. Nuncius, 16(2):897–903, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01216.

Anonymous:2001:IN

[2243] Anonymous. Indice dei nomi. Nuncius, 16(2):905–963, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01225.

Anonymous:2001:ISI

[2244] Anonymous. Indice dei siti Internet. Nuncius, 16(2):965–966, ???? 2001. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539101x01234.

Casati:2002:NLI

[2245] Stefano Casati. The new library of the Institute and Museum of History of Science, Florence. Nuncius, 17(1):3–8, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539102x00568.

Bellissima:2002:SAD

[2246] Fabio Bellissima. Il sistema assiomatico-deduttivo degli elementi armonici di Aristosseno. Nuncius, 17(1):9–44, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539102x00577.

Reynolds:2002:GGS

[2247] Anne Reynolds. Galileo Galilei and the satirical poem ≪contro il portar la toga≫: the literary foundations of science. Nuncius, 17(1):45–62, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00586.

Casati:2002:CLD

[2248] Stefano Casati and Giorgio Strano. Il ‘candore lunare’ e la difesa del sistema copernicano in due lettere galileiane conservate presso la Biblioteca Dell’Istituto e Museo di Storia della Scienza di Firenze. (Italian) [The ‘lunar whiteness’ and the Copernican defense system in two Galilean letters stored at the Library of the Institute and Museum of the History of Science of Florence]. Nuncius, 17(1):63–76, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00595.

Proverbio:2002:PAD

[2249] Edoardo Proverbio. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 17(1):77–84, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539102x00603.

Anonymous:2002:AIA

[2250] Anonymous. Alphabetical index of addressees. Nuncius, 17(1):89, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00612.

Anonymous:2002:AIS

[2251] Anonymous. Alphabetical index of senders. Nuncius, 17(1):90, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00621.

Vergari:2002:COT

[2252] Daniele Vergari. La corrispondenza di Ottaviano Targioni Tozzetti. Nuncius, 17(1):91–155, ???? 2002. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539102x00630.

Anonymous:2002:IDD

[2253] Anonymous. Indice delle date delle lettere. Nuncius, 17(1):156–159, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00649.

Anonymous:2002:ILP

[2254] Anonymous. Indice dei luoghi di partenza delle lettere. Nuncius, 17 (1):160–164, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00658.

Meschiari:2002:CGB

[2255] Alberto Meschiari. Corrispondenza di Giovanni Battista Amici con Franz Xaver Von Zach. Nuncius, 17(1):165–245, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539102x00667.

Bertini:2002:IF

[2256] Marta Bertini. Istituzioni e fonti. Nuncius, 17(1):247–262, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00676.

Marangon:2002:CAN

[2257] Martino Marangon. I codici astronomici nel Fondo Cicogna al Museo Correr di Venezia. Nuncius, 17(1):263–267, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539102x00685.

Mastrobisi:2002:MSA

[2258] Giorgio Jules Mastrobisi. Il ≪manoscritto di Singapore≫ (1923) di Albert Einstein. Per una teoria del ≪campo unificato≫ tra possibilità fisica e necessità matematica (Italian) [The ≪Singapore manuscript≫ (1923) of Albert Einstein. On a theory of the ≪unified field≫ between physical possibility and mathematical need]. Nuncius, 17(1):269–305, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00694.

Ferrari:2002:CFE

[2259] Graziano Ferrari. Census, filing and elaboration of scientific letters in the Earth sciences. Nuncius, 17(1):307–320, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539102x00702.

Ruffo:2002:BGN

[2260] Patrizia Ruffo. La bibliografia galileiana nell’archivio integrato galileothek. (Italian) [The Galilean bibliography of the Galileothek Integrated Archive]. Nuncius, 17(1):321–329, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539102x00711.

Lenzi:2002:ABD

[2261] Alessandra Lenzi. Attività bibliografica dell’Istituto e Museo di Storia della Scienza: lo Spoglio dei Periodici. Nuncius, 17(1):331–336, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00720.

Beretta:2002:MRC

[2262] Marco Beretta. The material realm of chemistry: Immutability vs. history? Nuncius, 17(1):337–347, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539102x00739.

Govoni:2002:QSS

[2263] Paola Govoni. Quando la scienza è sensazionale. Nuncius, 17(1):349– 355, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00748.

Polizzi:2002:RDP

[2264] Gaspare Polizzi. Una rivoluzione dimenticata. Per una nuova comprensione della scienza ellenistica. (Italian) [A forgotten revolution. Towards a new understanding of Hellenistic science]. Nuncius, 17(1):357–366, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00757.

Lippi:2002:BRG

[2265] Donatella Lippi. Book review: Gunther Von Hagens, Kórperwelte. Die Faszination des echten. Der Katalog zur Ausstellung, (German) [Bodyworld. The fascination of the real. The exhibition catalog], Heidelberg, Institut für Plastination Heidelberg, 2001, pp. 295. Nuncius, 17(1):367– 368, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00766.

Vista:2002:BRZ

[2266] Michele Vista. Book review: Zagli Andrea, Mineccia Francesco, e Giuntini Andrea, ≪Maladetti Beccari≫. Storia dei macellai fiorentini dal Cinquecento al Duemila, Firenze, Edizioni Polistampa, 2000, pp. 293, ill., 35.00 Euro, ISBN 88-8304-267-0. Nuncius, 17(1):368–370, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00775.

Abbri:2002:BRRa

[2267] Ferdinando Abbri. Book review: Robert Halleux, James Mcclellan, Daniela Berariu, Genevieve Xhayet, Les Publications de l’Académie Royale des Sciences de Paris (1666–1793). T. I. Description Bibliographique. T. II. Étude Statistique — Index Généraux, Turnhout, Brepols, 2001, 557 + 128 pp., ISBN 2-503-51197-X (≪De Diversis Artibus. Collection de travaux de l’académie internationale d’histoire des sciences≫. N.S. 15-1, 15-2). Nuncius, 17(1):370–371, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00784.

Guicciardini:2002:BRR

[2268] NiccolÒ Guicciardini. Book review: Richard Dalitz e Michael Nauenberg (eds.), The foundations of Newtonian scholarship, Singapore/ New Jersey/London/Hong Kong, World Scientific, 2000, xviii + 242 pp., s.i.p., ISBN 98-102-3920,3. Nuncius, 17(1):372–374, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00793.

Stancati:2002:BRI

[2269] Claudia Stancati. Book review: Ilaria Tani, l’Albero della mente. Sensi, pensiero, linguaggio in Herder, Roma, Carocci, 2000, pp. 218, lire 30.000. ISBN 88-430-1546-X. Nuncius, 17(1):374–377, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00801.

Ciardi:2002:BRE

[2270] Marco Ciardi. Book review: Enrico Giusti e Luigi Pepe (a cura di), La matematica in Italia (1800–1950), Firenze, Edizioni Polistampa, 2001, 182 pp., ill., EUR12,91, ISBN 88-8304-321-9, (≪Il Giardino di Archimede. un museo per la matematica≫ ). Nuncius, 17(1):377, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00810.

Miniati:2002:BRA

[2271] Mara Miniati. Book review: Alessandro Paoletti, Nadia Robotti, Fisica meteorologia e sismologia nell’Ottocento. Il contributo del seminario Vescovile di Chiavari, Genova, Università degli Studi di Genova, Dipartimento di Fisica, 2000, 138 pp., ill. Nuncius, 17(1):378, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00829.

Pogliano:2002:BRL

[2272] Claudio Pogliano. Book review: Lorenzo Calabi, I quaderni metafisici di Darwin. Teleologia ≪Metafisica≫ Causa Finale, Pisa, Edizioni Ets, 2001, 134 pp., EUR10.33, ISBN 88-467-0441-X. Nuncius, 17(1):379–381, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00838.

Scarpelli:2002:BRB

[2273] Giacomo Scarpelli. Book review: Bernadette Bensaude-Vincent e Christine Blondel (edd.), Des savants face à l’occulte, 1870–1940, Paris, Éditions La Découverte, 2002, 233, pp., 17,50, ISBN 2-7071-3616-6 (≪Sciences et Société≫ ). Nuncius, 17(1):381–384, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00847.

Shea:2002:BRR

[2274] William R. Shea. Book review: Raffaella Simili and Giovanni Paoloni (eds.), Per una storia del Consiglio Nazionale delle Ricerche. Rome, Edizioni Laterza, 2001. 2 vols. xi + 668 and xi + 876 pages. Price: EUR36,15 per volume. ISBN 88-420-5929-3 and 88-420-6225-1. Nuncius, 17(1):384– 385, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00856.

Lippi:2002:BRS

[2275] Donatella Lippi. Book review: Silvano Franco, Legislazione e politica sanitaria del fascismo, Roma, Editrice Apes, 2001, pp. 209, L. 30.000, ISBN 88-7233-016-5. Nuncius, 17(1):385–386, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00865.

Ciardi:2002:BRH

[2276] Marco Ciardi. Book review: Helmuth Plessner, Il riso e Il pianto. Una ricerca sui limiti del comportamento umano, a cura di Vallori Rasini, Milano, Bompiani, 2000, 261 pp., EUR17,56, ISBN 88-452-4665-5 (≪Filosofia≫, a cura di Pier Aldo Rovatti). Nuncius, 17(1):387–389, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00874.

Laurenza:2002:BRF

[2277] Domenico Laurenza. Book review: De Figura Umana: Fisiognomia, Anatomia e Arte in Leonardo, 2001, 241 pp., Olschki, Firenze. Nuncius, 17(1):391–392, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00883.

Santorio:2002:BRM

[2278] Santorio Santorio. Book review: La medicina statica, Edited by Giuseppe Ongaro, 2001, 187 pp., Giunti, Firenze. Nuncius, 17(1):392, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00892.

Cavalieri:2002:BRS

[2279] Bonaventura Cavalieri. Book review: Lo specchio ustorio, Edited by Enrico Giusti, 2001, 187 pp., Giunti, Firenze. Nuncius, 17(1):392, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00900.

Neri:2002:BRL

[2280] Antonio Neri. Book review: L’arte vetraria, Edited by Ferdinando Abbri, 2001, 127 pp., Giunti, Firenze. Nuncius, 17(1):393, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00919.

Toaldo:2002:BRC

[2281] Giuseppe Toaldo. Book review: Dei conduttori per preservare gli edifizj da’ fulmini, Edited by Stefano Casati, 2001, 229 pp., Giunti, Firenze. Nuncius, 17(1):393, ???? 2002. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539102x00928.

Mosso:2002:BRF

[2282] Angelo Mosso. Book review: La fatica, Edited by Michele Nani, 2001, 291 pp. Giunti, Firenze. Nuncius, 17(1):394, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539102x00937.

Cattaneo:2002:BRF

[2283] Carlo Cattaneo and Giovanni Milani. Book review: Ferdinandea. Scritti sulla ferrovia da Venezia a Milano 1836–1841, Edited by Pietro Redondi 2001, 2 voll., 542 pp., Giunti, Firenze. Nuncius, 17(1):394–395, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00946.

Borelli:2002:BRS

[2284] Giovanni Alfonso Borelli. Book review: Storia e meteorologia dell’eruzione dell’Etna del 1669, Edited by Nicoletta Morello, 2001, 270 pp., Giunti, Firenze. Nuncius, 17(1):395–396, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539102x00955.

Rolando:2002:BRS

[2285] Luigi Rolando. Book review: Saggio sopra la vera struttura del cervello dell’uomo e degli animali e spora le funzioni del sistema nervoso, Edited by Alessandro Dini, 2001, 93 pp. Giunti, Firenze. Nuncius, 17(1):396, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00964.

Rosa:2002:BRO

[2286] Daniele Rosa. Book review: Ologenesi, Edited by Antonello La Vergata, 2001, 446 pp., Giunti, Firenze. Nuncius, 17(1):396–397, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00973.

Luciani:2002:BRL

[2287] Luigi Luciani and Giuseppe Seppilli. Book review: Le localizzazioni funzionali del cervello, Edited by Carmela Morabito, 2001, 325 pp., Giunti, Firenze. Nuncius, 17(1):397, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539102x00982.

Genovesi:2002:BRD

[2288] Antonio Genovesi. Book review: Dissertatio physico-historica de rerum origine et constitutione, Edited by Sara Bonechi and Maurizio Torrini, 2001, 284 pp., Giunti, Firenze. Nuncius, 17(1):398–400, ????

2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00991.

Torrini:2002:LG

[2289] Maurizio Torrini. Letture Galileiane. Nuncius, 17(2):409–422, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00018.

Castellani:2002:LS

[2290] Vittorio Castellani. L’astrofisica stellare. Nuncius, 17(2):423–446, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00027.

Meskens:2002:SNB

[2291] Ad Meskens. Some new biographical data about Michiel Coignet. Nuncius, 17(2):447–454, ???? 2002. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539102x00036.

Zik:2002:GOA

[2292] Yaakov Zik. Galileo and optical aberrations. Nuncius, 17(2):455–465, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00045.

Bnoli:2002:TOM

[2293] Fabrizio Bnoli, Mara Miniati, Vincenzo Greco, and Giuseppe Molesini. Telescope optics of Montanari, Cellio, Campani and Bruni at the Museo della Specola in Bologna. Nuncius, 17(2):467–474, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00054.

Talas:2002:TEC

[2294] Sofia Talas. Thermometers in the Eighteenth Century: J. B. Micheli Du Crest’s works and the cooperation with G. F. Brander. Nuncius, 17 (2):475–496, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00063.

Brenni:2002:CSI

[2295] Paolo Brenni. 19th century scientific instrument advertising. Nuncius, 17(2):497–513, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00072.

Lombardi:2002:CWL

[2296] Anna Maria Lombardi. The constitution of white light, a source for the new questions of quantum mechanics. Nuncius, 17(2):515–545, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00081.

Arecco:2002:IF

[2297] Davide Arecco. Istituzioni e fonti. Nuncius, 17(2):547–565, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00090.

Casati:2002:CDT

[2298] Stefano Casati. La collezione delle tesi di medicina sostenute a Parigi dal 1798 al 1860. Nuncius, 17(2):567–579, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539102x00108.

Anonymous:2002:SIC

[2299] Anonymous. Scientific instrument commission. Nuncius, 17(2):581, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00117.

Bigg:2002:SH

[2300] Charlotte Bigg and Klaus Staubermann. Spectroscope histories. Nuncius, 17(2):583–587, ???? 2002. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539102x00126.

Hentschel:2002:SS

[2301] Klaus Hentschel. Spectroscopy or spectroscopies? Nuncius, 17(2):589– 614, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00135.

Aubin:2002:OOL

[2302] David Aubin. Orchestrating observatory, laboratory, and field: Jules Janssen, the spectroscope, and travel. Nuncius, 17(2):615–633, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00144.

Gamble:2002:ACS

[2303] Susan Gamble. An appealing case of spectra: Photographs on display at the Royal society, London 1891. Nuncius, 17(2):635–651, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00153.

Widmalm:2002:QSK

[2304] Sven Widmalm. Quantifying sunshine: Knut Ångstrom and the standardisation of pyrheliometry. Nuncius, 17(2):653–671, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00162.

Freiburger:2002:BJR

[2305] Dana A. Freiburger. Building a Japanese research tradition in physics: Hantaro Nagaoka and the spectroscope. Nuncius, 17(2):673–689, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00171.

Parrini:2002:FSE

[2306] Paolo Parrini. Filosofia e scienza in Enriques. Nuncius, 17(2):691–701, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00180.

Miniati:2002:BRI

[2307] Mara Miniati. Book review: Isaia Iannaccone, Storia e civilt della Cina. Cinque lezioni, Napoli, Libreria Dante Descartes, 2000, 145 pp. Nuncius, 17(2):703, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00199.

Grasso:2002:BRD

[2308] Roberto Grasso. Book review: Dag Nikolaus Hasse, Avicenna’s De anima in the Latin West. The formation of a peripatetic philosophy of the soul, London/Turin, The Warburg Institute, Nino Aragno editore, 2000, 350 pp. Lire 63.000, ISBN 0-85481-125-7 1471-5841. Nuncius, 17(2):703–707, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00207.

Calizzi:2002:BRR

[2309] Chiara Calizzi. Book review: Révolution scientifique et libertinage, études réunies par A. Mothu avec la collaboration d’A. Del Prete, Tumhout, Brepols, 2000, 312 pp. ISBN 2-503-51008-6. Nuncius, 17(2): 707–709, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00216.

Miniati:2002:BRG

[2310] Mara Miniati. Book review: Gerard l’E. Turner, Elizabethan instrument makers. The origins of the London trade in precision instrument making, Oxford, Oxford University Press, 2000, 290 pp., ill. £79.50, ISBN 0-19-856566-6. Nuncius, 17(2):710–711, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539102x00225.

Ciardi:2002:BRS

[2311] Marco Ciardi. Book review: Scott Mandelbrote, Footprints of the lion: Isaac Newton at work. Exhibition at Cambridge University Library (9 October 2001–23 March 2002), Cambridge, Cambridge University Library, 2001, 142 pp., 42 ill. ISBN 0-902205-58-7. Nuncius, 17(2):711– 712, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00234.

Abbri:2002:BRRb

[2312] Ferdinando Abbri. Book review: Renate Wilson, Pious Traders in Medicine. A German Pharmaceutical Network in Eighteenth-Century North America, University Park, Pennsylvania, The Pennsylvania State University Press, 2000, pp. xiv + 258. $37.50 cloth ISBN 0-271-02052-0. Nuncius, 17(2):712–713, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539102x00243.

Burdet:2002:BRGa

[2313] Carlo A. M. Burdet. Book review: Gianni Losano e Giuseppe Slaviero (a cura di) Michele Buniva, Memoria al figlio unigenito, Torino, Università degli Studi di Torino, Archivio Scientifico e Tecnologico dell’Università di Torino, 2000, pp. 166, ill. (Strumenti per la storia della scienza subalpina, [s. n., ma n. 1]). Nuncius, 17(2):714–715, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00252.

Burdet:2002:BRGb

[2314] Carlo A. M. Burdet. Book review: Gabriella Ballesio (a cura di), Carte Buniva, Inventario, Torino, Archivio Storico della Citt di Torino, 2000, pp. 127. 15,49, ISBN 88-86685-45-9. Nuncius, 17(2):715–717, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00261.

Ciardi:2002:BRL

[2315] Marco Ciardi. Book review: Luigi Blanco (a cura di), Amministrazione, formazione e professione: gli ingegneri in Italia tra Sette e Ottocento, Bologna, Il Mulino, 2000, pp. 538 pp. 30,99, ISBN 88-15-07665-4 (Annali dell’Istituto storico italo-germanico in Trento, Quaderni, 52). Nuncius, 17(2):717–718, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00270.

Rasini:2002:BRS

[2316] Vallori Rasini. Book review: Salvatore Lazzara, Claude Bernard. Tra Scienza e filosofia, Napoli, Cuen, 1999, pp. 85. 7,23, ISBN 88-7146-437-0. Nuncius, 17(2):718–719, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539102x00289.

Miniati:2002:BRF

[2317] Mara Miniati. Book review: Franco Soresini, Segnali sopra e sotto l’Atlantico, La Spezia, s.n.t., 2001, 79 pp., ill. Nuncius, 17(2):720, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00298.

Abbri:2002:BRY

[2318] Ferdinando Abbri. Book review: Yvette Conry, L’évolution créatrice d’Henry Bergson. Investigations critiques. Préface de François Dagognet, Paris, l’Harmattan, 2000, 328 pp. ISBN 2-7384-9903-1. Nuncius, 17(2): 721–723, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00306.

Miniati:2002:BRS

[2319] Mara Miniati. Book review: Salvo D’Agostino, Arcangelo Rossi (a cura di), Enrico Fermi e l’Enciclopedia Italiana, Roma, Istituto della Enciclopedia Italiana, 2001, 191 pp., ill. Nuncius, 17(2):723–724, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00315.

Bertini:2002:BRR

[2320] Silvia Bertini. Book review: Riccardo Luccio, con la collaborazione di Simonetta Gori Savellini, La psicologia: un profilo storico, Roma/Bari, Laterza, 2000, 307 pp., 19,63, ISBN 88-420-6139-5. Nuncius, 17(2):724– 726, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00324.

Ciardi:2002:BRM

[2321] Marco Ciardi. Book review: Mario Neve, Virtus loci. Lineamenti fondamentali di una teoria dell’informazione spaziale, Urbino, QuattroVenti, 1999, 439 pp., 68 ill. 24,8, ISBN 88-392-0518-7. Nuncius, 17(2):726–727, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00333.

Visconti:2002:BRP

[2322] Agnese Visconti. Book review: Peter Davis, Musei e ambiente naturale. Il ruolo del museo di storia naturale nella conservazione della biodiversità, Bologna, Clueb, 2001, xv + 350 pp. 28,41, ISBN 88-491-1284-X. Nuncius, 17(2):727–728, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00342.

Iluminati:2002:BRF

[2323] Augusto Iluminati. Book review: Felice Cimatti, La mente silenziosa. Come pensano gli animali non umani, Roma 2002, Editori Riuniti, pp. 239. 12,00, ISBN 88-359-5160-7. Nuncius, 17(2):729–732, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00351.

Toscano:2002:BRD

[2324] Anna Toscano. Book review: Daniel J. Siegel, La mente relazionale. Neurobiologia dell’esperienza interpersonale, tr. it. a cura di Luisa Madeddu, Milano, Raffaello Cortina Editore, 2001, pp. 362 + xiv, 32,54, ISBN 88-7078-693-5. Nuncius, 17(2):732–736, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00360.

Anonymous:2002:IB

[2325] Anonymous. The IMSS bookshelf. Nuncius, 17(2):737, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00379.

Anonymous:2002:BRDa

[2326] Anonymous. Book review: 42. Domenico Laurenza, De Figura Umana: Fisiognomia, Anatomia e Arte in Leonardo 2001, 241 pp., Olschki, Firenze. Nuncius, 17(2):737–738, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539102x00388.

Anonymous:2002:BRS

[2327] Anonymous. Book review: 23. Santorio Santorio, La medicina statica. Edited by Giuseppe Ongaro 2001, 187 pp., Giunti, Firenze. Nuncius, 17(2):738, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00397.

Anonymous:2002:BRB

[2328] Anonymous. Book review: 24. Bonaventura Cavalieri, Lo specchio ustorio, Edited by Enrico Giusti 2001, 187 pp., Giunti, Firenze. Nuncius, 17(2):738, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00405.

Anonymous:2002:BRAa

[2329] Anonymous. Book review: 25. Antonio Neri l’arte vetraria. Edited by Ferdinando Abbri 2001, 127 pp., Giunti, Firenze. Nuncius, 17(2):739, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00414.

Anonymous:2002:BRGa

[2330] Anonymous. Book review: 26. Giuseppe Toaldo Dei conduttori per preservare gli edifizj da’ fulmini. (Italian) [The conduits for preserving buildings from lightning]. Edited by Stefano Casati 2001, 229 pp., Giunti, Firenze. Nuncius, 17(2):739, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539102x00423.

Anonymous:2002:BRAb

[2331] Anonymous. Book review: 28. Angelo Mosso La fatica. Edited by Michele Nani 2001, 291 pp. Giunti, Firenze. Nuncius, 17(2):740, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00432.

Anonymous:2002:BRC

[2332] Anonymous. Book review: 29. Carlo Cattaneo/Giovanni Milani Ferdinandea. Scritti sulla ferrovia da Venezia a Milano 1836–1841, (Italian) [Writings on the railway from Venice to Milan 1836–1841]. Edited by Pietro Redondi 2001, 2 voll., 542 pp., Giunti, Firenze. Nuncius, 17(2): 740–741, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539102x00441.

Anonymous:2002:BRGb

[2333] Anonymous. Book review: 30. Giovanni Alfonso Borelli, Storia e meteorologia dell’eruzione dell’Etna del 1669. Edited by Nicoletta Morello 2001, 270 pp., Giunti, Firenze. Nuncius, 17(2):741–742, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00450.

Anonymous:2002:BRLa

[2334] Anonymous. Book review: 31. Luigi Rolando: Saggio sopra la vera struttura del cervello dell’uomo e degli animali e spora le funzioni del sistema nervoso. Edited by Alessandro Dini. 2001, 93 pp. Giunti, Firenze. Nuncius, 17(2):742, ???? 2002. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539102x00469.

Anonymous:2002:BRDb

[2335] Anonymous. Book review: 32. Daniele Rosa Ologenesi. Edited by Antonello La Vergata 2001, 446 pp., Giunti, Firenze. Nuncius, 17(2):742– 743, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00478.

Anonymous:2002:BRLb

[2336] Anonymous. Book review: 33. Luigi Luciani/Giuseppe Seppilli, Le localizzazioni funzionali del cervello. Edited by Carmela Morabito 2001, 325 pp., Giunti, Firenze. Nuncius, 17(2):743, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00487.

Anonymous:2002:BRAc

[2337] Anonymous. Book review: 34. Antonio Genovesi, Dissertatio physicohistorica de rerum origine et constitutione. Edited by Sara Bonechi and Maurizio Torrini 2001, 284 pp., Giunti, Firenze. Nuncius, 17(2):744, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00496.

Anonymous:2002:BRLc

[2338] Anonymous. Book review: L’arte della sperimentazione nell’Accademia Galileiana del Cimento (1656–1667), (Italian) [The art of experimentation in the Galilean Academy of Chemistry]. Edited by Paolo Galluzzi. 2001, 159 pp., Sillabe, Livorno Catalogue of the Exhibition Held at the Uffizi, Florence (From March 18 2001). Nuncius, 17(2):744–745, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00504.

Anonymous:2002:BRN

[2339] Anonymous. Book review: Nel segno di Masaccio. l’invenzione della prospettiva. (Italian) [In the sign of Masaccio. The invention of perspective]. Edited by Filippo Camerota, 2001, 311 pp., Giunti, Firenze. Nuncius, 17(2):745–746, ???? 2002. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539102x00513.

Anonymous:2002:IAX

[2340] Anonymous. Indici (anno XVII, 2002). Nuncius, 17(2):747–749, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00522.

Anonymous:2002:IM

[2341] Anonymous. Indice dei manoscritti. Nuncius, 17(2):751–754, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00531.

Anonymous:2002:IN

[2342] Anonymous. Indice dei nomi. Nuncius, 17(2):755–781, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00540.

Anonymous:2002:ISI

[2343] Anonymous. Indice dei siti Internet. Nuncius, 17(2):783, ???? 2002. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539102x00559.

Ciardi:2003:AAa

[2344] Marco Ciardi. Annuncio/announcement. Nuncius, 18(1):1, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00521.

Rosen:2003:DMP

[2345] Mark Rosen. Don Miniato Pitti and the second life of a scientist’s tools in Cinquecento Florence. Nuncius, 18(1):3–24, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00530.

Camerota:2003:TNA

[2346] Filippo Camerota. Two new attributions: a refractive dial of Guidobaldo Del Monte and the Roverino compass of Fabrizio Mordente. Nuncius, 18 (1):25–38, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00549.

Dupre:2003:DRD

[2347] Sven Dupre. The dioptrics of refractive dials in the sixteenth century. Nuncius, 18(1):39–67, ???? 2003. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539103x00558.

Donato:2003:LDP

[2348] Maria Pia Donato. L’onere della prova. Il Sant’Uffizio, l’atomismo e i medici romani. Nuncius, 18(1):69–87, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539103x00567.

Monti:2003:LSN

[2349] Maria Teresa Monti. I lombrichi di Spallanzani. notomi e inedite contro I garbugli e pasticci dell’argomento di analogia. Nuncius, 18(1):89–125, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00576.

Bandinelli:2003:LLA

[2350] Angela Bandinelli. 1783 — Lavoisier and Laplace: Another crucial year. Antiphlogistic chemistry and the investigation on Living beings between the Eighteenth and the Nineteenth Centuries. Nuncius, 18(1):127–139, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00585.

Babini:2003:PLD

[2351] Valeria P. Babini. Paola Lombroso, una donna nelle scienze dell’uomo. Nuncius, 18(1):141–165, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00594.

Minelli:2003:LR

[2352] Alessandro Minelli. L’omologia rivisitata. Nuncius, 18(1):167–200, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00602.

Meschiari:2003:PAD

[2353] Alberto Meschiari. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 18(1):201–242, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00611.

Anonymous:2003:DLW

[2354] Anonymous. Le due lettere di W. H. Fox Talbot alla Literary Gazette presentate da G. B. Amici all’Accademia Dei Georgofili in Firenze il 13 febbraio 1842. (Italian) [The two letters of W. H. Fox Talbot presented from the Literary Gazette from G. B. Amici of the Academy of Georgofili in Florence on February 13 1842]. Nuncius, 18(1):243–247, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00620.

Bino:2003:IF

[2355] Carla Bino. Istituzioni e fonti. Nuncius, 18(1):249–268, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00639.

Carnemolla:2003:PNN

[2356] Stefania Elena Carnemolla. Pedro Nuñes, matematico e cosmografo portoghese del XVI Secolo, e la sua Defesa Do Tratado Da Rumao Do Globo Para a Arte De Navegar. (Italian) [Pedro Nuñes, Portuguese mathematician and cartographer of the sixteenth century, and his Defense of the Global Treaty of Rome for the Art of Navigating]. Nuncius, 18(1):269– 286, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00648.

Nunes:2003:DDT

[2357] Pedro Nunes. Defenso do tratado da rumao do globo para a arte de navegar. (Portuguese) [Defense of the Global Treaty of Rome for the art of navigating]. Nuncius, 18(1):287–317, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00657.

Meli:2003:R

[2358] Domenico Bertoloni Meli. Ricordo. Nuncius, 18(1):319–325, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00666.

Caroti:2003:PS

[2359] Stefano Caroti. Pierre Souffrin (1935–2002). Nuncius, 18(1):327–329, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00675.

Raynaud:2003:LPM

[2360] Dominique Raynaud. Linear perspective in Masaccio’s Trinity fresco: Demonstration or self-persuasion? Nuncius, 18(1):331–344, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00684.

Ferrari:2003:LES

[2361] Graziano Ferrari. Letters in the Earth sciences: Their historic value and present-day scientific relevance. Nuncius, 18(1):345–354, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00693.

LaVergata:2003:BRD

[2362] Antonello La Vergata. Book review: David Millar, Ian Millar, John Millar, Margaret Millar, The Cambridge Dictionary of Scientists, second edition, Cambridge, Cambridge University Press, 2002, xii + 428 pp., ISBN 0-521-80602-X (paperback), 14.95 (US$50.00). Nuncius, 18(1): 355–357, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00701.

Miniati:2003:BRV

[2363] Mara Miniati. Book review: Viaggio nelle aree sismiche Liguria, basso Piemonte, Toscana, Emilia Romagna coste e Appennino dal 91 a.C. al 2000, (Italian) [Travel in seismic areas of Liguria, southern Piedmont, Tuscany, Emilia Romagna coast and Apennines from 91 A.D. to 2000]. Roma, Servizio Sismico Nazionale-Ufficio Formazione e Informazione, 2001, 494, pp., ill. (+DVD ROM). Nuncius, 18(1):357–358, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00710.

Bertini:2003:BRR

[2364] Silvia Bertini. Book review: Rudolph M. Bell, La santa anoressia. Digiuno e misticismo dal Medioevo ad oggi. (Italian) [Holy anorexia.

Fasting and mysticism from the Middle Ages to today], Editori Laterza, 2002, 281 pp., 18 ill., 8.00, ISBN 88-420-5629-4 (Economica 155). Nuncius, 18(1):358–359, ???? 2003. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539103x00729.

Lippi:2003:BRM

[2365] Donatella Lippi. Book review: Martin Dinges ed., Patients in the History of Homoeopathy, Sheffield 2002, 434 pp., ISBN 0-9536522-4-6. Nuncius, 18(1):359–361, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00738.

Pogliano:2003:BRB

[2366] Claudio Pogliano. Book review: Les Back and John Solomos (eds.), Theories of Race and Racism. A Reader, London/New York, Routledge, 2002, 646 pp., ISBN 0-415-15672-6. Nuncius, 18(1):361–362, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00747.

Visconti:2003:BRA

[2367] Agnese Visconti. Book review: Annamaria Galbani (a cura di), Donne politecniche, Milano, Scheiwiller, 2001, 198 pp., 13,43 ISBN 88-7644-298-7. Nuncius, 18(1):362–363, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00756.

Miniati:2003:BRLa

[2368] Mara Miniati. Book review: Lucia Fornari Schianchi (a cura di), l’Antica Spezieria di San Giovanni Evangelista in Parma, Reggio Emilia, Saetti Maestri, Fondazione Banca Monte di Parma, 2001, 277 pp., ill. Nuncius, 18(1):363–365, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00765.

Miniati:2003:BRK

[2369] Mara Miniati. Book review: Koenraad Van Cleempoel, A ‘catalogue raisonné’ of scientific instruments from the Louvain school, 1530 to 1600, Turnhout, Brepols, 2002, 284 pp., ill., ISBN 2-503-51218-6 (De diversis artibus t. 65). Nuncius, 18(1):365–366, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00774.

Abbri:2003:BRG

[2370] Ferdinando Abbri. Book review: Giovan Battista della Porta, Teatro. Tomo Secondo e Terzo — Commedie. A cura di Raffaele Sirri, Napoli, Edizioni Scientifiche Italiane, 2002, pp. xxvi + 550, pp. viii + 516, ISBN 88-495-0256-7, ISBN 88-495-0414-4 (Edizione Nazionale delle opere di Giovan Battista della Porta, 15-II-III), 42,00 ciascuno. Nuncius, 18(1): 366–368, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00783.

Ciancio:2003:BRD

[2371] Luca Ciancio. Book review: David Freedberg, The Eye of the Lynx. Galileo, His Friends, and the Beginnings of Modern Natural History, Chicago, The University of Chicago Press, 2002, 528 pp., 83 ill. a colori, 89 b/n, $50.00, ISBN 0-226-26147-6. Nuncius, 18(1):368–373, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00792.

Sciaccaluga:2003:BRB

[2372] Nicoletta Sciaccaluga. Book review: Benedino Gemelli, Isaac Beeckman atomista e lettore di Lucrezio, Firenze, Olschki, 2002, xiii + 132 pp., 16,00, ISBN 88-222-5075-3 (Subsidia, 1). Nuncius, 18(1):373–376, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00800.

Lippi:2003:BRL

[2373] Donatella Lippi. Book review: Luigi Guerrini, Antonio Cocchi naturalista e filosofo, Firenze, Polistampa 2002, pp. 339, 12,39, ISBN 88-8304-398-7. Nuncius, 18(1):376–377, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00819.

Rasini:2003:BRI

[2374] Vallori Rasini. Book review: Immanuel Kant, Bemerkungen. Note per un diario filosofico, a cura di Katrin Tenenbaum, Roma, Meltemi, 2001, 272 pp., 20,14, ISBN 88-8353-111-6. Nuncius, 18(1):378, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00828.

Lippi:2003:BRO

[2375] Donatella Lippi. Book review: Othmar Keel, L’avènement de la médecine clinique moderne en Europe, 1750–1815, Montréal, Les Presses de l’Université de Montréal. Georg éditeur, 2001, 542 pp., $59,95, ISBN 2-7606-1822-6. Nuncius, 18(1):378–381, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00837.

Burdet:2003:BRG

[2376] Carlo A. M. Burdet. Book review: Giuseppe Slaviero (a cura di) Michele Buniva, introduttore della vaccinazione in Piemonte, Scienza e sanit tra rivoluzione e restaurazione, Atti del Convegno di studi, Pinerolo, 14 ottobre 2000, Torino, Università degli Studi di Torino, Archivio Scientifico e Tecnologico dell’Università di Torino, 2002, 230 pp. (Strumenti per la storia della scienza subalpina, 2). Nuncius, 18(1):381–384, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00846.

Govoni:2003:BRK

[2377] Paola Govoni. Book review: Kathryn A. Neeley, Mary Somerville: Science, Illumination, and the Female Mind, Cambridge, Cambridge University Press, 2001, xvi + 263 pp., ISBN 0-521-62672-2 (Cambridge Science Biographies). Nuncius, 18(1):384–386, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00855.

Scarpelli:2003:BRA

[2378] Giacomo Scarpelli. Book review: Annabel Robinson, The Life and Work of Jane Ellen Harrison, Oxford, Oxford University Press, 2002, xvi + 332 pp., 17 ill., £45.000, ISBN 0-19-924233-X. Nuncius, 18(1):386–388, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00864.

DelSanto:2003:BRS

[2379] Paolo Del Santo. Book review: Sabino Maffeo, S.J., The Vatican Observatory: in the Service of Nine Popes, Roma, Vatican Observatory Pubblications, 2001, pp. xix-429, ISBN 88-209-7242-5. Nuncius, 18(1): 388–391, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00873.

Polizzi:2003:BRO

[2380] Gaspare Polizzi. Book review: Ornella Pompeo Faracovi e Luca Maria Scarantino (a cura di), Federigo Enriques. Matematiche e filosofia. Lettere inedite. Bibliografia degli scritti, Livorno, Belforte C. Editori, 2001, pp. 211, 18.07 (L 35.000) ISBN 88-7997-063-1. Nuncius, 18(1):391–394, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00882.

Ciardi:2003:BRK

[2381] Marco Ciardi. Book review: Kenneth Beauchamp, A history of telegraphy: its technology and application, London, The Institution of Electrical Engineers, 2001, 408 pp., ill., 55.00, ISBN 0-85296-792-6 (History of Technology Series, 26). Russell W. Burns, John Logie Baird: TV pioneer, London, The Institution of Electrical Engineers, 2001, 55.00, ISBN 0-85296-797-7 (History of Technology Series, 28). Nuncius, 18(1):394– 395, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00891.

Pogliano:2003:BRJ

[2382] Claudio Pogliano. Book review: Jonathan Marks, What It Means to Be 98 Chimpanzee. Apes, People, and Their Genes, Berkeley, University of California Press, 2002, pp. 314 (ISBN 0-520-24064-2). Nuncius, 18 (1):395–398, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00909.

Contardi:2003:BRM

[2383] Simone Contardi. Book review: Michela Nacci, Pensare la tecnica. Un secolo di incomprensioni, presentazione di Gianni Vattimo, Roma/Bari, Laterza, 2000, pp. 344, ISBN 88-420-5973-0 (Biblioteca di cultura moderna). Nuncius, 18(1):398–399, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00918.

Rasini:2003:BRH

[2384] Vallori Rasini. Book review: Hans Jonas, Sull’orlo dell’abisso. Conversazioni sul rapporto tra uomo e natura, a cura di Paolo Becchi, Torino, Einaudi, 2000, 150 pp., 11.36; ISBN 88-06-15051-0. Nuncius, 18(1):399– 400, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00927.

Anonymous:2003:IB

[2385] Anonymous. The IMSS bookshelf. Nuncius, 18(1):401, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00936.

Anonymous:2003:BRS

[2386] Anonymous. Book review: 43. Simone Contardi La casa di Salomone a Firenze. l’Imperiale e Reale Museo di fisica e storia naturale (1775– 1801) 2002, xx + 324 pp. con 8 taw. f.t., Olschki, Firenze. Nuncius, 18(1):401, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00945.

Anonymous:2003:BRGa

[2387] Anonymous. Book review: 44. Giambattista Riccioli e il merito scientifico dei Gesuiti nell’et barocca. A cura di Maria Teresa Borgato 2002, xviii-484 pp. con 29 figg. n.t., Olschki, Firenze. Nuncius, 18(1):401–402, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00954.

Anonymous:2003:BRGb

[2388] Anonymous. Book review: 45. Guido Alliney, Time and Soul in Fourteenth Century Theology. Three Questions of William of Alnwick on the Existence, the Ontological Status and the Unity of Time 2002, lxiv + 194 pp., Olschki, Firenze. Nuncius, 18(1):402–403, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00963.

Anonymous:2003:BRGc

[2389] Anonymous. Book review: 46. G. B. Hodierna e il Secolo Cristallino. Atti del convegno di Ragusa (22–24 ottobre 1997). A cura di Mario Pavone e Maurizio Torrini 2002, x-346 pp. con 3 figg. e 2 taw. n.t., Olschki, Firenze. Nuncius, 18(1):403, ???? 2003. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539103x00972.

Ciardi:2003:AAb

[2390] Marco Ciardi. Annuncio/announcement. Nuncius, 18(2):iii–410, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00017.

Zucchi:2003:BFL

[2391] Luca Zucchi. Brunfels e Fuchs: l’illustrazione botanica quale ritratto della singola pianta o immagine della specie. (Italian) [Brunfels and Fuchs: Botanical illustration: Portrait of the plant or single image of species?]. Nuncius, 18(2):411–465, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00026.

Levi-Donati:2003:FHT

[2392] Gemma Rosa Levi-Donati. Four hundred and twenty-five years later: Egnazio Danti’s anemoscope. Nuncius, 18(2):467–479, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00035.

Zik:2003:KT

[2393] Yaakov Zik. Kepler and the telescope. Nuncius, 18(2):481–514, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00044.

Sergio:2003:BHL

[2394] Emilio Sergio. Bacon, Hobbes e l’idea di ≪philosophia naturalis≫: Due modelli di riforma della scienza. Nuncius, 18(2):515–547, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00053.

Giudice:2003:LMA

[2395] Franco Giudice and Gianni Bonera. Lorenzo Mascheroni e Alessandro Volta: un dialogo tra scienza e poesia. Nuncius, 18(2):549–576, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00062.

Maison:2003:OIE

[2396] Laetitia Maison. Les observatoires italiens en 1875: un exemple pour le renouveau de l’astronomie française?. (French) [The Italian observatories in 1875: an example for the renewal of French astronomy?]. Nuncius, 18 (2):577–602, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00071.

Milighetti:2003:PAD

[2397] Maria Chiara Milighetti. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 18(2):603–617, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539103x00080.

DeCarolis:2003:CIT

[2398] Stefano De Carolis, Conti Andrea A., and Donatella Lippi. Un carteggio inedito tra Antonio Cocchi e Giovanni Bianchi. Nuncius, 18(2):619–636, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00099.

Scalva:2003:VSA

[2399] Giuse Scalva. Un viaggio scientifico alla metà del Secolo XVIII. Inventario del carteggio intorno al viaggio in Levante di Vitaliano Donati. Nuncius, 18(2):637–690, ???? 2003. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539103x00107.

Ciancio:2003:SCA

[2400] Luca Ciancio. Secondo contributo all’inventario del carteggio di Alberto Fortis. (Italian) [Second contribution to the inventory of correspondence of Alberto Fortis]. Nuncius, 18(2):691–703, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00116.

Pelucchi:2003:HP

[2401] Stéphane Pelucchi. Histoires Parallèles. Nuncius, 18(2):705–735, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00125.

Bigg:2003:I

[2402] Charlotte Bigg and Klaus Staubermann. Introduction. Nuncius, 18(2): 737–739, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00134.

Hennig:2003:BKS

[2403] Jochen Hennig. Bunsen, Kirchhoff, Steinheil and the elaboration of analytical spectroscopy. Nuncius, 18(2):741–754, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539103x00143.

Staubermann:2003:IVR

[2404] Klaus Staubermann. Investigating vision and the reversion spectroscope: Early astronomical colour studies in experimental psychology. Nuncius, 18(2):755–764, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00152.

Bigg:2003:SM

[2405] Charlotte Bigg. Spectroscopic metrologies. Nuncius, 18(2):765–777, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00161.

Staley:2003:ISM

[2406] Richard Staley. The interferometer and the spectroscope: Michelson’s standards and the spectroscopic community. Nuncius, 18(2):779–801, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00170.

Johnston:2003:UTM

[2407] Sean F. Johnston. An unconvincing transformation? Michelson’s interferential spectroscopy. Nuncius, 18(2):803–823, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00189.

Lombardi:2003:BSB

[2408] Anna M. Lombardi. The bolometer and the spectro-bolometer as steps towards the black-body spectrum. Nuncius, 18(2):825–840, ????

2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00198.

Mayring:2003:SHS

[2409] Eva A. Mayring. Sources for the history of spectroscopy at the deutsches museum. Nuncius, 18(2):841–852, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00206.

Galluzzi:2003:R

[2410] Paolo Galluzzi. Ricordo. Nuncius, 18(2):853–857, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539103x00215.

Ballerini:2003:SF

[2411] Arnaldo Ballerini. Storia e follia. Nuncius, 18(2):859–866, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00224.

Pogliano:2003:BRM

[2412] Claudio Pogliano. Book review: Mario Blagioli & Peter Galison (eds.), Scientific Authorship. Credit and Intellectual Property in Science, New York and London, Routledge, 2003, 384 pp., ISBN 0-415-94293-4. Nuncius, 18(2):867–868, ???? 2003. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539103x00233.

Lippi:2003:BRA

[2413] Donatella Lippi. Book review: Alessandra Parodi, Storie della Medicina, Torino, Edizioni di Comunita, 2002, pp. 309, EUR22,00. ISBN 88-245-0639-9. Nuncius, 18(2):868–869, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00242.

Miniati:2003:BRG

[2414] Mara Miniati. Book review: Gualtiero Bellucci, Gigliola Terenna, Francesca Vannozzi (a cura di), La collezione degli strumenti di anestesiologia, Siena, Nuova immagine, 2001, 186 pp., ill., ISBN 88-7145-180, EUR9,29 (Patrimonio storico-scientifico dell’Università degli studi di Siena. Materiali; 5) Gigliola Terenna, Francesca Vannozzi (a cura di), La collezione degli strumenti di ginecologia, ostetricia e pediatria, Siena, Nuova immagine, 2002, 223 pp., ill., ISBN 88-7145-188-0, EUR9,00 (Patrimonio storico-scientifico dell’Università degli studi di Siena. Materiali; 6). Nuncius, 18(2):870–871, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00251.

Miniati:2003:BRE

[2415] Mara Miniati. Book review: Erasmo Recami, Sergio Pizzigalli, Ettore Parigi, Francesco De Vincentiis, Virgilio Borlotti, Gli strumenti scientifici di interesse storico del ≪Lussana≫, del ≪Vittorio Emanuele≫, e del ≪Quarenghi≫ di Bergamo, Bergamo, Dipartimento di Ingegneria Università degli Studi di Bergamo, 2002, 79 pp., ill., ISBN e prezzo non indicati. Nuncius, 18(2):871–872, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00260.

diPasquale:2003:BRC

[2416] Giovanni di Pasquale. Book review: Carlo Santini (direzione e coordinamento), Ida Mastrorosa, Antonino Zumbo (a cura di), Letteratura scientifica e tecnica di Grecia e Roma, Roma, Carocci, 2003, 604 pp., ISBN 88-430-2142-7, EUR40,00. Nuncius, 18(2):872–874, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00279.

diPasquale:2003:BRG

[2417] Giovanni di Pasquale. Book review: Georgia Irby Massie, Paul Keyser, Greek Science of the Hellenistic Era, London-New York, Routledge, 2002, 392 pp., ISBN 0-415-23847-1. Nuncius, 18(2):874–876, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00288.

Acerbi:2003:BRA

[2418] Fabio Acerbi. Book review: Aeaomena, Euclid’s Data or the Importance of Being Given. The Greek Text translated and explained by Christian Marinus Taisbak. Acta Historica Scientiarum Naturalium et Medicinalium, vol. 45. Copenhagen, Museum Tusculanum Press, 2003, 271 pp., EUR48,00, ISBN 87-7289-815-1. Nuncius, 18(2):876–879, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00297.

Laurenza:2003:BRJ

[2419] Domenico Laurenza. Book review: Jole Agrimi, Ingeniosa scientia nature. Studi sulla fisiognomica medievale, Firenze, Sismel, Edizioni del Galluzzo, 2002, 180 pp., ISBN 88-8450-074-5, EUR34,00. Nuncius, 18 (2):879–881, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00305.

Abbri:2003:BRC

[2420] Ferdinando Abbri. Book review: Chiara Crisciani, Il papa e l’alchimia. Felice V, Guglielmo Fabri e l’elixir, Roma, Viella, 2002, 217 pp., ISBN 88-8334-079-5, 19,00 (La corte dei papi, 10). Nuncius, 18(2):881–883, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00314.

Abbri:2003:BRA

[2421] Ferdinando Abbri. Book review: Andrea Ubrizsy Savoia, Rapporti Italo-Ungheresi nella nascita della botanica in Ungheria, Pecs, Bornus Nyomda-Roma, Università degli Studi ≪La Sapienza≫, 2002, 413 pp., ill., ISBN 963-641-906-X. Nuncius, 18(2):883–885, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00323.

Toscano:2003:BRM

[2422] Anna Toscano. Book review: Michel Blay, La science du mouvement. De Galilée à Lagrange, Paris, Éditions Berlin, 2002, 332 pp. + ill., ISBN 2-7011-3076-X (≪Collection Berlin Sup Histoire des Sciences≫). Nuncius, 18(2):885–887, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00332.

Miniati:2003:BRLb

[2423] Mara Miniati. Book review: Lewis Pyenson, Jean–François Gauvin (eds.), The art of teaching physics. The Eighteenth-Century demonstration apparatus of Jean Antoine Nollet, Sillery (Québec), Les éditions du Septentrion, 2002, 221 pp., ill., ISBN 2-89448-320-1, senza indicazione di prezzo. Nuncius, 18(2):887–889, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00341.

Abbri:2003:BRJ

[2424] Ferdinando Abbri. Book review: Julia D. Douthwaite, The Wild Girl, Natural Man and the Monster. Dangerous Experiments in the Age of Enlightenment, Chicago and London, The University of Chicago Press, 2002, xm+314 pp., ISBN 0-226-16056-4 paper, prezzo non indicato. Nuncius, 18(2):889–891, ???? 2003. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539103x00350.

Bandinelli:2003:BRR

[2425] Angela Bandinelli. Book review: Raffaella Seligardi, Lavoisier in Italia. La comunità scientifica italiana e la rivoluzione chimica, (Italian) [Lavoisier in Italy. The Italian Scientific Community and the Chemical Revolution], Firenze, Leo Olschki, 2002, 414 pp., ISBN 88-222-5150-4. Nuncius, 18(2):891–894, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00369.

Pogliano:2003:BRD

[2426] Claudio Pogliano. Book review: Delia Frigessi, Cesare Lombroso, Torino, Einaudi, 2003, 426 pp., ISBN 88-06-13866-9. Nuncius, 18(2):894–897, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00378.

LaVergata:2003:BRG

[2427] Antonello La Vergata. Book review: Gregory Moore, Nietzsche, Biology and Metaphor, Cambridge, Cambridge University Press, 2002, viii, 228 pp., ISBN 0-521-81230-5 (hardback), EUR37.50 (US$55.00). Nuncius, 18(2):897–899, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00387.

Polizzi:2003:BRC

[2428] Gaspare Polizzi. Book review: Claudia Stancati, Donata Chiricò, Federica Vercillo (eds.), Henri Bergson: esprit et langage, Sprimont (Belgique), Pierre Mardaga editeur, 2001, 317 pp., s.i.p., ISBN 2-87009-765-4 (≪Philosophie et Langage≫ ). Nuncius, 18(2):899–902, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00396.

Polizzi:2003:BRM

[2429] Gaspare Polizzi. Book review: Maria Rita Abramo, Il raxionalismo di Gaston Bachelard, Messina, Armando Siciliano Editore, 2000, 156 pp., L. 20.000 (≪Interazioni≫, 3). Gaston Bachelard, l’Esperienza dello spazio nella fisica contemporanea, a cura di Maria Rita Abramo, Messina, Armando Siciliano Editore, 2002, lxxxviii + 91 pp., EUR10,50 (≪Interazioni≫, 9). Maria Rita Abramo, Gaston Bachelard e le fisiche del Novecento, Napoli, Alfredo Guida Editore, 2002, 237 pp., EUR15,00, ISBN 88-7188-593-7 (≪Transazioni≫, 2). Nuncius, 18(2):902–907, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00404.

Lippi:2003:BRE

[2430] Donatella Lippi. Book review: Eugenia Tognotti, La ≪ spagnola≫ in Italia. Storia dell’influenza che fece temere La fine del mondo (1918– 1919), Milano, Franco Angeli, 2002, pp. 165, EUR17,50. ISBN 88-464-3692-X. Nuncius, 18(2):908–909, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00413.

Bertini:2003:BRA

[2431] Silvia Bertini. Book review: Antonio Semerari, Storia, teorie e tecniche della psicoterapia cognitiva, Roma/Bari, Editori Laterza, 2002, 291 pp., EUR20,00, ISBN 88-420-6155-7 (≪Scienze della mente≫ ). Nuncius, 18 (2):909–911, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539103x00422.

Rasini:2003:BRK

[2432] Vallori Rasini. Book review: Karl Lowith, Dio, uomo e mondo nella metafisica da Cartesio a Nietzsche. A cura di O. Franceschelli, Roma, Donzelli, 2000, ISBN 88-7989-595-8, EUR19,63. Nuncius, 18(2):911–912, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00431.

Toscano:2003:BRG

[2433] Anna Toscano. Book review: George Gheverghese Joseph, C’era una volta un numero. La vera storia della matematica, (Italian) [There was a once a number. The true history of mathematics], tr. it. a cura di Barbara Mussini, Milano, il Saggiatore, 2000, 448 pp. + ill., ISBN 88-428-0778-8, EUR20,14. Nuncius, 18(2):913–916, ????

2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00440.

Anonymous:2003:BRB

[2434] Anonymous. Book review: Biblioteca di Nuncius 47. Domenico Laurenza, La ricerca dell’armonia. Rappresentazioni anatomiche nel Rinascimento 2003, x + 142 pp., Olschki, Firenze. Nuncius, 18(2):917, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00459.

Anonymous:2003:BRMa

[2435] Anonymous. Book review: 48. Marco Piccolino, The Taming of the Ray. Electric Fish Research in the Enlightenment, from John Walsh to Alessandro Volta 2003, xiv + 222 pp., Olschki, Firenze. Nuncius, 18(2): 918, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00468.

Anonymous:2003:BRMb

[2436] Anonymous. Book review: Musa Musaei. Studies on Scientific Instruments and Collections in Honour of Mara Miniati. Edited by Marco Beretta, Paolo Galluzzi, Carlo Triarico 2003, viii-488 pp., Olschki, Firenze. Nuncius, 18(2):918–919, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539103x00477.

Anonymous:2003:IAX

[2437] Anonymous. Indici (anno XVIII, 2003). Nuncius, 18(2):921–924, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00486.

Anonymous:2003:IM

[2438] Anonymous. Indice dei manoscritti. Nuncius, 18(2):925–929, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00495.

Anonymous:2003:IN

[2439] Anonymous. Indice dei nomi. Nuncius, 18(2):931–958, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00503.

Anonymous:2003:ISI

[2440] Anonymous. Indice dei siti Internet. Nuncius, 18(2):959, ???? 2003. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539103x00512.

Casi:2004:MAT

[2441] Fausto Casi. A medieval astrolabe in the tradition of Jean Fusoris. Nuncius, 19(1):3–29, ???? 2004. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539104x00016.

Bedin:2004:FRA

[2442] Silvio A. Bedin. Falconi, Renaissance astrologer and astronomical clock and instrument maker. Nuncius, 19(1):31–76, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00025.

Snchez:2004:AFC

[2443] Jos Ramn Bertomeu Snchez and Antonio Garca Belmar. Atoms in French chemistry textbooks during the first half of the Nineteenth Century. Nuncius, 19(1):77–119, ???? 2004. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539104x00034.

Cicenia:2004:GLF

[2444] Salvatore Cicenia. La geometria Lobacevskiana e la filosofia della matematica di Auguste Comte. Nuncius, 19(1):121–141, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00043.

Omodeo:2004:SSP

[2445] Pietro Omodeo. Selezione, storia di una parola e di un concetto. Nuncius, 19(1):143–170, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00052.

Pogliano:2004:SWC

[2446] Claudio Pogliano. Sciences at war and the cybernetic dream. Nuncius, 19(1):171–204, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00061.

Degani:2004:PAD

[2447] Franca Cattelani Degani and Mario Umberto Lugli. Per un archivio della corrispondenza degli scienziati italiani. Nuncius, 19(1):205–223, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00070.

Barbieri:2004:OCS

[2448] Francesco Barbieri and Marina Zuccoli. Odoardo corsini e la scienza delle acque: Opuscoli. Nuncius, 19(1):225–243, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539104x00089.

Gatto:2004:SRC

[2449] Romano Gatto, Iodice Lisa, and Emilia Olostro Cirella. Sul ritrovamento di un carteggio scientifico di Emanuele Fergola. Nuncius, 19(1):245–252, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00098.

Manara:2004:BRG

[2450] Alessandro Manara and Franca Chlistovsky. Book review: Giovanni Virginio Schiaparelli, Percival Lowell Scambi Epistolari Inediti (1896– 1910). Nuncius, 19(1):253–255, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539104x00106.

Mandrino:2004:CE

[2451] Agnese Mandrino. Criteri di edizione. Nuncius, 19(1):256–296, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00115.

Bartola:2004:AOM

[2452] Alberto Bartola. Alle origini del museo del collegio romano. Nuncius, 19(1):297–356, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00124.

Casati:2004:BPO

[2453] Stefano Casati. La biblioteca di Pietro Omodeo. Nuncius, 19(1):357– 373, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00133.

Casati:2004:NM

[2454] S. Casati and S. Contardi. Nova media. Nuncius, 19(1):375–383, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00142.

Ciardi:2004:DC

[2455] Marco Ciardi. Discussioni critiche. Nuncius, 19(1):385–423, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00151.

Triarico:2004:BRS

[2456] Carlo Triarico. Book review: Svante Lindqvist (ed.), Museums of Modern Science (Nobel Symposium 112), Canton MA, Science History Publications, 2000, 216 pp., illus., $39.95/$24.95 paperback. Nuncius, 19 (1):425–427, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00160.

Romizi:2004:BRM

[2457] Donata Romizi. Book review: Margaret Morrison, Unifying Scientific Theories. Physical Concepts and Mathematical Structures, Cambridge, Cambridge University Press, 2000, pp. 272, ISBN 0-521-65216-2. Nuncius, 19(1):427–430, ???? 2004. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539104x00179.

Strano:2004:BRF

[2458] Giorgio Strano. Book review: Francesco Bertola, Via Lactea, Cittadella (Pd), Biblos, 2003, pp. 216 ill., s.i.p. (ISBN 88-88064-22-2). Nuncius, 19 (1):430–432, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00188.

Acerbi:2004:BRC

[2459] Fabio Acerbi. Book review: C. J. Tuplin, T. E. Rihll (eds.), Science and Mathematics in Ancient Greek Culture, Oxford, Oxford University Press, 2002, xvi + 379 pp., £60, ISBN 0-19-815248-5. Nuncius, 19(1): 432–434, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00197.

Strano:2004:BRL

[2460] Giorgio Strano. Book review: Luigi Pepe (a cura di), Copernico e lo Studio di Ferrara: Università, dottori e studenti, Bologna, Clueb, 2003, 152 pp. ill., 13,50 (ISBN 88-491-2126-1). Nuncius, 19(1):435–437, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00205.

Fiocca:2004:BRD

[2461] Alessandra Fiocca. Book review: Della scienza et dell’arte del ben regolare le acque di Gio. Battista Aleotti detto l’Argenta architetto del Papa, et del publico ne lacitt di Ferrara, a cura di Massimo Rossi, Modena, Franco Cosimo Panini (Testi, Istituto di Studi Rinascimentali, Ferrara), 2000, pp. 909. ISBN 88-7686-980-8, 155. Nuncius, 19(1):437–441, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00214.

Bandinelli:2004:BRA

[2462] Angela Bandinelli. Book review: Antonio Clericuzio, Elements, Principles and Corpuscles. A Study of Atomism and Chemistry in the Seventeenth Century, Kluwer Academic Publishers, Dordrecht, Boston, London, 2000, pp. xi + 223, ISBN 0-7923-6782-0. Nuncius, 19(1):442–443, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00223.

Abbri:2004:BRGa

[2463] Ferdinando Abbri. Book review: Giuliano Pancaldi, Volta. Science and Culture in the Age of Enlightenment, Princeton and Oxford, Princeton University Press, 2003, pp. xx + 381, £22.95, ISBN 0-691-09685-6. Nuncius, 19(1):444–446, ???? 2004. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539104x00232.

Abbri:2004:BRR

[2464] Ferdinando Abbri. Book review: Roy Porter (ed. ), The Cambridge History of Science. Volume 4. Eighteenth-century Science, Cambridge, Cambridge University Press, 2003, pp. xxx + 912. £65.00. ISBN 0-521-57243-6. Nuncius, 19(1):446–450, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539104x00241.

Segala:2004:BRT

[2465] Marco Segala. Book review: Thomas Bach, Biologie und Philosophie bei C. F. Kielmeyer und F. W. J. Schelling, Stuttgart-Bad Cannstatt, Frommann-Holzboog, 2001, x + 347 pp., ISBN 3-7728-2034-4. Nuncius, 19(1):450–451, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00250.

Segala:2004:BRM

[2466] Marco Segala. Book review: Martin Kirschke, Liebigs Lehrer Karl W. G. Kastner (1783–1857). Eine Professorenkarriere in Zeiten naturwissenschaftlichen Umbruchs, Berlin-Diepholz, Verlag für Geschichte der Naturwissenschaften und der Technik, 2001, 450 pp., 38.50, ISBN 3-928186-56-6. Nuncius, 19(1):451–453, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539104x00269.

Abbri:2004:BRGb

[2467] Ferdinando Abbri. Book review: Guido Bonino, T. H. Green e il mito dell’empirismo britannico, Firenze, Leo S. Olschki, 2003, pp. 224 (Facolt di Lettere e Filosofia. Fondo di studi Parini-Chirio. Università degli Studi di Torino. Filosofia n. 3). ISBN 88-222-5214-4. Nuncius, 19(1):453–454, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00278.

Lippi:2004:BRD

[2468] Donatella Lippi. Book review: A. Droescher, Le Facolt medicochirurgiche italiane (1860–1915). Repertorio delle cattedre e degli stabilimenti annessi, dei docenti, dei liberi docenti e del personale scientifico, Bologna, Clueb, 2002, pp. 518. 73,50. ISBN 88-491-1940-2. Nuncius, 19 (1):454–456, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00287.

Polizzi:2004:BRJ

[2469] Gaspare Polizzi. Book review: Jean-François Stoffel, Le phénomnalisme problèmatique de Pierre Duhem, Préface de Jean Ladrire, Bruxelles, Académie royale de Belgique, 2002, 391 pp., 32, ISBN 2-8031-0190-4 (Mémoire de la Classe des Lettres, XXVII). Nuncius, 19(1):456–458, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00296.

Galavotti:2004:BRD

[2470] Maria Carla Galavotti. Book review: David Howie, Interpreting Probability. Controversies and Developments in the Early Twentieth Century, Cambridge, Cambridge University Press, 2002, pp. 262, ISBN 0-521-81251-8 (no price on the jacket). Nuncius, 19(1):458–460, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00304.

Bertini:2004:BRH

[2471] Silvia Bertini. Book review: Helmut E. Lck, Breve storia della psicologia, Il Mulino, Universale Paperbacks, 2002, 216 pp., 11,00, ISBN 88-15-08676-5. Nuncius, 19(1):461–462, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539104x00313.

Bond:2004:BRJ

[2472] Roberto Bond. Book review: James Lovelock, Omaggio a Gaia. La vita di uno scienziato indipendente, Torino, Bollati Boringhieri, 2002, 473 pp., ISBN 88-339-1435-6. Nuncius, 19(1):463–470, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00322.

Carolino:2004:SPE

[2473] Luís Miguel Carolino. Scienza, politica ed escatologia nella formazione dello ‘scienziato’ nell’Europa del XVII secolo: Il caso di Manuel Bocarro Francês — Jacob Rosales. Nuncius, 19(2):iii–506, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00331.

Capecchi:2004:LRD

[2474] Danilo Capecchi. Leibniz e la resistenza delle travi. Nuncius, 19(2): 507–544, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00340.

Piccardi:2004:UFH

[2475] Giovanni Piccardi. Uberto Francesco Hoefer e la tabula affinitatum. Nuncius, 19(2):545–568, ???? 2004. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539104x00359.

Giudice:2004:RSD

[2476] Franco Giudice. Ricerca, strumenti, didattica e pubblica utilità: Carlo Barletti e la fisica sperimentale a Pavia prima di Volta. Nuncius, 19 (2):569–599, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00368.

Logan:2004:WBR

[2477] Gabriella Berti Logan. Women and botany in risorgimento Italy. Nuncius, 19(2):601–628, ???? 2004. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539104x00377.

Brenni:2004:MHM

[2478] Paolo Brenni. Mechanical and hydraulic models for illustrating electromagnetic phenomena. Nuncius, 19(2):629–657, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539104x00386.

Casati:2004:DGF

[2479] Stefano Casati. La donazione ’Galileo Ferraris’. Nuncius, 19(2):659–666, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00395.

Omodeo:2004:B

[2480] Pietro Omodeo. La mia biblioteca. Nuncius, 19(2):667–670, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00403.

Anonymous:2004:P

[2481] Anonymous. Proceedings. Nuncius, 19(2):671–673, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539104x00412.

Rentetzi:2004:CVS

[2482] Maria Rentetzi. From Cambridge to Vienna: the scintillation counter in female hands. Nuncius, 19(2):675–689, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539104x00421.

Kinsley:2004:DID

[2483] Shaw Kinsley. A. E. Douglass and the instruments of dendrochronology. Nuncius, 19(2):691–699, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539104x00430.

Batllo:2004:RBO

[2484] J. Batllo, C. Clemente, F. Perez-Blanco, and F. vidal. Reconstruction of a bosch-omori seismograph. Nuncius, 19(2):701–711, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00449.

Singh:2004:DMU

[2485] Rajinder Singh. Designing, making and using scientific instruments in India: the example of Jagadis Chunder Bose. Nuncius, 19(2):713–723, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00458.

Paselk:2004:CR

[2486] Richard A. Paselk. The chemical refractometer. Nuncius, 19(2):725–732, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00467.

Segre:2004:CGQ

[2487] Michael Segre. Il caso Galileo: Una questione chiusa? Nuncius, 19 (2):733–747, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00476.

Abbri:2004:BRB

[2488] Ferdinando Abbri. Book review: Bernardino Ochino, Laberinti del Libero Arbitrio. A cura di Marco Bracali, Firenze, Leo S. Olschki Editore, 2004, pp. xcvi + 180, (≪Immagini della Ragione. 9≫ ). ISBN 88-222-5319-1, EUR19,00. Nuncius, 19(2):749–750, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00485.

Sargolini:2004:BRA

[2489] Federica Sargolini. Book review: Ann E. Moyer, The Philosophers’ Game. Rithmomachia in Medioeval and Renaissance Europe, University of Michigan Press, 2001, 205 pp., ISBN 0-472-11228-7. Nuncius, 19(2): 750–753, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00494.

Ciardi:2004:BRP

[2490] Marco Ciardi. Book review: Paolo Bensi, La vita del colore. Tecniche della pittura veneta dal Cinquecento al Settecento, Genova, Neos Edizioni, 2002, pp. 105, ill., ISBN 88-87262-14-4, EUR14,50. Nuncius, 19 (2):753–755, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00502.

Bernardoni:2004:BRA

[2491] Andrea Bernardoni. Book review: Alessandra Fiocca, Daniela Lamberini, Cesare Maffioli (a cura di), Arte e scienza delle acque nel Rinascimento, Venezia, Marsilio Editori, 2003, pp. 301, ISBN 88-317-8329-7, EUR25,00. Nuncius, 19(2):755–759, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539104x00511.

Guerrini:2004:BRG

[2492] Luigi Guerrini. Book review: Giulio Cesare Vanini dal tardo Rinascimento al ’libertinisme érudit’. Atti del convegno di studi, Lecce-Taurisano 24–26 ottobre 1985, a cura di Francesco Paolo Raimondi, Galatina (Le), Congedo, 2003, 528 pp., ISBN 88-8086-497-1, EUR23,00. Nuncius, 19 (2):759–762, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00520.

Iannaccone:2004:BRN

[2493] Isaia Iannaccone. Book review: Noël Golvers, Ferdinand Verbiest, S.J. (1623–1688) and the Chinese Heaven, University Press, 2003, 676 pp., 46 ill., 4 mappe. ISBN 90-5867-293-X. Nuncius, 19(2):762–763, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00539.

Jalon:2004:BRA

[2494] Mauricio Jalón. Book review: Andrés Galera, Ciencia a la sombra del Vesubio. Ensayo sobre el conocimiento de la naturaleza, Madrid, Consejo Superior de Investigaciones Cientificas, 2003, 250 pp., ISBN 84-00-08137-

4. Nuncius, 19(2):764–767, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539104x00548.

Abbri:2004:BRJ

[2495] Ferdinando Abbri. Book review: James E. McClellan III, Specialist Control. The Publications Committee of the Académie Royale des Sciences (Paris) 1700–1793, Philadelphia, American Philosophical Society, 2003, xii + 134 pp. (≪Transactions of the American Philosophical Society Held at Philadelphia for Promoting Useful Knowledge, volume 92, Part 3≫ ), ISBN 0-87169-933-8. Nuncius, 19(2):767–768, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00557.

Abbri:2004:BRH

[2496] Ferdinando Abbri. Book review: Hjalmar Fors, Mutual Favours. The Social and Scientific Practice of Eighteenth-Century Swedish Chemistry, Uppsala, Universitetstryckeriet, 2003, pp. 227 (≪Institutionen for ide- och lärdomshistoria. Uppsala Universitet. Skrifter n. 30≫ ). ISBN 91-506-1669-2. Nuncius, 19(2):768–770, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539104x00566.

Ciardi:2004:BRR

[2497] Marco Ciardi. Book review: Raimondo De Sangro, Lettera Apologetica. Introduzione, note e appendice documentaria a cura di Leen Spruit, Napoli, alóς edizioni, 2002, 303 pp. ill., ISBN 88-88247-09-2, EUR28. Nuncius, 19(2):771–773, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539104x00575.

Toscano:2004:BRP

[2498] Anna Toscano. Book review: Paola Bertucci and Giuliano Pancaldi (eds.), Electric Bodies. Episodes in the History of Medical Electricity, Universita di Bologna, Cis — Dipartimento di Filosofia, 2001, 298 pp., ISBN 88-900162-2-1 (≪Bologna Studies in History of Science≫ 9). Nuncius, 19(2):774–776, ???? 2004. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539104x00584.

Polizzi:2004:BRM

[2499] Gaspare Polizzi. Book review: Maine De Biran, Osservazioni sulle divisioni organiche del cervello, traduzione e introduzione a cura di Marco Piazza, Firenze, Olschki, 2002, 110 pp., ISBN 88-222-5168-7 (≪Biblioteca di Physis≫, 6), EUR13,00. Nuncius, 19(2):777–779, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00593.

Abbri:2004:BRM

[2500] Ferdinando Abbri. Book review: Maja Larsson, Den moraliska kroppen. Tolkningar av kön och individualitet i 1800-talets populärmedicin, Hedemora, Gidlunds forlag 2002, 232 pp., ill., ISBN 91-7844-617-1. Nuncius, 19(2):779–782, ???? 2004. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539104x00601.

Tatasciore:2004:BRG

[2501] Carlo Tatasciore. Book review: G. Polizzi, Leopardi e le ≪ ragioni della verità≫. Scienze e filosofia della natura negli scritti leopardiani, con una Prefazione di R. Bodei, Roma, Carocci, 2003, pp. Ixix, 1-288, EUR19,60, ISBN 88-430-2818-9. Nuncius, 19(2):782–785, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00610.

Gattinara:2004:BRV

[2502] Enrico Castelli Gattinara. Book review: Vallori Rasini, Teorie della Realtà Organica. Helmuth Plessner e Viktor von Weizsäcker, Modena, Edizioni Grafiche Sigem, 2002, 176 pp., EUR10 (privo di ISBN). Nuncius, 19(2):785–788, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00629.

Toscano:2004:BRR

[2503] Anna Toscano. Book review: Robert Kanigel, l’Uomo che vide l’infinito. La vita breve di Srinivasa Ramanujan, genio della matematica, tr. it. di Maddalena Mendolicchio, Milano, Rizzoli, 2003, 462 pp. + ill., ISBN 88-17-87169-9, EUR19,00. Nuncius, 19(2):788–791, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00638.

Ciardi:2004:BRL

[2504] Marco Ciardi. Book review: Luigi Cerruti, Bella e potente: La chimica del Novecento fra scienza e società, Roma, Editori Riuniti, 2003, 507 pp., ill., ISBN 88-359-5396-0, EUR26,00. Nuncius, 19(2):791–795, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539104x00647.

Manzocco:2004:BRD

[2505] Roberto Manzocco. Book review: Debora Hammond, The Science of Synthesis — Exploring the Social Implications of General Systems Theory, Boulder, University Press of Colorado, 2003, pp. 312, ISBN 0-87081-722-1. Nuncius, 19(2):795–796, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539104x00656.

Pogliano:2004:BRJ

[2506] Claudio Pogliano. Book review: Joseph Dumit, Picturing Personhood. Brain Scans and Biomedical Identity, Princeton and Oxford, Princeton University Press, 2004, 252 pp., ISBN 0-691-11398-X. Nuncius, 19(2): 797–799, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00665.

Scarpelli:2004:BRE

[2507] Giacomo Scarpelli. Book review: Enrico Castelli Gattinara, Strane alleanze. Storici, filosofi e scienziati a confrontonel Novecento, Milano, Mimesis, 2003; 213 pp., ISBN 88-8483-186-5, EUR16,00. Nuncius, 19 (2):799–807, ???? 2004. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539104x00674.

Beretta:2005:E

[2508] Marco Beretta. Editorial. Nuncius, 20(1):7–9, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00727.

Egmond:2005:CCS

[2509] Florike Egmond. Clusius, Cluyt, Saint Omer. Nuncius, 20(1):11–67, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00736.

Caudet:2005:TAC

[2510] Amparo Sebastian Caudet and Ruth Lopez Campos. Two astronomical compendia in the Spanish national museum of science and technology. Nuncius, 20(1):69–94, ???? 2005. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539105x00745.

DalPrete:2005:FBP

[2511] Ivano Dal Prete. Francesco Bianchini e il pianeta Venere. Nuncius, 20 (1):95–152, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00754.

Bertucci:2005:SCJ

[2512] Paola Bertucci. Sparking controversy: Jean Antoine Nollet and medical electricity south of the Alps. Nuncius, 20(1):153–187, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00763.

Logan:2005:CSA

[2513] Gabriella Berti Logan. Caterina Scarpellini: Astronomy and meteorology in Risorgimento Rome. Nuncius, 20(1):189–217, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00772.

Conforti:2005:HPM

[2514] Maria Conforti. Healing practices and medical professions in early modern Europe. Nuncius, 20(1):219–229, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00781.

Neve:2005:AMA

[2515] Mario Neve. Antiquity and middle ages. Nuncius, 20(1):231–232, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00790.

Acerbi:2005:BRC

[2516] Fabio Acerbi. Book review: Cleomedes Lectures on Astronomy: a Translation of The Heavens with an Introduction and Commentary by Alan C. Bowen and Robert B. Todd. Berkeley and Los Angeles: University of California Press, 2004. xvi + 238 pp., ISBN 0-520-23325-5. Nuncius, 20(1):232–233, ???? 2005. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539105x00808.

Beretta:2005:BRCa

[2517] Marco Beretta. Book review: Cristina Viano (ed.), Aristotele Chemicus. Il IV libro dei Meteorologica nella tradizione antica e medievale. Sankt Augustin: Academia Verlag, 2002. 206 pp. ISBN 3-89665-174-5. Nuncius, 20(1):233–234, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00817.

Galante:2005:BRR

[2518] Davide Galante. Book review: Robert Wardy, Aristotele in China. Language, Categories and Translation. Cambridge: Cambridge University Press, 2000. 182 pp., ISBN 0-521-77118-8. Nuncius, 20(1):234–235, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00826.

DeCallatay:2005:BRJ

[2519] Godefroid De Callatay. Book review: Jacques Sesiano, Les carrés magiques dans le pays islamiques. Lausanne: Presses polytechniques et universitaire romandes, 2004. 277 pp., ISBN 2-88074-571-3. Nuncius, 20 (1):235–236, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00835.

Craig:2005:BRC

[2520] Martin Craig. Book review: Cecilia Trifogli, Liber Tertius Physicorum Aristotelis. Repertorio delle Questioni. Commenti inglesi ca. 1250–1270.

Firenze: Sismel-Edizioni del Galluzzo, 2004. 393 pp., ISBN 88-8450-039-7. Nuncius, 20(1):237–238, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00844.

Guerrini:2005:REM

[2521] Luigi Guerrini. Renaissance and early modem science. Nuncius, 20(1): 238–239, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00853.

Bernardoni:2005:BRF

[2522] Andrea Bernardoni. Book review: Francesco Paolo Fiore (a cura di), Francesco di Giorgio alla corte di Federico da Montefeltro, Atti del Convegno Internazionale di Studi, Urbino, 11–13 ottobre 2001 (Biblioteca dell’‘Archivum Romanicum’. Serie I, vol. 317). Firenze: Olschki, 2004. 2 vols, xxiv + 710 pp., ill., ISBN 88-222-5322-1. Nuncius, 20(1):239–241, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00862.

Camerota:2005:BRN

[2523] Filippo Camerota. Book review: Nicoletta Marconi, Edificando Roma barocca. Macchine, apparati, maestranze e cantieri tra XVI e XVIII secolo (Studi e strumenti per la storia di Roma). Citta di Castello: Edimond, 2004. 350 pp., ill., ISBN 88-500-237-8. Nuncius, 20(1):241–242, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00871.

Bresadola:2005:BRM

[2524] Marco Bresadola. Book review: Medicina nei Secoli. Arte e scienza, vol. 15, fasc. 2, 2003. 401 pp., ill., ISSN 0394-9001. Nuncius, 20(1): 242–243, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00880.

Maffioli:2005:BRM

[2525] Cesare S. Maffioli. Book review: On the measurement of running water, a facsimile edition of Della misura dell’acque correnti of Dom Benedetto Castelli, together with English translation and commentary by Deane R. Blackman. Firenze: Leo S. Olschki Editore, 2004. 172 pp., ISBN 88-222-5335-3. Nuncius, 20(1):243–244, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00899.

Banks:2005:E

[2526] Terence Banks. Enlightenment. Nuncius, 20(1):245–246, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00907.

Abbri:2005:BRP

[2527] Ferdinando Abbri. Book review: Patricia Fara, Sex, Botany and Empire. The story of Carl Linnaeus and Joseph Banks (Revolutions in Science). Cambridge: Icon Books, 2004, 168 pp., ISBN 1-84046-573-5. Nuncius, 20(1):246–247, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00916.

Borgato:2005:BRA

[2528] Maria Teresa Borgato. Book review: Antoine Deparcieux, Essai sur les probabilités de la durée de la vie humaine (1746), Addition à l’Essai (1760), réédition en fac-simile avec une introduction et des notes de Cem Behar, accompagnée de contributions de G. Gallais-Hamonno, C. Rietsch, J. Berthon. Paris: Ined, 2003. xvii + ccxi + 187 pp., ill., ISBN 2-7332-1026-2. Nuncius, 20(1):247–249, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00925.

Monti:2005:BRR

[2529] Maria Teresa Monti. Book review: Repertorium zu Albrecht von Hallers Korrespondenz 1724–1777, hrsg von Urs Boschung, Barbara Braun-Bucher, Stephan Hachler, Anne Kathrin Ott, Hubert Steinke, Martin Stuber (‘Studia Halleriana’, VII/1–2). Basel: Schwabe, 2002. Vol. I: xlviii + 634 pp., ill.; vol. II: 352 pp., ISBN 3-7965-1325-5. Nuncius, 20(1):249– 251, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00934.

Beretta:2005:BRCb

[2530] Ilva Beretta. Book review: Carlo Allioni, Flora Pedemontana, con introduzioni di Lucia Tongiorgi Tomasi, Rosanna Caramiello, Giuliana Forneris (Rariora et Mirabilia 5), Firenze: Olschki, 2003. 2 vols, xl + 850 pp., ill., ISBN 88-222-5284-5. Nuncius, 20(1):251–252, ????

2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00943.

DiMitri:2005:BRR

[2531] Gino L. Di Mitri. Book review: Rosanna Caramiello e Giuliana Forneris (ed.), Le opere minori di Carlo Allioni. Dal‘Rariorum Pedemontii Stirpium’ all’‘Auctarium ad Floram Pedemontanam’. Firenze: Olschki, 2004. 264 pp., ill., ISBN 88-222-5378-7. Nuncius, 20(1):252– 253, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00952.

Heering:2005:BRO

[2532] Peter Heering. Book review: Oliver Hochadel, Öffentliche Wissenschaft: Elektrizität in der deutschen Aufklärung. Göttingen: Wallstein, 2003. 364 pp., ISBN 3-89244-629-6. Nuncius, 20(1):253–255, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00961.

Kawashima:2005:BRJ

[2533] Keiko Kawashima. Book review: Jean-Pierre Poirier, La Science et l’Amour, Madame Lavoisier. Paris: Pygmalion, 2004. 246 pp., ISBN 2-85704-926-9. Nuncius, 20(1):255–256, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00970.

Vaccari:2005:BRR

[2534] Valentina Vaccari. Book review: Robert E. Schofield, The Enlightened Joseph Priestley. A Study of His Life and Work from 1773 to 1804. University Park, Pennsylvania: the Pennsylvania State University Press, 2004. xv + 461 pp., ill., ISBN 0-271-02459-3. Nuncius, 20(1):256–257, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00989.

Clardi:2005:BRP

[2535] Marco Clardi. Book review: Paolo Mazzarello, Costantinopoli 1786: la congiura e la beffa. l’intrigo Spallanzani, Torino: Bollati Boringhieri, 2004. 327 pp., ill., ISBN 88-339-1573-5. Nuncius, 20(1):257–259, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00998.

DalPrete:2005:BRD

[2536] Ivano Dal Prete. Book review: Davide Arecco, Mongolfiere, Scienze e Lumi nel tardo Settecento. Cultura accademica e conoscenze tecniche dalla vigilia della Rivoluzione Francese all’età Napoleonica. Bari: Cacucci Editore, 2003. 247 pp., ill., ISBN 88-8422-257-5. Nuncius, 20(1):259–260, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01005.

Pepe:2005:BRR

[2537] Luigi Pepe. Book review: Roger Hahn, Le Système du monde. Pierre Simon Laplace. Un itinéraire dans la science. Paris: Gallimard, 2004. 306 pp., ISBN 2-07-072936-2. Nuncius, 20(1):260–261, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01014.

Beretta:2005:BRCc

[2538] Marco Beretta. Book review: Charles Coulston Gillispie, Science and Polity in France. The End of the Old Regame. Princeton and Oxford: Princeton University Press, 2004. 601 pp. ISBN 0-691-11849-3. Charles Coulston Gillispie, Science and Polity in France. The Revolutionary and Napoleonic Years. Princeton and Oxford: Princeton University Press, 2004. xii + 752 pp., ISBN 0-691-11541-9. Nuncius, 20(1):261–264, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01023.

Ratcliff:2005:BRM

[2539] Marc Ratcliff. Book review: Marc-Auguste Pictet 1752–1825. Correspondance sciences et techniques. Tome IV. Les Correspondants suisses, italiens, allemands, établi par René Sigrist. Genève: Slatkine, 2004. xii + 1121 pp., ISBN 2-05-101951-7. Nuncius, 20(1):264–265, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01032.

Ciancio:2005:MCS

[2540] Luca Ciancio. Modern and contemporary science. Nuncius, 20(1):265– 266, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01041.

Miniati:2005:BRJ

[2541] Mara Miniati. Book review: Jean-Paul Poirier, Anthony Turner, Antoine d’Abbadie (Académie des Sciences. Mémoire de la science 2). S. I.: Académie des Sciences & les auteurs, 2002. 127 pp., ill., ISBN 2-909344-09-06. Nuncius, 20(1):266–268, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x01050.

Carletti:2005:BRI

[2542] Christian Carletti. Book review: Iwan Rhys Morus, Michael Faraday and the Electrical Century (Revolutions in Science). Cambridge: Icon Books, 2004. 229 pp., ill., ISBN 1-84046-540-9. Nuncius, 20(1):268–269, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01069.

Bondio:2005:BRV

[2543] Mariacarla Gadebusch Bondio. Book review: Valeria P. Babini, Il caso Murri. Una storia italiana. Bologna: Il Mulino, 2004. 309 pp., ISBN 88-15-09730-9. Nuncius, 20(1):269–271, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x01078.

Sommer:2005:BRN

[2544] Marianne Sommer. Book review: Natalie Janine Chaoui, Mit Messzirkel und Schrotflinte. Das primatologische Werk von Adolph Hans Schultz (1891–1976). Marburg: Basilisken-Presse, 2004. 234 pp., ISBN 3-925347-73-9. Nuncius, 20(1):271–272, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x01087.

Turchetti:2005:BRR

[2545] Simone Turchetti. Book review: Robert Marc Friedman, The Politics of Excellence. Behind the Nobel Prize in Science. New York: Times Books, 2001. xv + 379 pp., ISBN 0-7167-3103-7. Nuncius, 20(1):272– 273, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01096.

Dennis:2005:BRJ

[2546] Michael Aaron Dennis. Book review: Jeff Hughes, The Manhattan Project: Big Science and the Atom Bomb (Revolutions in Science). Cambridge: Icon Books, 2002. vi + 170 pp., ill., ISBN 1-84046-504-2. Nuncius, 20(1):273–275, ???? 2005. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539105x01104.

Tournes:2005:G

[2547] Dominique Tournès. General. Nuncius, 20(1):275–276, ???? 2005. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x01113.

Minonzio:2005:BRE

[2548] Franco Minonzio. Book review: Émile Biémont, Ritmi del tempo. Astronomia e calendari, traduzione di Barbara Fujani. Bologna: Zanichelli, 2002. vii + 344 pp., ISBN 88-08-08931-2. Nuncius, 20(1):276–281, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01122.

Strano:2005:BRF

[2549] Giorgio Strano. Book review: Fred Watson, Stargazer: the Life and Times of the Telescope. Crows Nest: Allen & Unwin, 2004. x + 342 pp., ill., ISBN 1-86508-658-4. Nuncius, 20(1):281–282, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01131.

Dolza:2005:BRG

[2550] Luisa Dolza. Book review: I. Gouzévitch, A. Grelon, A. Karvar (eds.), La formation des ingégnieurs en perspective. Modèles de référence et résaux de médiation XVIIe–XXe siècle. Rennes: Presses Universitaire de Rennes, 2004. 176 pp., ISBN 2-86847-996-0. Nuncius, 20(1):282–284, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01140.

Capocci:2005:BRGa

[2551] Mauro Capocci. Book review: L. Garwin and T. Lincoln (eds.), A century of Nature: twenty-one discoveries that changed science and the world. Chicago: University of Chicago Press, 2003. xvm + 360 pp., ISBN 0-226-28415-8. Nuncius, 20(1):284–285, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x01159.

Passione:2005:BRD

[2552] Roberta Passione. Book review: David B. Baker (ed.), Thick Description and Fine Texture. Studies in the History of Psychology. Akron: The University of Akron Press, 2003. xxii + 249 pp., ISBN 1-931968-02-0. Nuncius, 20(1):285–286, ???? 2005. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539105x01168.

Ceregato:2005:BRH

[2553] Alessandro Ceregato. Book review: Helen M. Rozwadowski and David K. Van Keuren (eds.), The Machine in Neptune’s Garden — Historical Perspectives on Technology and the Marine Environment. Sagamore Beach, MA: Science History Publications/USA, 2004. xxviii + 371 pp., ISBN 0-88135-372-8. Nuncius, 20(1):286–287, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01177.

Meschiari:2005:BRM

[2554] Alberto Meschiari. Book review: Marian Fournier, Early Microscopes. A descriptive catalogue. Leiden: Museum Boerhaave, 2003. 236 pp., ill., ISBN 90-6292-143-4. Nuncius, 20(1):287–289, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01186.

Mantovani:2005:BRS

[2555] Roberto Mantovani. Book review: Science in the Provinces. A Descriptive Catalogue of the ‘Deventer’ Collection. Leiden: Museum Boerhaave, 2002. 214 pp., ill., ISBN 90-6292-139-6. Nuncius, 20(1):289–290, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01195.

Vannozzi:2005:BRK

[2556] Francesca Vannozzi. Book review: K. S. Gross, Early surgical instruments; B. W. J. Grob, The anatomical models of Dr. Louis Auzuox; B. W. J. Grob, The laboratory equipment of Willem Einthoven. A descriptive catalogue. Leiden: Museum Boerhaave, 2004. 183 pp., ill., ISBN 90-6292-148-5. Nuncius, 20(1):290–291, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x01203.

Govoni:2005:BRL

[2557] Paola Govoni. Book review: Liborio Dibattista, Storia della scienza e didattica delle discipline scientifiche. Roma: Armando Editore, 2004. 175 pp., ISBN 88-8358-546-1. Nuncius, 20(1):291–292, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x01212.

Capocci:2005:BRGb

[2558] Mauro Capocci. Book review: Giuliano Pancaldi (a cura di), Natura, cultura, identità. Le università e l’identità europea. Bologna: Studies in History of Science, 10: 2004. 213 pp., ISBN 88-900162-3-X. Nuncius, 20 (1):292–293, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x01221.

Pogliano:2005:BRL

[2559] Claudio Pogliano. Book review: Lorraine Daston and Fernando Vidal (eds.), The Moral Authority of Nature. Chicago and London: The University of Chicago Press, 2004. 520 pp., ISBN 0-226-13681-7. Nuncius, 20(1):293–294, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x01230.

Abbri:2005:BRA

[2560] Ferdinando Abbri. Book review: Alex Owen, The Place of Enchantment. British Occultism and the Culture of Modern. Chicago and London: The University of Chicago Press, 2004. xiv + 355 pp., ISBN 0-226-64201-1. Nuncius, 20(1):294–296, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x01249.

Hochadel:2005:BRD

[2561] Oliver Hochadel. Book review: Dan Burton and David Grandy, Magic, Mystery, and Science. The Occult in Western Civilization. Bloomington: Indiana University Press, 2004. 390 pp., ill., ISBN 0-253-21656-7. Nuncius, 20(1):296–297, ???? 2005. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539105x01258.

Werrett:2005:BRD

[2562] Simon Werrett. Book review: David N. Livingstone, Putting Science in Its Place: Geographies of Scientific Knowledge. Chicago and London:

University of Chicago Press, 2003. xii + 234 pp., ill., ISBN 0-226-48722-

9. Nuncius, 20(1):297–298, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x01267.

Strano:2005:Na

[2563] Giorgio Strano. Newsletter. Nuncius, 20(1):299–317, ???? 2005. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x01276.

Ferriello:2005:LHB

[2564] Giuseppina Ferriello. The lifter of heavy bodies of Heron of Alexandria in the Iranian World. Nuncius, 20(2):327–345, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00015.

Tognoni:2005:NDF

[2565] Federico Tognoni. Nature described: Fabio colonna and natural history illustration. Nuncius, 20(2):347–370, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00024.

Iannaccone:2005:DRL

[2566] Isaia Iannaccone. Documents relatifs l’histoire de l’astronomie chinoise et aux rapports scientifiques entre l’Europe et la Chine (XVIIE–XIXE siècles) conservés la Bibliothèque de l’Observatoire de Paris. Nuncius, 20(2):371–396, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00033.

DAgostino:2005:DRD

[2567] Salvo D’Agostino. Il difficile ricupero dell’Anschaulichkeit di Goethe nell’opera di Helmholtz. Nuncius, 20(2):401–414, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00042.

Lthy:2005:MSS

[2568] Christoph Lthy. Museum spaces and spaces of science. reflections on the explanatory possibilities of history of science collections. Nuncius, 20 (2):415–429, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00051.

Shea:2005:NBG

[2569] William R. Shea. A new biography of Galileo. Nuncius, 20(2):431–448, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00060.

Acerbi:2005:AMA

[2570] Fabio Acerbi. Antiquity and middle ages. Nuncius, 20(2):449–450, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00079.

Zavattero:2005:BRP

[2571] Irene Zavattero. Book review: Paloma Llorente Megas, con la colaboración de Merc Otero Vidal (eds.), Liber de Causis. Indice y Concordancia (Lessico Intellettuale Europeo. XCIII). Firenze: Olschki, 2004. x + 304, 30, ISBN 88-222-5351-5. Nuncius, 20(2):450–451, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00088.

DiPasquale:2005:BRF

[2572] Giovanni Di Pasquale. Book review: Francesco Prontera (ed.), Tabula Peutingeriana: le antiche vie del mondo. Firenze: Olschki, 2003. 66 pp., ill., 150, ISBN 88-222-5269-1. Nuncius, 20(2):451–452, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00097.

Acerbi:2005:BRR

[2573] Fabio Acerbi. Book review: R. Rashed and H. Bellosta, Ibrhm ibn Sinn. Logique et géomtrie au Xe siècle. Leiden/Boston/Köln: Brill, 2000. x + 809 pp., 225, ISBN 90-04-11804-7. Nuncius, 20(2):452–453, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00105.

Minonzio:2005:BRS

[2574] Franco Minonzio. Book review: Sinesio, La mia fortunosa navigazione da Alessandria a Cirene, a cura di Pietro Janni (Biblioteca di Geographia Antiqua, 1). Firenze: Leo S. Olschki Editore, 2003. 133 pp., ISBN 88-222-5187-3. Nuncius, 20(2):453–455, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00114.

Soldano:2005:REM

[2575] Adriano Soldano. Renaissance and early modem science. Nuncius, 20 (2):455–456, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00123.

Pepe:2005:BRB

[2576] Luigi Pepe. Book review: Biographica Copernica, Die Copernicus-Biographien des 16. bis 18. Jahrhunderts, Texte und Übersetzungen, Bearbeitet von Andreas Kh?ne und Stefan Kirschner, mit einem Katalog der F?rben Copernicus—Portraits von Godula Metze. Berlin: Akademie Verlag, 2004. xxx + 508 pp., ISBN 3-05-003848-9. Nuncius, 20(2):456– 457, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00132.

Cndito:2005:BRR

[2577] Cristina Cndito. Book review: Richard Bsel, Orazio Grassi, architetto e matematico gesuita (Biblioteca Blu/Saggi). Roma: Argos, 2004. 352 pp., ill., ISBN 88-88690-02-6. Nuncius, 20(2):457–459, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00141.

deCeglia:2005:BRS

[2578] Francesco Paolo de Ceglia. Book review: Sarah Carvallo (ed.), La controverse entre Stahl et Leibniz sur la vie, l’organisme et le mixte. Doutes concernant la Vraie Théorie Médicale du célébre Stahl, avec les répliques de Leibniz aux observations stahliennes. Texte introduit, traduit et annoté par Sarah Carvallo, Préface de Michel Serres. Paris: Librairie Philosophique J. Vrin, 2004. 216 pp., 28, ISBN 2-7116-1629-0. Nuncius, 20(2):459–460, ???? 2005. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539105x00150.

Clericuzio:2005:BRG

[2579] Antonio Clericuzio. Book review: A. G. Debus (ed.), Alchemy and Early Modern Chemistry: Papers from Ambix, [s.l.]: Jeremy Mills Publishing for the Society for the History of Alchemy and Chemistry, 2004. xv + 543 pp., ISBN 0-9546484-1-2. Nuncius, 20(2):460–461, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00169.

Boccadoro:2005:BRB

[2580] Brenno Boccadoro. Book review: Benedetto Di Falco, l’Origine degli alfabeti e dei numeri, a cura di Gino L. Di Mitri. Galatina: Congedo Editore, 2004. x + 154 pp., 15, ISBN 88-8086-556-0. Nuncius, 20(2): 461–462, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00178.

Tognoni:2005:BRD

[2581] Federico Tognoni. Book review: Daniela Fausti (ed.), La complessa scienza dei Semplici. Atti delle celebrazioni per il V centenario della nascita di Pietro Andrea Mattioli, 12 marzo–19 novembre 2001. Atti Accademia Fisiocritici di Siena, 2001, 20 (suppl.). xv + 278 pp., ill. Nuncius, 20(2):462–463, ???? 2005. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539105x00187.

Terrall:2005:BRM

[2582] Mary Terrall. Book review: Mary E. Fissell, Vernacular Bodies: the Politics of Reproduction in Early Modern England. Oxford and New York: Oxford University Press, 2004. x + 284 pp., ill., ISBN 0-19-926988-2. Nuncius, 20(2):463–464, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00196.

Parodi:2005:BRR

[2583] Massimo Parodi. Book review: R. L. Friedman and L. O. Nielsen (eds.), The Medieval Heritage in Early Modern Metaphysics and Modal Theory, 1400–1700. Dordrecht: Kluwer Academic Publishers, 2003. 346 pp., ISBN 1-4020-1631-X. Nuncius, 20(2):464–466, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00204.

Negruzzo:2005:BRP

[2584] Simona Negruzzo. Book review: Paul F. Grendler, The Universities of the Italian Renaissance. Baltimore and London: John Hopkins University Press, 2002. xx + 592 pp., ISBN 0-8018-8055-6. Nuncius, 20(2):466–467, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00213.

Redondi:2005:BRM

[2585] Pietro Redondi. Book review: Marcus Hellyer, Catholic Physics. Jesuit Natural Philosophy in Early Modern Germany. Notre Dame: University of Notre Dame Press, 2005. x + 337 pp., ill., $50, ISBN 0-268-03071-

5. Nuncius, 20(2):467–468, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00222.

Abbri:2005:BRJ

[2586] Ferdinando Abbri. Book review: Jamie C. Kassler, The Beginnings of the Modern Philosophy of Music in England. Francis North’s A Philosophical Essay of Musick (1677) with comments of Isaac Newton, Roger North and in the Philosophical Transactions. Aldershot: Ashgate, 2004. xv + 258 pp., ISBN 0-7546-0139-0. Nuncius, 20(2):468–469, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00231.

Sweet:2005:BRN

[2587] Rosemary Sweet. Book review: Neil Kenny, The Uses of Curiosity in Early Modern France and Germany. Oxford and New York: Oxford University Press, 2004. xv + 484 pp., ill., £68.00, ISBN 0-19-927136-4. Nuncius, 20(2):470–471, ???? 2005. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539105x00240.

Camerota:2005:BRO

[2588] Michele Camerota. Book review: Oddone Longo, Scritti su Galileo e il suo tempo. Padova: Esedra Editrice, 2004. 176 pp., ill., 16, ISBN 88-86413-88-2. Nuncius, 20(2):471–472, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00259.

Ciseri:2005:BRL

[2589] Lorenzo Montemagno Ciseri. Book review: Laura Lunger Knoppers and Joan B. Landes (eds.), Monstrous Bodies/Political Monstrosities in Early Modern Europe. Ithaca: Cornell University Press, 2004. xi + 304 pp., ill., $22.50, ISBN 0-8014-8901-6. Nuncius, 20(2):472–473, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00268.

Lamberini:2005:BRF

[2590] Daniela Lamberini. Book review: Francesco Martelli (ed.), Il viaggio in Europa di Pietro Guerrini (1682–1686). Edizione della corrispondenza e dei disegni di un inviato di Cosimo III dei Medici (2 vols.: vol. 1: Carteggio con Apollonio Bassetti; vol. 2: Disegni e indici). Firenze: Olschki, 2005. cxxv + 508 pp., 133 plates, 69,00, ISBN 88-222-5401-5. Nuncius, 20(2):473–475, ???? 2005. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539105x00277.

Shackelford:2005:BRB

[2591] Jole Shackelford. Book review: Bruce T. Moran, Distillling Knowledge: Alchemy, Chemistry, and the Scientific Revolution. Cambridge, Mass.: Harvard University Press, 2005. 210 pp., $24.95, ISBN 0-674-01495-2. Nuncius, 20(2):475–476, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00286.

Ciappi:2005:BRC

[2592] Silvia Ciappi. Book review: Cesare Moretti, Carlo Stefano Salerno and Sabina Tommasi Ferroni, Ricette vetrarie muranesi. Gasparo Brunoro e il manoscritto di Danzica. Firenze: Nardini, 2004. 208 pp., ill., ISBN 88-404-5066-1. Nuncius, 20(2):476–477, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00295.

Sebastin:2005:BRV

[2593] Amparo Sebastin. Book review: Victor Navarro (ed.), Jerónimo Muñoz. Introduccióna a la Astronomía y la Geografía. Paterna: Collecio Oberta, La Imprenta, 2004. 354 pp., ill., ISBN 84-482-3709-9. Nuncius, 20(2): 477–478, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00303.

Dear:2005:BRW

[2594] Peter Dear. Book review: William R. Shea, Designing Experiments and Games of Chance: the Unconventional Science of Blaise Pascal (Science History Publications). Canton: Watson Publishing International, 2003. 354 pp., ISBN 0-88135-376-0. Nuncius, 20(2):478–479, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00312.

Favino:2005:BRW

[2595] Federica Favino. Book review: William R. Shea and Mariano Artigas, Galileo in Rome. The Rise and Fall of a Troublesome Genius. Oxford: Oxford University Press, 2003. x + 226 pp., plates, ISBN 0-19-516598-5. Nuncius, 20(2):479–481, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00321.

Hyrup:2005:BRM

[2596] Jens Hyrup. Book review: Maryvonne Spiesser (ed.), Une arithmétique commerciale du XVe siècle. Le Compendy de la praticque des nombres de Barthlemy de Romans. (De Diversis artibus, 70). Turnhout: Brepols, 2004. 762 pp., ISBN 2-503-51196-1. Nuncius, 20(2):481–482, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00330.

Rignani:2005:BRT

[2597] Orsola Rignani. Book review: L. Tampellini, Ruggero Bacone. Un passaggio nodale all’origine della scienza moderna. Siena: Cantagalli, 2004. 154 pp., ill., ISBN 88-8272-182-5. Nuncius, 20(2):482–483, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00349.

Laurenza:2005:BRR

[2598] Domenico Laurenza. Book review: Ralf Vollmuth, Das anatomische Zeitalter. Die Anatomie der Renaissance von Leonardo da Vinci bis Andreas Vesal. München: Verlag, 2004. 185 pp., ill., ISBN 3-929360-70-5. Nuncius, 20(2):483–484, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00358.

Redi:2005:E

[2599] Carlo Alberto Redi. Enlightenment. Nuncius, 20(2):484–485, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00367.

Fara:2005:BRM

[2600] Patricia Fara. Book review: Mordechai Feingold, The Newtonian Moment: Isaac Newton and the Making of Modern Culture. New York Oxford: New York Public Library/Oxford University Press, 2004. xvi + 224 pp., ill., ISBN 0-19-517734-7. Nuncius, 20(2):485–486, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00376.

Abbri:2005:BRM

[2601] Ferdinando Abbri. Book review: Martin Fitzpatrick, Peter Jones, Christa Knellwolf and Lain McCalman (eds.), The Enlightenment World, London and New York: Routledge, 2004. xx + 714 pp., ill. ISBN 0-415-21575-7. Nuncius, 20(2):487–488, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00385.

Ciancio:2005:BRC

[2602] Luca Ciancio. Book review: Cinzio Gibin (a cura di), Lettere di Giuseppe Olivi (1769–1795) naturalista. Conselve (PD): Edizioni Think Adv, 2004. 180 pp., ill., ISBN 88-8354-002-6. Nuncius, 20(2):488–489, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00394.

Seligardi:2005:BRA

[2603] Raffaella Seligardi. Book review: Antoine-Laurent Lavoisier, Opuscoli fisici e chimici, a cura di Marco Ciardi e Marco Taddia. Bologna: Bononia University Press, 2005. xv + 286 pp., ill., 45,00, ISBN 88-7395-012-4. Nuncius, 20(2):489–490, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00402.

Vaccari:2005:BRB

[2604] Valentina Vaccari. Book review: Ben Marsden, Watt’s Perfect Engine. Steam and the Age of Invention (Revolutions in Science). Royston: Icon Books, 2004. x + 214 pp., ill., ISBN 1-84046-546-8. Nuncius, 20(2): 490–491, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00411.

Ciardi:2005:BRI

[2605] Marco Ciardi. Book review: Immanuel Kant, Geografia fisica. Riproduzione anastatica dell’edizione Silvestri, Milano, 1807–1811. Introduzione di Franco Farinelli. Bergamo: Leading Edizioni, 2004. 3 vols.; 1 vol.: xxiv + xxxv + 581 pp., ill., 2 vol.: 400 pp., 3 vol.: 508 pp., ISBN 88-86996-10-1. Nuncius, 20(2):491–493, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00420.

Bandinelli:2005:BRR

[2606] Angela Bandinelli. Book review: Russell McCormmach, Speculative Truth. Henry Cavendish, Natural Philosophy and the Rise of Modern Theoretical Science. Oxford, New York: Oxford University Press, 2004. x + 258 pp., ill., ISBN 0-19-516004-5. Nuncius, 20(2):493–494, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00439.

Amengual:2005:BRA

[2607] Rubn Amengual. Book review: Alessandro Nuvolari, The Making of Steam Power Technology. A Study of Technical Change during the British Industrial Revolution. Eindhoven: Eindhoven University Press, 2004. x + 211 pp., ISBN 90-386-2077-2. Nuncius, 20(2):494–495, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00448.

Ratcliff:2005:BRC

[2608] Marc J. Ratcliff. Book review: Concetta Pennuto (ed.), Antonio Vallisneri, Quaderni di osservazioni (Sezione I, Volume 1 dell’Edizione nazionale delle opere di Antonio Vallisneri). Firenze: Olschki, 2004. cv + 258 pp., ISBN 88-222-5331-0. Nuncius, 20(2):495–497, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00457.

Rousseau:2005:BRN

[2609] George Rousseau. Book review: Neil Vickers, Coleridge and the doctors, 1795–1806 (Oxford English Monographs). Oxford: Clarendon Press, 2004. 198 pp., ISBN 0-19-927117-8. Nuncius, 20(2):497–498, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00466.

Gourvish:2005:MCS

[2610] Terry Gourvish. Modem and contemporary science. Nuncius, 20(2): 498–499, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00475.

Casati:2005:BRJ

[2611] Stefano Casati. Book review: Jean-Baptiste Biot et la météorite de l’Aigle, Paris: Académie des sciences, 2003. Nuncius, 20(2):499–501, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00484.

Barsanti:2005:BRC

[2612] Giulio Barsanti. Book review: Carlo Luciano Bonaparte, Iconografia della Fauna Italica per le quattro classi dei Vertebrati (a cura di Alessandro Minelli e Augusto Vigna Taglianti). Treviso: Canova, 2003. 3 vols., 460 + 1116 pp., ill., ISBN 88-8409-068-7. Nuncius, 20(2):501–502, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00493.

Turchetti:2005:BRA

[2613] Simone Turchetti. Book review: Amy Dahan and Dominique Pestre (eds.), Les sciences pour la guerre, 1940–1960. Paris: Editions de l’École des hautes tudes en sciences sociales, 2004. 401 pp., ISBN 2-7132-2015-7. Nuncius, 20(2):502–503, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00501.

Galuzzi:2005:BRA

[2614] Massimo Galuzzi. Book review: Andrea Del Centina e Alessandra Fiocca, l’Archivio di Guglielmo Libri dalla sua dispersione ai fondi della Biblioteca Moreniana, con il Catalogo del Fondo Carte Libri, a cura di G. Adini e M. L. Tanganelli (Collana Cultura e Memoria). Firenze: Olschki, 2004. xvi + 407 pp., ill., ISBN 88-222-5394-9. Nuncius, 20(2):503–504, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00510.

McClellan:2005:BRM

[2615] James E. McClellan. Book review: Marie Jaisson, Fondations, Prix et Subventions de l’Académie des Sciences (1916–1996). Turnhout: Brepols, 2003. 2 vols., 1363 pp., ISBN 2-503-52160-6. Nuncius, 20(2):504– 506, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00529.

Capocci:2005:BRC

[2616] Mauro Capocci. Book review: Claudia Mantovani, Rigenerare la società. l’Eugenetica in Italia dalle origini ottocentesche agli anni Trenta. Prefazione di Ernesto Galli della Loggia. Soveria Mannelli: Rubbettino, 2004. 404 pp., 25, ISBN 88-498-0769-4. Nuncius, 20(2):506–507, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00538.

Opitz:2005:BRJ

[2617] Donald L. Opitz. Book review: Jack Meadows, The Victorian Scientist: the Growth of a Profession. London: the British Library, 2004. vi + 202 pp., ill., £16.95, ISBN 0-7123-0894-6. Nuncius, 20(2):507–508, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00547.

Pogliano:2005:BRC

[2618] Claudio Pogliano. Book review: Carmela Morabito, La mente nel cervello. Un’introduzione storica alla neuropsicologia cognitiva. Roma/ Bari: Laterza, 2004. xvi + 172 pp., ill., ISBN 88-420-7472-1. Nuncius, 20(2):508–510, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00556.

Ciancio:2005:BRF

[2619] Luca Ciancio. Book review: Fabrizio Rasera (ed.), Le et del museo. Storia, uomini, collezioni del Museo Civico di Rovereto. Rovereto: Edizioni Osiride, 2004. 413 pp., ISBN 88-7498-028-0. Nuncius, 20(2):510–511, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00565.

Borrelli:2005:BRS

[2620] Antonio Borrelli. Book review: Salvatore Tommasi, Il rinnovamento della medicina in Italia. Introduzione di Marco Segala. Bibliografia a cura di Francesco Masedu (Cultura medica, 1). l’Aquila: Facolt di Medicina e Chirurgia, 2003. l + 313 pp. Nuncius, 20(2):511–512, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00574.

Santi:2005:G

[2621] Raffaella Santi. General. Nuncius, 20(2):512–513, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00583.

Tucci:2005:BRG

[2622] Pasquale Tucci. Book review: Graziella Buccellati and Benedetta Manetti (eds.), Energia, luce, movimento. Le collezioni storiche del Dipartimento di Elettrotecnica. Le collezioni storiche del Dipartimento di Elettrotecnica, a cura di Andrea Silvestri. Milano: Politecnico di Milano, 2002. 228 pp., ill., ISBN 88-7398-001-5. Nuncius, 20(2):513–515, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00592.

Guerrini:2005:BRR

[2623] Luigi Guerrini. Book review: Romano Paolo Coppini and Alessandro Tosi (a cura di), La Sapienza di Pisa. Pisa: Edizioni Plus: 2004. 208 pp., ill., ISBN 88-8492-238-0. Nuncius, 20(2):515–516, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00600.

Monti:2005:BRG

[2624] Maria Teresa Monti. Book review: Gilberto Corbellini, Breve storia delle idee di salute e malattia (Le Bussole). Roma: Carocci, 2004. 144 pp., 9,00. ISBN 88-430-3108-2. Nuncius, 20(2):516–517, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00619.

Strano:2005:BRJ

[2625] Giorgio Strano. Book review: John Mary Gribbin, The Men Who Measured the Universe. Crows Nest (NSW): Allen Unwin, 2004. x + 134 pp., ill., UK £6.99, ISBN 1-84046-536-0. Nuncius, 20(2):517–518, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00628.

Sankaran:2005:BRF

[2626] Neeraja Sankaran. Book review: Frederic Lawrence Holmes, Investigative Pathways. Patterns and Stages in the Careers of Experimental Scientists (New Haven and London: Yale Univ. Press, 2004), pp. xx + 225, ISBN 0-300-10075-2. Nuncius, 20(2):518–519, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00637.

Vista:2005:BRP

[2627] Michele Vista. Book review: Pierre Laszlo, Storia del sale. Miti, cammini, saperi. Roma: Donzelli, 2004. v + 183 pp., ill., 24,50, ISBN 88-7989-891-4. Nuncius, 20(2):519–521, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00646.

Vico:2005:BRM

[2628] Romana Martorelli Vico. Book review: Micrologus. Natura, Scienze e Società Medievali, XIII. La pelle umana/The Human Skin. Firenze: Sismel — Edizioni del Galluzzo, 2005. 772 pp., plates, 60. ISSN 1123-2560, ISBN 88-8450-143-1. Nuncius, 20(2):521–522, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00655.

Iannaccone:2005:BRJ

[2629] Isaia Iannaccone. Book review: Joseph Needham, Science and Civilisation in China, vol. VII:2, General Conclusions and Reflections, ed. K. G. Robinson. Cambridge: Cambridge University Press, 2004. + 283 pp., ISBN 0-521-08732-5. Nuncius, 20(2):522–524, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539105x00664.

Gibin:2005:BRM

[2630] Cinzio Gibin. Book review: Marco Piccolino, Lo zufolo e la cicala. Divagazioni galileiana tra la scienza e la sua storia. Torino: Bollati Boringhieri, 2005. 359 pp., ill., 26,00, ISBN 88-339-1612-X. Nuncius, 20 (2):524–525, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00673.

Bresadola:2005:BRR

[2631] Marco Bresadola. Book review: Roy Porter, Breve ma veridica storia della medicina occidentale (trad. it. Gian Carlo Brioschi e Melania Mascarino). Roma: Carocci, 2004. 229 pp., ill., 16,50, ISBN 88-430-3114-7. Nuncius, 20(2):525–526, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539105x00682.

Fano:2005:BRA

[2632] Vincenzo Fano. Book review: Antonio Sparzani, Relatività, quante storie. Un percorso scientifico-letterario tra relativo e assoluto. Torino: Bollati Boringhieri, 2003. 321 pp., ill., 30, ISBN 88-339-1457-7. Nuncius, 20 (2):526–528, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539105x00691.

Friesen:2005:BRP

[2633] John Friesen. Book review: Paul Wood (ed.), Science and Dissent in England, 1688–1945. Aldershot: Ashgate, 2004. xiv + 274 pp., ISBN 0-7546-3718-2. Nuncius, 20(2):528–529, ???? 2005. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00709.

Strano:2005:Nb

[2634] Giorgio Strano. Newsletter. Nuncius, 20(2):531–538, ???? 2005. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539105x00718.

Belle:2006:COM

[2635] Riccardo Bellè. Il corpus ottico Mauroliciano: Origini e sviluppo. (Italian) [The Optical Corpus Mauroliciano: Origins and development]. Nuncius, 21(1):7–29, ???? 2006. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539106x00014.

Boner:2006:KOC

[2636] Patrick J. Boner. Kepler on the origins of comets: Applying earthly knowledge to celestial events. Nuncius, 21(1):31–47, ????

2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00023.

Diana:2006:BDS

[2637] Esther Diana. La Biblioteca dell’Ospedale di Santa Maria Nuova di Firenze. Nuncius, 21(1):49–100, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00032.

Lamy:2006:CDC

[2638] Jérôme Lamy. La carte du ciel et la création du ’Bureau des Dames’ à l’Observatoire de Toulouse. Nuncius, 21(1):101–120, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00041.

Camerota:2006:HMS

[2639] Michele Camerota. Historia magistra scientiae. Nuncius, 21(1):121–132, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00050.

Shackelford:2006:RBE

[2640] Jole Shackelford. Reading bodies for epistemology: Art and science in the scientific revolution. Nuncius, 21(1):133–150, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00069.

Ackermann:2006:AMA

[2641] Silke Ackermann. Antiquity and middle ages. Nuncius, 21(1):151, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00078.

Gazzaniga:2006:BRI

[2642] Valentina Gazzaniga. Book review: Isabella Andorlini and Arnaldo Marcone, Medicina, medico e società nel mondoantico. Grassina: Le Monnier Università, 2004. v + 268 pp., ill., ISBN 88-00-86088-5. Nuncius, 21(1): 152–153, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00087.

Petit:2006:BRK

[2643] Caroline Petit. Book review: Karl-Heinz Leven, Antike Medizin: Ein Lexikon. München: C. H. Beck, 2005. xv + 967 pp., ISBN 3-406-52891-

0. Nuncius, 21(1):153–154, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00096.

Viano:2006:BRI

[2644] Cristina Viano. Book review: Ilaria Ramelli (ed.), Corpus Hermeticum. Edizione e commento di A. D. Nock e A.-J. Festugière, edizione dei testi ermetici copti e commento di I. Ramelli, testo greco, latino e copto (Il pensiero occidentale). Milano: Bompiani, 2005. 1627 pp., ISBN 88-452-3370-7. Nuncius, 21(1):154–155, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00104.

Pereira:2006:BRA

[2645] Michela Pereira. Book review: Anna Rodolfi, Il concetto di materia nell’opera di Alberto Magno. Firenze: Sismel, 2004. xx + 220 pp., ISBN 88-8450-126-1. Nuncius, 21(1):155–156, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00113.

DiPasquale:2006:BRL

[2646] Giovanni Di Pasquale. Book review: Lucio Russo, The forgotten revolution: how science was born in 300 BC and why it had to be reborn, with the collaboration of the translator, Silvio Levy. Berlin: Springer, 2004. 487 pp., ill., ISBN 3-540-20068-1. Nuncius, 21(1):156– 157, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00122.

Acerbi:2006:BRJa

[2647] Fabio Acerbi. Book review: J. Schönbeck, Euklid, Vita Mathematica (Band 12). Basel/Boston/Berlin: Verlag, 2003. x + 264 pp., ISBN 3-7643-6584-6. Nuncius, 21(1):157–159, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x00131.

Cavarra:2006:BRC

[2648] Berenice Cavarra. Book review: Cristina Viano (ed.), l’Alchimie et ses racines philosophiques. La tradition grècque et arabe (Histoire des Doctrines de l’Antiquité classique, Directeurs: Jean Pépin et Luc Brisson). Paris: Vrin, 2005. 240 pp., ISBN 2-7116-1754-8. Nuncius, 21(1):159–160, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00140.

Camerota:2006:REM

[2649] Michele Camerota. Renaissance and early modern science. Nuncius, 21 (1):160–162, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00159.

Strano:2006:BRA

[2650] Giorgio Strano. Book review: Anna Cassini, Gio: Domenico Cassini. Uno Scienziato del Seicento. Pinerolo: Arti Grafiche Alzani, 2003. 420 pp., ill. + plates. Nuncius, 21(1):162–163, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x00168.

Giglioni:2006:BRA

[2651] Guido Giglioni. Book review: Antonio Clericuzio, La macchina del mondo. Teorie e pratiche scientifiche dal Rinascimento a Newton. Roma: Carocci, 2005. 378 pp., ill., ISBN 88-430-3414-6. Nuncius, 21(1):163–164, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00177.

Guerrini:2006:BRS

[2652] Luigi Guerrini. Book review: Stéphane Garcia, Élie Diodati et Galilée. Naissance d’un réseau scientifique dans l’Europe du XVIIe siècle. Preface d’Isabelle Pantin. Firenze: Olschki, 2004. xx + 448 pp., ill., ISBN 88-222-5416-3. Nuncius, 21(1):164–165, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00186.

Betti:2006:BRA

[2653] Gian Luigi Betti. Book review: Atlee Naess, Galileo Galilei — When the World Stood Still. Berlin-New York: Springer, 2005. v + 221 pp., ill., ISBN 3-540-21961-7. Nuncius, 21(1):165–166, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00195.

Galuzzi:2006:BRM

[2654] Massimo Galuzzi. Book review: Marco Panza, Newton et les origines de l’analyse: 1664–1666. Paris: Blanchard, 2005. xv + 548 pp., ill., ISBN 2-85367-230-1. Nuncius, 21(1):166–167, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00203.

Dodd:2006:BRG

[2655] William Dodd. Book review: Gail Kern Paster, Humoring the Body. Emotions and the Shakespearean Stage. Chicago and London: Chicago University Press, 2004. xvi + 274 pp., ill., ISBN 0-226-64847-8. Nuncius, 21(1):167–169, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00212.

Haskell:2006:BRG

[2656] Yasmin Haskell. Book review: Gianna Pomata and Nancy Siraisi (eds.), Historia: Empiricism and Erudition in Early Modern Europe. Cambridge (Massachusetts): The MIT Press, 2005. vm + 490 pp., ill., ISBN 0-262-16229-6. Nuncius, 21(1):169–170, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00221.

DeVos:2006:BRL

[2657] Paula De Vos. Book review: Londa Schiebinger and Claudia Swan (eds.), Colonial Botany: Science, Commerce, and Politics in the Early Modern World. Philadelphia: University of Pennsylvania Press, 2005. vi + 346 pp., ill., ISBN 0-8122-3827-3. Nuncius, 21(1):170–171, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00230.

Clericuzio:2006:BRJ

[2658] Antonio Clericuzio. Book review: Jole R. Shackelford, A Philosophical Path for Paracelsian Medecine. The Ideas, Intellectual Context, and Influence of Petrus Severinus (1540–1602). Copenhagen: Museum Tusculanum press, 2004. 520 pp., ill., ISBN 87-7289-817-8. Nuncius, 21 (1):171–173, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00249.

Lopez:2006:BRE

[2659] Susana Gómez Lopez. Book review: Emanuele Zinato, La scienza dissimulata nel Seicento (premessa di Paolo Rossi). Napoli: Liguori, 2005. 140 pp., ISBN 88-207-3704-3. Nuncius, 21(1):173–174, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00258.

Messbarger:2006:BRJ

[2660] Rebecca Messbarger. Book review: Judith P. Zinsser (ed.), Men, Women, and the Birthing of Modern Science. DeKalb: Northern Illinois University Press, 2005. v + 216 pp., ISBN 978-0-87580-340-1. Nuncius, 21(1):174–175, ???? 2006. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539106x00267.

Mandelbrote:2006:E

[2661] Scott Mandelbrote. Enlightenment. Nuncius, 21(1):176–177, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00276.

Daniel:2006:BRR

[2662] Margócsy Daniel. Book review: Renée E. Kistemaker, Natalya Kopaneva, Debora J. Meijers and Georgy V. Vilibankhov (eds.), The Paper Museum of the Academy of Sciences in St. Petersburgh c. 1725–1760: Introduction and Interpretation (History of Sciences and Scholarship in the Netherlands 6). Amsterdam: Royal Netherlands Academy of Arts and Sciences, 2005. xii + 348 pp., ill., Dvd. ISBN 906844265. Nuncius, 21(1):177–178, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00285.

Pirillo:2006:BRR

[2663] Nestore Pirillo. Book review: Raphael Lagier, Les races humaines selon Kant. Paris: Presses Universitaires de France, 2004. vii + 200 pp., ISBN 978-2-13-054657-3. Nuncius, 21(1):178–179, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00294.

Bertrams:2006:BRJ

[2664] Kenneth Bertrams. Book review: Janis Langins, Conserving the Enlightenment: French Military Engineering from Vauban to the Revolution. Cambridge, Mass.: The MIT Press, 2004. xv + 532 pp., ISBN 0-262-12258-8. Nuncius, 21(1):179–180, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00302.

Ciancio:2006:BRM

[2665] Luca Ciancio. Book review: Maria Toscano, Alberto Fortis nel Regno di Napoli: naturalismo e antiquaria 1783–1791, Bari: Cacucci Editore, 2004. 182 pp., ISBN 88-8422-362-8. Nuncius, 21(1):180–181, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00311.

Morus:2006:MCS

[2666] Iwan Rhys Morus. Modem and contemporary science. Nuncius, 21(1): 182, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00320.

Borrelli:2006:BRR

[2667] Antonio Borrelli. Book review: Raffaele Colapietra, Per la biografia di Salvatore Tommasi. Con un intervento di Ferdinando Di Orio (Deputazione Abruzzese di Storia Patria. Studi e testi, 28). l’Aquila: Edizioni Libreria Colacchi, 2004. 129 pp., ISBN 88-88676-19-8. Nuncius, 21(1): 182–184, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00339.

Zylberman:2006:BRG

[2668] Patrick Zylberman. Book review: Giorgio Cosmacini, Il medico materialista: vita e pensiero di Jakob Moleschott. Roma: Laterza, 2005. x + 189 pp., ISBN 88-420-7516-7. Nuncius, 21(1):184–185, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00348.

Battimelli:2006:BRJ

[2669] Giovanni Battimelli. Book review: James W. Cronin (ed.), Fermi Remembered. Chicago: University of Chicago Press, 2004. xii + 287 pp., ill., ISBN 0-226-12111-9. Nuncius, 21(1):185–186, ????

2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00357.

Carnemolla:2006:BRM

[2670] Stefania Elena Carnemolla. Book review: Marco Di Giovanni, Scienza e potenza. Miti della guerra moderna, istituzioni scientifiche e politica di massa nell’Italia fascista. 1935–1945. Torino: Zamorani, 2005. 308 pp., ISBN 88-7158-132-6. Nuncius, 21(1):186–187, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00366.

Capocci:2006:BRA

[2671] Mauro Capocci. Book review: Alan H. Goodman, Deborah Heath and M. Susan Lindee (eds.), Genetic Nature/Culture: Anthropology and Science Beyond the Two-Culture Divide. Berkeley: University of California Press, 2003. xv + 311 pp., ill., ISBN 0-520-23793-5. Nuncius, 21(1):187– 188, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00375.

Ferraresi:2006:BRP

[2672] Alessandra Ferraresi. Book review: Paola Govoni, Un pubblico per la scienza. La divulgazione scientifica nell’Italia in formazione. Roma: Carocci, 2002. 351 pp., ill., ISBN 88-430-2321-7. Nuncius, 21(1):188– 189, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00384.

Pogliano:2006:BRM

[2673] Claudio Pogliano. Book review: Michael Hagner, Geniale Gehirne. Zur Geschichte der Elitegehirnforschung. Göttingen: Wallstein Verlag, 2004. 376 pp., ill., ISBN 3-89244-649-0. Nuncius, 21(1):189–190, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00393.

Turchetti:2006:BRG

[2674] Simone Turchetti. Book review: Gerald Holton, Victory and Vexation in Science. Einstein, Bohr, Heisenberg and Others. Cambridge, Mass.: Harvard University Press, 2005. xiv + 230 pp., ill., ISBN 0-674-01519-3. Nuncius, 21(1):191–192, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00401.

Abbri:2006:BRO

[2675] Ferdinando Abbri. Book review: Olof Ljungström, Oscariansk antropologi. Etnografi, förhistoria och rasforskning under sent 1800tal, Hedemora, Gidlunds Förlag, 2004, 496 pp., ill., ISBN 91-7844-635-X. Nuncius, 21(1):192–193, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00410.

Pancaldi:2006:BRI

[2676] Giuliano Pancaldi. Book review: Iwan Rhys Morus, When Physics Became King. Chicago and London: The University of Chicago Press, 2005. xii + 303 pp., ill., ISBN 0-226-54202-5. Nuncius, 21(1):193–194, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00429.

Muller-Wille:2006:BRC

[2677] Staffan Müller-Wille. Book review: Claudio Pogliano, l’Ossesione della razza: Antropologia e genetica nel XX secolo. Pisa: Edizioni della Normale, 2005. 582 pp., ISBN 88-7642-146-7. Nuncius, 21(1):194–195, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00438.

Piccolino:2006:BRR

[2678] Marco Piccolino. Book review: Richard Rapport, Nerve endings, the discovery of the synapse. New York: Norton, 2005. 240 pp., ill., ISBN 0-393-06019-5; Elliot S. Valenstein, The war of the soups and the sparks, the discovery of neurotransmitters and the dispute over how nerves communicate. New York: Columbia University Press, 2005. xviii + 237 pp., ill., ISBN 0-231-13588-2. Nuncius, 21(1):196–198, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00447.

Riccini:2006:BRG

[2679] Raimonda Riccini. Book review: Giorgio Zanrosso, Dirigibili 1852– 1937. Dal Giffard n. 1 all’Hindenburg. Vicenza: Editrice Veneta, 2004. 254 pp., ill., ISBN 88-8449-225-4. Nuncius, 21(1):198–199, ????

2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00456.

Gingerich:2006:G

[2680] Owen Gingerich. General. Nuncius, 21(1):199–200, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x00465.

Heering:2006:BRH

[2681] Peter Heering. Book review: Hasok Chang, Inventing Temperature. Measurement and Scientific Progress (Oxford Studies in the Philosophy of Science). Oxford: Oxford University Press, 2004. 304 pp., ill., ISBN 0-19-517127-6. Nuncius, 21(1):200–202, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00474.

Borck:2006:BRM

[2682] Cornelius Borck. Book review: Mariacarla Gadebusch Bondio, Medizinische Ästhetik. Kosmetik und plastische Chirurgie zwischen Antike und früher Neuzeit. München: Fink Verlag, 2005. 238 pp., ill., ISBN 3-7705-4101-4. Nuncius, 21(1):202–203, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00483.

Meschiari:2006:BRB

[2683] Alberto Meschiari. Book review: Brian Bracegirdle, A catalogue of the Microscopy Collections at the Science Museum, London. CD-ROM available from Savona Books, 400 Seawall Lane, Haven Sands, North Cotes, Grimsby DN36 5XE, UK. Nuncius, 21(1):203–204, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00492.

Marr:2006:BRL

[2684] Alexander Marr. Book review: Lorraine Daston (ed.), Things that Talk: Object Lessons from Art and Science. New York: Zone Books, 2004. 448 pp., ISBN 1-890951-43-9. Nuncius, 21(1):204–205, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00500.

Candito:2006:BRM

[2685] Cristina Càndito. Book review: Marco Franciosi (ed.), Prospettiva e geometriadello spazio (Quaderni del Centro Studi Enriques, 2). Sarzana: Agorà edizioni, 2005. 280 pp., ill., ISBN 88-87218-98-6. Nuncius, 21 (1):205–207, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00519.

Motta:2006:BRY

[2686] Franco Motta. Book review: Yasmin A. Haskell, Loyola’s Bees. Ideology and Industry in Jesuit Latin Didactic Poetry. Oxford and New York: Oxford University Press — The British Academy, 2003 (British Academy Postdoctoral Fellowship Monographs). x + 354 pp., ill., ISBN 0-19-726284-8. Nuncius, 21(1):207–208, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00528.

Sergio:2006:BRT

[2687] Emilio Sergio. Book review: Thomas Holden, The Architecture of Matter. Galileo to Kant. Oxford: Clarendon Press, 2004. 305 pp., ISBN 0-19-926326-4. Nuncius, 21(1):208–209, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00537.

Wormbs:2006:BRT

[2688] Nina Wormbs. Book review: Thomas P. Hughes, Human-built world: how to think about technology and culture. Chicago: The University of Chicago Press, 2004. x + 223 pp., ill., ISBN 0-226-35933-6. Nuncius, 21 (1):209–210, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00546.

Buttner:2006:BRW

[2689] Jochen Büttner. Book review: Wolfgang Lefèvre (ed.), Picturing Machines, 1400–1700. Cambridge: The MIT Press, 2004. vi + 347 pp., ISBN 0-262-12269-3. Nuncius, 21(1):210–212, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00555.

Giacomoni:2006:BRM

[2690] Paola Giacomoni. Book review: Michel Meulders, Helmholtz. Dal secolo dei Lumi alle neuroscienze. Torino: Bollati Boringhieri, 2005.

304 pp., ill., ISBN 88-339-1579-4. Nuncius, 21(1):212–213, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00564.

Tomas:2006:BRG

[2691] José Pardo Tomás. Book review: Giuseppe Olmi (ed.), Rappresentare il corpo. Arte e anatomia da Leonardo all’Illuminismo. Bologna: Bononia University Press, 2004. 326 pp. ill., ISBN 88-7395-094-9. Nuncius, 21 (1):213–214, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00573.

deCeglia:2006:BRN

[2692] Francesco Paolo de Ceglia. Book review: Nico Pitrelli and Giancarlo Sturloni (eds.), La comunicazione della scienza. Atti del I e II Convegno Nazionale. Introduzione di Pietro Greco, Prolusione di Paolo Rossi. Roma: Zadigroma Editore, 2004. 368 pp., ISBN 88-88734-09-0. Nuncius, 21(1):214–215, ???? 2006. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539106x00582.

Micheletti:2006:BRC

[2693] Mario Micheletti. Book review: Carlo Augusto Viano, Le imposture degli antichi e i miracoli dei moderni. Torino: Einaudi, 2005. 157 pp., ISBN 88-06-16098-2. Nuncius, 21(1):215–217, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00591.

Strano:2006:Na

[2694] Giorgio Strano. Newsletter. Nuncius, 21(1):219–228, ???? 2006. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x00609.

Segre:2006:LE

[2695] Michael Segre. Letter to the Editor. Nuncius, 21(1):229, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00618.

Miniati:2006:P

[2696] Mara Miniati. Preface. Nuncius, 21(2):247–249, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x00627.

Contardi:2006:OSI

[2697] Simone Contardi. The origins of a scientific institution. Felice Fontana and the birth of the Real Museo Di Fisica e Storia Naturale Di Firenze. Nuncius, 21(2):251–263, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00636.

Cipriani:2006:FFF

[2698] Curzio Cipriani. Felice Fontana and the formation of the naturalistic collections of the Imperial Royal Museum of Physics and Natural History of Florence. Nuncius, 21(2):265–294, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00645.

Marker:2006:AMS

[2699] Anna Märker. The anatomical models of la specola: Production, uses, and reception. Nuncius, 21(2):295–321, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x00654.

Bertucci:2006:PUS

[2700] Paola Bertucci. Public utility and spectacular display: the physics cabinet of the Royal museum in Florence (1775). Nuncius, 21(2):323–336, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00663.

Mazzolini:2006:VFR

[2701] Renato G. Mazzolini. Visitors to Florence’s r. museum of physics and natural history from September 1784 to October 1785. Nuncius, 21(2): 337–348, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00672.

Olmi:2006:WCI

[2702] Giuseppe Olmi. ‘A wonderful collection indeed!’: the Royal Museum of Florence in the testimony of two travellers. Nuncius, 21(2):349–368, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00681.

Vaccari:2006:O

[2703] Ezio Vaccari. Obituary. Nuncius, 21(2):369–370, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x00690.

Kahn:2006:CDC

[2704] Didier Kahn. Le contexte des débats sur le clonage: Art et nature, de Zosime de Panopolis à Goethe. Nuncius, 21(2):371–378, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00708.

Maravela-Solbakk:2006:AMA

[2705] Anastasia Maravela-Solbakk. Antiquity and middle ages. Nuncius, 21 (2):379–380, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00717.

Doody:2006:BRM

[2706] Aude Doody. Book review: Mary Beagon (ed.), The Elder Pliny on the Human Animal (Natural History, book 7). Oxford: Clarendon Press, 2005. xxii + 515 pp., ISBN 0-19-815065-2. Nuncius, 21(2):380–381, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00726.

Bernardoni:2006:BRR

[2707] Andrea Bernardoni. Book review: Robert Bork, Scott Montgomery, Carol Neuman De Vegvar, Ellen Shortell and Steven Walton (eds.), De Re Metallica, the Uses of Metal in the Middle Ages. Aldershot/Burlington, VT: Ashgate, 2005. 442 pp., ill., ISBN 0-7546-5048-0. Nuncius, 21(2): 381–382, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00735.

Acerbi:2006:BRJb

[2708] Fabio Acerbi. Book review: Jean Christianidis (ed.), Classics in the History of Greek Mathematics (Boston Studies in the Philosophy of Science 240). Dordrecht/Boston/London: Kluwer, 2004. x + 461 pp., EUR176.80, ISBN 1-4020-0081-2. Nuncius, 21(2):382–383, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00744.

Hunt:2006:BRP

[2709] Anthony Hunt. Book review: Piers D. Mitchell, Medicine in the Crusades: Warfare, Wounds and the Medieval Surgeon. Cambridge: Cambridge University Press, 2004. ix + 293 pp., ill., ISBN 0-521-84455-X. Nuncius, 21(2):383–385, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00753.

Acerbi:2006:BRR

[2710] Fabio Acerbi. Book review: Roshdi Rashed, Geometry and Dioptrics in Classical Islam. London: Al-Furqän Islamic Heritage Foundation, 2005. xiv + 1178 + vi pp., ISBN 1-873992-99-8. Nuncius, 21(2):385–386, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00762.

Ralley:2006:REM

[2711] Robert Ralley. Renaissance and early modern science. Nuncius, 21(2): 386–387, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00771.

Gatto:2006:BRV

[2712] Romano Gatto. Book review: Veronica Gavagna e Carlotta Leone (eds.), Giovan Battista della Porta: Elementorum curvilineorum libri tres. Napoli e Roma: Edizioni scientifiche italiane, 2005. xxxiv + 178 pp., ill. ISBN 88-495-1060-8. Nuncius, 21(2):387–388, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00780.

Luthy:2006:BRL

[2713] Christoph Lüthy. Book review: Luigi Guerrini (ed.), Federico Cesi. Apiarium. Testo e traduzione. Vol. I: Roma: Accademia Nazionale dei Lincei, 2005. xx + 265 pp. ISBN 88-218-0950-1; Vol. II: Roma: Accademia Nazionale dei Lincei, 2006. xxiv + 135 pp. ISBN 88-218-0963-3. Nuncius, 21(2):388–390, ???? 2006. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539106x00799.

Trabucco:2006:BRL

[2714] Oreste Trabucco. Book review: Luigi Ingaliso (ed.), Giovan Filippo Ingrassia: Informatione del pestifero et contagioso morbo. Milano: Angeli, 2005. 650 pp., ISBN 88-464-7290-X. Nuncius, 21(2):390–392, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00807.

Abbri:2006:BRG

[2715] Ferdinando Abbri. Book review: Geert H. Janssen, Creaturen van de macht. Patronage bij Willem Frederik van Nassau (1613–1664). Amsterdam: Amsterdam University Press, 2005. 304 pp., ill., ISBN 50-5356-787-6. Nuncius, 21(2):392–393, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00816.

Nenci:2006:BRI

[2716] Elio Nenci. Book review: Ian Maclean (ed.), Girolamo Cardano: De libris propriis. The editions of 1544, 1550, 1557, 1562, with supplementary material. Milano: Franco Angeli, 2004. 400 pp., ISBN 88-464-5449-9. Nuncius, 21(2):393–394, ???? 2006. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539106x00825.

Alexander:2006:BRN

[2717] Amir R. Alexander. Book review: Noel Malcolm and Jacqueline Stedall (eds.), John Pell (1611–1685) and His Correspondence with Sir Charles Cavendish. The Mental World of an Early Modern Mathematician. Oxford: Oxford University Press, 2004. 664 pp., ISBN 0-19-856484-8. Nuncius, 21(2):394–395, ???? 2006. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539106x00834.

Henry:2006:BRW

[2718] John Henry. Book review: William R. Newman, Atoms and Alchemy: Chymistry and the Experimental Origins of the Scientific Revolution. Chicago: University of Chicago Press, 2006. xiii + 250 pp., ill., ISBN 0-226-57696-5. Nuncius, 21(2):395–396, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00843.

Egmond:2006:BRD

[2719] Florike Egmond. Book review: David Pegler and David Freedberg (eds.), The paper museum of Cassiano Dal Pozzo. Series B-Natural History. Part II: Fungi. London: Royal Collection Enterprises, 2005. 3 vols. viii + 1028 pp., ill., ISBN 1-905375-05-0. Nuncius, 21(2):396–398, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00852.

Guerrini:2006:BRM

[2720] Luigi Guerrini. Book review: Mauro Pesce, l’Ermeneutica biblica di Galileo e le due strade della teologia cristiana (Uomini e dottrine, 43). Roma: Edizioni di Storia e Letteratura, 2005. vii + 240 pp., ISBN 88-8498-207-3. Nuncius, 21(2):398–399, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00861.

Bresadola:2006:BRM

[2721] Marco Bresadola. Book review: Mauro Simonazzi, La malattia inglese. La melanconia nella tradizione filosofica e medica dell’Inghilterra moderna. Bologna: Il Mulino, 2004. 455 pp., ISBN 88-15-09759-7. Nuncius, 21(2):399–400, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00870.

Parigi:2006:BRM

[2722] Silvia Parigi. Book review: Monica Uguaglia (ed.), Leonardo Garzoni: Trattati della calamita. Milano: Angeli, 2005. 352 pp., ISBN 88-464-7132-6. Nuncius, 21(2):400–401, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00889.

Tognoni:2006:E

[2723] Federico Tognoni. Enlightenment. Nuncius, 21(2):401–403, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00898.

Vaccari:2006:BRM

[2724] Valentina Vaccari. Book review: Mary Ellen Bowden and Lisa Rosner, Joseph Priestley, Radical Thinker: A Catalogue to Accompany the Exhibit at the Chemical Heritage Foundation Commemorating the 200th Anniversary of the Death of Joseph Priestley. Philadelphia: Chemical Heritage Foundation, 2005. 72 pp., ill., ISBN 0-941901-38-6. Nuncius, 21(2):403–404, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00906.

Peiffer:2006:BRF

[2725] Jeanne Peiffer. Book review: Frans A. Cerulus (ed.), Die Werke von Daniel Bernoulli (Band 8. Technologie II). Basel & c.: Birkhäuser Verlag, 2004. xx + 360 pp., ill., ISBN 3-7643-7007-6. Nuncius, 21(2):404–405, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00915.

Zanier:2006:BRI

[2726] Sonia Zanier. Book review: Ivano Dal Prete (ed.), Antonio Vallisneri, Miglioramenti e correzioni d’alcune esperienze ed osservazioni del Signor Redi; Mauro De Zan (ed.), Carlo Francesco Cogrossi — Antonio Vallisneri, Nuova Idea del male contagioso de’ buoi (Sezione I, Volume 5 dell’Edizione Nazionale delle opere di Antonio Vallisneri). Firenze: Olschki, 2005. vii + 173 pp., ISBN 88-222-5484-8. Nuncius, 21(2):406– 407, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00924.

Ciardi:2006:BRR

[2727] Marco Ciardi. Book review: Raffaele Giura Longo e Pasquale Longo (eds.), La Expedición Malaspina, 1789–1794. Bari: Cacucci, 2002–2004. vol. I: Da Cadice ad Acapulco (2002), 186 pp.; vol. II: Dalle Filippine a Cadice (2004), 186 pp., ISBN 88-8422-180-3. Nuncius, 21(2):407–408, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00933.

Roche:2006:BRR

[2728] John Roche. Book review: Roger Hahn, Pierre Simon Laplace (1749– 1827): a determined scientist. Cambridge (Mass.): Harvard University Press, 2005. x + 310 pp., ISBN 0-674-01892-3. Nuncius, 21(2):408–410, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00942.

Ratcliff:2006:BRT

[2729] Marc J. Ratcliff. Book review: Thierry Hoquet, Buffon: histoire naturelle et philosophie. Paris: Honoré Champion, 2005. 809 pp., ISBN 2-7453-1246-4. Nuncius, 21(2):410–411, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x00951.

Bandinelli:2006:BRM

[2730] Angela Bandinelli. Book review: Mi Gyung Kim, Affinity, That Elusive Dream. A Genealogy of the Chemical Revolution. Cambridge, Massachusetts: The MIT Press, 2003. 599 pp., ISBN 0-262-11273-6. Nuncius, 21(2):411–412, ???? 2006. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539106x00960.

Ostwald:2006:BRB

[2731] Jamel Ostwald. Book review: Brett D. Steele and Tamera Dorland (eds.), The Heirs of Archimedes. Science and the Art of War through the Age of Enlightenment (Dibner Institute studies in the history of science and technology). Cambridge (Mass.): The MIT Press, 2005. 397 pp.: ill., ISBN 0-262-19516-X. Nuncius, 21(2):412–413, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00979.

Monti:2006:BRH

[2732] Maria Teresa Monti. Book review: Hubert Steinke, Irritating Experiments. Haller’s Concept and the European Controversy on Irritability and Sensibility. 1750–1790. Amsterdam: Rodopi, 2005. 354 pp., ill., ISBN 90-420-1852-6. Nuncius, 21(2):413–415, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x00988.

Navarro:2006:MCS

[2733] Jaume Navarro. Modern and contemporary science. Nuncius, 21(2): 415–416, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x00997.

Poggi:2006:BRT

[2734] Stefano Poggi. Book review: Thomas Bach and Olaf Breidbach (eds.), Naturphilosophie nach Schelling, Stuttgart-Bad Cannstatt: Verlag, 2005. xii + 836 pp., ISBN 3-7728-2255-X. Nuncius, 21(2):416–417, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01004.

Beck:2006:BRC

[2735] Naomi Beck. Book review: Claude Blanckaert, La nature de la société. Organicisme et sciences sociales au XIXe siècle. Paris: l’Harmattan, 2004. 158 pp., ISBN 2-7475-7691-4. Nuncius, 21(2):418–419, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01013.

Koehler:2006:BRC

[2736] Peter Koehler. Book review: Cornelius Borck, Hirnströme. Eine Kulturgeschichte der Elektroenzephalographie. Göttinger: Wallstein Verlag, 2005. 382 pp., ill., ISBN 3-89244-893-0. Nuncius, 21(2):419–420, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01022.

Beetham:2006:BRG

[2737] Margaret Beetham. Book review: Geoffrey Cantor, Gowan Dawson, Graeme Gooday, Richard Noakes, Sally Shuttleworth and Jonathan R. Topham, Science in the Nineteenth-Century periodical: reading the magazine of nature (Cambridge studies in Nineteenth Century literature and culture; 45). Cambridge: Cambridge University Press, 2004. xi + 329 pp.: ill., ISBN 0-521-83637-9. Nuncius, 21(2):420–421, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01031.

Pogliano:2006:BRF

[2738] Claudio Pogliano. Book review: Francesco Cassata, Molti, sani e forti. l’eugenetica in Italia, Torino: Bollati Boringhieri, 2006. 396 pp., ISBN 88-339-1644-8. Nuncius, 21(2):421–422, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x01040.

Turchetti:2006:BRH

[2739] Simone Turchetti. Book review: Harry Collins, Gravity’s Shadow. The Search for Gravitational Waves. Chicago & London: University of Chicago Press, 2004. xxii + 870 pp., ISBN 0-226-11378-7. Nuncius, 21 (2):422–423, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x01059.

Gattei:2006:BRS

[2740] Stefano Gattei. Book review: Steve W. Fuller, Kuhn vs Popper: the struggle for the soul of science (Revolutions in Science). Duxford: Icon Books, 2003. iv + 236 pp., ISBN 1-84046-468-2. Nuncius, 21(2):423–425, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01068.

Maiocchi:2006:BRG

[2741] Roberto Maiocchi. Book review: Giovanni Paoloni e Raffaella Simili (eds.), I Lincei nell’Italia unita. Roma: Bretschneider, 2004. 255 pp., ISBN 88-7689-192-7. Nuncius, 21(2):425–426, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x01077.

Faracovi:2006:BRP

[2742] Ornella Pompeo Faracovi. Book review: Paolo Parrini, Filosofia e scienza nell’Italia del Novecento. Figure, correnti, battaglie. Milano: Guerini e associati, 2004. 390 pp., ISBN 88-8335-539-3. Nuncius, 21(2):426–427, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01086.

Marchis:2006:BRP

[2743] Vittorio Marchis. Book review: Pietro Redondi e Paola Zocchi (eds.), Milano 1906. l’Esposizione internazionale del Sempione. La scienza, La cittä, la vita. Milano: Guerini e Associati, 2006. 268 pp., ISBN 88-8335-777-9. Nuncius, 21(2):427–428, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x01095.

Bensaude-Vincent:2006:G

[2744] Bernadette Bensaude-Vincent. General. Nuncius, 21(2):429–430, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01103.

Nielsen:2006:BRE

[2745] Henry Nielsen. Book review: Enrico Baraldi, Hjalmar Fors and Anders Houltz (eds.), Taking Place. The spatial contexts of science, technology and business. Sagamore Beach: Watson Publishing International, 2006. xxii + 402 pp., ISBN 0-88135-252-7. Nuncius, 21(2):430–431, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01112.

Abbri:2006:BRR

[2746] Ferdinando Abbri. Book review: Ruth Barcan, Nudity. A Cultural Anatomy (Dress, Body, Culture). Oxford-New York: Berg, 2004. xii + 308 pp., ill., ISBN 1-85973-872-9. Nuncius, 21(2):431–432, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01121.

Golinski:2006:BRP

[2747] Jan Golinski. Book review: Peter J. Bowler and Iwan Rhys Morus, Making modern science: A historical survey. Chicago: The University of Chicago Press, 2005. viii + 529 pp., ill., ISBN 0-226-06860-9. Nuncius, 21(2):432–433, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x01130.

Zacchini:2006:BRP

[2748] Simone Zacchini. Book review: Pierre Hadot, Il velo di Iside. Storia dell’i-dea di natura. Torino: Einaudi, 2006. xvii + 332 pp., ill., ISBN 88-06-17235-2. Nuncius, 21(2):433–435, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x01149.

Patiniotis:2006:BRE

[2749] Manolis Patiniotis. Book review: Ekmeleddin Ihsanoğlu, Science, technology and learning in the Ottoman Empire. Western Influence, local institutions, and the transfer of knowledge. Aldershot: Ashgate, 2004. 352 pp., ill., ISBN 0-86078-924-1. Nuncius, 21(2):435–436, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01158.

Borck:2006:BRB

[2750] Cornelius Borck. Book review: Bettina Von Jagow and Florian Steger (eds.), Literatur und Medizin: Ein Lexikon. Göttingen: Vandenhoek & Ruprecht 2005. 492 pp., ISBN 3-525-21018-3. Nuncius, 21(2):436–437, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01167.

Eddy:2006:BRF

[2751] Matthew Eddy. Book review: Frank A. J. L. James, ’The Common Purposes of Life’. Science and society at the Royal Institution of Great Britain. Aldershot: Ashgate, 2002. xiv + 456 pp., ill., ISBN 0-7546-0960-X. Nuncius, 21(2):437–438, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x01176.

Toscano:2006:BRK

[2752] Anna Toscano. Book review: Kennet F. Kiple (ed.), The Cambridge historical dictionary of disease. Cambridge: Cambridge University Press, 2003. xiii + 412 pp., ISBN 0-521-53026-1. Nuncius, 21(2):438–439, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01185.

Numerico:2006:BRJ

[2753] Teresa Numerico. Book review: Jeremy M. Norman (ed.), From Gutenberg to the Internet. A Sourcebook on the History of Information Technology. Novato (CA): historyofscience.com, 2005. xvi + 900 pp., ill., ISBN 0-930405-87-0. Nuncius, 21(2):439–440, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01194.

Laurenza:2006:BRG

[2754] Domenico Laurenza. Book review: Giuseppe Olmi e Giuseppe Papagno (eds.), La natura e il corpo: studi in memoria di Attilio Zanca. Firenze: L. S. Olschki, 2006. xiv + 244 pp., ISBN 88-222-5467-8. Nuncius, 21 (2):440–442, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539106x01202.

Conforti:2006:BRM

[2755] Maria Conforti. Book review: Margaret Pelling and Scott Mandelbrote (eds.), The Practice of Reform in Health, Medicine and Science, 1500–2000. Essays for Charles Webster. Aldershot: Ashgate, 2005. xv + 376 pp., ISBN 0-7546-3933-9. Nuncius, 21(2):442–443, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01211.

DalPrete:2006:BRM

[2756] Ivano Dal Prete. Book review: Martin J. S. Rudwick, The New Science of Geology. Studies in the Earth Sciences in the Age of Revolution. Aldershot: Ashgate: 2004. xviii + 316 pp., ISBN 0-86078-958-6. Nuncius, 21(2):443–444, ???? 2006. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539106x01220.

Gauvin:2006:BRA

[2757] Jean-François Gauvin. Book review: Anthony J. Turner, Après Gassendi. Son influence et sa reputation, essai, avec l’histoire des collections scientifiques et un catalogue des instruments et appareils concernant les sciences exactes appartenant au Musée Gassendi à Digne-les-Bains. Digne-les-Bains: Musée Gassendi, 2006. 323 pp., ISBN 2-912450-07-1. Nuncius, 21(2):444–445, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539106x01239.

Strano:2006:Nb

[2758] Giorgio Strano. Newsletter. Nuncius, 21(2):447–453, ???? 2006. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539106x01248.

Anonymous:2006:NC

[2759] Anonymous. Notes on contributors. Nuncius, 21(2):455–460, ???? 2006. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539106x01257.

Hannah:2007:MUB

[2760] Robert Hannah. The meridiana of Ulugh Beg in Hagia Sophia, Constantinople. Nuncius, 22(1):7–14, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00013.

Mertens:2007:CBH

[2761] Joost Mertens. Charles Bourgeois (1759–1832), his antinewtonian colour theory and the reconcilement of art and science. Nuncius, 22(1):15–48, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00022.

Locher:2007:LTS

[2762] Fabien Locher. L’Émergence d’une techno-science de l’environnement océanique (1860–1914). Nuncius, 22(1):49–68, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539107x00031.

Toscano:2007:ADH

[2763] Marco Toscano. An ‘abstruse’ doctrine: the historical–philosophical meaning of Einstein’s Relativity in Federigo Enriques. Nuncius, 22(1): 69–95, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00040.

DePalma:2007:WMN

[2764] Armando De Palma and Germana Pareti. The ways of metaphor in neuroscience, or being on the right or wrong track. Nuncius, 22(1):97–124, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00059.

Kleeberg:2007:DHM

[2765] Bernhard Kleeberg. Dr. Haeckel and Mr. Hyde. New works on 19th and 20th-Century German biology and anthropology. Nuncius, 22(1): 125–138, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00068.

Acerbi:2007:AMAa

[2766] Fabio Acerbi. Antiquity and middle ages. Nuncius, 22(1):139–140, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00077.

Perilli:2007:BRI

[2767] Lorenzo Perilli. Book review: Ilsetraut Hadot, Arts libraux et philosophie dans la pensé antique: contribution l’histoire de l’éducation et de la culture dans l’antiquité. Seconde édition revue et considérablement augmenté. Paris: Libr. philosophique J. Vrin, 2005. 578 pp., ISBN 2-7116-1823-4. Nuncius, 22(1):140–141, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00086.

Iliffe:2007:REM

[2768] Robert Iliffe. Renaissance and early modern science. Nuncius, 22(1): 142–143, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00095.

Caleca:2007:BRS

[2769] Antonino Caleca. Book review: Simona Cremante, Leonardo da Vinci. Artista, scienziato, inventore. Firenze-Milano: Giunti, 2005. 640 pp., ISBN 88-09-03811-8. Nuncius, 22(1):143–144, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00103.

Bernardoni:2007:BRM

[2770] Andrea Bernardoni. Book review: Marie-Claude Duprez-Masson, Technique, mot et Image. Le de re metallica d’Agricola. Turnhout: Brepols, 2006. 520 pp., ISBN 978-2-503-52365-1. Nuncius, 22(1):144–145, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00112.

Gmez:2007:BRS

[2771] Susana Gmez. Book review: Saul Fisher, Pierre Gassendi’s Philosophy and Science. Atomism for Empiricists. Leyden-Boston: Brill, 2005. xxviii + 436 pp., ISBN 90-04-11996-5. Nuncius, 22(1):145–146, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00121.

Watson:2007:BRM

[2772] Richard Watson. Book review: Matthew L. Jones, The Good Life in the Scientific Revolution. Descartes, Pascal, Leibniz, and the Cultivation of Virtue. Chicago and London: The University of Chicago Press, 2006. xiii + 384 pp., ISBN 0-226-40955-4. Nuncius, 22(1):146–148, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00130.

Guerrini:2007:BRL

[2773] Luigi Guerrini. Book review: Lorella Mangani e Giuseppe Martini, La biblioteca di Francesco Redi e della sua famiglia. Catalogo. Arezzo: L. P. Grafiche, 2006. 686 pp., ill., ISBN 88-86830-02-5. Nuncius, 22(1): 148–149, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00149.

Donato:2007:BRE

[2774] Maria Pia Donato. Book review: Ernan McMullin (ed.), The Church and Galileo. Notre Dame: University of Notre Dame Press, 2005. xii + 391 pp., ill., ISBN 0-268-03484-2. Nuncius, 22(1):149–150, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00158.

Margocsy:2007:BRB

[2775] Daniel Margocsy. Book review: Brian W. Ogilvie, The Science of Describing: Natural History in Renaissance Europe. Chicago and London: The University of Chicago Press. 2006. xvi + 385 pp., ISBN 0-226-62087-5. Nuncius, 22(1):150–152, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00167.

Abbri:2007:BRC

[2776] Ferdinando Abbri. Book review: Claude V. Palisca, Music and Ideas in the Sixteenth and Seventeenth Centuries. Urbana and Chicago: University of Illinois Press 2006. x + 302 pp., ISBN 0-252-03156-3. Nuncius, 22(1):152–153, ???? 2007. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539107x00176.

Bennett:2007:BRS

[2777] Jim Bennett. Book review: Stephen Pumfrey, Latitude and the Magnetic Earth. Duxford: Icon Books Ltd, 2003. 249 pp., ISBN 1-84046-486-0. Nuncius, 22(1):153–154, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00185.

Ratcliff:2007:E

[2778] Marc Ratcliff.  Enlightenment.  Nuncius, 22(1):154–155, ????

2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00194.

Vaccari:2007:BRM

[2779] Valentina Vaccari. Book review: Malcolm Dick (ed.), Joseph Priestley and Birmingham. Studley: Brewin Books, 2005. 128 pp., ill., ISBN 1-85858-269-5. Nuncius, 22(1):155–156, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539107x00202.

Bertucci:2007:BRG

[2780] Paola Bertucci. Book review: Gino L. Di Mitri, Storia biomedica del tarantismo nel XVIII secolo. Firenze: Olschki, 2006. xxiv + 322 pp., ISBN 88-222-5508-9. Nuncius, 22(1):156–157, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00211.

Guicciardini:2007:BRR

[2781] Niccol Guicciardini. Book review: Robert Iliffe, Milo Keynes and Rebekah Higgitt (eds.), Early Biographies of Isaac Newton, 1660–1885. Vol. 1: Eighteenth Century Biography of Isaac Newton: the Unpublished Manuscripts and Early Texts. Vol. 2: Nineteenth-Century Biography of Isaac Newton: Public Debate and Private Controversy. London: Pickering and Chatto, 2006. Vol. 1: lxxx + 387 pp., vol. 2: xliv + 420 pp., ISBN-10 1-85196-778-8. Nuncius, 22(1):157–159, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00220.

Zanzi:2007:BRJ

[2782] Luigi Zanzi. Book review: Jean Gaudant, Dolomieu et la géologie de son temps. Paris: Ecole des mines, 2005. 200 pp., no ISBN. Nuncius, 22 (1):159–161, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00239.

Graziosi:2007:BRR

[2783] Elisabetta Graziosi. Book review: Rebecca Messbarger and Paula Findlen (eds.), Maria Gaetana Agnesi et alia. The contest for knowledge. Chicago and London: Chicago University Press, 2005. xxxii + 182 pp., ill., ISBN 0-226-01055-4. Nuncius, 22(1):161–162, ????

2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00248.

Rupke:2007:BRP

[2784] Nicolaas A. Rupke. Book review: Peter Hanns Reill, Vitalizing Nature in the Enlightenment. Berkeley, Los Angeles, London: University of California Press, 2005. x + 388 pp., ISBN 0-520-24135-5. Nuncius, 22 (1):162–163, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00257.

Monti:2007:BRM

[2785] Maria Teresa Monti. Book review: Martin Stuber, Stefan Höchler and Luc Lienhard (eds.), Hallers Netz. Ein Europischer Gelehrtenbriefwechsel zur Zeit der Aufklärung (Studia Halleriana, Band IX). Basel: Verlag, 2005. x + 592 pp., ISBN 3-7965-1327-1. Nuncius, 22(1):163–165, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00266.

Roero:2007:BRE

[2786] Clara Silvia Roero. Book review: Edith Dudley Sylla (ed.), Jacob Bernoulli: the art of conjecturing. Together with Letter to a Friend on Sets in Court Tennis. Baltimore: The Johns Hopkins University Press, 2006. xx + 430 pp., ISBN 0-8018-8235-4. Nuncius, 22(1):165–166, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00275.

Giuli:2007:BRR

[2787] Paola Giuli. Book review: Rita Unfer Lukoschik (ed.), Lettere di Elizabetta Caminer (1751–1769): organizzatrice culturale. Conselve: Edizioni Think Adv, 2006. 367 pp., ISBN 88-8354-003-4. Nuncius, 22(1):166–167, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00284.

Grandin:2007:MCS

[2788] Karl Grandin. Modern and contemporary science. Nuncius, 22(1):168– 169, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00293.

Ciardi:2007:BRA

[2789] Marco Ciardi. Book review: Alessandra Ferraresi, Stato, scienza, amministrazione, saperi. La formazione degli ingegneri in Piemonte dall’antico regime all’Unit d’Italia. Bologna: Il Mulino, 2004, 411 pp., ISBN 88-15-09926-3. Nuncius, 22(1):169–170, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00301.

Mandelbrote:2007:BRA

[2790] Scott Mandelbrote. Book review: Andrew Janiak (ed.), Newton: Philosophical writings. Cambridge: Cambridge university press, 2004. xxxix + 148 pp., ISBN 0-521-83122-9. Nuncius, 22(1):170–171, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00310.

Capocci:2007:BRA

[2791] Mauro Capocci. Book review: Abigail Lustig, Robert J. Richards and Michael Ruse (eds.), Darwinian Heresies. Cambridge: Cambridge University Press, 2004. viii + 200 pp., ISBN 0-521-81516-9. Nuncius, 22 (1):171–172, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00329.

Morini:2007:BRT

[2792] Simona Morini. Book review: Theodore Porter, Karl Pearson: the Scientific Life in a Statistical Age. Princeton: Princeton University Press, 2005. 352 pp., ill., ISBN 0-691-12635-6. Nuncius, 22(1):172–174, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00338.

Abbri:2007:BRM

[2793] Ferdinando Abbri. Book review: Michael Ruse, The Evolution-Creation Struggle. Cambridge, Mass. and London: Harvard University Press, 2005. 327 pp., ISBN 0-674-01687-4. Nuncius, 22(1):174–175, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00347.

Fox:2007:BRJ

[2794] Daniel M. Fox. Book review: Jennifer Tucker, Nature Exposed: Photography as Eyewitness to Victorian Science. Baltimore, Maryland: The Johns Hopkins University Press, 2005. ix + 294 pp., ISBN 0-8018-7991-4. Nuncius, 22(1):175–176, ???? 2007. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539107x00356.

Polizzi:2007:G

[2795] Gaspare Polizzi. General. Nuncius, 22(1):176–177, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539107x00365.

Toscano:2007:BRL

[2796] Maria Toscano. Book review: Luca Ciancio, Teatro del mutamento. Immagini del ’Tempio di Serapide’ a Pozzuoli (1750–1900). Rovereto: Nicolodi, 2005. 184 pp., ill., ISBN 88-8447-218-0. Nuncius, 22(1):177–178, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00374.

Bresadola:2007:BRG

[2797] Marco Bresadola. Book review: Giovanni Cipriani, Il trionfo della ragione. Salute e malattia nella Toscana dell’Età Moderna. Firenze: Nicomp L. E., 2005. 293 pp., ISBN 88-87814-30-9. Nuncius, 22(1):178– 179, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00383.

Calaresu:2007:BRS

[2798] Melissa Calaresu. Book review: Silvia Collini and Antonella Vannoni (eds.), Les instructions scientifiques pour les voyageurs (XVIIe–XIXe siècle). Paris: l’Harmattan, 2005. 348 pp., ISBN 2-7475-8650-2. Nuncius, 22(1):179–181, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00392.

Mauskopf:2007:BRE

[2799] Seymour H. Mauskopf. Book review: Edward Grant, Science and Religion, 400 BC–AD 1550. From Aristotle to Copernicus. Baltimore: The Johns Hopkins University Press, 2006. xviii + 310 pp., ISBN 0-8018-8401-2. Richard Olson, Science and Religion, 1450–1900. From Copernicus to Darwin. Baltimore: The Johns Hopkins University Press, 2006. xviii + 294 pp., ISBN 0-8018-8400-4. Nuncius, 22(1):181–183, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00400.

Galuzzi:2007:BRI

[2800] Massimo Galuzzi. Book review: Ivar Ekeland, The Best of All Possible Worlds. Mathematics and Destiny. Chicago and London: The University of Chicago Press, 2006. 207 pp., ISBN 0-226-19994-0. Nuncius, 22(1): 183–185, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00419.

Mandelbrote:2007:BRR

[2801] Scott Mandelbrote. Book review: Robert J. W. Evans and Alexander Marr (eds), Curiosity and Wonder from the Renaissance to the Enlightenment. Aldershot: Ashgate, 2006. xvi + 265 pp., ISBN 0-7546-4102-3. Nuncius, 22(1):185–186, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00428.

Widmalm:2007:BRM

[2802] Sven Widmalm. Book review: Michael Rand Hoare, The Quest for the True Figure of the Earth: Ideas and Expeditions in Four Centuries of Geodesy. Aldershot: Ashgate, 2005. xii, 275 pp., ISBN 0-7546-5020-0. Nuncius, 22(1):186–187, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00437.

Rampelt:2007:BRD

[2803] Jason M. Rampelt. Book review: David M. Knight and Matthew D. Eddy (eds.), Science and Beliefs: From Natural Philosophy to Natural Science, 1700–1900. Aldershot: Ashgate, 2005. xiv + 286 pp., ill., ISBN 0-7546-3996-7. Nuncius, 22(1):187–188, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00446.

Rossiter:2007:BRJ

[2804] Margaret W. Rossiter. Book review: Jonathan Harwood, Technology’s Dilemma. Agricultural Colleges between Science and Practice in Germany, 1860–1934. Bern: Peter Lang, 2005. 288 pp., ISBN 3-03910-299-0. Nuncius, 22(1):188–189, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00455.

Bleichmar:2007:BRD

[2805] Daniela Bleichmar. Book review: David N. Livingstone and Charles W. J. Withers (eds.), Geography and Revolution. Chicago and London: The University of Chicago Press, 2006. viii + 433 pp., ISBN 0-226-48733-4. Nuncius, 22(1):190–191, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00464.

Segala:2007:BRP

[2806] Marco Segala. Book review: Peter Pesic, Sky in a bottle. Cambridge, Mass: The MIT Press, 2005. 262 pp., ISBN 978-0-262-16234-0. Nuncius, 22(1):191–192, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00473.

Porter:2007:BRM

[2807] Theodore M. Porter. Book review: Margaret Schabas, The Natural Origins of Economics. Chicago and London: University of Chicago Press, 2005. xii + 231 pp., ISBN 0-226-73569-9. Nuncius, 22(1):192–194, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00482.

Diogo:2007:BRM

[2808] Maria Paula Diogo. Book review: Manuel Silva Suárez (ed.), Técnica e ingenieríaen España. Zaragoza: Real academia de ingeniería, 2004. Vol. 1: El Renacimiento. Zaragoza: Real academia de ingeniería, 2004. v + 624 pp., ill., ISBN 84-7820-814-3, vol. II: El siglo de las luces. De la ingeniería a la nueva navegación. 622 pp., ill., ISBN 84-7820-815-1; vol. III: El siglo de las luces. De la industria al ámbito agroforestal. 574 pp., ill., ISBN 84-7820-816-X. Nuncius, 22(1):194–195, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00491.

Nenci:2007:BRM

[2809] Elio Nenci. Book review: Marco Veneziani (ed.), Machina. XI Colloquio Internazionale. Roma, 8–10 gennaio 2004. Atti. Firenze: Olschki, 2005. 543 pp., ISBN 88-222-5419-8. Nuncius, 22(1):195–196, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00509.

Toscano:2007:BRR

[2810] Anna Toscano. Book review: Richard J. Weiss, Breve storia della luce. Arte e scienza dal Rinascimento a oggi. (Storia e Civiltà, 59). Bari: Edizioni Dedalo, 2005. 328 pp., ISBN 88-220-0559-7. Nuncius, 22(1):196– 197, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00518.

Rolfo:2007:Na

[2811] Iolanda Rolfo and Giorgio Strano. Newsletter. Nuncius, 22(1):199–209, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00527.

Munns:2007:CVO

[2812] David P. D. Munns. The challenge of variations: the observational traditions of Ptolemy and Aristotle, and Copernicus’ heliocentric solution. Nuncius, 22(2):223–259, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00536.

Goldstein:2007:WNP

[2813] Bernard R. Goldstein. What’s new in Ptolemy’s almagest? Nuncius, 22 (2):261–285, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00545.

Fabbri:2007:GMH

[2814] Natacha Fabbri. Genesis of Mersenne’s harmonie universelle: the manuscript livre de la nature des sons. Nuncius, 22(2):287–308, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00554.

Tagliavini:2007:OCD

[2815] Maria Grazia Tagliavini. ‘Opere di Cristallo Delicatissime e Meravigliose’: Works of art for connoisseurs and scientific instruments for the Accademia Del Cimento. Nuncius, 22(2):309–333, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00563.

Turner:2007:SMM

[2816] Anthony Turner. Stagecraft and mathematical magic in early modern London. Nuncius, 22(2):335–349, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00572.

Pogliano:2007:ONM

[2817] Claudio Pogliano. Old and new museums in Turin. Nuncius, 22(2): 351–359, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00581.

Acerbi:2007:AMAb

[2818] Fabio Acerbi. Antiquity and middle ages. Nuncius, 22(2):361–362, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00590.

Egmond:2007:BRJ

[2819] Florike Egmond. Book review: Jean A. Givens, Karen M. Reeds, Alain Touwaide (eds.), Visualizing Medieval Medicine and Natural History, 1200–1550. Aldershot: Ashgate, 2006. xx + 278 pp., ISBN 0-7546-5296-3. Nuncius, 22(2):362–363, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00608.

Strano:2007:BRB

[2820] Giorgio Strano. Book review: Barbara Obrist, La cosmologie médiévale: Textes et images. I: Les fondements antiques. Florence: Sismel — Edizioni del Galluzzo, 2004. 384 pp., ISBN 88-8450-140-7. Nuncius, 22(2): 363–364, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00617.

Martin:2007:BRC

[2821] Craig Martin. Book review: Cristina Viano, La matière des choses. Le livre IV des Météorologiques d’Aristote et son interprétation par Olympiodore. Paris: Librairie Philosophique J. Vrin, 2006. 409 pp., ISBN 2-7116-1828-5. Nuncius, 22(2):364–366, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00626.

Beretta:2007:BRL

[2822] Marco Beretta. Book review: Leonid Zhmud, The Origin of the History of Science in Antiquity. Translated from the Russian by Alexander Chernoglazov. Berlin-New York: Walter de Gruyter, 2006. viii + 331 pp., ISBN-10: 3-11-017966-0. Nuncius, 22(2):366–367, ????

2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00635.

Fabbri:2007:REM

[2823] Natacha Fabbri. Renaissance and early modern science. Nuncius, 22 (2):367–368, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00644.

Clark:2007:BRA

[2824] Fiona Clark. Book review: Antonio Barrera-Osorio, Experiencing Nature. The Spanish American Empire and the Early Scientific Revolution. Austin: The University of Texas Press, 2006. xii + 211 pp., ISBN 0-292-70981-1. Nuncius, 22(2):368–369, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00653.

Laird:2007:BRD

[2825] W. R. Laird. Book review: Domenico Bertoloni Meli, Thinking with Objects: the Transformation of Mechanics in the Seventeenth Century. Baltimore: The Johns Hopkins University Press, 2006. 389 pp., ISBN 0-8018-8426-8. Nuncius, 22(2):369–371, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00662.

Candito:2007:BRM

[2826] Cristina Càndito. Book review: Marianne Cojannot-Le Blanc, Marisa Dalai Emiliani, Pascal Glatigny (eds.), l’cartista, l’opera e la sfida della prospettiva. Roma: École Française de Rome, 2006. 485 pp., ISBN 2-7283-0740-7. Nuncius, 22(2):371–372, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00671.

Egmond:2007:BRA

[2827] Florike Egmond. Book review: Alix Cooper, Inventing the Indigenous. Local Knowledge and Natural History in Early Modern Europe. Cambridge and New York: Cambridge University Press, 2007. 218 pp., ISBN 978-0-521-87087-0. Nuncius, 22(2):372–373, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00680.

Grell:2007:BRA

[2828] Ole Peter Grell. Book review: Allen G. Debus, The Chemical Promise. Experiment and Mysticism in the Chemical Philosophy: 1550–1800. Sagamore Beach: Science History Publication/Watson. Nuncius, 22(2):373– 374, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00699.

Galuzzi:2007:BRV

[2829] Massimo Galuzzi. Book review: Vincent Jullien, Philosophie naturelle et géométrie au XVIIe siècle. Paris: Honoré Champion, Paris, 2006. 477 pp., ISBN 2-7453-1363-0. Nuncius, 22(2):374–376, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00707.

Principe:2007:BRL

[2830] Lawrence M. Principe. Book review: Lauren Kassell, Medicine and Magic in Elizabethan London. Simon Forman: Astrologer, Alchemist, and Physician. Oxford: Oxford University Press, 2005. xviii + 281 pp., ISBN 0-19-927905-5 2005. Nuncius, 22(2):376–377, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00716.

Daniel:2007:BRW

[2831] Dane T. Daniel. Book review: Wilhelm Kohlmann, Joachim Telle (eds.), Der Frühparacelsismus. Zweiter Teil. Tübingen: Max Niemeyer Verlag, 2004. xii + 1090 pp., ISBN 3-484-36589-7. Nuncius, 22(2):377–378, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00725.

Hanafi:2007:BRK

[2832] Zakiya Hanafi. Book review: Kathleen P. Long, Hermaphrodites in Renaissance Europe: Women and Gender in the Early Modern World. Aldershot: Ashgate, 2006. x + 268 pp., ISBN 0-7546-5609-8. Nuncius, 22(2):378–379, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00734.

Palmerino:2007:BRA

[2833] Carla Rita Palmerino. Book review: Ad Meskens, Joannes della Faille S.J.: Mathematics, Modesty, Missed Opportunities. Brussels and Rome: Istituto Storico Belga di Roma (Commercial distribution Brepols Publishers), 2005. 177 pp., ISBN 90-74461-53-0. Nuncius, 22(2):379–381, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00743.

Clericuzio:2007:BRW

[2834] Antonio Clericuzio. Book review: William R. Newman, Lawrence M. Principe (eds.), George Starkey: Alchemical laboratory notebooks and correspondence. Chicago: The University of Chicago Press, 2004. xxxvi + 352 pp., ISBN 0-226-57701-5. Nuncius, 22(2):381–382, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00752.

Guerrini:2007:BRK

[2835] Anita Guerrini. Book review: Katharine Park, Secrets of Women. Gender, Generation, and the Origins of Human Dissection. New York: The MIT Press, 2006. 419 pp., ISBN 1-890951-67-6. Nuncius, 22(2):382–383, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00761.

Brambilla:2007:BRA

[2836] Elena Brambilla. Book review: Alessandro Pastore, Le regole dei corpi. Medicina e disciplina nell’Italia moderna. Bologna: Il Mulino, 2006. 246 pp., ISBN 88-15-11277-4. Nuncius, 22(2):383–384, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00770.

Bredekamp:2007:BRV

[2837] Horst Bredekamp. Book review: Volker Remmert, Widmung, Welterklärung und Wissenschaftslegitimierung. Titelbilderund ihre Funktionen in der Wissenschaftlichen Revolution. Wiesbaden: Harrassowitz Verlag, 2005. 267 pp., ISBN 3-447-05337-2. Nuncius, 22(2):384–385, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00789.

DeCallatay:2007:BRG

[2838] Godefroid De Callatay. Book review: George Saliba, Islamic Science and the Making of the European Renaissance. Cambridge, Mass.: The MIT Press, 2007. xi + 315 pp., ISBN-13: 978-0-262-19557-7. Nuncius, 22(2):386–387, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00798.

Fantoli:2007:BRW

[2839] Annibale Fantoli. Book review: William R. Shea, Mariano Artigas, Galileo Observed. Science and the Politics of Belief. Sagamore Beach: Science History Publications/Watson Publishing International, 2006. 212 pp., ISBN 0-88135-356-6. Nuncius, 22(2):387–388, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00806.

Pereira:2007:BRU

[2840] Michela Pereira. Book review: Urszula Szulakowska, The Sacrificial Body and the Day of Doom. Alchemy and Apocalyptic Discourse in the Protestant Reformation. Leiden-Boston: Brill, 2006. 180 pp., ISBN 90-04-15025-0. Nuncius, 22(2):388–390, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00815.

Pennuto:2007:BRL

[2841] Concetta Pennuto. Book review: Lynne Tatlock (ed.), Justine Siegemund: the Court Midwife. Chicago and London: The University of Chicago Press, 2005. xxxii + 260 pp., ISBN 0-226-75709-9. Nuncius, 22(2):390–391, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00824.

Mora:2007:E

[2842] Víctor Guijarro Mora. Enlightenment. Nuncius, 22(2):391–392, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00833.

Albury:2007:BRC

[2843] W. R. Albury. Book review: Claude Blanckaert, Michel Porret (eds.), l’Encyclopédie Methodique (1782–1832): des lumières au positivisme.

Genève: Droz, 2006. 830 pp., ISBN 2-600-00805-5. Nuncius, 22(2):392– 393, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00842.

Pancaldi:2007:BRJ

[2844] Giuliano Pancaldi. Book review: James Delbourgo, A most amazing scene of wonders. Electricity and enlightenment in early America. Cambridge (MA) and London: Harvard University Press, 2006. xii + 367 pp., ISBN 978-0-674-02299-7. Nuncius, 22(2):393–395, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00851.

Sanchez:2007:BRF

[2845] José María García Sánchez. Book review: Felix Driver, Luciana Martins (eds.), Tropical Visions in an Age of Empire. Chicago: The University of Chicago Press, 2005. 279 pp., ISBN 0-226-16472-1. Nuncius, 22(2): 395–396, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00860.

Torre:2007:BRH

[2846] Javier Torre. Book review: Harry Liebersohn, The Traveler’s World. Europe to the Pacific. Cambridge (Mass.) and London: Harvard University Press, 2006. xiii + 380 pp., ISBN 0-674-02185-1. Nuncius, 22 (2):396–397, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00879.

Lavenia:2007:MCS

[2847] Vincenzo Lavenia. Modern and contemporary science. Nuncius, 22(2): 397–399, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00888.

Whitaker:2007:BRA

[2848] Andrew Whitaker. Book review: Andrew Brown, J. D. Bernal. The sage of science. Oxford: Oxford University Press, 2005. xiv + 562 pp., ISBN 0-19-851544-8. Nuncius, 22(2):399–400, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00897.

Candela:2007:BRG

[2849] Andrea Candela. Book review: Guido Chiesura, Charles Darwin geologo. La formazione del giovane Darwin. Docenti e mentori, il viaggio iniziatico tra vulcani e atolli. Le opere geologiche. Benevento: Hevelius Edizioni, 2002. 207 pp., ISBN 88-86977-28-X; Guido Chiesura (ed.), Charles Darwin: opere geologiche. Benevento: Hevelius Edizioni, 2004. 431 pp., ISBN 88-86977-59-X. Nuncius, 22(2):400–401, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00905.

Brockliss:2007:BRW

[2850] Laurence Brockliss. Book review: William Clark, Academic Charisma and the Origins of the Research University. Chicago: The University of Chicago Press, 2006. 576 pp., Isnb 0-226-10921-6. Nuncius, 22(2): 401–403, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x00914.

Bergia:2007:BRJ

[2851] Silvio Bergia. Book review: Jeffrey Crelinsten, Einstein’s Jury. The Race to Test Relativity. Princeton and Oxford: Princeton University Press, 2006. xxix + 397 pp., ISBN 978-0-691-12310-3. Nuncius, 22(2):403–404, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00923.

Bergia:2007:BRA

[2852] Silvio Bergia. Book review: Albert Einstein, The Collected Papers of Albert Einstein, Volume 10: the Berlin Years: Correspondence, May–December 1920, and Supplementary Correspondence, 1909– 1920. Princeton and Oxford: Princeton University Press, 2006. lxix + 686 pp., ISBN 978-0-691-12825-2. Nuncius, 22(2):404–405, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00932.

Canadelli:2007:BRP

[2853] Elena Canadelli. Book review: Philipp Felsch, Laborlandschaften. Physiologische Alpenreisen im 19. Jahrhundert, Göttingen: Wallstein Verlag, 2007. 254 pp., ISBN 3-8353-0159-7. Nuncius, 22(2):405–406, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00941.

Heering:2007:BRE

[2854] Peter Heering. Book review: Elisabeth R. Neswald, Thermodynamik als Kultureller Kampfplatz. Zur Faszinationsgeschichte. Berling: Verlag, 2006. 475 pp., ISBN 3-7930-9448-0. Nuncius, 22(2):406–407, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00950.

Morabito:2007:BRR

[2855] Carmela Morabito. Book review: Roberta Passione (ed.), Ugo Cerletti. Scritti sull’elettroshock. Milano: Angeli, 2006. 236 pp., ISBN 88-464-7823-1. Nuncius, 22(2):407–409, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00969.

Sinatra:2007:BRM

[2856] Maria Sinatra. Book review: Marc J. Ratcliff, Martine Ruchat (eds.), Les Laboratoires de l’esprit. Une histoire de la psychologie à Genève, 1892–1965. Genève: Musée d’Histoire des Sciences de Genève, 2006. 181 pp., ISBN 2-606-01215-1. Nuncius, 22(2):409–410, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x00978.

Wieland:2007:BRI

[2857] Thomas Wieland. Book review: Inken Rebentrost, Das Labor in der Box. Technikentwicklung und Unternehmengründung in der frühen deutschen Biotechnologie. München: Verlag, 2006. 310 pp., ISBN 978-3-406-54403-3. Nuncius, 22(2):410–411, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00987.

Morus:2007:BRL

[2858] Iwan Rhys Morus. Book review: Laura J. Snyder, Reforming Philosophy: A Victorian Debate on Science and Society. Chicago and London: The University of Chicago Press, 2006. x + 386 pp., ISBN 0-226-76733-7. Nuncius, 22(2):411–412, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539107x00996.

Heering:2007:BRF

[2859] Peter Heering. Book review: Friedrich Steinle, Explorative Experimente. Ampère, Faraday und die Ursprünge der Elektrodynamik. Stuttgart: Franz Steiner Verlag, 2005. 450 pp., ISBN 3-515-08185-2. Nuncius, 22 (2):412–413, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x01003.

Chinnici:2007:G

[2860] Ileana Chinnici. General. Nuncius, 22(2):414, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539107x01012.

Sivasundaram:2007:BRG

[2861] Sujit Sivasundaram. Book review: Geoffrey Cantor, Quakers, Jews and science: religious responses to modernity and the sciences in Britain, 1650–1900. New York: Oxford University Press, 2005. xi + 420 pp., ISBN 0-19-927668-4. Nuncius, 22(2):414–416, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x01021.

Henry:2007:BRP

[2862] John Henry. Book review: Peter Dear, The Intelligibility of Nature. How Science Makes Sense of the World. Chicago and London: The University of Chicago Press, 2006. xii + 242 pp., ISBN 0-226-13948-4. Nuncius, 22 (2):416–417, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539107x01030.

Conforti:2007:BRM

[2863] Maria Conforti. Book review: Mariacarla Gadebusch Bondio (ed.), Blood in History and Blood Histories. Firenze: Sismel — Edizioni del Galluzzo, 2005. 385 pp., ISBN 88-8450-162-8. Nuncius, 22(2):417–419, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x01049.

Bernardoni:2007:BRA

[2864] Andrea Bernardoni. Book review: Arthur Greenberg, From Alchemy to Chemistry in Picture and Story. New York: John Wiley & Sons Ltd., 2007, xxiii + 637 pp., ISBN 0-471-75154-5. Nuncius, 22(2):419–420, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x01058.

Brogi:2007:BRT

[2865] Stefano Brogi. Book review: Tullio Gregory, Origini della terminologia filosofica moderna. Linee di ricerca. Firenze: Olschki, 2006. x + 118 pp., ISBN 88-222-5566-6. Nuncius, 22(2):420–422, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x01067.

Pogliano:2007:BRS

[2866] Claudio Pogliano. Book review: Steffan Müller-Wille, Hans-Jörg Rheinberger (eds.), Heredity Produced. at the Crossroads of Biology, Politics, and Culture, 1500–1870. Cambridge (MA): The MIT Press, 2007. x + 496 pp., ISBN 978-0-262-13476-7. Nuncius, 22(2):422–423, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x01076.

Pogliano:2007:BRL

[2867] Claudio Pogliano. Book review: Luc Pauwels (ed.), Visual Cultures of Science. Rethinking Representational Practices in Knowledge Building and Science Communication. Hanover (N.H.): Dartmouth College Press, 2006. xx + 300 pp., ISBN 1-58465-512-7. Nuncius, 22(2):423–424, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x01085.

Rolfo:2007:Nb

[2868] Iolanda Rolfo and Giorgio Strano. Newsletter. Nuncius, 22(2):425–431, ???? 2007. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539107x01094.

Raphael:2008:GDM

[2869] Renee Raphael. Galileo’s Discorsi and Mersenne’s Nouvelles Pensées: Mersenne as a reader of Galilean ’Experience’. Nuncius, 23(1):7–36, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00012.

Kardel:2008:NSN

[2870] Troels Kardel. Nicolaus Steno’s New Myology (1667): Rather than muscle, the motor fibre should be called animal’s organ of movement. Nuncius, 23(1):37–64, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00021.

Diana:2008:CSS

[2871] Esther Diana. Una collezione di strumentaria scientifica all’avvento dell’ospedale moderno: Gli strumenti fisico-matematici di Vincenzo Viviani e l’Ospedale di Santa Maria Nuova Di Firenze (1871–1895). Nuncius, 23(1):65–90, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00030.

Saint-Martin:2008:NAE

[2872] Arnaud Saint-Martin. The new astronomical Eldorado: the French understanding of American astrophysics, 1900–1920. Nuncius, 23(1):91– 113, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00049.

Clericuzio:2008:MFB

[2873] Antonio Clericuzio. The many facets of Boyle’s Natural Philosophy Michael Hunter (with contributions by Edward B. Davis, Harriet Knight, Charles Littleton and Lawrence M. Principe), the Boyle Papers. Understanding the Manuscripts of Robert Boyle, Aldershot: Ashgate, 2007, 674 + xi pp. $139.95/£70.00 ISBN 0-7546-5568-7. Boyle on Atheism. Transcribed and edited by John J. MacIntosh (Toronto: University of Toronto Press, 2005) 493 + xxv pp. £60.00 ISBN 0-8020-9018-4. Nuncius, 23(1):115–126, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00058.

DiPasquale:2008:BRA

[2874] Giovanni Di Pasquale. Book review: Antiquity and Middle Ages, Serafina Cuomo, Technology and culture in Greek and Roman Antiquity. Cambridge: Cambridge University Press, 2007. 212 pp., ISBN 978-0-521-00903-4. Nuncius, 23(1):127–128, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00067.

Pennuto:2008:BRL

[2875] Concetta Pennuto. Book review: Luke Demaitre, Leprosy in Premodern Medicine. A Malady of the Whole Body. Baltimore: The Johns Hopkins University Press, 2007. xvi + 323 pp., ISBN 978-0-8018-8613-3. Nuncius, 23(1):128–130, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00076.

DiPasquale:2008:BRI

[2876] Giovanni Di Pasquale. Book review: Ignacio Gonzlez Tascon and Isabel Velsquez, Ingeniera romana en Hispania. Historia y técnicas constructivas. Madrid: Edicin Fundacin Juanel Turriano, 2005. 542 pp., ISBN 89-95457-52-0. Nuncius, 23(1):130–131, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00085.

Montanari:2008:BRB

[2877] Massimo Montanari. Book review: Bruno Laurioux, Gastronomie, humanisme et société Rome au milieu du XVe siècle. Firenze: Sismel — Edizioni del Galluzzo, 2006. 653 pp., ISBN 88-8450-198-9. Nuncius, 23 (1):131–132, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00094.

Andretta:2008:BRD

[2878] Elisa Andretta. Book review: Daniel Le Bl?vec (ed.), l’Université de médecine de Montpellier et son rayonnement (XIIIe–XVe siècles). Actes du colloque international de Montpellier (Université Paul-Valry-Montpellier III), 17–19 mai 2001. Tournhout: Brepols, 2004. 356 pp., ISBN 2-503-51642-4. Nuncius, 23(1):132–134, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539108x00102.

Acerbi:2008:BRD

[2879] Fabio Acerbi. Book review: Daryn Lehoux, Astronomy, Weather, and Calendars in the Ancient World. Parapegmata and Related Texts in Classical and Near-Eastern Societies. Cambridge: Cambridge University Press, 2007. xiv + 566 pp., ISBN 978-0-521-85181. Nuncius, 23 (1):134–135, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00111.

Minervini:2008:BRR

[2880] Laura Minervini. Book review: Rita Librandi and Rosa Piro (eds.), Lo scaffale della biblioteca scientifica in volgare (secoli XIII-XVI). Atti del convegno (Matera, 14–15 ottobre 2004). Firenze: Sismel — Edizioni del Galluzzo, 2006. 526 pp., ISBN-10: 88-8450-210-1. Nuncius, 23(1): 135–136, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00120.

Acerbi:2008:BRR

[2881] Fabio Acerbi. Book review: Reviel Netz and William Noel, The Archimedes Codex. Revealing the Secrets of the World’s Greatest Palimpsest. London: Weidenfeld e Nicolson, 2007. xii + 305 pp., ISBN 978-0-297-64547-4 [tr. it.: Il codice perduto di Archimede. La storia di un libro ritrovato e dei suoi segreti matematici. Milano: Rizzoli, 2007. 428 pp., ISBN 978-88-17-01601-8]. Nuncius, 23(1):136–138, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00139.

Beretta:2008:BRR

[2882] Ilva Beretta. Book review: Renaissance and Early Modern Science, Alessandro Alessandrini and Alessandro Ceregato (eds.), Natura Picta. Ulisse Aldrovandi, Bologna: Editrici Compositori, 2007. 670 pp., ISBN 978-88-7794-490-0. Adriano Soldano (ed.), l’Erbario di Ulisse Aldrovandi, vols. I–XV, Venezia: Istituto Veneto di Scienze, Lettere e Arti. 2000–2005. 1369 pp., no ISBN. Nuncius, 23(1):138–140, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00148.

Clericuzio:2008:BRC

[2883] Antonio Clericuzio. Book review: Clara Bartocci (ed.), John Josselyn. Rarit della Nuova Inghilterra. New England’s Rarities. Perugia: Morlacchi Editore, 2006. xxxiv + 243 pp., ISBN 88-6074-044-4. Nuncius, 23 (1):140–141, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00157.

Margocsy:2008:BRF

[2884] Daniel Margocsy. Book review: Florike Egmond, Paul Hoftijzer and Robert Visser (eds.), Carolus Clusius. Towards a Cultural History of a Renaissance Naturalist. Amsterdam: Koninklijke Nederlandse Akademie van Wetenschappen, 2007. vi + 349 pp., ISBN 978-90-6984-506-7. Nuncius, 23(1):141–142, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00166.

Favino:2008:BRM

[2885] Federica Favino. Book review: Maurice A. Finocchiaro, Retrying Galileo, 1633–1992. Berkeley/Los Angeles/London: University of California Press, 2005. 485 pp., ISBN 978-0-520-24261-6. Nuncius, 23(1): 142–144, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00175.

Olmi:2008:BRF

[2886] Giuseppe Olmi. Book review: Fabio Garbari Lucia Tongiorgi Tomasi (eds.), Flora: the Erbario Miniato and other Drawings. London: Brepols, 2007. 664 pp. (2 vols.). Vol. 1: ISBN 978-1-905375-20-2, Vol. 2: ISBN 978-1-905375-21-9. Nuncius, 23(1):144–145, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00184.

Abbri:2008:BRA

[2887] Abbri. Book review: Anne Goldgar, Tulipmania. Money, Honor, and Knowledge in the Dutch Golden Age. Chicago and London: The University of Chicago Press, 2007. xx + 425 pp., ISBN 978-0-226-30125-9. Nuncius, 23(1):145–146, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00193.

Droscher:2008:BRD

[2888] Ariane Droscher. Book review: Dario Generali Marc Ratcliff (eds.), From Makers to Users. Microscopes, Markets,and Scientific Practices in the Seventeenth and Eighteenth Centuries / Dagli artigiani ai naturalisti. Microscopi, offerta dei mercati e pratiche scientifiche nei secoli XVII e XVIII. Firenze: Olschki, 2007. xv + 336 pp., ISBN 978-88-222-5669-0. Nuncius, 23(1):146–148, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00201.

Angelini:2008:BRJ

[2889] Annarita Angelini. Book review: James Hankins (ed.), The Cambridge Companion to Renaissance Philosophy. Cambridge: Cambridge University Press, 2007. xvi + 430 pp., ISBN 978-0-521-60893-0. Nuncius, 23 (1):148–149, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00210.

Smith:2008:BRV

[2890] A. Mark Smith. Book review: Vincent Ilardi, Renaissance Vision from Spectacles to Telescopes. Philadelphia: the American Philosophical Society, 2007. xv + 378 pp., ISBN 978-0-87169-259-7. Nuncius, 23(1):149– 150, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00229.

Bernardoni:2008:BRL

[2891] Andrea Bernardoni. Book review: Lawrence M. Principe (ed.), Chymists and Chymistry, Studies in the History of Alchemy and Early Modern Chemistry. Sagamore Beach: Chemical Heritage Foundation and Science History Publications, 2007. xiii + 274 pp., ISBN 978-0-88135-396-9. Nuncius, 23(1):150–152, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00238.

Guerrini:2008:BRS

[2892] Luigi Guerrini. Book review: Sylvie Taussig (ed.), Pierre Gassendi. Vie et mœurs d’épicure. Paris: Les Belles Lettres, 2006. 2 vols., I vol.: cxix + 279 pp., II vol.: 396 pp., ISBN 2-251-79983-4. Nuncius, 23(1):152–153, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00247.

Fregonese:2008:BRE

[2893] Lucio Fregonese. Book review: Enlightenment: Paola Bertucci, Viaggio nel paese delle meraviglie. Scienza e curiosit nell’Italia del Settecento. Torino: Bollati Boringhieri, 2007. 294 pp., ISBN 978-88-339-1802-0. Nuncius, 23(1):153–154, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00256.

Beretta:2008:BRAa

[2894] Marco Beretta. Book review: Annie Chassagne, La Bibliothèque de l’Académie Royale des Sciences au XVIIIe siècle. Paris: Editions du Comit des Travaux Historiques et Scientifiques, 2007. 305 pp., ISBN 978-2-7355-0637-8. Nuncius, 23(1):154–156, ????

2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00265.

DiMitri:2008:BRJ

[2895] Gino L. Di Mitri. Book review: José Correia Da Serra, Itinerrios Histrico-Naturis, Maria Paula Diogo, Ana Carneiro e Ana Simões (eds.), Porto: Porto Editora, 2003. viii + 104 pp., ISBN 972-0-34262-5. Ana Simões, Maria Paula Diogo e Ana Carneiro, Cidado do Mundo. Una Biografia Cientifica do Abade Correia da Serra. Porto: Porto Editora, 2006. v + 185 pp., ISBN 978-972-0-45090-6. Nuncius, 23(1):156–157, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00274.

Barsanti:2008:BRM

[2896] Giulio Barsanti. Book review: Michl Duchet (ed.), Buffon: De l’homme. Presentation par Michele Duchet. Paris: l’Harmattan, 2006. 467 pp., ISBN 2-296-01033-4. Nuncius, 23(1):157–158, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00283.

Bandinelli:2008:BRJ

[2897] Angela Bandinelli. Book review: Joseph S. Fruton, Fermentation. vital or Chemical Process? Leiden and Boston: Brill, 2006. xv + 141 pp., ISBN 978-90-04-15268-7. Nuncius, 23(1):158–159, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00292.

Beretta:2008:BRE

[2898] Marco Beretta. Book review: Etienne Guyon (ed.), l’École normale de l’an III. Leçons de physique, de chimie, d’histoire naturelle. Édition annotée des cours de Hay, Berthollet et Daubenton, avec introductions, notes et annexes. Paris: Editions rue d’Ulm, 2006. 645 pp., ISBN 2-7288-0356-0. Nuncius, 23(1):159–160, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00300.

Anderson:2008:BRR

[2899] Robert G. W. Anderson. Book review: Richard L. Hills, James Watt Volume 1: His Time in Scotland, 1736–1774. Ashbourne: Landmark Publishing, 2002. 480 pp., ISBN 1-84306-045-0. Richard L. Hills, James Watt Volume 2: the Years of Toil, 1775–1785. Ashbourne: Landmark Publishing, 2005. 256 pp., ISBN 1-84306-046-9. Richard L. Hills, James Watt Volume 3: Triumph Through Adversity, 1785–1819. Ashbourne: Landmark Publishing, 2006. 288 pp., ISBN-13 978-1-84306-193-9. Nuncius, 23(1):160–162, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00319.

Knight:2008:BRU

[2900] David Knight. Book review: Ursula Klein and Wolfgang Lefèvre, Materials in Eighteenth-Century Science. A Historical Ontology. Cambridge Ma: The MIT Press, 2007. x + 345 pp., ISBN 978-0-262-11306-9. Nuncius, 23(1):162–163, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00328.

Thebaud-Sorger:2008:BRM

[2901] Marie Thébaud-Sorger. Book review: Michael R. Lynn, Popular science and public opinion in Eighteenth-Century France. Manchester: Manchester University Press, 2006. ix + 177 pp., ISBN 0-7190-7373. Nuncius, 23(1):163–164, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00337.

Steinke:2008:BRM

[2902] Hubert Steinke. Book review: Maria Teresa Monti (ed.), Écriture et mémoire. Les carnets medico-biologiques de Vallisneri A. Wolff. Milano: Franco Angeli, 2006. 226 pp., ISBN 88-464-7937-8. Nuncius, 23(1):164– 165, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00346.

Stefani:2008:BRM

[2903] Marta Stefani. Book review: Maria Teresa Monti, Spallanzani e le rigenerazioni animali. l’inchiesta, la comunicazione, la rete. Firenze: Olschki, 2005. 424 pp., ISBN 88-222-5493-7. Nuncius, 23(1):165–166, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00355.

Giudice:2008:BRM

[2904] Franco Giudice. Book review: Marco Piccolino e Marco Bresadola, Rane, torpedini e scintille. Galvani, Volta e l’elettricit animale. Torino: Bollati Boringhieri, 2003. 723 pp., ISBN 88-339-1470-4. Nuncius, 23(1):167–168, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00364.

Galuzzi:2008:BRC

[2905] Massimo Galuzzi. Book review: Clara Silvia Roero (ed.), Alle origini dell’Accademia. I contributi di Leonhard Euler. Firenze: Olschki, 2007. x + 178 pp., ISBN 978-88-222-5658-4. Nuncius, 23(1):168–169, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00373.

Bresadola:2008:BRW

[2906] Marco Bresadola. Book review: Wayne Wild, Medicine-by-Post. The Changing Voice of Illness in Eighteenth-Century British Consultation Letters and Literature. Amsterdam, New York: Rodopi, 2006. 286 pp., ISBN 90-420-1868-2. Nuncius, 23(1):169–170, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00382.

Eddy:2008:BRM

[2907] Matthew D. Eddy. Book review: Modern and Contemporary Science: José Ramón Bertomeu-Sánchez Agustí Nieto-Galan (eds.), Chemistry, Medicine and Crime: Mateu J. B. Orfila (1787–1853) and His Times. Sagamore Beach: Science History Publications, 2006. xxvi + 306 pp., ISBN 0-88135-275-6. Nuncius, 23(1):170–171, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00391.

Canadelli:2008:BRO

[2908] Elena Canadelli. Book review: Olaf Breidbach, Visions of Nature. The Art and Science of Ernst Haeckel. München: Prestel Verlag, 2006. 304 pp., ISBN 3-7913-3664-9. Nuncius, 23(1):171–172, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00409.

Capocci:2008:BRW

[2909] Mauro Capocci. Book review: W. F. Bynum, A. Hardy, S. Jacyna and E. Tansey, The Western Medical Tradition. 1800–2000. Cambridge University Press, 2006. 628 pp., ISBN-13: 978-0-521-47565-5, ISBN-10: 0-521-47565-1. Nuncius, 23(1):172–173, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00418.

Turchetti:2008:BRV

[2910] Simone Turchetti. Book review: Valeria Del Gamba, Il Ragazzo di Via Panisperna. l’avventurosa vita del fisico Franco Rasetti. Torino: Bollati Boringhieri, 2007. 167 pp., ISBN 978-88-339-1746-7. Nuncius, 23(1): 174–175, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00427.

Sloan:2008:BRM

[2911] Phillip R. Sloan. Book review: Michael Friedman, Alfred Nordman (eds.), The Kantian Legacy in Nineteenth-Century Science. Cambridge, Ma: The MIT Press, 2006. vi + 370 pp., ISBN 978-0-262-06254-1. Nuncius, 23(1):175–176, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00436.

Krige:2008:BRK

[2912] John Krige. Book review: Klaus Hentschel, The Mental Aftermath. The Mentality of German Physicists 1945–1949. Oxford: Oxford University Press, 2007. 205 pp., ISBN 978-0-19-920566-0. Nuncius, 23(1):176–177, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00445.

Wittje:2008:BRM

[2913] Roland Wittje. Book review: Myles W. Jackson, Harmonious Triads Physicists, Musicians, and Instrument Makers in Nineteenth-Century Germany. Cambridge (MA): The MIT Press, 2006. 395 pp., ISBN 0-262-10116-5. Nuncius, 23(1):177–178, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00454.

Sodi:2008:BRA

[2914] Francesca Sodi. Book review: Alberto Meschiari, Edizione Nazionale delle Opere e della Corrispondenza di Giovanni Battista Amici. Napoli: Bibliopolis, 2006. 2 vols., 1085 pp., ISBN 88-7088-519-4. Nuncius, 23 (1):178–179, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00463.

Heering:2008:BRC

[2915] Peter Heering. Book review: Christine Nawa Christoph Meinel, Instrumente und Apparaturen in Heidelberger Laboratorien skizziert von Friedrich Veith (1817–1907). Regensburg: Universitt Regensburg, 2007. 127 pp., no ISBN. Nuncius, 23(1):179–180, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00472.

Abbri:2008:BRR

[2916] Ferdinando Abbri. Book review: Richard Newbury, John Maynard Keynes. vita pubblica e privata di un grande economista ed esteta trasgressivo. Milano: Boroli, 2007. 175 pp., ISBN 978-88-7493-113-2. Nuncius, 23(1):180–182, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00481.

Cerruti:2008:BRC

[2917] Luigi Cerruti. Book review: Carsten Reinhardt, Shifting and Rearranging: Physical Methods and the Transformation of Modern Chemistry. Sagamore Beach: Science History Publications, 2006. x + 428 pp., ISBN 0-88135-354-X. Nuncius, 23(1):182–183, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00490.

Maltese:2008:BRS

[2918] Giulio Maltese. Book review: Simone Turchetti, Il caso Pontecorvo. Fisica nucleare, politica e servizi di sicurezza nella guerra fredda. Sironi: Milano, 2007. 285 pp., ISBN 978-88-518-0081-9. Nuncius, 23(1):183–184, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00508.

Segala:2008:BRA

[2919] Marco Segala. Book review: Andrea Villa (ed.), Charles Babbage. Passaggi dalla vita di uno scienziato. Autobiografia dell’inventore del computer. Torino: Utet, 2007. xiii + 414 pp., ISBN 978-88-02-07715-4. Nuncius, 23(1):184–185, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00517.

Clark:2008:BRG

[2920] Fiona Clark. Book review: General Jorge Caizares-Esguerra, Nature, Empire, and Nation. Explorations of the History of Science in the Iberian World. Stanford: Stanford University Press, 2006. xiv + 230 pp., ISBN 978-0-8047-5544-3. Nuncius, 23(1):186–187, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00526.

Sodero:2008:BRF

[2921] Fernando Giuseppe Sodero. Book review: Francesco Paolo de Ceglia (ed.), Scienziati di Puglia. Secoli V a.C.–XXI d.C. Bari: Adda, 2007. 682 pp., ISBN 978-88-8082-681-1. Nuncius, 23(1):187–188, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00535.

Galuzzi:2008:BRA

[2922] Massimo Galuzzi. Book review: Alfred S. Posamentier and Ingmar Lehmann, The (Fabulous) Fibonacci Numbers. New York: Prometheus Books, 2007. 385 pp., ISBN 976-1-59102-475-0. Nuncius, 23(1):188–189, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00544.

Rolfo:2008:N

[2923] Iolanda Rolfo. Newsletter. Nuncius, 23(1):191–198, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539108x00553.

Nieto-Galan:2008:HSS

[2924] Agusti Nieto-Galan. The history of science in Spain a critical overview. Nuncius, 23(2):211–236, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00562.

Simoes:2008:PCH

[2925] Ana Simões, Ana Carneiro, and Maria Paula Diogo. Perspectives on contemporary history of science in Portugal. Nuncius, 23(2):237–263, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00571.

Patiniotis:2008:OHM

[2926] Manolis Patiniotis. Origins of the historiography of modern Greek science. Nuncius, 23(2):265–289, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00580.

Kilinc:2008:AMA

[2927] Berna Kilinc. Ahmed Midhat and Adnan Adivar on history of science and civilizations. Nuncius, 23(2):291–308, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539108x00599.

Homburg:2008:BAN

[2928] Ernst Homburg. Boundaries and audiences of national histories of science: Insights from the history of science and technology of the Netherlands. Nuncius, 23(2):309–345, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00607.

Redondi:2008:RAM

[2929] Pietro Redondi. La rivincita di Aldo Mieli. Nuncius, 23(2):347–351, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00616.

Turner:2008:SB

[2930] Anthony Turner. Silvio A. Bedini. Nuncius, 23(2):353–371, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00625.

Vidoni:2008:MFM

[2931] Ferdinando Vidoni. In memoriam felice mondella. Nuncius, 23(2):373– 379, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00634.

Fabbri:2008:AMA

[2932] Natacha Fabbri. Antiquity and middle ages. Nuncius, 23(2):381–382, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00643.

Beretta:2008:BRAb

[2933] Marco Beretta. Book review: Aldo Brancacci, Pierre-Marie Morel (eds.), Democritus: Science, the Arts, and the Care of the Soul. Proceedings of the International Colloquium on Democritus (Paris, 18–20 September 2003). Leiden/Boston: Brill, 2007. 348 pp., ISBN 90-04-15160-5. M. Laura Gemelli Marciano, Democrito e l’Accademia. Studi sulla trasmissione dell’atomismo antico da Aristotele a Simplicio. Berlin/New York: Walter de Gruyter, 2007. xii + 376 pp., ISBN 978-3-11-018542-3. Nuncius, 23(2):382–384, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00652.

Fabbri:2008:BRM

[2934] Natacha Fabbri. Book review: Marta Cristiani, Cecilia Panti, Graziano Perillo (eds.), Harmonia Mundi. Musica mondana e musica celeste fra Antichità e Medioevo. Firenze: Sismel — Edizioni del Galluzzo, 2007. 228 pp., ISBN 978-88-8450-239-1. Nuncius, 23(2):384–385, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00661.

Sumpter:2008:BRM

[2935] Denise Sumpter. Book review: Michael A. B. Deakin, Hypatia of Alexandria. Mathematician and Martyr. Amherst: Prometheus Books, 2007. 231 pp., ISBN 978-1-59102-520-7. Nuncius, 23(2):385–386, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00670.

Beretta:2008:BRS

[2936] Marco Beretta. Book review: Stuart Gillespie, Philip Hardie (eds.), The Cambridge Companion to Lucretius. Cambridge: Cambridge University Press, 2007. 364 pp., ISBN 978-0-521-61266-1. Nuncius, 23(2):386–388, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00689.

Strano:2008:BRD

[2937] Giorgio Strano. Book review: David A. King, In Synchrony with the Heavens: Studies in Astronomical Timekeeping and Instrumentation in Medieval Islamic Civilization. Leiden/Boston: Brill, 2004– 2005 (2 vols.). Vol. 1: lvii + 930 pp., ISBN 90-04-12233-8; vol. 2: lxxvi + 1066 pp., ISBN 90-04-14188-X. Nuncius, 23(2):388–390, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00698.

Lehoux:2008:BRP

[2938] Daryn Lehoux. Book review: Philip Sabin, Hans Van Mees, Michael Whitby (eds.), The Cambridge History of Greek and Roman Warfare. Cambridge, Cambridge University Press: 2007. 2 vols, xiv + 608 pp., ISBN-13: 978-0-521-85779-6. Nuncius, 23(2):390–391, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00706.

Repici:2008:BRD

[2939] Luciana Repici. Book review: David Sider, Carl Wolfram Brunschön (eds.), Theophrastus of Eresus. On Weather Signs. Leiden/ Boston: Brill Academic Publishers, 2007. x + 263 pp., ISBN-10: 90-04-15593-7, ISBN-13: 978-90-04-15593-0. Nuncius, 23(2):391–392, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00715.

Maerker:2008:BRK

[2940] Anna Maerker. Book review: Ken Arnold, Cabinets for the Curious. Looking Back at Early English Museums, Aldershot: Ashgate, 2006. xii + 297 pp., ISBN 0-7546-0506-X. Nuncius, 23(2):392–393, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00724.

Donato:2008:BRA

[2941] Maria Pia Donato. Book review: Andrea Battistini, Gilberto De Angelis, Giuseppe Olmi (eds.), All’origine della scienza moderna: Federico Cesi e l’Accademia dei Lincei. Bologna: Il Mulino, 2007. 479 pp., ISBN 978-88-15-11986-5. Nuncius, 23(2):393–395, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00733.

Avalos:2008:BRM

[2942] Ana Avalos. Book review: Marco Bertozzi (ed.), Nello Specchio del cielo. Giovanni Pico della Mirandola e le‘Disputationes’ contro l’astrologia divinatoria. Firenze, Leo S. Olschki, 2008. vi + 165 pp., ISBN 978-88-222-5722-2. Nuncius, 23(2):395–396, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00742.

Dupre:2008:BRH

[2943] Sven Dupré. Book review: Horst Bredekamp, Galilei der Künstler: Der Mond, die Sonne, die Hand. Berlin: Akademie Verlag, 2007. 517 pp., ISBN 978-3-05-004319-7. Nuncius, 23(2):396–397, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00751.

Margocsy:2008:BRH

[2944] Daniel Margocsy. Book review: Harold J. Cook, Matters of Exchange: Commerce, Medicine and Science in the Dutch Golden Age. New Haven and London: Yale University Press, 2007. xiv + 562 pp., ISBN 978-0-300-11796-7. Nuncius, 23(2):397–398, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00760.

Latimer:2008:BRJ

[2945] Bonnie Latimer. Book review: Juliet Cummins, David Burchell (eds.), Science, Literature, and Rhetoric in Early Modern England. Aldershot: Ashgate, 2007. xiii + 241 pp., ISBN 978-0-7546-5781-1. Nuncius, 23 (2):398–399, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00779.

Finocchiaro:2008:BRE

[2946] Maurice A. Finocchiaro. Book review: Egidio Festa, Galileo: La lotta per la scienza. Bari: Laterza, 2007. xi + 344 pp., ISBN 978-88-420-8377-1. Nuncius, 23(2):399–400, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00788.

Goddu:2008:BRS

[2947] André Goddu. Book review: Stephen Gaukroger, The Emergence of a Scientific Culture. Science and the Shaping of Modernity, 1210–1685. New York: Oxford University Press, 2006. ix + 563 pp., ISBN 0-19-929644-8. Nuncius, 23(2):401–402, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00797.

Egmond:2008:BRD

[2948] Florike Egmond. Book review: Deborah E. Harkness, The Jewel House. Elizabethan London and the Scientific Revolution. New Haven & London: Yale University Press, 2007. 349 pp., ISBN 978-0-300-11196-5. Nuncius, 23(2):402–403, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00805.

Crippa:2008:BRF

[2949] Sabina Crippa. Book review: Fabrizio Meroi, Elisabetta Scapparone (eds.), La magia nell’Europa moderna. Tra antica sapienza e filosofia naturale. Atti del Convegno (Firenze, 2–4 ottobre 2003). Firenze: Olschki, 2007 (2 vols). xii + 786 pp., ISBN 978-88-222-5749-9. Nuncius, 23(2): 403–406, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00814.

Haskell:2008:BRN

[2950] Yasmin Haskell. Book review: Nancy G. Siraisi, History, Medicine, and the Traditions of Renaissance Learning. Ann Arbor: University of Michigan Press, 2007. 438 pp., ISBN-13: 978-0-472-11602-7; ISBN-10: 0-472-11602-9. Nuncius, 23(2):406–407, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00823.

Ricci:2008:BRD

[2951] Saverio Ricci. Book review: Dario Tessicini, I dintorni dell’infinito. Giordano Bruno e l’astronomia del Cinquecento, Pisa/Roma: Fabrizio Serra, 2007. 212 pp., ISBN 978-88-6227-097-7. Nuncius, 23(2):407–409, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00832.

Pennuto:2008:BRJ

[2952] Concetta Pennuto. Book review: Jacqueline Vons, Stéphane Velut (eds.), André Vésale: Résumé de ses livres sur la fabrique du corps humain. Paris: Les Belles Lettres, 2008. cxxi + 282 pp., ISBN 978-2-251-34510-9. Nuncius, 23(2):409–410, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00841.

Abbri:2008:BRE

[2953] Ferdinando Abbri. Book review: Edward G. Andrew, Patrons of Enlightenment. Toronto/Buffalo/London: University of Toronto Press, 2006. 284 pp., ISBN 0-8020-9064-8. Nuncius, 23(2):410–411, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00850.

Toscano:2008:BRB

[2954] Anna Toscano. Book review: Biancastella Antonino, Giuseppe Olmi, Maria Gioia Tavoni (eds.), Il libro illustrato a Bologna nel Settecento. Bologna: Dipartimento di Italianistica, 2007. 263 pp., ISBN assente. Nuncius, 23(2):411–413, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00869.

Fazzari:2008:BRM

[2955] Michela Fazzari. Book review: Marco Bresadola (ed.), Antonio Vallisneri: Quaderni di osservazioni, Volume II. Firenze: Leo S. Olschki Editore, 2007. xlvi + 153 pp., ISBN 978-88-222-5686-7. Nuncius, 23 (2):413–414, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00878.

Toscano:2008:BRN

[2956] Maria Toscano. Book review: Neil Chambers, Joseph Banks and the British Museum: the World of Collecting, 1770–1830. London: Pickering and Chatto, 2007. xiv + 195 pp., ISBN-10: 1-85196-858-X, ISBN-13: 978-1-85196-858-9. Nuncius, 23(2):414–415, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00887.

Ciancio:2008:BRD

[2957] Luca Ciancio. Book review: Dennis R. Dean, Romantic Landscapes. Geology and Its Cultural Influence in Britain, 1765–1835. Ann Arbor: Scholars’ Facsimiles & Reprint, 2007. 604 pp., ISBN 0-8201-1555-X. Nuncius, 23(2):415–416, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x00896.

Cecere:2008:BRS

[2958] Imma Cecere. Book review: Simone Dumont, Un Astronome des Lumières: Jèrôme Lalande. Paris: Vuibert, 2007. viii + 359 pp., ISBN 978-2-7117-4028-4. Nuncius, 23(2):416–417, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539108x00904.

Bresadola:2008:BRS

[2959] Marco Bresadola. Book review: Stanley Finger, Doctor Franklin’s medicine. Philadelphia: University of Pennsylvania Press, 2006. xiii + 379 pp., ISBN-13: 978-0-8122-3913-3. Nuncius, 23(2):417–418, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00913.

Zanier:2008:BRD

[2960] Sonia Zanier. Book review: Dario Generali, Antonio Vallisneri. Gli anni della formazione e le prime ricerche. Firenze: Olschki, 2007. xiv + 466 pp., ISBN 978-88-222-5621-8. Nuncius, 23(2):418–420, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00922.

Mason:2008:BRA

[2961] Peter Mason. Book review: Arthur Macgregor, Curiosity and Enlightenment.Collectors and Collections from the Sixteenth to the Nineteenth Century. London and New Haven: Yale University Press, 2007. 386 pp., ISBN 978030012493. Nuncius, 23(2):420–421, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00931.

Conforti:2008:BRO

[2962] Maria Conforti. Book review: Ole P. Grell, Andrew Cunningham, Bernd Roeck (eds.), Health Care and Poor Relief in 18th- and 19th-Century Southern Europe. Aldershot: Ashgate, 2005. vi + 326 pp., ISBN 0-7546-5156-8. Nuncius, 23(2):421–423, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00940.

Cavazza:2008:BRA

[2963] Marta Cavazza. Book review: Adriano Prosperi, Dare l’anima. Storia di un infanticidio. Torino: Einaudi, 2005. 373 pp., ISBN 88-06-16754-5. Nuncius, 23(2):423–424, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x00959.

Bleichmar:2008:BRS

[2964] Daniela Bleichmar. Book review: Susan Scott Parrish, American Curiosity: Cultures of Natural History in the Colonial British Atlantic World. Williamsburg, Virginia: The University of North Carolina Press for the Omohundro Institute of Early American History and Culture, 2006. xvi + 321 pp., ISBN 0-8078-5678-9. Nuncius, 23(2):424–426, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00968.

Leoni:2008:BRL

[2965] Simona Boscani Leoni. Book review: Luigi Zanzi, Enrico Rizzi (eds.), Déodat de Dolomieu, Viaggi nelle Alpi. Ornavasso: Fondazione Enrico Monti, Fondazione Maria Giussani Bernasconi, 2006. 271 pp., ISBN 88-85295-51-7. Nuncius, 23(2):426–427, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539108x00977.

LaVergata:2008:BRP

[2966] Antonello La Vergata. Book review: Peter J. Bowler, Monkey Trials and Gorilla Sermons. Evolution and Christianity from Darwin to Intelligent Design. Cambridge, MA and London: Harvard University Press, 2007. 256 pp., ISBN-13: 978-0-674-02615-5, ISBN-10: 0-674-02615-2. Nuncius, 23(2):427–429, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x00986.

Segala:2008:BRR

[2967] Marco Segala. Book review: Robert M. Brain, Robert S. Cohen, Ole Knudsen (eds.), Hans Christian Ørsted and the Romantic Legacy in Science. Ideas, Disciplines, Practices. Dordrecht: Springer, 2007. xx + 442 pp., ISBN 978-1-4020-2979-0. Nuncius, 23(2):429–430, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00995.

Zublena:2008:BRM

[2968] Paolo Zublena. Book review: Massimo Bucciantini, Italo Calvino e la scienza. Gli alfabeti del mondo. Roma: Donzelli, 2007. v + 184 pp., ISBN 978-88-6036-163-9. Nuncius, 23(2):430–431, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01002.

Carletti:2008:BRM

[2969] Christian Carletti. Book review: Maria Pia Casalena, Per lo Stato, per la Nazione. I congressi degli scienziati in Francia e in Italia (1830–1914). Roma: Carocci, 2007. 253 pp., ISBN 978-88-430-4270-8. Nuncius, 23 (2):431–432, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x01011.

Kessler:2008:BRJ

[2970] Elizabeth A. Kessler. Book review: James Elkins, Six Stories from the End of Representation. Images in Painting, Photography, Astronomy, Microscopy, Particle Physics and Quantum Mechanics, 1980– 2000. Stanford: Stanford University Press, 2008, xx + 274 pp. ISBN 978-0-8047-4147-7, ISBN 978-0-8047-4148-4. Nuncius, 23(2):432–433, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01020.

Yalda:2008:BRJ

[2971] Sepi Yalda. Book review: James Rodger Fleming, The Callendar Effect: the Life and Work of Guy Stewart Callendar (1898–1964), the Scientist Who Established the Carbon Dioxide Theory of Climate Change. Boston: American Meteorological Society, 2007. xv + 155 pp., ISBN13: 978-1-878220-76-9, ISBN-10: 1-878220-76-4. Nuncius, 23(2): 434, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01039.

Cahan:2008:BRK

[2972] David Cahan. Book review: Karl Grandin, Nina Wormbs, Sven Widmalm (eds.), The Science-Industry Nexus. History, Policy, Implications. Sagamore Beach: Science History Publications, 2004. xvii + 457 pp., ISBN 0-88135-365-5. Nuncius, 23(2):434–436, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01048.

Govoni:2008:BRN

[2973] Paola Govoni. Book review: Nicole Hulin, Les femmes, l’enseignement et Les sciences. Un long cheminement (XIXe–XXe siècle). Paris: l’Harmattan, 2008. 238 pp., ISBN 978-2-296-05606-0. Nuncius, 23(2): 436–437, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x01057.

Carletti:2008:BRC

[2974] Christian Carletti. Book review: C. Luperini, T. Paladini, Ho costruito il seguente apparecchietto. Antonio Pacinotti, manoscritti e strumenti. Pisa: Ets, 2007. m + 127 pp., ISBN 978-88-467-1895-2. Nuncius, 23 (2):437–438, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x01066.

Pantalony:2008:BRA

[2975] David Pantalony. Book review: Alison D. Morrison-Low, Making Scientific Instruments in the Industrial Revolution. Aldershot: Ashgate, 2007. xvi + 408 pp., ISBN 978-0-7546-5758-3. Nuncius, 23(2):438–439, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01075.

Oldroyd:2008:BRR

[2976] David Oldroyd. Book review: Ralph O’Connor, The Earth on Show: Fossils and the Poetics of Science, 1802–1856. Chicago and London: The University of Chicago Press, 2007. xiv + 541 pp., ISBN-13: 978-0-226-61668-1; 10: 0-226-61668-1. Nuncius, 23(2):439–441, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01084.

Kojevnikov:2008:BRD

[2977] Alexei Kojevnikov. Book review: David E. Rowe, Robert Schulmann (eds.), Einstein on Politics. His Private Thoughts and Public Stands on Nationalism, Zionism, War, Peace, and the Bomb. Princeton: Princeton University Press, 2007. xxix + 523 pp., ISBN-13: 978-0-691-12094-2. Nuncius, 23(2):441–442, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x01093.

Corbellini:2008:BRU

[2978] Gilberto Corbellini. Book review: Ulf Schmidt, Andreas Frewer (eds.), History and Theory of Human Experimentation. The Declaration of Helsinki and Modern Medical Ethics. Stuttgart: Franz Steiner Verlag, 2007. 370 pp., ISBN 978-3-515-08862-6. Nuncius, 23(2):442–443, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01101.

Canadelli:2008:BRJ

[2979] Elena Canadelli. Book review: Julia Voss, Darwins Bilder. Ansichten der Evolutionstheorie 1837–1874. Frankfurt am Main: Fischer Taschenbuch Verlag, 2007. 380 pp., ISBN 978-3-596-17627-4. Nuncius, 23(2):443–444, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01110.

Candito:2008:G

[2980] Cristina Càndito. General. Nuncius, 23(2):444–445, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539108x01129.

Bernardoni:2008:BRB

[2981] Andrea Bernardoni. Book review: Brenda J. Buchanan (ed.), Gunpowder, Explosives and the State: A Technological History. Burlington, VT: Ashgate, 2006. xxiii + 425 pp., ISBN 0-7546-5259-9. Nuncius, 23(2):446– 447, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01138.

Shea:2008:BRC

[2982] William Shea. Book review: Charles Coulston Gillispie, Essays and Reviews in History and History of Science. Philadelphia: American Philosophical Society, 2007. xix + 424 pp., ISBN 978-0-87169-965-7. Nuncius, 23(2):447–448, ???? 2008. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539108x01147.

Pogliano:2008:BRL

[2983] Claudio Pogliano. Book review: Lorraine Daston, Peter Galison, Objectivity. New York: Zone Books, 2007. 502 pp., ISBN 978-1-890951-78-8. Nuncius, 23(2):448–450, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x01156.

Raposo:2008:BRR

[2984] Pedro Raposo. Book review: Roger Hutchins, British University Observatories: 1772–1939. Aldershot: Ashgate, 2008. xxiii + 533 pp., ISBN 978-0-7546-3250-4. Nuncius, 23(2):450–451, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01165.

Pogliano:2008:BRM

[2985] Claudio Pogliano. Book review: Martin Kemp, The Human Animal in Western Art and Science. Chicago and London: The University of Chicago Press, 2007. xx + 308 pp., ISBN 978-0-226-43033-1. Nuncius, 23(2):451–452, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x01174.

Abbri:2008:BRJ

[2986] Ferdinando Abbri. Book review: Jacomien Prins, Mariken Teeuwen (eds.), Harmonisch labyrint. De muziek van de kosmos in de westerse wereld. Hilversum: Hilversum Verloren, 2007. 181 pp., ISBN 978-90-8550-974-1. Nuncius, 23(2):452–453, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539108x01183.

Maronne:2008:BRP

[2987] Sébastien Maronne. Book review: Peter S. Rudman, How Mathematics Happened. The First 50,000 Years. New York: Prometheus Books, 2007. 314 pp., ISBN 976-1-59102-477-4. Nuncius, 23(2):453–454, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01192.

Fara:2008:BRA

[2988] Patricia Fara. Book review: Alessandro Tosi, Portraits of Men and Ideas: Images of Science from the Renaissance to the Nineteenth Century. Pisa: Edizioni Plus, 2007. 183 pp., ISBN 978-88-8492-511-4. Nuncius, 23(2): 454–455, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539108x01200.

Bonaiuti:2008:N

[2989] Elisa Bonaiuti and Iolanda Rolfo. Newsletter. Nuncius, 23(2):457–465, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x01219.

Beretta:2009:JUT

[2990] Marco Beretta. Journals under threat: A joint response from history of science, technology and medicine Editors. Nuncius, 24(1):7–10, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00011.

Geymonat:2009:AGA

[2991] Mario Geymonat. Arithmetic and geometry in Ancient Rome: Surveyors, intellectuals, and poets. Nuncius, 24(1):11–34, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00020.

Omodeo:2009:GBN

[2992] Pietro Daniel Omodeo. Giordano Bruno and Nicolaus Copernicus. Nuncius, 24(1):35–59, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00039.

Mayer:2009:SIP

[2993] Thomas F. Mayer. The status of the Inquisition’s precept to Galileo (1616) in historical perspective. Nuncius, 24(1):61–95, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00048.

Smith:2009:PHT

[2994] S. Halikowski Smith. ‘The Physician’s Hand’: Trends in the evolution of the apothecary and his art across Europe (1500–1700). Nuncius, 24 (1):97–125, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00057.

Fiocca:2009:CDS

[2995] Alessandra Fiocca and Iolanda Nagliati. Le cours d’histoire des sciences de Guglielmo Libri au Collège de France (1833). Nuncius, 24(1):127– 171, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00066.

Repellini:2009:FLN

[2996] Ferruccio Franco Repellini. Formalismo in linguaggio naturale. Nuncius, 24(1):173–185, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00075.

Dekker:2009:AMA

[2997] Elly Dekker. Antiquity and middle ages. Nuncius, 24(1):187–188, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00084.

Romani:2009:BRC

[2998] Francesca Romana Romani. Book review: Charles Burnett (ed.), Ibn Baklarish’s Book of Simples. Medical Remedies between Three Faiths in Twelfth-Century Spain. Oxford: the Arcadian Library in association with Oxford University Press, 2008 (Studies in the Arcadian Library, 3) 216 pp., ISBN 978-0-19-954306-9. Nuncius, 24(1):188–189, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00093.

Giorgetti:2009:BRB

[2999] Dario Giorgetti. Book review: Bruce S. Eastwood, Ordering the Heavens. Roman Astronomy and Cosmology in the Carolingian Renaissance. Boston, Leiden: Brill, 2007 (History of Science and Medicine Library, 4). xiii + 452 pp., ISBN 978-90-04-16186-3. Nuncius, 24(1):190, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00101.

Acerbi:2009:BRJ

[3000] Fabio Acerbi. Book review: Jöran Friberg, Amazing Traces of a Babylonian Origin in Greek Mathematics. Hackensack, NJ, U.S.A.: World Scientific Press, 2007. xx + 476 pp., ISBN 978-981-270-452-8. Nuncius, 24(1):190–191, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00110.

Caroti:2009:BRT

[3001] Stefano Caroti. Book review: Tullio Gregory, Speculum naturale. Percorsi del pensiero medievale. Roma: Edizioni di Storia e Letteratura, 2007 (Storia e Letteratura. Raccolta di studi e testi, 235). x + 254 pp., ISBN 978-88-84 98338-1. Nuncius, 24(1):191–194, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00129.

Barreca:2009:BRJ

[3002] Francesco Barreca. Book review: Jens Høyrup, Jacopo da Firenze’s Tractatus Algorismi and Early Italian Abbacus Culture. Basel — Boston — Berlin: Birkhäuser Verlag AG, 2007. xii + 484 pp., ISBN 978-3-7643-8390-9. Nuncius, 24(1):194–195, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00138.

DiPasquale:2009:BRM

[3003] Giovanni Di Pasquale. Book review: Mary Jaeger, Archimedes and the Roman Imagination. Ann Arbor: The University of Michigan Press, 2008. ix + 230 pp., ISBN 978-0-472-11630-0. Nuncius, 24(1):195–196, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00147.

Tessicini:2009:BRM

[3004] Dario Tessicini. Book review: Michel-Pierre Lerner, Le Monde des Sphères. Tome I: Genèse et triomphe d’une représentation cosmique, xi + 401 pp., ISBN 978-2-251-42034-9. Tome II: La fin du cosmos classique, x + 529 pp., ISBN 978-2-251-42035-6. Paris: Les Belles Lettres, 2008. Nuncius, 24(1):196–198, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00156.

Belle:2009:BRJ

[3005] Riccardo Bellè. Book review: José Luis Mancha, Studies in Medieval Astronomy and Optics. Aldershot: Ashgate, 2006. x + 182 pp., ISBN 978-0-86078-996-3. Nuncius, 24(1):198–200, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00165.

Conforti:2009:BRM

[3006] Maria Conforti. Book review: Michael Mcvaugh, The Rational Surgery of the Middle Ages. Firenze: Sismel, Edizioni del Galluzzo, 2007. 296 pp., ISBN 88-8450-199-7. Nuncius, 24(1):200–201, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00174.

Bernardoni:2009:REM

[3007] Andrea Bernardoni. Renaissance and early modern science. Nuncius, 24 (1):201–202, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00183.

Palumbo:2009:BRM

[3008] Margherita Palumbo. Book review: Maria Rosa Antognazza, Leibniz. An Intellectual Biography. Cambridge/New York: Cambridge University Press, 2009. xxvii + 623 pp., ISBN 978-0-521-80619-0. Nuncius, 24(1): 202–203, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00192.

Brioist:2009:BRA

[3009] Pascal Brioist. Book review: Andrea Bernardoni, Leonardo e il monumento equestre a Francesco Sforza. Firenze: Giunti, 2007. 126 pp., ISBN 978-88-09-05396-0. Nuncius, 24(1):203–205, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00200.

Pennuto:2009:BRJ

[3010] Concetta Pennuto. Book review: Jol Coste, Représentations et comportements en temps d’épidémie dans la littérature imprimè de peste (1490– 1725). Contribution l’histoire culturelle de la peste en France l’époque moderne. Préface du professeur Yves-Marie Berc. Paris: Honor Champion Éditeur, 2007. 848 pp., ISBN 978-2-7453-1532-8. Nuncius, 24(1): 205–207, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00219.

Shea:2009:BRM

[3011] William R. Shea. Book review: Maurice A. Finocchiaro (ed. and transl.), The Essential Galileo. Indianapolis: Hackett Publishing Company, 2009. 380 pp., ISBN 978-0-87220-937. Nuncius, 24(1):207–208, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00228.

Diana:2009:BRJ

[3012] Esther Diana. Book review: John Henderson, The Renaissance Hospital. Healing the Body and healing the Soul. New Haven and London: Yale University Press, 2006. xxxiv + 458 pp., ISBN 978-0-300-10995-5. Nuncius, 24(1):208–209, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00237.

Coradeschi:2009:BRN

[3013] Gabriele Coradeschi. Book review: Nicholas Jardine et Alain-Philippe Segons (éds.), La Guerre des Astronomes. La querelleau sujet de l’origine du système géohéliocentrique à la fin du XVIe siècle. Voll. I–II. Paris: Les Belles Lettres, 2008. 14 xxiv + 286 pp., xix + 852 pp., ISBN 978-2-251-34513-0, 978-2-251-34512-3. Nuncius, 24(1):209–211, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00246.

Laurenza:2009:BRM

[3014] Domenico Laurenza. Book review: Martin Kemp, Leonardo da Vinci. The Marvellous Works of Nature and Man. Oxford: Oxford University Press, 2006. xxix + 381 pp., ISBN 978-0-19-280725-0. Nuncius, 24(1): 211–212, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00255.

Camerota:2009:BRD

[3015] Filippo Camerota. Book review: Daniela Lamberini, Il Sanmarino. Giovanni Battista Belluzzi architetto militare e trattatista del Cinquecento, 2 voll. (I. La vita e le opere, II. Gli scritti). Firenze: Leo S. Olschki Editore, 2007. xvi + 840 pp., ISBN 978-88-222-5660-7. Nuncius, 24(1):212–214, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00264.

Abbri:2009:BRS

[3016] Ferdinando Abbri. Book review: Steven Matthews, Theology and Science in the Thought of Francis Bacon. Aldershot: Ashgate, 2008. xiv + 150 pp., ISBN 978-0-7546-6252-5. Nuncius, 24(1):214–215, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00273.

Karrow:2009:BRM

[3017] Robert W. Karrow. Book review: Matthew Mclean, The Cosmographia of Sebastian Münster: Describing the World in the Reformation. St. Andrews Studies in Reformation History. Aldershot: Ashgate, 2007. 378 pp., ISBN 978-0-7546-5843-6. Nuncius, 24(1):215–216, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00282.

Long:2009:BRL

[3018] Pamela O. Long. Book review: Lissa Roberts, Simon Schaffer and Peter Dear (eds.), The mindful hand: Inquiry and invention from the late Renaissance to early industrialization. Amsterdam: Koninklijke Nederlandse Akademie van Wetenschappen, 2007. xxvii + 503 pp., ISBN 978-90-6984-483-1. Nuncius, 24(1):216–218, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00291.

Ruiz-Castell:2009:BRJ

[3019] Pedro Ruiz-Castell. Book review: Joan Vernet i Ramon Parés, La Ciéncia en la História dels Països Catalans. II Del naixement de la ciència moderna a La Il· lustració. Valencia: Institut d’Estudis Catalans Universitat de València, 2007. 774 pp., ISBN 978-84-7283-905-2. Nuncius, 24(1):218– 220, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00309.

Imbruglia:2009:E

[3020] Girolamo Imbruglia. Enlightenment. Nuncius, 24(1):220–221, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00318.

Candela:2009:BRD

[3021] Andrea Candela. Book review: Donata Brianta, Europa mineraria. Circolazione delle élites e trasferimento tecnologico (secoli XVIII–XIX). Milano: Franco Angeli, 2007. 447 pp., ISBN 978-88-464-8818-3. Nuncius, 24(1):221–222, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00327.

Clerici:2009:BRM

[3022] Luca Clerici. Book review: Marco Ciardi (ed.), Esplorazioni e viaggi scientifici nel Settecento. Milano: Rizzoli, 2008. 458 pp., ISBN 978-88-17-02334-4. Nuncius, 24(1):223–224, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00336.

Brockliss:2009:BRM

[3023] Laurence Brockliss. Book review: Maurice Crosland, Scientific Institutions and Practice in France and Britain c. 1700–c. 1870. Aldershot: Ashgate/Variorum, 2007. xvi + 270 pp., ISBN 978-0-7546-5913-6. Nuncius, 24(1):224–225, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00345.

Imbruglia:2009:BRAa

[3024] Girolamo Imbruglia. Book review: Adrien Paschoud, Le monde amérindien au miroir des Lettres édifiantes et curieuses, SVEC (Studies on Voltaire and XVIII century). Oxford: Voltaire Foundation, 2008. vi + 229 pp., ISBN 978-0-7294-0935-3. Nuncius, 24(1):225–226, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00354.

Bowden:2009:BRI

[3025] Mary Ellen Bowden. Book review: Isabel Rivers and David L. Wykes (eds.), Joseph Priestley, Scientist, Philosopher, and Theologian. Oxford: Oxford University Press, 2008. xii + 252 pp., ISBN 978-0-19-921530-0. Nuncius, 24(1):226–227, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00363.

deCeglia:2009:BRSa

[3026] Francesco Paolo de Ceglia. Book review: Stéphane Schmitt, Les forces vitales et leur distribution dans la nature. Un essai de ≪ systématique physiologique≫. Turnhout: Brepols 2006. 240 pp., ISBN 2-503-52442-7. Nuncius, 24(1):228–229, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00372.

Conforti:2009:BRA

[3027] Maria Conforti. Book review: Antonio Vallisneri, Consulti medici. Vol. I, a cura di Benedino Gemelli. Firenze: Olschki, 2006. clxxxi + 433 pp., ISBN 88-222-5554-2. Nuncius, 24(1):229–230, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00381.

Ottaviani:2009:MCS

[3028] Alessandro Ottaviani. Modern and contemporary science. Nuncius, 24 (1):230–231, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00390.

Ferraresi:2009:BRM

[3029] Alessandra Ferraresi. Book review: Marco Ciardi (ed.), Il fisico sublime. Amedeo Avogadro e la cultura scientifica del primo Ottocento. Bologna: Il Mulino, 2007. 271 pp., ISBN 978-88-15-11889-9. Nuncius, 24(1):231– 232, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00408.

Carneiro:2009:BRP

[3030] Ana Carneiro. Book review: Pietro Corsi, Fossils and Reputations. A Scientific Correspondence: Pisa, Paris, London, 1853–1857. Pisa: Edizioni Plus–Pisa University Press, 2008, ISBN 978-88-8492-564-0. Nuncius, 24(1):232–234, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00417.

Luciano:2009:BRA

[3031] Erika Luciano. Book review: Angelo Guerraggio e Pietro Nastasi, Roma 1908: il Congresso Internazionale dei Matematici. Torino: Bollati Boringhieri, 2008. 215 pp., ISBN 978-88-339-1866-2. Nuncius, 24(1):234–235, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00426.

Segala:2009:BRJ

[3032] Marco Segala. Book review: Jim Endersby, Imperial Nature. Joseph Hooker and the Practices of Victorian Science. Chicago and London:

The University of Chicago Press, 2008. xii + 429 pp., ISBN 978-0-226-20791-9. Nuncius, 24(1):235–237, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00435.

Hochadel:2009:BRA

[3033] Oliver Hochadel. Book review: Aileen Fyfe and Bernard Lightman (eds.), Science in the Marketplace. Nineteenth-Century Sites and Experiences. Chicago and London: University of Chicago Press, 2007. 410 pp., ISBN 978-0-226-27650-2. Bernard Lightman, Victorian Popularizers of Science. Designing Nature for New Audiences. Chicago and London: The University of Chicago Press, 2007. 545 pp., ISBN 978-0-226-48118-0. Nuncius, 24(1):237–239, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00444.

Polizzi:2009:BRN

[3034] Gaspare Polizzi. Book review: Noel Jackson, Science and Sensation in Romantic Poetry. Cambridge: Cambridge University Press, 2008. xiv + 288 pp., ISBN 978-0-521-86937-9. Nuncius, 24(1):239–241, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00453.

Capocci:2009:BRH

[3035] Mauro Capocci. Book review: Hannah Landecker, Culturing Life. How Cells Became Technologies. Cambridge (MA): Harvard University Press, 2007. x + 276 pp., ISBN 0-674-02328-5. Nuncius, 24(1):241–242, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00462.

Mauskopf:2009:BRU

[3036] Seymour Mauskopf. Book review: Ulf Larsson, Alfred Nobel. Networks of Innovation. Sagamore Beach: Science History Publications, 2008. 216 pp., ISBN 978-0-88135-399-0. Nuncius, 24(1):242–243, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00471.

Travis:2009:BRS

[3037] Anthony S. Travis. Book review: Stephan H. Lindner, Inside IG Farben: Hoechst During the Third Reich, transl. by Helen Schoop. Cambridge:

Cambridge University Press, 2008. xx + 388 pp., ISBN 978-0-521-88766-

3. Nuncius, 24(1):243–244, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00480.

Nasim:2009:BRG

[3038] Omar W. Nasim. Book review: Gavin Parkinson, Surrealism, Art and Modern Science: Relativity, Quantum Mechanics, Epistemology. New Haven London: Yale University Press, 2008. viii + 294 pp., ISBN 978-0-300-09887-7. Nuncius, 24(1):244–246, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00499.

Pogliano:2009:BRR

[3039] Claudio Pogliano. Book review: Roberta Passione, Ugo Cerletti. Il romanzo dell’elettroshock. Reggio Emilia: Aliberti editore, 2007. 206 pp., ISBN 978-88-7424-216-0. Nuncius, 24(1):246–247, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00507.

Beretta:2009:BRPa

[3040] Marco Beretta. Book review: Presenze femminili nella cultura del Novecento. Adriana Tramontano, Maria Luisa Righini Bonelli, Luisa Becherucci. Firenze: Centro Di, 2008. 71 pp., ISBN 978-88-7038-467-

3. Nuncius, 24(1):247–248, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00516.

Abbri:2009:BRR

[3041] Ferdinando Abbri. Book review: Roberto Scazzieri and Raffaella Simili (eds.), The Migration of Ideas. Sagamore Beach: Science History Publications/USA, 2008. x + 245 pp., ISBN 978-0-88135-397-6. Nuncius, 24 (1):248–249, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00525.

Focaccia:2009:G

[3042] Miriam Focaccia. General. Nuncius, 24(1):250–251, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539109x00534.

DalPrete:2009:BRP

[3043] Ivano Dal Prete. Book review: Pascal Richet, A Natural History of Time. Chicago and London: The University of Chicago Press, 2007. xiv + 471 pp., ISBN 978-0-226-71287-1. Nuncius, 24(1):251–252, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00543.

Rousseau:2009:BRS

[3044] George Rousseau. Book review: Sharon Ruston (ed.), Literature and Science: Essays and Studies 2008 for the English Association. Cambridge: D. S. Brewer, 2008. xi + 176 pp., ISBN 978-1-84384-178-4. Nuncius, 24 (1):252–253, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00552.

Rolfo:2009:N

[3045] Iolanda Rolfo. Newsletter. Nuncius, 24(1):255–263, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539109x00561.

Hagner:2009:I

[3046] Michael Hagner, Renato G. Mazzolini, and Claudio Pogliano. Introduction. Nuncius, 24(2):279–289, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00570.

Bernardoni:2009:BTP

[3047] Andrea Bernardoni. The biography of a technical picture: Biringuccio’s boring machine woodcut. Nuncius, 24(2):291–311, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539109x00589.

Kramer:2009:PIU

[3048] Fabian Krämer. The persistent image of an unusual centaur. A biography of Aldrovandi’s two-legged centaur woodcut’. Nuncius, 24(2):313–340, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00598.

Gattei:2009:EFK

[3049] Stefano Gattei. The engraved frontispiece of Kepler’s Tabulae Rudolphinae (1627): A preliminary study. Nuncius, 24(2):341–365, ????

2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00606.

Brusius:2009:IDS

[3050] Mirjam Brusius. Inscriptions in a double sense: the biography of an early scientific photograph of script. Nuncius, 24(2):367–392, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00615.

Nasim:2009:SIS

[3051] Omar W. Nasim. On seeing an image of a spiral nebula: From Whewell to Flammarion. Nuncius, 24(2):393–414, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539109x00624.

VonBrevern:2009:PLP

[3052] Jan Von Brevern. A picture’s life. The photograph of an erratic boulder. Nuncius, 24(2):415–438, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00633.

Canadelli:2009:SCD

[3053] Elena Canadelli. ‘Some curious drawings’. Mars through Giovanni Schiaparelli’s eyes: Between science and fiction. Nuncius, 24(2):439–464, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00642.

Solhdju:2009:MNI

[3054] Katrin Solhdju. Mapping nerves. an investigation into the vitality of an image. Nuncius, 24(2):465–488, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00651.

Beltrame:2009:WOB

[3055] Lorenzo Beltrame and Silvia Giovanetti. When the oocyte becomes an embryo. the social life of an ambiguous scientific image in Italian newspapers (1996–2007). Nuncius, 24(2):489–507, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00660.

Sannino:2009:AMA

[3056] Antonella Sannino. Antiquity and middle ages. Nuncius, 24(2):509–510, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00679.

Fausti:2009:BRS

[3057] Daniela Fausti. Book review: Sabrina Grimaudo, Difendere la salute. Igiene e disciplina del soggetto nel‘De sanitate tuenda’ di Galeno. Napoli: Bibliopolis, 2008. 280 pp., ISBN 978-88-7088-543-9. Nuncius, 24(2):510– 511, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00688.

Sannino:2009:BRD

[3058] Antonella Sannino. Book review: David Juste, Les Alcheandrana primitifs. Études sur les plus anciens traités astrologiques latins d’origine arabe (Xe siècle) (Brill’s Studies in Intellectual History, 152). Leiden: Brill Academic Publishing, 2007. xvi + 736 pp., ISBN 978-90-04-15827-6. Nuncius, 24(2):511–512, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00697.

Beretta:2009:BRPb

[3059] Marco Beretta. Book review: Paul T. Keyser and Georgia L. Irby-Massie (eds.), The Encyclopedia of Ancient Natural Scientists. The Greek tradition and its many heirs. London/New York: Routledge, 2008. ix + 1062 pp., ISBN 978-0-415-34020-5. Nuncius, 24(2):512–513, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00705.

DiPasquale:2009:BRJ

[3060] Giovanni Di Pasquale. Book review: Jo Marchant, Decoding the heavens. A 2.000 years old computer and the century long search to discover its secrets. Cambridge, Mass.: Da Capo Press Edition, 2009. 328 pp., ISBN 978-0-306-81742-7. Nuncius, 24(2):513–515, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00714.

DiPasquale:2009:BRC

[3061] Giovanni Di Pasquale. Book review: Cesare Rossi, Flavio Russo, and Ferruccio Russo (eds.), Ancient engineers’ inventions: precursors of the present. New York: Springer, c2009. xv + 339 pp., ISBN 978-90-481-2252-3. Nuncius, 24(2):515–516, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00723.

Rodolfi:2009:BRG

[3062] Anna Rodolfi. Book review: Giorgio Stabile, Dante e la filosofia della natura. Percezioni, linguaggi, cosmologie (Micrologus’ Library 20). Firenze: Sismel-Edizioni del Galluzzo, 2007. 416 pp., ISBN 978-88-8450-247-6. Nuncius, 24(2):516–517, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00732.

Beretta:2009:BRJ

[3063] Marco Beretta. Book review: James Warren (ed.), The Cambridge Companion to Epicureanism. Cambridge: Cambridge University Press, 2009. x + 342 pp., ISBN 978-0-521-69530-5. Nuncius, 24(2):517–520, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00741.

Barreca:2009:BRE

[3064] Francesco Barreca. Book review: Erhard Weigel, Analysis Aristotelica ex Euclide restituta (Clavis Pansophiae 3,3). Stuttgart–Bad Cannstatt: Frommann-Holzboog Verlag, 2008, 358, lxvii + 355 pp., ISBN 978-3-7728-2375-6. Nuncius, 24(2):520, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00750.

Angelini:2009:BRJ

[3065] Annarita Angelini. Book review: Jacopo Aconcio, Il pensiero scientifico e l’idea di tolleranza, a cura di P. Giacomoni e L. Dappiano. Trento: Editrice universitaria, 2005. 305 pp., ISBN 88-8443-115-8. Nuncius, 24 (2):521–522, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00769.

Conforti:2009:BRD

[3066] Maria Conforti. Book review: David Gentilcore, Medical Charlatanism in Early Modern Italy. Oxford: Oxford University Press, 2006. xii + 426 pp., ISBN 978-0-19-924535-2. Nuncius, 24(2):522–523, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00778.

Guicciardini:2009:BRAa

[3067] Niccolò Guicciardini. Book review: Andrew Janiak, Newton as Philosopher. Cambridge: Cambridge University Press, 2008. 296 pp., ISBN 978-0-521-86286-8. Nuncius, 24(2):523–524, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539109x00787.

Bernardoni:2009:BRT

[3068] Andrea Bernardoni. Book review: Tara Nummedal, Alchemy and Authority in the Holy Roman Empire. Chicago and London: The University of Chicago Press, 2007. 260 pp., ISBN 978-0-226-60856. Nuncius, 24 (2):524–525, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00796.

Long:2009:BRA

[3069] Pamela O. Long. Book review: Antonella Romano (éd.), Rome et la science moderne entre Renaissance et Lumières, (Collection de l’école Française de Rome, 403). Rome: Ecole française de Rome, 2008. 751 pp., ISBN 978-2-7283-0833-0. Nuncius, 24(2):525–527, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00804.

Favino:2009:BRP

[3070] Federica Favino. Book review: Pamela H. Smith and Benjamin Schmidt (eds.), Making knowledge in Early Modern Europe. Practices, Objects and Texts, 1400–1800. Chicago and London: Chicago University Press, 2007. 360 pp., ISBN 978-0-226-76329-3. Nuncius, 24(2):527–528, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00813.

Beretta:2009:BRPc

[3071] Marco Beretta. Book review: Philippe Véron, Verres d’optique et lunettes astronomiques. Firenze: Fondazione Ronchi, 2009. 541 pp., ISBN-13 978-88-88649-22-1. Nuncius, 24(2):528–529, ????

2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00822.

Bianchi:2009:BRC

[3072] Massimo L. Bianchi. Book review: Charles Webster, Paracelsus. Medicine, Magic and Mission at the End of Time. New Haven and London: Yale University Press, 2008. xiv + 326 pp., ISBN 978-0-300-13911-2. Nuncius, 24(2):529–531, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00831.

Visconti:2009:BRU

[3073] Agnese Visconti. Book review: Ulrich PäßLer, Ein ‘Diplomat aus den Wäldern des Orinoco’. Alexander von Humboldt als Mittler zwischen Preußen und Frankreich. Stuttgart: Franz Steiner Verlag, 2009. 244 pp., ISBN 978-3-515-09344-6. Nuncius, 24(2):531–532, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00840.

Imbruglia:2009:BRAb

[3074] Girolamo Imbruglia. Book review: Ann Thomson, Bodies of thought. Science, Religion and the Soul in the early Enlightenment. Oxford: Oxford University Press, 2008. viii + 293 pp., ISBN 978-0-19-923619-0. Nuncius, 24(2):532–533, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00859.

Simon:2009:BRN

[3075] Jonathan Simon. Book review: Nathalie Vuillemin, Les beautés de la nature à l’épreuve de l’analyse. Paris: Presses Sorbonne Nouvelle, 2008. 412 pp., ISBN 978-2-87854-466-4. Nuncius, 24(2):533–534, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00868.

Civello:2009:BRF

[3076] Andrea Civello. Book review: Frederick Burkhardt, Samantha Evans, and Alison Pearn (eds.), Evolution. Selected Letters of Charles Darwin. Cambridge: Cambridge University Press, 2008, xxii + 308 pp., ISBN 978-0-521-87412-0. Nuncius, 24(2):534–535, ????

2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00877.

Carli:2009:BRE

[3077] Alberto Carli. Book review: Elena Canadelli (a cura di), Milano scientifica 1875–1924. La rete del grande Politecnico. Nuncius, 24(2):535–538, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00886.

OConnor:2009:BRJ

[3078] Ralph O’Connor. Book review: Jane P. Davidson, A History of Paleontology Illustration. Bloomington, Indiana: Indiana University Press, 2008. xvi + 217 pp., ISBN 978-0-253-35175-3. Nuncius, 24(2):538–539, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00895.

Guicciardini:2009:BRAb

[3079] Niccolò Guicciardini. Book review: Alex D. D. Craik, Mr Hopkins’ Men: Cambridge Reform and British Mathematics in the 19th Century. London, Springer: 2007. xiv + 410 p., ISBN 978-1-84628-790-9. Nuncius, 24 (2):539–540, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00903.

Bonolis:2009:BRF

[3080] Luisa Bonolis. Book review: Francesco Guerra and Nadia Robotti, Ettore Majorana. Aspects of his Scientific and Academic Activity. Pisa: Edizioni Scuola Normale Superiore, 2008. 243 pp., ISBN 978-88-7642-331-4. Nuncius, 24(2):540–541, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00912.

Berlivet:2009:BRC

[3081] Luc Berlivet. Book review: Caroline Hannaway (ed.), Biomedicine in the Twentieth Century: Practices, Policies, and Politics. Amsterdam– Berlin–Oxford–Tokyo–Washington D.C.: IOS Press, 2008. x + 377 pp., ISBN 978-1-58603-832-8. Nuncius, 24(2):541–542, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539109x00921.

Canadelli:2009:BRP

[3082] Elena Canadelli. Book review: Pamela Kort and Max Hollein (eds.), Darwin. Art and the Search for Origins. Cologne: Wienand Verlag, 2009. 288 pp., ISBN 978-3-87909-973-3. Nuncius, 24(2):543–544, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00930.

Perez:2009:BRC

[3083] Liliane Pérez. Book review: Christine Macleod, Heroes of invention. Technology, liberalism and British identity 1750–1914. Cambridge: Cambridge University Press, 2007. xv + 458 pp., ISBN 978-0-521-87370-3. Nuncius, 24(2):544–545, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x00949.

Canadelli:2009:BRR

[3084] Elena Canadelli. Book review: Richard Milner, Darwin’s Universe. Evolution from A to Z. Berkeley: University of California Press, 2009. 488 pp., ISBN 978-0-520-24376-7. Nuncius, 24(2):545–546, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00958.

Pogliano:2009:BRS

[3085] Claudio Pogliano. Book review: Silvano Montaldo e Paolo Tappero (a cura di), Cesare Lombroso cento anni dopo. Torino: Utet, 2009. xvi + 410 pp., ISBN 978-88-02-08139-7. Nuncius, 24(2):546–547, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x00967.

Shea:2009:BRH

[3086] William R. Shea. Book review: Harry Nussbaumer and Lydia Bieri, Discovering the Expanding Universe. Cambridge: Cambridge University Press, 2009. xvii + 226 pp., ISBN 978-0-521-51484-2. Nuncius, 24(2): 547–548, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00976.

DiBartolo:2009:BRO

[3087] Alexander Di Bartolo. Book review: Olivier Orain, De plein-pied dans le monde. écriture et réalisme dans la géographie française au XXème siècle. Paris: l’Harmattan, 2009. 427 pp., ISBN 978-2-296-07957-1. Nuncius, 24(2):548–549, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x00985.

Muscatello:2009:BRG

[3088] Umberto Muscatello. Book review: Gaetana Silvia Rigo (ed.), Emilio Veratti e Camillo Golgi. Corrispondenza inedita e altri documenti. Firenze: Olschki, 2008. vii + 113 pp., ISBN 978-88-222-5859-5. Nuncius, 24(2): 549–551, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x00994.

deCeglia:2009:BRSb

[3089] Francesco Paolo de Ceglia. Book review: La scienza nel Mezzogiorno dopo l’Unità d’Italia. Soveria Mannelli (Catanzaro): Rubbettino Editore, 2008. 3 vols., 1570 pp., ISBN 978-88-498-1889-5. Nuncius, 24(2): 551–552, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x01001.

DeSio:2009:BRA

[3090] Fabio De Sio. Book review: Alessandro Volpone, Gli inizi della genetica in Italia. Bari: Cacucci, 2008. 334 pp., ISBN 978-88-8422-706-5. Nuncius, 24(2):552–553, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x01010.

Kim:2009:BRJ

[3091] Mi Gyung Kim. Book review: José Ramón Bertomeu-Sanchez, Duncan Thorburn Burns, and Brigitte Van Tiggelen (eds.), Neighbours and Territories. The Evolving Identity of Chemistry. Louvain-la-Neuve, Belgique: Memosciences asbl, 2008. 752 pp., ISBN 978-2-9600815-0-3. Nuncius, 24(2):553–554, ???? 2009. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539109x01029.

Taub:2009:BRG

[3092] Liba Taub. Book review: Görel Cavalli–Björkman and Svante Lindqvist (eds.), Research and Museums (RaM): Proceedings of an International Symposium in Stockholm, 22–25 May 2007 (Archives of the Nobel Museum 11, Nationalmusei skriftserie N. S. 20). Stockholm: National Museum Stockholm, the Nobel Museum, 2008. 236 pp., ISBN 978-91-7100-803-9. Nuncius, 24(2):554–555, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539109x01038.

Morabito:2009:BRP

[3093] Carmela Morabito. Book review: Paul Churchland, Neurophilosophy at Work. Cambridge: Cambridge University Press, 2007. xi + 249 pp., ISBN 978-0-521-69200-7. Nuncius, 24(2):556–557, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x01047.

Becchi:2009:BRL

[3094] Antonio Becchi. Book review: Luisa Dolza, Storia della tecnologia. Bologna: Il Mulino, 2008. 241 pp., ISBN 978-88-15-12574-3. Nuncius, 24(2):557–558, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x01056.

Stefani:2009:BRA

[3095] Marta Stefani. Book review: Ariane Droscher, Biologia. Storia e concetti. Roma: Carocci, 2008. 339 pp., ISBN 978-88-430-4535-8. Nuncius, 24(2): 558–559, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539109x01065.

Roberts:2009:BRM

[3096] Lissa Roberts. Book review: Massimo Mazzotti (ed.), Knowledge as Social Order. Rethinking the Sociology of Barry Barnes. Aldershot: Agate, 2008. xi + 184 pp., ISBN 978-0-7546-4863-5. Nuncius, 24(2):559–561, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x01074.

Anonymous:2009:N

[3097] Anonymous. Newsletter. Nuncius, 24(2):563–569, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539109x01083.

Anonymous:2009:NC

[3098] Anonymous. Notes on contributors. Nuncius, 24(2):571–575, ???? 2009. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539109x01092.

Anonymous:2010:USG

[3099] Anonymous. Under the sign of Galileo: the new Museo di Storia della Scienza. Nuncius, 25(1):vii–xii, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539110x00019.

Eddy:2010:ANC

[3100] Matthew D. Eddy. The alphabets of nature: Children, books and natural history in Scotland, circa 1750–1800. Nuncius, 25(1):xiii–22, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00028.

Segala:2010:LWL

[3101] Marco Segala. Lagrange: ‘Why I left Turin and never went back again’. Nuncius, 25(1):23–40, ???? 2010. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539110x00037.

Sommer:2010:DAH

[3102] Marianne Sommer. From descent to ascent — the human exception in the evolutionary synthesis. Nuncius, 25(1):41–67, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00046.

DePalma:2010:TCR

[3103] Armando De Palma and Germana Pareti. Tra cambiamenti radicali e conservazione: l’indagine sui fenomeni bioelettrici nella Germania del Secondo Ottocento. Nuncius, 25(1):69–100, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539110x00055.

Durand-Richard:2010:PIC

[3104] Marie-José Durand-Richard. Planimeters and integraphs in the 19th century. before the differential analyser. Nuncius, 25(1):101–124, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00064.

Luthy:2010:MEC

[3105] Christoph Lüthy. Magnifying Eighteenth-Century microscopy. Nuncius, 25(1):125–134, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539110x00073.

Rinaldi:2010:AMA

[3106] Massimo Rinaldi. Antiquity and middle ages. Nuncius, 25(1):135–136, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00082.

Barreca:2010:BRL

[3107] Francesco Barreca. Book review: Luca Bianchi, Pour une histoire de la ‘double vérité’. Paris: Vrin, 2008. 192 pp., ISBN 978-2-7116-2147-

7. Nuncius, 25(1):136–137, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539110x00091.

Fortuna:2010:BRM

[3108] Stefania Fortuna. Book review: Monica H. Green, Making Women’s Medicine Masculine. The Rise of Male Authority in Pre-Modern Gynaecology. Oxford: Univ. Press, 2008. xx + 409 pp., ISBN 978-0-19-921149-

4. Nuncius, 25(1):137–138, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539110x00109.

Beretta:2010:BRL

[3109] Marco Beretta. Book review: Lucretius, De Natuur van de dingen. Uitgegeven, vertaald, ingeleid en van aantekenineng voorzien door Piet Schrijvers. Groningen: Historische Uitgeverij, 2008. 612 pp., ISBN 978-90-6554-424-7. Nuncius, 25(1):138–140, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539110x00118.

Grimaudo:2010:BRS

[3110] Sabrina Grimaudo. Book review: Susan P. Mattern, Galen and the Rhetoric of Healing. Baltimore: The Johns Hopkins University Press, 2008. 279 pp., ISBN 978-0-8018-8835-9. Nuncius, 25(1):140–141, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00127.

Dekker:2010:BRD

[3111] Elly Dekker. Book review: David Pingree, Eastern Astrolabes. Historic Scientific Instruments of the Adler Planetarium & Astronomy Museum, Volume II. Chicago, 2009, 268 pp. ISBN 978-1-891220-02-0. Nuncius, 25 (1):141–142, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539110x00136.

Andretta:2010:REM

[3112] Elisa Andretta. Renaissance and early modern science. Nuncius, 25 (1):142–144, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539110x00145.

Antognazza:2010:BRD

[3113] Maria Rosa Antognazza. Book review: Daniel Garber, Leibniz: Body, Substance, Monad. Oxford: Oxford University Press, 2009. xxii + 428 pp., ISBN 978-0-19-956664-8. Nuncius, 25(1):144–145, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00154.

Ottaviani:2010:BRL

[3114] Alessandro Ottaviani. Book review: Leah Knight, Of Books and Botany in Early Modern England. Sixteenth-Century Plants and Print Culture. Aldershot: Ashgate, 2009. xvi + 163 pp., ISBN 978-0-7546-6586-1. Nuncius, 25(1):145–146, ???? 2010. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539110x00163.

Conforti:2010:BRA

[3115] Maria Conforti. Book review: Alessandro Pastore e Giovanni Rossi, Paolo Zacchia. Alle origini della medicina legale (1584–1659). Milano: F. Angeli, 2008. 351 pp., ISBN 978-88-568-0023-4. Nuncius, 25(1):146– 148, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00172.

Barras:2010:BRM

[3116] Vincent Barras. Book review: Marion Maria Ruisinger, Patientenwege. Die Konsiliarkorrespondenz Lorenz Heisters (1683–1758) in der Trew Sammlung Erlangen. Institut fur Geschichte der Medizin der Robert Bosch Stiftung, MedGG-Beihefte 28. Stuttgart: Franz Steiner Verlag, 2008. 308 pp., ISBN 978-3-515-08806-0. Nuncius, 25(1):148–149, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00181.

Abbri:2010:E

[3117] Ferdinando Abbri. Enlightenment. Nuncius, 25(1):149–150, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00190.

Panza:2010:BRJ

[3118] Marco Panza. Book review: J. B. Shank, The Newton Wars and the Beginning of the French Enlightenment. Chicago and London: University of Chicago Press, 2008. xvi + 572 pp., ISBN 978-0-226-74945-7. Nuncius, 25(1):150–152, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539110x00208.

Shea:2010:MCS

[3119] William R. Shea. Modern and contemporary science. Nuncius, 25(1): 152–153, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539110x00217.

Barreca:2010:BRP

[3120] Francesco Barreca. Book review: Paolo Bussotti (a cura di), Federigo Enriquese la cultura europea, pubblicazioni Centro Studi Enriques, 7. La Spezia: Agora edizioni, 2008. 298 + viii pp., ISBN 978-88-6067-089-2. Nuncius, 25(1):153–155, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539110x00226.

Beretta:2010:BRF

[3121] Marco Beretta. Book review: Fabio Minazzi (a cura di), Ludovico Geymonat. Un maestro del Novecento. Il filosofo, il partigiano e il docente. Milano: Edizioni Unicopli, 2009. 694 pp., illus., ISBN 978-88-4001-368-7. Nuncius, 25(1):155–156, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539110x00235.

Canadelli:2010:BRM

[3122] Elena Canadelli. Book review: M. J. S. Hodge, Darwin Studies. A Theorist and his Theories in their Contexts. Aldershot: Ashgate, 2009. 354 pp., ISBN 978-0-7546-5939-6. Nuncius, 25(1):156–157, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00244.

Canadelli:2010:BRK

[3123] Elena Canadelli. Book review: Karl Möbius, Asthetik der Tierwelt. Stuttgart: Franz Steiner Verlag, 2008. 188 pp., ISBN 978-3-515-09281-4. Nuncius, 25(1):157–158, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539110x00253.

Cerruti:2010:BRE

[3124] Luigi Cerruti. Book review: Eric R. Scerri, The Periodic Table. Its Story and Its Significance. Oxford: Oxford University Press, 2007. xxii + 346 pp., ISBN 978-0-19-530573-9. Nuncius, 25(1):158–159, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00262.

Peruzzi:2010:BRS

[3125] Giulio Peruzzi. Book review: Sandra Linguerri e Raffaella Simili (a cura di), Einstein parla italiano. Itinerari e polemiche. Bologna: Edizioni Pendragon, 2008. x + 359 pp., ISBN 978-88-8342-530-1. Nuncius, 25(1): 159–160, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539110x00271.

Sauer:2010:BRR

[3126] Tilman Sauer. Book review: Richard Staley, Einstein’s Generation. The Origins of the Relativity Revolution. Chicago and London: The University of Chicago Press, 2008. x + 494 pp., ISBN 978-0-226-77056-7. Nuncius, 25(1):160–161, ???? 2010. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539110x00280.

Creager:2010:BRB

[3127] Angela N. H. Creager. Book review: Bruno Strasser, La fabrique d’une nouvelle science: La biologie moléculaire à l’âge atomique (1945–1964). Firenze: Leo S. Olschki Editore, 2006. xxxiii + 450 pp., ISBN 978-88-222-5496-1. Nuncius, 25(1):161–163, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539110x00299.

Palumbo:2010:G

[3128] Giorgio G. C. Palumbo. General. Nuncius, 25(1):163–164, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00307.

Canadelli:2010:BRH

[3129] Elena Canadelli. Book review: Horst Bredekamp, Birgit Schneider, und Vera Dünkel (ed.), Das Technische Bild. Kompendium zu einer Stilgeschichte wissenschaftlicher Bilder. Berlin: Akademie Verlag, 2008. 231 pp., ISBN 978-3-05-004496-5. Nuncius, 25(1):164–165, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00316.

Pogliano:2010:BRM

[3130] Claudio Pogliano. Book review: Mariacarla Gadebusch Bondio (Hg.), Die Hand. Elemente einer Medizin- und Kulturgeschichte. Berlin: Lit Verlag, 2010. xviii + 480 pp., ISBN 978-3-8258-1958-3. Nuncius, 25 (1):165–166, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539110x00325.

Gallois:2010:BRP

[3131] William Gallois. Book review: Paul Glennie and Nigel Thrift, Shaping the Day: A History of Timekeeping in England and Wales 1300–1800. Oxford: Oxford University Press, 2009. xi + 456 pp., ISBN 978-0-19-927820-6. Nuncius, 25(1):166–167, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539110x00334.

Bonaiuti:2010:N

[3132] Elisa Bonaiuti and Iolanda Rolfo. Newsletter. Nuncius, 25(1):169–183, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539110x00343.

Rodriguez-Arribas:2010:DAI

[3133] Josefina Rodríguez-Arribas. Did Abraham ibn ezra maintain a circumsolar arrangement of Venus and mercury? Nuncius, 25(2):191–212, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-88000011.

Escobar:2010:KGL

[3134] Jorge M. Escobar. Kepler’s generalization of the law of reflection. Nuncius, 25(2):213–240, ???? 2010. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/18253911-88000012.

Milani:2010:SRS

[3135] Nausicaa Elena Milani. Il Système di Régis e le sue immagini tra nuova filosofia e censura. Nuncius, 25(2):241–297, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-88000013.

Raimondi:2010:CLA

[3136] Alessio Raimondi. Il contributo di Lamarck alla Fondazione della Meteorologia Moderna. Nuncius, 25(2):299–316, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/10.1163/18253911-
88000014.

Angeloni:2010:CRB

[3137] Roberto Angeloni. On the cultural relationship between Niels Bohr and Harald Høffding. Nuncius, 25(2):317–356, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/10.1163/18253911-
88000015.

Bianchi:2010:SGB

[3138] Simone Bianchi. Gli strumenti di Giovan Battista Amici dalla vecchia specola di Firenze al Nuovo Osservatorio di Arcetri. Nuncius, 25(2): 357–382, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/18253911-88000016.

Gattei:2010:CTA

[3139] Stefano Gattei. Copernicus’ terse argument scientific and philosophical foundations of the heliocentric hypothesis. Nuncius, 25(2):383–396, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-88000017.

Monti:2010:ASG

[3140] Maria Teresa Monti. “archaeology” of sources for a genuine “genetics” of theories. in memory of Carlo Castellani. Nuncius, 25(2):397–404, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-88000019.

Anonymous:2010:BR

[3141] Anonymous. Book reviews. Nuncius, 25(2):405–443, ???? 2010. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/10.1163/18253911-
88000020.

vanDijk:2011:PN

[3142] Arjan van Dijk. Publisher’s note. Nuncius, 26(1):1, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539111x569720.

Beretta:2011:E

[3143] Marco Beretta. Editorial. Nuncius, 26(1):3–6, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539111x569739.

deCeglia:2011:WVT

[3144] Francesco Paolo de Ceglia and Claudio Pogliano. Ways of voyaging through the human body. Nuncius, 26(1):7–19, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539111x569748.

DeSantis:2011:MV

[3145] Dario De Santis. Maiden voyage. Nuncius, 26(1):21–49, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569757.

Pogliano:2011:LOW

[3146] Claudio Pogliano. Lumen in obscuris: the winding road of modern endoscopy. Nuncius, 26(1):50–82, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x569766.

deCeglia:2011:IBF

[3147] Francesco Paolo de Ceglia. The importance of being Florentine: A journey around the world for wax anatomical Venuses. Nuncius, 26(1):83– 108, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569775.

Podgorny:2011:MEC

[3148] Irina Podgorny. Modern embalming, circulation of fluids, and the voyage through the human arterial system: Carl L. Barnes and the culture of immortality in America. Nuncius, 26(1):109–131, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569784.

Braun:2011:VVM

[3149] Ramona Braun. Voyaging in the vein: Medical experimentation with heart catheters in the Twentieth Century. Nuncius, 26(1):132–158, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569793.

Canadelli:2011:SPS

[3150] Elena Canadelli. ‘Scientific Peep Show’ the human body in contemporary science museums. Nuncius, 26(1):159–184, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539111x569801.

Cassou-Nogus:2011:EBL

[3151] Pierre Cassou-Nogus. Exploring the brain, looking for thoughts: on Asimov’s second fantastic voyage. Nuncius, 26(1):185–200, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569810.

Brodesco:2011:IGY

[3152] Alberto Brodesco. I’ve got you under my skin: Narratives of the inner body in cinema and television. Nuncius, 26(1):201–221, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569829.

Olsn:2011:BVV

[3153] Jan Eric Olsn. The body voyage as visual representation and art performance. Nuncius, 26(1):222–241, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x569838.

Chinnici:2011:MJR

[3154] Ileana Chinnici and Donatella Randazzo. A manuscript by Jesse Ramsden at Palermo Observatory. Nuncius, 26(1):243–264, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569847.

Anonymous:2011:NCa

[3155] Anonymous. Notes on contributors. Nuncius, 26(1):265–267, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x570016.

Anonymous:2011:EMA

[3156] Anonymous. The 2011 European Museum Academy Prize to the Museo Galileo. Nuncius, 26(1):269, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x579396.

Martelli:2011:GAW

[3157] Matteo Martelli. Greek alchemists at work: ’alchemical laboratory’ in the Greco–Roman Egypt. Nuncius, 26(2):271–311, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x596630.

Barreca:2011:AWJ

[3158] Francesco Barreca. The allegory of war in Johannes Kepler’s Astronomia Nova. Nuncius, 26(2):312–333, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x596649.

Luzzini:2011:MCV

[3159] Francesco Luzzini. Multa curiosa Vallisneri’s early studies on Earth sciences. Nuncius, 26(2):334–354, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x596658.

Beretta:2011:RDL

[3160] Marco Beretta. Rinman, Diderot, and Lavoisier: New evidence regarding Guillaume François Rouelle’s private laboratory and chemistry course. Nuncius, 26(2):355–379, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x596667.

Micheli:2011:SSC

[3161] Gianni Micheli. Sistemi di sapere a confronto. Nuncius, 26(2):381–389, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x596676.

Bernardoni:2011:ALM

[3162] Andrea Bernardoni. Atomism in late medieval philosophy and theology. Nuncius, 26(2):391–392, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x569856.

Acerbi:2011:APC

[3163] Fabio Acerbi. Apollonius de Perge, Coniques. Tome 2.1: Livres II et III. Nuncius, 26(2):393–394, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x569865.

Acerbi:2011:PAB

[3164] Fabio Acerbi. Pappus of Alexandria: Book 4 of the collection. Nuncius, 26(2):395–396, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539111x569874.

Sannino:2011:OPI

[3165] Antonella Sannino. Opus praeclarum de imaginibus astrologicis. Nuncius, 26(2):397–398, ???? 2011. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539111x569883.

Martelli:2011:AHP

[3166] Matteo Martelli. The Arabic Hermes: From pagan sage to prophet of science. Nuncius, 26(2):399–400, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x569892.

Acerbi:2011:AEC

[3167] Fabio Acerbi. Arquimedes, sobre l’esfera i el cilindre. Nuncius, 26(2): 401–402, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539111x596685.

Giglioni:2011:L

[3168] Guido Giglioni. Lettere. Nuncius, 26(2):403–404, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539111x596694.

Gualandi:2011:S

[3169] Andrea Gualandi. On sunspots. Nuncius, 26(2):405–406, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x596702.

deCeglia:2011:CME

[3170] Francesco Paolo de Ceglia. Centres of medical excellence? medical travel and education in Europe, 1500–1789. Nuncius, 26(2):407–408, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569900.

Bernardoni:2011:WOF

[3171] Andrea Bernardoni. The worlds of oronce fine: Mathematics, instruments and print in Renaissance France. Nuncius, 26(2):409–410, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569919.

Bernardoni:2011:GLN

[3172] Andrea Bernardoni. Glossario Leonardiano: Nomenclatura delle macchine nei codici Madrid e Atlantico. Nuncius, 26(2):411–412, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x596711.

Clericuzio:2011:PAR

[3173] Antonio Clericuzio. Philosophie et alchimie la Renaissance et l’Age Classique. scolastique et alchimie (xvie–xviie siècles); Andreas Libavius and the transformation of alchemy: Separating chemical cultures with polemical fire. Nuncius, 26(2):413–415, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x569928.

Fabbri:2011:MUA

[3174] Natacha Fabbri. La musurgia universalis di Athanasius Kircher. Contenuti, fonti, terminologia. Nuncius, 26(2):416–417, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539111x596720.

Barreca:2011:RWN

[3175] Francesco Barreca. Reconfiguring the world: Nature, god, and human understanding from the Middle Ages to Early Modern Europe. Nuncius, 26(2):418–419, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539111x596739.

Conforti:2011:WMS

[3176] Maria Conforti. The world makers: Scientists of the restoration and the search for the origins of the Earth. Nuncius, 26(2):420–421, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569937.

Ciuti:2011:RRS

[3177] Francesco Ciuti. Regolamento dei Regi Spedali di Santa Maria Nuova e di Bonifazio. Nuncius, 26(2):422–423, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x596748.

Mazur:2011:CIV

[3178] Peter A. Mazur. Corpi invasi e viaggi dell’anima: Santit, possessioni, esorcismo dalla teologia barocca alla medicina illuminista. Nuncius, 26 (2):424–425, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539111x569946.

Monti:2011:FTD

[3179] Maria Teresa Monti. La di ferenza tra la dottrina di Stahl e la mia in patologia e terapia. Nuncius, 26(2):426–427, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569955.

Strazzoni:2011:LLD

[3180] Andrea Strazzoni. L’arbre et le labyrinthe, Descartes selon l’ordre des Lumières. Nuncius, 26(2):428–431, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x569964.

Roscioni:2011:LTM

[3181] Lisa Roscioni. Liberi tutti: Manicomi e psichiatri in Italia: una storia del novecento. Nuncius, 26(2):432–433, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x596757.

Canadelli:2011:MMF

[3182] Elena Canadelli. Mirrors of memory: Freud, photography, and the history of art. Nuncius, 26(2):434–435, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x596766.

Pogliano:2011:CHS

[3183] Claudio Pogliano. The Cambridge history of science. vol. 6: the modern biological and Earth sciences. Nuncius, 26(2):436–437, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x569973.

Bernardoni:2011:SCH

[3184] Andrea Bernardoni. Spaces and collections in the history of science: the laboratorio chimico overture. Nuncius, 26(2):438–440, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x596775.

Mazzarello:2011:LDR

[3185] Paolo Mazzarello. Laboratory disease: Robert Koch’s medical bacteriology. Nuncius, 26(2):441–442, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x569982.

Schmitt:2011:GDT

[3186] Stphane Schmitt. Über Galvanismus und deutsche Trumereien. Zur Rezeption romantischer Naturforschung in Frankreich zwischen 1800 und 1820. Nuncius, 26(2):443–445, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x596784.

Abbri:2011:MMS

[3187] Ferdinando Abbri. The making of modern science. science, technology, medicine, and modernity: 1789-1914. Nuncius, 26(2):446–447, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x596793.

Gattei:2011:SHR

[3188] Stefano Gattei. Science and its history: A reassessment of the historiography of science. Nuncius, 26(2):448–450, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539111x569991.

Gosseye:2011:RLA

[3189] Lise Gosseye. Representing light across arts and sciences: Theories and practices. Nuncius, 26(2):451–452, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539111x570007.

Spataro:2011:SSD

[3190] Stefano Spataro. Scorci di storia della scienza. Nuncius, 26(2):453–454, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x596801.

Maddaluno:2011:SLM

[3191] Lavinia Maddaluno. The scientific life: A moral history of a late modern vocation; never pure: Historical studies of science as if it was produced by people with bodies, situated in time, space, culture, and society, and struggling for credibility and authority. Nuncius, 26(2):455–458, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x596810.

Anonymous:2011:NCb

[3192] Anonymous. Notes on contributors. Nuncius, 26(2):459–463, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x598520.

Anonymous:2011:CV

[3193] Anonymous. Contents volume 26 (2011). Nuncius, 26(2):465–468, ???? 2011. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539111x600120.

Abbri:2012:PRD

[3194] Ferdinando Abbri. Paolo Rossi: 30 December 1923–14 January 2012. Nuncius, 27(1):1–10, ???? 2012. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539112x636319.

Bernardoni:2012:LCA

[3195] Andrea Bernardoni. Leonardo and the “chemical arts”. Nuncius, 27 (1):11–55, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539112x637165.

Maddaluno:2012:UNW

[3196] Lavinia Maddaluno. Unveiling nature: Wonder and deception in Eighteenth-Century London shows and exhibitions. Nuncius, 27(1):56– 80, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539112x637174.

Menin:2012:JJR

[3197] Marco Menin. Jean-Jacques Rousseau vitaliste: La moralisation de l’hygiène médicale entre régime diététique et éthique alimentaire. Nuncius, 27(1):81–109, ???? 2012. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539112x636328.

Donato:2012:SFE

[3198] Maria Pia Donato. Science on the fringe of the Empire: the Academy of the Linceans in the early Nineteenth Century. Nuncius, 27(1):110–140, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539112x637183.

Pogliano:2012:PHO

[3199] Claudio Pogliano. Penfield’s homunculus and other grotesque creatures from the Land of If. Nuncius, 27(1):141–162, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539112x637192.

Beretta:2012:IER

[3200] Marco Beretta. Imaging the experiments on respiration and transpiration of Lavoisier and Séguin: Two unknown drawings by Madame Lavoisier. Nuncius, 27(1):163–191, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/182539112x637200.

Reeves:2012:GDM

[3201] Eileen Reeves. Grand designs: Mutio Oddi in exile from Urbino. Nuncius, 27(1):193–205, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539112x636337.

Pardo-Tomas:2012:REM

[3202] José Pardo-Tomás. Renaissance and early modern science. Nuncius, 27 (1):207–209, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539112x636346.

Shea:2012:BRN

[3203] Willliam Shea. Book review: Natacha Fabbri, De l’utilité de l’harmonie. Filosofia, scienza e musica in Mersenne, Descartes e Galileo. Pisa: Edizioni della Normale, 2008. xvi + 314 pp., ISBN 978-88-7642-321-5. Nuncius, 27(1):210–211, ???? 2012. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/182539112x636355.

Camerota:2012:BRA

[3204] Filippo Camerota. Book review: Anthony Gerbino and Stephen Johnston, Compass and Rule: Architecture as Mathematical Practice in England. New Haven and London: Yale University Press, 2009. 208 pp., ISBN 978-0-300-15093-3. Nuncius, 27(1):212–214, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539112x636364.

Corsi:2012:BRN

[3205] Elisabetta Corsi. Book review: Noël Golvers and Efthymios Nicolaidis, Ferdinand Verbiest and Jesuit Science in 17th century China. An annotated edition and translation of the Constantinople manuscript (1676). Athens-Leuven: Ferdinand Verbiest Institute; Katholieke Universiteit Leuven, Leuven Chinese Studies — XIX; Institute for Neohellenic Research, National Hellenic Research Foundation — 108; Sources of Modern Greek Literature and Learning, 2009. 381 pp., 117 ill., ISBN 978-960-7916-83-9, 978-90-801833-9-1. Nuncius, 27(1):215–217, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539112x636373.

Sanchez:2012:BRA

[3206] José Ramón Bertomeu Sánchez. Book review: Alessandro Pastore, Veleno. Credenze, crimini, saperi nell’Italia moderna. Bologna: Il Mulino, 2010. 254 pp., ill., ISBN 978-88-15-13873-6. Nuncius, 27(1):218–219, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539112x636382.

Vagelli:2012:MCS

[3207] Matteo Vagelli. Modern and contemporary science. Nuncius, 27(1): 220–221, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539112x636391.

Meschiari:2012:BRB

[3208] Alberto Meschiari. Book review: Brian Bracegirdle, A History of Photography with the Light Microscope. Quekett Microscopical Club, 2010, viii + 221 pp., ill., ISBN 978-0-9564591-1-4. Nuncius, 27(1):222–223, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539112x636409.

Conforti:2012:BRL

[3209] Maria Conforti. Book review: Luca Ciancio, Le colonne del Tempo. Il “Tempio di Serapide” a Pozzuoli nella storia della geologia, dell’archeologia e dell’arte (1750–1900). Firenze: Edifir, 2009. 342 pp., ill., ISBN 978-88-7970-416-8. Nuncius, 27(1):224–227, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/182539112x636418.

Nuvolari:2012:BRR

[3210] Alessandro Nuvolari. Book review: Ross Thomson, Structures of Change in the Mechanical Age. Technological Innovation in the United States 1790–1865. Baltimore: Johns Hopkins University Press. Nuncius, 27 (1):228–230, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/182539112x636427.

Bresadola:2012:G

[3211] Marco Bresadola. General. Nuncius, 27(1):231–233, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/content/10.1163/182539112x636436.

deCeglia:2012:I

[3212] Francesco Paolo de Ceglia and Irina Podgorny. Introduction. Nuncius, 27(2):235–240, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/18253911-02702001.

deCeglia:2012:SSC

[3213] Francesco Paolo de Ceglia. The science of Santa Claus: Discussions on the manna of Nicholas of Myra in the Modern Age. Nuncius, 27 (2):241–269, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/18253911-02702002.

Redworth:2012:GGS

[3214] Glyn Redworth. God’s gift? Sacred relics, gift giving, and Luisa de Carvajal’s preparation of the Holy during the Long Reformation. Nuncius, 27(2):270–288, ???? 2012. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/18253911-02702003.

Podgorny:2012:CDS

[3215] Irina Podgorny. Changing the dead to statues of stone: the synthesis of fossils, petrifaction, photography, and the chemistry of the Gorgonean arts. Nuncius, 27(2):289–308, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/18253911-02702004.

Marinozzi:2012:EAM

[3216] Silvia Marinozzi. The embalming art in the Modern Age: the mummies of Caroline, Letizia and Joachim-Napoleon Agar as examples of funerary rites in the Napoleonic Empire. Nuncius, 27(2):309–329, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02702005.

Dias:2012:NCF

[3217] Nélia Dias. Nineteenth-Century French collections of skulls and the cult of bones. Nuncius, 27(2):330–347, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/18253911-02702006.

Constantino:2012:HLL

[3218] María Eugenia Constantino and Antonio Lafuente. The hidden logistics of Longinos’s novohispanic cabinet. Nuncius, 27(2):348–370, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02702007.

Ruve:2012:FER

[3219] Gerlind Rüve. Focusing on the exception to the rule for a time: on the border between life and death in a case of Scheintod in 1833. Nuncius, 27(2):371–392, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/18253911-02702008.

Knoblauch:2012:DCD

[3220] Hubert Knoblauch. The decline of clinical dissections and the “culture of death”. Nuncius, 27(2):393–410, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/18253911-02702009.

Grazzini:2012:DI

[3221] Maria Grazia Grazzini. Documenta inedita. Nuncius, 27(2):411–467, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02702010.

Bresadola:2012:WAM

[3222] Marco Bresadola. Wax anatomy and the material culture of enlightenment science. Nuncius, 27(2):469–476, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/10.1163/18253911-
02702011.

Ipsen:2012:LWI

[3223] Carl Ipsen. The last word on Italian science? Nuncius, 27(2):477–488, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02702012.

Nenci:2012:BRM

[3224] Elio Nenci. Book review: Matteo Valleriani, Galileo Engineer. Boston Studies in the Philosophy of Science vol. 269, Dordrecht: Springer, 2010, 320 pp., ISBN 978-90-481-8644-0. Nuncius, 27(2):489–490, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02702013.

Gattei:2012:BRH

[3225] Stefano Gattei. Book review: Horst Bredekamp (ed.), Galileo’s O. Vol. I: Galileo’s Sidereus Nuncius. A Comparison of the Proof Copy (New York) with Other Paradigmatic Copies, ed. by Irene Brückle and Oliver Hahn; vol. II: Paul Needham, Galileo Makes a Book. The First Edition of Sidereus Nuncius, Venice 1610. Berlin: Akademie Verlag, 2011, 192, 249 pp.; ISBN 978-3-05-005095-9. Nuncius, 27(2):491–495, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02702014.

Pogliano:2012:BRR

[3226] Claudio Pogliano. Book review: Renato G. Mazzolini and Hans–Jörg Rheinberger (eds.), Differing Routes to Stem Cell Research: Germany and Italy. Bologna-Berlin: Il Mulino-Duncker & Humblot, 2012, 272 pp., ISBN 978-88-15-23878-8; 978-3-428-13849-4. Nuncius, 27(2):496– 499, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02702015.

Anonymous:2012:CV

[3227] Anonymous. Contents volume 27 (2012). Nuncius, 27(2):501–503, ???? 2012. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02702016.

Beretta:2013:TSR

[3228] Marco Beretta and Alessandro Tosi. Tennis and the scientific revolution. Nuncius, 28(1):1–4, ???? 2013. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/18253911-02801001.

Angelini:2013:PCC

[3229] Annarita Angelini. Praecisio and conjecture: Cusanus’ ball game and the ‘Learned Ignorance’ of the world. Nuncius, 28(1):5–18, ????

2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02801002.

Beretta:2013:TTP

[3230] Marco Beretta. Training tennis players through natural philosophy: From Scaino’s trattato to Garsault’s art du paumier. Nuncius, 28(1): 19–42, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02801003.

Pennuto:2013:JPH

[3231] Concetta Pennuto. Jeu de la paume: Health of the body and the mind in early modern medicine. Nuncius, 28(1):43–65, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02801004.

Gattei:2013:GTR

[3232] Stefano Gattei. Galileo and tennis: Reconciling the new physics with commonsense. Nuncius, 28(1):66–84, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/10.1163/18253911-
02801005.

Tosi:2013:TEM

[3233] Alessandro Tosi. Tennis in early modern visual culture. Nuncius, 28 (1):85–114, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/18253911-02801006.

Dijksterhuis:2013:JPJ

[3234] Fokko Jan Dijksterhuis. Jeu de paume & jeux de la raison in Seventeenth-Century optics. Nuncius, 28(1):115–141, ???? 2013. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/10.1163/18253911-
02801007.

Sylla:2013:JBM

[3235] Edith Dudley Sylla. Jacob Bernoulli and the mathematics of tennis. Nuncius, 28(1):142–163, ???? 2013. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/10.1163/18253911-02801008.

Lehman:2013:ARM

[3236] Christine Lehman. Alchemy revisited by the mid-eighteenth century chemists in France: An unpublished manuscript by Pierre-Joseph Macquer. Nuncius, 28(1):165–216, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/18253911-02801010.

Brenni:2013:GET

[3237] Paolo Brenni. Gerard L’Estrange Turner 22 January 1926–19 July 2012. Nuncius, 28(1):217–222, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/10.1163/18253911-02801009.

Saito:2013:AMA

[3238] Ken Saito. Antiquity and the middle ages. Nuncius, 28(1):223–228, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02801016.

Ugaglia:2013:BRT

[3239] Monica Ugaglia. Book review: Tony Roark, Aristotle on Time. A Study of the Physics. Cambridge: Cambridge University Press, 2011. 248 pp., ISBN 978-1-107-00262-3. Nuncius, 28(1):229–231, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02801011.

Abbri:2013:REM

[3240] Ferdinando Abbri. Renaissance and early modern science. Nuncius, 28 (1):232–234, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/18253911-02801012.

Pugliano:2013:BRJ

[3241] Valentina Pugliano. Book review: James Shaw and Evelyn Welch, Making and Marketing Medicines in Renaissance Florence (Clio Medica 89/ The Wellcome Series in the History of Medicine). Amsterdam and New York: Rodopi, 2011. 356 pp., ISBN 978-90-420-3156-2. Nuncius, 28(1): 235–237, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/18253911-02801013.

deCeglia:2013:MCS

[3242] Francesco Paolo de Ceglia. Modern and contemporary science. Nuncius, 28(1):238–239, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/10.1163/18253911-02801014.

Kragh:2013:BRK

[3243] Helge Kragh. Book review: Kostas Gavroğlu, Ana Simões, Neither Physics nor Chemistry: A History of Quantum Chemistry. Cambridge, Mass.: Mit Press, 2012, pp. xiv + 351. Nuncius, 28(1):240–242, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
10.1163/18253911-02801015.

Keller:2013:RET

[3244] Vera Keller. Re-entangling the thermometer: Cornelis Drebbel’s description of his self-regulating oven, the regiment of fire, and the early history of temperature. Nuncius, 28(2):243–275, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802001.

Pecontal:2013:MML

[3245] Emmanuel Pécontal. Les mires méridiennes lointaines de l’observatoire de Lyon: Recherches bibliographiques, archivistiques et archéologiques. (French) [distant meridian patterns of the Observatory of Lyon: bibliographic, archival, and archaeological searches]. Nuncius, 28(2):276–312, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802002.

Pogliano:2013:EMH

[3246] Claudio Pogliano. Eye, mind, hand: Filippo Pacini’s microscopy. Nuncius, 28(2):313–344, ???? 2013. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-02802003.

Gattei:2013:WST

[3247] Stefano Gattei. The wandering Scot: Thomas Seget’s album amicorum. Nuncius, 28(2):345–463, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02802026.

Dupre:2013:TGT

[3248] Sven Dupré. The transnational Galileo: a telescopic view from somewhere. Nuncius, 28(2):465–476, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02802004.

Ecca:2013:BR

[3249] Giulia Ecca. Book reviews. Nuncius, 28(2):477–479, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-02802005.

Strano:2013:BRD

[3250] Giorgio Strano. Book review: David A. King, Astrolabes from Medieval Europe. Aldershot: Ashgate, Variorum, 2011. xvi + 404 pp., ill., ISBN 978-1-4094-2593-9. Nuncius, 28(2):480–481, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802006.

Martelli:2013:BRM

[3251] Matteo Martelli. Book review: Michaël Martin, La magie dans l’antiquité. Paris: Ellipses Édition, 2012. 159 pp., ISBN 978-2-7298-7304-2. Nuncius, 28(2):482–483, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02802008.

Nenci:2013:REM

[3252] Elio Nenci. Renaissance and early modern science. Nuncius, 28(2):484– 486, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802009.

Grant:2013:BRS

[3253] Florence Grant. Book review: Sven Dupré and Christoph Lüthy (eds.), Silent Messengers. The Circulation of Material Objects of Knowledge in the Early Modern Low Countries. Münster: LIT Verlag, 2011. 387 pp., ISBN 978-3-8258-1635-3. Nuncius, 28(2):487–489, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802010.

Bernardoni:2013:BRP

[3254] Andrea Bernardoni. Book review: Pamela O. Long, Artisan / Practitioners and the Rise of the New Sciences, 1400–1600. Corvallis: Oregon State University Press, 2011. xii + 196 pp., Ill., ISBN 978-0-87071-609-6. Nuncius, 28(2):490–491, ???? 2013. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-02802011.

Bernardoni:2013:BRC

[3255] Andrea Bernardoni. Book review: Cesare S. Maffioli, La via delle acque (1500–1700): Appropriazione delle arti e trasformazioni delle matematiche. Firenze: Olschki, 2010, ill., 394 pp., ISBN 978-88-222-6008-6. Nuncius, 28(2):492–493, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02802012.

Conforti:2013:BRC

[3256] Maria Conforti. Book review: Craig Martin, Renaissance Meteorology, Pomponazzi to Descartes. Baltimore, The Johns Hopkins University Press, 2011. 213 pp., ISBN 978-1-4214-0187-4. Nuncius, 28(2):494–495, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802013.

Milighetti:2013:BRM

[3257] Maria Chiara Milighetti. Book review: Michel Paoli e Monica Preti, L’Arioste et les Arts, Milano, Officina Libraria-Paris, Musée du Louvre, 2012. 336 pp., ISBN Musée du Louvre: 978-2-35031-368-9, ISBN Officina Libraria: 978-88-89854-68-6. Nuncius, 28(2):496–497, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802014.

Lillie:2013:BRA

[3258] Lisa Lillie. Book review: Alina Payne, The telescope and the compass. Teofilo Gallaccini and the Dialogue between architecture and science in the age of Galileo. Firenze, Leo S. Olschki Editore, 2012. 240 pp., ISBN 978-88-222-6122-9, ISSN 0066-6815. Nuncius, 28(2):498–499, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802015.

Verardi:2013:BRO

[3259] Donato Verardi. Book review: Ornella Pompeo Faracovi, Lo specchio alto. Astrologia e filosofia fra Medioevo e prima età moderna, Bruniana & Campanelliana, Supplementi, XXXII — Studi, 11, Fabrizio Serra Editore, Pisa 2012. 205 pp., ISBN 978-88-6227-449-4. Nuncius, 28(2): 500–502, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-02802016.

Beretta:2013:E

[3260] Marco Beretta. Enlightenment. Nuncius, 28(2):503–505, ???? 2013. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-02802017.

Spary:2013:BRD

[3261] E. C. Spary. Book review: Daniela Bleichmar, Visible Empire: Botanical Expedition and Visual Culture in the Hispanic Enlightenment. Chicago and London: The University of Chicago Press, 2012. 286 pp., ISBN 978-0-226-05853-5. Nuncius, 28(2):506–508, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802018.

Adamo:2013:MCS

[3262] Angelo Adamo. Modern and contemporary science. Nuncius, 28(2): 509–510, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-02802019.

Clarke:2013:BRH

[3263] Imogen Clarke. Book review: Helge Kragh, Niels Bohr and the Quantum Atom. The Bohr Model of Atomic Structure 1913–1925. Oxford: Oxford University Press, 2012. vi + 410 pp., ISBN 978-0-19-965498-7. Nuncius, 28(2):511–512, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-02802020.

deCeglia:2013:BRI

[3264] Francesco Paolo de Ceglia. Book review: Irina Podgorny, Charlatanes. Crónicas de remedios incurables. Buenos Aires: Eterna Cadencia Editora, 2012. 352 pp., ISBN 978-987-1673-67-4. Nuncius, 28(2):513–514, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802021.

Guerra:2013:BRL

[3265] Corinna Guerra. Book review: Leslie Tomory, Progressive Enlightenment. The Origins of the Gaslight Industry, 1780-1820. Cambridge: The MIT Press, 2012. ISBN 978-0-262-01675-9. Nuncius, 28(2):515–517, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802022.

Maddaluno:2013:G

[3266] Lavinia Maddaluno. General. Nuncius, 28(2):518–521, ???? 2013. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-02802023.

Martelli:2013:BRE

[3267] Matteo Martelli. Book review: Efthymios Nicolaidis, Science and Eastern Orthodoxy. From the Greek Fathers to the Age of Globalization, translated by Susan Emanuel. Baltimore, John Hopkins University Press (“Medicine, Science, and Religion in Historical Context”), 2011. xviii + 252 pp., ISBN-13: 978-1-4214-0298-7. Nuncius, 28(2):522–523, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802024.

Canadelli:2013:BRT

[3268] Elena Canadelli. Book review: Theodore W. Pietsch, Trees of Life. A Visual History of Evolution. Baltimore: The Johns Hopkins University Press, 2012. 358 pp., ISBN 978-1-4214-0479-0. Nuncius, 28(2):524–526, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802025.

Anonymous:2013:CV

[3269] Anonymous. Contents volume 28 (2013). Nuncius, 28(2):527–529, ???? 2013. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02802027.

Valleriani:2014:I

[3270] Matteo Valleriani. Introduction. Nuncius, 29(1):1–8, ???? 2014. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-02901001.

Roby:2014:EGR

[3271] Courtney Roby. Experiencing geometry in Roman surveyors’ texts. Nuncius, 29(1):9–52, ???? 2014. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-02901002.

vanLeeuwen:2014:TLD

[3272] Joyce van Leeuwen. Thinking and learning from diagrams in the Aristotelian mechanics. Nuncius, 29(1):53–87, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-02901003.

Siebert:2014:TEO

[3273] Harald Siebert. Transformation of Euclid’s optics in Late Antiquity. Nuncius, 29(1):88–126, ???? 2014. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-02901004.

Valleriani:2014:APT

[3274] Matteo Valleriani. Ancient pneumatics transformed during the Early Modern Period. Nuncius, 29(1):127–173, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-02901005.

McLaughlin:2014:E

[3275] Peter McLaughlin. Epilogue. Nuncius, 29(1):175–177, ???? 2014. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-02901006.

Bennett:2014:CCA

[3276] Jim Bennett and Giorgio Strano. The so-called ‘Chaucer Astrolabe’ from the Koelliker Collection, Milan. Nuncius, 29(1):179–229, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02901007.

Mertens:2014:BRP

[3277] Michèle Mertens. Book review: Pseudo-Democrito, Scritti alchemici: con il commentario di Sinesio, edizione critica del testo greco, trad. e commento di Matteo Martelli, pref. di Tiziano Dorandi (Textes et Travaux de Chrysopoeia, 12). Paris: Séha-Milano: Archè, 2011. 523 pp. ISBN 978-88-7252-319-3. Nuncius, 29(1):231–234, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-02901008.

Tessicini:2014:BRM

[3278] Dario Tessicini. Book review: Monica Azzolini, The Duke and the Stars: Astrology and Politics in Renaissance Milan (I Tatti Studies in Italian Renaissance History). Cambridge (MA) — London: Harvard University Press, 2013. 370 pp., ISBN 978-0-674-06663-2. Nuncius, 29(1):235–238, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02901009.

Cambi:2014:BRE

[3279] Maurizio Cambi. Book review: Eva del Soldato, Simone Porzio: Un aristotelico tra natura e grazia. Roma: Edizioni di Storia e Letteratura, 2010. xvi + 352 pp., ISBN 978-88-6372-275-8. Nuncius, 29(1):239–243, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02901010.

Bernardoni:2014:BRA

[3280] Andrea Bernardoni. Book review: Aline Durand (ed.), Des outils des machines et des hommes: études offertes à Georges Comet. With the collaboration of Laurent Maggiori. Aix-en-Provence: Presses universitaires de Provence, 2012, 236 p.: ill., ritr.; 21cm, Cahiers d’histoire des techniques, 1158-8403; 8, ISBN 978-2-85399-797-3. Nuncius, 29(1):244–247, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02901011.

Gattei:2014:BRS

[3281] Stefano Gattei. Book review: Stephen Greenblatt, The Swerve: How the World Became Modern. New York-London: W.W. Norton & Company, 2011. 356 pp., ISBN 978-0-393-34340-3; $26.95. Nuncius, 29(1):248–250, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02901012.

Bernardoni:2014:BRM

[3282] Andrea Bernardoni. Book review: Matteo Valleriani (ed.), Metallurgy, Ballistics and Epistemic Instruments. The Nova scientia of Nicolò Tartaglia. A New Edition (Sources 6, Edition open access). Berlin: Max Planck Research Library for the History and Development of Knowledge, 2013. 56+304 pp., ISBN 978-3-8442-5258-3. Nuncius, 29(1):251– 253, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02901013.

Barreca:2014:BRE

[3283] Francesco Barreca. Book review: Erhard Weigel, Philosophia Mathematica, Theologia Naturalis Solida (Clavis Pansophiae 3,4). Stuttgart-Bad Cannstatt: Fromman-Holzboog Verlag, 2013, 2 voll. lxvii + 372 pp.; lvii + 436 pp., ISBN 978-3-7728-2540-8. Nuncius, 29(1):254–255, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02901014.

Pepe:2014:BRR

[3284] Luigi Pepe. Book review: Roger Hahn (ed.), Correspondance de Pierre Simon Laplace (1749–1727). 2 vols. Turnhout: Brepols, 2013. 1416 p., ISBN 978-2-503-54129-7. Nuncius, 29(1):256–259, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02901015.

deCeglia:2014:BRE

[3285] Francesco Paolo de Ceglia. Book review: Elena Canadelli, “Più positivo dei positivisti”: Antropologia, psicologia, evoluzionismo in Tito Vignoli. Pisa: ETS, 2013. 188 pp., ISBN 978-88-467-3603-1. Nuncius, 29(1): 260–262, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-02901016.

Barreca:2014:BRA

[3286] Francesco Barreca. Book review: Andrea Del Centina and Alessandra Fiocca, Guglielmo Libri matematico e storico della matematica: L’irresistibile ascesa dall’Ateneo Pisano all’Institut de France. Firenze:

Leo S. Olschki, 2010. XXI, 320, xxii pp. + CD-ROM, ISBN 978-88-222-6003-1. Nuncius, 29(1):263–264, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02901017.

Micheli:2014:BRP

[3287] Gianni Micheli. Book review: Pierre Souffrin, Écrits d’histoire des sciences. Paris: Les Belles Lettres, 2012. 497 pp., ISBN 978-2-251-42047-9. Nuncius, 29(1):265–272, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02901018.

Fregonese:2014:BRP

[3288] Lucio Fregonese. Book review: Peter Heering and Roland Wittje (eds.), Learning by Doing: Experiments and Instruments in the History of Science Teaching. Stuttgart, Germany: Franz Steiner Verlag, 2011. 362, pp. ISBN 978-3-515-09842-7. Nuncius, 29(1):273–275, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02901019.

Fregonese:2014:BRK

[3289] Lucio Fregonese. Book review: Klaus Staubermann (ed.), Reconstructions: Recreating Science and Technology of the Past. Edinburgh: National Museum of Scotland, 2011. xii + 276 pp., ISBN 978-1-905267-48-4. Nuncius, 29(1):276–278, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02901020.

Martines:2014:IC

[3290] Giangiacomo Martines. Isidore’s compass. Nuncius, 29(2):279–311, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02902005.

Yassi:2014:MJO

[3291] Yousef Yassi. The magic jar and other devices from 9th Century Persia. Nuncius, 29(2):312–323, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02902006.

Raynaud:2014:TAD

[3292] Dominique Raynaud. A tentative astronomical dating of Ibn al-Haytham’s solar eclipse record. Nuncius, 29(2):324–358, ???? 2014. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-02902003.

Omodeo:2014:AAV

[3293] Pietro T. Omodeo. The authenticity of Amerigo Vespucci’s Mundus Novus and information untold about his third journey. Nuncius, 29(2): 359–388, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-02902001.

Sanchez:2014:OIW

[3294] Antonio Sánchez. An official image of the world for the Hispanic monarchy. Nuncius, 29(2):389–438, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02902002.

Suay:2014:KRF

[3295] Juan Miguel Suay and David Teira. Kites: The rise and fall of a scientific object. Nuncius, 29(2):439–463, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02902004.

Julich:2014:CHL

[3296] Solveig Jülich. Colouring the human landscapes. Nuncius, 29(2):464– 497, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-02902007.

Ricciardi:2014:DCC

[3297] Massimo Ricciardi and Maria Laura Castellano. Domenico Cirillo’s collections. Nuncius, 29(2):499–530, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02902008.

Cuomo:2014:BRD

[3298] Serafina Cuomo. Book review: Domninus of Larissa, Encheiridion and Spurious Works. Introduction, critical text, English translation, and commentary by Peter Riedlberger (Mathematica Graeca Antiqua 2), Pisa and Roma: Fabrizio Serra Editore, 2013. 285 pp., ISBN 978-88-6227-567-5. Nuncius, 29(2):531–534, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-02902009.

Abbri:2014:BRA

[3299] Ferdinando Abbri. Book review: Angela Bandinelli, Le origini chimiche della vita. Legami tra la Rivoluzione di Lavoisier e la Biologia di Lamarck. Firenze: Olschki, 2013. 252 pp., ISBN 978-88-222-6270-7. Nuncius, 29 (2):535–538, ???? 2014. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-02902010.

Borrelli:2015:I

[3300] Arianna Borrelli, Michael Korey, and Volker R. Remmert. Introduction. Nuncius, 30(1):1–7, ???? 2015. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03001001.

Perkins:2015:PI

[3301] Emma L. Perkins and Liba Taub. “perhaps irrelevant”. Nuncius, 30 (1):9–36, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03001002.

Gaulke:2015:PES

[3302] Karsten Gaulke. Perfect in every sense. Nuncius, 30(1):37–74, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03001003.

Gessner:2015:HN

[3303] Samuel Gessner. Heavenly networks. Nuncius, 30(1):75–95, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03001004.

Mersmann:2015:MSM

[3304] Jasmin Mersmann. Moving shadows, moving Sun. Nuncius, 30(1):96– 123, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03001005.

Ribouillault:2015:AHG

[3305] Denis Ribouillault. Atlas and Hercules in the garden. Nuncius, 30(1): 124–160, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03001006.

Behrmann:2015:MJ

[3306] Carolin Behrmann. Metrics of justice. Nuncius, 30(1):161–194, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03001007.

Bianchi:2015:OAI

[3307] Simone Bianchi and Daniele Galli. Les observatoires astronomiques en Italie. Nuncius, 30(1):195–227, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03001008.

Chinnici:2015:PME

[3308] Ileana Chinnici and Paolo Brenni. The Palermo Merz Equatorial Telescope. Nuncius, 30(1):228–279, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03001009.

Ottaviani:2015:CD

[3309] Alessandro Ottaviani. The coral of death. Nuncius, 30(2):281–319, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03002001.

Candela:2015:RT

[3310] Andrea Candela. The radium terrors. Nuncius, 30(2):320–344, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03002002.

Timmermann:2015:AC

[3311] Anke Timmermann. Alchemy in Cambridge. Nuncius, 30(2):345–511, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03002003.

Ecca:2015:BRL

[3312] Giulia Ecca. Book review: L’Empirisme d’Épicure, written by Julie Giovacchini. Nuncius, 30(2):513–516, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03002004.

Martelli:2015:BRS

[3313] Matteo Martelli. Book review: The Secrets of Alchemy, written by Lawrence M. Principe. Nuncius, 30(2):517–519, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03002005.

Nenci:2015:BRS

[3314] Elio Nenci. Book review: Le strade della tecnica. Tecnologia e pratica della scienza nel mondo antico, written by Giovanni Di Pasquale. Nuncius, 30(2):520–522, ???? 2015. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03002006.

Strano:2015:BRP

[3315] Giorgio Strano. Book review: I Pronostici di Domenico Maria da Novara, edited by Fabrizio Bònoli, Giuseppe Bezza, Salvo De Meis, Cinzia Colavita. Nuncius, 30(2):523–524, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03002007.

Brockmann:2015:BRB

[3316] Sophie Brockmann. Book review: The Bishop’s Utopia. Envisioning Improvement in Colonial Peru, written by Emily Berquist Soule. Nuncius, 30(2):525–527, ???? 2015. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03002008.

Conforti:2015:I

[3317] Maria Conforti and Jeanne Peiffer. Introduction. Nuncius, 30(3):529– 541, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03003018.

Doherty:2015:GLN

[3318] Meghan C. Doherty. Giving light to narrative. Nuncius, 30(3):543–569, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03003004.

Conforti:2015:IPF

[3319] Maria Conforti. Illustrating pathologies in the first years of the Miscellanea Curiosa, 1670–1687. Nuncius, 30(3):570–609, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03003006.

Gessner:2015:JIM

[3320] Samuel Gessner. The journals and the instrument maker. Nuncius, 30 (3):610–636, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03003005.

Canadelli:2015:AI

[3321] Elena Canadelli. Authoritative images. Nuncius, 30(3):637–674, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03003001.

Bigg:2015:TSC

[3322] Charlotte Bigg. Travelling scientist, circulating images and the making of the modern scientific journal. Nuncius, 30(3):675–698, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03003002.

Barreca:2015:CA

[3323] Francesco Barreca. Chicago–Arcetri 1933. Nuncius, 30(3):699–707, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03003003.

Clericuzio:2015:SCR

[3324] Antonio Clericuzio. Science in 17th Century Rome. Nuncius, 30(3): 709–714, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03003007.

Strano:2015:BRI

[3325] Giorgio Strano. Book review: Illustrating the Phaenomena: Celestial Cartography in Antiquity and the Middle Ages, written by Elly Dekker. Nuncius, 30(3):715–717, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03003008.

Battistini:2015:BRG

[3326] Andrea Battistini. Book review: Galileo. Scienziato e umanista, written by John L. Heilbron. Nuncius, 30(3):718–724, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03003009.

Lopez:2015:BRA

[3327] Susana Gómez López. Book review: Arte y ciencia en el barroco español, written by José Ramón Marcaida López. Nuncius, 30(3):725– 728, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03003010.

Baldassarri:2015:BRC

[3328] Fabrizio Baldassarri. Book review: Communicating Observations in Early Modern Letters (1500–1675). Epistolography and Epistemology in the Age of the Scientific Revolution, edited by Dirk van Miert. Nuncius, 30(3):729–733, ???? 2015. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03003011.

Tonetti:2015:BRG

[3329] Luca Tonetti. Book review: Giorgio Baglivi, Della tarantola. Lo studio di un medico nel Salento del XVII secolo, edited by Concetta Pennuto. Nuncius, 30(3):734–737, ???? 2015. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03003012.

Segala:2015:BRP

[3330] Marco Segala. Book review: Paris savant. Parcours et rencontres au temps des Lumières, written by Bruno Belhoste. Nuncius, 30(3):738– 740, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03003013.

Guerra:2015:BRE

[3331] Corinna Guerra. Book review: Émilie du Châtelet et Marie-Anne Lavoisier. Science et genre au XVIII siècle, written by Keiko Kawashima. Nuncius, 30(3):741–744, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03003014.

Schroter:2015:BRS

[3332] Harm Schröter. Book review: Solvay. A History of a Multinational Family Firm, written by Kenneth Bertrams, Nicolas Coupain, Ernst Homburg. Nuncius, 30(3):745–747, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03003015.

Mossman:2015:BRA

[3333] Susan Mossman. Book review: Aluminum Dreams. The Making of Light Modernity, written by Mimi Sheller. Nuncius, 30(3):748–750, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03003016.

Segala:2015:BRS

[3334] Marco Segala. Book review: Scientific Babel. How Science Was Done Before And After Global English, written by Michael Gordin. Nuncius, 30(3):751–753, ???? 2015. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03003017.

Vaccari:2015:PAD

[3335] Ezio Vaccari. Per un archivio della corrispondenza degli scienziati italiani. (Italian) [Toward an archive of correspondence of Italian scientists]. Nuncius, 30(3):??, ???? 2015. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/182539190x00697.

Feola:2016:I

[3336] Vittoria Feola. Introduction. Nuncius, 31(1):1–9, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03101007.

Suitner:2016:RRM

[3337] Riccarda Suitner. Radical reformation and medicine in the late Renaissance. Nuncius, 31(1):11–31, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03101003.

Neven:2016:RRA

[3338] Sylvie Neven. Recording and reading alchemy and art-technology in medieval and premodern German recipe collections. Nuncius, 31(1):32–49, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03101001.

Schilt:2016:IUH

[3339] Cornelis J. (Kees-Jan) Schilt. “To improve upon hints of things”. Nuncius, 31(1):50–77, ???? 2016. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03101004.

vandenHeuvel:2016:CCC

[3340] Charles van den Heuvel, Scott B. Weingart, Nils Spelt, and Henk Nellen. Circles of confidence in correspondence. Nuncius, 31(1):78–106, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03101002.

Feola:2016:PRV

[3341] Vittoria Feola. Paris, Rome, Venice, and Vienna in Peter Lambeck’s network. Nuncius, 31(1):107–128, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03101005.

Gulizia:2016:PIM

[3342] Stefano Gulizia. Printing and instrument making in the early modern Atlantic, 1520–1600. Nuncius, 31(1):129–162, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03101006.

Savoia:2016:BSS

[3343] Paolo Savoia. The book of the sick of Santa Maria della Morte in Bologna and the medical organization of a hospital in the Sixteenth-Century. Nuncius, 31(1):163–235, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03101008.

Egmond:2016:BRE

[3344] Florike Egmond. Book review: Echt Tierisch! Die Menagerie des Fürsten, written by Sabine Haag (ed.). Nuncius, 31(1):237–239, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03101009.

Egmond:2016:BRF

[3345] Florike Egmond. Book review: Flora: Federico Cesi’s Botanical Manuscripts. Vol. I — II — III. The Paper Museum of Cassiano dal Pozzo. Series B — Natural History, Part Seven, written by Brent Elliott, Luigi Guerrini and David Pegler. Nuncius, 31(1):240–242, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03101010.

Fiorentini:2016:BRO

[3346] Erna Fiorentini. Book review: Observing by hand: sketching the nebulae in the nineteenth century, written by Omar W. Nasim. Nuncius, 31 (1):243–245, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03101011.

Abbri:2016:BRL

[3347] Ferdinando Abbri. Book review: The Limits of Matter. Chemistry, Mining and Enlightenment, written by Hjalmar Fors. Nuncius, 31(1):246– 249, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03101012.

Chan:2016:BS

[3348] Eleanor Chan. Beautiful surfaces. Nuncius, 31(2):251–287, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03102001.

Lopez:2016:EET

[3349] Susana Gómez López. The encounter of the emblematic tradition with optics. Nuncius, 31(2):288–331, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03102002.

Hendriksen:2016:NS

[3350] Marieke M. A. Hendriksen. Necessary, not sufficient. Nuncius, 31(2): 332–360, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03102003.

Lehman:2016:WTN

[3351] Christine Lehman. What is the “true” nature of diamond? Nuncius, 31 (2):361–407, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03102004.

Campanella:2016:AS

[3352] Sara Campanella. The “abyssal society”. Nuncius, 31(2):408–438, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03102005.

Canadelli:2016:MBF

[3353] Elena Canadelli. Marble busts and fish fossils. Nuncius, 31(2):439–483, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03102006.

Smith:2016:OF

[3354] Pamela H. Smith. Other foreword. Nuncius, 31(3):485–488, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103007.

Skogh:2016:IVR

[3355] Lisa Skogh. Introduction: The varied role of the amateur. Nuncius, 31 (3):489–498, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03103006.

Taylor:2016:BA

[3356] Paul Taylor. The birth of the amateur. Nuncius, 31(3):499–522, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103001.

Keller:2016:ALS

[3357] Vera Keller. Art lovers and scientific virtuosi? Nuncius, 31(3):523–548, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103005.

Oosterhoff:2016:LP

[3358] Richard J. Oosterhoff. Lovers in paratexts. Nuncius, 31(3):549–583, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103002.

Valeriani:2016:LGF

[3359] Simona Valeriani. Lovers, gentlemen and farmers. Nuncius, 31(3):584– 609, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103003.

Brenni:2016:ITM

[3360] Paolo Brenni. Instruments, tools and materials. Nuncius, 31(3):611–646, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103004.

Tramelli:2016:BRD

[3361] Barbara Tramelli. Book review: Daughters of Alchemy: Women and Scientific Culture in Early Modern Italy, written by Meredith K. Ray. Nuncius, 31(3):647–650, ???? 2016. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03103008.

Beretta:2016:BRS

[3362] Marco Beretta. Book review: Sabliers d’Autrefois — Hourglasses of the Past — Clessidre d’Altri Tempi, edited by Anna Dominique, Éric Delalande. Nuncius, 31(3):651–652, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03103009.

Wragge-Morley:2016:BRO

[3363] Alexander Wragge-Morley. Book review: Observing the World through Images: Diagrams and Figures in the Early-Modern Arts and Science, edited by Nicholas Jardine and Isla Fay. Nuncius, 31(3):653–655, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103010.

Tonetti:2016:BRR

[3364] Luca Tonetti. Book review: Realdo Colombo, De re anatomica libri XV. Anatomia, edited by Gianluigi Baldo. Nuncius, 31(3):656–659, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103011.

Dubald:2016:BRM

[3365] Déborah S. Dubald. Book review: Le Musée de l’Homme. Histoire d’un musée-laboratoire, edited by Claude Blankaert. Nuncius, 31(3):660–662, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103012.

Ecca:2016:BRA

[3366] Giulia Ecca. Book review: The Aristotelian Problemata Physica. Philosophical and Scientific Investigations, edited by Robert Mayhew. Nuncius, 31(3):663–666, ???? 2016. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03103013.

deCeglia:2016:BRS

[3367] Francesco Paolo de Ceglia. Book review: Lo scudo contro il vaiolo. Antonio Miglietta e la profilassi nel Regno di Napoli (1801–1826), written by Caterina Tisci. Nuncius, 31(3):667–668, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03103014.

Abbri:2016:BRO

[3368] Ferdinando Abbri. Book review: Œuvres de Lavoisier. Correspondance. Avant-propos de Henri Kagan. Édition dirigée par Patrice Bret. Volume VII 1792–1794, edited by Patrice Bret. Nuncius, 31(3):669–672, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103015.

Canadelli:2016:BRW

[3369] Elena Canadelli. Book review: World’s Fairs on the Eve of War. Science Technology & Modernity 1937–1942, written by Robert H. Kargon, Karen Fiss, Morris Low, & Arthur P. Molella. Nuncius, 31(3):673–675, ???? 2016. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03103016.

Yassi:2017:AJW

[3370] Yousef Yassi. Al-Jazari’s water clocks. Nuncius, 32(1):1–24, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03201001.

Bondio:2017:FUF

[3371] Mariacarla Gadebusch Bondio. On the function, utility, and fragility of the nose. Nuncius, 32(1):25–51, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03201002.

Cafagna:2017:ITR

[3372] Fabio Cafagna. Images of transparency and resurrection from Leonardo da Vinci to Crisóstomo Martínez. Nuncius, 32(1):52–84, ???? 2017. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03201003.

Bruno-Chomin:2017:LCM

[3373] Giuseppe Bruno-Chomin. “… Che i matti dicano spropositi”. A discussion of cometary theory and superstition in Seventeenth Century Italy. Nuncius, 32(1):85–110, ???? 2017. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03201004.

Gomez-Ferrer:2017:VVG

[3374] Mercedes Gómez-Ferrer and Yolanda Gil. Vincenzo Vincenzi, “geometer and engineer to Cardinal Borja”. Nuncius, 32(1):111–145, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03201005.

Morabito:2017:DFE

[3375] Carmela Morabito. David Ferrier’s experimental localization of cerebral functions and the anti-vivisection debate. Nuncius, 32(1):146–165, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03201006.

Aahren:2017:FT

[3376] Eva Åhrén. Figuring things out. Nuncius, 32(1):166–211, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03201007.

Ottaviani:2017:REA

[3377] Alessandro Ottaviani. Race, ethnicity and Antisemitism in George Alexis Montandon between Kulturkreislehre, hologenesis and Mendelism. Nuncius, 32(1):212–252, ???? 2017. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03201008.

Rocca:2017:DHC

[3378] Mattia Della Rocca and Claudio Pogliano. Different histories from 20th Century neuroscience. Nuncius, 32(2):253–260, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03202001.

Abraham:2017:BMT

[3379] Tara H. Abraham. Between the mind twist and the brain spot. Nuncius, 32(2):261–285, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03202002.

Borck:2017:SSR

[3380] Cornelius Borck. Soups and sparks revisited. Nuncius, 32(2):286–329, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03202003.

Pogliano:2017:LTB

[3381] Claudio Pogliano. Lucky triune brain. Nuncius, 32(2):330–375, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03202004.

Rocca:2017:ANT

[3382] Mattia Della Rocca. Of the artistic nude and technological behaviorism. Nuncius, 32(2):376–411, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03202006.

Lysen:2017:IBI

[3383] Flora Lysen. It blinks, it thinks? Nuncius, 32(2):412–439, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03202007.

Casini:2017:BNR

[3384] Silvia Casini. Beyond the neuro-realism fallacy. Nuncius, 32(2):440–471, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03202005.

Morabito:2017:PMN

[3385] Carmela Morabito. Plastic maps: The new brain cartographies of the 21st-Century neurosciences. Nuncius, 32(2):472–500, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03202008.

Struhal:2017:II

[3386] Eva Struhal. Introduction introduction. Nuncius, 32(3):501–513, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03203001.

Ribouillault:2017:SCR

[3387] Denis Ribouillault. Seeing Christ, reading nature. Nuncius, 32(3):514– 555, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03203002.

Oy-Marra:2017:NCC

[3388] Elisabeth Oy-Marra. “Nel conoscere le cose tutte della natura”. Nuncius, 32(3):556–582, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03203003.

Fehrenbach:2017:CF

[3389] Frank Fehrenbach. Cut flowers. Nuncius, 32(3):583–614, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03203004.

vanGastel:2017:CF

[3390] Joris van Gastel. Campania felix? Nuncius, 32(3):615–639, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03203005.

Sapir:2017:SR

[3391] Itay Sapir. Shadowy realism. Nuncius, 32(3):640–657, ???? 2017. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03203006.

Gage:2017:AWS

[3392] Frances Gage. Air, wind, sun, spirit. Nuncius, 32(3):658–682, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03203007.

Struhal:2017:NPN

[3393] Eva Struhal. Natural painting and the new science in seventeenth-century Florence. Nuncius, 32(3):683–708, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03203008.

Favino:2017:OGB

[3394] Federica Favino. Other Gaspare Berti’s legacy. Nuncius, 32(3):709–773, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03202009.

Antonelli:2017:BRF

[3395] Francesca Antonelli. Book review: Femmes de sciences de l’Antiquité au XIXe siècle. Réalités et représentations, edited by Adeline Gargam. Nuncius, 32(3):775–777, ???? 2017. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03203010.

Salvia:2017:BRF

[3396] Stefano Salvia. Book review: La filosofia naturale di Giovanni Ciampoli, written by Federica Favino. Nuncius, 32(3):778–781, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03203011.

Cooper:2017:BRS

[3397] Irene Galandra Cooper. Book review: Il segreto di san Gennaro. Storia naturale di un miracolo napoletano, written by Francesco Paolo de Ceglia. Nuncius, 32(3):782–784, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03203012.

Saldarriaga:2017:BRF

[3398] Gragorio Saldarriaga. Book review: Food and Health in Early Modern Europe. Diet, Medicine and Society, 1450–1800, written by David Gentilcore. Nuncius, 32(3):785–787, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03203013.

Gauvin:2017:BRC

[3399] Jean-François Gauvin. Book review: Curious & Modern Inventions. Instrumental Music as Discovery in Galileo’s Italy, written by Rebecca Cypess. Nuncius, 32(3):788–790, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03203014.

Dibattista:2017:BRH

[3400] Liborio Dibattista. Book review: Haeckel’s Embryos: Images, Evolution and Fraud, written by Nick Hopwood. Nuncius, 32(3):791–793, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03203015.

Segala:2017:BRI

[3401] Marco Segala. Book review: The Invention of Nature: Alexander von Humboldt’s New World, written by Andrea Wulf. Nuncius, 32(3):794– 796, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03203016.

Luciano:2017:BRV

[3402] Erika Luciano. Book review: Vannevar Bush da ingegnere a tecnologo. La nascita della società dell’informazione, written by Benedetta Campanile. Nuncius, 32(3):797–799, ???? 2017. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03203017.

Dorin:2018:ST

[3403] Alan Dorin and Eva Anagnostou-Laoutides. The silver triton. Nuncius, 33(1):1–24, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03301001.

Golan:2018:PK

[3404] Nurit Golan. A portal to knowledge. Nuncius, 33(1):25–55, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03301002.

Giacomoni:2018:LE

[3405] Paola Giacomoni. The light of the emotions. Nuncius, 33(1):56–87, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03301003.

Staubermann:2018:IV

[3406] Klaus Staubermann. Investigating vision. Nuncius, 33(1):88–103, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03301004.

Leporiere:2018:MSE

[3407] Lorenzo Leporiere. Mediums and science in Early 20th Century Europe. Nuncius, 33(1):104–136, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03301005.

Ciocci:2018:BRT

[3408] Argante Ciocci. Book review: The Twelfth Chapter of Fibonacci’s Liber Abaci in its 1202 version, Bollettino di Storia delle Scienze Matematiche, by Enrico Giusti (ed.). Nuncius, 33(1):137–139, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03301006.

Egmond:2018:BRM

[3409] Florike Egmond. Book review: Medicine and Humanism in Late Medieval Italy. The Carrara Herbal in Padua, by Sarah R. Kyle. Nuncius, 33(1): 140–143, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03301007.

Savoia:2018:BRO

[3410] Paolo Savoia. Book review: The Optics of Giambattista Della Porta (ca. 1535–1615): A Reassessment, by Arianna Borrelli, Giora Hon, Yaakov Zik (eds.). Nuncius, 33(1):144–147, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03301008.

Fransen:2018:BRE

[3411] Sietske Fransen. Book review: Eye for Detail: Images of Plants and Animals in Art and Science, 1500–1630, by Florike Egmond. Nuncius, 33(1):148–150, January 2018. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03301009.

Heintzman:2018:BRP

[3412] Kit Heintzman. Book review: The Paper Museum of Cassiano dal Pozzo. Series B, Parts IV–V: Natural History: Birds, Other Animals and Natural Curiosities, by Henrietta McBurney, Ian Rolfe, Caterina Napoleone, and Paula Findlen (eds.), with Carlo Violani, Onno Wijnands, Arturo Morales-Muñiz, Eufrasia Roselló-Izquierdo, Arthur McGregor, and Kathie Way. Nuncius, 33(1):151–154, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03301010.

Traetta:2018:BRI

[3413] Luigi Traetta. Book review: Italian Psychology and Jewish emigration under Fascism. From Florence to Jerusalem and New York, by Patrizia Guarnieri. Nuncius, 33(1):155–157, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03301011.

Campanile:2018:BRS

[3414] Benedetta Campanile. Book review: Science without Frontiers. Cosmopolitanism and National Interests in the World of Learning, 1870– 1940, by Robert Fox. Nuncius, 33(1):158–160, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03301012.

Droscher:2018:BRB

[3415] Ariane Dröscher. Book review: The Beautiful Brain: The Drawings of Santiago Ramón y Cajal, by Eric A. Newman, Alfonso Araque, and Janet M. Dubinsky (eds.). Nuncius, 33(1):161–162, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03301013.

Spataro:2018:BRM

[3416] Stefano Spataro. Book review: Il mistero degli antichi astronauti, by Marco Ciardi. Nuncius, 33(1):163–165, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03301014.

Petrocelli:2018:BRT

[3417] Carla Petrocelli. Book review: The Turing Guide, by Jack Copeland, Jonathan Bowen, Mark Sprevak, and Robin Wilson. Nuncius, 33(1): 166–168, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03301015.

Dunn:2018:BRN

[3418] Richard Dunn. Book review: Navigation: A Very Short Introduction, by Jim Bennett. Nuncius, 33(1):169–170, January 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http:/

/booksandjournals.brillonline.com/content/journals/10.1163/
18253911-03301016.

Vermij:2018:FDE

[3419] Rienk Vermij. On the function of dedications in early modern scientific books. Nuncius, 33(2):171–197, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/content/journals/10.1163/18253911-03302001.

Hagendijk:2018:LCB

[3420] Thijs Hagendijk. Learning a craft from books. Nuncius, 33(2):198–235, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03302002.

Chapron:2018:TS

[3421] Emmanuelle Chapron. Tools for science. Nuncius, 33(2):236–264, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03302003.

Zuidervaart:2018:EDA

[3422] Huib Zuidervaart and Tiemen Cocquyt. The early development of the achromatic telescope revisited. Nuncius, 33(2):265–344, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03302004.

Durey:2018:TPC

[3423] Guillaume Durey, Denis Beaudouin, André Pierre Legrand, and Brigitte Leridon. Le télescope pyrhéliométrique de Charles Féry et la mesure de la température du Soleil. (French) [The pyrheliometric telescope of Charles Fery and the measurement of the temperature of the Sun]. Nuncius, 33 (2):345–371, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03302005.

Pogliano:2018:CO

[3424] Claudio Pogliano. “Conscience d’outil”. Nuncius, 33(2):373–378, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03302010.

Antonelli:2018:BRS

[3425] Francesca Antonelli. Book review: Sentimental Savants: Philosophical Families in Enlightenment France, by Meghan K. Roberts. Nuncius, 33 (2):379–381, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
content/journals/10.1163/18253911-03302006.

Marinucci:2018:BRP

[3426] Ludovica Marinucci. Book reviews: Paolo Savoia, Cosmesi e chirurgia. Bellezza, dolore e medicina nell’Italia moderna. Milano: Editrice Bibliografica, 2017, 232 pp., 19.50EUR, ISBN: 978-88-7075-982-2. & Annarita Angelini, Matematica e immaginazione nel Rinascimento. Milano: Editrice Bibliografica, 2017, 198 pp., 18.50EUR, ISBN: 978-88-7075-951-8. & Carlo Bovolo, I cattolici italiani e la scienza. Il discorso apologetico sulla stampa clericale nell’età del positivismo. Milano: Editrice Bibliografica, 2017, 144 pp., 16.50EUR, ISBN: 978-88-7075-986-0. & Francesco Barreca, La scienza che fu. Idee e strumenti di teorie abbandonate. Milano: Editrice Biblio-grafica, 2017, 128 pp., 16.50EUR, ISBN: 978-88-7075-952-5. & Claudio Pogliano, Storie di cervelli. Dall’antichità al Novecento. Milano: Editrice Bibliografica, 2017, 208 pp., 19.50EUR, ISBN: 978-88-7075-981-5. Nuncius, 33(2):382– 385, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03302007.

Leporiere:2018:BRL

[3427] Lorenzo Leporiere. Book review: Los límites de la ciencia. Espiritismo, hipnotismo y el estudio de los fenómenos paranormales (1850– 1930), by Annette Müllberger (ed.). Nuncius, 33(2):386–388, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03302008.

Yale:2018:BRS

[3428] Elizabeth Yale. Book review: Science in the Archives: Pasts, Presents, Futures, by Lorraine Daston (ed.). Nuncius, 33(2):389–392, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03302009.

Segala:2018:BRA

[3429] Marco Segala. Book review: Alfred Wegener. Science, Exploration, and the Theory of Continental Drift, by Mott T. Greene. Nuncius, 33(2):393– 396, June 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/content/
journals/10.1163/18253911-03302011.

Ciancio:2018:VRE

[3430] Luca Ciancio and Domenico Laurenza. Visual representation in Earth sciences history after “The Emergence”: Introduction. Nuncius, 33(3): 397–414, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
33/3/article-p397_1.xml.

DalPrete:2018:REL

[3431] Ivano Dal Prete. The ruins of the Earth: Learned meteorology and artisan expertise in Fifteenth-Century Italian landscapes. Nuncius, 33 (3):415–441, November 2018. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/abstract/journals/
nun/33/3/article-p415_2.xml.

Laurenza:2018:ITS

[3432] Domenico Laurenza. Images and theories: The study of fossils in Leonardo, Scilla and Hooke. Nuncius, 33(3):442–463, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/33/
3/article-p442_3.xml.

Ciancio:2018:IFA

[3433] Luca Ciancio. “Immoderatus fervor ad intra coërcendus”: Reactions to Athanasius Kircher’s central fire in Jesuit science and imagination. Nuncius, 33(3):464–504, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/abstract/

journals/nun/33/3/article-p464_4.xml.

Roos:2018:BRL

[3434] Anna Marie Roos and Edwin D. Rose. Book review: Lives and Afterlives of the Lithophylacii Britannici ichnographia (1699), the First Illustrated Field Guide to English Fossils. Nuncius, 33(3):505–536, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/33/
3/article-p505_5.xml.

Dekker:2018:BRA

[3435] Elly Dekker. Book review: The Astronomical Framework Underlying the Sala di Galatea. Nuncius, 33(3):537–561, November 2018. CODEN ????

ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/abstract/journals/nun/33/3/article-p537_6.xml.

Bell:2018:TMA

[3436] Janis Bell. A treatise on mirrors attributed to Matteo Zaccolini. Nuncius, 33(3):563–584, November 2018. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/abstract/journals/
nun/33/3/article-p563_7.xml.

Antonelli:2018:LMA

[3437] Francesca Antonelli. Lavoisier and music: An analysis of Notes sur la Musique (1778) within its cultural context. Nuncius, 33(3):585–630, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/33/
3/article-p585_8.xml.

Haskell:2018:BRS

[3438] Yasmin Haskell. Book review: Science Writing in Greco-Roman Antiquity, by Liba Taub. Nuncius, 33(3):631–633, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/abstract/journals/nun/33/3/article-p631_9.xml.

Maddaluno:2018:BRR

[3439] Lavinia Maddaluno. Book review: Ruins Past, Modernity in Italy (1744– 1836), by Sabrina Ferri. Nuncius, 33(3):634–636, November 2018. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/abstract/journals/nun/33/3/article-p634_10.xml.

Daniele:2018:BRA

[3440] Stefano Daniele. Book review: Alchemy, Medicine, and Commercial Book Production. A Codicological and Linguistic Study of the Voigts– Sloane Manuscript Group, by Alpo Honkapohja. Nuncius, 33(3):637– 640, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
33/3/article-p637_11.xml.

Ciocci:2018:BRM

[3441] Argante Ciocci. Book review: Mathematical Practitioners and the Transformation of Natural Knowledge in Early Modern Europe, by Lesley B. Cormack, Steven A. Walton, John A. Schuster, ed. Nuncius, 33(3):641– 644, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
33/3/article-p641_12.xml.

Bouley:2018:BRP

[3442] Bradford Bouley. Book review: Professors, Physicians, and Practices in the History of Medicine: Essays in Honor of Nancy Siraisi, by Gideon Manning and Cynthia Klestinec, ed. Nuncius, 33(3):645–647, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/33/
3/article-p645_13.xml.

Pepe:2018:BRL

[3443] Chiara Pepe. Book review: Logica e magica. Giovan Battista Della Porta e i segreti della natura, by Donato Verardi. Nuncius, 33(3):648– 650, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
33/3/article-p648_14.xml.

Ryan:2018:BRE

[3444] William J. Ryan. Book review: The Experiential Caribbean: Creating Knowledge and Healing in the Early Modern Atlantic, by Pablo F. Gómez. Nuncius, 33(3):651–653, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

abstract/journals/nun/33/3/article-p651_15.xml.

Tessicini:2018:BRA

[3445] Dario Tessicini. Book review: After the Scientific Revolution: Thinking Globally about the Histories of the Modern Sciences, by J. B. Shank, ed. Nuncius, 33(3):654–656, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/abstract/

journals/nun/33/3/article-p654_16.xml.

Prkachin:2018:BRB

[3446] Yvan Prkachin. Book review: Being Brains: Making the Cerebral Subject, by Fernando Vidal and Francisco Ortega. Nuncius, 33(3):657–660, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/33/
3/article-p657_17.xml.

Tocchetti:2018:BRS

[3447] Sarah Tocchetti. Book review: Synthetic: How Life Got Made, by Sophia Roosth. Nuncius, 33(3):661–663, November 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

abstract/journals/nun/33/3/article-p661_18.xml.

Pouderon:2019:PVC

[3448] Bernard Pouderon. La procréation des vierges chez les gnostiques et leurs contemporains orthodoxes (iie s.): De l’empreinte du regard à la parthénogénèse. (French) [The procreation of virgins among the Gnostics and their Orthodox contemporaries (2nd century): From the impression of the gaze to parthenogenesis]. Nuncius, 34(1):1–26, February 2019. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/abstract/journals/nun/34/1/article-p1_1.xml.

Davis:2019:MEA

[3449] John Davis. A Medieval English astrolabe now in Innsbruck, linked to the Lancastrian Court and with a Chaucer connection. Nuncius, 34(1): 27–68, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
34/1/article-p27_2.xml.

Ampollini:2019:CGP

[3450] Ilaria Ampollini. Celestial globes and popular astronomy: Giuseppe Toaldo translator of Jérôme Lalande. Nuncius, 34(1):69–98, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/34/
1/article-p69_3.xml.

Valdes:2019:LIW

[3451] Catalina Valdés and Magdalena Montalbán. “… It was highly desirable they should be illustrated”: Images from the U.S. Navy Astronomical Expedition in Chile (1849–1852). Nuncius, 34(1):99–127, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/34/
1/article-p99_4.xml.

Valentines-Alvarez:2019:FTM

[3452] Jaume Valentines-Álvarez and Jaume Sastre-Juan. The failed Technology Museum of Catalonia: Engineers and the politics of the musealization of technology in Barcelona (1929–1939). Nuncius, 34(1):128– 154, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/34/
1/article-p128_5.xml.

Raggetti:2019:BRA

[3453] Lucia Raggetti. Book review: Apollonius of Tyana’s Great Book of Talismans. Nuncius, 34(1):155–182, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

abstract/journals/nun/34/1/article-p155_6.xml.

Andrews:2019:BRD

[3454] Noam Andrews. Book review: Diagnosis: Melencolia I. Nuncius, 34(1): 183–187, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
34/1/article-p183_7.xml.

Klestinec:2019:BRP

[3455] Cynthia Klestinec. Book review: Pious Postmortems: Anatomy, Sanctity, and the Catholic Church in Early Modern Europe, by Bradford A. Bouley. Nuncius, 34(1):189–191, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

abstract/journals/nun/34/1/article-p189_8.xml.

Nenci:2019:BRJ

[3456] Elio Nenci. Book review: Janello Torriani and the Spanish Empire. A Vitruvian Artisan at the Dawn of the Scientific Revolution, by Cristiano Zanetti. Nuncius, 34(1):192–194, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

abstract/journals/nun/34/1/article-p192_9.xml.

Ciocci:2019:BRC

[3457] Argante Ciocci. Book review: Cieli in contraddizione. Giovanni Battista Riccioli e il terzo sistema del mondo, by Flavia Marcacci. Nuncius, 34(1): 195–197, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
34/1/article-p195_10.xml.

Ladee:2019:BRI

[3458] Charley Ladee. Book review: Interactive and Sculptural Printmaking in the Renaissance, by Suzanne Karr Schmidt. Nuncius, 34(1):198– 200, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/34/
1/article-p198_11.xml.

Werrett:2019:BRA

[3459] Simon Werrett. Book review: Artisanal Enlightenment: Science and the Mechanical Arts in Old Regime France, by Paola Bertucci. Nuncius, 34(1):201–202, February 2019. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/abstract/journals/
nun/34/1/article-p201_12.xml.

Savoia:2019:BRC

[3460] Paolo Savoia. Book review: Creatures Born of Mud and Slime: The Wonder and Complexity of Spontaneous Generation, by Daryn Lehoux. Nuncius, 34(1):203–206, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/abstract/

journals/nun/34/1/article-p203_13.xml.

DeCeglie:2019:BRM

[3461] Rossella De Ceglie. Book review: Le monde dans un carnet. Alexander von Humboldt en Italie (1805), by Marie-Noëlle Bourguet. Nuncius, 34 (1):207–210, February 2019. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/abstract/journals/
nun/34/1/article-p207_14.xml.

Dibattista:2019:BRV

[3462] Liborio Dibattista. Book review: Vivarium. Experimental, Quantitative, and Theoretical Biology at Vienna’s Biologische Versuchsanstalt, by Gerd B. Müller, ed. Nuncius, 34(1):211–213, February 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/abstract/journals/nun/34/1/article-p211_15.xml.

Beretta:2019:PAS

[3463] Marco Beretta and Ulf Larsson. Prizes and awards in science before Nobel: Introduction. Nuncius, 34(2):215–217, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/abstract/journals/nun/34/2/article-p215_1.xml.

Beretta:2019:NRA

[3464] Marco Beretta. Names as rewards: The ambiguous role of eponyms in the history of science. Nuncius, 34(2):219–235, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/abstract/journals/nun/34/2/article-p219_2.xml.

Cuomo:2019:RSA

[3465] Serafina Cuomo. Rewarding science in Ancient Greece and Rome. Nuncius, 34(2):236–245, June 2019. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/abstract/journals/
nun/34/2/article-p236_3.xml.

Taub:2019:CSA

[3466] Liba Taub. Celebrating science in ancient Greece and Rome. Nuncius, 34(2):246–257, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
34/2/article-p246_4.xml.

Azzolini:2019:TWN

[3467] Monica Azzolini. There were no medals to be won: Scientific duels in the Italian Renaissance. Nuncius, 34(2):258–283, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/abstract/journals/nun/34/2/article-p258_5.xml.

Higgitt:2019:SSB

[3468] Rebekah Higgitt. “In the Society’s strong box”: A visual and material history of the Royal Society’s Copley Medal, c. 1736–1760. Nuncius, 34 (2):284–316, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
34/2/article-p284_6.xml.

Bret:2019:FDR

[3469] Patrice Bret. Funding and directing research or rewarding scientific achievements?: Two centuries of prizes at the Academy of Sciences in Paris. Nuncius, 34(2):317–355, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/abstract/

journals/nun/34/2/article-p317_7.xml.

Widmalm:2019:NSP

[3470] Sven Widmalm. Normal science: Prizes in the Swedish Academy of Sciences before Nobel. Nuncius, 34(2):356–376, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/abstract/journals/nun/34/2/article-p356_8.xml.

Beckman:2019:SOG

[3471] Jenny Beckman. “Sections of one great institution”: Prizes, publications, and transnational communication at the Royal Swedish Academy of Sciences in the Nineteenth Century. Nuncius, 34(2):377– 391, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/34/
2/article-p377_9.xml.

Brenni:2019:PMH

[3472] Paolo Brenni. Prizes, medals and honourable mentions: How instrument makers were rewarded at 19th-Century exhibitions. Nuncius, 34 (2):392–420, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
34/2/article-p392_10.xml.

Lindqvist:2019:SR

[3473] Svante Lindqvist. Salon des refusés. Nuncius, 34(2):421–426, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL https://brill.com/abstract/journals/nun/34/2/
article-p421_11.xml.

Savoia:2019:CSR

[3474] Paolo Savoia. Cheesemaking in the scientific revolution: A Seventeenth-Century Royal Society report on dairy products and the history of European knowledge. Nuncius, 34(2):427–455, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/abstract/journals/nun/34/2/article-p427_12.xml.

Beretta:2019:BRL

[3475] Marco Beretta. Book review: For the Love of Science: The Correspondence of J. H. de Magellan (1722–1790), by Roderick W. Home, Isabel M. Malaquias, Manuel F. Thomaz (ed.). Nuncius, 34(2):457– 459, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL https://brill.com/abstract/journals/nun/34/
2/article-p457_13.xml.

Bucciarelli:2019:BRI

[3476] Lucia Bucciarelli. Book review: Isaac Newton and Natural Philosophy, by Niccolò Guicciardini. Nuncius, 34(2):460–462, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/abstract/journals/nun/34/2/article-p460_14.xml.

Akopyan:2019:BRS

[3477] Ovanes Akopyan. Book review: Sing Aloud Harmonious Spheres: Renaissance Conceptions of Cosmic Harmony, by Jacomien Prins and Maude Vanhaelen (eds.). Nuncius, 34(2):463–464, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/abstract/journals/nun/34/2/article-p463_15.xml.

Tan:2019:BRM

[3478] Wei Yu Wayne Tan. Book review: Making Time: Astronomical Time Measurement in Tokugawa Japan, by Yulia Frumer. Nuncius, 34(2): 465–467, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
34/2/article-p465_16.xml.

Zanetti:2019:BRE

[3479] Cristiano Zanetti. Book review: Engineering the Eternal City: Infrastructure, Topography, and the Culture of Knowledge in Late Sixteenth-Century Rome, by Pamela O. Long. Nuncius, 34(2):468–471, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL https://brill.com/abstract/journals/nun/34/2/
article-p468_17.xml.

Lester:2019:BRD

[3480] Gustave Lester. Book review: Darwin’s Evolving Identity: Adventure, Ambition, and the Sin of Speculation, by Alistair Sponsel. Nuncius, 34 (2):472–474, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/abstract/journals/nun/
34/2/article-p472_18.xml.

Lusito:2019:BRJ

[3481] Fabio Lusito. Book review: John von Neumann, Matematica hilbertiana e fondamenti della meccanica quantistica, by Rossella Lupacchini (ed.). Nuncius, 34(2):475–477, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/abstract/

journals/nun/34/2/article-p475_19.xml.

Basso:2019:BRH

[3482] Elisabetta Basso. Book review: The Human Body in the Age of Catastrophe. Brittleness, Integration, Science, and the Great War, by Stefanos Geroulanos, Todd Meyers. Nuncius, 34(2):478–480, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/abstract/journals/nun/34/2/article-p478_20.xml.

Campanile:2019:BRL

[3483] Benedetta Campanile. Book review: L’élite sous la mitraille. Les normaliens, les mathématiques et la Grande Guerre, 1900-1925, by David Aubin. Nuncius, 34(2):481–483, June 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

abstract/journals/nun/34/2/article-p481_21.xml.

Bettineschi:2019:DTA

[3484] Cinzia Bettineschi, Giulia Deotto, Donald James Ian Begg, Gianmario Molin, Paola Zanovello, Alessia Fassone, Christian Greco, and Matilde Borla, Ivana Angelini. A dig through archives and depots: Rediscovering the inlay workshop of Tebtynis and its materials. Nuncius, 34(3):485–516, December 2019. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
34/3/article-p485_1.xml.

Nothaft:2019:BRO

[3485] C. Philipp E. Nothaft. Book review: An Overlooked Construction Manual for the Quadrans Vetustissimus. Nuncius, 34(3):517–534, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/34/3/article-p517_2.xml.

Fix:2019:BRE

[3486] Adam Fix. Book review: Esperienza, Teacher of All Things: Vincenzo Galilei’s Music as Artisanal Epistemology. Nuncius, 34(3):535– 574, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/34/
3/article-p535_3.xml.

Fagnani:2019:ASN

[3487] Martino Lorenzo Fagnani. Agricultural science in Napoleonic universities: Didactic and research in Pavia, Bologna and Padua. Nuncius, 34(3):575–601, December 2019. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
34/3/article-p575_4.xml.

Pogliano:2019:UVR

[3488] Claudio Pogliano. Unconventional views of racial brains in the 19th Century. Nuncius, 34(3):602–634, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/34/3/article-p602_5.xml.

SanchezdeLollanoPrieto:2019:CAC

[3489] Joaquín Sánchez de Lollano Prieto and Alicia Sánchez Ortiz. Contributions to animal ceroplastics: The sculptor Cristóbal Garrigó de Nis (1800–1863) and the anatomical cabinet of the Royal Veterinary School in Madrid. Nuncius, 34(3):635–660, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/34/3/article-p635_6.xml.

Maddaluno:2019:FUL

[3490] Lavinia Maddaluno. Four unpublished letters from Nicolas Fatio de Duillier to Isaac Newton: Networks and alchemical knowledge. Nuncius, 34 (3):661–702, December 2019. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
34/3/article-p661_7.xml.

Ecca:2019:BRG

[3491] Giulia Ecca. Book review: Galien de Pergame ou la rhétorique de la Providence. Médecine, littérature et pouvoir à Rome, by Caroline Petit. Nuncius, 34(3):703–705, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/34/3/article-p703_8.xml.

Morar:2019:BRK

[3492] Florin-Stefan Morar. Book review: Knowledge of Nature and the Nature of Knowledge in Early Modern Japan, by Federico Marcon. Nuncius, 34(3):706–708, December 2019. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
34/3/article-p706_9.xml.

Savoia:2019:BRR

[3493] Paolo Savoia. Book review: Recipes and Everyday Knowledge: Medicine, Science, and the Household in Early Modern England, by Elaine Leong. Nuncius, 34(3):709–712, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/34/3/article-p709_10.xml.

Hanss:2019:BRE

[3494] Stefan Hanß. Book review: Exotic Animals in the Art and Culture of the Medici Court in Florence, by Angelica Groom. Nuncius, 34(3):713– 716, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/34/
3/article-p713_11.xml.

Chen:2019:BRI

[3495] Jessie Wei-Hsuan Chen. Book review: Insect Artifice: Nature and Art in the Dutch Revolt, by Marisa Anne Bass. Nuncius, 34(3):717– 719, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/34/
3/article-p717_12.xml.

Loiodice:2019:BRC

[3496] Eleonora Loiodice. Book review: Cronaca di una cometa non annunciata. Astronomia e comunicazione della scienza nel XVIII secolo, by Ilaria Ampollini. Nuncius, 34(3):720–721, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/34/3/article-p720_13.xml.

Gambaro:2019:BRD

[3497] Ivana Gambaro. Book review: I Diari Berlinesi (1857–1859) di Giovanni Virginio Schiaparelli, by Clara Silvia Roero, Pasquale Tucci, ed. Nuncius, 34(3):722–724, December 2019. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
34/3/article-p722_14.xml.

Lusito:2019:BRL

[3498] Fabio Lusito. Book review: L’inferno sulla vetta, by Paolo Mazzarello. Nuncius, 34(3):725–728, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/34/3/article-p725_15.xml.

Baca:2019:BRL

[3499] Katie Baca. Book review: A Lab of One’s Own: Science and Suffrage in the First World War, by Patricia Fara. Nuncius, 34(3):729– 731, December 2019. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/34/
3/article-p729_16.xml.

Cesteros:2020:NCA

[3500] Miriam Blanco Cesteros. ‘OδντΓ ιζω: New clues on ancient book production from alchemical papyri. Nuncius, 35(1):1–19, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/35/1/article-p1_1.xml.

Chen:2020:WCT

[3501] Jessie Wei-Hsuan Chen. A woodblock’s career: Transferring visual botanical knowledge in the Early Modern Low Countries. Nuncius, 35 (1):20–63, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
1/article-p20_2.xml.

Buckley:2020:PEC

[3502] Cali Buckley. Pathos, Eros, and curiosity: The history and reception of ivory anatomical models from the Seventeenth Century to today. Nuncius, 35(1):64–89, April 2020. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
35/1/article-p64_3.xml.

Luciano:2020:TPR

[3503] Erika Luciano and Chiara Pizzarelli. “the theory of practice”: The role of collections in technical and scientific teaching in Turin (1845–1861). Nuncius, 35(1):90–113, April 2020. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
35/1/article-p90_4.xml.

Xavier:2020:MPM

[3504] Manuel Xavier. Meanings of practice, morals of error: The case of Gago Coutinho’s Salmoiraghi Theodolites. Nuncius, 35(1):114–140, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/35/1/article-p114_5.xml.

Sirgiovanni:2020:MDM

[3505] Elisabetta Sirgiovanni and Alessandro Aruta. From the madhouse to the docu-museum: The enigma surrounding the Cerletti-Bini ECT apparatus prototype. Nuncius, 35(1):141–164, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/35/1/article-p141_6.xml.

Schober:2020:BRN

[3506] Sarah-Maria Schober. Book review: A New Order of Medicine: The Rise of Physicians in Reformation Nuremberg, by Hannah Murphy. Nuncius, 35(1):165–168, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
1/article-p165_7.xml.

Marcus:2020:BRR

[3507] Hannah Marcus. Book review: Reading Galileo: Scribal Technologies and the Two New Sciences, by Renée Raphael. Nuncius, 35(1):169–171, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/35/1/article-p169_8.xml.

Abbri:2020:BRD

[3508] Ferdinando Abbri. Book review: Dutch Cartesianism and the Birth of Philosophy of Science. From Regius to ’s Gravesande 1640–1750, by Andrea Strazzoni. Nuncius, 35(1):172–174, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/35/1/article-p172_9.xml.

Andrews:2020:BRP

[3509] Noam Andrews. Book review: The Philosophy Chamber: Art and Science in Harvard’s Teaching Cabinet, 1766–1820, by Ethan W. Lasser (ed.). Nuncius, 35(1):175–177, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/1/article-p175_10.xml.

Brixius:2020:BRW

[3510] Dorit Brixius. Book review: Worlds of Natural History, by Helen A. Curry, Nicholas Jardine, James A. Secord and Emma Spary (eds.). Nuncius, 35(1):178–180, April 2020. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
35/1/article-p178_11.xml.

Fontani:2020:BRP

[3511] Marco Fontani. Book review: The Politics of Chemistry. Science and Power in Twentieth-Century Spain, by Agustí Nieto-Galan. Nuncius, 35 (1):181–182, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
1/article-p181_12.xml.

Candela:2020:BRG

[3512] Andrea Candela. Book review: Greening the Alliance. The Diplomacy of NATO’s Science and Environmental Initiatives, by Simone Turchetti. Nuncius, 35(1):183–185, April 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/35/1/article-p183_13.xml.

Raposo:2020:OWW

[3513] Pedro M. P. Raposo. Other worlds within a hand’s reach: Planetary models from the Sixteenth Century through the Early Twentieth Century. Nuncius, 35(S1):187–190, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/2/article-p187_1.xml.

Gessner:2020:AS

[3514] Samuel Gessner, Michael Korey, and Karsten Gaulke. The anomalous Sun. Nuncius, 35(S1):191–234, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/2/article-p191_2.xml.

McEvoy:2020:GGO

[3515] Rory McEvoy. George Graham and the orrery. Nuncius, 35(S1): 235–250, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
2/article-p235_3.xml.

Tirapicos:2020:FBO

[3516] Luís Tirapicos and Thomas Horst. Francesco Bianchini (1662–1729) and the origins of planetary globes. Nuncius, 35(S1):251–273, ???? 2020.

CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/35/2/article-p251_4.xml.

Raposo:2020:RO

[3517] Pedro M. P. Raposo. Recounting the orbs. Nuncius, 35(S1):274–299, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/35/2/article-p274_5.xml.

Devoy:2020:LCM

[3518] Louise E. Devoy. Lunar crater models. Nuncius, 35(S1):300–332, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/35/2/article-p300_6.xml.

Swan:2020:EMG

[3519] Claudia Swan. Early modern geometries. Nuncius, 35(S1):333–340, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/35/2/article-p333_7.xml.

Zorach:2020:SSI

[3520] Rebecca Zorach. Stones, snowflakes, and insect eggs. Nuncius, 35(S1): 341–363, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
2/article-p341_8.xml.

Schmidt:2020:PF

[3521] Suzanne Karr Schmidt. Pfinzing and friends. Nuncius, 35(S1):364– 386, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
2/article-p364_9.xml.

Chen-Morris:2020:GMU

[3522] Raz Chen-Morris. Geometry and the making of utopian knowledge in Early Modern Europe. Nuncius, 35(S1):387–412, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/35/2/article-p387_10.xml.

Lipking:2020:JKT

[3523] Lawrence Lipking. Johannes Kepler and Twenty-First-Century science. Nuncius, 35(S1):413–428, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/2/article-p413_11.xml.

Clericuzio:2020:DSQ

[3524] Antonio Clericuzio. Descartes, Stensen and the quest for visible mechanisms. Nuncius, 35(S1):429–440, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/35/2/article-p429_12.xml.

Pierucci:2020:BRH

[3525] Federico Pierucci. Book review: Hobbes and Galileo: Method, Matter and the Science of Motion, by Gregorio Baldin. Cham: Springer, 2020. XXIII, 216 pp., 31 b/w ill. 79,99 EUR. ISBN: 978-3-030-41413-9. Nuncius, 35 (S1):441–443, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
2/article-p441_13.xml.

Borrelli:2020:BRP

[3526] Arianna Borrelli. Book review: Perspective as Practice. Renaissance Cultures of Optics, by Sven Dupré (ed.). Turnhout, Belgium: Brepols, 2019. 512 p., 143 b/w ill. + 46 colour ill. 90 EUR. ISBN: 978-2-503-58107-1. Nuncius, 35(S1):444–447, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/35/2/article-p444_14.xml.

Tomassetti:2020:BRM

[3527] Stefano Tomassetti. Book review: Materia medica. Savoirs et usages des médicaments aux époques médiévales et modernes, by Philip Rieder, François Zanetti (eds.). Geneva: Librairie Droz, 2018.312 pp. $45.60. ISBN: 978-2-600-05878-0. Nuncius, 35(S1):448–450, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/35/2/article-p448_15.xml.

Savoia:2020:BRF

[3528] Paolo Savoia. Book review: Food, Social Politics and the Order of Nature in Renaissance Italy, by Allen J. Grieco. Milan: Officina Libraria (for Europe); Cambridge: Harvard University Press (rest of the world), 2019. 327 pp., 35 color plates. 28 EUR. ISBN 978-88-336-7039-3. Nuncius, 35 (S1):451–454, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
2/article-p451_16.xml.

Maddaluno:2020:BRN

[3529] Lavinia Maddaluno. Book review: Newton the Alchemist: Science, Enigma, and the Quest for Nature’s “Secret Fire”, by William R. Newman. Princeton: Princeton University Press, 2019. 560 pp. $39.95/ £34.00. ISBN: 978-0-691-17487-7. Nuncius, 35(S1):455–457, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
2/article-p455_17.xml.

Egmond:2020:BRO

[3530] Florike Egmond. Book review: Ökonomien botanischen Wissens. Praktiken der Gelehrsamkeit in Basel um 1600, by Davina Benkert. [Basler Beiträge zur Geschichtswissenschaft 188], Basel/Berlin: Schwabe Verlag, 2020. 352 pp., 21 illustrations of which 12 in colour, 58 EUR. ISBN (printed): 978-3-7965-4073-8; ISBN eBook (PDF): 978-3-7965-4155-1; DOI 10.24894/978-3-7965-4155-1. Nuncius, 35(S1):458– 460, ???? 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
2/article-p458_18.xml.

Dacome:2020:UOM

[3531] Lucia Dacome, Meghan C. Doherty, Dahlia Porter, and Courtney Roby. Unruly objects — material entanglements in the arts and sciences. Nuncius, 35(3):461–470, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/3/article-p461_1.xml.

Rider:2020:EUO

[3532] Robin E. Rider. Equations as unruly objects. Nuncius, 35(3):471– 505, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
3/article-p471_2.xml.

Netz:2020:WWG

[3533] Reviel Netz. Why were Greek mathematical diagrams schematic? Nuncius, 35(3):506–535, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/3/article-p506_3.xml.

Anderson:2020:PBB

[3534] Emily R. Anderson. Printing the bespoke book. Nuncius, 35(3):536– 560, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
3/article-p536_4.xml.

Basham:2020:RBS

[3535] Sarah Basham. The Reader’s body in seventeenth-century Chinese statecraft texts. Nuncius, 35(3):561–605, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/35/3/article-p561_5.xml.

Virdi:2020:MTD

[3536] Jaipreet Virdi. Material traces of disability. Nuncius, 35(3):606– 631, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
3/article-p606_6.xml.

Neswald:2020:TDT

[3537] Elizabeth Neswald. Things that don’t talk much and things that feel. Nuncius, 35(3):632–659, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/3/article-p632_7.xml.

Scripps:2020:DD

[3538] Sarah Michel Scripps. The downfall of the diorama. Nuncius, 35(3): 660–684, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
3/article-p660_8.xml.

Learner:2020:ENS

[3539] Mary Learner. Embroidering the new science. Nuncius, 35(3):685– 717, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
3/article-p685_9.xml.

Pantin:2020:MDS

[3540] Isabelle Pantin. Le monument du savoir galiléen dans le mouvement continu de la recherche collaborative. (French) [The Galilean monument of knowledge in the continuous movement of collaborative research]. Nuncius, 35(3):719–736, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/3/article-p719_10.xml.

Bernardoni:2020:BRM

[3541] Andrea Bernardoni. Book review: Le macchine nel mondo antico. Dalle civiltà mesopotamiche a Roma imperiale, by Giovanni Di Pasquale. Roma:Carocci, 2019. 241pp. 18 EUR. ISBN: 978-88-430-9589-6. Nuncius, 35(3):737–739, December 2020. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
35/3/article-p737_11.xml.

Mackenzie:2020:BRA

[3542] Pamela Mackenzie. Book review: Aesthetic Science: Representing Nature in the Royal Society of London, 1650–1720, by Alexander Wragge-Morley. Chicago: University of Chicago Press, 2020. 243pp. $40. ISBN: 978-0-226-68086-6. Nuncius, 35(3):740–742, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/35/3/article-p740_12.xml.

Azzolini:2020:BRA

[3543] Monica Azzolini. Book review: After the Flood. Imagining the Global Environment in Early Modern Europe, by Lydia Barnett. Baltimore: Johns Hopkins University Press, 2019. 264 pp. $49,95. ISBN: 978-1-4214-2951-9. Nuncius, 35(3):743–745, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/3/article-p743_13.xml.

Cerruti:2020:BRS

[3544] Luigi Cerruti and Elena Ghibaudi. Book review: La scienza Impossibile. Percorsi dell’alchimia in Francia tra Ottocento e Novecento, by Leonardo Anatrini, Marco Ciardi. Roma: Carocci, 2020. 212 pp. 22 EUR. ISBN: 978-88-430-9620-6. Nuncius, 35(3):746–749, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/35/3/article-p746_14.xml.

Bovolo:2020:BRD

[3545] Carlo Bovolo. Book review: Donne e scienza nella Roma dell’Ottocento, by Federica Favino. Roma: Carocci, 2020. 212 pp. 22 EUR. ISBN: 978-88-430-9620-6. Nuncius, 35(3):750–752, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/view/journals/nun/35/3/article-p750_15.xml.

Lusito:2020:BRO

[3546] Fabio Lusito. Book review: Ombre nella mente. Lombroso e lo scapigliato, by Maria Antonietta Grignani, Paolo Mazzarello. Torino: Bollati Boringhieri, 2020. 174 pp. 15 EUR. ISBN: 978-88-399-3411-2. Nuncius, 35(3):753–755, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/35/3/article-p753_16.xml.

Schettini:2020:BRD

[3547] Laura Schettini. Book review: Donne delinquenti. Il genere e la nascita della criminologia, by Silvano Montaldo Roma: Carocci, 2019. 356 pp. 32 EUR. ISBN: 978-88-430-9827-9. Nuncius, 35(3):756–758, December 2020. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/35/
3/article-p756_17.xml.

Canadelli:2021:E

[3548] Elena Canadelli. Editorial. Nuncius, 36(1):1–3, April 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/36/1/article-p1_1.xml.

Danna:2021:F

[3549] Raffaele Danna. Figuring out. Nuncius, 36(1):5–48, April 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/36/1/article-p5_2.xml.

Yassi:2021:AKB

[3550] Yousef Yassi and Reza Yassi. Al-Khazini’s balance of wisdom. Nuncius, 36(1):49–66, April 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
1/article-p49_3.xml.

Ackermann:2021:HCR

[3551] Silke Ackermann and Louise Devoy. Humfrey Cole revisited. Nuncius, 36(1):67–94, April 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
1/article-p67_4.xml.

Ciancio:2021:PEK

[3552] Luca Ciancio. Promoting empirical knowledge in Habsburg Europe. Nuncius, 36(1):95–118, April 2021. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
36/1/article-p95_5.xml.

Marcacci:2021:SG

[3553] Flavia Marcacci. Seeing at a glance. Nuncius, 36(1):119–142, April 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/36/1/article-p119_6.xml.

Sandoval:2021:SIA

[3554] Omar Olivares Sandoval. The scientific images of the axolotl by José María Velasco and their role in Nineteenth-Century evolutionary thinking. Nuncius, 36(1):143–166, April 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/36/1/article-p143_7.xml.

Doran:2021:IVC

[3555] Connemara Doran. Instrumentalizing and visualizing the cosmic first light. Nuncius, 36(1):167–192, April 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/36/1/article-p167_8.xml.

Minuzzi:2021:PMK

[3556] Sabrina Minuzzi. Printing medical knowledge. Nuncius, 36(2):193– 197, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p193_1.xml.

Minuzzi:2021:CPM

[3557] Sabrina Minuzzi. 15th-century practical medicine in print. Nuncius, 36 (2):199–263, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p199_2.xml.

Cavallo:2021:EVM

[3558] Sandra Cavallo. Early vernacular medical advice books and their popular appeal in early modern Italy. Nuncius, 36(2):264–303, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/36/2/article-p264_3.xml.

Brenner:2021:RCB

[3559] Elma Brenner. The reception, consumption and broader context of a French vernacular plague tract printed in 1495. Nuncius, 36(2): 304–324, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p304_4.xml.

Storey:2021:LM

[3560] Tessa Storey. The lament of the melons. Nuncius, 36(2):325–355, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/36/2/article-p325_5.xml.

vanLeerdam:2021:PPI

[3561] Andrea van Leerdam. Popularising and personalising an illustrated herbal in Dutch. Nuncius, 36(2):356–393, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/view/journals/nun/36/2/article-p356_6.xml.

Barreiros:2021:MMH

[3562] Bruno Barreiros and Palmira Fontes da Costa. Materia medica and the history of the book in Seventeenth-Century Portugal. Nuncius, 36 (2):394–430, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p394_7.xml.

DiGennaroSplendore:2021:TT

[3563] Barbara Di Gennaro Splendore. The triumph of theriac. Nuncius, 36 (2):431–470, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p431_8.xml.

Savani:2021:BRP

[3564] Giacomo Savani. Book review: Public Baths and Bathing Habits in Late Antiquity. A Study of the Evidence from Italy, North Africa and Palestine A.D. 285–700, by Sadi Maréchal. Nuncius, 36(2):471–473, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/36/2/article-p471_9.xml.

Wingard:2021:BRT

[3565] Tim Wingard. Book review: The Thirteenth-Century Animal Turn: Medieval and Twenty-First-Century Perspectives, by Nigel Harris. Nuncius, 36(2):474–476, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p474_10.xml.

Zhao:2021:BRP

[3566] Lu Zhao. Book review: Physiognomy in Ming China: Fortune and the Body, by Xing Wang. Nuncius, 36(2):477–479, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/view/journals/nun/36/2/article-p477_11.xml.

Storey:2021:BRF

[3567] Tessa Storey. Book review: Forgotten Healers: Women and the Pursuit of Health in Late Renaissance Italy, by Sharon Strocchia. Nuncius, 36 (2):480–482, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p480_12.xml.

Giusti:2021:BRF

[3568] Enrico Giusti. Book review: Fermat et les débuts modernes de la géométrie, by Roshdi Rashed. Nuncius, 36(2):483–486, June 2021. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/view/journals/nun/36/2/article-p483_13.xml.

Abbri:2021:BRB

[3569] Ferdinando Abbri. Book review: Burchard de Volder and the Age of the Scientific Revolution, by Andrea Strazzoni. Nuncius, 36(2): 487–490, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p487_14.xml.

Baldassarri:2021:BRC

[3570] Fabrizio Baldassarri. Book review: Crocologia. A Detailed Study of Saffron, the King of Plants, by Sally Francis and Maria Teresa Ramandi, eds. Nuncius, 36(2):491–492, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/36/2/article-p491_15.xml.

deCeglia:2021:BRP

[3571] Francesco Paolo de Ceglia. Book review: La philosophie de la biologie avant de la biologie. Une histoire du vitalisme, by Charles Wolfe & The Gestation of German Biology. Philosophy and Physiology from Stahl to Schelling, by John H. Zammito. Nuncius, 36(2):493–496, June 2021. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/view/journals/nun/36/2/article-p493_16.xml.

Conforti:2021:BRP

[3572] Maria Conforti. Book review: Plague and the city, by Lukas Engelmann, John Henderson, and Christos Lynteris, eds. Nuncius, 36(2): 497–499, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p497_17.xml.

Cassata:2021:BRB

[3573] Francesco Cassata. Book review: Blood Relations. Transfusion and the Making of Human Genetics, by Jenny Bangham. Nuncius, 36(2):

500–502, June 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
2/article-p500_18.xml.

Capocci:2021:SR

[3574] Mauro Capocci and Claudio Pogliano. Science and race. Nuncius, 36 (3):503–515, November 2021. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
36/3/article-p503_1.xml.

Muller-Wille:2021:CTL

[3575] Staffan Müller-Wille. Corners, tables, lines. Nuncius, 36(3):517– 531, November 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
3/article-p517_2.xml.

Sera-Shriar:2021:CJT

[3576] Efram Sera-Shriar. Constructing the “Jewish Type”. Nuncius, 36(3): 532–567, November 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
3/article-p532_3.xml.

Pogliano:2021:TS

[3577] Claudio Pogliano. A tricky start. Nuncius, 36(3):568–610, November 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/36/3/article-p568_4.xml.

Iannuzzi:2021:LCP

[3578] Lucas Orlando Iannuzzi. Lidio Cipriani (1892–1962), the photographs in his popular science literature. Nuncius, 36(3):611–645, November 2021. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/36/3/article-p611_5.xml.

Limeira-DaSilva:2021:ARW

[3579] Victor Rafael Limeira-DaSilva and Juanma Sánchez Arteaga. Alfred Russel Wallace and the models of Amazonian “Indians” displayed at the Crystal Palace Ethnological Exhibition. Nuncius, 36(3):646– 675, November 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
3/article-p646_6.xml.

Falcucci:2021:VCP

[3580] Beatrice Falcucci. Visualizing colonial power. Nuncius, 36(3):676– 722, November 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
3/article-p676_7.xml.

Roque:2021:LSW

[3581] Ricardo Roque. The logic of skull writing. Nuncius, 36(3):723– 753, November 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
3/article-p723_8.xml.

Ariffin:2021:SD

[3582] Dayana Ariffin. Spatializing differences. Nuncius, 36(3):754–778, November 2021. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/36/
3/article-p754_9.xml.

Kortekangas:2021:PSS

[3583] Otso Kortekangas. Putting the Suenjel Sámi on the map. Nuncius, 36(3):779–804, November 2021. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
36/3/article-p779_10.xml.

Anonymous:2021:BM

[3584] Anonymous. Back matter. Nuncius, 36(3):805–809, November 2021. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/view/journals/nun/36/3/article-p805_11.xml.

Anonymous:2022:FM

[3585] Anonymous. Front matter. Nuncius, 37(1):i–iv, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/37/1/article-pi_0.xml.

Laube:2022:VM

[3586] Stefan Laube and Sergei Zotov. Vial movies. Nuncius, 37(1):1– 41, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
1/article-p1_1.xml.

Turner:2022:REM

[3587] Anthony Turner. Recycling early modern mathematical instruments. Nuncius, 37(1):42–58, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/37/1/article-p42_2.xml.

Seip:2022:GCT

[3588] Oscar Seip. Giulio Camillo’s theatre of knowledge revisited. Nuncius, 37 (1):59–83, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
1/article-p59_3.xml.

deCeglia:2022:FG

[3589] Francesco Paolo de Ceglia and Fabio Lusito. In the footsteps of Galileo. Nuncius, 37(1):84–117, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/37/1/article-p84_4.xml.

Hedesan:2022:APC

[3590] Georgiana D. Hedesan. Alchemy and Paracelsianism at the Casino di San Marco in Florence. Nuncius, 37(1):119–143, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/view/journals/nun/37/1/article-p119_5.xml.

Favino:2022:GAB

[3591] Federica Favino. Giovanni Alfonso Borelli’s last will (1679, December 31st). Nuncius, 37(1):144–173, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/1/article-p144_6.xml.

Beretta:2022:PB

[3592] Marco Beretta. Paolo Brenni (1954–2021). Nuncius, 37(1):175– 206, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
1/article-p175_7.xml.

Freguglia:2022:NEL

[3593] Paolo Freguglia. A new edition of Leonardo Fibonacci’s Liber Abbaci. Nuncius, 37(1):207–213, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/37/1/article-p207_8.xml.

Duranti:2022:BRH

[3594] Tommaso Duranti. Book review: Healers in the Making. Students, Physicians, and Medical Education in Medieval Bologna (1250–1550), by Kira Robison. Nuncius, 37(1):215–217, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/1/article-p215_9.xml.

Tomassetti:2022:BRS

[3595] Stefano Tomassetti. Book review: Saints, Infirmity, and Community in the Late Middle Ages, by Jenni Kuuliala. Nuncius, 37(1):218– 220, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
1/article-p218_10.xml.

Zanetti:2022:BRL

[3596] Cristiano Zanetti. Book review: Leonardo ingegnere, by Andrea Bernardoni. Nuncius, 37(1):221–223, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/1/article-p221_11.xml.

Leonard:2022:BRF

[3597] Marie-Louise Leonard. Book review: Florence Under Siege: Surviving Plague in an Early Modern City, by John Henderson. Nuncius, 37(1): 224–226, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
1/article-p224_12.xml.

Abbri:2022:BRT

[3598] Ferdinando Abbri. Book review: The Transmutations of Chymistry. Wilhelm Homberg and the Académie Royale des Sciences, by Lawrence M. Principe. Nuncius, 37(1):227–230, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/1/article-p227_13.xml.

Giovannetti-Singh:2022:BRT

[3599] Gianamar Giovannetti-Singh. Book review: Translation at Work: Chinese Medicine in the First Global Age, by Harold J. Cook, ed. Nuncius, 37(1):231–233, February 2022. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
37/1/article-p231_14.xml.

Nall:2022:BRM

[3600] Joshua Nall. Book review: Maps of the Moon. Lunar Cartography from the Seventeenth Century to the Space Age, by Thomas Haddad. Nuncius, 37(1):234–236, February 2022. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
37/1/article-p234_15.xml.

Mandrioli:2022:BRT

[3601] Mauro Mandrioli. Book review: A Tale of Two Viruses. Parallels in the Research Trajectories of Tumor and Bacterial Viruses, by Neeraja Sankaran. Nuncius, 37(1):237–239, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/1/article-p237_16.xml.

Ceccarelli:2022:BRC

[3602] David Ceccarelli. Book review: Catastrophic Thinking. Extinction and the Value of Diversity from Darwin to the Anthropocene, by David Sepkoski. Nuncius, 37(1):240–242, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/1/article-p240_17.xml.

Baldassarri:2022:STF

[3603] Fabrizio Baldassarri. The seed, the tree, the fruit, the juice. Nuncius, 37(2):243–253, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
2/article-p243_1.xml.

Baldassarri:2022:CO

[3604] Fabrizio Baldassarri. A clockwork orange. Nuncius, 37(2):255–283, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/37/2/article-p255_2.xml.

Niermeier-Dohoney:2022:MCT

[3605] Justin Niermeier-Dohoney. “To multiply corn two-hundred-fold”. Nuncius, 37(2):284–314, June 2022. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
37/2/article-p284_3.xml.

Jalobeanu:2022:ST

[3606] Dana Jalobeanu and Oana Matei. Spiritual technologies. Nuncius, 37 (2):315–345, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
2/article-p315_4.xml.

Eriksen:2022:PSP

[3607] Christoffer Basse Eriksen. Picturing seeds of poppies. Nuncius, 37(2): 346–373, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
2/article-p346_5.xml.

Mackenzie:2022:CV

[3608] Pamela Mackenzie. Contested vision. Nuncius, 37(2):374–396, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/37/2/article-p374_6.xml.

Turner:2022:ABC

[3609] Anthony Turner. The astrolabe and the book in 16th Century France. Nuncius, 37(2):397–420, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/37/2/article-p397_7.xml.

Ottaviani:2022:CDC

[3610] Alessandro Ottaviani. A chapter in the debate on chordate phylogeny. Nuncius, 37(2):421–478, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/37/2/article-p421_8.xml.

DiGennaroSplendore:2022:BRT

[3611] Barbara Di Gennaro Splendore. Book review: La Thériaque. Histoire d’ un remède millénaire, by Véronique Boudon-Millot and Françoise Micheau, eds. Nuncius, 37(2):479–481, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/2/article-p479_9.xml.

Rutkin:2022:BRA

[3612] H. Darrel Rutkin. Book review: Astrology, Almanacs, and the Early Modern English Calendar, by Phebe Jensen. Nuncius, 37(2):482– 484, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
2/article-p482_10.xml.

Tongiorgi:2022:BRB

[3613] Lucia Tomasi Tongiorgi. Book review: La biblioteca di Leonardo, by Carlo Vecce, ed. Nuncius, 37(2):485–488, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/view/journals/nun/37/2/article-p485_11.xml.

Ottaviani:2022:BRN

[3614] Alessandro Ottaviani. Book review: Natural desiderio di sapere. Roma barocca fra vecchi e nuovi mondi, by Sabina Brevaglieri. Nuncius, 37 (2):489–491, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
2/article-p489_12.xml.

Begley:2022:LRI

[3615] Justin Begley. Lady ranelagh: The incomparable life of Robert Boyle’s sister, by michelle DiMeo. Nuncius, 37(2):492–494, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/37/2/article-p492_13.xml.

Serrano:2022:BRM

[3616] Elena Serrano. Book review: Material Lives. Women Makers and Consumer Culture in the 18th Century, by Serena Dyer. Nuncius, 37(2): 495–497, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
2/article-p495_14.xml.

Serviant-Fine:2022:BRS

[3617] Thibaut Serviant-Fine. Book review: Strange Blood. The Rise and Fall of Lamb Blood Transfusion in 19th Century Medicine and Beyond, by Boel Berner. Nuncius, 37(2):498–500, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/2/article-p498_15.xml.

Mason:2022:BRF

[3618] Peter Mason. Book review: Florentino Ameghino y Hermanos. Empresa Argentina de Paleontología Ilimitada, by Irina Podgorny & Los argentinos vienen de los peces. Ensayo de filogenia nacional, by Irina Podgorny. Nuncius, 37(2):501–504, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/2/article-p501_16.xml.

Mazzarello:2022:BRB

[3619] Paolo Mazzarello. Book review: Brain and Race. A History of Cerebral Anthropology, by Claudio Pogliano. Nuncius, 37(2):505– 507, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
2/article-p505_17.xml.

Giorgione:2022:BRF

[3620] Claudio Giorgione. Book review: La Forma del Tempo (The Shape of Time), by Lavinia Galli. Nuncius, 37(2):509–511, June 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/37/2/article-p509_18.xml.

Pizzato:2022:OI

[3621] Fedra A. Pizzato. Objects of inquiry. Nuncius, 37(3):513–527, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/37/3/article-p513_1.xml.

Coltofean-Arizancu:2022:DAC

[3622] Laura Coltofean-Arizancu. Displaying archaeology and circulating knowledge. Nuncius, 37(3):529–563, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/3/article-p529_2.xml.

Pizzato:2022:BHE

[3623] Fedra A. Pizzato. Between the history of the Earth and the history of men. Nuncius, 37(3):564–589, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/3/article-p564_3.xml.

Mallart:2022:AAP

[3624] Lucila Mallart. Archaeology, archival practices and national identities in the iconographic repertoire of Spain (1915–1929). Nuncius, 37(3): 590–614, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
3/article-p590_4.xml.

Troilo:2022:VE

[3625] Simona Troilo. Visions of the empire. Nuncius, 37(3):615–642, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/37/3/article-p615_5.xml.

Hochadel:2022:FOA

[3626] Oliver Hochadel. Facing our ancestors. Nuncius, 37(3):643–673, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
3/article-p643_6.xml.

Raggetti:2022:AWR

[3627] Lucia Raggetti. Alchemical waters run deep. Nuncius, 37(3):675– 692, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
3/article-p675_7.xml.

Steinsiepe:2022:LVV

[3628] Klaus F. Steinsiepe. Leonardo da Vinci, the ventricles of the brain, and the foramen of Monro. Nuncius, 37(3):693–709, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/37/3/article-p693_8.xml.

Vignieri:2022:DRL

[3629] Valentina Vignieri. For the digital reconstruction of Leonardo’s Codex Atlanticus. Nuncius, 37(3):711–720, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/3/article-p711_9.xml.

Danna:2022:BRA

[3630] Raffaele Danna. Book review: Anachronisms in the History of Mathematics: Essays on the Historical Interpretation of Mathematical Texts, by Niccolò Guicciardini, ed. Nuncius, 37(3):721–723, December 2022. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/view/journals/nun/37/3/article-p721_10.xml.

Beneduce:2022:BRM

[3631] Chiara Beneduce. Book review: Le Moyen Âge et les sciences, by Danielle Jacquart and Agostino Paravicini Bagliani, eds. Nuncius, 37 (3):724–730, December 2022. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
37/3/article-p724_11.xml.

Taglialagamba:2022:BRS

[3632] Sara Taglialagamba. Book review: Lo specchio di Leonardo. Scritture e libri del genio universale, by Marco Cursi. Nuncius, 37(3):731– 735, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
3/article-p731_12.xml.

Mazzella:2022:BRS

[3633] Biagio Mazzella. Book review: The Shape of Sex. Nonbinary Gender from Genesis to the Renaissance, by Leah DeVun. Nuncius, 37(3):736– 738, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
3/article-p736_13.xml.

Salomoni:2022:BRN

[3634] David Salomoni. Book review: New World Objects of Knowledge. A Cabinet of Curiosities, by Mark Thurner and Juan Pimentel, eds. Nuncius, 37(3):739–741, December 2022. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
37/3/article-p739_14.xml.

Lange:2022:BRP

[3635] Carolin Lange. Book review: Painting with Fire: Sir Joshua Reynolds, Photography, and the Temporally Evolving Chemical Object, by Matthew C. Hunter. Nuncius, 37(3):742–743, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/37/3/article-p742_15.xml.

Mandrioli:2022:BRG

[3636] Mauro Mandrioli. Book review: Getting Under Our Skin. The Cultural and Social History of Vermin, by Lisa T. Sarasohn. Nuncius, 37(3): 744–746, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/37/
3/article-p744_16.xml.

Chapman:2022:BRS

[3637] Mary Chapman. Book review: Science Periodicals in Nineteenth-Century Britain: Constructing Scientific Communities, by Gowan Dawson, Bernard Lightman, Sally Shuttleworth, and Jonathan R. Topham, eds. Nuncius, 37(3):747–749, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/37/3/article-p747_17.xml.

McIvor:2022:BRG

[3638] Arthur McIvor. Book review: Germs in the English Workplace, c1880– 1945, by Laura Newman. Nuncius, 37(3):750–752, December 2022. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/view/journals/nun/37/3/article-p750_18.xml.

GrafvonHardenberg:2022:BRN

[3639] Wilko Graf von Hardenberg. Book review: Nature’s Diplomats. Science, Internationalism, and Preservation, 1920–1960, by Raf De Bont. Nuncius, 37(3):753–755, December 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/37/3/article-p753_19.xml.

Anonymous:2022:BM

[3640] Anonymous. Back matter. Nuncius, 37(3):757–763, December 2022. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/view/journals/nun/37/3/article-p757_20.xml.

Anonymous:2022:BCT

[3641] Anonymous. E-book collections title lists and MARC records. Nuncius, 37(3):??, December 2022. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/page/ebooktitles/
ebook-collections-title-lists-and-marc-records.

Anonymous:2023:FM

[3642] Anonymous. Front matter. Nuncius, 38(1):i–iv, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/38/1/article-pi_0.xml.

Veronesi:2023:LW

[3643] Umberto Veronesi. “The lute of wisdom”. Nuncius, 38(1):1–31, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/38/1/article-p1_1.xml.

Golvers:2023:DPE

[3644] Noël Golvers. A Dutch pharmacist in Early Modern Rome. Nuncius, 38 (1):32–71, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
1/article-p32_2.xml.

Gauvin:2023:HC

[3645] Jean-François Gauvin. Huygens’s carriole. Nuncius, 38(1):72–102, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL https://brill.com/view/journals/nun/38/1/article-p72_3.
xml.

Primbault:2023:HLR

[3646] Simon Dumas Primbault. How Leibniz read Pascal’s geometry. Nuncius, 38(1):103–136, February 2023. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
38/1/article-p103_4.xml.

Ezra:2023:CC

[3647] Ruth Ezra. Corpuscular conchology. Nuncius, 38(1):137–164, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/38/1/article-p137_5.xml.

deJonge:2023:DR

[3648] Hans Laurens de Jonge. Depicting race. Nuncius, 38(1):165–192, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
1/article-p165_6.xml.

He:2023:BRT

[3649] Yingtian He. Book review: Thinking in Cases: Ancient Greek and Imperial Chinese Case Narratives, by Markus Asper, ed. Nuncius, 38(1): 193–195, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
1/article-p193_7.xml.

Vinkesteijn:2023:BRP

[3650] Robert Vinkesteijn. Book review: Pseudo-Galenica. The Formation of the Galenic Corpus from Antiquity to the Renaissance, by Caroline Petit, Simon Swain, and Klaus-Dietrich Fischer, eds. Nuncius, 38(1): 196–198, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
1/article-p196_8.xml.

Celati:2023:FKM

[3651] Alessandra Celati. Forbidden knowledge: Medicine, science, and censorship in early modern Italy, by hannah Marcus. Nuncius, 38(1):199– 201, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
1/article-p199_9.xml.

Pardo-Tomas:2023:BRN

[3652] José Pardo-Tomás. Book review: Natura ed esattezza all’alba della scienza galileiana. Le ‘Observationes’ di Fabio Colonna, by Alessandro Ottaviani. Nuncius, 38(1):202–204, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/38/1/article-p202_10.xml.

Bowers:2023:BRH

[3653] Kristy Wilson Bowers. Book review: Health and Healing in the Early Modern Iberian World: A Gendered Perspective, by Margaret E. Boyle and Sarah E. Owens, eds. Nuncius, 38(1):205–207, February 2023. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/view/journals/nun/38/1/article-p205_11.xml.

Skelton:2023:BRM

[3654] Kimberley Skelton. Book review: The Moving Statues of Seventeenth-Century Amsterdam: Automata, Waxworks, Fountains, Labyrinths, by Angela Vanhaelen. Nuncius, 38(1):208–210, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/38/1/article-p208_12.xml.

Storni:2023:BRG

[3655] Marco Storni. Book review: Geminiano Montanari e altri studi di storia della scienza nella prima età moderna, by Salvatore Rotta. Nuncius, 38(1):211–213, February 2023. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
38/1/article-p211_13.xml.

Hegele:2023:BRK

[3656] Arden Hegele. Book review: Kept from All Contagion: Germ Theory, Disease, and the Dilemma of Human Contact in Late Nineteenth-Century Literature, by Kari Nixon. Nuncius, 38(1):214–216, February 2023. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/view/journals/nun/38/1/article-p214_14.xml.

Preveraud:2023:BRA

[3657] Thomas Preveraud. Book review: Au bureau de la revue. Une histoire de la publication scientifique (XIXe–XXe siècle), by Valérie Tesnière. Nuncius, 38(1):217–219, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/38/1/article-p217_15.xml.

Doran:2023:BRB

[3658] Connemara Doran. Book review: Beyond Sputnik and the Space Race. The Origins of Global Satellite Communications, by Hugh R. Slotten. Nuncius, 38(1):220–222, February 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/38/1/article-p220_16.xml.

Anonymous:2023:BCTa

[3659] Anonymous. E-book collections title lists and MARC records. Nuncius, 38(1):??, February 2023. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/page/ebooktitles/
ebook-collections-title-lists-and-marc-records.

Canadelli:2023:E

[3660] Elena Canadelli. Editorial. Nuncius, 38(2):223, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/view/journals/nun/38/2/article-p223_1.xml.

Didouan:2023:DBH

[3661] Amandine Victoria Didouan. “To draw a body, human or beast, one must study anatomy”. Nuncius, 38(2):225–249, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/38/2/article-p225_2.xml.

Bhaumik:2023:PS

[3662] Rahul Bhaumik. Picturing the snakes. Nuncius, 38(2):251–277, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/38/2/article-p251_3.xml.

Kaemmer:2023:DD

[3663] Hannah Kaemmer. Description or design. Nuncius, 38(2):278–310, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/38/2/article-p278_4.xml.

Wu:2023:STR

[3664] Chia-Yun Wu. The search for the Tang royal domain (Wangji [Chinese characters]). Nuncius, 38(2):311–339, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/38/2/article-p311_5.xml.

Kahn:2023:QVS

[3665] Didier Kahn. Les quatre versions successives du Cours de Chimie d’Étienne de Clave. Nuncius, 38(2):341–395, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/view/journals/nun/38/2/article-p341_6.xml.

Raggetti:2023:HD

[3666] Lucia Raggetti. The humours and the dyes. Nuncius, 38(2):397–443, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/38/2/article-p397_7.xml.

Azzolini:2023:US

[3667] Monica Azzolini. Under the surface. Nuncius, 38(2):445–463, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/38/2/article-p445_8.xml.

Luthy:2023:BRL

[3668] Christoph Lüthy. Book review: Lines of Thought: Branching Diagrams and the Medieval Mind, by Ayelet Even-Ezra. Nuncius, 38(2): 465–468, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
2/article-p465_9.xml.

Strano:2023:BRE

[3669] Giorgio Strano. Book review: An Early-Fourteenth-Century English Astrolabe in Milan and Contemporary Anglo-Italian Politics, by John Davis. Nuncius, 38(2):469–471, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/38/2/article-p469_10.xml.

Higgitt:2023:BRR

[3670] Rebekah Higgitt. Book review: Renaissance Invention. Stradanus’s Nova Reperta, by Lia Markey, ed. Nuncius, 38(2):472–474, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/38/2/article-p472_11.xml.

Hedesan:2023:BRP

[3671] Georgiana D. Hedesan. Book review: Pseudo-Paracelsus. Forgery and Early Modern Alchemy, Medicine and Natural Philosophy, by Didier Kahn and Hiro Hirai, eds. Nuncius, 38(2):475–477, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/38/2/article-p475_12.xml.

Chavarria:2023:BRM

[3672] Elisa Novi Chavarria. Book review: Medicina sotto il vulcano. Corpi e salute a Napoli in età moderna, by Maria Conforti. Nuncius, 38(2): 478–480, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
2/article-p478_13.xml.

Strazzoni:2023:BRM

[3673] Andrea Strazzoni. Book review: Il metodo al tavolo anatomico. Descartes e la medicina, by Fabrizio Baldassarri. Nuncius, 38(2): 481–484, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
2/article-p481_14.xml.

Beretta:2023:BRS

[3674] Marco Beretta. Book review: Spaces of Enlightenment Science, by Gordon McOuat and Larry Stewart, eds. Nuncius, 38(2):485– 487, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
2/article-p485_15.xml.

Dubald:2023:BRU

[3675] Déborah Dubald. Book review: Urban Histories of Science: Making Knowledge in the City, 1820–1940, by Oliver Hochadel and Agustí Nieto-Galan, eds. Nuncius, 38(2):488–490, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/38/2/article-p488_16.xml.

Mukherjee:2023:BRC

[3676] Nivedita Mukherjee. Book review: Cancer Virus Hunters: A History of Tumor Virology, by Gregory J. Morgan. Nuncius, 38(2): 491–494, June 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
2/article-p491_17.xml.

Anonymous:2023:BCTb

[3677] Anonymous. E-book collections title lists and MARC records. Nuncius, 38(2):??, June 2023. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/page/ebooktitles/
ebook-collections-title-lists-and-marc-records.

Gehrke:2023:EOF

[3678] Louisa-Dorothea Gehrke. Enclosing the open fields. Nuncius, 38(3):495– 519, November 2023. ISSN 0394-7394 (print), 1825-3911 (electronic).

URL https://brill.com/view/journals/nun/38/3/article-p495_
1.xml.

Campani:2023:OP

[3679] Francesca Campani. Objects of pleasure. Nuncius, 38(3):520–552, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
3/article-p520_2.xml.

Cristalli:2023:PSP

[3680] Claudia Cristalli. From postal scale to psychological apparatus. Nuncius, 38(3):553–582, November 2023. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
38/3/article-p553_3.xml.

Ceccarelli:2023:HPF

[3681] David Ceccarelli. Handling and photographing fossils. Nuncius, 38(3): 583–607, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
3/article-p583_4.xml.

Pogliano:2023:OTI

[3682] Claudio Pogliano. “Other than as illustration”. Nuncius, 38(3):608– 650, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
3/article-p608_5.xml.

Daniele:2023:ATP

[3683] Stefano Daniele. The angel and the thief with the pearl earrings. Nuncius, 38(3):651–689, November 2023. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
38/3/article-p651_6.xml.

Tripepi:2023:IPF

[3684] Alessandro Tripepi. International perspectives on the Florentine edition of Apollonius’ Conics. Nuncius, 38(3):690–710, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/38/3/article-p690_7.xml.

Pogliano:2023:NTM

[3685] Claudio Pogliano. New teche at the Museo Galileo. Nuncius, 38(3): 711–714, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
3/article-p711_8.xml.

Hagmann:2023:UHV

[3686] Johannes-Geert Hagmann. Unimaginable. How Van Leeuwenhoek’s microscope changed the world. Nuncius, 38(3):715–717, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/38/3/article-p715_9.xml.

Pantin:2023:BRL

[3687] Isabelle Pantin. Book review: Lynceorum Historia. Le ‘Schede Lincee’ di Martin Fogel, by Michele Camerota, Alessandro Ottaviani, and Oreste Trabucco, eds. Nuncius, 38(3):719–721, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.

com/view/journals/nun/38/3/article-p719_10.xml.

Baker:2023:BRG

[3688] Alexi Baker. Book review: Giuseppe Campani, “Inventor Romae,” an Uncommon Genius, by Silvio A. Bedini. Nuncius, 38(3):722– 724, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
3/article-p722_11.xml.

Antonelli:2023:BRL

[3689] Francesca Antonelli. Book review: Laura Bassi. Donne, genere e scienza nell’Italia del Settecento, by Marta Cavazza. Nuncius, 38(3): 725–727, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
3/article-p725_12.xml.

Orland:2023:BRB

[3690] Barbara Orland. Book review: Bodily Fluids, Chemistry and Medicine in the Eighteenth Century Boerhaave School, by Robert E. Verwaal. Nuncius, 38(3):728–730, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/38/3/article-p728_13.xml.

Minelli:2023:BRP

[3691] Alessandro Minelli. Book review: Plants and Politics in Padua during the Age of Revolution, 1820-1848, by Ariane Dröscher. Nuncius, 38(3): 731–733, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
3/article-p731_14.xml.

Cozzoli:2023:BRP

[3692] Daniele Cozzoli. Book review: Pasteur’s Empire: Bacteriology and Politics in France, its Colonies, and the World, by Aro Velmet. Nuncius, 38(3):734–737, November 2023. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
38/3/article-p734_15.xml.

Mandrioli:2023:BRQ

[3693] Mauro Mandrioli. Book review: Quel che resta. Scheletri e altri resti umani come beni culturali, by Maria Giovanna Belcastro, Giorgio Manzi, and Jacopo Moggi Cecchi, eds. Nuncius, 38(3):738–740, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/38/
3/article-p738_16.xml.

Bertozzi:2023:BRH

[3694] Eugenio Bertozzi. Book review: Historical Scientific Instruments in Contemporary Education, by Elizabeth Cavicchi and Peter Heering, eds. Nuncius, 38(3):741–743, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/38/3/article-p741_17.xml.

Madruga:2023:BRM

[3695] Catarina Madruga. Book review: Mobile Museums Collections in Circulation, by Felix Driver, Mark Nesbitt, and Caroline Cornish, eds. Nuncius, 38(3):744–746, November 2023. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/38/3/article-p744_18.xml.

Anonymous:2023:BM

[3696] Anonymous. Back matter. Nuncius, 38(3):747–752, November 2023. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https:

//brill.com/view/journals/nun/38/3/article-p747_19.xml.

Anonymous:2023:BCTc

[3697] Anonymous. E-book collections title lists and MARC records. Nuncius, 38(3):??, November 2023. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/page/ebooktitles/
ebook-collections-title-lists-and-marc-records.

Anonymous:2024:FM

[3698] Anonymous. Front matter. Nuncius, 39(1):i–iv, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/39/1/article-pi_0.xml.

Guignard:2024:ASW

[3699] Laurence Guignard. The amateur scientist’s workshop (1800–1950). Nuncius, 39(1):1–9, March 2024. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
39/1/article-p1_1.xml.

Henriquez:2024:PDO

[3700] Maria Angelica Esmeral Henriquez. To press, dry, organise, and represent. Nuncius, 39(1):11–34, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/39/1/article-p11_2.xml.

Baridon:2024:EVD

[3701] Laurent Baridon. Eugène Viollet-le-Duc, an amateur orologist on the cusp between romanticism and positivism. Nuncius, 39(1):35–50, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL

https://brill.com/view/journals/nun/39/1/article-p35_3.xml.

Richard:2024:WDG

[3702] Nathalie Richard. The work of a dilettante or a grand amateur? Nuncius, 39(1):51–75, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/39/
1/article-p51_4.xml.

Fages:2024:EBM

[3703] Volny Fages. Émile Belot’s models, or an engineer’s quest for scientific legitimacy. Nuncius, 39(1):76–98, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/39/1/article-p76_5.xml.

Serina:2024:PJS

[3704] Florent Serina. Pierre Janet’s secret garden, or the botanical collection of a psychopathologist. Nuncius, 39(1):99–115, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/39/1/article-p99_6.xml.

Guignard:2024:AAI

[3705] Laurence Guignard. The amateur astronomer as image producer. Nuncius, 39(1):116–140, March 2024. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
39/1/article-p116_7.xml.

Guillemain:2024:PDC

[3706] Hervé Guillemain. The pendulum that diagnoses and cures. Nuncius, 39 (1):141–161, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/39/
1/article-p141_8.xml.

Gigante:2024:MIA

[3707] Federica Gigante. A medieval Islamic astrolabe with Hebrew inscriptions in Verona. Nuncius, 39(1):163–192, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/39/1/article-p163_9.xml.

Luciano:2024:RMG

[3708] Erika Luciano. Retaining and mobilising geometric knowledge. Nuncius, 39(1):193–216, March 2024. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
39/1/article-p193_10.xml.

Pizzato:2024:BMU

[3709] Fedra A. Pizzato. Botanical Museum, University of Padua. Nuncius, 39 (1):217–221, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/39/
1/article-p217_11.xml.

Scuro:2024:BRU

[3710] Rachele Scuro. Book review: Uomini e fiumi. Per una storia idraulica ed agraria della bassa pianura del Po (1450–1620), by Franco Cazzola. Nuncius, 39(1):223–227, March 2024. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/view/journals/nun/
39/1/article-p223_12.xml.

Storni:2024:BRU

[3711] Marco Storni. Book review: Underground Mathematics. Craft Culture and Knowledge Production in Early Modern Europe, by Thomas Morel. Nuncius, 39(1):228–231, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/39/1/article-p228_13.xml.

Omodeo:2024:BRM

[3712] Pietro D. Omodeo. Book review: Mathematics and Society. Numbers and Measures in Early Modern South India, by Senthil Babu D. Nuncius, 39 (1):232–237, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/39/
1/article-p232_14.xml.

Mandrioli:2024:BRH

[3713] Mauro Mandrioli. Book review: In the Herbarium. The Hidden World of Collecting and Preserving Plants, by Maura C. Flannery. Nuncius, 39 (1):238–240, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/39/
1/article-p238_15.xml.

DeVos:2024:BRM

[3714] Paula DeVos. Book review: Merchants of Medicines. The Commerce and Coercion of Health in Britain’s Long Eighteenth Century, by Zachary Dorner. Nuncius, 39(1):241–243, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/39/1/article-p241_16.xml.

Osborne:2024:BRV

[3715] Michael A. Osborne. Book review: Victims of Fashion: Animal Commodities in Victorian Britain, by Helen Louise Cowie. Nuncius, 39(1): 244–246, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/39/
1/article-p244_17.xml.

Lefkaditou:2024:BRB

[3716] Ageliki Lefkaditou. Book review: Bones and Bodies. How South African Scientists Studied Race, by Alan G. Morris. Nuncius, 39(1): 247–249, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/39/
1/article-p247_18.xml.

DiLieto:2024:BRI

[3717] Paola Bernadette Di Lieto. Book review: The Italian Genius on Display. The First National Exhibition of History of Science (Florence, 1929) and the Preservation of Scientific Heritage in Fascist Italy, by Francesco Barreca. Nuncius, 39(1):250–251, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/

view/journals/nun/39/1/article-p250_19.xml.

Ienna:2024:BRS

[3718] Gerardo Ienna. Book review: Scienziati e guerra fredda. Tra collaborazione e diritti umani, by Elisabetta Bini and Elisabetta Vezzosi, eds. Nuncius, 39(1):252–254, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://brill.com/view/

journals/nun/39/1/article-p252_20.xml.

Johnston:2024:BRC

[3719] Stephen Johnston. Book review: Catalogue of Surveying and Related Instruments: Museo Galileo, by Jim Bennett. Nuncius, 39(1): 255–257, March 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL https://brill.com/view/journals/nun/39/
1/article-p255_21.xml.

Anonymous:2024:BCTa

[3720] Anonymous. E-book collections title lists and MARC records. Nuncius, 39(1):??, March 2024. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/page/ebooktitles/
ebook-collections-title-lists-and-marc-records.

Hsiung:2024:I

[3721] Hansun Hsiung. Introduction. Nuncius, 39(2):259–274, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/view/journals/
nun/39/2/article-p259_1.xml;   https://brill.com/view/journals/
nun/39/2/article-p259_1.xml.

Montero:2024:TFM

[3722] Elena Paulino Montero. Touching female memories in the purification funerary chapel in Burgos (c. 1482–1531). Nuncius, 39(2): 275–304, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
view/journals/nun/39/2/article-p275_2.xml;   https://brill.
com/view/journals/nun/39/2/article-p275_2.xml.

Cooley:2024:LAN

[3723] Mackenzie Cooley. Likeness across nature. Nuncius, 39(2):305– 337, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
view/journals/nun/39/2/article-p305_3.xml;   https://brill.
com/view/journals/nun/39/2/article-p305_3.xml.

Serrano:2024:TS

[3724] Elena Serrano. Touching the soul. Nuncius, 39(2):338–361, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/view/journals/
nun/39/2/article-p338_4.xml;   https://brill.com/view/journals/
nun/39/2/article-p338_4.xml.

Moscardo:2024:BOH

[3725] Ester García Moscardó. Bodies and objects in the Hispanic occupation of Tahiti. Nuncius, 39(2):362–389, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/view/journals/nun/39/2/article-p362_5.xml;
https://brill.com/view/journals/nun/39/2/article-p362_5.xml.

Hsiung:2024:GDH

[3726] Hansun Hsiung. The gentleman, the detective, and the housewife. Nuncius, 39(2):390–416, June 2024. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL http://booksandjournals.brillonline.
com/view/journals/nun/39/2/article-p390_6.xml;   https://brill.
com/view/journals/nun/39/2/article-p390_6.xml.

Palfreyman:2024:MMS

[3727] Harriet Palfreyman. Medical masculinity and the sensation of suffering in Leonard Portal Mark’s Acromegaly (1912). Nuncius, 39(2): 417–443, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-

3911 (electronic). URL http://booksandjournals.brillonline.com/
view/journals/nun/39/2/article-p417_7.xml;   https://brill.
com/view/journals/nun/39/2/article-p417_7.xml.

Jordanova:2024:TV

[3728] Ludmilla Jordanova. Touching visions. Nuncius, 39(2):444–450, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic).  URL http://booksandjournals.brillonline.com/
view/journals/nun/39/2/article-p444_8.xml;   https://brill.
com/view/journals/nun/39/2/article-p444_8.xml.

deCeglia:2024:SPE

[3729] Francesco Paolo de Ceglia. In search of the phoenix in Eighteenth-Century Naples. Nuncius, 39(2):451–485, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/view/journals/nun/39/2/article-
p451_9.xml;   https://brill.com/view/journals/nun/39/2/article-
p451_9.xml.

Dekker:2024:DSR

[3730] Elly Dekker. On dating the stars on the rete of the Verona astrolabe. Nuncius, 39(2):486–490, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/view/journals/nun/39/2/article-p486_10.xml;
 https://brill.com/view/journals/nun/39/2/article-p486_10.xml.

Kahn:2024:ERH

[3731] Didier Kahn. An early Renaissance Hebrew alchemical treatise in French translation (ca. 1540)? Nuncius, 39(2):491–528, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/view/

journals/nun/39/2/article-p491_11.xml;   https://brill.com/
view/journals/nun/39/2/article-p491_11.xml.

Strano:2024:MJJ

[3732] Giorgio Strano. In memory of James (Jim) Arthur Bennett (1947– 2023). Nuncius, 39(2):529–536, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/view/journals/nun/39/2/article-p529_12.xml;
 https://brill.com/view/journals/nun/39/2/article-p529_12.xml.

Scalas:2024:BRB

[3733] Giulia Scalas. Book review: Bodily Fluids in Antiquity, by Mark Bradley, Victoria Leonard, and Laurence Totelin, eds. Nuncius, 39(2):537–539, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/view/
journals/nun/39/2/article-p537_13.xml;   https://brill.com/
view/journals/nun/39/2/article-p537_13.xml.

Klerk:2024:BRM

[3734] Saskia Klerk. Book review: Making Physicians. Traditions, Teaching, and Trials at Leiden University, 1575-1639, by Evan R. Ragland. Nuncius, 39(2):540–542, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/view/journals/nun/39/2/article-p540_14.xml;
 https://brill.com/view/journals/nun/39/2/article-p540_14.xml.

Baldassarri:2024:BRS

[3735] Fabrizio Baldassarri. Book review: Skepticism Pictures: Figuring Descartes’s Natural Philosophy, by Melissa Lo. Nuncius, 39(2):543–544, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/view/
journals/nun/39/2/article-p543_15.xml;   https://brill.com/
view/journals/nun/39/2/article-p543_15.xml.

Bicak:2024:BRD

[3736] Ivana Bicak. Book review: Dangerous Drugs: The Self-Presentation of the Merchant-Poet Joannes Six van Chandelier (1620–1695), by Ronny Spaans. Nuncius, 39(2):545–546, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/view/journals/nun/39/2/article-p545_16.xml;
 https://brill.com/view/journals/nun/39/2/article-p545_16.xml.

Rizzi:2024:BRC

[3737] Elena Maria Rita Rizzi. Book review: Collective Wisdom: Collecting in the Early Modern Academy, by Anna Marie Roos and Vera Keller, eds. Nuncius, 39(2):547–549, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/view/journals/nun/39/2/article-p547_17.xml;
 https://brill.com/view/journals/nun/39/2/article-p547_17.xml.

Kambas:2024:BRL

[3738] Elizabeth Kambas. Book review: Ladies of Honor and Merit: Gender, Useful Knowledge, and Politics in Enlightened Spain, by Elena Serrano. Nuncius, 39(2):550–553, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/view/journals/nun/39/2/article-p550_18.xml;
 https://brill.com/view/journals/nun/39/2/article-p550_18.xml.

Abbri:2024:BRA

[3739] Ferdinando Abbri. Book review: The Arsenal of the Eighteenth-Century Chemistry. The Laboratories of Antoine Laurent Lavoisier (1743-1794), by Marco Beretta and Paolo Brenni. Nuncius, 39(2):554– 557, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/view/
journals/nun/39/2/article-p554_19.xml;   https://brill.com/
view/journals/nun/39/2/article-p554_19.xml.

Raposo:2024:BRA

[3740] Pedro M. P. Raposo. Book review: The Astronomer’s Chair: A Visual and Cultural History, by Omar W. Nasim. Nuncius, 39(2):558–560, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (elec-

tronic). URL http://booksandjournals.brillonline.com/view/
journals/nun/39/2/article-p558_20.xml;   https://brill.com/
view/journals/nun/39/2/article-p558_20.xml.

Schettini:2024:BRL

[3741] Laura Schettini. Book review: L’università delle donne. Accademiche e studentesse dal Seicento a oggi, by Andrea Martini e Carlotta Sorba, eds. Nuncius, 39(2):561–563, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://booksandjournals.

brillonline.com/view/journals/nun/39/2/article-p561_21.xml;
 https://brill.com/view/journals/nun/39/2/article-p561_21.xml.

Fabbri:2024:BRF

[3742] Natacha Fabbri. Book review: Femmes de science. Quatre siècles de conquêtes, entre langue et littérature, by Natasa Raschi and Cristina Trinchero, eds. Nuncius, 39(2):564–566, June 2024. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL http://

booksandjournals.brillonline.com/view/journals/nun/39/2/article-
p564_22.xml;   https://brill.com/view/journals/nun/39/2/article-
p564_22.xml.

Anonymous:2024:BCTb

[3743] Anonymous. E-book collections title lists and MARC records. Nuncius, 39(2):??, June 2024. CODEN ???? ISSN 0394-7394 (print),

1825-3911 (electronic). URL https://brill.com/page/ebooktitles/
ebook-collections-title-lists-and-marc-records.

Abramo:2002:GBF

[3744] Maria Rita Abramo. Gaston Bachelard e le fisiche del Novecento. (Italian) [Gaston Bachelard and the physics of the Twentieth Century], volume 2 of Transazioni. Guida, Napoli, Italia, 2002. ISBN 88-7188-593-7. 235 pp. LCCN B2430.B254 A22 2002. 15.00 EUR.

Ackermann:1998:HCM

[3745] Silke Ackermann, editor. Humphrey Cole: Mint, measurement, and maps in Elizabethan England, volume 126 of Occasional paper / British Museum. British Museum, London, UK, 1998. ISBN 0-86159-126-7. ISSN 0142-4815. vi + 106 pp. LCCN Q185 .H85 1998.

Acot:1998:EOS

[3746] Pascal Acot, Patrick Blandin, B. P. Hamm, Jean-Marc Drouin, et al., editors. The European origins of scientific ecology (1800–1901). History of science, technology and medicine. Gordon and Breach Science, Amsterdam, The Netherlands, 1998. ISBN 90-5699-103-5 (hardcover + CD-ROM), 90-5699-104-3 (softcover). ISSN 1027-4030. xix + 932 pp. LCCN QH540.8 .E87 1998. Translated from the original French by B.P. Hamm.

Agnoletti:1998:SFN

[3747] Mauro Agnoletti. Segherie e foreste nel Trentino: dal Medioevo ai giorni nostri. (Italian) [Sawmills and forests in Trentino from the Middle Ages to the present da]. Monografie etnografiche trentine. Museo degli usi e costumi della gente trentina, San Michele all’Adige, Italy, 1998. ISBN 88-85352-12-X. 233 pp. LCCN SD202.I82 T743 1998.

Agrimi:2002:ISN

[3748] Jole Agrimi. Ingeniosa scientia nature: studi sulla fisiognomica medievale (Latin and Italian) [Ingenious knowledge of nature: studies on medieval physiognomy], volume 36; 8 of Millennio medievale Studi. SIS-MEL edizioni del Galluzzo, Tavarnuzze (Firenze), Italia, 2002. ISBN 88-8450-074-5. xii + 180 pp. LCCN BF854 .A34 2002. URL http://

www.gbv.de/dms/casalini/00/00974097.pdf.

Albanese:1996:NSN

[3749] Denise Albanese. New science, new world. Duke University Press, Durham, NC, USA, 1996. ISBN 0-8223-1759-1 (hardcover), 0-8223-1768-0 (paperback). xi + 244 pp. LCCN PR438.S35 A43 1996. URL http:/

/www.gbv.de/dms/bowker/toc/9780822317685.pdf.

Angelini:1993:AAV

[3750] Annarita Angelini and Walter Tega. Anatomie accademiche, Vol. 3, L’Istituto delle Scienze e l’Accademia. (Italian) [Academic Anatomy]. Cultura e vita civile nel Settecento. Società editrice il Mulino, Bologna, Italia, 1993. ISBN 88-15-04157-5. 573 pp. LCCN ????

Anonymous:2005:PUH

[3751] Anonymous. La pelle umana = The human skin, volume 13 of Micrologus. SISMEL. Società internazionale per lo studio del medioevo latino, edizioni del Galluzzo, Firenze, Italia, 2005. ISBN 88-8450-143-1. 772 + 86 pp. LCCN GN191 .P45 2005. URL http://books.google.com/

books?id=DrHaAAAAMAAJ.

Anonymous:2008:SNM

[3752] Anonymous. La scienza nel Mezzogiorno dopo l’Unità d’Italia. (Italian) [Science in the South after the Unification of Italy]. SuperSaggi Storiografia. Rubbettino, Soveria Mannelli, Italy, 2008. ISBN 88-498-1889-0. 300 pp. LCCN Q127.I8. 58.00 EUR. URL http://www.gbv.de/dms/

casalini/08/08650241.pdf.

Ash:1995:GPG

[3753] Mitchell G. Ash. Gestalt psychology in German culture, 1890–1967: holism and the quest for objectivity. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-47540-6 (hardcover). xii + 513 pp. LCCN BF203 .A84 1995.

URL http://www.loc.gov/catdir/description/cam026/94036273.
html;   http://www.loc.gov/catdir/toc/cam026/94036273.html.

Ash:1998:GPG

[3754] Mitchell G. Ash. Gestalt psychology in German culture, 1890–1967: holism and the quest for objectivity. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-64627-8. xii + 513 pp. LCCN BF203 .A84 1995.

URL http://www.loc.gov/catdir/description/cam0210/99162510.
html;   http://www.loc.gov/catdir/toc/cam025/99162510.html.

Ausejo:1993:MME

[3755] Elena Ausejo and Mariano Hormigón. Messengers of mathematics: European mathematical journals (1800–1946). Ciencia. Siglo Veintiuno, Madrid, Spain, 1993. ISBN 84-323-0802-1 (paperback). xxiv + 297 pp. LCCN QA41.7 .M47 1993.

Azzolini:2012:DSA

[3756] Monica Azzolini. The duke and the stars: astrology and politics in Renaissance Milan. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-06663-4. 370 pp. LCCN BF1729.P6 A99 2012.

Babbage:2007:PDV

[3757] Charles Babbage and Andrea Villa, editors. Passaggi dalla vita di uno scienziato: autobiografia dell’inventore del computer. (Italian) [Passages from the Life of a Philosopher: autobiography of the inventor of the computer]. Frontiere / [UTET libreria]. UTET libreria, Torino, Italia, 2007. ISBN 88-02-07715-0. xvi + 414 pp. LCCN ???? With an introduction by Vittorio Marchis.

Back:2000:TRR

[3758] Les Back and John Solomos, editors. Theories of race and racism: a reader. Routledge readers in sociology; Routledge student readers. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2000. ISBN 0-415-15671-8 (hardcover), 0-415-15672-6 (paperback). xxiv + 646 pp. LCCN HT1521 .T473 2000. URL http://site.

ebrary.com/lib/academiccompletetitles/home.action;   http:/
/site.ebrary.com/lib/alltitles/docDetail.action?docID=10017837;
 http://www.gbv.de/dms/bowker/toc/9780415156721.pdf;   http:
//www.gbv.de/dms/hbz/toc/ht010508114.pdf;   http://www.gbv.
eblib.com/patron/FullRecord.aspx?p=165045;   http://www.loc.
gov/catdir/enhancements/fy0649/99016826-d.html.

BaguenaCervellera:1985:EMC

[3759] Ma. José (María José) Báguena Cervellera et al. Estudios sobre la Medicina y la ciencia valencianas (siglos XVI–XIX). (Spanish) [Medical studies and Valencian science (16th–19th Centuries)], volume 28 of Cuadernos valencianos de historia de la medicina y de la ciencia; Serie A, Monografías. Cátedra de Historia de la Medicina, Universidad de Valencia, Valencia, España, 1985. ISBN 84-600-4005-4. 193 pp. LCCN R557.V3 E88 1985.

Bandinelli:2013:OCD

[3760] Angela Bandinelli. Le origini chimiche della vita: legami tra la rivoluzione di Lavoisier e la biologia di Lamarck. (Italian) [The chemical origins of life: links between the revolution of Lavoisier and the biology of Lamarck], volume 69 of Biblioteca di Nuncius. L. S. Olschki, Firenze, Italia, 2013. ISBN 88-222-6270-0. ISSN 1122-0910. ix + 252 pp. LCCN QP511.8.L38 B26 2013.

Banfield:1993:IIE

[3761] Edwin Banfield. The Italian influence on English barometers from 1780. Baros, Trowbridge, UK, 1993. ISBN 0-948382-07-4. 160 pp. LCCN QC886 .B366 1993.

Barr:1997:THH

[3762] Alan P. Barr, editor. Thomas Henry Huxley’s place in science and letters: centenary essays. The University of Georgia Humanities Center Series on science and the humanities. University of Georgia Press, Athens, GA, USA, 1997. ISBN 0-8203-1865-5. xiii + 353 pp. LCCN Q125 .T427 1997.

URL http://www.gbv.de/dms/bowker/toc/9780820318653.pdf.

BassoRicci:1997:DSS

[3763] Maria Basso Ricci et al. Due secoli di strumenti geomagnetici in Italia (1740–1971) = Two centuries of geomagnetic instruments in Italy (1740– 1971). Editrice Compositori, Bologna, Italia, 1997. ISBN 88-7794-091-

3. 234 pp. LCCN QC819 .D84 1997. URL http://www.gbv.de/dms/

goettingen/241534852.pdf.

Beauchamp:2001:HTT

[3764] Kenneth G. Beauchamp. History of Telegraphy: Its Technology and Applications, volume 26 of IEE history of technology series. Institution of Electrical Engineers, London, UK, 2001. ISBN 0-85296-792-6, 1-55022-407-7. xxii + 413 pp. LCCN TK5115

.B43 2001. URL http://gbv.eblib.com/patron/FullRecord.aspx?
p=432576;   http://site.ebrary.com/lib/alltitles/docDetail.
action?docID=10276620;   http://www.gbv.de/dms/bowker/toc/9780852967928.
pdf;   http://www.gbv.de/dms/bs/toc/328848360.pdf;   http:/
/www.gbv.eblib.com/patron/FullRecord.aspx?p=432576;   http:
//www.loc.gov/catdir/enhancements/fy0631/2004533187-d.html;

 http://www.loc.gov/catdir/enhancements/fy0631/2004533187-t.
html.

Becquemont:1998:CSR

[3765] Daniel Becquemont and Laurent Mucchielli. Le cas Spencer: religion, science et politique. (French) [The case of Spencer: religion, science, and politics]. Science, histoire et société. Presses universitaires de France, Paris, France, 1998. ISBN 2-13-049107-3. ISSN 1242-5087. ix + 358 pp. LCCN MLCS 2001/07197 (B).

Bedini:1999:PAI

[3766] Silvio A. Bedini. Patrons, artisans, and instruments of science, 1600–1750, volume 635 of Variorum collected studies series. Ashgate/Variorum, Aldershot, Hampshire, Great Britain, 1999. ISBN 0-86078-781-8. 280 (est.) pp. LCCN Q185 .B425 1999. URL http://www.

gbv.de/dms/bowker/toc/9780860787815.pdf.

Belhoste:2011:PSP

[3767] Bruno Belhoste. Paris savant: parcours et rencontres au temps des Lumières. (French) [Scholarly Paris: courses and meetings in the Enlightenment]. Colin, Paris, France, 2011. ISBN 2-200-25563-2. 310

pp. LCCN Q127.F8 B44 2011. URL http://www.loc.gov/catdir/
toc/fy12pdf03/2012428139.html.

Beltrami:1998:DGN

[3768] Eugenio Beltrami, Jules Hoüel, L. (Luciano) Boi, Livia Giacardi, and Rossana Tazzioli. La découverte de la géométrie non euclidienne sur la pseudosphère: les lettres d’Eugenio Beltrami à Jules Hoüel (1868–1881). (French) [The discovery of non-Euclidean geometry on the pseudosphere: the letters of Eugenio Beltrami to Jules Hoüel (1868–1881)]. Collection Sciences dans l’histoire. Librairie Scientifique et Technique Albert Blanchard, Paris, France, 1998. ISBN 2-85367-208-5. ISSN 1258-0996. iii + 278 pp. LCCN QA29.B512 A4 1998. URL http://www.zentralblatt-

math.org/zmath/en/search/?an=0921.01011.

Bensaude-Vincent:2002:SFO

[3769] Bernadette Bensaude-Vincent and Christine Blondel. Des savants face à l’occulte: 1870–1940. (French) [Scientists against the occult: 1870–1940]. Sciences et société. Éditons La Découverte, Paris, France, 2002. ISBN 2-7071-3616-6. 231 pp. LCCN BF1409.5 .D47 2002.

Bertomeu-Sanchez:2006:CMC

[3770] José Ramón Bertomeu-Sanchez and Agustí Nieto-Galan, editors. Chemistry, medicine, and crime: Mateu J. B. Orfila (1787–1853) and his times. Science History Publications/USA, Sagamore Beach, MA, 2006. ISBN 0-88135-275-6. xxv + 306 pp. LCCN RA1025.O6 C44 2006.

Bertrams:2013:SHM

[3771] Kenneth Bertrams, Nicolas Coupain, and Ernst Homburg. Solvay: History of a Multinational Family Firm. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-02480-3, 1-139-17605-6. 650 pp. LCCN ????

Betri:1996:RSO

[3772] Maria Luisa Betri and Duccio Bigazzi. Ricerche di storia: in onore di Franco Della Peruta. (Italian) [Historical research: in honor of Franco Della Peruta]. F. Angeli, Milano, Italy, 1996. ISBN 88-204-9854-5 (vol. 1), 88-204-9853-7 (vol. 2). ???? pp. LCCN DG551 .R5 1996.

Bevilacqua:2000:NVSa

[3773] Fabio Bevilacqua and Lucio Fregonese. Nuova Voltiana: studies on Volta and his times. Vol. 2, …. Collana di storia della scienza. Università degli studi di Pavia; U. Hoepli, Pavia / Milano, Italia, 2000. ISBN 88-203-2790-2. 113 pp. LCCN ????

Bevilacqua:2000:NVSb

[3774] Fabio Bevilacqua and Lucio Fregonese. Nuova Voltiana: studies on Volta and his times. Vol. 1, …. Collana di storia della scienza. Università degli studi di Pavia; U. Hoepli, Pavia / Milano, Italia, 2000. ISBN 88-203-2789-9. 157 pp. LCCN ????

Bevilacqua:2000:NVSc

[3775] Fabio Bevilacqua and Lucio Fregonese, editors. Nuova Voltiana: studies on Volta and his times. Collana di storia della scienza. Università degli studi di Pavia, Pavia / Milano, Italia, 2000–2003. ISBN 88-203-2789-9 (vol. 1), 88-203-2790-2 (vol. 2), 88-203-2791-0 (vol. 3), 88-203-3079-2 (vol. 4), 88-203-3273-6 (vol. 5). 157 + 113 + 151 + ??? + ??? (5 volumes) pp. LCCN QC515.V8 N86 2000.

Bevilacqua:2001:NVS

[3776] Fabio Bevilacqua and Lucio Fregonese. Nuova Voltiana: studies on Volta and his times. Vol. 3, …. Collana di storia della scienza. Università degli studi di Pavia; U. Hoepli, Pavia / Milano, Italia, 2001. ISBN 88-203-2791-0. 151 pp. LCCN ????

Bimbi:1998:BBP

[3777] Bartolomeo Bimbi, Silvia Meloni Trkulja, Lucia Tongiorgi Tomasi, et al. Bartolomeo Bimbi: un pittore di piante e animali alla corte dei Medici. (Italian) [Bartolomeo Bimbi: a painter of plants and animals in the Court of the Medici]. Edifir, Firenze, Italia, 1998. ISBN 88-7970-099-5, 88-7970-053-7. 223 pp. LCCN ????

Blay:2002:SDM

[3778] Michel Blay. La science du mouvement: de Galilé à Lagrange. (French) [The science of motion: from Galileo to Lagrange]. Belin sup histoire des sciences. Belin, Paris, France, 2002. ISBN 2-7011-3076-X. 332 pp. LCCN ????

Bleichmar:2012:VEB

[3779] Daniela Bleichmar. Visible empire: botanical expeditions and visual culture in the Hispanic Enlightenment. University of Chicago Press, Chicago, IL, USA, 2012. ISBN 0-226-05853-0 (hardcover), 0-226-05855-7 (e-book), 1-299-19219-X (MyiLibrary). xii + 286 pp. LCCN QK21.S7 B54 2012.

Blue:1999:CHC

[3780] Gregory Blue and Timothy Brook, editors. China and historical capitalism: genealogies of sinological knowledge. Studies in modern capitalism. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-64029-6 (hardcover), 2-7351-0789-2 (hardback, France only). xii + 291 pp. LCCN HC427. URL http://site.ebrary.

com/lib/alltitles/docDetail.action?docID=10461817;   http:/
/www.gbv.de/dms/bowker/toc/9782735107896.pdf.

Bolzano:1993:PIF

[3781] Bernard Bolzano, Hourya Sinaceur, and Bernard Bolzano, Paradoxien des Unendlichen. Les paradoxes de l’infini. (French) [Paradoxes of infinity]. Sources du savoir. Éd. du Seuil, Paris, France, 1993. ISBN 2-02-015919-8. 191 pp. LCCN QA9 .B685. URL http://www.zentralblatt-

math.org/zmath/en/search/?an=0772.01015.

Bonoli:1998:LII

[3782] Fabrizio Bònoli. L’astronomia in Italia. (Italian) [Astronomy in Italy]. Arte tipografica, Napoli, Italia, 1998. ISBN 88-87375-00-3. 156 pp. LCCN QB33.I8 A55 1998.

Bonoli:2001:LAP

[3783] Fabrizio Bònoli and Daniela Piliarvu. I lettori di astronomia presso lo studio di Bologna dal XII al XX secolo. (Italian) [Readers of the study of astronomy at Bologna from the Twelfth to the Twentieth Century], volume 7 of Musei e archivi dello Studio bolognese. CLUEB, Bologna, Italia, 2001. ISBN 88-491-1651-9. 282 pp. LCCN QB35 .B66 2001. URL

http://digital.casalini.it/8849116519.

Bougard:1999:CNL

[3784] Michel Bougard. La chimie de Nicolas Lemery. (French) [The chemistry of Nicolas Lemery], volume 40 = N.S., 3 of De diversis artibus. Brepols, Turnhout, Belgium, 1999. ISBN 2-503-50739-5. 524 pp.

LCCN QD22.L5255 B68 1999. URL http://www.gbv.de/dms/bs/toc/
265542081.pdf.

Bret:1997:OLC

[3785] Patrice Bret. Oeuvres de Lavoisier: correspondance, Vol. 6, 1789–1791. (French) [Works of Lavoisier: correspondence, Vol. 6, 1789–1791]. Acad. des Sciences, Paris, France, 1997. ISBN 2-85367-205-0. xviii + 491 pp. LCCN ????

Brian:1996:EL

[3786] Denis Brian. Einstein: a life. John Wiley, New York, NY, USA, 1996. ISBN 0-471-11459-6 (hardcover), 0-471-19362-3 (paper). xiv + 509 pp. LCCN QC16.E5 B737 1996. URL http:/

/www.gbv.de/dms/bowker/toc/9780471114598.pdf;   http://www.
gbv.de/dms/ilmenau/toc/183614305.PDF;   http://www.loc.gov/
catdir/enhancements/fy0706/95012075-b.html;   http://www.loc.
gov/catdir/enhancements/fy0706/95012075-d.html;   http://www.
loc.gov/catdir/toc/onix04/95012075.html;   http://www.zentralblatt-
math.org/zmath/en/search/?an=0849.01023.

Brook:1999:CHC

[3787] Timothy Brook and Gregory Blue, editors. China and historical capitalism: genealogies of sinological knowledge. Studies in modern capitalism = Études sur le capitalisme moderne. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-64029-6 (hardcover), 2-7351-0789-2. xii + 291 pp. LCCN HC427 .C444 1999.

URL http://www.gbv.de/dms/bowker/toc/9782735107896.pdf;
http://www.loc.gov/catdir/description/cam029/98038596.html;
 http://www.loc.gov/catdir/samples/cam032/98038596.html;   http:
//www.loc.gov/catdir/toc/cam022/98038596.html.

Brugnatelli:1997:DVS

[3788] Luigi Vincenzo Brugnatelli and Alberto Gigli Berzolari. Diario del viaggio in Svizzera e in Francia con Alessandro Volta nel 1801. (Italian) [Diary of the journey in Switzerland and in France by Alessandro Volta in 1801], volume 28 of Fonti e studi per la storia dell’Università di Pavia. Cisalpino, Bologna, Italia, 1997. ISBN 88-205-0798-6. 301 pp. LCCN QD22.B83 A3 1997.

Brush:1995:NEO

[3789] Stephen G. Brush and Stephen G. Brush. Nebulous Earth: the origin of the solar system and the core of the earth from Laplace to Jeffreys, volume 1 of A history of modern planetary physics. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-44171-4. xii + 312 pp. LCCN

???? URL http://www.gbv.de/dms/bowker/toc/9780521441711.pdf.

Brush:1996:FEO

[3790] Stephen G. (Stephen George) Brush. Fruitful encounters: the origin of the solar system and of the Moon from Chamberlin to Apollo, volume 3 of History of modern planetary physics. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55214-1 (hardback). xii + 354 pp. LCCN QB601 .B89 1996 vol.

3. URL http://www.gbv.de/dms/bowker/toc/9780521552141.pdf;
http://www.loc.gov/catdir/description/cam027/95032973.html;
 http://www.loc.gov/catdir/toc/cam026/95032973.html.

Brush:1996:HMPa

[3791] Stephen G. Brush. A history of modern planetary physics. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-44171-4 (vol. 1), 0-521-55213-3 (vol. 2), 0-521-55214-1 (vol. 3). ???? pp. LCCN QB601 .B89 1996.

Brush:1996:HMPb

[3792] Stephen G. Brush. A history of modern planetary physics, Vol. 1, Nebulous Earth. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-44171-4 (hardcover). xii + 312 pp. LCCN QB503; QB601. URL

http://www.gbv.de/dms/bowker/toc/9780521441711.pdf.

Brush:1996:HMPc

[3793] Stephen G. Brush. A history of modern planetary physics, Vol. 2, Transmuted past. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55213-3 (Hardback). x + 134 pp. LCCN QB601. URL http://

www.gbv.de/dms/bowker/toc/9780521552134.pdf.

Brush:1996:HMPd

[3794] Stephen G. Brush. A history of modern planetary physics, Vol. 3, Fruitful encounters. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55214-1 (hardcover). xii + 354 pp. LCCN QB601. URL http://

www.gbv.de/dms/bowker/toc/9780521552141.pdf.

Brush:1996:NEO

[3795] Stephen G. Brush. Nebulous Earth: the origin of the solar system and the core of the Earth from Laplace to Jeffreys, volume 1 of A history of modern planetary physics. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-44171-4 (hardcover). xii + 312 pp. LCCN

QB601 .B89 1996 vol. 1. URL http://www.gbv.de/dms/bowker/toc/
9780521441711.pdf;   http://www.loc.gov/catdir/description/
cam027/95032975.html;   http://www.loc.gov/catdir/toc/cam026/
95032975.html.

Brush:1996:TPA

[3796] Stephen G. Brush. Transmuted past: the age of the Earth and the evolution of the elements from Lyell to Patterson, volume 2 of A history of modern planetary physics. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-55213-3 (Hardback). x + 134 pp. LCCN

QB601 .B89 1996 vol. 2. URL http://www.gbv.de/dms/bowker/toc/
9780521552134.pdf;   http://www.loc.gov/catdir/description/
cam027/95032974.html;   http://www.loc.gov/catdir/toc/cam023/
95032974.html.

Burns:2000:JLB

[3797] Russell W. Burns. John Logie Baird, Television Pioneer, volume 28 of IEE history of technology series. Institution of Electrical Engineers, London, UK, 2000. ISBN 0-85296-797-7. xxv + 417 pp. LCCN TK6635.B3

B87 2000. URL http://gbv.eblib.com/patron/FullRecord.aspx?
p=411110;   http://site.ebrary.com/lib/alltitles/docDetail.
action?docID=10271862;   http://www.gbv.de/dms/bsz/toc/bsz091893976inh.
pdf;   http://www.gbv.eblib.com/patron/FullRecord.aspx?p=411110.

Calabi:2001:QMD

[3798] Lorenzo (Lorenzo) Calabi. I quaderni metafisici di Darwin: teleologia “metafisica” causa finale. (Italian) [The metaphysical notebooks of Darwin: final cause of metaphysical teleology], volume 44 of Filosofia. ETS, Pisa, Italia, 2001. ISBN 88-467-0441-X. 132 pp. LCCN BD544

.C35 2001. 10.33 EUR. URL http://www.gbv.de/dms/goettingen/
341853526.pdf.

Canadelli:2013:PPP

[3799] Elena Canadelli. “Più positivo dei positivisti”: antropologia psicologia, evoluzionismo in Tito Vignoli. (Italian) [More “positive” positivists: anthropology, psychology, evolution in Tito Vignoli], volume 111 of Philosophica. ETS, Pisa, Italia, 2013. ISBN 88-467-3603-6. 178 pp. LCCN B3656.V544 C36 2013. 18.00 EUR. URL http://www.loc.gov/catdir/

toc/casalini12/2877430.pdf.

Capshew:1999:PMS

[3800] James H. Capshew. Psychologists on the march: science, practice, and professional identity in America, 1929–1969. Cambridge studies in the history of psychology. Cambridge University Press, Cambridge, UK, 1999. ISBN 0-521-56267-8 (hardcover), 0-521-56585-5 (paperback). xii + 276 pp. LCCN BF108.U5 C36 1999. URL http://site.ebrary.

com/lib/alltitles/docDetail.action?docID=10459774;   http:/
/www.gbv.de/dms/bowker/toc/9780521562676.pdf;   http://www.
loc.gov/catdir/description/cam028/97047526.html;   http://www.
loc.gov/catdir/toc/cam025/97047526.html.

Carlino:1994:FCL

[3801] Andrea Carlino. La fabbrica del corpo: libri e dissezione nel Rinascimento. (Italian) [The body factory: books and dissection in the Renaissance], volume 622 of Piccola biblioteca Einaudi. Giulio Einaudi editore, Torino, Italia, 1994. ISBN 88-06-13519-8, 88-06-13519-8. xxiv + 267 pp. LCCN ????

CastelnuovoFrigessi:2003:CL

[3802] Delia Castelnuovo Frigessi. Cesare Lombroso, volume 241 of Biblioteca di cultura storica. Einaudi, Torino, Italia, 2003. ISBN 88-06-13866-9. 426 + 16 pp. LCCN HV6023.L6 C37 2003.

Charles-Daubert:1998:LEF

[3803] Françoise Charles-Daubert. Les libertins érudits en France au XVIIe siècle. (French) [Libertines scholars in France in the Seventeenth Century], volume 106 of Philosophies. Presses universitaires de France, Paris, France, 1998. ISBN 2-13-049345-9. ISSN 0766-1398. 126 pp. LCCN ????

Cimatti:2002:MSC

[3804] Felice Cimatti. La mente silenziosa: come pensano gli animali non umani. (Italian) [The silent mind: how the non-human animals think]. Futura. Editori riuniti, Roma, Italia, 2002. ISBN 88-359-5160-7. 239 pp. LCCN QL785 .C447 2002.

Conry:2000:LCH

[3805] Yvette Conry. L’évolution créatrice d’Henri Bergson: investigations critiques. (French) [The creative evolution of Henri Bergson: critical investigations]. Collection Epistémologie et philosophie des sciences. L’Harmattan, Paris, France, 2000. ISBN 2-7384-9903-1. 328 pp. LCCN B2430.B43 C645 2000.

Cook:1998:EHC

[3806] Alan H. Cook. Edmond Halley: Charting the Heavens and the Seas. Clarendon Press, New York, NY, USA, 1998. ISBN 0-19-850031-

9. xvi + 540 pp. LCCN QB36.H25 C66 1998. URL http:/

/www.gbv.de/dms/bowker/toc/9780198500315.pdf;   http://www.
gbv.de/dms/goettingen/232124825.pdf;   http://www.loc.gov/catdir/
enhancements/fy0635/97012596-d.html;   http://www.loc.gov/catdir/
enhancements/fy0635/97012596-t.html;   http://www.loc.gov/catdir/
enhancements/fy0723/97012596-b.html.

Coyaud:2000:GAM

[3807] Sylvie Coyaud and Matteo Merzagora. Guida ai musei della scienza e della tecnica. (Italian) [Guide to the museums of science and technology], volume 3 of Guide tematiche. ClupGuide, Torino, Italia, 2000. ISBN 88-8344-008-0. 285 pp. LCCN Q105.E85 C69 2000; MLCS 2006/40212 (Q).

Crisciani:2002:PAF

[3808] Chiara Crisciani. Il Papa e l’alchimia: Felice V, Guglielmo Fabri e l’elixir. (Italian) [The Pope and alchemy: Felice V, Guglielmo Fabri and the elixir], volume 10 of La corte dei papi. Viella, Roma, Italia, 2002. ISBN 88-8334-079-5. 217 pp. LCCN BX1295 .C75 2002; QD26 .C7426

2002. URL http://www.gbv.de/dms/casalini/02/02472775.pdf.

Crowe:1998:CCS

[3809] Michael J. Crowe, David R. Dyck, and James J. Kevin, editors. A calendar of the correspondence of Sir John Herschel, 1792–1871. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-63149-1. 828 pp. LCCN QB36.H59 C35 1998. URL http:/

/www.gbv.de/dms/bowker/toc/9780521631495.pdf;   http://www.
gbv.de/dms/goettingen/244500126.pdf;   http://www.loc.gov/catdir/
description/cam029/98022928.html;   http://www.loc.gov/catdir/
toc/cam026/98022928.html.

Cuomo:2007:TCG

[3810] S. (Serafina) Cuomo. Technology and culture in Greek and Roman antiquity. Key themes in ancient history. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-81073-6 (hardcover), 0-521-00903-0 (paperback). xi + 212 pp. LCCN T16 .C78 2007. URL http://www.loc.

gov/catdir/enhancements/fy0805/2007008985-b.html;   http://
www.loc.gov/catdir/enhancements/fy0805/2007008985-d.html;
http://www.loc.gov/catdir/toc/ecip0712/2007008985.html.

Curtis:1991:GSP

[3811] Robert Irvin Curtis. Garun and salsamenta: production and commerce in materia medica, volume 3 of Studies in ancient medicine. Brill, Leiden, The Netherlands, 1991. ISBN 90-04-09423-7 (hardcover). ISSN 0925-1421. xv + 226 + 8 pp. LCCN RS160.7 .C87 1991. URL http://www.

gbv.de/dms/bs/toc/02115130X.pdf.

DalPra:1983:SDF

[3812] Mario Dal Pra. Storia della filosofia. (Italian) [History of philosophy]. Vallardi, Milano, Italy, second edition, 1983–1998. ISBN 88-299-0752-9 (vol. 10), 88-299-1487-8 (vol. 11, part 1–2). ???? pp. LCCN ????

Dameri:1997:BOC

[3813] Debora Dameri, Achille Lodovisi, and Giulia Luppi. La buona opinione: cultura, scienza e misure negli stati estensi, 1598–1860. (Italian) [The good opinion: culture, science and measures in the Estensi states, 1598– 1860]. Grandezze e misure nella storia. Museo della bilancia, Centro di documentazione, Campogalliano (Modena), Italia, 1997. ISBN 88-86143-08-7. 291 pp. LCCN Q185 .B57 1997.

Darwin:1990:CCD

[3814] Charles Robert Darwin and Frederick Henry Burkhardt. The correspondence of Charles Darwin (1856–1857), volume 6 of The correspondence of Charles Darwin. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-25586-4. xxix + 673 pp. LCCN ???? URL http://www.

gbv.de/dms/bowker/toc/9780521255868.pdf.

Darwin:1991:SC

[3815] Charles Robert Darwin and Frederick Henry Burkhardt. 1858–1859: Supplement to the correspondence 1821–1857, volume 7 of The correspondence of Charles Darwin. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-38564-4. xxxv + 671 pp. LCCN ???? URL

http://www.gbv.de/dms/bowker/toc/9780521385640.pdf.

Darwin:1997:CCD

[3816] Charles Robert Darwin and Frederick Henry Burkhardt. The correspondence of Charles Darwin: 1862, volume 10 of The correspondence of Charles Darwin. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-59032-9. xxxvii + 936 pp. LCCN ???? URL http://www.

gbv.de/dms/goettingen/230319017.pdf.

Davey:2002:HLD

[3817] Frank Davey. How Linda died. ECW Press, Toronto, ON, Canada, 2002. ISBN 1-55022-497-2. 344 pp. LCCN RC280.B7. URL http://site.

ebrary.com/lib/alltitles/docDetail.action?docID=10279811;
http://www.gbv.de/dms/bowker/toc/9780852967928.pdf;   http:
//www.gbv.eblib.com/patron/FullRecord.aspx?p=423076.

Davis:2001:MAN

[3818] Peter Davis. Musei e ambiente naturale: il ruolo dei musei di storia naturale nella conservazione della biodiversità. (Italian) [Museums and the natural environment: the role of museums of natural history in the conservation of biodiversity], volume 9 of Lexis 4: MuseoPoli luoghi per il sapere. CLUEB, Bologna, Italia, 2001. ISBN 88-491-1284-X. xv + 350

pp. LCCN Q. URL http://digital.casalini.it/884911284X.

Dean:2007:RLG

[3819] Dennis R. Dean. Romantic landscapes: geology and its cultural influence in Britain, 1765–1835, volume 555 of History of earth sciences series; Scholars’ Facsimiles and Reprints, 0161-7729. Scholars’ Facsimiles and Reprints, Ann Arbor, MI, USA, 2007. ISBN 0-8201-1555-X. 426

pp. LCCN QE261 .D43 2007. URL http://www.loc.gov/catdir/toc/
ecip078/2007000331.html.

DeBenedictis:2008:PFN

[3820] Cristina De Benedictis et al. Presenze femminili nella cultura del Novecento: Adriana Tramontano, Maria Luisa Bonelli Righini, Luisa Becherucci. Centro Di, Firenze, Italia, 2008. ISBN 88-7038-467-5. 71 pp. LCCN HQ1645.F56 P74 2008.

Debus:2001:CMD

[3821] Allen G. Debus. Chemistry and medical debate: van Helmont to Boerhaave. Science History Publications, Canton, MA, USA, 2001. ISBN 0-88135-292-6. xvii + 277 pp. LCCN R148. URL http://www.gbv.de/

dms/bowker/toc/9780881352924.pdf.

deCeglia:2000:IAF

[3822] Francesco Paolo de Ceglia and Georg Ernst Associated name Stahl. Introduzione alla fisiologia di Georg Ernst Stahl. (Italian) [Introduction to the physiology of Georg Ernst Stahl]. Edizioni Pensa Multimedia, Lecce, Italia, 2000. ISBN 88-8232-044-8. 236 pp. LCCN QP.

deClercq:1997:LCP

[3823] Peter R. de (Peter Richard) Clercq. The Leiden Cabinet of Physics: a descriptive catalogue, volume 271 of Museum Boerhaave communication. Museum Boerhaave, Leiden, The Netherlands, 1997. ISBN 90-6292-114-0. 198 pp. LCCN QC9.N4 D43 1997.

DelCentina:2010:GLM

[3824] Andrea Del Centina and A. Fiocca. Guglielmo Libri matematico e storico della matematica: l’irresistibile ascesa dall’Ateneo pisano all’Institut de France. (Italian) [Guglielmo Libri, mathematician and historian of mathematics: the irresistible rise from the University of Pisa to the Institute of France], volume 47; 2 of Cultura e memoria / Provincia di Firenze; Fondi scientifici della Biblioteca Moreniana. L. S. Olschki, Firenze, Italia, 2010. ISBN 88-222-6003-1. xxi + 320 + xxii pp. LCCN QA29.L589 D45 2010.

DellaPorta:2002:C

[3825] Giovan Battista Della Porta and Raffaele Sirri. Commedie, volume 15 of Edizione nazionale delle opere Teatro; T. 3. Edizioni Scientifiche Italiane, Napoli, Italia, 2002. ISBN 88-495-0094-7 (vol. 1), 88-495-0256-7 (vol. 2), 88-495-0414-4 (vol. 3), 88-495-0415-2 (vol. 4). viii + 516 pp. LCCN ????

DelSoldato:2010:SPA

[3826] Eva Del Soldato. Simone Porzio: un aristotelico tra natura e grazia. (Italian) [Simone Porzio: an Aristotelian between nature and grace], volume 6 of Centuria. Edizioni di storia e letteratura, Roma, Italy, 2010. ISBN 88-6372-275-7. xvi + 352 pp. LCCN B785.P684 D45 2010. 42.00

EUR. URL http://www.loc.gov/catdir/toc/casalini09/11004541.
pdf.

Desmond:1997:HEH

[3827] Adrian Desmond. Huxley: evolution’s high priest. Michael Joseph, London, UK, 1997. ISBN 0-7181-3882-1. xiv + 370 + 16 pp. LCCN AC-QUISITION IN PROCESS (COPIED); QH31.H9 D45 1997. URL http:

//www.gbv.de/dms/bs/toc/22277102X.pdf.

deTocqueville:1997:VIT

[3828] Alexis de Tocqueville and Umberto Coldagelli. Viaggi. (Italian) [Travels]. Pantheon. Bollati Boringhieri, Torino, Italia, 1997. ISBN 88-339-1055-5. lxxxviii + 811 pp. LCCN ????

Diasio:1999:SIA

[3829] Nicoletta Diasio. La science impure: anthropologie et médecine en France, Grande-Bretagne, Italie, Pays-Bas. (French) [Impure science: anthropology and medicine in France, Great Britain, Italy, the Netherlands]. Sciences sociales et sociétés. Presses universitaires de France, Paris, France, 1999. ISBN 2-13-050092-7. ISSN 1281-6256. 286 pp. LCCN GN296.5.E85 D53 1999.

Dickson:1998:QCN

[3830] William Michael Dickson. Quantum chance and non-locality: probability and non-locality in the interpretations of quantum mechanics. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-61947-5, 0-521-58127-3. xix + 244 pp. LCCN QC174.12 .D53 1998. URL http://site.

ebrary.com/lib/alltitles/docDetail.action?docID=10447653;
http://www.gbv.de/dms/bowker/toc/9780521581271.pdf;   http:
//www.loc.gov/catdir/description/cam028/97008813.html;   http:
//www.loc.gov/catdir/toc/cam021/97008813.html;   http://www.
zentralblatt-math.org/zmath/en/search/?an=0909.60087.

DiTrocchio:1993:BDS

[3831] Federico Di Trocchio. Le bugie della scienza: perchè e come gli scienziati imbrogliano. (Italian) [The lies of science: why and how scientists cheat]. Saggi. A. Mondadori, Milano, Italy, 1993. ISBN 88-04-35111-X. 364 pp. LCCN Q175.37 .D5 1993. L34000.

Dolby:1996:UKI

[3832] Riki G. A. Dolby. Uncertain knowledge: an image of science for a changing world. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56004-7 (hardcover), 0-521-89262-7 (paperback). xi + 365 pp. LCCN Q175 .D688 1996. URL http://site.ebrary.

com/lib/alltitles/docDetail.action?docID=10443911;   http:/
/www.gbv.de/dms/bowker/toc/9780521560047.pdf;   http://www.
gbv.de/dms/goettingen/213323524.pdf;   http://www.loc.gov/catdir/
description/cam027/96022588.html;   http://www.loc.gov/catdir/
samples/cam031/96022588.html;   http://www.loc.gov/catdir/toc/
cam022/96022588.html.

dosReis:1997:MEP

[3833] António Estc̀io dos Reis and Peter Ingham. Medir estrelas. (Portuguese) [Measuring stars]. CTT Correios, Lisboa, Portugal, 1997. ISBN 972-9127-40-9. 151 pp. LCCN ????

Douthwaite:2002:WGN

[3834] Julia V. Douthwaite. The wild girl, natural man, and the monster: dangerous experiments in the Age of Enlightenment. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-16055-6, 0-226-16056-4 (paperback). xiii + 314 pp. LCCN BD450. URL http://site.ebrary.

com/lib/alltitles/docDetail.action?docID=10431279;   http:/
/www.gbv.de/dms/bowker/toc/9780226160559.pdf;   http://www.
gbv.eblib.com/patron/FullRecord.aspx?p=616032.

Dupre:2011:SMC

[3835] Sven Dupré and Christoph (Christoph Herbert) Lüthy, editors. Silent messengers: the circulation of material objects of knowledge in the Early Modern Low Countries, volume 1 of Low Countries studies on the circulation of natural knowledge. LIT Verlag, Berlin, Germany, 2011. ISBN 3-8258-1635-4 (paperback). ii + 387 pp. LCCN GN406 .S5826 2011.

Dyck:1998:CCS

[3836] David R. Dyck, James J. Kevin, and Michael J. Crowe, editors. A calendar of the correspondence of Sir John Herschel. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-63149-1. 828 pp. LCCN QB36.H59 C35 1998.

Ellegaard:1990:DGR

[3837] Alvar Ellegård. Darwin and the general reader: a reception of Darwin’s Theory of Evolution in the British periodical press, 1859–1872. University of Chicago Press, Chicago, IL, USA, 1990. ISBN 0-226-20487-1. 394 pp. LCCN QH365.O8 E45 1990. URL http:/

/www.gbv.de/dms/bowker/toc/9780226204871.pdf;   http://www.
loc.gov/catdir/enhancements/fy0608/89078233-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0608/89078233-t.html.

Enriques:2001:FEM

[3838] Federigo Enriques, Manlio Mazziotti, and Ornella Pompeo Faracovi. Federigo Enriques: matematiche e filosofia; lettere inedite; bibliografia degli scritti. (Italian) [Federigo Enriques: mathematics and philosophy; unpublished letters; bibliography of writings], volume 6 of Dimensionilibri. Belforte, Livorno, Italia, 2001. ISBN 88-7997-063-1. 210 pp. LCCN QA.

Ernesti:1997:FMA

[3839] Giulio Ernesti, Andrea Locatelli, Antonello Negri, Silvia Bignami, and Alberto Lagomaggiore. La fortuna del moderno: architetture della produzione e dei servizi in area milanese negli anni Venti e Trenta. (Italian)

[The fortune of the modern: the architecture of production and services in the Milan area in the 1920s and 1930s]. Edifir, Firenze, Italia, 1997. ISBN 88-7970-053-7. 223 pp. LCCN NA6403.I8 E76 1997.

Euclides:2003:DED

[3840] Euclides and Christian Marinus Taisbak. Dedomena: Euclid’s data or the importance of being given, volume 45 of Acta historica scientiarum naturalium et medicinalium. Museum Tusculanum Press, Copenhagen, Denmark, 2003. ISBN 87-7289-815-1. 271 pp. LCCN QA31. EUR 45,00.

URL http://www.gbv.de/dms/goettingen/362001758.pdf;   http:
//www.zentralblatt-math.org/zmath/en/search/?an=1032.01005.

Fara:1996:SAM

[3841] Patricia Fara. Sympathetic attractions: magnetic practices, beliefs, and symbolism in Eighteenth-Century England. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-01099-4 (hardcover). x + 326 pp. LCCN QC750.4 .F37 1996. URL http:/

/www.gbv.de/dms/bowker/toc/9780691010991.pdf;   http://www.
loc.gov/catdir/description/prin031/96010785.html;   http://
www.loc.gov/catdir/toc/prin031/96010785.html.

Favrholdt:1992:NBP

[3842] David Favrholdt. Niels Bohr’s philosophical background, volume 63 of Historisk-filosofiske meddelelser. Royal Danish Academy of Sciences and Letters, Copenhagen, Denmark, 1992. ISBN 87-7304-228-5. 147 pp. LCCN QC16.B63 F3 1992; AS281 .C79h; AS281 .D201.

Feingold:1990:BNL

[3843] Mordechai Feingold, editor. Before Newton: the life and times of Isaac Barrow. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-30694-9. xi + 380 pp. LCCN QA29.B37 B44 1990.

URL http://www.gbv.de/dms/bowker/toc/9780521306942.pdf;
http://www.loc.gov/catdir/description/cam024/89038508.html;
 http://www.loc.gov/catdir/enhancements/fy0632/89038508-t.html.

Ferri:1997:PAM

[3844] Sara Ferri. Pietro Andrea Mattioli: Siena, 1501–Trento, 1578: la vita, le opere: con l’identificazione delle piante. (Italian) [Pietro Andrea Mattioli: Siena, 1501–Trento, 1578: life and works on the identification of plants]. Quattroemme, Ponte San Giovanni, Perugia, Italia, 1997. ISBN 88-85962-27-0. 405 pp. LCCN R520.M38 P54 1997.

Fibonacci:2020:LA

[3845] Leonardo Fibonacci, Enrico Giusti, and Paolo D’Alessandro. Liber Abbaci. Leo S. Olschki, Firenze, Italia, 2020. ISBN 88-222-6658-7. cxvii + 822 + 22 pp. LCCN QA32 .F47 2020.

Finocchiaro:1997:GWS

[3846] Maurice A. Finocchiaro. Galileo on the World Systems: a New Abridged Translation and Guide. University of California Press, Berkeley, CA, USA, 1997. ISBN 0-520-20548-0 (hardcover), 0-520-20646-0 (paperback). xi + 425 pp. LCCN QB41 .G173 1997.

URL http://www.gbv.de/dms/bowker/toc/9780520206465.pdf;
http://www.gbv.eblib.com/patron/FullRecord.aspx?p=223236;
http://www.loc.gov/catdir/bios/ucal051/96004256.html;   http:
//www.loc.gov/catdir/description/ucal041/96004256.html.

FoderaSerio:1997:LAP

[3847] Giorgia Foderà Serio and Ileana Chinnici. L’Osservatorio astronomico di Palermo: la storia e gli strumenti. (Italian) [The Astronomical Observatory of Palermo: history and tools]. Flaccovio, Palermo, Italia, 1997. ISBN 88-7804-144-0. 175 pp. LCCN QB82.I85 P354 1997.

Fortino:1997:CRS

[3848] Mirella Fortino. Convenzione e razionalità scientifica in Henri Poincaré. (Italian) [Convention and scientific rationality in Henri Poincaré]. Collana di storia delle idee. Rubbettino, Soveria Mannelli, Italia, 1997. ISBN 88-7284-597-1. 268 pp. LCCN B.

Fox:1995:IHS

[3849] Christopher Fox, Roy Porter, and Robert Wokler, editors. Inventing Human Science: Eighteenth-Century Domains. University of California Press, Berkeley, CA, USA, 1995. ISBN 0-520-20010-1. xv + 357

pp. LCCN H51 .I58 1995. URL http://www.loc.gov/catdir/bios/
ucal051/95006242.html;   http://www.loc.gov/catdir/description/
ucal041/95006242.html;   http://www.loc.gov/catdir/toc/ucal041/
95006242.html.

Fox:1999:NDI

[3850] Robert Fox and Agustí Nieto-Galan, editors. Natural dyestuffs and industrial culture in Europe, 1750–1880, volume 2 of European studies in science history and the arts. Science History Publications, Canton, MA, USA, 1999. ISBN 0-88135-273-X. xxix + 354 + 4 pp. LCCN TP919 .N38 1999.

Franco:2001:LPS

[3851] Silvano Franco. Legislazione e politica sanitaria del fascismo. (Italian) [Legislation and health policy of Fascism], volume 11 of Collana di saggi storici. APES, Roma, Italia, 2001. ISBN 88-7233-016-5. 209 pp. LCCN

KKH3075. 15.49 EUR. URL http://www.gbv.de/dms/casalini/02/
02605163.pdf.

Frangsmyr:2000:SIBa

[3852] Tore Frängsmyr. Svensk idéhistoria: bildning och vetenskap under tusen år, Del 2, 1809–2000. (Swedish) [Swedish History of Ideas: Education and Science for a Thousand Years. Part 2, 1809–2000]. Natur och Kultur, Stockholm, Sweden, 2000. ISBN 91-27-08105-2. 422 pp. LCCN ????

Frangsmyr:2000:SIBb

[3853] Tore Frängsmyr. Svensk idéhistoria: bildning och vetenskap under tusen år, Del 1, 1000–1809. (Swedish) [Swedish History of Ideas: Education and Science for a Thousand Years. Part 1, 1000–1809]. Natur och Kultur, Stockholm, Sweden, 2000. ISBN 91-27-07793-4. 448 pp. LCCN ????

Frangsmyr:2000:SIBc

[3854] Tore Frängsmyr. Svensk idehistoria: bildning och vetenskap under tusen år. (Swedish) [Swedish History of Ideas: Education and Science for a Thousand Years]. Natur och kultur, Stockholm, Sweden, 2000. ISBN 91-27-07793-4 (vol. 1), 91-27-08105-2 (vol. 2). ???? pp. LCCN DL635 .F73 2000.

Freedberg:2002:ELG

[3855] David Freedberg. The eye of the Lynx: Galileo, his friends, and the beginnings of modern natural history. University of Chicago Press, Chicago, IL, USA, 2002. ISBN 0-226-26147-6 (hardcover), 0-226-26148-4 (paper). xii + 513 pp. LCCN Q127.I8 F74 2002. US$50.00; US$30.00. URL http:

//site.ebrary.com/lib/academiccompletetitles/home.action;
http://site.ebrary.com/lib/alltitles/docDetail.action?docID=
10364135;   http://www.e-streams.com/es0606/es0606_2546.html;
 http://www.gbv.de/dms/bowker/toc/9780226261478.pdf;   http:
//www.h-net.org/review/hrev-a0d4h7-aa;   http://www.loc.gov/
catdir/bios/uchi051/2002000361.html;   http://www.loc.gov/catdir/
description/uchi051/2002000361.html;   http://www.loc.gov/catdir/
toc/fy037/2002000361.html;   http://www.myilibrary.com?id=250419.

Frezzotti:2000:CDS

[3856] Renato Frezzotti, Gigliola Terenna, and Francesca Vannozzi. La collezione degli strumenti di oculistica. (Italian) [The collection of the tools of ophthalmology], volume 4 of Materiali / Patrimonio Storico-Scientifico dell’Università degli Studi di Siena. Nuova immagine editrice, Siena, Italia, 2000. ISBN 88-7145-167-8 (paperback), 88-7145-167-1. 157 pp. LCCN ????

GadebuschBondio:2005:MAK

[3857] Mariacarla Gadebusch Bondio. Medizinische Ästhetik: Kosmetik und plastische Chirurgie zwischen Antike und früher Neuzeit. (German) [Medical Aesthetics: Cosmetic and Plastic Surgery Between Antiquity and Early Modern Times], volume 56 of Humanistische Bibliothek Reihe 1, Abhandlungen. Fink, München, Germany, 2005. ISBN 3-7705-4101-4 (paperback). 238 pp. LCCN ????

Galilee:1997:DMS

[3858] Galilée and Maurice A. Finocchiaro. Dialogo dei massimi sistemi: Galileo on the world systems: a new abridged translation and guide. University of California Press, Berkeley, CA, USA, 1997. ISBN 0-520-20548-0, 0-520-20646-0 (paperback). xi + 425 pp. LCCN QB41. URL http:/

/www.gbv.de/dms/goettingen/193658291.pdf;   http://www.loc.
gov/catdir/bios/ucal051/96004256.html;   http://www.loc.gov/
catdir/description/ucal041/96004256.html.

Galison:2008:O

[3859] Peter Galison and Lorraine Daston. Objectivity. Zone Books, New York, NY, USA, 2008. ISBN 1-890951-78-1. 501 pp. LCCN BD220 .D37 2007.

Garratt:1994:EHR

[3860] Gerald Reginald Mansel Garratt. The Early History of Radio: from Faraday to Marconi, volume 20 of IEE history of technology series. Institution of Electrical Engineers, in association with the Science Museum, London, UK, 1994. ISBN 0-85296-845-0. 105 pp. LCCN TK6547

.G37 1994. URL http://gbv.eblib.com/patron/FullRecord.aspx?
p=407985;   http://site.ebrary.com/lib/alltitles/docDetail.
action?docID=10263855;   http://www.gbv.eblib.com/patron/FullRecord.
aspx?p=407985.

Gassendi:2006:VME

[3861] Pierre Gassendi. Vie et mœurs d’épicure. Les Belles lettres, Paris, France, 2006. ISBN 2-251-79983-4. cxix + 278 + 396 (two volumes) pp. LCCN ???? Translation, introduction, and annotations by Sylvie Taussig.

Gaukroger:1995:DIB

[3862] Stephen Gaukroger. Descartes: an intellectual biography. Clarendon Press, New York, NY, USA, 1995. ISBN 0-19-823994-7, 0-19-823724-3 (paperback). xviii + 499 pp. LCCN ???? URL http://www.gbv.de/

dms/bowker/toc/9780198237242.pdf.

Gemelli:2002:IBA

[3863] Benedino Gemelli. Isaac Beeckman: atomista e lettore critico di Lucrezio. (Italian) [Isaac Beeckman: atomist and critical reader of Lucretius], volume 1 of Le corrispondenze letterarie, scientifiche ed erudite dal Rinascimento all’età moderna Subsidia. L. S. Olschki, Firenze, Italia, 2002. ISBN 88-222-5075-3. xiii + 132 pp. LCCN Q143.B4155 G454 2002. 16.00 EUR.

Giglioni:2000:IMS

[3864] Guido Giglioni. Immaginazione e malattia: saggio su Jan Baptiste van Helmont. (Italian) [Imagination and disease: essay on Jan Baptiste van Helmont], volume 53 of Filosofia e scienza nel Cinquecento e nel Seicento. Serie I, Studi. FrancoAngeli, Milano, Italy, 2000. ISBN 88-464-2286-4. 187 pp. LCCN R524.H4 G53 2000.

Giovacchini:2012:LD

[3865] Julie Giovacchini. L’empirisme d’Epicure, volume 11 of Les anciens et les modernes, études de philosophie. Classiques Garnier, Paris, France, 2012. ISBN 2-8124-0504-X. ISSN 2109-8042. 228 pp. LCCN B512 .G46 2012.

Giusti:2001:MII

[3866] Enrico Giusti and Luigi Pepe. La matematica in Italia (1800–1950). (Italian) [Mathematics in Italy (1800–1950)]. Polistampa, Firenze, Italia, 2001. ISBN 88-8304-321-9. 182 pp. LCCN QA27.I8 M368 2001.

GomezReig:1994:ESP

[3867] Constantino Gómez Reig, José María López Piänero, and Jorge Navarro-Pérez, editors. Los estudios sobre la salud pública en la ciudad de Valencia; 1880–1900. (Spanish) [Studies on public health in the city of Valencia: 1880–1900], volume 1 of Científicos Valencianos. Ayuntamiento de Valencia, Valencia, España, 1994. ISBN 84-88639-32-5. 235 pp. LCCN RA516.V34 G66 1994.

Gordin:2015:SBH

[3868] Michael D. Gordin. Scientific Babel: how science was done before and after global English. University of Chicago Press, Chicago, IL, USA, 2015. ISBN 0-226-00029-X (hardcover), 0-226-00032-X (e-book). 415 pp. LCCN Q223 .G67 2015.

Gouk:1999:MSN

[3869] Penelope Gouk. Music, science and natural magic in Seventeenth-Century England. Yale University Press, New Haven, CT, USA, 1999. ISBN 0-300-07383-6 (hardcover). xii + 308 pp. LCCN ML3800 .G66 1999. US$40,00, DM 90,80, Eur 46,43. URL http://www.gbv.de/dms/

bowker/toc/9780300073836.pdf.

Greenblatt:2012:SHW

[3870] Stephen Greenblatt. The swerve: how the world became modern. W. W. Norton, New York, NY, USA, 2012. ISBN 0-393-34340-5 (paperback). 356 + 8 pp. LCCN ????

Grooss:2004:ESI

[3871] K. S. Grooss and B. W. J. Grob. Early surgical instruments. Museum Boerhaave, Leiden, The Netherlands, 2004. ISBN 90-6292-148-5. 183 pp. LCCN RC71 .G76 2004.

Guerrini:2002:ACN

[3872] Luigi Guerrini. Antonio Cocchi, naturalista e filosofo. (Italian) [Antonio Cocchi, naturalist and philosopher], volume 2 of Biblioteca di medicina e storia. Polistampa, Firenze, Italia, 2002. ISBN 88-8304-398-7. 338 + 4 pp. LCCN R520.C636 .G84 2002.

Hafliger:1981:TDM

[3873] Ernst Häfliger and Josef Burn-Hool. Tavole delle malerbe: una rappresentazione sinottica delle flora che accompagna le colture agrarie. (Italian) [Tables of weeds: a synoptic representation of the flora accompanying agricultural crops]. Ciba-Geigy, Basel, Switzerland, 1981. xxi + 49 pp. LCCN ????

Hagner:2005:GGG

[3874] Michael Hagner. Geniale Gehirne: zur Geschichte der Elitegehirnforschung. (German) [Ingenious brains. On the elite brain research history]. Wissenschaftsgeschichte. Wallstein, Göttingen, Germany, second edition, 2005. ISBN 3-89244-649-0. 375 pp. LCCN QP376. Eur 38.00.

URL http://www.gbv.de/dms/faz-rez/F18200410062494890.pdf;
http://www.gbv.de/dms/faz-rez/F18200411262554992.pdf;   http:
//www.gbv.de/dms/faz-rez/FD1200410012493160.pdf;   http://www.
gbv.de/dms/hbz/toc/ht014841825.pdf.

Hahn:2013:CPS

[3875] Roger Hahn, Patricia Radelet-de Grave, et al., editors. Correspondance de Pierre-Simon Laplace (1749–1827). (French) [Letters of Pierre-Simon Laplace (1749–1827)], volume 90(1–2) of De diversis artibus; De diversis artibus. Brepols, Turnhout, Belgium, 2013. ISBN 2-503-54846-6 (t. 1), 2-503-54847-4 (t. 2), 2-503-54129-1 (set). xiv + 1416 (two volumes) pp. LCCN Q143.L36 A3 2013.

Halleux:1987:EHS

[3876] Robert Halleux and Pascal Lefebvre. Étudier l’histoire des sciences et des techniques. (French) [Studying the history of science and techniques], volume 3 of Collection belge d’histoire des sciences, des techniques et de l’industrie. Éd. APPS, Bruxelles, Belgium, 1987. ISBN 2-87192-006-0. 153 pp. LCCN ????

Halleux:2001:PAR

[3877] Robert Halleux et al. Les publications de l’Académie royale des sciences de Paris (1666–1793). (French) [Publications of the Royal Academy of Sciences of Paris (1666–1793)], volume t. 52 = n.s. 15 of De diversis artibus. Brepols, Turnhout, Belgium, 2001. ISBN 2-503-51197-X (set). ???? pp. LCCN Z7407.F7 P83 2001.

Hannaford:1996:RHI

[3878] Ivan Hannaford. Race: the history of an idea in the West. Woodrow Wilson Center Press, Washington, DC, USA, 1996. ISBN 0-8018-5222-6 (hardcover), 0-8018-5223-4 (paperback). xviii + 448 pp. LCCN HT1507

.H36 1996. URL http://www.gbv.de/dms/hbz/toc/ht007224601.
pdf;   http://www.loc.gov/catdir/bios/jhu051/96004162.html;
 http://www.loc.gov/catdir/description/jhu051/96004162.html.

Haslam:1996:HRM

[3879] Fiona Haslam. From Hogarth to Rowlandson: medicine in art in Eighteenth-Century Britain. Liverpool University Press, Liverpool, UK, 1996. ISBN 0-85323-630-5 (cased), 0-85323-640-2 (paperback). xv + 336 pp. LCCN N8223 .H36 1996. URL http:/

/www.gbv.de/dms/bowker/toc/9780853236405.pdf;   http://www.

loc.gov/catdir/enhancements/fy0642/96156381-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0642/96156381-t.html.

Hasse:2000:AAL

[3880] Dag Nikolaus Hasse. Avicenna’s “De anima” in the Latin West: the formation of a peripatetic philosophy of the soul: 1160–1300, volume 1 of Warburg Institute studies and texts. Warburg Institute, London, UK, 2000. ISBN 0-85481-125-7. x + 350 pp. LCCN ????

Holmes:1998:ALN

[3881] Frederic Lawrence Holmes. Antoine Lavoisier, the next crucial year, or The sources of his quantitative method in chemistry. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-01687-9 (hardcover). vii + 184 pp. LCCN QD22.L4 H65 1998. URL http:/

/www.gbv.de/dms/bowker/toc/9780691016870.pdf;   http://www.
loc.gov/catdir/description/prin031/97013566.html;   http://
www.loc.gov/catdir/toc/prin031/97013566.html.

Holton:2005:VVS

[3882] Gerald James Holton. Victory and vexation in science: Einstein, Bohr, Heisenberg, and others. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01519-3. xi + 229 pp. LCCN Q180.A3 H65 2005.

Hunter:2000:RBS

[3883] Michael Cyril William Hunter. Robert Boyle: (1627–1691); scrupulosity and science. Boydell Press, Woodbridge, Suffolk, England, 2000. ISBN 0-85115-798-X. ix + 293 pp. LCCN Q143.B77 R59 2000. URL http://

www.gbv.de/dms/bowker/toc/9780851157986.pdf.

Huxley:1997:MPT

[3884] Thomas Henry Huxley and Alan P. Barr, editors. The major prose of Thomas Henry Huxley. University of Georgia Humanities Center series on science and the humanities. University of Georgia Press, Athens, GA, USA, 1997. ISBN 0-8203-1864-7. xxii + 366 pp. LCCN Q127.G7 H892

1997. URL http://www.gbv.de/dms/goettingen/197703127.pdf.

Ing:1997:DSS

[3885] Istituto nazionale di geofisica (Italy), M. (Maria) Basso Ricci, et al. Due secoli di strumenti geomagnetici in Italia (1740–1971) = Two centuries of geomagnetic instruments in Italy (1740–1971). Editrice compositori, Bologna, Italy, 1997. ISBN 88-7794-091-3. 234 pp. LCCN ????

Irby-Massie:2002:GSH

[3886] Georgia Lynette Irby-Massie and Paul Turquand Keyser. Greek science of the Hellenistic era: a sourcebook. Routledge Sourcebooks for the Ancient World. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2002. ISBN 0-415-23847-1 (hardcover), 0-415-23848-X (paperback). xxxv + 392 pp. LCCN Q127.G7 I72 2002. URL http://site.ebrary.com/

lib/alltitles/docDetail.action?docID=5004261;   http://www.
gbv.de/dms/bowker/toc/9780415238472.pdf;   http://www.gbv.eblib.
com/patron/FullRecord.aspx?p=180470;   http://www.loc.gov/catdir/
enhancements/fy0650/2001041999-d.html.

Irwin:1992:PEC

[3887] Geoffrey Irwin. The prehistoric exploration and colonisation of the Pacific. Cambridge University Press, Cambridge, UK, 1992. ISBN 0-521-40371-5 (hardcover), 0-521-47651-8 (paperback). vii + 240 pp.

LCCN GN871 .I78 1992. URL http://www.gbv.de/dms/bowker/toc/
9780521403719.pdf;   http://www.loc.gov/catdir/description/
cam024/91023105.html;   http://www.loc.gov/catdir/toc/cam028/
91023105.html.

Jacob:1976:NER

[3888] Margaret C. Jacob. The Newtonians and the English Revolution, 1689– 1720. Classics in the history and philosophy of science. Gordon and Breach, New York, NY, USA, 1976. ISBN 2-88124-400-9. ISSN 1042-4628. 288 pp. LCCN BL245 .J3 1990.

Jammer:1999:ERP

[3889] Max Jammer. Einstein and religion: physics and theology. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-00699-7, 0-691-10297-X (paperback). 279 pp. LCCN QC16.E5 J36 1999. URL http:

//site.ebrary.com/lib/academiccompletetitles/home.action;
http://site.ebrary.com/lib/alltitles/docDetail.action?docID=
10496617;   http://www.gbv.de/dms/bowker/toc/9780691006994.
pdf;   http://www.loc.gov/catdir/description/prin021/99024124.
html;   http://www.loc.gov/catdir/samples/prin031/99024124.
html;   http://www.loc.gov/catdir/toc/prin032/99024124.html.

Jardine:2008:GAQ

[3890] Nicholas Jardine and A. Ph. Segonds. La guerre des astronomes: la querelle au sujet de l’origine du système géo-héliocentrique à la fin du XVIe siècle, volume 9.–10 of Science et humanisme. Belles lettres, Paris, France, 2008. ISBN 2-251-34513-2 (v. 1), 2-251-34512-4 (v. 2/1–2). v. 1–2 in 3 pp. LCCN QB603.M6 J37 2008.

Jones:1997:ACK

[3891] James H. (James Howard) Jones. Alfred C. Kinsey: a public / private life. W. W. Norton, New York, NY, USA, 1997. ISBN 0-393-04086-0. xx + 937 + 16 pp. LCCN HQ18.32.K56 J65 1997.

Jungnickel:1996:C

[3892] Christa Jungnickel and Russell McCormmach. Cavendish, volume 220 of Memoirs of the American Philosophical Society held at Philadelphia for Promoting Useful Knowledge. American Philosophical Society, Philadelphia, PA, USA, 1996. ISBN 0-87169-220-1. 414 pp. LCCN Q11 .P612 vol. 220; QD22.C4 J86 1996.

Kanz:1997:NIZ

[3893] Kai Torsten Kanz. Nationalismus und internationale Zusammenarbeit in den Naturwissenschaften: die deutsch-französischen Wissenschaftsbeziehungen zwischen Revolution und Restauration, 1789–1832; mit einer Bibliographie der Übersetzungen naturwissenschaftlicher Werke. (German) [Nationalism and international cooperation in the natural sciences: German–French scientific relations between Revolution and Restoration, 1789–1832, with a bibliography of translations of scientific works], volume 39 of Boethius. Steiner, Stuttgart, Germany, 1997. ISBN 3-515-07079-6. 352 pp. LCCN Q127.G3 K36 1997. DM 124.00, sfr 124.00, S

905.00. URL http://www.gbv.de/dms/bs/toc/235897418.pdf.

Kawashima:2013:EDC

[3894] Keiko Kawashima. Émilie du Châtelet et Marie-Anne Lavoisier: science et genre au XVIIIe siècle. (French) [Émilie du Châtelet and Marie-Anne Lavoisier: Science and type in the Eighteenth Century], volume 21 of Champion essais. Champion, Paris, France, 2013. ISBN 2-7453-2492-6. 322 pp. LCCN ???? Translation to French by Ayako Lecaille-Okamura.

Keel:2001:LMC

[3895] Othmar Keel. L’avènement de la médecine clinique moderne en Europe, 1750–1815: politiques, institutions et savoirs. (French) [The arrival of modern clinical medicine in Europe, 1750–1815: politics, institutions, and knowledge]. Bibliothèque d’histoire de la médecine et de la santé. Presses de l’Université de Montréal, Montréal, Québec, Canada, 2001. ISBN 2-7606-1822-6, 2-8257-0762-7. 542 pp. LCCN

R484 .K44 2001. URL http://site.ebrary.com/lib/alltitles/
docDetail.action?docID=10176884.

Kennedy:1999:MSA

[3896] E. S. (Edward Stewart) Kennedy, Paul Kunitzsch, and R. P. Lorch, editors. The melon-shaped astrolabe in Arabic astronomy, volume 43 of Boethius. Steiner, Stuttgart, Germany, 1999. ISBN 3-515-07561-5. vi + 235 pp. LCCN QB85.A283 K46 1999. DM 88.00, sfr 88.00, S

642.00. URL http://www.zentralblatt-math.org/zmath/en/search/
?an=0938.01006.

Kim:2000:NPC

[3897] Yung Sik Kim. The natural philosophy of Chu Hsi (1130–1200), volume 235 of Memoirs of the American Philosophical Society held at Philadelphia for promoting useful knowledge. American Philosophical Society, Philadelphia, PA, USA, 2000. ISBN 0-87169-235-X. ISSN 0065-9738. xii + 380 pp. LCCN B128.C54 K55 2000. URL http://www.gbv.de/dms/

bowker/toc/9780871692351.pdf.

King:2011:AME

[3898] David A. (David Anthony) King. Astrolabes from Medieval Europe, volume CS977 of Variorum collected studies series. Ashgate Publishing Ltd, London, UK, 2011. ISBN 1-4094-2593-2 (hardcover), 1-4094-2788-9, 1-4094-2326-3, 1-4094-3331-5, 1-4094-3686-1, 1-4094-4458-9, 1-4094-2401-4. xvi + 404 pp. LCCN QB85 .K564 2011. US$85.00.

Koerner:1999:LNN

[3899] Lisbet Koerner. Linnaeus: nature and nation. Harvard University Press, Cambridge, MA, USA, 1999. ISBN 0-674-09745-9, 0-674-00565-1. viii + 298 pp. LCCN QH44 .K58 1999.

URL http://www.gbv.de/dms/bowker/toc/9780674005655.pdf;
http://www.gbv.de/dms/bowker/toc/9780674097452.pdf;   http:
//www.gbv.de/dms/bs/toc/24079107x.pdf.

Kragh:1999:QGH

[3900] Helge Kragh. Quantum Generations: a History of Physics in the Twentieth Century. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-09552-3 (paperback), 0-691-01206-7 (hardcover). xiv + 494 pp. LCCN QC7 .K7 1999. URL http://www.loc.gov/

catdir/description/prin021/99017903.html;   http://www.loc.
gov/catdir/samples/prin031/99017903.html;   http://www.loc.
gov/catdir/toc/prin032/99017903.html.

Kragh:2012:NBQ

[3901] Helge Kragh. Niels Bohr and the Quantum Atom: The Bohr Model of Atomic Structure 1913–1925. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-965498-0 (hardcover), 0-19-163045-4 (e-book), 0-19-174169-8 (e-book). vi + 410 pp. LCCN QC774.A2.

Kunkel:1999:WZR

[3902] Natalie Kunkel. Wissenschaftsaustausch zwischen Rußland und Westeuropa, insbesondere Deutschland, in der Botanik und Pharmazie vom 18. bis zum frühen 20. Jahrhundert. (German) [Scientific exchange between Russia and Western Europe, especially Germany, in botany and pharmacy of the 18th until the early 20th Century], volume 28 of Algorismus; Münchener Universitätsschriften. Institut für Geschichte der Naturwissenschaften, München, Germany, 1999. ISBN 3-89241-029-1. 255 + 14

pp. LCCN Ql21.G3 G86 1999. URL http://www.gbv.de/dms/bs/toc/
254241158.pdf.

Landecker:2007:CLH

[3903] Hannah Landecker. Culturing life: how cells became technologies. Harvard University Press, Cambridge, MA, USA, 2007. ISBN 0-674-02328-5. x + 276 pp. LCCN QH585.2 .L36 2007.

Lazzara:1999:CBT

[3904] Salvatore Lazzara. Claude Bernard: tra scienza e filosofia. (Italian) [Claude Bernard: between science and philosophy], volume 55 of Tessere: testi di divulgazione. CUEN, Napoli, Italia, 1999. ISBN 88-7146-437-0. 85 pp. LCCN RA. L10000.

Lesle:1990:ZNT

[3905] Lutz Lesle and Christian Berg. Zwischen Notenhälsen und Tonköpfen: 75 Jahre Musikbibliothek Hamburg; mit einem Verzeichnis der in der Musikbibliothek und in der Staats- und Universitätsbibliothek Hamburg Carl von Ossietzky vorhandenen Vertonungen von Gedichten Matthias Claudius. (German) [Between note stems and tape heads: 75 years of the Music Library of Hamburg, with a directory in the Music Library and the Carl von Ossietzky State and University Library of Hamburg: existing settings of poems by Matthias Claudius], volume 3 of Bibliothemata. T. Bautz, Herzberg, Germany, 1990. ISBN 3-88309-030-1. 118 pp. LCCN ML136.H3 H329 1990. URL http://www.dandelon.com/

servlet/download/attachments/dandelon/ids/AT0012F5B2.pdf.

Lindberg:1992:BWS

[3906] David C. Lindberg. The beginnings of Western science: the European scientific tradition in philosophical, religious, and institutional context, 600 B.C. to A.D. 1450. University of Chicago Press, Chicago, IL, USA, 1992. ISBN 0-226-48230-8 (hardcover), 0-226-48231-6 (paperback). xviii + 455 pp. LCCN Q124.95 .L55 1992. URL http:/

/www.gbv.de/dms/bowker/toc/9780226482316.pdf;   http://www.
loc.gov/catdir/description/uchi051/91037741.html;   http://
www.loc.gov/catdir/enhancements/fy0608/91037741-t.html.

Lindeboom:1984:DMB

[3907] Gerrit Arie Lindeboom. Dutch medical biography: a biographical dictionary of Dutch physicians and surgeons, 1475–1975. Rodopi, Amsterdam, The Netherlands, 1984. ISBN 90-6203-676-7. xxxi + 2244 columns pp. LCCN R529.A1 L56 1984. fl 295.00.

Linguerri:2008:EPI

[3908] Sandra Linguerri and Raffaella Simili. Einstein parla italiano: itinerari e polemiche. (Italian) [Einstein speaks Italian: itineraries and controversy]. Pendragon, Bologna, Italia, 2008. ISBN 88-8342-530-8. x + 359

pp. LCCN ???? 20.00 EUR. URL http://www.gbv.de/dms/bsz/toc/
bsz283148705inh.pdf.

Long:2011:APR

[3909] Pamela Olivia Long. Artisan/practitioners and the rise of the new sciences, 1400–1600. The OSU Press Horning visiting scholars publication series. Oregon State University Press, Corvallis, OR, USA, 2011. ISBN 0-87071-609-3 (paperback), 0-87071-647-6 (e-book). xii + 196

pp. LCCN Q124.97 .L66 2011. URL http://muse.jhu.edu/books/
9780870716478.

Luccio:2009:PPS

[3910] Riccardo Luccio. La psicologia: un profilo storico. (Italian) [Psychology: a historical profile], volume 5 of Scienze della mente. GLF, Ed. Laterza, Roma, Italia, 2009. ISBN 88-420-6139-5. 307 pp. LCCN BF. URL

http://www.gbv.de/dms/casalini/00/00670294.pdf.

Lupacchini:2018:JNM

[3911] Rossella Lupacchini. John von Neumann: Metamatematica hilbertiana e fondamenti della meccanica quantistica. (Italian) [John von Neumann: Hilbertian metamathematics and foundations of quantum mechanics], volume 11 of Saggi di scienze e filosofia naturale. Bibliopolis, Napoli, Italia, 2018. ISBN 88-7088-661-1. 178 pp. LCCN ????

Luthy:2011:SMC

[3912] Christoph Lüthy and Sven Dupré. Silent Messengers The Circulation of Material Objects of Knowledge in the Early Modern Low Countries, volume 1 of Low countries studies on the circulation of natural knowledge. LIT Verlag, Münster, Westfalia, Germany, 2011. ISBN 3-8258-1635-4.

400 pp. LCCN ???? URL http://d-nb.info/1009621440/04.

MacGregor:2000:MCE

[3913] Arthur MacGregor, Melanie Mendonça, Julia White, Moira Hook, et al. Manuscript catalogues of the early museum collections, 1683–1886 [Ashmolean Museum Oxford], volume 907, 1569 of BAR International series. Archaeopress, Oxford, UK, 2000–2006. ISBN 1-84171-206-X (part 1: paperback), 1-85444-153-1 (part 1: paperback: Ashmolean Museum reprint), 1-84171-937-4 (part 2: paperback). ???? pp. LCCN Z6621.O95 M33 2000.

Maffioli:1994:GSW

[3914] Cesare Sergio Maffioli. Out of Galileo: the science of waters 1628–1718, volume 49 of Nieuwe Nederlandse bijdragen tot de geschiedenis der geneeskunde en der natuurwetenschappen. Erasmus Publ., Rotterdam, The Netherlands, 1994. ISBN 90-5235-071-X. xx + 509 pp. LCCN ????

Maffioli:2010:DAA

[3915] Cesare S. Maffioli. La via delle acque (1500–1700): appropriazione delle arti e trasformazione delle matematiche, volume 4 of Hyperchen. Leo S. Olschki, Firenze, Italia, 2010. ISBN 88-222-6008-2 (paperback). xxii + 394 + 1 pp. LCCN TC145 .M3155 2010. URL http://

catdir.loc.gov/catdir/toc/casalini09/10482326.pdf;   http:/
/www.ilibri.casalini.it/toc/10482326.pdf.

Maines:1998:TOH

[3916] Rachel P. Maines. The technology of orgasm: “hysteria,” the vibrator, and women’s sexual satisfaction, volume 24 (N.S.) of Johns Hopkins studies in the history of technology. The Johns Hopkins University Press, Baltimore, MD, USA, 1998. ISBN 0-8018-5941-7, 0-8018-6646-4 (paperback). xviii + 181 pp. LCCN HQ29 .M35 1998. URL http:

//www.gbv.de/dms/hbz/toc/ht010399283.pdf;   http://www.loc.
gov/catdir/bios/jhu051/98020213.html;   http://www.loc.gov/
catdir/description/jhu052/98020213.html;   http://www.loc.gov/
catdir/toc/jhu051/98020213.html.

Mandelbrote:2001:FLI

[3917] Scott Mandelbrote. Footprints of the lion: Isaac Newton at work: exhibition at Cambridge University Library, 9 October 2001–23 March 2002. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-902205-58-7. 142 pp. LCCN ????

Mangani:1998:MAO

[3918] Giorgio Mangani. Il “mondo” di Abramo Ortelio: misticismo, geografia e collezionismo nel Rinascimento dei Paesi Bassi. (Italian) [The “world” by Abraham Ortelius: mysticism, geography and collecting in the Renaissance of the Netherlands], volume 5 of Saggi. Franco Cosimo Panini, Modena, Italia, 1998. ISBN 88-7686-977-8. 332 + 123 pp. LCCN CB401 .M37 1998.

Marchant:2010:DHY

[3919] Josephine Marchant. Decoding the heavens: a 2,000-year-old-computer — and the century-long search to discover its secrets. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 2010. ISBN 0-306-81861-2 (paperback). 328 + 8 pp. LCCN QB107 .M38 2010. URL

http://www.decodingtheheavens.com/reviews.aspx.

Markowski:1990:AAC

[3920] Mieczysław Markowski. Astronomica et astrologica Cracoviensia ante annum 1550. (Latin) [Astronomical and astrological Krakow after the year 1550], volume 20 of Studi e testi / Istituto nazionale di studi sul Rinascimento. L. S. Olschki, Firenze, Italia, 1990. ISBN 88-222-3743-9. ISSN 0394-4409. xxxv + 377 pp. LCCN QB41; Z5152 M37 1990. L92000.

Marks:1997:PES

[3921] Harry M. Marks. The progress of experiment: science and therapeutic reform in the United States, 1900–1990. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1997. ISBN 0-521-58142-7, 0-521-78561-8 (paperback). xi + 258 pp.

LCCN RM47.U6 M37 1997. URL http://www.gbv.de/dms/bs/toc/
315530383.pdf;   http://www.gbv.de/dms/hbz/toc/ht008274711.
pdf;   http://www.loc.gov/catdir/description/cam028/96038997.
html;   http://www.loc.gov/catdir/toc/cam026/96038997.html.

Marks:2003:WIM

[3922] Jonathan (Jonathan M.) Marks. What It Means to be 98% Chimpanzee: Apes, People, and Their Genes. University of California Press, Berkeley, CA, USA, 2003. ISBN 0-520-24064-2 (paperback). xxi + 312 pp. LCCN ????

Martin:2011:RMP

[3923] Craig Martin. Renaissance meteorology: Pomponazzi to Descartes. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 1-4214-0187-8 (hardcover), 1-4214-0244-0 (e-book). viii + 213

pp. LCCN QC855 .M37 2011. URL http://muse.jhu.edu/books/
9781421402444/.

Martin:2012:MLF

[3924] Michaël Martin. La magie dans l’Antiquité. (French) [Magic in Antiquity]. Ellipses, Paris, France, 2012. ISBN 2-7298-7304-X (paperback). 159 pp. LCCN BF1591 M339.2012. URL http://catalogue.bnf.fr/

ark:/12148/cb43618320h.

Masani:1996:CNS

[3925] Alberto Masani. La cosmologia nella storia: fra scienza, religione e filosofia. (Italian) [Cosmology in history: between science, religion and philosophy]. Analisi e sintesi. La Scuola, Brescia, Italy, 1996. ISBN 88-350-9212-4. 410 pp. LCCN QB981 .M267 1996.

Mazumdar:1995:SSI

[3926] Pauline M. H. (Pauline Margaret Hodgson) Mazumdar. Species and specificity: an interpretation of the history of immunology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-43172-7 (hardcover). xiii + 457 pp. LCCN QR182 .M39 1995.

URL http://www.gbv.de/dms/bowker/toc/9780521431729.pdf;
http://www.loc.gov/catdir/description/cam025/93031219.html;
 http://www.loc.gov/catdir/samples/cam034/93031219.html;   http:
//www.loc.gov/catdir/toc/cam023/93031219.html.

Mazzarello:1998:GAV

[3927] Paolo Mazzarello. Il genio e l’alienista: la visita di Lombroso a Tolstoj. (Italian) [The genius and the psychiatrist: a visit of Lombroso to Tolstoy], volume 59 of Saggi Bibliopolis. Bibliopolis, Napoli, Italia, 1998. ISBN 88-7088-301-9. 94 pp. LCCN PG3400 .M38 1998.

McConnell:1998:LTJ

[3928] Anita McConnell. The life and times of J. J. Hicks (1837–1916): ‘King of the Clinicals’. William Sessions, York, UK, 1998. ISBN 1-85072-204-8. ix + 150 pp. LCCN ????

Meyerson:1998:DRI

[3929] Mile Meyerson and Carlo Vinti. La deduzione relativistica. (Italian) [Relativistic deduction]. Istituti editoriali e poligrafici internazionali, Pisa, Italia, 1998. ISBN 88-8147-136-1. LCCN ????

Michel:1992:CTM

[3930] Alain Michel. Constitution de la théorie moderne de l’intégration. (French) [Constitution of the modern theory of integration]. Mathesis. J. Vrin, Paris, France, 1992. ISBN 2-7116-1064-0. 338 pp. LCCN ????

Moore:2002:NBM

[3931] Gregory Moore and Friedrich Wilhelm Nietzsche. Nietzsche, biology, and metaphor. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-81230-5. viii + 228 pp. LCCN B3318.M4 M66 2002. URL http:

//site.ebrary.com/lib/academiccompletetitles/home.action;
http://site.ebrary.com/lib/alltitles/docDetail.action?docID=
10014986;   http://www.gbv.de/dms/bowker/toc/9780521812306.
pdf;   http://www.gbv.eblib.com/patron/FullRecord.aspx?p=202300;
 http://www.loc.gov/catdir/description/cam022/2001043121.html;
 http://www.loc.gov/catdir/enhancements/fy0731/2001043121-b.
html;   http://www.loc.gov/catdir/samples/cam031/2001043121.
html;   http://www.loc.gov/catdir/toc/fy02/2001043121.html.

Moquin-Tandon:1999:NPN

[3932] Alfred Moquin-Tandon and Jean-Louis Fischer. Un naturaliste à Paris, 1834: notes sur mon séjour à Paris, mon premier voyage à Paris en 1834, du 4 septembre au 23 octobre. (French) [A naturalist in Paris, 1834: notes on my stay in Paris, my first trip to Paris in 134, from 4 September to 23 October]. Collection SenS de l’histoire. Sciences en situation, Chilly-Mazarin, France, 1999. ISBN 2-908965-11-9. xxxiv + 163 pp. LCCN QH31.M55 A3 1999.

Morabito:1996:CCP

[3933] Carmela Morabito. La cartografia del cervello: il problema delle localizzazioni cerebrali nell’opera di David Ferrier, fra fisiologia, psicologia e filosofia. (Italian) [The cartography of the brain: the problem of cerebral localization in the work of David Ferrier from physiology, psychology and philosophy], volume 35 of Collana di psicologia Ricerche. F. Angeli, Milano, Italy, 1996. ISBN 88-204-9708-5 (L45000). 234 pp. LCCN R.

Moretti:2004:RVM

[3934] Cesare Moretti, Carlo Stefano Salerno, and Sabina Tommasi Ferroni. Ricette vetrarie muranesi: Gasparo Brunoro e il manoscritto di Danzica. (Italian) [Recipes of Murano glassmaking: Gasparo Brunora and the manuscript of Gdansk]. Arte e restauro Fonti. Nardini. ICR, Instituto Centrale per il Restauro, Firenze/Roma, Italia, 2004. ISBN 88-404-5066-

1. 206 pp. LCCN TP856. 22.00 EUR.

Morrell:1997:SCP

[3935] Jack Morrell. Science, culture and politics in Britain, 1750–1870, volume 567 of Variorum collected studies series. Variorum [u.a.], Aldershot, Great Britain, 1997. ISBN 0-86078-633-1. ???? pp. LCCN Q127.G4 M67

1997. URL http://www.gbv.de/dms/bowker/toc/9780860786337.
pdf;   http://www.gbv.de/dms/goettingen/227326482.pdf.

Mothu:2000:RSL

[3936] Alain Mothu and Antonella Del Prete. Révolution scientifique et libertinage. (French) [Scientific Revolution and libertinage], volume 48 = N.S., 11 of De diversis artibus. Brepols, Turnhout, Belgium, 2000. ISBN 2-503-51008-6. 312 pp. LCCN Q125.2 .R49 2000.

Mucciarelli:2000:VII

[3937] Roberta Mucciarelli, Laura Vigni, and Donatella Fabbri. Vergognosa immunditia: igiene pubblica e privata a Siena dal Medioevo all’età contemporanea. (Italian) [Disgraceful uncleanliness: public and private hygiene in Siena from the Middle Ages to the present]. Sienambiente, Siena, Italia, 2000. ISBN 88-7145-169-4. 223 pp. LCCN RA508.S56 M83 2000.

Muller-Jahncke:1996:GAG

[3938] Wolf-Dieter Müller-Jahncke, Christoph Friedrich, and Julian Paulus. Geschichte der Arzneimitteltherapie. (German) [History of drug therapy]. Deutscher Apotheker-Verlag, Stuttgart, Germany, 1996. ISBN 3-7692-2038-2. 295 pp. LCCN RM42 .M85 1996. DM 78.00, sfr 78.00, S

569.00. URL http://www.gbv.de/dms/bs/toc/216750520.pdf.

Munoz:2004:IAG

[3939] Jeronimo Munoz and Victor Navarro. Introducción a la Astronomia y la Geografia. Collectio oberta, Valencia, Spain, 2004. ISBN 84-482-3709-9. 354 pp. LCCN ????

Nauert:2012:HRC

[3940] Charles G. Nauert. Humanism and Renaissance Civilization. Variorum collected studies series. Ashgate Publishing Ltd, Farnham, Surrey, UK, 2012. ISBN 1-4094-2593-2, 1-4094-2326-3, 1-4094-3331-5. ???? pp. LCCN ????

NavarroBrotons:1999:BPM

[3941] Víctor Navarro Brotons et al. Bibliographia physico-mathematica hispanica (1475–1900) (Spanish) [Physical-mathematical Spanish bibliography], volume 56 of Cuadernos Valencianos de historia de la medicina y de la ciencia. Serie C, Repertorios bio-bibliográficos. Instituto de Historia de la Ciencia y Documentación “López Piänero,” Universitat de València-C.S.I.C., Valencia, España, 1999. ISBN 84-370-4219-4 (vol. 1). ???? pp. LCCN Z6651 .B53 1999; Z7141 .B53 1999.

NavarroBrotons:1999:LFS

[3942] Víctor Navarro Brotons. Libros y folletos, 1475–1600. (Spanish) [Books and brochures], volume 1; 56 of Bibliographia physico-mathematica hispanica (1475 - 1900); Cuadernos valencianos de historia de la medicina y de la ciencia Repertorios bio-bibliogrf̀icos. Instituto de Historia de la Ciencia y Documentación “López Piñero,” Universitat de València-C.S.I.C, Valencia, España, 1999. ISBN 84-370-4219-4. 374 pp. LCCN

???? URL http://www.zentralblatt-math.org/zmath/en/search/?
an=1002.01013.

Neeley:2001:MSS

[3943] Kathryn Angelyn Neeley. Mary Somerville: science, illumination, and the female mind. Cambridge science biographies. Cambridge University Press, Cambridge, UK, 2001. ISBN 0-521-62299-9, 0-521-62672-2 (paperback). xvi + 263 pp. LCCN Q143.S7 N44 2001. URL http://site.

ebrary.com/lib/academiccompletetitles/home.action;   http:/
/site.ebrary.com/lib/alltitles/docDetail.action?docID=10005053;
 http://www.gbv.de/dms/bowker/toc/9780521626729.pdf;   http:
//www.gbv.de/dms/goettingen/320849988.pdf;   http://www.loc.
gov/catdir/description/cam021/00049363.html;   http://www.loc.
gov/catdir/toc/cam025/00049363.html.

Neu:1997:ICBa

[3944] John Neu. Isis cumulative bibliography 1986–1995: a bibliography of the history of science formed from the annual Isis current bibliographies, Vol. 3, Subjects. Time periods: antiquity through 18th century. Science History Publications, Canton, MA, USA, 1997. ISBN 0-88135-133-4. xv + 620 pp. LCCN ????

Neu:1997:ICBb

[3945] John Neu, editor. Isis cumulative bibliography, 1986–1995: a bibliography of the history of science formed from the annual Isis current bibliographies. Science History Publications, Canton, MA, USA, 1997. ISBN 0-88135-131-8 (vol. 1), 0-88135-132-6 (vol. 2), 0-88135-133-4 (vol. 3), 0-88135-134-2 (vol. 4). xxx + 1362 (4 volumes) pp. LCCN Z7405.H6 I2 1997 Suppl. 3.

Newman:1991:SPP

[3946] William Royall Newman. The Summa perfectionis of Pseudo-Geber: a critical edition, translation, and study, volume 35 of Collection de travaux de l’Académie Internationale d’Histoire des Sciences. Brill, Leiden, The Netherlands, 1991. ISBN 90-04-09464-4. ISSN 0169-7897. iii + 785 pp. LCCN QD25 .G36713 1991. hfl 375.00. URL http://www.gbv.de/dms/

bs/toc/016109562.pdf.

Nicolaidis:2011:SEO

[3947] Efthymios Nicolaïdis. Science and Eastern Orthodoxy: from the Greek fathers to the age of globalization. Medicine, science, and religion in historical context. The Johns Hopkins University Press, Baltimore, MD, USA, 2011. ISBN 1-4214-0298-X (hardcover), 1-4214-0426-5 (e-book). xviii + 252 + 16 pp. LCCN BX342.9.S35 N55 2011. URL http://muse. jhu.edu/books/9781421404264/. Translated from the original Greek by Susan Emanuel.

Nicolini:1999:CVS

[3948] Nicoletta Nicolini and Gigliola Terenna. La collezione di vetreria scientifica. (Italian) [The collection of scientific glassware], volume 3 of Materiali / Patrimonio Storico-Scientifico dell’Università degli Studi di Siena. Nuova Immagine Editrice, Siena, Italia, 1999. ISBN 88-7145-160-0. 192 pp. LCCN QD53 .C54 1999.

Paganini:1998:FDS

[3949] Gianni Paganini. La filosofia della seconda metà del Novecento. (Italian) [The philosophy of the second half of the Twentieth Century]. Storia della filosofia. Piccin, Padova, Italia, second edition, 1998. ISBN 88-299-1487-8. xxxviii + 1762 (2 volumes) pp. LCCN B87 .S76 vol. 11, etc.

Pais:1991:NBT

[3950] Abraham Pais. Niels Bohr’s times: in physics, philosophy, and polity. Clarendon Press, New York, NY, USA, 1991. ISBN 0-19-852048-4 (paperback), 0-19-852049-2. xvii + 565 pp. LCCN QC773 .P35 1991. US$35.00.

URL http://www.loc.gov/catdir/enhancements/fy0635/90027248-
d.html.

Panconesi:1992:LMS

[3951] Emiliano Panconesi, Lorenzo Marri Malacrida, and Eugenio Garin. Lorenzo il Magnifico in salute e in malattia. Loggia Rucellai. A. Bruschi, Firenze, Italia, 1992. 65 pp. LCCN DG737.9 .P34 1992. Introduction by Eugenio Garin.

Paoli:2012:LAA

[3952] Michel Paoli and Monica Preti Hamard. L’Arioste et les arts: actes du colloque d’études. Louvre Éditions, Paris, France, 2012. ISBN 2-35031-368-9 (Musée du Louvre), 88-89854-68-5 (Officina libraria). 335

pp. LCCN PQ4586.A7 A75 2012. URL http://www.ilibri.casalini.
it/toc/11978067.pdf.

Payne:2012:TCT

[3953] Alina Alexandra Payne. The telescope and the compass: Teofilo Gallaccini and the dialogue between architecture and science in the age of Galileo, volume 393 of Biblioteca dell’ “Archivum Romanicum”. Serie 1, Storia, letteratura, paleografia. L. S. Olschki, Firenze, Italia, 2012. ISBN 88-222-6122-4. ISSN 0066-6815. xix + 240 pp. LCCN NA2543.S35 P39 2012.

PicaCiamarra:1996:LSI

[3954] Leonardo Pica Ciamarra. L’antropologia di Schopenhauer. (Italian) [The anthropology of Schopenhauer], volume 14 of Definizioni. Loffredo, Napoli, Italia, 1996. ISBN 88-8096-448-8. 191 pp. LCCN B3149.M3 P53 1996.

Piccolino:2003:RTS

[3955] Marco Piccolino and Marco Author Bresadola. Rane, torpedini e scintille: Galvani, Volta e l’elettricità animale. (Italian) [Frogs, torpedoes and sparks: Galvani, Volta and animal electricity]. Saggi, Scienze. Bollati Boringhieri, Torino, Italia, 2003. ISBN 88-339-1470-4. 722 + 8 pp. LCCN QP341. 42.00 EUR.

Piccolino:2005:ZCD

[3956] Marco Piccolino. Lo zufolo e la cicala: divazioni galileiane tra la scienza e la sua storia. (Italian) [The flute and the grasshopper: divisions between Galilean science and its history]. Saggi. Scienze. Bollati Boringhieri, Torino, Italia, 2005. ISBN 88-339-1612-X. 359 + 4 pp. LCCN ????

Pick:1989:FDE

[3957] Daniel Pick. Faces of degeneration: a European disorder, c. 1848–c. 1918, volume 15 of Ideas in context. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-36021-8, 0-521-45753-X (paperback). viii + 275 pp. LCCN HM111 .P54 1989.

URL http://www.gbv.de/dms/bowker/toc/9780521360210.pdf;
http://www.loc.gov/catdir/description/cam023/89031452.html;
 http://www.loc.gov/catdir/toc/cam022/89031452.html.

Pietsch:2012:TLV

[3958] Théodore W. Pietsch. Trees of Life — a Visual History of Evolution. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0479-6 (hardcover). xi + 358 pp. LCCN QH361 .P53 2012.

Podgorny:2012:CCR

[3959] Irina Podgorny. Charlatanes: crónicas de remedios incurables. Colección Nuestra América. Eterna Cadencia Editora, Buenos Aires, Argentina, 2012. ISBN 987-1673-67-1 (paperback). 348 pp. LCCN R730 .C43 2012.

URL http://core.cambeiro.com.ar/0-168920-0.pdf.

Poirier:2002:AA

[3960] Jean-Paul Poirier and Anthony Turner. Antoine d’Abbadie, volume 2 of Mémoire de la science. Académie des Sciences, Paris, France, 2002. ISBN 2-909344-09-6. 127 pp. LCCN ????

PompeoFaracovi:2012:SAA

[3961] Ornella Pompeo Faracovi. Lo specchio alto: astrologia e filosofia fra Medioevo e prima età moderna, volume 32 of Bruniana and Campanelliana / Supplementi / Studi 11. F. Serra, Pisa, Italia, 2012. ISBN 88-6227-449-1 (paperback), 88-6227-471-8, 88-6227-544-7. ISSN 1127-6045. 205 pp. LCCN BF1676 .P66 2012. URL http://

catdir.loc.gov/catdir/toc/casalini10/12614904.pdf;   http:/
/www.ilibri.casalini.it/toc/12614904.pdf.

Pressman:1998:LRP

[3962] Jack David Pressman. Last resort: psychosurgery and the limits of medicine. Cambridge history of medicine. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-35371-8 (hardcover). xv + 555 pp. LCCN RD594 .P655 1998. URL http:

//www.gbv.de/dms/hbz/toc/ht008603175.pdf;   http://www.loc.
gov/catdir/description/cam028/97014074.html;   http://www.loc.
gov/catdir/samples/cam031/97014074.html;   http://www.loc.gov/
catdir/toc/cam026/97014074.html.

Principe:2013:SA

[3963] Lawrence Principe. The Secrets of Alchemy. Synthesis. University of Chicago Press, Chicago, IL, USA, 2013. ISBN 0-226-68295-1 (hardcover), 0-226-92378-9 (e-book). v + 281 + 8 pp. LCCN QD13 .P75 2013.

Pseudo-Democritus:2011:SAC

[3964] Pseudo-Democritus and Matteo Martelli. Scritti alchemici: con il commentario di Senesio. (Italian) [Alchemical writings: with commentary by Senesio], volume 12 of Textes et travaux de Chrysopoeia. S.É.H.A., Paris, France, 2011. ISBN 88-7252-319-2. xvi + 523 pp. LCCN ????

Puccini:1999:ALV

[3965] Sandra Puccini. Andare lontano: viaggi ed etnografia nel secondo Ottocento. (Italian) [Going away: travel and ethnography in the second nineteenth century], volume 63 of Ricerche; Antropologia. Carocci, Roma, Italia, 1999. ISBN 88-430-1448-X. 297 pp. LCCN GN308.3.I8 P828 1999.

Pyenson:2002:ATP

[3966] Lewis Pyenson and Jean-François Gauvin. The art of teaching physics: the Eighteenth Century demonstration apparatus of Jean Antoine Nollet. Septentrion, Sillery, Québec, Canada, 2002. ISBN 2-89448-320-1. xvii + 220 + 16 pp. LCCN QC47.F7.

Raimondi:2003:GCV

[3967] Francesco Paolo Raimondi. Giulio Cesare Vanini dal tardo Rinascimento al libertinisme érudit: atti del convegno di studi, Lecce, Taurisano, 24-26 ottobre 1985. Testi e saggi. Congedo, Galatina, Lecce, Italias, 2003. ISBN 88-8086-497-1. 528 pp. LCCN B3578.V3 G58 2003.

Rebentrost:2006:LBT

[3968] Inken Rebentrost. Das Labor in der Box: Technikentwicklung und Unternehmensgründung in der frühen deutschen Biotechnologie. (German) [The laboratory in the box: Technology and development company founded in early German biotechnology], volume 16 of Schriftenreihe zur Zeitschrift für Unternehmensgeschichte. C. H. Beck, München, Germany, 2006. ISBN 3-406-54403-7. 309 pp. LCCN HD9999.B443 G375 2006.

Ricci:1996:LTF

[3969] Giacomo Ricci and David Rovai. L’Istituto Tecnico “Francesco Carrara” 1896–1996: cento anni di vita. (Italian) [The “Francesco Carrara” Technical Institute 1896–1996: one hundred years of life], volume 18 of Collana di cultura e storia lucchese. Maria Pacini Fazzi, Lucca, Italia, 1996. ISBN 88-7246-211-8. 246 pp. LCCN ????

Riedlberger:2013:ESW

[3970] Peter Riedlberger. Encheiridion and spurious works, volume 2 of Mathematica graeca antiqua. Fabrizio Serra editore, Pisa, Italia, 2013. ISBN 88-6227-567-6 (paperback). 279 pp. LCCN B517 .D6513 2013. URL

http://www.ilibri.casalini.it/toc/13016439.pdf.

Robinson:2002:LWJ

[3971] Annabel Robinson. The life and work of Jane Ellen Harrison. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2002. ISBN 0-19-924233-X. xvi + 332 pp. LCCN PA85.H33 R63 2002. URL

http://www.gbv.de/dms/bowker/toc/9780199242337.pdf;   http:
//www.loc.gov/catdir/enhancements/fy0611/2001055731-d.html;
 http://www.loc.gov/catdir/enhancements/fy0611/2001055731-t.
html;   http://www.loc.gov/catdir/enhancements/fy0725/2001055731-
b.html.

Rotzoll:2000:PSV

[3972] Maike Rotzoll. Pierleone da Spoleto: vita e opere di un medico del Rinascimento. (Italian) [Pierleone of Spoleto: the life and works of a doctor of the Renaissance], volume 187 of Studi / Accademia Toscana di Scienze e Lettere “La Colombaria”. L. S. Olschki, Firenze, Italia, 2000. ISBN 88-222-4842-2. ISSN 0065-0781. 127 pp. LCCN R154.P54 R68 2000.

Rovida:1999:DSA

[3973] Edoardo Rovida. Dallo scalpello al mouse: evoluzione e storia del disegno tecnico. (Italian) [From the chisel to the mouse: evolution and history of technical drawing]. Paravia scriptorium, Torino, Italia, 1999. ISBN 88-395-8300-9. 141 pp. LCCN TS174.4 .R68 1999.

Rowe:2007:EPH

[3974] David E. Rowe and Robert J. Schulmann, editors. Einstein on politics: his private thoughts and public stands on nationalism, Zionism, war, peace, and the bomb. Princeton University Press, Princeton, NJ, USA, 2007. ISBN 0-691-12094-3. xxxiv + 523 pp. LCCN QC16.E5 E5157 2007. URL http://www.loc.

gov/catdir/enhancements/fy0704/2006100303-b.html;   http://
www.loc.gov/catdir/enhancements/fy0704/2006100303-d.html;
http://www.loc.gov/catdir/toc/ecip076/2006100303.html.

Ruse:1996:MMC

[3975] Michael Ruse. Monad to man: the concept of progress in evolutionary biology. Harvard University Press, Cambridge, MA, USA, 1996. ISBN 0-674-58220-9. x + 628 pp. LCCN QH360.5 .R87 1996. URL http://

www.gbv.de/dms/bowker/toc/9780674582200.pdf.

Savoia:2002:RIU

[3976] Andrea Ubrizsy Savoia and Anes Farkas. Rapporti italo-ungheresi nella nascita della botanica in Ungheria. (Italian) [Italian–Hungarian reports on the birth of botany in Hungary]. Università degli studi di Roma “La Sapienza”, Roma, Italia, 2002. ISBN 963-641-906-X (paperback). 374 + 28 pp. LCCN ????

Scheifele:1996:FVD

[3977] Irene Scheifele. Festschriften der Versammlungen Deutscher Naturforscher und Ärzte 1822 bis 1920. (German) [Festschrifts of the Collections of German Scientists and Physicians 1822 to 1920], volume 6 of Schriftenreihe zur Geschichte der Versammlungen Deutscher Naturforscher und Ärzte. Wissenschaftliche Verlagsgesellschaft, Stuttgart, Germany, 1996. ISBN 3-8047-1050-6. 106 pp. LCCN QH149 .S328 1996. DM 48.00, sfr 48.00, S 375.00. URL http://www.gbv.de/dms/bs/

toc/09006190X.pdf.

Schnitzler:1981:TGJ

[3978] Arthur Schnitzler, Peter Michael Braunwarth, Werner Welzig, et al. Tagebuch 1879–1931. (German) [Journal 1879–1931]. Verlag der Österreichischen Akademie der Wissenschaften, Wien, Austria, 1981-2000. ISBN 3-7001-0395-6 (set), 3-7001-2120-2 (1927–1930), 3-7001-2121-0 (1931). ???? pp. LCCN PT2638.N5 Z47 1981. URL http://purl.ox.

ac.uk/uuid/28423c8f9cbf43038e6be57b41bada2d.

Schofield:1997:EJP

[3979] Robert E. Schofield. The enlightenment of Joseph Priestley: a study of his life and work from 1733 to 1773. Pennsylvania State University Press, University Park, PA, USA, 1997. ISBN 0-271-01662-0. xii + 305 pp. LCCN BX9869.P8 S36 1997. URL http://www.psupress.org/

books/titles/0-271-01662-0.html.

Segonds:2008:GAQ

[3980] Alain Philippe Segonds and Nicholas Jardine. La guerre des astronomes: la querelle au sujet de l’origine du système géo-héliocentrique à la fin du XVIe siècle, volume 9–10 of Science et humanisme. Les Belles Lettres, Paris, France, 2008. ISBN 2-251-34512-4 (II), 2-251-34513-2 (I). ???? (3 vol. en 2 t.) pp. LCCN ????

Seibold:1992:WBJ

[3981] Ilse Seibold and Johannes Associated name Walther. Der Weg zur Biogeologie: Johannes Walther (1860–1937); ein Forscherleben im Wandel der deutschen Universität. (German) [The road to biogeology: Johannes Walther (1860–1937); a researcher’s life in the changing German university]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-55341-X. viii + 196 pp. LCCN ????

Seligardi:2002:LIC

[3982] Raffaella Seligardi. Lavoisier in Italia: la comunità scientifica italiana e la rivoluzione chimica. (Italian) [Lavoisier in Italy: the Italian scientific community and the Chemical Revolution], volume 46 of Biblioteca di storia della scienza. L. S. Olschki, Firenze, Italia, 2002. ISBN 88-222-5150-4. ISSN 0394-5065. 410 pp. LCCN QD18.I8 S45 2002. URL

http://www.gbv.de/dms/casalini/02/02708914.pdf.

Shapin:1996:SR

[3983] Steven Shapin. The Scientific Revolution. Science.culture. University of Chicago Press, Chicago, IL, USA, 1996. ISBN 0-226-75020-5 (hardcover), 0-226-75021-3 (paperback). xiv + 218 pp. LCCN Q125 .S5166 1996.

URL http://site.ebrary.com/lib/academiccompletetitles/home.
action;   http://site.ebrary.com/lib/alltitles/docDetail.action?
docID=10230048;   http://www.h-net.org/review/hrev-a0a1e6-aa;
 http://www.loc.gov/catdir/description/uchi051/96013196.html;
 http://www.loc.gov/catdir/toc/uchi051/96013196.html.

Shapin:2008:SR

[3984] Steven Shapin. The Scientific Revolution. science * culture. University of Chicago Press, Chicago, IL, USA, 2008. ISBN 0-226-75022-1, 0-226-75020-5. 236 pp. LCCN Q125.S5166 1996.

URL http://gbv.eblib.com/patron/FullRecord.aspx?p=408583;
http://www.gbv.eblib.com/patron/FullRecord.aspx?p=408583.

Sheller:2014:ADM

[3985] Mimi Sheller. Aluminum dreams: the making of light modernity. Mobility studies. MIT Press, Cambridge, MA, USA, 2014. ISBN 0-262-02682-1 (hardcover). xii + 367 pp. LCCN TA480.A6 .S48 2014.

Simili:2001:PSCa

[3986] Raffaella Simili and Giovanni Paolini. Per una storia del Consiglio nazionale delle ricerche. (Italian) [A history of the National Research Council], volume 20–21 of Storia delle istituzioni scientifiche in Italia e in Europa; Percorsi. GLF editori Laterza, Roma, Italia, 2001. ISBN 88-420-5929-3 (vol. 1), 88-420-6225-1 (vol. 2). ???? pp. LCCN Q180.I8 P47 2001.

Simili:2001:PSCb

[3987] Raffaella Simili and Giovanni Paoloni. Per una storia del Consiglio nazionale delle ricerche, vol. 1, …. (Italian) [A history of the National Research Council, vol. 1], volume 20 of Percorsi; Storia delle istituzioni scientifiche in Italia e in Europa. Editori Laterza, Roma, Italia, 2001. ISBN 88-420-5929-3. x + 669 pp. LCCN ???? URL http://www.gbv.

de/dms/goettingen/358614791.pdf.

Simili:2001:PSCc

[3988] Raffaella Simili and Giovanni Paoloni. Per una storia del Consiglio nazionale delle ricerche, vol. 2, …. (Italian) [A history of the National Research Council, vol. 2], volume 21 of Percorsi; Storia delle istituzioni scientifiche in Italia e in Europa. Editori Laterza, Roma, Italia, 2001. ISBN 88-420-6225-1. x + 675 pp. LCCN ???? URL http://www.gbv.

de/dms/goettingen/358615232.pdf.

Simili:2001:TDN

[3989] Raffaella Simili and Ulisse Associated name Aldrovandi. Il teatro della natura di Ulisse Aldrovandi. (Italian) [The theater of nature of Ulisse Aldrovandi]. Compositori, Bologna, Italia, 2001. ISBN 88-7794-272-X. 143 pp. LCCN QH46.5 .T43 2001.

SistoEdreira:1999:PHC

[3990] Rafael Sisto Edreira. O patrimonio histórico-científico do Instituto Xelmírez I (Santiago de Compostela): inventario e catalogación, unha ollada ós gabinetes de ciencias do vello instituto de Santiago. (Galician) [The historical-scientific heritage of the Xelmírez I (Santiago de Compostela): inventory and cataloging, a look at the cabinets of the old Science Institute of Santiago]. Deputación Provincial da Coruña, Coruña, Spain, 1999. ISBN 84-89652-87-2. 191 pp. LCCN Q185.7 .S57 1999.

Solomos:2000:TRR

[3991] John Solomos and Les Back, editors. Theories of race and racism: a reader. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2000. ISBN 0-415-15671-8 (hardcover), 0-415-15672-6 (paperback). xxiv + 646 pp. LCCN HT1521 .T473 1999.

Souffrin:2012:ECH

[3992] P. Souffrin, Michel Blay, Francesco Furlan, and Michela Malpangotto. Écrits choisis d’histoire des sciences. (French) [Selected writings from the history of science], volume 38 of L’âne d’or. Les Belles Lettres, Paris, France, 2012. ISBN 2-251-42047-9 (paperback). ISSN 1158-9035. 497 pp. LCCN Q126.8 .S68 2012.

Soule:2014:BUE

[3993] Emily Berquist Soule. The Bishop’s Utopia: envisioning improvement in colonial Peru. The early modern Americas. University of Pennsylvania Press, Philadelphia, 2014. ISBN 0-8122-4591-1 (hardcover). 287 + 14 pp. LCCN F3611.T8 S68 2014.

Stafford:1999:VAC

[3994] Barbara Maria Stafford. Visual analogy: consciousness as the art of connecting. MIT Press, Cambridge, MA, USA, 1999. ISBN 0-262-19421-X. xvii + 219 pp. LCCN BD190 .S78 1999. URL http://www.gbv.de/

dms/bowker/toc/9780262194211.pdf.

Stafford:2001:VAC

[3995] Barbara Maria Stafford. Visual analogy: consciousness as the art of connecting. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-19421-X, 0-262-69267-8 (paperback). xvii + 219 pp. LCCN BD190. URL

http://www.gbv.de/dms/bowker/toc/9780262194211.pdf.

Staubermann:2011:RRS

[3996] K. B. (Klaus B.) Staubermann, editor. Reconstructions: recreating science and technology of the past. NMS Enterprises, Edinburgh, Scotland, 2011. ISBN 1-905267-48-7 (paperback). xii + 276 pp. LCCN T15 .R33 2011.

Stephenson:2000:UUI

[3997] Bruce Stephenson, Marvin Bolt, and Anna Felicity Friedman. The Universe unveiled: instruments and images through history. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-79143-X (hardcover). 152 pp. LCCN QB86 .S74 2000. URL

http://www.loc.gov/catdir/description/cam021/00059883.html;
 http://www.loc.gov/catdir/enhancements/fy0731/00059883-b.html;
 http://www.loc.gov/catdir/toc/cam021/00059883.html.

Stern:1999:EGW

[3998] Fritz Richard Stern. Einstein’s German world. Princeton University Press, Princeton, NJ, USA, 1999. ISBN 0-691-05939-x. ix + 335 pp. LCCN DD239 .S74 1999. URL http:/

/www.gbv.de/dms/bowker/toc/9780691059396.pdf;   http://www.
loc.gov/catdir/description/prin021/99020128.html;   http://
www.loc.gov/catdir/toc/prin032/99020128.html.

Stichweh:1984:EMS

[3999] Rudolf Stichweh. Zur Entstehung des modernen Systems wissenschaftlicher Disziplinen: Physik in Deutschland, 1740–1890. (German) [On the emergence of the modern system of scientific disciplines: physics in Germany, 1740–1890]. Suhrkamp, Frankfurt am Main, West Germany, 1984. ISBN 3-518-57688-7. 558 pp. LCCN QC9.G3 S75 1984.

Stoffel:1996:PDS

[4000] Jean-François Stoffel. Pierre Duhem et ses doctorands: bibliographie de la littérature primaire et secondaire. (French) [Pierre Duhem and his doctoral candidates: bibliography of the primary and secondary literature]. Centre Interfacultaire d’Étude en Histoire des Sciences, Louvain-la-Neuve, Belgium, 1996. ISBN 2-930175-00-1. 325 pp. LCCN QC16.D925 S76 1996.

Sutton:1995:SPS

[4001] Geoffrey V. Sutton. Science for a polite society: gender, culture, and the demonstration of enlightenment. Westview Press, Boulder, CO, USA, 1995. ISBN 0-8133-1575-1 (hardcover), 0-8133-1576-X (paperback). xiii + 391 pp. LCCN Q125 .S982 1995.

URL http://www.gbv.de/dms/goettingen/18561518X.pdf;   http:
//www.loc.gov/catdir/enhancements/fy0832/95020157-b.html;
http://www.loc.gov/catdir/enhancements/fy0832/95020157-d.html.

Taisbak:2003:DED

[4002] C. M. (Christian Marinus) Taisbak and Euclid. Dedomena–Euclid’s Data; or, The importance of being given, volume 45 of Acta historica scientiarum naturalium et medicinalium. Museum Tusculanum Press, Copenhagen, Denmark, 2003. ISBN 87-7289-815-1. ISSN 0065-1311. 271 pp. LCCN QA31 .T349 2003.

Taton:1989:PAR

[4003] René Taton and Curtis Wilson, editors. Planetary astronomy from the Renaissance to the rise of astrophysics, volume 2 of General history of astronomy. Cambridge University Press, Cambridge, UK, 1989-1995. ISBN 0-521-24254-1 (pt. A), 0-521-35168-5 (pt. B). ???? pp. LCCN QB15 .G38 1984 vol. 2. URL

http://www.loc.gov/catdir/description/cam023/88025817.html;
 http://www.loc.gov/catdir/toc/cam022/88025817.html.

Taton:1995:GHA

[4004] René Taton, Curtis Wilson, and Michael Hoskin. The general history of astronomy, Vol. 2. Planetary astronomy from the Renaissance to the rise of astrophysics. Part B. The Eighteenth and Nineteenth Centuries, volume 2. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-35168-5. xiii + 281 pp. LCCN ???? URL http:/

/www.gbv.de/dms/bowker/toc/9780521351683.pdf;   http://www.
zentralblatt-math.org/zmath/en/search/?an=0859.01001.

Tenon:1996:MPH

[4005] Jacques Tenon and Dora B. Weiner, editors. Memoirs on Paris hospitals. Resources in medical history. Science History Publications, USA, Canton, MA, USA, 1996. ISBN 0-88135-074-5. ISSN 1044-9477. xxxiii + 405 pp. LCCN RA989.F74 P3732813 1996.

Terenna:1997:SMS

[4006] Gigliola Terenna and Francesca Vannozzi. Strumenti medici dei secoli XVIII–XIX. (Italian) [Medical tools of the 18th and 19th Centuries], volume 1 of Mater.Iali; Materiali / Patrimonio Storico-Scientifico dell’Università degli Studi di Siena. Nuova Immagine Editrice, Siena, Italia, 1997. ISBN 88-7145-141-4. 78 + 47 pp. LCCN R.

Terenna:1998:CDS

[4007] Gigliola Terenna and Francesca Vannozzi. La collezione degli strumenti di psicologia. (Italian) [The collection of tools of psychology], volume 2 of Materiali / Patrimonio Storico-Scientifico dell’Università degli Studi di Siena. Nuova Immagine Editrice, Siena, Italia, 1998. ISBN 88-7145-146-5. 78 pp. LCCN ????

Tobias:1991:OGVa

[4008] Phillip V. Tobias. Olduvai Gorge. Volume 2. The cranium and maxillary dentition of australopithecus. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-20072-5. 263 pp. LCCN ???? URL http://www.

gbv.de/dms/bowker/toc/9780521200721.pdf.

Tobias:1991:OGVb

[4009] Phillip V. (Phillip Vallentine) Tobias. Olduvai Gorge. Volume 4. The skulls, endocasts and teeth of Homo habilis. Cambridge University Press, Cambridge, UK, 1991. ISBN 0-521-20072-5. xxxv + 921 + 104 (two parts) pp. LCCN GN283.9; GN282 O43 vol. 4. URL http://www.gbv.

de/dms/bowker/toc/9780521200721.pdf.

Tognotti:2002:SIS

[4010] E. (Eugenia) Tognotti. La spagnola in Italia: storia dell’influenza che fece temere la fine del mondo: 1918–1919. (Italian) [The Spanish flu in Italy: the history of influenza that made one fear the end of the world:

1918–1919], volume 303 of FrancoAngeli storia. F. Angeli, Milano, Italy, 2002. ISBN 88-464-3692-X. 165 pp. LCCN RA644.I6 T64 2002.

Tomory:2012:PEO

[4011] Leslie Tomory. Progressive enlightenment: the origins of the gaslight industry, 1780–1820. Transformations: studies in the history of science and technology. MIT Press, Cambridge, MA, USA, 2012. ISBN 0-262-01675-3 (hardcover), 0-262-30124-5 (e-book). xi + 348 pp. LCCN TP733.G6 .T66 2012. US$19.95. URL http://mitpress-ebooks.mit.

edu/product/progressive-enlightenment.

Tropp:1993:AFM

[4012] Eduard Abramovich Tropp, Viktor Yakovlevich Frenkel’, and Artur Davidovich Chernin. Alexander A. Friedmann: the man who made the universe expand. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-38470-2 (hardback), 0-521-02588-5 (paperback). x + 267 pp. LCCN QC16.F73 T7613 1993. URL http://site.ebrary.

com/lib/alltitles/docDetail.action?docID=10459828;   http:/
/www.gbv.de/dms/bowker/toc/9780521384704.pdf;   http://www.
loc.gov/catdir/description/cam025/92028315.html;   http://www.
loc.gov/catdir/toc/cam028/92028315.html.

Turner:2000:EIM

[4013] Gerard L’Estrange Turner. Elizabethan instrument makers: the origins of the London trade in precision instrument making. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-856566-6. xiv + 305 + 4 pp. LCCN Q185 .T854 2000. URL http:/

/www.gbv.de/dms/goettingen/067605869.pdf;   http://www.loc.
gov/catdir/enhancements/fy0611/2001268813-d.html;   http://
www.loc.gov/catdir/enhancements/fy0611/2001268813-t.html;
http://www.loc.gov/catdir/enhancements/fy0723/2001268813-b.
html.

Valerio:1998:PVN

[4014] Vladimiro Valerio and Ermanno Bellucci. Piante e vedute di Napoli dal 1486 al 1599: l’origine dell’iconografia urbana europea (Italian) [Plants and views of Naples from 1486 to 1599: the origin of the European urban iconography], volume 1 of Imago urbis. Electa Napoli, Napoli, Italia, 1998. ISBN 88-435-8725-0. 95 pp. LCCN DG842.5 .V35 1998.

Valleriani:2013:MBE

[4015] Matteo Valleriani. Metallurgy, Ballistics and Epistemic Instruments: The Nova scientia of Nicolò Tartaglia. A New Edition, volume 6 of Max Planck Research Library for the History and Development of Knowledge / Sources. Ed. Open Access, Berlin, Germany, 2013. ISBN 3-8442-5258-4 (e-book). 350 pp. LCCN ???? URL http://

d-nb.info/1034002716/34;   http://nbn-resolving.de/urn:nbn:
de:101:1-2013042611605;   http://www.edition-open-access.de/
sources/6/index.html;   http://www.epubli.de/shop/isbn/9783844252583.

Vallisnieri:2004:QO

[4016] Antonio Vallisnieri, Concetta Pennuto, Dario Generali, and Andrea Castellani. Quaderni di osservazioni, volume sezione I, v. 1 of Edizione nazionale delle opere di Antonio Vallisneri. L.S. Olschki, Firenze, Italia, 2004. ISBN 88-222-5331-0. ???? pp. LCCN ????

vanDamme:1997:ICS

[4017] Jacques van Damme, Koenraad van Cleempoel, and Gerard L’Estrange Turner. Instrumentos cientificos del siglo XVI: la Corte española y la Escuela de Lovaina. (Spanish) [Scientific instruments of the 16th Century: the Spanish Court and the School of Louvain]. Fundación Carlos de Amberes, Madrid, Spain, 1997. ISBN 84-87369-03-0. 248 pp. LCCN ????

vanMiert:2013:COE

[4018] Dirk van Miert, editor. Communicating observations in early modern letters (1500–1675): epistolography and epistemology in the age of the scientific revolution, volume 23 of Warburg Institute Colloquia. Warburg Institute, London, UK, 2013. ISBN 1-908590-46-7 (paperback). ISSN 1352-9986. ix + 268 pp. LCCN Q223.2 .C66 2013.

Vetrano:1996:GFD

[4019] Flavio Vetrano. Il Gabinetto di fisica dell’Università di Urbino: la sua storia, il suo museo. (Italian) [The old “cabinet of physics” of the University of Urbino]. Istituto Poligrafico e Zecca dello Stato, Libreria dello Stato, Roma, Italia, 1996. ISBN 88-240-3867-0. 263 pp. LCCN QC9.I8 G33 1996.

Vickers:1996:FB

[4020] Brian Vickers, editor. Francis Bacon. Oxford authors. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1996. ISBN 0-19-254198-6 (hardcover), 0-19-282025-7 (paperback). xii + 813 pp. LCCN PR2205

.V53 1996. URL http://www.loc.gov/catdir/enhancements/fy0638/
95016588-d.html;   http://www.loc.gov/catdir/enhancements/fy0724/
95016588-b.html.

Vinti:1997:SQG

[4021] Carlo Vinti. Il soggetto qualunque: Gaston Bachelard fenomenologo della soggettività epistemica (Italian) [The subject of whatever: Gaston Bachelard’s phenomenology of epistemic subjectivity], volume 6 of Ricerche filosofiche. Ed. Scientifiche Italiane, Napoli, Italia, 1997. ISBN 88-8114-538-3. 1010 pp. LCCN B.

Volta:1999:EEM

[4022] Alessandro Volta and Joseph Banks. On the electricity excited by the mere contact of conducting substances of different kinds. Collana di storia della scienza. Università degli studi di Pavia, Pavia, Italia, 1999. ISBN 88-203-2762-7. 99 pp. LCCN QC517 .V925 1999. Bicentenary in French, English, German and Italian of letter to Sir Joseph Banks of the 20th of March 1800.

vonHaller:2000:CFC

[4023] Albrecht von Haller and Maria Teresa Monti. Commentarius de formatione cordis in ovo incubato. (Latin) [The history of the formation of the heart in the incubated egg], volume 6 of Studia Halleriana. Schwabe, Basel, Switzerland, 2000. ISBN 3-7965-1324-7. clxxvii + 677 pp. LCCN ????

vonSchelling:1988:IPN

[4024] Friedrich Wilhelm Joseph von Schelling, Errol E. Harris, and Peter Lauchlan Heath. Ideas for a philosophy of nature as introduction to the study of this science 1797, second edition 1803. Cambridge University Press, Cambridge, UK, 1988. ISBN 0-521-32102-6, 0-521-35733-0. xxvi + 294 pp. LCCN BD581.

Vucinich:1988:DRT

[4025] Alexander Vucinich. Darwin in Russian thought. University of California Press, Berkeley, CA, USA, 1988. ISBN 0-520-06283-3. x + 468 pp.

LCCN B1623 .V83 1989. URL http://ark.cdlib.org/ark:/13030/
ft5290063h;   http://ark.cdlib.org/ark:/13030/ft5290063h/;
http://www.gbv.de/dms/hbz/toc/ht004355431.PDF;   http://www.
loc.gov/catdir/bios/ucal051/88002054.html;   http://www.loc.
gov/catdir/description/ucal041/88002054.html.

Weigel:2003:PMT

[4026] Erhard Weigel and Thomas Behme. Philosophia Mathematica, Theologia Naturalis Solida. (Latin) [Philosophy of Mathematics, Strong Natural Theology], volume 3 of Clavis pansophiae. Frommann-Holzboog, Stuttgart-Bad Cannstatt, Germany, 2003–2013. ISBN 3-7728-1625-8 (vol. 1), 3-7728-2540-0 (vol. 4, pts. 12). ???? pp. LCCN B785 .W3 2003.

Weigel:2008:AAE

[4027] Erhard Weigel and Thomas Behme. Analysis Aristotelica ex Euclide restituta. (Latin) [Analysis of Aristotle and Euclid restored], volume 3,3 of Werke; Clavis Pansophiae. Frommann-Holzboog, Stuttgart/Bad Cannstatt, Germany, 2008. ISBN 3-7728-2375-0. lxv + 355 pp. LCCN

???? URL http://www.zentralblatt-math.org/zmath/en/search/?
an=1143.01007.

Weigel:2013:PMS

[4028] Erhard Weigel and Thomas Behme. Philosophiae mathematicae, secunda disciplina specialis, purae matheseos pars altera: videlicet archimetria, seu expressioris usus communissimi geometria. (Latin) [Philosophy of mathematics, the second discipline, a part of pure mathematics that was altered, that is to say, archimetria, or expressing the most common uses of geometry], volume 4,2; 3,4,2 of Werke / Erhard Weigel; Clavis Pansophiae. Frommann-Holzboog, Stuttgart-Bad Cannstatt, Germany, 2013. ISBN 3-7728-2540-0. 436 pp. LCCN ???? URL http://d-nb.

info/1041725892/04.

Weiner:1993:CPR

[4029] Dora B. Weiner. The citizen-patient in revolutionary and imperial Paris. The Henry E. Sigerist series in the history of medicine. The Johns Hopkins University Press, Baltimore, MD, USA, 1993. ISBN 0-8018-4483-5 (hardcover), 0-8018-7002-X (paperback). xvi + 444 pp. LCCN R506.P3 W45 1993. URL http:/

/www.gbv.de/dms/bowker/toc/9780801844836.pdf;   http://www.
loc.gov/catdir/bios/jhu051/92049007.html;   http://www.loc.
gov/catdir/description/jhu051/92049007.html.

Wellman:1992:MMP

[4030] Kathleen Anne Wellman. La Mettrie: medicine, philosophy, and enlightenment. Duke University Press, Durham, NC, USA, 1992. ISBN 0-8223-1204-2 (hardcover). xiv + 342 pp. LCCN R507.L193 W46 1992.

URL http://www.gbv.de/dms/bowker/toc/9780822312048.pdf.

Williams:1994:PMH

[4031] Mari E. W. Williams. The precision makers: a history of the instruments industry in Britain and France, 1870–1939. Comparative and international business. Modern histories series. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1994. ISBN 0-415-03732-8. viii + 216 pp. LCCN HD9706.6.G72 W55 1994. URL http://www.gbv.de/

dms/bowker/toc/9780415037327.pdf.

Wilson:1990:CBS

[4032] David B. Wilson, editor. The Correspondence between Sir George Gabriel Stokes and Sir William Thomson, Baron Kelvin of Largs. Volume 1. 1846–1869. Volume 2. 1870–1901. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-32831-4. lvi + ix + 783 pp. LCCN QC16.K3.

Wilson:1993:SMA

[4033] John Elbert Wilson. Schellings Mythologie: zur Auslegung der Philosophie der Mythologie und der Offenbarung. (German) [Schelling’s mythology: the interpretation of the philosophy of mythology and revelation], volume 31 of Spekulation und Erfahrung. Abteilung II, Untersuchungen. Frommann-Holzboog, Stuttgart, Germany, 1993. ISBN 3-7728-1483-2. 292 pp. LCCN B2899.M9 W55 1993. DM 138.00, sfr 134.20, S 1076.40.

Wilson:2000:PTM

[4034] Renate Wilson. Pious traders in medicine: a German pharmaceutical network in Eighteenth-Century North America. Pennsylvania State University Press, University Park, PA, USA, 2000. ISBN 0-271-02052-0 (hardcover). xiv + 258 pp. LCCN R151 .W54 2000. URL http://www.

gbv.de/dms/hbz/toc/ht012905173.pdf.

Wolf:1996:APJ

[4035] Evemarie Wolf. Über die Anfänge der Pharmaziegeschichtsschreibung: von Johannes Ruellius (1529) bis David Peter Hermann Schmidt (1835). (French) [On the beginnings of writing of the history of pharmacy: John Ruellius (1529) to David Peter Hermann Schmidt (1835)], volume 72 of Quellen und Studien zur Geschichte der Pharmazie. Wiss. Verl.-Ges., Stuttgart, Germany, 1996. ISBN 3-8047-1447-1. x + 109 pp. LCCN ???? ca. DM 28.00, ca. sfr 28.00, ca. S 219.00. URL http://www.gbv.

de/dms/bs/toc/191629731.pdf.

Wolfschmidt:1995:MNG

[4036] Gudrun Wolfschmidt. Milchstraße, Nebel, Galaxien: Strukturen im Kosmos von Herschel bis Hubble. (German) [Milky Way, nebula, galaxies: structure in the Cosmos from Herschel to Hubble], volume N.F., 11 of Abhandlungen und Berichte / Deutsches Museum. R. Oldenbourg, München, Germany, 1995. ISBN 3-486-26308-0. 186 pp. LCCN AM101

.M9743 n.F., Bd. 11. URL http://www.gbv.de/dms/goettingen/
183368347.pdf.

Yabuuti:2000:HMC

[4037] Kiyosi Yabuuti. Une histoire des mathématiques chinoises. (French) [A history of Chinese mathematics]. Regards sur la science. Belin, Pour la science, Paris, France, 2000. ISBN 2-7011-2404-2. ISSN 0993-4812. 191 pp. LCCN ????

Zagli:2000:MBS

[4038] Andrea Zagli, Francesco Mineccia, and Andrea Giuntini. Maledetti beccari: storia dei macellai fiorentini dal Cinquecento al Duemila. (Italian) [Damn Beccari: history of butchers from the Florentine 500s to 2000s]. Polistampa, Firenze, Italia, 2000. ISBN 88-8304-267-0. 293 pp. LCCN HD9425.I83 F59 2000.

Abir-Am:1998:MMS

[4039] Pnina G. Abir-Am, editor. La mise en mémoire de la science: pour une ethnographie historique des rites commémoratifs. (French) [The storage of science: toward a historical ethnography of memorial rites], Histoire des sciences, des techniques et de la médecine. Éditions des archives contemporaines, Amsterdam, The Netherlands, 1998. ISBN 90-5709-007-4. ISSN 1027-4030. LCCN Q175.4 .M57 1998.

Albertazzi:1999:FSL

[4040] Liliana Albertazzi, Guido Cimino, and Simonetta Gori-Savellini, editors. Francesco De Sarlo e il laboratorio fiorentino di psicologia: [atti del Convegno “Francesco De Sarlo e il laboratorio fiorentino di psicologia”, tenuto a Pisa presso la Domus Galilaeana nei giorni 18–20 aprile 1994]. (Italian) [Francesco De Sarlo and the Florentine workshop of psychology: [proceedings of the conference held in Pisa at the Domus Galilaeana on 18–20 April 1994]], volume 5 of Collana del Seminario di storia della scienza. G. Laterza, Bari, Italia, 1999. ISBN 88-8231-075-2. LCCN BF41 .C67 1994.

Amaldi:1998:CPE

[4041] Edoardo Amaldi, Giovanni Battimelli, and Giovanni Paoloni, editors. 20th Century Physics: Essays and Recollections: a Selection of Historical Writings, volume 3 of Edoardo Amaldi Foundation series. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1998. ISBN 981-02-2369-2. xiii + 747 pp. LCCN Q127.I8

A53 1998. URL http://www.worldscientific.com/worldscibooks/
10.1142/2852.

Anonymous:1993:SMA

[4042] Anonymous, editor. Les scientifiques et la montagne: actes du 116e Congrès national des sociétés savantes, Chambéry, 29 avril–4 mai 1991, Section d’histoire des sciences et des techniques. (French) [Scientists and the mountain], Actes du 116e Congrès national des sociétés savantes. Section d’histoire des sciences et des techniques. Editions du CTHS, Paris, France, 1993. ISBN 2-7355-0272-4. ISSN 0998-5654. LCCN Q127.E8 C663 1991.

Anonymous:1999:SRC

[4043] Anonymous, editor. Sciences et religions de Copernic à Galilée, 1540– 1610: actes du colloque international organisé par l’Ecole française de Rome, en collaboration avec l’Ecole nationale des chartes et l’Istituto italiano per gli studi filosofici, avec la participation de l’Università di Napoli “Federico II”, Rome, 12–14 décembre 1996. (French) [Sciences and religions from Copernicus to Galileo, 1540–1610], volume 260 of Collection de l’École Française de Rome. École Française de Rome, Rome, Italy, 1999. ISBN 2-7283-0545-5. LCCN BL241 .S345 1999. URL http://

digital.casalini.it/2728305455.

BassoPeressut:1997:SDM

[4044] Luca Basso Peressut, editor. Stanze della meraviglia: i musei della natura tra storia e progetto. (Italian) [Rooms of wonder: the museums of nature between history and project], volume 6 of Lexis. IV, MuseoPoli. CLUEB, Bologna, Italia, 1997. ISBN 88-8091-519-3. LCCN QH70.A1 S735 1997.

Battistini:2001:SCT

[4045] Pierluigi Battistini and Bartholomaeus de Parma, editors. Seventh centenary of the teaching of astronomy in Bologna, 1297–1997: proceedings of the meeting held in Bologna at the Accademia delle Scienze on June 21, 1997, volume 8 of Musei e archivi dello Studio bolognese. CLUEB, Bologna, Italia, 2001. ISBN 88-491-1650-0. LCCN QB23 .S48 2001.

L55000. URL http://digital.casalini.it/8849116500;   http://
www.gbv.de/dms/goettingen/337289905.pdf.

Bauer:1995:RNT

[4046] Martin W. Bauer, editor. Resistance to new technology: nuclear power, information technology, and biotechnology. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-45518-9 (hardback), 0-521-59948-2 (paperback). LCCN T174.5 .R48 1995. URL http://site.

ebrary.com/lib/alltitles/docDetail.action?docID=10461834;
http://www.gbv.de/dms/bowker/toc/9780521455183.pdf;   http:

//www.gbv.de/dms/hbz/toc/ht006664769.pdf;   http://www.loc.
gov/catdir/description/cam026/94026745.html;   http://www.loc.
gov/catdir/toc/cam021/94026745.html. Papers from a Three-Day

Conference held at the Science Museum, London, 5–7 April 1993.

Beaune:1992:BAD

[4047] Jean-Claude Beaune and Jean Gayon, editors. Buffon 88: actes du Colloque international pour le bicentenaire de la mort de Buffon: (Paris, Montbard, Dijon, 14–22 juin 1988), Science, histoire, philosophie. J. Vrin, Paris, France, 1992. ISBN 2-7116-9755-X. LCCN QH31.B88 B84 1988.

Biagioli:2003:SAC

[4048] Mario Biagioli and Peter Galison, editors. Scientific authorship: credit and intellectual property in science. Routledge & Kegan Paul, London, UK and New York, NY, USA, 2003. ISBN 0-415-94292-6 (hardcover), 0-415-94293-4 (paperback). LCCN K1401 .S35 2003.

URL http://www.gbv.de/dms/goettingen/360109691.pdf;   http:
//www.loc.gov/catdir/enhancements/fy0651/2003276190-d.html.

Blanco:2000:AFP

[4049] Luigi Blanco, editor. Amministrazione, formazione e professione: gli ingegneri in Italia tra Sette e Ottocento. (Italian) [Administration, training and profession: the engineers of Italy from the 18th to 19th Centuries], volume 52 of Annali dell’Istituto Storico Italo–Germanico in Trento. Quaderni. Soc. Ed. Il Mulino, Bologna, Italia, 2000. ISBN 88-15-07665-4. LCCN TA79 .A66 2000. Atti del Convegno “Ingegneri, Publica Amministrazione e Istruzione Tecnico-Scientifica in Italia dall’Eta Napoleonica all’Unificazione Nazionale”; Trento, 24–25 novembre 1995.

Bonoli:2007:PDM

[4050] Fabrizio Bònoli, Giuseppe Bezza, Salvo De Meis, and Cinzia Colavita, editors. I pronostici di Domenico Maria da Novara, volume 69 of Biblioteca di Nuncius. Studi e testi. Leo S. Olschki, Firenze, 2007. ISBN 88-222-6216-6. ISSN 1122-0910. vii + 317 pp. LCCN BF1680 .N68 2012.

Bowers:1991:CPS

[4051] Brian Bowers and Lenore Symons, editors. Curiosity perfectly satisfyed: Faraday’s travels in Europe 1813–1815, volume 16 of IEE history of technology series. Peter Peregrinus in association with the Science Museum, London, UK, 1991. ISBN 0-86341-234-3. xvii + 168 pp. LCCN QC16.F2 C87 1991.

Bresadola:1999:LGI

[4052] Marco Bresadola and Giuliano Pancaldi, editors. Luigi Galvani International Workshop: Proceedings; Bologna, 9 October 1998, volume 7 of Bologna studies in history of science. Università di Bologna, Dipartimento di Filosofia, Centro Internazionale per la Storia delle Università e della Scienza, Bologna, Italia, 1999. ISBN 88-900162-1-3. LCCN MLCS 2004/03284.

Bret:1999:LEE

[4053] Patrice Bret, editor. L’expedition d’Egypte, une entreprise des Lumières 1798–1801: actes du colloque international. (French) [The Egyptian Expedition: an enterprise of the Lights 1798–1801: Proceedings of the international congress], Actes de colloques de l’Académie des Sciences. Technique and Documentation, Paris, France, 1999. ISBN 2-7430-0319-7. LCCN PJ1060.F7 E97 1999.

Brown:1989:PQPa

[4054] Laurie M. Brown, Max Dresden, and Lillian Hoddeson, editors. Pions to quarks: particle physics in the 1950s; based on a Fermilab symposium; [based on the lectures … at the 2. International Symposium on the History of Particle Physics, held at Fermilab on May 1–4, 1985], volume 2 of International symposium on the history of particle physics. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-30984-0. LCCN QC793.16 .P56 1989.

URL http://www.gbv.de/dms/bowker/toc/9780521309844.pdf;
http://www.loc.gov/catdir/description/cam023/88025644.html;
 http://www.loc.gov/catdir/toc/cam028/88025644.html.

Brown:1989:PQPb

[4055] Laurie M. Brown, Lillian Hoddeson, Max Dresden, and May West, editors. Pions to quarks: particle physics in the 1950s; based on a Fermilab symposium. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-30984-0. LCCN QC793.16. URL http://site.ebrary.

com/lib/alltitles/docDetail.action?docID=10460129;   http:/
/www.gbv.de/dms/bowker/toc/9780521309844.pdf.

Burkhardt:1985:CCD

[4056] Frederick Burkhardt and Sydney Smith, editors. The correspondence of Charles Darwin. Cambridge University Press, Cambridge, UK, 1985– 2010. ISBN 0-521-25587-2 (vol. 1), 0-521-25588-0 (vol. 2), 0-521-25589-9 (vol. 3), 0-521-25590-2 (vol. 4), 0-521-25591-0 (vol. 5), 0-521-25586-4 (vol. 6), 0-521-84459-2 (vol. 14), 0-521-85931-X (vol. 15: hardback), 0-521-88195-1 (vol. 16, pt. 1: hardback), 0-521-88196-X (vol. 16, pt. 2: hardback), 0-521-51836-9 (vol. 16, 2 vol. hardback set), 0-521-19030-4 (vol. 17), 0-521-76889-6 (vol. 18), 0-521-44241-9. ???? pp. LCCN QH31.D2 A4 1985.

Costantini:1997:PDC

[4057] Domenico Costantini and Maria Carla Galavotti, editors. Probability, dynamics, and causality: essays in honour of Richard C. Jeffrey. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4361-1. LCCN QA273.A1 P7634 1997. URL

http://www.gbv.de/dms/bowker/toc/9780792343615.pdf;   http:
//www.loc.gov/catdir/enhancements/fy0822/97165191-d.html;
http://www.loc.gov/catdir/enhancements/fy0822/97165191-t.html;
 http://www.zentralblatt-math.org/zmath/en/search/?an=0899.
00035.

deClercq:2000:SIO

[4058] P. R. de (Peter Richard) Clercq, editor. Scientific instruments: originals and imitations; the Mensing connection: proceedings of a symposium, held at the Museum Boerhaave, Leiden, 15–16 October 1999, Communication 286. Museum Boerhaave, Leiden, The Netherlands, 2000. ISBN 90-6292-129-9. LCCN ????

DeGennaro:2002:QUD

[4059] Giuseppe De Gennaro, editor. I quattro universi di discorso: atti del Congresso internazionale Orationis millennium: L’Aquila, 24–30 giugno 2000. (Italian) [The four universes of discourse: proceedings of the International Congress Orationis millennium: L’Aquila, 24–30 June 2000], Collana di mistica. Libreria editrice vaticana, Città del Vaticano, 2002. ISBN 88-209-7242-5. LCCN MLCM 2006/03330 (B).

deMedici:1990:L

[4060] Lorenzo de’ Medici and Bernard Toscani, editors. Laude, volume 21 of Studi e testi / Istituto nazionale di studi sul Rinascimento. L. S. Olschki, Firenze, Italia, 1990. ISBN 88-222-3743-9. ISSN 0394-4409. 122 pp. LCCN PQ4630.M3 L38 1990.

Dilg:1993:RDP

[4061] Peter Dilg and Hartmut Rudolph, editors. Resultate und Desiderate der Paracelsus-Forschung. (German) [Results and desiderata of the Paracelsus research], volume 31 of Sudhoffs Archiv. Beihefte. Steiner, Stuttgart, Germany, 1993. ISBN 3-515-06096-0. ISSN 0341-0773. LCCN R147.P2 R46 1993. DM 78.00. URL http://

www.dandelon.com/servlet/download/attachments/dandelon/ids/
AT001A033B5C3CFB4B3E0C125709D003A61CF.pdf.

Droscher:1996:ZII

[4062] Ariane Dröscher and Christoph J. Scriba, editors. Die Zellbiologie in Italien im 19. Jahrhundert: mit 18 Tabellen. (German) [Cell Biology in Italy in the 19th Century: with 18 tables], volume 26 of Acta historica Leopoldina. Deutsche Akademie der Naturforscher Leopoldina, Halle (Saale), German, 1996. ISBN 3-335-00500-7. ISSN 0001-5857. 184 pp. LCCN QH578.I8 D76 1996. ca. DM 28.00. URL http://www.gbv.de/

dms/bs/toc/216750008.pdf.

Durand:2012:OMH

[4063] Aline Durand, editor. Des outils, des machines et des hommes: études offertes à Georges Comet. (French) [Tools, machines, and men: studies offered to Georges Comet], volume 8 of Cahier d’Histoire des Techniques. Presses universitaires de Provence, Aix-en-Provence, France, 2012. ISBN 2-85399-797-9. ISSN 1158-8403. 236 pp. LCCN T14.5 .D477 2012.

Fiocca:1998:GAI

[4064] Alessandra Fiocca, editor. Giambattista Aleotti e gli ingegneri del Rinascimento. (Italian) [Giambattista Aleotti and the engineers of the Renaissance], volume 31 of Biblioteca di “Nuncius”. Studi e testi. L. S. Olschki, Firenze, Italia, 1998. ISBN 88-222-4660-8. ISSN 1122-0910.

LCCN TA140.A413 G53 1998. URL http://www.zentralblatt-math.
org/zmath/en/search/?an=0922.01014.

Fric:1955:OLC

[4065] René Fric, editor. Oeuvres de Lavoisier: correspondance. (French) [Works of Lavoisier: correspondence]. Albin Michel, Paris, France, 1955– 1997. ISBN 2-7011-1085-8 (vol. 4), 2-85367-184-4 (vol. 5), 2-85367-205-0 (vol. 6). ???? pp. LCCN QD22.L4 A4 1997.

Fritscher:1998:THM

[4066] Bernhard Fritscher and Fergus Henderson, editors. Toward a history of mineralogy, petrology, and geochemistry: proceedings of the International Symposium on the History of Mineralogy, Petrology, and Geochemistry, Munich, March 8–9, 1996, volume 23 of Algorismus; Münchener Universitätsschriften. Institut für Geschichte der Naturwissenschaften, München, Germany, 1998. ISBN 3-89241-024-0 (paperback). LCCN QE359.A1 I58 1996. URL http://www.gbv.de/dms/

goettingen/243174950.pdf.

Galbani:2001:DPA

[4067] Annamaria Galbani, editor. Donne politecniche: atti del Convegno e catalogo della Mostra: Milano, 22 maggio 2000 (Italian) [Polytechnic women: proceedings of the Conference and Catalog of the Show of Milan, 22 May 2000], volume 7 of Piccola biblioteca del Politecnico di Milano. Libri Scheiwiller, Milano, Italy, 2001. ISBN 88-7644-298-7. LCCN T36 .D66 2001.

Harman:1990:SLP

[4068] P. M. (Peter Michael) Harman, editor. The Scientific Letters and Papers of James Clerk Maxwell. Vol. I: 1846–1862, volume I. Cambridge University Press, Cambridge, UK, 1990. ISBN 0-521-25625-9 (hardcover). xxviii + 748 pp. LCCN QC21.2; QC670 .M385 1990.

URL http://www.gbv.de/dms/bowker/toc/9780521256254.pdf;
http://www.loc.gov/catdir/description/cam023/89000452.html;
 http://www.loc.gov/catdir/enhancements/fy0732/89000452-b.html;
 http://www.loc.gov/catdir/toc/cam021/89000452.html;   http:
//www.zentralblatt-math.org/zmath/en/search/?an=0744.01034.

Heering:2011:LDE

[4069] Peter Heering and Roland Wittje, editors. Learning by doing: experiments and instruments in the history of science teaching, Geschichte (Franz Steiner Verlag). Steiner, Stuttgart, Germany, 2011. ISBN 3-515-09842-9 (paperback). LCCN Q181 .L4954 2009.

Hessenbruch:2000:RGH

[4070] Arne Hessenbruch, editor. Reader’s guide to the history of science. Fitzroy Dearborn, London, UK, 2000. ISBN 1-884964-29-X. xxix + 934 pp. LCCN Q125 .R335 2000. URL http:/

/www.gbv.de/dms/bs/toc/215809416.pdf;   http://www.loc.gov/
catdir/enhancements/fy0653/2001270888-d.html;   http://www.
zentralblatt-math.org/zmath/en/search/?an=0977.01001.

Hoffmann:1999:PXI

[4071] Dieter Hoffmann, editor. Proceedings of the XXth International Congress of History of Science (Liège, 20–26 July 1997), volume t. 41–42, ¡44–46, 49–51, 53, 55–57, 60–64¿ = (n.s. 4–5, ¡7–9, 12–14, 16, 18–21, 23–27¿) of De diversis artibus. Brepols, Turnhout, Belgium, 1999–2002. ISBN 2-503-50890-1 (vol. 1), 2-503-50887-1 (vol. 2), 2-503-50891-X (vol. 4), 2-503-50992-4 (vol. 5), 2-503-51095-7 (vol. 6), 2-503-51158-9 (vol. 7), 2-503-51198-8 (vol. 8), 2-503-51268-2 (vol. 9), 2-503-51360-3 (vol. 10), 2-503-51197-X (vol. 12), 2-503-51199-6 (vol. 13), 2-503-51409-X (vol. 17), 2-503-51412-X (vol. 18), 2-503-51413-8 (vol. 19), 2-503-51414-6 (vol. 20), 2-503-51415-4 (vol. 21). LCCN Q124.6 .I67 1997.

Hoppe:1997:BIS

[4072] Brigitte Hoppe, editor. Biology integrating scientific fundamentals: contributions to the history of interrelations between biology, chemistry, and physics from the 18th to the 20th Centuries, volume 21 of Algorismus; Münchener Universitätsschriften. Institut für Geschichte der Naturwissen-Schaften, München, Germany, 1997. ISBN 3-89241-022-

4. LCCN QH315 .B62 1997. URL http://www.gbv.de/dms/bs/toc/
226181952.pdf.

James:1991:CMF

[4073] Frank A. J. L. James, editor. The correspondence of Michael Faraday. [History of technology series]. Institution of Electrical Engineers, London, UK, 1991–2012. ISBN 0-86341-248-3 (vol. 1), 0-86341-249-1 (vol. 2), 0-86341-250-5 (vol. 3), 0-86341-251-3 (vol. 4), 0-86341-823-6 (vol. 5), 0-86341-957-7 (vol. 6). ???? pp. LCCN QC16.F2 A4 1991.

James:1996:CMF

[4074] Frank A. J. L. James, editor. The correspondence of Michael Faraday, Vol. 3, 1841–December 1848, volume 3. Inst. of Electrical Engineers, London, UK, 1996. ISBN 0-86341-250-5. lvii + 835 pp. LCCN ????

URL http://www.gbv.de/dms/goettingen/220535655.pdf.

Knight:1998:MCS

[4075] David M. Knight and Helge Kragh, editors. The making of the chemist: the social history of chemistry in Europe, 1789–1914. Cambridge University Press, Cambridge, UK, 1998. ISBN 0-521-58351-9 (hardcover). xxi + 353 pp. LCCN QD39.7 .M34 1998. URL http://www.gbv.de/dms/bs/

toc/237428288.pdf;   http://www.loc.gov/catdir/description/
cam028/97043009.html;   http://www.loc.gov/catdir/toc/cam021/
97043009.html.

Maggioli:1998:ASS

[4076] Lidia Maggioli, editor. Alessandro Serpieri: scienziato riminese: Liceo scientifico statale “A. Serpieri”, Rimini, 22–23 marzo 1997. (Italian) [Alessandro Serpieri: Rimini scientist: scientific state high school “A. Serpieri”, Rimini, 22–23 March 1997]. Raffaelli e Luisè, Rimini, Italia, 1998. ISBN 88-85050-93-X. LCCN Q143.S48 A38 1998.

Margolin:1993:APR

[4077] Jean Claude Margolin and Sylvain Matton, editors. Alchimie et philosophie à la Renaissance: actes du colloque international de Tours, 4–7 décembre 1991. (French) [Alchemy and philosophy in the Renaissance: proceedings of the international conference in Tours, 4–7 December 1991], volume 57 of De Pétrarque à Descartes. J. Vrin, Paris, France, 1993. ISBN 2-7116-1172-8. ISSN 0418-4459. LCCN QD23.3 .A45 1993. 270F.

Meinel:2000:IEH

[4078] Christoph Meinel, editor. Instrument — Experiment: historische Studien; [… hatte die Deutsche Gesellschaft für Geschichte der Medizin, Naturwissenschaft und Technik zu ihrer 80. Jahrestagung im September 1997 nach Regensburg eingeladen]. (German) [Instrument — experiment: historical studies; the German Association for the History of Medicine, Science and Technology had their 80th annual meeting in September 1997 in Regensburg]. Verlag für Geschichte der Naturwissenschaft und der Technik, Berlin, Germany, 2000. ISBN 3-928186-51-

5. LCCN Q180.A3 I57 2000. URL http://www.gbv.de/dms/hebis-
darmstadt/toc/84296224.pdf.

Moore:1989:HHE

[4079] James R. (James Richard) Moore, editor. History, humanity, and evolution: essays for John C. Greene. Cambridge University Press, Cambridge, UK, 1989. ISBN 0-521-33511-6. xii + 429 pp. LCCN QH331

.H54 1989. URL http://www.dandelon.com/servlet/download/
attachments/dandelon/ids/AT0012585F66B7F1C9730C12570F6003DCE5D.
pdf;   http://www.gbv.de/dms/bowker/toc/9780521335119.pdf;
http://www.loc.gov/catdir/description/cam023/89032583.html;
 http://www.loc.gov/catdir/samples/cam031/89032583.html;   http:
//www.loc.gov/catdir/toc/cam028/89032583.html.

Moran:1991:PIS

[4080] Bruce T. Moran, editor. Patronage and institutions: science, technology and medicine at the European court, 1500–1750. Boydell Press, Woodbridge, UK; Rochester, NY, USA, 1991. ISBN 0-85115-285-6 (hardback). LCCN Q127.E8 P385 1991. US$73.00. URL http://www.gbv.de/dms/

hbz/toc/ht003777810.PDF.

Morini:1996:BMI

[4081] Luigina Morini, editor. Bestiari medievali. (Italian) [Medieval bestiary]. I millenni. Einaudi, Torino, Italia, 1996. ISBN 88-06-12446-3. xxx + 644 pp. LCCN ????

Olshausen:1998:SKH

[4082] Eckart Olshausen and Holger Sonnabend, editors. Stuttgarter Kolloquium zur Historischen Geographie des Altertums 6, 1996: “Naturkatastrophen in der Antiken Welt”. (German) [Stuttgart Conference on the Historical Geography of the Ancients 6, 1996: “natural disasters in the Ancient World”], volume 10 of Geographica historica. F. Steiner, Stuttgart, Germany, 1998. ISBN 3-515-07252-7. ISSN 1381-0472. LCCN D21.5 .S88 1996; GB5001 .S78 1998.

PompeoFaracovi:2001:FEM

[4083] Ornella Pompeo Faracovi, Luca M. Scarantino, Manlio Mazziotti, et al., editors. Federigo Enriques: matematiche e filosofia: lettere inedite, bibliografia degli scritti. (Italian) [Federigo Enriques: mathematics and philosophy: unpublished letters, bibliography of writings], volume 6 of Dimensionilibri. Belforte and C. editori, Livorno, Italia, 2001. ISBN 88-7997-063-1. LCCN QA29.E57 F44 2001.

Ruggiu:1998:FTI

[4084] Luigi Ruggiu, Luigi Alici, et al., editors. Filosofia del tempo. (Italian) [Philosophy of time], Sintesi. B. Mondadori, Milano, Italy, 1998. ISBN 88-424-9341-4. LCCN BD638 .F45 1998.

Santini:2002:LST

[4085] Carlo Santini, Ida Mastrorosa, and Antonino Zumbo, editors. Letteratura scientifica e tecnica di Grecia e Roma. (Italian) [Scientific and technical literature of Greece and Rome], volume 401 of Studi superiori; Lettere classiche. Carocci, Roma, Italia, 2002. ISBN 88-430-2142-7. xix + 604 pp. LCCN Q225.5 .L48 2002. 35.20 EUR. URL http://www.gbv.de/

dms/casalini/02/02470527.pdf.

Shea:2000:SVI

[4086] William R. Shea, editor. Science and the visual image in the Enlightenment, volume 4 of European studies in science history and the arts. Science History Publications, Canton, MA, USA, 2000. ISBN 0-88135-285-3. viii + 232 pp. LCCN Q222 .S35 2000; Q222 SCI. URL http://

www.gbv.de/dms/goettingen/318156385.pdf.

Simon:2001:OAP

[4087] Gérard Simon and Suzanne Débarbat, editors. Optics and astronomy: proceedings of the XXth International Congress of History of Science (Liège, 20–26 July 1997), volume 12 of Proceedings of the XXth International Congress of History of Science; De diversis artibus. Brepols, Turnhout, Belgium, 2001. ISBN 2-503-51197-X. LCCN ????

Stancati:2001:HBE

[4088] Claudia Stancati, Donata Chiricò, and Federica Vercillo, editors. Henri Bergson: esprit et langage. (French) [Henri Bergson: spirit and language], Philosophie et langage. Mardaga, Liège, Belgique, 2001. ISBN 2-87009-765-4. LCCN B2430.B43 H46 2001. Colloque … que l’Universitéde Calabre a organisé du 21 au 23 avril 1998.

Stoffel:1996:MGL

[4089] Jean-François Stoffel, editor. Mgr. Georges Lemaître, savant et croyant: actes du colloque commémoratif du centième anniversaire de sa naissance (Louvain-la-Neuve, le 4 novembre 1994): suivi de La physique d’Einstein: texte inédit. (French) [Monsignor Georges Lemaître, scholar and believer: proceedings of the commemorative conference on the hundredth anniversary of his birth (Louvain-la-Neuve, 4 November 1994), followed by the physics of Einstein: unpublished text]. Centre Interfacultaire d’Étude en Histoire des Sciences, Louvain-la-Neuve, Belgium, 1996. ISBN 2-930175-02-8 (paperback). LCCN ????

Stolzenberg:2001:GAK

[4090] Daniel Stolzenberg, editor. The great art of knowing: the baroque encyclopedia of Athanasius Kircher; published on the occasion of the exhibition at the Stanford University Libraries. Stanford University Press, Stanford, CA, USA, 2001. ISBN 0-911221-23-9. xii + 160 pp. LCCN Q125.2; Q143.K44 G74 2001. URL http://site.

ebrary.com/lib/academiccompletetitles/home.action;   http:/
/site.ebrary.com/lib/alltitles/docDetail.action?docID=10078367.

VanCleempoel:2002:CRS

[4091] Koenraad Van Cleempoel, editor. A catalogue raisonné of scientific instruments from the Louvain School, 1530 to 1600, volume 65 = N.S., 28 of Collection des travaux de l’Académie internationale d’histoire des sciences; De diversis artibus. Brepols, Turnhout, Belgium, 2002. ISBN 2-503-51218-6. LCCN Q185 .C58 2002. URL http://www.gbv.de/dms/

bsz/toc/bsz098173243inh.pdf.